A Bibliography of Publications in Revue d’Histoire des Sciences et de Leurs

Applications and Revue d’Histoire des Sciences

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

04 April 2024 Version 1.11

Title word cross-reference

#15882 [2400].

17 [2411]. ^{∘} [1890]. ^{e} [231, 349, 369, 896, 931, 1055, 1427, 1493, 1533, 1462, 1641, 1744, 1770, 1860, 1855, 1861, 1825, 1862, 3226, 4549, 1987, 3572, 4481, 4524, 4477, 1774, 76, 197, 1538, 3227, 386, 4512, 618, 279, 3983, 1849, 748, 4013, 658, 4468, 4410, 925, 1282, 2409, 2812, 3793, 1740, 2379, 4165, 14, 90, 25, 4173, 4003, 2295, 2396, 2422, 485, 2754, 496, 4482, 890, 1207, 1631, 1835, 2512, 2497, 625, 897, 1141, 2909, 892, 4411, 4434, 1463, 193, 257, 280, 499, 1461, 4028, 3929, 975, 1698, 647, 757, 1850, 2238, 1456, 4369, 3161, 3290]. ^{e} [2365, 2486, 4286, 4491, 3875, 4005, 4289, 4415, 4457, 3881, 4413, 3273, 3951, 4205, 759, 3569, 2900, 3462, 3867, 1379, 1500, 1662, 1699, 1788, 2045, 2218, 2577, 2044, 2266, 2267, 2251, 3993, 1402, 2014, 2472, 2582, 643, 570, 797, 838, 963, 964, 996, 1000, 1115, 1116, 1148, 1170, 1245, 1305, 1306, 1309, 1579, 1638, 1704, 1756, 2032, 2066, 2086, 2188, 2448, 2845, 2354, 2915, 2974, 3043, 3228, 3818, 4048, 433, 3779, 1772, 2469, 2828, 62, 792, 986, 1544, 291, 4337, 4403, 79, 4159, 954, 3428, 3048, 2160, 4362, 594, 2858, 2962, 1337, 1762, 3213, 4281, 3374, 4009, 4034, 2693]. ^{e} [3780, 3197, 2705, 1513, 1655, 756, 3624, 3157, 4319, 2260, 3173, 1224, 1821, 3918, 3376, 1990, 1008, 4397, 4510, 4530, 3695, 4154, 2101, 1268, 1346, 1735, 4146, 4283, 4062, 3295, 3215, 3345, 4338, 4452, 2583, 4523, 1661, 2454, 1739, 1658, 3293, 3113, 572, 568, 928, 1607, 1283, 2013, 2127, 2206, 2738, 3261, 4166, 867, 1206, 3453, 2315, 4493, 4565, 1319, 3623, 3827, 4070, 4051, 4124, 3239, 4023, 2669, 573, 1430, 1431, 1851, 2180, 333, 353, 160, 375, 654, 790, 791, 839, 860, 969, 1077, 1541, 4539, 815, 1014, 979, 1176, 1532, 1691, 2087, 2394, 869]. ^{e} [621, 848, 1052, 3013, 3244, 4558, 4333, 3530, 1615, 2939, 1584, 2246, 75, 2677, 2854, 656]. ^{e}r [897, 101, 279, 4003, 4081, 2319, 3664, 2105, 2405]. ^{me} [107]. ^{O} [1997, 1267, 1265, 1247, 1266, 164, 1973, 2308, 2320, 2868, 2870, 2942, 1752, 3359, 2189, 2275, 2276, 2343, 2345, 2422, 949, 968, 1434, 1640, 1975, 2037, 2117, 1214, 359, 1270, 1736, 3083, 443, 465, 1208, 1524, 397, 514, 2637, 2109, 797, 1003, 1061, 1302, 1306, 1307, 1311, 2323, 2393, 1847, 2603, 2453, 151, 1215, 1327, 1328, 1634, 1673, 1982, 2250, 2285, 1472, 2811, 4239, 945, 1268, 1271, 1351, 1352, 1354, 1355, 1413, 1735, 1751, 1018, 2583, 2758, 1058, 2756, 944, 3387, 2386, 2373, 353, 160, 184, 192, 240, 278, 421, 531, 836, 841, 1326, 1543]. o [1566, 1572, 2781, 221, 246, 247, 363, 1014, 1024, 1076, 2070, 2882, 3665, 953, 2246, 2614, 2854, 2855]. ^{os} [1594, 1595, 795, 592, 617, 843, 973, 1004, 1033, 1121, 1315, 1444, 1508, 1796, 2038, 2042, 2116, 2133, 328, 461, 2170, 1578, 3582, 1356, 1421, 1230, 383]. ^{r} [1459, 800, 1494, 2410, 1225]. ^{rd} [836]. ^{re} [3654, 3863, 3791, 4457, 883, 62, 3629, 4025, 1268, 3962, 4172, 4011, 1390]. _{e} [174, 386, 2584, 177, 388]. _{r} [1495]. ϵ́βδoμάδων [2320]. Δ [3148]. Πϵρὶ [2320]. S [3410]. XVI^{e} [550]. XVIII^{e} [550].

//cnum.cnam.fr [3987]. /Manuscrits [4090].

1 [163, 1035, 101, 221, 2191, 3501, 3460, 3970, 2807]. 1/2 [866, 974, 1122, 1035]. 10 [3525, 3432, 2788]. 100 [2779]. 106 [3083]. 106th [3083]. 11 [1973, 2117, 1997, 3665]. 11.-18 [769]. 111 [3968]. 111^{≫} [949]. 1177 [465]. 119-120 [2422]. 12 [3921]. 128 [4005]. 13 [696, 3964, 3871]. 134 [2583]. 136 [3533]. 137 [941]. 14 [3497]. 14-9-1769 [942]. 143 [3500]. 144 [4164]. 146 [4510]. 148 [1302]. 14th [4570]. 15 [797, 1775]. 1514- [2462]. 152 [4207]. 153 [3699]. 1550 [3438]. 1593 [2852]. 15th [1854]. 16 [4351, 1946, 945]. 1626 [3535]. 1633 [109]. 1638 [1232]. 1639 [1467]. 1641 [1667]. 1642 [1955]. 1643 [2107]. 1645 [2527]. 1646 [2658]. 1647 [2964]. 1648 [3124]. 1650 [2710]. 1662 [1522]. 1666 [1669]. 1670 [1670]. 1673 [1929]. 168 [3826]. 1684 [2057, 2096]. 1687 [1142]. 1695 [2492]. 16th [2506, 3653, 2697]. 17 [1311, 4036]. 1700 [3484]. 1709 [1671]. 1713 [2493].

1722 [2742]. 1730 [2782]. 175 [3949]. 1750 [418, 3947]. 1752 [4615]. 1762 [876]. 1763 [1332]. 1764 [1609]. 1772 [4196, 1389]. 1774 [2355]. 1783 [12]. 1786 [3122]. 1789 [3456]. 179 [3880]. 1791 [3937]. 1792 [280]. 1793 [1923, 2268]. 1794 [277]. 17th [3653, 2697]. 1800 [2478, 2131, 2840, 276, 2875]. 1800-talets [748]. 1805 [4053, 3753]. 1807 [2155]. 1809 [2808]. 1810 [3585]. 1815 [2445]. 182 [1507]. 1824 [4647]. 1827 [4080, 2879]. 1832 [3166]. 1835 [2337, 1755, 1834]. 1836 [3060]. 1838 [1869, 1936]. 1847 [4230]. 1850 [402, 1591, 2805]. 1859 [946]. 1860 [3301, 1734]. 1866 [2171, 903]. 1873 [3792]. 1875 [3003]. 1878 [2112]. 1883 [2993]. 1885 [2853]. 1889 [3797]. 1894 [1910, 731]. 1899 [381]. 18th [359, 3574, 2811]. 19 [2276, 727, 3816]. 1900 [3319, 3832, 771]. 1902 [3270]. 1906 [3384]. 1914 [3591, 3520, 3746, 2975, 2672, 3216, 4169]. 1915 [2960]. 1917 [3865]. 1918 [3567]. 1919 [2410, 3946]. 1920 [2378]. 1921 [1867]. 1927 [2750]. 1930s [4591]. 1931 [4393]. 1933 [3524]. 1936 [2313]. 1939 [4456, 3435]. 1940 [2920, 2582]. 1941 [279]. 1942 [4168]. 1946 [3666, 62, 205]. 1948 [165]. 1949 [165]. 1950 [3486, 1779, 3000, 3841, 422, 283, 477, 613]. 1951 [404, 359]. 1952 [488, 489, 459]. 1954 [667, 877]. 1955 [617, 939, 857]. 1956 [747, 665]. 1957 [941, 836, 1014]. 1958 [1079]. 1960 [4311, 1502, 1148, 3318]. 1960s [3902]. 1961 [1053, 1152, 1912, 1432, 1238]. 1961/1963 [1912]. 1962 [1294, 1208, 1303, 1246]. 1963 [1318, 1912, 1524, 1482]. 1964 [1484, 1509]. 1965 [1486, 1510, 1461, 1421]. 1966 [1594, 1835]. 1967 [1710, 1798, 2034]. 1968 [1889, 1942]. 1969 [1744, 1748, 1784, 1913]. 1970 [1822, 1823, 1843, 4616, 3580]. 1970s [3902]. 1971 [1861, 1825, 1862, 2043, 2115, 2319]. 1972 [1860, 2070]. 1973 [2323, 2604]. 1974 [2377]. 1975 [2324, 2250]. 1976 [2393]. 1977 [2448]. 1978 [2452]. 1982 [2842, 2844, 2745]. 1983 [3004, 2847]. 1987 [3299]. 1988 [3386]. 1989 [3664, 3368]. 1990 [3340, 3409]. 1993 [4261, 3965]. 1994 [4026, 3723]. 1997 [4077]. 1998 [4266]. 19th [2506, 3839, 4020, 4203, 1854]. 1re [4616, 666]. 1st [3743].

2 [2422, 1035, 1120, 2042, 2133, 96, 1749, 2192, 2041, 218, 4267, 3482, 1565, 1605, 4122, 160]. 20 [2039, 3990, 3536]. 2000 [1334]. 2001 [4317]. 2001/3 [4317]. 208 [1914]. 20th [2979]. 214 [3830]. 218 [1271]. 22 [1617]. 226/226 [1325]. 243 [4344]. 25 [2659, 613]. 26 [3870]. 27th [3965]. 29 [2400]. 2940 [3195]. 2A [3503]. 2B [3503]. 2e [4616, 1922]. 2nd [3315, 4203, 3743, 3041].

3 [972, 974, 1122, 2453, 2854, 384, 1061, 4317]. 3/4 [866, 972, 974, 1122]. 30 [2275]. 306 [4096]. 318 [3563]. 33 [3359, 4232]. 3495 [4032]. 35 [1270, 4345]. 350 [4645]. 37 [3596]. 38 [1018]. 39 [1685].

4 [888, 972, 974, 1122, 1511, 1837, 2040, 2246, 2190, 2296, 2132, 3417, 3363, 4529, 3583, 2444, 246, 2712]. 41 [843, 4006]. 42 [3755, 4098]. 436 [3754]. 44 [3497, 3677, 1263]. 45 [3035, 4198]. 49 [3679, 1033, 3860]. 4DE [2855].

5 [2262, 3584]. 50 [3769]. 50/51 [1121]. 51 [3846, 3519, 4391]. 52 [3906, 1121]. 52/53 [1121]. 53 [3974, 1121, 1259]. 54 [4038]. 55 [4100]. 56 [4148]. 57 [4212, 2400]. 58 [4273]. 5th [247, 609].

6 [164, 592, 2250, 3503, 4494, 1578, 3843]. 6-5-1859 [942]. 6/2A [3503]. 6/2B [3503]. 60 [1003]. ’600 [4083]. 61 [1315]. 63 [3411]. 65 [1444]. 651-1650 [1175]. 69 [3626]. 6th [631].

7 [3491, 3307]. ’700 [4083, 248]. 73 [1508, 3741, 1975]. 76 [1076]. 783 [4271]. 789 [4269]. 7A [3674].

8 [1709, 795, 3587]. 81 [1796].

9 [2037, 3420, 842, 2495]. 90th [2939]. 92 [3412, 4455]. 93 [2038]. 936 [4018]. 93g [3497]. 95 [4369]. 97c [3933, 3932].

= [748, 2252, 2568, 2830, 606, 2453, 601, 1105, 971]. ālāt [2372].

Âge [4491, 3462, 2267, 2828, 3842, 3924, 3626, 2272, 3741, 969, 2592, 2929, 476, 4089, 2076, 1954, 4646, 4066, 4486, 4126, 1916, 4129, 3873, 1794]. âme [4369]. être [4658]. êtres [919, 2108].

A-C [2974]. a-t-il [3355]. A-V.D.3 [3626]. A. [2441, 575, 2417, 1992, 2179, 2122, 1555, 2, 2879, 2613]. A.-C.-M [530]. A.-J [58]. A.-L [2837, 3279]. A.-M. [2879]. A.-P [4643]. A.-P. [1992]. A.D. [247]. aṡar [2372]. A/Part [3077]. år [2475]. Aarab [4614]. Aaron [1576, 821]. Aaserud [4079]. Abatage [748]. abbé [2895, 2435, 3660, 2900, 367, 777, 454, 775, 927]. Abbri [4076, 3786]. Abd [1388]. Abdelkader [751]. Abdou [3922]. Abdülhamid [1388]. abeilles [717]. Abel [828, 150, 4251, 82, 4321]. Abelian [3330]. Abendland [2843]. Aberdeen [1435]. Abetti [531, 531]. Abh [2392, 278]. Abhandlungen [3489, 3201, 2835, 1693, 1935]. able [4586]. above [359]. Abragam [2999]. Abraham [3371, 3125, 455]. abrégée [398]. abridged [3988]. absolus [241]. absolute [4348]. Abt [3503]. abus [976]. ac [4110, 529]. Academiae [1710]. Academic [2262]. académicien [4381, 3784]. académicienne [811]. Académiciens [1923]. Academie [991, 202, 1839, 4411, 1463, 1523, 1306, 1307, 1443, 1705, 151, 1084, 1826, 2954, 895, 2967, 3268, 1840]. Academies [186, 2328, 3224, 2220]. académique [4383, 4565]. Academy [3903, 3304, 3520, 2837, 608, 2954, 2837]. Acc [4076, 4435]. Accademia [2976, 2606, 4088, 4076, 2358, 4083]. accademica [4309]. accademici [3562]. accélératrice [3065, 3512]. acceptance [4669]. acceptation [4669]. accès [3987, 4505]. Accessibility [4631, 4650]. accompagnées [19]. according [4020, 2937]. accords [4567]. acédie [4199]. achievement [89]. Achille [1023, 2073]. acht [1216]. Acid [747]. acier [1588, 3208, 1558, 1558]. Acker [4291]. Ackerknecht [2031]. Acot [2104, 4360, 4032]. Acoustical [164]. Acoustics [2547]. Acoustique [78, 2995]. acque [2914]. acquisition [1880]. Acta [1710, 3536, 3755, 536, 2233, 690, 861, 628]. Actas [3592, 4409]. acte [1738]. Actes [3631, 4003, 3569, 3970, 3664, 2668, 3299, 786, 4510, 3938, 4363, 2583, 3759, 941, 3722, 3633, 2044, 3157, 160]. Acteurs [4614, 4292, 4441, 3374]. action [4447, 4357, 3600]. activité [5, 1741, 2299, 1736, 464]. activity [3814]. Actualisme [2860]. actualité [4284, 3934]. Actualités [944, 208, 465, 1024]. actuelle [2995, 3272, 3924]. actuelles [270, 366]. actuels [3547, 3047]. acupuncture [4654]. acupuncturist [4654]. acute [361]. acutis [3921]. AD [1270, 3012, 204, 2631, 2873, 4635]. Adam [742, 1056, 1233, 4502, 585]. Adanson [4538, 1624, 1084]. addenda [4198]. Addison [2127]. Addison-Wesley [2127]. Additions [3195, 1541, 627]. Adelard [4198]. Adèle [3863]. Adeline [4620]. Ader [1895, 1405]. adhibuerunt [3148]. Adieu [4639]. Adler [3996]. adli [1388]. Administrateurs [3695]. administration [2901, 1175]. administratives [4030]. admirable [595]. Adolf [2637, 2354]. Adolfo [3339]. Adolphe [10, 3345, 926, 4193]. Adoptées [349]. Adour [2583]. adressée [1188]. Adriaan [1517]. Adrian [3634, 3003]. Adrien [790]. Adrien-Marie [790]. Advance [190]. Advancement [2309]. Advances [3347]. advent [3257]. aeronomy [3347]. Aether [2918]. AFAS [4065]. afd [397]. affaire [1819, 4465, 3350, 2861]. affinité [3413]. Affinity [4316]. affranchie [2508]. Afhandling [747]. Afrique [1822]. After [3840]. afterword [4268]. against [216, 3120]. Agassiz [3389, 2498]. Agazzi [3824, 3824]. Age [3519, 2547, 4457, 2641, 2226, 2757, 4269, 3967, 1418, 2758, 3431, 4172]. Ages [1539, 3810, 1417, 3480, 2777, 589, 2692, 3275, 2113, 4238]. aggiornata [248]. Agnès [4291, 3458]. agonie [2858]. Agora [3412]. agrégat [3354]. Agricola [794, 3560, 1890, 178, 2120]. agricole [4155]. agriculture [95, 4169, 2273, 3177]. Agriculturist [3008]. Agrikultura [1597, 1355]. Agrimi [3626]. agronomes [2494, 2758]. agronomie [1380]. agronomiques [3896, 4169]. Ahjar [3650]. Ahmad [2372]. Ahmed [4614, 4164]. ai [2395, 248]. aimait [3213]. Ainsi [2738, 4166]. Ainsworth [3209]. Air [4598, 3192, 3654, 3928, 1437, 747]. Air-Pump [3192]. Aires [62, 154, 566]. Airs [3125, 2408, 3883]. Aiton [2869]. Akad [278]. Akademie [3223, 397, 942, 2392, 1323, 1332, 2355, 1028]. Akhbâr [246]. akkadien [419]. Akustik [3816]. Al [2684, 1605, 3042, 3650, 3994, 1388, 2372, 2570, 1436, 1675, 1818, 4493, 4023, 246, 2372, 1818, 4330, 3239, 2230]. al-Ahjar [3650]. Al-Akhbâr [246]. Al-Bahir [2230]. Al-Beruni [1605]. al-Dīn [2372]. al-Fārisī [1818]. al-Hasan [4330, 2372]. al-Haytham [1675, 1818, 4493, 4023, 4330]. al-hiyal [2570]. al-Jwarizmi [1436]. al-Khazin [4023]. al-Khwarizmi [3042]. Al-Kindi [2684]. al-Mutanna [1436]. al-qarn [2372]. al-Quhi [4023]. al-rūhāniyya [2372]. al-sadis [2372]. al-Samh [4023]. al-saniyya [2372]. al-Turuq [2372]. Àla [2588]. Alain [4179, 4615, 3103, 3676, 4033, 2770, 1000, 1579, 1869, 2035, 3299, 4643, 3300, 4296, 4145, 3038, 3049]. Alain-Philippe [4643]. Alan [3836, 3012]. Albert [3270, 3663, 3026, 3793, 9, 35, 95, 212, 4527, 1665, 138, 465, 3761, 3348, 3749, 4090, 1505, 2975, 964, 1997, 1328, 1683, 3096, 2105, 2405, 2444, 1145, 1632, 1684, 1242, 1243, 3865, 3032, 3033, 3031, 4590, 3029, 3256, 4638, 1519, 2542, 2746, 1391]. Albert-I [2105, 2405]. Alberti [4088, 4015]. Alberto [3368]. Alberts [3443]. Albertus [1908]. Albin [301, 4611]. albores [3732]. Albrecht [3833, 4243, 4202, 3418, 1413, 4566, 2449]. Alchemia [3316]. Alchemie [4121, 3650]. alchemiegeschichtliche [2695]. alchemist [3650]. alchemistische [2473]. Alchemists [3652]. Alchemy [1004, 1912, 2472, 3655, 3651, 3080, 746, 3653]. Alchimie [3645, 4491, 4555, 3641, 4337, 3644, 3643]. alchimique [3309, 3648]. alchimiques [3648, 2592]. alchimistes [3309, 945, 3962]. Alchimists [532]. Alcmaeon [2172]. alcohol [4364]. alcoolique [890]. alcooliques [1061]. Alcouffe [3103]. Aldini [2653]. Aldini-Valeriani [2653]. Aldo [1630, 417, 566]. Alec [3528]. Aleid [2534]. alejandrina [1907]. Alejandro [1108, 1108]. Alembert [4120]. Alessandri [3562]. Alessandro [3941, 831]. Alex [4254]. Alexander [3610, 1997, 942, 3570, 1171, 2829, 4354, 4358, 1236, 942, 3570, 1060, 1106, 1107, 1238]. Alexander-von-Humboldt-Forschung [2829]. Alexandra [3577]. Alexandre [585, 1288, 785, 1536, 1607, 926, 3375, 2532, 1396, 1009, 1393, 3920, 1395]. Alexandre-Édouard [3920]. Alexandrie [4003, 3294]. Alexandrov [2058]. Alexeï [3721]. Alexis [2410, 2894, 3658]. Alexis-Claude [3658]. Alexis-Marie [2894]. Alfred [89, 65, 4022, 2065, 3864, 3762, 1472, 2292, 4541, 4445, 4056]. Alfred-Louis-Armand-Marie [2065]. Alfredo [4081]. Algarotti [3365]. Algebra [1863, 3042, 1097, 1388, 2875, 4203, 4635, 830, 3078, 1672, 2214]. algèbre [3198, 1683, 4635, 867, 2222, 3133, 2469, 1683, 3239, 2230, 42]. algebriche [189]. algébrique [2213, 3328]. algébriques [375]. Algérie [2599, 1002, 3935]. Algorithme [2404, 2210]. algorithmique [4545]. algorithmisation [4013]. Algunos [62]. Alhazen [1675, 4015]. Aliénisme [4173]. alimentaire [4531]. Alix [2186]. alkahest [3646, 3646]. alla [3941, 3339, 565, 3454, 1291, 2739, 248]. All’alba [4122]. Allan [3341]. Allard [2316]. Allemagne [4169, 183, 410, 1139, 1285, 1427, 3401, 3376]. allemand [4512, 2110, 4172]. allemande [3529, 4338]. Allen [2564, 3083, 1685, 3074, 3743]. Aller [3074]. Aller-retour [3074]. Allgemeinen [4256]. Allibone [3036]. Alliez [4176]. allo [2629]. Allodi [1300]. allométrie [3956]. allures [3113]. Almagest [3014]. Almagia [989, 989]. almanac [3758]. Almanach [165, 165]. Almanacs [2696]. Almquist [747]. Àl’occasion [3072]. alone [4658]. Aloys [2574]. Alpers [3148]. Alpes [1349, 3948]. Alphabétique [1652]. Alphabétiques [2424]. Alphandéry [1353]. Alphonse [59, 3224]. alphonsine [4397]. Alphonsines [3097]. Alpinolo [602]. Alquié [234]. Alsace [1079]. Alt [1935, 1693]. Alt-Augsburg [1935]. Alt-Augsburger [1693]. alta [1907]. Altar [3362]. altdeutschen [2695]. alte [2849]. alternatives [4402]. alterquinus [4286].

Altertumswissenschaften [3416]. Althin [1267]. altindischer [536]. altri [4068]. aluminium [146]. Aly [476]. am [3489, 1107]. Amalia [1333]. amas [2157]. Amateur [3971]. Ambix. [1004, 1912]. Ambroise [2097, 3167, 460, 3146]. Amedeo [3985, 2988]. Ameliese [211]. Aménagement [4633]. America [4458, 4230, 2686, 3580, 164]. américaine [3740, 4521, 216, 4292]. américains [2101]. American [3486, 4458, 743, 3579, 1317, 1485, 1259, 1415, 2920, 2321, 3363, 2336, 2641, 947, 4206, 2921, 3860, 2271, 1893, 3830, 512, 610, 953, 2336]. americana [3735]. americanas [3735]. Amérique [948, 3854]. ami [2088, 1775, 381]. Amigues [4285, 3736]. amiral [887]. Amis [1744, 193]. Amistad [1171]. ammoniac [692]. Amodeo [189, 189]. among [3981]. Amor [4163, 4288]. amour [4624, 3286]. Amouretti [3896]. Ampère [2879, 2366, 2303, 3835, 2301, 2347, 2304, 3507, 1980, 858]. ampliata [248]. Amsterdam [349, 513, 2659]. Amundsen [3744]. Amy [2775, 4311, 3723, 4188]. Analecta [1909, 4510]. Analisi [830]. analogy [3110]. analomiques [1146]. Analyse [4085, 2169, 1882, 3079, 4407, 1322, 2458, 3133, 2255, 4571, 4504, 4287, 3199, 2724, 3379, 2454, 944, 1184, 3101, 4552, 1295, 208, 462, 2012, 3720, 3189, 3328]. Analyses [2663, 3319, 3486, 3563, 3950, 3972, 2503, 1898, 1267, 1593, 1594, 4576, 4585, 4594, 4605, 4659, 4667, 4672, 4681, 4690, 4409, 4453, 4458, 2458, 3225, 384, 671, 726, 749, 3464, 3694, 4189, 89, 165, 162, 163, 164, 241, 2827, 3010, 2421, 3631, 3519, 4557, 3635, 3767, 3942, 3270, 3663, 4142, 4264, 3901, 4168, 3591, 4526, 3939, 4310, 4433, 4450, 4477, 4508, 4511, 4515, 4546, 4569, 3752, 3747, 3836, 743, 1144, 1164, 939, 1774, 213, 689, 1538, 1539, 1540, 3516, 2374, 2490, 2524, 2547, 2564, 2976, 3458, 3562, 3459, 3632, 3634, 3748, 587, 2901, 2905]. Analyses [3900, 3697, 3866, 4050, 4512, 4513, 3578, 3485, 4092, 4390, 203, 279, 402, 1749, 1785, 1849, 1972, 1973, 2069, 2166, 2289, 2290, 2308, 2309, 2320, 2331, 2332, 2341, 2526, 2527, 2604, 2605, 2606, 2628, 2630, 2631, 2632, 2657, 2658, 2684, 2865, 2868, 2870, 2964, 3124, 3278, 3604, 3654, 967, 3833, 3810, 3815, 4097, 4175, 4240, 692, 695, 746, 748, 793, 794, 795, 876, 885, 3903, 3904, 4123, 4143, 3440, 3898, 3944, 3610, 3003, 3260, 3506, 2504, 2505, 2529, 2530, 2753, 2772, 2831, 2871, 2873, 2925, 2942, 3223]. Analyses [3336, 3489, 3969, 2337, 2338, 2834, 478, 487, 3579, 3533, 4254, 4410, 1752, 3359, 3896, 1633, 2409, 2478, 2536, 2812, 2885, 4365, 1779, 4200, 3340, 3349, 3863, 135, 158, 3343, 3880, 3721, 3700, 3793, 3826, 4144, 4179, 4233, 4242, 4263, 4266, 4314, 4475, 3785, 4312, 2252, 2262, 2509, 2775, 2783, 4165, 1522, 2877, 2878, 828, 7, 39, 40, 41, 43, 44, 45, 46, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 58, 59, 60, 63, 64, 65, 66]. Analyses [67, 70, 71, 72, 90, 91, 93, 94, 95, 137, 139, 159, 212, 1102, 4079, 3483, 157, 1345, 3934, 1016, 1020, 1021, 1022, 1025, 1030, 1545, 3119, 2629, 1974, 4004, 4173, 4232, 3592, 2212, 2228, 2467, 2544, 2627, 2656, 2683, 2803, 2826, 3236, 3303, 3304, 3504, 3811, 3882, 4003, 4088, 2190, 2191, 2192, 2189, 2193, 2275, 2276, 2274, 2295, 2342, 2344, 2296, 2297, 2343, 2345, 2395, 2396, 2411, 2422, 2052, 2061, 2104, 2310, 2532, 4488, 4201, 419, 446, 485, 490, 4077, 4367, 2913, 4297, 3501, 2167, 2168, 2169, 2183, 2227, 2248]. Analyses [2253, 2314, 2754, 611, 3943, 3178, 3207, 3971, 4243, 2750, 3384, 3411, 3517, 3701, 4202, 3522, 3566, 3703, 3727, 4078, 4125, 4170, 4258, 4271, 4344, 4345, 4351, 4352, 4368, 4527, 4081, 1977, 2078, 2215, 2384, 2507, 2531, 2609, 2610, 2664, 3053, 535, 507, 511, 515, 537, 605, 615, 592, 593, 616, 617, 632, 633, 652, 693, 696, 729, 730, 753, 801, 906, 909, 843, 844, 845, 846, 866, 888, 949, 968, 970, 989, 972, 973, 974, 1004, 1033, 1034, 1035, 1078, 1120, 1121, 1122, 1123, 1149, 1150, 1177, 1151, 1152, 1178]. Analyses [1313, 1314, 1334, 1315, 1316, 1317, 1318, 1434, 1435, 1444, 1445, 1477, 1482, 1506, 1525, 1511, 1484, 1485, 1486, 1507, 1508, 1509, 1510, 1630, 1631, 1639, 1640, 1618, 1619, 1665, 1690, 1709, 1710, 1796, 1797, 1798, 1829, 1835, 1836, 1868, 1837, 1975, 1979, 1980, 2028, 2037, 2041, 2038, 2039, 2040, 2042, 2043, 2114, 2115, 2117, 2131, 2116, 2132, 2133, 3487, 2451, 2512, 3245, 2059, 3360, 69, 444, 448, 2263, 3558, 3948, 119, 156, 398, 1214, 3103, 4411, 4434, 1389, 3902, 3992, 4311, 459, 74, 354, 358, 359, 360, 403, 956, 958, 959, 960, 961, 1576, 1577, 1863]. Analyses [3862, 4026, 2776, 3009, 3433, 2378, 3040, 3220, 3238, 3271, 3315, 3341, 3361, 3369, 3371, 3372, 3410, 3436, 3441, 3581, 3693, 3696, 3792, 3797, 3899, 3991, 4507, 138, 140, 218, 281, 282, 304, 400, 401, 450, 451, 452, 453, 586, 728, 767, 768, 864, 1258, 1259, 1269, 1270, 1347, 1349, 1353, 1415, 1416, 1417, 1420, 1422, 1590, 1591, 1727, 1728, 1729, 1736, 2711, 3121, 3598, 2261, 2388, 2389, 2546, 2569, 3000, 3005, 3083, 3169, 3212, 3262, 3282, 3673, 3940, 2957, 3311, 3461, 3528, 3594, 3927, 3945, 3649, 3153, 3235, 3272, 612, 3809, 4360]. Analyses [73, 186, 191, 204, 211, 214, 235, 275, 280, 301, 327, 328, 330, 371, 391, 399, 418, 443, 455, 456, 461, 465, 508, 581, 630, 651, 704, 940, 1032, 1096, 1191, 1208, 1523, 1524, 2423, 2774, 3152, 2058, 2928, 3520, 3628, 3669, 3671, 3676, 3677, 3746, 3761, 3791, 4261, 4393, 3990, 3075, 3145, 2661, 2662, 2779, 2924, 3042, 237, 302, 303, 325, 326, 373, 377, 396, 397, 506, 513, 514, 1754, 1866, 1867, 4196, 694, 3348, 3832, 3514, 3807, 3385, 4320, 3567, 1019, 2470, 2568, 2608, 2636, 2637, 2830, 509, 629]. Analyses [672, 948, 957, 962, 1113, 1850, 2842, 2914, 3766, 3949, 3963, 4033, 4230, 4369, 4436, 4437, 4438, 4093, 3515, 3922, 4022, 4236, 3123, 3557, 4260, 4315, 2770, 3011, 3302, 3370, 3536, 3763, 3787, 4053, 4141, 4298, 2445, 3122, 3233, 4024, 4056, 4174, 3755, 3014, 3097, 3118, 4286, 4491, 2667, 3733, 3786, 3936, 3937, 3941, 4005, 4008, 4018, 4076, 4195, 4255, 4257, 4259, 4267, 4270, 4289, 4415, 4416, 4435, 4457, 4478, 4514, 2109, 3824, 4098, 3414, 3672, 3457, 536, 1831, 2027, 2110, 2111, 2128, 2143, 2318, 2319, 2573, 2575, 2576, 2647, 2805, 3057, 3058, 3059, 3060, 3061, 3062, 3455]. Analyses [3460, 3885, 3886, 4287, 2063, 2294, 3652, 3881, 4343, 4413, 4019, 351, 352, 357, 3650, 2674, 3309, 3731, 2506, 3273, 2565, 2638, 4234, 3590, 4321, 4021, 3870, 3869, 3301, 3337, 3730, 3751, 3834, 4032, 4205, 2502, 3839, 4020, 4049, 4203, 4516, 305, 1348, 3935, 4309, 1610, 3841, 2312, 2313, 4094, 3521, 997, 602, 3154, 3155, 3276, 3277, 3569, 3585, 3749, 3947, 4090, 4091, 4441, 3666, 2639, 2640, 2642, 2837, 2838, 2900, 2988, 3208, 3242, 3279, 3342, 3386, 2709, 2920, 3054, 3105, 3313, 3314, 3362, 3408, 3417, 3431, 3462, 3480, 3482, 3488, 3503]. Analyses [3505, 3526, 3531, 3564, 3584, 3589, 3611, 3637, 3640, 3744, 3867, 2741, 3994, 4082, 4171, 2780, 3925, 4095, 1773, 1978, 3413, 983, 999, 1162, 1432, 1442, 1476, 1499, 1500, 1501, 1502, 1503, 1504, 1505, 1526, 1614, 1635, 1662, 1688, 1694, 1788, 1833, 2031, 2045, 2077, 2097, 2129, 2130, 2142, 2172, 2173, 2217, 2218, 2321, 2447, 2450, 2550, 2574, 2577, 2578, 2581, 2591, 2660, 2670, 2840, 2841, 2843, 2846, 2847, 2884, 2886, 2903, 2917, 2961, 2963, 2975, 2989, 3125, 3148, 3150, 3209, 3211, 3240, 3274, 3312, 3317, 3338, 3363, 3416, 3418, 3456, 3532, 3534, 3593, 3630, 3636, 3753, 3757]. Analyses [350, 1780, 2029, 2030, 2036, 2044, 2266, 2267, 2316, 2325, 2339, 2643, 2883, 4313, 2251, 2876, 2522, 3670, 3788, 3989, 3993, 4451, 3079, 3214, 2991, 1169, 2491, 2174, 2170, 2233, 2336, 2358, 2452, 2472, 2473, 2474, 2475, 2479, 2533, 2582, 2641, 2692, 2695, 2706, 2707, 3675, 3639, 1611, 1143, 161, 2420, 690, 747, 861, 862, 3381, 2214, 2477, 2534, 606, 666, 773, 607, 665, 673, 675, 691, 727, 744, 750, 751, 772, 796, 797, 798, 799, 800, 818, 819, 838, 879, 881, 882, 883, 884, 963, 964, 965, 995, 996, 998, 1000, 1001, 1002, 1003, 1029]. Analyses [1059, 1061, 1114, 1115, 1116, 1148, 1168, 1170, 1173, 1174, 1175, 1245, 1302, 1303, 1305, 1306, 1307, 1308, 1309, 1310, 1311, 1312, 1441, 1443, 1480, 1481, 1565, 1569, 1571, 1579, 1613, 1638, 1689, 1692, 1701, 1702, 1703, 1704, 1705, 1706, 1707, 1708, 1755, 1756, 1757, 1834, 1848, 1869, 1983, 2032, 2033, 2034, 2035, 2064, 2065, 2066, 2067, 2068, 2186, 2188, 2234, 2268, 2269, 2270, 2322, 2323, 2324, 2340, 2393, 2410, 2448, 2535, 2579, 2644, 2645, 2646, 2666, 2671, 2685, 2708, 2785, 2787, 2845, 1847, 1997, 2094, 2354, 2603, 3438, 2672, 2853, 2915, 2974, 3043, 3216, 210, 236, 324]. Analyses [466, 2510, 2548, 1685, 1771, 1772, 2453, 2469, 2828, 4071, 42, 109, 61, 62, 68, 87, 88, 92, 117, 131, 132, 134, 136, 150, 151, 152, 238, 239, 242, 245, 300, 355, 374, 395, 445, 447, 481, 482, 567, 603, 628, 653, 670, 770, 789, 792, 984, 985, 986, 987, 988, 1015, 1072, 1097, 1098, 1163, 1215, 1218, 1219, 1232, 1296, 1297, 1327, 1328, 1330, 1386, 1387, 1388, 1467, 1544, 1568, 1573, 1608, 1609, 1628, 1634, 1663, 1666, 1667, 1668, 1669, 1670, 1671, 1672, 1673, 1681, 1682, 1683, 1750, 1753, 1776, 1777, 1778, 1789, 1790, 1791, 1792, 1793, 1794, 1795, 1844]. Analyses [2025, 2054, 2055, 2056, 2057, 2093, 2095, 2096, 2107, 2919, 3039, 3126, 3586, 3587, 614, 991, 1263, 2987, 3008, 3218, 525, 532, 1468, 2806, 3729, 2230, 2306, 1982, 3648, 3651, 3653, 3655, 4316, 2773, 3745, 4398, 3523, 3525, 3561, 3704, 3725, 4432, 4492, 4494, 3970, 3316, 4121, 3339, 4474, 3868, 3965, 2372, 2570, 3638, 874, 942, 947, 954, 3706, 4207, 4439, 4531, 4096, 4206, 3383, 582, 583, 584, 591, 601, 604, 626, 2733, 2777, 2921, 3608, 1440, 3664, 4362, 565, 2962, 3259, 3789, 3840, 379, 1337, 1338, 1617, 1853, 2539, 2686, 3213, 4347, 4392, 4396, 4414, 4456]. Analyses [4528, 3559, 3560, 2076, 2999, 2922, 2508, 3149, 3389, 3966, 4009, 4034, 4178, 4269, 4292, 3814, 4120, 4532, 3580, 4176, 2693, 2696, 3754, 3795, 3796, 3842, 3864, 3924, 585, 1664, 3241, 1057, 1616, 1636, 2250, 2511, 3055, 1236, 1237, 2140, 2782, 3012, 3197, 3280, 3347, 3409, 3420, 4080, 3756, 3723, 3794, 1578, 3605, 3835, 4295, 4349, 3582, 2287, 2668, 2705, 2839, 3629, 3762, 886, 1331, 4199, 4084, 3860, 3535, 4010, 4025, 2182, 2293, 2538, 2916, 2352, 4231, 3299, 1786, 1787, 3106, 3206, 4139, 3743, 2833, 2835, 2918, 2958, 2959, 2960, 2972, 2977, 2979, 3002, 3007]. Analyses [3016, 3035, 3037, 3074, 3077, 3156, 3157, 3210, 3222, 3243, 3257, 3275, 4241, 4366, 4454, 4490, 4529, 249, 4319, 2247, 3828, 1471, 3439, 3518, 2231, 2260, 2992, 3344, 3500, 3699, 3702, 3844, 3080, 86, 524, 1017, 1216, 1288, 4417, 3454, 3490, 4412, 3529, 1472, 1473, 3484, 3412, 3627, 3481, 3524, 3606, 3724, 950, 992, 1109, 1117, 785, 786, 857, 859, 865, 900, 907, 937, 981, 982, 990, 1103, 1104, 1142, 1217, 1234, 1235, 1239, 1240, 2822, 2823, 4066, 2813, 2844, 2851, 3166, 3364, 3597, 3764, 4167, 1537, 1845, 1976, 2225]. Analyses [2226, 2271, 2523, 2537, 2810, 2811, 4177, 2244, 3432, 4188, 826, 832, 3565, 3570, 3607, 3612, 3758, 4290, 4397, 4486, 4487, 4510, 3626, 3365, 3742, 4530, 2333, 205, 206, 215, 216, 217, 4294, 3583, 4239, 2371, 3938, 3281, 3434, 3437, 3695, 3698, 4204, 4229, 4052, 4055, 4350, 3728, 901, 945, 1268, 1271, 1264, 1344, 1346, 1350, 1351, 1352, 1354, 1356, 1355, 1412, 1413, 1414, 1418, 1419, 1421, 1592, 1597, 1732, 1733, 1734, 1735, 1737, 2848, 2849, 2902, 3004, 3052, 3076, 3096, 3194, 3221, 3366, 1751, 2778, 3443, 3736, 3879, 3883, 3921, 3962, 3964, 4007, 4146, 4147]. Analyses [4285, 3568, 3813, 3923, 4072, 4363, 4346, 3946, 3165, 3234, 3318, 3705, 3527, 3491, 3735, 2024, 2060, 2106, 2126, 2154, 2157, 2158, 2216, 2232, 2288, 2334, 2419, 2501, 2528, 3195, 3577, 3224, 2385, 3147, 3829, 4348, 277, 2941, 2993, 1244, 745, 769, 887, 904, 905, 938, 1018, 1075, 1100, 1101, 1230, 1231, 1262, 1436, 1470, 1521, 1604, 1693, 1758, 1759, 1827, 1846, 1854, 2105, 2356, 2405, 2406, 2525, 3082, 3196, 3215, 3345, 3595, 3603, 3734, 3812, 3884, 3996, 4054, 4291, 3816, 4452, 155, 3307, 1695, 1981, 2113, 2272, 2494, 2514, 2583, 2648, 2758, 2815, 2850, 2926, 3104, 3177]. Analyses [3871, 4172, 3831, 3837, 4262, 2702, 2746, 2749, 3861, 4140, 4210, 1605, 1661, 2229, 2545, 4083, 2377, 3843, 3300, 3435, 3463, 3107, 3668, 3741, 3759, 3308, 3407, 4208, 875, 185, 209, 234, 274, 376, 392, 393, 504, 742, 1056, 1058, 1233, 1437, 1478, 1536, 1606, 1607, 1775, 863, 941, 1289, 1574, 2175, 2633, 2786, 2809, 4006, 4391, 4476, 3732, 2213, 2127, 2292, 2353, 2370, 2376, 2404, 2408, 2443, 2444, 2571, 2611, 2673, 2675, 2698, 2700, 2734, 2735, 2736, 2738, 2744, 2745, 2748, 2751, 2752, 2756, 2824, 2825, 2832, 2879, 3823, 3825, 479]. Analyses [483, 356, 484, 510, 674, 946, 3261, 3419, 4166, 3146, 731, 752, 754, 902, 944, 1392, 1483, 2375, 2990, 3056, 3098, 3167, 3168, 3306, 3367, 3387, 3453, 1575, 1686, 1687, 1998, 2026, 2062, 2108, 2141, 2171, 2185, 2264, 2315, 2476, 2549, 2572, 3442, 3444, 4035, 4036, 4493, 4394, 4440, 4318, 3667, 4161, 1479, 2387, 4163, 4235, 4296, 4317, 4361, 4455, 3827, 880, 380, 1438, 3897, 4162, 2869, 3120, 2771, 3036, 4145, 4238, 154, 417, 3722, 3726, 4068, 4070, 2307, 2694, 2697, 2699, 2732, 2743, 3790, 4209, 3038, 3258, 3995, 4051, 4124, 4164]. Analyses [4364, 4489, 2053, 2155, 2386, 3239, 3310, 3609, 4023, 3588, 3633, 3765, 4237, 4293, 4197, 3988, 4122, 1730, 1731, 2669, 2392, 1145, 1295, 1300, 1333, 1430, 1431, 1433, 1567, 1580, 1632, 1674, 1680, 1684, 1851, 1852, 1865, 1999, 2373, 2567, 3078, 4265, 4288, 2184, 2245, 4256, 243, 333, 306, 353, 362, 422, 833, 96, 118, 130, 133, 153, 160, 141, 184, 188, 189, 192, 207, 208, 240, 219, 250, 276, 283, 284, 278, 332, 307, 331, 372, 375, 394, 421, 383, 404, 442, 449, 463, 464, 477, 480, 491, 505, 526, 527, 528, 529, 530, 531, 488, 489]. Analyses [566, 627, 654, 667, 668, 669, 766, 784, 787, 788, 790, 791, 820, 821, 822, 823, 824, 825, 827, 829, 830, 831, 834, 836, 839, 840, 841, 842, 855, 856, 858, 860, 877, 903, 943, 969, 1070, 1071, 1073, 1074, 1077, 1099, 1105, 1118, 1290, 1291, 1292, 1293, 1298, 1299, 1323, 1324, 1325, 1326, 1329, 1332, 1335, 1336, 1294, 1403, 1469, 1541, 1542, 1543, 1564, 1566, 1570, 1572, 1612, 1629, 1826, 1828, 1864, 2291, 2311, 2355, 2407, 2755, 2781, 2852, 2866, 2867, 2872, 2874, 2875, 2880, 2881, 2904, 2927, 3041, 3305, 2112, 2159, 3750, 221, 246, 247, 361, 363, 454, 457]. Analyses [534, 588, 589, 590, 771, 835, 837, 878, 952, 993, 994, 1014, 1023, 1024, 1026, 1027, 1031, 1005, 1060, 1079, 1106, 1107, 1108, 1111, 1112, 1119, 1076, 1146, 1147, 1165, 1166, 1167, 1171, 1172, 1176, 1238, 1242, 1243, 1304, 1391, 1691, 1761, 1830, 2070, 2187, 2265, 2317, 2449, 2513, 2580, 2649, 2665, 2784, 2829, 2882, 2978, 3015, 3368, 2965, 3502, 4011, 4169, 4395, 3968, 2156, 2211, 2249, 2273, 2390, 2391, 2394, 2446, 2457, 2471, 2592, 2607, 2612, 2634, 2635, 2701, 2703]. Analyses [2704, 2739, 2740, 2757, 966, 1110, 1241, 4067, 3760, 2710, 2807, 2808, 2929, 3217, 3830, 3013, 3151, 3244, 3926, 4495, 4509, 3388, 3382, 3415, 2814, 4268, 3674, 476, 378, 3665, 4198, 971, 244, 4190, 3865, 1260, 1261, 3530, 3596, 3967, 248, 329, 381, 382, 420, 458, 460, 462, 486, 512, 533, 608, 609, 610, 613, 631, 774, 951, 953, 955, 1028, 1301, 1390, 1615, 1637, 1832, 4089, 1246, 2492, 2493, 2742, 3237, 3720, 2242, 2243, 2521, 2737, 2747, 2836, 2923, 2973, 3001, 3006, 3081]. Analyses [3219, 187, 190, 220, 2210, 2224, 2246, 2335, 2403, 1439, 1474, 1475, 3346, 2468, 2495, 2551, 2566, 2613, 2593, 2614, 2659, 2677, 2712, 2788, 2816, 2854, 2855, 4069, 1653]. Analysis [2506, 2025, 2333, 278, 89, 3169, 4571, 4659, 4667, 4672, 4681, 4690]. analytic [3839]. Analytical [1863]. analytics [3423]. Analytique [1654, 1544, 2705, 3606, 3453, 2056]. Analytiques [4559]. Analytische [3201]. Anamorphoses [825]. Anastasios [4641, 4509]. Anastassiadou [4486]. Anastassiadou-Dumont [4486]. Anastassiou [3148]. anastatica [2808]. Anatole [2999]. anatoliennes [4486]. Anatomia [1020, 2644, 3901]. Anatomica [1848]. Anatomical [3755, 1301]. Anatomie [496, 1689, 2916, 91, 4405, 4587, 4412, 1146, 4007]. anatomique [279, 1986, 356]. Anatomist [4067]. anatomiste [495, 750]. anatomistes [1703, 2066]. Anatomy [1020, 2321, 951, 4587, 2128]. Ancestors [3003]. Ancien [3753, 3621, 1115, 1708, 2234, 2448, 588, 606, 4662, 2852, 759]. Ancienne [1032, 2119, 2254, 3721, 677, 2494, 144, 112, 2273, 655, 3054]. Anciennes [1409, 2394, 2827, 2379, 2839, 1075]. Anciens [704, 4644, 10, 22, 255, 1581, 1341]. Ancient [3303, 3118, 1162, 2172, 2961, 2472, 1789, 784, 4627, 3417, 3482, 3564, 3584, 3589, 3637, 3640, 3744, 3921, 3964, 4036, 3830]. andalous [2272, 2758]. Anderson [3603]. Andler [2999]. Andrade [3041]. Andral [2652]. André [3950, 4410, 693, 3693, 864, 1729, 301, 957, 962, 2597, 3639, 1168, 324, 242, 670, 1215, 1634, 3409, 981, 4291, 4452, 3107, 2376, 669, 1172, 1304, 2156, 3151, 187, 2613, 4561, 858, 3164, 10, 4617, 3487, 3835]. André-Louis [981, 4617]. André-Marie [858, 3835]. Andrea [2914]. Andreae [939]. Andreas [2663, 36, 1020]. Andrée [2774, 3152, 1167, 40]. Andrew [3338, 3481, 3597, 4197]. Andrieu [4343]. Andrillat [3557]. Anecdota [4494, 4529, 4296]. anesthesia [1481]. anesthésie [753]. Anesthesiology [1505]. Anfänge [1926, 2053]. Anfängen [2885, 2671, 1325]. Angèle [4171]. Angeles [2733]. Angeles-London [2733]. Angelo [4161]. angina [800]. Angiolo [603]. anglais [2110, 1950, 1339, 2101]. anglaise [3206, 2820]. anglaises [3287]. Angle [2291]. Angleterre [507, 4205, 4124, 3198, 492]. Anglican [3301]. Anglicus [3587]. Anglo [1442, 3860, 3756]. Anglo-American [3860]. Anglo-Saxon [1442, 3756]. Ango [2102]. Angus [1237]. Anima [2287]. Animal [1476, 1110, 89, 2220, 923, 4454, 3984, 247, 3256]. animal-book [247]. animale [2194, 3983, 3984, 2045]. Animalium [2534, 1021, 389, 1676]. animals [247, 3534]. animateur [101]. animaux [4524, 4614, 2487, 1337, 1838, 2839, 1458, 4395, 356]. animaux-machines [2487]. Anita [4514, 2850]. anlässlich [1107]. Anleitung [1097]. Ann [3885, 3636, 2266, 2267]. Anna [32]. Annales [2657, 459, 1174, 2324, 1605, 2854, 1836, 3963]. Annali [2344]. Annals [727]. Anne [4512, 3533, 4243, 3360, 3121, 3311, 3153, 4261, 3060, 4413, 3670, 3989, 4176, 3407, 4265, 3750, 4011]. Anne-Claire [4261]. Anne-Francoise [4413, 3989]. Anne-Marie [4512, 3533, 3360, 3121, 3407, 4265, 3750]. Annecy [1148]. année [279, 5, 2422, 592, 843, 1033, 1121, 1445, 1507, 1510, 1727, 353, 1170]. anneés [2038, 2150, 432, 3687, 1315, 1316, 1318, 1708, 2095, 4591, 422, 283]. Annette [1313]. annexe [2754]. annexes [4049]. Annibale [3727]. Annick [3729]. Annie [4266, 4411, 4509]. Anniversaire [979, 1493, 4159, 3048, 2160, 145, 2180, 2939]. Anniversaires [772]. annonce [1204]. annotated [3903, 508, 2784, 4581]. annotée [559, 3648, 4581]. Annuaire [616, 801, 1150]. Annual [163, 2262, 1152, 1317, 1485, 1246, 1484, 1913, 2115, 384, 218]. annuel [217]. annum [3438]. anomalies [3493]. anonyme [4570]. anonymes [2455]. Anonymi [3921]. anonymous [4570]. Años [1171]. ans [4645, 461, 2842, 3949, 2910, 3554, 3439, 878, 1691, 3530]. Ansätze [2603]. anschaulichen [3416]. Anselm [4234]. ante [2805, 3438]. Antécédents [1]. anteriores [2644]. anthologie [3412]. anthology [4242]. Anthony [4315, 3206, 3345]. Anthropologe [2785, 2645]. Anthropologie [3123, 4166, 3654, 4624, 2538, 4476, 2352]. anthropologique [279, 4538, 4454, 2973]. Anti [3411, 4670]. anti-plague [4670]. Anti-realist [3411]. antibiotiques [648]. Antichi [989]. antidotarii [4206]. antigues^{≫} [904]. Antike [2989, 1689, 3237, 3416]. antiken [2467, 2849]. Antilles [3753]. antipesteux [4670]. antiprogressionnisme [2860]. antiqua [3482, 3637, 3640]. antique [3294, 4091, 818, 2288, 1661, 3366, 4285, 4035].

Antiquitätenfreunde [1262]. Antiquité [2656, 4620]. Antiquity [2547, 1432, 3432, 1071]. antisepsis [609]. antithesis [609]. Anto [3153]. Antoine [692, 1102, 2411, 3348, 3670, 4492, 2999, 585, 3723, 3305, 2298, 1556, 79, 3920, 1045, 1084, 2180, 714, 453, 4267, 1775, 1687, 2245]. Antoine-Francois [1687]. Antoine-Laurent [1045, 1084]. Antoine-Louis [2298]. Antoinette [2336, 1674, 2767]. Anton [2803, 3059, 2479, 2406, 3568]. Antonella [4511, 4096, 4362]. Antonello [3153]. Antoni [3903, 3078, 358, 326]. Antoniadès [1682]. Antonino [565, 3829]. Antonio [4255, 3653, 3454, 4083, 4068, 3960]. Antony [4507, 4267]. Antrittsrede [1028]. Anusia [2269]. anxio [4199]. anxio-dépression [4199]. août [349, 1861, 1825, 1862, 162, 1918, 795, 1461, 1699, 2491, 1467, 1667, 1201, 786, 3686, 1206, 220]. Ape [589]. Apercu [251, 956, 2238, 1456, 464]. Apes [589]. Apfel [2547]. aphasies [4412]. Aphorismes [1525]. Aphorismi [2669]. apices [76]. Apokryphen [3650]. Apollon [380]. Apollonius [3420]. Apologia [4068]. apologie [460]. apothicaire [692]. Apothicaireries [1138]. apparatus [3516, 4198]. apparence [4332]. Apparences [4332]. apparire [4475]. apparition [4013]. Appartenance [1919]. Appel [2547]. Appendice [2436, 4643]. appendicectomie [2451]. Appendix [1863, 4348]. Appl [2400]. application [3552, 1851]. Applications [3009, 2568, 3052, 2347, 3009]. appliqué [274]. appliquée [4349, 2100, 1270]. appliquées [795, 4003]. apport [4165, 4279, 1611]. apports [78]. Appreciation [784]. Apprentice [2109]. approach [4520, 4584, 3107]. approche [3426, 4195, 2666, 4584, 3379, 3920]. Approches [4545]. Approfondir [4219]. appropriation [4652]. appui [3654, 4436, 3608, 4166]. appui/Genre [4436]. apr [4003, 1105]. après [2256, 3343, 461, 4335, 3554, 3160, 2204, 2260, 83, 3530]. April [404, 489]. Aquila [4076]. Aquin [2473]. arabe [19, 1630, 4128, 3158, 1220, 4164, 52, 3595, 284]. arabe-hébreu [19]. arabes [2354, 3241, 3239]. Arabic [1792, 2372, 3420]. Arabische [1100]. Arabischen [2545, 2229]. ’arabiyya [2372]. Arabo [1539, 4130]. Arabo-Latin [1539]. arabo-latins [4130]. Arago [1524, 4504, 4568, 3092]. Arambourg [327, 2952]. Arato [3365]. Araujo [4232]. Arbeiten [2872]. arbeitende [2958]. arbitrage [4185]. Arbor [2661, 3214]. arbres [1222]. arbustes [1222]. arc [1818]. arc-en-ciel [1818]. Archaelogy [1261]. archaeology [4569, 1594, 1421]. archaïque [2919]. Archéol. [2266, 2267]. Archéologie [2902, 3460, 2654]. archéologique [1344, 1290]. archéologiques [1605]. Archetypes [3003]. Archibald [4172, 240]. Archimede [4602, 2055, 3882, 223, 1628, 2093, 1764, 3583, 1927]. Archimedes [689, 1539]. Archimedes. [628]. Archipelago [950]. Architectes [1541]. architectoniques [4685]. architectural [4685]. Architecture [4303, 4475]. Architettura [4088]. Archiv [4125, 2473]. Archive [2342, 1750, 4664, 4661, 2274, 1151]. Archives [2290, 2308, 571, 3969, 4242, 4228, 3948, 4434, 3922, 4259, 2491, 879, 1705, 1966, 3653, 4646, 682, 891, 4660, 2898, 4049, 4662, 4665, 4545, 2189, 2343, 617, 843, 1033, 1121, 1315, 1444, 1508, 1619, 1796, 2038, 141, 192, 307, 383, 488, 2070]. Archiving [4661]. Archivio [4351]. Arctic [3567]. Arcueil [1686]. ardente [382]. Arecco [4309]. arénaire [2055]. Arezzo [4076]. Argent [677]. Argentina [62, 566]. Argentine [127]. Argoud [4003]. argument [1361, 2936]. argumentatif [4636]. argumentative [4636]. arguments [2937]. Arias [3735]. arie [2914]. Ariew [3751]. Aris [3861]. Aristote [2839, 4668, 4058, 3111, 4138, 3643, 7]. Aristoteles [2467, 990, 3443]. Aristotelian [4402]. aristotélicien [3643]. Aristotelicum [1750]. Aristotelis [2287]. aristotelisch [3443]. aristotelisch-scholastische [3443]. Aristotle [3752, 4668, 395]. arithmetic [3888]. arithmetica [4012]. Arithmetical [2054]. arithmetice [3516]. arithmétique [3495, 1683, 1425, 1295, 1210, 1683, 3698, 1326]. Arithmétique. [3725]. arithmétiques [530, 1329]. Arithmetischer [1777]. Arithmos [2566]. Arithmos-Arrythmos [2566]. Armand [2658, 2964, 3124, 3278, 281, 3628, 2065, 239, 2514, 463, 795]. Armées [1002, 2884, 2975]. Armelle [4259, 3964]. Armengaud [4206]. Arménie [958]. Armitage [1237]. Armogathe [2631]. Armoise [279]. armourers [1345]. army [4617]. Arnaldo [2567]. Arnaud [213, 889]. Arnauld [4492, 382]. Arnold [653, 739, 651]. Aron [1386]. Árpád [1926, 2053]. arrêté [2540]. arrangements [4385]. Arrington [995]. Arrythmos [2566]. Ars [3503, 2740]. Arsbok [1150]. Art [1863, 3989, 3644, 3194, 1759, 840, 2689, 3294, 4470, 2660, 1168, 3779, 1558, 2993, 4225, 4035, 1761, 824]. arte [4303, 4006]. Arthur [3461, 2979, 3439, 3166, 3570, 1827, 2746, 667, 993, 3599, 3050]. arti [4088]. artibus [4394]. article [2466, 1205, 2285, 4581]. articles [342, 4223]. articolo [3941]. Artières [4004]. artificial [4478]. artificiale [2629]. artificielle [1136, 4174, 1787, 1986]. artillerie [3936]. Artis [2777, 3516]. Artisans [1541]. artist [2539]. Artistes [3463]. artistique [1146]. Arts [346, 892, 2452, 314, 181, 2583, 388, 3903, 4388, 3786, 3987, 4209, 760, 1535, 1840, 583, 1692]. Artsrouni [1407]. Arturo [248, 420]. Artz [1591]. Arzl [3123]. Arzneimittelforschung [2130]. arzte [3843]. Arztes [2670]. as-Samaw’al [2230]. Aschkâl [975]. Aschkâl-üt-teessis [975]. Aschoff [3739]. ascribed [689]. Ash [2244]. Ashcraft [3484]. Ashmole [1685]. Ashworth [1500]. asile [4662]. Asimov [2735, 2735]. Asklépios [380]. aspect [2517, 1579, 4379]. aspectibus [4331]. aspectos [4453, 62]. Aspects [1025, 2310, 4381, 3951, 1688, 1827, 1942, 4334, 658, 1960, 3952, 105, 3160, 3016, 1735]. Aspetti [4083]. aspirantes [1583]. assembly [3347]. Assis [3693]. Assistance [1175]. Association [3594, 2055, 3965, 1079, 1594, 2123]. associés [1375]. assyrische [1440]. astralabe [3812]. astrale [3873]. Astrarium [1893]. Astres [187, 4466]. Astrolabe [3082, 2761, 1721, 3812]. Astrolabes [3082, 4478, 619, 4291, 3217, 3996, 4054]. Astrolabica [2929]. astrolabists [745]. astrologer [2903]. astrologica [3438]. astrologie [8, 3884]. astrologue [3604]. Astrology [2696]. Astronom [2607, 876, 12]. astronome [3604, 659, 4119, 4381, 2894, 77, 2237]. astronomer [4290, 31]. astronomes [103, 2782, 3280, 3607, 3758, 4397, 4544]. Astronomia [531, 566, 3339]. Astronomica [2873, 3438]. Astronomical [3083, 3011, 1827, 833, 3169, 1101, 3282, 1933]. astronomicas [1436]. astronomico [4476]. Astronomie [1902, 397, 2603, 4493, 2485, 4378, 3273, 315, 2453, 4280, 2231, 2236, 4397, 4382, 2880, 2545, 2683, 3842, 2782, 3612, 187]. astronomii [2453]. astronomique [2727, 3660, 3607, 4382, 1395]. astronomiques [1346, 2929, 209]. astronomische [769, 1693, 3734]. astronomischen [1470]. Astronomy [2604, 2276, 2041, 2117, 3118, 938, 1758, 1827, 531, 1299, 3000, 2236, 3996, 3830, 3000, 1664, 3604, 2922, 4256, 3010, 1854, 4318]. Astrophysics [2992, 4668, 3000]. astrophysique [4668, 3612]. astros [1542]. Astruc [2019]. Asylum [2993, 4662]. Asylum-keeping [2993]. Atlantide [2925]. Atlas [3577]. atmosphériques [4184]. Atom [1472, 3036]. Atome [4, 2157, 3690]. atome/molécule [3690]. atomes [4170]. atomi [4072, 3238]. Atomic [1866, 1472, 3236]. atómico [1980]. atómico-molecular [1980]. atomique [3372, 3007, 1224, 1472]. Atomisme [4070, 4071, 4064, 4059, 122, 196, 270, 366]. atomismo [3523, 4070]. atomista [4068]. Atomistic [3139]. atomistique [4060]. Atomphysikers [3040]. Atran [3153]. Attali [2825]. attardés [3056]. Atti [4088, 4076, 2787, 4161, 4070, 4435, 4076, 4435]. attitude [761]. attraction [407]. attribuer [108, 227]. attribués [889]. Au-delà [4584]. aube [3288, 237, 2533, 2033, 2919, 1990]. Aubert [2122]. Aucante [4123]. audacieux [350]. audience [2605]. audition [1456]. Audouin [2575, 4392]. Auerbach [610]. Aufenthaltes [278]. Auffret [1704]. Aufklärung [2917, 954]. Auflage [506, 1470]. Aufsätze [1776, 1107, 3650]. Auftreten [1325]. Aufzeichnungen [1111]. Augenheilkunde [3505]. Auger [1384, 1335]. augmentée [834, 1828]. Augsburg [2479, 1693, 1935]. Augsburger [1693]. August [3386, 991, 2987, 667]. Auguste [32, 275, 3348, 2449, 4574, 3931, 2434, 519, 3920, 4266, 3705]. Augustin [2114, 1558, 681, 2017, 2003, 1348, 149]. Augustin-Normand [1348]. Augustine [3747]. Augustus [4431]. Aujac [3604, 2627, 2628, 1845, 3665]. aujourd’hui [3333, 3302, 2736, 766, 1242, 2999]. Aunes [1409]. Aurelianus [361]. aurora [4597]. aurore [2597]. aurores [4597]. Ausbau [787]. Ausdrucksformen [4093]. Auseinandersetzungen [787]. Ausejo [3666]. Ausgewählte [2835]. auspices [3903, 4608, 608]. Auspitz [2853]. aussi [2017, 2003]. Australasia [2686]. Australia [3455]. Australian [3432]. Austro [4056]. Austro-Polish [4056]. Auswahl [1666]. auteur [1102, 4570]. Auteurs [1652, 3408, 2101, 568]. authenticité [3496]. authentiques [3962]. author [3637, 4570]. authority [2936]. auto [4541]. auto-création [4541]. Autobiographical [2586]. autobiographies [4315]. autobiographiques [378, 1027]. autobiography [3005]. autographe [352]. autographes [1948, 754]. Autolycos [2627]. Automata [2710]. autonome [1342]. Autopsie [3233, 1978]. Autorenkollektiv [2058, 1684]. Autour [556, 3809, 4159, 3439, 1722, 3403, 4291, 3694, 3547, 4503, 1579, 4080, 3953, 4319]. Autre [1622, 4060, 4304, 3145]. autres [4546, 4641, 3924, 834]. Autriche [723, 1159, 1530]. Auvergne [3476]. Aux [3776, 3378, 4547, 4607, 1659, 3648, 4278, 3496, 3853, 2628, 2868, 4240, 571, 4109, 3555, 1102, 762, 2451, 3138, 3331, 1258, 1892, 280, 2622, 4685, 758, 2238, 3875, 3881, 4687, 3208, 2884, 2975, 3987, 879, 1148, 1579, 3043, 1340, 1966, 4403, 601, 2595, 468, 4633, 3157, 86, 3600, 2937, 4283, 3295, 2347, 2404, 2745, 270, 682, 4087, 4591, 1851, 668, 841, 1541, 2211, 891, 4127, 2293, 2538, 3407, 835]. Auxiliaire [4247]. auxiliaires [668]. av [4003]. avait [3168, 3530]. avancement [1949, 4065]. Avant [2430, 308, 4619, 3914, 948, 957, 1965, 2412, 1574, 3750, 878, 2260]. Avant-propos [2430, 308]. avatar [2851]. avatars [3290]. avènement [319]. avenir [212, 1058]. Aventur [2243]. aventure [16]. aventuriers [39]. averti [3971]. Avertissement [4109, 1651]. aviation [3004, 3076]. Avicenne [408]. Avignon [2218]. Avis [3208]. Avogadro [1980, 2988, 3985]. avr [192]. avril [1822, 3569, 941, 510, 370, 184, 463, 1784]. Awakening [653]. Axel [2406]. Axes [2775, 2783]. axiom [2876]. axiomatique [2313, 4129]. axiomes [2669]. Aydin [1388, 1018]. Azard [2962]. Azra [1294, 2311].

B [3972, 3278, 1025, 652, 3487, 1591, 3169, 327, 2661, 3042, 237, 325, 397, 2608, 4203, 1925, 2581, 2339, 2479, 2644, 2828, 653, 2306, 3664, 1853, 2539, 3535, 3243, 4319, 3344, 1240, 2385, 3463, 2611, 2476, 566, 4268, 1301, 190, 2551, 2614, 1101, 841]. B. [472, 2179, 519]. B.E.F.R.C.S [673]. Bâle [979]. bâtisseurs [968]. bête [4296]. ba [279]. Baader [2843]. Babbage [2725]. Babicza [2252]. Babini [566, 1628, 1778, 1907, 1324, 1542]. Babson [235]. babylonienne [237]. babyloniennes [251]. babylonisch [1440]. babylonisch-assyrische [1440]. Babylonische [237]. Bacchic [4364]. Bacciagaluppi [4507]. Bach [4645, 4361]. Bachelard [3633, 3032, 3033, 4657, 1281, 274, 376, 902, 2806, 4508]. Bachet [3, 986]. Bachir [3310]. Bachmann [4121]. Back [47, 48, 110, 97, 98, 168, 169, 170, 260, 261, 262, 263, 335, 336, 337, 424, 425, 426, 427, 539, 540, 541, 542, 635, 636, 637, 638, 706, 707, 708, 709, 803, 804, 805, 806, 911, 912, 913, 914, 1037, 1038, 1039, 1040, 1125, 1126, 1127, 1193, 1194, 1195, 1273, 1274, 1275, 1276, 1365, 1366, 1367, 1368, 1448, 1449, 1450, 1451, 1547, 1548, 1549, 1550, 1643, 1644, 1645, 1646, 1712, 1713, 1714, 1715, 1800, 1801, 1802, 1803, 1872, 1873, 1874, 1875, 2007, 2144, 2425, 2553, 2554, 2555, 2556]. Back [2616, 2617, 2618, 2714, 2715, 2716, 2717, 2790, 2791, 2792, 2888, 2889, 2890, 2943, 2944, 2945, 3017, 3018, 3019, 3020, 3085, 3086, 3087, 3127, 3128, 3129, 3180, 3181, 3182, 3246, 3247, 3248, 3320, 3321, 3322, 3391, 3392, 3393, 3465, 3466, 3467, 3539, 3540, 3541, 3615, 3616, 3617, 3680, 3681, 3682, 3770, 3771, 3772, 3847, 3848, 3849, 3907, 3908, 3909, 3975, 3976, 4039, 4040, 4101, 4102, 4103, 4104, 4149, 4150, 4213, 4214, 4274, 4275, 4323, 4324, 4372, 4373, 4419, 4420, 4460, 4461, 4498, 4499, 4534, 4535, 4630, 4649, 4666]. Background [2604, 3411, 1389, 1442, 1979, 1688]. backstaffs [4478]. Bacon [2925, 418, 940, 2336, 3126, 525, 3535, 4010, 4124, 4305, 3257, 1212, 9, 1468, 4025]. Baconiana [418, 940]. Bacq [2217]. bactériologie [63]. Bacteriology [882]. Baden [3301, 2391]. Baer [674]. Baeschlin [479]. Bahir [2230]. Bailey [2159]. Bain [4355, 4354, 4358, 4356, 4357]. Bainton [511]. Baire [2458]. baja [1907]. Bajosa [2644]. Baker [455, 3860, 2959, 2403]. balance [3694, 281, 1362]. balancerie [281]. Balaruc [4625]. Balashov [4254]. Balbinot [3344]. Baldasar [1020]. Balduini [2344]. Balestreri [2580]. Balibar [3302, 2611, 2972]. Ballistics [509]. Balmer [2296]. Balpe [3746, 4195]. Balthasar [2797]. Balthasar-Georges [2797]. Baltrusaitis [825]. Banû [2570]. Banû-Mûsa [2570]. Banach [3885]. Bancroft [2942]. Band [3223, 1928, 2536, 2661, 1440, 1216, 826, 1262, 2229, 4391, 787, 1325, 2265, 2513, 2468]. bande [1884, 1216]. Banesh [2751]. Bang [2845, 3610]. banh [279]. Bannerstone [2336]. Banu [3420, 4023]. Banyuls [1524]. Banyuls-sur-Mer [1524]. Baptiste [4475, 2108, 1186, 1187, 4552]. Baranowski [857]. Baratte [2648]. Barbara [4243, 3651, 2697]. Barbaras [4432]. Barbéra [1469]. Barberini [905]. Barcelone [897]. Barcia [3627]. Bardinet [3745]. Barjansky [3654]. Barjot [3558]. Barker [3751]. Barnabas [2733]. Barnes [4050]. Barock [875]. Barometer [1471]. barométrique [1918, 1204]. Baron [19, 1959]. baroque [2925]. Barrachi [2270]. barrages [1581]. barrages-voûtes [1581]. Barreau [3343, 3145, 3518, 2973]. Barrow [568]. Barry [3439]. Bart [3673]. Barthélemy [4286, 2256, 3514]. Barthez [2585, 1904]. Bartholin [2326]. Bartholomew [3381]. Bartolomeo [4543]. Bartomeu [904]. Bas [1850, 841, 413, 474, 440, 502, 523, 579, 664, 724, 782, 935, 1229, 1533, 1679, 10, 298]. Bas-Languedoc [1850]. basalt [4090]. basalte [4090]. Based [3165, 2155]. Basel [1928, 826]. Basis [3747]. Basler [1689]. Basque [2448]. Bassani [4161]. Bassermann [1420, 1262]. Bassermann-Jordan [1420, 1262]. Basson [2857]. bataille [3160]. batava [3488, 3611]. bateau [1896]. Bates [89]. Batsch [1783]. Battersby [2159]. Battista [4088, 1333]. Bau [1680]. Baudière [3767]. Baudrimont [3920]. Bauer [3372, 1383, 858]. Bauhin [387]. Bauingenieure [1734]. Baumé [692]. Baumeister [3489]. Baxter [216]. Bayer [894]. Bayet [3151, 4639]. Bayonne [2583]. Bazin [1558, 3884]. Bazin-Tacchella [3884]. Bd [3516, 3815, 4097, 4175, 3489, 2478, 2885, 3040, 2470, 2647, 3731, 4234, 3503, 2843, 3416, 3418, 3214, 2671, 2785, 3316, 2287, 2835, 2958, 3035, 3156, 2823, 3443, 1693, 2545, 3843, 4006, 3419, 2392, 2872, 2391, 2701, 4198, 3346]. be [3637, 4658]. Beagle [3455, 456]. Beam [2848, 4626]. beasts [3586]. Beatrice [2610]. Beatty [3434]. Beau [1730]. Beaudoin [1896]. Beaujouan [4286, 1231, 3741]. Beaulieu [3278, 2527, 2658, 2964, 3124, 3628]. Beaumont [3289]. Beaune [3515, 2291, 2089, 2588, 3722, 2087]. Beauregard [3055]. Beauvais [1409]. Beaux [346]. Beaux-Arts [346]. Bechetoille [2263]. Beck [1238]. Becker [394, 788, 791]. Becquemont [4556, 3382]. Becquerel [35, 88, 1472]. Bede [3754]. Bedel [3013]. bedeutender [2647]. Bedeutung [3734, 1106]. Bedini [1893, 1415, 2271]. bedside [3869]. Beeckman [3355, 3046, 2857, 3143, 2907]. Beekman [2883]. Beer [1263, 1578, 1827]. before [3377, 3637]. Béfort [4291]. Begegnung [1332]. Beginn [2661]. beginning [4573]. beginnings [2960, 2378]. Begriffes [4053]. Begründung [954]. Béguyer [1611]. Beiheft [2473]. Beihefte [4125, 188, 482]. Beijerinck [1109]. Beijing [4653]. being [2564]. Beitrag [3314, 2479]. Beiträge [2706, 690, 2287, 3346, 3815, 4097, 4175, 2478, 4125, 2661, 2674, 4234, 2479, 2707, 2685, 3443, 3444, 1332, 2829, 2566]. Béla [2130]. Belaval [3176, 1219, 393]. Belge [1433]. België [991, 991]. Belgique [3924, 362, 464, 231, 411, 438, 1161, 1744, 10, 128, 294, 662, 2266, 2267, 1604]. Belhoste [3746, 3723, 3825]. Belin [1406]. Bell [1025, 447]. Belle [3948]. Belles [892, 2916, 524, 4285, 2288, 4565]. Belles-Lettres [892, 524, 4565]. BellesLettres [2627]. Bellido [3960]. Bellivier [670]. Bellone [2471]. Belloni [1148]. Belmar [4255]. Belna [4139]. Belot [4140]. Ben [534, 3118]. benachbarter [2603]. Benavides [3735]. bench [3869]. Bendall [2957]. Benedetto [1614]. Bénédict [4485, 3749, 4090, 3418]. Bénédicte [4264]. Bénédiction [3993]. Benedictovitch [1970]. Benelux [896, 896]. Benfey [1439]. Bengt [876]. Benguigui [2900]. Bengy [285]. Bengy-Puyvallée [285]. Benitez [4296, 4066]. Benjamin [1468, 621, 2639]. Bennett [3603]. Benoît [3288]. benoeming [991]. Benoit [3515, 3525, 2352]. Bensaude [4477, 3786, 4255, 3559]. Bensaude-Vincent [4477, 3786, 4255, 3559]. Bentham [2643]. Bentley [946]. Benzenhöfer [3316]. beofening [1482]. Beppo [152]. berceau [311]. Berck [1896]. Berengario [3363, 951]. Beretta [3941, 3560]. Berg [2470]. Bergbau [2849, 1106]. Berge [3636]. Bergerac [2503, 4060]. Bergeron [3052]. Bergia [3344]. Bergier [283]. Bergman [793, 747]. Bergson [4192, 2061, 3938]. Bergwerk [211]. Bergwerk- [211]. Bericht [2479]. Berichte [3489]. Berio [2395]. Beris [372]. Berkeley [2733, 2959, 2881, 2982, 2875, 981, 955]. Berkeley-Los [2733]. Berkov [1332]. Berlin [4200, 1323, 3865, 942, 3093]. Berliner [1332, 2355]. Bernabé [4458]. Bernadette [4477, 3786, 4255, 3559]. Bernard [743, 16, 63, 1102, 1016, 1545, 4232, 3943, 632, 846, 973, 3487, 2470, 3118, 2576, 4343, 3935, 3342, 3313, 2472, 1689, 1667, 1955, 2107, 2306, 3725, 2696, 585, 4199, 3606, 907, 3698, 2528, 2752, 674, 3261, 4166, 3897, 2869, 2743, 3609, 1541, 3151, 3499, 1994, 2439, 3242, 2466, 1487, 92, 1087, 1381, 1322, 1460, 3927, 457, 2070, 33, 1243]. Bernardi [3062]. Bernardini [1571]. Bernardino [3153]. Bernardo [3950]. Bernd [4234]. Berner [2685]. Bernhard [3201, 1680]. Bernier [2293]. Berninger [2566]. Bernoulli [1928, 702, 4382, 4544, 826]. Berr [490, 1032, 421]. Berry [1144]. Bert [3119]. Bertele [1420, 1262]. Berthelot [138, 4610, 3233]. Berthollet [2446, 2704, 1991, 2100]. Berthoud [2997]. Berthoz [4033]. Bertin [484]. Bertoloni [4067]. Bertomeu [4255]. Bertotti [3344]. Bertrand [1347, 1544, 3577, 3388, 91, 1923, 2589]. Beruni [1605]. Berzelius [3485]. Besancon [892]. besoin [3168]. Besombes [1168]. Besomi [4078]. Bessel [3731]. Besso [2090]. Besson [4050]. bestiary [3586]. beta [4079]. Bétancourt [1585]. betrachtet [2603]. Betsy [512]. better [3926]. Betty [1936, 531]. Between [4588, 2935, 4660, 2877, 4653, 4663, 3749, 3418, 4430, 4656, 4670, 4604, 3869, 3588]. Bewegung [3443, 1564]. Beyond [4584, 1475]. Beziehungen [1332]. Bézout [4469]. Bezzola [4256]. Bharati. [2816]. Biagi [1614]. Biagioli [3781]. Bible [2809, 485]. biblical [3589, 4294]. Bibliografia [857]. bibliografici [2395]. Bibliographen [1070, 1298]. Bibliographia [1022]. Bibliographica [4066]. Bibliographical [1266]. Bibliographie [497, 2938, 3191, 4342, 4186, 3716, 3404, 1639, 1868, 2037, 3333, 3450, 3647, 1512, 4117, 1470, 2265, 2513, 1177, 699, 2842, 1303, 2646, 2157, 1922, 471, 822, 1060, 3349, 3244, 615, 3753, 2324, 153, 184, 1003, 1701]. Bibliographien [1564]. Bibliographies [2949, 3555, 2379]. bibliographique [4187, 1774, 2250, 2247]. bibliographiques [4358, 4522]. Bibliography [793, 1690, 418, 940, 1268, 2875, 1486, 1754, 2840, 2959, 2784, 2938, 2505, 1265, 1502, 2323, 2393]. Biblioteca [3694, 3901, 2606, 2395, 4344, 989, 3936, 3941, 4270, 4098, 2787, 603, 3523, 1907, 3583, 3923, 4072, 2106, 904, 4070, 2874, 4067, 4076]. Bibliotek [2475]. Biblioth [3703, 3664]. Bibliotheca [2131, 3488, 2233, 590, 2249, 3611]. Bibliothek [3416, 1262, 4361, 2872, 2740]. Bibliothéque [93, 95, 3793, 3785, 4411, 275, 2318, 357, 161, 445, 447, 4642, 3529, 1346, 1075, 1604, 2105, 2405, 376, 1536, 2213, 243, 372, 421, 1304, 1538, 3715, 92, 4601, 2204, 1906, 2291, 1538, 587, 4050, 4390, 693, 3121, 3761, 3385, 4441, 792, 3725, 3970, 4096, 2380, 1786, 3490, 1472, 865, 3096, 4647, 4346, 2288, 1230, 1231, 1695, 504, 1948, 2824, 902, 3453, 2732, 4237, 827, 839, 969, 378, 2243]. bibliothèques [1400]. Bicentenaire [124, 3167, 3168, 1808, 3072, 3554, 1223, 26]. Bichat [671, 461, 3968]. Bicquard [2745]. Bien [3824, 3582]. Biermann [2384, 1323, 2829]. Biezunski [2611, 2748]. bifurcation [2679]. big [3610]. bigness [2516]. Bigorre [1869]. Bigourdan [668]. bij [1793]. bijzonder [1793]. Bilan [283]. Bildanhang [2695]. Bildgedicht [2695]. Bildung [1680]. bilingual [4090]. Bilinski [2606]. Billard [3922]. Billroth [964]. Bilodeau [4264]. Bilson [2992]. binômes [1738]. Bineau [2108]. Binet [458, 66, 4444, 457]. Bính [381]. Bio [2840, 1774, 4358, 1303, 2646, 2973, 2247, 1003, 1701]. bio-bibliographie [1303, 2646, 1003, 1701]. bio-bibliographique [1774, 2247]. bio-bibliographiques [4358]. Bio-bibliography [2840]. bio-psychologique [2973]. biobibliographie [2066, 2067, 2065, 2269, 2270]. biobibliographique [1849]. Biochemical [2840]. Biochemist [3461]. Biochemistry [2264, 2063]. biochimie [3854]. biochimiques [3075]. bioénergétique [1947]. Bioesmat [4640]. Bioesmat-Martagon [4640]. Biogéographie [3374, 3376, 3373]. biogeography [3377]. Biographen [1070, 1298, 1564, 482]. Biographical [1268, 2110, 512]. Biographie [303, 1380, 2134, 2412, 1302, 2258, 4640, 2237, 2398]. Biographien [2647]. Biographies [3835, 2401, 1867]. biographique [560, 3216, 355, 3300, 3435, 1826]. biographiques [2139, 153]. Biographischer [1323]. Biography [2988, 2321, 3361, 3673, 1499, 3705, 2699, 2630, 3748, 3369, 3789]. Biohistoricae [2534]. Biologen [2647]. Biologia [3454, 1324, 566, 1324].

Biological [1025, 1688, 3743, 3177]. Biologie [3767, 1524, 1659, 2045, 2841, 1173, 1689, 4673, 341, 510, 3635, 226, 1477, 2497, 1813, 4679, 2667, 1815, 4674, 3951, 643, 2250, 4678, 2839, 2293, 3606, 3918, 4346, 1283, 122, 230, 196, 270, 366, 470, 3953, 1988, 4045, 454, 815, 4554, 2044, 1613, 835, 3767, 630, 4621, 1170]. biologique [4477, 3739, 2002, 2439, 1302, 4642, 4559, 385]. biologiques [3606, 3633]. biologiste [159, 1909, 2666]. Biology [2667, 4673, 3177, 3632, 4679, 3590, 3640, 3917, 4678, 3762, 3588, 4674, 2042, 2133, 2294, 2511, 37, 3528, 3945, 3590]. biophysique [63]. Biosphere [3972]. Biosphère. [3972]. Biot [4505, 4568, 4552]. biotechnology [3629]. Biran [4176]. Birembaut [3013, 3050]. Birgitta [793]. Birth [3165, 3120, 4572, 3578, 4475, 4183, 4197]. birthday [2939]. bis [2478, 2885, 1945, 3536, 2780, 2031, 2917, 1689, 2671, 3316, 954, 2229, 3843, 787, 1325]. Bischoffs [2479]. Bisher [278]. Bishop [675]. Biskup [2334]. bisontins [1917]. Bitbol [3578, 4616]. Bitschai [691]. Bitter [747]. Bitter-Selzer-Spa- [747]. Bizantini [565]. bizantinismo [565]. Björnbo [2406]. Bkouche [3433]. black [4182, 2390]. Blackmore [3826, 2248, 3500, 2835]. Blainville [2432, 2431, 2435, 2433, 2434, 1988, 2436]. Blaise [61, 4492, 4206, 1437, 1570, 3596, 4132, 3782, 4586, 4133]. Blaise-Pascal [3596]. Blake [2581]. Blamont [4629]. Blanc [2873, 401, 4236, 3756]. Blanchard [3793, 3761, 3096, 2542]. Blanché [651, 1785, 2331]. Blanckaert [4390, 3742, 4052, 3582]. Blaringhem [4155]. Blasius [689]. Blätter [1595, 1356]. Blaviani [989]. Blay [2913, 3807, 3302, 3787, 4005, 3561, 3453, 4141]. blé [4485, 4028]. bloc [4671]. block [4671]. Blois [2089]. Blondel [3694, 2879]. Blondiaux [4052]. Bloodletting [3125]. Blum [3077]. Blumenberg [1666]. Blundeville [492]. Blustein [3579]. Boas [2529, 1522, 2528, 3195, 862]. bobine [1411]. Bobinger [1935, 1693]. Bodenheimer [246, 247, 534, 929]. Bodet [3309]. bodies [1103]. Bodson [3460]. Boèce} [3495]. Boeren [2249]. Boerhaave [2342, 2141, 1679, 1022]. Boerhaaviana [1022]. Boethius [3516, 4006, 4391, 3419, 4238, 4198]. boeufs [2055]. Bogaert [2579]. Bogdan [2352, 2154]. Bohr [135, 3411, 3040, 3238, 3315, 3371, 3372, 3699, 3006, 2935, 2933, 2931, 3263, 2937]. Bohumil [1560]. Boirel [2736]. Bois [3059]. Bois-Reymond [3059]. Boisseau [1443]. Boissier [1782]. boissons [1061]. Boissy [2408]. Boistel [4487]. Boiteau [2319]. Bok [3673]. Boklund [747]. Bolens [2272, 2758]. Bölling [3667]. Bologna [1020, 3386, 1848, 420]. Bologne [4196]. Boltzmann [3826, 3384, 605, 2835]. Bolyai [4348]. Bolzano [2470, 3897, 3888, 3890, 3892, 3889, 3891, 1087, 1156, 1322, 3894, 2018, 3893, 3609]. bon [4016]. Bonah [3827]. Bonelli [2344, 989, 1469, 2224, 2690]. Boniface [4287]. bonne [924]. Bonner [1477]. Bonnet [3631, 378, 4543, 3738, 4027]. Bonnie [3579]. Bonser [1442]. Bontems [4508]. booc [2462]. Book [1931, 4627, 4614, 4623, 4626, 1928, 1779, 1943, 2826, 1923, 4617, 4615, 4616, 4624, 1912, 1946, 1911, 1915, 1945, 1913, 1944, 1914, 1892, 1887, 1888, 4645, 4644, 4628, 1909, 3650, 4641, 1960, 3838, 1925, 1956, 1941, 1953, 1938, 1939, 1926, 1927, 1955, 1929, 1958, 1959, 2570, 1940, 4625, 4642, 4619, 1933, 1908, 1907, 1954, 1936, 4640, 3808, 4643, 1889, 4646, 4629, 1906, 1910, 1924, 1896, 1895, 1894, 1890, 1893, 1891, 1897, 4647, 4618, 1922, 1935, 1934, 2804, 1930, 3497, 833, 4620, 1932, 1942, 4622, 1937, 1957, 4621, 3826, 4655, 3651, 4330, 247, 3497, 3652, 1908]. booklet [1758]. Books [485, 2581, 2452, 2054, 3420, 627, 2866, 4576, 4585, 4594, 4605, 4601, 4294, 608, 833, 1301]. booktitle [4600]. Boole [3990, 3310]. Booth [800]. bord [1116]. Bordeaux [3785, 2967]. bordelaise [3785]. Bordeu [2574]. Boréal [3591]. Borel [1673, 1697, 42, 4584, 243]. Borelli [364, 2]. Borelly [1839]. Borgato [2874]. Boris [478, 4393, 1580, 4395]. början [748]. Bornet [3990]. Borrelli [4068]. Borscheid [2391]. Bortolotti [829]. Bos [2659]. Boschung [4243]. Boscovich [2184, 4017, 1133, 1141, 3656, 3657, 3658, 1235]. Boscovitch [783, 1069]. Bosković [2184, 1105, 2854, 1969]. Bossi [4207]. Boston [1931, 3533, 3880, 3700, 3826, 3886, 3650, 2565, 4207, 3002, 3500, 3699]. Botanical [221, 1780]. Botanique [2413, 1378, 3161, 1762, 4030, 4285, 1319, 656]. botaniste [495, 3625, 1615, 735]. Botanistes [948]. Botany [2030, 3482, 3640]. Bothe [3091]. Bots [2308]. Bottazzini [4188]. Bouasse [4250]. Bouchard [304]. Boucher [3308]. Boudet [3884]. Bouguer [4469, 4468, 4016, 4464, 4465, 4467, 1514, 1512, 1513, 1423]. Bouillot [3863]. Bouleau [4189]. Bouligand [241, 68, 788, 832, 860]. Bouly [1250]. boundaries [3548]. boundless [4082]. bourbachique [2507]. Bourbaki [944, 1387]. Bourdier [484]. Bourgeois [4196, 3664, 2549]. Bourgeoisie [2853]. Bourgne [1294, 2311]. Bourgogne [1846, 3633]. Bourguet [3935]. bourguignon [79, 3433, 1677]. Boursier [4196]. Boursin [1978]. Bousicot [2268]. Boussingault [3008]. Boutaric [149]. bouteille [1489, 3428, 1221]. Boutlerov [926]. Bouton [2902]. Bouveresse [3343, 4231]. Bouvet [752]. bouwt [513]. Bowman [3060]. boxes [4182]. Boye [4008]. Boyer [4179, 3676, 1925]. Boyle [3654, 3649, 862, 3794, 569, 4085, 2487, 114, 3928, 3192]. Bozec [1000]. Bozzi [2914]. Brachay [1352]. Bradbury [2036]. Bradley [1414]. Bradwardine [4488, 789, 4132, 4133]. Braffort [1787]. Bragg [89, 3713]. Brague [3650]. Brahe [4402]. Braillard [1118]. brain [4055, 3125]. Brandi [506]. Brandli [4390]. Branly [3404, 3400, 3399, 3398, 3403]. Brannigan [3747]. Bras [1075]. Brassart [1480]. Braun [4243, 3401]. Braun-Bücher [4243]. Braunstein [4433, 4450, 4451]. Bray [3177]. Brazilian [2378]. breakthrough [3840]. Bréchignac [4645]. breeding [89]. bref [4306]. Breger [2958]. Bréhier [185, 370]. Brenner [4509, 4641]. Brésil [32]. Bresson [3458]. Bressou [2340]. Brest [963, 1708, 620]. Bret [3937, 4620, 3936]. Bretagne [3832, 4034, 1067, 1257, 1563, 1746, 1860, 618, 2014, 1498, 1532]. bretons [1703]. Bretz [964]. Breuken [1793]. Breve [154, 994]. breves [951]. brevets [1254, 967]. Breviarium [213]. brevium [2600]. Brezinski [4617].

Brian [3836, 3828, 3921, 3695]. Briefe [3223, 2449, 2460]. Briefen [2987]. Briefwechsel [3223, 826, 1332, 2355, 3059, 1609, 3667, 2829]. Brieux [3049]. Brigitte [4625]. Brill [2545]. Brinkley [3815]. Briot [2035]. Brisseau [3475]. Brisson [737]. Bristol [3459, 2753]. Britain [89, 402, 4230, 2445, 3743, 4489, 3967]. Britannicum [1685]. britannique [3929]. British [89, 2275, 1511, 1797, 1946, 2039, 2116, 3155, 673, 1237, 1578, 2226, 2112, 1261, 3377, 4456]. Brito [2287]. Britonis [2287]. Britton [3501]. Brittonia [221]. Brno [1843]. Broad [3307]. Broca [1303, 1051, 2646]. Brocas [881]. Brocbank [727]. Brock [1866, 2987]. Broda [605]. Brodie [1866]. Brodny [691]. Broglie [165, 135, 324, 1674, 165, 72, 448, 391, 2939]. Bromberg [3580]. Bronislaw [2606]. Brooke [3367]. Brooks [2113]. Brosse [2413]. brothers [3004]. Brotóns [2926, 3994]. Brouaut [890]. Brouillon [292]. Broussais [3242]. Broutet [215]. Brouwer [3100, 4581]. Brown [487, 3902, 3165, 4257, 2639, 3210]. Browne [3060, 3748]. Bru [3698, 3732]. Bruce [2342, 3436, 4348, 2836]. Bruhl [4366]. Bruijn [2185]. Bruins [1290, 1790]. Brullard [2512]. Brun [2613, 2518]. Brunés [1789]. Brunet [2627, 268]. Bruni [1216]. Bruning [940]. Brunner [651]. Bruno [508, 3746, 2077, 3723, 4490, 2244, 2060, 3224, 3825, 3151, 610, 4069, 1217]. Brunold [398, 117]. Brunot [2901]. Brunschvicg [4641]. Brunsdon [3586]. Brusch [3384]. Brush [2127]. Bruter [2227]. Bruxelles [1744, 1538, 333, 135]. Bruylants [2108]. Bruyns [4478, 3733]. Bryden [2807]. Brykman [4229]. bshs [2686]. Bub [3898]. Bucarest [1861, 2601]. Bucciantini [3523, 3923, 4344]. Bucciarelli [2664]. Buch [2287]. Bücher [1413, 4243, 1777]. Buchillustration [1639, 1868, 2037, 2265, 2513]. Buchwald [4197, 3696]. Bud [4018, 3629]. Budé [2055, 1079]. Buenos [62, 154, 566]. Buenos-Aires [62, 154]. Buffon [1338, 4317, 312, 407, 3573, 1063, 3717, 1573, 1062, 3515, 1574, 484, 4455]. Buffon-Cramer [1062]. Buffon^{≫} [182]. Bui [3970]. Buican [3104]. Builder [3194]. Building [3381, 3262]. Bulgarie [1821]. Bulletin [279, 1578, 3157, 249, 1271, 1354, 2583, 184, 528, 2816, 738]. Bullisiana [2614]. Buonamici [3414]. Bureau [3212]. Burgess [2129]. Burgière [4147]. burguesa [2635]. Burgundio [3503]. Burke [2641]. Burkhardt [3153, 3060, 3455, 3222, 2511]. Burndy [4348]. Burnet [2325]. Burning [2419]. Burov [3570]. Burr [695, 531]. Burrill [1109]. Burton [3083]. Busard [3516, 1163, 1792, 1793, 1794, 4006, 4391, 4198]. Buscaglia [3631]. Business [3655, 3861]. Bussières [1677]. Bustinza [1171]. Bustius [2479]. Butler [3083]. Butterfield [276]. Butts [1974]. Butzer [3668]. Buu [1305]. Bynum [3630]. Byzance [4510]. Byzantine [247, 1068]. byzantines [4510, 1532]. Byzantinisches [1795]. byzantins [3607, 3758].

C [1267, 1265, 1266, 89, 3767, 939, 2527, 2684, 3903, 3260, 3506, 13, 60, 2826, 4003, 2274, 2295, 3517, 1151, 1913, 2115, 3487, 359, 3433, 3361, 327, 2662, 373, 514, 2637, 2319, 3061, 3337, 2639, 3105, 3505, 3838, 1432, 2172, 3209, 2267, 2339, 2522, 2641, 2692, 800, 1059, 1148, 2340, 2974, 1959, 3126, 614, 525, 2773, 3868, 1338, 2539, 3149, 1616, 3206, 2977, 4643, 3439, 3702, 1268, 2024, 2158, 2216, 2113, 3308, 240, 530, 1942, 3041, 2580, 2249, 1301, 1246, 2242, 2243, 2468, 2816, 4489, 2855, 3211]. C. [2414]. C.-Fr [530]. C.N.A.M [1724]. C.N.R.S [948]. c/o [2855]. Côme [4081]. cône [1621, 1622]. côté [4133, 2324]. Côte-d’Ivoire [2324]. ca [2229, 608]. cabalistes [4058]. Cabanis [598, 3463, 2709]. Cabaret [3832]. Cabinet [1052, 4613, 1155]. Cabrera [238]. cachée [4066]. cachetés [2954]. Cacouros [4510]. Cadden [3480]. Cadeddu [3454]. cadrans [668]. cadre [3221]. caelestium [2656]. Caelius [361]. caelo [3557]. Cahan [3581]. Cahen [249]. Cahier [1247, 1545, 1270]. Cahiers [4477, 2657, 3359, 3721, 2396, 3433, 838, 4646, 3633, 2273, 4509, 2593, 3955, 1601]. Cahn [1831]. Cajal [2644, 4480]. Calascibetta [4076]. Calcidius [4549]. Calcoen [1604, 2105, 2405]. Calcolo [830]. Calcul [2568, 3427, 3777, 2982, 3025, 69, 1327, 1608, 1958, 3187, 1989, 3291, 2725, 1006, 2404, 2867, 41]. calculation [2725]. Calculus [1296, 787, 4020]. Caldani [2882]. Calendar [2334, 2686, 2881]. Calendars [3830]. calendrier [43]. calendriers [3197, 668]. calibres [1249]. California [3581, 3871]. Callandreau [236]. called [4198]. calling [3441]. Calmette [1243]. Calpe [62, 566]. Caltech [3590]. Calvert [1758]. Calzecchi [3402]. Cambodge [1309]. Cambridge [3810, 1863, 3480, 3526, 3630, 3864, 3275, 3565, 3367, 3459, 2309, 3579, 3274, 3868, 3835, 3762, 4055, 2993, 3967, 2564, 3037]. Cambridge-New [3864]. cameralism [4561]. Camerino [3368]. Camerota [4122, 4345]. Camille [4642, 837, 2952, 2729, 1225]. Camillo [3901]. Caminer [3861]. camouflet [4588]. Campanella [874]. Campanus [4391]. Campbell [608]. Camper [3015]. Campo [2256]. Campus [2958]. Canacci [603]. Canada [1502, 2136, 1320, 4539]. Canadian [3567]. canadienne [1424, 1612]. canal [4021, 3221]. canalizzazione [2174]. Canarios [1108]. canaux [3221]. cancer [3472, 3473, 3869, 3502]. Candaux [4615]. Candille [1175]. Candolle [3224, 2017, 2003, 3821, 1992]. Caneva [3605]. Canevari [2395]. Canguilhem [2318, 2666, 4642, 3914, 15, 3311, 4679, 4663, 4675, 3515, 4674, 3916, 3917, 1953, 4678, 4673, 4689, 4676, 3918, 4677, 4451]. Cannizzaro [3383]. Cannon [2831, 2523]. Canone [4069]. canons [3097]. Canterbury [4234]. Cantor [3337, 3689, 3203, 4512, 4513, 1682, 4139]. Cantù [2344]. Cap [1102]. Cape [1936]. Capek [2061, 1239]. CAPHÉS [4665]. capillarité [3137]. Capitaine [3409]. capitaliste [2711]. capitolo [4122]. Caplat [3216]. Capp [2696]. Capparelli [901]. Cappellari [2977]. Capponi [3562]. Capra [2874]. Caprina [4196]. captain [3262]. capture [3933]. caractère [4354, 585]. Caractéristique [3448]. Caractéristiques [1882]. Caravelli [985]. Carbonne [4022]. Carcavi [1918]. Carcavy [3782]. Cardan [3413]. Cardano [2903]. Cardini [134]. cardio [4521, 2845]. cardio-vasculaire [4521, 2845]. cardiologie [4293]. cardiologique [1457]. Cardon [4089]. Cardwell [903]. career [2137]. carence [3490]. Cariage [1756]. carie [4485]. Carinthia [1792]. Carl [2332, 13, 2384, 1887, 1925, 1028, 3697, 3642, 3970]. Carles [1478, 161]. Carlo [2406]. Carlsberg [730]. carme [4286]. Carmody [1101]. carnivores [4450]. Carnot [2509, 2637, 1958, 4349, 1891, 2377, 2593, 2139, 4547, 1491, 2239, 3900, 834]. Carobbi [2358]. Carole [4434]. Carolin [2534]. Caroline [4315, 3626]. Carolingian [3668]. Caron [945]. Caroti [3562]. Carozzi [3749, 3348, 4090]. Carpentier [3211]. Carpi [3363, 951]. Carra [19]. carré [4133]. carrefour [4169]. Carrel [2410]. carrés [3135, 52, 481]. carrière [2138, 1083]. Carro [3336]. Carrol [3060]. Carroll [1888, 4578]. carrosses [290]. Carroy [4546]. Carsten [4079, 4622]. Carstensen [2847]. Carta [2471]. Cartan [2168, 4429]. Carte [1742]. Carteggi [3240]. Carteggio [2490, 2882, 4351]. cartelle [4515]. Cartelli [2567]. cartes [2962, 4347, 4054, 1075]. cartésianisme [2240, 548, 234]. cartesianum [2631]. cartésien [2088]. cartésienne [4224, 2487, 924, 4051]. cartésiennes [3802, 3798]. cartographic [4653]. Cartographie [1744, 2252, 161]. cartographique [1819]. Cartography [3259]. Carvais [3151]. Carvalho [529]. Cas [4016, 3400, 4156, 179, 4028, 4518, 3293, 4564]. cas-témoins [4518]. casa [885, 1721]. Casanova [4196]. Case [4602, 631, 3220, 4406]. Caselli [1725]. Casimir [1909, 648]. Casimir-Joseph [1909]. Caspar [812]. Caspard [2974]. Cass [1973]. Cassar [1833]. Cassidy [2959, 3517]. Cassina [1426, 1074, 1099, 1291]. Cassini [4571]. Cassirer [4142, 3336, 4666, 4666]. Cassou [4474]. Cassou-Noguès [4474]. Castel [3809]. Castellani [2319]. Castiglioni [248, 420]. Castle [730]. Castres [1623]. catalan [4286]. catalana [904]. Catalogo [989, 2395, 989]. Catalogue [359, 235, 2581, 1175, 1940, 2006, 2380, 2216, 1230, 1521, 1854, 1079, 1538, 4478, 2027, 2902, 905, 2342, 1258, 1259, 1705, 1982, 1846, 3345, 840, 2677]. Catalogues [2949]. Catalogus [3143]. Catalyses [3606]. cataracte [3475]. catastrophisme [2860]. catégorie [2246]. catégories [4554]. Catelan [775, 927]. cathédrales [968]. Catherine [4557, 4168, 3533, 3703, 4434, 3372, 4645, 4415, 4416, 3933, 3932]. cathodiques [3688]. cathol [3924]. Catholic [2633]. catholique [3572, 184]. Catoptrique [988, 3496]. Cauchy [2458, 3009, 2568, 3326, 3331, 3333, 3328, 2152, 3330, 682, 1156, 2018, 3329, 1903, 3327, 2596, 1959]. Cauchyjev [2568]. Caudatus [2851]. Caullery [816]. causal [4519]. causalism [4604]. causalisme [4604]. Causalité [4550, 3873, 3676, 3138]. causation [4519]. cause [4134]. causes [3311]. causis [3119]. Caussat [1978]. Cavaillès [3026, 3031, 3027, 3662, 3028, 1386, 3856]. Cavalieri [3354]. Caveing [4310, 2868, 3876]. Cavellat [3215, 3215]. Cavendish [1144]. Cavini [3105]. Cazenave [2354]. Cazenobe [4162]. CCXXXIV [1028]. Ce [4218, 3724, 752]. Céard [2097]. cebri [1388]. Cecilia [1567, 3005]. Cédric [3966]. Célébration [1563, 1402, 1985]. Celebrations [2501]. Celebrazione [877]. Célèbres [236, 94, 305]. celebrity [4610]. Celeste [2344, 3522, 3566, 3187, 364, 3160, 1403]. celestes [1542, 3924, 4643]. celi [2094]. Céline [4618]. Cell [1636, 4587]. celle [3071, 3028, 2833, 2347]. celle-ci [2347]. celles [314]. cellulaire [1965, 4503, 4587, 1024]. Cellular [2069]. cellule [4073]. Cellules [3502]. Celse [3879]. Celsus [3505]. celui [4016, 4279]. Cena [2244]. Ceneri [2244]. Cenis [465]. censure [4686]. Census [3011]. Cent [461, 1824, 145]. Centaur. [1030]. Centaurus [537, 593, 888, 972, 1120, 332]. Centenaire [1563, 1743, 496, 1523, 1535, 2583, 2613, 549, 1402, 673, 1529, 105, 1742, 2653, 2611, 459, 1077]. Centenario [1469, 877]. centenary [4610]. Center [240, 4665]. centers [3282]. central [2204]. centrale [3715, 2318, 2380, 2851, 3168]. centrales [3874, 3622]. centrali [3901]. centralistes [4357]. Centre [4477, 5, 3504, 4003, 2396, 446, 2423, 2693, 2705, 886, 2260, 4530, 1735, 4363, 1846, 2405, 738, 3722, 3633, 2669, 1433, 4265, 766, 2407, 4665, 1463, 1341, 585]. centres [556]. centrifugation [4073]. Centro [2606, 2395]. cents [2842, 2910, 3554]. Centuries [3011, 4478, 3860, 1854, 2578, 2697, 3653]. Century [3459, 2753, 2310, 3000, 2294, 2506, 2581, 862, 2685, 2306, 2979, 2811, 2694, 1942, 247, 2457, 89, 4572, 3815, 3863, 4652, 1435, 359, 3581, 4664, 3567, 3839, 4020, 4203, 3521, 2838, 3636, 3868, 2372, 3574, 3860, 3659, 3728, 4006, 4317, 4573, 3476, 4654, 420, 4578, 4570, 186, 509, 1296, 3338, 3630, 3386]. céphalique [2687]. Ceramic [2452]. Ceramics [1940]. cercle [3352, 1927, 3268]. Cercles [2266, 2267]. cérébral [4480]. cérébrale [4480, 4618]. cérébrales [4355]. cerebrospinal [799]. Cérémonie [662]. CERN [3992]. Cernuschi [4615]. Ceroplastica [2787]. Cerruti [4435, 3383]. Certainty [2776]. certitude [3114]. cerveau [3594, 4033, 4343, 3518]. Ces [3280]. Ceskosclovensku [2394]. cétacés [2462]. cette [834, 2292]. Ch [90, 1757, 117, 1332, 2355, 3013, 1615, 2320]. Ch. [3099, 1555]. Ch.-Eug [90]. Château [4367, 3345]. Châtelet [3669, 1683, 4530, 1683]. Chabert [1320]. chacun [2243]. Chadwick [3967]. Chaia [3604]. chaire [3819, 3601, 3264]. chaires [273]. Chaldéens [1214, 197]. chaleur [2895]. chaleureux [3915]. Chaligne [1114]. challenges [4563, 4684]. Challinor [1025]. Chalmers [1774]. Chalubinski [3317]. Chambat [4615]. Chambdal [2634]. Chambéry [3569, 1148, 1941]. Chamonix [1349, 3749]. champ [4523, 2273, 4509]. Champagne [4296]. Champeix [1737]. Champion [3213]. Champion-Slatkine [3213]. champs [4445]. Chance [2069, 3434, 3434]. chancelier [2121]. Chancourtois [1611]. Chandra [3361]. Chandrasekhar [3361]. Chang [1777]. Change [2332, 1432, 2225, 2633]. changed [3434]. changing [4519, 4188]. Chanoine [2927]. Chantal [3345]. Chantier [4021]. chantiers [3547]. Chantilly [2423]. chap [3703, 944]. Chapdelaine [2614]. Chapelle [367]. Chapelon [3487]. Chaper [1023]. Chapiro [3345]. chapitre [3495]. Chapman [652]. Chappert [2376, 4452]. Chapront [3212]. Chaptal [4531, 1951, 660]. chapter [2848]. Chapters [3118]. Charaf [1112]. Charaf-ed-Din [1112]. Charbon [46]. Charbonnat [4526]. Charcot [611]. Chardin [327, 1478]. Chardon [963]. Chareix [4258]. Charitable [3868]. Charle [3043]. Charles [749, 89, 3939, 402, 2309, 2341, 3260, 2873, 2877, 535, 1078, 398, 191, 456, 2637, 2830, 3060, 3455, 4516, 3314, 1500, 3532, 3757, 2266, 1938, 984, 1608, 1663, 1927, 2055, 2056, 2093, 2773, 4178, 3754, 1057, 2810, 4239, 1895, 3004, 3076, 2528, 1230, 1521, 1846, 742, 1056, 1233, 356, 674, 4296, 1937, 4395, 378, 462, 4543, 923, 4381, 1857, 436, 2680, 4587, 2725, 3437, 3738, 920, 1225, 4027, 3631, 3748, 35, 2451]. Charles-C [2637]. Charles-Francois [4516]. Charles-Georges [4395]. Charlier [480]. Charlionet [4398]. Charlotte [3060]. Charmasson [4434, 4259, 3884, 2693]. Charrak [4410]. Chartres [4130]. Chassagne [4411, 2711]. chasseur [3353]. Chastellet [1828]. Chatelier [4349, 2026, 2026]. Chatelin [797]. Chauchard [93]. Chaudron [3834]. Chaunu [3244]. Chaussées [2901]. Chauvois [796, 1569, 1995, 510]. Chazal [3862]. che [773, 3039, 1390]. Chebyshev [4020]. check [667]. chef [1557]. Chefs [181]. Chefs-d’œuvre [181]. Chemical [1317, 2703, 3550, 4514, 3548, 4316, 4587, 3080, 4622, 1439]. Chemicum [1685]. Chemie [3214, 3489]. chemik [794]. Chemikers [1336]. chemin [3268]. Cheminement [4480, 2284, 4495]. chemins [4583]. Chemist [3008]. Chemistry [384, 1152, 1317, 1485, 1829, 1912, 1979, 218, 1867, 4609, 2063, 2988, 2472, 3560, 3080, 2476, 2390, 89, 4612, 3485, 37, 3786, 4514, 3386, 3653, 4611, 4607, 4608, 163, 1004, 1317, 1576, 400, 2640, 2474, 862, 532, 2062, 2572, 1485, 1475]. Chemists [2988, 2474, 4608]. Chemla [3525]. Chen [666, 3765]. Chene [4143]. chercheur [4189, 3490]. chercheurs [507]. Cherici [4618]. Cherni [4163, 4288]. Chernin [3570]. cheval [1892, 694]. Chevalley [1786, 4168, 3372]. Chevreul [4531]. chez [4549, 2431, 4598, 3422, 3425, 4000, 4332, 3426, 1747, 4014, 1456, 4356, 4657, 4431, 3199, 4550, 2433, 4305, 1571, 223, 645, 1218, 3998, 3114, 2432, 1811, 4587, 4673, 2839, 2327, 2236, 4429, 1360, 1841, 3449, 3964, 234, 568, 3803, 356, 3111, 1678, 3999, 4328, 3609, 3893, 4634, 869, 4554, 4302, 678, 4445, 2907, 2580]. Chiara [3626]. Chiche [1116]. Chick [2143]. chief [1665]. chiesa [3727]. chiffres [76, 197]. Childhood [2610]. Chilvers [4393]. Chimica [4161, 4076, 4270, 4435, 4435, 3383]. chimiche [4161]. Chimie [1822, 4085, 4416, 4608, 2745, 1901, 3552, 3874, 319, 4612, 658, 4312, 926, 320, 2439, 647, 4193, 4160, 4255, 4547, 4628, 4609, 2329, 2413, 2548, 177, 2896, 4337, 4531, 4611, 2348, 2302, 3920, 4169, 1474, 4607, 2783, 4092, 652, 450, 586, 3559, 1335, 4165, 451, 893]. chimique [3547, 20, 4298, 3413, 2014, 2367, 4587, 2108, 2141, 3507, 3690, 20, 90, 2704]. chimiques [2217, 1611, 1930]. Chimiste [4085, 794, 304, 1856, 409, 2446, 2797, 2897]. Chimistes [3969, 4257, 4547, 4608, 451]. Chin [1777]. China [665, 1565, 1297, 3080, 3177, 4652, 992]. Chine [3721, 2131, 3170, 3191, 3186, 3189, 3334, 2329, 3173, 1166]. Chinese [607, 3880, 3206]. chinois [2800]. Chinoise [1498, 3188, 3174, 643, 773, 1173, 3187, 3047, 952]. chinoises [3178, 3171, 2420, 3190, 2650, 3207]. Chirac [2799]. chirurgia [1503, 3735]. chirurgical [1707, 1112]. chirurgicale [4664]. chirurgici [2173]. Chirurgie [280, 1000, 1307, 1706, 4036, 279, 2218, 797, 1306, 1983, 1967, 3475, 1691, 1306, 1443, 1705, 2847]. chirurgien [1676, 1050, 2033, 3146]. chirurgiens [1115, 1579, 1708, 2268, 1114]. chirurgo [2077]. Chiti [1542]. Chladni [3816]. Chlorpromazine [2322]. choc [4111, 657]. choice [2876]. choisies [10, 1606]. Choix [601, 2507]. cholera [3166, 2537]. Cholesky [4617]. Chopin [631]. chose [4332]. choses [4157]. Chou [606]. Chouillet [4265]. Chr [303]. Christiaan [30, 303, 3674, 2659, 2819, 4110, 4111, 4116, 4114, 4115]. Christian [3747, 4082, 1692, 1609, 2054, 4414, 4528, 2404, 3827, 595, 657, 290]. Christiane [3059, 3828, 3490]. Christianidis [3700]. Christie [3337, 824]. Christine [3694, 2879, 1359]. Christoph [4527, 4053, 1668, 1791]. Christophe [3043, 998, 887]. Christopher [4393, 2694, 3967]. Christophore [3112]. Christus [3346]. Christus-Wahrheit [3346]. chromatique [4505]. Chronaxie [977]. chronic [361]. chroniis [3921]. chronique [1836, 3858, 1915]. Chronologie [2000, 3314, 1083, 714, 3823]. chronologies [3292]. chronologique [2996, 2953, 2969, 4119, 698, 1179, 1357, 2084, 2219]. Chronology [3012]. chronophotographie [3113]. Chrysopoeia [4286, 4491, 4084]. Chuong [1311]. Chuprov [2877]. Chuquet [2258]. church [3727]. chute [51]. Chymia [384, 218, 688, 163, 1152, 1317, 1485, 1798]. Chymici [2249]. CI [2295, 1472, 2347]. ci-dessus [1472]. Ciência [529]. Ciências [529]. Ciardi [4435, 4076]. Cibernetica [2739]. Cicenia [3829]. Cid [4453]. Ciel [2880, 534, 2254, 3121, 1997, 3280, 4054, 1818, 2820, 3787]. Cienc [1108]. Ciencia [3666, 2644, 1907, 566, 1324, 3592, 2805, 3994, 3735, 2926, 3732, 417, 1324, 300]. Ciencias [566, 4409, 885, 1324, 2635]. científica [529, 33, 3235]. cientifico [2805]. cieux [3160]. Cifoletti [3359]. cimeli [1469]. Cina [3339, 880]. cinéma [45, 3098]. cinématique [3354]. cinématographe [1250]. cinq [290, 458, 2842]. cinquante [461]. Cinquantenaire [977, 1824]. cinquantième [145]. cinquecento [2395, 4122, 4304]. Cioranescu [1108]. Ciphers [4238]. circa [2262]. Circle [4056, 240, 2477]. circles [4478]. Circular [2094]. circulari [2656]. Circulation [4656, 4653, 1702, 4597, 510]. citations [4423, 2350]. Cité [3408]. citoyen [3408, 3408]. città [4076]. Cittadino [3319]. Civic [1590]. civil [3142]. Civilisation [665, 3080, 2157, 3177, 1565, 2919, 3795, 1889, 1392, 144, 220]. Civilisations [4311, 1772, 272, 2904]. Civilita [565]. Civilization [3166, 3668, 162, 1888, 96, 250, 306]. Civiltà [880]. Cl [744, 1114, 1116, 1708, 1786]. Cl. [63]. Clagett [689, 1539, 1847, 982, 3830]. Claims [1895]. Clair [1849, 3413, 1504, 2247]. Clairaut [407, 2662, 3728, 471, 2400, 269, 287, 432, 557, 2219, 3658, 1828]. Claire [3896, 146, 4261, 4416, 1307, 4619, 3151, 4639]. claires [2502]. clairvoyant [2774]. clams [4236]. clandestine [4319]. clandestins [4066]. Clapeyron [1891]. Clara [4351, 693]. Clarac [350]. clarification [2516]. Clark [1942]. Clarke [3863, 2142]. Clasen [472]. clásicos [2965, 2644]. class [1919, 1919]. Classe [3934]. classement [2898]. Classer [3582]. Classic [1637]. classica [3488, 3611]. Classical [3501, 2608, 3432, 3281]. classicisme [4029]. classicos [2644]. Classics [1973, 1472, 2476, 1439, 2476, 1439].

Classification [1743, 3175, 2257, 1963, 2839, 1433, 2665, 2293, 3528]. classifications [4662]. classique [1858, 2868, 2656, 4491, 4005, 2919, 3608, 379, 4066, 3305]. Classiques [1437, 1730, 1474, 3302, 1663, 4492, 584, 3412, 1403, 4560]. Claude [3458, 4390, 3793, 1545, 4232, 4617, 2227, 3178, 3207, 1639, 1868, 2263, 3075, 4438, 3515, 2576, 4019, 4516, 3279, 3342, 1175, 1689, 986, 3648, 3970, 3965, 4292, 3582, 4084, 249, 3439, 3742, 4052, 2902, 4618, 3884, 2213, 3056, 3722, 3633, 457, 2446, 2704, 3760, 1615, 3, 2588, 2439, 2162, 2466, 1487, 3268, 1991, 2100, 1381, 2219, 3658, 2020, 1881, 1460, 2070, 3604, 16, 1016, 3305]. Claude-Bernard [2070]. Claude-Francois [2162]. Claude-Gaspar [986, 3]. Claude-Louis [2446, 2704, 1991, 2100]. Claude-Paul [2227]. Claude-Pierre [1881]. Claudette [3746, 4195]. Claudine [4264]. Claus [2037, 2265, 2513]. Clausius [1891]. Clave [4058]. clavecin [3809]. Clavelin [1858, 3704, 1954, 3790]. Clavius [3112]. Cle [4232]. Clément [4608, 1405, 1883, 1895]. Clémentel [4608]. Clendening [609, 631]. Cléomède [2656]. Cleonice [3359]. Clericuzio [3653]. Clerk [3792, 3991, 2918]. Clermont [3596]. clés [4433]. CLIII [3674]. climat [1695, 4360]. Climate [2113]. Climatologie [4539]. cliniche [4515]. Clinicien [3490]. clinique [4482, 962, 996, 1939, 4483, 4237, 4004]. Clio [3636, 3965, 4364, 3290, 1710]. Clocks [2710, 4143, 3830, 1894]. Clockwork [2710, 1017]. clos [3561, 1288]. closed [785]. Clouet [409]. Clusius [2196]. CLVI [1113]. CLXI [4083]. Clyde [487]. CNRS [2396, 2318, 4343, 4205, 4432, 4392, 4396, 3966, 3633]. cnum.cnam.fr [3987]. coaptations [1678]. Cocchi [4083]. Cockroft [3036]. Code [487, 3309]. codex [3674]. Codice [603]. Codices [2249, 2249]. Codification [3860]. coefficients [3331, 472]. cœlestium [4643]. cœlestium/Des [4643]. Coelho [3519]. coeli [4069]. Coenensoen [2462]. coeur [2916, 4007, 356, 4609]. cognitif [4575]. cognition [4033]. cognitive [4575, 3439]. cognitivisme [3439]. Cohen [1931, 3880, 652, 973, 2565, 3002, 2528, 3308, 743, 632, 846, 907, 674, 2869, 2743]. Cohérence [3263]. cohéreur [3405]. Cohn [1109]. cohorte [4521]. Coimbra [529]. Coladon [661]. Colbert [1186, 1187]. cold [4365]. Cole [872]. Colecã [4232]. colegios [1231]. Coleman [3235, 2846]. Coleridge [2838]. Colette [4196, 4487, 2749]. Coligny [3197]. Coll [2628, 1633, 3343, 2627, 2656, 753, 301, 948, 3302, 3521, 1662, 2843, 2603, 1954, 2972, 3736, 4007, 4146, 4147, 2673, 2675, 2736, 2756, 2669, 825, 2665, 2634, 2827, 2421, 4310, 2308, 2309, 2320, 2331, 2341, 4623, 2504, 2775, 2783, 2104, 2532, 2913, 2168, 2169, 2227, 2253, 2314, 4624, 2507, 968, 989, 4411, 4311, 327, 399, 3557, 4547, 4628, 2027, 2741, 1978, 1526, 2217, 2321, 2354, 2510, 1634, 1663, 2919, 4439, 2076, 3007, 4643, 981, 2822, 2823, 2813, 2225, 4629, 945, 4618, 2758, 3435, 2175, 2786, 2809, 2292, 2370, 2376, 2404, 2443]. coll [2751, 2752, 2824, 2832, 3098, 2264, 2771, 2732, 2386, 2373, 4620, 791, 1326, 1403, 1541, 1932, 2291, 1014, 2965, 2249, 2391, 2612, 2739, 2210, 1439, 1474]. collaborateur [4552]. collaborateurs [322]. collaboration [3476].

Colladon [814, 978]. Collana [4309, 3491, 4476, 4161]. Collectanea [4318]. Collected [3270, 3903, 3118, 3867, 3788, 3754, 3796, 3842, 3077, 3595, 3865, 4271, 4230, 4269, 4440, 4489, 3663, 4516]. collectifs [4525]. Collection [2870, 2529, 72, 1317, 138, 140, 2711, 73, 235, 3676, 2637, 324, 151, 1927, 2093, 1933, 4349, 2839, 2916, 2126, 905, 3300, 1606, 3442, 4296, 160, 667, 822, 246, 247, 2290, 3785, 22, 535, 444, 359, 371, 3585, 838, 1613, 242, 670, 1386, 1387, 4613, 863, 484, 421, 363, 476]. Collections [1940, 1854, 4612, 4644, 1169, 1304]. Collector [906]. College [2564, 1752, 1485, 727, 1933, 280, 759, 4195, 2859, 4033, 2576, 3819, 1706, 3601]. collèges [2915, 4289, 547, 550]. Collegium [2479]. collègue [4381]. collègues [2256]. Collet [1765]. Collier [1339]. Collini [4362]. Collins [3755, 3724]. Colloque [1055, 1161, 3631, 3578, 3343, 3700, 1207, 1659, 2582, 3043, 3970, 3664, 3299, 3759, 941, 3722, 3633, 1014, 3530, 4003, 4026, 4446, 2668, 3938, 4363, 1460, 1069, 1843, 1626, 1784, 1524, 2319, 1148, 3409, 1176]. Colloques [1053, 948, 3460, 3831, 3837]. Colloquia [2604, 2605]. Colnort [3309]. Colnort-Bodet [3309]. Colomb [887]. Colombaria [2976, 2358, 4083]. Colombe [444]. Colonel [89]. colonial [3592]. coloniale [4436]. coloniales [879]. colonialism [3638, 3527]. colonies [3753]. Colonna [2243]. colonnes [3064]. colorés [2895]. Colour [377, 2753]. Coloured [2826]. Colton [608]. combat [3966]. Combe [3261]. combinaison [3507, 3690]. combination [1439]. combinatoire [3494, 2404]. Combinatoria [2740]. comburentibus [3126]. Combustion [1389, 1254, 1730]. comentario [1436]. Comentarios [3994]. comète [432, 2994]. comètes [3924]. Comets [3012]. Coming [2757]. Comité [2295, 3208, 1943, 1835, 2879, 1230, 1521]. Comiti [2188]. Commandino [4602, 1948]. comme [4246, 2936, 4014, 4027, 4129, 2245]. commemorated [836]. Commemoration [2749, 1627, 295]. Commémorations [390]. Commémorative [662]. commencement [1739]. commencements [1727, 2680]. commensurabilitate [2094]. Comment [753, 4586, 1203, 1916, 3497, 3765]. Commentaire [3607, 2176, 86]. Commentaires [3063, 151, 2093, 2124, 1292, 1828, 984, 4120]. commentaries [1780]. Commentariolus [2231]. Commentarius [4202]. Commentary [3119, 3118, 3640, 3125, 3420, 3651]. commentateur [2164]. commentateurs [4304, 702, 4023]. Commentationes [844, 845, 219, 331, 2873]. commentators [3737]. commented [4581]. commentée [2326, 4581]. commentées [457, 10]. commentés [355]. Commento [4078]. Commerce [4608, 3362]. Commercial [1950]. commerciale [291]. Commercium [2662]. Commerson [2380]. Commis [193]. commissaire [4469, 4468]. Commission [3903, 4194, 3347]. Common [2642]. commun [4008]. communauté [3853]. communautés [4414]. Communication [2342, 949, 359, 836, 841, 3750, 2370, 865]. Communicationes [2534]. communications [89, 3578, 4182, 762, 3530]. Communis [2024]. communità [4270]. Community [3486, 3524]. como [238]. compact [1065]. Compagnie [161, 1114, 2177, 4645]. Companion [3337]. comparaison [3229]. comparaisons [19]. Comparative [3389, 3534, 3243]. comparatives [3721]. comparée [2212, 3819, 91]. comparées [3827]. compatible [4683]. compendium [4652]. Compère [4289, 2915]. Compétition [4670, 4670]. Compiègne [2665]. complément [3225, 1492]. complémentaire [3716, 4219]. complémentaires [3510]. Complémentarité [2934, 2938, 2933, 2931, 2936, 2937]. complementarity [2938, 2933, 2931, 2937, 2934, 2936]. complete [3903, 4689, 1544]. complètes [4615, 303, 4642, 4689, 1292, 1570, 2503, 30, 1102, 2591, 1753, 2140, 2156, 1781]. complets [2796]. complétude [3263]. Complexe [2532, 3958, 2760]. complexes [812]. complexité [4398]. composé [249]. composition [3494, 3688, 254]. compositions [387]. compréhension [570]. Comprehensive [2608, 2264]. Comprendre [4287, 4054, 3741]. Compte [1463, 459, 3408]. Comptes [1176, 1943, 2295, 1835, 2045, 1031]. Compton [2338]. computer [3861]. Computus [3754]. Comte [4450, 4266, 3705, 3931, 2434]. Con [3941, 1171, 2808, 3674]. Conan [330]. concentriques [2408]. Concept [4174, 1791, 4394, 2938, 2194, 3635, 3027, 4624, 1577, 3153, 2005, 4153, 3508, 4279, 4401, 2857, 3952, 4295, 2327, 3738, 3111, 3953, 4573, 3351, 2973]. conception [2017, 2003, 3431, 4452, 310]. conceptions [2017, 3046, 2750, 3888, 2003, 2347, 270]. Concepts [1487, 3449, 3139, 2486, 4519, 2127, 2738]. Conceptual [4592, 3524, 3486, 3389, 3794, 3808]. Conceptualisation [3188, 2913]. conceptuelle [4592]. concernant [3191, 194, 563, 3404, 15, 1738, 398, 3333, 1086, 1954, 756, 4116, 1254, 3790, 2015, 2349]. concerning [2938, 1665]. Concetta [4643]. concetti [4511]. concevait [1916]. Conciergerie [3221]. Concise [3206, 207]. conclusion [3986]. conclusione [2808]. Concorde [2410]. concours [273, 647, 2318]. condamnation [4669, 4058]. condemnation [4669]. Conditionnement [2973]. conditions [1432]. Condorcet [3698, 827, 1599, 2328, 4385, 313, 849, 870]. Condorcet. [2403]. Conférence [936, 1744, 3886, 279, 4439, 4507]. Conférences [1012, 933, 1093, 1823, 203, 1149, 1314, 1506, 1640, 1911, 403, 206, 491, 4367]. confidence [2993]. Configuration [4330]. configurationnelle [4592]. confins [3776]. Conflict [4592]. conflit [2532, 2975]. confœderatio [4061]. conforme [1102, 790]. congélation [1137, 1458]. Congourdeau [4510]. Congrés [231, 202, 349, 369, 896, 931, 1427, 1533, 1641, 1770, 1860, 1861, 1825, 1862, 2295, 1835, 625, 897, 499, 1461, 3569, 1699, 2044, 1206, 1031, 1176, 1532, 4510, 501, 764, 1462, 2601, 1943, 1463, 257, 2045, 2266, 2267, 1148, 1602, 786, 2583, 333, 160]. Congreso [4409, 3592]. Congresso [2787]. Conic [133]. Conics [3420]. Coniques [292, 1330, 3158, 552, 1181]. conjuntos [154]. connaître [395]. connaître^{≫} [138]. Connaissance [3604, 4679, 3372, 4675, 2681, 3292, 4164, 1105, 2196, 4142, 4229, 4476]. connaissances [2257, 1965, 3460]. connectionism [4024]. connections [4056, 3981]. connu [2221, 4113, 2266]. connue [1951]. conoïdes [1927]. conoscenza [4098]. conoscenze [4309]. conquête [3264, 2735]. conquérants [38]. conquise [3366]. Conrad [389, 2534]. Conry [2341, 3477, 2029, 2251]. consacré [459]. conscience [4405, 4357, 902]. conscripta [1216]. conseiller [386]. conseils [468]. consensus [4079]. conséquences [2260, 2347]. conservati [989]. conservation [4013, 3605]. Conservatoire [181, 388, 3987]. conservé [279, 1220]. conservée [879, 92]. conservées [571]. conservés [3715, 4434, 2318, 2801, 2006, 2380, 1211, 2135, 891]. consideration [3150]. consideratione [3316]. Considérations [2156, 1260, 516]. considéré [2245]. Consilia [3626]. Conspectus [3516, 4198]. conspiracy [4686]. conspirant [4223]. Constance [1977, 2078]. Constant [964]. Constant-Albert [964]. Constantin [4200, 2579]. Constantine [3651]. constants [3331]. Constantyn [3806]. Constitution [3874, 4096, 2973, 3354, 867, 3800, 4509]. Constitutions [487]. constitutive [4480]. constructeur [1346]. Constructing [4055]. construction [3410, 4280, 3142, 3854, 668, 2973]. Constructions [4287, 240, 3919]. constructivisme [3100]. construire [4054]. Consuegra [3388]. consummatio [2777]. contacts [1156]. contagions [1117]. containing [1863, 460]. Conté [1064]. contée [2648]. contemporain [585]. Contemporaine [1786, 587, 3043, 242, 376, 1536, 2213, 1430, 1431, 1865, 827, 376]. contemporaine^{≫} [902]. contemporains [2120, 1168, 1026]. contemporanea [3870]. Contemporary [4254, 2608, 3699, 1239, 784]. Contenau [1290]. Contet [1304]. Context [4609, 2190, 3786, 4061, 3434, 4317, 4190, 3810]. contexte [4609, 2651]. contextes [3199]. contextuelle [3512]. Contient [3408]. Continent [2772]. Continents [1615, 3277]. Contingence [4410]. Contingency [2633]. contingentia [4097, 4175]. continu [2868, 2195, 655, 3877]. continua [3327]. continuo [4070]. continuum [4070]. contrôle [387]. contraceptives [1764]. Contratación [1721]. contre [2170, 838, 4058, 3582, 4030, 2809, 4096]. Contrepoint [3403]. Contribuicão [529]. contributi [3339]. Contributie [959]. Contribution [279, 44, 51, 699, 4517, 2412, 2399, 964, 2368, 3147, 4523, 2135, 2196, 776, 2637, 758, 3749, 2463, 1821, 2681, 313, 692, 797, 881, 1245, 1302, 1303, 1308, 1310, 1311, 1480, 1702, 1704, 1706, 2032, 2064, 2066, 2067, 2646, 2708, 2845, 462]. Contributions [1265, 3439, 1109, 2811, 3700, 3408, 4252, 2760, 1593, 1271, 1354]. contributo [2629]. Control [3520, 4182]. Controriforma [4344, 4345]. controverse [228, 3983, 4567, 4562, 4583, 4580, 1738, 4196, 4551, 1047, 621]. Controverses [2813, 3694, 4310, 4433, 3882, 4352, 4566, 4670]. controversies [4563, 4670]. Controversy [1110, 4583, 4580, 3763, 3675, 4079]. Convegno [4088, 4076, 4435, 4070]. Convention [2050]. Converging [4653]. Conversation [1543]. Conversciences [3881]. convertir [1558]. convoyeurs [1884]. Conway [3368]. Cook [4578]. Cooke [3053, 3839, 4020, 1731]. Coolidge [118, 133]. coopération [3785, 4194]. Coordination [2476]. coordonnées [433]. Copans [2538]. Copenhagen [3755, 2934, 2932]. Copenhague [2934]. Copernic [2605, 582, 4643, 2231, 2260, 3258, 2134, 2136, 2561, 2140, 2135]. Copernican [3675, 4294, 938]. Copernicana [2604, 2605, 2770, 3011, 2334, 2501, 2525]. copernicanesimo [3727, 3923]. Copernicanism [3727]. copernicano [4078]. Copernico [2606, 1542]. Copernicu [2334]. Copernicus [2524, 2604, 2501, 2154]. copiste [1230, 1521, 1846]. Copping [512]. Copyflex [940]. coq [2734]. Coquand [2901]. Coquillages [4236, 1169]. Corbie [382]. Corboz [3950]. Cordemoy [1849]. cordes [3355, 3352]. cordis [4202]. Corfou [3700]. Corine [4050]. Cornelis [2658, 2964, 3278, 109, 1467, 1568, 1667, 1955, 2095, 2107, 1414, 1226, 3124, 1015, 1072, 1232, 626]. Cornell [3436, 3441]. Corner [512]. Cornu [4280]. Corollaires [3782]. coronary [4520]. Coronelli [1744, 989, 723]. Coronellienne [1159]. corporation [2268]. corporeæ [4688]. corporum [2656]. Corps [2901, 3964, 1282, 4033, 3875, 3881, 2055, 469]. Corpus [2631, 1750, 3607, 3758, 3224]. corpusculaire [3645, 2376]. corpuscularianism [4061]. Corrado [3583]. Corréa [216]. Corrections [3195]. corrélation [1811]. Correspondência [529]. Correspondance [4512, 34, 4615, 2662, 303, 3122, 3937, 2900, 1142, 2996, 2953, 2969, 1488, 4567, 926, 4383, 143, 2417, 2727, 3660, 3804, 2090, 519, 55, 2364, 4564, 1991, 2124, 557, 4566, 148, 891, 3858, 1062, 2073, 102, 2308, 2341, 2527, 2658, 2964, 3124, 3278, 4049, 2574, 109, 1015, 1072, 1232, 1467, 1568, 1667, 1955, 2095, 2107, 626, 742, 1056, 1233]. correspondances [4562]. correspondant [3806]. Correspondants [289, 10, 2064, 1826]. Correspondence [1522, 2877, 3418, 2452, 1936, 3195, 2171, 2493, 3663, 3060, 4563, 1671, 2881, 3455, 4290]. corrigé [790]. corrispondenza [4351]. corses [2188]. Corsi [3153, 3058, 3301]. corso [3383]. Cortambert [2257]. corte [565]. Cortes [245, 62, 87, 88, 136, 300]. Corvisier [3054]. Cosandey [73]. cose [3365]. cosmique [1598]. cosmogonie [4401]. cosmogonique [2205, 2285]. Cosmographicum [2869]. cosmographie [652, 3187]. cosmologia [4344, 3813, 3994]. cosmologie [3557, 1131, 2561, 1217, 2060, 2206]. cosmologique [3780, 2205, 2285]. cosmologiques [582, 4059]. Cosmologische [2460]. Cosmologists [3236]. Cosmology [3344, 4402, 4294, 2992]. cosmoplastic [4404]. cosmos [4082, 2927, 3514]. cosse [3039]. Costa [3055]. Costabel [2189, 2343, 2502, 3013, 3225, 1143, 1753, 2291, 3202]. Coste [3670]. costruzione [3238]. Cotes [2871]. Cotte [4140]. couchers [2627]. Couderc [43, 70]. couleur [2584, 2819, 2913, 1885]. couleurs [3711, 4141, 4008, 2680, 3510, 1483, 3427]. Coulomb [3694]. Couloubaritsis [2698]. Coulston [3939, 402, 1078, 1057, 3004, 2528]. Coumet [4253]. Count [2639, 3603]. Countries [456]. Coup [494, 286, 371]. coupe [4009]. Couples [2054]. cour [3781, 4081]. Courant [1977]. Courants [132]. courbe [3427]. courbes [3791]. Courcelle [4530]. Courcelles [963]. Courcier [530]. Cournot [2156, 2613, 3892, 470]. Courrier [1545]. Cours [650, 1013, 1094, 203, 4367, 5, 3495, 1821, 3168, 3623, 952, 1172, 2896, 4529]. court^{≫} [968, 945]. Coury [1757, 1938, 1937]. Couteau [4392, 3280]. Coutin [3831]. Couturat [4351, 2167, 4545]. Couty [4008]. Cowen [4230]. cowpox [2450]. Coyne [3727]. crânes [4208]. Cracovie [1461, 1658]. cracoviensia [3438]. Cracow [3011]. Craft [2550]. Cramer [557, 1062]. cranium [3363]. Crawford [2960]. creador [33]. create [4183]. creating [2979, 4514]. Creation [3696, 2622, 4637, 3006, 78, 4541, 1736, 1524, 3819, 2433, 4313, 3601, 2879, 4302]. Creation^{≫} [561]. créationnisme [2860]. Créativité [3252]. créatrice [1515]. creatures [3482, 3640]. Crédit [3948]. Crellin [1940]. Crépel [3698]. Créteil [4363]. crhesi [2812]. crible [4286]. Crisciani [3626]. cristallographe [3712]. cristallographie [34, 3713, 321]. cristaux [2315]. Cristiana [2174]. Critic [583]. critica [4511, 210, 1436, 2406, 189, 505, 1074]. Critical [3119, 3118, 3126, 3651, 3516, 4688, 3508, 2733, 4575, 3933, 3932, 982]. critici [3491]. criticism [3890]. critics [3236]. criticus [4198]. Critique [2443, 2612, 3563, 2220, 3689, 2164, 2868, 2982, 3336, 4473, 4688, 1396, 4002, 4575, 4014, 4286, 4448, 4470, 352, 1219, 553, 4525, 1201, 3607, 1675, 1294, 86, 2166, 44]. critiques [1361, 4064, 2407]. croise [4030]. croisée [4583, 3474]. Croisées [2822, 3606]. croissance [3956, 2014]. Croix [887, 1266]. Crombie [1913, 2115, 1246, 1432]. Crommelin [359]. Crommelinck [3943]. Crompton [89]. Crone [1299]. crónicas [3735]. Crook [3632]. Crookes [1472]. Cros [2680]. Crosby [789, 3864]. Crosland [1946, 3520, 1686, 2549, 2572, 4489]. Cross [3733, 4478]. Cross-staff [3733]. cross-staffs [4478]. Crossley [3206]. Crossroads [4583, 4507]. Crouvizier [4084]. Crowley [2848]. croyant [3844]. Crozon [3234]. Crucial [1389]. Crucibles. [2472]. Crustacés [1379]. Cruz [2378]. crystalli [2326]. Crystallography [3139]. Crystals [2294]. CS700 [4230]. CTHS [3831, 3837]. Cuadermos [3735]. Cuadernos [3994, 2644, 3735, 3732]. cubique [1738]. Cudworth [4404]. Cueillir [4291]. Cuénot [327]. Cues [4302]. cuestiones [4453]. Cugnetti [2812]. Cui [3148]. cuiuslibet [2777]. cuivre [748]. Cullen [2390]. Culpeper [1300]. Cultura [4309, 565, 3592, 4088, 4344, 4345]. Cultural [846, 821, 3485, 2345, 632, 973, 1123, 1316, 1445, 1507, 1914, 2477, 4318, 420]. culturales [2272]. Culture [3636, 2523, 1264, 3052, 2157, 3336, 4081, 4002, 4457, 4478, 3655, 4096, 4456, 3210, 3253, 75, 3480]. culturelles [4298, 884]. cultures [4670, 3838]. cum [3566]. cumulatif [2095]. Cunha [2012, 2414]. Cunningham [3338, 3481]. cura [2174, 830, 943]. Curas [3148]. Curbastro [877]. Curie [1472, 2571, 2415, 2463, 3091]. curieuses [825]. curiosité [3305]. Curious [4602]. Current [3464]. Curtis [3362]. Curtiu [1020]. Curtze [2106]. curve [189]. Cury [702]. Cushing [3410]. Cusset [791]. Cutro [4511]. Cuvier [3260, 2028, 2686, 1808, 4031, 2028, 1809, 1813, 2597, 1815, 1170, 1811, 254]. CVII [3149]. CXCV [4366]. CXCV/3 [4366]. cycle [2448]. Cycles [3796]. Cycloïde [1133]. cylindre [1927]. Cyrano [2503, 4060]. Cyril [211, 1732].

Cyrulnik [4395]. Cytochrome [1636]. Cytology [993]. Czartoryski [2525]. Czech [947]. Czechoslovak [1417].

D [3972, 3767, 1164, 2905, 3610, 2753, 4475, 58, 1545, 3304, 2189, 2343, 2052, 652, 1149, 1314, 1334, 1435, 1506, 1640, 1912, 1911, 2116, 3009, 1590, 2957, 3927, 3272, 327, 2058, 325, 514, 2568, 2111, 2576, 3834, 3105, 3362, 2550, 2961, 3148, 2066, 2410, 1997, 2974, 1669, 1670, 1929, 2057, 2096, 2919, 874, 2539, 3012, 2668, 4084, 4643, 1017, 1473, 2844, 3570, 4146, 3196, 2408, 2611, 3444, 4035, 491, 667, 856, 903, 1691, 2757, 2807, 608, 2492, 2742, 2521, 2737, 2923, 2551, 1151, 3354, 2857, 4120, 2074, 4573, 566, 4472, 1494, 1495, 1225]. D [800]. D-J [2974]. Dīn [2372, 2372]. Dîn [3239]. Dôme [1203]. d’Abbadie [4267]. d’Abbans [2902]. d’Abel [554]. d’Abidjan [2324]. d’acclimatation [4226]. d’action [1673, 215]. d’activité [3052]. dactylographie [1941]. d’Adam [558]. d’Adanson [1625]. Dadić [2184]. d’Adolphe [3934]. d’Adrien [2462]. d’Afrique [2851]. Dagen [1168, 4532]. d’agir [4357]. Dagnelie [1830]. Dagognet [2027, 2175, 3442, 3671, 4163]. d’agronomie [2272]. Dahan [4311, 3723, 4188, 2775]. Dain [59]. Dainian [3880]. Dainville [1075, 1483, 1920, 766]. d’air [4598]. d’Aix [273]. d’al [2454, 4602, 603, 248, 4614, 4560, 3454]. d’al-Khāzin [2454]. D’al-Khwarizmi [4560]. d’Albert [4134, 3030, 3429, 4638]. d’Alberuni [878]. d’alchimie [889]. d’Alcide [2860]. Dalechamp [3173]. D’Alembert [3704, 4615, 2662, 352, 1957, 4385, 2622, 3512, 1007, 309]. d’Alexandre [1373, 3162]. d’Alexandrie [3725, 2440, 2437]. d’Alexis [2219]. d’Alexis-Claude [2219]. d’Alfort [1155]. d’Alfred [4444, 3954]. d’Algebra [603]. d’algèbre [3099]. d’Alhazen [4331]. dalla [4309, 1291]. dalle [4515, 248]. Dallett [2479]. Dall’Europa [3339]. Dalmas [669]. Dalmedico [3723, 4188, 2775]. d’Alphonse [3821]. d’Alsace [1558]. Dalton [1980, 1754, 1563]. d’Alverny [1230]. Dame [1079]. d’aménagement [1583]. Dames [4196]. d’Ammonios [3495]. d’Ampère [3140, 2299, 2399, 3204, 2300, 293, 2347, 2302, 3690, 2398]. Dampier [445]. Dan [3411]. d’analyse [3722, 179]. d’anatomie [1050, 1155, 1319]. Danchin [2999]. d’Ancien [2681]. d’André [881, 2301, 3507]. d’André-Marie [2301, 3507]. D’Andrea [4068]. Dang [2845]. dangers [1598]. d’Angleterre [3596]. Daniel [3464, 31, 4615, 2609, 4434, 4438, 3935, 4392, 2999, 3756, 4556, 3442, 1564, 3382, 4387]. Danièle [4558, 3970]. Danielle [4312, 2754, 4147, 3884, 3837, 4035, 2669]. Daniil [4361]. Daninville [3013]. Danish [3150]. Danjon [3409, 187]. Danmark [876]. dans [3740, 85, 116, 147, 3289, 323, 2119, 2485, 2817, 1987, 3874, 3572, 4481, 3065, 4310, 4359, 3739, 3027, 3099, 2933, 3983, 3046, 4132, 2516, 2925, 3068, 3711, 3802, 4567, 4562, 1515, 3349, 3863, 4447, 2257, 2401, 10, 27, 57, 732, 115, 3687, 762, 2001, 1378, 4219, 3504, 3499, 3492, 4134, 4488, 4600, 4367, 2314, 496, 2934, 3645, 1630, 2497, 317, 2909, 4655, 1813, 3331, 320, 1349, 1736, 1892, 3294, 3153, 2622, 3188, 3045, 3328, 3746, 3145, 4128, 1698, 694, 4685, 4644, 342, 1850, 3622, 4136, 1884, 4609, 3354, 2005, 4287, 4153, 1815, 3855, 3919]. dans [4503, 4674, 3951, 3296, 2860, 3352, 3834, 3933, 4425, 4427, 1379, 2678, 3753, 2267, 3980, 2651, 3028, 4401, 2420, 321, 797, 884, 1173, 1708, 1869, 1939, 2234, 3228, 4048, 315, 4071, 117, 1663, 1683, 1790, 2919, 492, 4403, 3187, 3494, 3778, 1337, 4280, 4658, 3512, 55, 2693, 3795, 3924, 468, 1624, 677, 3158, 3756, 4158, 2916, 4454, 3291, 2205, 2285, 4676, 3529, 4114, 3286, 2844, 3376, 2074, 3070, 469, 3644, 4012, 2937, 2902, 3221, 4647, 6, 4613, 4060, 4062, 4363, 3601, 2288, 3295, 3292, 2494, 4635, 3738, 28, 3954, 234, 1437, 4133, 2102, 2304, 2443, 596]. dans [4166, 144, 1087, 3856, 2100, 4035, 4036, 4087, 4394, 4029, 4296, 380, 4221, 3894, 2983, 1252, 1916, 1656, 4129, 1730, 1433, 4566, 855, 1403, 1400, 2904, 1024, 1761, 2156, 2211, 2273, 4558, 4333, 3511, 3876, 2596, 655, 75, 547, 550, 1474, 719]. Dantes [1666]. d’Antoine [4156, 4058]. d’Apollonius [1330]. d’Application [1001]. d’aprés [4026, 4304, 926, 2207, 2318, 352, 1487, 161, 2048, 2329, 797, 1306, 1967, 519, 3224, 2272, 4036, 2073]. Darboux [3199, 1544]. d’Archeion [141]. d’Archimède [1292]. d’architecture [4009]. d’archives [4665]. d’Arcy [673]. Daremberg [1525, 4248]. Dareste [1225]. d’Ariane [4629]. d’Aristide [2270]. d’Aristote [2331, 3492, 1361, 1838, 2102, 3351, 2076, 2698]. d’arithmétisation [4287]. d’Arles [534]. Darling [695]. Darmois [586]. Darmon [3502]. d’Armurerie [249]. d’Arnaud [1538]. Darrel [3744]. Darrigol [3578]. D’Arsonval [63]. Darstellung [2479]. d’art [2253, 458]. d’articles [1027]. Darwin [3748, 2341, 3260, 456, 3060, 3455, 2477, 4241, 2633, 2809, 946, 878, 2723, 3377, 923, 2862, 922, 1811, 920, 2021, 921, 1025, 2029, 1578, 3382]. Darwinian [3319]. Darwinism [4406, 3632]. darwinisme [2251, 878, 3382]. d’Ascalon [1292]. Daston [3434, 3281]. d’astrologia [904]. d’astronomie [3241]. Data [86]. datant [1052]. datation [4138]. date [108, 1230, 1521, 1846, 227, 1425]. dater [619]. dates [1780]. d’Athènes [4096]. datis [2733]. d’attraction [2]. Dauben [2712, 2788]. Daudin [84]. d’Auguste [4450, 2269]. d’Augustin [3821, 2152, 1903]. d’Augustin-Pyramus [3821]. Daujat [157]. d’aujourd’hui [3594, 327, 651, 2738, 72, 1166]. Daumas [587, 20, 1889, 2756, 1392, 2863]. Dausset [2576]. d’Autheville [2665, 2665]. D’Autolycos [3665]. d’automne [935]. d’autorité [2936]. d’autrefois [486]. d’autres [3983, 3398, 4048]. d’Auvergne [1208]. Davaine [1909, 648]. David [4569, 2684, 3236, 3411, 3517, 1477, 2059, 3581, 3673, 2058, 4230, 3060, 3650, 2640, 3312, 2692, 799, 3126, 1636, 3743, 2959, 3243, 3364, 2024, 2158, 2216, 904, 3861, 4238, 823, 1851]. Davide [4309]. Davidson [487]. Daviel [3475]. Davies [2112]. Davis [3899, 1936, 1974]. d’Avogadro [3690]. Davy [692]. dawn [4172]. DD [3041]. ddr [3223]. Dear [3808]. Dearing [2310]. Death [2846]. Débarbat [3409, 4080]. débat [4446, 2858, 3780, 4578, 254, 3292, 2304, 4565, 1851, 3081]. Debate [3497, 3632, 3301, 4578, 3306]. Debates [1866, 4055]. Débats [3547, 4451, 4028, 3374, 4080, 3530]. Debeauvais [4291]. Deborah [4018].

Debru [3439, 3633, 3075, 4019, 2466, 3965, 3964]. Debus [2564, 1685]. début [3227, 386, 748, 25, 3272, 1456, 2218, 2413, 1000, 1756, 3345, 4452, 1964, 1930, 4573, 388]. débuts [1599, 2763, 4248, 3402, 2150, 10, 2441, 146, 3957, 3955, 3273, 2138, 1309, 3401, 4396, 1374, 2348, 3168, 1885]. Dec [2337, 2343]. decades [4656]. Décaillot [4512]. Decay [2112, 4079]. December [2834, 404]. décembre [279, 1659, 1602, 1467, 1667, 1826]. décennies [3555, 4656]. décès [2204]. Dechambre [484]. déchiffrement [1851]. decimum [490]. Decknamen [3642]. déclin [4589, 241]. decline [4589]. d’économie [2194]. déconstruction [4073]. d’Ecouen [3345]. Découvert [3784, 3091]. Découverte [2624, 554, 1911, 1728, 3717, 234, 292, 1817, 35, 4485, 4111, 4588, 3677, 1763, 3855, 1611, 2600, 1608, 2461, 1472, 2347, 1251, 2499, 3665, 1149, 1314, 1506, 1640, 491, 4392, 403]. Découvertes [3157, 3747, 399, 1456, 796, 324, 1472, 209, 283, 654, 2880, 764, 67]. découvreur [4157]. découvrir [2660]. Decree [3362]. décrites [435, 2267]. d’écrits [601, 1523]. d’écriture [4129]. dédales [2775]. Dedekind [3866, 2386, 3689, 2818, 1919, 2124, 2438, 3200]. d’Edition [2288]. d’Edo [3729]. d’Édouard [3403]. Dedron [1073]. d’Éducation [243, 2974]. déductibilité [3894]. déductif [4583]. deductive [4583]. déduite [858]. Dee [4137]. Deel [358, 326, 397]. deéllia [3062]. deep [2850]. deeper [3500]. defence [2963, 3120]. d’efficacité [3427]. Definition [3560, 4482, 3512]. Definitiones [2440, 2437]. définitions [1841]. degli [3901, 4351, 4076, 4301, 1441, 4072]. degré [1377, 2909, 42]. degrés [2895]. Dehergne [2131]. Deheuvels [2609]. d’Einstein [4142, 4347, 3844]. Dejiny [2394]. del [4409, 3901, 3592, 4088, 2395, 4078, 4076, 4435, 3058, 3994, 2077, 2173, 3150, 2991, 1441, 2787, 238, 3039, 3923, 3491, 1436, 2815, 3759, 4476, 2375, 4161, 154, 417, 4068, 4070, 4122, 566, 877, 1324, 2965, 2808, 3674, 248]. delà [4584]. Delachet [242, 1634]. Delacote [2611]. Delafond [780]. délais [2347]. Delambre [3212]. Delange [3213, 3057]. Delaporte [3166]. Delaunay [1242, 871, 363]. Deléage [3972]. délectables [986]. d’Électricité [731]. d’Elée [2868]. d’élément [2367]. Delga [4292]. Delhoume [401]. délia [4083]. d’Élien [59]. Delile [948]. délimitation [4445]. délire [4173]. Delisle [530, 1819, 2727, 2071, 530]. Delisle-Nolin [1819]. delivered [3386]. dell [2629, 3339, 1537, 531]. dell’800 [4161]. della [3901, 2344, 2297, 2395, 4344, 4345, 4351, 989, 1034, 1178, 1509, 1618, 3941, 4076, 4270, 4435, 4098, 4309, 2991, 2358, 1029, 2787, 210, 603, 2806, 901, 2232, 2406, 4476, 2375, 880, 4068, 4122, 1333, 189, 449, 505, 830, 877, 1074, 1469, 2874, 2814, 3674, 248, 2551]. dell’Algebra [602]. dell’anatomia [4083]. dell’Aquila [4076]. dell’aria [3941]. dell’Aritmetica [602]. dell’edizione [1099]. dell’Ematologia [994]. dell’embrione [3062]. dell’Istituto [2344]. dell’Oriente [880]. dell’Unione [830, 943]. dell’Università [2882]. dell’universitàdi [3674]. dell’VIII [4076]. Delom [1523]. delta [3332]. Deluc [3348]. démarches [2283]. d’embryogénie [3819]. d’émeraude [3648]. Demeulenaere [3828]. Demeure [527]. d’Émile [4476, 4472]. Démocrite [1598]. démographie [1836]. Demokrit [3488]. DeMolen [2804]. Demolon [95]. Demon [2918]. Demonferrand [4252]. demonic [874]. Démonstration [975, 4583, 3755, 3710, 4583, 4368, 3352]. démonstration-spectacle [3710]. Démonstrations [4012, 493, 3790]. démontré [3355]. d’Encouragement [388, 4612]. d’Endocrinologie [3252]. d’énergie [4279, 1728]. Deniel [1061]. Denis [1078, 3515, 3933, 3932, 4643, 4412, 3104, 588]. Denise [2677]. Denkansatzes [4234]. Denken [3416]. Denkens [4093, 3346]. denklemlerde [1388]. Dennis [4143, 3814]. Denomy [1997]. Denoville [4475]. d’enseignement [1938, 2974]. d’enseignements [3958]. d’ensemble [2017, 2003]. dentaire [1168]. d’entrepreneurs [3558]. d’entreprises [3558]. Denver [1659]. deny [4587]. dépôt [1571, 1048]. d’épaisseur [1249]. Deparis [4091]. départ [3093]. départements [1031]. Depau [10]. d’Éphèse [4147]. d’épistémologie [4179]. déplacements [2600]. déployé [3823]. déposé [2967]. dépression [4199]. depuis [3716, 3896, 1640, 1032, 1781, 1271, 1354, 3224, 1695]. d’équations [472, 2420]. d’équivalence [4554]. deras [747]. d’Erasme [2326]. Dercourt [4411]. Déré [4261]. Derek [906, 1572]. dérivées [3331, 2622]. dermatologie [1174]. d’Ernest [2754]. dernier [737, 3360, 1738, 2152, 1233]. dernières [432, 1708, 1079]. d’Ernst [4666]. Derrière [2571]. désaccords [4567]. desagüe [2815]. Desanti [3669]. Desargues [355, 1081, 292, 1157]. Desault [1967, 1983]. Desbats [68]. Descartes [3225, 2870, 4123, 4024, 2502, 3925, 4560, 584, 1337, 4532, 234, 742, 1056, 1233, 787, 199, 228, 3801, 1738, 4014, 4222, 4217, 4636, 4305, 3804, 1696, 3805, 365, 2950, 4635, 234, 3803, 230, 3800, 3798, 229, 3806, 227, 2631, 4143, 1218, 1219, 1325, 1569, 131, 3814, 392]. Descotes [4492]. Description [314, 4548]. descriptions [4198]. Descriptive [235, 4009, 1146, 1390, 359]. descubrimiento [3592, 88]. Déséquilibres [4624]. Desgranges [241, 486]. Desiderio [566, 1324]. design [1771]. Desmarest [3476]. Desmond [3634, 3003]. d’espace [568]. Despaigne [3244]. d’espèce [3400, 3153]. d’espèces [435]. despotisme [774]. dessalement [1732]. dessein [3686]. dessen [2479]. Dessins [1391, 1165, 1146]. dessus [1472]. Destaillats [1980]. Destin [1477, 46]. destinataire [227]. destiné [2595]. Destombes [279, 161]. Desurvire [3213]. détaille [4573]. details [4573]. d’État [2298]. Determinantentheorie [2661]. Déterminants [3998]. détermination [556, 3929]. déterminations [813]. Déterne [60]. d’Étienne [1831, 2001, 2004, 2005, 2000, 3819, 2006, 2284, 4677]. Detlef [3444]. détournement [2583]. detta [4076, 4435]. Dettonville [1187, 1186]. d’Étude [1341, 3460, 4530, 4265, 572, 928]. d’Etudes [1941, 2839, 886, 249, 1735, 766, 3227, 3555, 1808, 4441, 883, 4486, 4510, 2407]. d’Euclide [3225, 2312, 86, 3267, 975, 4128, 3296, 3493, 644, 4130, 4137, 4126, 3496, 4127, 151, 1329, 1629, 3499]. d’Euler [925, 3045, 1656, 2208]. d’Eutocius [2093, 1292]. Deutsch [3843, 3316, 1332]. deutsch-russische [1332]. Deutsche [4361, 769, 1323]. Deutschen [942, 2566, 1332]. Deutscher [3843]. Deutschland [1470]. Deux [3099, 2982, 926, 143, 3676, 778, 2625, 1085, 1883, 777, 2908, 3740, 4423, 4424, 813, 222, 159, 4383, 3267, 3071, 4368, 4426, 2910, 3554, 4609, 4425, 4427, 3867, 4428, 3621, 2188, 3717, 1781, 1954, 4283, 3094, 3224, 510, 3790, 409, 2135, 4604, 774, 889]. Deuxième [1563, 2969, 4222, 2110, 3376]. Devarieux [4176]. d’Évariste [2531, 375, 1657, 24, 1903]. Devaux [39, 4658]. developed [2452]. Development [3533, 729, 1591, 1476, 3125, 789, 1297, 2025, 2703, 3866, 1576, 3220, 3733, 3508, 2876, 938, 4440, 4656, 4670, 1976, 3388, 3841]. developmental [2294]. Développement [1698, 1610, 3767, 4542, 2605, 4468, 2622, 3951, 3960, 1340, 2367, 4490, 1821, 2724, 1350, 3954, 3953, 3623, 4670, 4346]. développements [4482, 1398, 693]. développer [834]. devenir [4288]. devenu [1709]. Devèze [1703]. déviation [1024]. device [4588, 4636]. devices [2570]. Devil [3634]. Deville [4416, 146]. Devillers [3149]. devint [4382]. dévoilé [3309]. Devoir [2634]. d’évolution [1312]. devoted [162, 2345, 632, 846, 973, 1123, 1316, 1445, 1486, 1507, 1914, 306, 96, 250, 220]. devriez [3724]. Dewhurst [1499, 2886, 2963]. d’exception [3922]. d’explication [2584]. d’exposition [2902]. d’extension [4445]. d’Extrême [203, 879]. d’Extrême-Orient [203, 879]. d’Harvey [1676]. d’Henri [1633, 342, 518, 4192, 4658]. d’Hercule [3064]. d’Héria [3197]. d’Hermès [8]. d’Hermies [652]. D’hier [766, 2026, 1166]. d’Hippocrate [3313, 130, 961]. d’hist [3700, 3703, 3664, 3530]. d’Histoire [202, 349, 415, 369, 475, 650, 896, 1013, 931, 980, 1094, 1055, 1053, 1257, 1285, 1427, 1429, 1462, 1641, 1561, 1744, 1767, 1770, 1823, 1860, 1825, 1862, 3170, 3186, 2374, 2870, 1, 1255, 182, 2050, 5, 1808, 2189, 2193, 2343, 2396, 446, 616, 617, 633, 696, 801, 843, 974, 1033, 1035, 1121, 1150, 1315, 1508, 1796, 2038, 2043, 2132, 625, 740, 895, 897, 1091, 1463, 3433, 1269, 257, 499, 1461, 2637, 3097, 2318, 2319, 1699, 2670, 2044, 2491, 2582, 1602, 2603, 3043, 151, 3651, 1338, 2380, 2705, 217, 1075, 1827, 2405, 1206]. d’Histoire [1045, 2771, 2386, 2669, 353, 160, 141, 192, 307, 383, 488, 247, 1005, 1052, 971, 3631, 2308, 3715, 3260, 3133, 3359, 3349, 2252, 2191, 2296, 1444, 1619, 3348, 3385, 3963, 4193, 3460, 3935, 4065, 3521, 3569, 3585, 3749, 3314, 3413, 838, 1115, 2548, 2453, 1387, 3653, 4337, 3525, 3725, 3970, 2598, 4592, 2693, 3924, 3157, 3490, 1897, 4613, 3601, 3741, 1536, 3823, 752, 3453, 3623, 3633, 4237, 4288, 184, 4509, 458]. d’historien [4306]. d’historiographie [3924]. d’holonomie [4429]. Dhombres [3900, 3832, 3900, 3207, 3832, 3233, 4034, 3167]. d’homéose [3953]. d’homo [3407]. D’Hondt [3664]. d’Humanisme [1538, 2373]. d’hybridation [1089]. d’hypothèses [919]. diable [4162]. Diagne [3310]. diagnosis [4656]. diagnostic [4656]. diagnostics [419]. diagonale [4133]. diagramma [4239]. dialectique [4368, 1984, 4398, 2404]. Dialog [1666]. Dialogo [4078]. Dialogue [4233, 4368, 1131, 3006, 3779, 1209, 3606, 1665]. dialoguent [4526]. Dialogues [1606]. Diara [3153]. Diaries [1936]. Diary [3455]. Diaz [3340]. d’Ibn [1675]. Dibner [4197]. Diccionario [2926]. Dichtung [2740]. Dickinson [89]. dictionary [3584]. Dictionnaire [2754, 345, 3936, 3216, 3884, 3435, 343, 2844, 652, 1353, 3456, 984, 1344, 2744]. dictionnaires [615, 3293]. Didact [4195]. Didactic [2474]. didactique [2781]. didattica [3829]. Diderot [3115, 1078, 4174, 352, 3408, 3407, 347, 2896, 3291, 310, 341, 351, 352, 4288]. Didier [3648]. Dieckmann [351]. Diego [1933]. Diepgen [1518, 954]. Diesel [1730]. diet [4236]. Dieter [2885, 2535, 2671, 3816]. dietetica [3039]. Dietlinde [2473]. Dietrich [3035, 3668]. Dietz [3675]. Dieu [1175, 1938, 380, 3572, 4222, 1115, 1967, 4114]. Dieudonné [1294, 2507, 2386, 2311]. Diez [1171]. Difference [3480, 3028]. différences [4354, 4020]. different [4598]. différentes [4598, 3427]. differential [4020, 2983]. différentielle [2983]. différentielles [2622, 1963, 1656, 2596]. Diffident [3794]. difformity [1847]. diffraction [3257]. diffusion [3710, 4046, 3359, 3288, 3290, 4065, 3342, 1662, 4096, 4062, 2748, 2135, 848, 3013, 3159]. Diffusione [3923]. d’Ignace [2240]. d’Ignace-Gaston [2240]. Digne [786]. Digression [704]. Dijksterhuis [937, 1496, 628, 244]. Dijon [3633, 4446]. dilemme [2365]. diluvianistes [2365]. dimensions [3824]. D’Imhotep [962]. Din [1112, 959, 975]. d’incomplétude [2937]. d’indivisibles [3354]. d’induction [1411]. d’industrie [3949]. d’industries [3948]. d’inertie [252]. Dinge [4095]. Dingle [1475]. d’insectologie [4027]. d’institutionnalisation [4484]. d’instruments [22, 255, 2178, 1346]. d’intégrité [1683]. d’intensité [3738]. d’interprétation [3133]. d’introduction [4155]. d’invention [967]. Diocles [2419]. diode [623]. Diophante [1527, 2098, 2149, 2222, 1295]. diophantienne [2454, 179]. Dioptrique [4051, 584]. Diouf [3922]. dipartimento [4076, 4068]. Diplôme [1941]. diplomacy [4610]. dir [3591]. Dirac [2935, 3271, 3369]. Directeur [1048]. Direction [3052, 2502]. directionem [2631, 4635]. Directoire [1064, 3044]. directory [2078]. Dirichlet [2829]. dirigada [529]. dirigée [2919]. Dirk [207]. d’Isaac [3046, 3143, 1404]. disant [197]. discepoli [2490]. discernement [3448]. disciple [1009, 3634]. disciples [3163]. disciplinaire [4509]. disciplinarité [2151]. Discipline [3441, 2151, 4589, 2550, 3950, 3808]. disciplines [4313, 3228, 4048]. discontinu [2195]. Discorsi [4359, 3352]. Discours [2168, 2227, 2314, 2507, 2217, 1600, 2397, 4356, 584, 4120, 1954, 1675, 3790]. Discourse [1103]. discovered [3091]. discoverer [3610]. Discoveries [2390, 3747]. Discovery [328, 1472, 3899, 4588, 1432, 3080, 1236, 900, 3904]. Discussion [1810, 3566, 1812, 1814, 1816, 114, 106]. Discussions [53]. disdiaclastici [2326]. Disease [3534, 4520, 2143, 2846, 3526, 2961, 3166]. Diseases [2578, 361, 361]. Disertage [960]. d’Islam [3147, 4164]. d’Isolda [3824]. disordered [4561]. Displacement [3464]. dispositif [4636]. dispositifs [2570]. Disproportion [4219]. disputation [3815]. disputes [3574, 4578]. Dissertatio [3566]. Dissertation [4599, 2771, 4074]. Dissertationsdrucke [2849]. dissolvant [3646]. Distillateurs [3309]. distinction [4337, 3690]. distinctions [4225]. distributions [2928]. Disziplingeschichte [2823]. Div [1485]. divers [460]. diverses [3138, 224]. diversis [4394]. divinatoire [2693]. divine [4014]. Division [931, 1053, 1561, 1317, 1091, 2336, 1206, 2411, 1033, 1444, 2054]. d’Ivoire [2324]. divulgazione [4262]. dix [4369, 4578, 2102, 4455]. dix-huitièmes [4369, 4455]. dix-neuvième [4578]. dix-septième [2102]. Dixhuitièmes [4619]. Dixon [2747]. Djebbar [4164]. DM [605]. DMF [3884]. d’Oberkampf [38]. Doblin [555]. d’obres [904]. d’observations [3596]. d’observer [3779, 4466]. docta [3346]. docteur [1775, 1995]. d’octobre [2161]. doctorat [157, 2448, 297, 462]. doctorats [3228]. Doctors [2110, 2129, 1504]. doctrine [3492, 2487, 379, 524, 2141, 2291]. Doctrines [2656, 2396, 2390, 2868, 582, 4510, 1930]. d’oculistique [2800]. document [1738, 1883, 1381]. documentaire [3785, 4665]. documentaires [2407]. documentary [3826, 4665]. Documentation [615, 4609, 738, 4609]. documented [420]. documenti [3150, 1469]. Documents [1247, 194, 4017, 2576, 2321, 756, 3500, 1410, 2758, 2867, 299, 1081, 2139, 3348, 4530, 2528, 480]. Doeblin [4228, 555]. d’oeil [371, 494, 286]. Doeming [1936]. Doerflinger [2392]. Doetsch [1109]. Doeuff [2925]. d’œuvre [181]. dogmatisme [4246]. Dohrn [3059]. doit [4218, 4658]. Dokouchaev [2763]. Dokumenten [1680]. d’Olivier [2494]. Dollfus [4392]. Dollo [3583]. Domain [747]. domaine [2516, 2401, 2314, 884, 1683, 1656]. Domandl [2479, 2479]. Domenico [4067]. domestication [1838]. Domingo [4271]. Domingue [3753, 3527]. Dominique [4617, 4352, 3558, 4311, 4094, 3408, 1707, 2915, 4492, 4241, 4556, 2851, 2771, 3995, 4089, 1600]. Dominis [2245, 2180]. Dompter [4646]. Donald [2570, 3439, 674]. Dondi [1893]. Donna [2294]. données [3171]. Donnini [3105]. Donnio [694]. Donovan [2390]. Doolittle [4178]. Doppler [4075]. d’Optique [604, 1423, 4051, 4124]. Dora [1853]. d’Orbigny [2860]. Dordain [1306, 1307]. d’organisation [3952]. Dorinda [2686]. Dorlin [4436]. Dorn [3219]. d’Orose [2827]. Dorothea [508, 3536]. Dorothée [2902]. Dorothy [1908, 955]. Dörries [3694]. d’Ørsted [1086]. Dortoman [4625]. Dortous [4596, 4598, 4600, 4601, 4597]. dossier [4643]. Dossiers [629]. Dostrovsky [2831]. Dott [901]. Douarin [4411]. double [1585]. doubles [3280]. douces [535]. Douglas [452, 453, 609]. d’Ouvrages [1267, 1593, 7, 2306, 2663, 3319, 3486, 3563, 3950, 3972, 2503, 1898, 4576, 4585, 4594, 4605, 4659, 4667, 4672, 4681, 4690, 4409, 4453, 4458, 2458, 3225, 384, 671, 726, 749, 3464, 3694, 4189, 89, 165, 241, 2827, 3010, 2421, 3631, 3519, 4557, 3635, 3767, 3942, 3270, 3663, 4142, 4264, 3901, 4168, 3591, 4526, 3939, 4310, 4433, 4450, 4477, 4508, 4511, 4515, 4546, 4569, 3752, 3747, 3836, 743, 1144, 1164, 939, 1774, 213, 689, 1538, 1539, 1540, 3516, 2374, 2490, 2524, 2547, 2564, 2976, 3458, 3562, 3459, 3632, 3634, 3748, 587, 2901, 2905, 3900, 3697]. d’ouvrages [3866, 4050, 4512, 4513, 3578, 3485, 4092, 4390, 203, 279, 402, 1749, 1785, 1849, 1972, 1973, 2069, 2166, 2289, 2290, 2308, 2309, 2320, 2331, 2332, 2341, 2526, 2527, 2604, 2605, 2606, 2628, 2630, 2631, 2632, 2657, 2658, 2684, 2865, 2868, 2870, 2964, 3124, 3278, 3604, 3654, 967, 3833, 3810, 3815, 4097, 4175, 4240, 692, 695, 746, 748, 793, 794, 795, 876, 885, 3903, 3904, 4123, 4143, 3440, 3898, 3944, 3610, 3003, 3260, 3506, 2504, 2505, 2529, 2530, 2753, 2772, 2831, 2871, 2873, 2925, 2942, 3223, 3336, 3489]. d’ouvrages [3969, 2337, 2338, 2834, 478, 487, 3579, 3533, 4254, 4410, 1752, 3359, 3896, 1633, 2409, 2478, 2536, 2812, 2885, 4365, 1779, 4200, 3340, 3349, 3863, 135, 158, 3343, 3880, 3721, 3700, 3793, 3826, 4144, 4179, 4233, 4242, 4263, 4266, 4314, 4475, 3785, 4312, 2252, 2262, 2509, 2775, 2783, 4165, 1522, 2877, 2878, 828, 39, 40, 41, 43, 44, 45, 46, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 58, 59, 60, 63, 64, 65, 66, 67, 70, 71]. d’ouvrages [72, 90, 91, 93, 94, 95, 137, 139, 159, 212, 1102, 4079, 3483, 157, 1345, 3934, 1016, 1020, 1021, 1022, 1025, 1030, 1545, 3119, 2629, 1974, 4004, 4173, 4232, 3592, 2212, 2228, 2467, 2544, 2627, 2656, 2683, 2803, 2826, 3236, 3303, 3304, 3504, 3811, 3882, 4003, 4088, 2190, 2191, 2192, 2275, 2276, 2295, 2342, 2344, 2395, 2396, 2411, 2422, 2052, 2061, 2104, 2310, 2532, 4488, 4201, 419, 446, 485, 4077, 4367, 2913, 4297, 3501, 2167, 2168, 2169, 2183, 2227, 2248, 2253, 2314, 2754, 611, 3943, 3178, 3207, 3971, 4243, 2750]. d’ouvrages [3384, 3411, 3517, 3701, 4202, 3522, 3566, 3703, 3727, 4078, 4125, 4170, 4258, 4271, 4344, 4345, 4351, 4352, 4368, 4527, 4081, 1977, 2078, 2215, 2384, 2507, 2531, 2609, 2610, 2664, 3053, 535, 507, 511, 515, 605, 615, 652, 693, 729, 730, 753, 906, 909, 949, 968, 970, 989, 1078, 1120, 1149, 1150, 1177, 1313, 1314, 1334, 1434, 1435, 1477, 1482, 1506, 1525, 1630, 1631, 1639, 1640, 1665, 1690, 1709, 1829, 1835, 1836, 1868, 1975, 1979, 1980, 2028, 2037, 2041, 2114, 2115, 2117, 2131, 3487, 2451, 2512, 3245, 2059, 3360, 69, 444, 448]. d’ouvrages [2263, 3558, 3948, 119, 156, 398, 1214, 3103, 4411, 4434, 1389, 3902, 3992, 4311, 459, 74, 354, 358, 359, 360, 403, 956, 958, 959, 960, 961, 1576, 1577, 1863, 3862, 4026, 2776, 3009, 3433, 2378, 3040, 3220, 3238, 3271, 3315, 3341, 3361, 3369, 3371, 3372, 3410, 3436, 3441, 3581, 3693, 3696, 3792, 3797, 3899, 3991, 4507, 138, 140, 218, 281, 282, 304, 400, 401, 450, 451, 452, 453, 586, 728, 767, 768, 864, 1258, 1259, 1269, 1347, 1349, 1353, 1415, 1416, 1417, 1420, 1422, 1590, 1591, 1727, 1728, 1729, 1736, 2711, 3121, 3598, 2261, 2388, 2389, 2546]. d’ouvrages [2569, 3000, 3005, 3083, 3169, 3212, 3262, 3282, 3673, 3940, 2957, 3311, 3461, 3528, 3594, 3927, 3945, 3649, 3153, 3235, 3272, 612, 3809, 4360, 73, 186, 191, 204, 211, 214, 235, 275, 280, 301, 327, 328, 330, 371, 391, 399, 418, 443, 455, 456, 461, 465, 508, 581, 630, 651, 704, 940, 1032, 1096, 1191, 1208, 1523, 1524, 2423, 2774, 3152, 2058, 2928, 3520, 3628, 3669, 3671, 3676, 3677, 3746, 3761, 3791, 4261, 4393, 3990, 3075, 3145, 2661, 2662, 2779, 2924, 3042, 237, 302, 303, 325, 326, 373, 377, 396, 397]. d’ouvrages [506, 513, 514, 1754, 1866, 1867, 4196, 694, 3348, 3832, 3514, 3807, 3385, 4320, 3567, 1019, 2470, 2568, 2608, 2636, 2637, 2830, 509, 629, 672, 948, 957, 962, 1113, 1850, 2842, 2914, 3766, 3949, 3963, 4033, 4230, 4369, 4436, 4437, 4438, 4093, 3515, 3922, 4022, 4236, 3123, 3557, 4260, 4315, 2770, 3011, 3302, 3370, 3536, 3763, 3787, 4053, 4141, 4298, 2445, 3122, 3233, 4024, 4056, 4174, 3755, 3014, 3097, 3118, 4286, 4491, 2667, 3733, 3786, 3936, 3937, 3941, 4005, 4008, 4018, 4076, 4195, 4255, 4257, 4259, 4267, 4270, 4289, 4415, 4416, 4435, 4457, 4478, 4514, 2109, 3824, 4098, 3414]. d’ouvrages [3672, 3457, 536, 1831, 2027, 2110, 2111, 2128, 2143, 2318, 2319, 2573, 2575, 2576, 2647, 2805, 3057, 3058, 3059, 3060, 3061, 3062, 3455, 3460, 3885, 3886, 4287, 2063, 2294, 3652, 3881, 4343, 4413, 4019, 351, 352, 357, 3650, 2674, 3309, 3731, 2506, 3273, 2565, 2638, 4234, 3590, 4321, 4021, 3870, 3869, 3301, 3337, 3730, 3751, 3834, 4032, 4205, 2502, 3839, 4020, 4049, 4203, 4516, 305, 1348, 3935, 4309, 1610, 3841, 2312, 2313, 4094, 3521, 997, 602, 3154, 3155, 3276, 3277, 3569, 3585, 3749, 3947, 4090, 4091, 4441, 3666, 2639, 2640, 2642, 2837, 2838, 2900]. d’ouvrages [2988, 3208, 3242, 3279, 3342, 3386, 2709, 2920, 3054, 3105, 3313, 3314, 3362, 3408, 3417, 3431, 3462, 3480, 3482, 3488, 3503, 3505, 3526, 3531, 3564, 3584, 3589, 3611, 3637, 3640, 3744, 3867, 2741, 3994, 4082, 4171, 2780, 3925, 4095, 1773, 1978, 3413, 983, 999, 1162, 1432, 1442, 1476, 1499, 1500, 1501, 1502, 1503, 1504, 1505, 1526, 1614, 1635, 1662, 1688, 1694, 1788, 1833, 2031, 2045, 2077, 2097, 2129, 2130, 2142, 2172, 2173, 2217, 2218, 2321, 2447, 2450, 2550, 2574, 2577, 2578, 2581, 2591, 2660, 2670, 2840, 2841, 2843, 2846, 2847, 2884, 2886, 2903, 2917]. d’ouvrages [2961, 2963, 2975, 2989, 3125, 3148, 3150, 3209, 3211, 3240, 3274, 3312, 3317, 3338, 3363, 3416, 3418, 3456, 3532, 3534, 3593, 3630, 3636, 3753, 3757, 350, 1780, 2029, 2030, 2036, 2044, 2266, 2267, 2316, 2325, 2339, 2643, 2883, 4313, 2251, 2876, 2522, 3670, 3788, 3989, 3993, 4451, 3079, 3214, 2991, 1169, 2491, 2174, 2233, 2336, 2358, 2452, 2472, 2473, 2474, 2475, 2479, 2533, 2582, 2641, 2692, 2695, 2706, 2707, 3675, 3639, 1611, 1143, 161, 2420, 690, 747, 861, 862, 3381, 2214, 2477, 2534, 606, 666, 773, 607, 665, 673, 675, 691, 727, 744, 750, 751, 772]. d’ouvrages [796, 797, 798, 799, 800, 818, 819, 838, 879, 881, 882, 883, 884, 963, 964, 965, 995, 996, 998, 1000, 1001, 1002, 1003, 1029, 1059, 1061, 1114, 1115, 1116, 1148, 1168, 1170, 1173, 1174, 1175, 1245, 1302, 1303, 1305, 1306, 1307, 1308, 1309, 1310, 1311, 1312, 1441, 1443, 1480, 1481, 1565, 1569, 1571, 1579, 1613, 1638, 1689, 1692, 1701, 1702, 1703, 1704, 1705, 1706, 1707, 1708, 1755, 1756, 1757, 1834, 1848, 1869, 1983, 2032, 2033, 2034, 2035, 2064, 2065, 2066, 2067, 2068, 2186, 2188, 2234, 2268, 2269, 2270, 2322, 2323, 2324, 2340, 2393, 2410, 2448, 2535, 2579, 2644, 2645, 2646, 2666]. d’ouvrages [2671, 2685, 2708, 2785, 2787, 2845, 1847, 1997, 2094, 2354, 2603, 3438, 2672, 2853, 2915, 2974, 3043, 3216, 210, 236, 324, 466, 2510, 2548, 1685, 1771, 1772, 2453, 2469, 2828, 4071, 42, 109, 61, 62, 68, 87, 88, 92, 117, 131, 132, 134, 136, 150, 151, 152, 238, 239, 242, 245, 300, 355, 374, 395, 445, 447, 481, 482, 567, 603, 628, 653, 670, 770, 789, 792, 984, 985, 986, 987, 988, 1015, 1072, 1097, 1098, 1163, 1215, 1218, 1219, 1232, 1296, 1297, 1327, 1328, 1330, 1386, 1387, 1388, 1467, 1544, 1568, 1573, 1608, 1609, 1628, 1634, 1663]. d’ouvrages [1666, 1667, 1668, 1669, 1670, 1671, 1672, 1673, 1681, 1682, 1683, 1750, 1753, 1776, 1777, 1778, 1789, 1790, 1791, 1792, 1793, 1794, 1795, 1844, 2025, 2054, 2055, 2056, 2057, 2093, 2095, 2096, 2107, 2919, 3039, 3126, 3586, 3587, 614, 991, 1263, 2987, 3008, 3218, 525, 532, 1468, 2806, 3729, 2230, 1982, 3648, 3651, 3653, 3655, 4316, 2773, 3745, 4398, 3523, 3525, 3561, 3704, 3725, 4432, 4492, 4494, 3970, 3316, 4121, 3339, 4474, 3868, 3965, 2372, 2570, 3638, 874, 942, 947, 954, 3706, 4207, 4439, 4531, 4096, 4206, 3383, 582, 583, 584, 591, 601, 604, 626, 2733, 2777, 2921]. d’ouvrages [3608, 1440, 3664, 4362, 565, 2962, 3259, 3789, 3840, 379, 1337, 1338, 1853, 2539, 2686, 3213, 4347, 4392, 4396, 4414, 4456, 4528, 3559, 3560, 2076, 2999, 2922, 2508, 3149, 3389, 3966, 4009, 4034, 4178, 4269, 4292, 3814, 4120, 4532, 3580, 4176, 2693, 2696, 3754, 3795, 3796, 3842, 3864, 3924, 585, 1664, 3241, 1057, 1616, 1636, 2250, 2511, 3055, 1236, 1237, 2140, 2782, 3012, 3197, 3280, 3347, 3409, 3420, 4080, 3756, 3723, 3794, 1578, 3605, 3835, 4295, 4349, 3582, 2287, 2668, 2705, 2839, 3629, 3762, 886, 1331, 4199, 4084, 3860, 3535, 4010, 4025, 2182, 2293, 2538, 2916]. d’ouvrages [2352, 4231, 3299, 1786, 1787, 3106, 3206, 4139, 3743, 2833, 2835, 2918, 2958, 2959, 2960, 2972, 2977, 2979, 3002, 3007, 3016, 3035, 3037, 3074, 3077, 3156, 3157, 3210, 3222, 3243, 3257, 3275, 4241, 4366, 4454, 4490, 4529, 249, 4319, 2247, 3828, 1471, 3439, 3518, 2231, 2260, 2992, 3344, 3500, 3699, 3702, 3844, 3080, 86, 524, 1017, 1216, 1288, 4417, 3454, 3490, 4412, 3529, 1472, 1473, 3484, 3412, 3627, 3481, 3524, 3606, 3724, 950, 992, 1109, 1117, 785, 786, 857, 859, 865, 900, 907, 937, 981, 982, 990, 1103, 1104, 1142]. d’ouvrages [1217, 1234, 1235, 1239, 1240, 2822, 2823, 4066, 2813, 2844, 2851, 3166, 3364, 3597, 3764, 4167, 1537, 1845, 1976, 2225, 2226, 2271, 2523, 2537, 2810, 2811, 4177, 2244, 3432, 4188, 826, 832, 3565, 3570, 3607, 3612, 3758, 4290, 4397, 4486, 4487, 4510, 3626, 3365, 3742, 4530, 2333, 205, 206, 215, 216, 217, 4294, 3583, 4239, 2371, 3938, 3281, 3434, 3437, 3695, 3698, 4204, 4229, 4052, 4055, 4350, 3728, 901, 945, 1268, 1271, 1344, 1346, 1350, 1351, 1352, 1354, 1356, 1412, 1413, 1414, 1418, 1419, 1592, 1597, 1732, 1733, 1734, 1735, 1737, 2848, 2849, 2902, 3004, 3052]. d’ouvrages [3076, 3096, 3194, 3221, 3366, 1751, 2778, 3443, 3736, 3879, 3883, 3921, 3962, 3964, 4007, 4146, 4147, 4285, 3568, 3813, 3923, 4072, 4363, 4346, 3946, 3165, 3234, 3318, 3705, 3527, 3491, 3735, 2024, 2060, 2106, 2126, 2154, 2157, 2158, 2216, 2232, 2288, 2334, 2419, 2501, 2528, 3195, 3577, 3224, 2385, 3147, 3829, 4348, 277, 2941, 2993, 1244, 745, 769, 887, 904, 905, 938, 1018, 1075, 1100, 1101, 1230, 1231, 1262, 1436, 1470, 1521, 1604, 1693, 1758, 1759, 1827, 1846, 1854, 2105, 2356, 2405, 2406, 2525, 3082, 3196, 3215, 3345, 3595, 3603, 3734, 3812, 3884, 3996, 4054]. d’ouvrages [4291, 3816, 4452, 155, 3307, 1695, 1981, 2113, 2272, 2494, 2514, 2583, 2648, 2758, 2815, 2850, 2926, 3104, 3177, 3871, 4172, 3831, 3837, 4262, 2702, 2746, 2749, 3861, 4140, 4210, 1605, 1661, 2229, 2545, 4083, 2377, 3843, 3300, 3435, 3463, 3107, 3668, 3741, 3759, 3308, 3407, 4208, 875, 185, 209, 234, 274, 376, 392, 393, 504, 742, 1056, 1058, 1233, 1437, 1478, 1536, 1606, 1607, 1775, 863, 941, 1289, 1574, 2175, 2633, 2786, 2809, 4006, 4391, 4476, 3732, 2213, 2127, 2292, 2353, 2370, 2376, 2404, 2408, 2443, 2444, 2571, 2611, 2673, 2675, 2698, 2700, 2734]. d’ouvrages [2735, 2736, 2738, 2744, 2745, 2748, 2751, 2752, 2756, 2824, 2825, 2832, 2879, 3823, 3825, 479, 483, 356, 484, 510, 674, 946, 3261, 3419, 4166, 3146, 731, 752, 754, 902, 944, 1392, 1483, 2375, 2990, 3056, 3098, 3167, 3168, 3306, 3367, 3387, 3453, 1575, 1686, 1687, 1998, 2026, 2062, 2108, 2141, 2171, 2185, 2264, 2315, 2476, 2549, 2572, 3442, 3444, 4035, 4036, 4493, 4394, 4440, 4318, 3667, 4161, 1479, 2387, 4163, 4235, 4296, 4317, 4361, 4455, 3827, 880, 380, 1438, 3897, 4162, 2869, 3120, 2771, 3036, 4145, 4238, 154]. d’ouvrages [417, 3722, 3726, 4068, 4070, 2307, 2694, 2697, 2699, 2732, 2743, 3790, 4209, 3038, 3258, 3995, 4051, 4124, 4164, 4364, 4489, 2053, 2155, 2386, 3239, 3310, 3609, 4023, 3588, 3633, 3765, 4237, 4293, 4197, 3988, 4122, 1730, 1731, 2669, 2392, 1145, 1295, 1300, 1333, 1430, 1431, 1433, 1567, 1580, 1632, 1674, 1680, 1684, 1851, 1852, 1865, 1999, 2373, 2567, 3078, 4265, 4288, 2184, 2245, 4256, 243, 333, 353, 362, 422, 833, 118, 130, 133, 153, 160, 184, 188, 189, 192, 207, 208, 240, 276, 283, 284, 278, 372, 375, 394, 421, 442, 449, 463, 464, 477, 480, 491]. d’ouvrages [505, 526, 527, 528, 529, 530, 531, 566, 627, 654, 667, 668, 669, 766, 784, 787, 788, 790, 791, 820, 821, 822, 823, 824, 825, 827, 829, 830, 831, 834, 836, 839, 840, 841, 842, 855, 856, 858, 860, 877, 903, 943, 969, 1070, 1071, 1073, 1074, 1077, 1099, 1105, 1118, 1290, 1291, 1292, 1293, 1298, 1299, 1323, 1324, 1325, 1326, 1329, 1332, 1335, 1336, 1294, 1403, 1469, 1541, 1542, 1543, 1564, 1566, 1570, 1572, 1612, 1629, 1826, 1828, 1864, 2291, 2311, 2355, 2407, 2755, 2781, 2852, 2866, 2867, 2872, 2874, 2875, 2880, 2881, 2904, 2927, 3041, 3305, 2112, 2159, 3750]. d’ouvrages [221, 246, 247, 361, 363, 454, 457, 534, 588, 589, 590, 771, 835, 837, 878, 952, 993, 994, 1014, 1023, 1024, 1026, 1027, 1031, 1060, 1079, 1106, 1107, 1108, 1111, 1112, 1119, 1146, 1147, 1165, 1166, 1167, 1171, 1172, 1176, 1238, 1242, 1243, 1304, 1391, 1691, 1761, 1830, 2187, 2265, 2317, 2449, 2513, 2580, 2649, 2665, 2784, 2829, 2882, 2978, 3015, 3368, 2965, 3502, 4011, 4169, 4395, 3968, 2156, 2211, 2249, 2273, 2390, 2391, 2394, 2446, 2457, 2471, 2592, 2607, 2612, 2634]. d’ouvrages [2635, 2701, 2703, 2704, 2739, 2740, 2757, 966, 1110, 1241, 4067, 3760, 2710, 2807, 2808, 2929, 3217, 3830, 3013, 3151, 3244, 3926, 4495, 4509, 3388, 3382, 3415, 2814, 4268, 3674, 476, 378, 3665, 4198, 244, 4190, 3865, 1260, 1261, 3530, 3596, 3967, 248, 329, 381, 382, 420, 458, 460, 462, 486, 512, 533, 608, 609, 610, 613, 631, 774, 951, 953, 955, 1028, 1301, 1390, 1615, 1637, 1832, 4089, 1246, 2492, 2493, 2742, 3237, 3720, 2242, 2243, 2521, 2737, 2747, 2836, 2923, 2973, 3001]. d’ouvrages [3006, 3081, 3219, 187, 190, 2210, 2224, 2246, 2335, 2403, 1439, 1474, 1475, 3346, 2468, 2495, 2551, 2566, 2613, 2659, 2677, 2712, 2788, 2816, 2854, 2855, 4069]. Douyère [3828]. Douze [3408]. Down [3194, 3155]. d’Oxford [3995, 1091]. Doyon [1729, 3164]. Dr [1765]. Drabkin [1104, 361]. Draelants [3924]. drafts [2477]. Drago [3829]. Dragoslav [3009, 2568]. Drake [1665, 2289, 2630, 1968, 1103, 1104]. Draparnaud [597]. Draperie [4089]. Drazen [1173, 1165, 1112]. Dream [1664, 4316]. dreams [2903]. Drebbel [1414]. Drei [1216, 1734]. Dresden [3902, 653]. Dreyfus [837]. Drifting [3277]. Dritte [3223]. Dritten [2674]. driven [3861]. droguerie [692]. droite [3451]. Dron [3570]. Drouin [3750]. Drouin-Hans [3750]. drugs [3531]. dualism [2836]. dualité [4308]. Duane [1164, 2052, 667]. Dubarle [2771]. Dubessy [4231]. Dubief [1391]. Dubosson [3970]. Dubost [1474]. Dubrueil [1050]. Duby [2514]. Duc [1541]. Ducassé [140]. Duchesneau [3457, 3925]. Duchesse [1359]. Duckworth [1301]. Duclos [2087]. Ducos [1886, 2680, 3884]. Ducrotay [2431, 2436]. Due [2921, 4078]. Duffin [3926]. Dufossé [1225]. Dufour [3265, 362, 479]. Dufrénoy [795]. Dugac [2386, 2458]. Dugald [3315]. Dugan [3880]. Dugas [448, 684]. Dugué [2609]. Duhamel [2316, 1222, 2897]. Duhem [4475, 3751, 4641, 2806, 2415, 582, 1465, 2350, 2633]. duizend [513]. Dujardin [1302]. Dujarric [214, 1832]. Dulieu [2218, 2755]. Dulong [2688]. Duluc [465]. Dumaître [1304, 3146]. Dumas [2108, 2179, 519]. Duminil [2916, 4007]. Dumont [4486]. d’un [3874, 1538, 2624, 4542, 563, 554, 159, 3687, 3189, 2689, 2215, 123, 1492, 2588, 1009, 4663, 3188, 1960, 472, 3932, 406, 2666, 3398, 2367, 2137, 1205, 2858, 3966, 4199, 3784, 4548, 4490, 1856, 1321, 1742, 2237, 3686, 4523, 2239, 4133, 2206, 4296, 292, 3305, 4656, 4509]. Duncan [2869]. D’Une [1429, 4524, 2164, 760, 3710, 2415, 10, 22, 4085, 4332, 4482, 2497, 2439, 3272, 3290, 3233, 4255, 3079, 643, 2033, 117, 614, 4337, 4096, 4396, 3044, 4597, 105, 4587, 4640, 4480, 2760, 4407, 4452, 2348, 3954, 122, 196, 270, 366, 867, 4138, 1134, 4296, 4589, 362, 375, 471, 4011, 2612, 4495, 4089]. Dunken [1323]. Dunn [4478]. Duns [4097, 4175]. Dunsheath [1419]. duodecimum [845]. Duong [279]. Duong-ba-banh [279]. Dupâquier [3054]. Dupont [4019, 4618, 392]. Dupréel [10]. Dupuy [2308]. Dupuytren [17]. Duquesne [1473, 1473]. durable [4542]. Durand [4292, 2687, 3310]. Durand-Delga [4292]. durante [2815]. Durbin [2522, 3307]. Dürer [3833, 2778]. Dürer. [1391]. Dürers [2778]. during [1296, 3347, 784, 3476]. Duris [3563, 4644]. Durkheim [4366]. Durling [3362, 3503, 3503, 3584]. Dürr [3077]. Durtain [391]. d’usage [1075]. Dusan [2245]. Dussert [4022]. Dussert-Carbonne [4022]. Dutch [3904, 1940]. Duthiers [4637]. d’utilisation [2204]. Dutra [4232]. Dutrochet [2317]. Duval [3197, 381, 1477]. Duval-Destin [1477]. Duveen [3555]. Dworsky [2664]. Dye [1263]. Dyk [2479]. Dyksterhuis [567]. Dynamica [4688, 4685]. Dynamics [2831, 2522, 4683, 4684, 4596, 4687]. Dynamique [4687, 2765, 4683, 1143, 1154, 4684, 3511, 3512]. dynamiques [4398, 4596]. dynamiste [4298]. Dynamistic [3139]. dynastie [1339]. Dziejow [2252].

E.-J [1894]. E.P.H.E [1735]. Eagleton [4557]. Earliest [771, 3236, 3749, 2932]. Early [3270, 1004, 443, 3332, 2271, 2523, 2746, 4557, 2228, 4271, 1912, 3110, 3235, 4514, 997, 2838, 2846, 4294, 4440, 2694, 4256, 2875, 3368, 2703, 1415]. Earth [2445, 4478, 3155, 3276, 3728, 3194, 4440, 2112, 4269, 3285]. East [3009]. Eastern [3668]. Easton [525]. eau [3366]. Eaux [3549, 535, 4633, 1732, 3883]. Eberhard [1522, 4125, 2661, 4204, 2778]. eccelent [4084]. Ecclesia [939]. ecclésiastique [1075]. échange [4663]. échangées [4049]. échanges [3854]. échappement [2623]. échec [4589]. Echeverria [3340]. echten [3314]. Eckehart [4056]. Éclairages [4530, 3549]. éclairé [774]. Eclipse [2491]. Eclipses [3012, 3924]. Eclosion [3585, 417]. Ecluse [2266]. École [203, 838, 1000, 1001, 3043, 792, 2693, 4529, 1346, 1231, 2377, 2745, 731, 3167, 3168, 839, 969, 1176, 2273, 4071, 2898, 2541, 2540, 3044, 1155, 1223, 2797, 3254]. Écoles [4096, 3874, 3622]. écologie [2723, 3740, 3853, 3624, 3606]. écologique [3378]. Ecology [3486, 3319]. Economic [729, 1592]. economics [4183]. économie [2431, 1270, 4502, 3296, 1811, 1735, 839, 969]. économique [1270, 4503]. économiques [3948, 4393, 1735, 2347, 2613]. economist [453]. économiste [214]. Economo [2579]. economy [2846]. écorché [1986]. écosystèmes [3853]. écrites [4367, 2212]. écrits [4046, 1294, 3648, 1786, 4123, 2311]. écriture [4004, 1230, 1521, 1846, 3800]. écritures [1851]. Écrivains [1208, 2573]. Ed [3767, 3731, 3756, 3724, 3863, 3315, 975, 4203, 4025, 3743, 901, 375, 830, 3041, 1112, 4011, 949, 4236, 4203, 4166, 3654, 4616, 3791, 4457, 1788, 2472, 666, 954, 3629, 1922, 4172, 1607, 2127, 790, 791, 834, 860]. Ed. [2440, 2437, 1631, 1500, 1661, 2738, 2315]. Edad [1907]. Eddington [3702]. Edgar [4204, 3812, 1317, 1485]. Edge [1057, 2633]. edición [1436]. Edidit [2233, 219]. Edit [154, 154, 90, 62]. édité [3877]. edited [689]. éditeur [4543, 680]. Edition [3119, 3536, 3118, 3126, 3651, 3316, 2287, 3607, 2364, 1752, 4688, 1499, 1685, 2733, 627, 4268, 1102, 1544, 1294, 971, 2164, 4616, 1408, 2417, 4286, 2110, 352, 986, 2460, 3648, 553, 530, 1637, 1201, 148, 474, 1665, 2528, 883, 1071, 1390]. Éditions [2157, 437, 2135]. editorial [4198, 154]. éditoriaux [4116]. edizione [4301, 2406, 3674, 248, 3240]. Edme [226]. Edmond [3372, 401, 858, 1383]. Edna [2924]. Édouard [909, 236, 987, 1098, 1295, 613, 3404, 3398, 3920, 1406]. Eduard [3570, 1438, 1496]. Eduardo [1436]. Education [2309, 1943, 1835, 1591, 903, 512, 279, 2905]. Edward [3810, 3904, 3899, 2477, 2094, 3275, 2525, 1543, 2159, 2450]. Edwards [4260]. Edwin [2142, 2030, 2479, 3610]. Eecke [219, 151, 481, 985, 988, 930, 1330, 1292]. Eerste [397, 991]. Eeuw [14]. eeuwse [1293]. effect [2338]. effectif [2074]. effects [3696]. effectué [3427]. effet [4075, 4134, 1585, 1252]. effets [4427]. effort [4573]. Efthymios [3700]. égaux [472]. égide [4160]. Egidio [4071, 4070]. Egill [3150]. egiziana [1291]. Église [1075, 3572]. Egret [4392]. Egypt [691]. Egypte [3745, 3935, 4036]. Egyptian [824, 3830]. égyptienne [237]. Egyptische [237]. Ehrenfest [2342]. Ehrenfried [2660, 2468]. Ehrlich [1109, 3877]. Eiffel [21]. Eighteenth [3459, 2310, 2581, 2685, 4652, 3338, 3868, 3860, 3728, 4317, 3476, 3386]. Eighteenth-Century [2310, 2685, 4652, 3868, 3728, 3386]. Eileen [4527]. Eilhard [3489]. Einbeziehung [2701]. einem [4125, 2058, 2695, 1470, 1684]. Einleitung [2287]. einst [1864]. Einstein [2935, 3270, 3663, 2783, 3245, 2972, 2746, 1632, 3865, 4590, 1665, 4671, 2090, 365, 2937, 1145, 4194, 2749, 130, 2335, 2532, 301, 2611, 1580]. Einsteins [1684]. Einzeldarstellungen [1890]. Einzelschriften [3757]. Eirenaeus [3642]. elaborando [3148]. Élaboration [3896, 4359, 3046, 736, 4656]. élasticité [3928]. élasticités [4598]. elasticities [4598]. Elasticity [2664]. élastique [657]. Eldridge [608]. Ele [2815]. Electric [1731, 89, 3696]. Electrical [1268, 1418, 1419, 89]. électricité [3983, 3710, 3985, 3987, 3157, 2832, 3980, 3986, 3984]. Electricity [966, 3982]. électrique [25, 2347, 1885]. électriques [157, 2900]. electrodynamical [3110]. electrodynamics [3835, 3693]. électrodynamique [2366, 3140, 3776, 4252, 2303, 2879]. électrodynamiques [858]. électroencéphalographiques [4483]. electromagnetic [3464]. électromagnétisme [4447, 1086]. electron [4197, 3899]. électroniques [4045]. électrons [3776]. électrophore [3428]. Electrophysicists [1268]. élégante [2997]. Elementa [4198, 4136]. élémentaire [2656, 3709, 1634, 1474, 1632]. élémentaires [482, 3318]. Elemente [482, 188]. elementis [3225, 3516]. Elementos [238]. Elements [2228, 2826, 4391, 627, 4198, 1792, 3497, 4142, 944, 3225, 4014, 1611, 2054, 2544, 3499, 151, 1663, 3725, 4130, 4001, 4145, 3873, 2258, 3497, 1922, 3267, 3811, 4128, 347, 3296, 3493, 1387, 4640, 4137, 4126, 243, 4127]. Elena [3592, 3666]. élève [737]. élèves [2072]. Éliane [1003]. Elias [1685]. Elie [2168, 1027, 4179, 4429]. élimination [3998]. Eliot [887]. Élisa [652]. Elisabeth [2411, 2491, 2960, 2977, 3760]. élisabéthaine [492]. Elisabethan [1019]. Elizabeth [2992]. elle [3672, 4683, 405]. Ellen [3579]. Ellenberger [4663, 3585, 2533]. elliptic [3330]. Ellis [4178]. Elliston [3743]. éloge [4016]. éloignement [779]. Els [3078]. Elsa [4436]. Else [824]. elusive [4316]. Emanuel [238]. Emanuele [1537]. embarras [2508]. Embassy [3362]. emblématique [4464]. emblèmes [144]. Embryological [4235]. embryologie [3960, 4163]. embryologiste [716]. embryology [3061]. Embryos [4235]. Emergence [1263, 2746, 4520, 1435, 2111, 3840, 2922, 4558, 4524, 4398]. Emerging [3593]. Emiel [3564]. Emil [4661]. Emiland [3670]. Emiland-Marie [3670]. Emile [1972, 2290, 2591, 4477, 612, 42, 1495, 185, 243, 1403, 4584, 370, 4473, 1673]. Emilie [3863, 4530]. Emilt [3059]. Emm [44]. Emma [3838]. Emmanuel [1538, 3121, 4298, 3755, 3097, 3259, 3842, 4397, 1346, 3168, 2929]. Emmy [3042]. émouvante [401]. Empêcheurs [4004, 3765, 4176, 4199, 4208, 3863, 3834, 3938, 4011]. Emperaire [1313]. Empéreur [3242]. Emperor [1637]. empiètements [3442]. Empire [3319, 2812, 2679, 3434, 3377, 1064, 3655]. Empires [3533]. Empiric [2550]. empirical [3659, 2836]. empiricism [4056]. Empirisme [3422]. emploi [76, 3045]. empreinte [1380]. Emptoz [4261]. enabling [4182]. énantiomorphes [469]. enceintes [751]. Enchaînements [3256]. enciclopedismo [3759]. encounter [4654]. Encouragement [4612]. Encyclopaedia [4664]. Encyclopedia [1078, 4018, 4600]. encyclopedias [4018]. Encyclopédie [3115, 446, 2585, 3936, 3408, 3292, 777, 3407, 4174, 3175, 2844, 527, 4600, 317, 342, 3291, 316, 318]. Encyclopédie. [2392]. encyclopédique [480]. encyclopédistes [777, 312]. end [2773, 4578, 3076, 4166]. Ende [161, 1564]. endocrinologie [3252]. énergétique [2415]. énergie [2461, 1472, 4558, 4558]. Energiebegriffs [2958]. énergies [3007]. Energy [1472, 3605]. enfante [4154]. enfants [4196, 2844]. Engelhardt [3536, 3276]. engendré [4296]. Engine [2848].

Engineering [1590, 89, 2905, 4182, 2372, 3259, 1419, 695]. Engineers [1268, 4633]. England [876, 507, 1442, 727, 3587, 2539, 3484, 903, 2855, 4557, 2310]. Englebert [605]. English [689, 3126, 3587, 1940, 2537, 903, 191, 2696, 2697]. enigma [3797, 245]. enjeu [4282]. Enjeux [3669, 4181, 4080, 4052, 4684]. enlightened [2872]. Enlightenment [4457, 2709, 3338, 3560, 3835, 4440, 2390, 3281, 4209, 1957, 3485]. énoncés [3189]. Enquête [521, 1253, 3924, 3098]. enregistrement [1726]. Enrico [3813, 2471]. Enrique [3994]. ENS [2593]. Enseignement [676, 3253, 841, 3552, 3226, 4264, 3874, 3227, 4553, 27, 732, 115, 4655, 2541, 3746, 962, 3622, 759, 996, 1967, 2086, 2666, 2679, 2853, 3043, 3216, 3231, 3818, 4048, 4249, 117, 367, 2595, 3044, 756, 1724, 4247, 3295, 3435, 4087, 766, 2211, 660, 547, 550, 1662, 3013]. enseignements [4679]. Enseigner [4195]. ensemble [2017, 2003, 3857]. Ensembles [3857, 2074, 944]. enseñanza [2635]. ensenyament [3078]. Entdeckung [3214]. Entdeckungsgeschichte [278]. entendement [4619]. enterprise [3386]. enthousiasme [4678]. enthusiasm [4678]. entier [1674]. entité [4482]. entomologie [2049]. entomologiques [1169]. entourage [3458]. Entre [936, 4423, 4588, 4604, 3740, 4226, 228, 4549, 2935, 4660, 4424, 618, 2341, 3654, 737, 4488, 3189, 3334, 4482, 4663, 4426, 4393, 3348, 2238, 4415, 4609, 2517, 2805, 2727, 3660, 3473, 4073, 4049, 4425, 4427, 2900, 3867, 3804, 1135, 1659, 4428, 3174, 884, 1939, 3779, 344, 4337, 3928, 2177, 4551, 3286, 3600, 4283, 4483, 254, 4564, 4029, 3984, 557, 3658, 4656, 3858, 3643, 4670, 159, 4413]. entre-deux-guerres [3740, 4609, 4425, 4427, 4604]. entrepreneur [2711]. entreprise [3824, 614]. Entrer [4170]. Entretiens [704, 3286]. Entstehung [2958]. Entwicklung [2636, 4093, 1350, 1734, 3816, 3444, 2740, 690]. Entwicklungsgeschichte [3419, 278]. Entwurf [3419]. envelopes [2954]. environment [4590]. environnement [4590, 2486]. environs [4633]. envisage [4586]. envisager [4586]. Enzymes [3461]. Eötvös [434]. épargnante [3600]. EPHE [2754, 2705, 2669]. Ephestia [3954]. Epic [1501]. epidemia [3946]. epidemics [3532]. Epidemienbuch [3757]. Epidemienbüchern [3314]. épidémies [1117, 4146]. epidemiologic [4519]. épidémiologie [4517, 4522, 4521, 4574, 4518]. Epidemiology [4574, 3235]. Épilepsie [4482]. épileptologie [4482]. Epiméthée [3225, 3790, 4123]. épinette [227]. Epistémé [3302, 2157]. epistemic [4095]. epistemische [4095]. Epistemologia [4511, 4475, 4232, 2806]. epistemological [4600, 4604]. epistemologici [3491]. Épistémologie [4266, 4660, 4512, 3027, 4284, 3793, 3677, 4641, 4340, 4162, 4573, 2781, 4163]. épistémologique [2657, 4600, 4282, 4051, 4124, 4604]. épistémologiques [2220, 4046, 4047, 2260, 2867]. epistemology [4660]. épistolaire [4663]. épistolaires [2517]. Epistolário [529, 1614, 1571]. Epistolicum [2662]. epistomology [4573]. epitaxy [4626]. épizooties [838]. época [2815]. Époque [3948, 3942, 3820, 2575, 3729, 1132, 1838, 517, 3308, 755, 3310, 1161, 3233].

Époques [1338]. épouse [4196]. Épures [4009]. équant [2856]. équation [1377, 2909]. equations [4020, 1388, 3331, 2622, 3328, 1963, 1656, 3329, 375, 2596]. équatorial [4118]. Équilibre [1437, 121, 2055]. équipages [2799]. Équipe [2396]. Equitation [3113]. Equivalence [4686, 3055, 1358]. Equivalent [2054]. Era [3485, 4655, 950]. Erasmus [2804]. ère [507, 4655, 2033]. Eresus [3482, 3640]. ereuthophobia [4656]. éreuthophobie [4656]. Erévan [958]. Erfinder [2607]. Erfindung [1733, 787]. Erhard [2780, 3222]. Eric [2388, 3828, 4290, 3695, 2872, 4436, 4257, 4176, 4293]. Erich [360, 506, 2449, 533]. Erik [4079, 394]. Erika [4351]. Eriksson [1020]. Erinnerungen [3156, 1111]. Erklärung [2989]. erläutert [2058]. ermete [3813]. Ermit [2830]. Erna [2536, 1526, 2645, 2785, 1028]. Ernest [689, 1592, 3491, 791, 4253, 3253]. Ernst [4142, 3336, 1420, 3153, 3515, 3059, 1933, 769, 1262, 1470, 674, 1299, 2566, 4579, 4666, 1931, 2248, 3500]. Errata [167, 2017]. Erratum [3542, 4399]. erreur [81, 775, 1964]. erreurs [2455, 1569]. Erstauflage [1470]. Erste [1070]. ersten [3757]. erster [1216]. Erträge [2701]. érudit [2089]. eruditionis [844, 845, 219, 331]. Erudizione [3562]. ervaring [14]. Erwin [3578, 4093, 2031, 583, 3156, 3578]. Escarpit [2370]. Esclangon [791]. Escudie [3487, 3261]. esferas [1542]. Esmond [631]. Espace [2972, 3055, 4640, 860]. Espaces [4414, 2796, 4415, 3373]. Espagne [576, 599, 4226, 4484, 897, 1601]. espagnole [3854]. España [2805, 3994, 3735, 2965, 2926]. español [2805]. española [4409]. españoles [2644, 2965]. Espasa [62, 566]. Espasa-Calpe [62, 566]. espèce [2857, 596]. Espèces [3373]. esperienza [1537]. Espinas [4445]. Esprit [371, 4141, 272, 4354, 4033, 3875, 4657, 3881, 2502, 2660, 3518, 1990, 941, 4017, 4119]. Esquisse [643, 1115, 122, 196, 270, 366, 2302, 471, 10, 32, 117, 277, 362]. Esquisses [457]. Essai [3225, 4189, 2996, 2953, 2969, 20, 137, 1185, 2313, 4592, 1423, 2698, 2108, 3995, 4051, 4124, 2704, 3777, 2598, 2895, 4179, 15, 2314, 2548, 2300, 1906]. Essais [2253, 1096, 3561, 2122, 3742, 3530, 4081, 3372, 4619, 1136]. Essay [2664, 4592, 1057, 4321, 512, 552, 3930]. Essays [3010, 3866, 3506, 1030, 4271, 452, 2992, 2810, 3603, 442, 4320, 3120, 3597, 65, 3751, 3965, 2804, 2700]. essentiae [3316]. essentiels [3612]. Essere [4475]. Essor [4089, 386, 150, 2236, 2377]. Est [1521, 1846, 3355, 4303, 3399, 1914, 4683, 4118, 4445, 4544]. est-elle [4683]. est-il [3399]. Establishments [2262]. estestvennonautchnye [601]. estestvennyh [983]. Esther [2917]. Esthétique^{≫} [909]. Estrange [3603]. estudio [1436]. Estudios [1076]. età [4309]. établissement [208]. établissements [3785, 1869]. était [3641]. Étalons [1917]. étape [3928, 1404]. étapes [1989, 1889, 2060, 2756, 385]. État [193, 4413, 3695, 971, 3188, 1965]. états [93, 1340, 4656, 1659]. États-Unis [1340, 4656, 1659]. étendue [3137]. étendues [3451]. éthérienne [3140]. ethical [3604]. éthico [4288]. éthico-politique [4288]. ethics [3482, 3860]. éthique [4542, 1756, 2634]. éthiques [3824]. ethnologie [3460]. ethnologique [3052]. Étienne [4003, 3153, 2774, 963, 4608, 4469, 2017, 2003, 3955, 4413, 4058, 1284, 4677, 3942, 2376]. Étienne-Louis [2376]. Étienvre [1302]. étoiles [3280, 4291]. étrange [2292]. étrangers [1307, 2072]. Etruschi [1441]. Ettore [829]. Étude [1774, 2868, 86, 902, 272, 1849, 692, 22, 91, 4354, 2027, 4521, 1487, 797, 881, 964, 1245, 1302, 1310, 1311, 1480, 1702, 1704, 1706, 2064, 3729, 3158, 2247, 2135, 76, 3661, 867, 1678, 524, 2436, 356]. Études [2827, 279, 2166, 2504, 3343, 2252, 2754, 3178, 2750, 4081, 1836, 4026, 4644, 3302, 4298, 4286, 3314, 1978, 2420, 3043, 792, 984, 4560, 2693, 3241, 3055, 2705, 886, 1346, 1735, 3147, 1231, 2758, 3741, 1775, 2786, 380, 2771, 839, 969, 2407, 1532, 2273, 3244, 2613, 2913, 728, 3676, 3079, 1579, 2354, 4518, 2376, 2386, 3239, 3609, 2291, 3326, 3025, 2970, 3231, 213, 3809, 4285, 1536, 1607, 3720]. Études-VI [792]. etwa [954]. eucharistique [4064]. Euclid [689, 2826, 1792, 2054, 4391, 3497, 627, 4198]. Euclide [2544, 988, 1663, 3725, 223, 4135]. Euclidean [2228, 4348]. euclidee [3829]. Euclides [238, 4198, 152, 4131]. euclidienne [4136, 1790, 3056, 2385]. euclidiennes [3147]. Euclidis [1163]. Eudoxe [2818]. Eudoxos [2467]. Eug [90]. Eugene [3319, 823, 1265, 2257, 46, 586]. eugenics [3756]. Eugenio [2406, 2926, 4069]. eugénisme [4438, 3756]. Euler [2662, 1097, 1608, 1609, 2025, 2056, 188, 1332, 4495, 4567, 3072, 1514, 1513, 3600]. Euleri [2873]. Eulers [1332, 2355]. Eurich [1771]. Europa [876, 1100]. Europäische [2823]. Europe [4165, 2451, 1707, 4207, 2686, 4647, 3827, 4237, 2050, 3173, 3854, 4656, 4514, 3796]. europea [4070]. European [3697, 3533, 3009, 3786, 3666, 3965, 4209, 3786, 4160, 1894]. européen [4160, 2651]. européenne [3929, 1379, 1026, 2273, 1830, 4089]. européennes [3896]. Europeo [2914, 2631]. Eva [4477, 4441, 1475]. evaluation [2061]. évaluations [3356]. évaluer [4415]. Evan [3485]. Evandro [3824]. Evangelista [947]. Evans [1936]. Évariste [375, 2311, 1294, 669]. Evelyn [4437, 1983, 4346]. événements [2156]. Everett [4077, 1476, 3726]. Evert [1790]. everyday [3434]. evidence [3749, 4183, 4131]. Évocation [145]. évolutif [137]. Evolution [2831, 2228, 2666, 89, 2933, 3506, 3528, 2036, 4676, 3368, 279, 1270, 357, 1058, 2933, 3260, 2002, 2001, 693, 1630, 2029, 2188, 3149, 4548, 4676, 2074, 1065, 23, 867, 3953, 1882, 2957]. Evolutionaries [4638]. evolutionary [2511, 4210]. évolutionnisme [4176, 3104, 3382, 3966]. évolutionnismes [4638]. évolutive [137]. Ewing [89]. Ex [4301, 3785, 4301]. ex-libris [3785]. Exact [2274, 1151, 1750, 3939, 1071]. Exactas [566, 1324]. exacte [371]. exactement [4466]. Exactes [1229, 935, 4003, 2582, 494, 841, 2612]. exakten [2384, 2058, 2612]. examens [3044]. exception [4582]. Excerpt [747]. exchange [4663]. exclure [3582]. exemplaire [161, 1425, 292]. exemplaires [4280, 1211, 437, 2135]. exemple [2584, 3263, 3188, 2486, 3512, 3190, 2204, 4551, 4062, 2454, 4027]. exemples [4128, 4338]. exercice [3494]. exercices [3875, 4591, 3225]. exercises [4591]. Exeter [3347]. Exhibition [2452, 833]. Exile [2977]. existence [3442]. exophtalmique [3919]. exotic [3638]. exotique [1305]. exotiques [279]. expand [3570]. Expansion [3533, 1889]. expédition [2910, 1320]. expeditions [4478, 1524, 3778]. experience [3808, 1918, 121, 114, 976, 4586, 4319, 1360, 4564, 1203, 858, 1204, 3001]. experiences [4687, 4687, 4166]. Experiencia [3735]. Experientia [4110]. Experiment [2452, 2224, 3312, 4586, 3794, 3341]. Experimenta [2326]. Experimental [4637, 3192, 2963, 4677, 3874, 386, 2005, 3954, 13]. experimental-fysiker [13]. expérimentale [1785, 226, 114, 4637, 3252, 3960, 1487, 4407, 196, 1284, 4443, 454, 4677, 1058]. expérimentales [155, 275]. Experimentalsysteme [4095]. expérimentateur [587]. Expérimentation [4483, 4027]. Experiments [1389, 4166, 3659]. expert [4469, 4468]. Explaining [3532]. explanation [3892]. Explication [606, 3711, 2463, 1818]. expliquent [4170, 4568]. expliquent-ils [4170]. Expliquer [4346]. exploitation [1584]. explorateurs [94]. exploration [3567, 1132]. explorations [3749]. explorer [798]. explosive [4588]. exponente [238]. exposé [1478, 1632]. Exposition [722, 279, 182, 181, 3221, 124, 1077, 1079, 1582, 1175, 815, 840]. Expositions [1011, 1530]. expression [4412]. extensions [2938]. externalism [4320]. Extr [4076, 4435]. Extrême [203, 3721]. Extrême-Orient [203]. Extrait [1605, 463, 529, 279, 1750, 249, 841, 1176]. Extraits [2436, 1271, 1354, 480, 1832, 1663, 1437, 1474]. Extraordinary [828]. Eyben [3564]. eye [3763, 3926]. eyes [3885]. Eyre [478]. Ezrahi [3726].

F [89, 4433, 3099, 137, 2411, 1317, 1485, 1459, 581, 651, 694, 3733, 4478, 536, 3455, 2312, 3526, 2741, 3630, 1780, 744, 2644, 132, 1671, 3008, 532, 2922, 3149, 1236, 2813, 215, 3603, 2926, 234, 3732, 2611, 479, 484, 2771, 184, 372, 766, 791, 820, 1332, 1542, 1612, 1629, 3041, 246, 247, 534, 1107, 2499, 2665, 3013, 1261, 2414, 2242, 1475]. F. [1992, 2896, 1321, 929]. F.R.S. [872]. Fārisī [1818]. fī [2372]. Fabien [3703, 4258]. Fabienne [3764]. Fabre [447, 3213, 3265, 2975, 2197]. Fabri [3458]. fabricants [255, 2178]. fabrication [3208]. Fabrice [4390]. Fabrics [2294]. Fac [3938, 906, 3924, 3722, 3596]. fac-similé [906]. Face [4066, 4355, 3400, 2002, 2422, 3348, 4687, 3502]. facing [4687]. facoltà [4076, 3674]. facsimile [1752, 4348, 1099, 1145]. facteurs [3311, 4517]. factor [4520]. factorielle [3379]. Faculdade [529]. Faculté [279, 956, 275, 1939, 1304, 4679, 3435, 279, 1729, 273, 2163, 2288, 1574, 2755, 3230, 458]. facultés [3229, 957]. Faculty [4679]. Fagot [3311]. Fagot-Largeault [3311]. failure [4589]. faire [3863]. fait [4490]. faite [279]. faith [3744]. Faits [2104]. Faivre [2538]. Falconer [3899]. fall [3728]. Fallstudie [3214]. Fälschung [3245]. false [4634]. famille [737, 193, 4647]. Fan [3880]. Fantini [3153, 2551]. Fantoli [3727]. Farabeuf [2067, 381]. Faraday [2171]. Farewell [4639]. Farriery [1244]. Farrington [1468]. Fasc [2592, 2191, 2193, 2296, 633, 989, 972, 973, 1035, 1120, 1122, 1123, 1316, 1445, 1511, 1507, 1837, 1946, 2037, 2039, 2040, 2043, 2132, 275, 3348, 1113, 4093, 3122, 3148, 2582, 4096, 4397, 3626, 1346, 1893, 2288, 1231, 3215, 941, 362, 307, 404, 489, 842, 3596, 2246, 1639]. Fasc. [3212]. fascicule [1945, 1914]. fascicules [866, 974, 1640, 1797]. Fasol [3384, 3384]. Fasol-Boltzmann [3384]. Fassin [4436]. father [4668]. Fauchard [1168]. Fauché [3881]. Faujas [2256]. Faulconer [1505]. Fauque [4312]. Fauser [2287]. fausse [1096]. fausses [1096]. Fauvergue [4619]. Faux [4567, 2623, 4634]. Favre [479]. Favrholdt [3411]. Faye [3411, 3587, 3699, 2523, 3463]. Febrer [3735]. Féd. [2266, 2267]. Fédérale [1427, 1285]. Fédération [1744, 1159]. Federico [4602, 565, 1300, 189, 1948]. Fedi [4233]. Feinberg [1472]. Feingold [3636]. Felice [2784, 4543, 2784, 2978]. Félins [1761]. Felip [4453]. Felix [1030, 2636, 1501, 953, 2215, 4252, 1302, 832, 1764]. Félix-Archimède [1764]. Fell [4165]. Felt [3840]. Feltz [3606, 3943]. femme [1172]. Femmes [4264, 751, 4147, 4620]. fer [3208, 1558, 1589, 21]. Ferchault [714]. Ferdinand [2313, 2157, 2997, 3401, 949]. Ferdinando [3786, 4076]. Ferguson [1265]. Feridun [961]. Fermat [3359, 3933, 3932, 770, 2137, 787, 1325, 1527, 1623, 346, 3804, 108, 175, 2137, 239]. Fermi [1472]. Fermilab [3165]. Fernand [4314, 35, 3258, 527]. Fernand-Demeure [527]. Fernande [1353]. Fernando [4369]. Fernex [2580]. Ferranti [89]. Ferrara [2874]. Ferrari [2567]. Ferre [4206, 3043]. Ferreiro [4475]. Ferréol [879]. Ferretti [2808]. ferreux [736, 3989]. Ferrier [4355, 1774]. Ferrner [876]. Ferro [2919]. Ferry [879]. Festa [4071, 4070]. Festgabe [2479]. Festschrift [36, 4177]. Festschriften [3843]. Festugière [8, 58]. feu [2509, 1723, 834, 2273, 225]. Feuer [2532]. feuilles [4283]. Fever [2537, 3235]. févr [162]. février [199, 1270, 3176, 3408, 744, 2950, 370]. Feza [4167]. fiables [4581]. fichiers [1579]. Fichtes [4053]. Ficin [4394]. Ficino [874]. fiction [4564]. field [3840]. Fields [2294, 1240]. Fierz [2903]. fifty [3169, 3945, 3257]. fig [3420]. Figala [2566]. figura [4069]. figure [4478, 4464, 4301, 3069, 2858, 483]. Figures [4508, 3787, 4474, 4139, 3862, 391, 2055, 3648, 66]. Figurovskij [1332, 1336]. filarete [4239]. Filipović [2184]. fille [4196]. Filliozat [2730, 379]. filosiofía [3062]. Filosofia [4511, 4344, 2806, 3454, 529, 566, 1324, 3562, 4076, 3414, 3870, 3383, 2387, 4068, 3674]. filosofica [4068]. filosofici [4069]. filosofico [2395]. Filosofo [2375]. filozofije [2184]. Fin [432, 366, 870, 4481, 386, 4512, 2812, 3793, 4286, 3621, 1309, 1704, 1756, 2448, 3779, 984, 1667, 3780, 4578, 4135, 1224, 1739, 1930]. fina [3901]. Final [2501, 3052]. Finance [4185]. Finanzmathematik [4204]. Finding [3933]. fine [4068, 4122, 248]. finies [3190]. finir [3608]. finite [4020]. Fink [3370]. Finlay [3235]. Finn [4079]. Finocchiaro [3988]. Finot [957, 962]. Fiocca [2874]. Fiolle [753, 753]. Fiore [2991]. Fiorentino [1216]. Fire [1891]. Firenze [3562, 3150, 1571, 603, 2551, 2976, 2787, 2344]. firme [1736]. Firmin [4252]. First [3632, 1020, 768, 1827, 1389, 2477, 4290, 3861, 442]. Fischer [3153, 4022, 2174, 3412, 3742, 3415, 2132, 3412, 3812, 3276, 2917]. Fischer-Homberger [2917]. Fisher [2043]. Fisica [566, 62, 3238]. fisiche [3941, 4076, 4435]. fisico [2808]. fisico-matematico [2808]. Fisk [955]. Fissel [3459]. Fission [2472]. FitzGerald [2171]. Five [3386]. fixed [240]. fixes [2254]. fixisme [3966]. Fizeau [4075]. Flachat [2020]. Flad [1327]. Flamel [3648]. Flamsteed [4290, 2388]. Flandre [1409]. Flavia [4516]. Fleck [3317, 4490]. Fleckenstein [875, 2519]. Fleischer [3148]. Fleming [3765, 1171]. Fletcher [2574, 3012]. Fleuves [2443]. flexibles [1282]. Flight [1895]. Flight-Claims [1895]. Flocon [1328]. Flore [948, 1390]. Florence [934, 4645, 2734, 1541, 1784, 1747, 2801, 1469]. Florencio [1171]. Florenz [1216, 2406]. Florey [3059]. Florian [73]. Florida [2109, 2539]. Florimond [2291, 2089, 2087]. Florkin [10, 2264, 1024, 1479, 1147, 2488, 774]. Florum [889]. Flos [889]. flottants [2055]. flou [3413]. Flowers [992]. Floyd [3262]. Fludd [2564]. fluid [799]. fluides [925, 3160, 1993]. fluviale [2174]. fluxions [1573]. foetus [751]. foi [924]. fois [151, 481]. Fokker [1590]. Folge [3815, 4097, 4175, 3489, 4234, 3416, 2287, 3443, 3346]. Folgen [3245]. Folie [3462, 3530]. Folio [3372]. Folio/essais [3372]. Folklore [1442]. folles [3502]. Folse [3699]. Fonction [2856, 2220]. fonctionnelle [3045, 1146]. fonctionnelles [1282, 3328]. fonctionnement [4389, 2163]. fonctions [2760]. Fond [2256]. fondamental [1322]. fondamentales [3448, 3234, 944]. Fondamenti [943, 4076, 4435]. fondateur [2498, 2797]. Fondateurs [4023, 4641]. Fondation [1429, 3115, 3336, 4094, 124]. Fondazione [2232, 1333, 3674]. Fondement [3747, 4014]. Fondements [2059, 3145, 2458, 2134, 2312, 3028, 2386, 2634, 2012, 4308]. Fondo [2395]. Fonds [351, 2984, 4434, 879, 1230, 573]. font [986]. Fontana [2784, 2978]. Fontenay [2593]. Fontenelle [704, 703, 720, 1488, 699, 698, 1090, 700, 4384, 701, 3286, 4001, 3285]. Fontenelle. [3299, 3300]. Fonteniers [1541]. Fonti [2882]. Forbes [1590, 4290, 2185, 2388, 513, 1998, 2872]. force [3065, 3512, 1209]. Forces [3985, 1947, 3600, 4573, 1240]. forces-de-vie [1947]. forerunner [3749]. Forerunners [946]. Forest [4412]. Forge [2247]. forgotten [4238, 610]. formalization [3860]. Forman [2262]. Formation [3723, 2194, 272, 3046, 157, 3236, 2120, 3381, 3398, 585, 4250, 3379, 3588]. formatione [4202]. formative [4144]. formazione [4262]. Forme [3894]. Forment [1271, 1354]. former [3154, 4662]. Formerly [1472]. formes [4628, 3722, 678, 3862]. Formio [2256]. formulaire [2817, 3189]. Formulario [1099]. formulation [1965]. formule [2366]. formules [4013, 3190]. Forscherleben [3035]. Forschung [2843, 2479, 2958, 2829, 2701, 1422]. Forschungen [4256]. Forschungsinstitut [2566]. Fort [1003]. Fortenbaugh [3482]. Forthomme [4199]. Fortier [4292]. Fortification [1266]. Fortino [4475, 4475]. Fortschritt [2849]. Fortun [4077]. Fortunato [4543]. Forzy [3938]. Fothergill [800]. fou [3408]. Foucault [295, 3530, 4569]. Fouchy [4383, 4381, 4388, 4387, 4386, 4378, 4380, 4389, 4384, 4385, 4379, 4382]. Foundation [730, 1867, 3590, 4160, 730]. Foundations [3810, 3524, 3194, 3486, 3866, 3604, 3581, 2025, 3389, 3794, 3808, 3439, 2747, 1505, 3219]. Founder [2159, 1752]. founders [532]. Founding [1417, 4514]. Fouquet [342]. four [4478]. Fourcroy [1687, 761, 1205, 1134]. Fourcy [3167]. Fourier [2155, 2207, 2595]. Fourmi [3730]. fourteenth [3815, 3521, 4006]. fourteenth-century [3521]. fous [3280]. Fowler [3304]. Fox [3939, 4437, 2672, 4346, 2275, 2116, 3788, 2509]. Foxon [4626]. Foyers [75]. Fr [1440, 1922, 2356, 154, 530]. Fracastor [4062]. Fracastoro [3813]. Fraccaro [1614]. fractions [3525]. fracture [3363]. fragment [1220]. Fragments [2868, 247, 2093, 3442]. Fragonard [1155]. Framingham [4520, 4521]. francais [4477, 3227, 967, 3487, 3746, 4196, 3348, 948, 4289, 2517, 2136, 1245, 2915, 151, 481, 2178, 4292, 4540, 4486, 3686, 3884, 863, 941, 920, 562, 1145, 184, 2797, 1052, 3244, 4604, 613, 2854, 4512, 304, 215, 1353, 484, 2211, 3289]. francaise [203, 4616, 4624, 3116, 1181, 2326, 3594, 3791, 3949, 559, 4641, 4065, 2975, 2651, 996, 1305, 1306, 1309, 1638, 1755, 1834, 2460, 3930, 3664, 1337, 1472, 4629, 2101, 2364, 3577, 3224, 277, 1675, 510, 3167, 3168, 2015, 4581, 1260, 2901, 95, 139, 391, 4436, 614, 2538]. francaises [4477, 4233, 2599, 1361, 4609, 3753, 4096, 3778, 644, 1349]. France [3563, 4512, 279, 2628, 692, 3260, 67, 4173, 2627, 2754, 3558, 3433, 1591, 1523, 2423, 4196, 4033, 4255, 2576, 1662, 3274, 2251, 797, 884, 1704, 1756, 1757, 2672, 2915, 3043, 466, 1772, 1927, 2093, 2773, 3259, 2539, 3966, 4366, 4397, 3052, 3736, 3879, 3883, 3962, 4007, 4146, 4147, 1075, 2494, 2514, 3104, 3825, 3261, 1930, 4489, 2155, 3633, 353, 2852, 4169, 2703, 3013, 3740, 1700, 85, 126, 116, 147, 201, 232, 258, 368, 412, 439, 498, 564, 578, 624, 649, 663, 763, 932, 1010, 936, 1054, 1066, 1363, 1385, 1428, 1589, 1641, 1660, 1745]. France [1768, 1859, 1823, 3572, 4424, 522, 2859, 618, 2766, 4248, 4506, 4580, 2150, 40, 3687, 3229, 1213, 685, 873, 898, 1160, 1228, 4637, 1409, 600, 1092, 1190, 1520, 1562, 1603, 2486, 3819, 4425, 4427, 4446, 4428, 1339, 348, 817, 852, 1140, 2459, 473, 3374, 4280, 3574, 3624, 3659, 1224, 1821, 1742, 1256, 1410, 1585, 4283, 3094, 3601, 2368, 3295, 1253, 2239, 1006, 1952, 4589, 4552, 4430, 2283, 83, 180, 296, 781, 1204, 1497, 2135, 721, 741, 1466, 1531, 4155, 1138, 1286, 286, 75, 547, 550, 4434, 3279, 2846, 3788, 4396, 1521, 1846, 2748, 3146, 1826]. Frances [186, 4290]. Francesca [4435, 3177, 4621]. Francesco [3414, 2358, 565, 3365, 4224]. francese [3870, 4309]. Francine [4526]. Francis [1774, 2925, 1468, 4010, 1893, 1101, 1759, 4291, 4061, 1340, 872, 418, 940, 2336, 3535, 4025]. Francis-Joseph [872]. Franciscan [2158]. Francisco [3791, 3753, 3735, 3388]. Franco [4076, 4435, 3365, 3568, 3081, 4426, 3529, 2368, 1842]. franco-allemande [3529]. franco-roumaines [2368]. franco-russes [4426, 1842]. Francois [1849, 2865, 3896, 21, 444, 4645, 3671, 3761, 4196, 4437, 3457, 2027, 4516, 4094, 3585, 3925, 1956, 3670, 4451, 2533, 1683, 4492, 1331, 3844, 3166, 1896, 2902, 3096, 1075, 2175, 484, 1483, 2026, 3442, 4163, 2771, 3633, 283, 855, 4395, 2273, 4433, 4450, 2996, 2953, 2969, 3349, 1677, 1782, 2121, 2162, 4086, 4503, 1529, 1088, 1856, 4646, 1920, 520, 849, 870, 3242, 1687]. Francois-Bourguignon [1677]. Francois-Vincent [4503]. Francoise [3863, 3302, 4413, 3989, 4432, 3970, 2076, 2972, 3435]. Frängsmyr [3485]. Frank [2069, 301, 3861, 2697]. Franklin [2421, 3341, 621, 743]. Frankrike [876]. Frans [577]. Franz [3608, 4654, 2645, 2785]. französischen [787]. Fraser [2242, 2521, 2737]. fraternel [2774]. Fraunhofer [3427]. Fréchet [1673, 2587]. Frederic [3386, 3827, 3588, 631, 4615, 4476, 271, 3686]. Frederick [2547, 695, 1591, 3060, 3455, 2457, 89, 2510]. Frédéricq [10]. Frederiksborg [730]. Fredholm [4049]. Fredrik [31]. Free [1289]. Free-Thought [1289]. Frege [4050, 3689, 4074, 4308, 3888, 3697]. Freiburger [2478]. Freiheit [4234]. French [2905, 186, 3520, 2109, 2709, 3638, 4188, 2633, 2611, 3306, 1686, 4364, 2938, 4576, 4585, 4594, 4605, 4659, 4667, 4672, 4681, 4690, 4574, 4572, 2935, 4598, 4660, 4612, 2933, 2954, 2931, 4669, 4583, 4590, 4580, 4600, 2934, 4688, 4655, 2936, 4637, 4575, 4679, 4664, 4663, 4571, 4588, 4675, 4671, 4636, 4685, 4609, 4607, 4657, 4674, 4683, 4687, 3933, 3338, 3932, 4638, 3646, 4601, 4593, 4592, 4611, 4639, 4597, 4599, 4586, 4658, 4608, 4587, 4678, 4673, 4689, 4633, 4578, 4584, 4676, 2236, 2725, 4603, 2937, 1722, 4613, 4662, 2151, 4665, 4635, 4684, 4668, 4570, 4582, 4591, 4596]. French [4589, 2983, 4573, 1919, 3091, 4666, 4634, 4656, 4677, 4581, 4604, 4670, 2939, 3636, 1289]. Frenkel [3570]. fréquence [3355]. Frère [9]. frères [1723]. Fresnel [88, 4568, 681]. Fresquet [3735]. Freud [4624]. Freudenthal [3521]. Fricii [939]. Fridolf [3362]. Friedmann [3570, 465]. Friedrich [2478, 2384, 3419, 4235, 791, 2607, 2701, 3731, 1111]. Friends [2942]. Frihetstiden [507]. frimaire [2540]. Frisi [2174]. Frison [949]. Fritz [3789]. froid [1212]. Frolow [3318]. Front [49, 50, 111, 99, 100, 171, 172, 173, 264, 265, 266, 267, 338, 339, 340, 428, 429, 430, 431, 543, 544, 545, 546, 639, 640, 641, 642, 710, 711, 712, 713, 807, 808, 809, 810, 915, 916, 917, 918, 1041, 1042, 1043, 1044, 1128, 1129, 1130, 1196, 1197, 1198, 1277, 1278, 1279, 1280, 1369, 1370, 1371, 1372, 1452, 1453, 1454, 1455, 1551, 1552, 1553, 1554, 1647, 1648, 1649, 1650, 1716, 1717, 1718, 1719, 1804, 1805, 1806, 1807, 1876, 1877, 1878, 1879, 2008, 2009, 2010, 2011, 2080, 2081, 2082]. Front [2083, 2145, 2146, 2147, 2148, 2200, 2201, 2202, 2203, 2278, 2279, 2280, 2281, 2360, 2361, 2362, 2363, 2426, 2427, 2428, 2429, 2481, 2482, 2483, 2484, 2557, 2558, 2559, 2560, 2619, 2620, 2621, 2718, 2719, 2720, 2721, 2793, 2794, 2795, 2891, 2892, 2893, 2946, 2947, 2948, 3021, 3022, 3023, 3024, 3088, 3089, 3090, 3130, 3131, 3132, 3183, 3184, 3185, 3249, 3250, 3251, 3323, 3324, 3325, 3394, 3395, 3396, 3468, 3469, 3470, 3543, 3544, 3545, 3618, 3619, 3620, 3683, 3684, 3685, 3773, 3774, 3775, 3850, 3851, 3852, 3910, 3911, 3912]. Front [3977, 3978, 3979, 4041, 4042, 4043, 4044, 4105, 4106, 4107, 4108, 4151, 4152, 4215, 4216, 4276, 4277, 4325, 4326, 4374, 4375, 4421, 4422, 4462, 4463, 4500, 4501, 4536, 4537, 4606, 4632]. frontières [3373]. frottement [1610, 1047]. FRS [3041]. frühe [4361]. frühen [3650, 3505, 1795, 3316]. Frühgeschichte [1260]. frühmittelalterlichen [2670]. Fruteau [4476]. Fruton [2840]. Fuat [2229, 2545]. Fuchs [3921]. Fueter [188]. fugitives [2909, 4296]. fuite [83]. Fujimura [3863]. Function [4020, 3839]. functions [3332]. Fundament [3489]. Fundamental [3693, 3749, 3702]. fundamentales [1542]. Funding [2262]. Funghi [3105]. Funktion [4053, 3734]. Furley [3236]. fusées [1063]. Fusoris [1346]. futur [3152]. futura [4097, 4175]. future [3945, 2182]. futures [4013]. Fys [3411]. fysiker [13].

G [89, 3752, 2564, 3562, 2628, 2657, 876, 3904, 1022, 2803, 3811, 3384, 1577, 3271, 2388, 3311, 211, 327, 651, 2423, 4053, 3014, 536, 3061, 3337, 2741, 4082, 2847, 3209, 2883, 2641, 772, 1170, 2324, 2666, 2685, 2603, 1685, 68, 355, 584, 3608, 3420, 3535, 3257, 4319, 1472, 1234, 832, 4290, 1910, 2157, 2419, 2941, 3603, 2749, 875, 234, 274, 1606, 2611, 484, 2185, 2771, 2697, 1580, 153, 375, 668, 788, 1290, 1332, 1469, 1612, 2781, 2872, 1060, 1110, 951, 2242, 2747, 3219, 1475]. G. [1808, 4688, 1281, 2896, 3141, 1225, 1294, 2311]. G.-B [4319]. G.-F. [2896]. Gaánita [2816]. Gaánita-Bharati. [2816]. Gabinetto [2941]. Gabriel [4546, 3971, 4176, 1075, 4558, 2652, 1405]. Gad [3521]. Gagnebin [204]. Gagnon [3648]. Gaissinovitch [1027]. galaxies [4506, 2157]. Galaxy [3083]. Galdeano [3994]. gale [998]. Galen [3362, 3503, 3584, 3125, 1689, 1301, 2172, 3125]. Galens [3737]. Galenus [3503]. Galien [2341, 1702, 3964]. Galilean [4686]. galileana [3523]. Galilée [134, 1954, 1606, 1373, 4359, 1858, 2164, 2290, 1358, 2689, 3781, 4258, 556, 1361, 1399, 1528, 2765, 1397, 629, 3255, 3354, 1131, 3353, 3352, 1402, 81, 976, 3350, 1209, 1360, 1362, 1401, 4059, 1359, 1559, 3790, 4051, 3351, 1357, 1403, 1400, 1541, 2861, 3858, 4071, 2164, 4368, 2972]. galiléen [2950]. galiléenne [1749, 3354, 1398]. galiléennes [1607]. Galilei [2490, 3522, 4078, 4125, 4345, 4527, 1665, 210, 62, 1666, 1103, 1104, 3988, 1469, 1542]. Galileiana [1469, 4122]. galileiani [3727, 4072]. Galileo [3519, 2490, 3562, 2289, 3522, 4078, 4271, 4527, 1665, 3414, 62, 1666, 1103, 1104, 3988, 2373, 1469, 1542, 1543, 1968, 3727, 2630, 3727, 4344, 4345, 210, 583, 3988]. Galison [4166, 4002, 4347]. Galissonière [1083, 1048]. Gall [2645, 2785]. Gallen [2479]. Gallia [3197]. gallico [4062]. Gallien [2027]. Gallo [3946]. Gallois [3938]. Galluzzi [2490, 2765]. Galois [375, 669, 1294, 2311, 1657, 24, 1903, 2531]. Galton [1089]. Galvani [3983]. galvaniques [3141]. galvanisme [3984]. Gand [294]. Gandavensi [845]. Gandavensis [219, 844, 331]. Gandt [2771]. Gani [2878]. Gapaillard [3672]. Garand [1230, 1521, 1846]. Garches [2340]. García [4255, 3960, 2824]. García-Bellido [3960]. Garcon [4413, 3989]. Gardair [2976]. garde [1722]. Garden [992]. Gardes [620]. Gardies [3882, 3385, 3498, 3497]. Gardner [2078]. Garfagnini [2406]. Gargam [4620]. Garin [2406]. Garland [4018, 3431]. Garma [2635, 2805]. Garnier [3342]. Garofalo [3921]. Garrabé [3591]. Garreau [9]. Garum [3362]. Garzon [2635]. Gascuel [4398]. Gaspar [986, 3]. Gaspard [26, 1490]. Gaspare [2991, 420]. Gasposchkin [3005]. Gassendi [585, 786, 1289, 4394, 2088, 3928, 568, 646, 1203, 4634]. gassendiste [468]. Gasser [2451]. Gaston [2341, 350, 1544, 376, 3633, 827, 4509, 3032, 2240, 3033, 4251, 1281, 4192, 847]. Gate [534]. Gatto [4070]. Gaubert [3336]. Gaubichon [1303]. gauches [3791]. Gaudin [3920]. Gaudon [1615]. Gaudry [3256]. Gauja [1523].

Gaujolle [2076]. Gaujolle-Zaslawsky [2076]. Gaukroger [3432]. Gaulois [3820]. gauloises [3197]. Gaultier [77]. Gauss [2384, 530, 3135, 2761, 2417, 559, 2414]. Gauthey [3670]. Gauthier [1294]. Gauthier-Villars [1294]. Gautier [2533]. Gavin [1578]. Gavroglu [3700, 3789]. Gay [253, 401, 2549]. Gay-Lussac [253, 401, 2549]. Gayon [4616, 4438, 3515, 3633, 4680]. gaz [720, 3356, 1352]. Gaza [247]. Gaziello [4434]. géant [392]. Gebändigte [1733]. Geber [3651]. Geburtstag [2479, 4177]. Gedenkschrift [942]. Gegenstandes [3444]. gehalten [2058]. Geheimnisse [4121]. Geisteswissenschaften [506]. Gelfand [2811]. Gellert [1107]. Géminos [2628]. géne [270, 366, 4437]. Généalogie [4436, 4350, 4592]. Genealogy [4592, 4316]. Geneeskunde [1482]. Genellis [694]. General [3000, 2333, 3347, 1018, 4635, 417, 566, 1324, 1059, 4643, 2408, 479, 486, 2151]. Generale [602, 166, 259, 334, 423, 538, 634, 705, 802, 910, 1036, 1124, 1192, 1272, 1364, 1447, 1546, 1642, 1711, 1799, 1871, 2079, 2199, 2277, 2359, 2480, 2552, 2615, 2713, 2789, 2887, 3538, 3614, 3679, 3769, 3846, 3906, 3974, 4038, 4100, 4148, 4212, 4273, 4322, 4371, 4418, 4459, 4497, 4533, 4648, 4186, 1949, 795, 4615, 909, 354, 3121, 2312, 818, 819, 1175, 792, 1889, 3601, 4635, 1661, 2370, 1392, 1430, 1431, 839, 860, 969, 1962, 2317]. généralisée [2151]. generation [4604, 158, 4441, 751, 1337, 4604]. generatione [1676]. Generationen [1734]. générations [2532]. généraux [2660, 3216]. généreuses [3600]. generous [2993]. gènes [3473]. Genèse [3782, 3198, 1065, 4523, 4624, 1736, 1348, 1581, 1724, 2315, 4472, 2973, 4542, 158, 2973, 3342, 4490]. genesi [2375]. Genesis [674, 402]. Genet [484]. Genet-Varcin [484]. généticien [716, 4155]. Genetics [3524, 4077, 3756, 4268]. génétique [3960, 2325, 3104, 3954, 3953, 3415, 3756]. Geneva [455]. Genève [3631, 3348, 3213, 1077, 3631, 3348, 3749, 814, 494, 4045, 378, 378]. Genevieve [1690, 1502, 4229, 2212]. genevois [4383, 2900, 573, 681, 718]. genevoise [3348]. Génie [124, 3487, 1135, 770, 3142, 1729]. Génin [2032]. Gennar [2739]. genom [876, 747]. Genova [2395]. Genre [4436]. Gent [991]. Genty [772]. Genuth [3996]. Géodésie [464]. géodésien [2237]. geodesy [4478]. géodynamiques [2486]. Geoffrey [1635, 1731, 460]. Geoffroy [2017, 89, 3942, 2001, 2003, 2004, 1831, 2005, 2000, 2006, 2284, 1284, 2002, 254]. Geoghegan [1912]. Geogr [1108]. Geograficos [1076]. Geografii [2252]. Geograph [2607]. géographe [3604]. géographes [1483]. Geographie [2392, 2827, 1740, 2252, 2257, 2401, 2379, 3375, 4091, 4282, 4030, 2085, 466]. Géographies [3950]. géographiques [2599, 433]. Geological [3155, 1910, 3476]. Géologie [44, 703, 3585, 2533, 4292, 4087]. géologique [178, 2651, 1742]. geologists [3947]. Geology [402, 1779, 456, 2445, 3277, 4269, 4440, 2159]. geomagnetism [3347]. géomagnétisme [4158]. géomancie [2693]. géomécanique [3194]. géométral [2508]. géomètre [669]. géomètres [3695]. Geometria [154, 189, 943, 1291, 3523]. Geometriæ [4136]. geometriam [1163]. Geometric [3303]. Geometrical [2333, 118, 240, 4006, 4636]. geometricum [1153]. Geometrie [3201, 2777, 2921, 2778, 2312, 1790, 517, 3829, 2817, 4549, 4074, 3872, 3811, 4655, 42, 1634, 4009, 4001, 1180, 2385, 3056, 4493, 3239, 120, 242, 4623, 2059, 584, 3801, 702, 4071, 3799, 1182]. Géométrie. [3833]. géométries [4640, 3147, 4492]. geometrii [2385]. géométrique [3448, 1515, 812, 4636, 984, 467, 4640]. géométriques [3451, 144]. Geometry [3839, 1847, 1789, 2777, 4655, 4348, 2228, 824]. géomorphologie [2150]. Geomorphology [2112]. Geophysical [3154, 3347]. géophysique [4091, 464]. Georg [2770, 3316, 4548, 4338, 3203, 1890]. Georg-Agricola-Gesellschaft [1890]. George [89, 4365, 2544, 3727, 970, 301, 899, 2779, 995, 2672, 245, 3586, 865, 2810, 4348, 2476, 442, 784, 903, 2982, 178, 864, 324, 574, 3990, 981]. Georges [241, 1749, 45, 10, 15, 2544, 4297, 611, 2028, 304, 330, 465, 3514, 3515, 2318, 1168, 1170, 1953, 2033, 1682, 1765, 2686, 4642, 585, 3197, 3844, 2157, 2514, 4208, 1687, 860, 4395, 1241, 1809, 1407, 4664, 4663, 2416, 3916, 4678, 4673, 4689, 4676, 3918, 3719, 2091, 2797, 2028]. Georgette [2410, 3478]. Georgio [845]. Georgius [844, 219, 331, 794]. geosciences [3347]. Gerald [2920, 2961, 1933, 3420, 2822, 2127, 2335]. Gerard [3603, 2700, 3862, 4261, 3990, 3938, 742, 1056, 1233, 2669, 187, 4342, 4339, 4335, 4449]. Geräte [1693]. Gerauld [1849]. Gerbert [2734, 29]. Gerd [3434]. Gereib [3245]. Gergonne [2967, 1984, 1841]. Gerhard [3153, 2843, 1323]. Gerhardt [2468]. Gericht [2917]. Gerichtsmedizin [2917]. Gericke [3237]. Gerlach [3035]. Germain [2664]. Germaine [3604, 2627, 1845, 3665]. Germains [421]. German [3319, 2109, 1263, 3524, 2710, 4561, 903]. Germany [2457, 2872]. Gernet [3207]. Gerrit [3218]. Gershon [534]. Gerson [3118]. Gersonides [3521]. Gert [3886]. Gesammelte [3650, 3077]. Gesamtausgabe [2470, 2835]. Gesch [3815, 4097, 3843]. Geschichte [1534, 4175, 2536, 2885, 1639, 1709, 2037, 506, 4234, 4095, 1526, 2130, 3214, 2479, 1689, 2685, 1617, 2287, 1413, 1733, 2849, 3443, 2106, 1693, 1935, 1438, 2392, 394, 1070, 1298, 1564, 2265, 2513, 2612, 3346, 2566, 1868, 3488, 2513, 1540, 360, 506, 2780, 2841, 2671, 954, 1412, 1890, 1470, 2229, 2545, 394, 787, 1325, 533]. Geschichtswissenschaft [506]. Geschiedenis [1482, 359, 514, 841, 396, 991]. Gesellschaft [1534, 1928, 2423, 2479, 1617, 826, 1890]. Gesner [2534, 389]. Gesnerus [2193, 2296, 633, 696, 866, 974, 1035, 1122, 1318, 1510, 1837, 1945, 2043, 2132, 1617]. Gesprach [1890]. gestation [594]. Getz [3587]. Gewerbe [1595, 1356]. Gewichten [514]. Geymonat [3357]. ghê [279]. Ghesquier [4558]. Ghesquier-Pourcin [4558]. Ghetaldi [1773]. Ghetaldus [1627]. Gholson [3439]. Giacomelli [210]. Giacomo [4196]. Gian [2406, 2646]. Gianfranco [3674]. Giard [3703, 3703]. Gibbs [2371, 1895, 3076]. Gibbs-Smith [1895, 3076]. Gibrat [1732]. Gibson [89, 418, 940, 1942]. Gift [2642]. Gigerenzer [3434]. Gignoux [614]. Gilain [3433]. Gilbert [4003, 4367, 906, 2065, 822, 2880, 1164]. Gilbertus [3587]. Gill [3752]. Gille [1347, 2589]. Gilles [3669, 2842, 1558, 2999, 827]. Gilles-Augustin [1558].

Gilles-Gaston [827]. Gillespie [402]. Gillis [991]. Gillispie [3939, 2637, 3004, 2528, 1078, 2773, 1057, 3260]. Gimignano [3759]. Gimpel [968]. Gingerich [1334, 3000]. Giordano [1217, 2244, 4069, 508]. Giorgio [4189, 629, 2232, 4083, 1333, 531, 2808]. Giornale [2976]. giorni [248]. Giovan [1333, 2808]. giovance [210]. giovanili [2375]. Giovanna [3359]. Giovanni [1893, 4083]. Girard [1557, 1157]. Girbal [1849]. Girolamo [2750, 3813, 2903]. Gispert [3433, 3825]. Giudice [4435]. Giuliano [4457]. Giunta [565]. Giuseppe [4351, 830, 831, 943, 1099, 2882]. given [240]. glace [4599]. glaciers [3749]. glaciology [3749]. Glaeser [2781]. Glass [3940, 946]. Glénisson [3309]. Glick [2926, 3732]. Glimpsing [2922]. Glisson [4061]. global [1515]. Globe [1744, 2119, 2657]. globes [1758]. Globi [989]. Globusfreude [723]. Gloden [153, 1519]. glorifié [4658]. glorified [4658]. Glory [3364]. Glosclaude [350]. Glossaire [2114]. glycogène [1381]. Glyn [2261]. Gmelig [2534]. Gmelig-Nijboer [2534]. Gnomonics [1854]. gnomonique [4379, 668]. Gnudi [420]. Goblot [4583]. God [4294]. Godart [2157]. Godeaux [10]. Gödel [3866, 4474]. Goebel [4234]. Goethe [4029, 1783, 3370, 941]. Goethezeit [3536]. Goff [4259, 3462, 2732]. Gohau [4558]. Goiffon [1986]. goitre [3919]. Gold [2992]. Goldammer [2479]. Goldbach [1609]. Goldbeck [1733]. Goldberg [3106]. Goldfarb [1936]. Goldhammer [3166]. Golding [2809]. Goldman [2918]. Goldschmidt [3953]. Goldstein [3118, 4675, 3118, 3933, 3932]. Goldstine [2506]. Golem [3724]. Golgi [3901]. Goltz [2473]. Golyer [667]. Gomella [1732]. Gomes [3791]. Gomez [4072]. Gomme [89]. Gonseth [2312, 2157, 2313]. Goodfield [1110]. Gooding [3312]. Goossens [380]. Gordon [3815, 2809, 2112]. górnik [794]. Gorrichon [2494]. Göschen [787, 1325]. Goschler [4200]. Gösta [4049]. Gotlob [4050]. Gottfried [3223, 3529, 4204, 2468]. Göttingen [2872]. Gouan [1556]. Goubert [3577]. Goudaroulis [3719]. Gould [4236, 3149, 2447]. Goulet [2656]. Goulven [3260, 3153]. Goupil [3122, 3413, 3556, 2446, 2704]. Gourevitch [4147, 4035]. Gouron [672]. Government [1910]. Govoni [4262]. Gow [1844]. Gowing [2871]. Goyanes [3627]. Goyard [447]. Grâce [1001]. Grace [235]. Grada [1105]. gradation [4008, 1423]. Graecae [3148]. Graecolatina [2826]. Graf [4204, 506]. Graham [2703, 3565]. Grainger [882]. grains [4028]. Grall [1312]. grammatica [830]. Gramme [10]. Granada [3780]. grand [4538, 2263, 304, 3628, 1909, 79, 1762, 2026, 3081, 3277, 2178]. Grande [3585, 4392, 1359, 121, 4319, 4011, 621, 4089, 1067, 1257, 1563, 1746, 1860, 618, 2014, 1498, 1532]. Grande-Bretagne [1067, 1257, 1563, 1746, 1860, 618, 2014, 1498, 1532]. Grande-Duchesse [1359]. grandes [2599, 67, 728, 1892, 391, 399, 1524, 3490, 1472, 2756, 1400, 40, 654, 2880]. grandeurs [3492, 3980]. Grandjean [4387, 4380, 4383, 4381, 4388, 4386, 4384, 4385, 4382]. Grands [2029, 132, 1663, 981, 1437, 863, 484, 1730, 1403, 1474, 4368, 193, 3348, 1673, 4531, 1761, 40]. Granger [2999, 827]. Grant [4203, 3810, 2094, 3275].

Grants [730]. granulaires [3405]. graphique [3113]. Grassé [327, 2319, 2318, 2575, 4638, 2955, 1241]. Grassmann [3886]. Grattan [3990, 2155, 3939, 2025]. Grattan-Guinness [3990, 2155, 3939, 2025]. gravée [161]. Graves [2389, 2569, 51]. Gravitation [3419, 4495, 4260]. gravité [556]. gravity [4260]. Gravium [976]. Gray [4348]. Grazyna [3011]. Great [402, 2110]. grec [151, 2919, 3070]. Grèce [380, 2119, 2254, 3054]. greci [3105]. greco [1907, 1892, 2705]. gréco-romain [2705]. gréco-romaine [1892]. greco-romano [1907]. grécoromaine [4036]. grecque [2485, 2817, 222, 3700, 1434, 237, 150, 3379, 3643, 3876]. grecques [2441, 3882, 4127]. Grecs [678, 197, 2705, 3962, 4036, 2236, 984, 379]. Greek [2320, 2228, 3236, 3431, 2172, 1844, 3608, 3420, 1863, 2371]. Greeks [2236]. Green [693]. Greenberg [3728]. Greene [2641, 3149, 3506]. Greenwich [89, 4478]. Greffe [4645]. Gregg [3486]. Grégoire [275, 380, 1153]. Gregor [1843]. Gregorio [877]. Gregory [2876, 2457]. Grenade [3753]. Grenoble [2812, 1729]. Gresham [1752, 2388]. Greslou [1941]. Greta [4647]. Gridley [695]. Griechisch [3503]. Griechisch-Lateinische [3503]. Griechischen [1926, 2053, 2603, 1605]. griego^{≫} [238]. Grieken [325]. Griekse [237]. Griffiths [2608]. Grimoult [3966]. Grinevald [3972]. gripe [3946]. Griset [4645]. Grison [3168]. Grmek [1545, 2189, 2343, 2576, 4035, 2551, 3927, 3272, 3834, 3961, 1173, 2668, 4146, 1112, 1691, 1165, 2919]. Grob [2920]. Groeben [3059]. groei [325, 396]. Grolier [1579]. Gros [2687]. Grosbras [4033]. Grosse [3040, 3035, 3156]. Grössenlehre [2470]. Grote [302]. Groult [4313]. groupe [1006]. Growth [156, 89, 2183, 2924, 2236, 1110]. Grubb [3940]. Gruby [610]. Grunblatt [1115]. grundad [1150]. Grunde [3201]. Grundlagen [1689]. Grundsätze [536]. Grundzüge [2535]. Gruner [1734]. Gruppo [4076]. Gruvan [748]. Gruvbrytning [748]. gruvbrytningen [748]. Gua [3172]. Guadeloupe [3753]. Guagno [2188]. Guaydier [399]. Gued [4558]. Guénard [2741]. Guénon [69]. guérison [3875]. Guerlac [2772, 1389, 2837]. Guéron [1787]. Guérout [4393]. Guerra [3753]. guerre [4574, 4562, 4201, 4311, 4427]. Guerres [1892, 3740, 4424, 4426, 4609, 4425, 4427, 4428, 4604]. Guespin [4398]. Guespin-Michel [4398]. Guettard [1424]. Guiart [139]. Guichet [4454]. Guide [4434, 3828, 442, 2826, 1780, 882, 1075, 3988]. Guidi [2412]. Guido [4507, 2412, 2358, 1343]. Guidobaldo [4602, 2991]. Guilaine [4201]. Guillain [611]. Guillaume [3348, 4286, 2055, 4231, 4319, 479, 356, 1079, 4086, 4158, 1856]. Guillaume-Antoine [3348]. Guillaume-Francois [4086, 1856]. Guillaume-Henri [479]. Guillaumin [3504, 4003, 3504]. Guillermand [2884]. Guillet [67]. Guillin [4433]. guillotine [4405]. Guinness [3990, 2155, 3939, 2025]. Guiod [3074]. Guitel [2212]. Guitton [2243]. Gulliver [4329]. Gundolf [2843]. Günergun [4167]. gunpowder [3551]. Günter [3505]. Günther [3734, 3886, 3416]. Gupta [2816]. Gurriá [2815]. Gusdorf [1749]. Gustav [690, 861, 1733, 2829]. Gustave [2340, 480, 860]. Gustave-Laubat [2340]. Guthrie [609]. Guy [4261, 4286, 3273, 2413, 3216, 4487, 1735, 3879, 1231, 3741, 3036]. Guyader [3942]. Guyane [3753]. Guye [90]. Guyotjeannin [462]. Gweneth [1021]. Gyldén [4049]. gymnastique [60]. gynécologie [2032]. Gyung [4316].

H [89, 1164, 3516, 2338, 37, 4079, 1022, 1025, 3304, 2193, 2296, 2052, 2310, 485, 2384, 511, 1631, 3487, 4026, 282, 3673, 3153, 327, 456, 3145, 1866, 4196, 3567, 2608, 2506, 2638, 4203, 3155, 3277, 2838, 2031, 2550, 2847, 2989, 2876, 1169, 751, 798, 799, 838, 882, 1312, 2034, 92, 789, 1163, 1672, 1792, 1793, 1794, 2987, 3218, 3655, 2508, 3535, 2182, 3077, 2231, 950, 992, 937, 1142, 1758, 1935, 1606, 4006, 2408, 479, 2264, 667, 589, 1111, 2665, 3926, 3415, 4198, 1301, 2242, 2566, 2659]. H. [1781, 1988, 2197, 2229]. H.-L.-L [1794]. H.-P [3077]. H.-V [327]. hôpital [1312, 1723, 486]. hôpitaux [1950, 1939, 3596]. Hôtel [1967]. ha-Me’ati [3125]. Habelts [2849]. Haber [2242]. Habert [2088, 2087]. Habicht [937]. habité [3604]. Hachette [527]. Hächler [4243]. Hacking [3434, 4263]. Hacohen [4144]. Hadamard [1515, 2314, 2312]. Haematology [837]. Hahn [3035, 1304, 2921, 3035, 2881]. Hai-Thuong [570]. Haïfa [4026]. Haim [4144]. Haines [903]. Halberstam [1672]. Halbertsma [377]. Haldane [1025, 327, 190]. Hall [2529, 1522, 1936, 3195, 2493, 2529, 509, 3195]. Halle [3536]. Haller [4202, 3418, 2064, 4163, 2873, 3573, 4611, 4566]. Halleriana [4243, 4202, 3418]. Hallers [4243, 2449]. Halleux [2307, 2288, 2592]. Halley [432, 2968, 2994]. Halls [2479]. Hallyn [4314, 3258]. Halperin [1980]. Halsted [4348]. Hamîd [1388]. Hamburg [1827]. Hamburgensis [3148]. Hamburger [2999]. Hamilton [1672]. Hamlin [3967]. Hammond [89, 3579]. Hammurapi [3107]. Hamon [2064]. Hamou [4051, 4124]. Hamraoui [4293]. Hanbury [3210]. handasa’l [2372]. Handbook [1019, 4006]. Handbuch [1262, 842]. handlist [2866, 443]. Hankins [1957]. Hanks [1574]. Hannaway [2474]. Hanno [1238]. Hannover [2423]. Hanoï [279]. Hans [2308, 3489, 2467, 2193, 2296, 605, 2043, 2132, 1420, 4095, 1788, 3148, 3214, 2473, 2034, 1666, 1262, 3960, 2785, 3203, 3750]. Hans-Joachim [3203]. Hans-Jörg [4095]. Hans-Werner [3214]. Hansen [3731, 3119]. Hansen-Matyssek [3731]. Hanson [900]. Haq [3650]. Haramundanis [3005]. Harant [3213]. Haraway [2294]. Hardouin [797]. Hareau [3764]. Harig [1234]. Haritasamhita [2186]. Harman [3792, 3991]. Harmonie [1331]. Harold [89, 2166, 1418, 2062, 3217]. Harré [1924]. Harriet [1934]. Harriette [2143]. Harriot [2526]. Harris [2172, 1414]. Harrison [726, 2530]. Harry [973, 1123, 2109, 3724, 2633, 4364, 3094]. Hart [2961]. Hartcup [3036]. Hartley [2062]. Hartlib [2309]. Hartmann [3156]. Hartner [2189, 2343]. Harvard [1895, 3076]. Harvey [1021, 1688, 796, 4083, 356, 510, 1676, 771, 1635, 1702]. Harwood [3524]. Hasan [4330, 2372]. Hasard [3855, 1978, 2678]. Haschmi [1605]. Hasquin [3809]. Hasselquist [271]. Hauniensis [2233]. Hauptschriften [4204]. Hauron [1886, 2680]. Haut [2407, 1916]. Haut-Rhin [2407]. Hautes [2754, 3178, 2420, 3043, 792, 2693, 2705, 886, 1346, 1231, 3741, 839, 969, 2273]. hauteur [4466]. hauteurs [4380]. Haüy [3222, 2017, 3716, 3715, 3711, 3708, 3710, 3712, 2003, 3713, 3709, 3714]. Hawkins [1976]. Haye [2502]. Haytham [1675, 1818, 4493, 4023, 4330]. Hayyan [3650]. Healing [3589, 3869, 3362, 3587, 2810]. Health [3967, 2846, 3965]. Hearnshaw [3169]. Heart [4609, 2172, 4520]. Heat [2837, 2155, 1476, 1110, 1476]. Heath [395, 627]. Heaven [534]. Heavenly [1017]. Heavens [3675, 4082]. Heaviside [89, 3797]. Heavy [2703]. hébraïque [4441, 3158]. hébreu [19]. Hebrew [3125, 4206]. heden [325]. Hedges [1592]. Hedley [3367]. Heft [1709]. Hegel [4298, 2771]. Heiberg [3725, 2440, 2437]. Heidelberg [3505]. Heidelberger [3316]. Heijenoort [1786]. Heilbron [2262]. Heilbrun [484]. Heilung [2479]. Heim [863, 484, 357, 863]. Heinrich [3696, 2423, 2479]. Heinsius [2308]. Heinz [3223, 2296, 4026, 3731]. Heisenberg [4168, 2959, 2977, 3077, 3599, 3517, 2977]. Helbing [4078, 3414, 4122]. Held [2452, 3965, 3347, 1827, 3368]. Helden [4527, 2444]. Helen [3262]. Hélène [2705, 1930, 3433, 4033, 4510, 3825, 2141, 2315, 4036]. helenistica [1907]. Helga [2903, 3840]. Helge [4079, 3369]. héliaques [2627]. hélice [1348]. héliocentrique [2605]. héliocentrisme [3159]. héliomètre [2728]. Helmholtz [2753, 3581, 3763, 2995, 4338, 4240]. Helmholtz-Hering [3763]. Helmut [3237]. helped [2477]. Helps [443]. Helvétique [765]. Helvétius [310]. Hendrik [3767]. Hendry [3037, 3006]. Heninger [1019]. Henk [3220, 4295]. Henri [4242, 1522, 17, 22, 71, 159, 2215, 3948, 74, 1258, 1416, 1032, 3557, 2770, 3937, 4008, 4049, 4545, 2533, 2033, 131, 670, 1217, 1759, 2356, 3831, 479, 2387, 380, 4430, 1632, 421, 2407, 2927, 1391, 2317, 1474, 2431, 35, 1516, 146, 4637, 4663, 2088, 3265, 549, 4658, 55, 4250, 84, 780, 2087, 949, 4416, 2316, 88, 670, 3213, 2317]. Henri-Ferdinand [949]. Henri-Louis [2088, 2087, 2316]. Henri-Mamer-Onésime [780]. Henrico [490]. Henrik [828]. Henrique [4232]. Henry [89, 2772, 2337, 2834, 1152, 1317, 1485, 1690, 1389, 2110, 2837, 999, 947, 3699, 857, 3195, 329, 486, 1855, 683, 1144, 4349, 2026]. Heou [606]. Heptateuchon [4130]. herausgegeben [506, 1617]. Herbert [400, 351, 2674, 2958, 276, 1475]. Herbier [2562, 4283]. herbiers [2267]. Herbrand [1786]. Herder [3529]. héréditaires [3574]. hérédité [3474, 3473]. heredity [3756]. hérétique [511, 3255]. Hering [3763]. Heritage [2524, 1974, 3411, 4085, 3601]. Herivel [2207]. Herman [2506]. Hermann [4641, 1792, 860, 4264, 4338, 3581, 3886]. Hermant [3863]. herméneutique [3336]. Hermès [58, 3648, 3757]. Hermeticism [2632]. Hermite [2830, 1555]. Hermon [3861]. Hernández [3735]. heroes [3981]. Héron [2440, 2437]. Hérophile [3294]. Herreman [4145]. Herschel [4315, 1237, 1936, 1936]. Hertz [3696]. Hervé [3809, 3213, 3884, 3942, 3343, 3518, 3345, 2973]. Herzen [909]. Heseler [1020]. Hess [3223]. Hesse [1240, 887]. Hessen [4393]. Heurck [949]. Heusser [3123]. heute [2479]. Hevelius [2364, 2283]. Hfl [1775]. Hiatus [4053]. Hideki [3156]. Hien [773]. hier [2736]. hiérarchiser [3582]. Hierholzer [3059]. hiéroglyphe [1851]. hiéroglyphique [4137]. hiéroglyphiques [3648]. Higgins [1268]. High [2777, 3840]. high-temperature [3840]. Hilaire [3942, 32, 4091, 2017, 2002, 2001, 2003, 2004, 2005, 2000, 2006, 2284, 254, 1284, 1831]. Hilbert [2059, 2058, 1977]. Hilbertschen [2058]. Hilfstein [2525]. Hilgevoord [3701]. Hill [2570]. Hillard [2677]. Hilton [2608]. him [235, 1754]. himii [2548]. Hindu [1566]. Hintikka [2333, 3866]. Hintzsche [2449]. Hio [773]. Hippocrate [1525, 2591, 3883, 4007, 4146, 3483, 4007, 4146]. Hippocrates [3362, 1005]. Hippocratic [2320, 3125, 3417]. Hippocratica [2668]. Hippocraticus [3148]. hippocratique [1113, 2842, 2668, 2916, 749]. Hippokrates [3757]. Hippokratische [3314, 3757]. Hippokratischen [3314]. Hirai [4394]. Hire [467]. Hiro [4394]. Hirsh [2922]. hispaniques [1231]. Hist [1485, 2266, 2267, 2392, 1108, 3994, 3735, 3732]. histochimique [1381]. Histoire [3563, 147, 497, 724, 782, 764, 765, 1229, 1589, 1822, 2123, 3635, 4477, 2901, 2166, 2868, 2504, 4410, 158, 2775, 2783, 4165, 676, 732, 1943, 1600, 2435, 2656, 496, 2750, 2689, 1149, 1207, 1314, 1506, 1640, 1835, 1911, 3360, 252, 3311, 1524, 3075, 233, 2796, 629, 3949, 4005, 4255, 1763, 4343, 4413, 357, 4441, 3413, 1135, 1662, 1659, 2045, 2044, 4451, 3079, 2533, 1143, 773, 1114, 1312, 1953, 2448, 3216, 4249, 1386, 1387, 1794, 582, 4625, 4362, 3559, 3149, 3966, 468, 3055, 4340, 4199, 3173, 1217, 4397, 4487, 4629, 4052, 2902, 3096, 1751, 6, 4618]. Histoire [3147, 3435, 28, 4208, 572, 928, 1606, 1607, 1574, 2175, 2786, 2764, 2376, 2734, 494, 738, 3056, 3497, 4455, 3827, 297, 4051, 4124, 4328, 2386, 333, 362, 4620, 208, 491, 791, 840, 1329, 2291, 2400, 3750, 1014, 1026, 1242, 4169, 2211, 1241, 3013, 294, 3382, 3530, 85, 116, 935, 1247, 3191, 4264, 4433, 4450, 4660, 272, 3099, 4246, 279, 1749, 2331, 2865, 563, 3260, 3025, 4334, 2812, 135, 3343, 4248, 3721, 4251, 44, 46, 10, 20, 27, 57, 115, 125, 1378, 4004, 2970, 2913, 4258, 3288, 3116, 51, 4386, 958, 1813, 3691, 281, 1720, 1887]. histoire [4664, 4663, 4360, 3761, 1963, 962, 4438, 3290, 3515, 3622, 4022, 4236, 4160, 4628, 4153, 1815, 4019, 4674, 4021, 4517, 4522, 3933, 4094, 4091, 3662, 3987, 643, 964, 1308, 1311, 1702, 2032, 2845, 2354, 3231, 1966, 551, 4560, 4439, 3608, 4392, 4396, 4642, 3952, 55, 1624, 2250, 3280, 4080, 3756, 4250, 4673, 4158, 4241, 1404, 3357, 1342, 558, 2074, 3742, 1906, 4247, 2368, 277, 1922, 3595, 4523, 23, 1283, 2751, 2824, 122, 196, 270, 366, 596, 3098, 4296, 4162, 4573, 4145, 4164, 3609, 3633, 3351, 1842, 2015, 2612, 717, 3858, 3244, 4509, 462, 2349, 2650, 3720]. histoire [286, 279, 3348, 1472, 4417, 726, 3635, 3767, 41, 18, 59, 139, 2212, 3207, 3703, 444, 448, 398, 140, 450, 586, 728, 1728, 3121, 3153, 3272, 4360, 581, 4261, 3460, 4019, 3585, 2884, 2975, 3639, 818, 819, 998, 1000, 1174, 1613, 1692, 2035, 445, 792, 1608, 2056, 614, 3525, 3559, 3828, 1889, 4412, 4052, 3194, 3366, 3224, 2385, 2514, 3104, 1661, 209, 1058, 1392, 3167, 1430, 1431, 788, 791, 588, 1304, 4011, 1241, 3415]. Histoires [3691, 3313, 4640, 3104, 2745, 2732, 2625, 2825]. Historia [2805, 3994, 2475, 861, 2534, 2644, 1907, 154, 566, 1324, 4409, 844, 845, 3735, 417, 219, 331, 529, 2191, 1778, 1907, 2712, 2788]. Historiae [1710, 389]. historian [2866]. historians [2078, 3947]. Historic [3996]. Historica [3536, 2233, 690, 3755, 536, 1909, 628]. Historical [3533, 2262, 4079, 3797, 3990, 1432, 1688, 3525, 4592, 4009, 1578, 3860, 3439, 1109, 3597, 2537, 3076, 3816, 4018, 799, 4185, 3603, 3367, 4263, 2192, 1944, 2572, 833, 2495]. Historici [2131]. historico [591, 2926, 3720]. historico-mathématiques [591, 3720]. historien [2379, 3297, 1988, 1395]. historienne [3571]. historiens [468, 3244]. Histories [4197]. Historii [971]. historiography [3700]. Historique [1884, 2459, 333, 2868, 4284, 1102, 2711, 1032, 1456, 984, 4592, 2902, 4407, 1075, 1678, 421, 464, 1836, 956, 4195, 355]. Historiques [1943, 2295, 1835, 728, 4287, 1579, 1398, 1775, 2879, 3387, 4565, 463, 2867, 2665]. historisch [3416]. historisch-vergleichender [3416]. historische [2287, 990]. historischen [1689, 2823]. History [1265, 1593, 4458, 384, 162, 163, 3010, 4557, 3459, 2309, 2831, 37, 1016, 1030, 2275, 2276, 2274, 2342, 2345, 2052, 3501, 4527, 2664, 537, 593, 632, 730, 846, 888, 972, 973, 1120, 1123, 1151, 1152, 1316, 1317, 1511, 1507, 1797, 1829, 1946, 1914, 1979, 2041, 2039, 2042, 2117, 2116, 2133, 359, 3436, 3441, 3696, 218, 400, 3000, 328, 443, 456, 2928, 377, 3118, 3733, 2110, 2143, 2063, 2506, 2642, 3386, 3480, 3526, 1925, 1432, 1442, 1500, 1502, 1694, 2129, 2142, 2840, 3274, 2876, 691, 744, 882, 965, 995, 1481, 2323, 2393, 1296, 1750]. History [1844, 1982, 3868, 1940, 4592, 1057, 1636, 1236, 3420, 1578, 3535, 2959, 3002, 3275, 1471, 950, 982, 2810, 2811, 3432, 3565, 1271, 1264, 1354, 1419, 2848, 2419, 2993, 1018, 1827, 2525, 2850, 3285, 2700, 561, 3306, 3367, 2062, 2264, 3120, 306, 96, 118, 133, 207, 250, 332, 442, 531, 667, 820, 2875, 2881, 2112, 771, 837, 993, 2457, 3967, 608, 631, 3006, 220, 2712, 2788, 2816, 89, 4660, 3632, 3810, 4182, 3579, 3880, 3826, 2190, 4271, 1445, 1485, 1486, 1577, 3581, 4664, 4663, 3042, 3370, 3786, 4018, 4514, 3061, 3652, 4674, 3590, 3337, 3751, 3933]. history [3531, 4095, 1833, 2450, 3209, 3211, 3532, 3534, 3630, 3636, 2883, 3917, 2477, 3586, 3587, 3965, 3347, 4673, 3629, 3743, 4188, 4055, 4613, 3196, 3871, 3107, 4318, 4317, 4573, 4209, 4197, 609, 1067, 695, 1486, 675, 798, 799, 800, 3206, 3565, 1592, 1895, 1759, 4458, 4626, 3838, 3506, 2190, 1484, 1913, 2115, 3992, 3154, 823, 953, 1246, 2551, 2855]. hiue [883]. hiyal [2570]. Hoai [2708]. Hoare [4478]. Hobbel [2905]. Hobbes [2166, 3654, 2353, 2282, 3046, 3422, 3425, 3423, 3192, 3424]. Hochschulschriften [2823]. Hocquet [3342]. Hoddeson [3902, 3165]. Hodge [3153, 3337]. Hoe [2420]. Hoeppli [997, 607]. Hoffmann [2751, 1332]. Hofmann [2987, 278, 394, 788, 2355, 2468, 2183, 1956, 787, 1325, 1097, 3835]. Hofmeier [4121]. Hohenheimius [2479]. Hohenheims [2479]. höheren [278]. Hohl [1175]. Holger [3755]. Holland [876]. hollandaises [161]. Hollande [233, 562]. Hölle [1666]. Holmes [3386, 3827, 3588]. Holmyard [746]. Holton [2822, 2127, 2335]. Holy [3655]. Homberger [2917]. hombre [3235]. Homet [2782]. Homine [2353]. Hommage [1393, 2667, 1399, 1459, 1281, 3742, 3922]. Homme [2538, 2421, 923, 4219, 2207, 3781, 3152, 1673, 3213, 3582, 4454, 4052, 234, 310, 4296, 3984, 2499, 3502, 516, 309]. Hommes [4477, 4416, 4222, 2851, 215, 2738, 2087]. Homo [2851]. homonymes [409]. Homosexualität [2989]. hongroise [1897]. Honnecourt [2623]. Honor [2525, 2992]. honoraires [1756]. Honoré [1155]. Honorer [4154]. Hooke [1975, 4124]. hookiennes [2584]. Hooykaas [14, 3048]. Hopwood [4235]. Hoquet [4621, 4455]. Horace [3749, 4090, 3418, 3749]. Horder [3061]. Horeau [3122]. Horiuchi [3729]. horizon [4454]. horizons [4478]. horlogerie [4468, 181, 3345]. horloges [2997, 4347, 1420]. Hormigon [3666]. Horoscopes [3196]. hors [1170, 1173, 2324, 4556, 777, 510, 855]. Horský [1854]. Horst [1266]. Horus [442]. Hoskin [3000, 4315, 1929, 2096, 2276, 1913, 2041, 2115, 2117, 2261, 4315, 1670, 1246]. hosoèdres [985]. Hospital [3868]. hospitalière [1312, 1704]. Hospitaliers [486]. Hospitals [2885, 2671]. Hou [666]. Houël [3199]. Hours [838]. Hourya [3897]. House [4607]. Houts [3439]. Houzel [3433, 2404]. Howard [3940, 328, 3726]. Howell [4294]. Hoyt [2389, 2569]. hrsg [3223]. Ht [2392]. http [3987]. Huard [279, 952, 1112, 1165, 1166, 1691, 2187, 203, 1146, 2768]. Hubbard [1731]. Hubble [3610]. Hubers [2785]. Hubert [4243, 2067, 4391, 421]. Hublin [3407, 3308]. Hud [4023]. Hudrisier [3831]. Hugeniorum [3674]. Hugh [2225]. Hughes [2733, 993]. Hughlings [2886]. Hugo [3083, 2475, 1436]. Hugonnard [2770, 2231]. Hugonnard-Roche [2770, 2231]. huguenienne [4117]. Huguenot [2539]. Hugues [2722]. Huguet [3435]. huitièmes [4369, 4455]. Hulin [4264, 3825]. Hulton [2539]. humain [4619]. humaine [1906, 596, 3372]. Humaines [1574, 1749, 2257, 1729, 4362, 886, 4490, 4052, 4346, 3750, 3635]. Human [3526, 4077, 3482, 3640, 3756, 3001]. Humanisme [492, 386, 3522, 3566, 4643, 3173, 4302, 185]. humanismo [3994]. humaniste [1538, 3]. Humanistische [2740]. Humanists [2373]. humanité [3695]. Humanity [1950, 3506]. Humbert [18, 61, 209, 504]. Humboldt [942, 1060, 1106, 1107, 1108, 1238, 2829, 3375, 2417, 1088]. Humboldt-Kommission [942]. Hume [2143]. Humphrey [1414]. Hunczovski [1707]. Hundert [2034]. Hundred [768, 3169]. Huneman [3968]. Hunt [2547, 3436]. Hunter [3649, 2128]. Husain [2990]. Husserl [3872]. Hutchins [4456]. Hutin [945]. Hutton [2159]. Huxley [3634]. Huygens [88, 1730, 3674, 2659, 2766, 2819, 30, 4110, 4111, 4112, 595, 657, 303, 548, 4687, 4113, 290, 4116, 4114, 3806, 4115, 303, 990]. Huygens-Fresnel [88]. hybridation [1007]. hydraulics [89]. hydraulique [1340, 3366]. hydrauliques [3052]. hydrodynamics [89]. hydrographe [4119]. hydrographique [1320]. hydrostatique [1362]. Hygiène [1116, 1172]. hylozoistic [4061]. hypothèse [3985, 2136, 2205, 2285, 4060, 2157, 3690]. Hypothesen [3201]. hypotheses [4686, 1358, 3138].

I.S.E.A [1270]. IAGA [3347]. Ian [4263, 3940, 3434]. Iannaccone [3339]. IAU [3083]. Ibañez [1501, 953, 1030]. Ibarra [3340]. Ibéro [2609]. ibid [279]. ibid. [2036]. Ibn [4493, 4023, 246, 689, 3650, 1388, 3241, 1436, 4330, 2669, 1566, 1818]. IC [2295]. Ice [4599]. Icona [4239]. iconographique [4643]. icy [4365]. id [566]. id. [566]. idea [4597]. Idéalisme [2169]. idéalité [3876]. Ideas [4478, 1688, 3434, 2749, 4190, 1778, 3653, 1016, 1030, 1057, 2743]. Idéaux [1683]. Idee [2740, 56, 2497, 2439, 4597, 2432, 2074, 1065, 1949]. idées [3563, 3740, 679, 1774, 1849, 2308, 3260, 158, 3821, 1456, 357, 4306, 1696, 3653, 3149, 2102, 366, 596, 4296, 1541, 2156, 3382, 656, 693, 2491, 4229, 279, 4091]. idées/International [3653]. Identification [123]. identité [4559, 3984, 4472]. identités [4628]. Idéographie. [4050]. Ideologie [2674]. idéologies [1117]. Ido [3797]. idraulica [2174]. Ignaz [3123]. ignorantia [3346]. Igor [3610]. Ihde [1576]. Ihre [2478, 1639, 1868, 2037, 2265, 2513, 3245]. ihren [2987]. II [1265, 1594, 3767, 939, 3516, 1749, 748, 2337, 44, 38, 2395, 4078, 1710, 1797, 2116, 3245, 1141, 3792, 2423, 3791, 2661, 2470, 4286, 4076, 4435, 2573, 3503, 1162, 1614, 2045, 3148, 3240, 2174, 2695, 665, 819, 1689, 2974, 984, 987, 1670, 1789, 2055, 2056, 3725, 954, 1440, 565, 3420, 3582, 4643, 1889, 1142, 3607, 1268, 3879, 1230, 2105, 3177, 2149, 3387, 2476, 4023, 192, 394, 521, 787, 790, 830, 856, 1105, 1332, 1014, 1238, 2265, 2513, 3830, 4198, 608, 774, 2551, 1700, 167, 166, 4409, 4303, 4571, 1344, 196, 4223, 550, 3516, 906, 1034]. II [358, 1892, 3076, 1998, 2264, 2242]. II/1 [954]. III [1265, 1533, 1594, 1744, 3767, 939, 213, 1749, 2604, 2834, 44, 4003, 1829, 1946, 3992, 3791, 3314, 1614, 3148, 3240, 3757, 2582, 109, 987, 1672, 1683, 1929, 2055, 3197, 2287, 2705, 4643, 1889, 3736, 2405, 2185, 417, 3722, 1430, 1431, 307, 787, 943, 1299, 4198, 231, 259, 896, 1744, 593, 1178, 2541, 3709, 2540, 270, 4382, 989, 3418, 3168, 880, 878, 2521, 333]. III.1 [4332]. Ijkwezen [514]. Ilana [4026, 3869, 3317, 4490]. Ilgauds [3203]. Illinois [1263]. Illness [2920]. Illustrated [820, 1758, 784, 1110]. Illustration [1199, 4399, 4376, 590]. illustrations [893, 4301, 2743, 958]. illustré [1353, 2675, 1112]. illustrée [1390]. ils [4170]. Ilse [3384, 3536, 2841]. iluminismo [1324]. Ilza [1637]. im [2663, 1540, 2536, 2636, 3650, 3505, 4095, 1526, 2843, 2847, 3757, 2603, 1750, 1890, 2849, 2106, 3429, 278, 1332, 2355]. Image [2109, 2119, 4209, 527, 4002]. Imagery [2979]. images [3161, 4205, 3814, 4188, 1587, 1885, 3873]. imaginaire [1987, 976, 4646, 2851, 2404, 3633, 3950]. Imagination [3814, 372]. imaginations [4332]. imaginé [4113]. Imagining [4256]. Imbert [750]. immenso [4069]. Immune [3762]. Immunologie [2724, 4559, 4574, 3360]. immunology [4574]. Immutabilité [4222, 4014]. Impact [1239, 2749, 3387, 1999, 3106]. Impacts [3943]. imperatore [565]. imperial [4652]. implications [4356]. implicit [4575]. implicite [4575]. implicites [1841]. importance [977, 1882]. important [563]. importante [2681, 4122, 1249]. impossible [3646]. imposture [1985]. impostures [4231]. impression [883]. imprimées [4119, 2238, 3987]. Imprimerie [1079, 1404]. imprimés [1102, 2677, 3753]. imprimeur [1538, 530]. imprimeur-libraire [530]. Imprimeurs [3215]. improvement [3551]. imprudente [3287]. impulsions [4297]. Imre [3703, 1750]. in-12 [10]. inaugurale [4074]. incandescent [89]. incertain [4154]. including [359, 1867]. inclus [1316, 1445, 1797]. inclusion [1919]. incommensurabilitate [2094]. incommensurable [2228]. incompleteness [2937]. inconnu [2416, 1046]. inconscient [234]. incubato [4202]. Inde [4670]. indépendant [1945]. Indes [161, 1114]. indéterminée [1295]. indéterministe [405]. Index [1323, 3810, 2095, 4643, 2631, 3148, 1826]. indexes [3420]. India [4670]. Indian [2816]. Indias [3735]. Indications [4358, 4522, 2188, 1230, 1521, 1846]. indice [2687, 861]. indienne [379]. Indies [2883]. indirect [3890]. indirecte [143]. individu [2857, 4296]. individualisation [4481]. Individualité [234, 3145]. individuelles [4354]. indivisibles [3423, 655]. Indochine [279]. indochinoises [279]. induction [3959, 1183]. inductive [2030]. Industrial [1594, 1421, 89, 2846]. industriale [528]. Industrie [1595, 345, 4608, 1356, 2391, 388, 4612, 692, 4165, 146, 3408, 2014, 4089, 4416]. industriel [2653, 2020]. Industrielle [2391, 1882, 2654]. industrielles [4165, 465, 3052, 208, 1024, 1416]. industrispion [876]. Industry [1263, 4608, 2703, 89, 3788, 4612, 1078]. inédit [1375, 1492, 2967, 4504, 1491, 3784, 2328, 1381, 1919]. Inedita [529]. inédite [3032, 1918, 3033, 2415, 926, 1377, 2179, 1967, 1088, 2595, 3044, 519, 1951, 293, 557, 83, 2073]. Inédites [700, 3031, 3265, 893, 2908, 718, 457]. inediti [4068]. inédits [1774, 2139, 4017, 143, 3348, 948, 2576, 351, 2455, 3698, 4296, 2386, 573, 4546, 2318]. infection [3473]. Infections [997]. Infectious [2578]. Infeld [2531]. inference [4519]. infigurabili [4069]. Infini [2167, 2637, 3561, 2404, 3101, 3897, 2124, 208, 1288, 4623]. infinies [1573]. infiniment [4012]. Infinite [508, 785]. infinitésimal [2982, 3025, 1006, 69, 1958]. infinitésimale [1184, 2012, 208]. infinitésimales [223, 4493, 4023]. infinitesimals [3332]. Influence [903, 1189, 3198, 2486, 2597, 3482, 3640, 2876, 1692, 3625, 585, 3379, 656, 2248, 1507, 632, 1445]. influences [1914, 2477, 846, 973, 1123, 1316, 2345]. influx [2217]. Information [2357, 2807, 2370]. Informations [1899, 1905, 1921, 1996, 2016, 2051, 2075, 2092, 2125, 2153, 2165, 2198, 2209, 2241, 2286, 2305, 2351, 2369, 2402, 2418, 2442, 2500, 2520, 2602, 2626, 2655, 2682, 2731, 2821, 2864, 2899, 2912, 2956, 2971, 3095, 3102, 3117, 3144, 3193, 3232, 3269, 3298, 3335, 3358, 3380, 3406, 3452, 3479, 3513, 3576, 3602, 3822, 3859, 3878, 3895, 4408, 2383, 2940, 2489, 1971, 2998, 2181, 2691, 3205, 2769, 3073, 3430, 2223, 3034, 1952, 3051, 2103, 2330, 2023, 2259, 2456, 2465, 2543, 2590, 2802, 2986, 1700]. informatique [2738]. informatiques [3439]. ingegneria [4076]. Ingenhousz [328]. ingénierie [3194]. ingénieur [3670, 1557, 1541]. Ingenieurbureau [1734]. ingénieurs [1347, 3832, 2595, 4034, 4633, 4486, 2681, 3142, 1541]. ingénieux [2570]. ingenii [2631, 4635]. ingenious [2570]. Ingram [1672]. ingrédients [3777]. Inicios [3437]. initiateurs [1880]. initiation [1881]. Inner [2977]. innovateur [2680]. Innovation [3861, 3400, 4210, 3464]. innovations [1736, 3157]. innumerabilibus [4069]. inoculation [4669, 2813]. inquiry [4010, 4622]. inscriptions [3197, 4285, 4565]. Insectes [1456, 3213, 1137]. insospechada [3732]. inspecteur [1059]. Inspecteurs [3216]. inspirateurs [2494]. inspirée [345]. Inst [3938]. installation [3093]. Institut [2374, 279, 3178, 1270, 3097, 2420, 3043, 886, 4486, 217, 2155, 3633, 1826, 2854, 980, 3357, 2653, 1523]. Institute [235, 4056, 2129, 3636, 874, 1940, 2158, 3196, 4197, 3473, 4430, 2143]. Instituti [2131]. Institution [1255, 282, 3495, 4178, 3591, 2960, 3986]. institutional [3810]. institutionalisation [2151]. institutionalization [2151]. Institutionen [2841]. institutionnalisation [4523]. institutionnelle [2812, 3854]. Institutions [3475, 2811, 4609, 4489, 4237]. Instruction [2411, 4196, 3216, 1821, 620]. Instructions [4362, 3778]. Instrument [2807, 4380, 3665, 3733]. Instrument-Makers [2807]. Instrumente [769]. Instruments [1259, 1415, 1982, 1759, 1854, 3722, 563, 3863, 1882, 359, 4281, 2204, 3345, 3603, 3996, 4051, 4124, 2929, 3082, 1981, 2850, 4018]. Instruments.Part [3082]. Integers [2054]. intégrale [1408, 1294]. integrals [3330]. Integration [1976, 3331]. intellectual [3810, 2190, 4271, 1432, 3705]. intellectuale [2631]. intellectuel [3549]. intellectuelle [585, 2398, 4194]. intellectuelles [4231]. Intellettuale [2914]. intelligence [4174, 3439]. intelligibilité [1360]. intensités [3427]. intensities [1847]. inter [3785, 4604]. inter-établissements [3785]. inter-war [4604]. intérêt [4250]. interactionniste [3677]. Interactions [3919, 3234]. interchamp [4523]. interdisciplinaire [4363]. interdit [2136]. Interdivisional [3347]. interests [4653]. interfuerunt [3148]. intérieur [4091]. interioribus [3503]. interleukin [3869]. interleukin-2 [3869]. Interlingua [830]. intermediary [3588]. intern [3700]. internacional [3592]. Internal [1637]. internat [2068]. International [202, 349, 369, 764, 934, 931, 1055, 1069, 1285, 1427, 1533, 1641, 1744, 1770, 1860, 1861, 1825, 1862, 3340, 2345, 537, 593, 846, 888, 972, 973, 1120, 1123, 1316, 1507, 1914, 625, 783, 897, 1141, 3083, 257, 499, 3386, 2961, 1602, 3653, 1264, 1827, 1206, 306, 96, 332, 1532, 1784, 3368, 953, 1341, 2712, 2788, 1462, 4226, 162, 3631, 5, 4003, 446, 632, 1445, 1463, 4026, 1524, 1461, 4609, 2319, 4446, 1659, 1699, 2044, 585, 2260, 4510, 4530, 3938, 4363, 1230, 1521, 941, 4265, 160, 250, 1014, 220, 221]. Internationale [349, 415, 475, 931, 1053, 1561, 1744, 1767, 740, 895, 1091, 257, 151, 3651, 1206, 4256, 160, 192, 246, 2870, 2191, 2422, 617, 1033, 1444, 1618, 2637, 3521, 2603, 4184, 865, 1827, 2386, 307, 383, 488, 247, 4194, 2246]. Internationalen [2058, 2835]. Internationales [202, 141, 192, 2308, 2189, 2343, 617, 843, 1033, 1121, 1315, 1444, 1508, 1619, 1796, 2038, 2491, 3653, 307, 383, 488, 2070, 3714]. internationalis [1710]. Internationaux [1053, 948]. internazionale [4088, 2297, 2787, 4070]. interne [4153, 1254, 1730]. interprétait [1203]. interpretare [4475]. Interpretation [3944, 2746, 2934, 4294, 137, 2934, 4471, 2822, 4221, 2608]. Interpretations [3307]. interpretazione [2643]. interpreter [2903].

Interpreting [3898]. Interprets [3285]. Interrogation [4171, 3457]. Intersciences [4628]. intervention [468, 2304]. Interwar [4609, 4144]. intill [748, 2475]. intorno [3941]. Introd [3420]. Introduccion [154, 2965]. introducteur [1006]. Introduction [4479, 3421, 3708, 1515, 4577, 3547, 4595, 91, 4377, 1435, 1863, 4682, 3817, 4607, 3446, 3951, 3471, 1487, 984, 3126, 4651, 3651, 3641, 2250, 1513, 4643, 4167, 3284, 4057, 3030, 4126, 3798, 4164, 3887, 4198, 951, 3740, 689, 812, 3534, 151, 355, 3016, 1410, 1585, 4348, 1145, 4660, 4424, 4607, 4673, 2628, 22, 3209, 1953, 1473, 3096, 2126, 2127]. Introductions [2231]. Introductive [1131]. introductory [442]. introduit [3070]. introspection [4443]. Intrusions [4231]. intuitive [570]. invariants [4047]. Invasions [1892]. Inventaire [4119, 437, 2135, 2084, 2219, 2291, 2400, 891, 2204, 2411, 351, 2705, 1604, 2105, 2405]. inventé [4380, 3765, 1221]. Inventeur [2680, 3399, 1258, 1064]. inventeur-innovateur [2680]. Inventeur-savant [2680]. inventeurs [1353, 305, 754]. Invention [2444, 4221, 1731, 89, 623, 2314, 1588, 1432, 1489, 3080, 3004, 145, 1730, 1726, 1210, 2951, 1249, 1981]. inventions [89, 1737, 40]. Inventory [2534]. inverse [3066, 4341]. inversement [3355]. Invertébrés [1678]. investigation [3648, 3001]. investigative [3386]. Invisible [3194, 4205, 2922]. Invited [2659]. Ioannes [1655]. ionisants [2170]. ionosphère [56]. Ions [3776]. Iosif [1970]. ippocratica [2914]. Ir [325]. ira [4083]. Iran [1290, 933]. Irena [2352]. Irene [3589, 3843, 4615, 4265]. Irigoin [4510]. Iris [2685]. Irma [526]. Irmer [3148]. Iron [2757]. irradiation [1598]. irrationalis [4053]. irrationalité [4133]. irrationnels [3334]. Irwin [3836]. Isaac [2529, 2530, 1258, 1416, 3598, 235, 3302, 2900, 1296, 1573, 1669, 1670, 1671, 1929, 2057, 2096, 604, 1142, 2735, 2699, 1572, 1828, 2492, 2493, 2742, 3355, 4485, 4502, 3069, 2857, 2907, 87, 907, 2528, 2375]. Isaac-Bénédict [4485]. Isabel [3946]. Isabelle [3522, 3566, 3922, 4022, 3559, 3924, 4643, 3244]. Isagogae [951]. Isaia [3339]. Ischer [204]. Isdatel’stvo [3731]. Ishizaka [4654]. Isis [162, 2345, 632, 846, 973, 1123, 1316, 1445, 1507, 1914, 306, 96, 250, 404, 489, 220]. Islam [3650, 1018]. Islamic [2233, 3764, 1345, 745]. islamique [3764]. Islamiques [3764]. Islandici [2326]. Isledovanija [591]. Isles [1261]. Isobel [3899]. isochronisme [3352]. Isomorphismus [3214]. Israel [4189, 499]. Issekutz [2130]. issued [3903, 2942, 608]. issus [2411]. Istit [3813]. Istituto [989, 3438, 4068, 528]. Istoria [2385]. Istoricul [959]. istorii [2548, 2453]. Istoriko [591]. Istoriko-matematitchesikja [591]. Istoriografiya [983]. Italia [989, 4270, 3923, 1567]. Italian [3727, 2373, 531, 4301]. Italiana [830, 943, 3901, 3941, 4270, 1503]. italiani [4351]. italiano [2077]. Italie [213, 500, 686, 722, 934, 3983, 3402, 1207, 2798, 625, 1741, 1747, 4062, 1784]. Italien [876]. italienne [1972]. Italy [2694, 1266]. Itard [2464, 1295, 1329, 1629, 1932, 986, 1073, 1326]. Ithaca [1206]. Ithaca-Philadelphie [1206]. itinéraire [3549, 2461]. Itinéraires [1117]. Iu [606]. IV [1267, 1265, 213, 2605, 1509, 1946, 3122, 1614, 3240, 1565, 987, 1387, 1671, 2057, 2093, 626, 4642, 2693, 2705, 1346, 3962, 4146, 4147, 944, 2185, 417, 4209, 2669, 383, 1077, 2156, 2855, 334, 1266, 1861, 4549, 496, 1402, 643, 366, 971, 1829, 1565, 886, 2737, 2583, 1469]. IV-1 [1565]. IV-2 [1565]. Ivan [3921]. Ivanovitch [3972]. Ivo [4204]. Ivor [3939, 3990, 2025, 2155]. IX [4409, 279, 843, 1486, 3240, 3214, 1689, 2828, 1568, 3218, 3725, 3924, 2959, 4493, 4023, 1324, 2808, 3530, 634, 897, 588]. Izaac [2735]. Izbrannye [601]. Izdatel’stvo [4203]. Izquierdo [885, 33].

J [89, 1931, 3663, 4433, 1144, 2069, 2320, 2631, 2632, 746, 478, 4017, 1633, 3340, 2878, 7, 35, 58, 1025, 2544, 2627, 2803, 3504, 4003, 2189, 2343, 2310, 3207, 2531, 906, 845, 1829, 1979, 1459, 1576, 3433, 3271, 3436, 3899, 1590, 3083, 3169, 3945, 3153, 275, 327, 330, 575, 3669, 4393, 2662, 302, 513, 3100, 2608, 3557, 3370, 4053, 3014, 3733, 4008, 4478, 2319, 2576, 3061, 3455, 2063, 2565, 3337, 3154, 3362, 3503, 3584, 1978, 999, 3211, 1992, 2030, 2266, 2473, 1611, 665, 675, 691, 796, 881, 1059, 1148, 1170, 1245, 1569, 1613, 1703, 1708]. J [1755, 1756, 1834, 2579, 2644, 1997, 2974, 68, 567, 628, 986, 1544, 1668, 1671, 1791, 614, 991, 1263, 3008, 947, 1940, 3664, 2999, 2508, 3149, 4269, 1936, 3420, 3756, 1787, 2835, 3074, 4643, 3439, 2231, 3699, 1017, 1472, 3412, 937, 990, 2813, 4294, 3657, 4204, 1894, 2419, 4348, 1101, 3082, 3603, 2702, 2749, 2545, 3308, 875, 1775, 1289, 2408, 2611, 2744, 484, 3167, 1998, 2185, 3497, 2869, 3722, 2155, 1295, 1865, 1969, 118, 133, 207, 372, 566, 856, 1105, 1293, 1294, 1542, 1543, 2311, 3041, 3658, 878, 1107, 2965, 2635, 2757, 1110, 2807, 4581, 244]. J [3865, 2242, 2521, 2737, 190, 2659, 2712, 2788]. J. [78, 2622, 3198, 2179, 2367, 1781, 364, 519, 3688, 1321, 2, 2197, 660, 2012, 2414, 611, 3899, 1059]. J.-A. [2]. J.-B [2611]. J.-B. [2179, 519]. J.-B.-S [327]. J.-C [4003, 3439]. J.-Cl [1708]. J.-F [4433, 1459, 2813]. J.-H [2508, 2408]. J.-H. [1781, 2197]. J.-J [3308]. J.-J. [3688]. J.-L [2631, 1978, 1756, 1544, 3412, 3497, 1865]. J.-M [1245]. J.-Matthias [4204]. J.-Moulin-Lyon [3722]. J.-P [2627, 3433, 2319, 1059, 1703, 4643, 2744, 1294, 2311]. J.-R [2631]. J.-V.-F. [1321]. J.-Y [3504]. Jâhiz. [4614]. Jaakko [3866, 2333]. Jaap [3637]. jaar [513]. Jabir [3650]. Jacalyn [3926]. Jacinto [529]. Jackson [2886, 4355]. Jacob [1863, 3589, 3150, 2030, 2999, 3149, 4437, 3589, 4414, 4528, 240]. Jacobean [1019]. Jacobi [4438]. Jacoboeus [3755]. Jacopo [951]. Jacquart [3884, 2669, 2754, 3867, 3837]. Jacqueline [4546, 4287, 2677]. Jacquemont [1049, 1023]. Jacques [3972, 2503, 3631, 2308, 2409, 3483, 2314, 3971, 652, 3487, 1728, 3153, 3152, 1850, 3515, 3233, 4174, 2667, 4628, 3672, 2573, 2312, 3054, 3462, 2577, 4494, 1337, 1338, 2539, 4231, 1786, 4629, 3742, 3883, 4146, 2060, 2648, 3407, 1437, 2404, 2825, 3261, 2732, 1730, 283, 3151, 3596, 1515, 3687, 1839, 4471, 597, 1064, 1950, 2204, 2327, 3659, 3173, 3297, 4596, 1729, 2033]. Jacques-Louis [2033]. Jacques-Philippe-Raymond [597]. Jacquot [2166, 2911]. Jaen [450]. Jaffe [2472]. Jahier [751]. Jahn [3536, 2841]. Jahre [2034]. Jahren [4053]. Jahrgang [1709]. Jahrgänge [1945]. Jahrhundert [2536, 1526, 1890, 278, 2663, 2636, 2031, 1332]. Jahrhunderts [1795, 3316, 769, 3734]. jährige [2479]. Jahrtausenden [1680]. Jaisson [3730]. Jaki [2546, 2244]. Jakob [1928, 3222, 4551, 188]. Jakov [2184]. Jal [2114]. Jallabert [2900]. James [89, 3752, 3506, 156, 3410, 3792, 3991, 3838, 999, 3381, 1844, 3835, 2918, 216, 1268, 3527, 608, 1339, 3262, 2159]. Jami [3533]. Jamin [2538]. Jan [3411, 3701, 2470, 4056, 1496, 3218, 3970, 2076, 3699, 3609, 328, 947]. Jane [3039]. Janet [3748, 3060]. Janine [4398, 1304]. Janis [3168]. János [4348]. Janson [589]. Janssen [4280]. January [2834]. Janus [3218]. janv [354, 192]. janv.-juin [354]. Janvier [1232, 1032, 2827, 2324]. Janvier-juillet [1232]. Jaouiche [2354, 3147]. Japan [3381, 4655, 4656, 1297, 4167]. Japanese [4654]. Japon [1174, 1615, 4655, 4656]. japonaise [3174, 2035, 2187, 1701]. japonaises [3729]. Jaquel [2407]. Jardin [2852, 3161, 2598]. Jardine [3838, 3120]. Jaren [1482]. Jarrige [4646]. Jarry [3922]. Jas [4490, 4169]. Jasmin [2339]. Jason [1309]. Jaubert [777]. jaugeage [4465]. Jaume [3078]. Jauregg [3766]. Jaussaud [3969]. Javier [2387]. Jay [4236, 2447]. Jayawardene [2866]. Jaynes [2923]. je [1214, 140, 4032, 1613, 242, 1327, 1328, 1634, 3007, 1352, 2736, 2756, 1326, 1932, 363]. Jean [3972, 241, 3631, 4142, 3591, 2976, 3900, 795, 3336, 135, 158, 3700, 3793, 4233, 4475, 2783, 10, 29, 1102, 157, 3504, 3882, 4003, 4615, 4367, 4616, 3178, 3207, 3703, 2507, 753, 968, 1434, 4411, 956, 452, 1349, 1736, 3121, 4645, 3153, 280, 630, 1096, 3348, 3832, 3385, 948, 4438, 3515, 4022, 2770, 3302, 3233, 4174, 3097, 4008, 4018, 4267, 1909, 2575, 4019, 3309, 2502, 4516, 3841, 4094, 4091, 2900, 3242, 3279, 3342, 3054, 2097, 2660, 2884, 3456, 4451, 3675, 1303, 1443, 1494, 1701, 1704, 1707, 2068, 3970, 4625, 379, 4034, 4120, 4532, 3795, 2782, 4080, 4084, 2538]. Jean [4231, 3299, 1786, 4139, 2833, 4454, 4319, 4643, 3844, 3412, 3606, 2851, 3612, 4510, 3742, 1346, 3194, 4363, 3577, 4618, 3884, 4054, 3407, 1058, 2213, 2824, 3261, 3056, 2108, 4162, 3722, 3038, 4164, 2386, 3633, 2669, 130, 1073, 1326, 1329, 1570, 1629, 1932, 2407, 454, 1024, 1026, 1167, 4011, 4169, 4395, 3415, 3665, 1260, 2613, 4433, 4450, 4477, 3031, 3027, 2911, 3026, 3349, 2464, 2382, 1902, 4387, 2588, 702, 2019, 3265, 2597, 2381, 3498, 3662, 4252, 2730, 4680, 3779, 4625, 2515, 1132, 3625, 468, 2237, 2364, 924, 3856, 1901, 1186, 1187, 4596, 4552, 4382, 2020]. Jean [3783, 387, 1900, 2774, 3152, 3273, 3704, 3213, 1957]. Jean-Baptiste [4475, 2108, 1186, 1187, 4552]. Jean-Claude [3793, 3178, 3207, 3515, 4019, 4516, 3279, 3342, 3970, 4084, 4618, 2213, 3056, 3722, 2588, 2020]. Jean-Daniel [4615]. Jean-Dominique [2851]. Jean-Fabien [3703]. Jean-Firmin [4252]. Jean-Francois [4645, 4094, 4451, 3844, 3633, 4433, 4450, 3349]. Jean-Henri [4008, 2407, 3265, 3213]. Jean-Jacques [4174, 3407, 2404, 4596]. Jean-Louis [3882, 1736, 3153, 3385, 4022, 1443, 3412, 3742, 2404, 3415, 3498]. Jean-Luc [2502, 4454, 4169, 4395]. Jean-Marc [3261]. Jean-Marie [2782, 2673]. Jean-Michel [2976, 4233, 1701]. Jean-Nicolas [3054]. Jean-Noël [4054, 4477]. Jean-Palerne [3504, 4003, 4363]. Jean-Patrice [3884]. Jean-Paul [3972, 4261, 3302, 4267, 3841, 4091, 2660, 1327, 2538, 4387]. Jean-Pierre [1434, 2770, 4080, 4139, 4643, 3612, 3577]. Jean-Victor [2575].

Jean-Yves [3504, 4003, 4367]. Jeanne [3833, 2775, 2294, 3318]. Jeannine [1706, 1473]. Jeans [156]. Jed [3696, 4197]. Jeff [4317]. Jefferson [2641, 3732]. Jeffrey [3898]. Jenicek [1417]. Jenkins [32]. Jenner [2450, 955]. Jensen [4079, 4170, 1109]. Jérôme [4487, 4062, 4487]. Jeremy [4348]. Jerne [3360]. Jerome [420]. Jerónimo [3994]. Jérusalem [4441]. Jérusalem. [4477]. Jerzy [2524]. Jesper [2928]. jésuite [4119, 4096]. Jesuiten [3429]. Jésuites [2131, 547, 550, 4017, 3429, 3071]. Jesuits [4652]. Jetter [2535, 2885, 2671]. jetzt [1864]. Jeu [825]. jeune [3549, 4590, 281]. jeunesse [34, 1881]. jeux [2404]. Jewish [3652, 3521]. Jiang [3880]. Jim [3525]. Jing [3175]. Jo [3094]. Joachim [4097, 4175, 2467, 2770, 1788, 2473, 2695, 4593, 3203, 2519]. Joan [3863, 4436, 3480, 3580]. João [529]. Joaquim [529]. Joaquin [885, 33]. job [3863]. Jody [3417]. Joël [3336, 4009]. Joëlle [3884]. Joh [1097]. Johan [1150, 1517, 13]. Johann [3222, 826, 2307, 4551, 2364, 2283, 3529]. Johanne [3316]. Johannes [4097, 4175, 3566, 730, 1793, 2869, 2392, 4006, 4198]. Johannesburg [1822]. John [89, 3270, 2905, 1973, 2526, 4626, 3506, 2530, 2831, 1752, 3826, 2262, 1974, 3501, 2248, 3992, 2388, 455, 4024, 3749, 3531, 2581, 2641, 2420, 800, 1668, 1263, 1936, 3206, 3016, 3037, 3500, 4290, 3434, 3728, 2158, 3196, 4172, 3861, 4210, 3367, 3041, 2703, 2740, 3006, 1563, 1949, 2207, 3645, 1339, 947, 4137, 4578, 1754, 1499, 2024]. Johns [1933]. Johnson [89]. Joint [1827]. Jole [3626]. Joliot [3091, 3686, 3091]. Joly [1113]. Jombert [4647]. Jonathan [4050, 3524]. Jonckheere [577]. Jones [89, 2261, 3754, 2166, 3934]. Jordan [1420, 1262, 2729]. Jordani [1216]. Jordanus [689, 2733, 3516]. Jörg [4095, 3276]. Jorge [2815]. Jos [1097, 788]. Jos. [278]. Jos.-E [788]. Jose [33, 1778, 4458, 885, 3235, 4255, 1628, 1907, 3627, 4363, 3735, 2926, 3732, 3310, 566, 1324]. Josef [3148, 3259, 394, 1070, 1298, 1564, 3427]. Joseph [89, 2337, 2834, 1923, 2183, 2131, 1909, 1956, 2840, 2479, 1565, 2645, 2785, 1544, 3074, 3080, 1235, 3177, 787, 1325, 2317, 2665, 2390, 2468, 2859, 1158, 1766, 2585, 1050, 1904, 2727, 2329, 1529, 77, 2071, 2760, 872, 2155]. Joseph-Francois [1529]. Joseph-Marie [1050]. Joseph-Nicolas [2727, 2071]. Josephinismus [1438]. Josette [887, 1230, 1521, 1846]. Josiah [2452, 2452]. Josua [161]. Jouan [788]. Jouanna [4146, 3483, 3883]. Jouffroy [4529, 2902]. Joule [2461]. jour [3310, 752]. Jourdin [2962]. Journal [1594, 164, 2275, 1004, 1511, 1797, 1912, 1946, 2039, 2116, 559, 673, 1421, 480, 2712, 2788, 1486, 2276, 2041, 2042, 2117, 2133, 456, 1707, 1045, 3596]. journalisme [1245]. journals [3666, 1499]. journaux [3456]. Journée [3157, 1131, 3460]. Journées [1808]. jours [435, 4297, 958, 2035, 2494, 4165, 1757, 2755, 2187, 3502]. Jousse [4387]. Jovan [3009, 2568]. jovis [3566]. Jozefa [2252]. Jr [512]. Jr. [1888]. Ju [1332, 2355]. Juan [4458, 4173, 3627, 904]. Jubiläumskongress [2479]. Judith [2322, 4620]. juill [192]. juill. [353]. juill.-sept. [353]. juillet [795, 1843, 461, 2319, 1232, 4366, 250, 1014, 1079, 1460]. juillet-août [795]. juin [3700, 4003, 2189, 459, 354, 1747, 1843, 4063, 3043, 3664, 3157, 3221, 2377, 3633, 333, 1490, 1079, 1460]. juin-juillet [1079]. juives [4441]. Jules [139, 1218, 1478, 2387, 1820, 1465, 4292]. Julia [2915]. Juliana [1339]. Julie [4624, 4010]. Julien [1351]. Juliette [792]. Julius [3766]. Jullien [4071, 3811]. July [1432, 3195, 489]. June [2837, 404]. Jung [3642]. Jungmayr [3316]. Jupiter [3566]. Jürgen [2612, 3663, 3223, 3443]. Jurgus [825]. juridiques [3449]. jurisprudence [3860]. Juschkewitsch [1540]. Juskevic [2662, 1332, 2355]. jusqu’à [958, 3194, 208]. jusqu’au [748, 1456, 643, 1000, 2394]. jusqu’aux [3713]. Just [3222, 3716, 3715, 3708, 3712, 3709, 3714]. Juste [4415]. justice [3967]. Justification [2332, 761]. Justin [3012]. Justus [3218, 2987]. Juvet [860]. Jwarizmi [1436].

K.-R [2384]. Kabbale [74]. Kaden [3505]. Kagan [3937]. Kahan [3007]. Kahane [3433]. Kahn [3648, 3648]. Kaiserzeit [3505]. Kaluza [3885]. Kameshwar [3361]. Kandel [4480]. Kant [3121, 4098, 3891, 2820, 869, 4233]. Kany [4363]. Kany-Turpin [4363]. k’ao [1390]. Kardel [3755]. Karin [2566]. Karine [3525]. Karl [3676, 506, 3370, 4055, 1350, 1734, 674, 3667, 2701, 3343, 4144]. Karli [2666]. Karten [4204]. Kartografii [2252]. Kasimir [3731]. Kasner [372]. Kastler [2292]. Katahiro [3729]. Katherine [3005, 2853]. Kathleen [3880]. Kathrin [4243]. Kathryn [3441]. Katisik [1388]. Katja [3731]. Kaucher [4647]. Kauffman [278, 2476]. Kay [3590]. Kayas [2544, 2544]. Kearney [2225]. Keating [3591]. Keckić [3009, 2568]. Keel [4237]. keeping [2993]. Keepsakes [2942]. Keferstein [2124]. Keil [2843]. Keilin [1636]. Keith [2403]. Kekulé [991]. Keller [4437, 4346]. Kelvin [3243]. Kemiens [2475]. Kempf [2580]. Kenneth [89, 2191, 2078, 2261, 3526, 1499, 2886, 2963, 3211, 3605, 4294, 4440]. Kepler [4314, 3566, 2869, 3120, 2307, 3038, 3258, 1543, 2335, 3136, 645, 2684, 1664, 1543]. Keplero [4344]. Kerisel [3194, 3194]. Kersaint [1687]. Kerszberg [3121]. Kessen [970]. Ketzerei [3650]. Kevin [3282]. Kevles [3756]. Key [2564, 3702]. Keynes [3455, 1635, 460]. Keys [1505, 1481]. Kh [2877]. Khāzin [2454]. Khalil [3147]. Khamsone [1310]. Khazin [4023]. Kholodilin [2749]. Khwarizmi [3042, 4560]. K’i [666, 883, 883, 1390]. Ki-Tsiun [1390]. Kiao [666]. Kielland [824]. Kildebæk [4514]. Kilgour [695]. Kilmister [3702]. Kim [4316]. Kindi [2684]. Kinematics [2094, 2694]. King [4238]. Kinnon [2749]. Kiple [3526]. Kirby [695]. Kirkbride [2993]. Kirtley [1779]. Kirwan [3930]. Kisch [610]. Kisin [3610]. Kitāb [2372]. Kitâb [2570]. Kitab [3650]. Kl [278]. Klado [2355]. Klasse [2392, 1028]. Klassiker [2384, 2058, 2785]. klassischen [3416]. Klaus [2647, 3059, 4204, 2710]. Klein [3663, 4477, 2666, 4622, 3865, 1863, 2636, 4441]. Kleine [1788, 1413]. Kleinert [2663]. Klickstein [400]. Klinckowstroem [1887]. Kline [2776]. Klinik [2034]. Klüver [2612]. Knight [2640]. knjiga [1105]. Knobloch [4204, 4125, 2661]. Knoefel [2784, 2978]. Knorr [2228, 3303]. Knowledge [4679, 3868, 4675, 4633, 3001, 4320]. Knowles [1694, 1982, 1471, 1981]. known [1847]. Koch [64, 3693, 1109]. Köcher [1440]. Kochina [2610]. Koecher [1440]. Köhler [4056]. Kohn [3060]. Kollros [479, 188]. Kolmogorov [3839, 4020, 4203, 2469]. Kölving [4530, 4265]. Kommentar [3757, 1795, 2287, 3443]. Kommission [942, 2392]. Kompassmacher [1693]. Königsberg [3441]. Koning [514]. Koninklijke [397, 991]. konkret [3416]. konkret-anschaulichen [3416]. Konopka [884]. Konrad [2841]. konst [747]. Kontext [2603]. Kontingenz [4097, 4175]. kontroversen [3444]. Konvitz [3259]. Kopaczewski [63]. Kopelevic [2355]. Kopernikowska [857]. Kopernikus [1234]. Kopf [246]. Kopolevic [1332]. Köpp [2670]. Kopparberg [748]. kopparhantering [748]. Kopparhanteringen [748]. Kordig [2332]. Kornberg [3461]. Korrespondenz [4243]. Korrespondenzen [2355]. Kosmologische [1680]. Kosmos [3734]. Kosmos-Modells [3734]. Kostant [3885]. Kostas [3700, 3789]. Kou [606]. K’ouo [666, 773]. Kouznetsov [1580]. Kovalevskaya [2610, 3053]. Kovalevsky [1009]. Kowalevskaya [2610]. Kowalewskaja [3667]. Kox [3663, 3865]. Koyré [585, 1288, 1373, 1396, 1393, 785, 1536, 1607, 3162, 1395, 1542]. Kraepelin [4661]. Krafft [2451]. Kraft [1733]. Kragelsky [1610]. Kragh [4079, 3369]. Kramer [2924]. Krankenbehandlung [3416]. Krankengeschichten [3757]. Krankenhausgeschichte [2535]. Kranksein [2479]. Kranzberg [1888]. Kraus [3650]. Krayer [3429]. Kremer [4171]. Kremer-Marietti [4171]. Krieglstein [3505]. Krige [3992]. Krisciunas [3282]. kristallographischen [3222]. Kritik [2612]. Kritische [2287]. kritischer [3316]. Krivomazov [2510]. Krohn [2522]. Kroll [3484]. Kronland [1655]. Krüger [3434]. Krysicki [1864]. Krystalle [3222]. Kubbinga [4295]. Kucharski [524]. Kudlien [3362]. Kues [3346]. Kuhn [2528, 3148, 3536, 938, 3954]. Kultur [2479, 1332]. Kumaniecki [939]. Kümmel [2478]. kungl [12]. Kunitzsch [1100]. Kunst [2778, 1262, 1422]. Kunst- [1262]. kunsten [991]. künstlerisches [2778]. Kunstuhrmacher [1935]. Kuriyama [4167]. Kurt [2423, 2479, 1777, 1795, 1323, 2829, 2985]. Kurt-R [1323]. Kurzbiographien [2841]. Kurze [482]. Kushyar [1566]. Kützing [1111]. KVA [747]. Kwartalnik [971]. KYA [747].

L [89, 3516, 3562, 2069, 2631, 2657, 3903, 2262, 1345, 1025, 3592, 485, 3703, 2664, 729, 958, 960, 2546, 3528, 327, 2662, 3042, 237, 397, 1754, 4230, 3276, 2837, 3279, 4082, 1978, 2321, 3363, 1170, 1569, 1756, 1848, 2067, 2234, 2579, 2644, 653, 1163, 1544, 1681, 1792, 1794, 2921, 2539, 4178, 3814, 3605, 3412, 992, 1235, 2244, 2371, 3728, 1414, 2126, 1244, 745, 2702, 2804, 4006, 4391, 2408, 479, 484, 510, 946, 2171, 3497, 4440, 3827, 1865, 1999, 362, 118, 133, 188, 627, 903, 1332, 2874, 2112, 246, 2390, 2635, 4198, 951, 1301, 1615, 1957]. L [2224, 2614, 2437, 3115, 2585, 3292, 777]. L. [3100, 4581]. l’être [234]. L’A.F.A.S. [501]. Labaste [4624]. Labat [419]. Labban [1566]. L’abbé [621, 2574]. l’Abeille [2649]. Laberrigue [3318]. Laberrigue-Frolow [3318]. labo [3591]. Laboratoire [1524, 4343, 3955, 4637]. laboratoires [1524, 1319, 4045]. Laboratory [730, 3481, 3319, 4637]. Laborde [1472]. Labosne [986]. Labrousse [2491]. l’absurde [3385]. labyrinthe [4532]. L’Académie [659, 892, 289, 142, 1535, 1376, 848, 1599, 2870, 1488, 1375, 4386, 4645, 2637, 758, 3521, 314, 2603, 4389, 3651, 1762, 756, 2177, 3828, 1374, 1200, 3657, 4285, 1739, 682, 4565, 1319, 2386, 82, 160, 246, 247, 891, 1923]. Lacaille [3660]. Lacassagne [4156]. Lacaze [4637]. Lacaze-Duthiers [4637]. L’accélération [137]. Lacepède [2138]. lachender [3488]. Lacki [3970]. Laclau [3577]. Laclos [4476]. Lacoin [1382]. Lacroix [1088, 2815, 520, 849, 870]. L’acte [20]. L’activité [376]. Ladina [4256]. Ladislao [1942]. Ladislav [1417]. L’admission [1869]. Ladrière [3606]. Ladurie [3259, 1695]. Laennec [2072, 3926, 382]. L’Affaire [3098, 2290]. Laforcade [2574]. Lafuente [3592]. Lagrange [1544, 194, 3136, 64, 612, 3891, 2084, 3093, 2662]. Lagrée [3993]. L’agriculture [2273]. l’agronomie [2175]. Laharie [3462]. Laird [3812]. Laissus [2318, 1766]. Lakatos [3703]. Lalande [4487, 4544]. L’Alchimie [4286, 4081]. Lallemand [2162]. Lalouère [1188]. Lamar [789]. Lamarck [3260, 2318, 2319, 3058, 2511, 2786, 2562, 2002, 2049, 2798, 2862, 922, 2048, 2195, 4225, 921, 3057, 2786]. Lamard [4140]. Lambarde [1594]. Lambert [4008, 2508, 2408, 2407, 311, 2460, 1781, 4256]. Lambrigot [1725]. Lambros [2698]. L’ambulance [2410]. l’Amérique [948, 2238, 1132]. Laming [1313]. Laming-Emperaire [1313]. Lamontagne [2316]. Lamouroux [1321]. lamp [89]. Lamure [1677]. Lamy [4319, 4487]. Lan [2708, 3190]. Lana [4196]. l’anatomie [279]. lancées [4296]. Lancisi [4083]. Land [2154, 3861]. Landelle [1405]. Landes [2905]. Landine [2270]. Landriani [3941]. Langage [4208, 2817, 2934, 3360, 4229, 1483, 865]. Langevin [4246, 1852]. Langins [3168]. Langlès [1615]. Langmuir [3972]. Language [970, 2572, 2934, 3786, 2477, 3642]. Langue [2370, 3116, 3594, 3224, 3884]. Langue-Linguistique-Communication [2370]. Languedoc [1850]. L’animalité [4454]. Lankester [2030]. Lantéri [4664, 4208]. Lantéri-Laura [4664, 4208]. l’Antiquité [2868, 958, 1892, 2884, 4129, 3502, 3896, 3834, 1790]. l’Antiquité. [3504, 4363]. Laos [1309]. laotienne [1310]. Lapérouse [2910]. Lapicque [977]. Laplace [3939, 2837, 4495, 2541, 3137, 778, 2455, 1085, 3044, 2205, 2285, 520, 2881]. laplacienne [3140]. L’Apogée [150]. Laporte [3746]. Lapouge [2687, 1225]. laquelle [3355, 3408]. L’Arbre [4532]. l’arc [3787]. l’arc-en-ciel [3787]. L’archéologie [1313]. L’architecture [70]. Lärdomshistoriska [1150]. Largeault [3311]. l’aristotélisme [2839]. L’Arithmétique [2382, 4308]. Larre [3760]. Larrie [4475]. Laser [3580, 1252]. Lashley [4055]. L’aspect [832]. l’Assemblée [1967]. l’Association [4065]. l’astrolabe [71, 2356]. l’Astronomie [1285, 791, 791]. Laszlo [4092]. late [3815, 2694]. Lateinische [3503]. Later [1301, 3826]. Latin [4458, 1539, 1792, 905, 481, 1230, 4458, 3842]. latine [3158, 1216, 1230, 1521, 1846]. latines [3267]. latini [4069]. latino [1436]. latins [4130, 2494]. Latinus [3503, 4131]. Latitudes^{≫} [3972]. l’atome [1472, 2170]. L’atomisme [630]. Latour [3151, 3224, 4490]. Lattin [1934]. Laubat [2340]. L’aube [753]. Laubry [356]. Laufer [2126]. Laugel [1681]. Launoy [468]. Laura [1472, 2493, 4664, 4208]. laureate [1867]. Laurence [89, 4415]. Laurent [3260, 4233, 4242, 3153, 4052, 519, 3920, 1045, 1084]. Lauria [4513]. Laurie [3902, 3165]. Lausanne [160]. Lautman [3032, 3033, 3031, 3026, 3029, 3028, 1786, 3030]. Lavagne [2677]. L’Avènement [4237, 2698]. L’Aventure [4629, 3056, 2832]. lavoisiennes [3555]. Lavoisier [453, 3122, 3937, 2837, 3279, 1474, 3552, 3549, 3550, 3982, 760, 3547, 195, 778, 893, 3554, 560, 3548, 3930, 1205, 3553, 3551, 4613, 1045, 1084, 1046, 1134, 2047, 2908, 148, 3560, 1575, 587, 3785, 4312, 214, 3786, 3936, 4270]. Lavoisier. [1389]. Law [1891]. Lawrence [3386, 1848, 3588, 2373, 2521, 2737, 3868, 2242]. Lawrynowicz [3731]. Lay [4487]. Layton [2522]. Lazare [1958, 3900, 2637]. Lazzaro [1614, 3240, 2882]. Lea [89]. leaders [610, 282]. Learned [4609, 4664]. Learning [2309, 3796]. least [4467]. Lebègue [10, 1567]. Leben [3040, 3316, 278, 1336]. Leber [1689]. Lebersforchung [1689, 1689]. Lebesgue [55, 2587, 2215, 1976]. Lebon [1730]. Leboucq [10]. Lecció [3627]. Leclercq [1058, 1830, 1026]. L’Ecluse [2266]. Lecointre [4231]. L’École [4413, 124, 2238]. Lecons [135, 2541, 3709, 860, 275, 4033, 2108]. Lecourt [4094, 3408]. l’écriture [4004]. l’Écrivain [2382]. lecteur [3447, 2412, 4570]. Lecteurs [4109]. lectual [727]. Lecture [2336, 4224, 4217]. Lectures [1020, 3384, 2388, 1867, 2128, 2063, 2336, 631, 3386, 609]. Lécuyer [3151]. Lederer [2611]. L’Édification [4452]. L’édition [4688]. Ledoux [4646]. Lee [89]. Leeuwenhoek [3903, 358, 326]. Leeuwenhoek-Commission [3903]. Lefebvre [2185, 131]. Leffler [4049]. Leflon [29]. Legacy [3411]. Legay [2673]. Legée [3478]. légendaires [2267]. légende [2688, 2571]. Legendre [2455, 886, 3107, 790]. legends [3417]. legibus [4688]. Légier [486]. Legrand [1308]. Legros [4091]. Legs [3927]. Leham [2319]. Lehrbuch [533]. Lehre [4097, 4175]. Leibnitiana [2423]. Leibniz [3223, 2423, 3925, 4619, 4204, 278, 2468, 4686, 1493, 3777, 1488, 4000, 4688, 143, 1189, 1181, 1492, 4685, 3447, 4683, 4687, 4550, 3998, 1201, 3449, 4012, 4603, 3163, 4564, 875, 3999, 4223, 393, 2183, 2423, 3457, 1143, 1219]. Leibniz-Gesellschaft [2423]. leibnizienne [3448, 4014, 3450]. leibniziens [3264]. Leibnizschen [278]. Leicester [1152, 1317, 1485]. Leiden [514]. Leidensis [2249]. Leikola [3153]. Leinen [605]. Lejeune [2829]. Lékarské [960]. lekarz [794]. L’électronique. [1118]. Lelièvre [2741]. Lemaître [2157, 2416, 3844, 2157, 3844]. L’émergence [4045, 4255]. Lemery [4193]. Lemire [3463]. lemma [3497]. lemmes [2055]. L’Empire [4347]. L’Enciclopedismo [3759]. L’Encyclopédie [343, 323, 320, 1408, 345, 4664, 311, 321, 314, 315, 344, 322, 313, 480, 1337, 4120, 4390, 1850, 350, 4313]. lendemain [4405]. l’énergie [398, 1732]. l’enfant [4241]. length [3599].

Lenihan [2702]. Lennox [3752]. Lenoble [11, 851]. Lenoir [1730, 1254]. L’enseignement [1757, 1939]. L’entrée [2491]. Leo [2976, 1034, 2787, 3668, 2551, 3098]. Leodiensis [590]. Léon [66, 67, 1575, 457, 458, 10, 401, 3265, 2565, 1457, 1003, 1384, 4088]. Léonard [481, 2409, 1747, 4236, 1601, 2577, 750, 1755, 1834, 1146, 3151, 527, 1165]. Leonardo [4236, 601, 526, 528, 4633, 528]. Leonello [3383]. Leonhard [2662, 1097, 1609, 188, 1332, 2355, 4567, 3072]. Leonhardi [2873]. Leopold [2531, 691]. Leopoldina [989, 3536]. Leopoldina-Meeting [3536]. Leopoldo [2882]. Lepetit [3935]. L’épistémologie [1281, 4616, 4476]. L’épopée [2340]. lèpre [1692]. Leprince [3289, 2832, 654]. Leprince-Ringuet [2832, 654]. L’équilibre [2339]. Lerner [2770, 4643]. Leroy [860, 4395, 981]. Lesky [2645, 2785, 1028, 2536, 1526]. Lesley [4290]. Leslie [1574]. l’espace [3074]. L’esprit [3075]. Lessico [2631, 2914]. lessicografia [2914]. Lesure [1331]. L’étape [34]. L’Éternel [4438]. léthargie [1136]. Leticia [1980]. L’étrange [2751]. l’étranger [4366, 3825]. L’étrenne [2307]. Letté [4349]. Letteràrio [2941]. Letterati [2976]. Letteratura [3365]. Lettere [2976, 2358, 4088, 4076, 4083, 3674]. letteren [991]. Letters [3903, 3792, 4679, 3362, 3218, 907, 2528, 3195, 3940, 2686, 4115, 1023]. Lettre [3032, 1918, 3033, 2466, 1359, 1949, 2415, 1188, 1377, 4387, 4063, 2179, 1967, 1088, 108, 2161, 1205, 2595, 3044, 1201, 1951, 2099, 1374, 293, 1134, 83, 4472, 227, 3782, 3750]. lettre-programme [1949]. Lettres [3031, 2687, 34, 2411, 892, 1729, 275, 984, 1514, 2916, 524, 4192, 3287, 4285, 2288, 2583, 1606, 1574, 1187, 718, 4544, 571, 4562, 2256, 4251, 4543, 157, 4679, 4644, 3265, 778, 4049, 2448, 1085, 3924, 1513, 1555, 510, 4565, 2908, 297, 1557, 2587, 3596, 309, 4477, 2374, 3458, 1186, 1145, 1147]. L’étudiant [213]. Leuckart [2647]. Leuilliot [1735]. leur [3933, 4065, 4510, 3244]. l’Europa [3568]. Leurquin [3607]. leurs [323, 4553, 2599, 1375, 3064, 3713, 1215, 3052, 3163, 3234]. leven [358, 326]. Levere [3567, 2838]. Levers [2627]. Levey [1566]. Levi [3118, 3957, 152, 4341]. Levi-Montalcini [3957]. Leviathan [3192, 3654]. L’Évolution [421, 3695, 212, 3676, 135, 90, 95]. Lévy [2609, 4366, 3673, 1960]. Lévy-Bruhl [4366]. Lewin [3871]. Lewis [2532, 3364, 4578, 234, 2694]. Lexell [659]. L’exemple [3142]. L’exercice [1756]. Lexikon [1070, 1298, 1564]. Lexique [3884]. L’expérience [2902, 1751]. L’expérimentation [3834]. L’Exploration [4618]. L’Exposition [1498]. Leyde [841, 1489, 3428, 1221, 841]. Leyden [359]. Leyendo [152]. lezioni [2375]. L’^{≪} [1949, 2400, 446]. L’Hôpital [1377, 2812]. l’Hôtel [1115, 1175, 1938]. L’hérédité [1477]. Lherminier [4614]. L’histoire [1472, 694, 2170, 4140, 4433, 4032, 4295, 3349, 1611, 1695, 2210]. l’Homme [2353, 2649, 212]. L’Huillier [3884]. L’humanisme [1079]. l’Humanité [421]. L’hypothèse [2157, 3995]. Li [3175, 3206, 3190, 3173, 992, 1227]. Liaigre [1729]. liaison [2348]. Lian [796]. Liba [3604]. libanaise [878]. libe [939]. Libelles [669]. Liber [2706, 3316, 4286, 389]. liberal [2965]. liberté [3950, 507]. libertins [1479]. Libois [3460]. libraire [530]. libraires [4647, 3215]. Librairie [485, 3167, 3168, 527, 669]. Library [1931, 3866, 1973, 2332, 2530, 2942, 3220, 235, 2581, 2225, 4348, 531, 1942, 4557, 4018, 4601, 2529, 1933]. Libre [4319, 135, 2086, 3938]. libres [1375, 3761]. Libreto [3039]. Libri [2801, 3505, 2991]. libris [3785]. Libro [1542, 3994, 3039]. librorum [3143]. librum [2287]. Licht [966]. Lichtenthaeler [3757, 749, 3314]. Lichts [2701]. Lick [3262]. Liebhaber [1262]. Liebig [2987, 3218, 1380]. Lieferung [1564, 1298, 1070]. Liège [774, 3460, 2288, 1915]. liegen [3201]. Liégeois [590]. liens [737, 4048]. lieu [3648, 1230, 1521, 1846, 4665]. lieux [3883, 4414, 4528]. Life [1144, 3579, 2248, 3517, 3598, 508, 2063, 3431, 1476, 1635, 859, 3192, 1998, 2185, 531, 2551, 2659, 3939, 3440, 1752, 3863, 3826, 3945, 4679, 4675, 3755, 3885, 3590, 3482, 3640, 3629, 2977, 3871, 2155, 2978, 3926, 420, 3434, 1590]. lifting [2233]. lifts [89]. light [4627, 3476, 3464]. ligne [4571, 3451]. Ligue [2853]. Liliane [3460, 1638]. Lille [1586]. Lillian [3902, 3165]. Lily [3590]. l’image [592]. L’imagerie [2253]. limaille [3402, 2221]. limite [1857]. Limites [4030, 4559, 3943, 1720, 3924, 678]. Limmer [3505]. Limoges [4642, 2339]. Lind [951, 2321, 3363, 1848]. Lindberg [2684, 876, 2692, 3126, 2024, 2158, 2216]. l’Inde [380]. Lindeboom [1022]. Lindroth [748, 477]. l’industrialisation [3989]. l’industrie [4349, 1352]. Line [4415, 4571]. linéaires [3331, 472, 1683, 4398]. linéaments [178]. Lineriis [1793]. Linet [2677]. l’infini [4001, 702]. Ling [1017, 665, 1565]. l’ingénieur [4189]. Linguae [3148]. Linguistique [2370]. Linguitti [2739]. Linnaeus [836]. Linné [2339, 1241, 3563]. linnéenne [3266]. Linneo [2643]. l’Institut [1808, 1242]. l’intellect [4528]. L’Intelligence [4395, 1787]. L’introduction [2251, 1182]. l’Intuitionnisme [2312]. L’inventeur [1416]. L’Invention [4367, 3935, 45, 2822]. Lion [2491]. Lionardo [417]. Lionnais [283, 132, 372, 855]. Liouville [2859, 2760]. Lipschitz [2596]. liqueur [4586]. Liqueurs [1437, 4586]. liqueurs^{≫} [121]. Lire [4173]. Lisa [3580]. Lise [3534, 4640, 1391, 3871, 4172]. list [667]. Liste [2949, 2930]. Lister [2143, 1109]. Literar [2695]. Literar- [2695]. literary [4317]. Literatur [1470, 278, 1060]. Literature [1266, 1913, 2115, 1933, 1246, 3550, 2190, 1484, 4230, 1780, 1019]. littéraire [3563, 4397, 4328, 353]. littéraires [698, 4196, 4036]. Litterarum [842]. Littérature [4205, 4173, 2651, 1692, 677, 4319, 3837, 655]. little [4669]. Littré [2591]. LIV [3482, 3994, 3879]. Livage [4628]. lives [1752, 3940, 3417, 767]. Livinec [1706]. living [3482, 3640, 729]. livraisons [2265]. Livre [4614, 3648, 3607, 2611, 2669, 588, 2624, 2689, 4655, 3296, 3493, 751, 1997, 151, 481, 2055, 2137, 2570, 4147, 944, 1304]. Livres [3725, 754, 222, 4255, 2862, 4601, 3070, 3736, 1329]. LIX [4288]. Lizak [2065]. Llasera [2925, 4205]. Lloyd [3608, 3608]. Lobacevskij [3829]. local [3497]. locale [3266, 752]. Locali [1021]. localisations [4355]. Locke [1499, 3198, 4229]. locomotion [3113]. locomotive [3261]. Locquin [3831].

L’Odyssée [2832]. L’œuvre [61, 355, 374, 479]. log [1633]. Logan [609, 631]. logarithmes [1608, 2056]. logic [3815, 4580, 4002, 3990, 4203, 3482, 2477, 4578, 4582, 4581, 4583, 4365]. Logica [830, 3105, 1074]. logical [4056, 1388]. logicisme [4580]. Logik [2740]. Logique [1533, 1861, 2722, 2469, 1433, 3777, 2331, 4623, 4583, 4000, 3029, 4428, 4545, 4578, 3357, 2404, 3856, 3894, 3893, 4581, 3336, 3609]. logiques [4578, 241, 1786]. logo [1537]. logoi [4401]. Logos [984, 2054, 822]. Lohrmann [3668]. Loi [4495, 2022, 3355, 3356, 51, 595, 2600, 3928, 2688, 1993]. Loïc [4052]. lois [4410, 4111, 4014, 4550, 645, 2303, 869]. Lola [4206]. Lomonosov [478]. Lomonossov [1852]. London [3003, 1752, 2452, 1685, 3868, 3789, 2855, 2733]. Londres [1860, 4393, 1498, 224, 701]. Long [631]. longitude [3929]. Longitudes [1334, 3212, 4017, 4468, 2997]. Longoburgo [2077]. longueur [3355]. look [3500]. Lopez [3235, 2644, 2965, 3735, 2926, 3732, 4072, 3735]. L’ophtalmologie [996]. L’Optique [988]. Lóránd [434]. Lorch [2030]. lore [589]. Loren [3565]. Lorenz [3434, 3416]. Lorenzo [2387]. L’Organisation [3882]. Lorgna [3568]. Lorient [1579]. Loris [1029]. Lorne [4438]. Lorraine [3281, 3434, 1359]. lors [3530]. Losada [154]. Loss [2776]. losses [744]. Lost [3420]. Lot [35]. Löther [2841]. Loty [4052]. louen [883]. Louis [165, 135, 4242, 72, 2683, 3882, 4351, 2664, 448, 1736, 3153, 391, 3385, 4022, 2218, 796, 1061, 1443, 1480, 1995, 2065, 2067, 2340, 324, 3965, 2247, 3412, 981, 3742, 3052, 2404, 2832, 284, 654, 2755, 2446, 2704, 3415, 2298, 1819, 977, 2088, 1886, 3660, 3498, 2498, 4545, 2680, 469, 1991, 2100, 4155, 2087, 2939, 4617, 2316, 2376, 1674, 2033]. Louis-Hubert [2067]. Louise [4546, 3752, 4196]. Loup [3648]. Louvain [2582, 3924]. Louvain-la-Neuve [2582]. louve [2734]. Louyat [2927]. Lovaniensia [4510]. Love [3461, 3579]. Loveland [4317]. Lowell [2389]. Lower [2963]. Lowitz [1336]. Löwy [3869, 3317, 4490]. l’U.N.E.S.C.O. [206]. Lubos [2214, 2394]. Luc [391, 2502, 4454, 4169, 4395, 2597]. Lucas [1295, 987]. Luce [3703, 1852, 449]. Luciano [4351]. Lucie [3753, 2272, 2758]. Lucien [10, 1729, 4398, 187, 1205]. Lucienne [2215, 832]. Lucrèce [4062]. Lucrèce. [2443]. ludis [4301]. Ludovico [3357, 2567]. Ludwig [3505, 2835, 4490, 3739, 3826, 3384, 605]. Ludwik [3317]. Lueurs [1724]. luftsyra [747]. Lugo [877]. Lui [4658]. Luig [4204]. Luigi [2629, 4435, 1148, 3383, 2739, 2808]. Luis [3735]. Luisa [2344, 989, 2690, 1469]. Luiz [4232]. Lukasiewicz [2076]. Luke [3945]. lumi [4309]. Lumière [3711, 4075, 2282, 4334, 4504, 4174, 4008, 3779, 645, 2833, 1423, 4452, 3803, 2102, 2408, 2752, 1304, 4333, 3302, 1675, 3098, 2504, 792, 2376, 3802]. Lumières [4390, 4619, 2844, 4647, 4240, 4538, 3288, 4470, 3621, 4282, 4265, 4390, 4562, 2711, 2573, 3670, 4454, 4066, 3070, 3379, 4647, 3571, 4410, 4532, 4141]. Luminescence [771]. lumineuse [2304, 3427]. lumineux [4452]. Lun [3206]. Luna [2695, 4643]. lunar [1664]. lunette [4392]. l’Univers [187, 70, 3557]. l’Université [135, 436, 3161, 2582, 2324, 217, 2288, 1231, 878, 4441]. luoghi [2914]. lus [2972]. Lussac [253, 401, 2549]. lutte [838]. Lützen [2928]. Luxembourg [501, 759]. Luxembourgeois [472]. Luz [3735, 136, 88, 245]. Luzio [989]. LVI [1893]. LVII [1538]. Lwoff [2667, 3151]. LXI [3637]. LXIII [948]. LXIV [795, 3640]. LXIX [2976, 354]. LXVIII [4390]. LXXXIX [1096]. LXXXV [1148]. LXXXVI [1176]. lycée [4195, 4249, 3825, 760]. lycées [4259]. Lychnos [2475, 515, 616, 801, 1150]. Lychnos-Bibliotek [2475]. Lycie [151]. Lynx [2580]. Lyon [280, 838, 2902, 3722, 4669, 1223]. Lyonet [1169]. lyonnais [3948, 2020]. lyonnaise [2902]. Lyons [1301].

M [3464, 3562, 3485, 279, 926, 1545, 3592, 2826, 2189, 2276, 2343, 611, 3384, 2078, 1152, 1317, 1477, 1485, 1946, 1913, 2041, 2115, 2117, 3433, 3441, 3792, 3991, 2261, 3000, 3153, 327, 456, 514, 694, 3122, 4008, 2576, 3455, 3337, 2640, 3105, 3531, 999, 1162, 2217, 2030, 2883, 107, 799, 1116, 1245, 2067, 2644, 2354, 1573, 1670, 1790, 1955, 1929, 2096, 3218, 3523, 2962, 3754, 3796, 1954, 2668, 2916, 3002, 4643, 2992, 1472, 1473, 215, 4055, 945, 3165, 2436, 2926, 2702, 2749, 2408, 2611, 2108, 2264, 380, 2869, 530, 566, 969, 1290, 1293]. M [2874, 952, 1024, 1691, 2580, 2882, 3368, 2965, 2211, 329, 1301, 1246, 2224, 2551, 2659]. M. [702, 4380, 2879, 2965]. M.-A [2962]. M.-F [694]. M.-M [2962]. M.-P [2916, 4643]. M.D [953]. mīkānīkiyya’l [2372]. même [3145, 4133]. Mûsa [2570]. ma [4196, 2372]. maître [381]. maîtres [2268, 604, 590]. maîtriser [3741]. Mabille [1230, 1521, 1846]. Macchine [2808, 2739]. Macciò [3105]. Macfarlane [2325, 2143]. Mach [1931, 2248, 3245, 3500, 3245]. Machabey [281, 239, 463]. Machinarum [2873]. Machine [4143, 3261, 1404, 1211, 1410, 1425, 1585, 1722, 1210]. machines [1582, 2487, 1407, 2372, 3439, 4346, 4162, 834]. machinisme [1889]. Mackintosh [999]. MacLean [990]. Macquer [777]. macroscopic [2747]. Madame [560, 3930]. Maddison [1759, 1893, 1759, 2356]. made [3570, 460]. Madeleine [4289, 2915, 1230, 1521, 1846, 1851]. Madrid [897, 3946]. Madrigal [4458]. madrileña [3946]. Maeyama [4318]. Magalhães [529]. Magasin [3289]. Magazine [537, 593, 888, 972, 1120, 332]. Magda [3766]. Magellan [571]. maggio [877]. Magic [874, 329]. magicien [21]. magie [447]. magiques [2267, 52]. Magistrat [1941]. Magistrats [1772]. Magnet [906]. Magnete [1164, 2176]. magnetic [3220]. magnétique [1726]. magnétiques [157]. Magnétisme [2415, 2220, 3987, 3984]. Magnitudes [2228]. Magnol [868]. Magnus [1908]. Magny [1586]. Mahoney [2137]. Mai [3536, 1107, 162, 279, 1808, 459, 3569, 2448, 3970, 3664, 3409, 4363, 333]. mai-3 [333]. Maillet [3288]. Maïmon [2983]. Main [1473]. Maine [4176]. mains [4528]. Mainz [278]. Mairan [4598, 4600, 4601, 4597, 4596]. mais [2017, 2003]. Maison [4607]. Maitte [2752]. Makers [3940, 2807, 514, 1415]. makes [3646]. Making [2905, 2445, 1418, 2878, 191, 4010, 3603]. mal [3591, 3824]. Malachi [4144]. malade [1869]. malades [2451]. maladie [1135]. Maladies [2844, 4035, 4481, 4477, 4028, 2799, 2234, 2448, 2919, 3574, 3490, 1172, 4147]. Malamoud [3703]. Malaria [607]. Malay [950]. Malebranche [1753, 811]. malebranchiste [1006, 1082]. Malet [3078, 1692]. Malgaigne [1459, 1529]. Malherbe [2374, 4120]. Malinas [4147]. Mallet [31]. Maloney [2842]. Malpangotto [4476]. Malpighi [3062, 4083, 4067]. Malta [1833]. Malus [2376]. Mamer [780]. Mamiani [2375]. Mammals [674]. Man [2905, 3673, 2961, 2579, 727, 3570, 674]. Mancosu [4623]. Mandelbrojt [2215]. Mandler [970]. Mandrou [1772]. Manetti [3105]. Manfred [3764]. Mani [1689]. Manier [2477]. Mann [2336]. Mannigfaltigkeitsbegriffs [2780]. Manning [1910]. männlichen [2989]. manoscritti [3150]. manquée [4407]. Mansfeld [2320, 3637]. mantikî [1388]. Mantoux [330]. Mantova [4088]. Manuel [4616, 4121, 822]. Manuels [3439, 4655, 4255, 3295]. manufactured [2452]. Manuscript [2564, 3011, 3749, 2372, 2686]. manuscripti [2249]. Manuscripts [2216, 4557, 3990, 4198, 4090]. Manuscrit [2328, 4168, 554, 123, 4504, 3936, 352, 1491, 682, 1087, 1112]. manuscrites [4119]. Manuscrits [3212, 4090, 573, 1538, 3715, 4228, 1741, 3103, 1258, 2318, 2801, 3717, 2006, 2380, 4548, 4066, 1230, 1521, 1604, 1846, 2105, 2405, 1948, 1294, 1231]. many [3737, 460]. manzano [3039]. mappemonde [161]. mapping [3476]. Maquet [3755]. Marat [3302]. Marbourg [3336]. Marc [3943, 3348, 750, 2666, 2919, 4052, 4208, 3261, 3426, 3920, 2180, 2263, 2245]. Marc-Antoine [3920, 2180, 2245]. Marc-Auguste [3348]. Marcel [10, 672, 4236, 161, 1175, 1958, 3756, 886, 3366, 3831, 2264, 1479, 1147, 774, 2488]. Marcelle [3213]. Marcellin [138, 381, 4610]. Marcello [4083, 880, 4067]. March [2452, 3965, 3368, 3599]. Marchal [40]. marche [2497, 2156]. Marchenay [2649]. marchesi [2991]. Marci [1655]. Marco [3793, 3941, 4076, 4435, 3560, 2926]. Marcohandschriften [2406]. Marconi [3402]. Marcou [4292]. Marcucci [3870, 2643]. Marcus [1655]. Maréchal [1260]. marées [2762, 4403, 1398, 1732]. Marganne [2705, 4036]. Margaret [2925, 2877, 2143, 4205, 2537, 4190]. Margerie [44]. marginalia} [4131]. Marginalien [1666]. margine [4301]. Margolin [4084]. Marguerite [1214, 581, 1775]. Maria [4224, 2344, 4202, 949, 989, 134, 4290, 3568, 4083, 1333, 841, 1469, 2808, 3946, 3735, 3732, 2690, 3735]. Mariafranca [4532]. Marian [2334]. Mariano [3666]. Marichal [1230, 1521, 1846]. Maricourt [2176, 4570]. Marie [4512, 2529, 3533, 3896, 1522, 2411, 3360, 1353, 3121, 4438, 4289, 3670, 862, 998, 1307, 2065, 2915, 3587, 4292, 2782, 3197, 2705, 4007, 2528, 3195, 3407, 2353, 2673, 3310, 1674, 4265, 790, 3750, 652, 2767, 1523, 1050, 4033, 2894, 3935, 2463, 2301, 4510, 1230, 1521, 4036, 3507, 858, 3835, 2571]. Marie-Antoinette [1674, 2767]. Marie-Claire [3896, 1307]. Marie-Claude [4438, 4292]. Marie-Élisa [652]. Marie-Elisabeth [2411]. Marie-Fernande [1353]. Marie-Madeleine [4289, 2915]. Marie-Paule [4007]. Marietti [4171, 4239]. Marin [2527, 2658, 2964, 3124, 3278, 109, 1015, 1072, 1232, 1467, 1667, 1955, 2095, 2107, 626, 1331, 1541, 1568]. Marine [2298, 963, 1059, 1708, 1755, 1048, 4469, 4468, 1524, 1834, 1050, 1723, 620, 2316]. mariner [4478]. marines [2997, 2962]. Marini [1773]. Marino [3631]. Marinus [2054, 86]. Mario [4078, 3414, 2988, 1503, 2077, 2173, 1441, 4122, 3781]. Marion [4434, 2502, 2631]. Mariotte [226, 3447, 3928, 3079]. Marisol [3437]. Maritime [1055, 4478, 839, 969]. maritimes [1524]. Marjorie [2143, 3996]. Mark [4627, 3706]. markets [4185]. Markoff [1555]. Markov [2877]. Marković [1105]. Markovits [4526]. Markovits-Pessel [4526]. Markowski [3438]. Markus [2903]. Markwart [3757]. Marlein [3611]. Marler [1025]. Marnef [3215]. Maroula [3935]. Marquis [827, 1377]. marquise [1828, 311]. mars [3343, 3460, 2582, 3221, 184, 250, 2389]. marseillais [4119]. Marshall [89, 689, 1539, 1847, 982, 3830]. Marsile [4394]. Marsilio [3941]. Marta [2344, 4368]. Martagon [4640]. Martha [2574, 2959, 420]. Marti [40, 1501, 953, 1030]. Marti-Ibañez [1501, 953, 1030]. Martin [3663, 4243, 2709, 2270, 4269, 4080, 2918, 3074, 1566, 3865, 1475, 1386]. Martine [4313, 2494]. Martinique [3753]. Martinus [2502, 2185, 1991, 1998, 2185]. martyr [511]. Martzloff [3207]. Marum [1998, 2185, 1991]. Marvels [3119]. Marvin [1566]. Marx [4393, 3103]. Mary [3752, 3459, 3389, 3794, 1240, 3094, 3705, 4190]. Masawayh [2669]. Maschine [2958]. Mason [581]. Masotti [2567]. Mass [1863, 235]. Massain [119, 451]. Masse [1437, 813, 3511]. Massignon [284]. massimi [4078]. Massimo [4344, 3923]. massive [2848]. Masson [3439, 79]. Mastai [2808]. masturbation [4011]. Mat [3411]. Matematica [830, 943, 2387, 1074, 1291, 566, 1778, 2567]. Matemáticas [3994]. Matematiche [829, 2874]. matematico [4239, 4476, 2808, 154]. matematitchesikja [591]. materia [3362]. Material [235, 4002]. Materiales [3735]. materiali [4515, 4076]. Materialism [2226, 2457]. matérialisme [2169, 2317, 2404]. Matérialité [3863]. materials [2686]. Matériaux [3884, 3721, 1105, 948]. materie [325]. matériels [2738]. math [1633, 278]. Mathematica [2191, 1098, 2106, 240, 2712, 2788, 3731, 1956, 1753]. Mathematical [1973, 3340, 1863, 4203, 4563, 789, 1670, 1672, 1929, 2025, 2057, 2096, 1572, 2742, 89, 2877, 2183, 4321, 3666, 3932, 1669, 2137, 3892, 3808, 3728, 3041, 3388, 2492]. mathematicarum [4301, 4301]. Mathematician [828, 4617, 3886, 2903]. Mathematicians [2373, 1098]. mathématicien [4512, 2215, 3, 1088, 770, 3268, 229, 734, 309, 1541, 1960]. mathématiciens [1698, 24, 1073]. mathematico [1099, 241]. mathématico-logiques [241]. Mathematics [4182, 3304, 3053, 3009, 2928, 2608, 3118, 2506, 2876, 395, 1098, 1297, 3420, 3206, 2419, 2373, 2712, 2788, 2831, 37, 4655, 2924, 3841, 3933, 3666, 3423, 4188, 3891, 4430, 1925, 1844, 207, 823, 856, 2403, 2816, 3866, 2078, 2779, 2776, 3839, 4020, 4203]. Mathematik [1540, 2636, 482, 1668, 2106, 2229, 2392, 278, 394, 787, 1325, 1926, 2053, 3429, 3237]. Mathematiker [278, 482]. Mathematiker-Biographen [482].

Mathematikerkongress [2058]. Mathematik^{≫} [188]. Mathématique [4513, 4189, 3027, 1633, 2314, 2909, 3426, 2967, 3433, 2981, 2637, 559, 4428, 1183, 3932, 518, 2469, 68, 81, 355, 1386, 4432, 2561, 4096, 1360, 3030, 4336, 3507, 3800, 2438, 3609, 3893, 4047, 208, 827, 858, 1490, 4554, 3876, 3720, 4310, 2167, 132, 150, 3561, 902]. Mathématiques [724, 782, 1823, 4132, 3026, 4000, 3669, 4305, 482, 1681, 3795, 4551, 3999, 4023, 208, 4302, 4304, 4189, 4424, 1949, 251, 10, 2441, 3882, 4303, 2191, 2227, 3207, 3189, 2507, 652, 4655, 3103, 1626, 4426, 2541, 4287, 3933, 4425, 2312, 2313, 3028, 3174, 2354, 3717, 2469, 367, 551, 987, 1387, 3175, 4560, 591, 55, 4137, 3190, 3286, 3600, 3595, 3884, 1082, 3264, 646, 3429, 3825, 493, 944, 3790, 2386, 3239, 316, 372, 375, 530, 834, 1290, 1294, 1828, 2311, 2781, 2349, 2650, 547, 550, 2169, 237, 1794, 3729, 3795, 4493, 3798, 2775, 22, 2741]. mathématiques [236, 832, 4145, 788, 3720, 3504, 1218, 3442, 1073]. mathématisation [3140, 3137, 3142, 3141, 3134, 4189]. mathématisations [3138]. Mathematische [2058, 1693]. mathematischen [2406]. Mathematischer [3223]. Mathematisches [2384]. Mather [1779]. Mathesis [4623, 4179, 3811, 3514, 3925, 3729, 4452, 4635, 4163, 4635, 3704, 4177]. Mathey [535, 73]. Mathieu [380]. matière [228, 4170, 398, 1379, 3160, 3692, 3234, 2292, 4394, 860]. Matières [166, 259, 334, 423, 538, 634, 705, 802, 910, 1036, 1124, 1192, 1272, 1364, 1447, 1546, 1642, 1654, 1711, 1799, 1871, 2079, 2199, 2277, 2359, 2480, 2552, 2615, 2713, 2789, 2887, 3538, 3614, 3679, 3769, 3846, 3906, 3974, 4038, 4100, 4148, 4212, 4273, 1962]. Matrice [4436]. Matricule [672]. Matrix [3410]. Matter [47, 48, 49, 50, 110, 97, 98, 111, 99, 100, 168, 169, 170, 171, 172, 173, 260, 261, 262, 263, 264, 265, 266, 267, 335, 336, 337, 338, 339, 340, 424, 425, 426, 427, 428, 429, 430, 431, 539, 540, 541, 542, 543, 544, 545, 546, 635, 636, 637, 638, 639, 640, 641, 642, 706, 707, 708, 709, 710, 711, 712, 713, 803, 804, 805, 806, 807, 808, 809, 810, 911, 912, 913, 914, 915, 916, 917, 918, 1037, 1038, 1039, 1040, 1041, 1042, 1043, 1044, 1125, 1126]. Matter [1127, 1128, 1129, 1130, 1193, 1194, 1195, 1196, 1197, 1198, 1273, 1274, 1275, 1276, 1277, 1278, 1279, 1280, 1365, 1366, 1367, 1368, 1369, 1370, 1371, 1372, 1448, 1449, 1450, 1451, 1452, 1453, 1454, 1455, 1547, 1548, 1549, 1550, 1551, 1552, 1553, 1554, 1643, 1644, 1645, 1646, 1647, 1648, 1649, 1650, 1712, 1713, 1714, 1715, 1716, 1717, 1718, 1719, 1800, 1801, 1802, 1803, 1804, 1805, 1806, 1807, 1872, 1873, 1874, 1875, 1876, 1877, 1878, 1879, 2007, 2008, 2009, 2010, 2011, 2080, 2081, 2082, 2083, 2144, 2145, 2146, 2147, 2148]. Matter [2200, 2201, 2202, 2203, 2278, 2279, 2280, 2281, 2360, 2361, 2362, 2363, 2425, 2426, 2427, 2428, 2429, 2481, 2482, 2483, 2484, 2553, 2554, 2555, 2556, 2557, 2558, 2559, 2560, 2616, 2617, 2618, 2619, 2620, 2621, 2714, 2715, 2716, 2717, 2718, 2719, 2720, 2721, 2790, 2791, 2792, 2793, 2794, 2795, 2888, 2889, 2890, 2891, 2892, 2893, 2943, 2944, 2945, 2946, 2947, 2948, 3017, 3018, 3019, 3020, 3021, 3022, 3023, 3024, 3085, 3086, 3087, 3088, 3089, 3090, 3127, 3128, 3129, 3130, 3131, 3132, 3180, 3181, 3182, 3183, 3184, 3185, 3246]. Matter [3247, 3248, 3249, 3250, 3251, 3320, 3321, 3322, 3323, 3324, 3325, 3391, 3392, 3393, 3394, 3395, 3396, 3465, 3466, 3467, 3468, 3469, 3470, 3539, 3540, 3541, 3543, 3544, 3545, 3615, 3616, 3617, 3618, 3619, 3620, 3680, 3681, 3682, 3683, 3684, 3685, 3770, 3771, 3772, 3773, 3774, 3775, 3847, 3848, 3849, 3850, 3851, 3852, 3907, 3908, 3909, 3910, 3911, 3912, 3975, 3976, 3977, 3978, 3979, 4039, 4040, 4041, 4042, 4043, 4044, 4101, 4102, 4103, 4104, 4105, 4106, 4107, 4108, 4149, 4150, 4151, 4152, 4213, 4214, 4215, 4216, 4274, 4275]. Matter [4276, 4277, 4323, 4324, 4325, 4326, 4372, 4373, 4374, 4375, 4419, 4420, 4421, 4422, 4460, 4461, 4462, 4463, 4498, 4499, 4500, 4501, 4534, 4535, 4536, 4537, 4630, 4649, 4606, 4632, 3560]. Mattew [2104]. Matthaeus [1020]. Matthew [4260, 3274]. Matthias [3694, 4177, 4204]. Matton [4529, 4319, 4296, 4491]. maturité [3318]. maturity [2183]. Matyssek [3731, 3731]. Maucuer [2073]. Maunoury [1736]. Maupertuis [3600, 3971]. Maupied [2435]. Maurepas [3625]. Maurice [4310, 587, 2868, 20, 3520, 1309, 1673, 3704, 1889, 86, 1473, 2756, 752, 1392, 1686, 2549, 2572, 3790, 4489, 3988, 1145, 1118, 2710, 1858, 1382, 2863, 816, 3876]. Maurin [2353]. Maurizio [2490, 4071, 4207, 3923, 2375]. Max [3902, 2318, 284]. Maximilian [2106, 1693, 1935]. Maximiliano [331]. Maxwell [3464, 2753, 3792, 4314, 3991, 2918]. Maxwellians [3436]. maxwellienne [3776]. May [836, 2712, 2788, 2191]. Mayaud [4169]. Mayenne [719]. Mayer [3605, 2872, 2710, 1345, 2727, 3660, 744]. Mayerhöfer [1070, 1298, 1564]. mayores [1231]. Mayot [1958]. Mayoux [3765]. Mayr [3153, 3515, 3149, 2710]. Mazaheri [476, 1630]. Mazaleigue [4624]. Mazaleigue-Labaste [4624]. Mazliak [4531]. Mazzolini [2685]. McAnally [667]. McClellan [3527]. McConnell [2850]. McCormmach [2262, 2192, 1944, 2495]. McCosh [3008]. McGuire [2632]. McKie [452, 453]. McVaugh [4206]. meaning [4321, 3932]. Meanings [3480]. Measure [3864]. measurement [3944, 3718]. Measuring [3082]. Me’ati [3125]. mécanicien [1729, 3514, 1881]. mécaniciens [1339]. mécanique [1817, 3578, 1858, 925, 1282, 4447, 4553, 652, 317, 1396, 3263, 3352, 1327, 364, 2160, 2725, 1993, 3453, 2, 1674, 4604, 1249, 448, 1544]. mécanisme [1696, 4638, 3113, 2736, 11]. Mécanismes [3677]. mécaniste [2584, 2317, 3067]. Meccanica [943, 4122, 3674]. mécènes [2283]. Mechanica [2873]. Mechanicae [2873]. mechanical [89, 2372, 2725, 1893]. Mechanics [3411, 3006, 3578, 3944, 1104, 3327, 4604, 4602, 567]. Mechanik [3419]. Mechaniques [1541]. mechanisering [244]. Mechanisierung [937]. mechanism [4638, 2226]. Mechanist [1110]. Mechanist-Vitalist [1110]. méconnu [4075, 2221, 2237, 2263]. méconnue [1404, 561]. méconnus [1737]. Mecquenem [1290]. Med [690, 861, 3994, 3735, 3732]. Médaille [899]. Medawar [3364]. Medd [3411]. médecin [794, 2409, 1909, 1059, 408, 1050]. Medecine [744, 724, 782, 1860, 279, 2193, 2296, 633, 696, 974, 1035, 2043, 2132, 956, 273, 436, 672, 2238, 1487, 773, 1498, 1005, 774, 765, 749, 4549, 4481, 4524, 4450, 4248, 4165, 3360, 958, 3252, 3294, 3927, 3188, 4255, 3460, 3314, 2412, 2884, 2975, 570, 964, 1115, 1173, 1305, 1309, 1310, 1638, 1701, 1756, 1757, 2035, 2086, 2340, 2448, 1966, 3745, 4625, 379, 3047, 3490, 3938, 3879, 3884, 3435, 1775, 4035, 4237, 4539, 952, 1304, 1937, 4169, 3244, 726, 279, 139, 535, 3272, 73, 257, 957, 2670, 2919, 779, 2163, 4490, 3764, 3577, 2187, 3151]. Médecine [458, 1427, 593, 2110, 2111, 3242, 1939, 2677, 2705, 3313, 2660, 884]. médecine-Inst [3938]. médecins [4405, 3071, 2577, 322, 4319, 1176, 613, 774, 4486, 1479]. Mededeling [514]. Médée [3834]. Media [1907]. Mediaevalia [2216]. mediators [4190]. Medica [3503, 1710, 3636, 3316, 3965, 4364, 4453, 3362]. Medical [1030, 2378, 2709, 1442, 3338, 3636, 798, 799, 800, 2811, 3584, 1499, 1504, 2450, 3209, 3274, 3860, 512, 4004, 2754, 461, 1245, 1480, 2666, 3295, 1304, 772, 1486, 1833, 675, 1940]. médicale [279, 60, 4482, 4196, 962, 3867, 1379, 3753, 1170, 1173, 1302, 1756, 2708, 92, 3293, 510, 1458, 4094]. médicales [935, 323, 280, 2238, 570, 883, 1579, 3633, 1229]. médicaux [4123, 3753, 606, 772, 3745, 3626, 2669, 419, 754]. mediche [2787]. medici [3921]. Medicina [1441, 2644, 33, 3505, 3454, 1324, 248, 1029, 3105]. Medicinae [1020, 1710]. médicinales [279, 387]. Medicinalium [2233, 3755, 536, 628]. Medicine [3459, 2342, 997, 3480, 1500, 1502, 2172, 3764, 2537, 2811, 2993, 1244, 1942, 631, 1637, 4557, 3579, 2190, 1486, 3417, 3584, 3589, 3274, 3317, 3630, 3636, 607, 3587, 3868, 3965, 3860, 4188, 3921, 3964, 4036, 4364, 3588, 3967, 608, 609, 610, 279, 888, 972, 1120, 999, 1501, 2129, 2142, 2581, 2323, 2393, 1940, 953, 3532, 3534, 3481, 537, 3744, 3597, 332]. Médicis [4081]. medico [3039, 459, 4664, 1707, 774, 4664]. médico-chirurgical [1707]. médico-chirurgicale [4664]. médico-psychologique [459]. médico-psychologiques [459]. médico-social [774]. Medico-surgical [4664]. médiétés [113]. Medieval [689, 3118, 1162, 1847, 1997, 2094, 1104, 2024, 2216, 3812, 784, 4557, 3516, 3744, 4082, 3586, 3587, 3796, 4131, 3737, 1566, 481, 2693, 3795, 2921, 905]. medievale [3759, 3097, 3924, 3158, 3644, 1661, 112, 565, 818]. Médiévales [1784, 3267, 4488, 2754, 2693, 3741]. médiévaux [619, 1231]. Medioevo [4078, 565, 4307, 1503]. médiques [1892]. Méditations [4636, 4636]. Méditerranée [3935]. Méditerranée. [969]. méditerranéenne [75]. mediterraneo [565]. Medium [4069]. Medizin [1534, 2478, 1709, 536, 3503, 1689, 2685, 2785, 954, 1440, 1617, 2392, 1709, 2843, 2917]. medizinische [2536, 1526, 2989]. Medizintheorie [3123]. Meeting [3536]. Megaterio [3732]. Mehrtens [2674]. Meier [729]. Meiji [4655]. Meitner [3871, 4172]. Melancholiker [3488]. mélancolie [3111]. Melanderhjelm [31]. Mélanges [3741, 4265]. Melbourne [3864]. Melchior [950]. Melchisédech [2099, 1374]. Melhado [3485]. Meli [4067]. Méliténiote [3607]. Melvin [1888]. Mém [2420]. membership [1919]. Membre [1839]. membres [1306, 1307, 1443, 1826]. Memoir [2837]. Mémoire [1986, 3162, 4307, 2967, 3819, 3855, 107, 2367, 3828, 4300, 4267, 2809, 1322, 4447, 281]. Mémoires [2455, 3413, 1290, 1294, 2311, 3631, 3260, 3504, 4003, 3178, 3348, 2575, 3749, 1338, 886, 2538, 1735, 3962, 4285, 4363, 2879, 378, 893, 4259]. Memoirs [743, 2321, 947, 3830]. Memorabilia [1098, 1119]. Memoria [3759, 831]. Memorial [1317]. memoriam [268, 4244, 4449]. Memorie [4076, 4435]. memory [4024]. Men [1237, 2271, 477, 784, 2957]. Mendel [1843, 814, 196, 661, 978]. Mendeleev [1743]. Mendelism [1616]. mendélisme [4155]. Mendelsohn [3726, 4077, 1476]. Mendoza [1891]. Mengal [4196, 3635, 4350]. Mengoli [4136]. Menjot [1775]. menor [566, 1324]. mens [513]. Mensajero [1542]. Mensch [605]. Menshutkin [478]. mensura [4069]. mensurandi [4006]. Mensuration [2921]. mental [1869, 2920]. mentale [2397]. mentales [4477]. Mentalités [1772, 3608]. mentioned [359]. Menut [1997]. Mer [1524, 4017, 3929, 4466, 887]. Méray [1857]. Mercator [278]. Mercier [3758]. Mercure [176, 4552]. Mercurialis [60]. mère [4450]. meridian [4571]. Méridien [4080, 2459]. Méridienne [4080, 4571, 4379]. mérites [4222]. merken [514]. Merleau [2060, 3557]. Merleau-Ponty [2060, 3557]. Merlin [4621]. Méropi [4486]. Mersenne [2527, 2658, 2964, 3124, 3278, 109, 1015, 1072, 1232, 1467, 1568, 1667, 1955, 2095, 2107, 626, 1331, 1541, 104, 2624, 2164, 1399, 1741, 4563, 103, 105, 468, 680, 106, 755, 101, 102, 11, 3628]. Mertens [3962]. Merton [2694]. mes [32]. Meslin [3345]. Meslin-Perrier [3345]. Mesmer [2421]. Mesnard [3309, 3299, 4319, 1775, 1570]. Mésopotamie [677]. Message [1403, 350, 2215]. Messager [3522, 3566]. Messeigneurs [4196]. Messenger [1543]. Messengers [3666]. Messina [3344]. Messmer [4411]. Messrs [2837]. Mésué [2669]. Mesurage [1409]. Mesure [3929, 3695, 3722, 3671, 4008, 4415, 1927, 1917]. mesurer [4415]. mesures [3549, 886, 3427]. mésusages [4663]. metabolism [3588]. Métadier [1472]. metafisiche [3062]. metafore [3238]. m’était [2648]. métal [3989, 1417]. métalliques [1350]. Metallographia [3645]. métallurgie [1351, 1260, 1260]. métallurgiste [794]. Metallurgy [1261, 1261]. metalurg [794]. métamorphose [1783, 1992, 4029]. métamorphoses [3853]. métaphisiques [4296]. metaphor [3762]. métaphores [4346]. Metaphors [2294]. metaphysical [4636, 4638, 4684]. Metaphysics [3611]. Métaphysique [1218, 4224, 3138, 2048, 1958, 3600, 234, 4221, 4333, 4526]. métaphysiques [4636, 4638, 4684]. metascience [3439]. metaulx [4084]. métaux [2120, 2288]. Metazoa [3368]. Metchnikov [1027]. météores [584]. Meteorological [1982, 1981]. Météorologie [362]. météorologiques [7]. Meteorology [1019, 1694]. Method [3009, 2333, 1110, 3362, 2990, 2963]. Méthode [2507, 2761, 1785, 2865, 3711, 569, 3359, 2415, 1378, 3135, 2227, 2314, 556, 2541, 3045, 1397, 3457, 2313, 472, 2217, 1625, 1569, 551, 1573, 2055, 4337, 584, 4466, 3113, 1058, 2438, 855, 2596, 2168, 371, 4417, 796]. Methoden [2841, 787]. méthodes [1720, 175, 223, 2272, 1283, 1764, 4433, 4547, 855]. méthodique [4390, 320, 3936, 3287]. méthodiques [2134]. Methodological [1974]. Méthodologie [1533, 1861, 3703, 3656]. méthodologique [2027]. Methodology [3703]. Methods [2142, 3608, 3235, 3423, 2372, 118]. métier [4306, 4382]. métiers [3987, 314, 181, 388]. Metman [1521, 1230, 1846]. Metodo [1628, 2375]. metodologica [3491]. Metrios [1682]. métrique [813, 3342]. métrologie [344, 463, 463]. Metzger [1930, 2141, 2315]. Meudon [4392]. Meulders [3943, 4240]. Mexico [566, 885, 2815]. Mexique [2599, 3552, 1429]. Meyer [745, 137]. Meyerhof [331]. Meyers [4625]. Meyerson [4477, 4476, 4473, 4472]. Mézières [124]. Méziriac [3]. Mgr [3844]. Mi [4316]. Mianet [1312]. Michael [4077, 4125, 3902, 3945, 3649, 4315, 4478, 3638, 4206, 3570, 2171, 3001, 2403, 2137]. Michajlovic [4361]. Michaux [86]. Michel [2976, 3578, 4390, 4240, 4233, 4616, 2913, 3943, 3178, 3971, 2609, 3832, 3807, 2770, 3302, 3787, 4141, 4005, 3242, 3993, 1302, 1308, 1701, 4398, 3561, 4439, 2962, 4396, 4292, 4120, 2782, 4349, 4643, 4510, 3234, 1759, 2356, 4140, 3463, 1606, 2404, 2443, 2748, 2751, 3453, 4122, 839, 1761, 4558, 3530, 167, 2982, 22, 71, 4538, 496, 1516, 301, 733, 1217, 1084, 511]. Michel-Pierre [2770, 4643]. Michela [4476]. Michelangelo [3759]. Michele [4345, 3039, 2941, 4122, 3962, 2925, 3949]. Michelet [1258, 1416]. Michelle [3212, 3413, 2446, 2704, 3556]. Micheloni [2814]. Michie [3439]. Michler [3757]. microbes [4670]. microbiologie [363]. Microbiology [1109]. Microcosme^{≫} [968]. micrographie [1148]. Microorganismes [535]. microphysics [4197, 4002]. microphysique [72]. microprocesseurs [2745]. Microscope [3904, 1586, 2036, 288, 815, 2700]. Microscopes [4124, 4045]. microscopi [1300]. microscopie [1148]. microscopiques [279]. mid [3659]. mid-seventeenth [3659]. Middeleeuwen [1793]. Middle [1539, 3810, 3480, 2692, 2777, 3275, 4238, 589, 3587, 4573]. Middleton [1694, 1982, 1471, 1981]. Midi [2915]. midst [4674]. Mieczyslaw [3438]. Mieli [1630, 566, 417]. mighty [1771]. Migliore [3562]. Miguel [4066, 4296, 3780]. Mikami [1297]. mil [1695]. milagro [238]. Milan [722, 4590, 4633]. milanaise [4590]. Milano [3674]. Milec [1239]. Milestones [837]. Milhaud [4509, 4251, 4192, 847, 742, 1056, 1233]. Milic [2061]. Milieu [1524, 2431, 2754, 4674, 1059, 585, 1513, 3345, 3261, 4573]. Milieux [3943, 3862, 3722, 3068, 3405, 2408]. Milieux^{≫} [3442]. militaire [1001]. Military [744]. Milky [3083, 3673]. Millás [1436]. mille [2734, 1691]. millennium [4188]. Miller [2870, 2870, 1690, 1502, 865, 2528, 2746]. Millet [16]. Milton [999]. min [2372]. Mind [2524, 3763, 4055]. mind-brain [4055]. mine [748, 3989]. Mineral [747]. minerali [2358]. mineralog [794]. mineralogia [2358]. Mineralogie [3214, 3645, 4401, 4087]. minéralogique [1424]. minéralogiste [794, 2797]. Minerals [2641, 1908]. mineralvattenområdet [747]. Minería [885]. Mines [1557, 4413, 2797]. mineur [794]. Ming [606, 1166, 2187, 1390]. ming-che [1390]. Mingshan [3880]. miniatures [1112]. minière [1584]. miniers [1884]. Minime [2527, 2658, 109, 1667, 2964, 1015, 1072, 1232, 1955, 2095, 2107, 626, 3628]. minimo [4069]. Minister [4608]. ministère [2411, 2316, 1943, 1835, 3052]. ministères [330]. ministre [4608]. minor [2491]. Minutier [2204]. mirabilium [3119]. miracle [284]. mirage [3958]. Mireille [1939]. Mirella [4475]. Mirko [3927, 3272, 3834, 1173, 2919, 4146, 4035, 1112, 1165, 3961]. miroir [4502, 4092]. miroirs [3496]. Mirrors [2419]. Miscellanea [4088, 3482]. misconceptions [4663]. mise [1722]. mises [273]. Missa [4477]. Mission [3052, 1290]. missionaries [2932]. missions [2599]. Misunderstanding [3836]. Mitcham [2522]. Mitglieder [1323]. Mitman [3486]. Mito [3454, 3627, 3583, 3058]. Mitrinović [3009, 2568]. Mitscherlich [3489]. Mittag [4049]. Mittag-Leffler [4049]. Mittelalter [2106, 1540]. mittelalterlichen [2843]. Mittelalters [3815, 4097, 4175, 4234, 2287, 3443, 3346]. Mittelstaedt [3944]. mixed [1388]. mixte [279]. Mlle [1523]. Mller [2979]. Mme [2845, 1955, 3289, 2908]. Mnemosyne [3488, 3611]. mobile [3669]. Mocrevii [939]. model [3902]. modèle [3348, 2678, 1818, 4129, 3643, 3067]. Modèles [4181, 3925, 4346, 1582, 3475, 4187, 4343]. modélisation [4189]. modélisations [4181]. Modéliser [4184]. Modells [3734]. Models [4182, 1259, 4185, 4235, 1758, 3501, 3220, 4190]. moderate [4686]. Modern [3697, 2061, 2664, 191, 1162, 2811, 2749, 1999, 2159, 2457, 3006, 4557, 4572, 4627, 3810, 1576, 3410, 2924, 3337, 3277, 2214, 2477, 532, 3423, 4167, 4294, 4256, 276, 2875, 3327, 610, 2142, 3344]. moderna [62, 2926, 1291]. modernas [1778]. moderne [726, 4572, 319, 926, 371, 1380, 3557, 4091, 2533, 1168, 1983, 2035, 3043, 594, 832, 1202, 112, 4237, 2634, 1394, 791, 4312, 399, 819]. moderne^{≫} [120]. modernen [2740]. Modernes [4644, 2754, 1683, 2693, 3741, 855, 2156, 704]. Modernité [4047]. modernités [3747]. moderno [3491]. modification [1375]. Mogenet [845]. Mohamed [1605]. Mohammad [4417]. Mohsen [4545]. moindre [3600]. moindres [3135]. moitié [177, 1821, 3376, 4062, 4523, 3623]. Molard [1881]. Moldenke [485, 485]. Moldova [959]. moléculaire [2667, 4678, 3490, 3606, 4045]. moléculaires [2857]. Molecular [3464, 2667, 3590, 1980, 4678, 4626]. molécule [4295, 3690]. Molecules [2957, 3985, 2833]. Molgo [2611]. Molinari [4261]. Mollat [2962, 969, 839]. moment [252]. momento [2765]. moments [3099, 2982]. Mommsen [1734]. mon [4022]. Monaco [1602, 1602]. monade [4137, 4069]. monadologie [4683]. monadology [4683]. monastique [4199]. Monceau [2316, 1222, 2897]. Mond [1693]. Mond- [1693]. Monde [301, 3256, 324, 582, 2705, 4060, 2292, 4304, 85, 116, 147, 679, 1817, 3604, 3802, 3721, 57, 407, 4170, 4368, 3929, 1480, 1997, 2919, 3561, 3756, 1288, 1845, 3070, 1732, 3366, 6, 28, 112, 3258, 1674, 1243, 3665, 483, 3038]. mondes [704]. mondial [2975]. Mondiale [1159, 1744, 4427, 1630]. mondiales [4426, 4428]. Mondo [2471, 4078, 1441]. Mondor [17, 17, 670]. Mong [666]. Monge [198, 26, 83, 1490, 374]. Mongrédien [585]. Monika [3706]. Monique [753, 2962, 1230, 1521, 1846]. monks [4238, 4557]. Monnerville [350]. Monnier [4381]. Monod [3922, 2667, 909, 2667]. Monod-Herzen [909]. Monograf [1108, 3735, 3732]. monografías [3994]. Monograph [2109, 2336, 2155]. Monographs [3590, 2686, 2537]. Monro [1504]. Monsieur [612]. Monspeliensis [1005]. monsters [4677]. monstres [2097, 4677]. Mont [465]. Montagne [15, 3569]. montagnes [2365]. Montalcini [3957]. Montanari [4072]. Montbard [3717]. Montbéliard [387]. Monte [4602, 2991]. Montée [4589]. Montel [4021]. Montgolfier [3004]. montgolfière [3004, 4309]. Montgomery [3060]. Monti [4202]. Montias [1474, 1632]. Montmor [2088, 2087]. Montpellier [956, 672, 2218, 2755, 2852, 1176, 273, 436, 1050, 2238, 2563, 1176]. montpelliérain [757]. montpelliéraine [758, 1005]. montpelliéraines [2238]. Montres [1420]. monument [3262]. monumentale [1969]. Moody [689]. Mooij [1633]. Moore [3506, 3940, 2876, 3440]. Moquin [4022]. Moquin-Tandon [4022]. Mora [3437]. Morale [130]. morality [3860]. Morando [2060]. Morange [4439]. morbis [3921]. morbo [4062]. Mordechaï [3636]. Moreau [4292, 1756, 2738]. Morée [2599, 3935]. Morel [1225]. Morelon [3241]. Morgagni [4083]. Morgan [4190, 4431, 196]. morganien [270]. Morgenthaler [4093]. Morgues [2539]. Morin [228]. Morison [887]. Moritz [1098]. Morley [1564]. Mormino [3674]. Moronia [2479]. morphologie [909, 1783, 1992, 4029]. Morris [2776, 2977, 859, 3368]. Morrison [4190]. Morse [3593]. Morsel [3948]. Morselli [2988]. Morsier [2033]. mort [199, 1493, 496, 3311, 461, 1190, 3072, 1529, 1657, 2180, 3968, 2613, 671]. mortels [3463]. mortes [1851]. Mortier [3809, 480]. morts [4450]. Mörzer [3733, 4478]. Moscou [1825, 1862, 3731, 3784]. Moscovici [2532, 1906, 1751]. Moskva [4203, 2510]. Mosnier [4319]. Moss [3675]. Mostra [3150, 1469]. Moström [793]. moteur [1254]. moteurs [3052, 1730]. Mothu [4296]. Motion [2094, 1104, 3752]. Motions [1847]. Motive [1891]. motoculture [2648]. Motors [1733]. motrice [2509, 834]. Motte [948]. Mottier [2410]. Motu [1021, 976, 2656, 733, 4122]. Motu^{≫} [210]. Motus [3674]. motuum [2094]. Moulin [3360, 3722, 3533]. Moulinier [4415]. Mounier [1707]. mountain [4236]. Mourelatos [2803]. Mourier [4297]. mouvement [2762, 1774, 1849, 2282, 3068, 3802, 4013, 2627, 4134, 757, 3672, 4287, 356, 3453, 1541, 774, 4223, 1751]. mouvements [4488]. movement [3659]. Moxa [279]. Moyen [3462, 2267, 2828, 1794, 3842, 3924, 3626, 2272, 2758, 3741, 144, 969, 2592, 2929, 476, 4089, 3884, 4379, 4126, 1916, 4129, 3873]. Moyen-Orient [144]. moyennes [3007]. Moyne [2539]. MR [2400]. MR1164428 [3497]. MR1351497 [3933, 3932]. ms [1848]. MSHE [4646]. Muccioli [880]. Mueller [2803]. Mugler [984, 1663, 1927, 2055, 2093]. Muhly [2757]. Mulder [3218]. mulhousien [2407]. Muller [2242, 3153, 3156, 2423, 536, 1332, 1111, 2499]. Multhauf [2345, 1259, 2642]. Multigrades [153]. multilingues [615]. multiplicatione [3126]. Multiplikation [2473]. München [278]. Mundialización [3592]. Mundini [1020]. Mundo [2166, 3592]. Mungo [798]. Muñoz [3994]. Münsters [3734]. Murdin [4290]. Murdoch [3315]. Muriel [3462, 4558]. Muris [4006]. Murphy [3945]. Murray [693]. Murs [2382, 2515]. Musa [3420, 4023]. musée [1586, 1589, 1255, 2085, 346, 181, 2902, 1079, 294, 841]. Museo [2344, 989]. Museología [4453]. Museum [1593, 1595, 2342, 730, 359, 768, 4478, 2452, 1498, 1982, 1940, 1578, 1271, 1354, 1356, 1758, 1854, 3082, 2807, 2050, 2318, 3996, 3715, 3820, 3969, 182, 2256, 2319, 1338, 2006, 2380, 2598, 3623, 3389]. Museums [2566, 1267].

Music [1590, 3482, 3519]. musicale [2995, 4112, 2907]. musiciens [755]. Musik [2478]. musiques [4265]. Musschenbroek [359]. must [4658]. Musulmans [476]. Mutanna [1436]. Mutation [4051, 4124, 2657]. Muzea [1355]. Muzeum [1854]. mycology [3209]. Mylon [3268]. mystères [1215]. Mysterium [2869]. Mysticism [2224]. myth [2450, 2306]. Mythe [4258, 1248, 3233, 3098]. Mythes [1434, 2580].

N [478, 12, 31, 485, 3337, 3839, 4020, 4203, 2920, 2510, 2469, 3409, 3206, 1109, 900, 1412, 2749, 4394, 3120, 1332, 1336, 2355, 3041, 1060, 3217, 2242, 2521, 2737, 1267, 1265, 1247, 1266, 1594, 1595, 164, 1973, 2308, 2320, 2868, 2870, 795, 2942, 1752, 3359, 2189, 2275, 2276, 2343, 2345, 2422, 592, 617, 843, 949, 968, 973, 1004, 1033, 1121, 1315, 1434, 1444, 1508, 1640, 1796, 1975, 2037, 2038, 2042, 2117, 2116, 2133, 1214, 359, 1270, 1736, 3083, 328, 443, 461, 465, 1208, 1524, 397, 514, 2637, 2109, 2170, 797, 1003, 1061, 1302]. n [1306, 1307, 1311, 2323, 2393, 1847, 1997, 2603, 2453, 151, 1215, 1327, 1328, 1634, 1673, 1982, 2250, 1578, 3582, 4643, 2285, 1472, 2811, 4239, 945, 1268, 1271, 1351, 1352, 1354, 1356, 1355, 1413, 1421, 1735, 1890, 1751, 1018, 1230, 2406, 2583, 2758, 1058, 2756, 944, 3387, 2386, 2373, 353, 160, 184, 192, 240, 278, 421, 383, 531, 488, 836, 841, 1326, 1543, 1566, 1572, 2781, 221, 246, 247, 363, 1014, 1024, 1076, 2070, 2882, 3665, 610, 953, 2246, 2614, 2854, 2855, 1964]. N.J. [2137]. N.M.H.S [836]. nôtre [2028]. Nabonnand [4049]. nach [3314, 2479]. Nachricht [1666]. Nachtrag [1470]. nacional [3592]. Nadal [2627]. Nadine [2574, 4055]. Nafiz [961]. Nagel [3491]. Naissance [1679, 4477, 4005, 4350, 3686, 3453, 4572, 3578, 1858, 4312, 11, 1808, 2765, 4196, 2678, 1402, 54, 3412, 4629, 3318, 4452, 2443, 3092, 2732, 1157, 1024]. Naldoni [1333]. Nam [1311, 279]. name [3508, 3756]. Namer [1972, 2290, 1403]. Names [3650]. Naming [3586]. Namur [275]. Nan [2921]. Nancy [2664, 2378, 2993]. Nantes [4261, 1115, 1579, 2268]. Naples [2517, 213]. Napoleon [1686, 4364]. napoleonica [4309]. Napoli [4068, 2551]. naquit [2738]. Narbonne [346]. Narodni [1854]. Narratio [3566, 2231, 2770]. Narrow [3307]. Nartov [1332]. nascita [877, 1469]. Nat [690, 861]. Nathalie [4021, 3408, 4490, 4052, 4169]. Nathan [2337, 2834]. nation [4436]. National [231, 4612, 2342, 730, 1835, 359, 4478, 2581, 1338, 181, 1271, 1354, 1854, 738, 1433, 3715, 1943, 2295, 2423, 2318, 2319, 3569, 2045, 3987, 3706, 2380, 2405, 3623, 333, 841, 3473]. Nationale [161, 1948, 1538, 1943, 1835, 1230, 388, 1967, 4612]. Nationales [571, 1966, 879]. nationalismes [4486]. nationalités [3820]. Natucci [602]. Natur [4093, 2958, 2701]. Natura [2375, 1537]. naturæ [4688, 4061, 4296]. Natural [402, 2871, 3384, 3991, 456, 3586, 1578, 907, 2226, 2528, 2403, 3604, 3994, 3838, 2903, 3312, 2883, 2477, 4613, 561, 4591, 4317, 4190]. Naturales [566]. naturali [4076, 4435, 4098, 3058, 2787]. naturalism [4404]. Naturalist [3794, 3743]. naturaliste [3942, 4538, 3922, 4022, 1059, 3213, 1321, 2897]. Naturalistes [863, 484, 386, 3820, 3969, 2598, 1176]. Naturalists [191]. Naturalium [2233, 3755, 536, 628]. Nature [3486, 3119, 592, 2533, 3126, 1338, 3389, 4542, 4410, 4583, 3492, 2227, 2497, 3121, 3528, 4014, 2924, 4550, 2838, 4171, 2339, 2120, 2463, 601, 2250, 3644, 3600, 4007, 4348, 3741, 2102, 3984, 869, 2907, 724, 782, 2123, 1906, 4558, 3319, 2865, 3638, 3742]. naturel [720]. Naturelle [182, 3631, 3715, 1600, 3288, 3311, 3348, 3622, 4236, 2318, 2319, 3749, 2380, 1811, 4084, 2833, 3173, 4613, 1574, 4455, 3623, 2184, 1828, 588, 1052, 4554, 2050, 2598, 1338, 1045]. Naturelles [1229, 935, 115, 2193, 2296, 633, 696, 974, 1035, 2043, 2132, 3266, 3569, 2670, 2044, 2582, 3818, 4048, 4030, 1253, 4591, 841, 878, 765, 1824, 73]. natures [3650]. Naturf [3843]. Naturfilosofi [3384]. Naturforschenden [1928, 826]. Naturforscher [3040, 2785, 3035, 3156, 3843, 2645]. Naturforschung [1106]. Naturkunde [3316]. Naturphilosophie [4551]. naturphilosophischen [4361]. naturwiss [278]. Naturwissenschaft [2674, 2391]. Naturwissenschaften [1709, 2685, 2785, 1890, 2392, 394, 1070, 1298, 1564, 2566, 1534, 1617]. naturwissenschaftlichen [690]. Naturwissenschaftlicher [3223, 3536]. Natuurkunde [325, 397]. Natuurwetenschap [14]. Natuurwetenschappen [1482, 359, 514, 841]. nauk [983]. Nauka [3731, 4203, 2548]. Nauki [1610, 971]. Nautchnobiografitcheskaïa [2510]. nautique [1721, 4470, 2114]. Nautonier [4158]. Naux [1608, 2056]. naval [4475, 1579]. navale [1305, 1638]. Navarro [3994, 2926]. Nave [3813]. navicula [4557]. navigants [1579, 1708]. navigateur [1059]. navigation [4475, 2902, 1299]. navigational [3733]. naviguer [4470]. navire [839, 969]. navires [1116]. naz [4076, 4435]. Nazionale [603, 4088, 4076, 4435, 3240, 3438, 3813]. Neapel [1216]. Nebulae [2261]. nécessaire [1722]. necessary [1722]. nécessité [4410, 869]. necessities [1388]. Neck [4011]. Neckam [191]. Necker [330, 779]. Nécrologie [2985, 285, 4593, 84, 3049, 149, 1920, 200, 370]. Nedeljković [2245]. Nederland [1482]. Nederlandse [397, 1293, 514]. Needham [665, 1565, 1017, 3177, 2063, 2329, 3080]. Néel [1708]. Néerlandaise [1229, 935]. n’eevklidovoy [2385]. négation [1809]. négociation [4184]. Negotiating [3786]. négresse [4296]. négriers [1116]. neige [2307]. Neimeyer [3439]. neither [3091]. nel [3562, 4309, 565, 4476, 4161, 4070, 3562]. Nélaton [2269]. Nell [1771]. nella [3870, 3150, 2787, 3523, 565, 3491, 4068, 4122]. nell’alto [1503]. nell’Arte [2787]. nell’età [4345, 4344]. nell’Europa [3150]. nell’Italia [4262]. nell’ottocento [4161]. nell’ultimo [4098]. Nemec [1560]. Nemore [689, 3516, 2733]. neohumanist [3886]. néolamarckien [4155]. néolamarckisme [4540]. néolamarckismes [4638]. neolamarckisms [4638]. Neper [2056, 1608]. Neptune [4588, 2642]. nerveuses [4481]. nerveux [2217]. nervoso [3901]. nés [751]. Netherlanders [1414]. Netherlands [3903, 3154]. network [3965]. Networks [3525, 3816, 4079, 3797, 3990, 4514, 4009]. Neubauer [2740]. Neuburger [4204]. Neue [3815, 4175, 3489, 4234, 3416, 2287, 3443, 3346, 4097]. neuen [1666, 787]. neuer [2706, 3757]. Neugebauer [3010, 1071]. Neun [1777]. neural [3959]. Neuro [3252]. Neuro-endocrinologie [3252]. neuroembryologie [3957]. neurologie [4483]. neurologist [3579]. Neurology [965]. Neurophilosophie [3439, 3938]. Neurophysiologie [977]. neurosciences [3938, 4240]. neurotransmission [4019]. neutres [4086]. neutron [3257, 3091]. Neuve [2582]. neuvième [4578]. Neuzeit [3316]. Never [2699]. Newcomen [1594]. Newman [3651, 372, 3749]. Newton [743, 2529, 235, 4393, 3302, 62, 1573, 1669, 1670, 1671, 1929, 2057, 2096, 2972, 907, 1142, 3728, 2528, 2375, 1572, 771, 4495, 2492, 2493, 2742, 4157, 3063, 3064, 2981, 2968, 4502, 3069, 1154, 54, 553, 604, 3114, 3163, 521, 437, 1828, 3067, 516, 3752, 2547, 2530, 1974, 3598, 3807, 87, 2699, 2505, 2772, 3514, 2141]. Newtonian [2743]. Newtoniana [2505]. newtoniano [2375]. newtonien [3068, 407]. newtonienne [3710, 2913, 3511, 3302]. Newtons [3419]. next [3945]. Neyman [2524]. Nguyên [381]. Nguyên-Bính [381]. Nguyen [1305, 2962]. Nguyen-Thanh [1305]. ni [3091]. Nicard [2436]. Niccolo [3150, 2606, 2567]. Nice [1349]. Niceron [680]. Nicholas [3455, 3838, 2182, 3455, 2501]. nicht [3314]. Nick [4235]. Nicol [2340, 2325]. Nicolaïdis [3700]. Nicolas [4189, 3458, 2308, 4415, 3054, 1941, 1387, 3648, 2140, 4643, 1542, 2134, 4193, 1064, 2727, 3660, 2258, 4529, 2071, 1883, 4302, 861, 4625]. Nicolas-Claude [3458]. Nicolas-Jacques [1064]. Nicolas-Louis [3660]. Nicolaus [278, 3150]. Nicole [4264, 3900, 4488, 4411, 3832, 2032, 1997, 1163, 1472, 3825, 4332, 772, 3119, 1847, 2094]. Nicole-Genty [772]. Nicolle [35]. Nicomaque [3495]. Nicquevert [4166]. Niderst [3299, 4296, 3300]. niederländische [769]. Niedersächsischen [2872]. Niels [3411, 3360, 3238, 3372, 3755, 690, 828, 3411, 3040, 3315, 3371, 3699]. Nielsen [4514]. Niesser [1097]. nihil [4303]. Nijboer [2534]. Nijhoff [2502]. Nikitina [277]. Nikolaus [1689, 3346]. Niman [2903]. Nineteenth [2753, 2578, 1296, 2457, 4572, 3581, 2838, 3630, 3636, 3860, 4573, 4654, 4578]. nineteenth-century [3581, 2838]. Nissen [1639, 1868, 2037, 2265, 2513]. Nitre [1355]. niveau [1306]. nivernaise [1735]. Nivikov [3610]. no [279, 2467, 3755, 2533, 3587, 4582, 3497, 529, 1932, 2400, 2593, 2850]. Nobel [1867, 2960, 205, 2702, 1335, 1867, 324]. noble [401]. Nock [58]. Noël [4054, 4477, 1059, 1243]. Noëlle [3935]. Noether [3042]. Noguès [4474]. Noland [821]. Nolani [1216]. Nolin [1819]. Nollet [2900, 3982, 621]. nom [3756, 3582]. Nomanul [3650]. nombre [3721, 435]. Nombres [175, 3689, 3721, 812, 3334, 1738, 4287, 4427, 2469, 447, 481, 986, 1215, 1087, 2124, 790, 1326, 1932, 2781, 3877]. nombreux [2386]. Nomenclature [221]. non [1880, 3989, 1306, 1307, 1790, 4398, 3190, 2385, 3829, 4348, 3056, 3101, 3750, 3147]. non-Euclidean [4348]. non-euclidiennes [3147]. non-ferreux [3989]. Nonancourt [4492]. n’ont [3091]. nonum [331]. nor [3091]. Nora [3060]. Nord [948, 2448, 1132, 3854, 4292]. nord-américaine [4292]. Nordenmark [12, 31]. Nordman [3935]. Nordon [3366]. Nordström [1150]. Norimberga [1300]. Normal [4676, 15, 3728]. normale [4529, 2541, 4495]. normand [1321, 1348]. normanna [565]. normanno [565]. normativité [4675]. normativity [4675]. North [4541, 65, 3235, 2189, 2343, 3196]. nos [4165, 435, 4297, 535, 1914, 958, 1757, 2035, 2755, 2187, 3502]. Nostalgie [234]. nostri [248]. notamment [568]. notation [1224, 4238]. Note [1266, 2895, 1409, 1131, 3352, 2801, 2729, 1762, 4525, 4064, 2285, 1007, 1559, 1046, 1988, 4219, 4301, 1490, 2914]. Notebooks [1578, 526]. Notes [1102, 1189, 2909, 3212, 4022, 1049, 3126, 4643, 733, 834, 2197, 2087, 689, 1020, 1545, 1181, 1499, 151, 1544, 3429, 480, 457, 962, 2576]. noti [4083]. Notice [1102, 560, 375, 2168]. notices [2139, 153]. Noticias [3735]. notion [4359, 135, 3138, 4663, 1857, 3354, 4061, 3639, 2367, 3114, 3055, 4480, 4429, 2, 1851, 4554, 2973]. notions [568]. Notre [2673, 1079, 2786, 1761, 2068]. Notre-Dame [1079]. Noumène [2983]. Noureddine [751]. nous [2580]. nouv [2583, 834, 3596]. Nouveau [1257, 3289, 3143, 2898, 2401, 1330, 3510, 978, 2141, 790, 751, 2114]. nouveau-nés [751]. Nouveaux [4619, 4181, 194, 2258, 1883, 4618, 4155, 4530]. nouvel [4380]. Nouvelle [2925, 4174, 3408, 3625, 1906, 1695, 3407, 2824, 4581, 2243, 3552, 1, 3359, 4553, 354, 1408, 3290, 4504, 4494, 3350, 2060, 510, 2820, 141, 2033, 1887]. Nouvelles [3109, 4181, 2134, 2650, 2573, 1954, 2368, 3790, 1541, 1079, 2164]. Nov [2337, 162]. novæ [4548]. Novara [4391]. Novecento [4515]. Novels [487]. Novelties [3675]. novembre [3631, 4088, 2423, 3176, 1659, 673, 220]. novembre-2 [1659]. novohispana [2815]. Nový [2214, 2394]. Nowotny [3840]. noyau [3007]. Nr [1151, 3536]. NRF [3561]. NS [2674]. NS-Ideologie [2674]. NTM [1709]. NTM. [1709]. Nuccetelli [4320]. nucléaire [3007, 1473, 3686]. Nuclear [4079, 2472, 3110, 3220, 4172, 2638]. Nueva [3735]. Nuevos [3735]. nuit [4454, 3310]. number [3516, 4203, 4238, 833, 1791, 3107]. numérations [2212]. numerical [2506]. numerico [830]. numérique [3987]. numériques [3356]. numeris [2733]. numero [4069, 4307, 2193, 459, 1170, 1173, 2324, 4366, 510, 528, 855, 2583, 2246]. numerorum [2382, 2515]. numéros [1005]. Nummern [1470]. Nuncio [3566]. Nuncius [3694, 3522, 3936, 1666, 3523, 3583, 3923, 4072, 4067, 1403, 4051, 134, 1559]. Nuncius/Le [3522]. Nuova [3829, 2406, 4345, 3238, 248]. Nye [3094]. Nystedt [3039].

O [3010, 2877, 2191, 3000, 4203, 2157, 277, 875, 2127, 969, 1071, 1439, 2712, 2788, 1610, 4643, 529, 2855]. O. [2499]. Oberkampf [2711]. oberta [3627]. Obituary [4593]. Objectivité [4308]. Objectivity [1057, 4010]. Objects [4622]. objet [1720, 2350]. objets [4310]. oblatum [219, 331, 490]. Obligationes [3815]. oblique [1621, 1622]. obra [62]. Obrist [3651]. obscurity [3797]. Observateurs [2538]. Observation [3118, 1236, 3566, 3280]. observations [3501, 4571]. Observatis [3566].

Observatoire [4118, 4281, 2968, 4571]. observatories [4456]. Observatory [89, 1018, 4571, 3262]. observée [4138]. observer [4380]. obstacle [4247]. obstacles [4046]. obstétrique [2032]. occasion [4013, 3048, 2653, 145, 2180, 2939]. Occasional [2866]. Occident [4415, 2693, 3795, 4318, 1691, 1916, 2683]. Occident. [3721]. occidental [3626, 2592, 2851]. occidentale [1749, 2909, 3867, 1173, 1707, 2919]. occidentales [2904]. occidentaux [3047]. occultation [4485, 4138]. occultes [8]. Occupation [4156]. océane [887]. Oceanographic [2850]. Océanographie [1602]. Ocerki [2548]. Oct [2337, 3759, 3722, 192]. oct.1948 [192]. octants [4478]. octavum [219]. octo [3505]. October [3195, 404, 489, 1942]. octobre [934, 2366, 783, 2381, 2730, 4439, 3299, 2583, 160, 4022]. oculaire [3809]. O’D [1236]. Odile [2064]. Odontoiatria [2814]. Odyssey [3461]. Oehser [282]. œil [3427]. Oesterreichische [1028]. Oesterreichischen [2392]. Oestmann [3734]. œther [4082]. Œuvre [4286, 1202, 3739, 271, 925, 1515, 4473, 269, 287, 407, 432, 2001, 4600, 4381, 4110, 977, 892, 345, 2004, 1831, 2005, 4503, 2860, 1086, 1457, 198, 881, 964, 1302, 518, 492, 3112, 1424, 4658, 585, 3509, 4116, 4114, 469, 1180, 4613, 3954, 1478, 561, 2026, 2100, 2302, 4379, 3953, 1969, 471, 467, 1105, 1357, 2084, 2219, 2400, 1024, 1079, 1243, 408, 516, 64, 611, 949, 3057, 1059, 2157, 1687, 1852, 2446]. oeuvres [553, 1753, 3224, 1629, 4119, 3116, 698, 2152, 644, 1292, 1400, 714, 2196, 598, 4689, 1832, 2503, 474, 4142, 1849, 30, 1102, 4615, 2609, 303, 3122, 3937, 2591, 1681, 1781, 4642, 3241, 2140, 2247, 3239, 375, 1570, 2156]. Öffentlichkeit [4200]. Offermann [3346]. offertes [1186, 1187]. offerts [3741, 4265]. officer [4617]. Official [3307]. officiers [1755, 1834]. Ohanessian [958]. Ohm [3141, 4338]. Ohya [2187, 1174]. oiseau [3310]. Ökologie [4093]. ökologischen [4093]. Ökonomie [4093]. Olby [3337, 1616]. Old [2773, 3527, 3431]. Oldenbourg [278]. Oldenburg [1522, 3195, 1855]. Olesko [3441]. olismo [4475]. Oliver [89, 3797, 2171, 4545]. Olivier [3950, 3578, 3055, 4530, 2411]. Olms [2468]. Olschki [2976, 1571, 2787, 2551, 1034]. Olsson [2475]. Oltre [3058]. om [747, 747]. O’Malley [674]. O’May [2078]. ombre [4008, 3779]. Omero [1537]. Omnes [3372]. Omnia [2873, 2662, 939, 1773, 2208]. omnium [3316]. omoiosis [901]. Ondar [2877]. onderrichtslaboratorium [991]. Ondes [4297]. ondulatoire [2160, 1674]. ondulatoria [88]. One [3169, 3826, 1979, 3482]. O’Neil [2104]. O’Neill [3945]. Onésime [780]. Onesti [3402]. Onfray [996]. Ong [2708]. Ongaro [2882]. ons [396]. ont [3070, 2408]. Ontario [3519]. ontogenèse [3635]. Ontogeny [2447]. ontologie [4513]. ontology [4263]. Ontwakende [237]. Ontwikkelingsgeschiedenis [3767]. Opera [2873, 939, 2662, 1773, 1956, 1216, 2208]. Opérationalisme [2612]. Operationalismus [2612]. Opere [2490, 830, 3240, 3150, 829, 830, 943, 2874]. opéron [2678]. Ophtalmology [995]. Opinion [402, 3224]. Opinions [3125, 367]. Opiologia [1985]. opium [1985]. opportunism [4242]. Opportunisme [4242]. opposant [4468]. oprichting [991]. Opsomer [2266, 2267]. Optical [2216, 3424]. opticien [2894]. Optics [2024, 2232, 4627]. Optique [4015, 2516, 3071, 3302, 4329, 2206, 2304, 4328, 3595, 2102]. Optique. [3302]. Opus [1153]. Opuscula [2826]. opuscule [2266, 2048, 4084]. orang [4296]. orang-outang [4296]. Orationes [939]. orbes [4643]. orbites [2771]. orbium [4643]. Orcel [2319, 795, 1102]. order [4636, 3788]. ordinaires [2596]. ordinary [4020]. ordonnée [1253]. ordre [4304, 4636, 3208, 4532, 4129, 4288]. Ordres [3442]. Ore [828]. Oresme [3119, 4488, 1847, 1997, 2094, 1163, 4332]. Orezza [2188]. organi [3901]. organiceskoj [2548]. organicism [2294]. organicisme [3853]. organique [4547, 177, 2548]. Organisation [2431, 903, 2435, 3493, 2540, 2122]. organisé [3700, 4026, 3043]. organisée [1175, 4510]. organisées [1808]. organisés [2108]. organism [4143]. organisme [2431, 4445]. organismes [469]. Organization [2672, 3497]. Organizzato [4076]. Orient [203, 4415, 2828, 4318, 3721, 144, 879, 3237]. orientale [4068]. Orientales [249, 161]. Orientalia [4510]. Orientalist [2607]. orig [3767, 4236, 3731, 4203, 3756, 3724]. orígens [3078]. Origin [3632, 1863, 3106, 2333, 89, 1389, 4090]. Original [3420, 3077, 292]. Originalarbeiten [3077]. originale [4166, 530, 1099]. originales [1400]. originalité [3171]. Originally [3727]. originaux [1102]. Origine [357, 2194, 919, 1809, 3272, 2856, 4090, 2459, 2738, 1249, 157, 189]. Origines [3348, 4521, 1749, 692, 2812, 53, 125, 25, 56, 762, 1434, 3378, 2728, 371, 2667, 4547, 4607, 2218, 2412, 1783, 643, 1000, 1757, 1340, 850, 2561, 379, 2538, 1889, 1411, 3366, 4278, 1661, 3308, 3407, 2213, 3261, 752, 1392, 3098, 3496, 3995, 2755, 454, 1242, 2187, 2394, 1241, 2596]. origini [248]. Origins [2667, 3125, 2474, 1976, 3076, 1999, 276, 2703, 3388, 4268, 2335, 2905, 3697, 4607, 2876, 2853, 2547, 2214, 1616]. Orla [1109]. Orla-Jensen [1109]. orléanaises [3621]. Orléans [3625]. orreries [1758]. Ortega [3592]. Orthogenesis [4406]. orthopédique [4523]. Orthoptères [1456]. Orton [4626]. Osborne [89, 3638]. Oseen [13]. Osiris [845, 219, 331]. Osiris. [490, 844]. Oskar [394, 788]. Osmowa [1412]. osservazioni [4301]. Ossia [4301]. ostataka [2568]. österreichischen [1438]. Ostoya [357]. Ostwalds [2384, 2058]. Oswaldo [2378]. other [89, 3005, 3532, 1891, 1758]. Othmar [4237]. Otilie [1854]. Ott [4243]. Ottavi [4241, 4556]. Ottavio [4078]. ottica [2375]. Ottiche [2808]. Otto [4515, 3414, 3418, 4122, 1730, 2519, 2710, 1733, 3035]. Otto-Motors [1733]. ottobre [4076, 4161]. ottomans [4486]. oublié [3971]. oubliée [1047]. oubliées [4064, 1737]. Ouest [2577, 797, 2234, 1846, 4544]. Ouguag [2269]. our [3701]. Oursel [3345]. outang [4296]. outils [3919, 4281, 4286, 4618]. outline [89, 400]. Outram [2686]. Outwikkeling [990]. Ouvert [2292]. ouverte [4189, 4092, 4170, 4236, 3724, 2675]. Ouverture [4513]. ouvrage [2240, 563, 2441, 3781, 4002, 2588, 975, 893, 2318, 1762, 3780, 4548, 1472, 2436, 3429, 1559, 1087, 3877, 3876]. Ouvrages [1620, 1653, 1760, 2346, 2594, 2949, 2930, 2677, 3071, 3378, 3554, 4193, 4159, 3439, 3094, 2015, 2349]. ouvriers [3208]. Ovaert [2404]. Overseas [3154]. ovo [4202]. Ovum [674]. Owen [1334, 2550, 2474, 178]. Owsei [946]. Oxford [1432, 4448, 4578, 2537, 4570, 2703]. oxygène [1739]. oxymore [4542]. Oystein [828]. Ozigova [2830].

P [2017, 89, 104, 1540, 279, 2527, 2658, 2870, 2964, 3124, 3278, 40, 8, 1025, 2627, 3811, 2189, 2343, 2345, 2104, 2532, 2610, 1435, 1946, 2116, 2003, 3433, 3792, 586, 1259, 3945, 327, 2058, 2608, 2637, 2830, 2111, 2318, 2319, 2575, 3839, 4020, 4203, 2642, 2837, 1978, 983, 2989, 2522, 798, 1059, 1703, 2234, 2322, 2666, 2354, 2974, 2469, 2828, 109, 447, 1015, 1072, 1232, 1467, 1568, 1667, 1955, 2095, 2107, 874, 584, 601, 626, 3389, 2916, 3077, 4643, 2231, 1472, 2822, 2941, 1244, 2525, 2113, 1606, 2102, 2408, 2611, 2744, 3306, 2549, 2572]. P [208, 566, 821, 1332, 1294, 1541, 2311, 2355, 952, 1691, 2580, 2249, 1241, 3013, 1615, 3219, 2437, 1459]. P. [2765, 1992, 105, 680]. P.-H [1606]. P.-J.-F. [1992]. P.-L [2234]. P.-P [327, 2319, 2575]. P.-S [2837]. P.V.A [1848]. pôle [2206]. Pablo [4170]. Pacaud [2243]. Pachter [329]. Padoue [3161, 3264]. Padova [2694]. Padovani [4516]. padovano [3039]. Pagel [1688]. pages [10]. Pailleur [1200]. Pais [3371]. Palaeontology [3003]. Palais [1149, 1314, 1506, 1640, 1911, 403, 491]. Palama [153]. Palatino [603]. paléographie [1230, 1521]. paléographique [76]. paléoichthyologie [2498]. paleontología [3732]. paléontologie [125, 3374, 3260, 327]. paléontologique [2597]. Palerne [3504, 4003, 4363]. Palissy [1102, 178]. palladiennes [4303]. Palm [3903]. Palmer [1577]. Palter [1162]. Pamela [3655]. panacée [890]. Panattoni [4515]. Pancaldi [4457]. Panchen [3528]. pandémie [3272]. Paneth [1475, 1475]. Pang [773]. Pannekoek [1299, 396]. Panofsky [583]. Panopolis [3962]. panorama [4187, 566, 2507, 417, 1324]. Panthéon [2613]. Pantin [4643, 3522, 3566]. Panza [3793]. Paola [4098, 4262]. Paolo [2490, 4623, 2174, 829]. Paoloni [3383]. paper [1919]. Paperbacks [2468]. Papers [3270, 3663, 2529, 2834, 3703, 3886, 2565, 1929, 2057, 2096, 2959, 3077, 907, 1891, 2334, 2528, 2476, 1942, 2875, 2881, 3865, 2492, 2742, 2659, 2933, 2337, 3792, 4661, 1669, 1670, 1672, 3347, 4318, 2923]. papiers [2411, 891]. Papin [4564]. papiri [3105]. Papp [566, 1324]. Papyrus [3745, 2705, 4036]. Paquot [4416]. para [3946, 529]. parabole [2055]. Paracelse [4401, 4058, 4159]. Paracelsus [2479, 2707, 329]. Paracelsus-Forschung [2479]. Paracelsusforschung [2479, 2706, 2707]. Paracelsustag [2707]. Paracelsustradition [2479]. Paradigm [3959]. paradigme [3510]. Paradís [3078]. paradossi [3238]. paradox [4364]. paradoxal [4678]. paradoxales [4687]. Paradoxes [3897]. paradoxical [4687, 4678]. paragraphe [4308]. Paraguay [32]. Parallèle [4384, 3829]. Parallelenproblem [1750]. parallèles [2828, 379, 3147, 3056]. parallélisme [3189]. paralysie [1916]. parameters [3501]. paraphrase [3516, 247]. parasite [2499]. Parasites [997]. parasitic [997]. paratonnerre [53]. parcours [4280]. Pardies [2240]. Pardo [3735]. Paré [2097, 3146, 460]. parfait [4031]. parfaite [227]. Paris [1538, 485, 2183, 2423, 2058, 4022, 2319, 2044, 1175, 1706, 1938, 1939, 2380, 3213, 4080, 2668, 2844, 3166, 2811, 4510, 3728, 3221, 4363, 2377, 3435, 1574, 3633, 278, 530, 766, 1294, 1031, 458, 486, 2613, 3715, 1582, 182, 2050, 4653, 1808, 3229, 1463, 2968, 4571, 1699, 2086, 2163, 3784, 2204, 3657, 1948, 3623, 3093, 3230, 891, 4118, 1460, 4670, 4289, 2318, 161, 217, 1304]. Paris-Province [3229]. Parisien [436, 1050]. parisienne [2412]. parisiennes [1400, 3714]. parisiens [1443, 4397, 4647, 3215]. Park [235, 798, 2521, 2737]. Parlement [4396]. parler [4219]. Parma [689]. Parme [4132, 4133]. Parmentier [4531]. parmi [3398]. Parot [4415]. Parrochia [3442]. Parsons [89]. Part [1265, 3516, 2190, 2640, 3482, 3077, 3177, 2476, 162, 2275, 2276, 1979, 2117, 3676, 4076, 4435, 3363, 4206, 2921, 3080, 1893, 3004, 2444, 250, 610, 220]. Parte [3240]. Partes [2921]. parthénogenèse [596]. participaient [3408]. Participation [4226, 1424]. Particle [3902, 3165, 2836]. particles [4598]. Particular [1019, 4663]. particules [4598, 3234, 3318]. particulier [179, 4663, 2064]. Partie [944, 2996, 2953, 2969, 4222, 2101, 1268, 3962, 4146, 2073, 4223]. partielles [3331, 2622]. Partington [1829, 1979]. partir [2600, 3427]. partnership [4315]. Parts [2264, 162, 2116, 96, 404, 489]. paru [2285]. parues [2265]. parus [1005]. parvenir [3711]. Pascal [4526, 4645, 4360, 4644, 4032, 4492, 4319, 1437, 1775, 1570, 3596, 121, 4218, 1189, 1188, 1181, 2255, 1190, 3782, 1183, 1182, 4586, 3928, 1200, 1180, 1211, 1425, 1722, 1184, 1202, 552, 1179, 61, 1191, 1208]. Pascale [4286]. pasigraphy [4579]. Pasquins [1479]. passage [4627, 2169, 2440, 2437, 1007, 176]. Passages [4142]. passé [3292, 2880]. Passermeetkunde [373]. Passeron [4615, 4265]. Passioni [4072]. passions [4080]. Past [3589, 2036, 2182, 3194, 4661]. Pasteur [34, 63, 3360, 3965, 3454, 1109, 1117, 1242, 17, 469, 4191, 2073, 34, 38, 864, 363, 3151, 1832]. Pasteur-Valléry [34]. pasteurienne [2340]. pastorienne [3151]. Pastoureau [1391]. Pasumot [3476]. Patai [3652]. patent [89]. Patents [89]. Paterson [2336]. path [4653]. pathologic [4676]. Pathologie [1689, 2397, 4199]. pathologique [3188, 2919, 4676, 15]. Pathology [2069]. Patients [3459]. Patologica [2644, 4083]. Patrice [2812, 3936, 3937, 3884, 4620]. Patricia [4172]. Patrick [3533, 693, 3940]. patrie [4065]. Patrimoine [3052, 4164, 4259]. Patriot [947]. patron [4566]. patronage [1079]. patrons [4425]. Pattaro [3491]. Patterns [729, 4662, 900]. Paty [4558, 2751]. Pau [1943]. Paul [3972, 3635, 3632, 2964, 3124, 3815, 2753, 43, 70, 91, 93, 1102, 2342, 2227, 2253, 2609, 2451, 2059, 959, 3271, 282, 304, 1892, 330, 4261, 4196, 3123, 3302, 3755, 4267, 352, 357, 3650, 3841, 1960, 4091, 1833, 2660, 1061, 1303, 1702, 2646, 109, 151, 481, 985, 988, 1015, 1072, 1327, 1330, 1467, 1667, 1955, 2107, 3525, 954, 4531, 626, 2539, 4392, 3197, 3280, 2538, 1787, 524, 1217, 4350, 1735, 4147, 1100, 4291, 3307, 3668, 392, 2353, 2373, 1292, 3015, 2634, 951, 4246, 4251, 1516, 4387, 2585, 1904, 2109, 1518, 4638, 871, 1051, 930, 3688, 3509, 3094].

Paul [2633, 4364, 2955]. Paul-André [4291]. Paul-Antoine [1102]. Paul-Henri [1217, 1516]. Paul-Joseph [2585, 1904]. Paul-Louis [1061]. Paul-Marie [3197, 2353]. Paula [4616]. Paule [4007, 3146, 1304]. Pauli [3006]. Paulo [219]. Pavia [2882]. Payen [3487, 3261, 1730, 1728, 2066, 1437]. Payne [3005]. Payne-Gasposchkin [3005]. Payot [3272, 1472]. Pays [2448, 841, 67, 1880, 1312, 3147, 4164, 774, 413, 474, 440, 502, 523, 579, 664, 724, 782, 935, 1229, 1533, 1679, 10, 298]. Pays-Bas [841, 413, 474, 440, 502, 523, 579, 664, 724, 782, 935, 1229, 1533, 1679, 10, 298]. paysannes [4028]. Peano [4351, 830, 1099, 830, 943, 831]. Pearce [2171, 1999]. peau [4283]. Pecham [2024, 2158]. Pecker [3178, 1172]. pectories [800]. pédagogie [117]. pédagogiques [3821, 766]. pédagogue [3213]. Pedersen [730]. Pédiatrie [2844]. pédologie [2763]. Pedro [3735]. Pehr [12]. Peiffer [2775, 3833]. Peirce [4239]. Peireisc [3458]. Peiresc [10, 4063, 2374, 386, 288]. Peirescel’ [1567]. Pékin [92]. Pélerin [3812]. Pell [1949]. Pellegrin [2839]. Pellet [1638]. Pelling [2537]. Pelseneer [10, 484, 135, 130]. Pelton [1340]. Pena [4402]. pendant [497, 4156, 507, 761, 1480, 4389, 177, 277, 1155, 286]. pendules [1420]. Penel [2851]. Penetralia [3674]. pénicilline [3765]. Pénisson [3529]. Pennsylvania [1317, 1485, 4572]. Pennsylvanie [4572]. pensable [3646]. pensant [4196]. Pensée [1217, 1014, 4310, 3739, 1749, 2925, 1515, 135, 2001, 3934, 1434, 444, 1396, 4026, 138, 1960, 1662, 2708, 132, 604, 1337, 1624, 4319, 2205, 2285, 3964, 3595, 4564, 23, 393, 1536, 3999, 4051, 1865, 1241, 3013, 678, 3876, 1143, 1386, 1387, 1751, 1606, 1607, 1329]. Pensées [2164, 4219, 4296]. Penser [3927, 3745, 4556, 4035, 3863, 4004, 3834, 4176, 4199, 3938, 4208, 3765, 4011, 2741, 3530]. Penseurs [981]. Pensiero [2395, 4475, 3491, 4122, 2739]. Pepe [2874]. peponthoton [3503]. percussion [1209]. Percy [1419]. perdus [1601, 2098, 2149]. père [4119, 3809, 1168, 4668, 2240]. pères [4017, 1158]. Perez [4025, 3484, 4625, 2635]. perfection [4031]. perfectionis [3651]. perfectionnement [1251]. performance [1934]. Perge [1330]. Peri [3503]. Pericle [3240, 1571, 994]. Périer [1201, 1723]. Perifano [4081]. péril [4564]. Period [4609, 4590, 2523]. période [497, 4590, 2412, 286, 196]. Périodique [1257, 1611]. Périodiques [1653, 1594, 162, 163, 164, 2189, 2193, 2274, 2296, 2297, 2343, 2345, 490, 537, 592, 593, 616, 617, 632, 633, 696, 801, 843, 844, 845, 846, 866, 888, 972, 973, 974, 1004, 1033, 1034, 1035, 1121, 1122, 1123, 1151, 1152, 1178, 1315, 1316, 1317, 1318, 1444, 1445, 1511, 1484, 1485, 1486, 1507, 1508, 1509, 1510, 1618, 1619, 1710, 1796, 1797, 1798, 1837, 2038, 2039, 2040, 2042, 2043, 2116, 2132, 2133, 1270, 2170, 1617, 1264, 1355, 1421, 306, 96, 141, 219, 250, 332, 307, 331, 383, 404, 488, 489, 1005, 1076, 2070, 971, 220]. périodiques [2593, 2614]. periodo [565, 1907]. périr [3623]. permanence [2904]. Permuda [1029]. perpétuels [4385]. Perrault [3305]. Perrier [3345]. Perrin [1902, 1901, 1900]. Perrot [3279]. Perry [3481, 2528]. persécutions [1775]. Personalisation [4053]. Personnel [2262, 3043, 3230]. personnels [1156]. Perspectiva [2024, 2158]. Perspective [2508, 4546, 2001, 1328, 517, 4015, 4333, 821, 3995]. Perspectives [4190, 4260, 3500, 825, 3367, 2052, 1026, 1830]. Perthes [3308]. Pertinentes [2873]. Peruzzi [3813]. pervaya [983]. perversion [4624]. Pesanteur [1437]. Peset [2635]. Pesic [4321]. Pessel [4526]. pessimism [3726]. peste [1941]. pestilentielles [2234]. Pestre [4311]. Peter [3944, 2505, 3991, 3123, 4321, 3751, 2450, 2670, 3125, 4347, 3808, 3364, 2849, 3861, 4166, 2784, 2829, 2978, 2391, 3591, 4002]. Pétersbourg [1739]. Petersburger [1332, 2355]. Petit [1170, 1241, 2675, 4266, 2122, 2091, 4509]. Petit-Thouars [2122]. Petite [3585, 1472, 4669, 279]. Petitjean [3533]. petits [4012]. Petra [3316]. Petrache [2368]. Petrarch [2373]. pétrole [2348]. Petruccioli [3238]. Petruck [1566]. Petrus [3015, 161]. peu [2266, 1951]. peuples [3529]. peur [2673, 4011]. Peurbach [4548]. Peut [619]. Peut-on [619]. Peyrard [1629]. Pezenas [4017, 4119]. Pfeffer [2580]. Pfeiffer [2508]. Ph [2268]. Pham [279, 2646]. Pham-van-ghê [279]. pharaonique [3745]. pharmaceutical [3587]. Pharmaceuticals [3589]. pharmaceutique [692, 3949]. Pharmacie [279, 1311, 752, 462]. pharmacien [661]. pharmaciens [4547, 3969]. Pharmacopoeias [4230]. Pharr [487, 487]. phénomène [135, 4452]. phénomènes [2819, 157, 2913, 3924, 3141, 858, 2628]. phénoménologique [902]. Phil [4097, 2392]. Phil.-Hist [2392]. Philadelphie [743, 1206]. Philalethes [3642]. Philibert [2380]. Philip [821, 3877]. Philipp [3153, 4654]. Philippe [4513, 4614, 3969, 3934, 4004, 2512, 301, 597, 4049, 4398, 4643, 4629, 3938, 1352, 4051, 4124, 2649, 3968, 3591, 777, 1557, 467]. Phillips [3501]. Philo [4229]. Philologie [2582, 2705, 2669, 2693]. Philologische [4093]. Philos [3815, 3700, 1485, 3699, 3664, 3924, 3722]. philosohy [4254]. Philosoph [605, 3123, 3488]. philosophe [3631, 3079, 1673, 1781, 3704, 585, 2407, 715]. Philosopher [455, 1931, 1499, 2903, 2871, 3521]. philosopher-scientist [3521]. Philosophes [651, 1478, 1541, 4641, 4319, 4191, 504]. Philosophia [3482, 3637, 3640]. philosophiae [4219, 4217, 4221, 4223]. philosophic [4582]. Philosophica [3677]. Philosophical [743, 1973, 2320, 3411, 3703, 1259, 2321, 2641, 947, 2921, 3439, 1893, 610, 3604, 4206, 3860, 3016, 3830, 1239, 2390]. Philosophicall [2564]. Philosophie [931, 1053, 1533, 1561, 1767, 1861, 4450, 4512, 4175, 2467, 2422, 4624, 1033, 1444, 740, 1091, 3433, 4679, 275, 2423, 3123, 4008, 3824, 4234, 4641, 1659, 4432, 2287, 1786, 3443, 2288, 1827, 376, 504, 1536, 2213, 2404, 1206, 3827, 1433, 827, 1784, 2612, 2246, 3346, 4549, 587, 3578, 1785, 1972, 4132, 3026, 4410, 3359, 1633, 3349, 3343, 4266, 20, 93, 95, 3872, 4616, 2253, 3645, 3116, 2967, 4222, 3963, 3515, 2455, 3313, 239, 445, 3970, 4444, 3664, 4642, 2999, 2195, 3241, 4673, 4084, 4529, 4556, 2300, 3357, 558, 3938, 3030, 3224, 1082, 1478]. philosophie [902, 3856, 2108, 4455, 385, 4162, 2983, 3239, 3633, 4666, 2184, 4620, 1828, 4509, 4558, 378, 4621, 3669, 4168, 4298, 4491, 4171, 1386, 3664, 4510, 4293]. Philosophies [1972, 4233, 4368, 3807, 2076, 1954, 2972, 3968, 3747, 3075, 3722, 4189]. philosophique [3457, 4171, 4199, 2301, 2352, 4366, 4114, 4552, 4129, 822, 4513]. philosophiques [598, 4050, 1849, 3121, 698, 4545, 4451, 1487, 4642, 2247, 4066, 378]. Philosophisch [2287]. Philosophisch-historische [2287]. philosophische [4234]. philosophischen [1684]. Philosophischhistorische [1028]. Philosophy [1931, 3519, 3697, 3578, 2320, 3533, 3501, 3384, 3411, 3696, 3650, 2565, 2709, 3363, 2522, 2336, 3126, 1468, 3860, 3002, 3500, 2226, 3432, 2528, 1827, 3307, 3891, 3120, 631, 2403, 2551, 4365, 3880, 3826, 4144, 3315, 3371, 3991, 3990, 3755, 3886, 3590, 3751, 3482, 3637, 3640, 3916, 3317, 4207, 3794, 4673, 3762, 3743, 3699, 2983, 4666, 3388, 609, 4679, 2870, 907, 65, 4254, 4024, 3484]. Phinney [216]. phlogiston [3930]. photographie [592, 2680]. photographies [1145]. photographique [530]. photométrie [1423, 4008]. Photosynthesis [328]. Phragmén. [4049]. phrénologie [4208]. phylogenèse [3635]. Phylogeny [2447]. Physical [2262, 2192, 1944, 2928, 3118, 2506, 2876, 3420, 2419, 2127, 2495, 2831, 156, 359, 1565, 3002, 3275]. Physician [2903, 1499, 4067]. physicien [571, 3708, 3687, 2207, 4141, 2900, 3398, 1655, 2204, 2897, 1900]. physiciens [4250, 119]. physicist [1931]. physico [90]. physico-chimique [90]. Physics [2061, 3696, 2638, 1565, 2979, 3002, 3037, 3165, 2127, 89, 37, 3902, 3371, 3410, 3693, 3482, 2151, 3871, 4666, 2338, 789, 3243, 1239, 2702, 2923, 3219, 3315, 3441, 2747, 2262, 3701, 3441]. Physik [2958, 3443, 2663]. Physiker [605]. Physikers [3156]. Physikgeschichtlichen [2701]. Physikunterricht [2701]. Physiocratie [330]. Physiological [1110, 3965, 3080]. Physiologie [1689, 533, 10, 63, 2576, 3919, 4483, 3954, 2317, 347, 533, 1569, 2916, 2665]. physiologie-outils [3919]. physiologiques [671, 4123, 4356, 3827]. Physiologist [328]. physiologiste [495]. Physiology [2685, 1110, 3482, 3640]. physiopathologiques [3919]. Physique [4332, 468, 3007, 3844, 3318, 2745, 405, 104, 3631, 3874, 2516, 3140, 135, 2783, 4297, 4218, 652, 2255, 399, 3669, 4502, 3348, 4195, 4448, 3749, 198, 3709, 2561, 4280, 4250, 1473, 4338, 2151, 376, 2443, 2698, 3825, 4166, 902, 1202, 1901, 3798, 1865, 4666, 834, 1335, 318, 4333, 2350, 1785, 3659, 2744, 72, 119, 3372, 4494]. physiques [732, 398, 1626]. Physis [2297, 1618, 2040]. Physis. [1034, 1178, 1509]. Pi [666]. Piaget [4241, 2824]. Picard [3273, 4094, 2237, 2364, 3633, 3783]. Pichevin [797]. Pichevin-Chatelin [797]. Pichot [3639]. Pickering [3705]. Picolet [3273]. Picon [3670, 3723, 3305]. Picone [3759]. Picot [1918]. Pictet [3348]. Pictorial [1078]. Pie [2808]. Pièces [700, 4296]. Pieere [3730]. Piélat [3724]. Pier [2808, 3841]. Pierce [420]. Pierre [2458, 3225, 4092, 203, 1849, 2873, 3969, 4475, 2169, 1434, 140, 728, 3121, 3153, 371, 399, 1523, 2770, 4628, 2318, 3751, 2502, 2591, 350, 2170, 1143, 1480, 2188, 2448, 61, 239, 770, 1753, 2137, 4474, 582, 2999, 585, 4080, 4139, 4319, 2247, 4643, 3529, 786, 3612, 3698, 1351, 3577, 4140, 209, 504, 2404, 2675, 2745, 2386, 3091, 1073, 1118, 1335, 2291, 2881, 1112, 1146, 1165, 1166, 1830, 2187, 1241, 3244, 458, 2210, 1474, 4307, 268, 4468, 2415, 1697, 4381, 2799, 4016, 4464, 4465, 4467, 4306, 3718, 2176, 4638, 108, 2137, 2178, 1512]. Pierre [1423, 4300, 4570, 3202, 2768, 2955, 1881, 2350, 3939, 1169, 1168]. Pierre-Charles [4381]. Pierre-Julien [1351]. Pierre-Paul [4638, 2955]. Pierre-Simon [2881, 3939]. pierres [4009]. Pietro [3153, 3058, 3301, 2806, 3306, 2855, 4136, 3240, 1571, 994]. pile [145]. Pillai [3306]. Pinault [3417, 4396, 3197]. Pinch [3312, 3724]. Pinel [3591]. Pinero [2644, 3735, 2926, 3732, 2965]. Ping [606]. Pingree [3650]. Pio [2808]. pionär [747]. Pioneer [747, 2872, 89, 950, 512]. Pioneers [2702]. pionnier [2451, 1983]. pionniers [926, 4531, 1612]. Pisa [3503, 3651, 3414]. Pise [481]. pistes [3474]. pistol [4663]. pitagorico [901]. Pitagorismo [2606]. Pitane [2627, 3665]. Pitot [518]. Piva [3568]. Piveteau [327, 484]. Piyo [3653]. Pizon [2170]. Pizzamiglio [2808]. Pla [88, 245, 62, 87, 136, 300]. Place [2257, 2837, 1895, 3856, 4506, 1016, 496, 3818, 4613, 1018, 4665, 4166, 2273]. placée [1079]. Places [3125, 460]. Placido [2814]. plagiat [1819]. plague [4670]. Plaies [4007]. plaisants [986]. plan [2508, 3952, 254, 4336]. Planches [1408, 314]. Plancius [161]. Planck [2022]. planes [3791, 2055]. Planet [1236]. planétaire [3842]. planétaires [3178]. Planetarians [2546]. Planetarium [3996]. planetarum [4548]. Planetary [1334, 938]. Planètes [2771]. Planets [2546, 2569]. Planning [4633]. plans [3227]. Plant [485, 328, 3319]. plantas [3735]. Plantes [73, 1762, 2852, 272, 279, 435, 3375, 2267, 2122, 3736]. Plants [485]. Plata [154]. plate [2657]. Plato [3304, 2803]. Platon [582, 2206]. Pleasing [2524]. Pledge [37, 1631]. plein [3310]. Plélat [3654]. Plessis [3591]. Plessis-Robinson [3591]. Pline [588]. Plinio [3994, 1614]. plis [2954]. ploasticité [4480]. Plon [2404]. plongée [1342]. Plumb [2310]. pluralité [704]. plus [584, 2852]. Plusieurs [1948, 2408, 1541]. plutôt [4546]. Pluto [2569]. pneumatics [4687]. pneumatique [4687]. poco [4083]. Pocock [3060]. Poczobut [4544]. Poemi [4069]. Poenaru [2368]. poésies [4296]. poetics [3482, 4256]. poétique [3258]. poétiques [4205]. Poetry [2838]. Poggi [4207]. Pogrebysski [1970]. poids [4157, 1917, 225, 3511]. Poincaré [670, 4553, 4242, 4580, 4049, 4545, 549, 4430, 4046, 1633, 4242, 4641, 2780, 4347, 2213, 2387, 4447]. point [2338, 923, 4134, 2856, 3508, 3174, 4336]. Points [2741, 2751, 2752, 4164, 4553, 1963]. Points-Sciences [2751, 2752, 4164]. Poirier [4267, 4091]. Poison [2883]. Pol [2436]. Polacca [2606]. polaires [4597]. Poland [2154]. polar [4597]. polarimétrie [3092]. polarisation [4505, 4568, 2833, 2206]. polemic [3424]. polemica [4072, 4068]. polemical [4317]. polémique [407]. polémiques [195]. Polge [2665]. police [4561]. Policy [2378]. Polish [4056, 3317]. politica [4068]. political [2846]. Politics [4144, 3535, 3482, 4611, 4010]. politique [4542, 4502, 2517, 4451, 3970, 4396, 4611, 3756, 4030, 4518, 3698, 2404, 2673, 4288, 3382, 3654, 4081]. Politiques [4237]. Polity [2773, 3371]. pollutions [4184]. Polnischen [1864]. Polnohospodarskeho [1355]. Pologne [884, 1090, 1952, 1557, 1287]. polonaise [1463]. polovina [983]. Polubarinova [2610]. Polubarinova-Kochina [2610]. polynômes [2455, 2420]. polytechnicienne [3723]. polytechnique [3254, 2377, 3168, 2898, 2540, 3044, 3167]. Poma [4196]. pompe [3654, 1723]. pompes [1583]. Poncetton [21]. ponctualisme [2952, 4446]. ponderibus [689]. Pont [4516, 3970, 3426, 3793, 2213, 3056]. Ponte [1118]. Ponts [2901]. Ponty [2060, 3557]. Poor [3459]. Popper [3343, 4144, 3676]. Popular [2696, 3274]. population [4268]. Populations [3474]. porc [2073]. Porras [3946]. Porret [4390]. Porset [4296]. Port [4492, 1191]. Port-Royal [4492, 1191]. Porta [1333]. portant [1230, 1521, 1846]. portatifs [3293]. Porte [534]. portée [4051, 4124]. Portela [2926]. Porter [3434, 2190, 2445]. Portrait [1856, 346, 2598, 700, 1321, 1046]. portrait-robot [2598]. portraits [613, 784, 2129]. Portsmouth [2529]. Portugal [529, 764, 1055]. Portulans [2962]. Poseïdon [154]. position [3028]. positive [4154]. positivism [4604]. positivisme [3601, 4604, 4390]. positiviste [4247]. positivistes [4266]. Possession [1259]. possible [3198, 661]. post [3788]. post-revolutionary [3788]. posthume [893, 585, 1087]. Posthumous [1975, 1664]. postmodernism [3726]. postremas [3148]. Postulat [975]. potentia [4688]. Potier [3969]. Pouchet [1764]. Poudres [3936, 3784, 2047]. Poulle [1538, 3097, 3842, 4397, 1346, 2929]. Poullet [530, 530]. Poullet-Delisle [530, 530]. Poulsen [3150]. Pour [724, 3349, 138, 4360, 553, 4241, 2211, 4469, 782, 765, 935, 1053, 1247, 1229, 2123, 3225, 4142, 4612, 279, 1949, 3711, 4468, 3721, 3356, 125, 1808, 3872, 1375, 556, 4311, 4664, 948, 893, 3936, 4160, 4380, 2576, 2502, 4065, 472, 151, 481, 4184, 2258, 4362, 2076, 1954, 4250, 3074, 249, 3047, 3357, 4247, 3884, 3986, 4296, 4145, 3633, 2283, 530, 552, 1105, 2398, 388, 3530, 2613, 3798, 395, 3608]. Pourcin [4558]. Pourquoi [753, 4587, 3091, 3943]. pourtant [3672]. Pouvoir [2634]. Powell [3301]. Power [3459, 673, 1891]. Poysson [106]. pp [278]. Pr [139, 513, 796, 797, 1241]. practicae [213]. Practical [4591, 2777]. prácticas [4453]. practice [4653, 3441, 3630, 3646, 4561, 4489, 3274]. Practices [4077, 4591]. Practique [3812]. practitioners [3868]. Pradeu [4559]. praeparationum [2706]. Prag [1929, 2096, 2742]. Prague [1748, 1626, 1854]. Praha [1854]. Prat [3625, 969]. praticiens [3047]. Pratike [2777]. pratique [1987, 4553, 4306, 792, 3646, 3494, 2693, 668, 839, 969, 3643]. Pratiques [3360, 4028, 4525, 4412, 3415, 4181, 3691, 4547, 4021, 1579, 4591]. Pratt [1934]. Praxis [2478]. Prazské [960]. Pre [2321]. Pre-Vesalian [2321]. préagronomiques [4028]. préambule [1840]. précartésiens [4026]. précédé [3302]. préceptes [2660]. Précis [4621]. Precision [3501, 3273, 3345, 1249]. Précisions [125, 409]. précolombienne [1937]. préconcues [656]. précurseur [4541, 226, 4663, 2952, 814, 4292, 4158, 661, 978, 2026]. précurseurs [920, 1225, 3056, 754]. precursor [4663]. prédarwinienne [2497]. prédécesseurs [4657]. predecessors [4657]. prédiction [4363]. Preface [3755, 1294]. préfets [774]. prefixed [1752]. Préhistoire [2022, 3254, 2813, 3147, 120, 3308, 1761]. préhistorique [1313, 2919, 1260, 4201]. Prehistory [1261, 2928]. préliminaire [4120]. prémices [3585]. premier [3334, 2397, 4022, 2412, 1602, 151, 4130, 2293, 1112]. Première [2110, 944, 2953, 592, 1131, 4441, 4427, 151, 481, 177, 2540, 1472, 1993, 4062, 3234, 4523, 1089, 467, 1490, 2797]. Premières [2194, 3356, 4484, 2366, 1282, 3178, 3980, 1154, 1608, 2857, 1989, 1889, 385, 437, 2135]. premiers [178, 4548, 2293, 1136, 1932, 3760]. préparation [1990, 2273, 747]. préparé [1762]. préparés [948]. Présences [4403]. Present [1432, 2182, 3194, 2036, 1208, 2880]. Présentation [4299, 4353, 4327, 4400, 2239, 3230, 1867, 4548, 3546, 3777, 3997, 3397, 4245, 4217, 3980, 1840, 3913]. presented [3603, 2155]. présentée [3221]. présentées [3578, 3700, 3530]. Preserve [3579]. preserved [2686]. presidente [3732]. presque [2950]. Press [1594, 2137, 3864, 2696]. presse [2974]. Presses [3933, 3932]. pression [3138]. presso [2808]. Prestet [924]. présumé [878]. présupposés [3045]. prétendue [81, 4337]. preuve [4174, 4027]. Prevention [2578]. Prévost [4485]. Prévot [2503]. Price [906, 1017]. Prigent [1245, 1579]. Prigogine [2593, 3413]. prima [2770, 3240, 2231]. primenama [2568]. primera [885]. Primeras [3735]. primero [1542]. primitif [2157]. Primitifs [135]. Primitiven [2031]. Primo [2490, 248]. prince [774]. Princeton [2834, 2137, 2137, 2924]. principî [1074]. Principal [1590, 567, 856, 1299]. principalement [797, 969]. Principales [2921, 3224, 474]. Principauté [1602]. Principaux [1373, 698, 1008, 3596]. principe [4308, 3820, 3353, 1811, 3600]. Principes [4222, 3199, 69, 4685, 2461, 2175, 4581, 909, 1828]. Principia [3807, 3065, 4224, 3063, 4219, 4218, 3064, 2981, 4217, 4393, 4220, 4221, 521, 437, 4223]. Principien [3384]. principles [4685, 4581, 2870, 2030, 1566]. Printed [2581, 2686]. priorités [195]. prirodne [2184]. prisme [2761]. prisons [3596]. Privas [2984]. Prix [853, 205, 1335, 1084]. Prize [2702]. Prizes [2960]. pro [3531, 2979]. Probabilidad [3437]. probabiliste [4584]. probabilistes [3449]. probabilistic [4584]. probabilités [2455, 3291, 2404, 2469, 155, 243]. Probability [3281, 2923, 2877, 4203, 3434]. probable [2856]. Probierbüchlein [211]. Problem [3728, 4582, 4676]. problema [901]. Problemata [2612, 1848]. problematic [3368]. Probleme [2058, 4310, 234, 2459, 3422, 3136, 2002, 179, 3066, 2997, 3028, 406, 2055, 3160, 4676, 1007, 2157, 2288, 106, 4336, 4582, 4341, 52, 256]. Problèmes [3170, 3186, 4138, 15, 1086, 236, 1673, 4116, 4059, 1283, 2408, 1026, 4310, 3882, 2453, 2813, 986]. problemi [2395]. Problems [3347, 982, 1110, 3608, 3303]. Procedures [1301]. Proceedings [3083, 3362, 3965, 3368]. procès [1819, 629, 2809]. process [3944, 4210]. Procissi [603]. Proclus [151, 4304]. prodiges [2097]. Production [3362, 3551]. Productivity [2262]. Products [2452]. produits [736]. professées [135]. professeur [436, 1050, 1227].

Professeurs [3435, 3229, 4441]. Professional [3274]. Professionalizing [2811]. professore [3414]. professors [1752]. profil [3687]. Profili [4345]. Prognostikon [3314]. programme [1949, 3929, 3200]. Programmes [3703, 3227, 1084]. progrès [2298, 279, 967, 446, 1270, 2576, 996, 2756, 4293, 656, 461, 772]. Progression [435]. Progymnasmata [3225]. Proiettiva [189]. proïzvedenijà [601]. project [4587, 292]. projectiles [3068]. projective [4074, 1180]. projet [1375, 4386, 4160, 4065, 4540, 4587, 3954]. Projetées [1095]. Prolegomena [3637]. Prolégomènes [2868]. prolongement [1696]. prolongements [2938]. Prométhée [3993, 4646]. promotion [2056]. promptitude [2516]. prononcée [4439]. pronostics [419]. Prooemium [3755]. proof [4321]. proofs [3890]. prop [3066]. Propagation [4452, 4598, 2155]. propensions [3676]. prophète [2774]. Proportion [3814]. Proportionibus [789]. proportionnalité [3499, 3492, 3511]. proportionnelle [3355]. Proportions [3725, 4136]. Propos [1738, 3761, 1579, 2430, 2240, 2164, 308, 563, 760, 2139, 2687, 222, 2415, 2441, 3267, 977, 496, 1676, 3573, 3071, 1968, 2689, 2588, 3378, 4663, 4222, 3554, 2466, 1986, 198, 1489, 406, 2679, 2367, 54, 2137, 4337, 1205, 2160, 3044, 105, 468, 4064, 3439, 1321, 558, 1742, 4001, 1425, 2654, 176, 3094, 4059, 1223, 3429, 661, 1134, 1047, 2208, 2861, 780, 2499]. proposition [4222, 3296, 4135, 3497, 3497]. propositions [3403]. proposto [2808]. Propounded [2054]. propres [834]. propriétés [2408]. proprioception [4153]. propulsive [1348]. prospecte [2368]. prospective [4521]. Prosper [4495]. prospettiva [2629, 2991]. protactinium [2600]. protagonistes [3348]. Protection [2170, 4542]. protéines [3075]. proteins [4095]. Proteinsynthese [4095]. protoplasme [4073]. Protozoaire [2499]. Protozoaires [1965]. Proust [4328, 1850, 2573]. provenance [3120]. provencal [77]. Provence [3266, 2782]. Province [3229, 2409, 957]. provinciale [1306]. Provine [4268]. provirus [3472]. provisional [905]. Provost [1003]. Prunelle [2563]. Prusse [3812]. Pruteanu [959]. pruténiques [3661]. Pseudepigraphic [3362]. Pseudo [2320, 2473, 3651, 1113, 2440, 2437]. Pseudo-Geber [3651]. pseudo-Héron [2440, 2437]. Pseudo-Hippocratic [2320]. pseudo-hippocratique [1113]. Pseudo-Thomas [2473]. pseudochemical [3642]. pseudohippokratische [3488]. psichiatrico [4515]. Psicoanalisi [2941]. Psicogénesis [154]. psychanalyse [693, 3530]. psyche [4573]. psychiatric [4662, 4656]. psychiatrie [3591, 4477, 4660, 4663, 4483, 3530]. psychiatrique [1987, 4662, 4656]. Psychiatry [4660, 2322, 4663, 4661, 2886]. psychisme [4573]. Psychogenèse [2824]. psychologie [2314, 4624, 4354, 3594, 4335, 4657, 4241, 4350, 4407, 3738, 4443, 4573, 4176, 4350, 3594]. psychologique [923, 459, 4356, 4343, 2973]. psychologiques [459, 1775]. Psychology [970, 3482, 1442, 3439, 4657, 4055, 4573]. psychophysiologie [2021]. Ptolemaeus [302]. Ptolémée [3604, 2856, 4015]. Ptolemy [3604, 3501, 3014]. pu [4586]. pubblica [2874]. Pubblicazioni [2232, 3674, 2551]. pubblico [4262]. Publ [2395, 3530, 3596, 3727, 3097].

Public [1020, 2846, 3836, 3967, 3208]. pública [2965]. Publication [1729, 1104, 760, 617, 1492, 2152, 4301, 4619, 1513, 192, 307, 383, 488]. Publications [697, 908, 1080, 1446, 1870, 1961, 2046, 2118, 2235, 2496, 2676, 2759, 2906, 2966, 2980, 3084, 3108, 3179, 3283, 3390, 3445, 3537, 3613, 3678, 3707, 3768, 3845, 3905, 3973, 4037, 4099, 4180, 4211, 4272, 4370, 4442, 4496, 2812, 4242, 1524, 4049, 2670, 4545, 1847, 1997, 2094, 886, 4530, 2024, 2158, 4348, 2405, 1574, 4265, 1566, 878, 2703, 3267, 4090, 1780, 3587, 2653, 3307, 1294, 2866]. publics [3558, 3168]. publié [2318, 3208]. publiée [2193, 2296, 633, 696, 974, 1033, 1035, 1444, 2043, 2132]. publiées [375]. publiés [1640, 355]. publique [2411, 1175, 3216]. publishing [4301]. Puech [3747]. Pugh [1244]. puissance [2509, 4550, 834]. Pump [3192]. pupils [3868]. pure [2100]. pures [795, 2507]. purgatoire [2732]. Purkert [3203]. Purkyna [1748]. Purkyne [947]. Pursell [1888]. Pushing [4260]. Puy [1203]. Puyvallée [285]. Pyke [3364]. Pyramus [3821]. Pyrmont [747]. Pyrmonter [747]. Pyrmonter-Valten [747]. Pythagore [4297]. Pythagoreer [397]. pythagoricienne [524].

qarn [2372]. Qing [4653]. QP.6 [4602]. Quaderni [210, 3813, 2406, 4068]. Quadrangles [2826]. Quadrans [2921]. quadrants [4478]. Quadratura [4012]. quadrature [1621, 1622, 2055]. Quadric [133]. Quadripartitum [2382, 2515]. quaestiones [2287, 1163]. qualitatif [1515]. Qualities [1847]. Quality [3501]. Quand [3641, 4573, 4382, 3646, 4030, 4526]. quanta [2751]. Quantification [3864]. Quantifier [4415]. quantique [3578, 137, 4471, 3263, 4666, 4604, 405, 3081]. quantiques [2750]. Quantum [3411, 3006, 3578, 3898, 3944, 3110, 4666, 4604, 4507]. Quarta [2662]. Quarterly [1264, 1486]. quaternions [1698]. Quatre [3265, 2587, 1579, 3146]. Quattrocento [4476]. Que [1214, 140, 4032, 1613, 242, 1327, 1328, 1634, 3007, 1352, 2736, 2756, 1326, 1932, 363, 4218, 3724, 4581]. Québec [3591]. Quelle [3047, 4644]. Quellen [4093, 1605]. Quelques [1081, 658, 3402, 255, 1523, 4128, 2862, 347, 4428, 4252, 223, 4159, 3952, 105, 1401, 3803, 1225, 755, 4047, 409, 1594, 162, 163, 164, 2139, 4334, 2379, 15, 3499, 2189, 2193, 2274, 2296, 2297, 2343, 2345, 490, 537, 592, 593, 616, 617, 632, 633, 696, 801, 843, 844, 845, 846, 866, 888, 972, 973, 974, 1004, 1033, 1034, 1035, 1121, 1122, 1123, 1151, 1152, 1178, 1315, 1316, 1317, 1318, 1444, 1445, 1511, 1484, 1485, 1486, 1507, 1508, 1509, 1510, 1618, 1619, 1710, 1796, 1797, 1798, 1837, 2038, 2039, 2040, 2042, 2043, 2116, 2132]. quelques [2133, 1270, 3252, 3378, 4193, 2517, 2573, 1086, 1581, 4427, 2267, 2170, 2066, 344, 1617, 4064, 4116, 558, 1264, 1355, 1421, 568, 4296, 306, 96, 153, 141, 219, 250, 332, 307, 331, 383, 404, 488, 489, 1400, 2015, 1005, 1076, 2070, 971, 3596, 2350, 220, 2593, 2614, 1960, 1794, 215]. Quénu [2068]. querelle [4405, 4644, 468, 2813]. Quest [4478]. Question [4332, 3425, 2858, 3803]. questioni [1074]. Questions [3063, 3133, 3637, 4337, 4059, 3349, 3499, 3252, 4439, 4052, 4132, 4133]. Quetelet [3934, 10]. queue [2851]. Quézel [4285]. Quhi [4023]. Qui [1221, 2411, 968, 357, 986, 3213, 3070, 945, 2347, 2408, 3722, 3530, 2673]. Quichoux [1702]. Quid [3531]. Quincentenary [2501]. Quinn [2539]. Quinon [4616, 4616]. quintae [3316]. quintants [4478]. quo [3531]. quoi [4439]. quotidienne [2409, 476]. Quoy [1059]. Qurra [689, 3241, 4023, 1101]. Qutayba [246].

R [89, 2547, 2901, 2332, 2631, 2632, 3506, 8, 14, 60, 1025, 2627, 2803, 2190, 1977, 2384, 652, 1829, 1979, 2116, 1141, 3433, 451, 1590, 3153, 214, 418, 455, 651, 940, 2662, 513, 509, 4260, 3118, 2565, 3337, 997, 3503, 3640, 1162, 2172, 2321, 3363, 3381, 607, 1059, 1148, 1848, 2354, 603, 1672, 3651, 2570, 3608, 2539, 1616, 3835, 2835, 3002, 3439, 2231, 3344, 1109, 900, 3565, 1924, 1893, 1891, 1758, 1759, 3603, 863, 484, 1998, 2185, 380, 4209, 2155, 3239, 240, 278, 788, 1105, 1290, 1323, 1111, 2665, 2829, 3926, 378, 1260]. R [1261, 631, 951, 2923, 2224, 1475, 2816]. R. [104, 4017, 1133, 285, 3657, 2102, 2021, 2124, 1969, 3658]. R.-C [240]. rūhāniyya [2372]. rêve [352, 3439, 3004, 1007]. Rêveries [4296]. rêves [4546, 2325]. Rôle [732, 4609, 3854, 1987, 3046, 4447, 4251, 1630, 3328, 2678, 2205, 2285, 4114, 4566, 2273]. Raalte [3611]. Rabbi [534]. rabbinic [3589]. Rabelais [495]. Rabinowitz [247]. race [4436]. racine [1738]. Racines [4393, 4224, 19, 4033, 3329, 3557]. racun [2568]. Radelet [2408]. Rädeuhren [1693]. Radiation [966]. Radical [2853]. radio [762, 2970]. radio-communications [762]. radioactifs [2600]. radioactive [2463]. Radioactividad [88]. radioactivité [3855, 35]. Radioactivity [1472]. radium [2745]. Radot [34, 1242, 458, 1117]. Radulphi [2287]. Radulphus [2287]. Raffaele [210]. Raffaella [4270]. Ragionamenti [603]. RAINS [4184]. raison [3529, 2186]. raisonnement [4583, 2689, 3385]. Raisons [3561, 3494]. Raissa [726, 1288]. Ralph [4404]. Ramazzini [1571]. Ramón [4255]. Ramsay [1472, 859]. Ramsey [3274]. Ramunni [2750]. Ranc [35, 138]. Ranchetti [2941]. Ranchin [2121]. Rand [4478]. Randell [89]. Randles [2657, 4082]. rang [2056]. Raphael [3652]. rapide [2657, 1588]. rapidités [4488]. Rappaport [3947]. Rapport [3566, 1222, 3052, 2576, 2408]. rapportant [2938, 601]. rapports [618, 4553, 4488, 3460, 4431, 1659, 445, 344, 2301, 2177, 4488, 4132, 4134, 4133]. rares [3698, 1541]. Rashed [4560, 3595, 4493, 4023, 3239]. Raspail [4503, 1853]. rassemblées [3343]. rassemblés [1523]. rat [380]. Rather [2069]. ratio [4110]. rationalisation [4608]. Rationalisierung [2849]. rationalisme [274]. rationalist [3888]. rationaliste [2397, 376]. rationalité [3171, 902, 3633]. rationalization [4608]. rationeque [844, 845, 331, 219]. rationnel [3422]. rationnelle [3703, 1984, 2973]. rattachent [357]. Rattansi [3653]. Ratzel [3376]. Rausky [2421]. Raven [191]. Ravetz [2155]. Ray [89, 191, 2922]. Rayet [4280]. Raymann [485]. Raymond [1931, 10, 16, 2842, 1386, 3758, 2404, 1567, 2169, 894, 597, 2813, 2404, 2210].

Raynaud [2962, 4352, 3995]. Raynaud-Nguyen [2962]. Rayon [1313, 1215, 1351, 3187, 4328, 1118]. rayonnement [2022, 4226]. rayonnements [2170]. Rayons [1964, 2895, 3688]. razionale [943]. razionalizzazione [1537]. razonamiento [154]. Razumovskij [1332]. Razvitie [1610]. Razzell [2450]. Re [4674, 2061, 3150]. re-consideration [3150]. re-evaluation [2061]. Re-register [4674]. Reaction [2633, 4580, 4580, 1381]. Read [2837, 1942]. reader [3550, 4570, 3933]. readers [3932]. Reading [3389]. readings [4254, 4254]. Reagenzglas [4095]. Real [885]. réalisations [53, 1781]. réaliser [4586]. Réalisme [3893]. realist [3411]. réaliste [4179]. réalité [2313, 2919, 4584, 4472, 3862, 2580]. Réalités [4620, 4001, 1865, 2325]. realities [4584]. reality [3864, 4584]. realize [4586]. realized [2838]. réanimation [2512]. Réapparition [688]. Reason [2226, 2224]. reasoning [4583]. Réaumur [776, 720, 737, 736, 1007, 716, 1136, 1137, 718, 734, 714, 717, 715, 891, 735, 719]. Reboul [155]. récapitulation [3635]. récente [760, 4618]. récentes [3267]. récents [222, 3071, 3378, 4159, 1401, 3094, 2015, 2349]. Reception [4602, 4580, 4600, 2151, 4604, 4580, 4600, 2151, 4126, 4604]. recettes [570]. Rechenberg [3077]. Rechenbuch [1795]. Rechenmaschine [2607]. Rechenpfennige [1776]. Recherche [2374, 3349, 2396, 2423, 3097, 738, 4477, 4506, 3943, 3703, 2588, 4416, 4205, 2502, 4094, 433, 3234, 2239, 3984, 4328, 3633, 3382, 102]. Recherches [299, 3872, 3145, 2919, 591, 2705, 4066, 2237, 3595, 1574, 2669, 1433, 208, 2781, 2613, 2194, 2134, 2366, 3721, 1258, 1416, 2765, 1135, 1169, 1154, 4396, 2693, 4278, 1656, 841, 2407, 2650, 3935, 3439, 1938, 3828, 671, 2396, 1113, 2573, 2693, 3736, 1295, 530]. Rechnen [1864, 3444]. réciproquement [4445]. Récit [4196]. Reckoning [1566]. Recollections [2977, 3005]. Réconcilier [4251]. Reconsidered [1025, 3649]. Reconsidering [4507]. reconstitution [4338]. reconstruction [3836, 3703, 3304]. recours [4685]. Récréations [987]. Rectification [1492]. rectificative [2285]. Rectitudo [4234]. Recueil [3578, 1177, 1408, 2582, 1027, 2401, 2453, 3197, 217]. recueilli [2896]. recueillies [457]. recueils [758]. Recues [697, 908, 1080, 1446, 1870, 1961, 2046, 2118, 2235, 2496, 2676, 2759, 2906, 2966, 2980, 3084, 3108, 3179, 3283, 3390, 3445, 3537, 3613, 3678, 3707, 3768, 3845, 3905, 3973, 4037, 4099, 4180, 4211, 4272, 4370, 4442, 4496]. Recus [1620, 1760, 2346, 2594, 2949, 2930]. rédacteurs [1084]. redaction [4198, 2466]. Redaktion [2058]. Rede [14]. redécouverte [117]. rédigé [3189, 1112]. rediscovered [4601]. rediuiua [529]. Redondi [3306, 2806, 2855]. réécriture [4221]. Reed [328]. réédition [2460]. Reeks [397]. réel [4189]. réels [3689, 4287, 4584, 1087, 3877]. Reese [2870]. Reeves [4527]. Refaire [3030]. Reference [1019, 2949, 3697, 4018, 2866]. reflecting [4478]. Reflections [3364, 1891]. reflets [1173]. réflexif [4187]. réflexion [3873]. Réflexions [1283, 3402, 224, 3429, 2441, 4236, 2347, 3986, 2509, 3573, 3671, 1958, 3439, 834]. réflexivité [4674]. reflexivity [4674]. réforme [3227, 1376, 2679, 1775, 4096].

Reformer [453, 1853]. réformes [3825]. réformiste [4331]. Reformkatholizismus [1438]. réfraction [645, 3067]. réfringens [2408]. refus [1358]. refuser [4587]. regard [4004, 3557, 4141, 4306]. Regards [4424, 3161, 3692, 2666, 3387]. Régent [1253]. Regesta [2334]. régie [2047]. Régime [1708, 3753, 3621, 2448, 2681, 2773, 3527, 1113, 2234]. Régime. [1115]. régimes [4525]. Régine [3043, 3607]. Regiomontano [4476]. Regiomontanus [2515]. région [2902]. regionale [2358]. Régis [3241]. régisseurs [3936]. Register [1945, 4674, 2468]. registers [4571]. registre [1741]. registres [4571]. règle [657]. règles [3802, 2502]. règne [1480]. Regnot [3608]. Regulae [2631, 4635, 3799]. réhabilitation [2365]. Reiches [2674]. Reichmann [1522]. Reid [1977, 2571]. Reihe [3223, 2661, 2470, 3416, 3214, 2823, 2849]. réimpr [3791]. réimpression [1390]. Reine [2470]. Reingold [2337, 2834]. Reinhard [3667]. Reinhardt [4622]. Reinhold [536]. Reino [1016, 2253]. Réinscrire [4674]. reinterpretation [2061]. Reisch [4365]. Réjane [2293]. rejection [1866]. Rekenboeken [1293]. rekenen [1793]. relación [3732]. relaciones [3735]. Related [3082, 1486, 4230, 2528]. relatifs [1258, 1410, 2781]. relating [235]. relation [3805]. Relations [936, 4481, 402, 2665, 3714, 3448, 1855, 1488, 4426, 2238, 2517, 1135, 884, 3158, 2368, 24, 1842, 1903, 3658]. relative [1918, 3956]. relatives [1189, 958, 2347, 596, 668]. relatividad [136]. relativistica [3674]. Relatività [3870]. Relativitätstheorie [3245, 1684]. Relativité [2312, 4142, 4046, 850, 4452, 2151, 2748, 1632, 860, 2972]. Relativity [3106, 2746, 2747, 2151, 1999]. relazione [2395]. relevance [3210]. reliable [4581]. Religieux [2527, 2658, 109, 1667, 2964, 1015, 1072, 1232, 1955, 2095, 2107, 626]. Religion [3301, 3993, 3484, 2683, 3482, 445, 4658, 3367, 3210]. religions [1434]. religious [3810]. remained [3497]. Remak [610]. remarks [4301, 4198]. remarquables [3791, 2408]. Remarque [1621, 1622]. Remarques [1209, 678, 347, 223, 3803]. remède [4246]. Remes [2333]. Rémi [3831, 3650]. Réminisciences [3844]. Renacimento [417]. Renacimiento [566]. Renaissance [1538, 2526, 4081, 1207, 1435, 1019, 4491, 1162, 2917, 2534, 750, 92, 4096, 2106, 2216, 1470, 3812, 4394, 2373, 784, 3739, 60, 1161, 1266, 4135, 4126, 1347, 4625, 2352, 589]. renaissances [3867]. Renan [4084, 3253]. Renard [2066]. Renard-Payen [2066]. Renate [2129]. Renato [2685]. Renault [4298]. Renck [4395]. rencontre [4393, 4407]. rend [3646]. Rendiconti [4076, 4435]. rendu [459, 1463, 3408]. rendus [1943, 2295, 1835, 2045, 1031, 1176]. René [3225, 3631, 213, 2631, 2870, 3278, 692, 41, 419, 652, 69, 448, 3348, 2502, 1662, 818, 819, 996, 1756, 355, 374, 445, 1327, 1328, 1959, 584, 2999, 3222, 1661, 1058, 2736, 2738, 2307, 1430, 1431, 3013, 1832, 2996, 2953, 2969, 3716, 3715, 3708, 3712, 4636, 4244, 3709, 1951, 684, 714, 3714]. René-Antoine [714]. René-Francois [2996, 2953, 2969]. René-Just [3222, 3716, 3715, 3708, 3712, 3709, 3714]. Renée [3343]. Renn [3663]. Renneberg [3706]. Rennes [1835, 797, 1303, 1306, 1307]. Renneville [4052, 4208]. renommée [3079]. réorganisation [4609]. Reorganization [4609]. répercussions [3428]. Repères [2398, 4428]. répertoire [2996, 2953, 2969, 1075, 4289, 2915, 984, 2131]. Repertorium [4243]. réplications [3694]. Réponse [3498, 1951]. Report [2501]. reporter [3885]. Reports [3340, 3482]. Repos [4223]. représentants [323]. representation [3889, 3604, 3929, 2260]. représentations [2254, 4620, 4205]. représentative [2973]. reprint [1685, 4235]. Reproduction [530, 940, 1763]. reproduites [1145]. Republic [3904, 2853]. république [4382, 4644, 4503, 4249, 3168, 4509]. rerum [4301, 3316]. Resa [876]. Research [1913, 2115, 2378, 3154, 1246, 2190, 3703, 1484, 3381, 3840, 3332, 2522]. Researches [456]. réseau [4567, 4566]. réseaux [1349, 4045]. résidents [1306, 1307]. Residues [3009]. résidus [2568]. Résistance [4642, 4467, 1137]. résistances [3624]. résistants [3068]. résolution [1377, 472]. Résolutions [349]. resonance [3220]. respirant [3964]. Respiration [1636, 1110]. respiratoire [2512]. responses [3891]. ressemblance [3286]. ressorts [2951]. Rest [2699]. Restaging [3694]. restera [405]. restera-t-elle [405]. Restless [3564]. restreinte [850, 1632]. résultats [2134]. Résumé [1188]. Résumés [2726]. Reti [1942]. réticences [4355]. reto [3946]. Retouches [700]. Retour [762, 4666, 2211, 432, 4438, 2994, 3074]. retrait [2937]. Retrospect [903, 2638, 3344]. retrouvé [1738, 2048]. retrouvée [4063]. retrouvés [4601]. Réunion [1463, 935]. Réunions [1095]. réunis [480]. réunissant [1544]. Rev [1973, 3497, 2400]. révélation [8]. Reverdin [2033]. Review [162, 1931, 4627, 4614, 4623, 4626, 1928, 1943, 1923, 2345, 4617, 4615, 4616, 4624, 632, 846, 973, 1123, 1316, 1507, 1912, 1946, 1911, 1915, 1945, 1913, 1944, 1914, 1892, 1887, 1888, 4645, 4644, 4628, 1909, 4641, 1960, 3838, 1925, 1956, 1941, 1953, 1938, 1939, 1926, 1927, 1955, 1929, 1958, 1959, 1940, 4625, 4642, 4619, 1933, 1908, 1907, 1954, 1936, 4640, 3808, 4643, 1889, 4646, 4629, 1906, 1910, 1924, 1896, 1895, 1894, 1890, 1893, 1891, 1897, 4647, 4618, 1922, 1935, 1934, 2804, 1930, 306, 4620, 96, 250, 1932, 1942, 4622, 1937, 1957, 1246, 4621, 220, 2190, 1445, 1484]. review [2115, 4575, 3933, 3932]. Reviews [4576, 4585, 4594, 4605]. revised [1665, 4203, 627]. révision [3403]. Revista [529]. revolución [2635]. Revoluciones [1542]. Revolution [4269, 3432, 4188, 938, 2743, 4624, 1626, 3664, 89, 4475, 3277, 3548, 3481, 2421, 1749, 3547, 1248, 4644, 3621, 2413, 105, 1395, 3151, 4495, 1064, 2975, 3577, 277, 1155, 3167, 3168, 2632, 2709, 3106, 787, 2224, 4316, 3808, 957, 2884, 1480]. revolutionary [3788]. revolutionibus [4643, 2135]. Révolutionnaire [4331, 497, 2541, 669]. révolutionnaires [3168]. revolutions [4655, 4655, 4643, 2175]. Revue [4187, 795, 3349, 4473, 2296, 2422, 974, 4002, 4575, 1032, 4448, 4470, 482, 3149, 4366, 463, 971, 2246, 4307, 2191, 2193, 633, 696, 1033, 1035, 1444, 1945, 2043, 2132, 2170, 986, 1544, 1897, 1005, 2472, 1, 354, 1269, 1032, 353, 2704]. Revues [3109]. Rex [1504]. Rey [3153, 1472, 3577, 4251, 3571, 3575, 150, 3742, 452]. Reyer [3048]. Reymond [651, 3059, 739, 651]. Reyneau [143]. Reynier [215]. Reynolds [89]. Rhône [3948]. Rheinau [2034]. Rheinberger [4095, 4095]. rhénanes [2407]. Rheticus [2770, 2231]. Rhetoric [3675, 3482, 4317]. rhétorique [4173, 4027, 4368]. rhétoriques [4314, 4567]. Rhin [2407]. rhinocéros [2764]. Rhoda [3947]. Ribot [4366]. Riccardo [4515]. Ricci [877]. Ricerca [4515]. Ricerche [3365, 2471, 3941]. Richard [726, 695, 2228, 2656, 3303, 729, 1975, 3598, 3262, 3153, 4478, 3455, 3362, 3503, 3584, 2922, 4292, 2511, 3484, 2804, 2699, 3953, 1919, 3815, 2963, 4052, 2386]. Richard-Harrison [726]. Richardson [2749]. riche [3958]. richesse [3279]. richesses [1253]. Richet [3971]. Richter [2674, 526]. ricostruzione [3491]. Riddle [3531]. Rider [2942, 2875]. Riedel [4053]. Riemann [3201, 2780, 3200, 1681, 2025]. Riese [965]. Rife [4172]. RIG [3197]. Rigal [1865]. Righini [2344, 2690, 2224]. Right [1973, 3863, 3341]. Rigoli [4173]. rigueur [2867]. Rijksmuseum [359, 841]. Rijksmusuem [514]. Rinascimento [3813, 3438]. Ringuet [2832, 654]. rinnovamento [4476]. Riolan [510, 4555]. Riordan [3902]. Riproduzione [1099, 2808]. Rise [3902, 2550, 3139, 3590, 3728, 4589]. risk [4520]. risque [4517, 4140, 3311]. Rist [613]. Rita [3957]. Ritter [3525]. Rivaz [1416, 1404, 1258]. Riverso [1537]. Rivière [214, 1832]. Rivista [2297, 1034, 1178, 1509, 1618, 2787]. Rivoluzione [4309, 4270, 2471]. Rizza [1567]. Robatel [3408]. Robert [1931, 3663, 3939, 1785, 2331, 3880, 2509, 10, 11, 1545, 1974, 2275, 2345, 535, 1975, 1259, 73, 704, 1113, 4018, 2642, 3362, 3431, 3788, 2672, 1772, 1098, 3794, 3629, 3860, 3439, 2226, 1732, 1737, 2288, 2528, 1230, 1521, 1846, 3812, 2702, 4140, 2370, 2571, 2307, 2697, 860, 1294, 2311, 3015, 2592, 3865, 569, 4085, 2487, 851, 1349, 3677, 4554, 2564, 64, 3649, 862, 3605]. Roberto [989, 4196]. Roberts [3934]. Roberts-Jones [3934]. Roberval [679, 179, 406, 2161, 3451, 3805, 680, 3811]. Robin [2942, 4010, 2875, 4587]. Robinet [4395, 2156, 2613]. Robinson [3591, 853]. robot [256, 2598]. Rocco [2629, 2991]. Rochas [1730]. Rochat [3760]. Roche [2770, 2231]. Rochefort [1000, 1723]. Rochelle [848]. Rochon [2895, 2894]. Rochot [3278, 585, 1994, 1667, 1955, 2107, 1541]. Rockefeller [4094]. Rockfeller [3590]. Roderick [3996]. Rodger [2870]. Rodolfo [2395]. Rodolico [2358]. Rodriguez [3388, 3994]. Rodriguez-Consuegra [3388]. Roemer [999, 4075, 2504]. Roero [4351]. Roger [2338, 9, 4079, 3053, 783, 3153, 235, 3515, 357, 2638, 3751, 3839, 4020, 3338, 1579, 1386, 4456, 2126, 1604, 2105, 2405, 2184, 2407, 2881, 3151, 159, 1337, 1338, 3149, 3297, 3742, 2871, 3126, 525, 1235]. Rogers [3487]. Rohault [4494, 2204, 2327, 3659]. Roi [1762, 83, 2598]. rois [3146]. Rokitansky [1028]. Roland [4097, 4175, 511, 3372, 3809, 3922, 3460, 2316, 4398, 480, 2382]. Rolando [2824]. Role [4609]. Rolf [2841]. Rolin [483]. Rolle [2982]. Roller [1164, 2052, 667]. Rollet [4242]. Rom [3505]. Roma [2606]. Romain [4319, 2705]. romaine [1892]. Roman [3885, 3655, 2371, 2531, 2233]. Romano [4070, 1907, 4096, 904]. Romantic [3485]. Romanticism [4207]. romantique [3820, 2575, 4646]. romantiques [3308]. romantisme [4029]. Rome [2976, 845, 4096, 2494, 1784, 3564]. Romein [1590]. Romein-Verschoor [1590]. Romer [1472]. Rommevaux [4488]. Ronald [2871, 3083]. Ronchi [792, 2232, 1333, 2232, 2353, 449]. Roncière [2962]. Rond [4615, 3704, 4120, 3863, 4004, 3834, 4176, 4199, 3938, 4208, 3765, 4011]. ronde [3408, 4510]. Rondelet [386]. Roofe [951]. Rooseboom [949, 841]. roots [3110, 2836, 821]. Rosa [3150]. Roscoff [4637]. Rose [3794, 2373]. Rose-Mary [3794]. Röseberg [3040]. Roselyne [3153, 3742, 3577, 3571, 3575]. Rosemary [2171]. Rosen [2231, 2525, 1664, 1543, 4554, 2810]. Rosenberg [4254, 3532, 2810]. Rosenfeld [2565, 2828, 3263]. Rosenkrantz [2923]. Roshdi [4560, 3595, 4493, 4023]. Rosińska [3011]. Rösler [1935]. Rosmorduc [2783, 2833, 4164, 2211]. Rosner [2479]. Rossetti [4072]. Rossier [2059]. Rossii [983]. Rostan [1457]. Rostand [35, 2774, 3152, 158, 630, 1096, 1909, 2381, 796, 1569, 1058, 454, 835, 1024, 1167]. Rosumek [2849]. Rota [4434]. Rota-Treguier [4434]. rotation [76]. rotatoire [4568, 2833]. Rothacker [506, 360, 394]. Rothmann [4402]. Rothschuh [533]. roue [1361]. Rouelle [4086, 2896, 1856]. Rouen [3299, 4586, 756]. Rouget [4196, 2073]. Rougier [4242, 4583, 2683]. Rouillé [2298]. roulement [1047]. Roulette [1185]. roumaines [2368]. Roumanie [129, 1861]. Rousseau [2318, 728, 1761, 4454]. Rousset [280]. Route [2408, 465]. Routes [2775]. Routine [4563]. Routledge [4254]. Roux [612]. Rouyer [2611]. Roviello [3121]. Rowan [1672]. Roy [3763, 2445, 3259, 1695, 436]. Royal [89, 280, 2837, 727, 4290, 2807, 3903, 2412, 4492, 1191]. Royale [1488, 289, 758, 142, 1307, 1443, 224, 701, 1599, 659, 1839, 1375, 4411, 1376, 1306, 1705, 1708, 779, 1762, 2177, 1374, 1604, 2105, 2405, 124, 1840, 4565]. Royaume [3279, 1557]. Royaumont [1014]. Royce [89]. Rozenfeld [2385]. Rozsivalova [960]. Ru [3704]. Ruben [1020]. Rubin [3417]. Rudaux [187]. Rüdiger [4177]. Rüdinger [4079]. Rudjer [2184, 2854]. Rudolf [2069, 188, 3739, 2647]. Rudolph [1111]. Rudra [380]. Rudwick [2659, 4269]. Rudzer [1105]. rue [485, 2340]. Ruestow [3904]. Ruffat [3949]. Ruffié [3971]. Ruffini [829]. Ruler [240]. Rules [221]. Rumford [2639]. Rumpf [2883]. Rumphius [2883]. Ruota [3870]. Rupert [2529, 1522, 3195, 2493]. Rupescissa [3316]. rupture [3174, 2904]. ruptures [4085]. rurale [2514]. russe [2548, 655]. Russel [2262, 3388]. Russell [2192, 1944, 3074, 2495, 4580, 2331]. russes [4426, 1698, 3428, 1842]. Russia [3565, 478]. Russian [2610, 3553, 4406]. Russie [936]. russische [1332]. russischen [1332, 1060]. Rußland [4361]. Russo [2865, 444, 3761, 3096, 1922, 184]. Ruth [2505, 3871]. Rutherford [1472]. Rutten [1290, 1214, 3488]. Ryan [3603]. Rydberg [507]. rythme [137].

S [1265, 89, 1931, 2976, 2905, 2632, 3610, 1025, 2190, 2532, 3384, 2215, 1034, 1975, 3009, 3361, 3369, 400, 3598, 3940, 2957, 3153, 204, 327, 328, 456, 581, 2058, 373, 1019, 2568, 2565, 3337, 1610, 2837, 2709, 3105, 2172, 2840, 3593, 2036, 2787, 238, 447, 3149, 4269, 1936, 3409, 2977, 3002, 3344, 1104, 945, 1412, 1592, 2528, 938, 2406, 3141, 1289, 4394, 2699, 566, 903, 2866, 246, 247, 534, 929, 2635, 4190, 1464, 190, 2551, 2659, 1151, 3867, 4634, 610]. S. [1464]. S.I [2131]. S008 [3212]. Sōtetsu [4654]. Sa [496, 949, 1169, 1059, 4613, 2026, 1852, 2446, 408, 3359, 4553, 64, 114, 301, 465, 2568, 2637, 3057, 4306, 4683, 2412, 3028, 796, 1569, 3648, 585, 3318, 124, 1478, 1687, 2291, 4472, 1260, 2350, 2613]. Saale [3536]. Sabine [4488]. sabio [417]. Sacha [4547]. s’achèvent [4166]. Sachsen [3443]. Sachwörter [1298, 1070, 1564]. Sacrobosco [3665]. Sadi [2509, 1891, 834, 2139, 1491, 2239, 2377]. sadis [2372]. Sadoul [45]. Sadoun [2446, 2704]. Sadoun-Goupil [2446, 2704]. saga [3262, 1504]. Sage [4260, 2797]. Sagesses [4488]. Saggi [2694]. Saggio [4475, 2358, 2808]. Saïgon [1312]. Saillard [4008]. Saint [3942, 32, 4003, 3933, 3753, 3932, 2270, 4319, 3527, 2017, 2002, 2256, 2001, 2003, 2004, 2005, 2000, 4413, 2722, 2006, 2284, 1153, 254, 1284, 588, 1831]. Saint-Denis [3933, 3932, 588]. Saint-Domingue [3753, 3527]. Saint-Étienne [4003, 4413]. Saint-Fond [2256]. Saint-Hilaire [3942, 32, 2017, 2002, 2001, 2003, 2004, 2005, 2000, 2006, 2284, 254, 1284, 1831]. Saint-Martin [2270]. Saint-Romain [4319]. Saint-Victor [2722]. Saint-Vincent [1153]. Sainte [4416, 3753, 146]. Sainte-Claire [4416]. Sainte-Claire-Deville [146]. Sainte-Lucie [3753]. sais [1214, 140, 4032, 1613, 242, 1327, 1328, 1634, 3007, 1352, 2736, 2756, 1326, 1932, 363]. sais-je [1214, 140, 4032, 1613, 242, 1327, 1328, 1634, 3007, 1352, 2736, 2756, 1326, 1932, 363]. Saison [1983]. Sakarovitch [4009]. Sakhri [4545]. Saladin [249]. Salamandre [1763]. Salamanque [1231]. Salanskis [4233]. Saldaña [4458]. Salem [1705]. Salhab [4417]. Sallantin [4174]. Sallmann [4415]. Salmon [4546]. Salomon [3151, 4639, 2983]. Salomon-Bayet [3151, 4639]. salon [311]. salpêtres [3936]. Salpêtriére [4481]. Salsamenta [3362]. Salt [2642]. Saltzer [2603]. salubrità [3941]. Salud [2965]. Salusbury [1103]. Salut [3208, 4432]. Salvatore [3829]. Salzburg [2479]. Salzburger [2706, 2707, 2479]. Samaran [1230, 1521, 1846]. Samarkandi [975]. Samaw’al [2230]. same [2837]. Samfundets [1150]. Samh [4023]. Samion [1304]. Samion-Contet [1304]. Sammler [1262]. Sammlung [787, 1325]. samt [747]. Sämtliche [3223]. Samuel [2838, 887, 2087, 2309]. San [3340, 1933, 3759, 883, 4662]. san-tchong [883]. Sanborn [2639]. Sanchez [3235, 4255]. Sanctallensis [1436]. Sanders [4239]. Sandra [3491]. Sandro [3238]. sang [1702, 2916, 356, 510]. sanguine [1089]. sanguinischer [3488]. sanitario [2965]. saniyya [2372]. sans [4638, 3373]. Sansuc [2404]. Santé [1001, 1002, 4524, 1116, 1755, 1834, 1941, 3293, 4237, 963, 1059, 1114, 3577, 3054]. Santiago [2805]. Santillana [629, 1606]. sapere [4476]. sapiens [3407]. Saporta [2341]. Sara [3996]. Sarah [2550]. Sargent [3794]. Sarkar [2990]. Sarl’ [2830]. Sarmant [2744]. Sarnowsky [3443]. Sarton [844, 845, 219, 331, 784, 899, 245, 442, 574, 784]. Sassady [1310]. satellite [3004]. satellites [3566, 3783]. satellitibus [3566]. satirische [2479]. Saturn [1236]. Saturne [2491, 3783]. Saumaise [3458]. Saunders [3217]. Saupin [4261]. Saussure [3749, 4090, 3418]. Sauvage [3098]. Sauvages [1782]. Sauveur [78]. savant [3631, 3942, 4543, 3971, 2089, 2263, 2637, 3233, 2894, 4464, 1064, 4555, 3079, 1663, 1781, 2680, 79, 585, 3844, 3308, 1437, 2675, 2026, 3310, 2245, 825, 1403, 2407, 1474]. Savantes [1835, 1176, 4165, 1943, 2295, 4664, 4028, 4609, 3569, 2045, 3621, 1148, 1031, 3569]. Savants [3793, 301, 324, 562, 1674, 40, 80, 4383, 3832, 2517, 4034, 1008, 3224, 504, 3168, 755, 681, 718, 130, 1243, 4544, 66, 1737]. savants-musiciens [755]. Savary [4252]. Savelli [2395]. Savie [611]. Savioz [378]. Savoie [4643, 2842]. Savoir [3863, 3439, 2175, 4168, 4508, 3833, 4485, 2722, 4474, 4414, 4528, 4633, 4139, 3724, 3897, 2735, 2634]. Savoir-faire [3863]. Savoir/sciences [3439]. Savoirs [3820, 4339, 4081, 4051, 3161, 4340, 3831, 4237]. Savonarola [3039]. Saxe [3097, 4134]. Saxon [1442, 3756]. Sayili [1388, 1018]. Sbornik [1355]. Sc/10 [2739]. scaphandre [1342]. Scattering [3257]. Scelte [830, 943]. scepticisme [4085]. sceptique [4085]. Schaffen [2778]. Schaffer [3654, 3312]. Schams [975]. Schams-ed-Din [975]. Schechner [3996]. Schedrov [1610]. Scheifele [3843]. scheikunde [991]. Scheiner [4527]. Scheler [1205]. Schelling [4361]. schéma [1986]. Schepers [2423]. Scherz [690, 861]. Scheurer [2822]. Schickard [2607]. Schid [960]. Schierbeek [358, 326]. Schiller [2317, 2665, 2317]. Schilperoort [2462]. Schiøler [2233]. Schipperges [2479]. Schlabach [3060]. Schlaudt [4545]. Schleiden [2030]. Schloesing [1109]. Schmidt [3148]. Schmitt [4437, 4346]. Schneider [4204, 46, 2706, 3415]. Schofield [2528, 2226]. scholar [3886]. scholastische [3443]. scholie [4135]. scholies [4128, 4127]. Scholtens [1775]. Scholz [2780, 3222]. Schön [4551]. schone [991]. School [3386, 3317]. Schools [3139, 3754]. Schopenhauer [493]. Schott [4026, 4026]. Schouten [302]. Schove [3012]. Schramm [4177]. Schreiber [2847]. Schrift [3843, 3346]. Schriften [3223, 3314, 2695, 4361]. Schriftenreihe [1709, 1693, 1935]. Schrifttums [2229, 2545]. Schrijvers [2989]. Schröder [3347, 4579, 2778]. Schrödinger [3578, 3440]. Schroeder [4495, 2823]. Schubring [3886]. Schuchardin [1412]. Schuder [842]. Schuhmann [2384]. Schule [2536, 1526, 3222]. Schulenburg [4204]. Schulman [3865]. Schulmann [3663]. Schumacher [1332]. Schütt [3489, 3214]. Schütz [533]. Schwann [1109, 1024, 1147]. Schweizer [3123]. Schweizerischen [1534, 1617]. Sci [3497, 2400]. Science [3486, 4304, 1267, 1429, 1861, 4189, 162, 1931, 2421, 3519, 4264, 3591, 4450, 4477, 4569, 174, 689, 2547, 3485, 4092, 203, 1973, 2942, 3533, 1752, 4079, 2190, 2275, 2342, 2345, 2052, 2532, 485, 3501, 3411, 3517, 3522, 3566, 4170, 4258, 4527, 2664, 537, 593, 632, 906, 909, 846, 888, 972, 973, 1120, 1123, 1207, 1313, 1316, 1507, 1946, 1913, 1914, 1975, 2115, 1214, 359, 1747, 2378, 3436, 3441, 3696, 3797, 282, 768, 1270, 1728, 391, 443, 3990, 4438, 3515, 4236, 2445, 4160, 4457, 4478, 2109, 4610, 4019, 3650, 2565, 3869, 4094, 3276]. Science [3386, 3480, 3867, 1162, 1432, 2030, 2522, 2336, 2452, 2641, 3675, 1847, 1997, 2094, 2672, 1685, 1215, 2306, 3655, 3525, 3970, 3259, 4396, 4456, 4592, 3389, 4009, 1237, 3756, 3794, 3835, 3535, 1787, 2959, 2960, 3002, 3275, 3808, 4643, 3500, 1472, 1473, 3484, 3524, 3724, 1104, 3364, 2811, 3432, 3565, 216, 1910, 1351, 1732, 3094, 3527, 2024, 2060, 4348, 1758, 3816, 2702, 2127, 2675, 3367, 3387, 1686, 2100, 2476, 3827, 3120, 243, 306, 96, 250, 276, 332, 442, 477, 531, 667, 784, 903, 1543, 1566, 1942, 2872, 2875, 2881, 1014, 1068, 1784, 2457, 2634]. Science [2703, 3830, 971, 1957, 1246, 3006, 220, 286, 2907, 3950, 4458, 89, 4557, 3572, 3836, 4614, 3422, 3654, 3810, 3802, 4013, 3710, 3579, 4254, 4365, 3349, 3863, 3343, 3880, 222, 3826, 4233, 4314, 4475, 3377, 38, 65, 178, 4085, 3943, 4354, 1445, 1630, 3581, 114, 3378, 4663, 371, 237, 3348, 4298, 3755, 3786, 4005, 4018, 4255, 4267, 4470, 3824, 3457, 3886, 4503, 1610, 3405, 4065, 4305, 3313, 3916, 4095, 3630, 3381, 2477, 818, 819, 150, 445, 3587, 492, 3641, 4432, 3638, 4207, 584, 594, 4525, 2076, 2195, 4269, 4292, 4658, 1954, 4349, 4633, 3743, 3016]. science [3210, 2327, 3439, 3357, 4561, 4167, 1845, 4188, 215, 3434, 4154, 3728, 2101, 2288, 277, 4452, 3871, 3837, 1661, 3986, 2673, 2698, 3453, 3856, 2315, 112, 4209, 4164, 4197, 1430, 1431, 1865, 1118, 1572, 1612, 2015, 2866, 2087, 4677, 4495, 4558, 1394, 3873, 1439, 1067, 1511, 1484, 1797, 2039, 2116, 156, 997, 1162, 525, 982, 2271, 820, 329, 3939, 3700, 39, 3581, 3370, 3337, 3751, 2838, 2030, 3608, 3724, 4294, 3728, 4190, 4458, 3533, 37, 2803, 592, 729, 1435, 1631, 767, 1096, 3520, 3567, 3301, 3630, 3788, 2692, 665, 744, 1565, 1771, 653, 2773]. Science [3706, 2182, 3080, 2225, 2523, 3565, 2525, 3177, 3668, 185, 3367, 4209, 4509, 190]. Science-Histoire-Philosophie [3827, 3343]. science/Biological [3743]. Science/History [4527]. Science/Physics [3696]. sciencefiction [3837]. Science^{≫} [212]. Sciences [147, 202, 349, 497, 650, 724, 782, 765, 1013, 931, 980, 1094, 1053, 1229, 1462, 1561, 1767, 1825, 1862, 2123, 3170, 3186, 165, 3747, 2166, 2870, 795, 2504, 2262, 5, 72, 57, 676, 732, 1943, 1808, 1923, 2192, 2193, 2296, 2396, 2411, 4616, 2750, 535, 617, 633, 696, 974, 1033, 1035, 1640, 1835, 1944, 2043, 2132, 897, 1091, 892, 4311, 1463, 3433, 1729, 1736, 73, 257, 275, 327, 1523, 1461, 1626, 2423, 2928, 3520, 3746, 233, 2637, 4369, 3118, 2109, 2575, 2506, 3171, 759, 2837, 2741, 1535, 2045, 2670, 2840, 2044, 2876]. Sciences [3079, 2452, 1953, 2603, 3043, 151, 1673, 1750, 1263, 3651, 1338, 3241, 3420, 886, 756, 3002, 217, 6, 2419, 3147, 1827, 2405, 2583, 1739, 28, 572, 928, 1058, 1437, 1574, 2786, 2751, 2752, 2879, 3825, 494, 738, 1206, 3306, 682, 3827, 2771, 297, 4164, 2386, 3239, 3633, 1730, 1430, 1431, 333, 160, 192, 307, 383, 491, 488, 791, 840, 841, 1071, 1826, 2755, 246, 247, 878, 1176, 2211, 2273, 2612, 848, 3244, 1824, 2246, 1474, 2495, 2613, 2677, 85, 116, 935, 3191, 323, 3635, 4481, 4526, 4433, 3227, 4246, 1749, 2605, 3260, 2831, 3336]. sciences [3711, 3359, 3349, 3863, 2050, 3172, 4266, 4251, 3785, 4165, 44, 8, 27, 115, 3934, 2435, 2803, 4003, 4339, 446, 4077, 2913, 3229, 2689, 3703, 652, 1435, 1444, 1486, 3116, 4411, 4386, 4645, 3153, 4663, 3145, 3266, 3348, 3385, 4644, 4033, 3787, 4141, 4005, 4018, 4195, 4255, 4416, 3571, 4343, 4032, 3935, 4065, 3521, 3569, 3585, 3408, 4171, 3413, 1662, 4451, 2582, 818, 819, 2354, 3818, 4048, 4249, 1663, 4389, 3725, 3970, 3965, 4439, 601, 4362, 1337, 3625, 4392, 4592, 4642, 3559, 2999, 3966, 4269, 1954, 3055, 3582, 3784, 3743, 4241, 4490, 3828, 3439, 3490, 558, 1845]. sciences [3070, 4510, 4629, 3657, 4052, 4346, 3601, 4247, 4618, 3224, 1922, 1661, 3297, 23, 3741, 2376, 2824, 4166, 3387, 3453, 4087, 4440, 4591, 4162, 4552, 3790, 4051, 4124, 3609, 4620, 855, 1403, 2291, 3230, 3750, 4169, 3858, 3013, 4509, 4558, 3382, 4445, 4621, 231, 202, 415, 369, 475, 896, 931, 1533, 1641, 1561, 1770, 1599, 2954, 659, 1488, 1839, 1375, 1149, 1314, 1506, 625, 740, 895, 1463, 193, 289, 499, 758, 142, 1699, 1376, 314, 1762, 2177, 1374, 1840, 1084, 1319, 1433, 82, 891, 294, 3268, 2289, 18, 2189, 2274, 2343, 617, 843, 1033, 1121]. Sciences [1150, 1151, 1315, 1508, 1796, 2038, 581, 2045, 333, 141, 192, 307, 383, 2551, 2865, 3903, 1444, 1619, 1096, 3761, 3312, 4231, 3823, 1433, 184, 2210, 1, 4003, 616, 801, 1911, 204, 155, 4265, 283, 766, 2300]. Scient [690, 861]. Scientia [689, 2979, 4303]. Scientiarum [2661, 3214, 2233, 844, 845, 3755, 536, 628, 219, 331]. Scientific [402, 2332, 2630, 2632, 2529, 3533, 730, 3792, 1415, 2111, 2550, 1057, 3106, 2979, 3524, 3432, 905, 2749, 2633, 2697, 821, 1110, 2807, 2224, 2335, 3747, 1144, 4590, 4475, 4653, 4600, 3703, 3220, 3369, 3696, 4657, 4301, 2988, 1432, 2030, 3789, 3796, 3077, 3808, 2236, 4055, 4613, 3996, 4668, 4318, 2743, 420, 3001, 4242, 3011, 1432, 1924, 1759, 4489, 2457]. scientifica [4345, 4270, 4262, 2471]. Scientifico [2941, 2629, 4301, 3491, 4122]. Scientifique [357, 971, 3740, 2298, 3226, 3289, 1538, 272, 256, 569, 4562, 135, 2256, 4553, 4590, 3793, 4242, 93, 95, 676, 2001, 3934, 4485, 4600, 4156, 2253, 1630, 444, 2299, 892, 1463, 3594, 1626, 2423, 3677, 3761, 1397, 4644, 757, 4657, 3935, 2138, 2413, 2666, 2679, 3818, 4249, 61, 374, 445, 447, 4444, 604, 2858, 2284, 105, 1624, 1513, 4490, 3509, 4417, 2236, 4114, 865, 2822, 1990, 3379, 4487, 1906, 3096, 3964, 4613, 3253, 2347, 3595, 3884, 1695, 4523, 1658, 23, 4668, 1964, 2206, 2824, 479, 738, 4164, 372, 464, 2211, 621, 4495, 3806, 774, 2243, 75]. Scientifique [590, 4264, 16, 3824, 117, 1536, 615]. scientifique^{≫} [504]. Scientifiques [936, 1943, 3569, 1855, 3747, 618, 563, 3349, 2401, 2599, 2168, 615, 1835, 4434, 255, 3212, 399, 465, 698, 1524, 4306, 4205, 3228, 2178, 3778, 4362, 4281, 4396, 3157, 2204, 1472, 3475, 2368, 1231, 1604, 2105, 2405, 2673, 2879, 754, 4566, 4045, 184, 208, 841, 1179, 1842, 1903, 1024, 1165, 4352, 3224, 3831, 2295]. scientifiques^{≫} [944]. Scientist [455, 2109, 947, 2549, 453, 3886, 1853, 2526, 2579, 3521]. Scientists [2129, 216, 3150, 3553, 3364]. Scienza [2344, 989, 1034, 2787, 3365, 3901, 3562, 2395, 3941, 4270, 4098, 2806, 3583, 2406, 4262, 4476, 4122, 2297, 1178, 1509, 210, 505, 1618]. Scienze [2976, 2606, 2358, 2787, 4088, 4076, 4435, 4098, 3058, 3062, 4309, 2941, 4083, 4161, 880]. scienziati [4351]. scineze [4161]. scolaire [4130]. scolaires [3295]. Scolaris [1020]. Scolastique [4491, 1394]. scolie [4222]. scope [3697]. Scott [3153, 4436, 3041, 1671]. Scottish [2390, 2807, 2807]. Scotus [4097, 4175]. Scriba [1668, 1791]. Scripta [240]. Scritti [2174, 505]. scuola [3523]. Scuole [2808]. scurvy [3211]. se [2938, 986, 3039, 601, 3623]. Sea [2850]. sealed [2954]. Sean [2137]. Séance [2123]. séances [459]. Search [2261, 525, 3599, 3702]. Searching [3508]. Sebastian [3340]. Sebastiano [505, 200]. Sébastien [2857]. Sebestik [3609]. sec [1610]. Secchi [3950]. Seccion [1108]. Seck [2607]. secolo [3562, 3039, 3365, 4070, 2173]. Second [2812, 2679, 1891, 1665, 42, 2528, 627, 1071, 4268]. seconda [2471]. secondaire [3227, 3746, 2679, 4087]. secondaires [1349]. seconde [4410, 2729, 1821, 2101, 3623]. secondo [3829, 829, 248]. Secord [3060, 3838]. Secret [2869, 3038]. secrétaire [2298]. secrétaires [4385]. secrétariat [1599, 4389]. Secretary [4178]. Secretos [3735]. Secrets [1789, 1185, 3651, 216]. Section [279, 3260, 1943, 535, 1835, 2575, 2693, 2879, 738, 878, 1176, 2551, 446, 3569, 3413, 792, 3924, 2705, 1346, 1735, 2669, 839, 1079, 5, 1014]. section^{≫} [969]. Sections [133, 2045, 3158]. secunda [2873]. secundus [939]. Sedacer [4286]. Sedacina [4286]. see [3926]. Seeger [1931]. Segal [3726]. Segno [3562, 4239]. Segonds [4643, 2770, 3038]. Segre [4125]. Seguin [2263, 2984, 3426]. segundo [3994]. sei [2991]. Seicento [4068, 248, 2395]. Seidengart [4142, 3336, 3121]. sein [3818, 278]. Seine [3123, 1106]. seiner [4093, 690, 2778, 3419]. seines [4234, 942, 1107]. seit [1734]. séjour [4022]. Seki [3729]. sekreterare [12]. sel [692]. Selbstbiographie [1028]. Selbstreflexion [2823]. Selected [3010, 3990, 1688, 2883, 3347, 2171, 2476, 4318, 2565]. selection [2529, 3410, 2477]. Selections [526]. selective [1780, 1924]. Séleucie [1220]. Séleucus [1220]. Self [3762, 4320, 3752]. self-knowledge [4320]. Self-motion [3752]. Seligardi [4270]. Selleri [3081]. selon [3355, 1858, 4251, 923, 4339, 3137, 2722, 4113, 2120, 4532, 4250, 2204, 3253, 234, 310, 316, 4395, 4509]. sels [4086]. Selsk [3411]. Seltzer [747]. Selzer [747]. Semaine [2260]. semantic [4320]. semence [4401, 4394]. semi [4296]. Séminaire [1823]. Seminar [3441, 1942]. seminari [4161]. Seminario [885, 4161]. sémiologie [2845]. sémiotique [4145]. Semita [889]. Semite [889]. sémitiques [19]. Senebier [3779]. Sénégal [4538]. Senglaub [2841]. Sens [2220, 4022, 3932, 2937]. sense [3697]. sensibilité [4153]. Sentier [4201]. Seoane [2965]. séparent [2347]. Sepp [2479]. Sepper [3814]. sept. [353]. September [3340, 2452, 489]. septembre [1744, 1748, 1860, 1861, 4409, 897, 1188, 1524, 1626, 4022, 2381, 2668, 4366, 3686, 1206, 1134, 101, 160]. septembre-19 [160]. septième [657, 2102]. ser [3735, 3732, 4435, 3608, 2406, 3596]. Sér. [3654]. Serbat [3879]. Serge [2711, 3677, 4393, 3515, 3881, 1115, 4529, 1906, 1751]. Sergescu [371, 208]. seria [2510]. Serie [2395, 1633, 1149, 4345, 3994, 566, 1324, 3359, 2411, 4615, 2169, 968, 1506, 1640, 1911, 354, 1270, 73, 4076, 2741, 2582, 879, 1170, 2324, 2644, 1338, 2432, 3582, 2583, 1574, 510, 141, 491, 855, 2854, 4477, 2812, 1314, 4441, 4625, 2693, 2613, 1032, 1173]. Series [3519, 2873, 2942, 485, 906, 3009, 3436, 3441, 2662, 3118, 3867, 2491, 2336, 3754, 3842, 1578, 3077, 3275, 3524, 2127, 2807, 4271, 4230, 3482, 3637, 3640, 3636, 3788, 4269, 3796, 3347, 3439, 2158, 4348, 1018, 3595, 4440, 4489, 608, 609, 631]. Serment [961]. Seron [3594]. Sérouya [74]. Serres [2443, 2494]. sert [4439]. serum [4670, 4670]. Servais [4395]. Servet [496, 511]. Service [3785, 963, 1001, 1002, 1059, 1114, 1742, 1817, 3572, 775, 3665]. services [2411]. servir [4296, 1105]. Servolo [4662]. Ses [1416, 3955, 379, 3158, 2938, 2254, 4660, 2256, 1676, 693, 1720, 289, 702, 4378, 3557, 4657, 3460, 3493, 3585, 4441, 2218, 2412, 3980, 1169, 1376, 2120, 796, 1000, 1168, 1569, 4048, 445, 4601, 4281, 2260, 2760, 2347, 1231, 4145, 2283, 409, 1294, 2755, 1832]. sesquicentennial [3886]. s’est [3722]. Seters [1169]. Sets [2054]. Settecento [4309]. Settembre [2808]. settimo [4161]. setting [3485]. settled [3637]. seu [4688, 529]. seul [4658]. seulement [4336]. Sève [4398]. Seventeenth [862, 2694, 1942, 4652, 3659]. Seventeenth-Century [862, 1942]. seventh [509]. Séville [1721]. Sex [3480]. sexangulaire [2307]. Sexl [2835]. sextants [4478, 4478]. sexualité [4436]. sexuelle [4624, 4436]. Sezgin [2545, 2229]. sfida [2567]. Sgard [3456]. ’sGravesande [359]. ’sGravesande-Musschenbroek [359]. Sguardo [4515]. Shaar [534]. Shaar-ha-Shama-yim [534]. Shackleton [704]. Shadish [3439]. Shallis [3001]. Shama [534]. Shan [3190]. Shanghaï [666]. Shape [3389, 2477, 3728]. Shapin [3654]. shaping [3904]. Shapiro [2697]. Sharaf [3239]. Shark [2836]. Sharlin [1418]. Sharon [2371]. Sharov [3610]. Sharples [3640]. Shea [4291, 4209, 2224]. Sheinin [4203]. Shelton [695]. Shen [3172]. Shenitzer [4203]. Sherman [2753]. Shermann [2753]. Sherrer [487]. Sherwood [89, 532, 820]. shifting [3277]. shifts [3959]. Shigehisa [4167]. Shippen [512]. Ships [4475]. Shírán [3206]. Shirley [2526]. Shizhen [3173]. Shlomoh [534]. Short [1500, 2581, 3565, 89, 4626, 37, 443, 1844, 951, 3046]. Show [2942]. Shryock [726]. Shu [1777, 3175]. Si [2648]. Sicht [3416, 2778]. Sicilia [565]. Sida [3360, 3272]. Sidereal [1543]. Sidereo [3566]. Sidereus [4051, 3522, 134, 1666, 1559, 1403]. Sidérurgie [1269]. Sidney [695, 966]. Sidot [3670]. Siebenthal [3795]. Siebold [4654]. siebzigsten [4177]. siècle [1855, 4549, 1987, 4572, 3572, 4481, 4524, 4477, 1774, 76, 174, 1538, 3227, 386, 4512, 618, 279, 3983, 1849, 1949, 967, 748, 2584, 658, 4468, 4410, 925, 1282, 2409, 2812, 1740, 2599, 4165, 25, 4003, 3189, 3334, 4482, 890, 2497, 2909, 3558, 892, 4434, 2711, 4664, 3809, 193, 1626, 4028, 3929, 1698, 647, 757, 1456, 4369, 3161, 3290, 2365, 2486, 4286, 4005, 4415, 4448, 3951, 4205, 759, 2900, 1379, 1662, 2014, 643, 570, 797, 838, 996, 1000, 1116, 1306, 1312, 1579, 1756, 2086, 2974, 3228, 3818, 4048, 433, 3779, 2469, 177, 291, 4337, 224, 79, 3428]. siècle [4362, 2858, 2962, 1337, 1762, 3374, 3780, 1513, 1655, 4578, 756, 3624, 4319, 3173, 1224, 1821, 3918, 3376, 1990, 1008, 3070, 4397, 4530, 2101, 1346, 4062, 3345, 4338, 4452, 4523, 2454, 1739, 1658, 3293, 3113, 572, 568, 928, 1283, 2013, 4570, 1964, 2102, 2206, 3825, 4166, 867, 1930, 4493, 4565, 1319, 3623, 3827, 4573, 4070, 4051, 4124, 3239, 4023, 573, 4265, 4620, 969, 1541, 4539, 815, 1014, 2087, 2394, 869, 621, 848, 1052, 388, 3244, 1615, 1584, 2246, 75, 547, 550, 2677, 656, 4538, 4437, 4470, 3571, 3621, 2809, 4286, 3793, 4173, 2396]. siècle [2754, 2512, 4411, 4645, 1850, 4413, 3273, 2218, 2577, 2251, 3993, 963, 964, 1115, 1245, 1309, 1638, 1704, 2032, 2066, 2188, 2845, 3043, 1772, 3213, 4009, 4034, 3695, 1735, 3215, 3261, 1930, 1430, 1431, 654, 3013]. siècle-1870 [2812]. siècle-1940 [2974]. siècles [3226, 197, 4013, 919, 2379, 159, 1207, 280, 2238, 4369, 4491, 3875, 4289, 1148, 1305, 1579, 1938, 2354, 2915, 2828, 1781, 4403, 594, 4281, 4619, 3924, 2293, 3157, 4283, 3295, 3224, 3453, 4455, 1851, 839, 4558, 4333, 458, 4644, 3881, 3462, 952, 1172, 1691]. Siegel [3464]. Siegfried [1117]. Siena [4076]. Sigalia [2831]. Sigaut [4395, 3896, 2273]. Sigerist [999, 2110, 683, 1690]. sight [4627]. sigla [4198]. Siglo [566, 3994, 1324]. siglorum [3516, 4198]. Siglos [3437, 1324]. signalétique [738]. signé [1425, 4363]. signes [1483, 4145]. Significance [789, 2228, 799, 3120, 3368, 2333]. Sigrist [3631, 3348]. Silverstein [905]. Silvester [2306]. Silvestro [3870, 2643]. Silvia [4351, 4362]. Silvio [1415, 2271, 1893]. simbòlica [3078]. Sime [3871]. similé [906]. similitude [3725, 1993]. Simon [3654, 10, 2629, 1590, 3312, 567, 4338, 856, 1299, 2881, 3368, 474, 4342, 4339, 4335, 4449, 1701, 3939]. Simondon [4367]. Simonetta [3368]. Simpson [327]. Simulation [4183]. Sinaceur [3897]. Sinan [4023]. Sinarellis [3935]. since [37, 1631, 2840]. Sinding [3490]. sine [4303, 1525]. Singe [2809]. Singer [1500, 508]. singolare [2643]. singularités [3624]. singulier [4464]. singuliers [1963]. Sinisgalli [2629, 2991]. sino [570, 1061]. sino-vietnamienne [570]. sino-vietnamiennes [1061]. Sinologica [607]. sinologiques [3760]. Sintesi [189]. Siou [606]. Sir [89, 1752, 235, 673, 1672, 859, 1171, 2925]. Sirmondian [487]. Sisco [211]. sistema [3901, 2965, 4098]. sistemi [4078]. Sitter [302]. Situation [2205, 2285, 1813, 1815, 83]. Sitzungsberichte [2392, 1028]. sive [4062]. Six [784, 3924]. Sixteenth [186, 2694, 1435, 2372, 420]. Siyuan [2420]. Skelton [2539]. Sketch [2586]. Skizzen [2566]. Skopová [1854]. skull [3363]. Slatkine [3213]. Sleinbuches [1605]. Sliousarev [584]. Sloan [3153]. slopes [4365]. Sluse [2996, 2969, 2953]. smallest [3599]. Smeur [1293]. Smintheus [380]. Smith [89, 1317, 1485, 3060, 3455, 3362, 3439, 1109, 4627, 211, 4502, 3655, 558, 1895, 3076, 823, 2703]. Smithsonian [282, 1982, 1255, 4178]. Smyth [1754]. SN [4009]. Snelders [3218, 2659]. Snorrason [3150]. so-called [4198]. sobre [3735]. Soc [3700, 1485, 3530]. sociabilités [4664]. Social [3486, 402, 2109, 1263, 1592, 3306, 2403, 3747, 2905, 2190, 453, 4018, 2517, 3933, 1432, 2846, 3274, 3788, 3743, 3016, 4646, 4190, 3967, 774]. sociale [3986, 827]. sociales [1736, 4393, 3043, 2347, 2273, 3382, 4445]. sociales^{≫} [3935]. socialism [3706]. socialismes [4646]. Socialized [2336]. sociaux [658]. Sociedad [4409, 3946, 300]. Société [935, 2193, 616, 633, 696, 801, 974, 1035, 1150, 2043, 2132, 3989, 2538, 1433, 4264, 4450, 4477, 4258, 3266, 4438, 4019, 4094, 3054, 3756, 3582, 3742, 2673, 4445, 724, 782, 765, 1229, 1429, 4226, 3631, 4612, 279, 2296, 4388, 459, 3433, 257, 3266, 3348, 3749, 2670, 224, 779, 3664, 701, 2288, 1075, 2583, 572, 928, 2673, 1005, 388, 1824, 3387]. Sociétés [2295, 1031, 1176, 4477, 4165, 4311, 4609, 3569, 3621, 3224, 3151, 1943, 1835, 3272, 2045, 1148, 2919]. Societies [4609, 4664, 4514]. Society [1594, 164, 743, 1004, 1317, 1912, 1259, 2920, 2321, 3363, 2641, 947, 2921, 2182, 1264, 1893, 3307, 2749, 1686, 610, 2816, 3744, 3587, 3638, 4206, 3864, 3597, 3830, 4612]. sociobiologiste [3730]. sociologie [4592, 4352]. Sociologies [4352, 3995]. Sociology [4077, 999, 4592, 3726, 4185]. Socrate [1537]. sodales [3148]. Soddy [2510, 1472]. Söder [4097, 4175]. sœur [4472]. Sofja [3667]. Sofya [2610]. soi [197, 4559, 4556]. soi-disant [197]. soif [1732]. soins [2451, 1869, 1938, 2844]. Soixantenaire [731]. Sol [2695]. solaire [3427]. solaires [1528, 668]. Solar [1334]. sold [3673]. Soleil [1161, 2491, 4634, 3661]. solid [4467, 4082]. solide [4628, 2291]. solides [3225, 1252]. solidorum [3225]. solitaire [4296, 392]. Solomon [3125, 3687, 4471, 4010]. Solovine [1145]. sols [290]. solution [2997, 2408, 52]. Solvay [4507]. solvent [3646]. som [747]. Some [1025, 3981, 3367, 2310, 1942, 837]. sommaire [1380, 2000, 344, 1179]. sommatoires [3190]. sommeil [4546, 93, 93]. Somnium [1664]. Son [3359, 46, 693, 1306, 4598, 3458, 2331, 3046, 737, 1515, 4473, 64, 137, 611, 949, 1630, 2028, 2255, 892, 1720, 4645, 301, 465, 962, 1380, 4298, 2597, 3057, 4674, 2651, 796, 1059, 1302, 1692, 103, 68, 976, 3952, 585, 2204, 1845, 4613, 2157, 145, 1478, 2786, 3098, 1687, 2026, 1322, 1852, 4566, 24, 2994, 3093, 3230, 2446, 408, 516, 2243, 2350]. Sona [4514]. sondern [3314]. Sonnen [1693]. Sonnen- [1693]. Sonntag [3418]. sonore [4359]. Sons [282]. sont [2411, 4581]. Sonya [3053]. Sophie [3765, 2664]. Sopra [4078]. Soranos [4147]. Sorbiers [2067]. Sorbonne [297, 2613]. Sorcerer [2109]. sorciers [1772]. sortes [883]. Soto [4271]. Soubeyran [3950]. Souchon [2611]. Souchon-Rouyer [2611]. souci [3029]. Souffrin [4307, 4306, 4300]. Souleymane [3310]. sound [2547, 4598]. source [3652, 1779]. Sources [4168, 4546, 3833, 1752, 3501, 1975, 4356, 3011, 3097, 3482, 3640, 2030, 1685, 1966, 4492, 3420, 2419, 3897, 1543, 1728, 4321, 3987, 2188, 3648, 2293, 1423, 3626, 3424, 463, 1572, 835, 2211, 2592]. souris [3473, 3958]. Soury [1820]. sous [2050, 1819, 2411, 330, 3557, 4160, 1064, 3753, 1115, 1480, 2234, 2324, 2679, 4249, 3648, 3044, 4608, 4565, 1079, 4509, 774, 3721]. sous-série [2411]. soutenue [2448]. soutenues [280]. Souvenirs [80, 159, 1523, 1009]. Souvenirs. [1027]. souveraine [670]. Soviet [3565]. soviétiques [2349]. Spa [747]. space [4348]. Spade [3815]. Spagyrical [3080]. Spallanzani [3062, 1614, 3240, 4532, 1109, 2882, 454]. Spalt [3444]. Spanien [2671]. Spary [3838]. spatiale [4335, 4629]. Special [2746, 2193, 4366, 2583, 528, 2246, 2747]. spéciales [3791]. spécialisation [4482]. specierum [3126]. species [3632, 1578]. speciosæ [4136]. spectacle [3710]. spectrale [3427]. spectre [3427]. spectroscopy [3169]. Speculations [3945, 1282]. speculis [3126]. Speech [3362]. speeches [1867]. Speiser [36, 2408]. spéléologie [4589]. speleology [4589]. Spemann [3960]. Spencer [89, 2262, 4556]. spermatikoi [4401]. sphère [2627, 1927, 1818, 4336, 3665]. sphérique [4493]. sphériques [2408]. sphéroïdes [1927]. sphéropée [1817]. SPHN [3348]. sphygmologie [1173]. Spinelli [4224]. Spink [2578, 1289]. Spinoza [114, 4220, 2422, 4432]. spinozienne [4221]. spirales [2055]. Spirit [2511, 4657, 2932]. spiritisme [4154]. Spirits [4143]. spiritual [2372, 874]. spiritualistes [4231]. Spitalebor [959]. Spitz [3703]. spontanée [158]. sprachlichen [4093]. Spranzi [4368]. Spriel [1472]. Spronsen [1611]. St [1504, 2479]. Staat [2391]. Staats [2872]. Staats- [2872]. Stachel [3270, 2565]. Stadt [1693, 1935]. Stadtarchivs [1693, 1935]. staff [3733]. staffs [4478]. Stafleu [1780]. Stahl [2141]. Stahlhochbaus [1350]. Stahlman [1334]. Stallybrass [2171]. Stampioen [1738]. standard [3902, 3101]. Stanislao [3383]. Stanislas [884]. Stanley [2546, 211, 3106, 2244, 1612]. Star [1934]. Stark [240]. starlight [3169]. Staroselskaïa [277]. State [3486, 1933, 4561]. statecraft [3259]. States [1502, 2686, 1271, 1354, 4656]. Station [2517]. Statique [2704, 2108]. Statistical [2923]. statisticians [2878]. Statistics [2923, 2877]. Statut [2839, 3422, 3359, 3138, 3856]. statuts [1375, 1376]. Staum [2709]. Stazione [2551]. STBS [2855]. steam [89, 2848]. Stearn [2539]. Stefan [4243]. Stefano [3562, 3150, 4207, 565]. Steffen [2674]. Stein [2877, 2877]. Steiner [188, 240]. Steinke [4243]. Steinle [3419]. Stelsels [514]. Sten [748, 876, 477]. Steneck [2182]. Stengers [4011, 3559]. Steno [3150, 861, 3755]. Stenone [3150]. Stensen [3755]. Stensens [690]. Stensiö [327]. Stepan [2378]. Stephan [3885]. Stéphane [4437, 4346]. Stephen [4236, 3060, 2447, 3593, 2777, 1936, 1472, 3432, 2127]. Sternen [1666]. Sternnamen [1100]. Sternuhren [1693]. sterrekunde [396]. sterrenkundigen [302]. Steven [3654, 3763]. Stevens [3754, 3796]. Stevenson [3060]. Stevin [10, 2629, 1590, 567, 856, 1299, 474]. Stewart [525]. Sticker [1680]. Stillman [2289, 2630, 2610, 1665, 1103, 1968]. Stipisić [2184]. Stock [2306]. Stockholm [747, 4388]. Stoffel [3844, 3349]. Stoic [4402, 4404]. stoïcien [3915]. stoïciennes [4403]. Stokes [3243]. Stomachion [2055]. Stones [3650]. Stootwetten [990]. Stora [748]. Storia [2344, 2297, 989, 1034, 1178, 1509, 2787, 3454, 531, 1618, 4076, 4270, 4435, 4098, 4309, 210, 2806, 3339, 3365, 2941, 2406, 4083, 4476, 4161, 4122, 505, 994, 2882, 1029, 449, 2814, 248]. storica [4344]. Storico [602, 189]. storico-critica [189]. Störig [1788]. Story [282, 2472, 1472, 2993, 2918, 1472]. Stotz [2264]. Strabon [1845]. Strasbourg [3343, 3965, 941, 462, 4679, 4291]. strasbourgeois [2666]. Strassburger [3734]. stratégies [4567]. stratigraphie [2597]. Straus [946]. Strauss [4341]. Strbánova [4514]. Street [2855]. stridulation [1456]. structurale [2255, 3920]. Structure [3027, 2163, 3258, 3801, 1965, 3495, 1704, 3175, 2833, 4523, 867, 2350]. Structures [4314, 944, 279, 3340]. Structures^{≫} [944]. Struik [207, 856]. Strumenti [2395, 3365, 2173]. Strumenti-Letteratura [3365]. Stuber [4243]. Stück [3419]. students [443]. Studi [3694, 4515, 2976, 2606, 2395, 3936, 2358, 3523, 3365, 3583, 3923, 4072, 4083, 2784, 2882, 4067, 4076, 3438, 3813, 3491, 2808, 3727, 3105]. Studia [2604, 2605, 4243, 4202, 2423, 2770, 3011, 3418, 2334, 2501, 2525, 2252]. Studie [3757, 2392, 3314]. Studien [2536, 2467, 4093, 1526, 3416, 2695, 3316, 3443, 1668]. Studienplan [3429]. studieresor [507]. Studies [384, 163, 1931, 2309, 2320, 2831, 3533, 3880, 3700, 3826, 2262, 3119, 2192, 3501, 2664, 1152, 1317, 1485, 1944, 218, 2928, 3118, 3786, 3886, 3650, 2506, 2565, 3417, 3531, 3584, 3589, 3867, 1162, 3312, 2876, 3788, 2477, 1263, 1982, 3525, 874, 4207, 4269, 3754, 3796, 3842, 3762, 3002, 3500, 3699, 3432, 4188, 3921, 3964, 2525, 3812, 3816, 3871, 2572, 4036, 4440, 4317, 2373, 240, 2457, 3006, 2495, 4458, 4079, 4271, 1435, 3581, 3797, 3990, 4230, 3482, 3637, 3640, 1432, 3532, 4009, 3860, 3016, 3439, 4055, 3595, 4209, 4489, 4197, 1968, 1866, 3521]. Studies [2321, 2062, 2659]. studio [4235]. Study [402, 3119, 2228, 1004, 1912, 191, 1442, 3125, 2477, 2306, 442, 2242, 2737, 4520, 1435, 3749, 3637, 2322, 3553, 3243, 420, 609, 2521, 3651]. Stuewer [2338, 2638, 4079]. Sturm [3099, 3329, 4516]. Sturtevant [2539]. Stüssi [479]. style [2981]. styles [4433, 3524]. Suan [1777]. suaréziennes [4224]. Suaudeau [60]. subjects [4662]. sublime [2372]. subreptice [4013]. Subsidia [2216]. Subsídios [529]. substantiel [2857]. substantielIe [2857]. subtile [228]. successeur [4469]. successeurs [3713]. such [4582]. Suchodolski [2352, 2154]. Suchowa [1060]. Sud [1822, 1309, 2234, 887, 1846]. Sud-Est [1846]. Sud-Ouest [2234, 1846]. Sud-Vietnam [1309]. Sudhoffs [4125, 2473]. Sudre [445]. sue [4301]. suédois [659, 507, 1615]. suédoise [507, 616, 801, 1150]. Suez [4021]. suicide [4529]. suisse [2193, 2296, 633, 696, 974, 1035, 2043, 2132, 2670, 414, 503, 580, 650, 725, 765, 1534, 854, 289, 4483, 1952, 687, 979, 1824]. Suite [407, 196, 270, 287, 480]. suites [1573]. suivants [3439]. suivi [2925, 3302, 3648, 4529, 4296, 668, 1260]. suivie [4286, 791]. suivies [375, 1292]. sujet [195, 1588, 4480]. sujets [1794, 4662, 834]. sukhoe [1610]. sul [4475, 3438, 3813]. sulla [2471, 3901]. Sulle [2914]. sull’eudiometria [3941]. Sully [2531]. Summa [3651]. Summaries [2726]. summary [2686]. Summer [3386]. Sun [4634]. sundials [4557, 2371]. sunspots [4527]. Sunt [2921]. Sunto [3383]. suo [3058, 3150, 210, 4122]. super [1163]. superconductivity [3840]. superficielle [3137]. supérieur [3043, 1724, 841]. supérieure [1821, 2745, 731]. supérieures [1941, 1978]. supérieurs [1458]. superstructures [1350]. Supper [2244]. Suppl [3488, 3218, 3611]. Supplement [940, 2400, 3197, 3555, 2754, 1524, 303, 3278]. Supplementa [2423]. supplémentaire [1914]. Suppléments [482]. surface [1621, 1622]. Surfaces [133]. surgeon [798, 4654, 420]. Surgeons [727, 2811]. Surgery [2550, 673, 608, 460]. surgical [1481, 4664]. surveillée [4454]. Survey [3155, 1910, 3076, 2155, 1924]. surveys [3196]. Survivance [886]. Susan [3868, 2523]. Susana [4320, 4072]. Susanna [4239]. Suse [1290]. Susiane [1290]. suspendu [3426]. Sutton [4024]. Suzanne [2411, 3309, 2325, 4080, 3736, 4285, 902]. Svedberg [36]. Sven [507]. Svenska [507]. Sverige [2475]. Sviluppi [3491]. Sviluppo [189, 2629, 602]. svolte [2806]. Swan [89]. Swanenburg [325]. Swazey [2322]. Swedish [477]. Swijtink [3434]. Swiss [3663]. Sworykin [1412]. s’y [357]. Sydney [3060, 3455]. Syed [3650]. syllogistique [2076]. Sylvain [4491, 4529, 4319, 4296]. Sylvestre [1088, 520, 849, 870]. Sylvestre-Francois [1088, 520, 849, 870]. Sylvie [3654, 3884]. symbiose [4278]. Symbolic [4578]. symbolique [3198, 1989, 4578, 3800, 3999]. symbolische [2740]. Symbolismus [2740]. symbolistes [3309]. Symmetrie [3222]. Symmetry [3139]. Symposia [953, 3347]. Symposion [4125]. Symposium [1744, 1822, 3340, 783, 1091, 1141, 1747, 3083, 3594, 2423, 3362, 1432, 1827, 2180, 1068, 3368, 2659, 3965, 934, 1285, 1748, 2961, 3165]. synchronismes [698]. syndicats [1756]. Synthé [3591]. Synthélabo [3949]. Synthese [1931, 3866, 2332, 3220, 446, 1032, 585, 2260, 3149, 3052, 3953, 5, 2260, 421, 354]. Synthèses [4628]. Synthesizing [4095]. Synthetical [747]. Syphilis [4062]. syphilographie [1174]. Syrie [1605]. System [2172, 2511, 89, 4652, 4404]. systematische [4053]. Système [1378, 679, 813, 407, 2722, 472, 3342, 1611, 582, 2724]. systèmes [4368, 2420, 4398, 4282, 3722, 4051, 4026]. Systémique [3606]. Systems [1665, 3988]. Szabó [2441, 1926, 2053]. Szczeciniarz [4174]. Szemle [1897].

T [1267, 1265, 2831, 37, 2248, 1631, 3410, 1727, 3669, 377, 3061, 2044, 2522, 447, 984, 1232, 1467, 1669, 1670, 1955, 1929, 2057, 2096, 3958, 1268, 2848, 1230, 1521, 1846, 3307, 1392, 3036, 1570, 1572, 2355, 2874, 3926, 2492, 2742, 2521, 2737, 2923, 384, 3355, 939, 1749, 2964, 3124, 3278, 748, 795, 44, 30, 1943, 2411, 617, 843, 1033, 1121, 1177, 1315, 1444, 1511, 1508, 1619, 1796, 354, 1892, 2423, 3746, 3791, 303, 4076, 4435, 3521, 3585, 2045, 2975, 665, 744, 818, 819, 1953, 2974, 109, 1072, 1568, 1608, 1753, 1789, 1927, 2055, 2056, 2095]. t [2107, 582, 1338, 4528, 4642, 3149, 249, 1142, 4397, 217, 1344, 3736, 3879, 3962, 4007, 4146, 4147, 1604, 2105, 2405, 1605, 1661, 742, 1574, 4394, 1430, 1431, 141, 192, 307, 383, 566, 790, 1290, 1324, 2156, 3720, 2242, 2973, 405, 488]. T. [1266, 2055, 2093]. Tân [2845]. têtes [4405]. Tabanelli [1503, 2077, 2173, 1441]. Tabarroni [2808]. Tabibito. [3156]. tablas [1436]. Table [1962, 166, 259, 334, 423, 538, 634, 705, 802, 910, 1036, 1124, 1192, 1272, 1364, 1447, 1546, 1642, 1652, 1654, 1711, 1653, 1799, 1871, 2079, 2199, 2277, 2359, 2480, 2552, 2615, 2713, 2789, 2887, 3538, 3614, 3679, 3769, 3846, 3906, 3974, 4038, 4100, 4148, 4212, 4273, 4322, 4371, 4418, 4459, 4497, 4533, 4648, 3648, 3408, 4510]. Tableau [698, 1179, 1357, 19]. tableaux [1169]. Tables [2424, 3011, 3097, 2095, 668, 1608, 3661]. Tabula [4206]. Tacchella [3884]. taches [1528]. tacticien [59]. Tagebuch [2384]. Tagliacozzi [420]. Tagung [2835]. Tai [606]. taijinkyōfushō [4656]. Taisbak [2826, 2054]. Tait [798]. Takakazu [3729]. Takebe [3729]. talets [748]. Talmy [1580]. Talon [1132]. Tamburello [3339]. Taming [3434]. T’an [666]. Tandon [4022]. Tannery [2527, 2658, 2964, 3124, 107, 109, 1015, 1072, 1467, 1667, 1955, 2107, 626, 4251, 551]. Tao [666]. Tapisseries [1409]. Taqī [2372]. Taqī’l [2372]. Taqī’l-Dīn [2372]. Tarde [4546, 4176]. tardo [4309]. Tardy [4054]. Tarr [726, 1288]. Tarski [4056]. Tartaglia [2567]. Taschenbuchreihe [1413]. Tasso [1666]. Tat [1234]. Taton [1327, 3013, 167, 3278, 41, 652, 2662, 4244, 1662, 818, 819, 2354, 355, 374, 792, 1328, 1959, 2231, 1661, 2307, 1430, 1431]. Taub [3604]. Tauber [3762]. taupe [380]. taxa [3368]. taxidermie [4283]. taxinomie [2027]. Taxonomic [1780]. taxonomique [1625]. Taylor [89, 2838, 3271, 532, 4440, 820]. Tche [1390]. Tche-wou [1390]. tchécoslovaque [682]. Tchécoslovaquie [1748, 1843, 1626, 1769, 2394]. Tchen [773]. Tcheng [666]. tchèque [1655]. tchong [883, 773]. tchou [883]. te [514, 991]. Teach [420]. Teaching [1913, 2115, 1246, 1484, 4655]. teachings [4679]. teaterhabitué [876]. Teboul [3178]. techn [3654, 3700]. Technical [1591, 1854, 1432]. Technick [1734]. technické [1854]. technico [2653]. technico-industriel [2653]. techniek [513]. Technik [1709, 1422, 2674, 1777, 1890, 2566, 1412]. Technikatörténeti [1897]. Technikgeschichte [1356, 1890, 1595, 1596]. Techniki [971]. Techniky [2394]. Technique [1747, 279, 3400, 1248, 1736, 3931, 343, 1306, 150, 2693, 1889, 3357, 1724, 277, 2756, 479, 660, 678, 1429, 1392, 967]. Techniques [1257, 217, 3096, 1437, 738, 1730, 283, 971, 1474, 2677, 1247, 3260, 3349, 3785, 4165, 4367, 4482, 615, 728, 1270, 1416, 1720, 1880, 3212, 4663, 4255, 3569, 3413, 2625, 1663, 4592, 3157, 1889, 1342, 3475, 2368, 4166, 1392, 3633, 1403, 1165, 4169, 2273, 2394, 3415, 446, 140, 1887, 4021, 1344, 1897, 1922, 2272, 3881]. Technischer [3223, 2849]. Technisches [1356, 1595]. techno [3824]. techno-scientifique [3824]. Technogenèse [4162]. technological [2905, 4210]. Technologie [1255, 4140, 2588, 3993, 1413, 1727]. Technologies [4646]. technologique [2239]. technologue [2020]. Technology [2522, 1565, 2672, 1982, 4592, 3307, 667, 2807, 4182, 3880, 2190, 4663, 4478, 3638, 3706, 4456, 3080, 4183, 4167, 4188, 3177, 4197, 1267, 1265, 2052, 2522, 1271, 1354, 2855, 3836, 3016, 1593, 1888, 1264, 3726]. Tecnica [1324]. técnicas [4409]. tecniche [4309, 4161, 2173]. tecnico [528]. tecnologie [4161]. tectonics [4572]. tectonique [4572]. teessis [975]. Teil [1216, 1332, 2355]. Teilen [1216]. Teilhard [327, 1478]. Teixeira [3791]. Telegraph [1731]. télégraphe [2347]. télégraphie [25]. Telescope [3940, 2444]. télescopes [4124]. Telescopio [1333]. télévision [4162, 1406]. Telkes [4477, 4441]. Telkes-Klein [4477, 4441]. Telle [2473, 2695, 4593]. Temkin [946]. témoignage [433]. témoignages [1242, 2868]. témoin [1741, 4130]. témoins [4655, 4518]. temperature [3840]. Temple [2779]. Tempo [3870, 3150, 3759, 3058]. temprana [1907]. Temps [968, 4415, 945, 3823, 1615, 3552, 1972, 2028, 301, 3557, 3313, 606, 796, 103, 4347, 3055, 106, 568, 1775, 2786, 4379, 1865, 24, 2156, 2243, 2973, 2673, 4298, 1845, 2786, 2825, 2026, 860]. Temps/Société [2673]. tenax [2499]. Tenerife [1108]. Tenon [1950, 3596]. tenore [901]. tentative [1089, 3953]. Tentatives [1583, 4484, 1282, 1416]. Tenth [2826]. tenu [3343, 2319, 3299, 3722]. teologia [4344]. teoria [88, 3437, 3829, 1980]. teoriche [2806]. teóricos [4453]. Teorija [2184]. Teplov [1332]. terapéutica [3735]. tératogenèse [1284]. tératologie [4677]. tératologique [2004, 2005]. teratology [4677]. Terence [1936]. Teresa [4202]. terme [2765, 2206]. termes [606, 1483]. terminologie [615, 984]. terms [3584]. Terracini [831]. Terrada [3735, 2644]. terrain [4155]. Terre [2657, 4091, 1338, 3948, 3348, 3924, 2762, 3069, 2858, 2013, 4087, 3672, 44]. terrestre [2119, 2657]. Terreur [761, 4405]. territoires [4155]. territoriale [3279]. Territories [3154]. Territory [1417]. tertium [2873, 2287]. terza [2739]. Terzian [2992]. terzo [829]. Tessier [4628]. test [4095]. Testart [3152]. Testi [2694, 3694, 3936, 3438, 3523, 3583, 3923, 4072, 4067]. Testimonia [2627]. Testo [4078, 4301]. tétrade [524]. Tétry [2774, 3152, 1167]. Tetty [2317]. Teubneriana [2106]. Text [3516, 2695, 1795, 2158, 3637, 3932, 3420]. Text- [2695]. Textbeleg [2706]. Textbook [2608]. textbooks [4655]. Texte [2661, 3536, 2473, 3241, 4643, 59, 2967, 3932, 1663, 3725, 1437, 2443, 1403, 1474, 2540]. Texten [1440]. Textes [4433, 2374, 2166, 3654, 2504, 2913, 2750, 3121, 4196, 4436, 3097, 4286, 4491, 4641, 2029, 4545, 3079, 2354, 355, 3316, 3608, 4120, 3055, 4080, 4084, 3612, 3698, 3147, 2786, 2376, 4166, 2771, 3239, 3609, 2291, 3244, 378, 4546, 1774, 4050, 2933, 1102, 4287, 3199, 1143, 4642, 4510, 1840, 4591, 4296, 2386, 1730, 2592, 1290]. Textiles [1341]. texto [1436]. textologie [2126]. Textos [2965, 2644]. Texts [2309, 2320, 3119, 3812, 4301, 3640, 3196, 4591]. Teysseire [2844]. Th [487, 2757]. Thabit [689, 3241, 1101]. Thalès [2783, 217, 1177, 422]. Thalès. [1727]. Thanh [1305]. Theater [3755]. Theatrum [1685]. Théétète [3882]. Theil [2591]. Their [3810, 1415, 1345, 4301, 747, 745]. Themas [2701]. Themata [2822]. Thematic [2335]. thématique [2898]. thème [3530]. Thémis [3824]. Thénard [304, 304]. Théo [3007]. Theodor [964, 1439]. Theodore [905, 3922, 3434, 1024, 1147, 4529, 3607]. Theodoric [608]. Théodoridès [1909, 1241, 2575, 1148, 1170, 1613, 2579]. Theodosian [487]. Theol [3815, 4097]. Theologie [4175, 4234, 2287, 3443, 3346, 1394]. théologiques [4059]. Theology [402]. Théophile [1831, 2574]. Théophraste [3611, 3736]. Theophrastus [3482, 3611, 3640]. theorem [3933, 3932]. théorème [3099, 3352, 2399, 3204, 1322]. Theoretical [1979, 4268]. théoria [154]. theoriam [2873]. Theoricæ [4548]. Théoricien [392, 587, 3687]. théoriciens [4604]. Theorie [2478, 3443, 4142, 4503, 944, 2184, 3857, 4554, 4549, 4355, 2605, 3422, 157, 4112, 2689, 4471, 3121, 2622, 3137, 1698, 2637, 4136, 4431, 2455, 4427, 3069, 1783, 1992, 175, 3646, 850, 3149, 4597, 4505, 4587, 2301, 1224, 3661, 2725, 1398, 2074, 2760, 3147, 4452, 1739, 2376, 2404, 2748, 1087, 3877, 3507, 3690, 3953, 3496, 1632, 243, 375, 860, 1326, 1024, 3382, 3643, 2596, 2350, 3081, 2907, 2169, 1965, 2029, 2469, 2013, 2820, 2656, 2828, 2370, 790, 858]. Theorien [2841, 2603]. Theories [2524, 3340, 3693, 3277, 3424, 2127, 4541, 2584, 2104, 3178, 398, 4028, 2486, 4357, 357, 4403, 2857, 2102, 4394, 3857, 3277, 3415, 2684, 2900]. théorique [1282, 3557, 106, 1840, 668]. théoriques [3360, 4412, 2634]. theorists [4604]. Theory [2753, 2831, 2228, 2664, 3009, 377, 1866, 3118, 1476, 2054, 1236, 3762, 1472, 1976, 2746, 2990, 1999, 3464, 89, 3516, 3697, 2877, 3236, 3110, 4507, 2928, 4260, 4020, 4203, 2477, 3646, 4597, 4587, 4185, 3702, 2725, 3889, 3107, 1439, 3839, 3410, 3276]. Therapeutic [2322, 966]. thérapeutique [2421, 890, 996]. thérapeutiques [3919, 2188]. Therapie [2031]. therapies [4364]. Therapy [631]. There [4582, 3659]. Thérèse [4434, 4259, 3884, 1523, 2693, 1230, 1521]. thermale [4625]. thermales [2188]. Thermodynamics [1891]. thermodynamique [2377]. Thermometer [1694]. thermométrie [776]. Thesauri [3148]. Thèse [279, 1306, 1307, 462, 157, 2136, 2729, 797, 1579, 1939, 2448, 1458, 279, 3499, 1303, 1305]. thèses [280, 2238, 1706, 4058]. Thévenin [3151]. Thévenot [2177, 2099, 1374]. Thi [2708, 2845]. Thiele [4177]. Thierry [3654, 3746, 3745, 3724, 4455, 4621, 4130, 2064]. Thillaud [2234, 2448]. thing [4582]. things [3650, 4095]. thinkable [3646]. Thinkers [2271]. thinking [3814]. Third [2853]. Thirring [605]. Thirteen [627]. Thirties [3037]. Thom [2999]. Thomas [2166, 1752, 4488, 3940, 3488, 1505, 2473, 727, 882, 1481, 395, 4121, 2992, 1103, 1976, 1910, 1268, 2528, 2993, 938, 2926, 3732, 2353, 4361, 627, 1957, 3046, 4132, 2586, 2397, 492, 3364, 1351, 4559, 2526, 789]. Thomasset [3884]. Thompson [693]. Thomson [3899, 2639, 4279, 3688]. Thorkild [2233]. Thouars [2122]. Thought [3486, 402, 1863, 508, 2979, 3524, 1924, 1289, 821, 2335, 3440, 3482, 3640, 938, 1999]. Threat [3364]. Three [1866, 3125, 833]. Thuillier [1735, 2675]. Thunberg [1615]. Thuong [570]. tiempo [3946]. Tierra [1324]. Tiger [2836]. till [507]. Tillier [3407]. Tilling [2493]. tilredande [747]. Time [3082, 2242, 2521, 2737, 3001, 3718, 3796, 3347, 216, 1686, 1388]. Timée [3030]. Times [771, 3371, 420, 3668]. Timm [1413]. Timmermans [450]. Timotheus [247]. Timpanaro [134, 200, 505]. Tin [1592]. Tinemue [4198]. Tinkers [2271]. Tintinnabulum [4296]. tirage [1330, 2141, 790]. tiré [3676, 1893]. tirées [2350]. Tissot [2449]. Title [2581, 3727, 2979]. titled [1472]. titre [2324, 1559]. Tobias [1336, 2727, 3660, 2872]. Toby [2811]. Tocco [1216]. Todestages [1107, 942]. Todhunter [1296]. together [1020]. tolemaico [4078]. Tollot [3348]. Tom [4626]. Tomás [3735]. Tome [2884, 1667, 2293, 1889, 1056, 1233, 1962, 167, 166, 259, 334, 423, 538, 634, 705, 802, 910, 1036, 1124, 1192, 1272, 1364, 1447, 1546, 1642, 1711, 1799, 1871, 2079, 2199, 2277, 2359, 2480, 2552, 2615, 2713, 2789, 2887, 3538, 3614, 3679, 3769, 3846, 3906, 3974, 4038, 4100, 4148, 4212, 4273, 279]. tomes [1596, 1798, 2324, 2993]. Tomic [4547]. Tomo [829]. ton [3503]. Tong [883, 883]. Tonnelat [2767, 1674]. Tons [1789]. too [2850]. Tools [3863]. Toomer [3420, 3014, 3420, 2419]. topographer [4617]. Topologie [4145, 3133, 3100, 2213]. topon [3503]. Torbern [747, 747, 793]. Tore [3485]. Torlais [1494]. Torres [3693]. Torricelli [3354, 2161]. Torrini [2490, 4071, 3523, 3923]. torsion [3694]. Torsten [1267]. tort [1682]. Torun [2604, 2605]. Toscana [2976, 2358, 4083]. Toscane [1359]. tot [302, 325, 990]. total [3213]. Totems [4045]. Tóth [1750]. Totok [2423]. t’ou [1390]. t’ou-k’ao [1390]. touchant [3499]. touche [4336]. Toulmonde [4643].

Toulouse [3594, 4281]. tour [4118]. tourbillon [4296]. Tourisme [466]. tournant [3226, 4013, 2150, 3453, 4558]. tourne [3672]. tournée [1950]. Tournefort [622, 863]. Tournès [4617]. Tournon [485]. Tours [2045]. tous [1972]. tout [4246, 4219, 3724]. toutes [4157]. toux [3188]. Touzé [3212]. Towers [1301]. tra [4515, 4072]. Tracasseries [4385]. Traces [4367, 4024]. trachéotomie [1308]. Tract [2320, 3046]. Tractat [904]. tractation [291]. Tractatus [2094, 789, 4122, 2158, 733]. Trad [3764, 3725, 1828, 726]. Trade [4608]. Trades [1078]. Tradition [3303, 2707, 4417, 3432, 2694, 4128, 3648, 3158, 3101, 442, 1539, 3381]. traditionnelle [3188, 1131]. traditionnelles [886]. traditions [3174, 3158, 2272, 4565, 3896]. tradizione [3583]. Traduction [1181, 2326, 975, 559, 2176, 1675, 3777, 4286, 2460, 3930, 3206, 4643, 1472, 3607, 2364, 510, 1559, 2820, 530, 4581, 1260]. traductions [3267, 1676, 644]. traduire [1407]. traduit [481]. traduites [1145, 860]. traduits [151]. tragique [753]. traitait [1916]. traitant [2367]. Traité [3711, 4545, 279, 3802, 890, 1113, 570, 2800, 3784, 249, 1362, 668, 4475, 3709, 985, 2353, 4027, 1474, 3659, 4132, 71, 4134, 4488, 419, 3791, 1558, 604, 1222, 3512, 2356, 4133]. traitement [748, 3068, 3499, 751]. traitements [4477, 3881]. traités [889, 2272, 1437, 3760, 58, 1185]. Trajectoire [4597, 4487]. Trajectoires [4052, 4266]. Trajectory [4597]. Traktat [2473]. Trames [4662]. tranchées [4405]. Transactions [3363, 2641, 4206, 2921, 1893, 610]. transatlantic [3916]. Transcendental [2640]. Transcription [2564]. transfini [4513]. transformation [919, 2463, 4199, 2743, 4569]. transformed [4365]. transformer [3623]. transformisme [357, 2284, 327]. transistors [1251]. transition [4519]. Translated [531]. Translation [3119, 508, 3118, 3125, 3126, 3651, 3420, 689, 3503, 1792, 3587, 4348, 3988, 4581, 4206, 2733]. Translations [2321]. Transmission [4131, 3831, 3896, 4028, 4510, 112, 1587, 1885]. Transmissions [4281, 2217]. Transmutation [1578, 3643, 1472]. transparence [4163]. transparente [1818]. transsubstantiation [4227, 4064]. transversalité [2304]. Trascendencia [62]. trattati [3562]. trattato [2914]. Traugott [1111]. Travail [1736, 855, 3448, 2123, 2416, 4279, 2239, 457]. Travailleurs [2673]. Travaux [1373, 1538, 279, 2870, 1943, 1835, 2637, 4033, 2582, 3651, 3924, 2436, 2879, 3258, 2386, 2373, 4288, 969, 4314, 2599, 2295, 2168, 1177, 3331, 1258, 2622, 3669, 1698, 4378, 4286, 4491, 1884, 2000, 3521, 4545, 797, 2603, 151, 1959, 3729, 4084, 558, 1401, 217, 2157, 2494, 2098, 2149, 752, 3168, 160, 1179, 246, 247, 3558]. travers [3348, 4441, 2927, 4472, 859]. Treatise [3125, 3516, 3815, 1847, 460, 747]. treatises [689, 4652]. Treatment [2578]. Treaty [1950]. Trebizond [3758]. Tree [2883]. Treguier [4434]. Treize [1938]. tremblements [3924]. Trembley [3631, 455]. Trendelenburg [533]. Trends [3306, 3700]. trenie [1610]. Trenii [1610]. trente [3687, 3530]. tres [4084]. tres-eccelent [4084]. Tresbens [904]. Treub [950]. Trevor [3567, 2838, 3312, 3724]. Trévoux [2844]. Tribiblos [3607]. tribulations [2206]. tribute [2667]. Tributes [3271]. Tricentenaire [1679, 1190, 199, 121, 54, 786]. Trichomonas [2499]. Tricot [7]. trigonométrie [4493]. Trillat [4574]. trimestre [279]. trimestrielle [2193, 2296, 617, 633, 696, 974, 1033, 1035, 1444, 1945, 2043, 2132, 192, 307, 383, 488, 971]. Trinh [2708]. Trinity [2564]. triomphe [4588]. triplici [4069]. Trismégiste [8, 58, 3648]. Triton [1763]. triumph [4588]. triviales [3190]. Trochet [2317]. Troels [3755]. Trois [919, 2256, 3555, 3348, 273, 883, 4531, 4280, 4319, 2653, 1557, 3403, 2060, 1145, 2087, 2349]. Troisième [1032, 4509, 2996, 3334, 1377, 2909, 105, 1523, 4249]. Trop [2221, 4219]. tropical [1312]. Tropp [3570]. Trouessart [2367]. Troupeau [2669]. Trousseau [962]. trouver [2283]. Troxler [3123]. Trudaine [193]. true [4478, 4634]. Truesdell [2274, 1151]. Trystam [2734]. Trystram [2734]. Tschernyschew [1412]. Tschirnhaus [2660]. TSF [3402, 3401]. Tsi [606]. Tsing [666]. Tsiun [1390]. Tu [2708]. tube [3402, 2221, 3399, 4095]. Tuberculosis [631]. Tugnoli [3491]. tumult [2937]. tumulte [2937]. tumultueuse [3805]. Tunis [886]. Tunisie [886]. tunnel [465]. turbine [89, 1340]. turbines [1340]. turbulences [2443]. turc [1112]. Turchini [956]. Turco [4435]. Turgot [330]. Turk [1388]. Turkey [4167]. Türk’ün [1388]. Turnbull [2803, 1142]. Turner [3345, 3763, 4267, 3082, 2700, 3603]. Turning [2338]. Turpin [1992, 4363]. Turuq [2372]. Tusî [3239]. tute [3039]. två [31]. Twelfth [2306]. Twentieth [3000, 2294, 2578, 3863, 4664, 4573]. Twentieth-Century [3000, 2294, 3863]. two [1665, 3482, 4294, 1414, 2289]. Tycho [3120]. Tychonic [4652]. Tylecote [1261]. Tyler [1894]. Tyndall [1109]. Type [4029, 3876]. types [1780]. typescript [940]. typographique [1404]. Typologie [3626, 2592]. Tyskland [876]. Tytus [3317].

U [2331, 2104, 2580, 278]. U.R.S.S. [936, 1770, 1825, 1862]. U.S [1910]. U.S.A [1848]. U.S.A. [416, 441, 688, 853, 899, 1068]. Übersetzung [1795, 1864]. Ucelli [1343]. Udo [3316]. Ugo [1074, 1099, 1291, 1426]. Uhr [3734]. Uhren [1262]. uit [514]. Ulla [4530, 4265]. Ullmann [3764, 2667, 3816]. Ulrich [3040, 3148, 3346]. Ulrike [4165, 3840]. Ultrajectinae [2534]. Ultraviolet [966]. umanesimo [4078]. Umberto [4188]. uncertainties [3347]. Uncertainty [3517]. undecimum [844]. Understanding [3106]. Underweysung [2778]. Underwood [1500]. Unendlichen [3444]. Unesco [1743]. ungebändigte [4125]. uniformity [1847]. unimaginable [4256]. Union [1767, 617, 1033, 1444, 3083, 3565, 1827, 184, 192, 307, 383, 488, 202, 349, 415, 475, 931, 1053, 1561, 740, 1091, 1206]. unique [161]. uniquement [858]. Unis [1659, 1340, 4656]. unità [4098]. unité [813, 2584, 2433, 2839, 254, 68]. United [1502, 2452, 2686, 1271, 1354, 4656]. Univ [3519, 1317, 1485, 3933, 3932, 2137, 3864, 2613, 3785, 3460, 3965, 3924, 3736, 3722, 3633, 3596]. Univers [2260, 3676, 4671, 3561, 1288, 4113]. univers-bloc [4671]. Universal [525, 3646]. universalia [3422]. universalis [4635]. Universalismus [875]. universalité [3934]. Universe [508, 2922, 1893, 2710, 3604, 4671, 4404, 3570, 785, 2869, 3610, 3702, 4330]. universel [3646, 3984]. universelle [4305, 4495, 1331]. Universidade [529]. università [4076]. universitario [4068]. Universitas [842]. Universität [2536, 1526]. Universitatis [2233, 590, 2249]. Universitäts [2872]. Universitätsgeschichte [2478]. Université [1808, 1915, 672, 2582, 4261, 4196, 2407, 979, 2665, 1077]. Universités [3883, 4007, 2628, 2627, 1927, 2093, 3879, 3962, 4146, 4147]. University [2529, 1435, 1432, 1848, 1847, 1997, 2094, 2225, 2024, 1827, 3368, 4456, 2109]. Unmaking [4082]. unpublished [1919, 2529]. unserer [3035]. unsolvability [4321]. Untersuchung [4234, 3346]. Untersuchungen [3201, 2473, 1440]. until [771]. Unto [2333]. unveränderte [1470]. upon [903, 460]. Uppsala [31]. Uppsala-astronomer [31]. ur [747]. Urban [1413]. urines [462]. Urkunden [2106]. Urology [691]. Urs [4243]. Ursula [4622]. Ursus [3120]. usage [4016, 4283]. Usages [3694, 4663, 4281]. Use [1694, 4685, 3332, 4131, 1789]. User [3861]. User-driven [3861]. Uses [3629, 3432, 3312, 3756, 4663]. usine [4413]. üt [975]. Utdrag [747]. Utérus [4196]. utiles [2379, 2502, 1541]. utilisation [1884]. utilisées [175]. utilité [2328]. Utopia [1771]. utopiste [2245]. Uvarov [652]. Uyûn [246]. Uzluk [961].

V [1055, 4433, 2763, 12, 31, 4615, 2754, 3727, 1509, 2039, 1590, 327, 2662, 3348, 4076, 4435, 3060, 1610, 3749, 4090, 1614, 3148, 3240, 1015, 2096, 3725, 601, 2693, 555, 3420, 2977, 3080, 1321, 3607, 1355, 3736, 4146, 1521, 3831, 2229, 3741, 2353, 3306, 3497, 4493, 1580, 160, 488, 566, 1332, 1691, 2493, 983, 2394, 423, 975, 3296, 3497, 2185]. V. [2763, 1851]. V.-David [1851]. V.D.3 [3626]. va [1477]. vaccination [2813]. vaccine [2450]. Vacher [1225, 2687]. Vaches [4450]. Vachon [2318]. Vademecum [2670]. Vagnetti [2629]. vaisseaux [2799, 2916]. Val [1001]. Val-de [1001]. valenc [3735, 3732]. valence [1577]. valencianos [3994, 2644]. Valentin [3971, 3242, 610, 2017, 2003]. Valentine [2870]. Valentini [4507]. Valeriani [2653]. Valéry [2253]. Valette [77]. Valeur [2995, 2056]. Validity [3125]. valle [2815]. Vallée [3760]. Valléry [34, 1242, 458, 1117]. Vallery-Radot [1242, 458, 1117]. Vallois [327]. Valore [4511]. Valson [1959]. Valten [747]. Vandel [212, 4638]. Vandeul [351]. Vanini [1990]. Vannereau [2962]. Vannoni [4362]. vapeur [1410, 1585, 2902, 3487]. Varcin [484]. Varet [822]. Varia [303, 1848, 3665]. variable [2760]. variation [3894]. Variations [1296, 4020]. Varie [943]. Variétés [3760]. Varignon [4227, 3718]. Variorum [4271, 4230, 4269, 4440, 4489]. various [456, 3482]. Varnoteaux [4629]. Varro [4122, 733]. Varsovie [1461]. Varsovie-Cracovie [1461]. Vasco [792, 2232, 1333, 449]. Vasconi [4098]. vasculaire [4521, 2845]. vasculaires [91]. Vascular [2172]. Vaucanson [1729]. Vaucouleurs [187]. Vaudoise [1824]. Vaux [19]. ve [1388]. vecteurs [1698]. Veen [373]. végétale [3853, 1783, 1992]. végétaux [91]. veil [89]. Veith [1637]. vel [2094]. Velde [575]. Veldkamp [3154]. Vellanskij [4361]. velléités [1781].

Velluz [1575]. Velocidad [136]. Velpeau [2065]. Vendrell [1436]. veneto [4161]. Venezia [4161]. Venice [4662]. Venise [4662]. Ver [151, 481, 985, 988, 930, 1330, 219, 1292]. Vera [154]. verbale [3750]. verbe [4565]. verdad [3235]. Verdet [2770, 4643, 2231, 3612]. verfasst [3314]. Vergata [3153]. Vergleichenden [4256]. vergleichender [3416]. Vergnaud [581]. Verhandelingen [397]. Verificazione [4475]. véritable [3711, 4196]. vérité [4416, 775, 2502]. Verlag [3505, 278]. Vermeer [2473]. Vermessung [1666]. Verna [4415]. Vernadsky [3972]. Vernant [1434]. Vernet [904, 1521]. Verneuil [2270]. Vernière [352]. Veröfentlichungen [1689]. Veröffentlichungen [2392]. vérole [4669]. Véronique [3704, 4395]. verovering [325]. Verrier [213]. Verriest [375]. Vers [1781, 3793, 3776, 2365, 4279, 2284, 2725, 4278, 4495]. versaillais [1480]. Versailles [2764]. Versamml [3843]. Versammlungen [3843]. Verschoor [1590]. Versicherungs [4204]. Versicherungs- [4204]. Version [3420, 1472, 1220, 1436, 4198]. Verso [4083]. versus [4406]. verte [3621]. vertes [2175]. Vésale [10, 881]. Vesalian [2321]. Vesalio [3627]. Vesaliu [1020]. Vestiges [561]. Vetenskapsakademiens [12]. vétérinaire [4524, 3460, 2340, 1223, 838]. Veterinary [1244, 3209]. Vetus [2921]. VI [349, 1931, 2536, 2275, 4202, 1004, 1509, 2116, 3937, 3240, 2420, 792, 1072, 1386, 3725, 591, 3607, 1735, 3736, 1846, 2545, 566, 839, 969, 1014, 538, 257, 3941, 742, 582]. VI. [407]. viajes [3735]. vibrantes [3355]. Vibration [2831, 4359, 2304]. Vibrations [4297]. Vicary [727]. vice [4199]. Victor [3519, 2575, 2926, 1023, 853, 2722, 1049, 2777, 3994]. Victor-Robinson [853]. Victorian [3003, 3243, 2523, 3940]. Vid [3411, 748]. Vidal [2243, 4369]. vide [3425, 4071]. Vidius [727, 2412]. Vie [73, 1524, 2436, 1167, 671, 271, 2409, 158, 64, 269, 287, 407, 432, 4381, 949, 401, 4679, 301, 698, 3145, 4675, 4644, 1831, 2000, 2027, 3057, 1947, 1696, 2412, 1169, 796, 881, 964, 1059, 1959, 3112, 1337, 585, 1513, 3412, 4346, 1658, 1478, 1687, 2026, 1852, 1105, 1179, 1357, 1243, 3968, 2446, 714, 476, 382, 408, 4445, 3863, 3153, 3639, 1575]. Vieille [3509]. vieillissement [1696]. Vienna [4056, 4056, 4144]. Vierge [4450]. Vierteljahrsschrift [1617]. Viet [1311, 279]. Viet-Nam [1311, 279]. Vieta [1863]. Viète [1956]. Vietnam [1309]. vietnamienne [570]. vietnamiennes [1061]. Vieusseux [2941]. vieux [159]. View [3411, 3701, 3916, 1686]. viewed [4315]. Viggo [2518]. vigilia [4309]. Vigneron [1892]. VII [2275, 4243, 1004, 4076, 3314, 3240, 1232, 3725, 591, 3420, 4146, 1018, 1262, 1056, 2476, 566, 2875, 2492, 369, 432, 499, 2061, 3758]. VII/1 [4243]. VII/2 [4243]. VIII [2275, 617, 1486, 1912, 4415, 3240, 2044, 2354, 1467, 2538, 4007, 2334, 1233, 1324, 2881, 1241, 2742, 591, 625]. Vijf [513]. Vijftig [1482]. Viktor [3570]. Villar [1600]. Villard [2623, 3688]. Villards [3197]. Villards-d’Héria [3197]. Villars [1294]. Villeneuve [213, 889]. Villon [4058]. Vincennes [3933, 3932]. Vincent [4477, 4508, 3815, 4123, 3811, 3786, 4255, 4091, 2188, 4071, 3559, 216, 4503, 1986, 1153]. Vincent-Pierre [2188]. Vincenzio [3562]. Vincenzo [989, 901]. Vinci [934, 601, 417, 527, 528, 1146, 1165, 1747, 1601, 4633, 750, 526]. Vindé [1225].

Vindicatio [2963]. vingt [2293]. Vinh [1311]. Vintéjoux [284]. Vinton [2547, 2310]. Viola [3626]. violence [4201, 4565]. Virchow [2069, 3739]. Viret [327]. virgiliana [4088]. Virieux [651]. Virieux-Reymond [651]. Virtanen [1016, 2253]. virtual [4603]. virtuel [4603]. Virtuosi [2697]. Virus [3472, 3473]. Viruses [3593]. Vis [2462]. Visages [4201]. visible [4051, 4124]. Vision [2684, 2753, 3140, 3763, 3590, 4549, 4335, 2021, 3737]. Visionary [3886]. visionnaire [3942]. visited [456]. Visser [3015]. Vistas [1827]. visual [4209]. visuel [4328]. Vita [3731, 3062, 901]. Vital [3123]. vitalisme [177]. Vitalist [1110]. Vitaly [3610]. vitamin [3211]. vitesse [4075, 2504, 4504, 3803, 3351]. Vito [985]. Vitrac [3499, 3725, 3313]. Vitruve [517]. Viva [2979]. vivant [1102, 3834, 3925, 3963, 4451]. vivante [1787, 1473, 1732]. vivantes [279]. vivants [4450, 919, 469]. vives [4052]. Viviane [3151]. vlaamse [991]. Vladimir [3972, 2184]. Vlastos [2803]. voûtes [1581]. vocation [670, 811]. Vocations [670]. Vogel [2985, 1777, 1795]. Vogt [3970]. voiture [2951]. Vol [1539, 2490, 3236, 4078, 1590, 3000, 3362, 1614, 567, 1671, 3197, 3080, 3195, 3082, 1299, 608, 2712, 3563, 1594, 4409, 4453, 162, 163, 164, 3631, 3519, 3767, 3270, 3663, 743, 3866, 4390, 3903, 2873, 2337, 2834, 3533, 2812, 3880, 3700, 3826, 1522, 10, 2826, 2190, 2191, 2192, 2189, 2193, 2275, 2276, 2274, 2296, 2297, 2343, 2345, 2061, 4615, 485, 4077, 3178, 3411, 3703, 3727, 4344, 593, 632, 633, 696, 846, 866, 888, 972, 973, 974, 1004, 1034, 1035, 1120, 1122, 1123, 1151, 1152, 1178, 1316, 1317, 1318, 1445, 1484, 1485, 1486, 1507]. vol [1509, 1510, 1710, 1797, 1829, 1837, 1912, 1946, 1913, 1944, 1914, 2041, 2039, 2040, 2042, 2043, 2115, 2117, 2131, 2116, 2132, 2133, 3992, 3693, 3792, 218, 1259, 328, 2423, 3990, 2662, 3963, 4286, 3937, 4076, 4435, 4098, 3060, 3455, 3886, 3650, 4020, 3417, 3482, 3584, 3589, 3637, 3640, 1162, 1526, 1662, 2321, 2670, 3240, 3363, 2522, 2233, 2641, 2420, 690, 861, 2477, 727, 1170, 1565, 2787, 628, 987, 1015, 1670, 1672, 1929, 2057, 2096, 3039, 1263, 3218, 3525, 3725, 947, 4207, 4206, 591, 626, 2921, 1617, 2140, 1578, 3582, 886, 3860, 3002, 4643, 3500]. vol [3699, 1472, 4066, 3607, 4290, 1421, 4647, 3921, 3964, 3705, 3577, 1827, 3812, 3996, 3307, 3177, 3871, 1605, 2444, 510, 3387, 1998, 2171, 2185, 2264, 2476, 4036, 4493, 4318, 880, 3726, 4051, 4124, 4023, 1430, 1431, 3078, 333, 96, 250, 278, 404, 529, 489, 830, 841, 856, 943, 1329, 1572, 2781, 221, 1108, 1238, 2965, 3830, 4558, 4198, 971, 3865, 610, 1246, 2492, 2493, 2742, 220, 1439, 2495, 2788, 2816, 2855]. Vol. [3362]. volantes [1063]. Volkov [3721]. Vollgraff [1517]. Vollständige [1097]. Vollzuges [4053]. volontaire [4573]. Volta [3941, 3982, 3983, 4457]. Voltaire [4317, 3287, 1289]. Volterra [2808]. Volume [3748, 2262, 1979, 2174, 2152, 1669, 248, 4090]. Volumen [490, 844, 845, 219, 331, 2873]. volumétrique [1882]. voluntary [4573]. vom [3536, 2706, 690]. voor [359, 514, 991, 841]. Voprossy [2453]. vor [3314, 2917]. Vorlesungen [4053, 2636]. vormen [514]. Vorstellungen [1689, 1680]. Vortices [3464]. Vortrag [2058]. Vorträge [2707, 1107]. vorzugsweise [278]. Vossiani [2249]. vous [3724]. Vovelle [2782]. Vowles [89]. Voyage [2925, 2256, 4538, 1707, 3778, 4091]. Voyages [1740, 32, 460, 4329]. voyageur [2020]. voyageurs [4362, 2598]. voyaging [3748]. Vrai [4634]. vrais [4567]. vraye [4084]. Vrin [2844, 275]. vu [1615]. Vue [344, 659, 4334, 4553, 923, 4134, 3348, 2101, 4407]. Vues [3787]. Vuillemin [1218]. vulgaires [4137]. Vulgarisation [4552, 3289, 3287, 3292, 3293]. vus [4137].

W [89, 2526, 2657, 4626, 13, 63, 1974, 2189, 2343, 1435, 2116, 4688, 1577, 1888, 3083, 3153, 418, 940, 2423, 302, 1866, 4436, 2109, 3455, 3105, 3482, 3640, 3744, 4082, 1694, 2578, 2847, 2522, 1169, 1611, 675, 727, 445, 2987, 3008, 1982, 3259, 3754, 3864, 2511, 3206, 3002, 3077, 1471, 3702, 859, 1142, 1412, 3094, 1759, 3603, 1981, 875, 2633, 2021, 2771, 4364, 589, 1301, 2747, 2712, 2788]. Waard [2527, 2658, 2964, 3124, 3278, 1667, 1955, 2107, 1226, 109, 1015, 1072, 1232, 1467, 1568, 2095, 626]. Waerden [653, 3042, 237, 397]. Waessenaer [1738]. Wagner [3766, 4551]. Wagner-Jauregg [3766]. während [278]. Wahrheit [4234, 3346]. Wai [3765]. Wakefield [4561]. wa’l [2372]. wa’l-handasa’l-mīkānīkiyya’l-’arabiyya [2372]. Walcott [4178]. Walden [586]. Waley [508]. Wali [3361]. Walker [874, 3706]. Wallace [4271, 922]. Wallis [2505, 1668, 3041, 2505]. Wallon [2514]. Walser [2296, 2193, 2034]. Walter [3440, 3062, 3589, 1688, 2603, 3035, 4204, 1111, 3203]. Walther [2660, 965, 3091]. Walusinski [2880]. Wandel [2847]. Wanderer [3156]. Wang [665, 1565, 1017]. Wangensteen [2550, 2550]. War [3076, 4574, 3632, 4365, 4604, 3632, 2143]. Warburg [874, 3196]. Ward [1752, 3424, 3869]. Wargentin [12]. Warington [1109]. Warner [4018]. warning [1722]. wars [4430]. Wartofsky [3002]. Warusfel [1215]. Warwick [4197]. Was [3659, 727, 4586]. Waschkies [2467]. Wasserstein [3125]. Water [2233, 4633, 1103]. Water-lifting [2233]. Waters [3125, 747, 747]. Watson [327]. Watt [89, 1410]. wave [2836]. wave-particle [2836]. waves [3696]. Wavre [483]. wax [4235]. Way [3083, 3431, 4686, 3673, 3808]. WC2N [2855]. Wear [3597]. Weart [2262]. Webb [4569]. Webber [235]. Weber [2702, 4083, 3693]. Webster [3996, 420, 2309, 3645]. Wechselbeziehungen [2478]. Wedgwood [2452]. Wednesday [2244]. Weg [2647, 2479]. Wege [2843, 690]. Weidman [4055]. Weierstraß [3667]. Weights [689]. Weil [3107]. Weill [4415]. Weill-Parot [4415]. Weiner [1853]. Weingart [2522]. Weingartner [3219]. Weinmann [2701]. Weiss [2905]. Weisskopf [2977]. Weisz [2672]. welche [3201]. Wellcome [2129, 3636, 1498, 1940]. Welling [3316]. Welt [2391]. Weltalls [1680]. Weltanschauung [1684]. Weltbildes [937]. Weltbund [723]. Weltgeschichte [1788]. Weltsysteme [1666]. Wen [3760]. were [3947]. wereld [513]. wereldbeeld [325, 396, 244]. Werk [3040, 2647, 2479, 3316, 1336]. Werke [1928, 3077]. werken [358, 326]. Werner [4168, 3489, 2478, 3517, 3214, 2977, 3077, 842, 3599, 2959]. Wertime [2757]. Wesley [3362, 2578, 3754, 3796, 2127]. Wessel [812]. Western [3519, 1888, 3864, 3860, 938, 3996, 3668]. Westfall [1975, 3598, 2699].

Westman [2632]. Wetenschap [237]. Wetenschappen [397, 991]. Weulersse [330]. Weyl [860]. Wheaton [2342, 2836]. Wheatstone [1731]. Wheeler [4172]. Wheels [2233]. which [1752, 2477]. Whilhelm [2607]. White [2166]. Whitehead [4541, 65]. Whitehouse [1025]. Whiteside [1670, 1929, 2096, 2742, 1669, 2057, 1572, 2492]. Whitmore [2166]. Whitrow [3766]. Whitteridge [1021]. Whitworth [89]. who [3673, 727, 3570]. Whyte [1235]. Wickersheimer [2754, 1495]. Widemer [651]. Wiederkehr [942, 1107]. Wien [2536, 1526, 2835, 2885]. Wiener [2536, 1526, 821, 2739]. Wightman [2116, 1435, 2111]. Wiksell [747]. Wilbur [2228, 3303]. Wilcke [13]. Wilczynski [878]. Wilfrid [1442]. Wilfried [3347, 2823]. Wilhelm [3223, 12, 2423, 3731, 2987, 4204, 533, 2468]. Wilkie [1025]. Wilkins [1973]. Wilkinson [3534]. Willem [3733, 4478, 3015]. William [1164, 3579, 1021, 2104, 4271, 906, 970, 1334, 2389, 2569, 3235, 2111, 3060, 3482, 1635, 2846, 3630, 1672, 3651, 2539, 3439, 859, 4291, 4083, 510, 946, 4209, 1942, 2390, 3415, 4268, 2224, 2855, 1774, 1688, 796, 1237, 512]. Williams [1477, 2171, 3481, 1999]. Willis [2397]. Willmoth [4290]. Wilma [3586]. Wilson [3155, 4578, 3243, 3815]. Wine [4364]. Winfried [2287]. Wings [784]. Winn [4196]. Winners [2702]. Winogradsky [1109]. Winsor [3389]. Winter [1332, 2355, 1336, 1438]. Wiriot [1939]. Wirken [278]. Wirkung [2479, 1684]. Wisconsin [1847, 1997, 2094, 2024, 3587]. wisdom [3210]. Wiskunde [1482, 237]. Wiss [278]. Wissen [4097, 4175]. Wissenchaften [2612]. Wissenschaft [1332, 1788, 4200]. Wissenschaften [3223, 2384, 2058, 506, 2603, 942, 2392, 394, 1323, 1332, 2355, 1028, 360]. Wissenschaftliche [3077, 1413, 2355, 2823]. wissenschaftlichen [1684]. Wissenschaftsforschung [2823]. Wissenschaftsgeschichte [2661, 2674, 2695, 2566, 1776]. wissenschaftsgeschichtliche [2392]. Wissenschaftskunde [842]. Wissenschaftslehre [4053]. wissenschaftsorganisatorische [2355]. Wissenschaftsund [2478]. Wit [3767]. witchcraft [609]. withdrawal [2937]. Withington [695]. without [4638]. Witkowski [3061]. Witnesses [4655]. Wittek [1350]. Wittgenstein [4582]. wittgensteinienne [4582]. Wittop [514]. Wlodzimierz [1864]. Woerden [3083]. Wohnlich [3244]. Wohnlich-Despaigne [3244]. Wojbor [3885]. Wojnar [2352]. Wolenski [4056]. Wolf [3536, 3276]. Wolff [3153, 2774, 3819, 3955, 4677]. Wolfgang [3697, 506, 2706, 4228]. Wolters [3245]. woman [4669]. Wong [952, 1166, 2187, 1464]. Woolf [973, 1123]. Woollam [799]. Word [2474]. Work [2630, 1025, 2248, 2539, 1998, 2185, 2659, 89, 3863, 4600, 3797, 508, 4658, 1664, 4613, 2155, 3041, 3015, 1144]. Works [1973, 1975, 1590, 235, 418, 940, 567, 3077, 2804, 856, 1299, 1572, 4659, 4667, 4672, 4681, 4690, 3903, 1754, 4516, 3553, 4689, 1101, 793, 3125, 1345, 745, 2978]. World [3632, 1665, 3526, 785, 2225, 3076, 3274, 3861, 3988, 191, 3898, 3701, 3282, 3589, 2078]. Worlds [508, 3744]. Wormald [3535]. worms [4236]. Woronoff [3515]. Wotschikowsky [2580]. Wou [1390, 1390]. Woyczyński [3885]. Wright [1504, 3262]. Wright-St [1504]. Writing [4095]. Writings [3011, 905, 3663, 1690, 3362, 3482, 3640, 2883, 3587, 3865, 460, 2959]. wrong [3341]. Wüest [3631]. Wunderlich [2647]. wurde [3314]. Wurtz [926, 4193, 2179, 2660]. Wussing [2384]. Wyckoff [1908]. Wylie [3061]. Wyngaert [535]. Wynne [3836].

X [931, 89, 76, 2631, 843, 2569, 4435, 3240, 1667, 2095, 1338, 2922, 1206, 1324, 1329, 2454, 705, 3901, 3523]. X-Ray [89, 2922]. Xavier [671, 3594, 2677]. XCI [1835]. XCIII [2045]. XCIV [1943]. XCVII [2295]. XI [2812, 3504, 1033, 1912, 328, 4435, 3462, 3240, 1955, 4493, 4023, 1324, 1691, 3013, 802, 1463, 1461, 3195, 1177, 1662]. XII [3516, 2754, 1033, 2040, 1727, 3867, 2044, 2107, 249, 2823, 3195, 3239, 1324, 910, 1462, 1641, 76, 197, 257, 3493, 1699]. XIII [1770, 1825, 1862, 2527, 2604, 3903, 2297, 1121, 3314, 3462, 2173, 2962, 3195, 1036, 4222, 4135, 1532, 4333, 3196]. XIIIe [3189]. XIV [2605, 2658, 1315, 4286, 2173, 2828, 4397, 1124, 174, 1819, 4333]. XIVe [4570]. XIVth [3011]. XIX [4512, 748, 2409, 3793, 4165, 4173, 3881, 4413, 983, 2218, 2577, 2251, 3993, 964, 1000, 1148, 1245, 1309, 1638, 2032, 2066, 2188, 2845, 3043, 2469, 4362, 3213, 4009, 3157, 1735, 4452, 3261, 3827, 1430, 1324, 3244, 4558, 1546, 3226, 1987, 4481, 1619, 2497, 1698, 2238, 3875, 2014, 2086, 3228, 3818, 4048, 177, 79, 4281, 3374, 3624, 1224, 1821, 3376, 1008, 4154, 2101, 4283, 4338, 4523, 1739, 1658, 3113, 2206, 3623, 815, 388]. XIXe [1756, 4620, 4572, 2599, 4573]. XL [4076, 4435, 3735, 362]. XLI [3735]. XLII [2358, 510, 3732]. XLIII [2266, 2267]. XLIV [3735, 4318]. XLV [3197]. XLVII [3538]. XLVIII [3614]. XV [3694, 1538, 2964, 2754, 1315, 3867, 3757, 2785, 2354, 1346, 1605, 1607, 839, 1192, 594, 656, 3039, 1108]. XVI [197, 279, 3124, 4003, 4081, 1207, 1444, 4434, 4644, 4369, 3755, 4491, 4289, 3994, 1480, 2032, 2915, 4009, 3437, 2525, 3215, 3345, 1014, 2677, 1272, 386, 2379, 890, 2909, 3161, 1379, 4403, 594, 3780, 3173, 4062, 867, 75]. XVIe [547]. XVII [1774, 3562, 1849, 3278, 2873, 2297, 2396, 1207, 1508, 280, 2423, 4491, 4005, 3273, 4205, 4441, 1305, 1579, 1772, 1753, 4362, 2962, 4319, 2501, 3345, 3453, 1930, 4070, 4051, 4124, 1851, 1324, 1541, 2249, 2246, 1364, 1855, 386, 2584, 4013, 658, 2238, 433, 291, 4337, 4403, 1762, 1655, 1990, 3295, 572, 568, 928, 1283, 2013, 1319, 566, 2087, 3437, 3726]. XVII-2 [1753]. XVIIe [1626, 1949, 224]. XVIII [4410, 2379, 14, 1508, 2512, 4411, 3809, 280, 2423, 1850, 4369, 4289, 4415, 3881, 2900, 983, 1662, 797, 838, 963, 1115, 1116, 1148, 1305, 1306, 1579, 1704, 2915, 2974, 1337, 4530, 3695, 1283, 3453, 1930, 1851, 839, 1324, 2394, 3013, 1615, 1447, 3572, 4524, 618, 3983, 4013, 4468, 925, 1282, 1740, 2909, 892, 193, 4028, 3929, 647, 757, 2238, 3290, 2365, 2486, 3875, 759, 570, 3779, 3428, 2858, 4281, 1513, 756, 4283, 3295, 1739, 3293, 4565, 573, 4539, 869, 621, 848, 1052, 1584, 75, 656, 58]. XVIIIe [4644, 4265, 969, 547]. XVIth [3011]. XX [4142, 4477, 1796, 4434, 397, 2770, 3993, 996, 1756, 2787, 3438, 3043, 4034, 3157, 4510, 4166, 1431, 529, 654, 4558, 1427, 1642, 3226, 3227, 25, 4482, 1456, 3951, 3918]. xxe [4664, 4573]. XXI [30, 4077, 2038, 4076, 2565, 1689, 2644, 1711, 3178, 3923]. XXII [2038, 303, 3011, 1799]. XXIII [2038, 3809, 4076, 1860, 1871, 4072]. XXIV [2189, 2343, 1574, 1962]. XXIX [4435, 2359, 4363]. XXVI [2343, 3419, 2079]. XXVII [4435, 4067]. XXVIII [485, 354, 2914, 3936, 2199]. XXX [2277]. XXXI [2260, 3813]. XXXII [1231, 2480]. XXXIII [2552]. XXXIV [1290, 2615]. XXXV [2713]. XXXVI [2789]. XXXVII [2131, 2887]. XXXVIII [3924].

Yûhannâ [2669]. Ya [3570]. Yahia [1605]. Yan [3206]. Yapp [3586]. Yaqubi [1705]. Yaron [3726]. Yasukatsu [4318]. Yates [186]. Yavetz [3797]. yazisi [1388]. Year [1389, 418, 940, 3755, 3758]. Yearbook [3726, 4077, 4056]. Years [1334, 3155, 3270, 3663, 4144, 3169, 3945, 2779, 3257, 3865, 2834, 768]. Years-2500 [1334]. Yellow [1637, 3235]. Yelnick [2611]. Yen [1227]. Yervant [2992]. yeux [4392]. yi [883]. yi-hiue-louen [883]. yim [534]. Yochelson [4178]. York [3864, 608, 3336]. York-Melbourne [3864]. Yoshio [1297]. Yougoslave [1535]. yougoslaves [1008]. Yougoslavie [980, 1069, 1627, 783, 1141, 1535]. Young [2753, 4590, 2477, 2586]. Youschkevitch [2354, 2637, 2469, 584, 2828]. youth [3564]. Yu [3175]. Yujian [2420]. Yukawa [3156, 3110]. Yun [606, 3175]. Yuri [4254]. Yushkevich [3839, 4020, 4203]. Yusuf [2372]. Yves [2827, 3504, 4003, 4367, 2318, 3057, 1306, 4147]. Yvette [2029, 2251, 3477]. Yvon [1219, 393, 3176].

Z [89, 3696, 2217, 4197, 878, 2252, 4257]. Zagorin [4025, 3484]. Zagreb [2854, 980]. Zahar [4179]. Zahlen [2124, 1864]. Zahlenlehre [2470]. zamanin [1388]. Zammit [4201]. Zandvoort [3220]. Zarko [2184]. zaruretler [1388]. Zaslawsky [2076]. Zaunick [1111]. Zawisza [2352]. Zdenek [1854]. Zecaire [4084]. Zeijlstra [950]. Zeit [2847, 3035, 1470]. Zeitgeschichte [2479]. Zeitschrift [1709, 482, 188]. Zeljko [1105]. Zellularpathologie [954]. Zeno [3434]. Zénobe [10]. Zénon [2868]. Zensetsou [1174, 2187]. Zermelo [2876]. Zerner [3433]. Zestiende [1293]. Zestiende-eeuwse [1293]. Zevenboom [514]. Zhi [3175]. zich [513]. Ziegler [4235, 4235]. Zielinska [4433]. Ziggelaar [2240]. Zijn [358, 326]. Zilberbogen [2614]. Zilberbogen-Chapdelaine [2614]. Ziman [3016, 4210]. Zimmermann [3276]. Zinguer [4026]. Zinner [1933, 769, 1470]. Zinsser [4620]. zodiaque [2485]. Zoologica [1119, 2551]. zoological [3015]. Zoologie [3572, 1241, 4484, 4637, 3460, 1026, 1830]. zoologique [4226, 2517]. zoologiques [3460]. zoologische [1639, 1868, 2037, 2265, 2513]. zoologiste [2575]. Zoology [4637, 2534, 3389]. Zosime [3962]. Zoubov [983, 601]. Zubov [1332]. zum [2031, 3416, 2479, 4177, 3443, 787, 1325]. Zur [2958, 1684, 278, 3815, 4097, 4175, 2478, 2536, 2467, 2661, 4093, 4053, 2674, 4234, 3314, 1526, 2917, 3214, 2479, 2695, 2706, 2707, 1689, 2685, 482, 1097, 1668, 1776, 3316, 942, 954, 2287, 3222, 4204, 3443, 2106, 1470, 1693, 1935, 3843, 3444, 4256, 787, 2829, 3346, 3536, 4053, 1260]. Zürich [3222]. zwei [1680]. Zweite [1298]. Zweiter [3223, 3314, 1216]. zwischen [3667, 1332, 2829]. zwölf [3757].

References

Berr:1947:ANR

[1] Henri Berr. Antécédents de la nouvelle Revue d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):5–8, juillet– septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904927.

Serrus:1947:MJB

[2] Ch. Serrus. La mécanique de J.-A. Borelli et la notion d’attraction. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):9–25, juillet– septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904929.

Collet:1947:MHC

[3] Claude-Georges Collet and Jean Itard. Un mathématicien humaniste, Claude-Gaspar Bachet de Méziriac (1581–1638). Revue d’Histoire des Sciences et de Leurs Applications, 1(1):26–50, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904931.

Metzger:1947:A

[4] Hélène Metzger. Atome. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):51–62, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904933.

Brunet:1947:LSH

[5] Pierre Brunet and Suzanne Delorme. L’activité de la Section d’Histoire des Sciences du Centre international de Synthèse au cours de l’année 1946. Revue d’Histoire des Sciences et de Leurs Applications, 1(1): 63–72, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904935.

Pelseneer:1947:LSM

[6] Jean Pelseneer and Q. Vetter. L’histoire des sciences dans le monde. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):72–77, juillet– septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904937.

Brunet:1947:AOBa

[7] Pierre Brunet. Analyses d’ouvrages: Aristote, Les météorologiques par J. Tricot. Revue d’Histoire des Sciences et de Leurs Applications, 1 (1):78–79, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904939.

Brunet:1947:AOBb

[8] Pierre Brunet. Analyses d’ouvrages: La révélation d’Hermès Trismégiste, I : L’astrologie et les sciences occultes par R. P. Festugière. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):79–81, juillet– septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904941.

Brunet:1947:AOBc

[9] Pierre Brunet. Analyses d’ouvrages: Frère Roger Bacon par Albert Garreau. Revue d’Histoire des Sciences et de Leurs Applications, 1(1): 81–82, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904943.

Brunet:1947:AOBd

[10] Pierre Brunet. Analyses d’ouvrages: André Vésale, 1944 par Georges Leboucq; Simon Stevin, 1942 par Robert Depau; Esquisse d’une histoire des mathématiques en Belgique, 1943 par Lucien Godeaux; Adolphe Quételet, pages choisies et commentées, 1942 par E. Dupréel; Zénobe Gramme, 1944 par Jean Pelseneer; Léon Frédéricq et les débuts de la physiologie en Belgique par Marcel Florkin; Les correspondants de Peiresc dans les anciens Pays-Bas, 1943, vol. in-12 par Raymond Lebègue. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):82–86, juillet– septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904945.

Brunet:1947:AOBe

[11] Pierre Brunet. Analyses d’ouvrages: Mersenne ou la naissance du mécanisme par Robert Lenoble. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):86–88, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904947.

Brunet:1947:AOBf

[12] Pierre Brunet. Analyses d’ouvrages: Pehr Wilhelm Wargentin, kungl. Vetenskapsakademiens sekreterare och astronom 1749–1783 par N. V. E.

Nordenmark. Revue d’Histoire des Sciences et de Leurs Applications, 1 (1):88–89, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904949.

Brunet:1947:AOBg

[13] Pierre Brunet. Analyses d’ouvrages: Johan Carl Wilcke, experimentalfysiker par C. W. Oseen. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):89, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904951.

Brunet:1947:AOBh

[14] Pierre Brunet. Analyses d’ouvrages: Rede en ervaring in de Natuurwetenschap der XVIII^{e} Eeuw par R. Hooykaas. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):89–90, juillet–septembre 1947. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904953.

Brunet:1947:AOBi

[15] Pierre Brunet. Analyses d’ouvrages: Essai sur quelques problèmes concernant le normal et le pathologique, ^{≪}La Montagne ^{≫}par Georges Canguilhem. Revue d’Histoire des Sciences et de Leurs Applications, 1(1): 90–91, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904955.

Brunet:1947:AOBj

[16] Pierre Brunet. Analyses d’ouvrages: Claude Bernard ou l’aventure scientifique Par Raymond Millet. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):91–92, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904957.

Brunet:1947:AOBk

[17] Pierre Brunet. Analyses d’ouvrages: Dupuytren par Henri Mondor; Pasteur par Henri Mondor. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):92–93, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904959.

Brunet:1947:AOBl

[18] Pierre Brunet. Analyses d’ouvrages: Histoire des Sciences par Pierre Humbert. Revue d’Histoire des Sciences et de Leurs Applications, 1 (1):93, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904961.

Brunet:1947:AOBm

[19] Pierre Brunet. Analyses d’ouvrages: Tableau des racines sémitiques (arabe-hébreu) accompagnées de comparaisons par Baron Carra de Vaux. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):94, juillet– septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904963.

Brunet:1947:AOBn

[20] Pierre Brunet. Analyses d’ouvrages: L’acte chimique. Essai sur l’histoire de la philosophie chimique par Maurice Daumas. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):94–95, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904965.

Brunet:1947:AOBo

[21] Pierre Brunet. Analyses d’ouvrages: Eiffel, le magicien du fer par François Poncetton. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):96, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904967.

Brunet:1947:AOBp

[22] Pierre Brunet. Analyses d’ouvrages: Introduction à l’étude d’une collection d’instruments anciens de mathématiques par Henri Michel. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):96, juillet– septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904969.

Reymond:1947:LPS

[23] Arnold Reymond. L’évolution de la pensée scientifique et l’histoire des sciences. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):97– 113, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903461.

Taton:1947:REG

[24] René Taton and Évariste Galois. Les relations d’Évariste Galois avec les mathématiciens de son temps. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):114–130, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903462.

Cahen:1947:TEO

[25] Louis Cahen. La télégraphie électrique des origines au début du XX^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):131– 161, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903463.

Sergescu:1947:BGM

[26] P. Sergescu. Le bicentenaire de Gaspard Monge. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):162–170, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903464.

Brunet:1947:LSE

[27] Pierre Brunet. L’histoire des sciences dans l’enseignement. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):171–174, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903465.

Reymond:1947:LSM

[28] Arnold Reymond and Pierre Brunet. L’histoire des sciences dans le monde. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):174– 177, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903466.

Brunet:1947:AOBq

[29] Pierre Brunet. Analyses d’ouvrages: Gerbert par Jean Leflon. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):178–179, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903467.

Brunet:1947:AOBr

[30] Pierre Brunet. Analyses d’ouvrages: Œuvres complètes, t. XXI par Christiaan Huygens. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):179–180, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903468.

Brunet:1947:AOBs

[31] Pierre Brunet. Analyses d’ouvrages: Fredrik Mallet och Daniel Melanderhjelm, två Uppsala-astronomer par N. V. E. Nordenmark. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):180–181, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903469.

Brunet:1947:AOBt

[32] Pierre Brunet. Analyses d’ouvrages: Esquisse de mes voyages au Brésil et Paraguay par Auguste de Saint-Hilaire; Anna E. Jenkins. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):181–182, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903470.

Brunet:1947:AOBu

[33] Pierre Brunet. Analyses d’ouvrages: Bernard creador de la medicina cientifica par Jose Joaquin Izquierdo. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):182–183, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903471.

Brunet:1947:AOBv

[34] Pierre Brunet. Analyses d’ouvrages: Pasteur, Correspondance. I : Lettres de jeunesse. L’étape de la cristallographie (1840–1857) par Pasteur-Valléry–Radot. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):183–184, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903472.

Brunet:1947:AOBw

[35] Pierre Brunet. Analyses d’ouvrages: Charles Nicolle par Fernand Lot; J. Rostand; Henri Becquerel et la découverte de la radioactivité par Albert Ranc. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):184– 185, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903473.

Brunet:1947:AOBx

[36] Pierre Brunet. Analyses d’ouvrages: The Svedberg ; Andreas Speiser (Festschrift). Revue d’Histoire des Sciences et de Leurs Applications, 1(2):185, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903474.

Brunet:1947:AOBy

[37] Pierre Brunet. Analyses d’ouvrages: Science since 1500. A short History of mathematics, physics, chemistry, biology par H. T. Pledge. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):185–186, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903475.

Brunet:1947:AOBz

[38] Pierre Brunet. Analyses d’ouvrages: Les conquérants de la science, II, d’Oberkampf à Pasteur. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):186–187, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903476.

Brunet:1947:AOBaa

[39] Pierre Brunet. Analyses d’ouvrages: Les aventuriers de la science par Pierre Devaux. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):187–188, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903477.

Brunet:1947:AOBab

[40] Pierre Brunet. Analyses d’ouvrages: Grandes inventions par Andrée Marti; Grands savants de France par P. Marchal. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):188, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903478.

Brunet:1947:AOBac

[41] Pierre Brunet. Analyses d’ouvrages: Histoire du calcul par René Taton. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):188–189, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903479.

Itard:1947:AOB

[42] Jean Itard. Analyses d’ouvrages: Algèbre et géométrie du second degré par Émile Borel. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):189, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903480.

Brunet:1947:AOBad

[43] Pierre Brunet. Analyses d’ouvrages: Le calendrier par Paul Couderc. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):190, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903481.

Brunet:1947:AOBae

[44] Pierre Brunet. Analyses d’ouvrages: Critique et Géologie. Contribution à l’histoire des sciences de la terre, t. I, t. II, t. III par Emm. de Margerie. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):190–191, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903482.

Brunet:1947:AOBaf

[45] Pierre Brunet. Analyses d’ouvrages: L’invention du cinéma (1832–1897) par Georges Sadoul. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):191–192, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903483.

Brunet:1947:AOBag

[46] Pierre Brunet. Analyses d’ouvrages: Le Charbon. Son histoire, son destin par Eugène Schneider. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):192, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903484.

Anonymous:1947:BMa

[47] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):??, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904971.

Anonymous:1947:BMb

[48] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):??, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903485.

Anonymous:1947:FMa

[49] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 1(1):??, juillet–septembre 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904925.

Anonymous:1947:FMb

[50] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 1(2):??, October 1947. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903460.

Costabel:1948:CHL

[51] Pierre Costabel. Contribution à l’histoire de la loi de chute des graves. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):193–205, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903876.

deVaux:1948:SAD

[52] Carra de Vaux. Une solution arabe du problème des carrés magiques. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):206–212, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903877.

Brunet:1948:ODP

[53] Pierre Brunet. Les origines du paratonnerre (discussions et réalisations). Revue d’Histoire des Sciences et de Leurs Applications, 1(3):213–253, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903878.

Itard:1948:PDT

[54] Jean Itard. À propos du tricentenaire de la naissance de Newton. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):254–257, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903879.

Leconte:1948:LMC

[55] Th. Leconte. L’histoire des mathématiques dans la correspondance de Henri Lebesgue. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):257–265, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903880.

Cahen:1948:OII

[56] Louis Cahen. Sur les origines de l’idée de ionosphère. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):266, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903881.

Brunet:1948:LSM

[57] Pierre Brunet. L’histoire des sciences dans le monde. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):266–270, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903882.

Brunet:1948:AOBa

[58] Pierre Brunet. Analyses d’ouvrages: Traités I à XVIII par Hermès Trismégiste; A. D. Nock; A.-J. Festugière. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):271–272, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903883.

Brunet:1948:AOBb

[59] Pierre Brunet. Analyses d’ouvrages: Histoire du texte d’Élien le tacticien par Alphonse Dain. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):272, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903884.

Brunet:1948:AOBc

[60] Pierre Brunet. Analyses d’ouvrages: La renaissance de la gymnastique médicale. Mercurialis par R. Suaudeau; C. Suaudeau–Déterne. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):272–273, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903885.

Itard:1948:AOBb

[61] Jean Itard. Analyses d’ouvrages: L’œuvre scientifique de Blaise Pascal par Pierre Humbert. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):273–274, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903886.

Itard:1948:AOBc

[62] Jean Itard. Analyses d’ouvrages: Algunos aspectos de la fisica moderna, Trascendencia de la obra de Galileo y Newton. Galileo Galilei, Buenos-Aires, Espasa-Calpe Argentina, 1^{re} édit., 1942; 2^{e} édit., 1946 par Cortès Pla. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):274– 275, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903887.

Brunet:1948:AOBd

[63] Pierre Brunet. Analyses d’ouvrages: Cl. Bernard et la physiologie, Pasteur et la bactériologie, D’Arsonval et la biophysique par W. Kopaczewski. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):275–278, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903888.

Brunet:1948:AOBe

[64] Pierre Brunet. Analyses d’ouvrages: Robert Koch, sa vie et son œuvre par E. Lagrange. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):278–279, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903889.

Brunet:1948:AOBf

[65] Pierre Brunet. Analyses d’ouvrages: Essays in science and philosophy par Alfred North Whitehead. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):279–280, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903890.

Brunet:1948:AOBg

[66] Pierre Brunet. Analyses d’ouvrages: Figures de savants par Léon Binet. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):280, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903891.

Brunet:1948:AOBh

[67] Pierre Brunet. Analyses d’ouvrages: La France, pays de grandes découvertes par Léon Guillet. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):280–281, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903892.

Itard:1948:AOBd

[68] Jean Itard. Analyses d’ouvrages: La mathématique et son unité par G. Bouligand; J. Desbats. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):281–282, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903893.

Costabel:1948:AOB

[69] Pierre Costabel. Analyses d’ouvrages: Les principes du calcul infinitésimal par René Guénon. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):282–283, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903894.

Brunet:1948:AOBi

[70] Pierre Brunet. Analyses d’ouvrages: L’architecture de l’univers par Paul Couderc. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):283–284, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903895.

Brunet:1948:AOBj

[71] Pierre Brunet. Analyses d’ouvrages: Traité de l’astrolabe par Henri Michel. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):284– 285, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903896.

Brunet:1948:AOBk

[72] Pierre Brunet. Analyses d’ouvrages: Physique et microphysique, Collection Sciences d’aujourd’hui par Louis de Broglie. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):286–287, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903897.

Delorme:1948:AOB

[73] Suzanne Delorme. Analyses d’ouvrages: Les Plantes et la Vie, Collection ^{≪}Sciences et Médecine ^{≫}, série des Sciences naturelles par Florian Cosandey; Robert Mathey. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):287–288, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903898.

D:1948:AOB

[74] S. D. Analyses d’ouvrages: La Kabbale par Henri Sérouya. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):288, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903899.

deDainville:1948:FCS

[75] François de Dainville. Foyers de culture scientifique dans la France méditerranéenne du XVI^{e} au XVIII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):289–300, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904136.

Beaujouan:1948:EPF

[76] Guy Beaujouan. Étude paléographique sur la ^{≪}rotation ^{≫}des chiffres et l’emploi des apices du x^{e} au XII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):301–313, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904137.

Humbert:1948:JGV

[77] Pierre Humbert. Joseph Gaultier de la Valette astronome provençal (1564–1647). Revue d’Histoire des Sciences et de Leurs Applications, 1(4):314–322, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904138.

Auger:1948:AJS

[78] Léon Auger. Les apports de J. Sauveur (1653–1716) à la création de l’acoustique. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):323–336, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904139.

Jovignot:1948:GSB

[79] Louis Jovignot. Un grand savant bourguignon du XIX^{e} siècle : Antoine Masson. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):337–350, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904140.

Brunet:1948:SS

[80] Pierre Brunet. Souvenirs des savants. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):351–354, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904141.

Itard:1948:PEM

[81] Jean Itard. Sur une prétendue erreur mathématique de Galilée. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):355–356, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904142.

Taton:1948:ALS

[82] R. Taton. Abel et l’Académie des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):356–358, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904143.

Taton:1948:LIM

[83] R. Taton. Une lettre inédite de Monge sur la situation en France, en 1791, après la fuite du roi. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):358–359, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904144.

Moreau:1948:NHD

[84] Joseph Moreau. Nécrologie : Henri Daudin. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):360–361, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904145.

Anonymous:1948:LSM

[85] Anonyme. L’histoire des sciences dans le monde: France. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):362–366, avril– juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904146.

Michel:1948:AOB

[86] Paul-Henri Michel. Analyses d’ouvrages: Le commentaire de Marinus aux Data d’Euclide. Étude critique par Maurice Michaux. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):367–368, avril– juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904147.

Itard:1948:AOBe

[87] Jean Itard. Analyses d’ouvrages: Isaac Newton par Cortès Pla. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):368–369, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904148.

Itard:1948:AOBf

[88] Jean Itard. Analyses d’ouvrages: Henri Becquerel, El descubrimiento de la Radioactividad par Cortès Pla; Huygens-Fresnel, La teoria ondulatoria de la luz. Revue d’Histoire des Sciences et de Leurs Applications, 1(4): 369, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904149.

Auger:1948:AOB

[89] Léon Auger. Analyses d’ouvrages: The Royal Observatory, Greenwich, Oliver Heaviside and the mathematical theory of electrical communications, James Watt and the industrial revolution, Osborne Reynolds and his work in hydraulics and hydrodynamics, The steam turbine and other inventions of Sir Charles Parsons, Sir Joseph Whitworth : A pioneer of mechanical engineering, Frédérick Henry Royce, an outline of his engineering achievement, Sir Alfred Ewing, a pioneer in physics and engineering, Colonel Crompton and the evolution of the electrical industry, Sir Joseph Swan and the invention of the incandescent electric lamp, S. Z. de Ferranti, The history of X-Ray analysis, A century of British chemistry, The science of animal breeding in Britain : a short history, Patents of invention origin and growth of the patent system in Britain; Science lifts the veil par Harold Spencer Jones; George Lee; H. W. Dickinson; H. P. Vowles; A. H. Gibson; R. H. Parsons; F. C. Lea; G. Geoffroy Smith; L. F. Bates; J. H. Johnson; W. L. Randell; Kenneth R. Swan; Laurence Bragg; F. Sherwood Taylor; F. H. A. Marshall; John Hammond; A. A. Gomme. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):369–372, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904150.

Brunet:1948:AOBl

[90] Pierre Brunet. Analyses d’ouvrages: L’évolution physico-chimique, 3^{e} édit. par Ch.-Eug. Guye. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):372, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904151.

Brunet:1948:AOBm

[91] Pierre Brunet. Analyses d’ouvrages: Les végétaux vasculaires. Introduction à l’étude de l’anatomie comparée par Paul Bertrand. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):372–373, avril– juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904152.

Itard:1948:AOBg

[92] Jean Itard. Analyses d’ouvrages: Une bibliothèque médicale de la Renaissance conservée à Pékin par H. Bernard. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):373, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904153.

Brunet:1948:AOBn

[93] Pierre Brunet. Analyses d’ouvrages: Le sommeil et les états de sommeil, (Bibliothéque de philosophie scientifique) par Paul Chauchard. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):373–374, avril– juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904154.

Brunet:1948:AOBo

[94] Pierre Brunet. Analyses d’ouvrages: Les explorateurs célèbres. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):374, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904155.

Brunet:1948:AOBp

[95] Pierre Brunet. Analyses d’ouvrages: L’évolution scientifique et l’agriculture française, (Bibliothéque de philosophie scientifique) par Albert Demolon. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):374–375, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904156.

Taton:1948:AQP

[96] R. Taton. Analyses de quelques périodiques: Isis, An International Review devoted to the History of Science and Civilization, vol. 38, 1947– 1948, parts 1 and 2. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):376, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904157.

Anonymous:1948:BMb

[97] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):??, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903900.

Anonymous:1948:BMc

[98] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):??, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904158.

Anonymous:1948:FMb

[99] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):??, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903875.

Anonymous:1948:FMc

[100] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 1(4):??, avril–juin 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904135.

Sergescu:1948:MAS

[101] Pierre Sergescu. Mersenne l’animateur (8 septembre 1588–1^{er} septembre 1648). Revue d’Histoire des Sciences et de Leurs Applications, 2 (1):5–12, juillet–décembre 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904492.

deWaard:1948:RCM

[102] C. de Waard. A la recherche de la correspondance de Mersenne. Revue d’Histoire des Sciences et de Leurs Applications, 2(1):13–28, juillet– décembre 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904493.

Humbert:1948:MAS

[103] Pierre Humbert. Mersenne et les astronomes de son temps. Revue d’Histoire des Sciences et de Leurs Applications, 2(1):29–32, juillet– décembre 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904494.

Auger:1948:RPM

[104] Léon Auger. Le R. P. Mersenne et la physique. Revue d’Histoire des Sciences et de Leurs Applications, 2(1):33–52, juillet–décembre 1948. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904495.

Lenoble:1948:QAR

[105] Robert Lenoble. Quelques aspects d’une révolution scientifique: A propos du troisième centenaire du P. Mersenne (1588–1648). Revue d’Histoire des Sciences et de Leurs Applications, 2(1):53–79, juillet–décembre 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904496.

Rochot:1948:DTA

[106] Bernard Rochot. Une discussion théorique au temps de Mersenne : Le problème de Poysson (1635–1636). Revue d’Histoire des Sciences et de Leurs Applications, 2(1):80–89, juillet–décembre 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904497.

Hochart:1948:MMT

[107] Hochart and C. de Waard. A la mémoire de m^{me} tannery. Revue d’Histoire des Sciences et de Leurs Applications, 2(1):90–94, juillet– décembre 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904498.

Itard:1948:DAL

[108] Jean Itard. Sur la date à attribuer à une lettre de Pierre Fermat. Revue d’Histoire des Sciences et de Leurs Applications, 2(1):95–98, juillet– décembre 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904499.

Itard:1948:AOBa

[109] Jean Itard. Analyses d’ouvrages: Correspondance du P. Marin Mersenne, Religieux Minime, t. III, 1631–1633 par Paul Tannery; Cornelis de Waard. Revue d’Histoire des Sciences et de Leurs Applications, 2(1):99–100, juillet–décembre 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904500.

Anonymous:1948:BMa

[110] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 2(1):??, juillet–décembre 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904501.

Anonymous:1948:FMa

[111] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 2(1):??, juillet–décembre 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904491.

Sarton:1949:TAM

[112] George Sarton. La transmission au monde moderne de la science ancienne et médiévale. Revue d’Histoire des Sciences et de Leurs Applications, 2(2):101–138, January 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904571.

Michel:1949:M

[113] Paul-Henri Michel. Les médiétés. Revue d’Histoire des Sciences et de Leurs Applications, 2(2):139–178, January 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904572.

Daudin:1949:SSE

[114] Henri Daudin. Spinoza et la science expérimentale : sa discussion de l’expérience de Boyle. Revue d’Histoire des Sciences et de Leurs Applications, 2(2):179–190, January 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904573.

Bugler:1949:LSN

[115] G. Bugler. L’histoire des sciences naturelles dans l’enseignement. Revue d’Histoire des Sciences et de Leurs Applications, 2(2):191–192, January 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904574.

Anonymous:1949:LSMa

[116] Anonyme. L’histoire des sciences dans le monde: France. Revue d’Histoire des Sciences et de Leurs Applications, 2(2):192–193, January 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904575.

Itard:1949:AOBa

[117] Jean Itard. Analyses d’ouvrages: Esquisse d’une pédagogie de la redécouverte dans l’enseignement scientifique par Ch. Brunold. Revue d’Histoire des Sciences et de Leurs Applications, 2(2):194, January 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904576.

Taton:1949:AOBa

[118] R. Taton. Analyses d’ouvrages: A History of Geometrical Methods par J. L. Coolidge. Revue d’Histoire des Sciences et de Leurs Applications, 2(2):195–196, January 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904577.

Courtin:1949:AOBa

[119] M. Courtin. Analyses d’ouvrages: Physique et physiciens par Massain. Revue d’Histoire des Sciences et de Leurs Applications, 2(2):196, January 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904578.

Taton:1949:PFG

[120] R. Taton. La préhistoire de la ^{≪}géométrie moderne^{≫}. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):197–224, mai–août 1949. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904762.

Brunold:1949:TFG

[121] Ch. Brunold. Le tricentenaire de ^{≪}la grande expérience de l’équilibre des liqueurs^{≫}de Pascal. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):225–240, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904763.

Rostand:1949:EHA

[122] Jean Rostand. Esquisse d’une histoire de l’atomisme en biologie. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):241–265, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904764.

Costabel:1949:IM

[123] P. Costabel. Identification d’un manuscrit. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):266–267, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904765.

Robinet:1949:LRD

[124] René Robinet. L’École royale du Génie de Mézières (1748–1794): L’exposition du bicentenaire de sa fondation. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):267–270, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904766.

Bugler:1949:PPH

[125] G. Bugler. Précisions pour une histoire des origines de la paléontologie. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):270–274, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904767.

Anonymous:1949:F

[126] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):274–275, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904768.

Anonymous:1949:A

[127] Anonyme. Argentine. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):276, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904769.

Pelseneer:1949:B

[128] J. Pelseneer and Paul Ver Eecke. Belgique. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):276–277, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904770.

Anonymous:1949:R

[129] Anonyme. Roumanie. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):277, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904771.

Taton:1949:AOBb

[130] R. Taton. Analyses d’ouvrages: Morale de savants d’Hippocrate à Einstein par Jean Pelseneer. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):278–279, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904772.

Itard:1949:AOBb

[131] Jean Itard. Analyses d’ouvrages: Descartes par Henri Lefebvre. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):279–280, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904773.

Itard:1949:AOBc

[132] Jean Itard. Analyses d’ouvrages: Les Grands Courants de la pensée mathématique par F. Le Lionnais. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):280–281, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904774.

Taton:1949:AOBc

[133] R. Taton. Analyses d’ouvrages: A History of the Conic Sections and Quadric Surfaces par J. L. Coolidge. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):282, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904775.

Itard:1949:AOBd

[134] Jean Itard. Analyses d’ouvrages: Sidereus nuncius par Galilée; Maria Timpanaro Cardini. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):283, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904776.

Bouvier:1949:AOBa

[135] Robert Bouvier. Analyses d’ouvrages: L’Évolution de la notion de phénomène physique, des Primitifs à Bohr et Louis de Broglie. Leçons sur l’histoire de la pensée scientifique, professées à l’Université libre de Bruxelles par Jean Pelseneer. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):283–286, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904777.

Itard:1949:AOBe

[136] Jean Itard. Analyses d’ouvrages: Velocidad de la luz y relatividad par Cortès Pla. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):286–287, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904778.

Bugler:1949:AOBa

[137] G. Bugler. Analyses d’ouvrages: L’accélération évolutive (Essai sur le rythme évolutif et son interprétation quantique) par F. Meyer. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):287–288, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904779.

Daumas:1949:AOBa

[138] Maurice Daumas. Analyses d’ouvrages: La pensée de Marcellin Berthelot, Collection ^{≪}Pour connaître^{≫} par Albert Ranc. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):288, mai–août 1949. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904780.

Bugler:1949:AOBb

[139] G. Bugler. Analyses d’ouvrages: Histoire de la Médecine française par Pr Jules Guiart. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):288–289, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904781.

Daumas:1949:AOBb

[140] Maurice Daumas. Analyses d’ouvrages: Histoire des techniques, Collection ^{≪}Que sais-je?^{≫}par Pierre Ducassé. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):289–290, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904782.

Taton:1949:AQP

[141] R. Taton. Analyses de quelques périodiques: Archives Internationales d’Histoire des Sciences, nouvelle série d’Archeion, t. I. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):290–292, mai–août 1949. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904783.

Gauja:1949:LRS

[142] Pierre Gauja. L’Académie Royale des Sciences (1666–1793). Revue d’Histoire des Sciences et de Leurs Applications, 2(4):293–310, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905425.

Costabel:1949:DIC

[143] P. Costabel. Deux inédits de la correspondance indirecte Leibniz– Reyneau. Revue d’Histoire des Sciences et de Leurs Applications, 2 (4):311–332, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905426.

Rutten:1949:EGC

[144] Marguerite Rutten. Les emblèmes géométriques dans la civilisation ancienne du moyen-Orient. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):333–339, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905427.

Polvani:1949:LPE

[145] Giovanni Polvani. L’invention de la pile: Évocation à l’occasion de son cent cinquantième anniversaire. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):340–351, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905428.

Daumas:1949:HSC

[146] Maurice Daumas. Henri Sainte-Claire-Deville et les débuts de l’industrie de l’aluminium. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):352–357, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905429.

Anonymous:1949:LSMb

[147] Anonyme. L’histoire des sciences dans le monde: France. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):358–360, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905430.

Taton:1949:ECL

[148] R. Taton. Édition de la correspondance de Lavoisier. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):360, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905431.

Rochot:1949:NAB

[149] Bernard Rochot. Nécrologie: Augustin Boutaric (1885–1949). Revue d’Histoire des Sciences et de Leurs Applications, 2(4):361, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905432.

Itard:1949:AOBf

[150] Jean Itard. Analyses d’ouvrages: L’Apogée de la science technique grecque. L’essor de la mathématique par Abel Rey. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):362–363, September 1949. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905433.

Itard:1949:AOBg

[151] Jean Itard. Analyses d’ouvrages: Les Commentaires sur le premier livre des Éléments d’Euclide, traduits pour la première fois du grec en français, avec une introduction et des notes, Collection des travaux de l’Académie Internationale d’Histoire des Sciences, n^{o} 1 par Proclus de Lycie; Paul Ver Eecke. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):363– 364, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905434.

Itard:1949:AOBh

[152] Jean Itard. Analyses d’ouvrages: Leyendo a Euclides par Beppo Levi. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):364–365, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905435.

Taton:1949:AOBd

[153] R. Taton. Analyses d’ouvrages: Bibliographie des Multigrades avec quelques notices biographiques par A. Gloden; G. Palama. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):365, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905436.

Sergescu:1949:AOB

[154] P. Sergescu. Analyses d’ouvrages: Psicogénesis del razonamiento matématico (Éditorial Poseïdon, Buenos-Aires, 1947) Introduccion a la théoria de conjuntos (Édit. del Plata, Buenos-Aires, 1948) Breve Historia de la Geometria (Edit. Losada, Buenos-Aires, 1948) par Fr. Vera. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):365–366, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905437.

Prevost:1949:AOB

[155] René Prévost. Analyses d’ouvrages: Sciences expérimentales et probabilités par Reboul. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):366–367, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905438.

Courtin:1949:AOBb

[156] M. Courtin. Analyses d’ouvrages: The Growth of physical Science par James Jeans. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):367–368, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905439.

Cahen:1949:AOB

[157] L. Cahen. Analyses d’ouvrages: Origine et formation de la théorie des phénomènes électriques et magnétiques (thèse de doctorat ès lettres) par Jean Daujat. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):368–370, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905440.

Bouvier:1949:AOBb

[158] Robert Bouvier. Analyses d’ouvrages: La Genèse de la vie. Histoire des idées sur la génération spontanée par Jean Rostand. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):370–373, September 1949. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905441.

Bugler:1949:AOBc

[159] G. Bugler. Analyses d’ouvrages: Entre deux siècles (Souvenirs d’un vieux biologiste) par Henri Roger. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):373–374, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905442.

Taton:1949:AOBe

[160] R. Taton. Analyses d’ouvrages: Actes du V^{e} Congrès international d’Histoire des Sciences (Lausanne, 30 septembre-19 octobre 1947), Collection de travaux de l’Académie Internationale d’Histoire des Sciences, n^{o} 2. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):374, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905443.

Herve:1949:AOB

[161] Roger Hervé. Analyses d’ouvrages: Carles hollandaises. La cartographie de la Compagnie des Indes orientales, 1593–1743, La mappemonde de Pétrus Plancius gravée par Josua Van den Ende, 1604, d’après l’unique exemplaire de la Bibliothèque Nationale de Paris par Marcel Destombes. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):374–378, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905444.

Auger:1949:AQPa

[162] Léon Auger. Analyses de quelques périodiques: Isis, An international Review devoted to the History of Science and Civilization, vol. 38, parts 3 et 4 (févr. 1948); vol. 39, parts 1 et 2 (mai 1948); part 3 (août 1948); part 4 (nov. 1948); vol. 40, part 1 (févr. 1949); part 2 (mai 1949). Revue d’Histoire des Sciences et de Leurs Applications, 2(4):378–379, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905445.

Auger:1949:AQPb

[163] L. Auger. Analyses de quelques périodiques: Chymia: Annual Studies in the History of Chemistry, vol. 1. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):380–382, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905446.

Auger:1949:AQPc

[164] L. Auger. Analyses de quelques périodiques: The Journal of the Acoustical Society of America, vol. 20, n^{o} 6. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):382, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905447.

Auger:1949:AOB

[165] Léon Auger. Analyses d’ouvrages: Almanach des Sciences 1948 par Louis de Broglie; Almanach des Sciences 1949 par Louis de Broglie. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):382–383, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905448.

Anonymous:1949:TGM

[166] Anonyme. Table générale des matières du Tome II. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):384–388, September 1949. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905449.

Anonymous:1949:EDT

[167] Anonyme. Errata du Tome II to Michel and Taton. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):388, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905450.

Anonymous:1949:BMa (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904579.

Anonymous:1949:BMb

[169] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):??, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904784.

Anonymous:1949:BMc

[170] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):??, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905451.

Anonymous:1949:FMa

[171] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 2(2):??, January 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904570.

Anonymous:1949:FMb

[172] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):??, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904761.

Anonymous:1949:FMc

[173] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):??, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905424.

Beaujouan:1950:SAX

[174] Guy Beaujouan. La science au XIV_{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):5–20, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903812.

Itard:1950:MUP

[175] Jean Itard. Les méthodes utilisées par Fermat en théorie des nombres. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):21–26, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903813.

Peiresc:1950:PDP

[176] Peiresc and Pierre Humbert. À propos du passage de Mercure 1631. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):27–31, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903814.

Jacques:1950:VCO

[177] J. Jacques. Le vitalisme et la chimie organique pendant la première moitié du XIX_{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):32–66, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903815.

Brunet:1950:PLS

[178] Pierre Brunet. Les premiers linéaments de la science géologique : Agricola, Palissy, George Owen. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):67–79, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903816.

Costabel:1950:PRC

[179] Pierre Costabel. Sur un problème de Roberval et un cas particulier d’Analyse Diophantienne. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):80–86, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903817.

Taton:1950:F

[180] R. Taton. France. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):86–88, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903818.

Loiseau:1950:LCO

[181] J. Loiseau. L’exposition des chefs-d’œuvre de l’horlogerie au Musée du Conservatoire National des Arts et Métiers. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):88–89, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903819.

Bourdier:1950:LFB

[182] Franck Bourdier. L’exposition ^{≪}Buffon^{≫}au Muséum d’Histoire Naturelle de Paris. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):

89–91, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903820.

Anonymous:1950:A

[183] Anonyme. Allemagne. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):91, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903821.

Taton:1950:AOBa

[184] R. Taton. Analyses d’ouvrages: Bibliographie d’histoire des sciences (Bulletin de l’Union catholique des scientifiques français, n^{o} 7, mars– avril 1948) par F. Russo. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):92, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903822.

Rochot:1950:AOBa

[185] B. Rochot. Analyses d’ouvrages: Science et Humanisme par Émile Bréhier. Revue d’Histoire des Sciences et de Leurs Applications, 3(1): 92–94, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903823.

Delorme:1950:AOBa

[186] Albert Delorme. Analyses d’ouvrages: The French Academies of the Sixteenth Century par Frances A. Yates. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):94, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903824.

deBengy-Puyvallee:1950:AOBa

[187] R. de Bengy-Puyvallée. Analyses d’ouvrages: Astronomie, Les Astres, l’Univers par Lucien Rudaux; Gérard de Vaucouleurs; André Danjon. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):95, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903825.

Taton:1950:AOBb

[188] René Taton. Analyses d’ouvrages: Jakob Steiner. Leonhard Euler, Beihefte zu der Zeitschrift ^{≪}Elemente der Mathematik^{≫}par L. Kollros; Rudolf Fueter. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):95–96, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903826.

Taton:1950:AOBc

[189] René Taton. Analyses d’ouvrages: Origine e Sviluppo della Geometria Proiettiva par Federico Amodeo; Sintesi storico-critica della Geometria delle curve algebriche par Federico Amodeo. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):96–97, January 1950. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903827.

deBengy-Puyvallee:1950:AOBb

[190] R. de Bengy-Puyvallée. Analyses d’ouvrages: Science Advance par J. B. S. Haldane. Revue d’Histoire des Sciences et de Leurs Applications, 3(1): 97–98, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903828.

Delorme:1950:AOBb

[191] Suzanne Delorme. Analyses d’ouvrages: English Naturalists from Neckam to Ray, A Study of the making of the Modern World par Charles E. Raven. Revue d’Histoire des Sciences et de Leurs Applications, 3(1): 98–99, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903829.

Taton:1950:AOBd

[192] R. Taton. Analyses d’ouvrages: Archives Internationales d’Histoire des Sciences, publication trimestrielle de l’Union Internationale d’Histoire des Sciences, t. II, n^{o} 5 (oct.1948), n^{o} 6 (janv. 1949), n^{o} 7 (avr. 1949), n^{o} 8 (juill. 1949), n^{o} 9 (oct. 1949). Revue d’Histoire des Sciences et de Leurs Applications, 3(1):99–100, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903830.

Delorme:1950:FGC

[193] Suzanne Delorme. Une famille de grands Commis de l’État amis des Sciences, au XVIII^{e} siècle Les Trudaine. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):101–109, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904641.

Beaujouan:1950:DNC

[194] Guy Beaujouan. Documents nouveaux concernant Lagrange. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):110–132, avril– juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904642. Avec un ap-

pendice de René Taton.

Daumas:1950:PAS

[195] Maurice Daumas. Les polémiques au sujet des priorités de Lavoisier. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):133–155, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904643.

Rostand:1950:EHA

[196] Jean Rostand. Esquisse d’une histoire de l’atomisme en biologie: II. — la Période expérimentale de Mendel à Morgan (suite). Revue d’Histoire des Sciences et de Leurs Applications, 3(2):156–169, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904644.

Beaujouan:1950:SDC

[197] Guy Beaujouan. Les soi-disant chiffres grecs ou chaldéens (XII^{e}–XVI^{e} siècles). Revue d’Histoire des Sciences et de Leurs Applications, 3(2):170– 174, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904645.

Hare:1950:POM

[198] Hare and René Taton. À propos de l’œuvre de Monge en physique. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):174–179, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904646.

Anonymous:1950:TMD

[199] Anonyme. Le tricentenaire de la mort de Descartes (11 février 1650– 11 février 1950). Revue d’Histoire des Sciences et de Leurs Applications, 3(2):179–180, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904647.

Schuhl:1950:NST

[200] Pierre-Maxime Schuhl. Nécrologie: Sebastiano Timpanaro (1888–1949). Revue d’Histoire des Sciences et de Leurs Applications, 3(2):180–181, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904648.

Anonymous:1950:Fa

[201] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):181–183, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904649.

Anonymous:1950:UAI

[202] Anonyme. Union et Académie Internationales d’Histoire des Sciences: Congrès International d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):183, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904650.

Bernard-Maitre:1950:AOB

[203] Henri Bernard-Maître. Analyses d’ouvrages: La Science et l’Extrême-Orient (Cours et Conférences de l’École française d’Extrême-Orient, 1948–1949) par Pierre Huard. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):184, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904651.

Delorme:1950:AOBc

[204] Suzanne Delorme. Analyses d’ouvrages: Sciences par Ad. Ischer; S. Gagnebin. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):184– 185, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904652.

P:1950:AOBa

[205] R. B. P. Analyses d’ouvrages: Les Prix Nobel en 1946. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):185–186, avril–juin 1950. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904653.

P:1950:AOBb

[206] R. B. P. Analyses d’ouvrages: Les Conférences de l’U.N.E.S.C.O. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):186, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904654.

Taton:1950:AOBe

[207] René Taton. Analyses d’ouvrages: A concise History of Mathematics par Dirk J. Struik. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):186–187, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904655.

Taton:1950:AOBf

[208] René Taton. Analyses d’ouvrages: Les Recherches sur l’infini mathématique jusqu’à l’établissement de l’analyse infinitésimale (Actualités scientifiques et industrielles, 1083, Histoire des Mathématiques, I) par P. Sergescu. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):187–189, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904656.

Rochot:1950:AOBb

[209] B. Rochot. Analyses d’ouvrages: Histoire des découvertes astronomiques par Pierre Humbert. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):189–190, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904657.

Humbert:1950:AOBa

[210] P. Humbert. Analyses d’ouvrages: Galileo Galilei giovance il suo ^{≪}De Motu^{≫}. Quaderni di storia e critica della Scienza par Raffaele Giacomelli. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):190– 191, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904658.

Delorme:1950:AOBd

[211] Albert Delorme. Analyses d’ouvrages: Bergwerk- und Probierbüchlein par Ameliese G. Sisco; Cyril Stanley Smith. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):191–192, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904659.

Bugler:1950:AOB

[212] G. Bugler. Analyses d’ouvrages: L’homme et l’évolution ^{≪}L’avenir de la Science^{≫}par Albert Vandel. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):192–193, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904660.

Beaujouan:1950:AOB

[213] Guy Beaujouan. Analyses d’ouvrages: Études sur Arnaud de Villeneuve, 1240(?)– 1311 : III Le Breviarium practicae ou Arnaud de Villeneuve et l’Italie; IV L’étudiant de Naples par René Verrier. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):193–194, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904661.

Delorme:1950:AOBe

[214] Albert Delorme. Analyses d’ouvrages: Lavoisier, économiste par R. Dujarric de la Rivière. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):194–195, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904662.

P:1950:AOBc

[215] R. B. P. Analyses d’ouvrages: Quelques Français, hommes de science et d’action par M. Reynier; F. Broutet. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):195, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904663.

P:1950:AOBd

[216] R. B. P. Analyses d’ouvrages: Les secrets de la Science américaine (Scientists against time) par James Phinney Baxter; Vincent Corréa. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):195, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904664.

P:1950:AOBe

[217] R. B. P. Analyses d’ouvrages: Recueil annuel des travaux de l’Institut d’Histoire des Sciences et des Techniques de l’Université de Paris, t. 5 par Thalès. Revue d’Histoire des Sciences et de Leurs Applications, 3(2): 196, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904665.

Daumas:1950:AOB

[218] Maurice Daumas. Analyses d’ouvrages: Annual Studies in the History of Chemistry, vol. 2 par Chymia. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):196–199, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904666.

Taton:1950:AQPa

[219] R. Taton. Analyses de quelques périodiques: Osiris, Commentationes de scientiarum et eruditionis historia rationeque. Edidit Georgius Sarton Gandavensis. Volumen octavum Paulo Ver Eecke oblatum. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):199–200, avril– juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904667.

deBengy-Puyvallee:1950:AQP

[220] R. de Bengy-Puyvallée. Analyses de quelques périodiques: Isis, An international Review devoted to the History of Science and Civilisation, vol. 40, part. 3 (août 1949), part. 4 (novembre 1949). Revue d’Histoire des Sciences et de Leurs Applications, 3(2):200, avril–juin 1950. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904668.

Tetry:1950:AOBa

[221] A. Tétry. Analyses d’ouvrages: International Rules of Botanical Nomenclature. Brittonia, vol. 6, n^{o} 1. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):200, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904669.

Brehier:1950:PDL

[222] Émile Bréhier. À propos de deux livres récents sur la science grecque. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):201– 209, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903703.

Itard:1950:QRM

[223] Jean Itard. Quelques remarques sur les méthodes infinitésimales chez Euclide et Archimède. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):210–213, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903704.

Jones:1950:SRL

[224] H. W. Jones. La Société Royale de Londres au XVIIe siècle : réflexions diverses. Revue d’Histoire des Sciences et de Leurs Applications, 3(3): 214–221, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903705.

Vassails:1950:PDF

[225] Gérard Vassails. Le poids du feu. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):222–241, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903706.

Bugler:1950:PBE

[226] G. Bugler. Un précurseur de la biologie expérimentale : Edme Mariotte. Revue d’Histoire des Sciences et de Leurs Applications, 3(3): 242–250, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903707.

deWaard:1950:DDA

[227] C. de Waard. Sur le destinataire et la date à attribuer à une lettre de Descartes sur une épinette parfaite. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):251–255, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903708.

Auger:1950:CEM

[228] Léon Auger. La controverse entre Morin et Descartes sur la matière subtile. Revue d’Histoire des Sciences et de Leurs Applications, 3(3): 255–262, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903709.

Sergescu:1950:DM

[229] Pierre Sergescu. Descartes mathématicien. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):262–265, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903710.

Rostand:1950:DB

[230] Jean Rostand. Descartes et la biologie. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):265–269, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903711.

Anonymous:1950:BIC

[231] Anonyme. Belgique: III^{e} Congrés National des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):269–270, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903712.

Anonymous:1950:Fb

[232] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):270–273, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903713.

Dijksterhuis:1950:HHS

[233] E. J. Dijksterhuis. Hollande: L’histoire des sciences en Hollande de 1940 à 1950. Revue d’Histoire des Sciences et de Leurs Applications, 3(3): 273–276, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903714.

Rochot:1950:AOBc

[234] Bernard Rochot. Analyses d’ouvrages: La Découverte métaphysique de l’homme chez Descartes, La Nostalgie de l’être par F. Alquié; Le Problème de l’inconscient dans le cartésianisme L’Individualité selon Descartes par G. Lewis. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):277–281, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903715.

Delorme:1950:AOBf

[235] A. Delorme. Analyses d’ouvrages: A Descriptive Catalogue of the Grace K. Babson Collection of the Works of Sir Isaac Newton, and the Material relating to him in the Babson Institute Library, Babson Park, Mass par Roger Babson Webber. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):281–283, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903716.

Humbert:1950:AOBb

[236] P. Humbert. Analyses d’ouvrages: Célèbres problèmes mathématiques par Édouard Callandreau. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):283–284, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903717.

Dijksterhuis:1950:AOB

[237] E. J. Dijksterhuis. Analyses d’ouvrages: Ontwakende Wetenschap. Egyptische, Babylonische en Griekse Wiskunde (L’aube de la science. Mathématiques égyptienne, babylonienne et grecque) par B. L. Van der Waerden. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):

285–286, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903718.

Itard:1950:AOBa

[238] Jean Itard. Analyses d’ouvrages: Los Elementos de Euclides como exponente del ^{≪}milagro griego^{≫} par Emanuel S. Cabrera. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):286–287, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903719.

Itard:1950:AOBb

[239] Jean Itard. Analyses d’ouvrages: La philosophie de Pierre de Fermat par Armand Machabey. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):287–288, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903720.

Taton:1950:AOBg

[240] R. Taton. Analyses d’ouvrages: Jacob Steiner’s Geometrical Constructions with a Ruler given a fixed Circle with its Center …, (The Scripta Mathematica Studies n^{o} 4) par R.-C. Archibald; A. Stark. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):288, juillet– septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903721.

Auger:1950:AOB

[241] L. Auger. Analyses d’ouvrages: Le Déclin des absolus mathématicologiques par Georges Bouligand; Jean Desgranges. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):289, juillet–septembre 1950. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903722.

Itard:1950:AOBc

[242] Jean Itard. Analyses d’ouvrages: La Géométrie contemporaine, collection ≪Que sais-je?^{≫} par André Delachet. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):289–290, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903723.

T:1950:AOB

[243] R. T. Analyses d’ouvrages: Éléments de la théorie des probabilités, (Bibliothèque d’Éducation par la Science), par Émile Borel. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):290, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903724.

Vollgraff:1950:AOB

[244] J. A. Vollgraff. Analyses d’ouvrages: De mechanisering van het wereldbeeld par E. J. Dijksterhuis. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):290–291, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903725.

Itard:1950:AOBd

[245] Jean Itard. Analyses d’ouvrages: El enigma de la Luz par Cortes Pla; George Sarton. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):291–292, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903726.

Tetry:1950:AOBb

[246] A. Tétry. Analyses d’ouvrages: The Uyûn Al-Akhbâr of Ibn Qutayba. Collection de travaux de l’Académie Internationale des Sciences, n^{o} 4 par F. S. Bodenheimer; L. Kopf. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):293, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903727.

Tetry:1950:AOBc

[247] A. Tetry. Analyses d’ouvrages: Timotheus of Gaza on animals. Fragments of a Byzantine paraphrase of an animal-book of the 5th Century A.D. Collection des travaux de l’Académie internationale d’Histoire des Sciences, n^{o} 3 par F. S. Bodenheimer; A. Rabinowitz. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):293, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903728.

Wickersheimer:1950:AOB

[248] Ernest Wickersheimer. Analyses d’ouvrages: Storia della medicina (primo volume, dalle origini alla fine del Seicento; secondo volume, dal ’700 ai giorni nostri), nuova edizione ampliata e aggiornata par Arturo Castiglioni. Revue d’Histoire des Sciences et de Leurs Applications, 3(3): 294–297, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903729.

Mazaheri:1950:AOB

[249] Aly Mazaheri. Analyses d’ouvrages: Un traité d’Armurerie composé pour Saladin (extrait du Bulletin d’Études Orientales, t. XII, 1947–48) par Claude Cahen. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):297–300, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903730.

Taton:1950:AQPb

[250] R. Taton. Analyses de quelques périodiques: Isis, An international Review devoted to the History of Science and Civilization, vol. 41: part. 1 (mars 1950), part. 2 (juillet 1950). Revue d’Histoire des Sciences et de Leurs Applications, 3(3):300, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903731.

Bruins:1950:AMB

[251] E. M. Bruins. Aperçu sur les mathématiques babyloniennes. Revue d’Histoire des Sciences et de Leurs Applications, 3(4):301–314, October 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904205.

Costabel:1950:HDM

[252] Pierre Costabel. Histoire du moment d’inertie. Revue d’Histoire des Sciences et de Leurs Applications, 3(4):315–336, October 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904206.

Daumas:1950:GL

[253] Maurice Daumas. Gay-Lussac (1778–1850). Revue d’Histoire des Sciences et de Leurs Applications, 3(4):337–342, October 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904207.

Piveteau:1950:DEC

[254] Jean Piveteau. Le débat entre Cuvier et Geoffroy Saint-Hilaire sur l’unité de plan et de composition. Revue d’Histoire des Sciences et de Leurs Applications, 3(4):343–363, October 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904208.

Daumas:1950:QFI

[255] Maurice Daumas. Quelques fabricants d’instruments scientifiques anciens. Revue d’Histoire des Sciences et de Leurs Applications, 3(4):364– 370, October 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904209.

Bernard-Maitre:1950:PDR

[256] Henri Bernard-Maitre. Le Problème du robot scientifique. Revue d’Histoire des Sciences et de Leurs Applications, 3(4):370–375, October 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904210.

Delorme:1950:VCI

[257] Suzanne Delorme. VI^{e} Congrès international d’histoire des sciences et XII^{e} Congrès de la Société internationale d’histoire de la Médecine. Revue d’Histoire des Sciences et de Leurs Applications, 3(4):375–379, October 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904211.

Anonymous:1950:Fc

[258] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 3(4):379–381, October 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904212.

Anonymous:1950:TGM

[259] Anonyme. Table générale des matières du Tome III. Revue d’Histoire des Sciences et de Leurs Applications, 3(4):382–388, October 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904213.

Anonymous:1950:BMa

[260] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):??, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903831.

Anonymous:1950:BMb

[261] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):??, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904670.

Anonymous:1950:BMc

[262] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):??, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903732.

Anonymous:1950:BMd

[263] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 3(4):??, October 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904214.

Anonymous:1950:FMa

[264] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 3(1):??, January 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903811.

Anonymous:1950:FMb

[265] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):??, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904640.

Anonymous:1950:FMc

[266] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 3(3):??, juillet–septembre 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903702.

Anonymous:1950:FMd

[267] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 3(4):??, October 1950. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904204.

Berr:1951:MPB

[268] Henri Berr. In memoriam: Pierre Brunet. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):5–12, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904035.

Brunet:1951:VOCa

[269] Pierre Brunet. La vie et l’œuvre de Clairaut. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):13–40, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904036.

Rostand:1951:EHA

[270] Jean Rostand. Esquisse d’une histoire de l’atomisme en biologie: III. — du géne morganien aux conceptions actuelles (suite). Revue d’Histoire des Sciences et de Leurs Applications, 4(1):41–59, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904037.

Bodenheimer:1951:VOF

[271] F. S. Bodenheimer. La vie et l’œuvre de Frédéric Hasselquist (1722– 1752). Revue d’Histoire des Sciences et de Leurs Applications, 4(1): 60–77, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904038.

Belin-Milleron:1951:LPF

[272] J. Belin-Milleron. L’histoire des plantes, la formation de l’esprit scientifique et l’étude des civilisations. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):78–84, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904039.

Dulieu:1951:TCF

[273] L. Dulieu. Trois chaires de la Faculté de Médecine d’Aix mises au concours à Montpellier. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):85–90, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904040.

Rochot:1951:AOB

[274] Bernard Rochot. Analyses d’ouvrages: Le rationalisme appliqué par G. Bachelard. Revue d’Histoire des Sciences et de Leurs Applications, 4(1): 91–92, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904041.

Delorme:1951:AOBa

[275] Suzanne Delorme. Analyses d’ouvrages: Leçons de Philosophie des Sciences expérimentales, Bibliothèque de la Faculté de Philosophie et Lettres de Namur, fasc. 8 par Auguste Grégoire; J. Vrin. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):92–93, January 1951. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904042.

Taton:1951:AOBa

[276] R. Taton. Analyses d’ouvrages: The Origins of modern Science, 1300– 1800 par Herbert Butterfield. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):93–94, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904043.

Portal:1951:AOB

[277] R. Portal. Analyses d’ouvrages: Esquisse de l’histoire de la science et de la technique pendant la Révolution française, 1789–1794 par O. Staroselskaïa–Nikitina. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):94–95, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904044.

Taton:1951:AOJ

[278] R. Taton. Analyses d’ouvrages: Jos. E. Hofmann, Die Entwicklungsgeschichte der Leibnizschen Mathematik während des Aufenthaltes in Paris (1672–1676), Leibniz Verlag, München (Bisher R. Oldenbourg Verlag), 1949, 1 vol., ^{≪}Nicolaus Mercator (Kauffman): sein Leben und Wirken, vorzugsweise als Mathematiker ^{≫}(Akad. der Wiss. u. der Literatur, Abh. der math, naturwiss. Kl., n^{o} 3, Mainz, 1950, pp. 45–103); ≪Zur Entdeckungsgeschichte der höheren Analysis im 17. Jahrhundert ≫. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):95–97, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904045.

Bernard-Maitre:1951:AOB

[279] Henri Bernard-Maitre. Analyses d’ouvrages: Tome IX des Travaux de l’Institut anatomique de la Faculté de Médecine de l’Indochine (Section anthropologique); Contribution à l’histoire de l’anatomie, Thèse de la Faculté mixte de Médecine et de Pharmacie de l’Indochine, année 1947, no 149 par Pham-van-ghê; Les structures vivantes microscopiques. Évolution de la technique et des idées (extrait de l’Éducation, no 10 de mai 1948,) par P. Huard; La France et les progrès de la chirurgie, Conférence faite à l’Exposition médicale de Hanoï, décembre 1941 par P. Huard; Un traité des plantes médicinales exotiques du XVI^{e} siècle, conservé à Hanoï (extrait du Bulletin de la Société des Études indochinoises, 1^{er} trimestre 1948), par P. Huard; M. Destombes; Petite histoire du Moxa (Armoise) (ibid., 4^{e} trimestre 1948) par P. Huard; Histoire de la médicine du Viet-Nam (Thèse pour la Faculté mixte de Médecine et de Pharmacie de l’Indochine), par Duong-ba-banh. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):97–98, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904046.

Delorme:1951:AOBb

[280] S. Delorme. Analyses d’ouvrages: Les thèses médicales soutenues à Lyon aux XVII^{e} et XVIII^{e} siècles et le Collège Royal de Chirurgie de 1774 à 1792 par Jean Rousset. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):99, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904047.

Daumas:1951:AOBa

[281] Maurice Daumas. Analyses d’ouvrages: Mémoire sur l’histoire de la balance et de la balancerie par Armand Machabey Jeune. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):100–101, January 1951. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904048.

Daumas:1951:AOBb

[282] Maurice Daumas. Analyses d’ouvrages: Sons of Science. The Story of the Smithsonian Institution and its Leaders par Paul H. Oehser. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):101–102, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904049.

Taton:1951:AOBb

[283] René Taton. Analyses d’ouvrages: Sciences; Techniques, in 50 Années de découvertes, Bilan 1900–1950 par François Le Lionnais; Jacques Bergier. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):102–103, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904050.

Taton:1951:AOBc

[284] René Taton. Analyses d’ouvrages: Le miracle arabe par Max Vintéjoux; Louis Massignon. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):103, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904051.

D:1951:NRB

[285] S. D. Nécrologie: R. De Bengy-Puyvallée (1919–1951). Revue d’Histoire des Sciences et de Leurs Applications, 4(1):104, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904052.

deBroglie:1951:COH

[286] Louis de Broglie. Coup d’œil sur l’histoire de la science en France pendant la période 1799–1810. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):105–108, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904857.

Brunet:1951:VOCb

[287] Pierre Brunet. La vie et l’œuvre de Clairaut (suite). Revue d’Histoire des Sciences et de Leurs Applications, 4(2):109–153, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904858.

Humbert:1951:PM

[288] Pierre Humbert. Peiresc et le microscope. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):154–158, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904859.

Delorme:1951:LRS

[289] Suzanne Delorme. L’Académie Royale des Sciences : ses correspondants en Suisse. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):159–170, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904860.

Jansen:1951:CCS

[290] P. Jansen. Les carrosses à cinq sols et Christian Huygens. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):171–172, avril– juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904861.

Jansen:1951:TCA

[291] P. Jansen. Une tractation commerciale au XVII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):173–176, avril–juin 1951. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904862.

Taton:1951:DEO

[292] R. Taton. Découverte d’un exemplaire original du ^{≪}Brouillon project ≫sur les coniques de Desargues. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):176–181, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904863.

Pelseneer:1951:LIA

[293] Jean Pelseneer. Une lettre inédite d’Ampère. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):181–182, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904864.

VandeVelde:1951:BMH

[294] A. J. J. Van de Velde. Belgique: Le Musée de l’Histoire des Sciences à Gand. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):183– 184, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904865.

Taton:1951:CF

[295] R. Taton. Commémoration Foucault (1851–1951). Revue d’Histoire des Sciences et de Leurs Applications, 4(2):184–185, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23910683.

Taton:1951:F

[296] R. Taton. France. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):185–188, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904866.

Sergescu:1951:DLH

[297] P. Sergescu. Un doctorat ès lettres (histoire des sciences) à la sorbonne. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):188–191, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904867.

Burger:1951:PB

[298] D. Burger. Pays-Bas. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):191–192, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904868.

Anonymous:1951:RD

[299] Anonyme. Recherches de documents. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):192, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904869.

Itard:1951:AOB

[300] Jean Itard. Analyses d’ouvrages: Ciencia y Sociedad par Cortès Pla. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):193–194, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904870.

Delorme:1951:AOBc

[301] Suzanne Delorme. Analyses d’ouvrages: Einstein, sa vie et son temps, (Coll. Les Savants et le Monde) par Philippe Frank; André George; Albin Michel. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):194– 195, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904871.

Dijksterhuis:1951:AOBa

[302] E. J. Dijksterhuis. Analyses d’ouvrages: Grote sterrenkundigen van Ptolemaeus tot De Sitter par W. J. A. Schouten. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):195, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904872.

Dijksterhuis:1951:AOBb

[303] E. J. Dijksterhuis. Analyses d’ouvrages: Œuvres complètes de Christiaan Huygens, t. XXII. Supplément à la Correspondance. Varia. Biographie de Chr. Huygens. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):195–196, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904873.

Daumas:1951:AOBc

[304] Maurice Daumas. Analyses d’ouvrages: Un grand Français, le chimiste Thénard par Paul Thénard; Georges Bouchard. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):196–197, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904874.

Gille:1951:AOB

[305] Bertrand Gille. Analyses d’ouvrages: Les inventeurs célèbres. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):197–198, avril– juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904875.

T:1951:AQP

[306] R. T. Analyses de quelques périodiques: Isis, An International Review devoted to the History of Science and Civilization. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):198–199, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904876.

Taton:1951:AQPa

[307] R. Taton. Analyses de quelques périodiques: Archives internationales d’Histoire des Sciences, publication trimestrielle de l’Union internationale d’Histoire des Sciences, t. III (fasc. 10 à 13). Revue d’Histoire des Sciences et de Leurs Applications, 4(2):199–200, avril–juin 1951. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904877.

Berr:1951:AP

[308] Henri Berr. Avant-propos. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):201–203, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23905017.

deBroglie:1951:MHL

[309] Louis de Broglie. Un mathématicien, homme de lettres : d’Alembert. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):204– 212, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905018.

Rostand:1951:CHS

[310] Jean Rostand. La conception de l’homme selon Helvétius et selon Diderot. Revue d’Histoire des Sciences et de Leurs Applications, 4(3– 4):213–222, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905019.

Delorme:1951:SML

[311] Suzanne Delorme. Le salon de la Marquise de Lambert, berceau de l’Encyclopédie. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):223–227, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905020.

Bourdier:1951:BE

[312] Franck Bourdier and Yves François. Buffon et les Encyclopédistes. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):228–232, juillet– décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905021.

Sergescu:1951:CCL

[313] Pierre Sergescu. La contribution de Condorcet à l’Encyclopédie. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):233–237, juillet– décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905022.

Huard:1951:PLC

[314] Georges Huard. Les planches de l’Encyclopédie et celles de la description des arts et Métiers de l’Académie des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):238–249, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905023.

Humbert:1951:LL

[315] Pierre Humbert. L’astronomie dans l’Encyclopédie. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):250–254, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905024.

Taton:1951:MSB

[316] René Taton. Les mathématiques selon l’Encyclopédie. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):255–266, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905025.

Costabel:1951:MB

[317] Pierre Costabel. La mécanique dans l’Encyclopédie. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):267–293, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905026.

Vassails:1951:LP

[318] Gérard Vassails. L’Encyclopédie et la physique. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):294–323, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905027.

Bedel:1951:LCM

[319] Charles Bedel. L’avènement de la chimie moderne. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):324–333, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905028.

Daumas:1951:CLL

[320] Maurice Daumas. La chimie dans l’Encyclopédie et dans l’Encyclopédie méthodique. Revue d’Histoire des Sciences et de Leurs Applications, 4(3– 4):334–343, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905029.

Hooykaas:1951:CL

[321] R. Hooykaas. La cristallographie dans l’Encyclopédie. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):344–352, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905030.

Laignel-Lavastine:1951:MCL

[322] Maxime Laignel-Lavastine. Les médecins collaborateurs de l’Encyclopédie. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):353– 358, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905031.

Astruc:1951:SML

[323] Pierre Astruc. Les sciences médicales et leurs représentants dans l’Encyclopédie. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):359–368, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905032.

Humbert:1951:AOB

[324] Pierre Humbert. Analyses d’ouvrages: Nobel, Savants et découvertes, (Collection ^{≪}Les Savants et le Monde ^{≫}) par Louis de Broglie; André George. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4): 369–370, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905033.

Dijksterhuis:1951:AOBc

[325] E. J. Dijksterhuis. Analyses d’ouvrages: De verovering der materie. De groei van het wereldbeeld der Natuurkunde van de Grieken tot heden par Ir B. D. Swanenburg. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):370, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905034.

Dijksterhuis:1951:AOBd

[326] E. J. Dijksterhuis. Analyses d’ouvrages: Antoni van Leeuwenhoek. Zijn leven en werken. Deel I par A. Schierbeek. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):371, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905035.

Delorme:1951:AOBd

[327] A. Delorme. Analyses d’ouvrages: Paléontologie et Transformisme, (coll. ≪Sciences d’aujourd’hui ^{≫}) par G. G. Simpson; E. A. Stensiö; P. Teilhard de Chardin; H.-V. Vallois; J. Viret; D. M. S. Watson; C. Arambourg; L. Cuénot; P.-P. Grassé; J.-B.-S. Haldane; J. Piveteau. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):371–372, juillet– décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905036.

Delorme:1951:AOBe

[328] A. Delorme. Analyses d’ouvrages: Jan Ingenhousz, Plant Physiologist, with a History of the Discovery of Photosynthesis, vol. XI, n^{os} 5–6 par Howard S. Reed. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):372–373, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905037.

Wickersheimer:1951:AOB

[329] Ernest Wickersheimer. Analyses d’ouvrages: Paracelsus, Magic into Science par Henry M. Pachter. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):373–374, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905038.

Delorme:1951:AOBf

[330] A. Delorme. Analyses d’ouvrages: La Physiocratie sous les ministères de Turgot et de Necker (1774–1781) par Georges Weulersse; Paul Mantoux; J. Conan. Revue d’Histoire des Sciences et de Leurs Applications, 4 (3–4):375, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905039.

Taton:1951:AQPb

[331] R. Taton. Analyses de quelques périodiques: Osiris, Commentationes de scientiarum et eruditionis historia rationeque, Volumen nonum Maximiliano Meyerhof oblatum; Georgius Sarton Gandavensis. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):376, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905040.

Taton:1951:AQP

[332] R. Taton. Analyses de quelques périodiques: Centaurus, International Magazine of the History of Science and Medicine. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):376–377, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905041.

T:1951:AOB

[333] R. T. Analyses d’ouvrages: III^{e} Congrès national des Sciences, Bruxelles, 30 mai-3 juin 1950, vol. I : Historique du Congrès et Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):377, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905042.

Anonymous:1951:TGM

[334] Anonyme. Table générale des matières du Tome IV. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):378–384, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905043.

Anonymous:1951:BMa (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904053.

Anonymous:1951:BMb

[336] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):??, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904878.

Anonymous:1951:BMc

[337] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):??, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905044.

Anonymous:1951:FMa

[338] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 4(1):??, January 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904034.

Anonymous:1951:FMb

[339] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):??, avril–juin 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904856.

Anonymous:1951:FMc

[340] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):??, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905016.

Rostand:1952:DB

[341] Jean Rostand. Diderot et la biologie. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):5–17, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903552.

Dulieu:1952:AHF

[342] Louis Dulieu. Les articles d’Henri Fouquet dans l’Encyclopédie. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):18–25, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903553.

Gille:1952:LDT

[343] Bertrand Gille. L’Encyclopédie, dictionnaire technique. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):26–53, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903554.

Jeune:1952:VSQ

[344] Armand Machabey Jeune. Vue sommaire sur quelques rapports entre l’Encyclopédie et la métrologie. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):54–63, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903555.

Dautry:1952:OIL

[345] Jean Dautry. Une œuvre inspirée de l’Encyclopédie : le dictionnaire de l’industrie de 1776. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):64–72, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903556.

Cotard:1952:PFA

[346] Henri Cotard. Un portrait de Fermat au Musée des Beaux-Arts de Narbonne? Revue d’Histoire des Sciences et de Leurs Applications, 5(1): 73–76, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903557.

Francois:1952:QRF

[347] Y. François and T. François. Quelques remarques sur les ^{≪}Éléments de Physiologie^{≫}de Diderot. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):77–82, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903558.

Huard:1952:F

[348] Georges Huard. France. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):82–86, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903559.

Anonymous:1952:UIHa

[349] Anonyme. Union Internationale d’Histoire des Sciences: Résolutions Adoptées Par Le VI^{e} Congrès International d’Histoire des Sciences (Amsterdam, 14–21 août 1950). Revue d’Histoire des Sciences et de Leurs Applications, 5(1):86–88, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903560.

Grosclaude:1952:AOB

[350] P. Grosclaude and Henri Bernard-Maitre. Analyses d’ouvrages: Un audacieux message. L’Encyclopédie par Pierre Glosclaude; Gaston Monnerville; Pierre Clarac. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):89–91, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903561.

Francois:1952:AOBa

[351] Y. François. Analyses d’ouvrages: Inventaire du Fonds Vandeul et inédits de Diderot par Herbert Dieckmann. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):91, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903562.

Francois:1952:AOBb

[352] Y. François. Analyses d’ouvrages: Le rêve de d’Alembert …, édition critique d’après le manuscrit autographe de Diderot par Diderot; Paul Vernière. Revue d’Histoire des Sciences et de Leurs Applications, 5(1): 92–93, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903563.

T:1952:AOBa

[353] R. T. Analyses d’ouvrages: Revue d’Histoire littéraire de la France, 51^{e} année, n^{o} 3, juill.-sept. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):93, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903565.

D:1952:AOBa

[354] S. D. Analyses d’ouvrages: Revue de Synthèse, t. XXVIII de la nouvelle série, t. LXIX de la série générale (janv.-juin 1951). Revue d’Histoire des Sciences et de Leurs Applications, 5(1):93–94, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903566.

Itard:1952:AOBa

[355] Jean Itard. Analyses d’ouvrages: L’œuvre mathématique de G. Desargues. Textes publiés et commentés, avec une introduction biographique et historique par René Taton. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):94–95, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903567.

Rostand:1952:AOB

[356] Jean Rostand. Analyses d’ouvrages: Étude anatomique du mouvement du coeur et du sang chez les animaux par Guillaume Harvey; Charles Laubry. Revue d’Histoire des Sciences et de Leurs Applications, 5(1): 95–97, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903568.

Francois:1952:AOBc

[357] Y. François. Analyses d’ouvrages: Les théories de l’Évolution. (Origine et Histoire du transformisme et des idées qui s’y rattachent.) (Bibliothèque Scientifique) par Paul Ostoya; Roger Heim. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):97–98, January 1952. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903569.

D:1952:AOBb

[358] E. J. D. Analyses d’ouvrages: Antoni Van Leeuwenhoek. Zijn leven en zijn werken. Deel II par A. Schierbeek. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):98, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903570.

D:1952:AOBc

[359] E. J. D. Analyses d’ouvrages: Descriptive Catalogue of the physical instruments of the 18th century (including the collection ’sGravesande-Musschenbroek) in the Rijksmuseum voor de Geschiedenis der Natuurwetenschappen (National Museum of the History of Science) at Leyden, Communication n^{o} 81 from the above mentioned Museum, 1951 par C. A. Crommelin. Revue d’Histoire des Sciences et de Leurs Applications, 5 (1):98, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903571.

D:1952:AOBd

[360] E. J. D. Analyses d’ouvrages: Geschichte der Wissenschaften par Erich Rothacker. Revue d’Histoire des Sciences et de Leurs Applications, 5(1): 99, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903572.

Tetry:1952:AOBa

[361] A. Tétry. Analyses d’ouvrages: On acute diseases and on chronic diseases par Caelius Aurelianus; I. E. Drabkin. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):99–100, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903573.

T:1952:AOBb

[362] R. T. Analyses d’ouvrages: Esquisse d’une Histoire de la Météorologie en Belgique, fasc. XL par L. Dufour. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):100, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903575.

Tetry:1952:AOBb

[363] A. Tétry. Analyses d’ouvrages: Pasteur et la microbiologie, (collection ^{≪}Que sais-je? ^{≫}, n^{o} 467) par A. Delaunay. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):100, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903574.

Koyre:1952:MCJ

[364] Alexandre Koyré. La mécanique céleste de J. A. Borelli. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):101–138, avril–juin 1952. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905403.

LeLionnais:1952:DE

[365] F. Le Lionnais. Descartes et Einstein. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):139–154, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23905404.

Rostand:1952:EHA

[366] Jean Rostand. Esquisse d’une histoire de l’atomisme en biologie (fin) : IV. — les idées actuelles sur le Gène. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):155–170, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905405.

Itard:1952:OAC

[367] Jean Itard. Les opinions de l’abbé de la chapelle sur l’enseignement des mathématiques. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):171–175, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905406.

Anonymous:1952:Fa

[368] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):175–179, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905407.

Anonymous:1952:VCI

[369] Anonyme. VII^{e} Congrès International d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):179, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23910687.

Schuhl:1952:NEB

[370] Pierre-Maxime Schuhl. Nécrologie: Émile Bréhier (12 avril 1876–3 février 1952). Revue d’Histoire des Sciences et de Leurs Applications, 5(2):180, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905408.

Delorme:1952:AOBa

[371] Suzanne Delorme. Analyses d’ouvrages: Coup d’oeil sur les origines de la science exacte moderne (collection ^{≪}Esprit et Méthode ^{≫}) par Pierre Sergescu. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):181–182, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905409.

Taton:1952:AOBa

[372] R. Taton. Analyses d’ouvrages: Les mathématiques et l’imagination, (Bibliothèque scientifique) par E. Kasner; J. Newman; F. Beris; F. Le Lionnais. Revue d’Histoire des Sciences et de Leurs Applications, 5(2): 182, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905410.

Dijksterhuis:1952:AOBa

[373] E. J. Dijksterhuis. Analyses d’ouvrages: Passermeetkunde par S. C. Van Veen. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):183, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905411.

Itard:1952:AOBb

[374] Jean Itard. Analyses d’ouvrages: L’œuvre scientifique de Monge par René Taton. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):183– 185, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905412.

Taton:1952:AOBb

[375] R. Taton. Analyses d’ouvrages: Œuvres mathématiques d’Évariste Galois publiées en 1897, suivies d’une notice sur Évariste Galois et la théorie des équations algébriques, 2^{e} ed. par Évariste Galois; G. Verriest. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):185–186, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905413.

Rochot:1952:AOBa

[376] Bernard Rochot. Analyses d’ouvrages: L’activité rationaliste de la physique contemporaine, Bibliothèque de Philosophie contemporaine par Gaston Bachelard. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):186–187, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905414.

Dijksterhuis:1952:AOBb

[377] E. J. Dijksterhuis. Analyses d’ouvrages: A History of the Theory of Colour par K. T. A. Halbertsma. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):187–189, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23905415.

Virieux:1952:AOB

[378] Antoinette Virieux. Analyses d’ouvrages: La philosophie de Charles Bonnet de Genève par R. Savioz; Mémoires autobiographiques de Charles Bonnet de Genève. Bibliothèque des Textes philosophiques. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):189–190, avril– juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905416.

Laignel-Lavastine:1952:AOB

[379] M. Laignel-Lavastine. Analyses d’ouvrages: La doctrine classique de la médecine indienne. Ses origines et ses parallèles grecs par Jean Filliozat. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):190–192, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905417.

Schuhl:1952:AOB

[380] P.-M. Schuhl. Analyses d’ouvrages: Asklépios, Apollon Smintheus et Rudra, Études sur le dieu à la taupe et le dieu au rat dans la Grèce et dans l’Inde par Henri Grégoire; R. Goossens; M. Mathieu. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):192, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905418.

Wickersheimer:1952:AOBa

[381] Ernest Wickersheimer. Analyses d’ouvrages: Un maître et un ami de Farabeuf: Marcellin Duval, 1807–1899 par Nguyên-Bính. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):193, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905419.

Wickersheimer:1952:AOBb

[382] Ernest Wickersheimer. Analyses d’ouvrages: La vie ardente de Laënnec par Arnauld de Corbie. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):193–194, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905420.

Taton:1952:AQPa

[383] René Taton. Analyses de quelques périodiques: Archives internationales d’Histoire des Sciences, publication trimestrielle de l’Union internationale d’Histoire des Sciences, t. IV (n^{os} 14 à 17). Revue d’Histoire des Sciences et de Leurs Applications, 5(2):194–196, avril–juin 1952. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905421.

Aron:1952:AOB

[384] J.-P. Aron. Analyses d’ouvrages: Annual Studies in the History of Chemistry, t. 3 par Chymia. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):196, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905422.

Schuhl:1952:PEP

[385] Pierre-Maxime Schuhl. Les premières étapes de la philosophie biologique. Revue d’Histoire des Sciences et de Leurs Applications, 5(3): 197–221, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904387.

Belin-Milleron:1952:NEH

[386] J. Belin-Milleron. Les naturalistes et l’essor de l’humanisme expérimental (fin du XVI_{e}, début du XVII^{e} siècle): de Rondelet au conseiller Peiresc. Revue d’Histoire des Sciences et de Leurs Applications, 5(3): 222–233, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904388.

Wickersheimer:1952:JBC

[387] Ernest Wickersheimer. Jean Bauhin et le contrôle des compositions médicinales à Montbéliard. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):234–245, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904389.

Tresse:1952:CAM

[388] R. Tresse. Le conservatoire des arts et Métiers et la Société d’encouragement pour l’industrie nationale au début du XIX_{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):246– 264, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904390.

VandeVelde:1952:LFB

[389] A. J. J. Van de Velde. Le Liber I^{≪}Historiae Animalium^{≫}de 1551 de Conrad Gesner. Revue d’Histoire des Sciences et de Leurs Applications, 5(3): 265–269, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904391.

Anonymous:1952:C

[390] Anonyme. Commémorations. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):269–273, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904392.

Delorme:1952:AOBb

[391] Suzanne Delorme. Analyses d’ouvrages: Les grandes figures de la Science française par Luc Durtain; Louis de Broglie. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):274, juillet–septembre 1952. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904393.

Rochot:1952:AOBb

[392] B. Rochot. Analyses d’ouvrages: Descartes, Théoricien géant et solitaire par Paul Dupont. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):274–275, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904394.

Rochot:1952:AOBc

[393] B. Rochot. Analyses d’ouvrages: La pensée de Leibniz par Yvon Belaval. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):276, juillet– septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904395.

Taton:1952:AOBc

[394] R. Taton. Analyses d’ouvrages: Geschichte der Mathematik, (Geschichte der Wissenschaften, II Naturwissenschaften) par Oskar Becker; Josef E. Hofmann; Erik Rothacker. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):276, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904396.

Itard:1952:AOBc

[395] Jean Itard. Analyses d’ouvrages: Pour connaître: Mathematics in Aristotle par Thomas Heath. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):276–278, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904397.

Dijksterhuis:1952:AOBc

[396] E. J. Dijksterhuis. Analyses d’ouvrages: De groei van ons wereldbeeld. Een geschiedenis van de sterrekunde par A. Pannekoek. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):278–279, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904398.

Dijksterhuis:1952:AOBd

[397] E. J. Dijksterhuis. Analyses d’ouvrages: Die Astronomie der Pythagoreer. (Verhandelingen der Koninklijke Nederlandse Akademie van Wetenschappen, afd. Natuurkunde. Eerste Reeks, Deel XX, n^{o}. 1) par B. L. Van der Waerden. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):279–281, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904399.

Courtin:1952:AOB

[398] Madeleine Courtin. Analyses d’ouvrages: Histoire abrégée des théories physiques concernant la matière et l’énergie par Charles Brunold. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):281–282, juillet– septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904400.

Delorme:1952:AOBc

[399] A. Delorme. Analyses d’ouvrages: Les grandes découvertes de la physique moderne, coll. ^{≪}Les grandes découvertes scientifiques ^{≫}par Pierre Guaydier. Revue d’Histoire des Sciences et de Leurs Applications, 5(3): 282, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904401.

Daumas:1952:AOBa

[400] Maurice Daumas. Analyses d’ouvrages: Outline of the History of Chemistry par Herbert S. Klickstein. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):283–284, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904402.

Daumas:1952:AOBb

[401] Maurice Daumas. Analyses d’ouvrages: La vie émouvante et noble de Gay-Lussac par Edmond Blanc; Léon Delhoume. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):284–285, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904403.

Bernard-Maitre:1952:AOB

[402] Henri Bernard-Maitre. Analyses d’ouvrages: Genesis and Geology. A Study in the Relations of Scientific Thought, Natural Theology and Social Opinion in Great Britain, 1790–1850 par Charles Coulston Gillespie. Revue d’Histoire des Sciences et de Leurs Applications, 5(3): 285–286, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904404.

D:1952:AOBe

[403] S. D. Analyses d’ouvrages: Les Conférences du Palais de la Découverte. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):286– 287, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904405.

Taton:1952:AQPb

[404] René Taton. Analyses de quelques périodiques: Isis, vol. 42, parts 1, 2, 3, 4 (fasc. 127, 128, 129, 130), April–June–October–December 1951. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):288, juillet– septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904407.

deBroglie:1952:PQR

[405] Louis de Broglie. La physique quantique restera-t-elle indéterministe? Revue d’Histoire des Sciences et de Leurs Applications, 5(4):289–311, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905080.

Hofmann:1952:PPR

[406] Jos.-E. Hofmann and Pierre Costabel. À propos d’un problème de Roberval. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):312–333, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905081.

Brunet:1952:VOC

[407] Pierre Brunet. La vie et l’œuvre de clairaut: VI. — le système du monde newtonien la polémique avec Buffon sur l’attraction (suite). Revue d’Histoire des Sciences et de Leurs Applications, 5(4):334–349, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905082.

Yahia:1952:AMS

[408] B. Ben Yahia. Avicenne médecin: Sa vie, son œuvre. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):350–358, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905083.

Taton:1952:QPC

[409] René Taton. Quelques précisions sur le chimiste Clouet et deux de ses homonymes. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):359–367, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905084.

Anonymous:1952:A

[410] Anonyme. Allemagne. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):368, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905085.

Anonymous:1952:B

[411] Anonyme. Belgique. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):368, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905086.

Anonymous:1952:Fb

[412] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):368–370, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905087.

Anonymous:1952:PB

[413] Anonyme. Pays-Bas. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):370, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905088.

Anonymous:1952:S

[414] Anonyme. Suisse. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):370, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905089.

Anonymous:1952:UIHb

[415] Anonyme. Union Internationale d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):370–371, October 1952. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905090.

Anonymous:1952:U

[416] Anonyme. U.S.A. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):371, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905091.

Sergescu:1952:AOB

[417] P. Sergescu. Analyses d’ouvrages: Panorama general de historia de la ciencia. III. La Eclosion del Renacimento, IV. Lionardo da Vinci, sabio par Aldo Mieli. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):372–373, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905092.

Delorme:1952:AOBd

[418] Suzanne Delorme. Analyses d’ouvrages: Francis Bacon, a Bibliography of his Works and of Baconiana to the year 1750 par R. W. Gibson. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):373–374, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905093.

Chalus:1952:AOB

[419] Paul Chalus. Analyses d’ouvrages: Traité akkadien de diagnostics et pronostics médicaux par René Labat. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):374–377, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23905094.

Wickersheimer:1952:AOBc

[420] Ernest Wickersheimer. Analyses d’ouvrages: The life and times of Gaspare Tagliacozzi, surgeon of Bologna, 1545–1599, with a documented study of the scientific and cultural life of Bologna in the sixteenth century … par Martha Teach Gnudi; Jerome Pierce Webster; Arturo Castiglioni. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):377–379, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905095.

Taton:1952:AOBd

[421] Juliette Taton. Analyses d’ouvrages: Les Germains, (Bibliothèque de Synthèse historique, collection ^{≪}L’Évolution de l’Humanité ^{≫}, n^{o} 23) par Henri Hubert; Henri Berr. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):379–380, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905096.

T:1952:AOBc

[422] R. T. Analyses d’ouvrages: Thalès, années 1949–1950. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):380, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905097.

Anonymous:1952:TGM

[423] Anonyme. Table générale des matières du Tome V. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):381–388, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905098.

Anonymous:1952:BMa

[424] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):??, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903576.

Anonymous:1952:BMb

[425] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):??, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905423.

Anonymous:1952:BMc

[426] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):??, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904408.

Anonymous:1952:BMd

[427] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):??, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905099.

Anonymous:1952:FMa

[428] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 5(1):??, January 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903551.

Anonymous:1952:FMb

[429] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):??, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905402.

Anonymous:1952:FMc

[430] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):??, juillet–septembre 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904386.

Anonymous:1952:FMd

[431] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):??, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905079.

Brunet:1953:VOC

[432] Pierre Brunet. La vie et l’œuvre de Clairaut: VII. — le retour de la comète de Halley. les dernières années de Clairaut (fin). Revue d’Histoire des Sciences et de Leurs Applications, 6(1):1–17, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903498.

Humbert:1953:TRC

[433] Pierre Humbert. Un témoignage sur la recherche des coordonnées géographiques au XVII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):18–21, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903499.

Vassails:1953:LE

[434] Gérard Vassails. Lóránd Eötvös. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):22–41, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903500.

Cailleux:1953:PDN

[435] André Cailleux. Progression du nombre d’espèces de plantes décrites de 1500 à nos jours. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):42–49, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903501.

Dulieu:1953:PPL

[436] Louis Dulieu. Un parisien, professeur à l’Université de Médecine de Montpellier : Charles Le Roy (1726–1779). Revue d’Histoire des Sciences et de Leurs Applications, 6(1):50–59, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903502.

Taton:1953:IEP

[437] R. Taton. Inventaire des exemplaires des premières éditions des ≪Principia^{≫}de Newton. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):60–63, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903503.

Anonymous:1953:B

[438] Anonyme. Belgique. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):63, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903504.

Anonymous:1953:Fa

[439] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):63–66, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23910021.

Anonymous:1953:PBa

[440] Anonyme. Pays-Bas. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):66, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23910022.

Anonymous:1953:U

[441] Anonyme. U.S.A. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):67, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23910023.

Taton:1953:AOBa

[442] René Taton. Analyses d’ouvrages: Horus, A Guide to the History of Science; a first Guide for the Study of the History of Science with introductory Essays on Science and Tradition par George Sarton. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):68–69, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903505.

Delorme:1953:AOBa

[443] Suzanne Delorme. Analyses d’ouvrages: The Early History of Science. A short Handlist (Helps for students of History, n^{o} 52). Revue d’Histoire des Sciences et de Leurs Applications, 6(1):69, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903506.

Costabel:1953:AOBa

[444] Pierre Costabel. Analyses d’ouvrages: Histoire de la pensée scientifique, (collection ^{≪}La Colombe ^{≫}) par François Russo. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):69–70, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903507.

Itard:1953:AOBa

[445] Jean Itard. Analyses d’ouvrages: Histoire de la science et de ses rapports avec la philosophie et la religion (Bibliothèque scientifique) par W. Dampier; René Sudre. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):70–71, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903508.

Chalus:1953:AOBa

[446] Paul Chalus. Analyses d’ouvrages: L’^{≪}Encyclopédie ^{≫}et le progrès des sciences et des techniques, (Centre international de Synthèse, section d’Histoire des Sciences). Revue d’Histoire des Sciences et de Leurs Applications, 6(1):71, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903509.

Itard:1953:AOBb

[447] Jean Itard. Analyses d’ouvrages: La magie des nombres (Bibliothèque scientifique) par E. T. Bell; S. Fabre; P. Goyard. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):72, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903510.

Costabel:1953:AOBb

[448] Pierre Costabel. Analyses d’ouvrages: Histoire de la mécanique par René Dugas; Louis de Broglie. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):72–74, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903511.

Taton:1953:AOBb

[449] R. Taton. Analyses d’ouvrages: Storia della luce par Vasco Ronchi. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):74–75, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903512.

Daumas:1953:AOBa

[450] Maurice Daumas. Analyses d’ouvrages: Histoire de la chimie par Jaen Timmermans. Revue d’Histoire des Sciences et de Leurs Applications, 6 (1):75, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903513.

Daumas:1953:AOBb

[451] Maurice Daumas. Analyses d’ouvrages: Chimie et chimistes par R. Massain. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):75–77, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903514.

Daumas:1953:AOBc

[452] Maurice Daumas. Analyses d’ouvrages: The Essays of Jean Rey par Douglas McKie. Revue d’Histoire des Sciences et de Leurs Applications, 6(1): 77–79, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903515.

Daumas:1953:AOBd

[453] Maurice Daumas. Analyses d’ouvrages: Antoine Lavoisier, scientist, economist, social Reformer par Douglas McKie. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):79–80, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903516.

Tetry:1953:AOBa

[454] A. Tétry. Analyses d’ouvrages: Les origines de la biologie expérimentale et l’abbé Spallanzani par Jean Rostand. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):80–81, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903517.

Delorme:1953:AOBb

[455] Albert Delorme. Analyses d’ouvrages: Abraham Trembley of Geneva, Scientist and Philosopher (1710–1784) par John R. Baker. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):81–83, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903518.

Delorme:1953:AOBc

[456] A. Delorme. Analyses d’ouvrages: Journal of Researches into the Geology and Natural History of the various Countries visited par Charles Darwin; H. M. S. Beagle. Revue d’Histoire des Sciences et de Leurs Applications, 6 (1):83, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903519.

Tetry:1953:AOBb

[457] A. Tétry. Analyses d’ouvrages: Esquisses et notes de travail inédites de Claude Bernard, recueillies et commentées par Léon Binet. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):84, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903520.

Wickersheimer:1953:AOBa

[458] Ernest Wickersheimer. Analyses d’ouvrages: La Faculté de Médecine de Paris, cinq siècles d’art et d’histoire par Léon Binet; Pierre Vallery-Radot. Revue d’Histoire des Sciences et de Leurs Applications, 6(1): 84–86, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903521.

Cuvillier:1953:AOB

[459] Armand Cuvillier. Analyses d’ouvrages: Centenaire de la Société médicopsychologique, numéro des Annales médico-psychologiques consacré au compte rendu des séances des 25 et 26 mai 1952, juin 1952. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):87–88, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903522.

Wickersheimer:1953:AOBb

[460] Ernest Wickersheimer. Analyses d’ouvrages: The apologie and treatise … containing the voyages made into divers places, with many of his writings upon surgery par Ambroise Paré; Geoffrey Keynes. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):88, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903523.

Delorme:1953:AOBd

[461] A. Delorme. Analyses d’ouvrages: Le Progrès médical, n^{os} 13–14, 10– 24 juillet 1952 : Cent cinquante ans après la mort de Bichat. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):89, January 1953.

CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903524.

Wickersheimer:1953:AOBc

[462] Ernest Wickersheimer. Analyses d’ouvrages: Contribution à l’histoire de l’analyse des urines, (Thèse de doctorat en pharmacie de Strasbourg (2).) par Charles Guyotjeannin. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):89–90, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903525.

Taton:1953:AOBc

[463] R. Taton. Analyses d’ouvrages: Les sources historiques de la métrologie, (Extrait de la Revue de Métrologie, avril 1952.) par Armand Machabey. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):91, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903526.

Taton:1953:AOBd

[464] R. Taton. Analyses d’ouvrages: Aperçu historique de l’activité scientifique de la Belgique en Géodésie et en Géophysique. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):91, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903527.

Delorme:1953:AOBe

[465] Suzanne Delorme. Analyses d’ouvrages: Le Mont Cenis, sa route, son tunnel …, Actualités scientifiques et industrielles, n^{o} 1177 par Albert Duluc; Georges Friedmann. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):91–92, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903528.

Humbert:1953:AOB

[466] P. Humbert. Analyses d’ouvrages: La France : Géographie, Tourisme. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):92, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903529.

Taton:1953:POG

[467] René Taton. La première œuvre géométrique de Philippe de la Hire. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):93–111, avril– juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903902.

Lenoble:1953:HPP

[468] Robert Lenoble. Histoire et physique: À propos des conseils de Mersenne aux historiens et de l’intervention de Jean de Launoy dans la querelle gassendiste. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):112–134, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903903.

Nicolle:1953:FCE

[469] Jacques Nicolle. ^{≪}Corps énantiomorphes et organismes vivants ^{≫}dans l’œuvre de Louis Pasteur. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):135–149, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903904.

Rostand:1953:CB

[470] Jean Rostand. Cournot et la biologie. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):150–160, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903905.

Taton:1953:EBO

[471] René Taton. Esquisse d’une bibliographie de l’œuvre de clairaut. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):161–168, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903906.

Gloden:1953:MDL

[472] A. Gloden. La méthode du luxembourgeois B. I. Clasen pour la résolution d’un système d’équations linéaires (méthode des coefficients égaux). Revue d’Histoire des Sciences et de Leurs Applications, 6(2):168–170, avril– juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903907.

Laignel-Lavastine:1953:F

[473] M. Laignel-Lavastine. France. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):171–174, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903908.

Anonymous:1953:PBE

[474] Anonyme. Pays-Bas: édition des Œuvres Principales de Simon Stevin. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):174–175, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903912.

Anonymous:1953:UIH

[475] Anonyme. Union Internationale d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):175, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903913.

Vintejoux:1953:AOB

[476] M. Vintéjoux. Analyses d’ouvrages: La vie quotidienne des Musulmans au Moyen Âge, (collection ^{≪}La vie quotidienne ^{≫}) par Aly Mazaheri. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):176, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903914.

Taton:1953:AOBe

[477] René Taton. Analyses d’ouvrages: Swedish Men of Science 1650–1950 par Sten Lindroth. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):176–177, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903915.

Bodet:1953:AOBa

[478] Suzanne Bodet. Analyses d’ouvrages: Russia’s Lomonosov par Boris N. Menshutkin; J. Eyre. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):177–178, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903916.

Rossier:1953:AOBa

[479] P. Rossier. Analyses d’ouvrages: L’œuvre scientifique el technique du général Guillaume-Henri Dufour par F. Baeschlin; H. Favre; L. Kollros; F. Stüssi. Revue d’Histoire des Sciences et de Leurs Applications, 6(2): 178, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903917.

Taton:1953:AOBf

[480] R. Taton. Analyses d’ouvrages: Une suite de l’Encyclopédie: le Journal encyclopédique (1756–1793): notes, documents, et extraits réunis par Gustave Charlier; Roland Mortier. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):178–179, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903918.

Itard:1953:AOBc

[481] Jean Itard. Analyses d’ouvrages: Le livre des nombres carrés, traduit pour la première fois du latin médiéval en français par Léonard de Pise; Paul Ver Eecke. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):179–180, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903919.

Itard:1953:AOBd

[482] Jean Itard. Analyses d’ouvrages: Beihefte zur Zeitschrift ^{≪}Elemente der Mathematik ^{≫}, Suppléments à la ^{≪}Revue de Mathématiques élémentaires ≫par Kurze Mathematiker-Biographen. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):180–181, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903920.

Rossier:1953:AOBb

[483] P. Rossier. Analyses d’ouvrages: La figure du monde par Rolin Wavre. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):181, avril– juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903921.

Rostand:1953:AOBa

[484] Jean Rostand. Analyses d’ouvrages: Buffon, collection des ^{≪}Grands Naturalistes français ^{≫}par L. Bertin; F. Bourdier; E. Dechambre; Y. François; E. Genet-Varcin; G. Heilbrun; R. Heim; J. Pelseneer; J. Piveteau. Revue d’Histoire des Sciences et de Leurs Applications, 6 (2):181–183, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903922.

Chalus:1953:AOBb

[485] P. Chalus. Analyses d’ouvrages: Plants of the Bible, (Paris, Librairie Raymann, 17, rue de Tournon, 6^{e}) (vol. XXVIII de la ^{≪}New Series of Plant Science Books ^{≫}) par H. N. Moldenke; A. L. Moldenke. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):183, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903923.

Wickersheimer:1953:AOBd

[486] Ernest Wickersheimer. Analyses d’ouvrages: Hospitaliers d’autrefois, Hôpital général de Paris (1656–1790) par Henry Légier–Desgranges. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):183–185, avril– juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903924.

Bodet:1953:AOBb

[487] Suzanne Bodet. Analyses d’ouvrages: The Theodosian Code and Novels and the Sirmondian Constitutions par Clyde Pharr; Th. Sherrer Davidson; Brown Pharr. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):185, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903925.

Taton:1953:AQPa

[488] R. Taton. Analyses de quelques périodiques: Archives internationales d’Histoire des Sciences, publication trimestrielle de l’Union internationale d’Histoire des Sciences, t, . V (n os 18 à 21), 1952. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):185–186, avril– juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903926.

Taton:1953:AQPb

[489] René Taton. Analyses de quelques périodiques: Isis, vol. 43, parts 1, 2, 3, 4 (fasc. 131, 132, 133, 134), April–July–September–October 1952. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):186–187, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903927.

Chalus:1953:AQP

[490] Paul Chalus. Analyses de quelques périodiques: Osiris. Volumen decimum Henrico Berr oblatum. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):187–188, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903928.

Taton:1953:AOBg

[491] René Taton. Analyses d’ouvrages: Les Conférences du Palais de la Découverte, série D (Histoire des Sciences). Revue d’Histoire des Sciences et de Leurs Applications, 6(2):188, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903929.

Jacquot:1953:HSA

[492] Jean Jacquot. Humanisme et science dans l’Angleterre élisabéthaine: L’œuvre de Thomas Blundeville. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):189–202, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904110.

Rostand:1953:SDM

[493] François Rostand. Schopenhauer et les démonstrations mathématiques. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):203– 230, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904111.

Rossier:1953:COH

[494] Paul Rossier. Coup d’œil sur l’histoire des sciences exactes à Genève. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):231– 249, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904112.

Astruc:1953:RBA

[495] Pierre Astruc. Rabelais botaniste, anatomiste et physiologiste. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):250–261, juillet– septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904113.

Chauvois:1953:PDI

[496] L. Chauvois. À propos du IV^{e} centenaire de la mort de Michel servet: Sa place dans l’histoire de l’anatomie. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):262–268, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904114.

Anonymous:1953:BHS

[497] Anonyme. Bibliographie de l’histoire des sciences pendant la période révolutionnaire. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):269–271, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904115.

Anonymous:1953:Fb

[498] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):271–272, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904116.

Delorme:1953:IVC

[499] S. Delorme. Israël: VII^{e} Congrès International d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 6(3): 272–275, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904117.

Anonymous:1953:I

[500] Anonyme. Italie. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):275, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904118.

Anonymous:1953:LCF

[501] Anonyme. Luxembourg: Congrès de L’A.F.A.S. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):275–276, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904119.

Anonymous:1953:PBb

[502] Anonyme. Pays-Bas. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):276–277, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904120.

Anonymous:1953:S

[503] Anonyme. Suisse. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):277, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904121.

Rochot:1953:AOB

[504] Bernard Rochot. Analyses d’ouvrages: Philosophes et savants, ≪Bibliothèque de Philosophie scientifique^{≫} par Pierre Humbert. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):278–279, juillet– septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904122.

Taton:1953:AOBh

[505] René Taton. Analyses d’ouvrages: Scritti di storia e critica della scienza par Sebastiano Timpanaro. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):279, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904123.

Dijksterhuis:1953:AOBa

[506] E. J. Dijksterhuis. Analyses d’ouvrages: Geschichte der Geschichtswissenschaft. 2., Auflage (Geschichte der Wissenschaften, herausgegeben I. Geisteswissenschaften) par Karl Brandi; Erich Rothacker; Wolfgang Graf. Revue d’Histoire des Sciences et de Leurs Applications, 6(3): 279–280, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904124.

Colnort-Bodet:1953:AOBa

[507] Suzanne Colnort-Bodet. Analyses d’ouvrages: Svenska studieresor till England under Frihetstiden [Les chercheurs suédois en Angleterre pendant l’ère de liberté suédoise, 1718–1772] par Sven Rydberg. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):280, juillet– septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904125.

Delorme:1953:AOBf

[508] Suzanne Delorme. Analyses d’ouvrages: Giordano Bruno, His Life and Thought, with annotated Translation of his work On the Infinite Universe and Worlds par Dorothea Waley Singer. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):281, juillet–septembre 1953. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904126.

Dugas:1953:AOB

[509] R. Dugas. Analyses d’ouvrages: Ballistics in the seventh Century par A. R. Hall. Revue d’Histoire des Sciences et de Leurs Applications, 6(3): 281–282, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904127.

Rostand:1953:AOBb

[510] Jean Rostand. Analyses d’ouvrages: ^{≪}Les deux lettres de William Harvey à Jean Riolan sur la circulation du sang^{≫}(nouvelle traduction française), Biologie médicale, numéro hors série, vol. XLII, avril 1953 par L. Chauvois. Revue d’Histoire des Sciences et de Leurs Applications, 6(3): 282–284, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904128.

Colnort-Bodet:1953:AOBb

[511] S. Colnort-Bodet. Analyses d’ouvrages: Michel Servet, hérétique et martyr par Roland H. Bainton. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):284–285, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904129.

Wickersheimer:1953:AOBe

[512] Ernest Wickersheimer. Analyses d’ouvrages: William Shippen, Jr., pioneer in American medical education, a biographical essay … par Betsy Copping Corner. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):285–286, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904130.

Dijksterhuis:1953:AOBb

[513] E. J. Dijksterhuis. Analyses d’ouvrages: De mens bouwt zich een wereld. Vijf duizend jaar techniek. Amsterdam par Pr R. J. Forbes. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):287, juillet– septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904131.

Dijksterhuis:1953:AOBc

[514] E. J. Dijksterhuis. Analyses d’ouvrages: Nederlandse Gewichten. Stelsels, Ijkwezen, vormen, makers en merken, Mededeling n^{o} 86 uit het Rijksmusuem voor de Geschiedenis der Natuurwetenschappen te Leiden par K. M. C. Zevenboom; D. A. Wittop Koning. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):287–288, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904132.

Colnort-Bodet:1953:AOBc

[515] S. Colnort-Bodet. Analyses d’ouvrages: Lychnos. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):288, juillet–septembre 1953. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904133.

daCAndrade:1953:NCH

[516] E. N. da C. Andrade, F.R.S. Newton: Considérations sur l’homme et son œuvre. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):289– 307, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905530.

Luce:1953:GPE

[517] Jean-H. Luce. Géométrie de la perspective à l’époque de vitruve. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):308–321 (1954), October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905531.

Humbert:1953:LMH

[518] Pierre Humbert. L’œuvre mathématique d’Henri Pitot. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):322–328 (1954), October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905532.

Laurent:1953:ALJ

[519] Aug. Laurent, J. B. Dumas, and J. Jacques. Auguste Laurent et J.-B. Dumas d’après une correspondance inédite. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):329–349, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905533.

Taton:1953:LSF

[520] René Taton. Laplace et Sylvestre-François Lacroix. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):350–360, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905534.

Taton:1953:EFB

[521] R. Taton. Enquête sur les ^{≪}Principia ^{≫}de Newton (II). Revue d’Histoire des Sciences et de Leurs Applications, 6(4):360–361, October 1953. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905535.

Bedel:1953:F

[522] Charles Bedel. France. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):361–363, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905536.

Anonymous:1953:PBc

[523] Anonyme. Pays-Bas. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):363, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905537.

Michel:1953:AOB

[524] Paul-Henri Michel. Analyses d’ouvrages: Étude sur la doctrine pythagoricienne de la tétrade, ^{≪}Les Belles-Lettres ^{≫}par Paul Kucharski. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):364–365, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905538.

Jacquot:1953:AOBa

[525] Jean Jacquot. Analyses d’ouvrages: Roger Bacon and his Search for a Universal Science par Stewart C. Easton. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):365–367, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905539.

Taton:1953:AOBi

[526] René Taton. Analyses d’ouvrages: Selections from the Notebooks of Leonardo da Vinci par Irma A. Richter. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):367, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905540.

Taton:1953:AOBj

[527] René Taton. Analyses d’ouvrages: Léonard de Vinci (1452–1519), Librairie Hachette (^{≪}Encyclopédie par l’image ^{≫}) par Fernand-Demeure. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):367, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905541.

Taton:1953:AOBk

[528] René Taton. Analyses d’ouvrages: Leonardo, numéro spécial du Bulletin de l’^{≪}Istituto tecnico industriale Leonardo da Vinci ^{≫}. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):367–368, October 1953. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905542.

Taton:1953:AOBl

[529] René Taton. Analyses d’ouvrages: Correspondência científica dirigada a João Jacinto de Magalhães (1769–1789). Contribuição para o seu Epistolário (Extrait de la Revista da Faculdade de Ciências da Universidade de Coimbra, vol. XX. — Inedita ac rediuiua. Subsídios para a História da Filosofia e da Ciência em Portugal, no 3.) par Joaquim de Carvalho. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):368, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905543.

Taton:1953:AOBm

[530] René Taton. Analyses d’ouvrages: Recherches arithmétiques (Reproduction photographique de l’édition originale de la traduction de Poullet-Delisle, Paris, Courcier imprimeur-libraire pour les mathématiques, 1807) par C.-Fr. Gauss; A.-C.-M. Poullet-Delisle. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):368–370, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905545.

Taton:1953:AOBn

[531] René Taton. Analyses d’ouvrages: The History of Astronomy, Translated from the Italian ^{≪}Storia dell’ Astronomia ^{≫}(^{≪}The Life of Science Library ^{≫}, n^{o} 24) par Giorgio Abetti; Betty Burr Abetti. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):370, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905547.

Jacquot:1953:AOBb

[532] Jean Jacquot. Analyses d’ouvrages: The Alchimists founders of modern Chemistry par F. Sherwood Taylor. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):370–372, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905549.

Wickersheimer:1953:AOBf

[533] Ernest Wickersheimer. Analyses d’ouvrages: Geschichte der Physiologie (Lehrbuch der Physiologie) par K. E. Rothschuh; Wilhelm Trendelenburg; Erich Schütz. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):372–373, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905551.

Theodorides:1953:AOB

[534] J. Théodoridès. Analyses d’ouvrages: The Gate of Heaven (Shaar-ha-Shama-yim, La Porte du Ciel) par Rabbi Gershon Ben Shlomoh d’Arles; F. S. Bodenheimer. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):374–375, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905553.

Collet:1953:AOB

[535] Cl.-G. Collet. Analyses d’ouvrages: Microorganismes de nos eaux douces (collection ^{≪}Sciences et Médecine ^{≫}, Section des Sc.) par Charles de Wyngaert; Robert Mathey. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):375, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905555.

Filliozat:1953:AOB

[536] J. Filliozat. Analyses d’ouvrages: ^{≪}Grundsätze altindischer Medizin ^{≫}, Acta historica scientiarum naturalium et medicinalium par Reinhold F. G. Müller. Revue d’Histoire des Sciences et de Leurs Applications, 6(4): 375, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905557.

Colnort-Bodet:1953:AQP

[537] Suzanne Colnort-Bodet. Analyses de quelques périodiques: Centaurus, International Magazine of the History of Science and Medicine. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):376, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905559.

Anonymous:1953:TGM

[538] Anonyme. Table générale des matières du Tome VI. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):377–384, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905561.

Anonymous:1953:BMa

[539] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):??, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903530.

Anonymous:1953:BMb

[540] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):??, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903930.

Anonymous:1953:BMc

[541] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):??, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904134.

Anonymous:1953:BMd

[542] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):??, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23910054.

Anonymous:1953:FMa

[543] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):??, January 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903497.

Anonymous:1953:FMb

[544] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 6(2):??, avril–juin 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903901.

Anonymous:1953:FMc

[545] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 6(3):??, juillet–septembre 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904109.

Anonymous:1953:FMd

[546] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 6(4):??, October 1953. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905529.

deDainville:1954:LMCa

[547] François de Dainville. L’enseignement des mathématiques dans les Collèges Jésuites de France du XVIe au XVIIIe siècle. Revue d’Histoire des Sciences et de Leurs Applications, 7:6–21, 1954. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

Dugas:1954:CH

[548] René Dugas. Sur le cartésianisme de Huygens. Revue d’Histoire des Sciences et de Leurs Applications, 7:22–33, 1954. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

Hadamard:1954:CHP

[549] Jacques Hadamard. Le centenaire de Henri Poincaré. Revue d’Histoire des Sciences et de Leurs Applications, 7:101–108, 1954. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

deDainville:1954:LMCb

[550] François de Dainville. L’enseignement des mathématiques dans les Collèges Jésuites de France du XVI^{e} au XVIII^{e} siècle. II. Revue d’Histoire des Sciences et de Leurs Applications, 7:109–123, 1954. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

Itard:1954:MTH

[551] Jean Itard. Sur la méthode de Tannery en histoire des mathématiques. Revue d’Histoire des Sciences et de Leurs Applications, 7:326–332 (1955), 1954. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

Taton:1955:EPC

[552] René Taton. L’ “Essay pour les coniques” de Pascal. Revue d’Histoire des Sciences et de Leurs Applications, 8:1–18, 1955. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

Koyre:1955:PEC

[553] Alexandre Koyré. Pour une édition critique des oeuvres de Newton. Revue d’Histoire des Sciences et de Leurs Applications, 8:19–37, 1955. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

Brun:1955:DMA

[554] Viggo Brun. Découverte d’un manuscrit d’Abel. Revue d’Histoire des Sciences et de Leurs Applications, 8:103–106, 1955. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

Levy:1955:DVD

[555] Paul. W. Lévy. Doeblin (V. Doblin) (1915–1940). Revue d’Histoire des Sciences et de Leurs Applications, 8:107–115, 1955. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

Costabel:1955:AMG

[556] Pierre Costabel. Autour de la méthode de Galilée pour la détermination des centres de gravité. Revue d’Histoire des Sciences et de Leurs Applications, 8:116–128, 1955. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

Speziali:1955:CIE

[557] Pierre Speziali. Une correspondance inédite entre Clairaut et Cramer. Revue d’Histoire des Sciences et de Leurs Applications, 8:193–237, 1955. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

Moscovici:1956:PQT

[558] Serge Moscovici. À propos de quelques travaux d’Adam Smith sur l’histoire et la philosophie des sciences. Revue d’Histoire des Sciences et de Leurs Applications, 9(1):1–20, January 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904694.

Eymard:1956:JMG

[559] P. Eymard and J. P. Lafon. Le journal mathématique de Gauss: Traduction française annotée. Revue d’Histoire des Sciences et de Leurs Applications, 9(1):21–51, January 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904695.

Gillispie:1956:NBL

[560] Charles Gillispie. Notice biographique de Lavoisier par Madame Lavoisier. Revue d’Histoire des Sciences et de Leurs Applications, 9(1): 52–61, January 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904696.

Rostand:1956:OMF

[561] Jean Rostand. Une œuvre méconnue : les ^{≪}Vestiges of the natural History of Creation^{≫}. Revue d’Histoire des Sciences et de Leurs Applications, 9(1):62–73, January 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904697.

Sergescu:1956:SFH

[562] Pierre Sergescu. Savants français en Hollande. Revue d’Histoire des Sciences et de Leurs Applications, 9(1):74–78, January 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904698.

Birembaut:1956:PIO

[563] Arthur Birembaut. À propos d’un important ouvrage concernant l’histoire des instruments scientifiques. Revue d’Histoire des Sciences et de Leurs Applications, 9(1):78–87, January 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904699.

Anonymous:1956:Fa

[564] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 9(1):87–90, January 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904700.

Lacoin:1956:AOB

[565] M. Lacoin. Analyses d’ouvrages: La Cultura in Sicilia nel periodo normanno, — La Cultura alla corte di Federico II imperatore — Civilita Medievale par Antonino de Stefano; Bizantini e bizantinismo nella Sicilia normanna, — Medioevo mediterraneo par Francesco Giunta. Revue d’Histoire des Sciences et de Leurs Applications, 9(1):91–92, January 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904701.

Taton:1956:AOA

[566] René Taton. Analyses d’ouvrages: Aldo Mieli, Desiderio Papp et José Babini, Panorama general de Historia de la Ciencia, La ciencia del Renacimiento: t. V par A. M., Matemática y Ciencias Naturales, Buenos Aires Mexico, Espasa-Calpe Argentina S. A. (^{≪}Historia y Filosofia de la Ciencia ^{≫}, serie menor), s. d. (1952); t. VI (par D. P. et J. B.), Astronomia, Fisica y Biologia, id., s. d. (1952); t. VII (par D. P. et J. B.), Las Ciencias Exactas en el Siglo XVII, id. Revue d’Histoire des Sciences et de Leurs Applications, 9(1):93–94, January 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904702.

Itard:1956:AOBa

[567] Jean Itard. Analyses d’ouvrages: The Principal Works of Simon Stevin, Vol. I. Mechanics par E. J. Dyksterhuis. Revue d’Histoire des Sciences et de Leurs Applications, 9(1):95–96, January 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904703.

Rochot:1956:NTE

[568] B. Rochot. Sur les notions de temps et d’espace chez quelques auteurs du XVII^{e} siècle, notamment Gassendi et Barrow. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):97–104, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904975.

Boas:1956:MSR

[569] Marie Boas. La méthode scientifique de Robert Boyle. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):105–125, avril–juin 1956. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904976.

Huard:1956:TMS

[570] P. Huard and M. Durand. Un traité de médecine sino-vietnamienne du XVIII^{e} siècle : La compréhension intuitive des recettes médicales de Hai-Thuong. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):126–149, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904977.

Birembaut:1956:LDP

[571] Arthur Birembaut. Sur les lettres du physicien Magellan conservées aux archives nationales. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):150–161, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904978.

Rochot:1956:LSS

[572] B. Rochot. L’histoire des sciences à la Société d’Étude du XVII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):162–165, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904979.

Speziali:1956:MID

[573] Pierre Speziali. Manuscrits inédits du XVIII^{e} siècle (fonds genevois). Revue d’Histoire des Sciences et de Leurs Applications, 9(2):165–168, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904980.

Taton:1956:GS

[574] René Taton. George Sarton (1884–1956). Revue d’Histoire des Sciences et de Leurs Applications, 9(2):168–171, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904981.

Delorme:1956:JJV

[575] S. Delorme. A. J. J. Van de Velde (1871–1956). Revue d’Histoire des Sciences et de Leurs Applications, 9(2):171, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23910474.

Anonymous:1956:Ea

[576] Anonyme. Espagne. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):172, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23910475.

Taton:1956:FJ

[577] R. Taton. Frans Jonckheere (1903–1956). Revue d’Histoire des Sciences et de Leurs Applications, 9(2):172, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904982.

Anonymous:1956:Fb

[578] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):173–176, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23910476.

Anonymous:1956:PB

[579] Anonyme. Pays-Bas. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):176, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23910477.

Anonymous:1956:S

[580] Anonyme. Suisse. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):176, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23910478.

Delorme:1956:AOBa

[581] Suzanne Delorme. Analyses d’ouvrages: Histoire des Sciences par S. F. Mason; Marguerite Vergnaud. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):177–178, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904983.

Koyre:1956:AOBa

[582] A. Koyré. Analyses d’ouvrages: Le système du Monde, Histoire des doctrines cosmologiques, de Platon à Copernic, t. vi par Pierre Duhem. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):178–179, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904984.

Koyre:1956:AOBb

[583] Alexandre Koyré. Analyses d’ouvrages: Galileo as a Critic of the Arts par Erwin Panofsky. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):179–181, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904985.

Koyre:1956:AOBc

[584] A. Koyré. Analyses d’ouvrages: Discours de la méthode plus La dioptrique, Les météores, et La géométrie, (Les classiques de la science) par René Descartes; G. G. Sliousarev; A. P. Youschkevitch. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):181–183, avril–juin 1956. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904986.

Lenoble:1956:AOB

[585] R. Lenoble. Analyses d’ouvrages: Centre international de Synthèse, Pierre Gassendi, 1592–1655, sa vie et son œuvre. (La vie, le caractère et la formation intellectuelle, Le savant, Le philosophe, L’influence sur le milieu contemporain, L’influence posthume) par Bernard Rochot; Alexandre Koyré; Georges Mongrédien; Antoine Adam. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):183–185, avril–juin 1956. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904987.

Daumas:1956:AOB

[586] Maurice Daumas. Analyses d’ouvrages: Histoire de la chimie par P. Walden; Eugène Darmois. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):185–186, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904988.

Bedel:1956:AOB

[587] Charles Bedel. Analyses d’ouvrages: Lavoisier théoricien et expérimentateur, ≪Bibliothèque de philosophie contemporaine ^{≫}par Maurice Daumas. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):186–187, avril– juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904989.

Theodorides:1956:AOBa

[588] J. Théodoridès. Analyses d’ouvrages: Histoire naturelle. Livre IX par Pline l’Ancien; E. de Saint-Denis. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):187, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904990.

Theodorides:1956:AOBb

[589] J. Théodoridès. Analyses d’ouvrages: Apes and Ape lore in the Middle Ages and the Renaissance par H. W. Janson. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):187–188, avril–juin 1956. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904991.

Theodorides:1956:AOBc

[590] J. Théodoridès. Analyses d’ouvrages: Maîtres Liégeois de l’Illustration Scientifique. ^{≪}Bibliotheca Universitatis Leodiensis ^{≫}. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):189, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904992.

Koyre:1956:AOBd

[591] A. Koyré. Analyses d’ouvrages: Istoriko-matematitchesikja Isledovanija (Recherches historico-mathématiques), vol. VI, vol. VII, vol. VIII. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):189–191, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904993.

Colnort:1956:AQPa

[592] S. Colnort. Analyses de quelques périodiques: Science et Nature, par la photographie et par l’image, première année, 1954, n^{os} 1 à 6. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):191–192, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904994.

Colnort:1956:AQPb

[593] Suzanne Colnort. Analyses de quelques périodiques: Centaurus, International Magazine of the History of Science and Medecine. vol. III. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):192, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904995.

Lacoin:1956:GSM

[594] M. Lacoin. Sur la gestation de la science moderne (XV^{e} et XVI^{e} siècles). Revue d’Histoire des Sciences et de Leurs Applications, 9(3): 193–207, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905193.

Costabel:1956:FLA

[595] Pierre Costabel. La ^{≪}loi admirable ^{≫}de Christian Huygens. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):208–220, juillet– septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905194.

Rostand:1956:HIR

[596] Jean Rostand. Sur l’histoire des idées relatives à la parthénogenèse dans l’espèce humaine. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):221–235, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905195.

Dulieu:1956:JPR

[597] Louis Dulieu. Jacques-Philippe-Raymond Draparnaud. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):236–258, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905196.

Astruc:1956:COP

[598] P. Astruc. Cabanis, Œuvres philosophiques. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):259–266, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905197.

Anonymous:1956:Eb

[599] Anonyme. Espagne. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):266, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905198.

Delorme:1956:F

[600] S. Delorme. France. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):266–269, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905199.

Koyre:1956:AOBe

[601] A. Koyré. Analyses d’ouvrages: Izbrannye estestvennonautchnye proïzvedenijà (= Choix d’écrits se rapportant aux sciences de la nature) par Leonardo da Vinci; V. P. Zoubov. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):270–271, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905200.

Gloden:1956:AOB

[602] Albert Gloden. Analyses d’ouvrages: Sviluppo Storico dell’Aritmetica Generale e dell’Algebra par Alpinolo Natucci. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):271–273, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905201.

Itard:1956:AOBb

[603] Jean Itard. Analyses d’ouvrages: I ^{≪}Ragionamenti d’Algebra ^{≫}di R. Canacci (dal Codice Palatino 567 della Biblioteca Nazionale di Firenze) par Angiolo Procissi. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):273–274, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905202.

Koyre:1956:AOBf

[604] Alexandre Koyré. Analyses d’ouvrages: Traité d’Optique. ^{≪}Les maîtres de la pensée scientifique ^{≫} par Isaac Newton. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):274–275, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905203.

Colnort:1956:AOBa

[605] S. Colnort. Analyses d’ouvrages: Ludwig Boltzmann. Mensch, Physiker, Philosoph, (in Leinen DM. 11) par Englebert Broda; Hans Thirring. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):275, juillet– septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905204.

Huan:1956:AOB

[606] Nguyen Tran Huan. Analyses d’ouvrages: Kou Tai Tsi Ping Ming Heou Chou I (= Explication des termes médicaux de l’ancien temps) par Iu Yun Siou. Revue d’Histoire des Sciences et de Leurs Applications, 9(3): 275–276, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905205.

Huard:1956:AOB

[607] P. Huard. Analyses d’ouvrages: ^{≪}Malaria in Chinese medicine ^{≫}, Sinologica par R. Hoeppli. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):276, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905206.

Wickersheimer:1956:AOBa

[608] Ernest Wickersheimer. Analyses d’ouvrages: The surgery … ca. a. D 1267, Vol. I (books I and II). (History of medicine series issued under the auspices of the New York Academy of Medicine, vol. 12) par Theodoric; Eldridge Campbell; James Colton. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):277–278, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905207.

Wickersheimer:1956:AOBb

[609] Ernest Wickersheimer. Analyses d’ouvrages: From witchcraft to antisepsis : a study in antithesis … (Logan Clendening lectures on the history and philosophy of medicine, 5th series) par Douglas Guthrie. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):278–279, juillet– septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905208.

Wickersheimer:1956:AOBc

[610] Ernest Wickersheimer. Analyses d’ouvrages: Forgotten leaders in modern medicine, Valentin, Gruby, Remak, Auerbach. (Transactions of the American Philosophical Society, n. s., vol. 44, part. 2) par Bruno Kisch. Revue d’Histoire des Sciences et de Leurs Applications, 9(3): 279–280, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905209.

Chauvois:1956:AOB

[611] L. Chauvois. Analyses d’ouvrages: J. M. Charcot (1825–1893). Savie, son œuvre par Georges Guillain. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):280–282, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905210.

Delaunay:1956:AOB

[612] A. Delaunay. Analyses d’ouvrages: Monsieur Roux par Émile Lagrange. Revue d’Histoire des Sciences et de Leurs Applications, 9(3): 282–283, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905211.

Wickersheimer:1956:AOBd

[613] Ernest Wickersheimer. Analyses d’ouvrages: 25 portraits de médecins français, 1900–1950 par Édouard Rist. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):283–284, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905212.

Jacques:1956:AOB

[614] J. Jacques. Analyses d’ouvrages: Histoire d’une entreprise française par C. J. Gignoux. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):285–286, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905213.

Colnort:1956:AOBb

[615] S. Colnort. Analyses d’ouvrages: Bibliographie des dictionnaires scientifiques et techniques multilingues, Documentation et terminologie scientifique. Revue d’Histoire des Sciences et de Leurs Applications, 9(3): 286–287, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905214.

Colnort:1956:AQPc

[616] S. Colnort. Analyses de quelques périodiques: Lychnos, Annuaire de la Société suédoise d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):287, juillet–septembre 1956. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905215.

Colnort:1956:AQPd

[617] S. Colnort. Analyses de quelques périodiques: Archives internationales d’Histoire des Sciences, publication trimestrielle de l’Union internationale d’Histoire des Sciences, t. VIII, (n^{os} 30 à 33), 1955. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):287–288, juillet– septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905216.

Bernal:1956:RSE

[618] J. D. Bernal. Les rapports scientifiques entre la Grande-Bretagne et la France au XVIII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):289–300, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903793.

Poulle:1956:PDA

[619] E. Poulle. Peut-on dater les astrolabes médiévaux? Revue d’Histoire des Sciences et de Leurs Applications, 9(4):301–322, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903794.

deDainville:1956:LGM

[620] François de Dainville. L’instruction des Gardes de la Marine à Brest en 1692. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):323– 338, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903795.

Torlais:1956:GCS

[621] Jean Torlais. Une grande controverse scientifique au XVIII^{e} siècle: L’abbé Nollet et Benjamin Franklin. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):339–349, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903796.

Leroy:1956:T

[622] Jean-F. Leroy. Tournefort (1656–1708). Revue d’Histoire des Sciences et de Leurs Applications, 9(4):350–354, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903797.

Bouthillon:1956:LD

[623] Léon Bouthillon. L’invention de la diode. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):354–356, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903798.

Anonymous:1956:Fc

[624] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):357–359, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903799.

Costabel:1956:IVC

[625] Pierre Costabel. Italie: Viii^{e} Congrès International d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):360– 363, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903800.

Koyre:1956:AOBg

[626] Alexandre Koyré. Analyses d’ouvrages: Correspondance du P. Marin Mersenne, religieux minime, vol. IV (1634) par Paul Tannery; Cornélis de Waard. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):364–366, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903801.

Taton:1956:AOB

[627] R. Taton. Analyses d’ouvrages: The Thirteen Books of Euclid’s Elements, second edition revised with additions par Thomas L. Heath. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):366, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903802.

Itard:1956:AOBc

[628] Jean Itard. Analyses d’ouvrages: Archimedes. (^{≪}Acta historica scientiarum naturalium et medicinalium ^{≫}, vol. 12) par E. J. Dijksterhuis.

Revue d’Histoire des Sciences et de Leurs Applications, 9(4):366–371, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903803.

Dugas:1956:AOB

[629] René Dugas. Analyses d’ouvrages: Le procès de Galilée ^{≪}Dossiers de l’Histoire ^{≫} par Giorgio de Santillana. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):371–372, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903804.

Delorme:1956:AOBb

[630] Suzanne Delorme. Analyses d’ouvrages: L’atomisme en biologie par Jean Rostand. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):372–373, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903805.

Wickersheimer:1956:AOBe

[631] Ernest Wickersheimer. Analyses d’ouvrages: A History of the Therapy of Tuberculosis and the Case of Frederic Chopin … (Logan Clendening Lectures on the History and Philosophy of Medicine, 6th series) par Esmond R. Long. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):373–375, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903806.

Colnort:1956:AQPe

[632] S. Colnort. Analyses de quelques périodiques: Isis, an international Review devoted to the History of Science and its cultural influence, vol. 46 par Bernard Cohen. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):375, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903807.

Colnort:1956:AQPf

[633] S. Colnort. Analyses de quelques périodiques: Gesnerus, revue trimestrielle, publiée par la Société suisse d’Histoire de la Médecine et des Sciences naturelles, vol. 12 (fasc. 1–2 et 3–4). Revue d’Histoire des Sciences et de Leurs Applications, 9(4):376, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903808.

Anonymous:1956:TGM

[634] Anonyme. Table générale des matières du Tome IX. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):377–384, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903809.

Anonymous:1956:BMa

[635] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 9(1):??, January 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904704.

Anonymous:1956:BMb

[636] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):??, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904996.

Anonymous:1956:BMc

[637] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):??, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905217.

Anonymous:1956:BMd

[638] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):??, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903810.

Anonymous:1956:FMa

[639] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 9(1):??, January 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904693.

Anonymous:1956:FMb

[640] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 9(2):??, avril–juin 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904974.

Anonymous:1956:FMc

[641] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 9(3):??, juillet–septembre 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905192.

Anonymous:1956:FMd

[642] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 9(4):??, October 1956. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903792.

Huan:1957:EHB

[643] Nguyen Tran Huan. Esquisse d’une histoire de la biologie chinoise des origines jusqu’au IV^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 10(1):1–37, January 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904786.

Lacoarret:1957:TFO

[644] Marie Lacoarret. Les traductions françaises des œuvres d’Euclide. Revue d’Histoire des Sciences et de Leurs Applications, 10(1):38–58, January 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904787.

Itard:1957:LRL

[645] Jean Itard. Les lois de la réfraction de la lumière chez Kepler. Revue d’Histoire des Sciences et de Leurs Applications, 10(1):59–68, January 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904788.

Rochot:1957:GM

[646] B. Rochot. Gassendi et les mathématiques. Revue d’Histoire des Sciences et de Leurs Applications, 10(1):69–78, January 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904789.

Dulieu:1957:CCA

[647] Louis Dulieu. Un concours de chimie au XVIII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 10(1):79–85, January 1957. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904790.

Rostand:1957:CDA

[648] Jean Rostand. Casimir Davaine et les antibiotiques. Revue d’Histoire des Sciences et de Leurs Applications, 10(1):86–87, January 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904791.

Anonymous:1957:Fa

[649] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 10(1):87–92, January 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904792.

Anonymous:1957:SCH

[650] Anonyme. Suisse: Cours d’histoire des sciences. Revue d’Histoire des Sciences et de Leurs Applications, 10(1):92, January 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904793.

Delorme:1957:AOBa

[651] Suzanne Delorme. Analyses d’ouvrages: Arnold Reymond, ^{≪}Philosophes d’aujourd’hui ^{≫}par A. Virieux-Reymond; R. Blanché; G. Widemer; F. Brunner. Revue d’Histoire des Sciences et de Leurs Applications, 10(1): 93–94, January 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904794.

Colnort:1957:AOBa

[652] Suzanne Colnort. Analyses d’ouvrages: Dictionnaire des sciences mathématiques, mécanique, cosmographie, physique, chimie par E. B. Uvarov; D. R. Chapman; Jacques d’Hermies; Marie-Élisa Cohen; René Taton. Revue d’Histoire des Sciences et de Leurs Applications, 10(1):94, January 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904795.

Itard:1957:AOBa

[653] Jean Itard. Analyses d’ouvrages: Science Awakening par B. L. Van der Waerden; Arnold Dresden. Revue d’Histoire des Sciences et de Leurs Applications, 10(1):94–95, January 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904796.

Taton:1957:AOBa

[654] R. Taton. Analyses d’ouvrages: Grandes découvertes du XX^{e} siècle par Louis Leprince-Ringuet. Revue d’Histoire des Sciences et de Leurs Applications, 10(1):96, January 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904797.

Zoubov:1957:FIF

[655] V. Zoubov. Les ^{≪}indivisibles ^{≫}et le continu dans l’ancienne littérature russe. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):97– 109, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905485.

deVirville:1957:IIP

[656] Ad. Davy de Virville. De l’influence des idées préconçues sur les progrès de la botanique du XV^{e} au XVIII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):110–119, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905486.

Costabel:1957:SRD

[657] Pierre Costabel. La septième règle du choc élastique de Christian Huygens. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):120– 131, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905487.

Boas:1957:QAS

[658] Marie Boas. Quelques aspects sociaux de la chimie au XVII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):132–147, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905488.

Birembaut:1957:LRS

[659] Arthur Birembaut. L’Académie royale des Sciences en 1780 vue par l’astronome suédois Lexell (1740–1784). Revue d’Histoire des Sciences et de Leurs Applications, 10(2):148–166, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905489.

Tresse:1957:JCE

[660] René Tresse. J. A. Chaptal et l’enseignement technique de 1800 à 1819. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):167–174, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905490.

Rostand:1957:PDP

[661] Jean Rostand. À propos du pharmacien Coladon précurseur possible de Mendel. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):175–176, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905491.

VandeVelde:1957:BCC

[662] A. J. J. Van de Velde. Belgique: Cérémonie commémorative. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):176–177, avril– juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905492.

Anonymous:1957:Fb

[663] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):177–179, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905493.

Anonymous:1957:PB

[664] Anonyme. Pays-Bas. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):179, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905494.

Huard:1957:AOBa

[665] P. Huard and M. Wong. Analyses d’ouvrages: Science and Civilisation in China, t. I, 1954, t. II, 1956 par J. Needham; Wang Ling. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):180–182, avril– juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905495.

Huan:1957:AOB

[666] Nguyen Tran Huan. Analyses d’ouvrages: Mong K’i Pi T’an Kiao Tcheng. Shanghaï, 1re éd. par Chen K’ouo; Hou Tao Tsing. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):182–183, avril– juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905496.

Taton:1957:AOBb

[667] R. Taton. Analyses d’ouvrages: A check list of the E. De Golyer Collection in the History of Science and Technology as of August 1, 1954 par Arthur McAnally; Duane H. D. Roller. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):183, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905497.

Taton:1957:AOBc

[668] R. Taton. Analyses d’ouvrages: Gnomonique ou traité théorique et pratique de la construction des cadrans solaires suivi de tables auxiliaires relatives aux cadrans et aux calendriers par G. Bigourdan. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):183–184, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905498.

Taton:1957:AOBd

[669] René Taton. Analyses d’ouvrages: Évariste Galois révolutionnaire et géomètre, (^{≪}Librairie des Libelles ^{≫}) par André Dalmas. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):184–187, avril– juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905499.

Itard:1957:AOBb

[670] Jean Itard. Analyses d’ouvrages: Henri Poincaré ou la vocation souveraine, (collection ^{≪}Vocations ^{≫}) par André Bellivier; Henri Mondor. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):187, avril– juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905500.

Astruc:1957:AOB

[671] P. Astruc. Analyses d’ouvrages: Recherches physiologiques sur la vie et la mort par Xavier Bichat. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):188–189, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905501.

Dulieu:1957:AOB

[672] L. Dulieu. Analyses d’ouvrages: Matricule de l’Université de Médecine de Montpellier (1503–1599) par Marcel Gouron. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):189–190, avril–juin 1957. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905502.

Huard:1957:AOBb

[673] P. Huard and J. Brocas. Analyses d’ouvrages: ^{≪}Le centenaire de Sir d’Arcy Power, B.E.F.R.C.S. (1852–1941) ^{≫}(The British Journal of Surgery, novembre 1955). Revue d’Histoire des Sciences et de Leurs Applications, 10(2):190–191, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905503.

Rostand:1957:AOB

[674] Jean Rostand. Analyses d’ouvrages: On the Genesis of the Ovum of Mammals and of Man par Karl Ernst von Baer; Charles Donald O’Malley; I. Bernard Cohen. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):191–192, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905504.

Huard:1957:AOBc

[675] P. Huard and M. Wong. Analyses d’ouvrages: Medical History par W. J. Bishop. Revue d’Histoire des Sciences et de Leurs Applications, 10(2): 192, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905505.

Brunold:1957:ESH

[676] Charles Brunold. Enseignement scientifique et histoire des sciences. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):193– 196, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904216.

Levey:1957:LLA

[677] Martin Levey. L’argent dans la littérature ancienne de Mésopotamie. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):197– 204, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904217.

Vernant:1957:RFL

[678] J.-P. Vernant. Remarques sur les formes et les limites de la pensée technique chez les grecs. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):205–225, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904218.

Auger:1957:IRS

[679] Léon Auger. Les idées de Roberval sur le système du monde. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):226–234, juillet– septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904219.

Lenoble:1957:RFE

[680] Robert Lenoble. Roberval ^{≪}éditeur ^{≫}de Mersenne et du P. Niceron. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):235– 254, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904220.

Speziali:1957:AFS

[681] Pierre Speziali. Augustin Fresnel et les savants genevois. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):255–259, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904221.

Rychlik:1957:MCA

[682] Karel Rychlík. Un manuscrit de Cauchy aux archives de l’Académie tchécoslovaque des sciences. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):259–261, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904222.

Huard:1957:HS

[683] Pierre Huard. Henry Sigerist (1891–1957). Revue d’Histoire des Sciences et de Leurs Applications, 10(3):261–263, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904223.

Taton:1957:RD

[684] René Taton. René Dugas (1897–1957). Revue d’Histoire des Sciences et de Leurs Applications, 10(3):263–264, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904224.

Costabel:1957:F

[685] Pierre Costabel and S. D. France. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):264–270, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904225.

Anonymous:1957:I

[686] Anonyme. Italie. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):270–271, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904226.

Speziali:1957:S

[687] Pierre Speziali. Suisse. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):271–273, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904227.

Anonymous:1957:URB

[688] Anonyme. U.S.A.: Réapparition de Chymia. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):273, juillet–septembre 1957. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904228.

Beaujouan:1957:AOB

[689] G. Beaujouan. Analyses d’ouvrages: The Medieval Science of Weights (Scientia de ponderibus): treatises ascribed to Euclid, Archimedes, Thabit ibn Qurra, Jordanus de Nemore and Blasius of Parma edited with introduction, English translation and notes par Ernest A. Moody; Marshall Clagett. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):274–276, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904229.

Hooykaas:1957:AOB

[690] R. Hooykaas. Analyses d’ouvrages: Vom Wege Niels Stensens. Beiträge zu seiner naturwissenschaftlichen Entwicklung ^{≪}Acta Historica Scient. Nat. et Med., ^{≫}, vol. 14 par Gustav Scherz. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):276–279, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904230.

Huard:1957:AOBd

[691] P. Huard and J. Brocas. Analyses d’ouvrages: A History of Urology in Egypt par J. Bitschai; Léopold Brodny. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):279–281, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904231.

Birembaut:1957:AOBa

[692] Arthur Birembaut. Analyses d’ouvrages: Contribution à l’étude des origines de la droguerie pharmaceutique et de l’industrie du sel ammoniac en France : l’apothicaire Antoine Baumé (1728–1804) par René Davy. Revue d’Histoire des Sciences et de Leurs Applications, 10(3): 281–284, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904232.

Colnort:1957:AOBb

[693] S. Colnort. Analyses d’ouvrages: La psychanalyse. Son évolution, ses développements, ^{≪}Bibliothèque des Idées ^{≫}par Clara Thompson; Patrick Murray; André Green. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):284–285, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904233.

Drieux:1957:AOB

[694] Y. Drieux. Analyses d’ouvrages: Le cheval dans l’histoire par M.-F. de Genellis; Donnio. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):285–286, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904234.

Birembaut:1957:AOBb

[695] Arthur Birembaut. Analyses d’ouvrages: Engineering in History par Richard Shelton Kirby; Sidney Withington; Arthur Burr Darling; Frederick Gridley Kilgour. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):286–287, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904235.

Colnort:1957:AQP

[696] S. Colnort. Analyses de quelques périodiques: Gesnerus, revue trimestrielle, publiée par la Société suisse d’Histoire de la Médecine et des Sciences naturelles, vol. 13. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):287–288, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904236.

Anonymous:1957:PR

[697] Anonyme. Publications reçues. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):288, juillet–septembre 1957. CODEN RHSAAM.

ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904237.

Delorme:1957:TCV

[698] Suzanne Delorme. Tableau chronologique de la vie et des œuvres de Fontenelle: avec les principaux synchronismes littéraires philosophiques et scientifiques. Revue d’Histoire des Sciences et de Leurs Applications, 10(4):289–299, October 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904018.

Delorme:1957:CBF

[699] Suzanne Delorme. Contribution à la bibliographie de Fontenelle. Revue d’Histoire des Sciences et de Leurs Applications, 10(4):300–309, October 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904019.

Martin:1957:RAP

[700] Geneviève Martin. Retouches au portrait de Fontenelle: Pièces inédites. Revue d’Histoire des Sciences et de Leurs Applications, 10(4):310–333, October 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904020.

McKie:1957:FSR

[701] Douglas McKie. Fontenelle et la Société Royale de Londres. Revue d’Histoire des Sciences et de Leurs Applications, 10(4):334–338, October 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904021.

Delorme:1957:BIS

[702] Suzanne Delorme. La “Géométrie de l’infini” et ses commentateurs de Jean Bernoulli à M. de Cury. Revue d’Histoire des Sciences et de Leurs Applications, 10(4):339–359, October 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904022.

Birembaut:1957:FG

[703] Arthur Birembaut. Fontenelle et la géologie. Revue d’Histoire des Sciences et de Leurs Applications, 10(4):360–374, October 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904023.

Delorme:1957:AOBb

[704] Suzanne Delorme. Analyses d’ouvrages: Entretiens sur la pluralité des mondes. Digression sur les Anciens et les Modernes par Fontenelle; Robert Shackleton. Revue d’Histoire des Sciences et de Leurs Applications, 10(4):375–377, October 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904024.

Anonymous:1957:TGM

[705] Anonyme. Table générale des matières du Tome X. Revue d’Histoire des Sciences et de Leurs Applications, 10(4):378–384, October 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904025.

Anonymous:1957:BMa

[706] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 10(1):??, January 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904798.

Anonymous:1957:BMb

[707] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):??, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905506.

Anonymous:1957:BMc

[708] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):??, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904238.

Anonymous:1957:BMd

[709] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 10(4):??, October 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904026.

Anonymous:1957:FMa

[710] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 10(1):??, January 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904785.

Anonymous:1957:FMb

[711] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 10(2):??, avril–juin 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905484.

Anonymous:1957:FMc

[712] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 10(3):??, juillet–septembre 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904215.

Anonymous:1957:FMd

[713] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 10(4):??, October 1957. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904017.

Torlais:1958:CVO

[714] Jean Torlais. Chronologie de la vie et des œuvres de René-Antoine Ferchault de Réaumur. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):1–12, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905046.

Torlais:1958:RP

[715] Jean Torlais. Réaumur philosophe. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):13–33, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23905047.

Rostand:1958:REG

[716] Jean Rostand. Réaumur embryologiste et généticien. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):34–50, January 1958. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905048.

Torlais:1958:RHA

[717] Jean Torlais. Réaumur et l’histoire des abeilles. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):51–67, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905049.

Speziali:1958:RSG

[718] Pierre Speziali. Réaumur et les savants genevois: Lettres inédites. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):68–80, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905050.

deVirville:1958:RM

[719] A. Davy de Virville. Réaumur dans la Mayenne. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):81–82, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905052.

Birembaut:1958:FRG

[720] Arthur Birembaut. Fontenelle, Réaumur et le gaz naturel. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):82–84, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905053.

Theodorides:1958:Fa

[721] Jean Théodoridès, P. Costabel, and Lise Dubief. France. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):84–87, January 1958. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905054.

Anonymous:1958:IEM

[722] Anonyme. Italie: Exposition à Milan. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):87, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905055.

Anonymous:1958:ACW

[723] Anonyme. Autriche: Coronelli Weltbund Der Globusfreude. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):87–88, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905056.

Anonymous:1958:PBSa

[724] Anonyme. Pays-Bas: Société Pour L’Histoire de La Médecine, des Mathématiques et des Sciences de la Nature. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):88, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905057.

Anonymous:1958:S

[725] Anonyme. Suisse. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):88–89, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905058.

Astruc:1958:AOBa

[726] P. Astruc. Analyses d’ouvrages: Histoire de la Médecine moderne, trad par Richard-Harrison Shryock; Raissa Tarr. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):90–92, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905059.

Huard:1958:AOBa

[727] P. Huard. Analyses d’ouvrages: ^{≪}The man who was Vidius (1508–1569) ≫. Thomas Vicary lectual, Annals of the Royal College of Surgeons of England, 1956, vol. 19 par W. Brocbank. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):92–93, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905060.

Daumas:1958:AOBa

[728] Maurice Daumas. Analyses d’ouvrages: Histoire des techniques ^{≪}Les grandes études historiques ^{≫} par Pierre Rousseau. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):93–95, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905061.

Colnort:1958:AOBa

[729] Suzanne Colnort. Analyses d’ouvrages: Science and Economic Development : New Patterns of Living par Richard L. Meier. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):96, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905063.

Colnort:1958:AOBb

[730] Suzanne Colnort. Analyses d’ouvrages: The Carlsberg Foundation. The Carlsberg Laboratory, Scientific Grants, The Museum of National History at Frederiksborg Castle, The New Carlsberg Foundation par Johannes Pedersen. Revue d’Histoire des Sciences et de Leurs Applications, 11(1): 96, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905064.

Russo:1958:AOBa

[731] F. Russo. Analyses d’ouvrages: 1894–1954 : Soixantenaire de l’École supérieure d’Électricité. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):96, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905062.

Brunold:1958:RHE

[732] Charles Brunold. Rôle de l’histoire dans l’enseignement des sciences physiques. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):97–107, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905126.

Moscovici:1958:NBM

[733] Serge Moscovici. Notes sur le “De motu tractatus” de Michel Varro. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):108–129, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905127.

Taton:1958:RM

[734] René Taton. Réaumur mathématicien. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):130–133, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905128.

deVirville:1958:RB

[735] Ad. Davy de Virville. Réaumur botaniste. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):134–137, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905129.

Birembaut:1958:REP

[736] Arthur Birembaut. Réaumur et l’élaboration des produits ferreux. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):138–166, avril– juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905130.

Birembaut:1958:LFE

[737] Arthur Birembaut. Les liens de famille entre Réaumur et Brisson, son dernier élève. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):167–169, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905131.

Russo:1958:DHS

[738] F. Russo. La documentation en histoire des sciences et des techniques: La section ^{≪}histoire des sciences et des techniques ^{≫}du Bulletin signalétique du centre national de la recherche scientifique. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):169–171, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905132.

Delorme:1958:AR

[739] Suzanne Delorme. Arnold Reymond (1874–1958). Revue d’Histoire des Sciences et de Leurs Applications, 11(2):171–174, avril–juin 1958. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905133.

Costabel:1958:UIH

[740] P. Costabel. Union Internationale d’Histoire et de Philosophie des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):174– 175, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905134.

Theodorides:1958:Fb

[741] J. Théodoridès. France. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):175–178, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905136.

Rochot:1958:AOB

[742] B. Rochot. Analyses d’ouvrages: Correspondance, t. VI (1956) par Descartes; Charles Adam; Gérard Milhaud. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):179–180, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905143.

Bauer:1958:AOB

[743] Edmond Bauer. Analyses d’ouvrages: Franklin and Newton, Philadelphie, (^{≪}Memoirs of the American Philosophical Society ^{≫}, vol. 45) par I. Bernard Cohen. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):180–182, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905144.

Huard:1958:AOBb

[744] P. Huard. Analyses d’ouvrages: Science History and its recent losses (Military Medecine, février 1957, t. 42) par Cl. F. Mayer. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):182–184, avril– juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905145.

Poulle:1958:AOBa

[745] Emmanuel Poulle. Analyses d’ouvrages: Islamic astrolabists and their works par L. A. Meyer. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):184–185, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905146.

Birembaut:1958:AOBa

[746] Arthur Birembaut. Analyses d’ouvrages: Alchemy par E. J. Holmyard. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):185–186, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905147.

Hooykaas:1958:AOT

[747] R. Hooykaas. Analyses d’ouvrages: Torbern Bergman, On Acid of Air (Excerpt from KYA 1773); Treatise on Bitter, Seltzer, Spa and Pyrmont Waters and their Synthetical Preparation (Excerpt from KYA 1775). Uno Boklund, Torbern Bergman as Pioneer in the Domain of Mineral Waters, Stockholm, Almquist & Wiksell, 1956 Torbern Bergman, Om luftsyra (Utdrag ur KVA 1773); Afhandling om Bitter-Selzer-Spa-och Pyrmonter-Valten, samt deras tilredande genom konst (Utdrag ur KVA 1775). Uno Boklund, Torbern Bergman som pionär på mineralvattenområdet, Stockholm, Almquist & Wiksell, 1956. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):186–187, avril–juin 1958. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905148.

Birembaut:1958:AOBb

[748] Arthur Birembaut. Analyses d’ouvrages: Gruvbrytning och kopparhantering vid Stora Kopparberg intill 1800-talets början (= Abatage et traitement du cuivre au Stora Kopparberg jusqu’au début du XIX^{e} siècle), t. I : Gruvan och gruvbrytningen (= La mine et l’abatage), t. II : Kopparhanteringen (= Le traitement du cuivre) par Sten Lindroth. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):187–188, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905149.

Astruc:1958:AOBb

[749] P. Astruc. Analyses d’ouvrages: La médecine hippocratique par Charles Lichtenthaeler. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):188–189, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905150.

Huard:1958:AOBc

[750] P. Huard. Analyses d’ouvrages: Un anatomiste de la Renaissance : Léonard de Vinci par Marc Imbert. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):189, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905151.

Huard:1958:AOBd

[751] P. Huard and M. Wong. Analyses d’ouvrages: Le livre de la génération du foetus et le traitement des femmes enceintes et des nouveau-nés par H. Jahier; Abdelkader Noureddine. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):189–191, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23905152.

Russo:1958:AOBb

[752] F. Russo. Analyses d’ouvrages: Les travaux d’histoire locale de la pharmacie des origines à ce jour par Maurice Bouvet. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):191, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905153.

Colnort:1958:AOBc

[753] Suzanne Colnort. Analyses d’ouvrages: L’aube tragique de l’anesthésie, Coll. ^{≪}Pourquoi? Comment? ^{≫} par Jean Fiolle; Monique Fiolle. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):191–192, avril– juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905154.

Russo:1958:AOBc

[754] F. Russo. Analyses d’ouvrages: Précurseurs et inventeurs. Livres et autographes scientifiques et médicaux. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):192, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905155.

Senior:1958:QSM

[755] Armand Machabey Senior. Quelques savants-musiciens de l’époque de Mersenne. Revue d’Histoire des Sciences et de Leurs Applications, 11(3): 193–206, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903606.

Martin:1958:DLR

[756] Geneviève Martin. Documents de l’Académie de Rouen concernant l’enseignement des sciences au XVIII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):207–226, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903607.

Dulieu:1958:MSM

[757] Louis Dulieu. Le mouvement scientifique montpelliérain au XVIII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 11(3): 227–249, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903608.

Dulieu:1958:CMR

[758] Louis Dulieu. La contribution montpelliéraine aux recueils de l’Académie Royale des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):250–262, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903609.

Gloden:1958:LSA

[759] Albert Gloden. L’enseignement des sciences à l’Ancien Collège de Luxembourg au XVIII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):263–266, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903610.

Birembaut:1958:PPR

[760] Arthur Birembaut. À propos d’une publication récente sur Lavoisier et le Lycée des Arts. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):267–273, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903611.

Daumas:1958:JAF

[761] Maurice Daumas. Justification de l’attitude de Fourcroy pendant la terreur. Revue d’Histoire des Sciences et de Leurs Applications, 11(3): 273–274, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903612.

Cahen:1958:ROR

[762] Louis Cahen. Retour aux origines dans les radio-communications. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):274–276, juillet– septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903613.

Anonymous:1958:F

[763] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):276–277, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903614.

Anonymous:1958:PCI

[764] Anonyme. Portugal: Congrès International de l’Histoire des Découvertes. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):277– 278, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903615.

Anonymous:1958:SSH

[765] Anonyme. Suisse: Société Helvétique pour l’Histoire de la Médécine et des Sciences Naturelles. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):278, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903616.

Taton:1958:AOBa

[766] R. Taton. Analyses d’ouvrages: Sciences et enseignement. D’hier à aujourd’hui, Paris, Centre d’Études pédagogiques par F. de Dainville. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):279–280, juillet– septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903617.

Daumas:1958:AOBb

[767] M. Daumas. Analyses d’ouvrages: Lives in Science. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):280, juillet–septembre 1958. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903618.

Daumas:1958:AOBc

[768] M. Daumas. Analyses d’ouvrages: The Science Museum. The First Hundred Years. Revue d’Histoire des Sciences et de Leurs Applications, 11(3): 280–281, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903619.

Poulle:1958:AOBb

[769] Emmanuel Poulle. Analyses d’ouvrages: Deutsche und niederländische astronomische Instrumente des 11.-18. Jahrhunderts par Ernst Zinner. Revue d’Histoire des Sciences et de Leurs Applications, 11(3): 281–282, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903620.

Itard:1958:AOBa

[770] Jean Itard. Analyses d’ouvrages: Un mathématicien de génie. Pierre de Fermat (1601–1665). Revue d’Histoire des Sciences et de Leurs Applications, 11(3):282–283, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903621.

Theodorides:1958:AOB

[771] Jean Théodoridès. Analyses d’ouvrages: A History of Luminescence from the Earliest Times until 1900 par E. Newton Harvey. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):283–284, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903622.

Huard:1958:AOBe

[772] P. Huard. Analyses d’ouvrages: Anniversaires médicaux en 1957, Progrès médical par G. Nicole-Genty. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):284, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903623.

Huan:1958:AOB

[773] Nguyen Tran Huan. Analyses d’ouvrages: Tchong Kouo I Hio che [Histoire de la Médecine chinoise] par Tchen Pang Hien. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):284–286, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903624.

Wickersheimer:1958:AOB

[774] Ernest Wickersheimer. Analyses d’ouvrages: Un prince, deux préfets; le mouvement scientifique et médico-social au pays de Liège sous le despotisme éclairé (1771–1830) (Médecine et médecins au pays de Liège, II) par Marcel Florkin. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):286–288, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903625.

Robinet:1958:LCO

[775] André Robinet. L’abbé de Catelan: ou l’erreur au service de la vérité. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):289–301, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904880.

Birembaut:1958:CRT

[776] Arthur Birembaut. La contribution de Réaumur à la thermométrie. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):302–329, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904881.

Proust:1958:DEH

[777] Jacques Proust. Deux encyclopédistes hors de l’ “Encyclopédie” : Philippe Macquer et l’abbé Jaubert. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):330–336, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904882.

Duveen:1958:DLL

[778] Denis I. Duveen and Roger Hahn. Deux lettres de Laplace à Lavoisier. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):337–342, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904883.

Kersaint:1958:SRM

[779] G. Kersaint. La Société royale de Médecine et l’éloignement de Necker. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):343–344, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904884.

Theodorides:1958:PHM

[780] Jean Théodoridès. À propos de Henri-Mamer-Onésime Delafond (1805– 1861). Revue d’Histoire des Sciences et de Leurs Applications, 11(4):344– 345, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904885.

Taton:1958:F

[781] R. Taton and Franck Bourdier. France. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):345–350, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904886.

Anonymous:1958:PBSb

[782] Anonyme. Pays-Bas: Société pour l’Histoire de la Médecine, des Mathématiques et des Sciences de la Nature. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):350–351, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904887.

Costabel:1958:YSI

[783] Pierre Costabel. Yougoslavie: Le Symposium International Roger Boscovitch (23–25 octobre 1958). Revue d’Histoire des Sciences et de Leurs Applications, 11(4):351–353, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904888.

Taton:1958:AOBb

[784] René Taton. Analyses d’ouvrages: The Appreciation of Ancient and Medieval Science during the Renaissance (1450–1600) par George Sarton; Six Wings. Men of Science in the Renaissance, illustrated with contemporary portraits par George Sarton. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):354–356, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904889.

Moscovici:1958:AOBa

[785] Serge Moscovici. Analyses d’ouvrages: From the closed World to the Infinite Universe par Alexandre Koyré. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):356–358, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904890.

Moscovici:1958:AOBb

[786] Serge Moscovici. Analyses d’ouvrages: Tricentenaire de Pierre Gassendi (1655–1955), Actes du Congrès (Digne, 4–7 août 1955). Revue d’Histoire des Sciences et de Leurs Applications, 11(4):358–360, October 1958. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904891.

Taton:1958:AOBc

[787] R. Taton. Analyses d’ouvrages: Geschichte der Mathematik, II. Von Fermat und Descartes bis zur Erfindung des Calculus und bis zum Ausbau der neuen Methoden; III. Von den Auseinandersetzungen um den Calculus bis zur französischen Revolution. (^{≪}Sammlung Göschen ^{≫}, Band 875 et Band 882) par Joseph E. Hofmann. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):360, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904892.

Taton:1958:AOBd

[788] R. Taton. Analyses d’ouvrages: Histoire des mathématiques par Oskar Becker; Jos.-E. Hofmann; G. Bouligand; R. Jouan. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):361, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904893.

Itard:1958:AOBb

[789] Jean Itard. Analyses d’ouvrages: Thomas of Bradwardine. His Tractatus de Proportionibus. Its Significance for the Development of Mathematical Physics par H. Lamar Crosby. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):361–362, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904894.

Taton:1958:AOBe

[790] René Taton. Analyses d’ouvrages: Théorie des nombres, 4^{e} éd. conforme à la 3^{e}, nouveau tirage corrigé, t. I; t. II par Adrien-Marie Legendre. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):363–364, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904895.

Taton:1958:AOBf

[791] R. Taton. Analyses d’ouvrages: Histoire de l’astronomie, 2^{e} éd., suivie de l’Astronomie moderne, coll. ^{≪}Histoire des Sciences ^{≫}par Friedrich Becker; F. Cusset; Ernest Esclangon. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):364, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904896.

Itard:1958:AOBc

[792] Jean Itard. Analyses d’ouvrages: Histoire de la lumière, ^{≪}Bibliothèque générale de l’École pratique des Hautes Études-VI^{e} section ^{≫}par Vasco Ronchi; Juliette Taton. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):365, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904897.

Birembaut:1958:AOBc

[793] Arthur Birembaut. Analyses d’ouvrages: Torbern Bergman: A Bibliography of his Works par Birgitta Moström. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):366, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904898.

Birembaut:1958:AOBd

[794] Arthur Birembaut. Analyses d’ouvrages: Georgius Agricola 1494– 1555, górnik, metalurg, mineralog, chemik, lekarz (mineur, métallurgiste, minéralogiste, chimiste, médecin). Revue d’Histoire des Sciences et de Leurs Applications, 11(4):366–367, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904899.

Birembaut:1958:AOBe

[795] Arthur Birembaut. Analyses d’ouvrages: Armand Dufrénoy (1792– 1857), Revue générale des Sciences pures et appliquées, juillet-août 1957, t. LXIV, n^{os} 7–8 par Jean Orcel. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):367–369, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904900.

Huard:1958:AOBf

[796] P. Huard. Analyses d’ouvrages: William Harvey, sa vie et son temps, ses découvertes, sa méthode par Louis Chauvois; Pr Lian; J. Rostand. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):370–371, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904901.

Huard:1958:AOBg

[797] P. Huard. Analyses d’ouvrages: Contribution à l’étude de la chirurgie dans l’Ouest de la France au XVIII^{e} siècle et principalement à Rennes d’après les travaux du Pr Hardouin, thèse de Rennes, 1957, n^{o} 15 par Pichevin-Chatelin. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):371–372, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904902.

Huard:1958:AOBh

[798] P. Huard. Analyses d’ouvrages: Mungo Park surgeon and explorer, Medical History par H. P. Tait. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):372, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904903.

Huard:1958:AOBi

[799] P. Huard. Analyses d’ouvrages: The historical significance of the cerebrospinal fluid, Medical History par David H. M. Woollam. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):372–375, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904904.

Huard:1958:AOBj

[800] P. Huard. Analyses d’ouvrages: D^{r} John Fothergill and the angina pectories, Medical History par C. C. Booth. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):375, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904905.

Colnort:1958:AQP

[801] S. Colnort. Analyses de quelques périodiques: Lychnos, Annuaire de la Société suédoise d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):375–376, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904906.

Anonymous:1958:TGM

[802] Anonyme. Table générale des matières du Tome XI. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):377–384, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904907.

Anonymous:1958:BMa

[803] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):??, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905065.

Anonymous:1958:BMb

[804] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):??, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905156.

Anonymous:1958:BMc

[805] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):??, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903626.

Anonymous:1958:BMd

[806] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):??, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904908.

Anonymous:1958:FMa

[807] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 11(1):??, January 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905045.

Anonymous:1958:FMb

[808] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 11(2):??, avril–juin 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905125.

Anonymous:1958:FMc

[809] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 11(3):??, juillet–septembre 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903605.

Anonymous:1958:FMd

[810] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 11(4):??, October 1958. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904879.

Robinet:1959:VAM

[811] André Robinet. La vocation académicienne de Malebranche. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):1–18, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904503.

Brun:1959:CWI

[812] Viggo Brun. Caspar Wessel et l’introduction géométrique des nombres complexes. Revue d’Histoire des Sciences et de Leurs Applications, 12(1): 19–24, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904504.

Birembaut:1959:DDU

[813] Arthur Birembaut. Les deux déterminations de l’unité de masse du système métrique. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):25–54, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904505.

Hahn:1959:CGP

[814] Roger Hahn. Colladon de Genève, précurseur de Mendel. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):55–56, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904506.

Theodorides:1959:EMB

[815] J. Théodoridès. Une exposition sur le microscope et la biologie au XIX^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):57– 59, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904507.

Theodorides:1959:MC

[816] J. Théodoridès. Maurice Caullery (1868–1958). Revue d’Histoire des Sciences et de Leurs Applications, 12(1):60–62, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904508.

Huard:1959:Fa

[817] P. Huard and J. Théodoridès. France. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):63–70, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904509.

Huard:1959:AOBa

[818] P. Huard. Analyses d’ouvrages: Histoire générale des sciences, t. I, La science antique et médiévale par René Taton. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):71–72, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904511.

Huard:1959:AOBb

[819] P. Huard. Analyses d’ouvrages: Histoire générale des sciences, t. II : La science moderne par René Taton. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):72–74, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904512.

Taton:1959:AOBa

[820] R. Taton. Analyses d’ouvrages: An Illustrated History of Science par F. Sherwood Taylor. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):74–75, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904513.

Taton:1959:AOBb

[821] R. Taton. Analyses d’ouvrages: Roots of Scientific Thought. A Cultural Perspective par Philip P. Wiener; Aaron Noland. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):75, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904514.

Taton:1959:AOBc

[822] René Taton. Analyses d’ouvrages: Manuel de bibliographie philosophique, (Collection ^{≪}Logos ^{≫}) par Gilbert Varet. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):75–76, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904515.

Taton:1959:AOBd

[823] R. Taton. Analyses d’ouvrages: History of Mathematics par David Eugene Smith. Revue d’Histoire des Sciences et de Leurs Applications, 12 (1):77, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904516.

Taton:1959:AOBe

[824] R. Taton. Analyses d’ouvrages: Geometry in Egyptian Art par Else Christie Kielland. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):77, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904517.

Taton:1959:AOBf

[825] René Taton. Analyses d’ouvrages: Anamorphoses ou perspectives curieuses, (Coll. ^{≪}Jeu savant ^{≫}) par Jurgus Baltrusaitis. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):78, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904518.

Naux:1959:AOBa

[826] Charles Naux. Analyses d’ouvrages: Der Briefwechsel von Johann Bernoulli, Band I Naturforschenden Gesellschaft in Basel. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):79–80, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904519.

Taton:1959:AOBg

[827] René Taton. Analyses d’ouvrages: La mathématique sociale du Marquis de Condorcet, (^{≪}Bibliothèque de Philosophie contemporaine ^{≫}) par Gilles-Gaston Granger. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):80–81, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904520.

Brun:1959:AOB

[828] Viggo Brun. Analyses d’ouvrages: Niels Henrik Abel, Mathematician Extraordinary par Oystein Ore. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):81–82, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904521.

Taton:1959:AOBh

[829] René Taton. Analyses d’ouvrages: Opere Matematiche, Tomo secondo, Tomo terzo par Paolo Ruffini; Ettore Bortolotti. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):83–84, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904522.

Taton:1959:AOBi

[830] René Taton. Analyses d’ouvrages: Opere Scelte a cura dell’Unione Matematica Italiana …, vol. I : Analisi matematica. Calcolo numerico par Giuseppe Peano; Opere Scelte a cura dell’Unione Matematica Italiana …, vol. II : Logica matematica. Interlingua ed algebra della grammatica par Giuseppe Peano. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):84–85, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904523.

Taton:1959:AOBj

[831] R. Taton. Analyses d’ouvrages: In memoria di Giuseppe Peano par Alessandro Terracini. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):85, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904524.

Naux:1959:AOBb

[832] Ch. Naux. Analyses d’ouvrages: L’aspect moderne des mathématiques par Lucienne Félix; G. Bouligand. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):85–86, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904525.

T:1959:AOB

[833] R. T. Analyses d’ouvrages: Historical Astronomical Books, (Book Exhibition number three). Revue d’Histoire des Sciences et de Leurs Applications, 12(1):86, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904526.

Taton:1959:AOBk

[834] R. Taton. Analyses d’ouvrages: Réflexions sur la puissance motrice du feu et sur les machines propres à développer cette puissance, nouv. éd. augmentée des Notes sur les mathématiques, la physique et autres sujets par Sadi Carnot. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):86, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904527.

Theodorides:1959:AOBa

[835] Jean Théodoridès. Analyses d’ouvrages: Aux sources de la Biologie par Jean Rostand. Revue d’Histoire des Sciences et de Leurs Applications, 12 (1):87, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904528.

Taton:1959:AOBl

[836] R. Taton. Analyses d’ouvrages: Linnaeus commemorated, 1707–May 23^{rd} 1957, (Communication n^{o} 103 from the N.M.H.S.). Revue d’Histoire des Sciences et de Leurs Applications, 12(1):88, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904529.

Theodorides:1959:AOBb

[837] J. Théodoridès. Analyses d’ouvrages: Some Milestones in the History of Haematology par Camille Dreyfus. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):88–89, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904530.

Huard:1959:AOBc

[838] P. Huard. Analyses d’ouvrages: La lutte contre les épizooties et l’École Vétérinaire de Lyon au XVIII^{e} siècle, (collection des ^{≪}Cahiers d’histoire ≫) par H. Hours. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):89–90, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904531.

Taton:1959:AOBm

[839] R. Taton. Analyses d’ouvrages: Le navire et l’économie maritime du XV^{e} au XVIII^{e} siècles, (^{≪}Bibliothèque générale de l’École pratique des Hautes Études, VI^{e} section ^{≫}) par Michel Mollat. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):90–91, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904532.

Taton:1959:AOBn

[840] R. Taton. Analyses d’ouvrages: Exposition Art, Histoire et Sciences. Catalogue. Revue d’Histoire des Sciences et de Leurs Applications, 12(1): 91, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904533.

Taton:1959:AOBo

[841] R. Taton. Analyses d’ouvrages: Musée national des Sciences exactes et naturelles à Leyde (extrait de Enseignement supérieur et recherches scientifiques aux Pays-Bas, vol. I, n^{o} 3, 1957), (Communication 101 b du ^{≪}Rijksmuseum voor de Geschiedenis der Natuurwetenschappen ^{≫}, Leyde) par Maria Rooseboom. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):91, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904534.

Taton:1959:AOBp

[842] R. Taton. Analyses d’ouvrages: Universitas Litterarum, Handbuch der Wissenschaftskunde … fasc. 3, 4, 5, 6, 7, 8, 9 par Werner Schuder. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):91–92, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904535.

Colnort:1959:AQPa

[843] S. Colnort. Analyses de quelques périodiques: Archives internationales d’Histoire des Sciences, t. IX (année 1956) et X (année 1957), n^{os} 34 à 37 et 38 à 41. Revue d’Histoire des Sciences et de Leurs Applications, 12(1): 92–93, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904536.

Colnort:1959:AQPb

[844] S. Colnort. Analyses de quelques périodiques: Osiris. Commentationes de scientiarum et eruditionis historia rationeque Volumen undecimum par Georgius Sarton Gandavensis. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):93–94, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904537.

Colnort:1959:AQPc

[845] Suzanne Colnort. Analyses de quelques périodiques: Osiris, Commentationes de scientiarum et eruditionis historia rationeque, Volumen duodecimum par A. Georgio Sarton Gandavensi; A. Rome; J. Mogenet. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):94–95, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904538.

Colnort:1959:AQPd

[846] S. Colnort. Analyses de quelques périodiques: Isis, an International Review devoted to the History of Science and its Cultural Influences, vol. 47, vol. 48 par I. Bernard Cohen. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):95–96, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904539.

Nadal:1959:GM

[847] André Nadal. Gaston Milhaud (1858–1918). Revue d’Histoire des Sciences et de Leurs Applications, 12(2):97–110, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905572.

Torlais:1959:LRD

[848] Jean Torlais. L’Académie de la Rochelle et la diffusion des sciences au XVIII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):111–125, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905573.

Taton:1959:CSFa

[849] René Taton. Condorcet et Sylvestre-François Lacroix (I). Revue d’Histoire des Sciences et de Leurs Applications, 12(2):127–158, avril– juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905574.

Kahan:1959:OTR

[850] T. Kahan. Sur les origines de la théorie de la relativité restreinte. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):159–165, avril– juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905575.

Costabel:1959:RL

[851] P. Costabel. Robert Lenoble (1902–1959). Revue d’Histoire des Sciences et de Leurs Applications, 12(2):167–169, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905576.

Huard:1959:Fb

[852] P. Huard, J. Théodoridès, and P. Costabel. France. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):169–173, avril–juin 1959. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905577.

Anonymous:1959:UPV

[853] Anonyme. U.S.A.: Prix Victor-Robinson. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):174, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905579.

D:1959:S

[854] S. D. Suisse. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):174, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905578.

Taton:1959:AOBq

[855] René Taton. Analyses d’ouvrages: La méthode dans les sciences modernes (numéro hors série de ^{≪}Travail et Méthodes ^{≫}) par François Le Lionnais. Revue d’Histoire des Sciences et de Leurs Applications, 12 (2):175–176, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905580.

Taton:1959:AOBr

[856] René Taton. Analyses d’ouvrages: The Principal Works of Simon Stevin, vol. II : Mathematics par D. J. Struik. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):176–178, avril–juin 1959. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905581.

Moscovici:1959:AOBa

[857] Serge Moscovici. Analyses d’ouvrages: Bibliografia Kopernikowska, 1509–1955 par Henry K. Baranowski. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):178–179, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23905582.

Taton:1959:AOBs

[858] René Taton. Analyses d’ouvrages: Théorie mathématique des phénomènes électrodynamiques uniquement déduite de l’expérience par André-Marie Ampère; Edmond Bauer. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):179, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23905583.

Moscovici:1959:AOBb

[859] Serge Moscovici. Analyses d’ouvrages: A Life of Sir William Ramsay par Morris W. Travers. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):179–180, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905584.

Taton:1959:AOBt

[860] René Taton. Analyses d’ouvrages: Temps, espace, matière. Leçons sur la théorie de la relativité générale, traduites sur la 4^{e} éd par Hermann Weyl; Gustave Juvet; Robert Leroy; Georges Bouligand. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):180, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905585.

Hooykaas:1959:AOBa

[861] R. Hooykaas. Analyses d’ouvrages: Nicolas Steno and his Indice ^{≪}Acta Historia Scient. Nat. et Med. ^{≫}, vol. 15 par Gustav Scherz. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):181–186, avril– juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905586.

Hooykaas:1959:AOBb

[862] R. Hooykaas. Analyses d’ouvrages: Robert Boyle and Seventeenth-Century Chemistry par Marie Boas. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):186–189, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23905587.

Roger:1959:AOB

[863] Jacques Roger. Analyses d’ouvrages: Tournefort, (collection ^{≪}Les Grands Naturalistes français ^{≫}) par R. Heim; Roger Heim. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):189–190, avril– juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905588.

Daumas:1959:AOB

[864] Maurice Daumas. Analyses d’ouvrages: Pasteur par André George. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):191, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905589.

Moscovici:1959:AOBc

[865] Serge Moscovici. Analyses d’ouvrages: Langage et communication, ≪Bibliothèque scientifique internationale ^{≫}par George A. Miller. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):191–192, avril– juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905590.

Colnort:1959:AQPe

[866] S. Colnort. Analyses de quelques périodiques: Gesnerus, vol. 14 (fascicules 1/2 et 3/4). Revue d’Histoire des Sciences et de Leurs Applications, 12(2):192, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905591.

Russo:1959:CAA

[867] F. Russo. La constitution de l’algèbre au XVI^{e} siècle: Étude de la structure d’une évolution. Revue d’Histoire des Sciences et de Leurs Applications, 12(3):193–208, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905101.

Dulieu:1959:M

[868] Louis Dulieu. Les magnol. Revue d’Histoire des Sciences et de Leurs Applications, 12(3):209–224, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23905102.

Tonelli:1959:NLN

[869] Giorgio Tonelli. La nécessité des lois de la nature au XVIII^{e} siècle et chez Kant en 1762. Revue d’Histoire des Sciences et de Leurs Applications, 12(3):225–241, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905103.

Taton:1959:CSFb

[870] René Taton. Condorcet et Sylvestre-François Lacroix (fin). Revue d’Histoire des Sciences et de Leurs Applications, 12(3):243–262, juillet– septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905104.

Huard:1959:PD

[871] Pierre Huard and Jean Théodoridès. Paul Delaunay (1878–1958). Revue d’Histoire des Sciences et de Leurs Applications, 12(3):263–266, juillet– septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905105.

Theodorides:1959:FJC

[872] J. Théodoridès. Francis-Joseph Cole, F.R.S. (1872–1959). Revue d’Histoire des Sciences et de Leurs Applications, 12(3):266–268, juillet– septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905106.

Costabel:1959:Fa

[873] P. Costabel. France. Revue d’Histoire des Sciences et de Leurs Applications, 12(3):268–270, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905107.

Klein:1959:AOB

[874] Robert Klein. Analyses d’ouvrages: Spiritual and demonic Magic, from Ficino to Campanella, ^{≪}Studies of the Warburg Institute ^{≫}par D. P. Walker. Revue d’Histoire des Sciences et de Leurs Applications, 12(3): 271–274, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905109.

Robinet:1959:AOB

[875] André Robinet. Analyses d’ouvrages: Barock und Universalismus par J. O. Fleckenstein; G. W. Leibniz. Revue d’Histoire des Sciences et de Leurs Applications, 12(3):275–276, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905110.

Birembaut:1959:AOBa

[876] Arthur Birembaut. Analyses d’ouvrages: Resa i Europa. En astronom, industrispion och teaterhabitué genom Danmark, Tyskland, Holland, England, Frankrike och Italien 1758–1762 par Bengt Ferrner; Sten G. Lindberg. Revue d’Histoire des Sciences et de Leurs Applications, 12(3): 276–277, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905111.

Taton:1959:AOBu

[877] R. Taton. Analyses d’ouvrages: Celebrazione in Lugo del centenario della nascita di Gregorio Ricci Curbastro. 2 maggio 1954,^{≪}Gregorio Ricci Curbastro ^{≫}. Revue d’Histoire des Sciences et de Leurs Applications, 12(3): 277–278, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905112.

Theodorides:1959:AOBc

[878] J. Théodoridès. Analyses d’ouvrages: Sur le darwinisme présumé d’Alberuni 800 ans avant Darwin, ^{≪}Publications de l’Université libanaise, Section des Sciences naturelles ^{≫}, III par J. Z. Wilczynski. Revue d’Histoire des Sciences et de Leurs Applications, 12(3):278, juillet– septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905113.

Huard:1959:AOBd

[879] P. Huard. Analyses d’ouvrages: La série d’Extrême-Orient du fonds des Archives coloniales conservée aux Archives nationales par Ferréol de Ferry. Revue d’Histoire des Sciences et de Leurs Applications, 12(3): 278–279, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905114.

Schrimpf:1959:AOB

[880] Robert Schrimpf. Analyses d’ouvrages: Scienze della Cina, in Le Civiltà dell’Oriente, vol. III par Marcello Muccioli. Revue d’Histoire des Sciences et de Leurs Applications, 12(3):279–280, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905115.

Huard:1959:AOBe

[881] P. Huard. Analyses d’ouvrages: Contribution à l’étude de la vie et de l’œuvre d’André Vésale par J. Brocas. Revue d’Histoire des Sciences et de Leurs Applications, 12(3):280–281, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905116.

Huard:1959:AOBf

[882] P. Huard and J. Théodoridès. Analyses d’ouvrages: A guide to the History of Bacteriology par Thomas H. Grainger. Revue d’Histoire des Sciences et de Leurs Applications, 12(3):281–283, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905117.

Huard:1959:AOBg

[883] P. Huard and Wong. Analyses d’ouvrages: Tong Ki yi-hiue-louen tchou san-tchong (Trois sortes d’études médicales), 1^{re} impression, 1^{re} édition par Tong Ki. Revue d’Histoire des Sciences et de Leurs Applications, 12 (3):283–284, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905118.

Huard:1959:AOBh

[884] P. Huard. Analyses d’ouvrages: Les relations culturelles entre la France et la Pologne dans le domaine de la Médecine par Stanislas Konopka. Revue d’Histoire des Sciences et de Leurs Applications, 12(3): 284–285, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905119.

Birembaut:1959:AOBb

[885] Arthur Birembaut. Analyses d’ouvrages: La primera casa de las ciencias en México, el Real Seminario de Minería (1792–1811) par José Joaquin Izquierdo. Revue d’Histoire des Sciences et de Leurs Applications, 12 (3):285, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905120.

Machabey:1959:AOB

[886] Armand Machabey. Analyses d’ouvrages: Survivance des mesures traditionnelles en Tunisie, ^{≪}Publications de l’Institut des Hautes Études de Tunis, Mémoires du Centre d’Études des Sciences humaines ^{≫}, vol. IV par Marcel Legendre. Revue d’Histoire des Sciences et de Leurs Applications, 12(3):285–286, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905121.

Poulle:1959:AOBa

[887] Emmanuel Poulle. Analyses d’ouvrages: Christophe Colomb amiral de la mer océane, ^{≪}La Croix du Sud ^{≫} par Samuel Eliot Morison; Josette Hesse. Revue d’Histoire des Sciences et de Leurs Applications, 12(3): 286–287, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905122.

Colnort:1959:AQPf

[888] S. Colnort. Analyses de quelques périodiques: Centaurus, International Magazine of the History of Science and Medicine, vol. 4. Revue d’Histoire des Sciences et de Leurs Applications, 12(3):287–288, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905123.

Payen:1959:FFS

[889] Jacques Payen. Flos Florum et Semita Semite. Deux traités d’alchimie attribués à Arnaud de Villeneuve. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):289–300, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905653.

Colnort:1959:TTA

[890] Suzanne Colnort. Un traité de thérapeutique au XVI^{e} siècle : Brouaut et la panacée alcoolique. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):301–313, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905654.

Torlais:1959:ICP

[891] Jean Torlais. Inventaire de la correspondance et des papiers de Réaumur conservés aux archives de l’Académie des Sciences de Paris. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):315–326, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905655.

Cousin:1959:LSB

[892] Jean Cousin. L’Académie des Sciences, Belles-Lettres et Arts de Besançon au XVIII^{e} siècle et son œuvre scientifique. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):327–344, October 1959. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905656.

Duveen:1959:IIP

[893] Denis I. Duveen and Lucien Scheler. Des illustrations inédites pour les Mémoires de chimie ouvrage posthume de Lavoisier. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):345–353, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905657.

Delorme:1959:RB

[894] Suzanne Delorme. Raymond Bayer (1898–1959). Revue d’Histoire des Sciences et de Leurs Applications, 12(4):355–357, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905658.

Costabel:1959:AIH

[895] P. Costabel. Académie Internationale d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):357–359, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905659.

Anonymous:1959:BIC

[896] Anonyme. Benelux: III^{e} Congrès Benelux d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):359, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905660.

Costabel:1959:EIC

[897] Pierre Costabel and Suzanne Delorme. Espagne: IX^{e} Congrès International d’Histoire des Sciences (Barcelone–Madrid, 1^{e}r–7 septembre 1959). Revue d’Histoire des Sciences et de Leurs Applications, 12(4):359–363, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905661.

Costabel:1959:Fb

[898] P. Costabel. France. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):363–367, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905662.

Delorme:1959:UMG

[899] Suzanne Delorme. U.S.A. Médaille George Sarton. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):367–368, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905663.

Moscovici:1959:AOBd

[900] Serge Moscovici. Analyses d’ouvrages: Patterns of Discovery par N. R. Hanson. Revue d’Histoire des Sciences et de Leurs Applications, 12(4): 369, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905664.

Payen:1959:AOB

[901] J. Payen. Analyses d’ouvrages: Il tenore di vita pitagorico ed il problema della omoiosis par Dott. Vincenzo Capparelli. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):370, October 1959. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905666.

Russo:1959:AOB

[902] F. Russo. Analyses d’ouvrages: La conscience de la rationalité. Étude phénoménologique sur la physique mathématique, ^{≪}Bibliothèque de philosophie contemporaine^{≫}par Suzanne Bachelard. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):370, October 1959. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905665.

Taton:1959:AOBv

[903] R. Taton. Analyses d’ouvrages: German Influence upon English Education and Science, 1800–1866 par George Haines; The Organisation of Science in England. A Retrospect par D. S. L. Cardwell. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):371, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905667.

Poulle:1959:AOBb

[904] Emmanuel Poulle. Analyses d’ouvrages: Tractat d’astrologia,^{≪}Biblioteca catalana d’obres antigues^{≫}par Bartomeu de Tresbens; Juan Vernet; David Romano. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):372, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905668.

Poulle:1959:AOBc

[905] Emmanuel Poulle. Analyses d’ouvrages: Medieval Latin Scientific Writings in the Barberini Collection; a provisional Catalogue par Theodore Silverstein. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):372–373, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905669.

Colnort:1959:AOBa

[906] S. Colnort. Analyses d’ouvrages: On the Magnet, fac-similé, ^{≪}The Collector’s Series in Science ^{≫}, II par William Gilbert; Derek J. Price. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):373–374, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905670.

Moscovici:1959:AOBe

[907] S. Moscovici. Analyses d’ouvrages: Isaac Newton, Papers and Letters on Natural Philosophy par I. Bernard Cohen. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):374–376, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905671.

Anonymous:1959:PR

[908] Anonyme. Publications reçues. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):376, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905673.

Colnort:1959:AOBb

[909] S. Colnort. Analyses d’ouvrages: Principes de morphologie générale, ≪Science et Esthétique^{≫} par Édouard Monod-Herzen. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):376, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905672.

Anonymous:1959:TGM

[910] Anonyme. Table générale des matières du Tome XII. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):377–384, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905674.

Anonymous:1959:BMa

[911] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):??, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904540.

Anonymous:1959:BMb

[912] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):??, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905592.

Anonymous:1959:BMc 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905124.

Anonymous:1959:BMd

[914] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):??, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905675.

Anonymous:1959:FMa

[915] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 12(1):??, January 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904502.

Anonymous:1959:FMb

[916] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 12(2):??, avril–juin 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905571.

Anonymous:1959:FMc

[917] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 12(3):??, juillet–septembre 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905100.

Anonymous:1959:FMd

[918] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 12(4):??, October 1959. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905652.

Bourdier:1960:TSH

[919] Franck Bourdier. Trois siècles d’hypothèses sur l’origine et la transformation des êtres vivants (1550–1859). Revue d’Histoire des Sciences et de Leurs Applications, 13(1):1–44, January 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904028.

Rostand:1960:PFC

[920] Jean Rostand. Les précurseurs français de Charles Darwin. Revue d’Histoire des Sciences et de Leurs Applications, 13(1):45–58, January 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904029.

Vandel:1960:LD

[921] Albert Vandel. Lamarck et Darwin. Revue d’Histoire des Sciences et de Leurs Applications, 13(1):59–72, January 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904030.

Grasse:1960:LWD

[922] Pierre-P. Grassé. Lamarck, Wallace et Darwin. Revue d’Histoire des Sciences et de Leurs Applications, 13(1):73–79, January 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904031.

Canguilhem:1960:LAD

[923] Georges Canguilhem. L’homme et l’animal du point de vue psychologique selon Charles Darwin. Revue d’Histoire des Sciences et de Leurs Applications, 13(1):81–94, January 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904032.

Robinet:1960:JPO

[924] André Robinet. Jean Prestet ou la bonne foi cartésienne (1648–1691). Revue d’Histoire des Sciences et de Leurs Applications, 13(2):95–104, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904308.

Bouligand:1960:LEM

[925] Georges Bouligand. L’œuvre d’Euler et la mécanique des fluides au XVIII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):105–113, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904309.

Bykov:1960:DPC

[926] G. V. Bykov and J. Jacques. Deux pionniers de la chimie moderne, Adolphe Wurtz et Alexandre M. Boutlerov, d’après une correspondance inédite. Revue d’Histoire des Sciences et de Leurs Applications, 13 (2):115–134, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904310.

Robinet:1960:LC

[927] André Robinet. L’abbé de Catelan. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):135–137, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904311.

Rochot:1960:LSS

[928] B. Rochot. L’histoire des sciences à la Société d’Étude du XVII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):137–138, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904312.

Theodorides:1960:FB

[929] J. Théodoridès. F. S. Bodenheimer (1897–1959). Revue d’Histoire des Sciences et de Leurs Applications, 13(2):139–141, avril–juin 1960. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904313.

Itard:1960:PVE

[930] Jean Itard. Paul Ver Eecke (1867–1959). Revue d’Histoire des Sciences et de Leurs Applications, 13(2):141–143, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904314.

Anonymous:1960:UIH

[931] Anonyme. Union Internationale d’Histoire et de Philosophie des Sciences: Division d’Histoire des Sciences: X^{e} Congrès International d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 13(2): 144, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904315.

Anonymous:1960:Fa

[932] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):144–147, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904316.

Anonymous:1960:IC

[933] Anonyme. Iran: Conférences. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):147–148, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904317.

Anonymous:1960:ISI

[934] Anonyme. Italie: Symposium International (Florence–Vinci, 8–10 octobre 1960). Revue d’Histoire des Sciences et de Leurs Applications, 13(2): 148, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904318.

Anonymous:1960:PBS

[935] Anonyme. Pays-Bas: Société néerlandaise pour l’histoire des sciences exactes, naturelles et médicales (réunion d’automne). Revue d’Histoire des Sciences et de Leurs Applications, 13(2):148, avril–juin 1960. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904319.

Anonymous:1960:URC

[936] Anonyme. U.R.S.S.: Conférence sur les relations scientifiques entre la France et la Russie. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):148, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904320.

Moscovici:1960:AOBa

[937] Serge Moscovici. Analyses d’ouvrages: Die Mechanisierung des Weltbildes par E. J. Dijksterhuis; H. Habicht. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):149–151, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904321.

Poulle:1960:AOBa

[938] Emmanuel Poulle. Analyses d’ouvrages: The Copernican Revolution : Planetary Astronomy in the development of Western Thought par Thomas S. Kuhn. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):151–152, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904322.

Bayard-Pierlot:1960:AOB

[939] J. Bayard-Pierlot. Analyses d’ouvrages: Opera omnia, t. II : Orationes, 1954, t. III : De Ecclesia libe secundus, 1955 par Andreae Fricii Mocrevii; C. Kumaniecki. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):153, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904324.

Delorme:1960:AOBa

[940] Suzanne Delorme. Analyses d’ouvrages: Francis Bacon, A Bibliography of his Works and of Baconiana to the year 1750, Supplement, Bruning Copyflex reproduction of typescript par R. W. Gibson. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):153, avril–juin 1960. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904323.

Roger:1960:AOB

[941] Jacques Roger. Analyses d’ouvrages: Goethe et l’esprit français, Actes du Colloque international de Strasbourg, 23–27 avril 1957. fasc. 137. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):154–156, avril– juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904325.

Klein:1960:AOBa

[942] Marc Klein. Analyses d’ouvrages: Alexander von Humboldt (14-9-1769 / 6-5-1859), Gedenkschrift zur 100. Wiederkehr seines Todestages Alexander von Humboldt-Kommission der Deutschen Akademie der Wissenschaften zu Berlin. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):156–157, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904326.

Taton:1960:AOBa

[943] R. Taton. Analyses d’ouvrages: Opere Scelte a cura dell’Unione Matematica Italiana …, vol. III : Geometria e Fondamenti, Meccanica razionale, Varie par Giuseppe Peano. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):157–158, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904327.

Russo:1960:AOB

[944] F. Russo. Analyses d’ouvrages: ^{≪}Structures^{≫}, chap. IV du livre I (Théorie des ensembles) de la Première Partie (Structures fondamentales de l’analyse) des Éléments de mathématiques,^{≪}Actualités scientifiques^{≫}, n^{o} 1258 par N. Bourbaki. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):158, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904328.

Payen:1960:AOB

[945] J. Payen. Analyses d’ouvrages: Les alchimistes, coll. ^{≪}Le Temps qui court^{≫}, n^{o} 16 par M. Caron; S. Hutin. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):159, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904329.

Rostand:1960:AOB

[946] Jean Rostand. Analyses d’ouvrages: Forerunners of Darwin, 1745–1859 par Bentley Glass; Owsei Temkin; William L. Straus. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):159–160, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904330.

Klein:1960:AOBb

[947] Marc Klein. Analyses d’ouvrages: Jan Evangelista Purkyne : Czech Scientist and Patriot, (1787–1869) (Memoirs of the American Philosophical Society, vol. 49) par Henry J. John. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):160–163, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904331.

Dulieu:1960:AOBa

[948] Louis Dulieu. Analyses d’ouvrages: Matériaux inédits préparés par Delile pour une Flore de l’Amérique du Nord, in Les Botanistes français en Amérique du Nord avant 1850 …, Coll. ^{≪}Colloques internationaux du C.N.R.S. ^{≫}, LXIII par Jean Motte. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):163–164, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904332.

Colnort:1960:AOBa

[949] S. Colnort. Analyses d’ouvrages: Henri-Ferdinand Van Heurck (1838– 1909). Sa vie, son œuvre, ^{≪}Communication n^{o} 111^{≫}par Éd. Frison; Maria Rooseboom. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):164, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904333.

Moreau:1960:AOBa

[950] F. Moreau. Analyses d’ouvrages: Melchior Treub, pioneer of a new era in the History of the Malay Archipelago par H. H. Zeijlstra. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):164–165, avril– juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904334.

Wickersheimer:1960:AOBa

[951] Ernest Wickersheimer. Analyses d’ouvrages: A short Introduction to Anatomy (Isagogae breves) par Jacopo Berengario da Carpi; L. R. Lind; Paul G. Roofe. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):166–167, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904335.

Theodorides:1960:AOBa

[952] J. Théodoridès. Analyses d’ouvrages: La médecine chinoise au cours des siècles par P. Huard; M. Wong. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):167–169, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904336.

Wickersheimer:1960:AOBb

[953] Ernest Wickersheimer. Analyses d’ouvrages: History of American Medicine, (M.D. International Symposia, n^{o} 5) par Felix Marti-Ibañez. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):169–170, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904337.

Klein:1960:AOBc

[954] Marc Klein. Analyses d’ouvrages: Geschichte der Medizin, II/1, Von der Medizin der Aufklärung bis zur Begründung der Zellularpathologie (etwa 1740–etwa 1858). 2^{e} éd. par Paul Diepgen. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):170–171, avril–juin 1960. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904338.

Wickersheimer:1960:AOBc

[955] Ernest Wickersheimer. Analyses d’ouvrages: Jenner of Berkeley par Dorothy Fisk. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):172–173, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904339.

D:1960:AOBa

[956] L. D. Analyses d’ouvrages: La Faculté de Médecine de Montpellier. Aperçu historique par Jean Turchini. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):174, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904341.

Dulieu:1960:AOBb

[957] Louis Dulieu. Analyses d’ouvrages: Les Facultés de Médecine de province avant la Révolution par André Finot. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):174, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904340.

D:1960:AOBb

[958] L. D. Analyses d’ouvrages: Illustrations relatives à l’histoire de la médecine en Arménie, de l’Antiquité jusqu’à nos jours, Erévan par L. A. Ohanessian. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):175, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904342.

D:1960:AOBc

[959] L. D. Analyses d’ouvrages: Contributie la Istoricul Spitalebor din Moldova par Paul Pruteanu. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):175, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904343.

D:1960:AOBd

[960] L. D. Analyses d’ouvrages: Prazské Lékarské Disertage par L. Schid; E. Rozsivalova. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):175, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904344.

D:1960:AOBe

[961] L. D. Analyses d’ouvrages: Serment d’Hippocrate par Feridun Nafiz Uzluk. Revue d’Histoire des Sciences et de Leurs Applications, 13(2): 175, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904345.

Dulieu:1960:AOBc

[962] Louis Dulieu. Analyses d’ouvrages: Notes sur l’histoire de la clinique médicale et de son enseignement. D’Imhotep à Trousseau par André Finot. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):175– 176, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904346.

Huard:1960:AOBa

[963] P. Huard. Analyses d’ouvrages: Chardon de Courcelles et le Service de Santé de la Marine à Brest au XVIII^{e} siècle par Étienne Le Fur. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):176–177, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904347.

Huard:1960:AOBb

[964] P. Huard. Analyses d’ouvrages: La vie et l’œuvre de Theodor Billroth. Contribution à l’étude de l’histoire de la médecine au XIX^{e} siècle par Constant-Albert Bretz. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):177–178, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904348.

Huard:1960:AOBc

[965] P. Huard. Analyses d’ouvrages: A History of Neurology par Walther Riese. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):178– 179, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904349.

Torlais:1960:AOB

[966] Jean Torlais. Analyses d’ouvrages: Therapeutic, Electricity and Ultraviolet Radiation par Sidney Licht. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):179–180, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904350.

Bert:1960:AOB

[967] Maurice Bert. Analyses d’ouvrages: Brevets d’invention français (1791– 1902). Un siècle de progrès technique. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):180–183, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904351.

Colnort:1960:AOBb

[968] S. Colnort. Analyses d’ouvrages: Les bâtisseurs de cathédrales, coll. ≪Microcosme^{≫}, série ^{≪}Le Temps qui court^{≫}, n^{o} 11 par Jean Gimpel. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):183–184, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904352.

Taton:1960:AOBb

[969] R. Taton. Analyses d’ouvrages: Le navire et l’économie maritime du Moyen Âge au XVIIIe siècle, principalement en Méditerranée. Travaux … (^{≪}Bibliothèque générale de l’École pratique des Hautes Études, VI^{e} section^{≫}) par M. Mollat; O. de Prat. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):184, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904353.

Colnort:1960:AOBc

[970] S. Colnort. Analyses d’ouvrages: The Language of Psychology … par George Mandler; William Kessen. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):184–185, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904354.

Voise:1960:AQP

[971] W. Voisé. Analyses de quelques périodiques: Kwartalnik Historii Nauki i Techniki (= Revue trimestrielle d’Histoire de la Science et des Techniques), (Édition Scientifique de l’État), vol. I–IV. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):185–188, avril–juin 1960. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904355.

Colnort:1960:AQPa

[972] S. Colnort. Analyses de quelques périodiques: Centaurus, International Magazine of the History of Science and Medicine, vol. 5, fasc. 1, 2, 3/4. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):188, avril– juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904356.

Colnort:1960:AQPb

[973] S. Colnort. Analyses de quelques périodiques: Isis, an International Review devoted to the History of Science and its cultural Influences, (vol. 49), (vol. 50), (4 fasc.), n^{os} 155 à, (4 fasc.) n^{os} 159 à 162 inc. par I. Bernard Cohen; Harry Woolf. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):188–189, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904357.

Colnort:1960:AQPc

[974] S. Colnort. Analyses de quelques périodiques: Gesnerus, Revue trimestrielle, publiée par la Société suisse d’Histoire de la Médecine et des Sciences naturelles, vol. 15, 1958, fascicules 1/2, 3/4. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):190, avril–juin 1960. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904358.

Dilgan:1960:DDV

[975] Hâmid Dilgan. Démonstration du V^{e} postulat d’Euclide par Schamsed-Din Samarkandi: Traduction de l’ouvrage Aschkâl-üt-teessis de Samarkandi. Revue d’Histoire des Sciences et de Leurs Applications, 13 (3):191–196, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904608.

Koyre:1960:BMG

[976] Alexandre Koyré. Le de Motu Gravium de Galilée : de l’expérience imaginaire et de son abus. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):197–245, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904609.

Chauchard:1960:PDC

[977] Paul Chauchard. À propos du cinquantenaire de la chronaxie : L’importance de l’œuvre de Louis Lapicque en neurophysiologie. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):247–258, juillet– septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904610.

Rostand:1960:DNC

[978] Jean Rostand. Du nouveau sur colladon: précurseur de Mendel. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):259–262, juillet– septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904611.

Theodorides:1960:SAU

[979] J. Théodoridès. Suisse: 500^{e} anniversaire de l’Université de Bâle. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):263–264, juillet– septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904612.

Anonymous:1960:YIH

[980] Anonyme. Yougoslavie: L’Institut d’Histoire des Sciences de Zagreb. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):264, juillet– septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904613.

Moscovici:1960:AOBb

[981] Serge Moscovici. Analyses d’ouvrages: George Berkeley, coll. ^{≪}Les Grands Penseurs ^{≫}par André-Louis Leroy. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):265, juillet–septembre 1960. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904614.

Moscovici:1960:AOBc

[982] Serge Moscovici. Analyses d’ouvrages: Critical Problems in the History of Science par Marshall Clagett. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):266–268, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904615.

Grmek:1960:AOBa

[983] M. D. Grmek. Analyses d’ouvrages: Istoriografiya estestvennyh nauk v Rossii (XVIII v. pervaya polovina XIX v.) par V. P. Zoubov. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):268–269, juillet– septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904616.

Itard:1960:AOBa

[984] Jean Itard. Analyses d’ouvrages: Dictionnaire historique de la terminologie géométrique des Grecs. T. I : Introduction, Répertoire géométrique, Lettres A à K. T. II : Logos à la fin. ^{≪}Études et commentaires ^{≫}par Charles Mugler. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):269–271, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904617.

Itard:1960:AOBb

[985] Jean Itard. Analyses d’ouvrages: Le Traité des hosoèdres par Vito Caravelli; Paul Ver Eecke. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):271, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904618.

Itard:1960:AOBc

[986] Jean Itard. Analyses d’ouvrages: Problèmes plaisants et délectables qui se font par les nombres, 5^{e} édition revue par Claude-Gaspar Bachet; A. Labosne; J. Itard. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):271–272, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904619.

Itard:1960:AOBd

[987] Jean Itard. Analyses d’ouvrages: Récréations mathématiques. vol. I : vol. II : vol. III : vol. IV par Édouard Lucas. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):272, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904620.

Itard:1960:AOBe

[988] Jean Itard. Analyses d’ouvrages: L’Optique et la Catoptrique par Euclide; Paul Ver Eecke. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):272–273, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904621.

Colnort:1960:AOBd

[989] Suzanne Colnort. Analyses d’ouvrages: Catalogo dei Globi Antichi conservati in Italia, fasc. 1 : I Globi Blaviani. coll. ^{≪}Istituto e Museo di Storia della Scienza. Biblioteca ^{≫}, I par Leopoldina Luzio; Roberto Almagia; Catalogo dei Globi Antichi conservati in Italia, fasc. 2 : I Globi di Vincenzo Coronelli. coll. ^{≪}Istituto e Museo di Storia della Scienza. Biblioteca ≫, III par Maria Luisa Bonelli; Roberto Almagia. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):273, juillet–septembre 1960. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904622.

Moscovici:1960:AOBd

[990] Serge Moscovici. Analyses d’ouvrages: De historische Outwikkeling der Stootwetten van Aristoteles tot Huygens par J. MacLean. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):274, juillet– septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904623.

Jacques:1960:AOB

[991] J. Jacques. Analyses d’ouvrages: De geschiedenis van de benoeming van August Kekulé te Gent en de oprichting van het eerste onderrichtslaboratorium voor scheikunde in Belgie, ^{≪}Koninklijke vlaamse Academie voor wetenschappen, letteren en schone kunsten van België ^{≫}par J. Gillis. Revue d’Histoire des Sciences et de Leurs Applications, 13(3): 274–275, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904624.

Moreau:1960:AOBb

[992] F. Moreau. Analyses d’ouvrages: The Garden Flowers of China par H. L. Li. Revue d’Histoire des Sciences et de Leurs Applications, 13(3): 275–276, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904625.

Theodorides:1960:AOBb

[993] Jean Théodoridès. Analyses d’ouvrages: A History of Cytology par Arthur Hughes. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):276–277, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904626.

Theodorides:1960:AOBc

[994] J. Théodoridès. Analyses d’ouvrages: Breve storia dell’Ematologia par Pericle di Pietro. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):277, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904627.

Huard:1960:AOBd

[995] P. Huard. Analyses d’ouvrages: A History of Ophtalmology par George E. Arrington. Revue d’Histoire des Sciences et de Leurs Applications, 13(3): 277–278, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904628.

Huard:1960:AOBe

[996] P. Huard. Analyses d’ouvrages: L’ophtalmologie française au XX^{e} siècle, Les progrès de l’enseignement de la clinique et de la thérapeutique par René Onfray. Revue d’Histoire des Sciences et de Leurs Applications, 13 (3):278–279, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904629.

Glesinger:1960:AOB

[997] L. Glesinger. Analyses d’ouvrages: Parasites and parasitic Infections in early Medicine and Science par R. Hoeppli. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):280, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904630.

Huard:1960:AOBf

[998] P. Huard. Analyses d’ouvrages: Histoire de la gale par Marie Christophe. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):280, juillet– septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904631.

Grmek:1960:AOBb

[999] M. D. Grmek. Analyses d’ouvrages: On the Sociology of Medicine par Henry E. Sigerist; Milton J. Roemer; James M. Mackintosh. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):281–282, juillet– septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904632.

Huard:1960:AOBg

[1000] Pierre Huard. Analyses d’ouvrages: Histoire de l’École de Chirurgie de Rochefort, de ses origines jusqu’au début du XIX^{e} siècle (1722–1836) par Alain le Bozec. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):282–283, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904633.

Huard:1960:AOBh

[1001] P. Huard. Analyses d’ouvrages: Le Val-de–Grâce et l’École d’Application du Service de Santé militaire. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):283, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904634.

Huard:1960:AOBi

[1002] P. Huard. Analyses d’ouvrages: Le Service de Santé des Armées en Algérie, (1830–1958). Revue d’Histoire des Sciences et de Leurs Applications, 13(3):283–284, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904635.

Huard:1960:AOBj

[1003] P. Huard. Analyses d’ouvrages: Bio-bibliographie de Léon Le Fort (1829– 1893), n^{o} 60 par Éliane Provost. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):284–285, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904636.

Colnort:1960:AQPd

[1004] S. Colnort. Analyses de quelques périodiques: Ambix. The Journal of the Society for the Study of Alchemy and Early Chemistry (vol. VI, n^{os} 1, 2,

3) (vol. VII, n^{os} 1, 2, 3). Revue d’Histoire des Sciences et de Leurs Applications, 13(3):285, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904637.

Theodorides:1960:AQP

[1005] J. Théodoridès. Analyses de quelques périodiques: Monspeliensis Hippocrates, revue de la Société montpelliéraine d’Histoire de la Médecine, 9 numéros parus (1958–1960). Revue d’Histoire des Sciences et de Leurs Applications, 13(3):286, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904638.

Robinet:1960:GMI

[1006] André Robinet. Le groupe malebranchiste introducteur du calcul infinitésimal en France. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):287–308, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904729.

Mortier:1960:NPD

[1007] Roland Mortier. Note sur un passage du Rêve de d’Alembert : Réaumur et le problème de l’hybridation. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):309–316, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904730.

Nedelkovitch:1960:PSY

[1008] Duchan Nedelkovitch. Les principaux savants yougoslaves au XIX^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):317–323, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904731.

Davydoff:1960:AKS

[1009] Constantin Davydoff. Alexandre Kovalevsky (1840–1901) Souvenirs d’un disciple. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):325–348, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904732.

Anonymous:1960:Fb

[1010] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):349–350, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904733.

Huard:1960:E

[1011] P. Huard and Jean Théodoridès. Expositions. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):350–352, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904734.

Anonymous:1960:C

[1012] Anonyme. Conférences. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):352–353, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904735.

Anonymous:1960:CHS

[1013] Anonyme. Cours d’histoire des sciences. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):353–354, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904736.

Theodorides:1960:AOBd

[1014] J. Théodoridès. Analyses d’ouvrages: La Science au XVI^{e} siècle, Colloque international de Royaumont, 1–4 juillet 1957, coll. ^{≪}Histoire de la Pensée ^{≫}, n^{o} II, VI^{e} section. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):355, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904737.

Itard:1960:AOBf

[1015] Jean Itard. Analyses d’ouvrages: Correspondance du P. Marin Mersenne, religieux minime, vol. V (1635) par Paul Tannery; Cornélis de Waard. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):355–357, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904738.

Canguilhem:1960:AOBa

[1016] G. Canguilhem. Analyses d’ouvrages: Claude Bernard and his place in the History of Ideas par Reino Virtanen. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):357–358, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904739.

Michel:1960:AOB

[1017] Henri Michel. Analyses d’ouvrages: Heavenly Clockwork par J. Needham; Wang Ling; D. J. Price. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):358–359, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904740.

Poulle:1960:AOBb

[1018] Emmanuel Poulle. Analyses d’ouvrages: The Observatory in Islam and its place in the general History of the Observatory, series VII, n^{o} 38 par Aydin Sayili. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):359–360, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904741.

Dufour:1960:AOB

[1019] L. Dufour. Analyses d’ouvrages: A Handbook of Renaissance Meteorology with Particular Reference to Elisabethan and Jacobean Literature par S. K. Heninger. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):360–361, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904742.

Canguilhem:1960:AOBb

[1020] G. Canguilhem. Analyses d’ouvrages: Andreas Vesalius’ First Public Anatomy at Bologna (1540), Medicinae Scolaris, together with his notes on Matthaeus Curtius’ Lectures on Anatomia Mundini par Baldasar Heseler; Ruben Eriksson. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):361, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904743.

Canguilhem:1960:AOBc

[1021] G. Canguilhem. Analyses d’ouvrages: De Motu Locali Animalium (1627) par William Harvey; Gweneth Whitteridge. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):361–362, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904744.

Canguilhem:1960:AOBd

[1022] G. Canguilhem. Analyses d’ouvrages: Bibliographia Boerhaaviana par G. A. Lindeboom; H. Boerhaave. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):362, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904745.

Theodorides:1960:AOBe

[1023] J. Théodoridès. Analyses d’ouvrages: Letters to Achille Chaper par Victor Jacquemont. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):362–363, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904746.

Theodorides:1960:AOBf

[1024] J. Théodoridès. Analyses d’ouvrages: Naissance et déviation de la théorie cellulaire dans l’œuvre de Théodore Schwann, ^{≪}Actualités scientifiques et industrielles ^{≫}, n^{o} 1282 par M. Florkin; Jean Rostand. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):363–364, October 1960. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904747.

Canguilhem:1960:AOBe

[1025] G. Canguilhem. Analyses d’ouvrages: Darwin’s Biological Work, Some Aspects Reconsidered par P. R. Bell; J. Challinor; J. B. S. Haldane; P. Marler; H. L. K. Whitehouse; J. S. Wilkie. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):364–365, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904748.

Theodorides:1960:AOBg

[1026] J. Théodoridès. Analyses d’ouvrages: Perspectives de la zoologie européenne. Histoire, problèmes contemporains par Jean Leclercq. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):365–366, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904749.

Theodorides:1960:AOBh

[1027] J. Théodoridès. Analyses d’ouvrages: Souvenirs. Recueil d’articles autobiographiques par Elie Metchnikov; A. Gaissinovitch. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):366, October 1960. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904750.

Wickersheimer:1960:AOBd

[1028] Ernest Wickersheimer. Analyses d’ouvrages: Selbstbiographie und Antrittsrede, (Oesterreichische Akademie der Wissenschaften, Philosophischhistorische Klasse, Sitzungsberichte, CCXXXIV, 3) par Carl von Rokitansky; Erna Lesky. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):367–368, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904751.

Huard:1960:AOBk

[1029] Pierre Huard. Analyses d’ouvrages: Storia della Medicina par Loris Permuda. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):368– 369, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904752.

Canguilhem:1960:AOBf

[1030] G. Canguilhem. Analyses d’ouvrages: Centaur. Essays on the History of Medical Ideas par Felix Martí-Ibáñez. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):369, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904753.

Theodorides:1960:AOBi

[1031] J. Théodoridès. Analyses d’ouvrages: Comptes rendus du Congrès des Sociétés savantes de Paris et des départements. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):369–370, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904754.

Delorme:1960:AOBb

[1032] S. Delorme. Analyses d’ouvrages: Revue de Synthèse. Ancienne Revue de Synthèse historique, Troisième série (depuis janvier 1956) par Henri Berr. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):370–371, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904755.

Colnort:1960:AQPe

[1033] S. Colnort. Analyses de quelques périodiques: Archives internationales d’Histoire des Sciences, revue trimestrielle, publiée par la division d’Histoire des Sciences de l’Union internationale d’Histoire et de Philosophie des Sciences, t. XI (année 1958) et t. XII (année 1959), n^{os} 42 à 45 et 46 à 49. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):371–372, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904756.

Colnort:1960:AQPf

[1034] S. Colnort. Analyses de quelques périodiques: Physis. Rivista di Storia della Scienza, vol. I et II par Leo S. Olschki. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):372–373, October 1960. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904757.

Colnort:1960:AQPg

[1035] S. Colnort. Analyses de quelques périodiques: Gesnerus, revue trimestrielle, publiée par la Société suisse d’Histoire de la Médecine et des Sciences naturelles, vol. 16, fasc. 1/2. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):373, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904758.

Anonymous:1960:TGM

[1036] Anonyme. Table générale des matières du Tome XIII. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):374–382, October 1960. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904759.

Anonymous:1960:BMa

[1037] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 13(1):??, January 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904033.

Anonymous:1960:BMb

[1038] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):??, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904359.

Anonymous:1960:BMc

[1039] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):??, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904639.

Anonymous:1960:BMd

[1040] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):??, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904760.

Anonymous:1960:FMa

[1041] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 13(1):??, January 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904027.

Anonymous:1960:FMb

[1042] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 13(2):??, avril–juin 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904307.

Anonymous:1960:FMc

[1043] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 13(3):??, juillet–septembre 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904607.

Anonymous:1960:FMd

[1044] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 13(4):??, October 1960. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904728.

Scheler:1961:ALLa

[1045] Lucien Scheler. Antoine-Laurent Lavoisier et le “Journal d’Histoire naturelle”. Revue d’Histoire des Sciences et de Leurs Applications, 14(1): 1–9, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904998.

Scheler:1961:NPI

[1046] Lucien Scheler. Note sur un portrait inconnu de Lavoisier. Revue d’Histoire des Sciences et de Leurs Applications, 14(1):10–12, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904999.

Tabor:1961:PDF

[1047] David Tabor. À propos du frottement de roulement : Une controverse oubliée. Revue d’Histoire des Sciences et de Leurs Applications, 14(1): 13–18, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905000.

Lamontagne:1961:GDD

[1048] Roland Lamontagne. La Galissonière : Directeur du Dépôt de la marine. Revue d’Histoire des Sciences et de Leurs Applications, 14(1):19–26, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905001.

Humboldt:1961:NVJ

[1049] A. Humboldt and Jean Théodoridès. Notes sur Victor Jacquemont. Revue d’Histoire des Sciences et de Leurs Applications, 14(1):27–31, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905002.

Dulieu:1961:JMD

[1050] Louis Dulieu. Joseph-Marie Dubrueil: Médecin parisien, chirurgien de la Marine et professeur d’anatomie à Montpellier (1790–1852). Revue d’Histoire des Sciences et de Leurs Applications, 14(1):33–46, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905003.

Huard:1961:PB

[1051] Pierre Huard and S. Pozzi. Paul Broca (1824–1880). Revue d’Histoire des Sciences et de Leurs Applications, 14(1):47–86, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905004.

Torlais:1961:CHN

[1052] Jean Torlais. Un cabinet d’histoire naturelle français datant du XVIII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 14(1):87– 88, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905005.

Anonymous:1961:UIH

[1053] Anonyme. Union Internationale d’Histoire et de Philosophie des Sciences: Division d’Histoire des Sciences Colloques Internationaux pour 1961. Revue d’Histoire des Sciences et de Leurs Applications, 14(1):89, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905006.

Anonymous:1961:Fa

[1054] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 14(1):89–90, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905007.

Anonymous:1961:PVC

[1055] Anonyme. Portugal: V^{e} colloque international d’histoire maritime. Revue d’Histoire des Sciences et de Leurs Applications, 14(1):90, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905008.

Rochot:1961:AOBa

[1056] B. Rochot. Analyses d’ouvrages: Correspondance, Tome VII (1646– 1647) par Descartes; Charles Adam; Gérard Milhaud. Revue d’Histoire des Sciences et de Leurs Applications, 14(1):91, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905009.

Leroy:1961:AOB

[1057] A.-L. Leroy. Analyses d’ouvrages: The Edge of Objectivity, An Essay in the History of Scientific Ideas par Charles Coulston Gillispie. Revue d’Histoire des Sciences et de Leurs Applications, 14(1):91–92, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905010.

Rochot:1961:AOBb

[1058] B. Rochot. Analyses d’ouvrages: Histoire et avenir de la méthode expérimentale, ^{≪}Évolution des Sciences ^{≫}, n^{o} 17 par René Leclercq; Jean Rostand. Revue d’Histoire des Sciences et de Leurs Applications, 14(1): 92–93, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905011.

Huard:1961:AOBa

[1059] P. Huard. Analyses d’ouvrages: J. R. C. Quoy (1790–1869), inspecteur général du Service de Santé de la Marine, médecin, naturaliste, navigateur. Sa vie, son milieu, son œuvre par J.-P. Noël. Revue d’Histoire des Sciences et de Leurs Applications, 14(1):93–94, January 1961. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905012.

Theodorides:1961:AOBa

[1060] J. Théodoridès. Analyses d’ouvrages: Alexander von Humboldt in der russischen Literatur (bibliographie) par N. G. Suchowa. Revue d’Histoire des Sciences et de Leurs Applications, 14(1):94, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905013.

Huard:1961:AOBb

[1061] P. Huard and M. Wong. Analyses d’ouvrages: Les boissons alcooliques sino-vietnamiennes, n^{o} 3 par Paul-Louis Deniel. Revue d’Histoire des Sciences et de Leurs Applications, 14(1):95–96, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905014.

Weil:1961:CBC

[1062] Françoise Weil. La correspondance Buffon-Cramer. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):97–136, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905677.

Hanks:1961:BFV

[1063] Lesley Hanks. Buffon et les fusées volantes. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):137–154, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905678.

Ferenczi:1961:NJC

[1064] Rosemarie Ferenczi and R. J. Maras. Nicolas-Jacques Conté (1755–1805) : Un savant et un inventeur sous la Révolution, le Directoire et l’Empire. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):155–168, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905679.

Pier:1961:GEI

[1065] Jean-Paul Pier. Genèse et évolution de l’idée de compact. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):169–179, avril– juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905680.

Anonymous:1961:Fb

[1066] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):181–183, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905681.

Anonymous:1961:GBF

[1067] Anonyme. Grande-Bretagne: ^{≪}History of Science^{≫}. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):183, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905682.

Theodorides:1961:USS

[1068] J. Théodoridès. U.S.A.: Un symposium sur la science Byzantine. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):183–184, avril– juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905683.

Anonymous:1961:YCI

[1069] Anonyme. Yougoslavie: Colloque International Boscovitch. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):184, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905684.

Taton:1961:AOBa

[1070] René Taton. Analyses d’ouvrages: Lexikon der Geschichte der Naturwissenschaften, Biographen, Sachwörter und Bibliographen, Erste Lieferung par Josef Mayerhöfer. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):185–186, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905685.

Taton:1961:AOBb

[1071] R. Taton. Analyses d’ouvrages: The exact Sciences in Antiquity, second édition par O. Neugebauer. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):186, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905686.

Itard:1961:AOBa

[1072] Jean Itard. Analyses d’ouvrages: Correspondance du P. Marin Mersenne, religieux minime, t. VI (1636–1637) par Paul Tannery; Cornélis de Waard. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):186–187, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905687.

Taton:1961:AOBc

[1073] René Taton. Analyses d’ouvrages: Mathématiques et mathématiciens par Pierre Dedron; Jean Itard. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):187–188, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905688.

Taton:1961:AOBd

[1074] René Taton. Analyses d’ouvrages: Critica dei principî della matematica e questioni di logica par Ugo Cassina. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):188–189, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905689.

Poulle:1961:AOBa

[1075] E. Poulle. Analyses d’ouvrages: Cartes anciennes de l’Église de France; historique, répertoire, guide d’usage. (Bibliothèque de la Société d’Histoire ecclésiastique de la France) par François de Dainville; Gabriel Le Bras. Revue d’Histoire des Sciences et de Leurs Applications, 14 (2):189–190, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905690.

Theodorides:1961:AQP

[1076] J. Théodoridès. Analyses de quelques périodiques: Estudios Geograficos, n^{o} 76. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):190, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905691.

Taton:1961:AOBe

[1077] R. Taton. Analyses d’ouvrages: Université de Genève (1559–1959). Exposition du IV^{e} Centenaire. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):190–191, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905692.

Colnort:1961:AOBa

[1078] Suzanne Colnort. Analyses d’ouvrages: A Diderot Pictorial Encyclopedia of Trades and Industry par Denis Diderot; Charles Coulston Gillispie.

Revue d’Histoire des Sciences et de Leurs Applications, 14(2):191, avril– juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905694.

Theodorides:1961:AOBb

[1079] J. Théodoridès. Analyses d’ouvrages: L’humanisme en Alsace. Exposition placée sous le patronage de l’Association Guillaume Budé, section Alsace, Musée de l’œuvre Notre-Dame, juin-juillet 1958. Catalogue, Imprimerie des Dernières Nouvelles. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):191, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905693.

Anonymous:1961:PR

[1080] Anonyme. Publications reçues. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):192, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905695.

Birembaut:1961:QDD

[1081] Arthur Birembaut. Quelques documents sur Desargues. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):193–204, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903932.

Robinet:1961:PMM

[1082] André Robinet. La philosophie malebranchiste des mathématiques. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):205– 254, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903933.

Lamontagne:1961:CCG

[1083] Roland Lamontagne. Chronologie de la carrière de la Galissonière. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):255– 256, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903934.

Scheler:1961:ALLb

[1084] Lucien Scheler. Antoine-Laurent Lavoisier et Michel Adanson rédacteurs de programmes des Prix à l’Académie des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):257–284, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903935.

Laissus:1961:DLL

[1085] Yves Laissus. Deux lettres de Laplace. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):285–296, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903936.

Gerard:1961:QPC

[1086] Jean-Pierre Gérard. Sur quelques problèmes concernant l’œuvre d’ørsted en électromagnétisme. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):297–312, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903937.

Rychlik:1961:TNR

[1087] Karel Rychlík. La théorie des nombres réels dans un ouvrage posthume manuscrit de Bernard Bolzano. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):313–327, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903938.

Humboldt:1961:LIH

[1088] A. Humboldt and Jean Théodoridès. Une lettre inédite de Humboldt au mathématicien Sylvestre-François Lacroix. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):329–330, juillet–décembre 1961. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903939.

Rostand:1961:PTH

[1089] Jean Rostand. La première tentative d’hybridation sanguine (Galton, 1871). Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4): 331–332, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903940.

Delorme:1961:FP

[1090] Suzanne Delorme. Fontenelle et la Pologne. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):332, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903941.

Costabel:1961:UIH

[1091] P. Costabel and R. Taton. Union Internationale d’Histoire et de Philosophie des Sciences: Division d’Histoire des Sciences: Symposium d’Oxford. Revue d’Histoire des Sciences et de Leurs Applications, 14(3– 4):333–336, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903942.

Delorme:1961:F

[1092] S. Delorme and P. Huard. France. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):336–341, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903943.

Anonymous:1961:C

[1093] Anonyme. Conférences. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):341–343, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23910982.

Anonymous:1961:CHS

[1094] Anonyme. Cours d’histoire des sciences. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):343–345, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23910983.

Anonymous:1961:RP

[1095] Anonyme. Réunions projetées. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):345–346, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903944.

Delorme:1961:AOB

[1096] Suzanne Delorme. Analyses d’ouvrages: Science fausse et fausses sciences, (Les Essais, LXXXIX) par Jean Rostand. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):347–348, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903945.

Itard:1961:AOBb

[1097] Jean Itard. Analyses d’ouvrages: Vollständige Anleitung zur Algebra par Leonhard Euler; Joh. Niesser; Jos. E. Hofmann. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):348, juillet–décembre 1961. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903946.

Itard:1961:AOBc

[1098] Jean Itard. Analyses d’ouvrages: On Mathematics and Mathematicians (Memorabilia Mathematica) par Robert Édouard Moritz. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):348, juillet– décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903947.

Taton:1961:AOBf

[1099] René Taton. Analyses d’ouvrages: Formulario mathematico. Riproduzione in facsimile dell’edizione originale par Giuseppe Peano; Ugo Cassina. Revue d’Histoire des Sciences et de Leurs Applications, 14(3– 4):348–349, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903948.

Poulle:1961:AOBb

[1100] E. Poulle. Analyses d’ouvrages: Arabische Sternnamen in Europa par Paul Kunitzsch. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):349–351, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903949.

Poulle:1961:AOBc

[1101] Emmanuel Poulle. Analyses d’ouvrages: The astronomical works of Thabit b. Qurra par Francis J. Carmody. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):351–353, juillet–décembre 1961. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903950.

Bugler:1961:AOB

[1102] G. Bugler. Analyses d’ouvrages: Œuvres complètes. Édition conforme aux textes originaux imprimés du vivant de l’auteur; avec des Notes et une Notice historique par Bernard Palissy; Paul-Antoine Cap; Jean Orcel. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4): 353–355, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903951.

Moscovici:1961:AOBa

[1103] S. Moscovici. Analyses d’ouvrages: Discourse on bodies in water par Galileo Galilei; Thomas Salusbury; Stillman Drake. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):355, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903952.

Moscovici:1961:AOBb

[1104] S. Moscovici. Analyses d’ouvrages: On motion and on mechanics. (Publication in the Medieval Science, 5) par Galileo Galilei; I. E. Drabkin; S. Drake. Revue d’Histoire des Sciences et de Leurs Applications, 14 (3–4):355, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903953.

Taton:1961:AOBg

[1105] R. Taton. Analyses d’ouvrages: Rudzer Bosković. Grada knjiga, II (= ≪Matériaux pour servir à la connaissance de la vie et l’œuvre de R. J. Bosković ^{≫}, II) apr Zeljko Marković. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):356, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903954.

Theodorides:1961:AOBc

[1106] J. Théodoridès. Analyses d’ouvrages: Alexander von Humboldt (1769– 1859), Seine Bedeutung für den Bergbau und die Naturforschung. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):356– 357, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903955.

Theodorides:1961:AOBd

[1107] Jean Théodoridès. Analyses d’ouvrages: Alexander von Humboldt, Vorträge und Aufsätze anlässlich der 100. Wiederkehr seines Todestages am 6 Mai 1959 par J. F. Gellert. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):357–358, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903956.

Theodorides:1961:AOBe

[1108] J. Théodoridès. Analyses d’ouvrages: Alejandro de Humboldt en Tenerife, Canarios Monograf. Seccion I (Cienc. Hist. y Geogr.), vol. XV par Alejandro Cioranescu. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):358–359, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903957.

Moreau:1961:AOBa

[1109] F. Moreau. Analyses d’ouvrages: Microbiology, Historical Contributions from 1776 to 1908, by Spallanzani, Schwann, Pasteur, Cohn, Tyndall, Koch, Lister, Schloesing, Burrill, Ehrlich, Winogradsky, Warington, Beijerinck, Smith, Orla-Jensen par R. N. Doetsch. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):359–360, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903958.

Torlais:1961:AOB

[1110] Jean Torlais. Analyses d’ouvrages: The growth of Scientific Physiology. Physiological Method and the Mechanist-Vitalist Controversy illustrated par the Problems of Respiration and Animal Heat par G. J. Goodfield. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4): 361–363, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903959.

Theodorides:1961:AOBf

[1111] Jean Théodoridès. Analyses d’ouvrages: Friedrich Traugott Kützing (1807–1893), Aufzeichnungen und Erinnerungen par R. H. Walter Müller; Rudolph Zaunick. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):363, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903960.

Theodorides:1961:AOBg

[1112] Jean Théodoridès. Analyses d’ouvrages: Le premier manuscrit chirurgical turc rédigé par Charaf-ed-Din (1465), et illustré de 140 miniatures par Pierre Huard; Mirko Drazén Grmek. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):363–364, juillet–décembre 1961. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903961.

Dulieu:1961:AOB

[1113] L. Dulieu. Analyses d’ouvrages: Recherches sur le traité pseudohippocratique du Régime, fasc. CLVI par Robert Joly. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):364–365, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903962.

Huard:1961:AOBc

[1114] Pierre Huard. Analyses d’ouvrages: Chirurgiens de la Compagnie des Indes. Histoire du Service de Santé de la Compagnie (1664–1793) par Cl. Chaligne. Revue d’Histoire des Sciences et de Leurs Applications, 14(3– 4):365–367, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903963.

Huard:1961:AOBd

[1115] Pierre Huard. Analyses d’ouvrages: Les chirurgiens de l’Hôtel–Dieu de Nantes sous l’Ancien Régime. Esquisse d’histoire de la médecine à Nantes au XVIII^{e} siècle par Serge Grunblatt. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):367–368, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903964.

Huard:1961:AOBe

[1116] P. Huard. Analyses d’ouvrages: Hygiène et santé à bord des navires négriers au XVIII^{e} siècle par M. Cl. Chiche. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):368–369, juillet–décembre 1961. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903965.

Moreau:1961:AOBb

[1117] F. Moreau. Analyses d’ouvrages: Itinéraires des contagions, épidémies et idéologies par A. Siegfried; Pasteur Vallery-Radot. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):369–371, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903966.

Taton:1961:AOBh

[1118] R. Taton. Analyses d’ouvrages: L’électronique. (^{≪}Le rayon de la science ≫) par Maurice Ponte; Pierre Braillard. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):371, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903967.

Theodorides:1961:AOBh

[1119] J. Théodoridès. Analyses d’ouvrages: Memorabilia Zoologica (1958– 1961). Revue d’Histoire des Sciences et de Leurs Applications, 14(3– 4):371, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903968.

Colnort:1961:AOBb

[1120] S. Colnort. Analyses d’ouvrages: Centaurus, International Magazine of the History of Science and Medicine, vol. 6, 1959, fasc. 1, 2. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):372, juillet– décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903970.

Colnort:1961:AQPa

[1121] S. Colnort. Analyses de quelques périodiques: Archives internationales d’Histoire des Sciences, t. XIII, année 1960, n^{os} 50/51 et 52/53. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):372, juillet– décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903969.

Colnort:1961:AQPb

[1122] S. Colnort. Analyses de quelques périodiques: Gesnerus, vol. 17, 1960, fasc. 1/2 et 3/4. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):373, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903971.

Colnort:1961:AQPc

[1123] S. Colnort. Analyses de quelques périodiques: Isis, an International Review devoted to the History of Science and its cultural Influences, vol. 51, 1960, fasc. 163, 164, 165, 166 par Harry Woolf. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):373–374, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903972.

Anonymous:1961:TGM

[1124] Anonyme. Table générale des matières du Tome XIV. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):375–384, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903973.

Anonymous:1961:BMa

[1125] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 14(1):??, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905015.

Anonymous:1961:BMb

[1126] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):??, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905696.

Anonymous:1961:BMc

[1127] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):??, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903974.

Anonymous:1961:FMa

[1128] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 14(1):??, January 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904997.

Anonymous:1961:FMb

[1129] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 14(2):??, avril–juin 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905676.

Anonymous:1961:FMc

[1130] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 14(3–4):??, juillet–décembre 1961. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903931.

Galilei:1962:GCT

[1131] Galileo Galilei and Maurice Clavelin. Galilée et la cosmologie traditionnelle la première journée du dialogue: Note introductive. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):1–26, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904800.

Lamontagne:1962:LLD

[1132] Roland Lamontagne. L’exploration de l’Amérique du nord à l’époque de Jean talon. Revue d’Histoire des Sciences et de Leurs Applications, 15(1): 27–30, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904801.

Costabel:1962:CRB

[1133] Pierre Costabel. Le de Cycloïde de R. Boscovich. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):31–42, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904802.

Scheler:1962:PLF

[1134] Lucien Scheler. À propos d’une lettre de Fourcroy à Lavoisier du 3 septembre 1793. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):43–50, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904803.

Grmek:1962:HRR

[1135] M. D. Grmek. Histoire des recherches sur les relations entre le génie et la maladie. Revue d’Histoire des Sciences et de Leurs Applications, 15(1): 51–68, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904804.

Rostand:1962:RPE

[1136] Jean Rostand. Réaumur et les premiers essais de léthargie artificielle. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):69–71, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904805.

Rostand:1962:RRI

[1137] Jean Rostand. Réaumur et la résistance des insectes à la congélation. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):71–72, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904806.

Torlais:1962:AF

[1138] Jean Torlais. Apothicaireries de France. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):72–73, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904807.

Anonymous:1962:A

[1139] Anonyme. Allemagne. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):74, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904808.

Huard:1962:F

[1140] P. Huard. France. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):74–75, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904809.

Costabel:1962:YIS

[1141] Pierre Costabel. Yougoslavie: II^{e} Symposium International R. Boscovich. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):75–77, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904810.

Moscovici:1962:AOB

[1142] Serge Moscovici. Analyses d’ouvrages: The Correspondance of Isaac Newton, t. I : 1661–1675; t. II : 1676–1687 par I. Newton; H. W. Turnbull. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):78–81, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904811.

Herivel:1962:AOB

[1143] J. W. Herivel. Analyses d’ouvrages: Leibniz et la dynamique. Les textes de 1692, (Histoire de la Pensée, I) par Pierre Costabel. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):81–82, January 1962. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904812.

Bauer:1962:AOBa

[1144] E. Bauer. Analyses d’ouvrages: Henry Cavendish, His Life and scientific Work par A. J. Berry. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):82–83, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904813.

Speziali:1962:AOB

[1145] Pierre Speziali. Analyses d’ouvrages: Lettres à Maurice Solovine, reproduites en facsimilé et traduites en français avec une introduction et trois photographies par Albert Einstein. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):84–85, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904814.

Theodorides:1962:AOBa

[1146] J. Théodoridès. Analyses d’ouvrages: Dessins analomiques (anatomie artistique, descriptive et fonctionnelle) par Léonard de Vinci; Pierre Huard. Revue d’Histoire des Sciences et de Leurs Applications, 15(1): 85–86, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904815.

Theodorides:1962:AOBb

[1147] Jean Théodoridès. Analyses d’ouvrages: Lettres de Théodore Schwann par Marcel Florkin. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):86, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904816.

Huard:1962:AOBa

[1148] P. Huard. Analyses d’ouvrages: Colloque sur la microscopie et la micrographie aux XVIII^{e} et XIX^{e} siècles C. R. du LXXXV^{e} Congrès des Sociétés savantes, Chambéry–Annecy, 1960 par Luigi Belloni; J. Théodoridès. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):86–87, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904817.

Colnort:1962:AOBa

[1149] S. Colnort. Analyses d’ouvrages: Les Conférences du Palais de la Découverte. Série D: Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):87–90, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904818.

Colnort:1962:AOBb

[1150] S. Colnort. Analyses d’ouvrages: Lychnos, Lärdomshistoriska Samfundets Arsbok grundad, Annuaire de la Société suédoise d’Histoire des Sciences par Johan Nordström. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):91, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904820.

Colnort:1962:AQPa

[1151] S. Colnort. Analyses de quelques périodiques: Archive for the History of Exact Sciences, vol. I, s. d. (1960–1961), Nr. 1, 2, 3, 4 par C. Truesdell. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):91, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904819.

Colnort:1962:AQPb

[1152] S. Colnort. Analyses de quelques périodiques: Chymia, Annual Studies in the History of Chemistry, vol. 5, 1959; vol. 6, 1960 vol. 7, 1961 par Henry M. Leicester. Revue d’Histoire des Sciences et de Leurs Applications, 15 (1):92, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904821.

Naux:1962:LGG

[1153] Ch. Naux. L’Opus geometricum de Grégoire de Saint-Vincent. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):93–104, avril– juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903734.

Herivel:1962:PRN

[1154] J. W. Herivel. Sur les premières recherches de Newton en dynamique. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):105–140, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903735.

Poulle-Drieux:1962:HFC

[1155] Yvonne Poulle-Drieux, Thillaye, Fragonard, Delzeuzes, Landrieux, Flandrin, Peyrilhe, Mahon, Le Clerc, G. Cuvier, and Eti. Geoffroy. Honoré Fragonard et le cabinet d’anatomie de l’École d’Alfort pendant la Révolution. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):141–162, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903736.

Rychlik:1962:CPC

[1156] Karel Rychlík. Sur les contacts personnels de Cauchy et de Bolzano. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):163–164, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903737.

Taton:1962:NGD

[1157] R. Taton. Sur la naissance de Girard Desargues. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):165–166, avril–juin 1962. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903738.

Costabel:1962:JP

[1158] Pierre Costabel. Joseph Pérès (1890–1962). Revue d’Histoire des Sciences et de Leurs Applications, 15(2):167–169, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903739.

Anonymous:1962:AFM

[1159] Anonyme. Autriche: Fédération Mondiale Coronellienne. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):170, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903740.

Costabel:1962:F

[1160] P. Costabel. France. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):170–173, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903741.

Anonymous:1962:BCS

[1161] Anonyme. Belgique: Colloque sur le soleil à l’Époque de la Renaissance. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):174, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23911246.

Grmek:1962:AOB

[1162] M. D. Grmek. Analyses d’ouvrages: Toward Modern Science, vol. I : Studies in Ancient and Medieval Science; Vol. II : Studies in Renaissance Science par R. M. Palter. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):175, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903742.

Itard:1962:AOB

[1163] Jean Itard. Analyses d’ouvrages: Quaestiones super geometriam Euclidis par Nicole Oresme; H. L. L. Busard. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):176–177, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903743.

Bauer:1962:AOBb

[1164] E. Bauer. Analyses d’ouvrages: The De Magnete of William Gilbert par Duane H. D. Roller. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):177, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903744.

Theodorides:1962:AOBc

[1165] J. Théodoridès. Analyses d’ouvrages: Léonard de Vinci, dessins scientifiques et techniques par Pierre Huard; Mirko Drazen Grmek. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):178, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903745.

Theodorides:1962:AOBd

[1166] Jean Théodoridès. Analyses d’ouvrages: Chine d’hier et d’aujourd’hui par Pierre Huard; Ming Wong. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):178–179, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903746.

Theodorides:1962:AOBe

[1167] J. Théodoridès. Analyses d’ouvrages: La Vie par Jean Rostand; Andrée Tétry. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):179, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903747.

Huard:1962:AOBb

[1168] Pierre Huard. Analyses d’ouvrages: Pierre Fauchard, père de l’art dentaire moderne (1678–1761) et ses contemporains par André Besombes; Georges Dagen. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):179–180, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903748.

Guyenot:1962:AOB

[1169] E. Guyénot. Analyses d’ouvrages: Pierre Lyonet (1706–1789). Sa vie, ses collections de coquillages et de tableaux, ses recherches entomologiques par W. H. Van Seters. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):180–182, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903749.

Huard:1962:AOBc

[1170] P. Huard. Analyses d’ouvrages: Biologie médicale, numéro hors série, vol. L, 59^{e} année par G. Petit; J. Théodoridès; Georges Cuvier. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):182, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903750.

Theodorides:1962:AOBf

[1171] J. Théodoridès. Analyses d’ouvrages: Diez Años de Amistad con Sir Alexander Fleming par Florencio Bustinza. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):182–183, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903751.

Theodorides:1962:AOBg

[1172] J. Théodoridès. Analyses d’ouvrages: Hygiène et maladies de la femme au cours des siècles par André Pecker. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):183, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903752.

Huard:1962:AOBd

[1173] P. Huard. Analyses d’ouvrages: Les reflets de la sphygmologie chinoise dans la médecine occidentale, Biologie médicale, numéro hors série par Mirko Drazen Grmek. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):183–184, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903753.

Huard:1962:AOBe

[1174] P. Huard. Analyses d’ouvrages: Histoire de la dermatologie au Japon, in Annales de dermatologie et de syphilographie par Zensetsou Ohya. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):185–186, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903754.

Huard:1962:AOBf

[1175] P. Huard and J. Théodoridès. Analyses d’ouvrages: L’Hôtel– Dieu de Paris (651-1650). Catalogue de l’exposition organisée par l’administration générale de l’Assistance publique par Marcel Candille; Claude Hohl. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):186, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903755.

Theodorides:1962:AOBh

[1176] J. Théodoridès. Analyses d’ouvrages: Colloque sur les médecins et naturalistes de l’École de Montpellier, extrait des Comptes rendus du LXXXVI^{e} Congrès des Sociétés Savantes, Montpellier, 1961 (Section des Sciences). Revue d’Histoire des Sciences et de Leurs Applications, 15(2):186–187, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903756.

Colnort:1962:AOBc

[1177] S. Colnort. Analyses d’ouvrages: Thalès, Recueil des travaux et bibliographie, t. XI. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):187, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903757.

Colnort:1962:AQPc

[1178] S. Colnort. Analyses de quelques périodiques: Physis. Rivista di Storia della Scienza, vol. III. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):188, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903758.

Taton:1962:TCS

[1179] René Taton. Tableau chronologique sommaire de la vie et des travaux scientifiques de Pascal. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4):191–195, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904543.

Pascal:1962:LPG

[1180] René Taton. L’œuvre de Pascal en géométrie projective. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4):197–252, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904544.

Costabel:1962:TFN

[1181] Pierre Costabel. Traduction française de notes de Leibniz sur les Coniques de Pascal. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4):253–268, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904545.

Itard:1962:BGP

[1182] Jean Itard. L’introduction à la géométrie de Pascal. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4):269–286, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904546.

Hara:1962:PIM

[1183] Kokiti Hara. Pascal et l’induction mathématique. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4):287–302, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904547.

Russo:1962:PAI

[1184] François Russo. Pascal et l’analyse infinitésimale. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4):303–320, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904548.

Costabel:1962:ESB

[1185] Pierre Costabel. Essai sur les secrets des Traités de la Roulette. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4):321– 350, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904549.

Scheler:1962:BDO

[1186] Lucien Scheler. Les Lettres de Dettonville offertes à Jean-Baptiste Colbert. Revue d’Histoire des Sciences et de Leurs Applications, 15:351–365, 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

Scheler:1962:LDO

[1187] Lucien Scheler. Les lettres de Dettonville offertes à Jean-Baptiste Colbert. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4): 351–365, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904550.

Costabel:1962:RLA

[1188] Pierre Costabel. Résumé de la lettre adressée par Pascal à Lalouère le 4 septembre 1658. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4):367–369, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904551.

Costabel:1962:NRI

[1189] Pierre Costabel. Notes relatives à l’influence de Pascal sur Leibniz. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4): 369–374, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904552.

Delorme:1962:FTM

[1190] Suzanne Delorme. France: Tricentenaire de la mort de Pascal. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4):375– 382, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904553.

Delorme:1962:AOB

[1191] Suzanne Delorme. Analyses d’ouvrages: Pascal et Port-Royal. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4):383– 384, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904567.

Anonymous:1962:TGM

[1192] Anonyme. Table générale des matières du Tome XV. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4):385–392, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904568.

Anonymous:1962:BMa

[1193] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):??, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904822.

Anonymous:1962:BMb

[1194] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):??, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903759.

Anonymous:1962:BMc

[1195] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4):??, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904569.

Anonymous:1962:FMa

[1196] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):??, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904799.

Anonymous:1962:FMb

[1197] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 15(2):??, avril–juin 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903733.

Anonymous:1962:FMc

[1198] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4):??, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904541.

Anonymous:1962:I

[1199] Anonyme. [illustration]. Revue d’Histoire des Sciences et de Leurs Applications, 15(3–4):??, juillet–décembre 1962. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904542.

Mesnard:1963:PLP

[1200] Jean Mesnard. Pascal à l’Académie Le Pailleur. Revue d’Histoire des Sciences et de Leurs Applications, 16(1):1–10, January 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904071.

Leibniz:1963:ECL

[1201] Jean Mesnard and René Taton. Édition critique de la lettre de Leibniz à Périer du 30 août 1676. Revue d’Histoire des Sciences et de Leurs Applications, 16(1):11–22, January 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904072.

Sadoun-Goupil:1963:LPP

[1202] Michelle Sadoun-Goupil. L’œuvre de Pascal et la physique moderne. Revue d’Histoire des Sciences et de Leurs Applications, 16(1):23–52, January 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904073.

Rochot:1963:CGI

[1203] Bernard Rochot. Comment Gassendi interprétait l’expérience du puy de Dôme. Revue d’Histoire des Sciences et de Leurs Applications, 16(1): 53–76, January 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904074.

Taton:1963:LEB

[1204] René Taton. L’annonce de l’expérience barométrique en France. Revue d’Histoire des Sciences et de Leurs Applications, 16(1):77–83, January 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904075.

Kersaint:1963:PAL

[1205] G. Kersaint. À propos d’un article de Lucien Scheler sur une lettre de Fourcroy à Lavoisier. Revue d’Histoire des Sciences et de Leurs Applications, 16(1):83–84, January 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904076.

Russo:1963:UIH

[1206] F. Russo and J. Théodoridès. Union Internationale d’Histoire et de Philosophie des Sciences: Division d’Histoire des Sciences: X^{e} Congrès International d’Histoire des Sciences (Ithaca-Philadelphie, 24 août–2 septembre 1962). Revue d’Histoire des Sciences et de Leurs Applications, 16(1):85–92, January 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904077.

Colnort:1963:ICH

[1207] Suzanne Colnort. Italie: Colloque sur l’Histoire de la Science de la Renaissance (XVI^{e} et XVII^{e} siècles). Revue d’Histoire des Sciences et de Leurs Applications, 16(1):92–93, January 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904078.

Delorme:1963:AOB

[1208] Suzanne Delorme. Analyses d’ouvrages: Pascal présent, 1662–1962, (Écrivains d’Auvergne, n^{o} 3). Revue d’Histoire des Sciences et de Leurs Applications, 16(1):95–96, January 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904079.

Moscovici:1963:RDG

[1209] Serge Moscovici. Remarques sur le dialogue de Galilée de la force de la percussion. Revue d’Histoire des Sciences et de Leurs Applications, 16(2):97–137, avril–juin 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905067.

Taton:1963:IMA

[1210] René Taton. Sur l’invention de la machine arithmétique. Revue d’Histoire des Sciences et de Leurs Applications, 16(2):139–160, avril–juin 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905068.

Payen:1963:ECM

[1211] Jacques Payen. Les exemplaires conservés de la machine de Pascal. Revue d’Histoire des Sciences et de Leurs Applications, 16(2):161–178, avril– juin 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905069.

Rostand:1963:BF

[1212] Jean Rostand. Bacon et le froid. Revue d’Histoire des Sciences et de Leurs Applications, 16(2):179, avril–juin 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905070.

Colnort:1963:F

[1213] Suzanne Colnort. France. Revue d’Histoire des Sciences et de Leurs Applications, 16(2):180–182, avril–juin 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905071.

Courtin:1963:AOB

[1214] Madeleine Courtin. Analyses d’ouvrages: La Science des Chaldéens, (^{≪}Que sais-je? ^{≫}, n^{o} 893) par Marguerite Rutten. Revue d’Histoire des Sciences et de Leurs Applications, 16(2):183, avril–juin 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905072.

Itard:1963:AOBa

[1215] Jean Itard. Analyses d’ouvrages: Les nombres et leurs mystères, (Le Rayon de la Science, n^{o} 9) par André Warusfel. Revue d’Histoire des Sciences et de Leurs Applications, 16(2):184, avril–juin 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905073.

Michel:1963:AOB

[1216] Paul-Henri Michel. Analyses d’ouvrages: Opera latine conscripta. Neapel und Florenz, 1879–1891. Drei Bände in acht Teilen. Zweiter Band, erster Teil par Jordani Bruni Nolani; Tocco Fiorentino. Revue d’Histoire des Sciences et de Leurs Applications, 16(2):184–186, avril–juin 1963. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905074.

Moscovici:1963:AOBa

[1217] S. Moscovici. Analyses d’ouvrages: La cosmologie de Giordano Bruno, (Histoire de la Pensée) par Paul-Henri Michel. Revue d’Histoire des Sciences et de Leurs Applications, 16(2):186, avril–juin 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905075.

Itard:1963:AOBb

[1218] Jean Itard. Analyses d’ouvrages: Mathématiques et Métaphysique chez Descartes par Jules Vuillemin. Revue d’Histoire des Sciences et de Leurs Applications, 16(2):186–187, avril–juin 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905076.

Itard:1963:AOBc

[1219] Jean Itard. Analyses d’ouvrages: Leibniz critique de Descartes par Yvon Belaval. Revue d’Histoire des Sciences et de Leurs Applications, 16(2):187–192, avril–juin 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905077.

Pines:1963:FSS

[1220] S. Pinès. Un fragment de Séleucus de Séleucie conservé en version arabe. Revue d’Histoire des Sciences et de Leurs Applications, 16(3): 193–209, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905158.

Torlais:1963:QIB

[1221] Jean Torlais. Qui a inventé la bouteille de Leyde? Revue d’Histoire des Sciences et de Leurs Applications, 16(3):211–219, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905159.

Lamontagne:1963:RBA

[1222] Roland Lamontagne. Rapport sur le Traité des arbres et arbustes … de Duhamel du Monceau. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):221–225, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905160.

Poulle-Drieux:1963:PDB

[1223] Yvonne Poulle-Drieux. À propos du bicentenaire de l’École vétérinaire de Lyon. Revue d’Histoire des Sciences et de Leurs Applications, 16(3): 227–232, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905161.

Metz:1963:NAT

[1224] André Metz. La notation atomique et la théorie atomique en France à la fin du XIX^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):233–239, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905162.

Rostand:1963:QPC

[1225] Jean Rostand. Quelques précurseurs: Charles Morel de Vindé, Camille Dareste, le D^{r} Dufossé, G. Vacher de Lapouge. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):241–251, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905163.

Rochot:1963:CW

[1226] Bernard Rochot. Cornelis de waard 1879–1963. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):253–256, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905164.

Wong:1963:PLY

[1227] Ming Wong. Le professeur Li yen 1892–1963. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):256–257, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905165.

Costabel:1963:F

[1228] P. Costabel, S. Colnort, and Suzanne Delorme. France. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):258–266, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905166.

Anonymous:1963:PBS

[1229] Anonyme. Pays-Bas: Société Néerlandaise pour l’Histoire des Sciences Exactes Naturelles et Médicales. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):266, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905173.

Poulle:1963:AOBa

[1230] Emmanuel Poulle. Analyses d’ouvrages: Comité international de paléographie. Catalogue des manuscrits en écriture latine portant des indications de date, de lieu ou de copiste, T. II : Bibliothèque nationale, fonds latin (n^{os} 1 à 8000) par Charles Samaran; Robert Marichal; Marie– Thérèse d’Alverny; Monique Garand; Madeleine Mabille; Josette Metman. Revue d’Histoire des Sciences et de Leurs Applications, 16(3): 267–269, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905174.

Poulle:1963:AOBb

[1231] Emmanuel Poulle. Analyses d’ouvrages: Manuscrits scientifiques médiévaux de l’Université de Salamanque et de ses ^{≪}colegios mayores ^{≫}. ≪Bibliothèque de l’École des Hautes Études hispaniques ^{≫}, fasc. XXXII par Guy Beaujouan. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):269–271, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905175.

Itard:1963:AOBd

[1232] Jean Itard. Analyses d’ouvrages: Correspondance du P. Marin Mersenne, religieux minime, T. VII : Janvier-juillet 1638 par Cornélis de Waard. Revue d’Histoire des Sciences et de Leurs Applications, 16(3): 271–273, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905176.

Rochot:1963:AOB

[1233] Bernard Rochot. Analyses d’ouvrages: Correspondance, Tome VIII et dernier par Descartes; Charles Adam; Gérard Milhaud. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):273–274, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905177.

Moscovici:1963:AOBb

[1234] S. Moscovici. Analyses d’ouvrages: Die Tat des Kopernikus par G. Harig. Revue d’Histoire des Sciences et de Leurs Applications, 16(3): 274–275, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905178.

Moscovici:1963:AOBc

[1235] S. Moscovici. Analyses d’ouvrages: Roger Joseph Boscovich par L. L. Whyte. Revue d’Histoire des Sciences et de Leurs Applications, 16 (3):275, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905179.

Levy:1963:AOBa

[1236] J. Lévy. Analyses d’ouvrages: The Planet Saturn. A History of Observation, Theory and Discovery par A. F. O’D. Alexander. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):275–277, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905180.

Levy:1963:AOBb

[1237] Jacques Lévy. Analyses d’ouvrages: William Herschel. (British Men of Science) par Angus Armitage. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):277–279, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905181.

Theodorides:1963:AOBa

[1238] J. Théodoridès. Analyses d’ouvrages: Alexander von Humboldt. vol. I : 1959, vol. II : 1961 par Hanno Beck. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):279–281, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905182.

Moscovici:1963:AOBd

[1239] S. Moscovici. Analyses d’ouvrages: Philosophical Impact of contemporary Physics par Milec Capek. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):281, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905183.

Moscovici:1963:AOBe

[1240] S. Moscovici. Analyses d’ouvrages: Forces and Fields par B. Mary Hesse. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):281, juillet– septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905184.

Torlais:1963:AOB

[1241] Jean Torlais. Analyses d’ouvrages: Histoire de la Zoologie. Des origines à Linné, (Histoire de la pensée, VIII) par Georges Petit; Jean Théodoridès; Pr Pierre P. Grassé. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):282–283, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905185.

Theodorides:1963:AOBb

[1242] Jean Théodoridès. Analyses d’ouvrages: L’Institut Pasteur des origines à aujourd’hui. (Histoire et témoignages) par Albert Delaunay; Pasteur Vallery-Radot. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):283–284, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905186.

Theodorides:1963:AOBc

[1243] Jean Théodoridès. Analyses d’ouvrages: La vie et l’œuvre de Albert Calmette (1863–1933). (Les savants et le monde) par Noël Bernard. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):284–285, juillet– septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905187.

Poulle-Drieux:1963:AOB

[1244] Yvonne Poulle-Drieux. Analyses d’ouvrages: From Farriery to Veterinary Medicine, 1785–1795 … par L. P. Pugh. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):285–286, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905188.

Huard:1963:AOB

[1245] P. Huard. Analyses d’ouvrages: Contribution à l’étude du journalisme médical français au XIX^{e} siècle par J.-M. Prigent. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):286–287, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905189.

Youschkevitch:1963:AOB

[1246] A. Youschkevitch. Analyses d’ouvrages: History of Science. An Annual Review of Literature, Research and Teaching, vol. I, 1962 par A. C. Crombie; M. A. Hoskin. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):287–288, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23905190.

Anonymous:1963:DPH

[1247] Anonyme. Documents pour l’histoire des techniques: Cahier n^{o} 3. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):289–290, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903628.

Daumas:1963:MRT

[1248] Maurice Daumas. Le mythe de la révolution technique. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):291–302, October 1963. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903629.

Woodbury:1963:IIM

[1249] R. S. Woodbury. Une invention importante de la mécanique de précision: L’origine des calibres d’épaisseur. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):303–316, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903630.

Soulard:1963:CB

[1250] Robert Soulard. Le cinématographe Bouly. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):317–322, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903631.

Renard:1963:DPT

[1251] G. Renard. La découverte et le perfectionnement des transistors. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):323–358, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903632.

Sevin:1963:LLS

[1252] J. Sevin. L’effet laser dans les solides. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):359–372, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903633.

Pilisi:1963:LOP

[1253] Jean Pilisi. L’enquête ordonnée par le Régent sur les richesses naturelles de la France (1716–1717). Revue d’Histoire des Sciences et de Leurs Applications, 16(4):373–374, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903634.

Payen:1963:BLC

[1254] Jacques Payen. Les brevets de Lenoir concernant le moteur à combustion interne. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):374–380, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903635.

Bishop:1963:MHT

[1255] P. W. Bishop. Le Musée d’histoire et de technologie de la ^{≪}Smithsonian institution ^{≫}. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):380–386, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903636.

Payen:1963:F

[1256] Jacques Payen, J. Payen, and O. Popovitch. France. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):387–394, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903637.

Anonymous:1963:GBN

[1257] Anonyme. Grande-Bretagne: Un nouveau périodique d’histoire des techniques. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):394, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903640.

Daumas:1963:AOBa

[1258] M. Daumas. Analyses d’ouvrages: Catalogue des manuscrits relatifs aux recherches et aux travaux de l’inventeur Isaac de Rivaz par Henri Michelet. Revue d’Histoire des Sciences et de Leurs Applications, 16(4): 395, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903641.

Daumas:1963:AOBb

[1259] M. Daumas. Analyses d’ouvrages: Catalogue of Instruments and Models in the Possession of the American Philosophical Society, vol. 53 par Robert P. Multhauf. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):395–396, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903642.

Weill:1963:AOBa

[1260] Adrienne R. Weill. Analyses d’ouvrages: Zur Frühgeschichte der Metallurgie, suivi de sa traduction française : Considérations sur la métallurgie préhistorique par Jean R. Maréchal. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):396–397, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903643.

Weill:1963:AOBb

[1261] Adrienne R. Weill. Analyses d’ouvrages: Metallurgy in Archaelogy (A Prehistory of Metallurgy in the British Isles) par R. F. Tylecote. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):397–399, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903644.

Poulle:1963:AOBc

[1262] Emmanuel Poulle. Analyses d’ouvrages: Ein Handbuch für Sammler und Liebhaber. Uhren … ^{≪}Bibliothek für Kunst- und Antiquitätenfreunde^{≫}, Band VII par Ernst von Bassermann-Jordan; Hans von Bertele. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):399–400, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903645.

Jacques:1963:AOB

[1263] J. Jacques. Analyses d’ouvrages: The Emergence of the German Dye Industry, ^{≪}Illinois Studies in the Social Sciences ^{≫}, vol. 44 par John J. Beer. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):400– 401, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903646.

Payen:1963:AQP

[1264] J. Payen. Analyses de quelques périodiques: Technology and Culture, The International Quarterly of the Society for the History of Technology. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):401, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903647.

Anonymous:1963:CBH

[1265] Anonyme. Contributions to Bibliography in the History of Technology, T. and C., III, n^{o} 1, (Part I); III, n^{o} 2, (Part II); III, n^{o} 3, (Part III); IV, n^{o} 3, (Part IV) par Eugène S. Ferguson. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):401–402, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903648.

Anonymous:1963:LFR

[1266] Anonyme. The literature on fortification in Renaissance Italy. bibliographical note, T. and C., IV, n^{o} 1 par Horst de la Croix. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):402–403, October 1963. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903649.

Anonymous:1963:AOB

[1267] Anonyme. Analyses d’ouvrages: Museums of Science and Technology. T. and C., IV, n^{o} 1 par Torsten Althin. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):403, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903650.

Payen:1963:AOBa

[1268] Jacques Payen. Analyses d’ouvrages: A Biographical Bibliography of Electrical Engineers and Electrophysicists, T. and C., II, n^{o} 1, (I^{re} partie); II, n^{o} 2, (II^{e} partie) par Thomas James Higgins. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):403, October 1963. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903651.

Daumas:1963:AOBc

[1269] M. Daumas. Analyses d’ouvrages: Revue d’Histoire de la Sidérurgie. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):404–405, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903652.

Daumas:1963:AQP

[1270] Maurice Daumas. Analyses de quelques périodiques: Cahier de l’Institut de Science économique appliquée, n^{o} 134, février 1963, série AD, n^{o} 2 : Évolution des techniques et progrès de l’économie. I.S.E.A., 35. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):405–406, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903653.

Payen:1963:AOBb

[1271] J. Payen. Analyses d’ouvrages: Contributions from the Museum of History and Technology. Forment des extraits de : United States National Museum Bulletin, depuis le n^{o} 218. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):406–408, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903654.

Anonymous:1963:TGM

[1272] Anonyme. Table générale des matières du Tome XVI. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):409–416, October 1963. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903655.

Anonymous:1963:BMa

[1273] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 16(1):??, January 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904080.

Anonymous:1963:BMb

[1274] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 16(2):??, avril–juin 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905078.

Anonymous:1963:BMc

[1275] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):??, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905191.

Anonymous:1963:BMd

[1276] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):??, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903656.

Anonymous:1963:FMa

[1277] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 16(1):??, January 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904070.

Anonymous:1963:FMb

[1278] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 16(2):??, avril–juin 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905066.

Anonymous:1963:FMc

[1279] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 16(3):??, juillet–septembre 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905157.

Anonymous:1963:FMd

[1280] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 16(4):??, October 1963. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903627.

Hyppolite:1964:HGB

[1281] Jean Hyppolite. En hommage à Gaston Bachelard: L’épistémologie de G. Bachelard. Revue d’Histoire des Sciences et de Leurs Applications, 17(1): 1–12, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904910.

Bouligand:1964:MTC

[1282] G. Bouligand. La mécanique théorique des corps flexibles (1638–1788) et les premières tentatives de ^{≪}spéculations fonctionnelles ^{≫}au XVIII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):13– 24, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904911.

Roger:1964:RHB

[1283] Jacques Roger. Réflexions sur l’histoire de la biologie (XVII^{e}–XVIII^{e} siècle) : problèmes de méthodes. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):25–40, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904912.

Rostand:1964:EGS

[1284] Jean Rostand. Étienne Geoffroy Saint-Hilaire et la tératogenèse expérimentale. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):41–50, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904913.

Anonymous:1964:AFS

[1285] Anonyme. Allemagne Fédérale: Symposium International d’Histoire de l’Astronomie. Revue d’Histoire des Sciences et de Leurs Applications, 17 (1):51, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904914.

Wong:1964:F

[1286] M. Wong. France. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):51–53, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904915.

Taton:1964:P

[1287] René Taton. Pologne. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):54–56, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904917.

Michel:1964:AOB

[1288] Paul-Henri Michel. Analyses d’ouvrages: Du monde clos à l’univers infini par Alexandre Koyré; Raïssa Tarr. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):57–60, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904918.

Roger:1964:AOB

[1289] Jacques Roger. Analyses d’ouvrages: French Free-Thought from Gassendi to Voltaire par J. S. Spink. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):60–63, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904919.

Taton:1964:AOBa

[1290] René Taton. Analyses d’ouvrages: Textes mathématiques de Suse (Mémoires de la Mission archéologique en Iran), t. XXXIV: (Mission de Susiane) par E. M. Bruins; M. Rutten; G. Contenau; R. de Mecquenem. Revue d’Histoire des Sciences et de Leurs Applications, 17(1): 63–64, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904920.

Taton:1964:AOBb

[1291] René Taton. Analyses d’ouvrages: Dalla geometria egiziana alla matematica moderna par Ugo Cassina. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):64–65, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904921.

Taton:1964:AOBc

[1292] René Taton. Analyses d’ouvrages: Les œuvres complètes d’Archimède suivies des Commentaires d’Eutocius d’Ascalon par Paul Ver Eecke. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):65–67, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904922.

Taton:1964:AOBd

[1293] René Taton. Analyses d’ouvrages: De Zestiende-eeuwse Nederlandse Rekenboeken par A. J. E. M. Smeur. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):67–68, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904923.

Taton:1964:AOE

[1294] René Taton. Analyses d’ouvrages: Évariste Galois, Écrits et mémoires mathématiques d’É. G. Édition critique intégrale de ses manuscrits et publications par Robert Bourgne et J.-P. Azra. Préface de J. Dieudonné, Paris, Gauthier-Villars, 1962. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):68–72, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904924.

Speziali:1964:AOBa

[1295] Pierre Speziali. Analyses d’ouvrages: Recherches sur l’analyse indéterminée et l’arithmétique de Diophante par Édouard Lucas; J. Itard. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):72–73, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904926.

Itard:1964:AOBa

[1296] Jean Itard. Analyses d’ouvrages: A History of the Calculus of Variations during the Nineteenth Century par Isaac Todhunter. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):73–74, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904928.

Itard:1964:AOBb

[1297] Jean Itard. Analyses d’ouvrages: The Development of Mathematics in China and Japan par Yoshio Mikami. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):74, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904930.

Taton:1964:AOBe

[1298] René Taton. Analyses d’ouvrages: Lexikon der Geschichte der Naturwissenschaften, Biographen, Sachwörter und Bibliographen … Zweite lieferung par Josef Mayerhöfer. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):74–75, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904932.

Taton:1964:AOBf

[1299] René Taton. Analyses d’ouvrages: The Principal Works of Simon Stevin, Vol. III : Astronomy, Navigation, par A. Pannekoek; Ernst Crone. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):75–76, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904934.

Speziali:1964:AOBb

[1300] Pierre Speziali. Analyses d’ouvrages: I microscopi Culpeper di Norimberga par Federico Allodi. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):76–77, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904936.

Wickersheimer:1964:AOB

[1301] Ernest Wickersheimer. Analyses d’ouvrages: On Anatomical Procedures The Later Books par Galen; W. L. H. Duckworth; M. C. Lyons; B. Towers. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):77–78, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904938.

Huard:1964:AOBa

[1302] P. Huard. Analyses d’ouvrages: Contribution à la biographie de Félix Dujardin et étude de son œuvre biologique et médicale, n^{o} 148 par Michel Étienvre. Revue d’Histoire des Sciences et de Leurs Applications, 17(1): 78–79, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904940.

Huard:1964:AOBb

[1303] P. Huard. Analyses d’ouvrages: Contribution à la bio-bibliographie de Paul Broca. Thèse de Rennes, 1962 par Jean Le Gaubichon. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):80, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904942.

Theodorides:1964:AOB

[1304] Jean Théodoridès. Analyses d’ouvrages: Histoire de la médecine et du livre médical à la lumière des collections de la Bibliothèque de la Faculté de Médecine de Paris par André Hahn; Paule Dumaître; Janine Samion-Contet. Revue d’Histoire des Sciences et de Leurs Applications, 17(1): 80–82, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904944.

Huard:1964:AOBc

[1305] P. Huard. Analyses d’ouvrages: La médecine navale et exotique française au XVII^{e} et au XVIII^{e} siècles, Thèse par Nguyen-Thanh Buu. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):82–83, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904946.

Huard:1964:AOBd

[1306] P. Huard. Analyses d’ouvrages: La chirurgie provinciale française au XVIII^{e} siècle. Son niveau technique d’après les membres non résidents de l’Académie royale de Chirurgie. Thèse de Rennes, 1962, n^{o} 281 par Yves Dordain. Revue d’Histoire des Sciences et de Leurs Applications, 17(1): 84–85, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904948.

Huard:1964:AOBe

[1307] P. Huard. Analyses d’ouvrages: Les membres non résidents étrangers de l’Académie Royale de Chirurgie. Thèse de Rennes, 1964, n^{o} 397 par Marie-Claire Dordain. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):85–86, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904950.

Huard:1964:AOBf

[1308] P. Huard. Analyses d’ouvrages: Contribution à l’histoire de la trachéotomie par Michel Legrand. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):86–87, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904952.

Huard:1964:AOBg

[1309] P. Huard. Analyses d’ouvrages: Les débuts de la médecine française au Sud-Vietnam, au Cambodge et au Laos, à la fin du XIX^{e} siècle par Maurice Jason. Revue d’Histoire des Sciences et de Leurs Applications, 17(1): 87–88, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904954.

Huard:1964:AOBh

[1310] P. Huard. Analyses d’ouvrages: Contribution à l’étude de la médecine laotienne par Khamsone Sassady. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):88–89, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904956.

Huard:1964:AOBi

[1311] P. Huard. Analyses d’ouvrages: Contribution à l’étude de l’histoire de la pharmacie au Viet-Nam, n^{o} 17 par Chuong Van Vinh. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):89–90, January 1964. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904958.

Huard:1964:AOBj

[1312] P. Huard. Analyses d’ouvrages: Un siècle d’évolution hospitalière en pays tropical : Histoire de l’hôpital Grall de Saïgon (1861–1961) par H. Mianet. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):91, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904960.

Colnort:1964:AOBa

[1313] S. Colnort. Analyses d’ouvrages: L’archéologie préhistorique, ^{≪}Le Rayon de la Science ^{≫}par Annette Laming-Emperaire. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):91–92, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904962.

Colnort:1964:AOBb

[1314] S. Colnort. Analyses d’ouvrages: Les Conférences du Palais de la Découverte. Série D : Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):92, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904964.

Colnort:1964:AQPa

[1315] S. Colnort. Analyses de quelques périodiques: Archives internationales d’Histoire des Sciences, t. XIV et XV, années 1961 et 1962, n^{os} 54–55, 56–57, 58–59, 60–61. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):93, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904966.

Colnort:1964:AQPb

[1316] S. Colnort. Analyses de quelques périodiques: Isis, an International Review devoted to the History of Science and its cultural Influences, vol. 52 et 53, années 1961 et 1962, fasc. 167 à 174 inclus. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):93–94, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904968.

Colnort:1964:AQPc

[1317] S. Colnort. Analyses de quelques périodiques: Chymia, Annual Studies in the History of Chemistry, vol. 8, 1962, ^{≪}Edgar F. Smith Memorial Collection: Univ. of Pennsylvania, Division of History of Chemistry, American Chemical Society. ^{≫}par Henry M. Leicester. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):94–95, January 1964. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904970.

Colnort:1964:AQPd

[1318] S. Colnort. Analyses de quelques périodiques: Gesnerus, vol. 18, 19, 20, années 1961, 1962, 1963. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):95, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904972.

Schiller:1964:LAB

[1319] Joseph Schiller and Marainville. Les laboratoires d’anatomie et de botanique à l’Académie des Sciences au XVII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):97–114, avril–juin 1964. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904672.

Lamontagne:1964:LHC

[1320] Roland Lamontagne. L’expédition hydrographique de Chabert au Canada. Revue d’Histoire des Sciences et de Leurs Applications, 17 (2):115–119, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904673.

Moreau:1964:PPD

[1321] Fernand Moreau. À propos d’un portrait du naturaliste normand J.-V.- F. Lamouroux. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):121–128, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904674.

Sebestik:1964:BBS

[1322] Jan Sebestik and Bernard Bolzano. Bernard Bolzano et son Mémoire sur le théorème fondamental de l’analyse. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):129–164, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904675.

Taton:1964:AOBg

[1323] R. Taton. Analyses d’ouvrages: Deutsche Akademie der Wissenschaften zu Berlin. Biographischer Index der Mitglieder par Kurt-R. Biermann; Gerhard Dunken. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):165, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904676.

Taton:1964:AOBh

[1324] René Taton. Analyses d’ouvrages: Panorama general de Historia de la Ciencia, t. VIII: El siglo del iluminismo (^{≪}Historia y Filosofia de la Ciencia ^{≫}, serie menor) t. IX: Biologia y medicina en los siglos XVII y XVIII, t. X: Las ciencias exactas en el siglo XIX, t. XI: Biología y medicina del siglo XIX, t. XII: Ciencia de la Tierra y Tecnica del siglo XIX par Desiderio Papp; José Babini. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):166–167, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904677.

Taton:1964:AOBi

[1325] René Taton. Analyses d’ouvrages: Geschichte der Mathematik, I : Von den Anfängen bis zum Auftreten von Fermat und Descartes (^{≪}Sammlung Göschen ^{≫}, Band 226/226 a) par Joseph E. Hofmann. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):168, avril–juin 1964. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904678.

Taton:1964:AOBj

[1326] René Taton. Analyses d’ouvrages: Arithmétique et théorie des nombres (coll. ^{≪}Que sais-je? ^{≫}, n^{o} 1093) par Jean Itard. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):168–169, avril–juin 1964. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904679.

Itard:1964:AOBc

[1327] Jean Itard. Analyses d’ouvrages: Le calcul mécanique (^{≪}Que sais-je? ^{≫}, n^{o} 367) par René Taton; Jean-Paul Flad. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):169, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904680.

Itard:1964:AOBd

[1328] Jean Itard. Analyses d’ouvrages: La perspective (^{≪}Que sais-je? ^{≫}, n^{o} 1050) par Albert Flocon; René Taton. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):169–170, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904681.

Taton:1964:AOBk

[1329] René Taton. Analyses d’ouvrages: Les livres arithmétiques d’Euclide (Histoire de la Pensée, vol. X) par Jean Itard. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):170–171, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904682.

Itard:1964:AOBe

[1330] Jean Itard. Analyses d’ouvrages: Les coniques d’Apollonius de Perge [nouveau tirage] par Paul Ver Eecke. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):171–172, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904683.

Machabey:1964:AOB

[1331] A. Machabey, Sr. Analyses d’ouvrages: Harmonie universelle par Marin Mersenne; François Lesure. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):172–174, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904684.

Taton:1964:AOBl

[1332] René Taton. Analyses d’ouvrages: Die deutsch-russische Begegnung und Leonhard Euler. Beiträge zu den Beziehungen zwischen der deutschen und der russischen Wissenschaft und Kultur im 18. Jahrhundert par E. Winter; P. N. Berkov; N. A. Figurovskij; V. P. Zubov; Die Berliner und die Petersburger Akademie der Wissenschaften im Briefwechsel Leonhard Eulers : Teil I, Der Briefwechsel L. Eulers mit G. F. Müller, 1735– 1767, Teil II, Der Briefwechsel L. Eulers mit Nartov, Razumovskij, Schumacher, Teplov und der Petersburger Akademie, 1730–1763 par A. P. Juskevic; E. Winter; P. Hoffmann; Ju. Ch. Kopolevic. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):175–178, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904685.

Speziali:1964:AOBc

[1333] Pierre Speziali. Analyses d’ouvrages: De Telescopio, ^{≪}Fondazione Giorgio Ronchi ^{≫}par Giovan Battista della Porta; Vasco Ronchi; Maria Amalia Naldoni. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):178–179, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904686.

Colnort:1964:AOBc

[1334] S. Colnort. Analyses d’ouvrages: Solar and Planetary Longitudes for Years-2500 to 2000 par William D. Stahlman; Owen Gingerich. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):180, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904687.

Taton:1964:AOBm

[1335] René Taton. Analyses d’ouvrages: Les Prix Nobel de physique et de chimie par Pierre Auger. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):180–181, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904688.

Taton:1964:AOBn

[1336] René Taton. Analyses d’ouvrages: Leben und Werk des Chemikers Tobias Lowitz (1757–1804) par N. A. Figurovskij; E. Winter. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):181, avril–juin 1964. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904689.

Laissus:1964:AOBa

[1337] Yves Laissus. Analyses d’ouvrages: Les sciences de la vie dans la pensée française du XVIII^{e} siècle. La génération des animaux de Descartes à l’Encyclopédie par Jacques Roger. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):182–191, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904690.

Laissus:1964:AOBb

[1338] Yves Laissus. Analyses d’ouvrages: Les Époques de la Nature. Mémoires du Muséum National d’Histoire Naturelle, série C, Sciences de la Terre, t. X par Buffon; Jacques Roger. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):191–192, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904691.

Hemardinquer:1964:DMA

[1339] J.-J. Hémardinquer. Une dynastie de mécaniciens anglais en France : James, John et Juliana Collier (1791–1847). Revue d’Histoire des Sciences et de Leurs Applications, 17(3):193–208, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903532.

Hunter:1964:OTF

[1340] Louis C. Hunter. Les origines des turbines Francis et Pelton : développement de la turbine hydraulique aux États-Unis de 1820 à 1900. Revue d’Histoire des Sciences et de Leurs Applications, 17(3): 209–242, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903533.

deMicheaux:1964:CIE

[1341] R. de Micheaux and J. Payen. Le Centre International d’Étude des Textiles Anciens. Revue d’Histoire des Sciences et de Leurs Applications, 17(3):243–250, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903534.

Momot:1964:LTP

[1342] Jacques Momot. L’histoire des techniques et la plongée en scaphandre autonome. Revue d’Histoire des Sciences et de Leurs Applications, 17(3): 251–257, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903535.

Payen:1964:GU

[1343] J. Payen. Guido Ucelli (1885–1964). Revue d’Histoire des Sciences et de Leurs Applications, 17(3):257–258, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903536.

Payen:1964:AOBa

[1344] Jacques Payen. Analyses d’ouvrages: Dictionnaire archéologique des techniques, t. I, t. II. Revue d’Histoire des Sciences et de Leurs Applications, 17(3):259–261, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903537.

Cahen:1964:AOB

[1345] Claude Cahen. Analyses d’ouvrages: Islamic armourers and their works par L. A. Mayer. Revue d’Histoire des Sciences et de Leurs Applications, 17(3):261–262, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903538.

Payen:1964:AOBb

[1346] Jacques Payen. Analyses d’ouvrages: Un constructeur d’instruments astronomiques au XV^{e} siècle : Jean Fusoris (Bibliothèque de l’École des Hautes Études, IV^{e} section, fasc. 318) par Emmanuel Poulle. Revue d’Histoire des Sciences et de Leurs Applications, 17(3):262–267, juillet– septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903539.

Daumas:1964:AOBa

[1347] Maurice Daumas. Analyses d’ouvrages: Les ingénieurs de la Renaissance par Bertrand Gille. Revue d’Histoire des Sciences et de Leurs Applications, 17(3):267–275, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903540.

Gille:1964:AOB

[1348] Paul Gille. Analyses d’ouvrages: La genèse de l’hélice propulsive par Paul Augustin-Normand. Revue d’Histoire des Sciences et de Leurs Applications, 17(3):275–279, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903541.

Daumas:1964:AOBb

[1349] M. Daumas. Analyses d’ouvrages: De Nice à Chamonix. Les réseaux secondaires dans les Alpes françaises par Jean Robert. Revue d’Histoire des Sciences et de Leurs Applications, 17(3):280, juillet–septembre 1964. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903542.

Payen:1964:AOBc

[1350] J. Payen. Analyses d’ouvrages: Die Entwicklung des Stahlhochbaus (Le développement des superstructures métalliques) par Karl A. Wittek. Revue d’Histoire des Sciences et de Leurs Applications, 17(3): 280–281, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903543.

Payen:1964:AOBd

[1351] Jacques Payen. Analyses d’ouvrages: La métallurgie, (^{≪}Le Rayon de la Science ^{≫}, n^{o} 19) par Pierre-Julien Le Thomas. Revue d’Histoire des Sciences et de Leurs Applications, 17(3):281–282, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903544.

Payen:1964:AOBe

[1352] Jacques Payen. Analyses d’ouvrages: L’industrie du gaz, (^{≪}Que saisje? ^{≫}, n^{o} 239) par Philippe Brachay. Revue d’Histoire des Sciences et de Leurs Applications, 17(3):282, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903545.

Daumas:1964:AOBc

[1353] M. Daumas. Analyses d’ouvrages: Dictionnaire illustré des inventeurs français par Marie-Fernande Alphandéry. Revue d’Histoire des Sciences et de Leurs Applications, 17(3):283, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903546.

Payen:1964:AOBf

[1354] Jacques Payen. Analyses d’ouvrages: Contributions from the Museum of History and Technology. (forment des extraits de : United States National Museum Bulletin, depuis le n^{o} 218). Revue d’Histoire des Sciences et de Leurs Applications, 17(3):284–285, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903547.

Payen:1964:AQP

[1355] J. Payen. Analyses de quelques périodiques: Agrikultura, Sbornik Polnohospodarskeho Muzea V Nitre, n^{o} 1 (1962); n^{o} 2 (1963). Revue d’Histoire des Sciences et de Leurs Applications, 17(3):285–287, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903548.

Payen:1964:AOBg

[1356] J. Payen. Analyses d’ouvrages: Blätter für Technikgeschichte, n^{os} 24 (1962) et 25 (1963). Technisches Museum für Industrie und Gewerbe. Revue d’Histoire des Sciences et de Leurs Applications, 17(3):287– 288, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903549.

Taton:1964:TCV

[1357] René Taton. Tableau chronologique de la vie et de l’œuvre de Galilée. Revue d’Histoire des Sciences et de Leurs Applications, 17(4):293–304, October 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903783.

Clavelin:1964:GRE

[1358] Maurice Clavelin. Galilée et le refus de l’équivalence des hypothèses. Revue d’Histoire des Sciences et de Leurs Applications, 17(4):305–330, October 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903784.

Russo:1964:LGC

[1359] François Russo. Lettre de Galilée à Christine de Lorraine Grande-Duchesse de Toscane (1615). Revue d’Histoire des Sciences et de Leurs Applications, 17(4):331–368, October 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903785.

Namer:1964:LME

[1360] Emile Namer. L’intelligibilité mathématique et l’expérience chez Galilée. Revue d’Histoire des Sciences et de Leurs Applications, 17(4):369–384, October 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903786.

Costabel:1964:RAC

[1361] Pierre Costabel. La roue d’Aristote et les critiques françaises à l’argument de Galilée. Revue d’Histoire des Sciences et de Leurs Applications, 17(4):385–396, October 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903787.

Namer:1964:TFB

[1362] Émile Namer. Le traité de ^{≪}la balance hydrostatique ^{≫}de Galilée. Revue d’Histoire des Sciences et de Leurs Applications, 17(4):397–403, October 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903788.

Anonymous:1964:F

[1363] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 17(4):405–408, October 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903789.

Anonymous:1964:TGM

[1364] Anonyme. Table générale des matières du Tome XVII. Revue d’Histoire des Sciences et de Leurs Applications, 17(4):409–416, October 1964. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903790.

Anonymous:1964:BMa

[1365] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):??, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904973.

Anonymous:1964:BMb

[1366] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):??, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904692.

Anonymous:1964:BMc

[1367] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 17(3):??, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903550.

Anonymous:1964:BMd

[1368] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 17(4):??, October 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903791.

Anonymous:1964:FMa

[1369] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 17(1):??, January 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904909.

Anonymous:1964:FMb

[1370] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):??, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904671.

Anonymous:1964:FMc

[1371] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 17(3):??, juillet–septembre 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903531.

Anonymous:1964:FMd

[1372] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 17(4):??, October 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903780.

Anonymous:1964:PTA

[1373] Anonyme. Principaux travaux d’Alexandre Koyré sur Galilée. Revue d’Histoire des Sciences et de Leurs Applications, 17(4):??, October 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903781.

McKeon:1965:LMT

[1374] Robert M. McKeon. Une lettre de Melchisédech Thévenot sur les débuts de l’Académie royale des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):1–6, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903976.

Chapin:1965:ALL

[1375] Seymour L. Chapin. Les associés libres de l’Académie royale des Sciences: Un projet inédit pour la modification de leurs statuts (1788). Revue d’Histoire des Sciences et de Leurs Applications, 18(1):7–13, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903977.

Hahn:1965:LRS

[1376] Roger Hahn and de la Rochefoucauld. L’Académie royale des Sciences et la réforme de ses statuts en 1789. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):15–28, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903978.

Costabel:1965:LID

[1377] Pierre Costabel. Une lettre inédite du marquis de L’Hôpital sur la résolution de l’équation du troisième degré. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):29–43, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903979.

Callot:1965:SMH

[1378] Émile Callot. Système et méthode dans l’histoire de la botanique. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):45–53, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903980.

Grmek:1965:CMM

[1379] M. D. Grmek and D. Guinot. Les Crustacés dans la matière médicale européenne au XVI^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):55–71, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903981.

Dumas:1965:LSE

[1380] Jean-Louis Dumas. Liebig et son empreinte sur l’agronomie moderne: Biographie sommaire de Liebig. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):73–108, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903982.

Schiller:1965:CBR

[1381] Joseph Schiller. Claude Bernard et la réaction histochimique du glycogène: Un document inédit. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):109–112, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903983.

Russo:1965:ML

[1382] F. Russo. Maurice Lacoin (1877–1963). Revue d’Histoire des Sciences et de Leurs Applications, 18(1):113–114, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903984.

Taton:1965:EB

[1383] René Taton. Edmond Bauer (1880–1963). Revue d’Histoire des Sciences et de Leurs Applications, 18(1):114–116, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903985.

Rochot:1965:LA

[1384] B. Rochot. Léon Auger (1886–1964). Revue d’Histoire des Sciences et de Leurs Applications, 18(1):116–117, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903986.

Anonymous:1965:Fa

[1385] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):117–118, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903987.

Itard:1965:AOBa

[1386] Jean Itard. Analyses d’ouvrages: Philosophie mathématique, (collection ^{≪}Histoire de la Pensée ^{≫}, VI) par Jean Cavaillès; Raymond Aron; Roger Martin. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):119–120, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903988.

Itard:1965:AOBb

[1387] Jean Itard. Analyses d’ouvrages: Éléments d’histoire des mathématiques, (collection ^{≪}Histoire de la Pensée ^{≫}, IV) par Nicolas Bourbaki. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):120–123, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903989.

Itard:1965:AOBc

[1388] Jean Itard. Analyses d’ouvrages: Abdülhamid ibn Türk’ün Katisik denklemlerde mantikî zaruretler adli yazisi ve zamanin cebri (Logical necessities in mixed equations by Abd al H .amîd ibn Turk and the Algebra of his time) par Aydin Sayili. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):123–124, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903990.

Crosland:1965:AOB

[1389] Maurice Crosland. Analyses d’ouvrages: Lavoisier. The Crucial Year. The Background and origin of his first Experiments on Combustion in 1772 par Henry Guerlac. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):124–126, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903991.

Wong:1965:AOB

[1390] M. Wong. Analyses d’ouvrages: Tche-wou ming-che t’ou-k’ao (Flore descriptive et illustrée), 1^{re} réimpression de la 1^{re} édition par Wou Ki-Tsiun. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):126– 127, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903992.

Theodorides:1965:AOB

[1391] Jean Théodoridès. Analyses d’ouvrages: Albert Dürer. Dessins par Henri Pastoureau; Lise Dubief. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):127, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903993.

Russo:1965:AOB

[1392] F. Russo. Analyses d’ouvrages: Histoire générale des techniques, T. I : Les origines de la civilisation technique par Maurice Daumas. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):128, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903994.

Delorme:1965:HAK

[1393] Suzanne Delorme. Hommage à Alexandre Koyré. Revue d’Histoire des Sciences et de Leurs Applications, 18(2):129–131 133–139, avril–juin 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904160.

Vignaux:1965:TSS

[1394] Paul Vignaux. De la théologie scolastique à la science moderne. Revue d’Histoire des Sciences et de Leurs Applications, 18(2):141–146, avril– juin 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904161.

Taton:1965:AKH

[1395] René Taton. Alexandre Koyré historien de la ^{≪}révolution astronomique ≫. Revue d’Histoire des Sciences et de Leurs Applications, 18(2):147–154, avril–juin 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904162.

Costabel:1965:AKC

[1396] Pierre Costabel. Alexandre Koyré critique de la pensée mécanique. Revue d’Histoire des Sciences et de Leurs Applications, 18(2):155–159, avril– juin 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904163.

Dubarle:1965:MSG

[1397] Dominique Dubarle. La méthode scientifique de Galilée. Revue d’Histoire des Sciences et de Leurs Applications, 18(2):161–192, avril–juin 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904164.

Moscovici:1965:DHT

[1398] Serge Moscovici. Les développements historiques de la théorie galiléenne des marées. Revue d’Histoire des Sciences et de Leurs Applications, 18(2):193–220, avril–juin 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904165.

Costabel:1965:HMG

[1399] Pierre Costabel and Pierre Piveteau. Un hommage de Mersenne à Galilée. Revue d’Histoire des Sciences et de Leurs Applications, 18(2):221–226, avril–juin 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904166.

Taton:1965:OOG

[1400] R. Taton. Les œuvres originales de Galilée dans quelques grandes bibliothèques parisiennes. Revue d’Histoire des Sciences et de Leurs Applications, 18(2):227–230, avril–juin 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904167.

Namer:1965:QTR

[1401] Émile Namer. Quelques travaux récents sur Galilée. Revue d’Histoire des Sciences et de Leurs Applications, 18(2):230–234, avril–juin 1965. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904168.

Hahn:1965:CDI

[1402] Roger Hahn and René Taton. Célébration du IV^{e} centenaire de la naissance de Galilée (1564–1964). Revue d’Histoire des Sciences et de Leurs Applications, 18(2):234–238, avril–juin 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904169.

Taton:1965:AOB

[1403] René Taton. Analyses d’ouvrages: Sidereus nuncius (Le Message céleste). (coll. ^{≪}Grands classiques des sciences et des techniques. ^{≪}Un savant dans le texte ^{≫}^{≫}) par Galilée; Émile Namer. Revue d’Histoire des Sciences et de Leurs Applications, 18(2):239–240, avril–juin 1965. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904170.

Michelet:1965:MTI

[1404] Henri Michelet. La machine typographique d’Isaac de Rivaz: Une étape méconnue de l’histoire de l’imprimerie. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):241–247, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23906152.

Payen:1965:CAG

[1405] Jacques Payen and G. de la Landelle. Clément Ader et Gabriel de la Landelle (1883–1884). Revue d’Histoire des Sciences et de Leurs Applications, 18(3):249–264, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23906153.

Soulard:1965:EBT

[1406] Robert Soulard. Édouard Belin et la télévision. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):265–281, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23906154.

Daumas:1965:MTG

[1407] Maurice Daumas. Les machines à traduire de Georges Artsrouni. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):283–302, juillet– septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23906155.

Daumas:1965:NEI

[1408] Maurice Daumas. Une nouvelle édition intégrale du ^{≪}recueil de planches de l’Encyclopédie ^{≫}. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):303–304, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23906156.

Delesalle:1965:AFA

[1409] Hubert Delesalle. Aunes de France et aunes de Flandre: Note sur le mesurage des anciennes tapisseries de Beauvais. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):305–308, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23906157.

Payen:1965:DRI

[1410] Jacques Payen. Documents relatifs à l’introduction en France de la machine à vapeur de Watt. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):309–314, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23906158.

Payen:1965:OBI

[1411] Jacques Payen. Les origines de la bobine d’induction. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):315–318, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23906159.

Payen:1965:AOBa

[1412] Jacques Payen. Analyses d’ouvrages: Geschichte der Technik par A. A. Sworykin; N. I. Osmowa; W. I. Tschernyschew; S. W. Schuchardin. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):319–323, juillet– septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23906160.

Payen:1965:AOBb

[1413] Jacques Payen. Analyses d’ouvrages: Kleine Geschichte der Technologie. (Urban–Bücher, die Wissenschaftliche Taschenbuchreihe, n^{o} 78) par Albrecht Timm. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):323, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23906161.

Payen:1965:AOBc

[1414] J. Payen. Analyses d’ouvrages: The two Netherlanders. Humphrey Bradley and Cornelis Drebbel par L. E. Harris. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):323–324, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23906162.

Daumas:1965:AOBa

[1415] Maurice Daumas. Analyses d’ouvrages: Early American Scientific Instruments and Their Makers par Silvio A. Bedini. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):325–327, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23906163.

Daumas:1965:AOBb

[1416] Maurice Daumas. Analyses d’ouvrages: L’inventeur Isaac de Rivaz (1752–1828). Ses recherches techniques et ses tentatives industrielles par Henri Michelet. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):328–329, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23906164.

Daumas:1965:AOBc

[1417] Maurice Daumas. Analyses d’ouvrages: Metal Founding through the Ages on Czechoslovak Territory par Ladislav Jenicek. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):329–330, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23906165.

Payen:1965:AOBd

[1418] Jacques Payen. Analyses d’ouvrages: The Making of the Electrical Age par Harold I. Sharlin. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):330–332, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23906166.

Payen:1965:AOBe

[1419] Jacques Payen. Analyses d’ouvrages: A History of Electrical Engineering par Percy Dunsheath. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):332–333, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23906167.

Daumas:1965:AOBd

[1420] Maurice Daumas. Analyses d’ouvrages: Montres, pendules et horloges par Ernst von Bassermann-Jordan; Hans von Bertele. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):333–334, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23906168.

Payen:1965:AQP

[1421] Jacques Payen. Analyses de quelques périodiques: The Journal of Industrial Archaeology, vol. I, n^{os} 1 à 4, 1964–1965. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):335, juillet–septembre 1965. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23906169.

Daumas:1965:AOBe

[1422] Maurice Daumas. Analyses d’ouvrages: Forschung und Technik in der Kunst. Revue d’Histoire des Sciences et de Leurs Applications, 18 (3):336, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23906170.

Morere:1965:PSE

[1423] Jean-Édouard Morère. La photométrie : les sources de l’essai d’optique sur la gradation de la lumière de Pierre Bouguer, 1729. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):337–384, October 1965. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904581.

Lamontagne:1965:PCO

[1424] Roland Lamontagne. La participation canadienne à l’œuvre minéralogique de Guettard. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):385–388, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904582.

Payen:1965:MAP

[1425] Jacques Payen. La machine arithmétique de Pascal : À propos de l’exemplaire signé et daté de 1652. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):389–391, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904583.

Speziali:1965:UC

[1426] Pierre Speziali. Ugo Cassina (1897–1964). Revue d’Histoire des Sciences et de Leurs Applications, 18(4):391–393, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904584.

Anonymous:1965:AFX

[1427] Anonyme. Allemagne fédérale: XX^{e} Congrès International d’Histoire de la Médécine. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):393–394, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904585.

Anonymous:1965:Fb

[1428] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):394–396, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904586.

Anonymous:1965:MFS

[1429] Anonyme. Mexique: Fondation d’une Société d’Histoire de la Science et de la Technique. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):396, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904587.

Speziali:1965:AOBa

[1430] Pierre Speziali. Analyses d’ouvrages: Histoire générale des Sciences, t. III : La science contemporaine, vol. 1 : Le XIX^{e} siècle par René Taton. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):397–401, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904588.

Speziali:1965:AOBb

[1431] Pierre Speziali. Analyses d’ouvrages: Histoire générale des Sciences, t. III : La science contemporaine, vol. 2 : Le XX^{e} siècle par René Taton. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):401–406, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904589.

Grmek:1965:AOBa

[1432] M. D. Grmek. Analyses d’ouvrages: Scientific Change. Historical studies in the intellectual, social and technical conditions for scientific discovery and technical invention, from Antiquity to the Present, Symposium on the History of Science, University of Oxford, 9–15 July 1961 par A. C. Crombie. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):406–407, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904590.

Speziali:1965:AOBc

[1433] Pierre Speziali. Analyses d’ouvrages: La classification dans les sciences, Centre National Belge de Recherches de Logique, Société belge de Logique et de Philosophie des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):408–409, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904591.

Colnort:1965:AOBa

[1434] Suzanne Colnort. Analyses d’ouvrages: Les origines de la pensée grecque, (^{≪}Mythes et religions ^{≫}, n^{o} 45) par Jean-Pierre Vernant. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):409–410, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904592.

Colnort:1965:AOBb

[1435] Suzanne Colnort. Analyses d’ouvrages: Science and the Renaissance. An Introduction to the study of the emergence of the sciences in the sixteenth century …, (^{≪}Aberdeen University studies ^{≫}, 143, 144) par W. P. D. Wightman. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):410–411, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904593.

Poulle:1965:AOB

[1436] Emmanuel Poulle. Analyses d’ouvrages: El comentario de ibn al-Mutanna a las tablas astronomicas de al-Jwarizmi; estudio y edición critica del texto latino en la version de Hugo Sanctallensis par Eduardo Millás Vendrell. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):411–413, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904594.

Rochot:1965:AOB

[1437] B. Rochot. Analyses d’ouvrages: Extraits des Traités de l’Équilibre des Liqueurs el de la Pesanteur de la Masse de l’air. (Grands Classiques des Sciences et des Techniques : Un savant dans le texte, 1) par Blaise Pascal; Jacques Payen. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):413, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904595.

Sebestik:1965:AOB

[1438] Jan Sebestik. Analyses d’ouvrages: Der Josephinismus. Die Geschichte des österreichischen Reformkatholizismus 1740–1848 par Eduard Winter. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):414–416, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904596.

deDestaillats:1965:AOB

[1439] Leticia Halperin de Destaillats. Analyses d’ouvrages: Classics in the theory of chemical combination, (coll. ^{≪}Classics of science ^{≫}, vol. I) par O. Theodor Benfey. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):416–417, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904597.

Labat:1965:AOB

[1440] René Labat. Analyses d’ouvrages: Die babylonisch-assyrische Medizin in Texten und Untersuchungen, Band II par Fr. Koecher; Fr. Köcher. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):417–418, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904598.

Huard:1965:AOBa

[1441] P. Huard. Analyses d’ouvrages: La Medicina del mondo degli Etruschi par Mario Tabanelli. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):418–420, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904599.

Grmek:1965:AOBb

[1442] M. D. Grmek. Analyses d’ouvrages: The Medical Background of Anglo-Saxon England. A Study in History, Psychology and Folklore par Wilfrid Bonser. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):420–421, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904600.

Huard:1965:AOBb

[1443] P. Huard. Analyses d’ouvrages: Les membres parisiens de l’Académie Royale de Chirurgie par Jean-Louis Boisseau. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):421–422, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904601.

Colnort:1965:AQPa

[1444] S. Colnort. Analyses de quelques périodiques: Archives internationales d’histoire des sciences, revue trimestrielle, publiée par la division d’histoire des sciences de l’Union internationale d’Histoire et de Philosophie des sciences, t. XVI, 1963, n^{os} 62–65. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):422–423, October 1965. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904602.

Colnort:1965:AQPb

[1445] S. Colnort. Analyses de quelques périodiques: Isis, an international review devoted to the history of science and its cultural influence, vol.

54, année 1963, fasc. 175 à 178 inclus. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):423–424, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904603.

Anonymous:1965:PR

[1446] Anonyme. Publications reçues. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):424, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904604.

Anonymous:1965:TGM

[1447] Anonyme. Table générale des matières du Tome XVIII. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):425–432, October 1965. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904605.

Anonymous:1965:BMa

[1448] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):??, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903995.

Anonymous:1965:BMb

[1449] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 18(2):??, avril–juin 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904171.

Anonymous:1965:BMc

[1450] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):??, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23906171.

Anonymous:1965:BMd

[1451] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):??, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904606.

Anonymous:1965:FMa

[1452] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 18(1):??, January 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903975.

Anonymous:1965:FMb

[1453] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 18(2):??, avril–juin 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904159.

Anonymous:1965:FMc

[1454] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 18(3):??, juillet–septembre 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23906151.

Anonymous:1965:FMd

[1455] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 18(4):??, October 1965. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904580.

Dumortier:1966:SAC

[1456] Bernard Dumortier. La stridulation et l’audition chez les insectes Orthoptères : Aperçu historique sur les idées et les découvertes jusqu’au début du XX^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):1–28, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904824.

Grmek:1966:LCL

[1457] M. D. Grmek and A. Rousseau. L’œuvre cardiologique de Léon Rostan. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):29–52, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904825.

Rostand:1966:MCA

[1458] Jean Rostand. Une thèse médicale sur la congélation des animaux supérieurs (1811). Revue d’Histoire des Sciences et de Leurs Applications, 19(1):53–54, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904826.

D:1966:HAP

[1459] S. D. Un hommage au P^{r} J.-F. Malgaigne. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):54, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904827.

Virtanen:1966:CCB

[1460] Reino Virtanen. Le colloque Claude Bernard: Paris, 29 juin–2 juillet 1965. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):55– 58, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904828.

Delorme:1966:XCI

[1461] Suzanne Delorme. Le XI^{e} Congrès international d’Histoire des Sciences: Varsovie-Cracovie, août 1965. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):59–63, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904829.

Anonymous:1966:XCI

[1462] Anonyme. XII^{e} Congrès international d’histoire des sciences. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):64, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904831.

D:1966:RAC

[1463] S. D. Réunion au Centre scientifique de l’Académie polonaise des Sciences à Paris: Compte rendu du XI^{e} Congrès international d’histoire des sciences. Revue d’Histoire des Sciences et de Leurs Applications, 19(1): 64, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904830.

Wong:1966:W

[1464] M. Wong. S. S. Wong (1889–1965). Revue d’Histoire des Sciences et de Leurs Applications, 19(1):65–66, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904832.

Taton:1966:JD

[1465] René Taton. Jules Duhem (1888–1964). Revue d’Histoire des Sciences et de Leurs Applications, 19(1):66, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904833.

Theodorides:1966:Fa

[1466] J. Théodoridès and S. D. France. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):67–68, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904834.

Itard:1966:AOBa

[1467] Jean Itard. Analyses d’ouvrages: Correspondance du P. Marin Mersenne, T. VIII : août 1638–décembre 1639 par Paul Tannery; Cornelis De Waard. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):69–70, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904835.

Jacquot:1966:AOB

[1468] F. Jacquot. Analyses d’ouvrages: The Philosophy of Francis Bacon par Benjamin Farrington. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):70–73, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904836.

Taton:1966:AOBa

[1469] René Taton. Analyses d’ouvrages: IV Centenario della nascita di Galileo Galilei, 1564–1964, Mostra di documenti e cimeli Galileiana par Maria Luisa Bonelli; G. Barbéra Florence. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):73–74, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904837.

Poulle:1966:AOBa

[1470] Emmanuel Poulle. Analyses d’ouvrages: Geschichte und Bibliographie der astronomischen Literatur in Deutschland zur Zeit der Renaissance … 2., unveränderte Auflage der Erstauflage von 1941 mit einem Nachtrag von 622 Nummern par Ernst Zinner. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):74, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904838.

McKeon:1966:AOB

[1471] Robert M. McKeon. Analyses d’ouvrages: The History of the Barometer par W. E. Knowles Middleton. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):75–77, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904839.

Mongredien:1966:AOBa

[1472] André Mongredien. Analyses d’ouvrages: The Story of Atomic Theory and Atomic Energy (Formerly titled: The Atom Story) par J. G. Feinberg; Histoire de l’atome, ^{≪}Les grandes découvertes scientifiques ≫(traduction française de la première version de l’ouvrage ci-dessus) par J. G. Feinberg; Stephen Spriel; L’histoire de l’énergie atomique par Laura Fermi; Nicole Rey; La découverte de l’atome, (Petite Bibliothèque Payot, n^{o} 26) par Alfred Romer; J. Métadier; The Discovery of Radioactivity and Transmutation (Classics of Science, vol. 2) par Alfred Romer; Becquerel; Rutherford; Crookes; P. Soddy; M. Curie; Laborde; Ramsay. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):77–81, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904840.

Mongredien:1966:AOBb

[1473] André Mongredien. Analyses d’ouvrages: Introduction à la physique nucléaire …, (^{≪}La Science vivante ^{≫}) par D. M. Main; Maurice Duquesne; Jeannine Duquesne. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):81–82, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904841.

deDestaillats:1966:AOBa

[1474] Leticia Halperin de Destaillats. Analyses d’ouvrages: Extraits du Traité élémentaire de chimie (coll. Grands Classiques des Sciences et des Techniques, ^{≪}Un savant dans le texte ^{≫}) par Lavoisier; Henri Montias; Pierre Dubost. Revue d’Histoire des Sciences et de Leurs Applications, 19(1): 82–83, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904842.

deDestaillats:1966:AOBb

[1475] Leticia Halperin de Destaillats. Analyses d’ouvrages: Chemistry and Beyond par F. A. Paneth; Herbert Dingle; G. R. Martin; Eva Paneth. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):83–84, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904843.

Grmek:1966:AOBa

[1476] M. D. Grmek. Analyses d’ouvrages: Heat and Life. The Development of the Theory of Animal Heat par Everett Mendelsohn. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):85–87, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904844.

Colnort:1966:AOBa

[1477] S. Colnort. Analyses d’ouvrages: L’hérédité, (^{≪}Où va la biologie? ^{≫}) par David M. Bonner; Williams; Duval-Destin. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):87, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904845.

Rochot:1966:AOBa

[1478] B. Rochot. Analyses d’ouvrages: Teilhard de Chardin, sa vie, son œuvre, avec un exposé de sa philosophie (Philosophes) par Jules Carles. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):87–88, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904846.

Schiller:1966:AOB

[1479] J. Schiller. Analyses d’ouvrages: Médecins, libertins et Pasquins par Marcel Florkin. Revue d’Histoire des Sciences et de Leurs Applications, 19(1): 88–89, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904847.

Huard:1966:AOBa

[1480] P. Huard. Analyses d’ouvrages: Contribution à l’étude du monde médical versaillais sous le règne de Louis XVI et pendant la Révolution par Pierre Brassart. Revue d’Histoire des Sciences et de Leurs Applications, 19(1): 89–91, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904848.

Huard:1966:AOBb

[1481] P. Huard. Analyses d’ouvrages: The History of surgical anesthesia par Thomas E. Keys. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):91–93, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904849.

Colnort:1966:AOBb

[1482] S. Colnort. Analyses d’ouvrages: Vijftig Jaren beofening van de Geschiedenis der Geneeskunde, Wiskunde en Natuurwetenschappen in Nederland. 1913–1963. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):93, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904850.

Russo:1966:AOB

[1483] F. Russo. Analyses d’ouvrages: Le langage des géographes: termes, signes, couleurs (1500–1800) par François de Dainville. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):94–95, January 1966. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904851.

Colnort:1966:AQPa

[1484] Suzanne Colnort. Analyses de quelques périodiques: History of Science, An annual review of literature, research and teaching, vol. 2, 1963; vol. 3, 1964. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):95, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904852.

Colnort:1966:AQPb

[1485] S. Colnort. Analyses de quelques périodiques: Chymia, Annual Studies in the history of chemistry, vol. 9, ^{≪}Edgar F. Smith College, Univ. of Pennsylvania and Div. of Hist. of Chemistry, American Philos. Soc. ^{≫}par Henry M. Leicester. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):95–96, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904853.

Colnort:1966:AQPc

[1486] Suzanne Colnort. Analyses de quelques périodiques: Medical History, A quarterly journal devoted to the history and bibliography of medicine and the related sciences, vol. VIII, 1964; vol. IX, 1965. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):96, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904854.

Halpern:1966:CPC

[1487] Bernard Halpern. Concepts philosophiques de Claude Bernard d’après l’introduction à l’étude de la médecine expérimentale. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):97–114, avril–juin 1966. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903676.

Birembaut:1966:CLF

[1488] A. Birembaut, P. Costabel, S. Delorme, and Leibniz. La correspondance Leibniz–Fontenelle et les relations de Leibniz avec l’Académie Royale des Sciences en 1700–1701. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):115–132, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903677.

Heilbron:1966:PIB

[1489] John L. Heilbron. À propos de l’invention de la bouteille de Leyde. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):133–142, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903678.

Taton:1966:PNM

[1490] René Taton and Monge. La première note mathématique de Gaspard Monge (juin 1769). Revue d’Histoire des Sciences et de Leurs Applications, 19(2):143–149, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903679.

Gabbey:1966:MIS

[1491] W. A. Gabbey and J. W. Herivel. Un manuscrit inédit de Sadi Carnot. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):151–166, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903680.

Costabel:1966:RCP

[1492] Pierre Costabel. Rectification et complément à la publication d’un inédit de Leibniz. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):167–169, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903681.

Anonymous:1966:AML

[1493] Anonyme. Le 250^{e} anniversaire de la mort de Leibniz. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):170, avril–juin 1966. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903682.

Huard:1966:DJT

[1494] P. Huard. Le D^{r} Jean Torlais (1897–1964). Revue d’Histoire des Sciences et de Leurs Applications, 19(2):171–172, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903683.

Kayser:1966:DEW

[1495] Ch. Kayser. Le D_{r} Émile Wickersheimer (1880–1965). Revue d’Histoire des Sciences et de Leurs Applications, 19(2):173–176, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903684.

Hooykaas:1966:EJD

[1496] R. Hooykaas. Eduard Jan Dijksterhuis (1892–1965). Revue d’Histoire des Sciences et de Leurs Applications, 19(2):177–179, avril–juin 1966. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903685.

Taton:1966:F

[1497] René Taton. France. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):179–182, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903686.

Huard:1966:GBL

[1498] P. Huard. Grande-Bretagne: L’Exposition de Médecine Chinoise du Wellcome Museum de Londres. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):182, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903687.

Grmek:1966:AOBb

[1499] M. D. Grmek. Analyses d’ouvrages: John Locke (1632–1704), physician and philosopher. A medical biography, with an edition of the medical notes in his Journals par Kenneth Dewhurst. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):183–184, avril–juin 1966. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903688.

Grmek:1966:AOBc

[1500] M. D. Grmek. Analyses d’ouvrages: A Short History of Medicine, 2^{e} éd. par Charles Singer; E. Ashworth Underwood. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):184, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903689.

Grmek:1966:AOBd

[1501] M. D. Grmek. Analyses d’ouvrages: The Epic of Medicine par Felix Marti-Ibañez. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):185, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903690.

Grmek:1966:AOBe

[1502] M. D. Grmek. Analyses d’ouvrages: Bibliography of the History of Medicine of United States and Canada, 1939–1960 par Genevieve Miller. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):185, avril– juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903691.

Grmek:1966:AOBf

[1503] M. D. Grmek. Analyses d’ouvrages: La chirurgia italiana nell’alto medioevo par Mario Tabanelli. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):185–186, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903692.

Grmek:1966:AOBg

[1504] M. D. Grmek. Analyses d’ouvrages: Doctors Monro. A medical saga par Rex E. Wright-St. Clair. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):186–187, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903693.

Grmek:1966:AOBh

[1505] M. D. Grmek. Analyses d’ouvrages: Foundations of Anesthesiology par Albert Faulconer; Thomas E. Keys. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):187, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903694.

Colnort:1966:AOBc

[1506] S. Colnort. Analyses d’ouvrages: Les Conférences du Palais de la Découverte, série D : Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):187–188, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903695.

Colnort:1966:AQPd

[1507] S. Colnort. Analyses de quelques périodiques: Isis, an International Review devoted to the History of Science and its cultural Influence, vol. 55, année 1964, fasc. 179 à 182. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):188, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903696.

Colnort:1966:AQPe

[1508] S. Colnort. Analyses de quelques périodiques: Archives internationales d’Histoire des Sciences, t. XVII, 1964, n^{os} 66, 67, 68–69; t. XVIII, 1965, n^{os} 70–71, 72–73. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):188–189, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903697.

Colnort:1966:AQPf

[1509] S. Colnort. Analyses de quelques périodiques: Physis. Rivista di Storia della Scienza, vol. IV, 1962; vol. V, 1963; vol. VI, 1964. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):189–191, avril– juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903698.

Colnort:1966:AQPg

[1510] S. Colnort. Analyses de quelques périodiques: Gesnerus, vol. 21, année 1964; vol. 22, année 1965. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):191–192, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903699.

Colnort:1966:AQP

[1511] S. Colnort. Analyses de quelques périodiques: The British Journal for the History of Science, t. I, 1962–63, fasc. 1–4. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):192, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903700.

Maheu:1966:BPB

[1512] Gilles Maheu. Bibliographie de Pierre Bouguer (1698–1758). Revue d’Histoire des Sciences et de Leurs Applications, 19(3):193–205, juillet– septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904241.

Maheu:1966:VSA

[1513] Gilles Maheu. La vie scientifique au milieu du XVIII^{e} siècle : Introduction à la publication des lettres de Bouguer à Euler. Revue d’Histoire des Sciences et de Leurs Applications, 19(3):206–224, juillet–septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904242.

Lamontagne:1966:LBE

[1514] Roland Lamontagne. Lettres de Bouguer à Euler. Revue d’Histoire des Sciences et de Leurs Applications, 19(3):225–246, juillet–septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904243.

Bouligand:1966:IPC

[1515] Georges Bouligand. Introduction à la pensée créatrice de Jacques Hadamard (1865–1963): Le qualitatif et le global dans son œuvre géométrique. Revue d’Histoire des Sciences et de Leurs Applications, 19 (3):247–265, juillet–septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904244.

Costabel:1966:PHM

[1516] Pierre Costabel. Paul-Henri Michel (1894–1964). Revue d’Histoire des Sciences et de Leurs Applications, 19(3):267–269, juillet–septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904245.

Koning:1966:JAV

[1517] D. A. Wittop Koning. Johan Adriaan Vollgraff (1877–1965). Revue d’Histoire des Sciences et de Leurs Applications, 19(3):270, juillet– septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904246.

Grmek:1966:PD

[1518] M. D. Grmek. Paul Diepgen (1878–1966). Revue d’Histoire des Sciences et de Leurs Applications, 19(3):271–273, juillet–septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904247.

Taton:1966:AG

[1519] René Taton. Albert Gloden (1901–1966). Revue d’Histoire des Sciences et de Leurs Applications, 19(3):274, juillet–septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904248.

Delorme:1966:F

[1520] Suzanne Delorme. France. Revue d’Histoire des Sciences et de Leurs Applications, 19(3):275–279, juillet–septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904249.

Poulle:1966:AOBb

[1521] Emmanuel Poulle. Analyses d’ouvrages: Comité international de Paléographie, Catalogue des manuscrits en écriture latine portant des indications de date, de lieu ou de copiste, T. V : Est de la France par Charles Samaran; Robert Marichal; Monique Garand; Madeleine Mabille; Josette Metman; Marie–Thérèse Vernet. Revue d’Histoire des Sciences et de Leurs Applications, 19(3):280–281, juillet–septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904251.

Brown:1966:AOB

[1522] Harcourt Brown. Analyses d’ouvrages: The Correspondence of Henri Oldenburg, vol. I : 1641–1662 par A. Rupert Hall; Marie Boas Hall; Eberhard Reichmann. Revue d’Histoire des Sciences et de Leurs Applications, 19(3):281–283, juillet–septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904252.

Delorme:1966:AOBa

[1523] Suzanne Delorme. Analyses d’ouvrages: Institut de France. Académie des Sciences. Troisième Centenaire, 1666–1966. Quelques souvenirs rassemblés et décrits par Pierre Gauja; Mlle Marie–Thérèse Delom. Revue d’Histoire des Sciences et de Leurs Applications, 19(3):283–285, juillet– septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904253.

Delorme:1966:AOBb

[1524] Albert Delorme. Analyses d’ouvrages: Colloque international sur l’Histoire de la Biologie marine. Les grandes expéditions scientifiques et la création des laboratoires maritimes, Banyuls-sur-Mer, 2–6 septembre 1963. (Publications du Laboratoire Arago. Supplément n^{o} 19 à Vie et Milieu). Revue d’Histoire des Sciences et de Leurs Applications, 19(3): 285–287, juillet–septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904254.

Colnort:1966:AOBd

[1525] S. Colnort. Analyses d’ouvrages: Les Aphorismes par Hippocrate; Daremberg; Siné. Revue d’Histoire des Sciences et de Leurs Applications, 19(3):287, juillet–septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904255.

Grmek:1966:AOBi

[1526] M. D. Grmek. Analyses d’ouvrages: Die Wiener medizinische Schule im 19. Jahrhundert, (coll. ^{≪}Studien zur Geschichte der Universität Wien ^{≫}, vol. 6) par Erna Lesky. Revue d’Histoire des Sciences et de Leurs Applications, 19(3):288, juillet–septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904256.

Bachmakova:1966:DF

[1527] Isabelle Bachmakova. Diophante et Fermat. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):289–306, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905707.

Dame:1966:GTS

[1528] Bernard Dame. Galilée et les taches solaires (1610–1613). Revue d’Histoire des Sciences et de Leurs Applications, 19(4):307–370, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905708.

Huard:1966:CMJ

[1529] Pierre Huard. Le centenaire de la mort de Joseph-François Malgaigne (1806–1865). Revue d’Histoire des Sciences et de Leurs Applications, 19(4):371–382, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905709.

Theodorides:1966:AE

[1530] Jean Théodoridès. Autriche: Expositions. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):383, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905710.

Theodorides:1966:Fb

[1531] Jean Théodoridès and P. Huard. France. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):383–386, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905711.

Theodorides:1966:GBX

[1532] J. Théodoridès. Grande-Bretagne: XIII^{e} Congrès International des Études Byzantines. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):386–387, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905714.

Anonymous:1966:PBI

[1533] Anonyme. Pays-Bas: III^{e} Congrès International de Logique, Méthodologie et Philosophie des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):387, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905715.

Anonymous:1966:SSG

[1534] Anonyme. Suisse: Schweizerischen Gesellschaft für Geschichte der Medizin und der Naturwissenschaften. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):387, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905716.

Grmek:1966:YCL

[1535] M. D. Grmek. Yougoslavie: Le centenaire de L’Académie Yougoslave des Sciences et des Arts. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):388–389, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905717.

Rochot:1966:AOBb

[1536] B. Rochot. Analyses d’ouvrages: Études d’histoire de la pensée scientifique, (Bibliothèque de Philosophie contemporaine) par Alexandre Koyré. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):390– 391, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23905718.

Mugler:1966:AOB

[1537] Ch. Mugler. Analyses d’ouvrages: Natura e logo. La razionalizzazione dell’ esperienza da Omero a Socrate par Emanuele Riverso. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):391–393, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905719.

Beaujouan:1966:AOBa

[1538] Guy Beaujouan. Analyses d’ouvrages: La bibliothèque scientifique d’un imprimeur humaniste au XV^{e} siècle : catalogue des manuscrits d’Arnaud de Bruxelles à la Bibliothèque nationale de Paris, ^{≪}Travaux d’Humanisme et Renaissance ^{≫}, LVII par Emmanuel Poulle. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):393–394, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905720.

Beaujouan:1966:AOBb

[1539] Guy Beaujouan. Analyses d’ouvrages: Archimedes in the Middle Ages : Vol. I, The Arabo-Latin tradition par Marshall Clagett. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):394–396, October 1966. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905721.

Beaujouan:1966:AOBc

[1540] Guy Beaujouan. Analyses d’ouvrages: Geschichte der Mathematik im Mittelalter par A. P. Juschkewitsch. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):396–398, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23905722.

Taton:1966:AOBb

[1541] R. Taton. Analyses d’ouvrages: ^{≪}Les Mechaniques ^{≫}de Galilée, Mathématicien & Ingénieur du Duc de Florence, avec plusieurs additions rares et nouvelles, utiles aux Architectes, Ingénieurs, Fonteniers, Philosophes & Artisans. (coll. ^{≪}Le mouvement des idées au XVII^{e} siècle ≫) par P. Marin Mersenne; Bernard Rochot. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):398–399, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905723.

Taton:1966:AOBc

[1542] R. Taton. Analyses d’ouvrages: Las Revoluciones de las esferas celestes. Libro primero, (^{≪}Los fundamentales ^{≫}) par Nicolas Copernico; A. Koyré; J. F. Chiti; El Mensajero de los astros, (^{≪}Los fundamentales ^{≫}) par Galileo Galilei; J. Babini. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):399–400, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905724.

Taton:1966:AOBd

[1543] R. Taton. Analyses d’ouvrages: Kepler’s Conversation with Galileo’s Sidereal Messenger, (^{≪}The Sources of Science ^{≫}, n^{o} 5) par J. Kepler; Edward Rosen. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):400–403, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905725.

Itard:1966:AOBb

[1544] Jean Itard. Analyses d’ouvrages: Mécanique analytique. Édition complète réunissant les notes de la 3^{e} édition revue par J.-L. Lagrange; Joseph Bertrand; Gaston Darboux. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):403–404, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23905726.

Canguilhem:1966:AOB

[1545] G. Canguilhem. Analyses d’ouvrages: Cahier de notes (1850–1860) par Claude Bernard; M. D. Grmek; Robert Courrier. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):405–406, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905727.

Anonymous:1966:TGM

[1546] Anonyme. Table générale des matières du Tome XIX. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):407–415, October 1966. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905728.

Anonymous:1966:BMa

[1547] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):??, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904855.

Anonymous:1966:BMb

[1548] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):??, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903701.

Anonymous:1966:BMc

[1549] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 19(3):??, juillet–septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904257.

Anonymous:1966:BMd

[1550] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):??, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905729.

Anonymous:1966:FMa

[1551] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):??, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904823.

Anonymous:1966:FMb

[1552] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 19(2):??, avril–juin 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903675.

Anonymous:1966:FMc

[1553] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 19(3):??, juillet–septembre 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904240.

Anonymous:1966:FMd

[1554] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 19(4):??, October 1966. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905706.

Ogigova:1967:LCH

[1555] H. Ogigova. Les lettres de Ch. Hermite à A. Markoff, 1885–1899. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):1–32, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904082.

Dulieu:1967:AG

[1556] Louis Dulieu. Antoine Gouan (1733–1821). Revue d’Histoire des Sciences et de Leurs Applications, 20(1):33–48, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904083.

Szwejcerowa:1967:TLP

[1557] Aniela Szwejcerowa. Trois lettres de Philippe de Girard (1775–1845) ingénieur en chef des Mines du Royaume de Pologne. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):49–61, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904084.

Klein:1967:FBA

[1558] Marc Klein. ^{≪}Traité sur l’acier d’Alsace ou l’art de convertir le fer en acier ^{≫}(1737) de Gilles-Augustin Bazin (1681–1754). Revue d’Histoire des Sciences et de Leurs Applications, 20(1):63–66, January 1967. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904085.

Russo:1967:NTD

[1559] Fr. Russo. Note sur la traduction du titre de l’ouvrage de Galilée ≪Sidereus nuncius ^{≫}. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):67–69, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904086.

Klastersky:1967:BN

[1560] I. Klásterský. Bohumil Němec (1873–1966). Revue d’Histoire des Sciences et de Leurs Applications, 20(1):69–71, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904087.

Anonymous:1967:UIH

[1561] Anonyme. Union Internationale d’Histoire et de Philosophie des Sciences: Division d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):72, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904088.

Delorme:1967:Fa

[1562] Suzanne Delorme. France. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):72–76, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904089.

Anonymous:1967:GBC

[1563] Anonyme. Grande-Bretagne: Célébration du deuxième centenaire de John Dalton. Revue d’Histoire des Sciences et de Leurs Applications, 20 (1):76, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904090.

Taton:1967:AOBa

[1564] R. Taton. Analyses d’ouvrages: Lexikon der Geschichte der Naturwissenschaften, Biographen, Sachwörter und Bibliographien …, 4. und 5. Lieferung; Bewegung (Ende) par Josef Mayerhöfer; Daniel von Morley. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):77, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904091.

Huard:1967:AOBa

[1565] P. Huard and M. Wong. Analyses d’ouvrages: Science and civilisation in China. Physics and physical Technology, vol. IV-1 : vol. IV-2 par Joseph Needham; Wang Ling. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):78–81, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904092.

Taton:1967:AOBb

[1566] R. Taton. Analyses d’ouvrages: Principles of Hindu Reckoning. (^{≪}Publications in medieval Science ^{≫}, n^{o} 8) par Kushyar ibn Labban; Martin Levey; Marvin Petruck. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):82–83, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904093.

Speziali:1967:AOBa

[1567] Pierre Speziali. Analyses d’ouvrages: Peirescel’ Italia par Cecilia Rizza; Raymond Lebègue. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):83–87, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904094.

Itard:1967:AOBa

[1568] Jean Itard. Analyses d’ouvrages: Correspondance du P. Marin —, t. IX par Mersenne; Cornelis de Waard. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):88–90, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904095.

Huard:1967:AOBb

[1569] P. Huard. Analyses d’ouvrages: Descartes, sa méthode et ses erreurs en physiologie par L. Chauvois; J. Rostand. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):90–91, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904096.

Taton:1967:AOBc

[1570] René Taton. Analyses d’ouvrages: Œuvres complètes. T. I par Blaise Pascal; Jean Mesnard. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):91–93, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904097.

Huard:1967:AOBc

[1571] P. Huard. Analyses d’ouvrages: Epistolario, (en dépôt chez Olschki, à Firenze) par Bernardini Ramazzini; Pericle di Pietro. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):93–94, January 1967. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904098.

Taton:1967:AOBd

[1572] R. Taton. Analyses d’ouvrages: The Mathematical Works, vol. I, (^{≪}The sources of science ^{≫}, n^{o} 3) par Isaac Newton; Derek T. Whiteside. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):94–95, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904099.

Itard:1967:AOBb

[1573] Jean Itard. Analyses d’ouvrages: La méthode des fluxions et des suites infinies par Isaac Newton; M. de Buffon. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):95–97, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904100.

Roger:1967:AOB

[1574] Jacques Roger. Analyses d’ouvrages: Buffon avant l’ ^{≪}Histoire naturelle ≫, (Publications de la Faculté des Lettres et Sciences Humaines de Paris, série ^{≪}Recherches ^{≫}, t. XXIV) par Leslie Hanks. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):97–98, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904101.

Sadoun-Goupil:1967:AOB

[1575] Michelle Sadoun-Goupil. Analyses d’ouvrages: Vie de Lavoisier par Léon Velluz. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):99, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904102.

D:1967:AOBa

[1576] Leticia Halperin D. Analyses d’ouvrages: The development of modern Chemistry par Aaron J. Ihde. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):100, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904103.

D:1967:AOBb

[1577] Leticia Halperin D. Analyses d’ouvrages: A history of the concept of valence par W. G. Palmer. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):101, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904104.

Limoges:1967:AOB

[1578] Camille Limoges. Analyses d’ouvrages: Darwin’s Notebooks on Transmutation of species, Bulletin of the British Museum (Natural History), Historical Series, vol. 2, n^{os} 2, 3, 4, 5, 6 par Gavin de Beer. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):101–103, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904105.

Huard:1967:AOBd

[1579] P. Huard. Analyses d’ouvrages: Les chirurgiens navigants de Lorient et de Nantes au XVIII^{e} siècle. Propos historiques autour de quatre fichiers (thèse de Nantes, 1964–1965) par Roger Grolier; L’aspect naval des études médicales pratiques à Nantes aux XVII^{e} et XVIII^{e} siècles (thèse de Nantes, 1965–1966) par Alain Prigent. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):103–105, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904106.

Speziali:1967:AOBb

[1580] Pierre Speziali. Analyses d’ouvrages: Einstein par Boris G. Kouznetsov; V. Talmy. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):105–108, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904107.

Goblot:1967:QBA

[1581] Henri Goblot. Sur quelques barrages anciens et la genèse des barragesvoûtes. Revue d’Histoire des Sciences et de Leurs Applications, 20 (2):109–140, avril–juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903833.

Birembaut:1967:LMM

[1582] Arthur Birembaut. L’exposition de modèles de machines à Paris en 1683. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):141–158, avril–juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903834.

Costabel:1967:TAP

[1583] Pierre Costabel. Tentatives d’aménagement des pompes aspirantes en 1766. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):159– 168, avril–juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903835.

deBeaurepaire-Louvagny:1967:EMA

[1584] B. de Beaurepaire-Louvagny. Une exploitation minière au XVIII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):169–186, avril–juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903836.

Payen:1967:BIF

[1585] Jacques Payen and M. le chr. de Bétancourt. Bétancourt et l’introduction en France de la machine à vapeur à double effet (1789). Revue d’Histoire des Sciences et de Leurs Applications, 20(2):187–198, avril–juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903837.

Daumas:1967:MMA

[1586] Maurice Daumas. Un microscope de Magny au musée de Lille. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):199–201, avril– juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903839.

Soulard:1967:TI

[1587] Robert Soulard. Sur la transmission des images. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):202–206, avril–juin 1967. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903840.

Garanger:1967:ASI

[1588] A. Garanger. Au sujet de l’invention de l’acier rapide. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):207–208, avril–juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903841.

Anonymous:1967:FMH

[1589] Anonyme. France: Musée de l’Histoire du Fer. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):209–210, avril–juin 1967. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903842.

Daumas:1967:AOBa

[1590] Maurice Daumas. Analyses d’ouvrages: The Principal Works of Simon Stevin. Vol. V : Engineering, Music, Civic Life par R. J. Forbes; A. D. Fokker; A. Romein-Verschoor. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):211–212, avril–juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903843.

Daumas:1967:AOBb

[1591] Maurice Daumas. Analyses d’ouvrages: The Development of Technical Education in France, 1500–1850 par Frederick B. Artz. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):212–214, avril–juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903844.

Payen:1967:AOBa

[1592] Jacques Payen. Analyses d’ouvrages: Tin in Social and Economic History par Ernest S. Hedges. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):214–215, avril–juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903845.

Anonymous:1967:AOB

[1593] Anonyme. Analyses d’ouvrages: Contributions from the Museum of History and Technology. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):215, avril–juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903846.

Anonymous:1967:AQP

[1594] Anonyme. Analyses de quelques périodiques: The Journal of Industrial Archaeology, vol. II, n^{os} 1 et 2, 1965 : The Lambarde Press, in association with the Newcomen Society; vol. II, n^{os} 3 et 4, 1965; vol. III, n^{os} 1 à 4, 1966. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):215– 217, avril–juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903847.

Anonymous:1967:BTT

[1595] Anonyme. Blätter für Technikgeschichte, n^{os} 26 (1964), 27 (1965) et 28 (1966). Technisches Museum für Industrie und Gewerbe. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):217–218, avril– juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903848.

Anonymous:1967:TT

[1596] Anonyme. Technikgeschichte, tomes 32 (1965) et 33 (1966). Revue d’Histoire des Sciences et de Leurs Applications, 20(2):218–219, avril– juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903849.

Payen:1967:AOBb

[1597] Jacques Payen. Analyses d’ouvrages: Agrikultura. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):219–220, avril–juin 1967. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903850.

Mugler:1967:DDI

[1598] Charles Mugler. Démocrite et les dangers de l’irradiation cosmique. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):221– 228, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905785.

Baker:1967:DCA

[1599] K. M. Baker. Les débuts de Condorcet au secrétariat de l’Académie royale des Sciences (1773–1776). Revue d’Histoire des Sciences et de Leurs Applications, 20(3):229–280, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905786.

Callot:1967:DDV

[1600] Émile Callot. Un discours de Dominique Villar sur l’histoire naturelle. Revue d’Histoire des Sciences et de Leurs Applications, 20(3): 281–288, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905787.

Gillmor:1967:ECF

[1601] C. Stewart Gillmor. Espagne: Les cahiers^{≪}perdus^{≫}de Léonard de Vinci. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):289, juillet– septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905788.

Huard:1967:PMP

[1602] P. Huard. Principauté de Monaco: premier Congrès International d’Histoire de l’Océanographie (Monaco, 12–17 décembre 1966). Revue d’Histoire des Sciences et de Leurs Applications, 20(3):289–290, juillet– septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905789.

Delorme:1967:Fb

[1603] Suzanne Delorme and S. Colnort. France. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):290–296, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905790.

Poulle:1967:AOB

[1604] Emmanuel Poulle. Analyses d’ouvrages: Inventaire des manuscrits scientifiques de la Bibliothèque royale de Belgique, t. I par Roger Calcoen. Revue d’Histoire des Sciences et de Leurs Applications, 20(3): 297–298, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905796.

Rashed:1967:AOB

[1605] R. Rashed. Analyses d’ouvrages: Die griechischen Quellen des Sleinbuches von Al-Beruni. Extrait des Annales archéologiques de Syrie, vol. XV, t. 2 par Mohamed Yahia Haschmi. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):298–299, juillet–septembre 1967. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905797.

Rochot:1967:AOBa

[1606] B. Rochot. Analyses d’ouvrages: Dialogues et Lettres choisies. Collection ≪Histoire de la Pensée^{≫}par Galilée; P.-H. Michel; G. di Santillana. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):300–301, juillet– septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905798.

Rochot:1967:AOBb

[1607] B. Rochot. Analyses d’ouvrages: Études galiléennes, 2^{e} éd. (Histoire de la Pensée, XV) par Alexandre Koyré. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):301–302, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905799.

Itard:1967:AOBc

[1608] Jean Itard. Analyses d’ouvrages: Histoire des logarithmes de Néper à Euler, t. I: La découverte des logarithmes et le calcul des premières tables par Charles Naux. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):302, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905800.

Itard:1967:AOBd

[1609] Jean Itard. Analyses d’ouvrages: Briefwechsel 1729–1764 par Leonhard Euler; Christian Goldbach. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):303–306, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905801.

Gillmor:1967:AOB

[1610] C. Stewart Gillmor. Analyses d’ouvrages: Razvitie Nauki o Trenii — sukhoe trenie (Développement de la science du frottement — frottement sec) par I. V. Kragelsky; V. S. Schedrov. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):306–307, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905802.

Helperin-Donghi:1967:AOB

[1611] Leticia Helperin-Donghi. Analyses d’ouvrages: L’histoire de la découverte du système périodique des éléments chimiques et l’apport de Béguyer de Chancourtois par J. W. van Spronsen. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):307, juillet–septembre 1967. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23905803.

Taton:1967:AOBe

[1612] R. Taton. Analyses d’ouvrages: Les pionniers de la science canadienne par G. F. G. Stanley. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):308, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905804.

Huard:1967:AOBe

[1613] P. Huard. Analyses d’ouvrages: Histoire de la Biologie, collection ≪Que sais-je? ^{≫}par J. Théodoridès. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):309, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905805.

Grmek:1967:AOBa

[1614] M. D. Grmek. Analyses d’ouvrages: Epistolario, Vol. I (1750–1773), vol. II (1774–1781), vol. III (1782–1787), vol. IV (1788–1793), vol. V (1794– 1799) par Lazzaro Spallanzani; Benedetto Biagi; Plinio Fraccaro. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):309–310, juillet– septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23905806.

Wong:1967:AOBa

[1615] M. Wong. Analyses d’ouvrages: Le Japon du XVIII^{e} siècle vu par un botaniste suédois, Ch. P. Thunberg. (Temps et Continents) par Claude Gaudon; L. Langlès. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):310–312, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905807.

Leroy:1967:AOBa

[1616] J. F. Leroy. Analyses d’ouvrages: Origins of Mendelism par R. C. Olby. Revue d’Histoire des Sciences et de Leurs Applications, 20(3): 312–313, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23905808.

Laissus:1967:AQP

[1617] Yves Laissus. Analyses de quelques périodiques: Gesnerus, Vierteljahrsschrift, herausgegeben von der schweizerischen Gesellschaft für Geschichte der Medizin und der Naturwissenschaften, vol. 22. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):313–315, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905809.

Colnort:1967:AQPa

[1618] S. Colnort. Analyses de quelques périodiques: Physis, Rivista internationale di storia della scienza. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):315, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23905810.

Colnort:1967:AQPb

[1619] S. Colnort. Analyses de quelques périodiques: Archives internationales d’histoire des sciences, t. XIX. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):315–316, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23905811.

Anonymous:1967:OR

[1620] Anonyme. Ouvrages reçus. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):316, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905812.

Hara:1967:RQS

[1621] Kokiti Hara. Remarque sur la quadrature de la surface du cône oblique. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):317–332, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904706.

Itard:1967:ARQ

[1622] Jean Itard. Autre remarque sur la quadrature de la surface du cône oblique. Revue d’Histoire des Sciences et de Leurs Applications, 20 (4):333–335, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904707.

Chabbert:1967:FC

[1623] Pierre Chabbert. Fermat à Castres. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):337–348, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904708.

Leroy:1967:AHP

[1624] Jean François Leroy. Adanson dans l’histoire de la pensée scientifique. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):349–360, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904709.

Guedes:1967:MTA

[1625] Michel Guédès. La méthode taxonomique d’Adanson. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):361–386, October 1967. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904710.

Delorme:1967:TCP

[1626] Suzanne Delorme. Tchécoslovaquie: Colloque de Prague 25–29 septembre 1967 La Révolution scientifique du XVIIe siècle et les Sciences mathématiques et physiques. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):387–390, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904711.

Anonymous:1967:YCG

[1627] Anonyme. Yougoslavie: Commémoration de Ghetaldus. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):390, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904712.

Itard:1967:AOBe

[1628] Jean Itard. Analyses d’ouvrages: El ^{≪}Metodo ^{≫} par Archimède; José Babini. Revue d’Histoire des Sciences et de Leurs Applications, 20(4): 391, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904713.

Taton:1967:AOBf

[1629] R. Taton. Analyses d’ouvrages: Les Œuvres d’Euclide par F. Peyrard; Jean Itard. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):391–392, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904714.

Colnort:1967:AOBa

[1630] Suzanne Colnort. Analyses d’ouvrages: La science arabe et son rôle dans l’évolution scientifique mondiale par Aldo Mieli; A. Mazahéri. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):392, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904715.

Colnort:1967:AOBb

[1631] Suzanne Colnort. Analyses d’ouvrages: Science since 1500, 2^{e} éd. par H. T. Pledge. Revue d’Histoire des Sciences et de Leurs Applications, 20(4): 393, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904716.

Speziali:1967:AOBc

[1632] Pierre Speziali. Analyses d’ouvrages: La théorie de la relativité restreinte, exposé élémentaire par Albert Einstein; Henri Montias. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):393–394, October 1967. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904717.

Bouligand:1967:AOB

[1633] G. Bouligand. Analyses d’ouvrages: La philosophie mathématique d’Henri Poincaré. (Coll. log. math., Série A) par J. J. A. Mooij. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):394–395, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904718.

Itard:1967:AOBf

[1634] Jean Itard. Analyses d’ouvrages: La géométrie élémentaire, (coll. ^{≪}Que sais-je? ^{≫}, n^{o} 1211) par André Delachet. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):395–397, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904719.

Grmek:1967:AOBb

[1635] M. D. Grmek. Analyses d’ouvrages: The Life of William Harvey par Geoffrey Keynes. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):397–398, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904720.

Leroy:1967:AOBb

[1636] J. F. Leroy. Analyses d’ouvrages: The History of Cell Respiration and Cytochrome par David Keilin. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):398–399, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904721.

Wong:1967:AOBb

[1637] Ming Wong. Analyses d’ouvrages: The Yellow Emperor’s Classic of Internal Medicine. New Edition par Ilza Veith. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):399–401, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904722.

Huard:1967:AOBf

[1638] Pierre Huard. Analyses d’ouvrages: La médecine navale française au XIX^{e} siècle par Liliane Pellet. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):401–402, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904723.

Colnort:1967:AOBc

[1639] Suzanne Colnort. Analyses d’ouvrages: Die zoologische Buchillustration. Ihre Bibliographie und Geschichte, 3 fasc par Claude Nissen. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):402, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904724.

Colnort:1967:AOBd

[1640] S. Colnort. Analyses d’ouvrages: Les Conférences du Palais de la Découverte, série D: Histoire des Sciences, fascicules publiés depuis le n^{o} 105 (1966). Revue d’Histoire des Sciences et de Leurs Applications, 20(4): 402, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904725.

Anonymous:1967:FXC

[1641] Anonyme. France: XII^{e} Congrès International d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):403–404 390, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904726.

Anonymous:1967:TGM

[1642] Anonyme. Table générale des matières du Tome XX. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):405–412, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904727.

Anonymous:1967:BMa (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904108.

Anonymous:1967:BMb

[1644] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):??, avril–juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903851.

Anonymous:1967:BMc

[1645] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):??, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905813.

Anonymous:1967:BMd

[1646] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):??, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23911247.

Anonymous:1967:FMa

[1647] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 20(1):??, January 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904081.

Anonymous:1967:FMb

[1648] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 20(2):??, avril–juin 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903832.

Anonymous:1967:FMc

[1649] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 20(3):??, juillet–septembre 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23905784.

Anonymous:1967:FMd

[1650] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 20(4):??, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904705.

Delorme:1968:A

[1651] Suzanne Delorme. Avertissement. Revue d’Histoire des Sciences et de Leurs Applications, 21(1):1–2, January 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903578.

Anonymous:1968:TAA

[1652] Anonyme. Table alphabétique des auteurs. Revue d’Histoire des Sciences et de Leurs Applications, 21(1):3–32, January 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903579.

Anonymous:1968:TOP

[1653] Anonyme. Table des ouvrages et périodiques analysés. Revue d’Histoire des Sciences et de Leurs Applications, 21(1):33–74, January 1968. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903580.

Anonymous:1968:TAM

[1654] Anonyme. Table analytique des matières. Revue d’Histoire des Sciences et de Leurs Applications, 21(1):75–116, January 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903581.

Marek:1968:PTD

[1655] Jiri Marek. Un physicien tchèque du XVII^{e} siècle : Ioannes Marcus Marci de Kronland (1595–1667). Revue d’Histoire des Sciences et de Leurs Applications, 21(2):109–130, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904470.

Simonov:1968:RED

[1656] N. I. Simonov. Sur les recherches d’Euler dans le domaine des équations différentielles. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):131–156, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904471.

Infantozzi:1968:MEG

[1657] Carlos Alberto Infantozzi. Sur la mort d’Évariste Galois. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):157–160, avril– juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904472.

Rederowa:1968:VSC

[1658] Danuta Rederowa. La vie scientifique à cracovie au XIX^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):161–167, avril– juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904473.

Grmek:1968:EUC

[1659] M. D. Grmek. Aux États-unis, un colloque international sur les rapports entre la biologie, l’histoire et la philosophie Denver, 27 novembre-2 décembre 1967. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):169–171, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904474.

Anonymous:1968:Fa

[1660] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):171–172, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904475.

Rashed:1968:AOB

[1661] R. Rashed. Analyses d’ouvrages: Histoire générale des sciences, t. I : La science antique et médiévale (Des origines à 1450), 2^{e} éd. par René Taton. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):173– 178, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904476.

Grmek:1968:AOBa

[1662] M. D. Grmek. Analyses d’ouvrages: Enseignement et diffusion des sciences en France au XVIII^{e} siècle, Coll. ^{≪}Histoire de la pensée ^{≫}, vol. XI par René Taton. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):178–180, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904477.

Itard:1968:AOBa

[1663] Jean Itard. Analyses d’ouvrages: Extraits des Éléments, coll. ^{≪}Grands classiques des sciences et des techniques. Un savant dans le texte ^{≫} par Euclide; Charles Mugler. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):180, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904478.

Lerner:1968:AOB

[1664] Michel-Pierre Lerner. Analyses d’ouvrages: Kepler’s Somnium. The Dream, or posthumous work on lunar Astronomy par E. Rosen. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):180–182, avril– juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904479.

Colnort:1968:AOBa

[1665] S. Colnort. Analyses d’ouvrages: Dialogue concerning the two chief World Systems, second revised edition par Galileo Galilei; Stillman Drake; Albert Einstein. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):182, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904480.

Itard:1968:AOBb

[1666] Jean Itard. Analyses d’ouvrages: Sidereus Nuncius, Nachricht von neuen Sternen, Dialog über die Weltsysteme (Auswahl), Vermessung der Hölle Dantes, Marginalien zu Tasso par Galileo Galilei; Hans Blumenberg. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):182, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904481.

Itard:1968:AOBc

[1667] Jean Itard. Analyses d’ouvrages: Correspondance du P. Marin Mersenne, Religieux Minime. Tome X : Du 6 août 1640 à la fin décembre 1641 par Paul Tannery; Cornelis de Waard; Bernard Rochot. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):183–185, avril– juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904482.

Itard:1968:AOBd

[1668] Jean Itard. Analyses d’ouvrages: Studien zur Mathematik des John Wallis (1616–1703) par Christoph J. Scriba. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):185–186, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904483.

Itard:1968:AOBe

[1669] Jean Itard. Analyses d’ouvrages: The mathematical papers of Isaac Newton, volume I, 1664–1666 par D. T. Whiteside. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):186–188, avril–juin 1968. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904484.

Itard:1968:AOBf

[1670] Jean Itard. Analyses d’ouvrages: The Mathematical papers of Isaac Newton, vol. II, 1667–1670 par D. T. Whiteside; M. A. Hoskin. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):188–190, avril– juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904485.

Itard:1968:AOBg

[1671] Jean Itard. Analyses d’ouvrages: The correspondence of Isaac Newton, Vol. IV : 1694–1709 par J. F. Scott. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):190–191, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904486.

Itard:1968:AOBh

[1672] Jean Itard. Analyses d’ouvrages: The Mathematical papers of Sir William Rowan Hamilton, vol. III : Algebra par H. Halberstam; R. E. Ingram. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):191– 193, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904487.

Itard:1968:AOBi

[1673] Jean Itard. Analyses d’ouvrages: Émile Borel philosophe et homme d’action. ^{≪}Les grands problèmes des Sciences ^{≫}, n^{o} 19 par Maurice Fréchet. Revue d’Histoire des Sciences et de Leurs Applications, 21(2): 194, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904488.

Speziali:1968:AOBa

[1674] Pierre Speziali. Analyses d’ouvrages: Louis de Broglie et la mécanique ondulatoire, (Savants du monde entier) par Marie-Antoinette Tonnelat. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):194–196, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904489.

Rashed:1968:BLI

[1675] Roshdi Rashed. Le ^{≪}Discours de la Lumière ^{≫}d’Ibn al-Haytham (Alhazen): Traduction française critique. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):197–224, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23909912.

Chauvois:1968:PDG

[1676] L. Chauvois and P. Huard. À propos du De generatione animalium d’Harvey et de ses traductions: Harvey chirurgien. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):225–231, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23909913.

Dulieu:1968:FBB

[1677] Louis Dulieu. François-Bourguignon de Bussières de Lamure (1717– 1787). Revue d’Histoire des Sciences et de Leurs Applications, 21(3): 233–244, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23909914.

Sahuc:1968:EHC

[1678] L. J.-M. Sahuc. Étude historique des coaptations chez les Invertébrés. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):245– 264, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23909915.

Anonymous:1968:PBT

[1679] Anonyme. Pays-Bas: Tricentenaire de la naissance de Boerhaave. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):265, juillet– septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23909916.

Speziali:1968:AOBb

[1680] Pierre Speziali. Analyses d’ouvrages: Bau und Bildung des Weltalls. Kosmologische Vorstellungen in Dokumenten aus zwei Jahrtausenden par Bernhard Sticker. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):266–267, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23909917.

Itard:1968:AOBj

[1681] Jean Itard. Analyses d’ouvrages: Œuvres Mathématiques de Riemann par L. Laugel. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):267–268, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23909918.

Itard:1968:AOBk

[1682] Jean Itard. Analyses d’ouvrages: Cantor a tort par Georges Antoniadès Metrios. Revue d’Histoire des Sciences et de Leurs Applications, 21 (3):268, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23909919.

Itard:1968:AOBl

[1683] Jean Itard. Analyses d’ouvrages: Arithmétique et algèbre modernes. III. Idéaux dans un domaine d’intégrité. Algèbre et arithmétique linéaires par Albert Châtelet; François Châtelet. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):268, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23909920.

Speziali:1968:AOBc

[1684] Pierre Speziali. Analyses d’ouvrages: Relativitätstheorie und Weltanschauung. Zur philosophischen und wissenschaftlichen Wirkung Albert Einsteins (Von einem Autorenkollektiv). Revue d’Histoire des Sciences et de Leurs Applications, 21(3):268–271, juillet–septembre 1968. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23909921.

Hutin:1968:AOB

[1685] Serge Hutin. Analyses d’ouvrages: Theatrum Chemicum Britannicum, a reprint of the London edition (1652). ^{≪}The Sources of Science ^{≫}, 39 par Elias Ashmole; Allen G. Debus. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):271–272, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23909922.

Sadoun-Goupil:1968:AOBa

[1686] Michelle Sadoun-Goupil. Analyses d’ouvrages: The Society of Arcueil. A view of French Science at the time of Napoleon I par Maurice Crosland. Revue d’Histoire des Sciences et de Leurs Applications, 21(3): 272–274, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23909923.

Sadoun-Goupil:1968:AOBb

[1687] Michèle Sadoun-Goupil. Analyses d’ouvrages: Antoine-François de Fourcroy (1755–1809), sa vie, son œuvre par Georges Kersaint. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):274–275, juillet– septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23909924.

Grmek:1968:AOBb

[1688] M. D. Grmek. Analyses d’ouvrages: William Harvey’s Biological Ideas. Selected Aspects and Historical Background par Walter Pagel. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):275–276, juillet– septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23909925.

Huard:1968:AOBa

[1689] P. Huard. Analyses d’ouvrages: Die historischen Grundlagen der Lebersforchung: I. Die Vorstellungen über Anatomie, Physiologie und Pathologie der Leber in der Antike: II. Die Geschichte der Lebersforchung, von Galen bis Claude Bernard, ^{≪}Basler Veröfentlichungen zur Geschichte der Medizin und der Biologie ^{≫}, IX, 2, 1965; XXI, 1967 par Nikolaus Mani. Revue d’Histoire des Sciences et de Leurs Applications, 21(3): 276–278, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23909926.

Colnort:1968:AOBb

[1690] S. Colnort. Analyses d’ouvrages: A Bibliography of the writings of Henry E. Sigerist par Genevieve Miller. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):278, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23909927.

Theodorides:1968:AOB

[1691] J. Théodoridès. Analyses d’ouvrages: Mille ans de chirurgie en Occident: V^{e}–XI^{e} siècles par P. Huard; M. D. Grmek. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):278–279, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23909928.

Huard:1968:AOBb

[1692] P. Huard. Analyses d’ouvrages: Histoire de la lèpre et de son influence sur la littérature et les arts par Christian Malet. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):279–280, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23909929.

Poulle:1968:AOB

[1693] Emmanuel Poulle. Analyses d’ouvrages: Alt-Augsburger Kompassmacher; Sonnen-, Mond-, und Sternuhren, astronomische und mathematische Geräte, Rädeuhren. (Abhandlungen zur Geschichte der Stadt Augsburg, Schriftenreihe des Stadtarchivs Augsburg, Bd. 16) par Maximilian Bobinger. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):280–281, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23909930.

Grmek:1968:AOBc

[1694] M. D. Grmek. Analyses d’ouvrages: A History of the Thermometer and Its Use in Meteorology par W. E. Knowles Middleton. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):281–282, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23909931.

Pueyo:1968:AOB

[1695] Guy Pueyo. Analyses d’ouvrages: L’histoire du climat depuis l’an mil, (Nouvelle Bibliothèque scientifique) par E. Le Roy Ladurie. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):282–284, juillet– septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23909932.

Grmek:1968:IDP

[1696] M. D. Grmek. Les idées de Descartes sur le prolongement de la vie et le mécanisme du vieillissement. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):285–302, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904263.

Chabbert:1968:PB

[1697] Pierre Chabbert. Pierre Borel (1620?–1671). Revue d’Histoire des Sciences et de Leurs Applications, 21(4):303–343, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904265.

Dobrovolskij:1968:DTV

[1698] W. Dobrovolskij. Développement de la théorie des vecteurs et des quaternions dans les travaux des mathématiciens russes du XIX^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):345–349, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904266.

Grmek:1968:XCI

[1699] M. D. Grmek. XII^{e} Congrès international d’Histoire des Sciences Paris, 26–31 août 1968. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):351–355, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904268.

Anonyme:1968:IIF

[1700] Anonyme. II. Informations. France. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):356, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904270.

Huard:1968:AOBc

[1701] P. Huard. Analyses d’ouvrages: Bio-bibliographie de la médecine japonaise par Jean-Michel Simon. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):357, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904272.

Huard:1968:AOBd

[1702] P. Huard. Analyses d’ouvrages: Contribution à l’étude de l’histoire de la circulation du sang, de Galien à Harvey par Paul Quichoux. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):357–358, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904274.

Huard:1968:AOBe

[1703] P. Huard. Analyses d’ouvrages: Les anatomistes bretons par J.-P. Devèze. Revue d’Histoire des Sciences et de Leurs Applications, 21 (4):358–360, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904276.

Huard:1968:AOBf

[1704] P. Huard. Analyses d’ouvrages: Contribution à l’étude de la structure hospitalière en France à la fin du XVIII^{e} siècle par Jean Auffret. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):360, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904278.

Huard:1968:AOBg

[1705] Pierre Huard. Analyses d’ouvrages: Catalogue des Archives de l’Académie royale de Chirurgie par Salem Yaqubi. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):360–361, October 1968. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904280.

Huard:1968:AOBh

[1706] P. Huard. Analyses d’ouvrages: Contribution à l’étude des thèses du Collège de Chirurgie de Paris (1749–1789) par Jeannine Livinec. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):361, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904282.

Huard:1968:AOBi

[1707] P. Huard. Analyses d’ouvrages: Journal du voyage médico-chirurgical de Jean Hunczovski (1752–98) en Europe occidentale (1782) par Dominique Mounier. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):361–362, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904284.

Huard:1968:AOBj

[1708] P. Huard. Analyses d’ouvrages: Les chirurgiens navigants de la Marine royale à Brest dans les dernières années de l’Ancien Régime (1780–90) par J.-Cl. Le Néel. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):362–363, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904286.

Colnort:1968:AOBc

[1709] S. Colnort. Analyses d’ouvrages: NTM. Zeitschrift für Geschichte der Naturwissenschaften, Technik und Medizin, devenu, en 1965 : NTM. Schriftenreihe für Geschichte der Naturwissenschaften, Technik und Medizin, 1. Jahrgang, 1960–1964, Heft 1, 2, 3, 4; 2. Jahrgang, 1965, Heft 5, 6; 3. Jahrgang, 1966, Heft 7, 8. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):364–365, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904288.

Colnort:1968:AQP

[1710] S. Colnort. Analyses de quelques périodiques: Clio medica, Acta Academiae internationalis Historiae Medicinae, vol. I et II, 1965–1966 et 1967. Revue d’Histoire des Sciences et de Leurs Applications, 21(4): 366, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904290.

Anonymous:1968:TGM

[1711] Anonyme. Table générale des matières du Tome XXI. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):367–372, October 1968. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904292.

Anonymous:1968:BMa

[1712] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 21(1):??, January 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903582.

Anonymous:1968:BMb

[1713] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):??, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904490.

Anonymous:1968:BMc

[1714] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):??, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23909933.

Anonymous:1968:BMd

[1715] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):??, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904294.

Anonymous:1968:FMa

[1716] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 21(1):??, January 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903577.

Anonymous:1968:FMb

[1717] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):??, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904469.

Anonymous:1968:FMc

[1718] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 21(3):??, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23909911.

Anonymous:1968:FMd

[1719] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 21(4):??, October 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904261.

Daumas:1969:LTS

[1720] Maurice Daumas. L’histoire des techniques : son objet, ses limites, ses méthodes. Revue d’Histoire des Sciences et de Leurs Applications, 22(1): 5–32, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903761.

Destombes:1969:ANC

[1721] Marcel Destombes. Un astrolabe nautique de la Casa de Contratación (Séville, 1563). Revue d’Histoire des Sciences et de Leurs Applications, 22(1):33–64, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903762.

Payen:1969:AMP

[1722] Jacques Payen. Autour de la machine de Pascal: une mise en garde nécessaire. (French) [About Pascal’s machine: a necessary warning]. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):65–66, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903763.

Payen:1969:FPP

[1723] Jacques Payen. Les frères Périer et la pompe à feu de l’hôpital de la Marine à Rochefort (1783). Revue d’Histoire des Sciences et de Leurs Applications, 22(1):66–69, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903764.

Payen:1969:LGE

[1724] Jacques Payen. Lueurs sur la genèse de l’enseignement technique supérieur au c.n.a.m. (1819). Revue d’Histoire des Sciences et de Leurs Applications, 22(1):70–75, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903765.

Soulard:1969:CL

[1725] Robert Soulard. Caselli et Lambrigot. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):76–78, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903766.

Soulard:1969:IEM

[1726] Robert Soulard. Sur l’invention de l’enregistrement magnétique. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):78–80, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903767.

Daumas:1969:AOBa

[1727] Maurice Daumas. Analyses d’ouvrages: Thalès. T. XII, année 1966 : Les commencements de la technologie. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):81–83, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903768.

Daumas:1969:AOBb

[1728] Maurice Daumas. Analyses d’ouvrages: Histoire des sources d’énergie, (Découverte de la Science). par Jacques Payen. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):83, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903769.

Daumas:1969:AOBc

[1729] Maurice Daumas. Analyses d’ouvrages: Jacques Vaucanson, mécanicien de génie, (Publication de la Faculté des Lettres et Sciences humaines de Grenoble.) par André Doyon; Lucien Liaigre. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):84–87, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903770.

Soulard:1969:AOBa

[1730] Robert Soulard. Analyses d’ouvrages: Les moteurs à combustion interne. (Grands Classiques des Sciences et des Techniques : Une invention dans les textes) par Jacques Payen; Huygens; Lebon; Lenoir; Beau de Rochas; Otto; Diesel. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):87–88, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903771.

Soulard:1969:AOBb

[1731] Robert Soulard. Analyses d’ouvrages: Cooke and Wheatstone and the Invention of the Electric Telegraph par Geoffrey Hubbard. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):88–89, January 1969. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903772.

Payen:1969:AOBa

[1732] Jacques Payen. Analyses d’ouvrages: L’énergie des marées. La soif du monde et le dessalement des eaux, (La Science vivante) par Robert Gibrat; Cyril Gomella. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):89, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903773.

Payen:1969:AOBb

[1733] Jacques Payen. Analyses d’ouvrages: Gebändigte Kraft. Die Geschichte der Erfindung des Otto-Motors par Gustav Goldbeck. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):89–90, January 1969. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903774.

Payen:1969:AOBc

[1734] Jacques Payen. Analyses d’ouvrages: Drei Generationen Bauingenieure : das Ingenieurbureau Gruner und die Entwicklung der Technick seit 1860 par Karl Mommsen. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):91, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903775.

Payen:1969:AOBd

[1735] Jacques Payen. Analyses d’ouvrages: Aspects de l’économie nivernaise au XIX^{e} siècle, (E.P.H.E., VI^{e} section, Centre d’Études économiques, ≪Études et Mémoires ^{≫}, n^{o} 60) par Guy Thuillier; Paul Leuilliot. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):91–94, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903776.

Daumas:1969:AOBd

[1736] Maurice Daumas. Analyses d’ouvrages: La genèse des innovations. La création technique dans l’activité de la firme, (Sciences sociales du Travail, n^{o} 3) par Jean-Louis Maunoury. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):94–95, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903777.

Payen:1969:AOBe

[1737] Jacques Payen. Analyses d’ouvrages: Savants méconnus, inventions oubliées par Robert Champeix. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):96, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903778.

Costabel:1969:DRC

[1738] Pierre Costabel. Descartes et la racine cubique des nombres binômes: Propos sur un document retrouvé concernant le dernier acte de la controverse Stampioen–Waessenaer. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):97–116, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor. org/stable/23904174. Propos sur un document retrouvé concernant le dernier acte de la controverse Stampioen-Waessenaer.

Raskine:1969:TOL

[1739] Naum Mikhaïlovitch Raskine. La théorie de l’oxygène à l’Académie des Sciences de Pétersbourg à la fin du XVIII^{e} et au commencement du XIX^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):117– 136, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904175.

Broc:1969:VGA

[1740] Numa Broc. Voyages et géographie au XVIII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):137–154, avril–juin 1969. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904176.

Costabel:1969:RMT

[1741] Pierre Costabel. Un registre de manuscrits témoin de l’activité de Mersenne en Italie en 1645. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):155–162, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904177.

Nickles:1969:SCG

[1742] M. Nicklès. Le service de la carte géologique de la France: À propos d’un centenaire. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):163–166, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904178.

Anonymous:1969:UCC

[1743] Anonyme. Unesco: Centenaire de la classification de Mendeleev. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):166–167, avril– juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904179.

Anonymous:1969:BIS

[1744] Anonyme. Belgique: III^{e} Symposium International de la Fédération Mondiale Coronelli des Amis du Globe III^{e} Conférence Internationale d’Histoire de la Cartographie Bruxelles, 17–20 septembre 1969. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):167, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904180.

Anonymous:1969:Fa

[1745] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):167–168, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904181.

Anonymous:1969:GB

[1746] Anonyme. Grande-Bretagne. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):168–169, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904183.

D:1969:ISF

[1747] S. D. Italie: Symposium sur ^{≪}science et technique chez Léonard de Vinci ≫Florence–Vinci, 23–26 juin 1969. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):169, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904184.

Anonymous:1969:TSP

[1748] Anonyme. Tchécoslovaquie: Symposium Purkynǎ à Prague 7–10 septembre 1969. Revue d’Histoire des Sciences et de Leurs Applications, 22 (2):169–170, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904185.

Bernhardt:1969:AOBa

[1749] Jean Bernhardt. Analyses d’ouvrages: Les sciences humaines et la pensée occidentale, I. De l’histoire des sciences à l’histoire de la pensée, 1966, II. Les origines des sciences humaines, 1967, III. La révolution galiléenne, 1969, t. 1, t. 2 par Georges Gusdorf. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):171–179, avril–juin 1969. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904187.

Itard:1969:AOBa

[1750] Jean Itard. Analyses d’ouvrages: Das Parallelenproblem im Corpus Aristotelicum, extrait de Archive for History of Exact Sciences par Imre Tóth. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):179–181, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904188.

Pedersen:1969:AOB

[1751] Kurt Møller Pedersen. Analyses d’ouvrages: L’expérience du mouvement, (Histoire de la Pensée, n^{o} 16) par Serge Moscovici. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):181–182, avril–juin 1969. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904189.

Borghgraef:1969:AOB

[1752] Béatrice Borghgraef. Analyses d’ouvrages: The lives of the professors of Gresham College to which is prefixed the life of the founder Sir Thomas Gresham, a facsimile of the London edition of 1740 (The Sources of Science, n^{o} 71) par John Ward. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):182–183, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904190.

Itard:1969:AOBb

[1753] Jean Itard. Analyses d’ouvrages: Mathematica, Œuvres complètes, t. XVII-2 par Malebranche; Pierre Costabel. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):183–186, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904191.

Donghi:1969:AOB

[1754] Leticia Halperin Donghi. Analyses d’ouvrages: John Dalton, 1766–1844. A bibliography of works par and about him par A. L. Smyth. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):186–187, avril– juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904192.

Huard:1969:AOBa

[1755] P. Huard. Analyses d’ouvrages: Les officiers de santé de la Marine française de 1814 à à 1835 par J. Léonard. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):187–188, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904193.

Huard:1969:AOBb

[1756] P. Huard and J. Théodoridès. Analyses d’ouvrages: L’exercice de la médecine en France à la fin du XIXe siècle et au début du XX^{e} siècle (honoraires, syndicats, éthique médicale) par J.-L. Cariage; René Moreau. Revue d’Histoire des Sciences et de Leurs Applications, 22 (2):188–189, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904194.

Huard:1969:AOBc

[1757] P. Huard. Analyses d’ouvrages: L’enseignement de la médecine en France des origines à nos jours par Ch. Coury. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):189–190, avril–juin 1969. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904195.

Poulle:1969:AOBa

[1758] Emmanuel Poulle. Analyses d’ouvrages: A Science Museum illustrated booklet. Astronomy : globes, orreries and other models … par H. R.

Calvert. Revue d’Histoire des Sciences et de Leurs Applications, 22(2): 190, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904196.

Poulle:1969:AOBb

[1759] Emmanuel Poulle. Analyses d’ouvrages: Scientific Instruments in Art and History par Henri Michel; R. E. W. Maddison; Francis R. Maddison. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):191–192, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904197.

Anonymous:1969:OR

[1760] Anonyme. Ouvrages reçus. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):192, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904199.

Theodorides:1969:AOB

[1761] J. Théodoridès. Analyses d’ouvrages: Les grands Félins dans l’art de notre Préhistoire par Michel Rousseau. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):192, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904198.

Laissus:1969:PDR

[1762] Yves Laissus and Anne-Marie Monseigny. Les plantes du roi: Note sur un grand ouvrage de botanique préparé au XVII^{e} siècle par l’Académie royale des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):193–236, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903585.

Fischer:1969:HDR

[1763] Jean-Louis Fischer. Histoire de la découverte de la reproduction du Triton et de la Salamandre. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):237–256, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903586.

Rostand:1969:FAP

[1764] Jean Rostand. Félix-Archimède Pouchet et les méthodes contraceptives. Revue d’Histoire des Sciences et de Leurs Applications, 22(3): 257–258, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903587.

Itard:1969:DGC

[1765] Jean Itard. Dr Georges Collet (1878–1968). Revue d’Histoire des Sciences et de Leurs Applications, 22(3):259–260, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903588.

Delorme:1969:JL

[1766] Suzanne Delorme and Yves Laissus. Joseph Laissus (1900–1969). Revue d’Histoire des Sciences et de Leurs Applications, 22(3):261–263, juillet– septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903589.

Anonymous:1969:UIH

[1767] Anonyme. Union internationale d’histoire et de philosophie des sciences. Revue d’Histoire des Sciences et de Leurs Applications, 22(3): 264–265, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903590.

Anonymous:1969:Fb

[1768] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):265–266, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903591.

Delorme:1969:T

[1769] S. Delorme. Tchécoslovaquie. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):266–268, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903592.

Anonymous:1969:URX

[1770] Anonyme. U.R.S.S. XIII^{e} Congrès International d’Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):268, juillet– septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903593.

Hutin:1969:AOBa

[1771] Serge Hutin. Analyses d’ouvrages: Science in Utopia : a mighty design par Nell Eurich. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):269–270, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903594.

Hutin:1969:AOBb

[1772] Serge Hutin. Analyses d’ouvrages: Magistrats et sorciers en France au XVII^{e} siècle, (Civilisations et Mentalités) par Robert Mandrou. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):270–272, juillet– septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903595.

Grisard:1969:AOB

[1773] J. Grisard. Analyses d’ouvrages: Opera Omnia par Marini Ghetaldi. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):272– 273, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903596.

Beaude:1969:AOB

[1774] Joseph Beaude. Analyses d’ouvrages: William Chalmers (1596–1678). Étude bio-bibliographique avec des textes inédits, (Le mouvement des idées au XVII^{e} siècle) par Francis Ferrier. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):273–274, juillet–septembre 1969. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903597.

Rochot:1969:AOB

[1775] B. Rochot. Analyses d’ouvrages: Antoine Menjot, docteur en médecine, ami de Pascal, réformé au temps des persécutions. Études historiques et psychologiques. Hfl. 15 par Marguerite Scholtens; J. Mesnard. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):275–277, juillet– septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903598.

Itard:1969:AOBc

[1776] Jean Itard. Analyses d’ouvrages: Rechenpfennige, Aufsätze zur Wissenschaftsgeschichte. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):277–278, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903599.

Itard:1969:AOBd

[1777] Jean Itard. Analyses d’ouvrages: Chin Chang Suan Shu; Neun Bücher Arithmetischer Technik par Kurt Vogel. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):278–279, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903600.

Itard:1969:AOBe

[1778] Jean Itard. Analyses d’ouvrages: Historia de las ideas modernas en Matematica par Jose Babini. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):279–280, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903601.

Bourdier:1969:AOB

[1779] Franck Bourdier. Analyses d’ouvrages: Source Book in Geology 1900– 1950 par Kirtley E. Mather. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):280–282, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903602.

Guedes:1969:AOB

[1780] Michel Guédès. Analyses d’ouvrages: Taxonomic literature, a selective guide to botanical publications with dates, commentaries and types par F. A. Stafleu. Revue d’Histoire des Sciences et de Leurs Applications, 22(3): 282–284, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903603.

Jaquel:1969:VBC

[1781] Roger Jaquel. Vers les Œuvres complètes du savant et philosophe J.- H. Lambert (1728–1777): velléités et réalisations depuis deux siècles. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):285–302, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903997.

Dulieu:1969:FBS

[1782] Louis Dulieu. François Boissier de Sauvages (1706–1767). Revue d’Histoire des Sciences et de Leurs Applications, 22(4):303–322, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903998.

Guedes:1969:TMM

[1783] Michel Guédès. La théorie de la métamorphose en morphologie végétale : Des origines à Gœthe et Batsch. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):323–363, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903999.

Theodorides:1969:ICI

[1784] J. Théodoridès and K. Jaouiche. Italie: Colloque International sur la Philosophie et la Science Médiévales Rome–Florence, 9–15 avril 1969. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):365, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904000.

Bernhardt:1969:AOBb

[1785] Jean Bernhardt. Analyses d’ouvrages: La méthode expérimentale et la philosophie de la physique par Robert Blanché. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):366–368, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904001.

Martin:1969:AOBa

[1786] Roger Martin. Analyses d’ouvrages: Ecrits logiques (^{≪}Bibliothèque de Philosophie Contemporaine ^{≫}) par Jacques Herbrand; Jean van Heijenoort; Cl. Chevalley; A. Lautman. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):368–369, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904002.

Martin:1969:AOBb

[1787] Roger Martin. Analyses d’ouvrages: L’intelligence artificielle, (^{≪}La Science vivante ^{≫}) par Paul Braffort; J. Guéron. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):369, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904003.

Grmek:1969:AOB

[1788] M. D. Grmek. Analyses d’ouvrages: Kleine Weltgeschichte der Wissenschaft, 3^{e} éd. par Hans Joachim Störig. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):369–370, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904004.

Itard:1969:AOBf

[1789] Jean Itard. Analyses d’ouvrages: The Secrets of Ancient Geometry and its use, t. I, t. II par Tons Brunés. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):370, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904005.

Itard:1969:AOBg

[1790] Jean Itard. Analyses d’ouvrages: La Géométrie non euclidienne dans l’Antiquité par Evert M. Bruins. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):370–371, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23904006.

Itard:1969:AOBh

[1791] Jean Itard. Analyses d’ouvrages: The Concept of Number par Christoph J. Scriba. Revue d’Histoire des Sciences et de Leurs Applications, 22(4): 371, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904007.

Itard:1969:AOBi

[1792] Jean Itard. Analyses d’ouvrages: The translation of the Elements of Euclid from the Arabic into Latin by Hermann of Carinthia (?) par H. L. L. Busard. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):371–372, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904008.

Itard:1969:AOBj

[1793] Jean Itard. Analyses d’ouvrages: Het rekenen met Breuken in de Middeleeuwen, in het bijzonder bij Johannes de Lineriis par H. Busard. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):372–373, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904009.

Itard:1969:AOBk

[1794] Jean Itard. Analyses d’ouvrages: Quelques sujets de l’Histoire des Mathématiques au Moyen Âge par H.-L.-L. Busard. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):373–374, October 1969. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904010.

Itard:1969:AOBl

[1795] Jean Itard. Analyses d’ouvrages: Ein Byzantinisches Rechenbuch des frühen 14. Jahrhunderts. Text, Übersetzung und Kommentar par Kurt Vogel. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):374– 375, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904011.

Colnort:1969:AQPa

[1796] S. Colnort. Analyses de quelques périodiques: Archives internationales d’Histoire des Sciences, t. XX, 1967, n^{os} 78–79, 80 et 81. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):375, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904012.

Colnort:1969:AQPb

[1797] S. Colnort. Analyses de quelques périodiques: The British Journal for the History of Science, vol. II, 1964–1965, fascicules 5 à 8 inclus. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):375–376, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904013.

Colnort:1969:AQPc

[1798] S. Colnort. Analyses de quelques périodiques: Chymia, tomes 10, 1965; 11, 1966; 12, 1967. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):376, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904014.

Anonymous:1969:TGM

[1799] Anonyme. Table générale des matières du Tome XXII. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):377–384, October 1969. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904015.

Anonymous:1969:BMa

[1800] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):??, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903779.

Anonymous:1969:BMb

[1801] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):??, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904200.

Anonymous:1969:BMc 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903604.

Anonymous:1969:BMd

[1803] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):??, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904016.

Anonymous:1969:FMa

[1804] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 22(1):??, January 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903760.

Anonymous:1969:FMb

[1805] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 22(2):??, avril–juin 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904173.

Anonymous:1969:FMc

[1806] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 22(3):??, juillet–septembre 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903584.

Anonymous:1969:FMd

[1807] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 22(4):??, October 1969. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903996.

Canguilhem:1970:JEO

[1808] Georges Canguilhem. Journées d’études organisées par l’Institut d’Histoire des Sciences de l’Université de Paris les 30 et 31 mai 1969 pour le bicentenaire de la naissance de G. Cuvier. Revue d’Histoire des Sciences et de Leurs Applications, 23(1):6–8, January 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903487.

Courtes:1970:GCO

[1809] Francis Courtès. Georges Cuvier ou l’origine de la négation. Revue d’Histoire des Sciences et de Leurs Applications, 23(1):9–27, January 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903488.

Canguilhem:1970:D

[1810] Georges Canguilhem, François Dagognet, M. D. Grmek, and Bernard Balan. Discussion. Revue d’Histoire des Sciences et de Leurs Applications, 23(1):28–34, January 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903489.

Limoges:1970:LNP

[1811] Camille Limoges. L’économie naturelle et le principe de corrélation chez Cuvier et Darwin. Revue d’Histoire des Sciences et de Leurs Applications, 23(1):35–45, January 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903490.

Leroy:1970:D

[1812] J.-Fr. Leroy and M. D. Grmek. Discussion. Revue d’Histoire des Sciences et de Leurs Applications, 23(1):46–48, January 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903491.

Dagognet:1970:SCH

[1813] François Dagognet. La situation de Cuvier dans l’histoire de la biologie. Revue d’Histoire des Sciences et de Leurs Applications, 23(1):49–60, January 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903492.

Piveteau:1970:Da

[1814] Pierre Piveteau and Michel Foucault. Discussion. Revue d’Histoire des Sciences et de Leurs Applications, 23(1):61–62, January 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903493.

Foucault:1970:SCH

[1815] Michel Foucault. La situation de Cuvier dans l’histoire de la biologie. Revue d’Histoire des Sciences et de Leurs Applications, 23(1):63–69, January 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903494.

Piveteau:1970:Db

[1816] Pierre Piveteau, François Dagognet, J.-Fr. Leroy, Yvette Conry, Bernard Balan, Georges Canguilhem, Suzanne Delorme, Camille Limoges, ?. Saint-Sernin, M. D. Grmek, and ?. Salomon. Discussion. Revue d’Histoire des Sciences et de Leurs Applications, 23(1):70–92, January 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903495.

Aujac:1970:SOM

[1817] Germaine Aujac. La sphéropée, ou la mécanique au service de la découverte du monde. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):93–107, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903853.

Rashed:1970:MST

[1818] Roshdi Rashed. Le modèle de la sphère transparente et l’explication de l’arc-en-ciel: Ibn al-Haytham, al-Fārisī. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):109–140, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903854.

Broc:1970:APC

[1819] Numa Broc. Une affaire de plagiat cartographique sous Louis XIV : le procès Delisle-Nolin. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):141–153, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903855.

Huard:1970:JS

[1820] Pierre Huard and M.-J. Imbault-Huart. Jules Soury (1842–1915). Revue d’Histoire des Sciences et de Leurs Applications, 23(2):155–164, avril– juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903856.

Miloschev:1970:CFA

[1821] Boris Josifov Miloschev. La contribution de la France au développement de l’instruction supérieure en Bulgarie au cours de la seconde moitié du XIX^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):165–169, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903857.

Anonymous:1970:ADS

[1822] Anonyme. Afrique du Sud: Symposium sur l’Histoire de la Chimie Johannesburg, 13 et 14 avril 1970. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):170, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903858.

Anonymous:1970:FSH

[1823] Anonyme. France: Séminaire d’Histoire des Mathématiques Conférences 1970. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):170, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903859.

Virieux-Reymond:1970:SCC

[1824] Antoinette Virieux-Reymond. Suisse: Cent cinquantenaire de la Société Vaudoise des Sciences Naturelles. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):171, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903860.

Anonymous:1970:URXa

[1825] Anonyme. U.R.S.S. XIII^{e} Congrès International d’Histoire des Sciences Moscou, 18–24 août 1971. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):171–172, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903861.

Taton:1970:AOBa

[1826] R. Taton. Analyses d’ouvrages: Index biographique des membres et correspondants de l’Académie des Sciences, du 22 décembre 1666 au 15 décembre 1967 Institut de France. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):173–174, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903862.

Poulle:1970:AOBa

[1827] Emmanuel Poulle. Analyses d’ouvrages: New Aspects in the History and Philosophy of Astronomy. First Joint Symposium of the International Astronomical Union and the Union internationale d’Histoire et de Philosophie des Sciences, held at the University of Hamburg, (Vistas in Astronomy, vol. 9) par Arthur Beer. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):174–175, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903863.

Taton:1970:AOBb

[1828] René Taton. Analyses d’ouvrages: Principes mathématiques de la philosophie naturelle, trad. de la marquise du Chastellet augmentée des Commentaires de Clairaut par Isaac Newton. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):175–180, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903864.

Colnort:1970:AOBa

[1829] Suzanne Colnort. Analyses d’ouvrages: A History of Chemistry, vol. III, IV par J. R. Partington. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):180–181, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903865.

Theodorides:1970:AOB

[1830] Jean Théodoridès. Analyses d’ouvrages: Perspectives de la zoologie européenne par Jean Leclercq; Pierre Dagnelie. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):181–182, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903866.

Fischer:1970:AOB

[1831] Jean-Louis Fischer. Analyses d’ouvrages: La vie et l’œuvre d’Étienne Geoffroy Saint-Hilaire par Théophile Cahn. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):182–183, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903867.

Wrotrowska:1970:AOB

[1832] Denise Wrotnowska. Analyses d’ouvrages: Pasteur, extraits de ses œuvres par René Dujarric de la Rivière. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):183–184, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903868.

Grmek:1970:AOB

[1833] M. D. Grmek. Analyses d’ouvrages: Medical history of Malta par Paul Cassar. Revue d’Histoire des Sciences et de Leurs Applications, 23(2): 184, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23903869.

Huard:1970:AOBa

[1834] P. Huard. Analyses d’ouvrages: Les officiers de santé de la marine française de 1814 à 1835 par J. Léonard. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):184–185, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903870.

Colnort:1970:AOBb

[1835] S. Colnort. Analyses d’ouvrages: Comptes rendus du XCI^{e} Congrès National des Sociétés Savantes, Rennes, 1966, Section des Sciences, Y. I, Histoire des Sciences, ^{≪}Ministère de l’Education nationale. Comité des Travaux scientifiques et historiques ^{≫}. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):185–186, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903871.

Colnort:1970:AOBc

[1836] S. Colnort. Analyses d’ouvrages: Annales de démographie historique, 1965 (Études et chronique). Revue d’Histoire des Sciences et de Leurs Applications, 23(2):186–187, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903872.

Colnort:1970:AQP

[1837] S. Colnort. Analyses de quelques périodiques: Gesnerus, vol. 23, 1966, fasc. 1–2, 3–4; vol. 24, 1967, fasc. 1–2, 3–4; vol. 25, 1968, fasc. 1–2, 3–4. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):187–188, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903873.

Louis:1970:DAE

[1838] Pierre Louis. La domestication des animaux à l’époque d’Aristote. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):189–201, juillet– septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903658.

Chabbert:1970:JBM

[1839] Pierre Chabbert and M. Dissés. Jacques Borelly (16??–1689) Membre de l’Académie royale des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):203–227, juillet–septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23903659.

Salomon-Bayet:1970:PTA

[1840] Claire Salomon-Bayet. Un préambule théorique à une Académie des Arts: Académie royale des Sciences, 1693–1696. Présentation et textes. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):229–250, juillet– septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903660.

Otero:1970:DIC

[1841] Mario H. Otero. Les définitions implicites chez Gergonne. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):251–255, juillet– septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903661.

Taton:1970:HRS

[1842] René Taton. Sur l’histoire des relations scientifiques franco-russes. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):257–264, juillet– septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903662.

D:1970:TCG

[1843] S. D. Tchécoslovaquie: Colloque Gregor Mendel Brno, 28 juin–3 juillet 1970. Revue d’Histoire des Sciences et de Leurs Applications, 23(3): 265–266, juillet–septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903663.

Itard:1970:AOB

[1844] Jean Itard. Analyses d’ouvrages: A short History of Greek Mathematics par James Gow. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):267, juillet–septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903664.

Mugler:1970:AOB

[1845] Ch. Mugler. Analyses d’ouvrages: Strabon et la science de son temps, ≪Les sciences du monde ^{≫}par Germaine Aujac. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):267–268, juillet–septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903665.

Poulle:1970:AOBb

[1846] Emmanuel Poulle. Analyses d’ouvrages: Catalogue des manuscrits en écriture latine portant des indications de date, de lieu ou de copiste, T.

VI : Bourgogne, Centre, Sud-Est et Sud-Ouest de la France par Charles Samaran; Robert Marichal; Monique Garand; Madeleine Mabille; Josette Metman. Revue d’Histoire des Sciences et de Leurs Applications, 23(3): 268–269, juillet–septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903666.

Hugonnard-Roche:1970:AOB

[1847] Henri Hugonnard-Roche. Analyses d’ouvrages: Nicole Oresme and the Medieval Geometry of Qualities and Motions. A treatise on the uniformity and difformity of intensities known as ????, (The University of Wisconsin Publications in Medieval Science, n^{o} 12) par Marshall Clagett. Revue d’Histoire des Sciences et de Leurs Applications, 23(3): 269–272, juillet–septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903667.

Huard:1970:AOBb

[1848] P. Huard. Analyses d’ouvrages: Problemata Varia Anatomica (P.V.A.): ms. 1165, The University of Bologna, Lawrence (U.S.A.) par L. R. Lind. Revue d’Histoire des Sciences et de Leurs Applications, 23(3): 272–273, juillet–septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903668.

Bernhardt:1970:AOB

[1849] Jean Bernhardt. Analyses d’ouvrages: Œuvres philosophiques, avec une étude biobibliographique, (Le mouvement des idées au XVII^{e} siècle) par Gerauld de Cordemoy; Pierre Clair; François Girbal. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):273–276, juillet–septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23903669.

Dulieu:1970:AOB

[1850] Louis Dulieu. Analyses d’ouvrages: L’Encyclopédie dans le Bas-Languedoc au XVIII^{e} siècle par Jacques Proust. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):276–277, juillet–septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903670.

Speziali:1970:AOBa

[1851] Pierre Speziali. Analyses d’ouvrages: Le débat sur les écritures et l’hiéroglyphe aux XVII^{e} et XVIII^{e} siècles et l’application de la notion de déchiffrement aux écritures mortes par Madeleine V.-David. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):277–280, juillet– septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23903671.

Speziali:1970:AOBb

[1852] Pierre Speziali. Analyses d’ouvrages: Lomonossov, 1711–1765. Sa vie, son œuvre par Luce Langevin. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):280–283, juillet–septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23903672.

Laissus:1970:AOB

[1853] Yves Laissus. Analyses d’ouvrages: Raspail, scientist and reformer par Dora B. Weiner. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):283, juillet–septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903673.

Poulle:1970:AOBc

[1854] Emmanuel Poulle. Analyses d’ouvrages: Astronomy, Gnomonics; a Catalogue of Instruments of the 15th to the 19th centuries in the Collections of the National Technical Museum, Prague (Narodni technické Muzeum, Praha) par Zdenek Horský; Otilie Skopová. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):284, juillet–septembre 1970. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23911709.

Anonymous:1970:HOR

[1855] Anonyme. Henry Oldenburg et les relations scientifiques au XVII^{e} siècle. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):285–304, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904055.

Mayer:1970:PCG

[1856] Jean Mayer. Portrait d’un chimiste : Guillaume-François Rouelle (1703– 1770). Revue d’Histoire des Sciences et de Leurs Applications, 23(4):305– 332, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904056.

Dugac:1970:CMN

[1857] Pierre Dugac. Charles Méray (1835–1911) et la notion de limite. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):333–350, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904057.

Bernhardt:1970:GNM

[1858] Jean Bernhardt. Galilée et la naissance de la mécanique classique selon Maurice Clavelin. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):351–364, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904058.

Anonymous:1970:F

[1859] Anonyme. France. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):365–367, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904059.

Anonymous:1970:GBX

[1860] Anonyme. Grande-Bretagne: XXIII^{e} Congrès International d’Histoire de la Médecine, Londres, 2–09 septembre 1972. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):367, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23911710.

Anonymous:1970:RIC

[1861] Anonyme. Roumanie: IV^{e} Congrès International de Logique, Méthodologie et Philosophie de la Science Bucarest, 29 août–4 septembre 1971. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):367– 368, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23911711.

Anonymous:1970:URXb

[1862] Anonyme. U.R.S.S. XIII^{e} Congrès International d’Histoire des Sciences Moscou, 18–24 août 1971. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):368, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23911712.

Dadic:1970:AOB

[1863] Zarko Dadić. Analyses d’ouvrages: Greek Mathematical Thought and the Origin of Algebra, With an Appendix containing Vieta’s Introduction to the Analytical Art, Cambridge (Mass.) par Jacob Klein. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):369–370, October 1970. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904060.

Taton:1970:AOBc

[1864] R. Taton. Analyses d’ouvrages: Zählen und Rechnen einst und jetzt, übersetzung aus dem Polnischen par Wlodzimierz Krysicki. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):370, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904061.

Speziali:1970:AOBc

[1865] Pierre Speziali. Analyses d’ouvrages: Le temps et la pensée physique contemporaine, (^{≪}Réalités de la science ^{≫}) par J.-L. Rigal. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):370–372, October 1970. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904062.

Donghi:1970:AOBa

[1866] Leticia Halperin Donghi. Analyses d’ouvrages: The Atomic Debates. Brodie and the rejection of the Atomic Theory. Three Studies par W. H. Brock. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):372– 374, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23904063.

Donghi:1970:AOBb

[1867] Leticia Halperin Donghi. Analyses d’ouvrages: Nobel Foundation, Nobel Lectures, including presentation speeches and laureates’ biographies, Chemistry, 1901–1921. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):374, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904064.

Colnort:1970:AOBd

[1868] S. Colnort. Analyses d’ouvrages: Die zoologische Buchillustration. Ihre Bibliographie und Geschichte par Claude Nissen. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):374–375, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23904065.

Huard:1970:AOBc

[1869] P. Huard. Analyses d’ouvrages: L’admission du malade mental dans les établissements de soins, de 1789 à 1838 par Alain Bigorre. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):375, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904066.

Anonymous:1970:PR

[1870] Anonyme. Publications reçues. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):376, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904067.

Anonymous:1970:TGM

[1871] Anonyme. Table générale des matières du Tome XXIII. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):377–384, October 1970. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23904068.

Anonymous:1970:BMa

[1872] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 23(1):??, January 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903496.

Anonymous:1970:BMb

[1873] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):??, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903874.

Anonymous:1970:BMc

[1874] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):??, juillet–septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903674.

Anonymous:1970:BMd

[1875] Anonyme. Back matter. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):??, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904069.

Anonymous:1970:FMa

[1876] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 23(1):??, January 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903486.

Anonymous:1970:FMb

[1877] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 23(2):??, avril–juin 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23903852.

Anonymous:1970:FMc

[1878] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 23(3):??, juillet–septembre 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23903657.

Anonymous:1970:FMd

[1879] Anonyme. Front matter. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):??, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23904054.

Daumas:1971:LTP

[1880] Maurice Daumas. L’acquisition des techniques par les pays non initiateurs. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):5–11, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/24468283.

Tresse:1971:JID

[1881] René Tresse. La jeunesse et l’initiation du mécanicien Claude-Pierre Molard de 1759 à 1791. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):13–24, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

24468284.

Christophe:1971:LVC

[1882] R. Christophe. L’analyse volumétrique de 1790 à 1860: Caractéristiques et importance industrielle. Évolution des instruments. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):25–44, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/24468285.

Payen:1971:DND

[1883] Jacques Payen. Deux nouveaux document sur Nicolas Clément. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):45–60, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/24468286.

Faddeev:1971:HUC

[1884] B. V. Faddeev. Historique de l’utilisation des convoyeurs à bande dans les travaux miniers. Revue d’Histoire des Sciences et de Leurs Applications, 49(1):61–66, 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/24468287.

Soulard:1971:DTE

[1885] Robert Soulard. Sur les débuts de la transmission électrique des images en couleur. Revue d’Histoire des Sciences et de Leurs Applications, 24 (1):67–73, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/24468288.

Foiret:1971:LDD

[1886] Jacques Foiret. Louis Ducos du Hauron. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):74–78, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/24468289.

Daumas:1971:BRN

[1887] Maurice Daumas. Book review: Nouvelle histoire des techniques by Carl von Klinckowstroem. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):79–80, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

24468290.

Daumas:1971:BRT

[1888] Maurice Daumas. Book review: Technology in Western Civilization by Melvin Kranzberg, Carroll W. Pursell Jr. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):80–82, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/24468291.

Melderen:1971:BRH

[1889] Joseph-Adolphe Melderen. Book review: Histoire générale des techniques Tome I : Les origines de la civilisation technique, 1962, Tome II : Les premières étapes du machinisme, 1965, Tome III : L’expansion du machinisme, 1968 by Maurice Daumas. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):82–85, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/24468292.

Payen:1971:BRG

[1890] J. Payen. Book review: Geschichte der Naturwissenschaften und der Technik im 19. Jahrhundert., 6. Gesprach der Georg-Agricola-Gesellschaft (Technikgeschichte in Einzeldarstellungen, n^{∘} 16). Revue d’Histoire des Sciences et de Leurs Applications, 24(1):85–86, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/24468293.

Payen:1971:BRR

[1891] Jacques Payen. Book review: Reflections on the Motive Power of Fire by Sadi Carnot, and other Papers on the Second Law of Thermodynamics by E. Clapeyron, R. Clausius, E. Mendoza. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):86–87, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/24468294.

Daumas:1971:BRC

[1892] Maurice Daumas. Book review: Le cheval dans l’Antiquité gréco-romaine (Des Guerres médiques aux grandes Invasions), t. I, t. II by Paul Vigneron. Revue d’Histoire des Sciences et de Leurs Applications, 24(1): 87–89, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/24468295.

Payen:1971:BRM

[1893] Jacques Payen. Book review: Mechanical Universe: The Astrarium of Giovanni De’ Dondi, Transactions of the American Philosophical Society, LVI, 1966, fasc. 5 (tiré à part) by Silvio A. Bedini, Francis R. Maddison. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):89–91, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/24468296.

Payen:1971:BRE

[1894] J. Payen. Book review: European clocks by E.-J. Tyler. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):91, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/24468297.

Payen:1971:BRC

[1895] J. Payen. Book review: Clément Ader. His Flight-Claims and his Place in History by Charles Harvard Gibbs-Smith. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):91–92, ???? 1971. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/24468298.

Payen:1971:BRB

[1896] Jacques Payen. Book review: Le bateau de Berck by François Beaudoin. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):92– 93, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/24468299.

Payen:1971:BRT

[1897] J. Payen. Book review: Technikatörténeti Szemle, revue hongroise d’histoire des techniques. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):93–95, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

24468300.

Anonyme:1971:AO

[1898] Anonyme. Analyses d’ouvrages. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):96, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

Anonyme:1971:Ia

[1899] Anonyme. Informations. Revue d’Histoire des Sciences et de Leurs Applications, 24(1):96, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

deBroglie:1971:JPP

[1900] Louis de Broglie. Jean Perrin physicien. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):99–105, 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/24468260.

Sadron:1971:JPC

[1901] Charles Sadron. Jean Perrin et la chimie physique. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):107–116, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/24468262.

Couderc:1971:JPA

[1902] Paul Couderc. Jean Perrin et l’astronomie. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):117–122, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/24468264.

Taton:1971:RSA

[1903] René Taton. Sur les relations scientifiques d’Augustin Cauchy et d’Évariste Galois. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):123–148, 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/24468266.

Dulieu:1971:PJB

[1904] Louis Dulieu. Paul-Joseph Barthez. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):149–176, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/24468268.

Anonyme:1971:Ib

[1905] Anonyme. Informations. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):177–180, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

24468270.

Paul:1971:BRE

[1906] H. W. Paul. Book review: Essai sur l’histoire humaine de la nature, (Nouvelle bibliothèque scientifique) by Serge Moscovici. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):181, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/24468271.

Lejeune:1971:BRH

[1907] Albert Lejeune. Book review: Historia de la Ciencia: 6. Ciencia alejandrina; 7. Ciencia helenistica; 8. Ciencia en el periodo greco-romano; 9. Ciencia en la temprana Edad Media; 10. Ciencia en la alta Edad Media;

11. Ciencia en la baja Edad Media, (Biblioteca de Historia de la Ciencia, 1006 à 1011) by José Babini. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):181–182, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/24468272.

Lejeune:1971:BRB

[1908] Albert Lejeune. Book review: Book of Minerals by Albertus Magnus, Dorothy Wyckoff. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):182–183, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

24468273.

Fischer:1971:BRG

[1909] Jean-Louis Fischer. Book review: Un grand médecin et biologiste, Casimir-Joseph Davaine (1812–1882), (Analecta historica) by Jean Théodoridès, Jean Rostand. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):183–185, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/24468274.

Paul:1971:BRG

[1910] H. W. Paul. Book review: Government in Science. The U.S. Geological Survey, 1867–1894 by Thomas G. Manning. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):185, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/24468275.

Colnort:1971:BRCa

[1911] S. Colnort. Book review: Les Conférences du Palais de la Découverte, série D: Histoire des Sciences. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):185–186, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/24468276.

Colnort:1971:BRA

[1912] S. Colnort. Book review: Ambix. The Journal of the Society for the Study of Alchemy and early Chemistry, vol. VIII–XI, 1961/1963 by D. Geoghegan. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):186– 187, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/24468277.

Colnort:1971:BRHa

[1913] S. Colnort. Book review: History of Science. An annual review of Literature, Research and Teaching, vol. 4, 1965, vol. 5, 1966, vol. 6, 1967, vol. 7, 1968, vol. 8, 1969 by A. C. Crombie, M. A. Hoskin. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):187–188, ???? 1971. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/24468278.

Colnort:1971:BRI

[1914] S. Colnort. Book review: Isis, an International Review devoted to the History of Science and its cultural influences. vol. 56, 1965, nos 183– 186; vol. 57, 1966, nos 187–190; vol. 58, 1967, nos 191–195 (195 est un fascicule supplémentaire); vol. 59, 1968, nos 196–200 (200 est un fascicule supplémentaire); vol. 60, 1969, nos 201–205; vol. 61, nos 206– 208. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):188– 192, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/24468279.

Colnort:1971:BRCb

[1915] S. Colnort. Book review: Chronique de l’Université de Liège. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):192, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/24468280.

Sigal:1971:CCT

[1916] Pierre André Sigal. Comment on conçevait et on traitait la paralysie en Occident dans le haut Moyen Âge. Revue d’Histoire des Sciences et de Leurs Applications, 24(3):193–211, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/24468230.

Lavagne:1971:EBP

[1917] François-G. Lavagne. Étalons bisontins de poids et de mesure. Revue d’Histoire des Sciences et de Leurs Applications, 24(3):213–232, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/24468231.

Beaude:1971:LIP

[1918] Joseph Beaude. Lettre inédite de Picot à Carcavi relative à l’expérience barométrique (5 août 1649). Revue d’Histoire des Sciences et de Leurs Applications, 24(3):233–246, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/24468232.

Sinaceur:1971:AII

[1919] Mohammed A. Sinaceur. Appartenance et inclusion. Un inédit de Richard Dedekind. (French) [Class membership and class inclusion. An unpublished paper of Richard Dedekind]. Revue d’Histoire des Sciences et de Leurs Applications, 24(3):247–254, 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/24468234.

Russo:1971:NFD

[1920] François Russo. Nécrologie. François de Dainville (1909–1970). Revue d’Histoire des Sciences et de Leurs Applications, 24(3):255–256, ????

1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/24468235.

Anonyme:1971:Ic

[1921] Anonyme. Informations. Revue d’Histoire des Sciences et de Leurs Applications, 24(3):257–258, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

24468236.

Poulle:1971:BRE

[1922] Emmanuel Poulle. Book review: Eléments de bibliographie de l’histoire des sciences et des techniques, 2e éd. by Fr. Russo. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):260–261, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.

jstor.org/stable/24468238.

Cazenave:1971:BRL

[1923] M. Cazenave. Book review: L’Académie des Sciences et les Académiciens de 1666 à 1793 by Joseph Bertrand. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):261, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.jstor.org/

stable/24468239.

Paul:1971:BRS

[1924] H. W. Paul. Book review: Scientific Thought, 1900–1960, a selective Survey by R. Harré. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):261–262, ???? 1971. CODEN RHSAAM. ISSN 0048-7996,

0151-4105, 1969-6582. URL http://www.jstor.org/stable/24468240.

Grisard:1971:BRH

[1925] J. Grisard. Book review: A History of Mathematics by Carl B. Boyer. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):262–263, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582.

URL http://www.jstor.org/stable/24468241.

Itard:1971:BRAa

[1926] Jean Itard. Book review: Anfänge der Griechischen Mathematik by Árpád Szabó. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):263–266, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-

4105, 1969-6582. URL http://www.jstor.org/stable/24468242.

Itard:1971:BRAb

[1927] Jean Itard. Book review: Archimède, t. I er, De la sphère et du cylindre, La mesure du cercle, Sur les conoïdes et les sphéroïdes (Collection des Universités de France) by Charles Mugler. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):266–267, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.jstor.org/

stable/24468243.

Bos:1971:BRW

[1928] H. J. M. Bos. Book review: Die Werke von -, Band I by Jakob Bernoulli, Naturforschenden Gesellschaft in Basel. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):267–269, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.jstor.org/

stable/24468244.

Itard:1971:BRM

[1929] Jean Itard. Book review: The Mathematical Papers of Isaac Newton, vol. III, 1670–1673 by D. T. Whiteside, M. A. Hoskin, A. Prag. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):270–272, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL

http://www.jstor.org/stable/24468245.

Sadoun-Goupil:1971:BRD

[1930] Michelle Sadoun-Goupil. Book review: Les doctrines chimiques en France, du début du XVII e siècle à la fin du XVIII e siècle by Hélène Metzger. Revue d’Histoire des Sciences et de Leurs Applications, 24(2): 272–273, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105,

1969-6582. URL http://www.jstor.org/stable/24468246.

Avramesco:1971:BRE

[1931] A. Avramesco. Book review: Ernst Mach physicist and philosopher, <<Boston Studies in the Philosophy of Science >>, VI, <<Synthese Library >> by Robert S. Cohen, Raymond J. Seeger. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):273–275, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.

jstor.org/stable/24468247.

Taton:1971:BRN

[1932] René Taton. Book review: Les nombres premiers (coll. <<Que saisje? >>, no 571) by Jean Itard. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):275–276, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.jstor.org/

stable/24468248.

Lejeune:1971:BRA

[1933] Albert Lejeune. Book review: Astronomical Literature in the Ernst Zinner Collection San Diego State College Library by Gerald Johns. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):276, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http:

//www.jstor.org/stable/24468249.

Poulle:1971:BRS

[1934] Emmanuel Poulle. Book review: Star performance by Harriet Pratt Lattin. Revue d’Histoire des Sciences et de Leurs Applications, 24(2): 276–277, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105,

1969-6582. URL http://www.jstor.org/stable/24468250.

Poulle:1971:BRK

[1935] Emmanuel Poulle. Book review: Kunstuhrmacher in Alt-Augsburg, (Abhandlungen zur Geschichte des Stadt Augsburg. Schriftenreihe des Stadtarchivs Augsburg, 18) by Maximilian Bobinger, H. Rösler. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):277–278, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL

http://www.jstor.org/stable/24468251.

Levy:1971:BRH

[1936] Jacques R. Lévy. Book review: Herschel at the Cape. Diaries and Correspondence of -, 1834–1838 by John Herschel, Davis S. Evans, Terence J. Doeming, Betty Hall Evans, Stephen Goldfarb. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):278–279, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.

jstor.org/stable/24468252.

Theodorides:1971:BRM

[1937] Jean Théodoridès. Book review: La médecine précolombienne by Charles Coury. Revue d’Histoire des Sciences et de Leurs Applications, 24(2): 279–280, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105,

1969-6582. URL http://www.jstor.org/stable/24468253.

Huard:1971:BRLa

[1938] P. Huard. Book review: L’Hôtel–Dieu de Paris. Treize siècles de soins, d’enseignement et de recherches by Charles Coury. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):280–282, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.

jstor.org/stable/24468254.

Huard:1971:BRLb

[1939] P. Huard. Book review: L’enseignement clinique dans les hôpitaux de Paris entre 1794 et 1848, thèse de la Faculté de Médecine de Paris by Mireille Wiriot. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):282–283, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-

4105, 1969-6582. URL http://www.jstor.org/stable/24468255.

Koning:1971:BRM

[1940] D. A. Wittop Koning. Book review: Medical Ceramics. A Catalogue of the English and Dutch Collections in the Museum of the Wellcome Institute of the History of Medicine by J. K. Crellin. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):284, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.

jstor.org/stable/24468256.

Huard:1971:BRC

[1941] P. Huard. Book review: Chambéry, le Magistrat de santé, la peste (1564– 1724), Diplôme d’Etudes supérieures, dactylographie by Nicolas Greslou. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):284–285, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582.

URL http://www.jstor.org/stable/24468257.

Taton:1971:BRS

[1942] René Taton. Book review: Some Aspects of Seventeenth-Century Medicine and Science, Papers Read at a Clark Library Seminar October 12, 1968 by Ladislao Reti, William C. Gibson. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):285–286, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.

jstor.org/stable/24468259.

C:1971:BRC

[1943] S. C. Book review: Comptes rendus du XCIV e Congrès national des Sociétés savantes, Pau, 1969, Section des Sciences, t. I : Histoire des Sciences, <<Ministère de l’Education nationale. Comité des Travaux Scientifiques et Historiques. >>. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):286, ???? 1971. CODEN RHSAAM. ISSN 0048-7996,

0151-4105, 1969-6582. URL http://www.jstor.org/stable/24468261.

Colnort:1971:BRHb

[1944] S. Colnort. Book review: Historical Studies in the Physical Sciences, vol. I by Russell McCormmach. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):286–287, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.jstor.org/

stable/24468263.

Colnort:1971:BRG

[1945] S. Colnort. Book review: Gesnerus, revue trimestrielle, Register der Jahrgänge 1 bis 25, 1943–1944–1968, fascicule indépendant. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):287, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http:

//www.jstor.org/stable/24468265.

Colnort:1971:BRB

[1946] S. Colnort. Book review: The British Journal for the History of Science, vol. III, 1966–1967, fasc. 9–12; vol. IV, 1968–1969, fasc. 13–16 by M. P. Crosland. Revue d’Histoire des Sciences et de Leurs Applications, 24(2): 287–288, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105,

1969-6582. URL http://www.jstor.org/stable/24468267.

Florkin:1971:FVB

[1947] Marcel Florkin. Des forces-de-vie à la bioénergétique. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):289–298, ???? 1971. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/24467724.

Rose:1971:PMA

[1948] Paul Lawrence Rose. Plusieurs manuscrits autographes de Federico Commandino à la Bibliothèque Nationale de Paris. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):299–307, 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/24467725.

Bernhardt:1971:LPP

[1949] Jean Bernhardt. Une lettre-programme pour ^{≪}l’avancement des mathématiques ^{≫}au XVIIe siècle: L’^{≪}Idée générale des mathématiques ≫de John Pell. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):309–316, 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/24467726.

Greenbaum:1971:CTH

[1950] Louis S. Greenbaum. The commercial treaty of humanity. La tournée des hôpitaux anglais par Jacques Tenon en 1787. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):317–350, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/24467727.

Lemercier:1971:LPC

[1951] Léa Lemercier. Une lettre peu connue de Chaptal. Réponse à une lettre inédite de René. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):351–363, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

24467728.

Rosinska:1971:IFP

[1952] G. Rosińska and Jean Théodoridès. Informations: France, Pologne, Suisse. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):365– 368, ???? 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/24467729.

Huard:1971:BRI

[1953] P. Huard. Book review: Introduction à l’Histoire des Sciences, t. I by Georges Canguilhem. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):369, ???? 1971. CODEN RHSAAM. ISSN 0048-7996,

0151-4105, 1969-6582. URL http://www.jstor.org/stable/24467730.

Lerner:1971:BRD

[1954] M.-P. Lerner. Book review: Discours concernant deux sciences nouvelles, Coll. <<Philosophies pour l’âge de la science >> by Galilée, M. Clavelin. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):369–374, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582.

URL http://www.jstor.org/stable/24467731.

Itard:1971:BRC

[1955] Jean Itard. Book review: Correspondance du P. Marin Mersenne, religieux minime, T. XI : 1642 by Cornelis de Waard, Mme Paul Tannery, M. Bernard Rochot. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):374–377, ???? 1971. CODEN RHSAAM. ISSN 0048-7996,

0151-4105, 1969-6582. URL http://www.jstor.org/stable/24467732.

Grisard:1971:BRO

[1956] J. Grisard. Book review: Opera mathematica by François Viète, Joseph E. Hofmann. Revue d’Histoire des Sciences et de Leurs Applications, 24 (4):377–378, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105,

1969-6582. URL http://www.jstor.org/stable/24467733.

Yamazaki:1971:BRJ

[1957] Eizo Yamazaki. Book review: Jean d’Alembert, Science and the Enlightenment by Thomas L. Hankins. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):378–380, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.jstor.org/

stable/24467734.

Itard:1971:BRR

[1958] Jean Itard. Book review: Réflexions sur la métaphysique du calcul infinitésimal by Lazare Carnot, Marcel Mayot. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):380–381, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.

jstor.org/stable/24467735.

Itard:1971:BRV

[1959] Jean Itard. Book review: La vie et les travaux du baron Cauchy by C. A. Valson, René Taton. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):381–382, ???? 1971. CODEN RHSAAM. ISSN 0048-7996,

0151-4105, 1969-6582. URL http://www.jstor.org/stable/24467736.

Giuculescu:1971:BRQ

[1960] Alexandru Giuculescu. Book review: Quelques aspects de la pensée d’un mathématicien by Paul Lévy. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):382–384, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.jstor.org/

stable/24467737.

Anonymous:1971:PR

[1961] Anonymous. Publications reçues. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):384, ???? 1971. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582. URL http://www.jstor.org/

stable/24467738.

Anonyme:1971:TGM

[1962] Anonyme. Table génèrale des matières du tome XXIV. Revue d’Histoire des Sciences et de Leurs Applications, 24(4):385–392, ???? 1971. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

Dobrovolsky:1972:HCP

[1963] Viatceslav Alexevitch Dobrovolsky. Sur l’histoire de la classification des points singuliers des équations différentielles. Revue d’Histoire des Sciences, 25(1):3–11, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631712.

Rosmorduc:1972:ESA

[1964] Jean Rosmorduc. Une erreur scientifique au début du siècle : ^{≪}les rayons N ^{≫}. Revue d’Histoire des Sciences, 25(1):13–25, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23631713.

Faure-Fremiet:1972:ECS

[1965] E. Fauré-Frémiet and Jean Théodoridès. État des connaissances sur la structure des protozoaires avant la formulation de la Théorie cellulaire. Revue d’Histoire des Sciences, 25(1):27–44, January/March 1972. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631714.

Imbault-Huart:1972:SHM

[1966] M. J. Imbault-Huart. Sources de l’histoire de la médecine aux archives nationales de 1750 à 1822. Revue d’Histoire des Sciences, 25(1):45–53, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631715.

Huard:1972:LCH

[1967] Pierre Huard and M. J. Imbault-Huart. L’enseignement de la chirurgie à l’Hôtel–Dieu : d’après une lettre inédite de Desault à l’Assemblée Nationale (1791). Revue d’Histoire des Sciences, 25(1):55–63, January/ March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631716.

Clavelin:1972:PGS

[1968] Maurice Clavelin. À propos des Galileo Studies de Stillman Drake. Revue d’Histoire des Sciences, 25(1):65–68, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631717.

Stipanic:1972:OMR

[1969] Ernest Stipanić. Une œuvre monumentale sur R. J. Bošković. Revue d’Histoire des Sciences, 25(1):69–72, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631718.

Youschkevitch:1972:IBP

[1970] A. P. Youschkevitch and A. T. Grigorian. Iosif Benedictovitch Pogrebysski (1906–1971). Revue d’Histoire des Sciences, 25(1):73–74, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631719.

D:1972:I

[1971] S. D. Informations. Revue d’Histoire des Sciences, 25(1):75–80, January/ March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631720.

Bernhardt:1972:AOBa

[1972] Jean Bernhardt. Analyses d’ouvrages: La philosophie italienne, (^{≪}Philosophies de tous les temps ^{≫}) par Emile Namer. Revue d’Histoire des Sciences, 25(1):81–82, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631721.

Bernhardt:1972:AOBb

[1973] Jean Bernhardt. Analyses d’ouvrages: The Mathematical and Philosophical Works of the Right Rev. ^{≪}Cass Library of Science Classics ^{≫}, n^{o} 11 par John Wilkins. Revue d’Histoire des Sciences, 25(1):82, January/ March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631722.

Casini:1972:AOB

[1974] Paolo Casini. Analyses d’ouvrages: The Methodological Heritage of Newton par Robert E. Butts; John W. Davis. Revue d’Histoire des Sciences, 25(1):83–87, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631723.

Colnort:1972:AOBa

[1975] S. Colnort. Analyses d’ouvrages: The Posthumous Works of —, ^{≪}The Sources of Science ^{≫}, n^{o} 73 par Robert Hooke; Richard S. Westfall. Revue d’Histoire des Sciences, 25(1):87–88, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631724.

Murata:1972:AOB

[1976] Tamotsu Murata. Analyses d’ouvrages: Lebesgue’s Theory of Integration. Its Origins and Development par Thomas Hawkins. Revue d’Histoire des Sciences, 25(1):88–89, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631725.

Cohen:1972:AOB

[1977] E. Cohen. Analyses d’ouvrages: Hilbert par Constance Reid; R. Courant. Revue d’Histoire des Sciences, 25(1):89–90, January/March 1972. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631726.

Grisard:1972:AOB

[1978] J. Grisard. Analyses d’ouvrages: Autopsie du hasard, (coll. ^{≪}Études Supérieures ^{≫}) par J.-L. Boursin; P. Caussat. Revue d’Histoire des Sciences, 25(1):90–91, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631727.

Colnort:1972:AOBb

[1979] S. Colnort. Analyses d’ouvrages: A History of Chemistry, Volume one, part I : Theoretical Background par J. R. Partington. Revue d’Histoire des Sciences, 25(1):91, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631728.

Colnort:1972:AOBc

[1980] S. Colnort. Analyses d’ouvrages: Teoría atómico-molecular par Dalton; Avogadro; Ampère; Leticia Halperin de Destaillats. Revue d’Histoire des Sciences, 25(1):91–92, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631729.

Pueyo:1972:AOB

[1981] G. Pueyo. Analyses d’ouvrages: Invention of the Meteorological Instruments par W. E. Knowles Middleton. Revue d’Histoire des Sciences, 25 (1):92–94, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631730.

Jilek:1972:AOB

[1982] Frantisek Jílek. Analyses d’ouvrages: Catalogue of Meteorological Instruments in the Museum of History and Technology (Smithsonian Studies in History and Technology, n^{o} 2) par W. E. Knowles Middleton. Revue d’Histoire des Sciences, 25(1):94–95, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631731.

Huard:1972:AOB

[1983] P. Huard. Analyses d’ouvrages: Un pionnier de la chirurgie moderne : Desault par Evelyn Saison. Revue d’Histoire des Sciences, 25(1):95–96, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631732.

Giard:1972:FDR

[1984] Luce Giard. La ^{≪}dialectique rationnelle ^{≫}de Gergonne. Revue d’Histoire des Sciences, 25(2):97–124, juin 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631578.

Salomon-Bayet:1972:OIC

[1985] Claire Salomon-Bayet. Opiologia, imposture et célébration de l’opium. Revue d’Histoire des Sciences, 25(2):125–150, juin 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631579.

Guillerme:1972:EAS

[1986] Jacques Guillerme. De l’écorché au schéma anatomique: A propos de^{≪}La Mémoire artificielle ^{≫}de Goiffon et Vincent (1779). Revue d’Histoire des Sciences, 25(2):151–170, juin 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631580.

Balan:1972:RIP

[1987] Bernard Balan. Sur le rôle de l’imaginaire dans la pratique psychiatrique au XIX^{e} siècle. Revue d’Histoire des Sciences, 25(2):171–190, juin 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631581.

Shuster-Aziza:1972:NHB

[1988] Eveline Shuster-Aziza and Evelyne Shuster-Aziza. Note sur H. de Blainville historien de la biologie. Revue d’Histoire des Sciences, 25(2): 191–200, juin 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631582.

Lusternik:1972:PED

[1989] L. A. Lusternik and S. S. Petrova. Les premières étapes du calcul symbolique. Revue d’Histoire des Sciences, 25(3):201–206, juillet–septembre 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631698.

Namer:1972:VPE

[1990] Emile Namer. Vanini et la préparation de l’esprit scientifique à l’aube du XVII^{e} siècle. Revue d’Histoire des Sciences, 25(3):207–220, juillet– septembre 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631699.

Sadoun-Goupil:1972:CCL

[1991] Michelle Sadoun-Goupil. La correspondance de Claude-Louis Berthollet et Martinus Van Marum (1786–1805). Revue d’Histoire des Sciences, 25(3):221–252, juillet–septembre 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631700.

Guedes:1972:TMM

[1992] Michel Guédès. La théorie de la métamorphose en morphologie végétale: A.-P. de Candolle et P.-J.-F. Turpin. Revue d’Histoire des Sciences, 25 (3):253–270, juillet–septembre 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631701.

Noir:1972:PLS

[1993] Dominique Noir. La première loi de similitude de la mécanique des fluides. Revue d’Histoire des Sciences, 25(3):271–274, juillet–septembre 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631702.

Costabel:1972:BR

[1994] Pierre Costabel. Bernard Rochot (1900–1971). Revue d’Histoire des Sciences, 25(3):275–277, juillet–septembre 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631703.

Huard:1972:DLC

[1995] Pierre Huard. Le Docteur Louis Chauvois (1881–1972). Revue d’Histoire des Sciences, 25(3):278–282, juillet–septembre 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631704.

Anonymous:1972:I

[1996] Anonyme. Informations. Revue d’Histoire des Sciences, 25(3):283– 286, juillet–septembre 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631705.

Hugonnard-Roche:1972:AOB

[1997] Henri Hugonnard-Roche. Analyses d’ouvrages: Le livre du ciel et du monde, ^{≪}The University of Wisconsin Publications in Medieval Science ≫, n^{O} 11 par Nicole Oresme; Albert D. Menut; Alexander J. Denomy. Revue d’Histoire des Sciences, 25(3):287–288, juillet–septembre 1972. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631706.

Sadoun-Goupil:1972:AOB

[1998] Michelle Sadoun-Goupil. Analyses d’ouvrages: Martinus Van Marum, Life and Work, vol. I et II par R. J. Forbes. Revue d’Histoire des Sciences, 25(3):288–290, juillet–septembre 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631707.

Speziali:1972:AOB

[1999] Pierre Speziali. Analyses d’ouvrages: Relativity Theory : its Origins and Impact on Modern Thought par L. Pearce Williams. Revue d’Histoire des Sciences, 25(3):290–292, juillet–septembre 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631708.

Fischer:1972:CSV

[2000] J.-L. Fischer. Chronologie sommaire de la vie et des travaux d’Étienne Geoffroy Saint-Hilaire (1772–1844). Revue d’Histoire des Sciences, 25 (4):293–300, October 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631980.

Cahn:1972:LEG

[2001] Théophile Cahn. L’œuvre d’Étienne Geoffroy Saint-Hilaire dans une perspective de l’évolution de la pensée scientifique. Revue d’Histoire des Sciences, 25(4):301–310, October 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631981.

Bourdier:1972:LGS

[2002] Franck Bourdier. Lamarck et Geoffroy Saint-Hilaire face au problème de l’évolution biologique. Revue d’Histoire des Sciences, 25(4):311–325, October 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631982.

Dagognet:1972:VHE

[2003] Fr. Dagognet. Valentin Haüy, Étienne Geoffroy Saint-Hilaire, Augustin P. de Candolle : Une conception d’ensemble mais aussi un ensemble de conceptions. Revue d’Histoire des Sciences, 25(4):327–336, October 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631983. Voir errata

[2017].

Duhamel:1972:LTE

[2004] Bernard Duhamel. L’œuvre tératologique d’Étienne Geoffroy Saint-Hilaire. Revue d’Histoire des Sciences, 25(4):337–346, October 1972. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631984.

Fischer:1972:CEO

[2005] J. L. Fischer. Le concept expérimental dans l’œuvre tératologique d’Étienne Geoffroy Saint-Hilaire. Revue d’Histoire des Sciences, 25 (4):347–364, October 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631985.

Laissus:1972:CME

[2006] Yves Laissus. Catalogue des manuscrits d’Étienne Geoffroy Saint-Hilaire conservés au Muséum. Revue d’Histoire des Sciences, 25(4):365–390, October 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631986.

Anonymous:1972:BM

[2007] Anonyme. Back matter. Revue d’Histoire des Sciences, 25(4):??, October 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631987.

Anonymous:1972:FMa

[2008] Anonyme. Front matter. Revue d’Histoire des Sciences, 25(1):??, January/March 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631711.

Anonymous:1972:FMb

[2009] Anonyme. Front matter. Revue d’Histoire des Sciences, 25(2):??, juin 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631577.

Anonymous:1972:FMc

[2010] Anonyme. Front matter. Revue d’Histoire des Sciences, 25(3):??, juillet– septembre 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631697.

Anonymous:1972:FMd

[2011] Anonyme. Front matter. Revue d’Histoire des Sciences, 25(4):??, October 1972. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631979.

Youschkevitch:1973:JCF

[2012] A. P. Youschkevitch. J. A. da Cunha et les fondements de l’analyse infinitésimale. Revue d’Histoire des Sciences, 26(1):3–22, January 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23631795.

Roger:1973:TTA

[2013] Jacques Roger. La Théorie de la Terre au XVII^{e} siècle. Revue d’Histoire des Sciences, 26(1):23–48, January 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631796.

Hall:1973:CIC

[2014] Marie Boas Hall. La croissance de l’industrie chimique en Grande-Bretagne au XIX^{e} siècle. Revue d’Histoire des Sciences, 26(1):49–68, January 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631797.

Taton:1973:QOR

[2015] René Taton. Sur quelques ouvrages récents concernant l’histoire de la science française. Revue d’Histoire des Sciences, 26(1):69–90, January 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631798.

Anonymous:1973:Ia

[2016] Anonyme. Informations. Revue d’Histoire des Sciences, 26(1):91–96, January 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631799.

Anonymous:1973:EVH

[2017] Anonyme. Errata: Valentin Haüy, Étienne Geoffroy Saint-Hilaire, Augustin P. de Candolle : Une conception d’ensemble mais aussi un ensemble de conceptions. Revue d’Histoire des Sciences, 26(1):96, January 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631800. Voir [2003].

Sinaceur:1973:CB

[2018] Hourya Sinaceur. Cauchy et Bolzano. Revue d’Histoire des Sciences, 26(2):97–112, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631766.

Dulieu:1973:JA

[2019] Louis Dulieu. Jean astruc. Revue d’Histoire des Sciences, 26(2):113– 135, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631767.

Todericiu:1973:JCF

[2020] Doru Todériciu. Jean-Claude Flachat (1700–1775): voyageur, industriel et technologue lyonnais. Revue d’Histoire des Sciences, 26(2):137–143, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631768.

Rouquette:1973:RWD

[2021] M. L. Rouquette. R. W. Darwin et la psychophysiologie de la vision. Revue d’Histoire des Sciences, 26(2):145–151, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631769.

Acloque:1973:PLD

[2022] P. Acloque. Préhistoire de la loi du rayonnement de Planck. Revue d’Histoire des Sciences, 26(2):153–157, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631770.

Theodorides:1973:I

[2023] J. Théodoridès. Informations. Revue d’Histoire des Sciences, 26(2):158– 163, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631771.

Picolet:1973:AOBa

[2024] Guy Picolet. Analyses d’ouvrages: John Pecham and the Science of Optics : Perspectiva Communis, (^{≪}The University of Wisconsin Publications in Medieval Science ^{≫}, 14) par David C. Lindberg. Revue d’Histoire des Sciences, 26(2):164–166, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631772.

Itard:1973:AOBa

[2025] Jean Itard. Analyses d’ouvrages: The Development of the foundations of Mathematical Analysis from Euler to Riemann par Ivor Grattan-Guinness. Revue d’Histoire des Sciences, 26(2):166–169, avril 1973. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631773.

Sadoun-Goupil:1973:AOBa

[2026] Michelle Sadoun-Goupil. Analyses d’ouvrages: Henry Le Chatelier. Un grand savant d’hier. Un précurseur. Sa vie, son œuvre, son temps par François Le Chatelier. Revue d’Histoire des Sciences, 26(2):169–170, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631774.

Fischer:1973:AOB

[2027] Jean Louis Fischer. Analyses d’ouvrages: Le catalogue de la vie, étude méthodologique sur la taxinomie, coll. ^{≪}Gallien ^{≫} par François Dagognet. Revue d’Histoire des Sciences, 26(2):170–171, avril 1973. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631775.

Colnort:1973:AOBa

[2028] S. Colnort. Analyses d’ouvrages: Georges Cuvier, de son temps au nôtre par Georges Cuvier. Revue d’Histoire des Sciences, 26(2):171–172, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631776.

Guedes:1973:AOBa

[2029] Michel Guédès. Analyses d’ouvrages: Darwin : Théorie de l’évolution (Les Grands Textes) par Yvette Conry. Revue d’Histoire des Sciences, 26(2):172–174, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631777.

Guedes:1973:AOBb

[2030] Michel Guédès. Analyses d’ouvrages: Principles of scientific Botany or Botany as an inductive science. ^{≪}The Sources of Science, 40 ^{≫}par M. J. Schleiden; Edwin Lankester; Jacob Lorch. Revue d’Histoire des Sciences, 26(2):174–175, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631778.

Grmek:1973:AOBa

[2031] M. D. Grmek. Analyses d’ouvrages: Therapie von den Primitiven bis zum 20. Jahrhundert par Erwin H. Ackerknecht. Revue d’Histoire des Sciences, 26(2):176, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631779.

Huard:1973:AOBa

[2032] P. Huard. Analyses d’ouvrages: Contribution à l’histoire de l’obstétrique et de la gynécologie, du XVI^{e} au XIX^{e} siècle par Nicole Génin. Revue d’Histoire des Sciences, 26(2):176–177, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631780.

Huard:1973:AOBb

[2033] P. Huard. Analyses d’ouvrages: Jacques-Louis Reverdin (1842–1929). Un chirurgien à l’aube d’une ère nouvelle par Henri Reverdin; Georges de Morsier. Revue d’Histoire des Sciences, 26(2):177–179, avril 1973. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631781.

Huard:1973:AOBc

[2034] P. Huard. Analyses d’ouvrages: Hundert Jahre Klinik Rheinau, 1867– 1967 par Hans H. Walser. Revue d’Histoire des Sciences, 26(2):179, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631782.

Huard:1973:AOBd

[2035] P. Huard. Analyses d’ouvrages: Histoire de la médecine japonaise moderne (de 1868 à nos jours) par Alain Briot. Revue d’Histoire des Sciences, 26(2):179–180, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632859.

Guedes:1973:AOBc

[2036] Michel Guédès. Analyses d’ouvrages: The evolution of the microscope, The microscope. Past and present, ibid. par S. Bradbury. Revue d’Histoire des Sciences, 26(2):180–182, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631783.

Colnort:1973:AOBb

[2037] S. Colnort. Analyses d’ouvrages: Die zoologische Buchillustration. Ihre Bibliographie und Geschichte, 1 fasc., n^{o} 9 par Claus Nissen. Revue d’Histoire des Sciences, 26(2):183, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631784.

Colnort:1973:AQPa

[2038] S. Colnort. Analyses de quelques périodiques: Archives internationales d’Histoire des Sciences anneés XXI, XXII et XXIII, 1968, 1969, 1970, n^{os} 82–83, 84–85, 86–87, 88–89, 90–91, 92–93. Revue d’Histoire des Sciences, 26(2):183–184, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631785.

Colnort:1973:AQPb

[2039] S. Colnort. Analyses de quelques périodiques: The British Journal of the History of Science, vol. V, 1970–1971, fasc. 17, 18, 19, 20. Revue d’Histoire des Sciences, 26(2):184, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631786.

Colnort:1973:AQPc

[2040] S. Colnort. Analyses de quelques périodiques: Physis, vol. XII, 1970, fasc. 1, 2, 3, 4. Revue d’Histoire des Sciences, 26(2):184–185, avril 1973. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631787.

Colnort:1973:AOBc

[2041] S. Colnort. Analyses d’ouvrages: Journal for the History of Astronomy, vol. 1, vol. 2 par M. A. Hoskin. Revue d’Histoire des Sciences, 26(2):185– 186, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631788.

Colnort:1973:AQPd

[2042] S. Colnort. Analyses de quelques périodiques: Journal of the History of Biology, vol. 1, 1968, n^{os} 1, 2; vol. 2, 1969, n^{os} 1, 2; vol. 3, 1970, n^{os} 1, 2. Revue d’Histoire des Sciences, 26(2):186–187, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631789.

Colnort:1973:AQPe

[2043] S. Colnort. Analyses de quelques périodiques: Gesnerus, revue trimestrielle publiée par la Société suisse d’Histoire de la Médecine et des Sciences naturelles, vol. 26, 1969, fasc. 1–2, 3–4; vol. 27, 1970, fasc. 1–2, 3–4; vol. 28, 1971 par Hans Fisher. Revue d’Histoire des Sciences, 26(2):187–188, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631790.

Guedes:1973:AOBd

[2044] Michel Guédès. Analyses d’ouvrages: Actes du XII^{e} Congrès international d’Histoire des Sciences (Paris, 1968). T. VIII : Histoire des Sciences naturelles et de la Biologie. Revue d’Histoire des Sciences, 26(2): 189–191, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631791.

Grmek:1973:AOBb

[2045] M. D. Grmek. Analyses d’ouvrages: Comptes rendus du XCIII^{e} Congrès national des Sociétés savantes (Tours, 1968), sections des Sciences, t. II : Histoire des Sciences et Biologie animale. Revue d’Histoire des Sciences, 26(2):191, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631792.

Anonymous:1973:PR

[2046] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 26(2): 192, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631793.

Scheler:1973:LRP

[2047] Lucien Scheler. Lavoisier et la régie des poudres. Revue d’Histoire des Sciences, 26(3):193–222, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631989.

Hodge:1973:MLA

[2048] J. M. S. Hodge and M. J. S. Hodge. La métaphysique de Lamarck d’après un opuscule retrouvé. Revue d’Histoire des Sciences, 26(3):223–229, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631990.

Bremond:1973:LE

[2049] J. Brémond and J. Lessertisseur. Lamarck et l’entomologie. Revue d’Histoire des Sciences, 26(3):231–250, juillet 1973. CODEN RHSAAM.

ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631991.

Boyer:1973:MHN

[2050] Ferdinand Boyer. Le Muséum d’Histoire naturelle à Paris et l’Europe des sciences sous la Convention. Revue d’Histoire des Sciences, 26(3):251– 257, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631992.

Anonymous:1973:Ib

[2051] Anonyme. Informations. Revue d’Histoire des Sciences, 26(3):259–260, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631993.

Cazenobe:1973:AOBa

[2052] J. Cazenobe. Analyses d’ouvrages: Perspectives in the History of Science and Technology par Duane H. D. Roller. Revue d’Histoire des Sciences, 26(3):261–262, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631994.

Sinaceur:1973:AOB

[2053] Mohammed-Allal Sinaceur. Analyses d’ouvrages: Anfänge der Griechischen Mathematik par Árpád Szabó. Revue d’Histoire des Sciences, 26(3): 262–264, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631995.

Itard:1973:AOBb

[2054] Jean Itard. Analyses d’ouvrages: Division and Logos, A Theory of Equivalent Couples and Sets of Integers Propounded par Euclid in the Arithmetical Books of the Elements par Christian Marinus Taisbak. Revue d’Histoire des Sciences, 26(3):265, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631996.

Itard:1973:AOBc

[2055] Jean Itard. Analyses d’ouvrages: t. II : Des spirales, De l’équilibre des figures planes, L’arénaire, La quadrature de la parabole; t. III : Des corps flottants, Stomachion, La méthode, Le livre des lemmes, Le problème des boeufs; (Association Guillaume–Budé) par Archimède; Charles Mugler. Revue d’Histoire des Sciences, 26(3):266–267, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631997.

Itard:1973:AOBd

[2056] Jean Itard. Analyses d’ouvrages: Histoire des logarithmes de Neper à Euler; t. II : La promotion des logarithmes au rang de valeur analytique par Charles Naux. Revue d’Histoire des Sciences, 26(3):267–268, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631998.

Itard:1973:AOBe

[2057] Jean Itard. Analyses d’ouvrages: The Mathematical Papers of Isaac Newton, vol. IV : 1674–1684 par D. T. Whiteside. Revue d’Histoire des Sciences, 26(3):269–271, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631999.

Demidov:1973:AOB

[2058] S. S. Demidov. Analyses d’ouvrages: Die Hilbertschen Probleme. Vortrag ^{≪}Mathematische Probleme ^{≫}von D. Hilbert gehalten auf dem 2. Internationalen Mathematikerkongress Paris 1900 erläutert von einem Autorenkollektiv unter der Redaktion ^{≪}Ostwalds Klassiker der exakten Wissenschaften ^{≫}, 252 par David Hilbert; P. S. Alexandrov. Revue d’Histoire des Sciences, 26(3):271–273, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632000.

Contro:1973:AOB

[2059] Walter S. Contro. Analyses d’ouvrages: Les Fondements de la géométrie par David Hilbert; Paul Rossier. Revue d’Histoire des Sciences, 26(3): 273–275, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632001.

Picolet:1973:AOBb

[2060] Guy Picolet. Analyses d’ouvrages: Les trois étapes de la cosmologie, (^{≪}Science nouvelle ^{≫}) par Jacques Merleau-Ponty; Bruno Morando. Revue d’Histoire des Sciences, 26(3):275–276, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632002.

Cazenobe:1973:AOBb

[2061] J. Cazenobe. Analyses d’ouvrages: Bergson and Modern Physics : a reinterpretation and a re-evaluation, vol. VII par Milic Capek. Revue d’Histoire des Sciences, 26(3):276–279, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632003.

Sadoun-Goupil:1973:AOBb

[2062] Michelle Sadoun-Goupil. Analyses d’ouvrages: Studies in the History of Chemistry par Harold Hartley. Revue d’Histoire des Sciences, 26(3):279– 280, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632004.

Florkin:1973:AOB

[2063] Marcel Florkin. Analyses d’ouvrages: The Chemistry of Life. Lectures on the History of Biochemistry par J. Needham. Revue d’Histoire des Sciences, 26(3):280–281, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632005.

Huard:1973:AOBe

[2064] P. Huard. Analyses d’ouvrages: Contribution à l’étude des correspondants de Haller et, en particulier, de Thierry par Odile Hamon. Revue d’Histoire des Sciences, 26(3):281–282, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632006.

Huard:1973:AOBf

[2065] P. Huard. Analyses d’ouvrages: Biobibliographie de Alfred-Louis-Armand-Marie Velpeau (1795–1867) par Gilbert Lizak. Revue d’Histoire des Sciences, 26(3):282–283, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632007.

Huard:1973:AOBg

[2066] Pierre Huard. Analyses d’ouvrages: Contribution à la biobibliographie de quelques anatomistes du XIX^{e} siècle par D. Renard-Payen. Revue d’Histoire des Sciences, 26(3):283–284, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632008.

Huard:1973:AOBh

[2067] P. Huard. Analyses d’ouvrages: Contribution à la biobibliographie de Louis-Hubert Farabeuf (1841–1910) par M. L. du Sorbiers. Revue d’Histoire des Sciences, 26(3):285–286, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632009.

Huard:1973:AOBi

[2068] P. Huard. Analyses d’ouvrages: Notre internat par Jean Quénu. Revue d’Histoire des Sciences, 26(3):286, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632010.

Bernhardt:1973:AOB

[2069] Jean Bernhardt. Analyses d’ouvrages: Cellular Pathology par Rudolf Virchow; Frank Chance; L. J. Rather. Revue d’Histoire des Sciences, 26(3): 286–287, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632011.

Theodorides:1973:AQP

[2070] J. Théodoridès. Analyses de quelques périodiques: Archives internationales Claude-Bernard, n^{o} 1, 1971, n^{o} 2, 1972. Revue d’Histoire des Sciences, 26(3):287–288, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632012.

Nevskaja:1973:JND

[2071] Nina I. Nevskaja. Joseph-Nicolas Delisle (1688–1768). Revue d’Histoire des Sciences, 26(4):289–313, October 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631648.

Huard:1973:EEL

[2072] Pierre Huard and Mirko D. Grmek. Les élèves étrangers de Laennec. Revue d’Histoire des Sciences, 26(4):315–337, October 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631649.

Wrotnowska:1973:FRD

[2073] D. Wrotnowska. Le ^{≪}rouget du porc ^{≫}Pasteur et Achille Maucuer: d’après une correspondance en partie inédite. Revue d’Histoire des Sciences, 26(4):339–364, October 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631650.

Murata:1973:EIE

[2074] Tamotsu Murata. Sur l’évolution de l’idée d’ ^{≪}effectif ^{≫}dans l’histoire de la théorie des ensembles. Revue d’Histoire des Sciences, 26(4):365–368, October 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631651.

Anonymous:1973:Ic

[2075] Anonyme. Informations. Revue d’Histoire des Sciences, 26(4):369–370, October 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631652.

Largeault:1973:AOB

[2076] J. Largeault. Analyses d’ouvrages: La syllogistique d’Aristote. coll. ≪Philosophies pour l’âge de la science ^{≫}par Jan Lukasiewicz; Françoise Gaujolle-Zaslawsky. Revue d’Histoire des Sciences, 26(4):371–375, October 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631653.

Grmek:1973:AOBc

[2077] M. D. Grmek. Analyses d’ouvrages: Un chirurgo italiano del 1200 : Bruno da Longoburgo par Mario Tabanelli. Revue d’Histoire des Sciences, 26(4):375–376, October 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631654.

Cohen:1973:AOB

[2078] E. Cohen. Analyses d’ouvrages: World directory of historians of mathematics par Kenneth O’May; Constance M. Gardner. Revue d’Histoire des Sciences, 26(4):376, October 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631655.

Anonymous:1973:TGM

[2079] Anonyme. Table générale des matières du Tome XXVI. Revue d’Histoire des Sciences, 26(4):377–384, October 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631656.

Anonymous:1973:FMa

[2080] Anonyme. Front matter. Revue d’Histoire des Sciences, 26(1):??, January 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631794.

Anonymous:1973:FMb

[2081] Anonyme. Front matter. Revue d’Histoire des Sciences, 26(2):??, avril 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631765.

Anonymous:1973:FMc

[2082] Anonyme. Front matter. Revue d’Histoire des Sciences, 26(3):??, juillet 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631988.

Anonymous:1973:FMd

[2083] Anonyme. Front matter. Revue d’Histoire des Sciences, 26(4):??, October 1973. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631647.

Taton:1974:ICO

[2084] René Taton. Inventaire chronologique de l’œuvre de Lagrange. Revue d’Histoire des Sciences, 27(1):3–36, January 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632077.

Broc:1974:MG

[2085] Numa Broc. Un Musée de Géographie en 1795. Revue d’Histoire des Sciences, 27(1):37–43, January 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632078.

Huard:1974:LLM

[2086] Pierre Huard and M. J. Imbault-Huart. L’enseignement libre de la médecine à Paris au XIX^{e} siècle. Revue d’Histoire des Sciences, 27(1): 45–62, January 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632079.

Todericiu:1974:NTH

[2087] Doru Todériciu, Suzanne Delorme, and Pierre Costabel. Notes sur trois hommes de science du XVII^{e} siècle : Samuel Duclos, Henri-Louis Habert de Montmor et Florimond de Beaune. Revue d’Histoire des Sciences, 27(1):63–67, January 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632080.

Delorme:1974:CAG

[2088] Suzanne Delorme. Un cartésien ami de Gassendi: Henri-Louis Habert de Montmor. Revue d’Histoire des Sciences, 27:68–72, 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

Costabel:1974:FBE

[2089] Pierre Costabel. Florimond de Beaune, érudit et savant de Blois. Revue d’Histoire des Sciences, 27:73–75, 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

Kouznetsov:1974:CEB

[2090] B. Kouznetsov and A. Frenk. La correspondance Einstein–Besso. Revue d’Histoire des Sciences, 27(1):77–82, January 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632081.

Theodorides:1974:GP

[2091] Jean Théodoridès. Georges Petit (1892–1973). Revue d’Histoire des Sciences, 27(1):83–85, January 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632082.

Anonymous:1974:Ia

[2092] Anonyme. Informations. Revue d’Histoire des Sciences, 27(1):86, January 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632083.

Itard:1974:AOBa

[2093] Jean Itard. Analyses d’ouvrages: t. IV, Commentaires d’Eutocius et fragments (Collection des Universités de France) par Archimède; Charles Mugler. Revue d’Histoire des Sciences, 27(1):87–88, January 1974. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632084.

Hugonnard-Roche:1974:AOB

[2094] Henri Hugonnard-Roche. Analyses d’ouvrages: Nicole Oresme and the Kinematics of Circular Motion, Tractatus de commensurabilitate vel incommensurabilitate motuum celi, (The University of Wisconsin Publications in Medieval Science, 15) par Edward Grant. Revue d’Histoire des Sciences, 27(1):88–91, January 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632085.

Itard:1974:AOBb

[2095] Jean Itard. Analyses d’ouvrages: Correspondance du P. Marin Mersenne, religieux minime, Tables et index cumulatif des t. I à X (années 1617–1641) par Cornelis De Waard. Revue d’Histoire des Sciences, 27(1):91–92, January 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632086.

Itard:1974:AOBc

[2096] Jean Itard. Analyses d’ouvrages: The Mathematical Papers of Isaac Newton, vol. V : 1683–1684 par D. T. Whiteside; M. A. Hoskin; A. Prag. Revue d’Histoire des Sciences, 27(1):92–94, January 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632087.

Grmek:1974:AOBa

[2097] M. D. Grmek and Jacques Roger. Analyses d’ouvrages: Des monstres et prodiges par Ambroise Paré; Jean Céard. Revue d’Histoire des Sciences, 27(1):95–96, January 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632088.

Rashed:1974:TPD

[2098] Roshdi Rashed. Les travaux perdus de Diophante. I. Revue d’Histoire des Sciences, 27(2):97–122, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631676.

McClaughlin:1974:LMT

[2099] Trevor McClaughlin. Une lettre de Melchisédech Thévenot. Revue d’Histoire des Sciences, 27(2):123–126, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631677.

Sadoun-Goupil:1974:SPS

[2100] Michelle Sadoun-Goupil. Science pure et science appliquée dans l’œuvre de Claude-Louis Berthollet. Revue d’Histoire des Sciences, 27(2):127– 145, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631678.

Paul:1974:SFS

[2101] Harry W. Paul. La science française de la seconde partie du XIX^{e} siècle vue par les auteurs anglais et américains. Revue d’Histoire des Sciences, 27(2):147–163, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631679.

Rosmorduc:1974:TNL

[2102] Jean Rosmorduc. Les théories sur la nature de la lumière au dix-septième siècle: les idées d’Aristote dans l’^{≪}Optique ^{≫}du R. P. Ango. Revue d’Histoire des Sciences, 27(2):165–168, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631680.

Taton:1974:I

[2103] R. Taton and R. Hahn. Informations. Revue d’Histoire des Sciences, 27(2):169–174, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631681.

Cazenobe:1974:AOB

[2104] J. Cazenobe. Analyses d’ouvrages: Faits et théories, (coll. ^{≪}U 2 ^{≫}) par William Mattew O’Neil; P. Acot. Revue d’Histoire des Sciences, 27(2): 175, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631682.

Poulle:1974:AOB

[2105] Emmanuel Poulle. Analyses d’ouvrages: Inventaire des manuscrits scientifiques de la Bibliothèque royale Albert-I^{er}, t. II par Roger Calcoen. Revue d’Histoire des Sciences, 27(2):175–176, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631683.

Picolet:1974:AOBa

[2106] Guy Picolet. Analyses d’ouvrages: Urkunden zur Geschichte der Mathematik im Mittelalter und der Renaissance (Biblioteca Mathematica Teubneriana, 45) par Maximilian Curtze. Revue d’Histoire des Sciences, 27(2):176–177, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631684.

Itard:1974:AOBd

[2107] Jean Itard. Analyses d’ouvrages: Correspondance du P. Marin Mersenne, religieux minime, t. XII : 1643 par Paul Tannery; Cornelis De Waard; Bernard Rochot. Revue d’Histoire des Sciences, 27(2):177– 178, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631685.

Sadoun-Goupil:1974:AOBa

[2108] Michelle Sadoun-Goupil. Analyses d’ouvrages: Leçons sur la philosophie chimique, L’Essai de statique chimique des êtres organisés par Jean-Baptiste Dumas; M. Bineau; M. Dumas; A. Bruylants. Revue d’Histoire des Sciences, 27(2):179–182, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631686.

Faure:1974:AOB

[2109] Françoise Faure. Analyses d’ouvrages: The Sorcerer’s Apprentice. The French Scientist’s Image of German Science, 1840–1919, (^{≪}University of Florida Social Sciences Monograph ^{≫}, n^{o} 44) par Harry W. Paul. Revue d’Histoire des Sciences, 27(2):182–183, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631687.

Fischer:1974:AOBa

[2110] Jean-Louis Fischer. Analyses d’ouvrages: The Great Doctors, a biographical History of Medecine, (Première édition en allemand, 1931, deuxième édition en anglais, 1933) par Henry E. Sigerist. Revue d’Histoire des Sciences, 27(2):183–184, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631688.

Fischer:1974:AOBb

[2111] Jean-Louis Fischer. Analyses d’ouvrages: The emergence of Scientific Medecine par William P. D. Wightman. Revue d’Histoire des Sciences, 27(2):184–185, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631689.

Taylor:1974:AOB

[2112] Kenneth L. Taylor. Analyses d’ouvrages: The Earth in Decay : A History of British Geomorphology, 1578–1878 par Gordon L. Davies. Revue d’Histoire des Sciences, 27(2):185–188, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631690.

Pueyo:1974:AOB

[2113] Guy Pueyo. Analyses d’ouvrages: Climate through the Ages par C. E. P. Brooks. Revue d’Histoire des Sciences, 27(2):188–189, avril 1974. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631691.

Colnort:1974:AOBa

[2114] S. Colnort. Analyses d’ouvrages: Nouveau Glossaire nautique par Augustin Jal. Revue d’Histoire des Sciences, 27(2):189, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631692.

Colnort:1974:AOBb

[2115] S. Colnort. Analyses d’ouvrages: History of Science, An annual review of Literature, Research and Teaching, vol. 9, 1970, vol. 10, 1971 par A. C. Crombie; M. A. Hoskin. Revue d’Histoire des Sciences, 27(2):189– 190, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631693.

Colnort:1974:AQPa

[2116] S. Colnort. Analyses de quelques périodiques: The British Journal for the History of Science, vol. VI, parts I et II, n^{os} 21, 22 par W. P. D. Wightman; R. Fox. Revue d’Histoire des Sciences, 27(2):190, avril 1974. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631694.

Colnort:1974:AOBc

[2117] S. Colnort. Analyses d’ouvrages: Journal for the History of Astronomy, vol. 3, 1972, part 1, n^{o} 6, part 2, n^{o} 7, part 3, n^{o} 8; vol. 4, 1973, part 1, n^{o} 9, part 2, n^{o} 10, part 3, n^{o} 11 par M. A. Hoskin. Revue d’Histoire des Sciences, 27(2):190–191, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631695.

Anonymous:1974:PR

[2118] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 27(2):191– 192, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631696.

Aujac:1974:LDG

[2119] Germaine Aujac. L’image du globe terrestre dans la Grèce ancienne. Revue d’Histoire des Sciences, 27(3):193–210, juillet 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631659.

Halleux:1974:NFM

[2120] Robert Halleux. La nature et la formation des métaux selon Agricola et ses contemporains. Revue d’Histoire des Sciences, 27(3):211–222, juillet 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631660.

Dulieu:1974:CFR

[2121] Louis Dulieu. Le chancelier François Ranchin. Revue d’Histoire des Sciences, 27(3):223–239, juillet 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631661.

Legee:1974:EOP

[2122] Georgette Legée. Essais sur l’organisation des plantes par A. Aubert Du Petit-Thouars. Revue d’Histoire des Sciences, 27(3):241–249, juillet 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631662.

Anonymous:1974:APH

[2123] Anonyme. Association pour l’Histoire des Sciences de la Nature: Séance de travail. Revue d’Histoire des Sciences, 27(3):250, juillet 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631663.

Sinaceur:1974:LNC

[2124] Mohammed-A. Sinaceur and R. Dedekind. L’infini et les nombres: Commentaires de R. Dedekind à ^{≪}Zahlen ^{≫}la correspondance avec Keferstein. Revue d’Histoire des Sciences, 27(3):251–278, juillet 1974. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631664.

Anonymous:1974:Ib

[2125] Anonyme. Informations. Revue d’Histoire des Sciences, 27(3):279–280, juillet 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631665.

Picolet:1974:AOBb

[2126] Guy Picolet. Analyses d’ouvrages: Introduction à la textologie, (^{≪}Collection L ^{≫}) par Roger Laufer. Revue d’Histoire des Sciences, 27 (3):281, juillet 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631666.

Rosmorduc:1974:AOB

[2127] Jean Rosmorduc. Analyses d’ouvrages: Introduction to concepts and theories in Physical Science, 2^{e} éd. (Addison-Wesley Series in Physics) par Gerald Holton; Stephen O. Brush. Revue d’Histoire des Sciences, 27(3):

281–283, juillet 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631667.

Fischer:1974:AOBc

[2128] Jean-Louis Fischer. Analyses d’ouvrages: Hunter’s Lectures of Anatomy. Revue d’Histoire des Sciences, 27(3):283–284, juillet 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631668.

Grmek:1974:AOBb

[2129] M. D. Grmek. Analyses d’ouvrages: Portraits of Doctors and Scientists in the Wellcome Institute of the History of Medicine par Renate Burgess. Revue d’Histoire des Sciences, 27(3):284–285, juillet 1974. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631669.

Grmek:1974:AOBc

[2130] M. D. Grmek. Analyses d’ouvrages: Die Geschichte der Arzneimittelforschung par Béla Issekutz. Revue d’Histoire des Sciences, 27(3):285– 286, juillet 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631670.

Colnort:1974:AOBd

[2131] S. Colnort. Analyses d’ouvrages: Répertoire des Jésuites de Chine de 1552 à 1800, ^{≪}Bibliotheca Instituti Historici S.I. ^{≫}, vol. XXXVII par Joseph Dehergne. Revue d’Histoire des Sciences, 27(3):286, juillet 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23631671.

Colnort:1974:AQPb

[2132] S. Colnort. Analyses de quelques périodiques: Gesnerus, revue trimestrielle publiée par la Société suisse d’Histoire de la Médecine et des Sciences naturelles, vol. 29, 1972, fasc. 1–2, 3–4 par Hans Fischer. Revue d’Histoire des Sciences, 27(3):286–287, juillet 1974. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631672.

Colnort:1974:AQPc

[2133] S. Colnort. Analyses de quelques périodiques: Journal of the History of Biology, vol. 4, 1971, n^{os} 1, 2; vol. 5, 1972, n^{os} 1, 2. Revue d’Histoire des Sciences, 27(3):287–288, juillet 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631673.

Biskup:1974:NRB

[2134] Marian Biskup. Nouvelles recherches sur la biographie de Nicolas Copernic (fondements méthodiques et résultats). Revue d’Histoire des Sciences, 27(4):289–306, October 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631745.

Taton:1974:CED

[2135] René Taton and Maylis Cazenave. Contribution à l’étude de la diffusion du “De Revolutionibus” de Copernic: Inventaire des exemplaires des deux premières éditions conservés en France. Revue d’Histoire des Sciences, 27(4):307–328, October 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631746.

Galarneau:1974:CAC

[2136] Claude Galarneau. Copernic au Canada français : l’interdit, l’hypothèse et la thèse. Revue d’Histoire des Sciences, 27(4):329–333, October 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23631747.

Itard:1974:PLP

[2137] Jean Itard. À propos d’un livre sur Pierre Fermat: (The mathematical career of Pierre de Fermat (1601–1665), Princeton Univ. Press, Princeton, N.J., 1973) by Michael Sean Mahoney. Revue d’Histoire des Sciences, 27(4):335–346, October 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631748.

Hahn:1974:DCS

[2138] Roger Hahn. Sur les débuts de la carrière scientifique de Lacepède. Revue d’Histoire des Sciences, 27(4):347–353, October 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631749.

Birembaut:1974:PNB

[2139] Arthur Birembaut. À propos des notices biographiques sur Sadi Carnot : quelques documents inédits. Revue d’Histoire des Sciences, 27(4):355– 370, October 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631750.

Levy:1974:AOB

[2140] Jacques Lévy. Analyses d’ouvrages: Œuvres complètes, vol. I par Nicolas Copernic. Revue d’Histoire des Sciences, 27(4):371–372, October 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631751.

Sadoun-Goupil:1974:AOBb

[2141] Michelle Sadoun-Goupil. Analyses d’ouvrages: Newton, Stahl, Boerhaave et la doctrine chimique (nouveau tirage) par Hélène Metzger. Revue d’Histoire des Sciences, 27(4):372–374, October 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631752.

Grmek:1974:AOBd

[2142] M. D. Grmek. Analyses d’ouvrages: Modern Methods in the History of Medicine par Edwin Clarke. Revue d’Histoire des Sciences, 27(4):374– 375, October 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631753.

Fischer:1974:AOBd

[2143] Jean-Louis Fischer. Analyses d’ouvrages: War on disease, a History of the Lister Institute par Harriette Chick; Margaret Hume; Marjorie Macfarlane. Revue d’Histoire des Sciences, 27(4):375–376, October 1974. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631754.

Anonymous:1974:BM

[2144] Anonyme. Back matter. Revue d’Histoire des Sciences, 27(4):??, October 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631755.

Anonymous:1974:FMa

[2145] Anonyme. Front matter. Revue d’Histoire des Sciences, 27(1):??, January 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632076.

Anonymous:1974:FMb

[2146] Anonyme. Front matter. Revue d’Histoire des Sciences, 27(2):??, avril 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631675.

Anonymous:1974:FMc

[2147] Anonyme. Front matter. Revue d’Histoire des Sciences, 27(3):??, juillet 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631658.

Anonymous:1974:FMd

[2148] Anonyme. Front matter. Revue d’Histoire des Sciences, 27(4):??, October 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631744.

Rashed:1975:TPD

[2149] Roshdi Rashed. Les travaux perdus de Diophante (II). Revue d’Histoire des Sciences, 28(1):3–30, January 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631812.

Broc:1975:DGF

[2150] Numa Broc. Les débuts de la géomorphologie en France : le tournant des années 1890. Revue d’Histoire des Sciences, 28(1):31–60, January 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631813.

Pyenson:1975:RRG

[2151] Lewis Pyenson. La réception de la relativité généralisée : disciplinarité et institutionalisation en physique. (French) [The reception of General Relativity: discipline and institutionalization in physics]. Revue d’Histoire des Sciences, 28(1):61–73, January 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631814.

Dugac:1975:PDD

[2152] Pierre Dugac. Sur la publication du dernier volume des œuvres d’Augustin Cauchy. Revue d’Histoire des Sciences, 28(1):75–83, January 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631815.

Anonymous:1975:Ia

[2153] Anonyme. Informations. Revue d’Histoire des Sciences, 28(1):84–86, January 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631816.

Picolet:1975:AOBa

[2154] Guy Picolet. Analyses d’ouvrages: Poland, The Land of Copernicus par Bogdan Suchodolski. Revue d’Histoire des Sciences, 28(1):87–88, January 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631817.

Sinaceur:1975:AOB

[2155] Hourya Sinaceur. Analyses d’ouvrages: Joseph Fourier, 1768–1830, A survey of his life and work, based of his monograph on the propagation of heat, presented to the Institut de France in 1807 par Ivor Grattan-Guinness; J. R. Ravetz. Revue d’Histoire des Sciences, 28(1):88–90, January 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631818.

Todericiu:1975:AOB

[2156] Doru Todériciu. Analyses d’ouvrages: Œuvres complètes, t. IV : Considérations sur la marche des idées et des événements dans les temps modernes par A. A. Cournot; André Robinet. Revue d’Histoire des Sciences, 28(1):90–91, January 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631819.

Picolet:1975:AOBb

[2157] Guy Picolet. Analyses d’ouvrages: L’hypothèse de l’atome primitif. L’hypothèse de l’atome primitif et le problème des amas de galaxies, Georges Lemaître et son œuvre; bibliographie des travaux de Georges Lemaître Éditions ^{≪}Culture et Civilisation ^{≫}(^{≪}Epistèmè ^{≫}) par Georges Lemaître; Ferdinand Gonseth; G. Lemaître; O. Godart. Revue d’Histoire des Sciences, 28(1):91–93, January 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631820.

Picolet:1975:AOBc

[2158] Guy Picolet. Analyses d’ouvrages: Tractatus de perspectiva, (^{≪}Franciscan Institute Publications ^{≫}, Text series, 16) par John Pecham; David C. Lindberg. Revue d’Histoire des Sciences, 28(1):93–95, January 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23631821.

Taylor:1975:AOB

[2159] K. L. Taylor. Analyses d’ouvrages: James Hutton — the Founder of Modern Geology par Edward Battersby Bailey. Revue d’Histoire des Sciences, 28(1):96, January 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631822.

Kubli:1975:PDA

[2160] Fritz Kubli. À propos du 50^{e} anniversaire de la mécanique ondulatoire. Revue d’Histoire des Sciences, 28(2):97–112, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632014.

Itard:1975:LTR

[2161] Jean Itard. La lettre de Torricelli à Roberval d’octobre 1643. Revue d’Histoire des Sciences, 28(2):113–124, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632015.

Dulieu:1975:CFL

[2162] Louis Dulieu. Claude-François Lallemand (1790–1854). Revue d’Histoire des Sciences, 28(2):125–138, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632016.

LeRoux:1975:SFF

[2163] J.-J. Le Roux, P. Huard, and M.-J. Imbault-Huart. Structure et fonctionnement de la Faculté de Médecine de Paris en 1813. Revue d’Histoire des Sciences, 28(2):139–168, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632017.

Bernhardt:1975:MCG

[2164] Jean Bernhardt. Mersenne, commentateur de Galilée: à propos d’une édition critique des ^{≪}Nouvelles Pensées de Galilée ^{≫}. Revue d’Histoire des Sciences, 28(2):169–177, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632019.

Anonymous:1975:Ib

[2165] Anonyme. Informations. Revue d’Histoire des Sciences, 28(2):178–181, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632020.

Bernhardt:1975:AOB

[2166] Jean Bernhardt. Analyses d’ouvrages: Critique du De Mundo de Thomas White, (^{≪}L’Histoire des Sciences, Textes et Études ^{≫}) par Thomas Hobbes; Jean Jacquot; Harold Whitmore Jones. Revue d’Histoire des Sciences, 28(2):182–183, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632021.

Charbonneau:1975:AOBa

[2167] Louis Charbonneau. Analyses d’ouvrages: De l’Infini mathématique par Louis Couturat. Revue d’Histoire des Sciences, 28(2):183–184, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632022.

Charbonneau:1975:AOBb

[2168] Louis Charbonneau. Analyses d’ouvrages: Notice sur les travaux scientifiques, (coll. ^{≪}Discours de la Méthode ^{≫}) par Elie Cartan. Revue d’Histoire des Sciences, 28(2):184–186, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632023.

Charbonneau:1975:AOBc

[2169] Louis Charbonneau. Analyses d’ouvrages: Le passage au matérialisme : Idéalisme et matérialisme, Mathématiques et matérialisme, (coll. ≪Théorie ^{≫}, série ^{≪}Analyse ^{≫}) par Pierre Raymond. Revue d’Histoire des Sciences, 28(2):186–187, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632024.

Halleux:1975:AQP

[2170] Robert Halleux. Analyses de quelques périodiques: L’atome et l’histoire, (n^{os} 45 et 46 de la revue Protection contre les rayonnements ionisants) par Pierre Pizon. Revue d’Histoire des Sciences, 28(2):187, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632025.

Sadoun-Goupil:1975:AOBa

[2171] Michelle Sadoun-Goupil. Analyses d’ouvrages: The Selected Correspondence of Michael Faraday, vol. 1 : 1812–1848 vol. 2 : 1849–1866 par L. Pearce Williams; Rosemary FitzGerald; Oliver Stallybrass. Revue d’Histoire des Sciences, 28(2):187–188, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632026.

Grmek:1975:AOBa

[2172] M. D. Grmek. Analyses d’ouvrages: The Heart and the Vascular System in Ancient Greek Medicine. From Alcmaeon to Galen par C. R. S.

Harris. Revue d’Histoire des Sciences, 28(2):189–190, avril 1975. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632027.

Grmek:1975:AOBb

[2173] M. D. Grmek. Analyses d’ouvrages: Tecniche e strumenti chirurgici del XIII e XIV secolo par Mario Tabanelli. Revue d’Histoire des Sciences, 28(2):190, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632028.

Halleux:1975:AOB

[2174] Robert Halleux. Analyses d’ouvrages: Scritti di idraulica fluviale e di canalizzazione, a cura di Cristiana Fischer, Volume II par Paolo Frisi. Revue d’Histoire des Sciences, 28(2):191, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632029.

Roger:1975:AOB

[2175] Jacques Roger. Analyses d’ouvrages: Des révolutions vertes. Histoire et principes de l’agronomie, coll. ^{≪}Savoir ^{≫}par François Dagognet. Revue d’Histoire des Sciences, 28(2):191–192, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632030.

Grave:1975:BMP

[2176] P. Radelet de Grave and D. Speiser. Le “De magnete” de Pierre de Maricourt: Traduction et commentaire. Revue d’Histoire des Sciences, 28(3): 193–234, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632032.

McClaughin:1975:REC

[2177] Trevor McClaughin and Trevor McClaughlin. Sur les rapports entre la compagnie de Thévenot et l’Académie royale des Sciences. Revue d’Histoire des Sciences, 28(3):235–242, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632033.

Kniajetskaia:1975:PGF

[2178] E. A. Kniajetskaia and V. L. Chenakal. Pierre le Grand et les fabricants français d’instruments scientifiques. Revue d’Histoire des Sciences, 28(3): 243–258, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632034.

Hjelm-Hansen:1975:LIW

[2179] Nils Hjelm-Hansen. Une lettre inédite de A. Wurtz à J.-B. Dumas. Revue d’Histoire des Sciences, 28(3):259–265, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632035.

Stipanic:1975:SOD

[2180] Ernest Stipanić. Symposium à l’occasion du 350^{e} anniversaire de la mort de Marc-Antoine Dominis. Revue d’Histoire des Sciences, 28(3):267–269, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632036.

Delorme:1975:I

[2181] Suzanne Delorme. Informations. Revue d’Histoire des Sciences, 28(3): 270–273, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632037.

Mandelbaum:1975:AOB

[2182] Jonathan Mandelbaum. Analyses d’ouvrages: Science and Society : Past, Present, and Future par Nicholas H. Steneck. Revue d’Histoire des Sciences, 28(3):274–275, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632038.

Charbonneau:1975:AOBd

[2183] Louis Charbonneau. Analyses d’ouvrages: Leibniz in Paris (1672– 1676). His growth to mathematical maturity par Joseph E. Hofmann. Revue d’Histoire des Sciences, 28(3):275–277, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632039.

Stipanic:1975:AOB

[2184] Ernest Stipanić. Analyses d’ouvrages: Teorija prirodne filozofije [Théorie de la philosophie naturelle] par Rudjer Bosković; Roger Boscovich; Jakov Stipisić; Zarko Dadić; Vladimir Filipović. Revue d’Histoire des Sciences, 28(3):277–279, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632040.

Sadoun-Goupil:1975:AOBb

[2185] Michelle Sadoun-Goupil. Analyses d’ouvrages: Martinus Van Marum, Life and Work, vol. III par R. J. Forbes; Martinus Van Marum, Life and Work, vol. IV et V par E. Lefebvre; J. G. de Bruijn. Revue d’Histoire des Sciences, 28(3):279–281, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632041.

Huard:1975:AOBa

[2186] P. Huard. Analyses d’ouvrages: La Haritasamhita par Alix Raison. Revue d’Histoire des Sciences, 28(3):281–282, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632042.

Theodorides:1975:AOB

[2187] Jean Théodoridès. Analyses d’ouvrages: La Médecine japonaise des origines à nos jours par Pierre Huard; Zensetsou Ohya; Ming Wong. Revue d’Histoire des Sciences, 28(3):282–283, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632043.

Huard:1975:AOBb

[2188] P. Huard. Analyses d’ouvrages: Orezza, Guagno : évolution des indications thérapeutiques de deux sources thermales corses au XIX^{e} siècle par Vincent-Pierre Comiti. Revue d’Histoire des Sciences, 28(3):283–285, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632044.

Cazenave:1975:AQPa

[2189] Maylis Cazenave. Analyses de quelques périodiques: Archives internationales d’Histoire des Sciences, vol. XXIV, juin 1974, n^{o} 94 par W. Hartner; P. Costabel; M. D. Grmek; J. D. North. Revue d’Histoire des Sciences, 28(3):285–286, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632045.

Cazenave:1975:AOBa

[2190] Maylis Cazenave. Analyses d’ouvrages: History of Science, A review of literature and research in the history of science, medicine and technology in its intellectual and social context, vol. 11, 1973,Part 1, 2, 3, 4 par R. S. Porter. Revue d’Histoire des Sciences, 28(3):286, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632924.

Cazenave:1975:AOBb

[2191] Maylis Cazenave. Analyses d’ouvrages: Historia Mathematica, revue internationale d’histoire des mathématiques, vol. 1, 1974, fasc. 1, 2, 3, 4, vol. 2, 1975, fasc. 1 par Kenneth O. May. Revue d’Histoire des Sciences, 28(3):286–287, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632925.

Cazenave:1975:AOBc

[2192] Maylis Cazenave. Analyses d’ouvrages: Historical Studies in the Physical Sciences, vol. 2 par Russell McCormmach. Revue d’Histoire des Sciences, 28(3):287, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632926.

Cazenave:1975:AQPb

[2193] Maylis Cazenave. Analyses de quelques périodiques: Gesnerus, revue trimestrielle publiée par la Société suisse d’Histoire de la Médecine et des Sciences naturelles, vol. 31, 1974, fasc. 1–2, 3–4, vol. 32, 1975, fasc. 1–2 (numéro spécial) par Hans H. Walser. Revue d’Histoire des Sciences, 28(3):287–288, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632927.

Balan:1975:PRO

[2194] Bernard Balan. Premières recherches sur l’origine et la formation du concept d’économie animale. Revue d’Histoire des Sciences, 28(4):289– 326, October 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632090.

Laurent:1975:LPD

[2195] Goulven Laurent. Lamarck : de la philosophie du continu à la science du discontinu. Revue d’Histoire des Sciences, 28(4):327–360, October 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632091.

Ubrizsy:1975:CCO

[2196] Andrea Ubrizsy. Contribution à la connaissance des œuvres de Clusius. Revue d’Histoire des Sciences, 28(4):361–370, October 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632092.

Theodorides:1975:NJH

[2197] Jean Théodoridès. Notes sur J.-H. Fabre. Revue d’Histoire des Sciences, 28(4):371–373, October 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632093.

Anonymous:1975:Ic

[2198] Anonyme. Informations. Revue d’Histoire des Sciences, 28(4):374–377, October 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632094.

Anonymous:1975:TGM

[2199] Anonyme. Table générale des matières du Tome XXVIII. Revue d’Histoire des Sciences, 28(4):378–384, October 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632095.

Anonymous:1975:FMa

[2200] Anonyme. Front matter. Revue d’Histoire des Sciences, 28(1):??, January 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631811.

Anonymous:1975:FMb

[2201] Anonyme. Front matter. Revue d’Histoire des Sciences, 28(2):??, avril 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632013.

Anonymous:1975:FMc

[2202] Anonyme. Front matter. Revue d’Histoire des Sciences, 28(3):??, juillet 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632031.

Anonymous:1975:FMd

[2203] Anonyme. Front matter. Revue d’Histoire des Sciences, 28(4):??, October 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632089.

McClaughlin:1976:EUD

[2204] Trevor McClaughlin and Guy Picolet. Un exemple d’utilisation du Minutier central de Paris : La bibliothèque et les instruments scientifiques du physicien Jacques Rohault selon son inventaire après décès. Revue d’Histoire des Sciences, 29(1):3–20, January 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632047.

Merleau-Ponty:1976:SRH

[2205] Jacques Merleau-Ponty. Situation et rôle de l’hypothèse cosmogonique dans la pensée cosmologique de Laplace. Revue d’Histoire des Sciences, 29(1):21–49, January 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632048.

Rosmorduc:1976:TTS

[2206] Jean Rosmorduc. Les tribulations d’un terme scientifique : ^{≪}pôle ^{≫}… ≪polarisation ^{≫}, de la cosmologie de platon à l’optique du XIX^{e} siècle. Revue d’Histoire des Sciences, 29(1):51–62, January 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632049.

Charbonneau:1976:FHP

[2207] Louis Charbonneau. Fourier, l’homme et le physicien d’après John Herivel. Revue d’Histoire des Sciences, 29(1):63–72, January 1976. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632050.

Taton:1976:PBO

[2208] René Taton. À propos des ^{≪}Opera omnia ^{≫}d’Euler. Revue d’Histoire des Sciences, 29(1):73–77, January 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632051.

Anonymous:1976:Ia

[2209] Anonyme. Informations. Revue d’Histoire des Sciences, 29(1):78, January 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632052.

deBuzon:1976:AOBa

[2210] C. de Buzon. Analyses d’ouvrages: L’histoire et les sciences, coll. ≪Algorithme ^{≫}par Pierre Raymond. Revue d’Histoire des Sciences, 29 (1):79, January 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632053.

Todericiu:1976:AOBa

[2211] Doru Todériciu. Analyses d’ouvrages: Le Retour aux sources. Pour l’Histoire des Sciences dans l’enseignement scientifique français par M. Rosmorduc. Revue d’Histoire des Sciences, 29(1):80–81, January 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632054.

Caveing:1976:AOBa

[2212] M. Caveing. Analyses d’ouvrages: Histoire comparée des numérations écrites par Geneviève Guitel. Revue d’Histoire des Sciences, 29(1):81–86, January 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632055.

Rosenfeld:1976:AOB

[2213] B. A. Rosenfeld and A. P. Youschkevitch. Analyses d’ouvrages: La topologie algébrique, des origines à Poincaré (Bibliothèque de Philosophie contemporaine) par Jean-Claude Pont. Revue d’Histoire des Sciences, 29(1):86–89, January 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632056.

Horowitz:1976:AOB

[2214] Philip Horowitz. Analyses d’ouvrages: Origins of modern algebra par Lubos Nový. Revue d’Histoire des Sciences, 29(1):89–91, January 1976. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632057.

Cohen:1976:AOB

[2215] E. Cohen. Analyses d’ouvrages: Message d’un mathématicien : Henri Lebesgue par Lucienne Félix; S. Mandelbrojt. Revue d’Histoire des Sciences, 29(1):92, January 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632058.

Picolet:1976:AOBa

[2216] Guy Picolet. Analyses d’ouvrages: A Catalogue of Medieval and Renaissance Optical Manuscripts, (^{≪}Subsidia Mediaevalia ^{≫}, 4) par David C. Lindberg. Revue d’Histoire des Sciences, 29(1):92–94, January 1976. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632059.

Grmek:1976:AOBa

[2217] M. D. Grmek. Analyses d’ouvrages: Les transmissions chimiques de l’influx nerveux, (coll. ^{≪}Discours de la méthode ^{≫}) par Z. M. Bacq. Revue d’Histoire des Sciences, 29(1):94–95, January 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632060.

Grmek:1976:AOBb

[2218] M. D. Grmek. Analyses d’ouvrages: La chirurgie à Montpellier de ses origines au début du XIX^{e} siècle, Avignon par Louis Dulieu. Revue d’Histoire des Sciences, 29(1):95–96, January 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632061.

Taton:1976:ICO

[2219] René Taton. Inventaire chronologique de l’œuvre d’Alexis-Claude Clairaut (1713–1765). Revue d’Histoire des Sciences, 29(2):97–122, avril 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631962.

Azouvi:1976:SFE

[2220] François Azouvi. Sens et fonction épistémologiques de la critique du magnétisme animal par les Académies. Revue d’Histoire des Sciences, 29(2):123–142, avril 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631963.

Cazenobe:1976:TCM

[2221] J. Cazenobe. Trop connu, méconnu : le tube à limaille (1890). Revue d’Histoire des Sciences, 29(2):143–165, avril 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631964.

Sinaceur:1976:AD

[2222] Mohammed-Allal Sinaceur. Sur l’algèbre de Diophante. Revue d’Histoire des Sciences, 29(2):167–170, avril 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631965.

Jaouiche:1976:I

[2223] K. Jaouiche. Informations. Revue d’Histoire des Sciences, 29(2):171– 174, avril 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631966.

deBuzon:1976:AOBb

[2224] Catherine de Buzon. Analyses d’ouvrages: Reason, Experiment and Mysticism in the Scientific Revolution par M. L. Righini Bonelli; William R. Shea. Revue d’Histoire des Sciences, 29(2):175–176, avril 1976. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631967.

Murphy:1976:AOBa

[2225] Terence Murphy. Analyses d’ouvrages: Science and Change, 1500–1700 (coll. ^{≪}World University Library ^{≫}) par Hugh Kearney. Revue d’Histoire des Sciences, 29(2):176–178, avril 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631968.

Murphy:1976:AOBb

[2226] Terence Murphy. Analyses d’ouvrages: Mechanism and Materialism : British Natural Philosophy in an Age of Reason par Robert E. Schofield. Revue d’Histoire des Sciences, 29(2):178–179, avril 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631969.

Charbonneau:1976:AOBa

[2227] Louis Charbonneau. Analyses d’ouvrages: Sur la nature des mathématiques, (coll. ^{≪}Discours de la méthode ^{≫}) par Claude-Paul Bruter. Revue d’Histoire des Sciences, 29(2):179–180, avril 1976. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631970.

Caveing:1976:AOBb

[2228] Maurice Caveing. Analyses d’ouvrages: The Evolution of the Euclidean Elements : A Study of the Theory of incommensurable Magnitudes and its significance for early Greek Geometry par Wilbur Richard Knorr. Revue d’Histoire des Sciences, 29(2):180–183, avril 1976. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631971.

Rashed:1976:AOB

[2229] Roshdi Rashed. Analyses d’ouvrages: Geschichte des arabischen Schrifttums, Band V, Mathematik, bis ca. 430 H. par Fuat Sezgin. Revue d’Histoire des Sciences, 29(2):183–184, avril 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631972.

Jaouiche:1976:AOB

[2230] K. Jaouiche. Analyses d’ouvrages: Al-Bahir en algèbre par as-Samaw’al. Revue d’Histoire des Sciences, 29(2):184–186, avril 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631973.

Merleau-Ponty:1976:AOBa

[2231] Jacques Merleau-Ponty. Analyses d’ouvrages: Introductions à l’astronomie de Copernic. Le Commentariolus de Copernic. La Narratio prima de Rheticus par H. Hugonnard-Roche; E. Rosen; J. P. Verdet; R. Taton. Revue d’Histoire des Sciences, 29(2):186–187, avril 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631974.

Picolet:1976:AOBb

[2232] Guy Picolet. Analyses d’ouvrages: New Optics, (^{≪}Pubblicazioni della Fondazione ^{≪}Giorgio Ronchi ^{≫}^{≫}, 14) par Vasco Ronchi. Revue d’Histoire des Sciences, 29(2):188, avril 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631975.

Halleux:1976:AOB

[2233] Robert Halleux. Analyses d’ouvrages: Roman and Islamic Waterlifting Wheels, vol. 28 (Acta Historica Scientiarum Naturalium et Medicinalium. Edidit Bibliotheca Universitatis Hauniensis, vol. 28.) Thorkild Schiøler. Revue d’Histoire des Sciences, 29(2):189–190, avril 1976. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631976.

Huard:1976:AOBa

[2234] P. Huard. Analyses d’ouvrages: Les maladies pestilentielles dans le Sud-Ouest, sous l’Ancien Régime par P.-L. Thillaud. Revue d’Histoire des Sciences, 29(2):190, avril 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631977.

Anonymous:1976:PR

[2235] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 29(2):191– 192, avril 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631978.

Moreau:1976:LAS

[2236] Joseph Moreau. L’essor de l’astronomie scientifique chez les Grecs. (French) [The growth of scientific astronomy with the Greeks]. Revue d’Histoire des Sciences, 29(3):193–212, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631825.

Olmsted:1976:RBA

[2237] J. W. Olmsted. Recherches sur la biographie d’un astronome et géodésien méconnu : Jean Picard (1620–1682). Revue d’Histoire des Sciences, 29(3):213–222, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631826.

Dulieu:1976:ARE

[2238] Louis Dulieu. Aperçu sur les relations entre l’Amérique et l’École de Médecine de Montpellier : Les thèses médicales montpelliéraines imprimées aux XVII^{e}, XVIII^{e} et XIX^{e} siècles. Revue d’Histoire des Sciences, 29(3):223–242, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631827.

Redondi:1976:SCR

[2239] Pietro Redondi. Sadi Carnot et la recherche technologique en France de 1825 à 1850: Présentation d’un travail de recherche. Revue d’Histoire des Sciences, 29(3):243–259, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631828.

Beaude:1976:CIG

[2240] Joseph Beaude. Sur le cartésianisme d’Ignace-Gaston Pardies: A propos de l’ouvrage du Père A. Ziggelaar. Revue d’Histoire des Sciences, 29(3): 261–267, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631829.

Anonymous:1976:Ib

[2241] Anonyme. Informations. Revue d’Histoire des Sciences, 29(3):268–272, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631830.

deBeauregard:1976:AOBa

[2242] O. Costa de Beauregard. Analyses d’ouvrages: The Study of Time, t. I, t. II par J. T. Fraser; F. C. Haber; G. H. Muller; N. Lawrence. Revue d’Histoire des Sciences, 29(3):273, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631831.

deBeauregard:1976:AOBb

[2243] O. Costa de Beauregard. Analyses d’ouvrages: A chacun son temps, (Nouvelle Bibliothèque scientifique) par A. Pacaud; C. Vidal; Colonna; Aventur; Guitton. Revue d’Histoire des Sciences, 29(3):273–274, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631832.

Namer:1976:AOB

[2244] Emile Namer. Analyses d’ouvrages: The Ash Wednesday Supper, La Cena de le Ceneri par Giordano Bruno; Stanley L. Jaki. Revue d’Histoire des Sciences, 29(3):274–276, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631833.

Stipanic:1976:AOB

[2245] Ernest Stipanić. Analyses d’ouvrages: Marc-Antoine de Dominis considéré comme savant et utopiste par Dusan Nedeljković. Revue d’Histoire des Sciences, 29(3):276–278, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631834.

deBuzon:1976:AOBc

[2246] C. de Buzon. Analyses d’ouvrages: ^{≪}La catégorie de XVII^{e} siècle; Philosophie et Sciences ^{≫}. Numéro spécial de la Revue internationale de Philosophie, n^{o} 114, 1975, fasc. 4. Revue d’Histoire des Sciences, 29(3): 278–279, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631835.

McClaughlin:1976:AOB

[2247] Trevor McClaughlin. Analyses d’ouvrages: Œuvres philosophiques avec une étude bio-bibliographique par Louis de la Forge; Pierre Clair. Revue d’Histoire des Sciences, 29(3):279, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631836.

Charbonneau:1976:AOBb

[2248] Louis Charbonneau. Analyses d’ouvrages: Ernst Mach, His Work, Life and Influence par John T. Blackmore. Revue d’Histoire des Sciences, 29(3):279–280, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631837.

Todericiu:1976:AOBb

[2249] Doru Todériciu. Analyses d’ouvrages: Codices Vossiani Chymici, (coll. Bibliotheca Universitatis Leidensis. Codices manuscripti, XVII) par P. C. Boeren. Revue d’Histoire des Sciences, 29(3):281, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631838.

Leroy:1976:AOB

[2250] Jean François Leroy. Analyses d’ouvrages: ^{≪}Introduction bibliographique à l’histoire de la biologie. ^{≫}Histoire et nature, 1974–1975, n^{o} 5–6. Revue d’Histoire des Sciences, 29(3):281–282, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631839.

Guilhot:1976:AOB

[2251] Joëlle Guilhot. Analyses d’ouvrages: L’introduction du darwinisme en France au XIX^{e} siècle par Yvette Conry. Revue d’Histoire des Sciences, 29(3):282–285, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631840.

Broc:1976:AOB

[2252] Numa Broc. Analyses d’ouvrages: Studia z Dziejow Geografii i Kartografii (= Études d’histoire de la géographie et de la cartographie) par Jozefa Babicza. Revue d’Histoire des Sciences, 29(3):285–288, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631841.

Charbonneau:1976:AOBc

[2253] Louis Charbonneau. Analyses d’ouvrages: L’imagerie scientifique de Paul Valéry, (coll. ^{≪}Essais d’art et de philosophie ^{≫}) par Reino Virtanen. Revue d’Histoire des Sciences, 29(3):288, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631842.

Aujac:1976:CFS

[2254] Germaine Aujac. Le ciel des fixes et ses représentations en Grèce ancienne. Revue d’Histoire des Sciences, 29(4):289–307, October 1976. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632097.

Costabel:1976:PPS

[2255] Pierre Costabel. La physique de Pascal et son analyse structurale. Revue d’Histoire des Sciences, 29(4):309–324, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632098.

Boyer:1976:VSA

[2256] Ferdinand Boyer. Un voyage scientifique après Campo Formio: Trois lettres de Barthélemy Faujas de Saint-Fond à ses collègues du Muséum. Revue d’Histoire des Sciences, 29(4):325–336, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632099.

Broc:1976:ECP

[2257] Numa Broc. Eugène Cortambert et la place de la géographie dans la classification des connaissances humaines (1852). Revue d’Histoire des Sciences, 29(4):337–345, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632100.

LHuillier:1976:ENP

[2258] Hervé L’Huillier. Éléments nouveaux pour la biographie de Nicolas Chuquet. Revue d’Histoire des Sciences, 29(4):347–350, October 1976. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632101.

Theodorides:1976:I

[2259] Jean Théodoridès. Informations. Revue d’Histoire des Sciences, 29 (4):351–356, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632102.

Merleau-Ponty:1976:AOBb

[2260] Jacques Merleau-Ponty. Analyses d’ouvrages: Avant, avec, après Copernic. La représentation de l’Univers et ses conséquences épistémologiques (Centre international de Synthèse : XXXI^{e} Semaine de Synthèse). Revue d’Histoire des Sciences, 29(4):357–359, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632103.

Debarbat:1976:AOB

[2261] Suzanne Débarbat. Analyses d’ouvrages: The Search for the Nebulae par Kenneth Glyn Jones; M. A. Hoskin. Revue d’Histoire des Sciences, 29(4):359–360, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632104.

Brouzeng:1976:AOB

[2262] Paul Brouzeng. Analyses d’ouvrages: Physics circa 1900. Personnel, Funding, and Productivity of the Academic Establishments, in Historical Studies in the Physical Sciences, Annual Volume 5 par Paul Forman; John L. Heilbron; Spencer Weart; Russel McCormmach. Revue d’Histoire des Sciences, 29(4):360–361, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632105.

Costabel:1976:AOB

[2263] Pierre Costabel. Analyses d’ouvrages: Marc Seguin, grand savant méconnu par Claude Bechetoille. Revue d’Histoire des Sciences, 29 (4):361–362, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632106.

Sadoun-Goupil:1976:AOB

[2264] Michelle Sadoun-Goupil. Analyses d’ouvrages: A History of Biochemistry, Parts I and II, (coll. ^{≪}Comprehensive Biochemistry ^{≫}, vol. 30) par Marcel Florkin; M. Florkin; E. H. Stotz. Revue d’Histoire des Sciences, 29(4):362–363, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632107.

Theodorides:1976:AOB

[2265] Jean Théodoridès. Analyses d’ouvrages: Die zoologische Buchillustration. Ihre Bibliographie und Geschichte, Band I, Bibliographie, Band II, Geschichte (6 livraisons parues) par Claus Nissen. Revue d’Histoire des Sciences, 29(4):363–364, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632108.

Guedes:1976:AOBa

[2266] Michel Guédès. Analyses d’ouvrages: Sur un opuscule peu connu de Charles de L’Ecluse, Ann. du XLIII^{e} Congrès de la Féd. des Cercles Archéol. Hist. Belgique par J. E. Opsomer; Charles de L’Ecluse. Revue d’Histoire des Sciences, 29(4):365, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632109.

Guedes:1976:AOBb

[2267] Michel Guédès. Analyses d’ouvrages: Sur quelques plantes magiques ou légendaires décrites dans les ^{≪}herbiers ^{≫}du Moyen Âge, Ann. du XLIII^{e} Congrès de la Féd. des Cercles Archéol. Hist. Belgique par C. Opsomer. Revue d’Histoire des Sciences, 29(4):365, October 1976. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632110.

Huard:1976:AOBb

[2268] P. Huard. Analyses d’ouvrages: La corporation des maîtres chirurgiens de Nantes, de 1752 à 1793 par Ph. Bousicot. Revue d’Histoire des Sciences, 29(4):366–367, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632111.

Huard:1976:AOBc

[2269] P. Huard. Analyses d’ouvrages: Biobibliographie d’Auguste Nélaton (1807–1873) par Anusia Ouguag. Revue d’Histoire des Sciences, 29 (4):367–368, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632112.

Huard:1976:AOBd

[2270] P. Huard. Analyses d’ouvrages: Biobibliographie d’Aristide Verneuil de Saint-Martin (1823–1895) par Landine Barrachi. Revue d’Histoire des Sciences, 29(4):368–370, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632114.

Murphy:1976:AOBc

[2271] Terence Murphy. Analyses d’ouvrages: Thinkers and Tinkers, Early American Men of Science par Silvio A. Bedini. Revue d’Histoire des Sciences, 29(4):370–371, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632115.

Pueyo:1976:AOB

[2272] Guy Pueyo. Analyses d’ouvrages: Les méthodes culturales au Moyen Âge d’après les traités d’agronomie andalous : traditions et techniques par Lucie Bolens. Revue d’Histoire des Sciences, 29(4):371–372, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632116.

Todericiu:1976:AOBc

[2273] Doru Todériciu. Analyses d’ouvrages: L’agriculture et le feu. Rôle et place du feu dans les techniques de préparation du champ de l’ancienne agriculture européenne, (École des Hautes Études en Sciences sociales, Cahiers des Études Sociales, I) par François Sigaut. Revue d’Histoire des Sciences, 29(4):372–373, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632117.

Cazenave:1976:AQP

[2274] Maylis Cazenave. Analyses de quelques périodiques: Archive for History of Exact Sciences, vol. 11–15 par C. Truesdell. Revue d’Histoire des Sciences, 29(4):373–375, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632118.

Cazenave:1976:AOBa

[2275] Maylis Cazenave. Analyses d’ouvrages: The British Journal for the History of Science, vol. VI, 1973, part 3, n^{o} 23, part 4, n^{o} 24; vol. VII, 1974, part. 1, n^{o} 25, part 2, n^{o} 26, part 3, n^{o} 27; vol. VIII, 1975, part 1, n^{o} 28, part 2, n^{o} 29, part 3, n^{o} 30 par Robert Fox. Revue d’Histoire des Sciences, 29(4):375–376, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633019.

Cazenave:1976:AOBb

[2276] Maylis Cazenave. Analyses d’ouvrages: Journal for the History of Astronomy, vol. 5, 1974, part 1, n^{o} 12, part 2, n^{o} 13, part 3, n^{o} 14; vol. 6, 1975, part 1, n^{o} 15, part 2, n^{o} 16, part 3, n^{o} 17; vol. 7, 1976, part 1, n^{o} 18, part 2, n^{o} 19 par M. A. Hoskin. Revue d’Histoire des Sciences, 29(4): 376, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633020.

Anonymous:1976:TGM

[2277] Anonyme. Table générale des matières du Tome XXX. Revue d’Histoire des Sciences, 29(4):377–384, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632119.

Anonymous:1976:FMa

[2278] Anonyme. Front matter. Revue d’Histoire des Sciences, 29(1):??, January 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632046.

Anonymous:1976:FMb

[2279] Anonyme. Front matter. Revue d’Histoire des Sciences, 29(2):??, avril 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631961.

Anonymous:1976:FMc

[2280] Anonyme. Front matter. Revue d’Histoire des Sciences, 29(3):??, juillet 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631824.

Anonymous:1976:FMd

[2281] Anonyme. Front matter. Revue d’Histoire des Sciences, 29(4):??, October 1976. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632096.

Bernhardt:1977:HML

[2282] Jean Bernhardt. Hobbes et le mouvement de la lumière. Revue d’Histoire des Sciences, 30(1):3–24, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631559.

Targosz:1977:JHS

[2283] Karolina Targosz. Johann Hevelius et ses démarches pour trouver des mécènes en France. Revue d’Histoire des Sciences, 30(1):25–41, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631560.

Laurent:1977:CEG

[2284] Goulven Laurent. Le cheminement d’Étienne Geoffroy Saint-Hilaire (1772–1844) vers un transformisme scientifique. Revue d’Histoire des Sciences, 30(1):43–70, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631561.

Merleau-Ponty:1977:ESR

[2285] Jacques Merleau-Ponty. Situation et rôle de l’hypothèse cosmogonique dans la pensée cosmologique de Laplace (note rectificative à l’article paru dans le n^{o} 1 de 1976). Revue d’Histoire des Sciences, 30(1):71–72, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631562.

Anonymous:1977:Ia

[2286] Anonyme. Informations. Revue d’Histoire des Sciences, 30(1):73–76, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631563.

Louis:1977:AOB

[2287] Pierre Louis. Analyses d’ouvrages: Der Kommentar des Radulphus Brito zu Buch III, De Anima, Radulphi Britonis quaestiones in Aristotelis librum tertium De Anima. Kritische Edition und Philosophisch-historische Einleitung, (Beiträge zur Geschichte der Philosophie und Theologie des Mittelalters, Neue Folge, Bd 12) par Winfried Fauser. Revue d’Histoire des Sciences, 30(1):77–78, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631564.

Picolet:1977:AOBa

[2288] Guy Picolet. Analyses d’ouvrages: Le problème des métaux dans la science antique, Société d’Edition ^{≪}Les Belles Lettres ^{≫}, 1974, (^{≪}Bibliothèque de la Faculté de Philosophie et Lettres de l’Université de Liège ^{≫}, fasc. 209) par Robert Halleux. Revue d’Histoire des Sciences, 30(1):78–80, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631565.

Bernhardt:1977:AOBa

[2289] Jean Bernhardt. Analyses d’ouvrages: Two New Sciences par Galileo; Stillman Drake. Revue d’Histoire des Sciences, 30(1):80–81, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631566.

Bernhardt:1977:AOBb

[2290] Jean Bernhardt. Analyses d’ouvrages: L’affaire Galilée, (collection ≪Archives ^{≫}) par Emile Namer. Revue d’Histoire des Sciences, 30(1): 81–82, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631567.

Taton:1977:AOBa

[2291] René Taton. Analyses d’ouvrages: Doctrine de l’Angle solide. Inventaire de sa bibliothèque. (coll. ^{≪}L’Histoire des sciences. Textes et études ^{≫}) par Florimond de Beaune; Pierre Costabel. Revue d’Histoire des Sciences, 30(1):82–84, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631568.

Rosmorduc:1977:AOBa

[2292] J. Rosmorduc. Analyses d’ouvrages: Cette étrange matière, (coll. ≪Le Monde Ouvert ^{≫}) par Alfred Kastler. Revue d’Histoire des Sciences, 30(1):84–87, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631569.

Mandelbaum:1977:AOB

[2293] Jonathan Mandelbaum. Analyses d’ouvrages: Aux sources de la biologie. Tome premier: Les vingt premiers siècles. La classification par Réjane Bernier. Revue d’Histoire des Sciences, 30(1):87–89, January 1977. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631570.

Florkin:1977:AOB

[2294] Marcel Florkin. Analyses d’ouvrages: Crystals, Fabrics and Fields. Metaphors of organicism in Twentieth-Century developmental Biology par Donna Jeanne Haraway. Revue d’Histoire des Sciences, 30(1):89–90, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631571.

Cazenave:1977:AOBa

[2295] M. Cazenave. Analyses d’ouvrages: Comptes rendus du XCVII^{e}, IC^{e}, C^{e} et CI^{e} Congrès national des Sociétés savantes Comité des travaux historiques et scientifiques. Revue d’Histoire des Sciences, 30(1):90–93, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631572.

Cazenave:1977:AQPa

[2296] M. Cazenave. Analyses de quelques périodiques: Gesnerus, Revue trimestrielle publiée par la Société suisse d’histoire de la Médecine et des Sciences naturelles, vol. 32, 1975, fasc. 3–4; vol. 33, 1976, fasc. 1–2, 3–4 par Hans H. Walser; Heinz Balmer. Revue d’Histoire des Sciences, 30 (1):93, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631573.

Cazenave:1977:AQPb

[2297] M. Cazenave. Analyses de quelques périodiques: Physis, Rivista internazionale di Storia della Scienza, vol. XIII (1971)–vol. XVII (1975). Revue d’Histoire des Sciences, 30(1):94–96, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631574.

Allard:1977:ALR

[2298] Michel Allard. Antoine-Louis Rouillé secrétaire d’État à la marine (1749–1754) : progrès scientifique et marine. Revue d’Histoire des Sciences, 30(2):97–103, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631600.

Costabel:1977:LSA

[2299] Pierre Costabel. L’activité scientifique d’Ampère. Revue d’Histoire des Sciences, 30(2):105–112, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631601.

Merleau-Ponty:1977:LPS

[2300] Jacques Merleau-Ponty. L’Essai sur la philosophie des sciences d’Ampère. Revue d’Histoire des Sciences, 30(2):113–118, avril 1977. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631602.

Marcovich:1977:TPR

[2301] André Marcovich. La théorie philosophique des rapports d’André-Marie Ampère. Revue d’Histoire des Sciences, 30(2):119–123, avril 1977. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631603.

Sadoun-Goupil:1977:EOA

[2302] Michelle Sadoun-Goupil. Esquisse de l’œuvre d’Ampère en chimie. Revue d’Histoire des Sciences, 30(2):125–141, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631604.

Kastler:1977:ALE

[2303] Alfred Kastler. Ampère et les lois de l’électrodynamique. Revue d’Histoire des Sciences, 30(2):143–157, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631605.

Rosmorduc:1977:AOI

[2304] Jean Rosmorduc. Ampère et l’optique : une intervention dans le débat sur la transversalité de la vibration lumineuse. Revue d’Histoire des Sciences, 30(2):159–167, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631606.

Anonymous:1977:Ib

[2305] Anonyme. Informations. Revue d’Histoire des Sciences, 30(2):169–172, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631607.

Jeauneau:1977:AOB

[2306] Édouard Jeauneau. Analyses d’ouvrages: Myth and Science in the Twelfth Century. A Study of Bernard Silvester par B. Stock. Revue d’Histoire des Sciences, 30(2):173–174, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631608.

Shea:1977:AOB

[2307] William R. Shea. Analyses d’ouvrages: L’étrenne ou la neige sexangulaire par Johann Kepler; Robert Halleux; René Taton. Revue d’Histoire des Sciences, 30(2):174–175, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631609.

Bernhardt:1977:AOBc

[2308] Jean Bernhardt. Analyses d’ouvrages: Correspondance de Jacques Dupuy et de Nicolas Heinsius (1646–1656), (coll. ^{≪}Archives internationales d’histoire des idées ^{≫}, n^{o} 40) par Hans Bots. Revue d’Histoire des Sciences, 30(2):175–176, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631610.

Bernhardt:1977:AOBd

[2309] Jean Bernhardt. Analyses d’ouvrages: Samuel Hartlib and the Advancement of Learning, (coll. ^{≪}Cambridge Texts and Studies in the History of Education ^{≫}) par Charles Webster. Revue d’Histoire des Sciences, 30(2): 176–178, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631611.

Cazenobe:1977:AOB

[2310] J. Cazenobe. Analyses d’ouvrages: Some Aspects of Eighteenth-Century England par J. H. Plumb; Vinton A. Dearing. Revue d’Histoire des Sciences, 30(2):178, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631612.

Taton:1977:AOBb

[2311] René Taton. Analyses d’ouvrages: Écrits et mémoires mathématiques d’E. G. par Évariste Galois; Robert Bourgne; J.-P. Azra; J. Dieudonné. Revue d’Histoire des Sciences, 30(2):179, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631613.

Giuculescu:1977:AOBa

[2312] A. Giuculescu. Analyses d’ouvrages: Les fondements des mathématiques. De la Géométrie d’Euclide à la Relativité générale et à l’Intuitionnisme par F. Gonseth; Jacques Hadamard. Revue d’Histoire des Sciences, 30(2): 179–180, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631614.

Giuculescu:1977:AOBb

[2313] A. Giuculescu. Analyses d’ouvrages: Les mathématiques et la réalité, Essai sur la méthode axiomatique, 1936 par Ferdinand Gonseth. Revue d’Histoire des Sciences, 30(2):180–181, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631615.

Charbonneau:1977:AOB

[2314] Louis Charbonneau. Analyses d’ouvrages: Essai sur la psychologie de l’invention dans le domaine mathématique, (coll. ^{≪}Discours de la méthode ^{≫}) par Jacques Hadamard. Revue d’Histoire des Sciences, 30(2): 181, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631616.

Sadoun-Goupil:1977:AOB

[2315] Michelle Sadoun-Goupil. Analyses d’ouvrages: La genèse de la science des cristaux, 2^{e} éd. par Hélène Metzger. Revue d’Histoire des Sciences, 30(2):182, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631617.

Guedes:1977:AOBa

[2316] Michel Guédès. Analyses d’ouvrages: Henri-Louis Duhamel du Monceau et le ministère de la Marine par Michel Allard; Roland Lamontagne. Revue d’Histoire des Sciences, 30(2):182–183, avril 1977. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631618.

Theodorides:1977:AOB

[2317] Jean Théodoridès. Analyses d’ouvrages: Henri Dutrochet (Henri du Trochet 1776–1847), le matérialisme mécaniste et la physiologie générale par Joseph Schiller; Tetty Schiller. Revue d’Histoire des Sciences, 30(2):183– 184, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631619.

Fischer:1977:AOBa

[2318] Jean-Louis Fischer. Analyses d’ouvrages: Inédits de Lamarck, d’après les manuscrits conservés à la Bibliothèque centrale du Muséum national d’Histoire naturelle de Paris, (ouvrage publié avec le concours du CNRS) par Max Vachon; Georges Rousseau; Yves Laissus; Georges Canguilhem; Pierre P. Grassé. Revue d’Histoire des Sciences, 30(2):184–186, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631620.

Fischer:1977:AOBb

[2319] Jean-Louis Fischer. Analyses d’ouvrages: Colloque international Lamarck, tenu au Muséum national d’Histoire naturelle de Paris les 1^{er}, 2 et 3 juillet 1971 par P.-P. Grassé; J. Orcel; J.-P. Leham; P. Boiteau; C. Castellani. Revue d’Histoire des Sciences, 30(2):186–187, avril 1977. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631621.

Bernhardt:1977:AOBe

[2320] Jean Bernhardt. Analyses d’ouvrages: The Pseudo-Hippocratic Tract Πϵρὶ ϵ́βδoμάδων, ch. 1–11 and Greek Philosophy, (coll. ^{≪}Philosophical Texts and Studies ^{≫}, n^{o} 20) par J. Mansfeld. Revue d’Histoire des Sciences, 30(2):187–188, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631622.

Grmek:1977:AOB

[2321] M. D. Grmek. Analyses d’ouvrages: Studies in Pre-Vesalian Anatomy. Biography, Translations, Documents, (coll. ^{≪}Memoirs of the American Philosophical Society ^{≫}, vol. 104) par L. R. Lind. Revue d’Histoire des Sciences, 30(2):188, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631623.

Huard:1977:AOBa

[2322] P. Huard. Analyses d’ouvrages: Chlorpromazine in Psychiatry. A study of Therapeutic par Judith P. Swazey. Revue d’Histoire des Sciences, 30(2):189–190, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631624.

Huard:1977:AOBb

[2323] P. Huard. Analyses d’ouvrages: Bibliography of the History of Medicine, n^{o} 9, 1973. Revue d’Histoire des Sciences, 30(2):190–191, avril 1977. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631625.

Huard:1977:AOBc

[2324] P. Huard. Analyses d’ouvrages: Bibliographie de la Côte-d’Ivoire, 3 tomes, sous le titre des Annales de l’Université d’Abidjan, numéro hors série, 1972, 1973, 1975 par G. Janvier. Revue d’Histoire des Sciences, 30(2):191–192, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631626.

Guedes:1977:AOBb

[2325] Michel Guédès. Analyses d’ouvrages: La génétique : rêves et réalités par Macfarlane Burnet; Suzanne Nicol. Revue d’Histoire des Sciences, 30(2): 192, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631627.

Cuvelier:1977:BCI

[2326] Pierre Cuvelier. Les Experimenta crystalli Islandici disdiaclastici d’Erasme Bartholin: Traduction française commentée. Revue d’Histoire des Sciences, 30(3):193–224, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631865.

McClaughlin:1977:CSC

[2327] Trevor McClaughlin. Le concept de science chez Jacques Rohault. Revue d’Histoire des Sciences, 30(3):225–240, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631866.

McClellan:1977:MIC

[2328] James E. McClellan. Un manuscrit inédit de Condorcet : Sur l’utilité des académies. Revue d’Histoire des Sciences, 30(3):241–253, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23631867.

Hominal:1977:CCA

[2329] François Hominal. Sur la chimie en Chine d’après Joseph Needham. Revue d’Histoire des Sciences, 30(3):255–261, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631868.

Taton:1977:I

[2330] R. Taton. Informations. Revue d’Histoire des Sciences, 30(3):262–264, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631869.

Bernhardt:1977:AOBf

[2331] Jean Bernhardt. Analyses d’ouvrages: La logique et son histoire d’Aristote à Russell, (coll. ^{≪}U ^{≫}) par Robert Blanché. Revue d’Histoire des Sciences, 30(3):265, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631870.

Bernhardt:1977:AOBg

[2332] Jean Bernhardt. Analyses d’ouvrages: The Justification of Scientific Change, (^{≪}Synthese Library ^{≫}) par Carl R. Kordig. Revue d’Histoire des Sciences, 30(3):266–267, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631871.

Otero:1977:AOB

[2333] Mario H. Otero. Analyses d’ouvrages: The Method of Analysis. Its Geometrical Origin and Its General Significance par Jaakko Hintikka; Unto Remes. Revue d’Histoire des Sciences, 30(3):267–272, juillet 1977. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631872.

Picolet:1977:AOBb

[2334] Guy Picolet. Analyses d’ouvrages: Regesta Copernicana (Calendar of Copernicus’ Papers), (^{≪}Studia Copernicana ^{≫}, VIII) par Marian Biskup. Revue d’Histoire des Sciences, 30(3):272–274, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631873.

deBuzon:1977:AOB

[2335] Catherine de Buzon. Analyses d’ouvrages: Thematic Origins of Scientific Thought. Kepler to Einstein par Gerald Holton. Revue d’Histoire des Sciences, 30(3):275–276, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631874.

Halleux:1977:AOBa

[2336] Robert Halleux. Analyses d’ouvrages: Francis Bacon and Socialized Science, (^{≪}a Monograph in The Bannerstone Division of American Lectures in Philosophy ^{≫}. ^{≪}American Lecture Series, 96 ^{≫}) par Antoinette Mann Paterson. Revue d’Histoire des Sciences, 30(3):276, juillet 1977. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631875.

Blondel:1977:AOBa

[2337] Christine Blondel. Analyses d’ouvrages: The papers of Joseph Henry, vol. I: Dec. 1797–Oct. 1822; vol. II: Nov. 1832–Dec. 1835 par Nathan Reingold. Revue d’Histoire des Sciences, 30(3):277, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631876.

Blondel:1977:AOBb

[2338] Christine Blondel. Analyses d’ouvrages: The Compton effect. Turning point in Physics par Roger H. Stuewer. Revue d’Histoire des Sciences, 30 (3):278, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631877.

Guedes:1977:AOBc

[2339] Michel Guédès. Analyses d’ouvrages: L’équilibre de la nature … par C. Linné; B. Jasmin; C. Limoges. Revue d’Histoire des Sciences, 30(3):278– 280, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631878.

Huard:1977:AOBd

[2340] Pierre Huard. Analyses d’ouvrages: L’épopée pasteurienne et la médecine vétérinaire, (21, rue Gustave-Laubat, 92380 Garches) par Louis Nicol; C. Bressou. Revue d’Histoire des Sciences, 30(3):280–282, juillet 1977. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631879.

Bernhardt:1977:AOBh

[2341] Jean Bernhardt. Analyses d’ouvrages: Correspondance entre Charles Darwin et Gaston de Saporta, (coll. ^{≪}Galien ^{≫}) par Y. Conry. Revue d’Histoire des Sciences, 30(3):282, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631880.

Cazenave:1977:AOBb

[2342] M. Cazenave. Analyses d’ouvrages: Catalogue of the Paul Ehrenfest Archive at the Museum Boerhaave, (Communication 151 of the National Museum for the History of Science and Medicine — Museum Boerhaave) par Bruce Wheaton. Revue d’Histoire des Sciences, 30(3):282–283, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631881.

Cazenave:1977:AQPc

[2343] Maylis Cazenave. Analyses de quelques périodiques: Archives internationales d’Histoire des Sciences, vol. XXIV, n^{o} 95 (déc. 1974) à vol. XXVI, n^{o} 99 (déc. 1976) par W. Hartner; P. Costabel; M. D. Grmek; J. D. North. Revue d’Histoire des Sciences, 30(3):283, juillet 1977. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631882.

Cazenave:1977:AOBc

[2344] Maylis Cazenave. Analyses d’ouvrages: Annali dell’Istituto e Museo di Storia della Scienza di Firenze par Maria Luisa Righini Bonelli; Celeste Cantù; Marta Balduini. Revue d’Histoire des Sciences, 30(3):284, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631883.

Cazenave:1977:AQPd

[2345] Maylis Cazenave. Analyses de quelques périodiques: Isis, an International Review devoted to the History of Science and its cultural influences, vol. 61, n^{o} 209, 1970 à vol. 67, n^{o} 240, 1976 par Robert P. Multhauf. Revue d’Histoire des Sciences, 30(3):284–287, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631884.

Anonymous:1977:OR

[2346] Anonyme. Ouvrages reçus. Revue d’Histoire des Sciences, 30(3):288, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631885.

Ponte:1977:ATE

[2347] Maurice Ponte. Ampère et le télégraphe électrique: Réflexions sur les conceptions d’Ampère relatives aux conséquences économiques et sociales de la découverte scientifique et sur les délais qui séparent celle-ci de ses applications. Revue d’Histoire des Sciences, 30(4):289–302, October 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631904.

Rabkin:1977:CPD

[2348] Yakov M. Rabkin. La chimie et le pétrole : les débuts d’une liaison. Revue d’Histoire des Sciences, 30(4):303–336, October 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631905.

Youschevitch:1977:TOS

[2349] A. P. Youschkevitch, René Taton, and Pierre Dugac. Sur trois ouvrages soviétiques récents concernant l’histoire des mathématiques. Revue d’Histoire des Sciences, 30(4):337–359, October 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631906.

deBeauregard:1977:QCT

[2350] Olivier Costa de Beauregard. Sur quelques citations tirées de ^{≪}la théorie physique, son objet, sa structure ^{≫}de Pierre Duhem. Revue d’Histoire des Sciences, 30(4):361–366, October 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631907.

Anonymous:1977:Ic

[2351] Anonyme. Informations. Revue d’Histoire des Sciences, 30(4):367–368, October 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631908.

Margolin:1977:AOB

[2352] Jean-Claude Margolin. Analyses d’ouvrages: Anthropologie philosophique de la Renaissance par Bogdan Suchodolski; Irena Wojnar; Benoit Zawisza. Revue d’Histoire des Sciences, 30(4):369–370, October 1977. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631909.

Rosmorduc:1977:AOBb

[2353] J. Rosmorduc. Analyses d’ouvrages: De Homine. Traité de l’Homme par Thomas Hobbes; Paul-Marie Maurin; V. Ronchi. Revue d’Histoire des Sciences, 30(4):370–371, October 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631910.

Hugonnard-Roche:1977:AOB

[2354] Henri Hugonnard-Roche. Analyses d’ouvrages: Les mathématiques arabes (VIII^{e}–XV^{e} siècles), (coll. ^{≪}L’histoire des sciences, Textes et études ^{≫}) par Adolf P. Youschkevitch; M. Cazenave; K. Jaouiche; R. Taton. Revue d’Histoire des Sciences, 30(4):371–373, October 1977. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631911.

Taton:1977:AOBc

[2355] René Taton. Analyses d’ouvrages: Die Berliner und die Petersburger Akademie der Wissenschaften im Briefwechsel Leonhard Eulers : Teil 3, Wissenschaftliche und wissenschaftsorganisatorische Korrespondenzen, 1726–1774 par A. P. Juskevic; E. Winter; P. Hofmann; T. N. Klado; Ju. Ch. Kopelevic. Revue d’Histoire des Sciences, 30(4):373–375, October 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631912.

Poulle:1977:AOB

[2356] Emmanuel Poulle. Analyses d’ouvrages: Traité de l’astrolabe par Henri Michel; Fr. Maddison. Revue d’Histoire des Sciences, 30(4):375–376, October 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631913.

Anonymous:1977:Id

[2357] Anonyme. Information. Revue d’Histoire des Sciences, 30(4):376, October 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631915.

Halleux:1977:AOBb

[2358] Robert Halleux. Analyses d’ouvrages: I minerali della Toscana. Saggio di mineralogia regionale, (Accademia Toscana di Scienze e Lettere ^{≪}La Colombaria ^{≫}, Studi, XLII) par Guido Carobbi; Francesco Rodolico. Revue d’Histoire des Sciences, 30(4):376, October 1977. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631914.

Anonymous:1977:TGM

[2359] Anonyme. Table générale des matières du Tome XXIX. Revue d’Histoire des Sciences, 30(4):377–384, October 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631916.

Anonymous:1977:FMa

[2360] Anonyme. Front matter. Revue d’Histoire des Sciences, 30(1):??, January 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631558.

Anonymous:1977:FMb

[2361] Anonyme. Front matter. Revue d’Histoire des Sciences, 30(2):??, avril 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631599.

Anonymous:1977:FMc

[2362] Anonyme. Front matter. Revue d’Histoire des Sciences, 30(3):??, juillet 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631864.

Anonymous:1977:FMd

[2363] Anonyme. Front matter. Revue d’Histoire des Sciences, 30(4):??, October 1977. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631903.

Picolet:1978:CJP

[2364] Guy Picolet. La correspondance de Jean Picard avec Johann Hevelius (1671–1679): Edition et traduction française. Revue d’Histoire des Sciences, 31(1):3–42, January 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631926.

Ellenberger:1978:DMA

[2365] François Ellenberger. Le dilemme des montagnes au XVIII^{e} siècle : vers une réhabilitation des diluvianistes? Revue d’Histoire des Sciences, 31(1):43–52, January 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631927.

Blondel:1978:PRF

[2366] Christine Blondel. Sur les premières recherches de formule électrodynamique par Ampère (octobre 1820). Revue d’Histoire des Sciences, 31(1):53–65, January 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631928.

Hurwic:1978:PMP

[2367] Józef Hurwic. À propos d’un mémoire de 1854 par J. Trouessart traitant du développement de la notion d’élément chimique. Revue d’Histoire des Sciences, 31(1):67–72, January 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631929.

Petrescou:1978:CHR

[2368] Aurel D. Petrescou. Contribution à l’histoire des relations scientifiques franco-roumaines : Petrache Poenaru prospecte en France les nouvelles techniques (1825–1831). Revue d’Histoire des Sciences, 31(1):73–75, January 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631930.

Anonymous:1978:Ia

[2369] Anonyme. Informations. Revue d’Histoire des Sciences, 31(1):76–78, January 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631931.

Rosmorduc:1978:AOBa

[2370] Jean Rosmorduc. Analyses d’ouvrages: Théorie générale de l’information et de la communication, (coll. ^{≪}Langue-Linguistique-Communication ^{≫}) par Robert Escarpit. Revue d’Histoire des Sciences, 31(1):79–80, January 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631932.

Pares:1978:AOB

[2371] Jean Parès. Analyses d’ouvrages: Greek and Roman Sundials par Sharon L. Gibbs. Revue d’Histoire des Sciences, 31(1):80–81, January 1978. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631933.

Khawam:1978:AOB

[2372] René R. Khawam. Analyses d’ouvrages: Taqī’l-Dīn wa’l-handasa’l-mīkānīkiyya’l-’arabiyya, ma’ Kitāb al-Turuq al-saniyya fī ālāt alrūh.āniyya, min al-qarn al-sadis aṡar, (Taqī al-Dīn and Arabic mechanical engineering with the sublime methods of spiritual machines. An Arabic manuscript of the sixteenth century) par Ahmad Yusuf al-H .asan. Revue d’Histoire des Sciences, 31(1):81–82, January 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631934.

Speziali:1978:AOB

[2373] Pierre Speziali. Analyses d’ouvrages: The Italian Renaissance of Mathematics. Studies on Humanists and Mathematicians from Petrarch to Galileo, (coll. ^{≪}Travaux d’Humanisme et Renaissance ^{≫}, n^{o} 145) par Paul Lawrence Rose. Revue d’Histoire des Sciences, 31(1):82–86, January 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631935.

Beaulieu:1978:AOB

[2374] Armand Beaulieu. Analyses d’ouvrages: Lettres à Malherbe (1606– 1628), (Institut de Recherche et d’Histoire des Textes) par Peiresc. Revue d’Histoire des Sciences, 31(1):86–87, January 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631936.

Russo:1978:AOB

[2375] F. Russo. Analyses d’ouvrages: Isaac Newton Filosofo della Natura : Le lezioni giovanili di ottica e la genesi del metodo newtoniano par Maurizio Mamiani. Revue d’Histoire des Sciences, 31(1):87–88, January 1978. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631937.

Rosmorduc:1978:AOBb

[2376] Jean Rosmorduc. Analyses d’ouvrages: Étienne-Louis Malus (1775– 1812) et la théorie corpusculaire de la lumière, (coll. ^{≪}L’Histoire des sciences. Textes et études^{≫}) par André Chappert. Revue d’Histoire des Sciences, 31(1):88–90, January 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631938.

Redondi:1978:AOB

[2377] Pietro Redondi. Analyses d’ouvrages: Sadi Carnot et l’essor de la thermodynamique. Paris, École polytechnique, 11–13 juin 1974. Revue d’Histoire des Sciences, 31(1):90–94, January 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631939.

Dantes:1978:AOB

[2378] Maria Amelia Dantes. Analyses d’ouvrages: Beginnings of Brazilian Science. Oswaldo Cruz, Medical Research and Policy, 1890–1920 par Nancy Stepan. Revue d’Histoire des Sciences, 31(1):95–96, January 1978. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631940.

Broc:1978:QBA

[2379] Numa Broc. De quelques bibliographies anciennes utiles à l’historien de la géographie (XVI^{e}–XVIII^{e} siècles). Revue d’Histoire des Sciences, 31(2): 97–130, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631845.

Laissus:1978:CMP

[2380] Yves Laissus. Catalogue des manuscrits de Philibert Commerson (1727– 1773) conservés à la Bibliothèque centrale du Muséum national d’Histoire naturelle (Paris). Revue d’Histoire des Sciences, 31(2):131–162, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631846.

Fischer:1978:JRO

[2381] Jean-Louis Fischer. Jean Rostand (30 octobre 1894–4 septembre 1977). Revue d’Histoire des Sciences, 31(2):163–172, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631847.

Charmasson:1978:BRE

[2382] Thérèse Charmasson. “L’Arithmétique” de Roland l’Écrivain et le “Quadripartitum numerorum” de Jean de Murs. Revue d’Histoire des Sciences, 31(2):173–176, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631848.

Bernhardt:1978:I

[2383] Jean Bernhardt. Informations. Revue d’Histoire des Sciences, 31(2):177– 179, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631849.

Cohen:1978:AOB

[2384] E. Cohen. Analyses d’ouvrages: Mathematisches Tagebuch, 1796–1814, (^{≪}Ostwalds Klassiker der exakten Wissenschaften ^{≫}, 256) par Carl Friedrich Gauss; E. Schuhmann; K.-R. Biermann; H. Wussing. Revue d’Histoire des Sciences, 31(2):180, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631850.

Pont:1978:AOB

[2385] Jean-Claude Pont. Analyses d’ouvrages: Histoire de la géométrie non euclidienne (Istoria n’eevklidovoy geometrii) par B. A. Rozenfeld. Revue d’Histoire des Sciences, 31(2):180–181, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631851.

Sinaceur:1978:AOB

[2386] M. Sinaceur. Analyses d’ouvrages: Richard Dedekind et les fondements des mathématiques (avec de nombreux textes inédits), (coll. ^{≪}Histoire des Sciences. Textes et études ^{≫}, Coll. des Travaux de l’Académie internationale d’Histoire des Sciences, n^{o} 24) par Pierre Dugac; Jean Dieudonné. Revue d’Histoire des Sciences, 31(2):182–183, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23631852.

Schmid:1978:AOB

[2387] Anne-Françoise Schmid. Analyses d’ouvrages: La filosofia de la matematica de Jules Henri Poincaré par Javier de Lorenzo. Revue d’Histoire des Sciences, 31(2):183–185, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631853.

Debarbat:1978:AOBa

[2388] S. Débarbat. Analyses d’ouvrages: The Gresham Lectures of John Flamsteed par Eric G. Forbes. Revue d’Histoire des Sciences, 31(2):185–186, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631854.

Debarbat:1978:AOBb

[2389] S. Débarbat. Analyses d’ouvrages: Lowell and Mars par William Graves Hoyt. Revue d’Histoire des Sciences, 31(2):186–187, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631855.

Todericiu:1978:AOBa

[2390] Doru Todericiu. Analyses d’ouvrages: Philosophical Chemistry in the Scottish Enlightenment. The Doctrines and Discoveries of William Cullen and Joseph Black par A. L. Donovan. Revue d’Histoire des Sciences, 31(2):187–188, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631856.

Todericiu:1978:AOBb

[2391] Doru Todericiu. Analyses d’ouvrages: Naturwissenschaft, Staat und Industrie in Baden (1848–1914), (coll. ^{≪}Industrielle Welt ^{≫}, Bd 17) par Peter Borscheid. Revue d’Histoire des Sciences, 31(2):189, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23631857.

Specklin:1978:AOB

[2392] Robert Specklin. Analyses d’ouvrages: Die Geographie in der Encyclopédie. Eine wissenschaftsgeschichtliche Studie (Sitzungsberichte der Oesterreichischen Akademie der Wissenschaften, Phil.-Hist. Klasse, 304. Bd, 1. Abh. — Veröffentlichungen der Kommission für Geschichte der Mathematik, Naturwissenschaften und Medizin, 17. Ht) par Johannes Doerflinger. Revue d’Histoire des Sciences, 31(2):189–190, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23631858.

Huard:1978:AOBa

[2393] P. Huard. Analyses d’ouvrages: Bibliography of the History of Medicine, n^{o} 12, 1976. Revue d’Histoire des Sciences, 31(2):190–191, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23631859.

Todericiu:1978:AOBc

[2394] Doru Todericiu. Analyses d’ouvrages: Dejiny Techniky v Ceskosclovensku (Anciennes techniques de Tchécoslovaquie, des origines jusqu’au XVIII^{e} siècle) par Lubos Nový. Revue d’Histoire des Sciences, 31(2): 191–192, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23631860.

Cazenave:1978:AOBa

[2395] M. Cazenave. Analyses d’ouvrages: Catalogo del Fondo Canevari della Biblioteca Berio di Genova, (Publ. del ^{≪}Centro di Studi del Pensiero filosofico del cinquecento e del seicento in relazione ai problemi della scienza ^{≫}, Serie II : Strumenti bibliografici) par Rodolfo Savelli. Revue d’Histoire des Sciences, 31(2):192, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631861.

Cazenave:1978:AOBb

[2396] M. Cazenave. Analyses d’ouvrages: Recherches sur le XVII^{e} siècle, I (Cahiers de l’Équipe de Recherche 75 du CNRS, Centre d’Histoire des Sciences et des Doctrines). Revue d’Histoire des Sciences, 31(2):192, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631862.

Conry:1978:TWO

[2397] Yvette Conry. Thomas Willis ou le premier discours rationaliste en pathologie mentale. Revue d’Histoire des Sciences, 31(3):193–231, juillet 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631888.

Taton:1978:RPB

[2398] René Taton. Repères pour une biographie intellectuelle d’Ampère. Revue d’Histoire des Sciences, 31(3):233–248, juillet 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631889.

Hamamdjian:1978:CAA

[2399] Pierre-Gérard Hamamdjian. Contribution d’Ampère au ^{≪}théorème d’Ampère ^{≫}. Revue d’Histoire des Sciences, 31(3):249–268, juillet 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23631890.

Taton:1978:SFI

[2400] René Taton. Supplément à l’^{≪}Inventaire de l’œuvre de Clairaut (1713– 1765)^{≫}(I) [Rev. Histoire Sci. Appl. 29 (1976), no. 2, 97–122; MR 57 #15882]. Revue d’Histoire des Sciences, 31(3):269–271, juillet 1978. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631891.

Broc:1978:NRB

[2401] Numa Broc. Un nouveau recueil de biographies scientifiques dans le domaine de la géographie. Revue d’Histoire des Sciences, 31(3):272–273, juillet 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631892.

Anonymous:1978:Ib

[2402] Anonyme. Informations. Revue d’Histoire des Sciences, 31(3):274–276, juillet 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631893.

deBuzon:1978:AOB

[2403] Catherine de Buzon. Analyses d’ouvrages: Condorcet. From Natural Philosophy to Social Mathematics par Keith Michael Baker. Revue d’Histoire des Sciences, 31(3):277–280, juillet 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631894.

Rosmorduc:1978:AOBc

[2404] Jean Rosmorduc. Analyses d’ouvrages: De la combinatoire aux probabilités par Pierre Raymond; La théorie des jeux : une politique imaginaire par Michel Plon; Philosophie et calcul de l’infini par Christian Houzel; Jean-Louis Ovaert; Pierre Raymond; Jean-Jacques Sansuc; Matérialisme dialectique et logique, (coll. ^{≪}Algorithme ^{≫}) par Pierre Raymond. Revue d’Histoire des Sciences, 31(3):281–282, juillet 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631895.

Poulle:1978:AOBa

[2405] Emmanuel Poulle. Analyses d’ouvrages: Inventaire des manuscrits scientifiques de la Bibliothèque royale Albert-I^{er}, t. III, (Publications du Centre national d’Histoire des Sciences, I, 3) par Roger Calcoen. Revue d’Histoire des Sciences, 31(3):282–283, juillet 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631896.

Poulle:1978:AOBb

[2406] Emmanuel Poulle. Analyses d’ouvrages: Die mathematischen S. Marcohandschriften in Florenz. Nuova edizione. (^{≪}Quaderni di storia e critica della scienza ^{≫}, n. sér., 8) par Axel Anton Björnbo; Gian Carlo Garfagnini; Eugenio Garin. Revue d’Histoire des Sciences, 31(3):283– 284, juillet 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631897.

Taton:1978:AOB

[2407] R. Taton. Analyses d’ouvrages: Le savant et philosophe mulhousien Jean-Henri Lambert (1728–1777). Études critiques et documentaires, (^{≪}Université du Haut-Rhin, Centre de recherches et d’études rhénanes ≫) par Roger Jaquel. Revue d’Histoire des Sciences, 31(3):284, juillet 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631898.

Rosmorduc:1978:AOBd

[2408] Jean Rosmorduc. Analyses d’ouvrages: Les propriétés remarquables de la Route de la lumière, par les airs et en général par plusieurs milieux réfringens sphériques et concentriques, avec la solution des problèmes qui y ont du rapport … par J.-H. Lambert; P. Radelet; D. Speiser; M. L. Boissy. Revue d’Histoire des Sciences, 31(3):285–286, juillet 1978. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23631899.

Boulle:1978:AOB

[2409] Lydie Boulle. Analyses d’ouvrages: La vie quotidienne du médecin de province au XIX^{e} siècle, (^{≪}La vie quotidienne ^{≫}) par Jacques Léonard. Revue d’Histoire des Sciences, 31(3):286–287, juillet 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631900.

Huard:1978:AOBb

[2410] P. Huard. Analyses d’ouvrages: L’ambulance du D^{r} Alexis Carrel, 1914– 1919 ^{≪}La Concorde ^{≫}par Georgette Mottier. Revue d’Histoire des Sciences, 31(3):287–288, juillet 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631901.

Cazenave:1978:AOBc

[2411] M. Cazenave. Analyses d’ouvrages: Inventaire des papiers de la division des Sciences et Lettres du ministère de l’Instruction publique et des services qui en sont issus (sous-série F^{17}), t. I, par Marie-Elisabeth Antoine; Suzanne Olivier. Revue d’Histoire des Sciences, 31(3):288, juillet 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631902.

Grmek:1978:CBV

[2412] Mirko Drazen Grmek. Contribution à la biographie de Vidius (Guido Guidi) premier lecteur royal de médecine : ses origines et sa vie avant la période parisienne. Revue d’Histoire des Sciences, 31(4):289–299, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632063.

Howard:1978:GBB

[2413] Rio Howard. Guy de la brosse : Botanique et chimie au début de la révolution scientifique. Revue d’Histoire des Sciences, 31(4):301–326, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632064.

Youschkevitch:1978:CFG

[2414] A. P. Youschkevitch. C. F. Gauss et J. A. da Cunha. Revue d’Histoire des Sciences, 31(4):327–332, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632065.

Brouzeng:1978:MEM

[2415] Paul Brouzeng. Magnétisme et énergétique: La méthode de Duhem: À propos d’une lettre inédite de Pierre Curie. Revue d’Histoire des Sciences, 31(4):333–344, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632066.

Godart:1978:TIG

[2416] O. Godart, M. Heller, and Georges Lemaître. Un travail inconnu de Georges Lemaître. Revue d’Histoire des Sciences, 31(4):345–359, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632067.

Dugac:1978:ECG

[2417] Pierre Dugac. Sur une édition de la correspondance de Gauss avec A. von Humboldt. Revue d’Histoire des Sciences, 31(4):361–366, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632068.

Anonymous:1978:Ic

[2418] Anonyme. Informations. Revue d’Histoire des Sciences, 31(4):367–370, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632069.

Picolet:1978:AOB

[2419] Guy Picolet. Analyses d’ouvrages: On Burning Mirrors, (^{≪}Sources in the History of Mathematics and Physical Sciences ^{≫}, 1) par Diocles; G. J. Toomer. Revue d’Histoire des Sciences, 31(4):371–373, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632070.

Hominal:1978:AOB

[2420] François Hominal. Analyses d’ouvrages: Les systèmes d’équations polynômes dans le Siyuan Yujian (1303) (Mém. de l’Institut des Hautes Études chinoises, vol. VI) par John Hoe. Revue d’Histoire des Sciences, 31(4):374–377, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632071.

Azouvi:1978:AOB

[2421] François Azouvi. Analyses d’ouvrages: Mesmer ou la révolution thérapeutique, (coll. ^{≪}Science de l’homme ^{≫}) par Franklin Rausky. Revue d’Histoire des Sciences, 31(4):377–378, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632072.

Cazenave:1978:AOBd

[2422] M. Cazenave. Analyses d’ouvrages: Spinoza (1632–1677), Revue internationale de Philosophie, 31^{e} année, n^{o} 119-120, 1977, face. 1–2. Revue d’Histoire des Sciences, 31(4):378, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632073.

Delorme:1978:AOB

[2423] Suzanne Delorme. Analyses d’ouvrages: Leibniz à Paris, t. I :Les Sciences, t. II : La Philosophie de Leibniz. Symposium de la G. W. Leibniz-Gesellschaft (Hannover) et du Centre national de la Recherche scientifique (Paris) à Chantilly (France) du 14 au 18 novembre 1976, (Studia Leibnitiana, Supplementa, vol. XVII et vol. XVIII.) par Kurt Müller; Heinrich Schepers; Wilhelm Totok. Revue d’Histoire des Sciences, 31(4): 378, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632074.

Anonymous:1978:TA

[2424] Anonyme. Tables alphabétiques. Revue d’Histoire des Sciences, 31 (4):379–387, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632075.

Anonymous:1978:BM

[2425] Anonyme. Back matter. Revue d’Histoire des Sciences, 31(4):??, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633047.

Anonymous:1978:FMa

[2426] Anonyme. Front matter. Revue d’Histoire des Sciences, 31(1):??, January 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631925.

Anonymous:1978:FMb

[2427] Anonyme. Front matter. Revue d’Histoire des Sciences, 31(2):??, avril 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631844.

Anonymous:1978:FMc

[2428] Anonyme. Front matter. Revue d’Histoire des Sciences, 31(3):??, juillet 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631887.

Anonymous:1978:FMd

[2429] Anonyme. Front matter. Revue d’Histoire des Sciences, 31(4):??, October 1978. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632062.

Bachelard:1979:AP

[2430] Suzanne Bachelard. Avant-propos. Revue d’Histoire des Sciences, 32(1): 3, January 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631942.

Balan:1979:OOE

[2431] Bernard Balan. Organisation, organisme, économie et milieu chez Henri Ducrotay de Blainville. Revue d’Histoire des Sciences, 32(1):5–24, January 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631943.

Lessertisseur:1979:LSC

[2432] J. Lessertisseur and F. K. Jouffroy. L’idée de série chez blainville. Revue d’Histoire des Sciences, 32(1):25–42, January 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631944.

Gohau:1979:LCC

[2433] Gabriel Gohau. L’unité de la création chez Blainville. Revue d’Histoire des Sciences, 32(1):43–58, January 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631945.

Gouhier:1979:BAC

[2434] Henri Gouhier. Blainville et Auguste Comte. Revue d’Histoire des Sciences, 32(1):59–72, January 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631946.

Canguilhem:1979:FHS

[2435] Georges Canguilhem. L’ ^{≪}histoire des sciences de l’organisation ^{≫}de Blainville et l’abbé Maupied. Revue d’Histoire des Sciences, 32(1):73–91, January 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631947.

Prevost:1979:AEO

[2436] C. Prévost. Appendice: Extraits de l’ouvrage de Pol Nicard: “Étude sur la Vie et les Travaux de M. Ducrotay de Blainville”. Revue d’Histoire des Sciences, 32(1):93–96, January 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633131.

Federspiel:1979:PBP

[2437] Michel Federspiel. Sur un passage des Definitiones pseudo-Héron d’Alexandrie (Ed. Heiberg, p. 156, l. 1–5). Revue d’Histoire des Sciences, 32(2):97–106, 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

Sinaceur:1979:MMD

[2438] Mohamed A. Sinaceur and Richard Dedekind. La méthode mathématique de Dedekind. Revue d’Histoire des Sciences, 32(2):107–142, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23631632.

Debru:1979:CBI

[2439] Claude Debru. Claude Bernard et l’idée d’une chimie biologique. Revue d’Histoire des Sciences, 32(2):143–162, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631633.

Federspiel:1979:PBD

[2440] Michel Federspiel. Sur un passage des Definitiones du pseudo-Héron d’Alexandrie (Ed. Heiberg, 1. 1–5). Revue d’Histoire des Sciences, 32(2): 156, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631631.

Caveing:1979:PDM

[2441] Maurice Caveing. À propos des débuts des mathématiques grecques: Réflexions sur l’ouvrage de A. Szabó. Revue d’Histoire des Sciences, 32(2):163–168, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631634.

Anonymous:1979:I

[2442] Anonyme. Informations. Revue d’Histoire des Sciences, 32(2):169–173, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631635.

Rosmorduc:1979:AOBa

[2443] Jean Rosmorduc. Analyses d’ouvrages: La naissance de la physique dans le texte de Lucrèce. Fleuves et turbulences, (coll. ^{≪}Critique ^{≫}) par Michel Serres. Revue d’Histoire des Sciences, 32(2):174–176, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631636.

Rosmorduc:1979:AOBb

[2444] Jean Rosmorduc. Analyses d’ouvrages: The Invention of the Telescope, vol. 67, part. 4 par Albert Van Helden. Revue d’Histoire des Sciences, 32(2):176–177, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631637.

Ellenberger:1979:AOB

[2445] François Ellenberger. Analyses d’ouvrages: The Making of Geology : Earth Science in Britain, 1660–1815 par Roy Porter. Revue d’Histoire des Sciences, 32(2):177–182, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23631638.

Todericiu:1979:AOBa

[2446] D. Todericiu. Analyses d’ouvrages: Le chimiste Claude-Louis Berthollet (1748–1822). Sa vie, son œuvre par Michelle Sadoun-Goupil. Revue d’Histoire des Sciences, 32(2):182–186, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23631639.

Grmek:1979:AOBa

[2447] M. D. Grmek. Analyses d’ouvrages: Ontogeny and Phylogeny par Stephen Jay Gould. Revue d’Histoire des Sciences, 32(2):186–187, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631640.

Huard:1979:AOB

[2448] P. Huard. Analyses d’ouvrages: Les maladies et la médecine en Pays Basque Nord à la fin de l’Ancien Régime (1690–1789), thèse de doctorat ès lettres (3^{e} cycle, Histoire de la médecine), soutenue en mai 1977 par Pierre Thillaud. Revue d’Histoire des Sciences, 32(2):187–189, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631641.

Theodorides:1979:AOB

[2449] Jean Théodoridès. Analyses d’ouvrages: Albrecht von Hallers Briefe an Auguste Tissot (1754–1777) par Erich Hintzsche. Revue d’Histoire des Sciences, 32(2):189, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631642.

Grmek:1979:AOBb

[2450] M. D. Grmek. Analyses d’ouvrages: Edward Jenner’s cowpox vaccine : the history of a medical myth par Peter Razzell. Revue d’Histoire des Sciences, 32(2):190, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631643.

Comiti:1979:AOB

[2451] V. P. Comiti. Analyses d’ouvrages: Charles Krafft (1863–1921); un pionnier en Europe de l’appendicectomie et des soins aux malades par Paul Gasser. Revue d’Histoire des Sciences, 32(2):191, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631644.

Halleux:1979:AOBa

[2452] Robert Halleux. Analyses d’ouvrages: Josiah Wedgwood : ^{≪}the Arts and Sciences United ^{≫}. An Exhibition of Josiah Wedgwood Correspondence, Experiment Books and the Ceramic Products he developed and manufactured. Held at the Science Museum, London, 21 March to 24 September 1978. Revue d’Histoire des Sciences, 32(2):192, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23631645.

Ilic:1979:AOBa

[2453] Miriana Ilić. Analyses d’ouvrages: Voprossy istorii astronomii [= Problèmes d’histoire de l’astronomie], recueil n^{o} 3. Revue d’Histoire des Sciences, 32(2):192, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23631646.

Rashed:1979:LDA

[2454] Roshdi Rashed. L’analyse diophantienne au x^{e} siècle : l’exemple d’al-Khāzin. Revue d’Histoire des Sciences, 32(3):193–222, juillet 1979. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632334.

Gillispie:1979:MIO

[2455] Charles C. Gillispie. Mémoires inédits ou anonymes de Laplace sur la théorie des erreurs, les polynômes de Legendre, et la philosophie des probabilités. Revue d’Histoire des Sciences, 32(3):223–279, juillet 1979. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632335.

Theodorides:1979:Ia

[2456] Jean Théodoridès. Informations. Revue d’Histoire des Sciences, 32(3): 281–284, juillet 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632336.

Todericiu:1979:AOBb

[2457] Doru Todericiu. Analyses d’ouvrages: Scientific materialism in Nineteenth Century Germany, (^{≪}Studies in the History of Modern Science ≫, 1) par Frederick Gregory. Revue d’Histoire des Sciences, 32(3):285– 286, juillet 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632337.

Arboleda:1979:AOB

[2458] Luis Carlos Arboleda. Analyses d’ouvrages: Les fondements de l’analyse de Cauchy à Baire par Pierre Dugac. Revue d’Histoire des Sciences, 32(3):286–288, juillet 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632338.

Lagarde:1979:HDP

[2459] Lucie Lagarde. Historique du probléme du Méridien origine en France. Revue d’Histoire des Sciences, 32(4):289–304, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632397.

Jaki:1979:ERT

[2460] Stanley L. Jaki. Sur l’édition et la réédition de la traduction française des Cosmologische Briefe de Lambert. Revue d’Histoire des Sciences, 32(4):305–314, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632398.

Merleau-Ponty:1979:DPE

[2461] Jacques Merleau-Ponty. La découverte des principes de l’énergie : l’itinéraire de Joule. Revue d’Histoire des Sciences, 32(4):315–331, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632399.

Arvy:1979:CDF

[2462] Lucie Arvy and Bernard Heuvelmans. Les cétacés du ^{≪}vis booc ^{≫}(1577– 1578) d’Adrien Coenensoen Van Schilperoort (1514-?). Revue d’Histoire des Sciences, 32(4):333–338, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632400.

Hurwic:1979:CMC

[2463] Józef Hurwic. La contribution de Marie Curie à l’explication de la nature de la transformation radioactive. Revue d’Histoire des Sciences, 32(4):339–343, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632401.

Caveing:1979:JI

[2464] M. Caveing, R. Rashed, and R. Taton. Jean Itard (1902–1979). Revue d’Histoire des Sciences, 32(4):345–350, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632402.

Theodorides:1979:Ib

[2465] Jean Théodoridès. Informations. Revue d’Histoire des Sciences, 32 (4):351–355, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632403.

Grmek:1979:LRP

[2466] M. D. Grmek. Lettre à la rédaction: [À propos de l’article de Claude Debru sur Claude Bernard]. Revue d’Histoire des Sciences, 32(4):356, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632404.

Caveing:1979:AOB

[2467] M. Caveing. Analyses d’ouvrages: Von Eudoxos zu Aristoteles, (^{≪}Studien zur antiken Philosophie ^{≫}, no 8) par Hans Joachim Waschkies. Revue d’Histoire des Sciences, 32(4):357–360, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632405.

deValence:1979:AOB

[2468] M. de Valence. Analyses d’ouvrages: Register zu Gottfried Wilhelm Leibniz. (^{≪}Olms Paperbacks Band ^{≫}, 49) par Joseph Ehrenfried Hofmann; C. I. Gerhardt. Revue d’Histoire des Sciences, 32(4):360–361, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632406.

Ilic:1979:AOBb

[2469] Miriana Ilic. Analyses d’ouvrages: Les mathématiques au XIX^{e} siècle : Logique mathématique; Algèbre; Théorie des nombres; Théorie des probabilités par A. N. Kolmogorov; A. P. Youschkevitch. Revue d’Histoire des Sciences, 32(4):361–362, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632407.

Dugac:1979:AOB

[2470] Pierre Dugac. Analyses d’ouvrages: Reine Zahlenlehre, (Gesamtausgabe, Reihe II, Bd. 2, Grössenlehre II) par Bernard Bolzano; Jan Berg. Revue d’Histoire des Sciences, 32(4):362–366, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632408.

Todericiu:1979:AOBc

[2471] Doru Todericiu. Analyses d’ouvrages: Il Mondo di Carta, Ricerche sulla seconda rivoluzione scientifica par Enrico Bellone. Revue d’Histoire des Sciences, 32(4):366–367, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632409.

Halleux:1979:AOBb

[2472] Robert Halleux. Analyses d’ouvrages: Crucibles. The Story of Chemistry. From Ancient Alchemy to Nuclear Fission, 4^{e} éd. revue par Bernard Jaffe. Revue d’Histoire des Sciences, 32(4):367, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632410.

Halleux:1979:AOBc

[2473] Robert Halleux. Analyses d’ouvrages: Der alchemistische Traktat ^{≪}Von der Multiplikation^{≫}von Pseudo-Thomas von Aquin. Untersuchungen und Texte, (^{≪}Sudhoffs Archiv ^{≫}, Beiheft 19) par Dietlinde Goltz; Joachim Telle; Hans J. Vermeer. Revue d’Histoire des Sciences, 32(4):367–368, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632411.

Halleux:1979:AOBd

[2474] Robert Halleux. Analyses d’ouvrages: The Chemists and the Word. The Didactic Origins of Chemistry par Owen Hannaway. Revue d’Histoire des Sciences, 32(4):368–369, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632412.

Halleux:1979:AOBe

[2475] Robert Halleux. Analyses d’ouvrages: Kemiens Historia i Sverige intill år 1800, (^{≪}Lychnos-Bibliotek ^{≫}, 17, 4) par Hugo Olsson. Revue d’Histoire des Sciences, 32(4):369–370, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632413.

Sadoun-Goupil:1979:AOB

[2476] Michelle Sadoun-Goupil. Analyses d’ouvrages: Classics in Coordination Chemistry, Part. II, Selected Papers (1798–1899), (^{≪}Classics of Science ^{≫}, vol. VII) par George B. Kauffman. Revue d’Histoire des Sciences, 32(4):370–371, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632414.

Howard:1979:AOB

[2477] Rio Howard. Analyses d’ouvrages: The young Darwin and his cultural circle. A Study of influences which helped shape the language and logic of the first drafts of the theory of natural selection, (^{≪}Studies in the history of modern science, vol. 2 ^{≫}) par Edward Manier. Revue d’Histoire des Sciences, 32(4):371–373, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632415.

Boulle:1979:AOB

[2478] Lydie Boulle. Analyses d’ouvrages: Musik und Medizin. Ihre Wechselbeziehungen in Theorie und Praxis von 800 bis 1800, (^{≪}Freiburger– Beiträge zur Wissenschaftsund Universitätsgeschichte ^{≫}, Bd. 2) par Werner Friedrich Kümmel. Revue d’Histoire des Sciences, 32(4):373–375, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632416.

Halleux:1979:AOBf

[2479] Robert Halleux. Analyses d’ouvrages: Salzburger Beiträge zur Paracelsusforschung, 13–18 : Paracelsus, Werk und Wirkung. Festgabe für Kurt Goldammer zum 60. Geburtstag par Sepp Domandl; Paracelsus in Moronia. Bischoffs Halls satirische Darstellung von ^{≪}Bustius Hohenheimius ^{≫}und dessen Collegium par Joseph B. Dallett; Sepp Domandl; Paracelsus-Forschung heute, Bericht über den Jubiläumskongress 1976 und die 25 jährige Geschichte der Gesellschaft par Kurt Goldammer; Anton Dyk; Sepp Domandl; Hohenheims Weg von St Gallen nach Augsburg (1531–1536) par Edwin Rosner; Sepp Domandl; Paracelsus und die Paracelsustradition in Salzburg (1524–1976). Ein Beitrag zur Kultur und Zeitgeschichte par Sepp Domandl; Kranksein und Heilung bei Paracelsus par Heinrich Schipperges; Sepp Domandl. Revue d’Histoire des Sciences, 32(4):375–376, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632417.

Anonymous:1979:TGM

[2480] Anonyme. Table générale des matières du Tome XXXII. Revue d’Histoire des Sciences, 32(4):377–384, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632418.

Anonymous:1979:FMa

[2481] Anonyme. Front matter. Revue d’Histoire des Sciences, 32(1):??, January 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631941.

Anonymous:1979:FMb

[2482] Anonyme. Front matter. Revue d’Histoire des Sciences, 32(2):??, avril 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23631630.

Anonymous:1979:FMc

[2483] Anonyme. Front matter. Revue d’Histoire des Sciences, 32(3):??, juillet 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632333.

Anonymous:1979:FMd

[2484] Anonyme. Front matter. Revue d’Histoire des Sciences, 32(4):??, October 1979. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632396.

Aujac:1980:ZAG

[2485] Germaine Aujac. Le zodiaque dans l’astronomie grecque. Revue d’Histoire des Sciences, 33(1):3–32, January 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632161.

Ellenberger:1980:IEC

[2486] François Ellenberger. De l’influence de l’environnement sur les concepts : l’exemple des théories géodynamiques au XVIII^{e} siècle en France. Revue d’Histoire des Sciences, 33(1):33–68, January 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632162.

Conry:1980:RBD

[2487] Yvette Conry. Robert Boyle et la doctrine cartésienne des animauxmachines. Revue d’Histoire des Sciences, 33(1):69–74, January 1980. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632163.

Theodorides:1980:MF

[2488] Jean Théodoridès. Marcel Florkin (1900–1979). Revue d’Histoire des Sciences, 33(1):75–76, January 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632164.

Costabel:1980:I

[2489] Pierre Costabel. Informations. Revue d’Histoire des Sciences, 33(1):77– 82, January 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632165.

Beaulieu:1980:AOBa

[2490] Armand Beaulieu. Analyses d’ouvrages: Le Opere dei discepoli di Galileo Galilei. Vol. Primo. Carteggio (1642–1648) par Paolo Galluzzi; Maurizio Torrini. Revue d’Histoire des Sciences, 33(1):83–85, January 1980. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632166.

Halbronn:1980:AOB

[2491] J. Halbronn. Analyses d’ouvrages: L’entrée de Saturne au Lion. L’Eclipse de Soleil du 12 août 1654, (^{≪}Archives internationales d’Histoire des Idées ≫, Series minor, 14) par Elisabeth Labrousse. Revue d’Histoire des Sciences, 33(1):85–87, January 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632167.

Youschkevitch:1980:AOBa

[2492] A. Youschkevitch. Analyses d’ouvrages: The mathematical Papers of Isaac Newton, vol. VII : 1691–1695 par D. T. Whiteside. Revue d’Histoire des Sciences, 33(1):87–91, January 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632168.

Youschkevitch:1980:AOBb

[2493] A. Youschkevitch. Analyses d’ouvrages: The Correspondence of Isaac Newton, vol. V : 1709–1713 par A. Rupert Hall; Laura Tilling. Revue d’Histoire des Sciences, 33(1):91–93, January 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632169.

Pueyo:1980:AOBa

[2494] Guy Pueyo. Analyses d’ouvrages: Les travaux et les jours à Rome et dans l’ancienne France. Les agronomes latins inspirateurs d’Olivier de Serres par Martine Gorrichon. Revue d’Histoire des Sciences, 33(1):93– 94, January 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632170.

deValence:1980:AOBa

[2495] M. de Valence. Analyses d’ouvrages: Historical Studies in the Physical Sciences, vol. 8, vol. 4–9 par Russell McCormmach. Revue d’Histoire des Sciences, 33(1):94–95, January 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632171.

Anonymous:1980:PR

[2496] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 33(1):96, January 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632172.

Conry:1980:LFM

[2497] Yvette Conry. L’idée d’une ^{≪}marche de la nature ^{≫}dans la biologie prédarwinienne au XIX^{e} siècle. Revue d’Histoire des Sciences, 33(2):97– 149, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632279.

Gaudant:1980:LAF

[2498] Jean Gaudant. Louis Agassiz (1807–1873), fondateur de la paléoichthyologie. Revue d’Histoire des Sciences, 33(2):151–162, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632280.

Theodorides:1980:PDT

[2499] Jean Théodoridès and Jean-Jacques Rousset. À propos de la découverte de Trichomonas tenax, protozoaire parasite de l’homme, par O. F.

Müller. Revue d’Histoire des Sciences, 33(2):163–164, avril 1980. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632281.

Anonymous:1980:Ia

[2500] Anonyme. Informations. Revue d’Histoire des Sciences, 33(2):165–166, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632282.

Picolet:1980:AOBa

[2501] Guy Picolet. Analyses d’ouvrages: Nicholas Copernicus, Quincentenary Celebrations Final Report, (^{≪}Studia Copernicana ^{≫}, XVII). Revue d’Histoire des Sciences, 33(2):167–168, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632283.

Giard:1980:AOB

[2502] Luce Giard. Analyses d’ouvrages: Règles utiles et claires pour la direction de l’esprit en la recherche de la vérité par René Descartes; Jean-Luc Marion; Pierre Costabel; La Haye; Martinus Nijhoff. Revue d’Histoire des Sciences, 33(2):168–170, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632284.

Akagi:1980:AOB

[2503] S. Akagi. Analyses d’ouvrages: Œuvres complètes, par Cyrano de Bergerac; Jacques Prévot. Revue d’Histoire des Sciences, 33(2):170–172, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632285.

Blay:1980:AOBa

[2504] Michel Blay. Analyses d’ouvrages: Roemer et la vitesse de la lumière. (coll. ^{≪}Histoire des Sciences. Textes et Études ^{≫}). Revue d’Histoire des Sciences, 33(2):172–175, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632286.

Blay:1980:AOBb

[2505] Michel Blay. Analyses d’ouvrages: Newton and Newtoniana, 1672–1975, A Bibliography par Peter Wallis; Ruth Wallis. Revue d’Histoire des Sciences, 33(2):176, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632287.

Fubini:1980:AOB

[2506] Silvia Fubini. Analyses d’ouvrages: A History of numerical Analysis, from the 16th through the 19th Century, (^{≪}Studies in the History of Mathematics and Physical Sciences ^{≫}, 2) par Herman H. Goldstine. Revue d’Histoire des Sciences, 33(2):176–178, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632288.

Cohen:1980:AOBa

[2507] E. Cohen. Analyses d’ouvrages: Panorama des mathématiques pures. Le choix bourbachique, (coll. ^{≪}Discours de la Méthode ^{≫}) par Jean Dieudonné. Revue d’Histoire des Sciences, 33(2):178–179, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632289.

Laurent:1980:AOB

[2508] Roger Laurent. Analyses d’ouvrages: La Perspective affranchie de l’embarras du plan géométral par J.-H. Lambert; H. Pfeiffer. Revue d’Histoire des Sciences, 33(2):179–181, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632290.

Brouzeng:1980:AOB

[2509] Paul Brouzeng. Analyses d’ouvrages: Réflexions sur la puissance motrice du feu par Sadi Carnot; Robert Fox. Revue d’Histoire des Sciences, 33(2): 181–183, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632291.

Hurwic:1980:AOBa

[2510] Józef Hurwic. Analyses d’ouvrages: Frederick Soddy, 1877–1956, Moskva, (coll. ^{≪}Nautchnobiografitcheskaïa seria ^{≫}) par A. N. Krivomazov. Revue d’Histoire des Sciences, 33(2):183–186, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632292.

Leroy:1980:AOB

[2511] Jean-F. Leroy. Analyses d’ouvrages: The Spirit of System : Lamarck and evolutionary Biology par Richard W. Burkhardt. Revue d’Histoire des Sciences, 33(2):186–188, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632293.

Comiti:1980:AOB

[2512] Vincent-Pierre Comiti. Analyses d’ouvrages: La réanimation respiratoire au XVIII^{e} siècle par Philippe Brullard. Revue d’Histoire des Sciences, 33(2):188–190, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632294.

Theodorides:1980:AOB

[2513] Jean Théodoridès. Analyses d’ouvrages: Die zoologische Buchillustration. Ihre Bibliographie und Geschichte, Band II : Geschichte par Claus Nissen. Revue d’Histoire des Sciences, 33(2):190–191, avril 1980. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632295.

Pueyo:1980:AOBb

[2514] Guy Pueyo. Analyses d’ouvrages: Histoire de la France rurale par Georges Duby; Armand Wallon. Revue d’Histoire des Sciences, 33(2): 191–192, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632296.

LHuillier:1980:RBN

[2515] Ghislaine L’Huillier. Regiomontanus et le “Quadripartitum numerorum” de Jean de Murs. Revue d’Histoire des Sciences, 33(3):193–214, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632237.

Blay:1980:CDO

[2516] Michel Blay. Une clarification dans le domaine de l’optique physique : “bigness” et promptitude. Revue d’Histoire des Sciences, 33(3):215–224, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632238.

Fischer:1980:LSP

[2517] Jean-Louis Fischer. L’aspect social et politique des relations épistolaires entre quelques savants français et la station zoologique de Naples de 1878 à 1912. Revue d’Histoire des Sciences, 33(3):225–251, juillet 1980. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632239.

Taton:1980:VB

[2518] René Taton. Viggo Brun (1885–1978). Revue d’Histoire des Sciences, 33(3):253–254, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632240.

Taton:1980:JOF

[2519] René Taton. Joachim Otto Fleckenstein (1914–1980). Revue d’Histoire des Sciences, 33(3):255–258, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632241.

Anonymous:1980:Ib

[2520] Anonyme. Informations. Revue d’Histoire des Sciences, 33(3):259–260, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632242.

deBeauregard:1980:AOB

[2521] O. Costa de Beauregard. Analyses d’ouvrages: The study of Time III par J. T. Fraser; N. Lawrence; D. Park. Revue d’Histoire des Sciences, 33(3): 261, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632243.

Guillerme:1980:AOB

[2522] J. Guillerme. Analyses d’ouvrages: Research in Philosophy and Technology (vol. I) par P. T. Durbin; C. Mitcham; The Dynamics of Science and Technology par W. Krohn; E. T. Layton; P. Weingart. Revue d’Histoire des Sciences, 33(3):262–263, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632244.

Murphy:1980:AOBa

[2523] Terence Murphy. Analyses d’ouvrages: Science in Culture. The Early Victorian Period par Susan Faye Cannon. Revue d’Histoire des Sciences, 33(3):264, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632245.

Beaulieu:1980:AOBb

[2524] A. Beaulieu. Analyses d’ouvrages: The Heritage of Copernicus : Theories ^{≪}More Pleasing to the Mind ^{≫} par Jerzy Neyman. Revue d’Histoire des Sciences, 33(3):264–265, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632246.

Poulle:1980:AOB

[2525] Emmanuel Poulle. Analyses d’ouvrages: Science and History, Studies in Honor of Edward Rosen, (^{≪}Studia Copernicana ^{≫}, XVI) par P. Czartoryski; E. Hilfstein. Revue d’Histoire des Sciences, 33(3):265–267, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632247.

Bernhardt:1980:AOBa

[2526] Jean Bernhardt. Analyses d’ouvrages: Thomas Harriot. Renaissance Scientist par John W. Shirley. Revue d’Histoire des Sciences, 33(3):267– 270, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632248.

Bernhardt:1980:AOBb

[2527] Jean Bernhardt. Analyses d’ouvrages: Correspondance du P. Marin Mersenne, Religieux Minime, XIII : 1644–1645 par P. Tannery; C. De Waard; A. Beaulieu. Revue d’Histoire des Sciences, 33(3):270–272, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632249.

Picolet:1980:AOBb

[2528] Guy Picolet. Analyses d’ouvrages: Isaac Newton’s Papers and Letters on Natural Philosophy and related documents, second edition par I. Bernard Cohen; Robert E. Schofield; Marie Boas; Charles Coulston Gillispie; Thomas S. Kuhn; Perry Miller. Revue d’Histoire des Sciences, 33(3):272– 274, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632250.

Blay:1980:AOBc

[2529] Michel Blay. Analyses d’ouvrages: Unpublished Scientific Papers of Isaac Newton. A selection from the Portsmouth Collection in the University Library par A. Rupert Hall; Marie Boas Hall. Revue d’Histoire des Sciences, 33(3):274–275, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632251.

Blay:1980:AOBd

[2530] Michel Blay. Analyses d’ouvrages: The Library of Isaac Newton par John Harrison. Revue d’Histoire des Sciences, 33(3):275–276, juillet 1980. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632252.

Cohen:1980:AOBb

[2531] E. Cohen. Analyses d’ouvrages: Le roman d’Évariste Galois par Leopold Infeld; J. Sully. Revue d’Histoire des Sciences, 33(3):276–277, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632253.

Cazenobe:1980:AOB

[2532] J. Cazenobe. Analyses d’ouvrages: Einstein et le conflit des générations, ≪Complexe ^{≫}, (coll. ^{≪}de la Science ^{≫}) par Lewis S. Feuer; S. Moscovici; P. Alexandre. Revue d’Histoire des Sciences, 33(3):277–278, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632254.

Halleux:1980:AOB

[2533] Robert Halleux. Analyses d’ouvrages: A l’aube de la géologie moderne : Henri Gautier (1660–1737), (^{≪}Histoire et Nature ^{≫}, no 7, 1975, et 9–10, 1976–1977) par François Ellenberger. Revue d’Histoire des Sciences, 33 (3):279, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632255.

Howard:1980:AOB

[2534] Rio Howard. Analyses d’ouvrages: Conrad Gesner’s ^{≪}Historia Animalium ^{≫}. An Inventory of Renaissance Zoology, (^{≪}Communicationes Biohistoricae Ultrajectinae ^{≫}, 72) par Carolin Aleid Gmelig-Nijboer. Revue d’Histoire des Sciences, 33(3):280, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632256.

Huard:1980:AOB

[2535] Pierre Huard and Marie-José Imbault-Huart. Analyses d’ouvrages: Grundzüge der Krankenhausgeschichte (1800–1900) par Dieter Jetter. Revue d’Histoire des Sciences, 33(3):280–282, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632257.

Boulle:1980:AOB

[2536] Lydie Boulle. Analyses d’ouvrages: Die Wiener medizinische Schule im 19. Jahrhundert, (^{≪}Studien zur Geschichte der Universität Wien ^{≫}, Band VI) par Erna Lesky. Revue d’Histoire des Sciences, 33(3):283–285, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632258.

Murphy:1980:AOBb

[2537] Terence Murphy. Analyses d’ouvrages: Cholera, Fever and English Medicine 1825–1865, (^{≪}Oxford Historical Monographs ^{≫}) par Margaret Pelling. Revue d’Histoire des Sciences, 33(3):285–286, juillet 1980. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632259.

Mandelbaum:1980:AOB

[2538] Jonathan Mandelbaum. Analyses d’ouvrages: Aux origines de l’anthropologie française. Les Mémoires de la Société des Observateurs de l’Homme en l’an VIII par Jean Copans; Jean Jamin; Jean-Paul Faivre. Revue d’Histoire des Sciences, 33(3):286–287, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632260.

Laissus:1980:AOB

[2539] Yves Laissus. Analyses d’ouvrages: The Work of Jacques Le Moyne de Morgues, a Huguenot artist in France, Florida and England … par Paul Hulton; D. B. Quinn; R. A. Skelton; William C. Sturtevant; William L. Stearn. Revue d’Histoire des Sciences, 33(3):288, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632261.

Langins:1980:POE

[2540] Janis Langins. Sur la première organisation de l’École polytechnique. Texte de l’arrêté du 6 frimaire an III. Revue d’Histoire des Sciences, 33(4):289–313, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632264.

Dhombres:1980:LMP

[2541] Jean Dhombres. L’enseignement des mathématiques par la ^{≪}méthode révolutionnaire ^{≫}. les leçons de Laplace à l’École Normale de l’an III. Revue d’Histoire des Sciences, 33(4):315–348, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632265.

Taton:1980:AB

[2542] René Taton. Albert Blanchard (1891–1980). Revue d’Histoire des Sciences, 33(4):349–350, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632266.

Theodorides:1980:I

[2543] Jean Théodoridès and François Ellenberger. Informations. Revue d’Histoire des Sciences, 33(4):351–360, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632267.

Caveing:1980:AOB

[2544] Maurice Caveing. Analyses d’ouvrages: Les Eléments par Euclide; George J. Kayas; Georges J. Kayas. Revue d’Histoire des Sciences, 33(4):361–364, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632268.

Rashed:1980:AOB

[2545] Roshdi Rashed. Analyses d’ouvrages: Geschichte des Arabischen Schrifttums, Bd. VI, Astronomie par Fuat Sezgin; E. J. Brill. Revue d’Histoire des Sciences, 33(4):364–366, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632269.

Debarbat:1980:AOB

[2546] Suzzane Débarbat. Analyses d’ouvrages: Planets and Planetarians par Stanley L. Jaki. Revue d’Histoire des Sciences, 33(4):366–367, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632270.

Beaulieu:1980:AOBc

[2547] Armand Beaulieu. Analyses d’ouvrages: Origins in Acoustics. The Science of sound from Antiquity to the Age of Newton par Frederick Vinton Hunt; R. E. Apfel; R. E. Appel. Revue d’Histoire des Sciences, 33(4):367– 368, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632271.

Hurwic:1980:AOBb

[2548] Józef Hurwic. Analyses d’ouvrages: Essai d’histoire de la chimie organique [en russe : Ocerki po istorii organiceskoj himii]. ^{≪}Nauka ^{≫}. Revue d’Histoire des Sciences, 33(4):368–372, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632272.

Sadoun-Goupil:1980:AOB

[2549] Michelle Sadoun-Goupil. Analyses d’ouvrages: Gay-Lussac Scientist and Bourgeois par Maurice P. Crosland. Revue d’Histoire des Sciences, 33(4):372–375, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632273.

Grmek:1980:AOB

[2550] M. D. Grmek. Analyses d’ouvrages: The Rise of Surgery. From Empiric Craft to Scientific Discipline par Owen H. Wangensteen; Sarah D. Wangensteen. Revue d’Histoire des Sciences, 33(4):375–376, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632274.

deValence:1980:AOBb

[2551] M. de Valence. Analyses d’ouvrages: History and Philosophy of the Life Sciences (^{≪}Pubblicazioni della Stazione Zoologica di Napoli ^{≫}, Section II) par M. D. Grmek; Leo S. Olschki Firenze; B. Fantini. Revue d’Histoire des Sciences, 33(4):376, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632275.

Anonymous:1980:TGM

[2552] Anonyme. Table générale des matières du Tome XXXIII. Revue d’Histoire des Sciences, 33(4):377–384, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632276.

Anonymous:1980:BMa

[2553] Anonyme. Back matter. Revue d’Histoire des Sciences, 33(1):??, January 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632173.

Anonymous:1980:BMb

[2554] Anonyme. Back matter. Revue d’Histoire des Sciences, 33(2):??, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632297.

Anonymous:1980:BMc

[2555] Anonyme. Back matter. Revue d’Histoire des Sciences, 33(3):??, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632262.

Anonymous:1980:BMd

[2556] Anonyme. Back matter. Revue d’Histoire des Sciences, 33(4):??, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632858.

Anonymous:1980:FMa

[2557] Anonyme. Front matter. Revue d’Histoire des Sciences, 33(1):??, January 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632160.

Anonymous:1980:FMb

[2558] Anonyme. Front matter. Revue d’Histoire des Sciences, 33(2):??, avril 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632278.

Anonymous:1980:FMc

[2559] Anonyme. Front matter. Revue d’Histoire des Sciences, 33(3):??, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632236.

Anonymous:1980:FMd

[2560] Anonyme. Front matter. Revue d’Histoire des Sciences, 33(4):??, October 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632263.

Kerszberg:1981:CCO

[2561] Pierre Kerszberg. La cosmologie de Copernic et les origines de la physique mathématique. Revue d’Histoire des Sciences, 34(1):3–23, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632299.

Aymonin:1981:LL

[2562] Gérard G. Aymonin. L’herbier de Lamarck. Revue d’Histoire des Sciences, 34(1):25–58, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632300.

Dulieu:1981:PM

[2563] Louis Dulieu. Prunelle à Montpellier. Revue d’Histoire des Sciences, 34(1):59–69, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632301.

Beaulieu:1981:AOB

[2564] A. Beaulieu. Analyses d’ouvrages: Robert Fludd and His Philosophicall Key, being a Transcription of the Manuscript at Trinity College, Cambridge par Allen G. Debus. Revue d’Histoire des Sciences, 34(1):71–72, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632302.

Galles:1981:AOB

[2565] Carlos Delfino Galles. Analyses d’ouvrages: Selected Papers of Léon Rosenfeld, (^{≪}Boston Studies in the Philosophy of Science ^{≫}, XXI) par R. S. Cohen; J. J. Stachel. Revue d’Histoire des Sciences, 34(1):72–74, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632303.

deValence:1981:AOBa

[2566] M. de Valence. Analyses d’ouvrages: Arithmos-Arrythmos : Skizzen aus der Wissenschaftsgeschichte (^{≪}Wissenschaftsgeschichte. Beiträge aus dem Forschungsinstitut des Deutschen Museums für die Geschichte der Naturwissenschaften und der Technik ^{≫}) par Karin Figala; Ernst H. Berninger. Revue d’Histoire des Sciences, 34(1):74–75, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632304.

Speziali:1981:AOB

[2567] Pierre Speziali. Analyses d’ouvrages: Ludovico Ferrari e Niccolò Tartaglia. Cartelli di sfida matematica par Arnaldo Masotti. Revue d’Histoire des Sciences, 34(1):75–78, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632305.

Dugac:1981:AOBa

[2568] Pierre Dugac. Analyses d’ouvrages: Cauchyjev racun ostataka sa primenama [ = Calcul des résidus de Cauchy, avec applications] par Dragoslav S. Mitrinović; Jovan D. Keckić. Revue d’Histoire des Sciences, 34(1):79– 80, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632306.

Debarbat:1981:AOB

[2569] Suzanne Débarbat. Analyses d’ouvrages: Planets X and Pluto par William Graves Hoyt. Revue d’Histoire des Sciences, 34(1):80–81, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632307.

Khawam:1981:AOB

[2570] René R. Khawam. Analyses d’ouvrages: Kitâb al-hiyal, The Book of ingenious devices [le livre des dispositifs ingénieux] par Banû-Mûsa; Donald R. Hill. Revue d’Histoire des Sciences, 34(1):81, January 1981. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632308.

Rosmorduc:1981:AOBa

[2571] Jean Rosmorduc. Analyses d’ouvrages: Marie Curie. Derrière la légende par Robert Reid. Revue d’Histoire des Sciences, 34(1):82–83, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632309.

Sadoun-Goupil:1981:AOB

[2572] Michelle Sadoun-Goupil. Analyses d’ouvrages: Historical Studies in the Language of Chemistry par Maurice P. Crosland. Revue d’Histoire des Sciences, 34(1):83, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632310.

Fischer:1981:AOBa

[2573] Jean-Louis Fischer. Analyses d’ouvrages: Recherches nouvelles sur quelques écrivains des Lumières (II) par Jacques Proust. Revue d’Histoire des Sciences, 34(1):83–84, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632311.

Grmek:1981:AOBa

[2574] M. D. Grmek. Analyses d’ouvrages: Correspondance par Théophile de Bordeu; Martha Fletcher; Nadine Labbé; Aloys de Laforcade. Revue d’Histoire des Sciences, 34(1):84–85, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632312.

Fischer:1981:AOBb

[2575] Jean-Louis Fischer. Analyses d’ouvrages: Un zoologiste de l’époque romantique, Jean-Victor Audouin (1797–1841), (^{≪}Mémoires de la Section des Sciences ^{≫}, 6) par Jean Théodoridès; P.-P. Grassé. Revue d’Histoire des Sciences, 34(1):85–87, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632313.

Fischer:1981:AOBc

[2576] Jean-Louis Fischer. Analyses d’ouvrages: Notes pour le rapport sur les progrès de la physiologie, (^{≪}Documents inédits du Collège de France ^{≫},

2) par Claude Bernard; M. D. Grmek; J. Dausset. Revue d’Histoire des Sciences, 34(1):87–89, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632314.

Grmek:1981:AOBb

[2577] M. D. Grmek. Analyses d’ouvrages: Les médecins de l’Ouest au XIX^{e} siècle par Jacques Léonard. Revue d’Histoire des Sciences, 34(1):89, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632315.

Grmek:1981:AOBc

[2578] M. D. Grmek. Analyses d’ouvrages: Infectious Diseases. Prevention and Treatment in the Nineteenth and Twentieth Centuries par Wesley W. Spink. Revue d’Histoire des Sciences, 34(1):90, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632316.

Huard:1981:AOB

[2579] Pierre Huard. Analyses d’ouvrages: Constantin von Economo, the Man and the Scientist par L. Van Bogaert; J. Théodoridès. Revue d’Histoire des Sciences, 34(1):91–93, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632317.

Theodorides:1981:AOB

[2580] Jean Théodoridès. Analyses d’ouvrages: Chez nous, le Lynx? Mythes et réalité par C. Kempf; A. Balestreri; U. Wotschikowsky; M. Fernex; P. Pfeffer. Revue d’Histoire des Sciences, 34(1):93–94, January 1981. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632318.

Grmek:1981:AOBd

[2581] M. D. Grmek. Analyses d’ouvrages: A Short Title Catalogue of Eighteenth Century Printed Books in the National Library of Medicine par John B. Blake. Revue d’Histoire des Sciences, 34(1):94–95, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632319.

Halleux:1981:AOB

[2582] Robert Halleux. Analyses d’ouvrages: Les sciences exactes et naturelles à l’Université de Louvain de 1835 à 1940. Colloque d’Histoire des Sciences, III (Louvain-la-Neuve, 17 mars 1977), (Université de Louvain, Recueil de Travaux d’Histoire et de Philologie, 6^{e} série, fasc. 15). Revue d’Histoire des Sciences, 34(1):95, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632320.

Pueyo:1981:AOB

[2583] Guy Pueyo. Analyses d’ouvrages: IV^{e} Centenaire du détournement de l’Adour, 1578–1978, Actes du Congrès de Bayonne, 28–29 octobre 1978, Numéro spécial du Bulletin de la Société des Sciences, Lettres et Arts de Bayonne, nouv. série, n^{o} 134. Revue d’Histoire des Sciences, 34(1):95–96, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632321.

Blay:1981:EEM

[2584] Michel Blay. Un exemple d’explication mécaniste au XVII_{e} siècle : l’unité des théories hookiennes de la couleur. Revue d’Histoire des Sciences, 34(2):97–121, avril 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632611.

DiTrocchio:1981:PJB

[2585] Federico Di Trocchio. Paul-Joseph Barthez et l’ “Encyclopédie”. Revue d’Histoire des Sciences, 34(2):123–136, avril 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632612.

Chappert:1981:AST

[2586] André Chappert and Thomas Young. L’ “autobiographical sketch” de Thomas Young. Revue d’Histoire des Sciences, 34(2):137–147, avril 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632613.

Taylor:1981:QLL

[2587] Angus E. Taylor, Pierre Dugac, and H. Lebesgue. Quatre lettres de Lebesgue à Fréchet. Revue d’Histoire des Sciences, 34(2):149–169, avril 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632614.

Daumas:1981:RTP

[2588] Maurice Daumas. Àla recherche de la technologie: À propos d’un ouvrage de Jean-Claude Beaune. Revue d’Histoire des Sciences, 34(2):171–176, avril 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632615.

Poulle:1981:BG

[2589] Emmanuel Poulle. Bertrand Gille (1920–1980). Revue d’Histoire des Sciences, 34(2):177–178, avril 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632616.

Theodorides:1981:I

[2590] Jean Théodoridès. Informations. Revue d’Histoire des Sciences, 34(2): 179–186, avril 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632617.

Grmek:1981:AOBe

[2591] Mirko D. Grmek. Analyses d’ouvrages: Œuvres complètes par Hippocrate; Emile Littré; Pierre Theil. Revue d’Histoire des Sciences, 34(2): 187–188, avril 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632618.

Todericiu:1981:AOBa

[2592] Doru Todericiu. Analyses d’ouvrages: Les textes alchimiques (^{≪}Typologie des sources du Moyen Âge occidental ^{≫}, Fasc. 32) par Robert Halleux. Revue d’Histoire des Sciences, 34(2):188–190, avril 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632619.

deValence:1981:AQPa

[2593] M. de Valence. Analyses de quelques périodiques: Les Cahiers de Fontenay (ENS), no 18, ^{≪}de Carnot à Prigogine ^{≫}. Revue d’Histoire des Sciences, 34(2):190, avril 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632620.

Anonymous:1981:OR

[2594] Anonyme. Ouvrages reçus. Revue d’Histoire des Sciences, 34(2):191– 192, avril 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632621.

Langins:1981:LIF

[2595] Janis Langins and J.-B. Fourier. Une lettre inédite de Fourier sur l’enseignement destiné aux ingénieurs en 1797. Revue d’Histoire des Sciences, 34(3–4):193–207, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632457.

Youschkevitch:1981:OFM

[2596] A. P. Youschkevitch. Sur les origines de la ^{≪}méthode de Cauchy– Lipschitz ^{≫}dans la théorie des équations différentielles ordinaires. Revue d’Histoire des Sciences, 34(3–4):209–215, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632458.

Ellenberger:1981:ASP

[2597] François Ellenberger and Gabriel Gohau. A l’aurore de la stratigraphie paléontologique : Jean–André De Luc, son influence sur Cuvier. Revue d’Histoire des Sciences, 34(3–4):217–257, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632459.

Laissus:1981:VND

[2598] Yves Laissus. Les voyageurs naturalistes du Jardin du roi et du Muséum d’histoire naturelle : essai de portrait-robot. Revue d’Histoire des Sciences, 34(3–4):259–317, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632460.

Broc:1981:GMS

[2599] Numa Broc. Les grandes missions scientifiques françaises au XIXe siècle (Morée, Algérie, Mexique) et leurs travaux géographiques. Revue d’Histoire des Sciences, 34(3–4):319–358, juillet–octobre 1981. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632461.

Hurwic:1981:DDB

[2600] Józef Hurwic. La découverte du brevium (protactinium 234) à partir de la loi des déplacements radioactifs. Revue d’Histoire des Sciences, 34(3– 4):359–361, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632462.

Beaujouan:1981:CB

[2601] Guy Beaujouan. Le Congrès de Bucarest. Revue d’Histoire des Sciences, 34(3–4):362–366, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632463.

Anonymous:1981:I

[2602] Anonyme. Informations. Revue d’Histoire des Sciences, 34(3–4):367–368, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632464.

Hugonnard-Roche:1981:AOB

[2603] Henri Hugonnard-Roche. Analyses d’ouvrages: Theorien und Ansätze in der griechischen Astronomie — im Kontext benachbarter Wissenschaften betrachtet, (^{≪}Coll. des travaux de l’Académie internationale d’Histoire des Sciences ^{≫}, n^{o} 23) par Walter G. Saltzer. Revue d’Histoire des Sciences, 34(3–4):369–370, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632465.

Bernhardt:1981:AOBa

[2604] Jean Bernhardt. Analyses d’ouvrages: Colloquia Copernicana III, Astronomy of Copernicus and its Background, Torun, 1973, (^{≪}Studia Copernicana ^{≫}, XIII). Revue d’Histoire des Sciences, 34(3–4):370–371, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632466.

Bernhardt:1981:AOBb

[2605] Jean Bernhardt. Analyses d’ouvrages: Colloquia Copernicana IV, L’audience de la théorie héliocentrique; Copernic et le développement des sciences …, Torun, 1973, (^{≪}Studia Copernicana ^{≫}, XIV). Revue d’Histoire des Sciences, 34(3–4):371–373, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632467.

Bernhardt:1981:AOBc

[2606] Jean Bernhardt. Analyses d’ouvrages: Il Pitagorismo di Niccolò Copernico, (Accademia Polacca delle Scienze, Biblioteca e Centro di Studi a Roma) par Bronislaw Bilinski. Revue d’Histoire des Sciences, 34 (3–4):373, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632468.

Todericiu:1981:AOBb

[2607] Doru Todericiu. Analyses d’ouvrages: Whilhelm Schickard (1592– 1635), Astronom, Geograph, Orientalist, Erfinder der Rechenmaschine par Friedrich Seck. Revue d’Histoire des Sciences, 34(3–4):374, juillet– octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632469.

Dugac:1981:AOBb

[2608] Pierre Dugac. Analyses d’ouvrages: A Comprehensive Textbook of Classical Mathematics. A Contemporary Interpretation par H. B. Griffiths; P. J. Hilton. Revue d’Histoire des Sciences, 34(3–4):374–375, juillet– octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632470.

Cohen:1981:AOBa

[2609] Eliott Cohen. Analyses d’ouvrages: Œuvres par Paul Lévy; Daniel Dugué; Paul Deheuvels; Michel Ibéro. Revue d’Histoire des Sciences, 34(3–4):375, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632471.

Cohen:1981:AOBb

[2610] Eliott Cohen. Analyses d’ouvrages: A Russian Childhood par Sofya Kovalevskaya; Sofya Kowalevskaya; Beatrice Stillman; P. Y. Polubarinova-Kochina. Revue d’Histoire des Sciences, 34(3–4):376, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632472.

Rosmorduc:1981:AOBb

[2611] Jean Rosmorduc. Analyses d’ouvrages: Einstein, le Livre du Centenaire par A. P. French; G. Delacote; J. Souchon-Rouyer; J.-B. Yelnick; D. Lederer; F. Balibar; J. Molgo; M. Biezunski. Revue d’Histoire des Sciences, 34(3–4):376–377, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632473.

Todericiu:1981:AOBc

[2612] Doru Todericiu. Analyses d’ouvrages: Operationalismus, Kritik und Geschichte einer Philosophie der exakten Wissenchaften [L’^{≪}Opérationalisme ^{≫}, Critique et histoire d’une philosophie des Sciences exactes], (coll. ^{≪}Problemata ^{≫}, 2) par Jürgen Klüver. Revue d’Histoire des Sciences, 34(3–4):377–378, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632474.

deValence:1981:AOBb

[2613] M. de Valence. Analyses d’ouvrages: A. Cournot Études pour le Centenaire de sa mort (1877–1977). (^{≪}Recherches Panthéon–Sorbonne, Univ. de Paris I, Série : Sciences économiques ^{≫}) par Jean Brun; André Robinet. Revue d’Histoire des Sciences, 34(3–4):378, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632475.

deValence:1981:AQPb

[2614] M. de Valence. Analyses de quelques périodiques: Bullisiana, n^{o} 1, ^{≪}B ≫par E. L. Zilberbogen-Chapdelaine. Revue d’Histoire des Sciences, 34 (3–4):378, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632476.

Anonymous:1981:TGM

[2615] Anonyme. Table générale des matières du Tome XXXIV. Revue d’Histoire des Sciences, 34(3–4):379–384, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632477.

Anonymous:1981:BMa

[2616] Anonyme. Back matter. Revue d’Histoire des Sciences, 34(1):??, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632322.

Anonymous:1981:BMb

[2617] Anonyme. Back matter. Revue d’Histoire des Sciences, 34(2):??, avril 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632622.

Anonymous:1981:BMc

[2618] Anonyme. Back matter. Revue d’Histoire des Sciences, 34(3–4):??, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632478.

Anonymous:1981:FMa

[2619] Anonyme. Front matter. Revue d’Histoire des Sciences, 34(1):??, January 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632298.

Anonymous:1981:FMb

[2620] Anonyme. Front matter. Revue d’Histoire des Sciences, 34(2):??, avril 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632610.

Anonymous:1981:FMc

[2621] Anonyme. Front matter. Revue d’Histoire des Sciences, 34(3–4):??, juillet–octobre 1981. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632456.

Demidov:1982:CDT

[2622] Serghei S. Demidov. Création et développement de la théorie des équations différentielles aux dérivées partielles dans les travaux de J. d’Alembert. Revue d’Histoire des Sciences, 35(1):3–42, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632624.

Daumas:1982:FEV

[2623] Maurice Daumas. Le faux échappement de Villard de Honnecourt. Revue d’Histoire des Sciences, 35(1):43–54, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632625.

Beaulieu:1982:DLM

[2624] Armand Beaulieu. Découverte d’un livre de Mersenne. Revue d’Histoire des Sciences, 35(1):55–56, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632626.

Herlea:1982:DHT

[2625] Alexandre Herlea. Deux histoires des techniques. Revue d’Histoire des Sciences, 35(1):57–63, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632627.

Anonymous:1982:Ia

[2626] Anonyme. Informations. Revue d’Histoire des Sciences, 35(1):64–66, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632628.

Caveing:1982:AOBa

[2627] Maurice Caveing. Analyses d’ouvrages: La sphère en mouvement, Levers et couchers héliaques, Testimonia, ^{≪}Les BellesLettres ^{≫}, (Coll. des Universités de France) par Autolycos de Pitane; Germaine Aujac; J.-P. Brunet; R. Nadal. Revue d’Histoire des Sciences, 35(1):67–69, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632629.

Bernhardt:1982:AOBa

[2628] Jean Bernhardt. Analyses d’ouvrages: Introduction aux Phénomènes, (Coll. des Universités de France) par Géminos; G. Aujac. Revue d’Histoire des Sciences, 35(1):70, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632630.

Casali:1982:AOB

[2629] Serge Casali and Samuela Etter. Analyses d’ouvrages: Il contributo di Simon Stevin allo sviluppo scientifico dell prospettiva artificiale par Rocco Sinisgalli; Luigi Vagnetti. Revue d’Histoire des Sciences, 35(1):70–71, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632631.

Bernhardt:1982:AOBb

[2630] Jean Bernhardt. Analyses d’ouvrages: Galileo at Work, his Scientific Biography par Stillman Drake. Revue d’Histoire des Sciences, 35(1):71– 73, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632632.

Bernhardt:1982:AOBc

[2631] Jean Bernhardt. Analyses d’ouvrages: Index des ^{≪}Regulae ad directionem ingenii ^{≫}de René Descartes, (^{≪}Lessico intellectuale europeo, X, Corpus cartesianum, 1 ^{≫}) par J.-R. Armogathe; J.-L. Marion. Revue d’Histoire des Sciences, 35(1):73–74, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632633.

Bernhardt:1982:AOBd

[2632] Jean Bernhardt. Analyses d’ouvrages: Hermeticism and the Scientific Revolution par R. S. Westman; J. E. McGuire. Revue d’Histoire des Sciences, 35(1):74–75, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632634.

Roger:1982:AOB

[2633] Jacques Roger. Analyses d’ouvrages: The Edge of Contingency. French Catholic Reaction to Scientific Change from Darwin to Duhem par Harry W. Paul. Revue d’Histoire des Sciences, 35(1):75–76, January 1982. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632635.

Todericiu:1982:AOBa

[2634] Doru Todericiu. Analyses d’ouvrages: Savoir, Devoir, Pouvoir. La Science moderne et les fondements de l’éthique, (Coll. ^{≪}Théoriques ^{≫}) par Paul Chambdal. Revue d’Histoire des Sciences, 35(1):76–77, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632636.

Todericiu:1982:AOBb

[2635] Doru Todericiu. Analyses d’ouvrages: Ciencias y enseñanza en la revolución burguesa par J. L. Peset; S. Garma; J. S. Pérez Garzon. Revue d’Histoire des Sciences, 35(1):77–78, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632637.

Dugac:1982:AOBa

[2636] Pierre Dugac. Analyses d’ouvrages: Vorlesungen über die Entwicklung der Mathematik im 19. Jahrhundert par Felix Klein. Revue d’Histoire des Sciences, 35(1):78–79, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632638.

Dugac:1982:AOBb

[2637] Pierre Dugac. Analyses d’ouvrages: Lazare Carnot savant et sa contribution à la théorie de l’infini mathématique, (Collection des Travaux de l’Académie internationale d’Histoire des Sciences ^{≫}, n^{o} 20) par Charles-C. Gillispie; Adolf P. Youschkevitch. Revue d’Histoire des Sciences, 35(1):79–82, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632639.

Galles:1982:AOB

[2638] Carlos Delfino Galles. Analyses d’ouvrages: Nuclear Physics in Retrospect par Roger H. Stuewer. Revue d’Histoire des Sciences, 35(1):82–83, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632640.

Goupil-Sadoun:1982:AOBa

[2639] Michelle Goupil-Sadoun. Analyses d’ouvrages: Benjamin Thomson, Count Rumford par Sanborn C. Brown. Revue d’Histoire des Sciences, 35 (1):84, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632641.

Goupil-Sadoun:1982:AOBb

[2640] Michelle Goupil-Sadoun. Analyses d’ouvrages: The Transcendental Part of Chemistry par David M. Knight. Revue d’Histoire des Sciences, 35(1): 85, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632642.

Halleux:1982:AOBa

[2641] Robert Halleux. Analyses d’ouvrages: The Science of Minerals in the Age of Jefferson, (Transactions of the American Philosophical Society, vol. 68, 4) par John C. Greene; John G. Burke. Revue d’Histoire des Sciences, 35(1):85–86, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632643.

Goupil-Sadoun:1982:AOBc

[2642] Michelle Goupil-Sadoun. Analyses d’ouvrages: Neptune’s Gift, A History of Common Salt par Robert P. Multhauf. Revue d’Histoire des Sciences, 35(1):87–88, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632644.

Guedes:1982:AOB

[2643] Michel Guédès. Analyses d’ouvrages: Bentham e Linneo, una interpretazione singolare par Silvestro Marcucci. Revue d’Histoire des Sciences, 35(1):88, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632645.

Huard:1982:AOBa

[2644] Pierre Huard. Analyses d’ouvrages: Clasicos españoles de la Anatomia Patologica anteriores a Cajal, (^{≪}Cuadernos valencianos de Historia de la Medicina et de la Ciencia ^{≫}, XXI, série B, textos classicos) par J. M. Lopez Pinero; F. Bajosa; M. L. Terrada. Revue d’Histoire des Sciences, 35(1):88–89, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632646.

Huard:1982:AOBb

[2645] Pierre Huard. Analyses d’ouvrages: Franz Joseph Gall, naturforscher und anthropologe par Erna Lesky. Revue d’Histoire des Sciences, 35(1):

89–90, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632647.

Huard:1982:AOBc

[2646] Pierre Huard. Analyses d’ouvrages: Contribution à la bio-bibliographie de Paul Broca par Pham Van Gian. Revue d’Histoire des Sciences, 35(1): 91–92, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632648.

Fischer:1982:AOB

[2647] Jean-Louis Fischer. Analyses d’ouvrages: Rudolf Leuckart, Weg und Werk, (^{≪}Biographien bedeutender Biologen ^{≫}, Bd. 2) par Klaus Wunderlich. Revue d’Histoire des Sciences, 35(1):92–94, January 1982. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632649.

Pueyo:1982:AOB

[2648] Guy Pueyo. Analyses d’ouvrages: Si la motoculture m’était contée … par Jacques Baratte. Revue d’Histoire des Sciences, 35(1):94–95, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632650.

Theodorides:1982:AOBa

[2649] Jean Théodoridès. Analyses d’ouvrages: L’Homme et l’Abeille par Philippe Marchenay. Revue d’Histoire des Sciences, 35(1):96, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632651.

Youschkevitch:1982:NRH

[2650] Adolf P. Youschkevitch. Nouvelles recherches sur l’histoire des mathématiques chinoises. Revue d’Histoire des Sciences, 35(2):97–110, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632129.

Halleux:1982:LGF

[2651] Robert Halleux. La littérature géologique française de 1500 à 1650 dans son contexte européen. Revue d’Histoire des Sciences, 35(2):111–130, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632130.

Huard:1982:GA

[2652] Pierre Huard and Marie-Josè Imbault-Huart. Gabriel Andral (1797– 1876). Revue d’Histoire des Sciences, 35(2):131–153, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632131.

Payen:1982:TPO

[2653] Jacques Payen. Trois publications à l’occasion du centenaire de l’Institut technico-industriel Aldini-Valeriani. Revue d’Histoire des Sciences, 35(2): 155–157, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632132.

Payen:1982:PAI

[2654] Jacques Payen. À propos de l’archéologie industrielle. Revue d’Histoire des Sciences, 35(2):158–162, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632133.

Anonymous:1982:Ib

[2655] Anonyme. Informations. Revue d’Histoire des Sciences, 35(2):163–164, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632134.

Caveing:1982:AOBb

[2656] Maurice Caveing. Analyses d’ouvrages: Théorie élémentaire (De motu circulari corporum caelestium), (Coll. ^{≪}Histoire des Doctrines de l’Antiquité classique ^{≫}) par Cléomède; Richard Goulet. Revue d’Histoire des Sciences, 35(2):165–167, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632135.

Bernhardt:1982:AOBe

[2657] Jean Bernhardt and L. Lagarde. Analyses d’ouvrages: De la Terre plate au globe terrestre, une mutation épistémologique rapide, 1480–1520, (^{≪}Cahiers des Annales ^{≫}, 38) par W. G. L. Randles. Revue d’Histoire des Sciences, 35(2):167–168, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632136.

Bernhardt:1982:AOBf

[2658] Jean Bernhardt. Analyses d’ouvrages: Correspondance du P. Marin Mersenne, Religieux Minime, XIV : 1646 par P. Tannery; Cornelis De Waard; Armand Beaulieu. Revue d’Histoire des Sciences, 35(2):169–170, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632137.

deValence:1982:AOBa

[2659] Maylis de Valence. Analyses d’ouvrages: Studies on Christiaan Huygens, Invited Papers from the Symposium on the Life and Work of Christiaan Huygens, Amsterdam, 22–25 par H. J. M. Bos; M. J. S. Rudwick; H. A. M. Snelders. Revue d’Histoire des Sciences, 35(2):170–171, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632138.

Grmek:1982:AOBa

[2660] M. D. Grmek. Analyses d’ouvrages: Médecine de l’esprit ou préceptes généraux de l’art de découvrir par Ehrenfried Walther von Tschirnhaus; Jean-Paul Wurtz. Revue d’Histoire des Sciences, 35(2):171, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632139.

Dieudonne:1982:AOBa

[2661] Jean Dieudonné. Analyses d’ouvrages: Der Beginn der Determinantentheorie, (Arbor Scientiarum, Beiträge zur Wissenschaftsgeschichte, Reihe B : Texte Band II) par Eberhard Knobloch. Revue d’Histoire des Sciences, 35(2):171–173, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632140.

Dieudonne:1982:AOBb

[2662] Jean Dieudonné. Analyses d’ouvrages: Opera Omnia, Series Quarta A : Commercium Epistolicum, vol. V, Correspondance de Leonhard Euler avec A. C. Clairaut, J. d’Alembert et J. L. Lagrange par Leonhard Euler; A. Juskevic; R. Taton. Revue d’Histoire des Sciences, 35(2):173–174, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632141.

Acloque:1982:AOB

[2663] Paul Acloque. Analyses d’ouvrages: Physik im 19. Jahrhundert par Andreas Kleinert. Revue d’Histoire des Sciences, 35(2):174–175, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632142.

Cohen:1982:AOB

[2664] Eliott Cohen. Analyses d’ouvrages: Sophie Germain, An Essay in the History of the Theory of Elasticity, (^{≪}Studies in the History of Modern Science ^{≫}, 6) par Louis L. Bucciarelli; Nancy Dworsky. Revue d’Histoire des Sciences, 35(2):175–176, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632143.

Theodorides:1982:AOBb

[2665] Jean Théodoridès. Analyses d’ouvrages: Physiologie et classification. Relations historiques, (Coll. ^{≪}Université de Compiègne ^{≫}) par Joseph Schiller; F. d’Autheville; R. H. Polge d’Autheville. Revue d’Histoire des Sciences, 35(2):176–177, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632144.

Huard:1982:AOBd

[2666] Pierre Huard. Analyses d’ouvrages: Regards d’un biologiste. Evolution de l’approche scientifique. L’enseignement médical strasbourgeois par Marc Klein; G. Canguilhem; P. Karli. Revue d’Histoire des Sciences, 35(2): 177–178, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632145.

Fantini:1982:AOB

[2667] Bernardino Fantini. Analyses d’ouvrages: Les origines de la biologie moléculaire. Hommage à Jacques Monod, Origins of Molecular Biology. A tribute to Jacques Monod par A. Lwoff; A. Ullmann. Revue d’Histoire des Sciences, 35(2):178–180, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632146.

Louis:1982:AOBa

[2668] Pierre Louis. Analyses d’ouvrages: Hippocratica, Actes du colloque hippocratique de Paris (4–9 septembre 1978) par M. D. Grmek. Revue d’Histoire des Sciences, 35(2):181–182, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632147.

Sournia:1982:AOB

[2669] J.-C. Sournia. Analyses d’ouvrages: Le Livre des axiomes médicaux (Aphorismi), (Coll. du Centre de Recherches d’Histoire et de Philologie de la IV^{e} section de l’EPHE) par Yûhannâ ibn Masawayh; Jean Mésué; Danielle Jacquart; Gérard Troupeau. Revue d’Histoire des Sciences, 35(2):182–183, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632148.

Grmek:1982:AOBb

[2670] M. D. Grmek. Analyses d’ouvrages: Vademecum eines frühmittelalterlichen Arztes, (^{≪}Publications de la Société suisse d’Histoire de la Médecine et des Sciences naturelles, ^{≫}vol. 34) par Peter Köpp. Revue d’Histoire des Sciences, 35(2):183–184, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632149.

Huard:1982:AOBe

[2671] Pierre Huard. Analyses d’ouvrages: Geschichte des Hospitals, Bd. 4 : Spanien von den Anfängen bis um 1500 par Dieter Jetter. Revue d’Histoire des Sciences, 35(2):184–185, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632150.

Hulin:1982:AOB

[2672] Nicole Hulin. Analyses d’ouvrages: The Organization of Science and Technology in France, 1808–1914 par Robert Fox; George Weisz. Revue d’Histoire des Sciences, 35(2):185–188, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632151.

Rosmorduc:1982:AOBa

[2673] Jean Rosmorduc. Analyses d’ouvrages: Qui a peur de la science? Travailleurs scientifiques, politique et société, (Coll. ^{≪}Notre Temps/Société ≫) par Jean-Marie Legay. Revue d’Histoire des Sciences, 35(2):188–189, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632152.

Freudenthal:1982:AOB

[2674] Gad Freudenthal. Analyses d’ouvrages: Naturwissenschaft, Technik und NS-Ideologie. Beiträge zur Wissenschaftsgeschichte des Dritten Reiches par Herbert Mehrtens; Steffen Richter. Revue d’Histoire des Sciences, 35(2):189–190, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632153.

Rosmorduc:1982:AOBb

[2675] Jean Rosmorduc. Analyses d’ouvrages: Le petit savant illustré, (Coll. ≪Science ouverte ^{≫}) par Pierre Thuillier. Revue d’Histoire des Sciences, 35(2):190–191, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632154.

Anonymous:1982:PR

[2676] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 35(2): 192, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632157.

deValence:1982:AOBb

[2677] Maylis de Valence. Analyses d’ouvrages: Catalogue des Ouvrages imprimés au XVI^{e} siècle. Sciences — Techniques — Médecine par Jacqueline Linet; Denise Hillard; Xavier Lavagne. Revue d’Histoire des Sciences, 35(2):192, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632155.

Grmek:1982:RDH

[2678] Mirko D. Grmek and Bernardino Fantini. Le rôle du hasard dans la naissance du modèle de l’opéron. Revue d’Histoire des Sciences, 35(3): 193–215, juillet 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632421.

Hulin:1982:PES

[2679] Nicole Hulin. À propos de l’enseignement scientifique : une réforme de l’enseignement secondaire sous le second empire, la ^{≪}bifurcation ^{≫}(1852– 1864). Revue d’Histoire des Sciences, 35(3):217–245, juillet 1982. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632422.

Jongh:1982:ISI

[2680] Ariane Isler de Jongh. Inventeur-savant et inventeur-innovateur : Charles Cros et Louis Ducos du Hauron : Les commencements de la photographie en couleurs. Revue d’Histoire des Sciences, 35(3):247–273, juillet 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632423.

Pernot:1982:ICC

[2681] J.-F. Pernot. Une importante contribution à la connaissance des ingénieurs d’ancien Régime. Revue d’Histoire des Sciences, 35(3):275– 280, juillet 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632424.

Anonymous:1982:Ic

[2682] Anonyme. Informations. Revue d’Histoire des Sciences, 35(3):281–282, juillet 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632425.

Caveing:1982:AOBc

[2683] Maurice Caveing. Analyses d’ouvrages: Astronomie et religion en Occident par Louis Rougier. Revue d’Histoire des Sciences, 35(3):283–284, juillet 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632426.

Bernhardt:1982:AOBg

[2684] Jean Bernhardt. Analyses d’ouvrages: Theories of Vision from Al-Kindi to Kepler par David C. Lindberg. Revue d’Histoire des Sciences, 35(3): 284–285, juillet 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632427.

Huard:1982:AOBf

[2685] Pierre Huard. Analyses d’ouvrages: The Iris in Eighteenth-Century Physiology, (^{≪}Berner Beiträge zur Geschichte der Medizin und der Naturwissenschaften ^{≫}) par Renato G. Mazzolini. Revue d’Histoire des Sciences, 35(3):285–287, juillet 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632428.

Laissus:1982:AOB

[2686] Yves Laissus. Analyses d’ouvrages: The letters of Georges Cuvier : A summary calendar of manuscript and printed materials preserved in Europe, the United States of America, and Australasia, (bshs Monographs,

2) par Dorinda Outram. Revue d’Histoire des Sciences, 35(3):288, juillet 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632429.

Boissel:1982:PIC

[2687] Jean Boissel and J.-P. Durand. À propos de l’indice céphalique: Lettres de Durand de Gros à Vacher de Lapouge. Revue d’Histoire des Sciences, 35(4):289–319, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632175.

Medard:1982:LLD

[2688] Louis Médard. La légende de la loi de Dulong. Revue d’Histoire des Sciences, 35(4):321–330, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632177.

Clavelin:1982:HST

[2689] Maurice Clavelin. Histoire des sciences et théorie du raisonnement: À propos d’un livre récent sur Galilée et l’art du raisonnement. Revue d’Histoire des Sciences, 35(4):331–340, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632178.

Shea:1982:MLR

[2690] William R. Shea. Maria Luisa Righini Bonelli (1917–1981). Revue d’Histoire des Sciences, 35(4):341–342, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632179.

Delorme:1982:I

[2691] Suzanne Delorme and Jean Théodoridès. Informations. Revue d’Histoire des Sciences, 35(4):343–348, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632180.

Halleux:1982:AOBb

[2692] Robert Halleux. Analyses d’ouvrages: Science in the Middle Ages par David C. Lindberg. Revue d’Histoire des Sciences, 35(4):349–350, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632181.

Lejbowicz:1982:AOBa

[2693] Max Lejbowicz. Analyses d’ouvrages: Recherches sur une technique divinatoire : la géomancie dans l’Occident médiéval, (Centre de recherches d’histoire et de philologie de la IV^{e} Section de l’École pratique des Hautes Études, Série V : ^{≪}Hautes études médiévales et modernes ^{≫}, 44) par Thérèse Charmasson. Revue d’Histoire des Sciences, 35(4):350–352, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632182.

Shea:1982:AOBa

[2694] William R. Shea. Analyses d’ouvrages: The Merton Tradition and Kinematics in late Sixteenth and early Seventeenth Century Italy, Padova, (^{≪}Saggi e Testi ^{≫}, 15) par Christopher Lewis. Revue d’Histoire des Sciences, 35(4):352–354, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632183.

Halleux:1982:AOBc

[2695] Robert Halleux. Analyses d’ouvrages: Sol und Luna. Literar- und alchemiegeschichtliche Studien zu einem altdeutschen Bildgedicht. Mit Text- und Bildanhang, (^{≪}Schriften zur Wissenschaftsgeschichte ^{≫}, II) par Joachim Telle. Revue d’Histoire des Sciences, 35(4):354–355, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632184.

Lejbowicz:1982:AOBb

[2696] Max Lejbowicz. Analyses d’ouvrages: English Almanacs, 1500–1800. Astrology and the Popular Press par Bernard Capp. Revue d’Histoire des Sciences, 35(4):355–357, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632185.

Shea:1982:AOBb

[2697] William R. Shea. Analyses d’ouvrages: English Scientific Virtuosi in the 16th and 17th Centuries par Barbara Shapiro; Robert G. Frank. Revue d’Histoire des Sciences, 35(4):357–358, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632186.

Rosmorduc:1982:AOBc

[2698] J. Rosmorduc. Analyses d’ouvrages: L’avènement de la science physique. Essai sur la Physique d’Aristote par Lambros Couloubaritsis. Revue d’Histoire des Sciences, 35(4):358, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632187.

Shea:1982:AOBc

[2699] William R. Shea. Analyses d’ouvrages: Never at Rest. A biography of Isaac Newton par Richard S. Westfall. Revue d’Histoire des Sciences, 35(4):358–359, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632188.

Rosmorduc:1982:AOBd

[2700] J. Rosmorduc. Analyses d’ouvrages: Essays on the History of the Microscope par Gerard L’E. Turner. Revue d’Histoire des Sciences, 35(4):360, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632190.

Todericiu:1982:AOBc

[2701] D. Todericiu. Analyses d’ouvrages: Die Natur des Lichts. Einbeziehung eines Physikgeschichtlichen Themas in den Physikunterricht, (^{≪}Erträge der Forschung ^{≫}, Bd. 128) par Karl Friedrich Weinmann. Revue d’Histoire des Sciences, 35(4):360, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632189.

Ramunni:1982:AOB

[2702] Girolamo Ramunni. Analyses d’ouvrages: Pioneers of Science, Nobel Prize Winners in Physics par Robert L. Weber; J. M. A. Lenihan. Revue d’Histoire des Sciences, 35(4):361, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632191.

Todericiu:1982:AOBd

[2703] D. Todericiu. Analyses d’ouvrages: The Origins and early Development of the Heavy Chemical Industry in France, (^{≪}Oxford Science Publications ≫) par John Graham Smith. Revue d’Histoire des Sciences, 35(4):361– 362, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632192.

Todericiu:1982:AOBe

[2704] D. Todericiu. Analyses d’ouvrages: Revue de l’Essai de Statique chimique par Claude-Louis Berthollet; Michelle Sadoun-Goupil. Revue d’Histoire des Sciences, 35(4):362–363, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632193.

Louis:1982:AOBb

[2705] Pierre Louis. Analyses d’ouvrages: Inventaire analytique des papyrus grecs de médecine, (Centre de Recherches d’Histoire et de Philologie de la IV^{e} section EPHE, III Hautes Études du Monde gréco-romain, 12) par Marie–Hélène Marganne. Revue d’Histoire des Sciences, 35(4):364, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632194.

Halleux:1982:AOBd

[2706] Robert Halleux. Analyses d’ouvrages: Ein neuer Textbeleg vom Liber praeparationum, (^{≪}Salzburger Beiträge zur Paracelsusforschung ^{≫}, 19) par Wolfgang Schneider. Revue d’Histoire des Sciences, 35(4):364–365, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632195.

Halleux:1982:AOBe

[2707] Robert Halleux. Analyses d’ouvrages: Paracelsus in der Tradition. Vorträge Paracelsustag 1978, (^{≪}Salzburger Beiträge zur Paracelsusforschung ^{≫}, 21). Revue d’Histoire des Sciences, 35(4):365, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632196.

Huard:1982:AOBg

[2708] Pierre Huard. Analyses d’ouvrages: Contribution à la pensée médicale de Lan Ong (1720–1791) par Thi Hoai Tu Trinh. Revue d’Histoire des Sciences, 35(4):365–367, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632197.

Gourevitch:1982:AOB

[2709] Michel Gourevitch. Analyses d’ouvrages: Cabanis, Enlightenment and Medical Philosophy in the French Revolution par Martin S. Staum. Revue d’Histoire des Sciences, 35(4):367–369, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632198.

Turner:1982:AOB

[2710] A. J. Turner. Analyses d’ouvrages: The Clockwork Universe : German Clocks and Automata 1550–1650 par Otto Mayer; Klaus Maurice; Otto Mayr. Revue d’Histoire des Sciences, 35(4):369–370, October 1982. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632199.

Daumas:1982:AOB

[2711] Maurice Daumas. Analyses d’ouvrages: Oberkampf, un entrepreneur capitaliste au siècle des Lumières, (^{≪}Collection historique ^{≫}) par Serge Chassagne. Revue d’Histoire des Sciences, 35(4):370–373, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632200.

deValence:1982:AOBc

[2712] Maylis de Valence. Analyses d’ouvrages: Historia Mathematica. International Journal of History of Mathematics. Vol. 2–4 par K. O. May; J. W. Dauben. Revue d’Histoire des Sciences, 35(4):373–374, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632201.

Anonymous:1982:TGM

[2713] Anonyme. Table générale des matières du Tome XXXV. Revue d’Histoire des Sciences, 35(4):375–384, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632202.

Anonymous:1982:BMa

[2714] Anonyme. Back matter. Revue d’Histoire des Sciences, 35(1):??, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632652.

Anonymous:1982:BMb

[2715] Anonyme. Back matter. Revue d’Histoire des Sciences, 35(2):??, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632158.

Anonymous:1982:BMc

[2716] Anonyme. Back matter. Revue d’Histoire des Sciences, 35(3):??, juillet 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632430.

Anonymous:1982:BMd

[2717] Anonyme. Back matter. Revue d’Histoire des Sciences, 35(4):??, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632203.

Anonymous:1982:FMa

[2718] Anonyme. Front matter. Revue d’Histoire des Sciences, 35(1):??, January 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632623.

Anonymous:1982:FMb

[2719] Anonyme. Front matter. Revue d’Histoire des Sciences, 35(2):??, avril 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632128.

Anonymous:1982:FMc

[2720] Anonyme. Front matter. Revue d’Histoire des Sciences, 35(3):??, juillet 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632420.

Anonymous:1982:FMd

[2721] Anonyme. Front matter. Revue d’Histoire des Sciences, 35(4):??, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632174.

Giard:1983:LSD

[2722] Luce Giard. Logique et système du savoir selon Hugues de Saint-Victor. Revue d’Histoire des Sciences, 36(1):3–32, January 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632326.

Acot:1983:DE

[2723] Pascal Acot. Darwin et l’écologie. Revue d’Histoire des Sciences, 36(1): 33–48, January 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632327.

Moulin:1983:AAS

[2724] Anne-Marie Moulin. De l’analyse au système : le développement de l’immunologie. Revue d’Histoire des Sciences, 36(1):49–67, January 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632328.

Mosconi:1983:CBV

[2725] Jean Mosconi. Charles Babbage: vers une théorie du calcul mécanique. (French) [Charles Babbage: toward a theory of mechanical calculation]. Revue d’Histoire des Sciences, 36(1):69–107, January 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632329.

Anonymous:1983:RS

[2726] Anonyme. Résumés et summaries. Revue d’Histoire des Sciences, 36(1):109–110, January 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632330.

Forbes:1983:CAE

[2727] Eric G. Forbes. La correspondance astronomique entre Joseph-Nicolas Delisle et Tobias Mayer. Revue d’Histoire des Sciences, 36(2):113–151, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632363.

Delisle:1983:OH

[2728] Delisle, Danielle Fauque, and M. Folkes. Les origines de l’héliomètre. Revue d’Histoire des Sciences, 36(2):153–171, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632364.

Hirano:1983:NSC

[2729] Yoīchi Hirano and J.-M.-C. Duhamel. Note sur la seconde thèse de Camille Jordan. Revue d’Histoire des Sciences, 36(2):173–174, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632365. With an ap-

pendix by J.-M.-C. Duhamel.

Huard:1983:JFO

[2730] Pierre Huard. Jean Filliozat (4 octobre 1906–27 octobre 1982). Revue d’Histoire des Sciences, 36(2):175–176, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632366.

Anonymous:1983:I

[2731] Anonyme. Informations. Revue d’Histoire des Sciences, 36(2):177–178, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632367.

Shea:1983:AOBa

[2732] William R. Shea. Analyses d’ouvrages: La naissance du purgatoire, (coll. ≪Bibliothèque des Histoires ^{≫}) par Jacques Le Goff. Revue d’Histoire des Sciences, 36(2):179–180, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632368.

LHuillier:1983:AOBa

[2733] Hervé L’Huillier. Analyses d’ouvrages: De numeris datis, A critical edition and translation par Jordanus de Nemore; Barnabas Hughes; Berkeley-Los Angeles-London. Revue d’Histoire des Sciences, 36(2):180– 181, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632369.

Rosmorduc:1983:AOBa

[2734] Jean Rosmorduc. Analyses d’ouvrages: Le coq et la louve. Histoire de Gerbert et l’an mille par Florence Trystram; Florence Trystam. Revue d’Histoire des Sciences, 36(2):181, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632370.

Rosmorduc:1983:AOBb

[2735] J. Rosmorduc. Analyses d’ouvrages: La conquête du savoir par Isaac Asimov; Izaac Asimov. Revue d’Histoire des Sciences, 36(2):182, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632371.

Rosmorduc:1983:AOBc

[2736] J. Rosmorduc. Analyses d’ouvrages: Le mécanisme hier et aujourd’hui, (Coll. ^{≪}Que sais-je? ^{≫}) par René Boirel. Revue d’Histoire des Sciences, 36(2):182–183, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633392.

deBeauregard:1983:AOBa

[2737] O. Costa de Beauregard. Analyses d’ouvrages: The Study of Time IV par J. T. Fraser; N. Lawrence; D. Park. Revue d’Histoire des Sciences, 36(2): 183–184, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632372.

Rosmorduc:1983:AOBd

[2738] J. Rosmorduc. Analyses d’ouvrages: Ainsi naquit l’informatique. Les hommes, les matériels, à l’origine des concepts de l’informatique d’aujourd’hui, 2^{e} éd. par René Moreau. Revue d’Histoire des Sciences, 36(2):184, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632373.

Todericiu:1983:AOBa

[2739] Doru Todericiu. Analyses d’ouvrages: Macchine e pensiero. Da Wiener alla terza Cibernetica, (coll. Sc/10, 95) par Gennar Luigi Linguitti. Revue d’Histoire des Sciences, 36(2):184–185, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632374.

Todericiu:1983:AOBb

[2740] D. Todericiu. Analyses d’ouvrages: Symbolismus und symbolische Logik. Die Idee der Ars Combinatoria in der Entwicklung der modernen Dichtung, (^{≪}Humanistische Bibliothek ^{≫}) par John Neubauer. Revue d’Histoire des Sciences, 36(2):185, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632375.

Gourlet:1983:AOB

[2741] V. Gourlet. Analyses d’ouvrages: Penser les mathématiques, (coll. ≪Points ^{≫}, série ^{≪}Sciences ^{≫}) par F. Guénard; G. Lelièvre. Revue d’Histoire des Sciences, 36(2):186, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632376.

Youschkevitch:1983:AOB

[2742] A. P. Youschkevitch. Analyses d’ouvrages: The Mathematical Papers of Isaac Newton, vol. VIII : 1697–1722 par D. T. Whiteside; A. Prag. Revue d’Histoire des Sciences, 36(2):186–188, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632377.

Shea:1983:AOBb

[2743] William R. Shea. Analyses d’ouvrages: The Newtonian Revolution, with illustrations of the transformation of scientific ideas par I. Bernard Cohen. Revue d’Histoire des Sciences, 36(2):188–190, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632378.

Rosmorduc:1983:AOBe

[2744] J. Rosmorduc. Analyses d’ouvrages: Dictionnaire de physique par J.-P. Sarmant. Revue d’Histoire des Sciences, 36(2):191, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632379.

Rosmorduc:1983:AOBf

[2745] J. Rosmorduc. Analyses d’ouvrages: Du radium aux microprocesseurs. Histoires de l’École supérieure de Physique et de Chimie, 1882–1982 par Pierre Bicquard. Revue d’Histoire des Sciences, 36(2):191, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632380.

Ramunni:1983:AOBa

[2746] Girolamo Ramunni. Analyses d’ouvrages: Albert Einstein’s Special Theory of Relativity. Emergence (1905) and Early Interpretation (1905– 1911) par Arthur I. Miller. Revue d’Histoire des Sciences, 36(2):191– 192, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632381.

deBeauregard:1983:AOBb

[2747] O. Costa de Beauregard. Analyses d’ouvrages: Special relativity. The foundations of macroscopic physics par W. G. Dixon. Revue d’Histoire des Sciences, 36(2):192–194, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632382.

Rosmorduc:1983:AOBg

[2748] Jean Rosmorduc. Analyses d’ouvrages: La diffusion de la théorie de la relativité en France par Michel Biezunski. Revue d’Histoire des Sciences, 36(2):194–195, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632383.

Ramunni:1983:AOBb

[2749] Girolamo Ramunni. Analyses d’ouvrages: The Impact of Modern Scientific Ideas on Society. In Commemoration of Einstein par Colette M. Kinnon; A. N. Kholodilin; J. G. Richardson. Revue d’Histoire des Sciences, 36(2):195, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632384.

Chevalley:1983:AOB

[2750] Catherine Chevalley. Analyses d’ouvrages: Les conceptions quantiques de 1911 à 1927, (^{≪}L’Histoire des Sciences Textes et Études ^{≫}) par Girolamo Ramunni. Revue d’Histoire des Sciences, 36(2):196–198, avril 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632385.

Rosmorduc:1983:AOBh

[2751] J. Rosmorduc. Analyses d’ouvrages: L’étrange histoire des quanta, (coll. ^{≪}Points-Sciences ^{≫}) par Banesh Hoffmann; Michel Paty. Revue d’Histoire des Sciences, 36(2):198, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632386.

Rosmorduc:1983:AOBi

[2752] J. Rosmorduc. Analyses d’ouvrages: La lumière, (coll. ^{≪}Points-Sciences ≫, 528) par Bernard Maitte. Revue d’Histoire des Sciences, 36(2):198– 199, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632387.

Blay:1983:AOBa

[2753] Michel Blay. Analyses d’ouvrages: Colour Vision in the Nineteenth Century. The Young–Helmholtz–Maxwell Theory, Bristol par Paul D. Shermann; Paul D. Sherman. Revue d’Histoire des Sciences, 36(2):199–201, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632388.

Charmasson:1983:AOB

[2754] Thérèse Charmasson. Analyses d’ouvrages: Le milieu médical en France du XII^{e} au XV^{e} siècle. En annexe, 2^{e} Supplément au ^{≪}Dictionnaire ≫d’Ernest Wickersheimer, (EPHE, V, ^{≪}Hautes Études médiévales et modernes ^{≫}, 46) par Danielle Jacquart. Revue d’Histoire des Sciences, 36(2):201–203, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632389.

Taton:1983:AOBa

[2755] René Taton. Analyses d’ouvrages: La Faculté des Sciences de Montpellier de ses origines à nos jours par Louis Dulieu. Revue d’Histoire des Sciences, 36(2):203–204, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632390.

Rosmorduc:1983:AOBj

[2756] J. Rosmorduc. Analyses d’ouvrages: Les grandes étapes du progrès technique, (Coll. ^{≪}Que sais-je? ^{≫}, n^{o} 1960) par Maurice Daumas. Revue d’Histoire des Sciences, 36(2):204, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632391.

Todericiu:1983:AOBc

[2757] D. Todericiu. Analyses d’ouvrages: The Coming of the Age of Iron par Th. A. Wertime; J. D. Muhly. Revue d’Histoire des Sciences, 36(2): 205, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632392.

Pueyo:1983:AOB

[2758] Guy Pueyo. Analyses d’ouvrages: Agronomes andalous du Moyen Age, (coll.^{≪}Études et Documents^{≫}, n^{o} 13) par Lucie Bolens. Revue d’Histoire des Sciences, 36(2):205–206, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632393.

Anonymous:1983:PR

[2759] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 36(2):207– 208, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632394.

Peiffer:1983:JLS

[2760] Jeanne Peiffer. Joseph Liouville (1809–1882): ses contributions à la théorie des fonctions d’une variable complexe. Revue d’Histoire des Sciences, 36(3–4):209–248, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632205.

Debarbat:1983:MGA

[2761] Suzanne Débarbat. Méthode de Gauss et astrolabe à prisme. Revue d’Histoire des Sciences, 36(3–4):249–263, juillet–décembre 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632206.

Acloque:1983:MMT

[2762] Paul Acloque. Les marées et le mouvement de la Terre. Revue d’Histoire des Sciences, 36(3–4):265–283, juillet–décembre 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632207.

Boulaine:1983:VVD

[2763] Jean Boulaine. V. V. Dokouchaev et les débuts de la pédologie. Revue d’Histoire des Sciences, 36(3–4):285–306, juillet–décembre 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632208.

Rookmaaker:1983:HDR

[2764] L. C. Rookmaaker. Histoire du rhinocéros de Versailles (1770–1793). Revue d’Histoire des Sciences, 36(3–4):307–318, juillet–décembre 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632209.

DeGandt:1983:GND

[2765] François De Gandt. Galilée et la naissance de la dynamique : recherches de P. Galluzzi sur le terme ^{≪}momento ^{≫}. Revue d’Histoire des Sciences, 36(3–4):319–324, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632210.

Blay:1983:HF

[2766] Michel Blay. Huygens et la France. Revue d’Histoire des Sciences, 36(3– 4):325–328, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632211.

Costabel:1983:MAT

[2767] Pierre Costabel. Marie-Antoinette Tonnelat (1912–1980). Revue d’Histoire des Sciences, 36(3–4):329–331, juillet–décembre 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632212.

Theodorides:1983:PH

[2768] Jean Théodoridès. Pierre Huard (1901–1983). Revue d’Histoire des Sciences, 36(3–4):332–334, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632213.

Gohau:1983:I

[2769] Gabriel Gohau. Informations. Revue d’Histoire des Sciences, 36(3–4): 335–344, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632214.

Elena:1983:AOB

[2770] Alberto Elena. Analyses d’ouvrages: Narratio prima, (^{≪}Studia Copernicana ^{≫}, XX) par Georg Joachim Rheticus; Henri Hugonnard-Roche; Jean-Pierre Verdet; Michel-Pierre Lerner; Alain Segonds. Revue d’Histoire des Sciences, 36(3–4):345–346, juillet–décembre 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632215.

Seidengart:1983:AOB

[2771] Jean Seidengart. Analyses d’ouvrages: Les orbites des Planètes (Dissertation de 1801), (coll. ^{≪}Textes et Études d’Histoire des Sciences ^{≫}) par G. W. F. Hegel; François De Gandt; Dominique Dubarle. Revue d’Histoire des Sciences, 36(3–4):346–348, juillet–décembre 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632216.

Blay:1983:AOBb

[2772] Michel Blay. Analyses d’ouvrages: Newton on the Continent par Henry Guerlac. Revue d’Histoire des Sciences, 36(3–4):348–350, juillet– décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632217.

Jorland:1983:AOB

[2773] Gérard Jorland. Analyses d’ouvrages: Science and Polity in France at the end of the Old Regime par Charles C. Gillispie. Revue d’Histoire des Sciences, 36(3–4):350–353, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632218.

Delorme:1983:AOB

[2774] Suzanne Delorme. Analyses d’ouvrages: Jean Rostand, prophète clairvoyant et fraternel par Andrée Tétry; Étienne Wolff. Revue d’Histoire des Sciences, 36(3–4):353–355, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632219.

Brouzeng:1983:AOBa

[2775] Paul Brouzeng. Analyses d’ouvrages: Routes et dédales — Histoire des mathématiques, (coll.^{≪}Axes^{≫}) par Amy Dahan–Dalmédico; Jeanne Peiffer. Revue d’Histoire des Sciences, 36(3–4):355–356, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632220.

Dahan-Dalmedico:1983:AOB

[2776] Amy Dahan-Dalmédico. Analyses d’ouvrages: Mathematics, The Loss of Certainty par Morris Kline. Revue d’Histoire des Sciences, 36(3–4): 356–358, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632221.

LHuillier:1983:AOBb

[2777] Hervé L’Huillier. Analyses d’ouvrages: Practical Geometry in the High Middle Ages. ^{≪}Artis cuiuslibet consummatio ^{≫}and the ^{≪}Pratike de Geometrie ^{≫} par Stephen K. Victor. Revue d’Histoire des Sciences, 36(3– 4):358–360, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632222.

Peiffer:1983:AOB

[2778] Jeanne Peiffer. Analyses d’ouvrages: Dürer — Kunst und Geometrie, Dürers künstlerisches Schaffen aus der Sicht seiner ^{≪}Underweysung ≫ par Eberhard Schröder. Revue d’Histoire des Sciences, 36(3–4):

360–361, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632223.

Dieudonne:1983:AOB

[2779] Jean Dieudonné. Analyses d’ouvrages: 100 years of mathematics par George Temple. Revue d’Histoire des Sciences, 36(3–4):361–364, juillet– décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632224.

Grattan-Guinness:1983:AOB

[2780] I. Grattan-Guinness. Analyses d’ouvrages: Geschichte des Mannigfaltigkeitsbegriffs von Riemann bis Poincaré par Erhard Scholz. Revue d’Histoire des Sciences, 36(3–4):364–365, juillet–décembre 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632225.

Taton:1983:AOBb

[2781] René Taton. Analyses d’ouvrages: Epistémologie des nombres relatifs, Recherches en didactique des mathématiques, vol. 2, n^{o} 3 par G. Glaeser. Revue d’Histoire des Sciences, 36(3–4):365, juillet–décembre 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632226.

Levy:1983:AOB

[2782] Jacques Lévy. Analyses d’ouvrages: Astronomie et astronomes en Provence, 1680–1730 par Jean-Marie Homet; Michel Vovelle. Revue d’Histoire des Sciences, 36(3–4):366–368, juillet–décembre 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632227.

Brouzeng:1983:AOBb

[2783] Paul Brouzeng. Analyses d’ouvrages: De Thalès à Einstein — Histoire de la physique et de la Chimie, (coll. ^{≪}Axes ^{≫}) par Jean Rosmorduc. Revue d’Histoire des Sciences, 36(3–4):368, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632228.

Theodorides:1983:AOB

[2784] Jean Théodoridès. Analyses d’ouvrages: Felice Fontana 1730–1805. An annotated bibliography, (^{≪}Studi su Felice Fontana ^{≫}, 1) par Peter K. Knoefel. Revue d’Histoire des Sciences, 36(3–4):369, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632229.

Huard:1983:AOBa

[2785] Pierre Huard. Analyses d’ouvrages: Franz Joseph Gall, 1758–1828, Naturforscher und Anthropologe. (^{≪}Hans Hubers Klassiker der Medizin und der Naturwissenschaften ^{≫}, Bd. XV) par Erna Lesky. Revue d’Histoire des Sciences, 36(3–4):369–370, juillet–décembre 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632230.

Roger:1983:AOB

[2786] Jacques Roger. Analyses d’ouvrages: Lamarck et son temps, Lamarck et notre temps, (coll. ^{≪}L’Histoire des Sciences. Textes et Études ^{≫}). Revue d’Histoire des Sciences, 36(3–4):370–371, juillet–décembre 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632231.

Huard:1983:AOBb

[2787] P. Huard. Analyses d’ouvrages: La Ceroplastica nella Scienza e nell’Arte, Atti del I Congresso internazionale, (^{≪}Biblioteca della Rivista di Storia delle Scienze mediche e naturali ^{≫}, vol. XX) par Leo S. Olschki Firenze. Revue d’Histoire des Sciences, 36(3–4):371–373, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632232.

deValence:1983:AOB

[2788] M. de Valence. Analyses d’ouvrages: Historia Mathematica. International Journal of History of Mathematics, vol. 5–10 par K. O May; J. W. Dauben. Revue d’Histoire des Sciences, 36(3–4):373–376, juillet– décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632233.

Anonymous:1983:TGM

[2789] Anonyme. Table générale des matières du Tome XXXVI. Revue d’Histoire des Sciences, 36(3–4):377–384, juillet–décembre 1983. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632234.

Anonymous:1983:BMa

[2790] Anonyme. Back matter. Revue d’Histoire des Sciences, 36(1):??, January 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632331.

Anonymous:1983:BMb

[2791] Anonyme. Back matter. Revue d’Histoire des Sciences, 36(2):??, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632395.

Anonymous:1983:BMc

[2792] Anonyme. Back matter. Revue d’Histoire des Sciences, 36(3–4): ??, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632235.

Anonymous:1983:FMa

[2793] Anonyme. Front matter. Revue d’Histoire des Sciences, 36(1):??, January 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632325.

Anonymous:1983:FMb

[2794] Anonyme. Front matter. Revue d’Histoire des Sciences, 36(2):??, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632362.

Anonymous:1983:FMc

[2795] Anonyme. Front matter. Revue d’Histoire des Sciences, 36(3–4): ??, juillet–décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632204.

Dugac:1984:HEC

[2796] Pierre Dugac. Histoire des espaces complets. Revue d’Histoire des Sciences, 37(1):3–28, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632340.

Todericiu:1984:BGS

[2797] Doru Todericiu. Balthasar-Georges Sage (1740–1824) chimiste et minéralogiste français: fondateur de la première École des Mines (1783). Revue d’Histoire des Sciences, 37(1):29–46, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632341.

Corsi:1984:LI

[2798] Pietro Corsi. Lamarck en Italie. Revue d’Histoire des Sciences, 37(1): 47–64, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632342.

Dulieu:1984:PCM

[2799] Louis Dulieu. Pierre Chirac et les maladies des équipages des vaisseaux. Revue d’Histoire des Sciences, 37(1):65–70, January 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632343.

Huard:1984:TCO

[2800] Pierre Huard. Un traité chinois d’oculistique. Revue d’Histoire des Sciences, 37(1):71–74, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632344.

Grattan-Guinness:1984:NML

[2801] I. Grattan-Guinness. Note sur les manuscrits de libri conservés à Florence. Revue d’Histoire des Sciences, 37(1):75–76, January 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632345.

Theodorides:1984:I

[2802] Jean Théodoridès. Informations. Revue d’Histoire des Sciences, 37(1): 77–81, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632346.

Caveing:1984:AOBa

[2803] Maurice Caveing. Analyses d’ouvrages: Science and the sciences in Plato par J. Anton; G. Vlastos; A. Mourelatos; R. Turnbull; I. Mueller. Revue d’Histoire des Sciences, 37(1):82–83, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632347.

Reeds:1984:BRB

[2804] Karen Reeds. Book review: Essays on The Works of Erasmus par Richard L. DeMolen. Revue d’Histoire des Sciences, 37(1):84, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632348.

Fischer:1984:AOB

[2805] Jean-Louis Fischer. Analyses d’ouvrages: El cientifico español ante su Historia. La ciencia en España entre 1750–1850 par Santiago Garma. Revue d’Histoire des Sciences, 37(1):84–85, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632349.

Jaki:1984:AOB

[2806] Stanley L. Jaki. Analyses d’ouvrages: Epistemologia e storia della scienza. Le svolte teoriche da Duhem a Bachelard, (^{≪}Filosofia della Scienza ^{≫}, 18) par Pietro Redondi. Revue d’Histoire des Sciences, 37(1): 85–87, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632350.

Turner:1984:AOBa

[2807] A.-J. Turner. Analyses d’ouvrages: Scottish Scientific Instrument-Makers 1600–1900. Royal Scottish Museum. Information Series, ≪Technology ^{≫}, 1 par D. J. Bryden. Revue d’Histoire des Sciences, 37 (1):87, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632351.

Turner:1984:AOBb

[2808] A.-J. Turner. Analyses d’ouvrages: Pio IX e le Macchine Ottiche, con riproduzione anastatica del saggio fisico-matematico proposto da Giovan Maria Mastai Ferretti a conclusione dei studi presso le Scuole Pie di Volterra il 7 Settembre 1809 par Pier Luigi Pizzamiglio; Giorgio Tabarroni. Revue d’Histoire des Sciences, 37(1):87–88, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632352.

Roger:1984:AOB

[2809] Jacques Roger. Analyses d’ouvrages: Le procès du Singe. La Bible contre Darwin, (coll. ^{≪}La Mémoire du Siècle ^{≫}) par Gordon Golding. Revue d’Histoire des Sciences, 37(1):88–89, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632353.

Murphy:1984:AOBa

[2810] Terence D. Murphy. Analyses d’ouvrages: Healing and History. Essays for George Rosen par Charles E. Rosenberg. Revue d’Histoire des Sciences, 37(1):89–90, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632354.

Murphy:1984:AOBb

[2811] Terence D. Murphy. Analyses d’ouvrages: Professionalizing Modern Medicine. Paris Surgeons and Medical Science and Institutions in the 18th Century, (^{≪}Contributions in Medical History ^{≫}, n^{o} 6) par Toby Gelfand. Revue d’Histoire des Sciences, 37(1):90–91, January 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632355.

Boulle:1984:AOBa

[2812] Lydie Boulle. Analyses d’ouvrages: L’hôpital de Grenoble des origines à la fin du Second Empire (XI^{e} siècle-1870), (Publications du crhesi de Grenoble, ^{≪}Série histoire institutionnelle ^{≫}, vol. 8) par Patrice Cugnetti. Revue d’Histoire des Sciences, 37(1):91–92, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632356.

Moulin:1984:AOBa

[2813] Anne-Marie Moulin. Analyses d’ouvrages: Querelle de l’inoculation ou préhistoire de la vaccination, (coll. ^{≪}Problèmes et Controverses ^{≫}) par J.-F. de Raymond. Revue d’Histoire des Sciences, 37(1):93, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632357.

Vidal:1984:AOB

[2814] F. Vidal. Analyses d’ouvrages: Storia della Odontoiatria par Placido Micheloni. Revue d’Histoire des Sciences, 37(1):94–95, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632358.

Pueyo:1984:AOBa

[2815] Guy Pueyo. Analyses d’ouvrages: Ele desagüe del valle de México durante la época novohispana par Jorge Gurriá Lacroix. Revue d’Histoire des Sciences, 37(1):95–96, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632359.

deValence:1984:AOBa

[2816] M. de Valence. Analyses d’ouvrages: Gaán.ita-Bharati. Bulletin of the Indian Society for History of Mathematics, vol. 1–3 par R. C. Gupta. Revue d’Histoire des Sciences, 37(1):96, January 1984. CODEN RHSAAM.

ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632360.

Aujac:1984:LFG

[2817] Germaine Aujac. Le langage formulaire dans la géométrie grecque. Revue d’Histoire des Sciences, 37(2):97–109, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632537.

Gardies:1984:ED

[2818] Jean-Louis Gardies. Eudoxe et Dedekind. Revue d’Histoire des Sciences, 37(2):111–125, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632538.

Blay:1984:CHP

[2819] Michel Blay. Christiaan Huygens et les phénomènes de la couleur. Revue d’Histoire des Sciences, 37(2):127–150, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632539.

Seidengart:1984:NTA

[2820] Jean Seidengart. Une nouvelle traduction anglaise de la ^{≪}Théorie du ciel ≫de Kant. Revue d’Histoire des Sciences, 37(2):151–155, avril 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632540.

Anonymous:1984:Ia

[2821] Anonyme. Informations. Revue d’Histoire des Sciences, 37(2):156–158, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632541.

Moscovici:1984:AOBa

[2822] Serge Moscovici. Analyses d’ouvrages: L’invention scientifique. Themata et interprétation, (coll. ^{≪}Croisées ^{≫}) par Gerald Holton; P. Scheurer. Revue d’Histoire des Sciences, 37(2):159, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632542.

Moscovici:1984:AOBb

[2823] Serge Moscovici. Analyses d’ouvrages: Disziplingeschichte als wissenschaftliche Selbstreflexion der historischen Wissenschaftsforschung, (coll. ^{≪}Europäische Hochschulschriften ^{≫}, Reihe XII, Bd. 7) par Wilfried Schroeder. Revue d’Histoire des Sciences, 37(2):160, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632543.

Rosmorduc:1984:AOBa

[2824] J. Rosmorduc. Analyses d’ouvrages: Psychogenèse et histoire des sciences, (coll. ^{≪}Nouvelle Bibliothèque scientifique ^{≫}) par Jean Piaget; Rolando Garcia. Revue d’Histoire des Sciences, 37(2):160–161, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632544.

Rosmorduc:1984:AOBb

[2825] Jean Rosmorduc. Analyses d’ouvrages: Histoires du temps par Jacques Attali. Revue d’Histoire des Sciences, 37(2):161, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632545.

Caveing:1984:AOBb

[2826] Maurice Caveing. Analyses d’ouvrages: Coloured Quadrangles. A guide to the Tenth Book of Euclid’s ^{≪}Elements ^{≫}, (^{≪}Opuscula Graecolatina ^{≫}, vol. 24) par C. M. Taisbak. Revue d’Histoire des Sciences, 37(2):161– 163, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632546.

Aujac:1984:AOB

[2827] G. Aujac. Analyses d’ouvrages: La géographie d’Orose, (coll. ^{≪}Études anciennes ^{≫}) par Yves Janvier. Revue d’Histoire des Sciences, 37(2):163– 164, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632547.

Ilic:1984:AOB

[2828] Myriana Ilic. Analyses d’ouvrages: Théorie des parallèles en Orient au Moyen Âge, IX–XIV^{e} siècles par B. A. Rosenfeld; A. P. Youschkevitch. Revue d’Histoire des Sciences, 37(2):164–165, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632548.

Theodorides:1984:AOBa

[2829] Jean Théodoridès. Analyses d’ouvrages: Briefwechsel zwischen Alexander von Humboldt und Peter Gustav Lejeune Dirichlet, (^{≪}Beiträge zur Alexander-von-Humboldt-Forschung ^{≫}, 7) par Kurt R. Biermann. Revue d’Histoire des Sciences, 37(2):165, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632549.

Dugac:1984:AOB

[2830] Pierre Dugac. Analyses d’ouvrages: Sarl’ Ermit [= Charles Hermite] par E. P. Ozigova. Revue d’Histoire des Sciences, 37(2):165–168, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632550.

Blay:1984:AOBa

[2831] Michel Blay. Analyses d’ouvrages: The Evolution of Dynamics; Vibration Theory from 1687 to 1742, (^{≪}Studies in the History of mathematics and physical sciences ^{≫}, 6) par John T. Cannon; Sigalia Dostrovsky. Revue d’Histoire des Sciences, 37(2):168–169, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632551.

Rosmorduc:1984:AOBc

[2832] J. Rosmorduc. Analyses d’ouvrages: L’aventure de l’électricité, (coll. ≪L’Odyssée ^{≫}) par Louis Leprince-Ringuet. Revue d’Histoire des Sciences, 37(2):169–170, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632552.

Mathieu:1984:AOBa

[2833] J.-P. Mathieu. Analyses d’ouvrages: La polarisation rotatoire naturelle, de la structure de la lumière à celle des molécules par Jean Rosmorduc. Revue d’Histoire des Sciences, 37(2):170–171, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632553.

Blondel:1984:AOB

[2834] Christine Blondel. Analyses d’ouvrages: The Papers of Joseph Henry, vol. III : January 1836–December 1837, The Princeton Years par Nathan Reingold. Revue d’Histoire des Sciences, 37(2):171, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632554.

Mathieu:1984:AOBb

[2835] J.-P. Mathieu. Analyses d’ouvrages: Gesamtausgabe, Bd. 8 : Ausgewählte Abhandlungen der internationalen Tagung, Wien, 1981 par Ludwig Boltzmann; R. Sexl; J. Blackmore. Revue d’Histoire des Sciences, 37(2):172, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633434.

deBeauregard:1984:AOB

[2836] O. Costa de Beauregard. Analyses d’ouvrages: The Tiger and the Shark, empirical roots of wave-particle dualism par Bruce Wheaton. Revue d’Histoire des Sciences, 37(2):173–174, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632555.

Goupil:1984:AOBa

[2837] Michelle Goupil. Analyses d’ouvrages: Memoir on Heat, Read to the Royal Academy of Sciences, June 28, 1783 by Messrs. Lavoisier & de la Place of the same Academy par A.-L. Lavoisier; P.-S. Laplace; Henry Guerlac. Revue d’Histoire des Sciences, 37(2):174–175, avril 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632556.

Goupil:1984:AOBb

[2838] Michelle Goupil. Analyses d’ouvrages: Poetry realized in nature; Samuel Taylor Coleridge and early nineteenth-century science par Trevor H. Levere. Revue d’Histoire des Sciences, 37(2):175–176, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632557.

Louis:1984:AOB

[2839] Pierre Louis. Analyses d’ouvrages: La classification des animaux chez Aristote. Statut de la biologie et unité de l’aristotélisme, (^{≪}Collection d’Études anciennes ^{≫}) par Pierre Pellegrin. Revue d’Histoire des Sciences, 37(2):176–177, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632558.

Grmek:1984:AOBa

[2840] Mirko D. Grmek. Analyses d’ouvrages: A Bio-bibliography for the History of the Biochemical Sciences since 1800 par Joseph S. Fruton. Revue d’Histoire des Sciences, 37(2):177–178, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632559.

Grmek:1984:AOBb

[2841] Mirko D. Grmek. Analyses d’ouvrages: Geschichte der Biologie, Theorien, Methoden, Institutionen und Kurzbiographien par Ilse Jahn; Rolf Löther; Konrad Senglaub. Revue d’Histoire des Sciences, 37(2):178–179, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632560.

Duminil:1984:AOB

[2842] M.-P. Duminil. Analyses d’ouvrages: Cinq cents ans de bibliographie hippocratique, 1473–1982 par Gilles Maloney; Raymond Savoie. Revue d’Histoire des Sciences, 37(2):179–180, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632561.

Grmek:1984:AOBc

[2843] Mirko D. Grmek. Analyses d’ouvrages: Medizin im mittelalterlichen Abendland, (Coll. ^{≪}Wege der Forschung ^{≫}, Bd. 363) par Gerhard Baader; Gundolf Keil. Revue d’Histoire des Sciences, 37(2):180–181, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632563.

Moulin:1984:AOBb

[2844] Anne-Marie Moulin. Analyses d’ouvrages: Pédiatrie des Lumières. Maladies et soins des enfants dans l’Encyclopédie et le Dictionnaire de Trévoux, Paris, Vrin, 1982 par D. Teysseire. Revue d’Histoire des Sciences, 37(2):181, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632565.

Huard:1984:AOB

[2845] Pierre Huard. Analyses d’ouvrages: Contribution à l’histoire de la sémiologie cardio-vasculaire au XIX^{e} siècle par Mme Thi Bang Tân Dang. Revue d’Histoire des Sciences, 37(2):181–183, avril 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632567.

Grmek:1984:AOBd

[2846] Mirko D. Grmek. Analyses d’ouvrages: Death is a social disease, Public health and political economy in early industrial France par William Coleman. Revue d’Histoire des Sciences, 37(2):183–184, avril 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632569.

Grmek:1984:AOBe

[2847] Mirko D. Grmek. Analyses d’ouvrages: Chirurgie im Wandel des Zeit 1945–1983 par H. W. Schreiber; G. Carstensen. Revue d’Histoire des Sciences, 37(2):184–185, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632571.

Payen:1984:AOBa

[2848] Jacques Payen. Analyses d’ouvrages: The Beam Engine. A massive chapter in the History of Steam par T. E. Crowley. Revue d’Histoire des Sciences, 37(2):185–186, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632573.

Payen:1984:AOBb

[2849] Jacques Payen. Analyses d’ouvrages: Technischer Fortschritt und Rationalisierung im antiken Bergbau, (^{≪}Habelts Dissertationsdrucke. Reihe alte Geschichte ^{≫}) par Peter Rosumek. Revue d’Histoire des Sciences, 37(2):186–187, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632575.

Pueyo:1984:AOBb

[2850] Guy Pueyo. Analyses d’ouvrages: No Sea too deep. The History of Oceanographic Instruments par Anita McConnell. Revue d’Histoire des Sciences, 37(2):187–189, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632577.

Moulin:1984:AOBc

[2851] Anne-Marie Moulin. Analyses d’ouvrages: Homo Caudatus, Les hommes à queue d’Afrique centrale : un avatar de l’imaginaire occidental par Jean-Dominique Penel. Revue d’Histoire des Sciences, 37(2):189, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632580.

Taton:1984:AOBa

[2852] René Taton. Analyses d’ouvrages: Le Jardin des Plantes de Montpellier, le plus ancien jardin de France, 1593. Revue d’Histoire des Sciences, 37 (2):189, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632578.

Hulin:1984:AOB

[2853] Nicole Hulin. Analyses d’ouvrages: The Radical Bourgeoisie. The Ligue de l’enseignement and the origins of the Third Republic 1866–1885 par Katherine Auspitz. Revue d’Histoire des Sciences, 37(2):190–191, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632582.

deValence:1984:AOBb

[2854] M. de Valence. Analyses d’ouvrages: ^{≪}Rudjer Bosković ^{≫}, Annales de l’Institut français de Zagreb, 3^{e} série, n^{o} 3. Revue d’Histoire des Sciences, 37(2):191, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632584.

deValence:1984:AOBc

[2855] M. de Valence. Analyses d’ouvrages: History and Technology, (c/o STBS, 42 William IV Street, London WC2N 4DE, England), vol. 1, n^{o} 1 par Pietro Redondi. Revue d’Histoire des Sciences, 37(2):191–192, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632586.

Evans:1984:FOP

[2856] James Evans. Fonction et origine probable du point équant de Ptolémée. Revue d’Histoire des Sciences, 37(3–4):193–213, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632654.

Kubbinga:1984:PTF

[2857] H. H. Kubbinga. Les premières théories ^{≪}moléculaires ^{≫}: Isaac Beeckman (1620) et Sébastien Basson (1621): Le concept d’ ^{≪}individu substantiel ^{≫}et d’ ^{≪}espèce substantielIe ^{≫}. Revue d’Histoire des Sciences, 37(3–4):215–233, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632655.

Lafuente:1984:QFT

[2858] Antonio Lafuente and Josè L. Peset. La question de la figure de la Terre: L’agonie d’un débat scientifique au XVIII^{e} siècle. Revue d’Histoire des Sciences, 37(3–4):235–254, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632656.

Belhoste:1984:JLC

[2859] Bruno Belhoste and Jesper Lützen. Joseph Liouville et le Collège de France. Revue d’Histoire des Sciences, 37(3–4):255–304, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632657.

Gaudant:1984:AAC

[2860] Jean Gaudant. Actualisme, antiprogressionnisme, catastrophisme et créationnisme dans l’œuvre d’Alcide d’Orbigny (1802–1857). Revue d’Histoire des Sciences, 37(3–4):305–312, juillet–décembre 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632658.

Taton:1984:PAG

[2861] René Taton, François Russo, Pierre Costabel, and Pietro Redondi. À propos de l’affaire Galilée. Revue d’Histoire des Sciences, 37(3–4): 313–324, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632659.

Fischer:1984:QLL

[2862] Jean-Louis Fischer. Quelques livres sur Lamarck et Darwin. Revue d’Histoire des Sciences, 37(3–4):325–333, juillet–décembre 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632660.

Taton:1984:MD

[2863] René Taton. Maurice Daumas (1910–1984). Revue d’Histoire des Sciences, 37(3–4):334–338, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632661.

Anonymous:1984:Ib

[2864] Anonyme. Informations. Revue d’Histoire des Sciences, 37(3–4):339– 346, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632662.

Bernhardt:1984:AOBa

[2865] Jean Bernhardt. Analyses d’ouvrages: Nature et méthode de l’histoire des sciences par François Russo. Revue d’Histoire des Sciences, 37(3–4): 347–349, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632663.

Taton:1984:AOBb

[2866] René Taton. Analyses d’ouvrages: Reference Books for the historian of science : a handlist, (^{≪}Occasional publications ^{≫}, 2) par S. A. Jayawardene. Revue d’Histoire des Sciences, 37(3–4):350, juillet–décembre 1984.

CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632664.

Taton:1984:AOBc

[2867] René Taton. Analyses d’ouvrages: La rigueur et le calcul — Documents historiques et épistémologiques. Revue d’Histoire des Sciences, 37(3–4): 350–351, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632665.

Bernhardt:1984:AOBb

[2868] Jean Bernhardt. Analyses d’ouvrages: Zénon d’Elée, Prolégomènes aux doctrines du continu, Étude historique et critique des Fragments et Témoignages. (^{≪}Histoire des Doctrines de l’Antiquité classique ^{≫}, n^{o} 7) par Maurice Caveing. Revue d’Histoire des Sciences, 37(3–4): 352–353, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632666.

Segonds:1984:AOB

[2869] A. Ph. Segonds. Analyses d’ouvrages: Mysterium Cosmographicum. The Secret of the Universe par Johannes Kepler; A. M. Duncan; E. J. Aiton; I. Bernard Cohen. Revue d’Histoire des Sciences, 37(3–4):354–355, juillet– décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632667.

Bernhardt:1984:AOBc

[2870] Jean Bernhardt. Analyses d’ouvrages: Principles of Philosophy, (^{≪}Collection des Travaux de l’Académie internationale d’Histoire des Sciences ^{≫}, n^{o} 30) par René Descartes; Valentine Rodger Miller; Reese P. Miller. Revue d’Histoire des Sciences, 37(3–4):355, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632668.

Blay:1984:AOBb

[2871] Michel Blay. Analyses d’ouvrages: Roger Cotes, Natural Philosopher par Ronald Gowing. Revue d’Histoire des Sciences, 37(3–4):355–356, juillet– décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632669.

Taton:1984:AOBd

[2872] René Taton. Analyses d’ouvrages: Tobias Mayer (1723–62), Pioneer of enlightened Science in Germany, (^{≪}Arbeiten aus der Niedersächsischen Staats- und Universitäts–Bibliothek Göttingen ^{≫}, Bd. 17) par Eric G. Forbes. Revue d’Histoire des Sciences, 37(3–4):357–358, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632670.

Blay:1984:AOBc

[2873] Michel Blay. Analyses d’ouvrages: Commentationes Mechanicae ad theoriam Machinarum Pertinentes, volumen tertium, (Leonhardi Euleri Opera Omnia, Series secunda, Opera Mechanica et Astronomica, vol. XVII) par Leonhardi Euleri; Charles Blanc; Pierre de Haller. Revue d’Histoire des Sciences, 37(3–4):358–360, juillet–décembre 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632671.

Taton:1984:AOBe

[2874] René Taton. Analyses d’ouvrages: Opere matematiche della pubblica Biblioteca di Ferrara (1753–1815) par M. T. Borgato; L. Capra; A. Fiocca; L. Pepe. Revue d’Histoire des Sciences, 37(3–4):360–361, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632672.

Taton:1984:AOBf

[2875] René Taton. Analyses d’ouvrages: A Bibliography of early modern Algebra, 1500–1800, (^{≪}Berkeley Papers in History of Science ^{≫}, VII) par Robin E. Rider. Revue d’Histoire des Sciences, 37(3–4):361–363, juillet– décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632673.

Guillemot:1984:AOB

[2876] Michel Guillemot. Analyses d’ouvrages: Zermelo’s axiom of choice : Its origins, development, and influence, (^{≪}Studies in the History of Mathematics and Physical Sciences ^{≫}, 8) par Gregory H. Moore. Revue d’Histoire des Sciences, 37(3–4):363–364, juillet–décembre 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632674.

Bru:1984:AOBa

[2877] Bernard Bru. Analyses d’ouvrages: The Correspondence between A. A. Markov and A. A. Chuprov on the theory of probability and mathematical statistics par Kh. O. Ondar; Charles Stein; Margaret Stein. Revue d’Histoire des Sciences, 37(3–4):365, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632675.

Bru:1984:AOBb

[2878] Bernard Bru. Analyses d’ouvrages: The making of statisticians par J. Gani. Revue d’Histoire des Sciences, 37(3–4):365–366, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632676.

Rosmorduc:1984:AOBd

[2879] Jean Rosmorduc. Analyses d’ouvrages: A.-M. Ampère et la création de l’électrodynamique, 1820–1827, (^{≪}Comité des Travaux historiques et scientifiques. Mémoires de la Section des Sciences^{≫}, 10) par Christine Blondel. Revue d’Histoire des Sciences, 37(3–4):366–368, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632677.

Taton:1984:AOBg

[2880] René Taton. Analyses d’ouvrages: Grandes découvertes de l’astronomie. Ciel passé, présent par Gilbert Walusinski. Revue d’Histoire des Sciences, 37(3–4):368, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632678.

Taton:1984:AOBh

[2881] René Taton. Analyses d’ouvrages: Calendar of the correspondence of Pierre-Simon Laplace, (^{≪}Berkeley Papers in History of Science ^{≫}, VIII) par Roger Hahn. Revue d’Histoire des Sciences, 37(3–4):369, juillet– décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632679.

Theodorides:1984:AOBb

[2882] Jean Théodoridès. Analyses d’ouvrages: Leopoldo M. A. Caldani — Lazzaro Spallanzani, Carteggio (1768–1798), (^{≪}Fonti e Studi per la storia dell’Università di Pavia ^{≫}, n^{o} 4) par Giuseppe Ongaro. Revue d’Histoire des Sciences, 37(3–4):369–370, juillet–décembre 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632680.

Guedes:1984:AOB

[2883] Michel Guédès. Analyses d’ouvrages: The Poison Tree. Selected writings of Rumphius on the natural history of the Indies par G. E. Rumpf; E. M. Beekman. Revue d’Histoire des Sciences, 37(3–4):370–371, juillet– décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632681.

Grmek:1984:AOBf

[2884] M. D. Grmek. Analyses d’ouvrages: Histoire de la médecine aux armées, Tome 1 : De l’Antiquité à la Révolution par Jean Guillermand. Revue d’Histoire des Sciences, 37(3–4):371–373, juillet–décembre 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632682.

Boulle:1984:AOBb

[2885] Lydie Boulle. Analyses d’ouvrages: Wien von den Anfängen bis um 1900, (^{≪}Geschichte des Hospitals ^{≫}, Bd. 5) par Dieter Jetter. Revue d’Histoire des Sciences, 37(3–4):373–374, juillet–décembre 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632683.

Grmek:1984:AOBg

[2886] M. D. Grmek. Analyses d’ouvrages: Hughlings Jackson on psychiatry par Kenneth Dewhurst. Revue d’Histoire des Sciences, 37(3–4):374, juillet– décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632684.

Anonymous:1984:TGM

[2887] Anonyme. Table générale des matières du Tome XXXVII. Revue d’Histoire des Sciences, 37(3–4):375–384, juillet–décembre 1984. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632685.

Anonymous:1984:BMa

[2888] Anonyme. Back matter. Revue d’Histoire des Sciences, 37(1):??, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632361.

Anonymous:1984:BMb

[2889] Anonyme. Back matter. Revue d’Histoire des Sciences, 37(2):??, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632588.

Anonymous:1984:BMc

[2890] Anonyme. Back matter. Revue d’Histoire des Sciences, 37(3–4): ??, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632686.

Anonymous:1984:FMa

[2891] Anonyme. Front matter. Revue d’Histoire des Sciences, 37(1):??, January 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632339.

Anonymous:1984:FMb

[2892] Anonyme. Front matter. Revue d’Histoire des Sciences, 37(2):??, avril 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632536.

Anonymous:1984:FMc

[2893] Anonyme. Front matter. Revue d’Histoire des Sciences, 37(3–4): ??, juillet–décembre 1984. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632653.

Fauque:1985:AMR

[2894] Danielle Fauque. Alexis-Marie Rochon (1741–1817), savant astronome et opticien. Revue d’Histoire des Sciences, 38(1):3–36, January 1985. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632596.

Blay:1985:NFB

[2895] Michel Blay. Note sur l’^{≪}Essai sur les degrés de chaleur des rayons colorés ^{≫}de l’abbé Rochon. Revue d’Histoire des Sciences, 38(1):37–42, January 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632597.

Jacques:1985:FCC

[2896] Jean Jacques. Le ^{≪}Cours de chimie de G.-F. Rouelle recueilli par Diderot ≫. Revue d’Histoire des Sciences, 38(1):43–53, January 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632598.

Viel:1985:DDM

[2897] Claude Viel. Duhamel du Monceau, naturaliste, physicien et chimiste. Revue d’Histoire des Sciences, 38(1):55–71, January 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632599.

Bayle:1985:NCT

[2898] Nadia Bayle and Claudine Billoux. Le nouveau classement thématique des archives de l’École Polytechnique. Revue d’Histoire des Sciences, 38(1):73–82, January 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632600.

Anonymous:1985:Ia

[2899] Anonyme. Informations. Revue d’Histoire des Sciences, 38(1):83–86, January 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632601.

Goupil:1985:AOB

[2900] Michelle Goupil. Analyses d’ouvrages: Théories électriques du XVIII^{e} siècle. Correspondance entre l’abbé Nollet (1700–1770) et le physicien genevois Jean Jallabert (1712–1768) par Isaac Benguigui. Revue d’Histoire des Sciences, 38(1):87–88, January 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632602.

Belhoste:1985:AOBa

[2901] Bruno Belhoste. Analyses d’ouvrages: Le Corps des Ponts et Chaussées, Histoire de l’administration française par A. Brunot; R. Coquand. Revue d’Histoire des Sciences, 38(1):88–89, January 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632603.

Payen:1985:AOB

[2902] Jacques Payen. Analyses d’ouvrages: L’expérience de Jouffroy d’Abbans en 1783 et la navigation à vapeur dans la région lyonnaise (catalogue d’exposition) par Musée historique de Lyon; Claude François Dorothée de Jouffroy d’Abbans, 1751–1832, (^{≪}Histoire et Archéologie ^{≫}) par E. Bouton. Revue d’Histoire des Sciences, 38(1):89–91, January 1985. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632604.

Grmek:1985:AOBa

[2903] M. D. Grmek. Analyses d’ouvrages: Girolamo Cardano, 1501–1576. Physician, natural philosopher, mathematician, astrologer and interpreter of dreams par Markus Fierz; Helga Niman. Revue d’Histoire des Sciences, 38(1):91, January 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632605.

Taton:1985:AOBa

[2904] R. Taton. Analyses d’ouvrages: Rupture et permanence dans les civilisations occidentales. Revue d’Histoire des Sciences, 38(1):91–92, January 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632606.

Belhoste:1985:AOBb

[2905] Bruno Belhoste. Analyses d’ouvrages: The Making of technological Man, the social origins of French engineering education par John Hobbel Weiss; D. S. Landes. Revue d’Histoire des Sciences, 38(1):92–94, January 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632607.

Anonymous:1985:PR

[2906] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 38(1):95– 96, January 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632608.

deBuzon:1985:SNT

[2907] Frédéric de Buzon. Science de la nature et théorie musicale chez Isaac Beeckman (1588–1637). Revue d’Histoire des Sciences, 38(2):97–120, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632510.

Scheler:1985:DLI

[2908] Lucien Scheler. Deux lettres inédites de Mme Lavoisier. Revue d’Histoire des Sciences, 38(2):121–130, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632511.

Costabel:1985:NFE

[2909] Pierre Costabel. Notes fugitives sur l’équation du troisième degré dans la mathématique occidentale du XVI^{e} au XVIII^{e} siècle. Revue d’Histoire des Sciences, 38(2):131–147, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632512.

Fauque:1985:DCA

[2910] Danielle Fauque. Il y a deux cents ans : l’expédition Lapérouse. Revue d’Histoire des Sciences, 38(2):149–160, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632513.

Bernhardt:1985:JJ

[2911] Jean Bernhardt. Jean Jacquot (1909–1983). Revue d’Histoire des Sciences, 38(2):161–162, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632514.

Anonymous:1985:Ib

[2912] Anonyme. Informations. Revue d’Histoire des Sciences, 38(2):163–165, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632515.

Chappert:1985:AOB

[2913] André Chappert. Analyses d’ouvrages: La conceptualisation newtonienne des phénomènes de la couleur, (coll.^{≪}L’histoire des sciences, Textes et études ^{≫}) par Michel Blay. Revue d’Histoire des Sciences, 38(2):167– 168, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632516.

Duminil:1985:AOB

[2914] Marie-Paule Duminil. Analyses d’ouvrages: Note di lessicografia ippocratica. Il trattato ^{≪}Sulle arie, le acque, i luoghi ^{≫}, (^{≪}Lessico Intellettuale Europeo ^{≫}, XXVIII) par Andrea Bozzi. Revue d’Histoire des Sciences, 38(2):168–169, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632517.

Hulin:1985:AOB

[2915] Nicole Hulin. Analyses d’ouvrages: Les collèges français, XVI^{e}–XVIII^{e} siècles. Répertoire 1: France du Midi par Marie-Madeleine Compère; Dominique Julia. Revue d’Histoire des Sciences, 38(2):169–170, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632518.

Marganne:1985:AOB

[2916] Marie-Hélène Marganne. Analyses d’ouvrages: Le sang, les vaisseaux, le coeur dans la Collection hippocratique. Anatomie et physiologie, ^{≪}Les Belles Lettres ^{≫}par M.-P. Duminil. Revue d’Histoire des Sciences, 38 (2):170–172, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632519.

Grmek:1985:AOBb

[2917] M. D. Grmek. Analyses d’ouvrages: Medizin vor Gericht. Gerichtsmedizin von der Renaissance bis zur Aufklärung par Esther Fischer-Homberger. Revue d’Histoire des Sciences, 38(2):172–173, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632520.

Mathieu:1985:AOB

[2918] J.-P. Mathieu. Analyses d’ouvrages: The Demon in the Aether. The story of James Clerk Maxwell par Martin Goldman. Revue d’Histoire des Sciences, 38(2):173, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632521.

Jacquart:1985:AOB

[2919] Danielle Jacquart. Analyses d’ouvrages: Les maladies à l’aube de la civilisation occidentale, Recherches sur la réalité pathologique dans le monde grec préhistorique, archaïque et classique, (^{≪}Médecine et Sociétés ≫, coll. dirigée par Marc Ferro) par Mirko D. Grmek. Revue d’Histoire des Sciences, 38(2):173–175, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632522.

Gourevitch:1985:AOB

[2920] Michel Gourevitch. Analyses d’ouvrages: Mental Illness and American Society, 1875–1940 par Gerald N. Grob. Revue d’Histoire des Sciences, 38(2):175–176, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632523.

LHuillier:1985:AOB

[2921] Hervé L’Huillier. Analyses d’ouvrages: Medieval Mensuration : ≪Quadrans Vetus ^{≫}and ^{≪}Geometrie Due Sunt Partes Principales … ^{≫}, (^{≪}Transactions of the American Philosophical Society ^{≫}, vol. 72, part 8, 1982) par Nan L. Hahn. Revue d’Histoire des Sciences, 38(2):176–177, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632524.

Lasota:1985:AOB

[2922] J.-P. Lasota. Analyses d’ouvrages: Glimpsing an invisible Universe. The emergence of X-ray astronomy par Richard F. Hirsh. Revue d’Histoire des Sciences, 38(2):177–179, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632525.

deBeauregard:1985:AOB

[2923] O. Costa de Beauregard. Analyses d’ouvrages: Papers on Probability, Statistics and Statistical Physics par E. T. Jaynes; R. D. Rosenkrantz. Revue d’Histoire des Sciences, 38(2):179–180, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632526.

Dieudonne:1985:AOB

[2924] J. Dieudonné. Analyses d’ouvrages: The nature and growth of modern mathematics, Princeton par Edna E. Kramer. Revue d’Histoire des Sciences, 38(2):180–181, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632527.

Blay:1985:AOBa

[2925] Michel Blay. Analyses d’ouvrages: Sir Francis Bacon. La Nouvelle Atlantide suivi de Voyage dans la pensée baroque par Michèle Le Doeuff; Margaret Llasera. Revue d’Histoire des Sciences, 38(2):181–182, avril– juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632528.

Pueyo:1985:AOB

[2926] Guy Pueyo. Analyses d’ouvrages: Diccionario histórico de la ciencia moderna en España par José M. López Piñero; Thomas F. Glick; Victor Navarro Brotóns; Eugenio Portela Marco. Revue d’Histoire des Sciences, 38(2):183–184, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632529.

Taton:1985:AOBb

[2927] R. Taton. Analyses d’ouvrages: A travers le Cosmos par Chanoine Henri Louyat. Revue d’Histoire des Sciences, 38(2):184, avril–juin 1985. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632530.

Demidov:1985:AOB

[2928] S. Demidov. Analyses d’ouvrages: The prehistory of the theory of distributions, (^{≪}Studies in the History of Mathematics and Physical Sciences ≫, 7) par Jesper Lützen. Revue d’Histoire des Sciences, 38(2):185–187, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632531.

Turner:1985:AOB

[2929] A. J. Turner. Analyses d’ouvrages: Les instruments astronomiques du Moyen Âge, (^{≪}Astrolabica ^{≫}, 3) par Emmanuel Poulle. Revue d’Histoire des Sciences, 38(2):187–188, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632532.

Anonymous:1985:LOR

[2930] Anonyme. Liste des ouvrages reçus. Revue d’Histoire des Sciences, 38(2):189–190, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632533.

Blay:1985:BC

[2931] Michel Blay. Bohr et la complémentarité. (French) [Bohr and complementarity]. Revue d’Histoire des Sciences, 38(3–4):193–194, juillet–décembre 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632496;   http://www.persee.fr/web/revues/home/prescript/
article/rhs_0151-4105_1985_num_38_3_4004.

Heilbron:1985:EMC

[2932] John L. Heilbron. The earliest missionaries of the Copenhagen spirit. Revue d’Histoire des Sciences, 38(3–4):195–230, juillet–décembre 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632497;
http://www.persee.fr/web/revues/home/prescript/article/rhs_
0151-4105_1985_num_38_3_4005.

Bensaude-Vincent:1985:LCT

[2933] Bernadette Bensaude-Vincent. L’évolution de la complémentarité dans les textes de Bohr (1927–1939). (French) [The evolution of complementarity in Bohr’s papers (1927–1939)]. Revue d’Histoire des Sciences, 38(3–4):231–250, juillet–décembre 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632498;   http:
//www.persee.fr/web/revues/home/prescript/article/rhs_0151-
4105_1985_num_38_3_4006.

Chevalley:1985:CLI

[2934] Catherine Chevalley. Complémentarité et langage dans l’interprétation de Copenhague. (French) [Complementarity and language in the Copenhagen interpretation]. Revue d’Histoire des Sciences, 38(3–4):251– 292, juillet–décembre 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:/

/www.jstor.org/stable/23632499;   http://www.persee.fr/web/
revues/home/prescript/article/rhs_0151-4105_1985_num_38_3_4007.

Balibar:1985:BEE

[2935] Françoise Balibar. Bohr entre Einstein et Dirac. (French) [Bohr between Einstein and Dirac]. Revue d’Histoire des Sciences, 38(3–4): 293–307, juillet–décembre 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:/

/www.jstor.org/stable/23632500;   http://www.persee.fr/web/
revues/home/prescript/article/rhs_0151-4105_1985_num_38_3_4008.

Darrigol:1985:CCA

[2936] Olivier Darrigol. La complémentarité comme argument d’autorité (1927– 1934). (French) [Complementarity as an argument of authority (1927– 1934)]. Revue d’Histoire des Sciences, 38(3–4):309–323, juillet–décembre 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632501;   http://www.persee.fr/web/revues/home/prescript/
article/rhs_0151-4105_1985_num_38_3_4009.

Paty:1985:ECA

[2937] Michel Paty. Einstein et la complémentarité au sens de Bohr: du retrait dans le tumulte aux arguments d’incomplétude. (French) [Einstein and complementarity according to Bohr: withdrawal in the tumult of the incompleteness arguments]. Revue d’Histoire des Sciences, 38(3–4):325–351, juillet–décembre 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632502;   http:
//www.persee.fr/web/revues/home/prescript/article/rhs_0151-
4105_1985_num_38_3_4010.

Anonymous:1985:BSR

[2938] Anonyme. Bibliographie se rapportant au concept de complémentarité et à ses prolongements. (French) [Bibliography concerning the concept of complementarity and its extensions]. Revue d’Histoire des Sciences, 38(3–4):353–363, juillet–décembre 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632503;   http:
//www.persee.fr/web/revues/home/prescript/article/rhs_0151-
4105_1985_num_38_3_4012.

deBeauregard:1985:ODA

[2939] O. Costa de Beauregard. A l’occasion du 90^{e} anniversaire de Louis de Broglie. (French) [On the occasion of the 90th birthday of Louis de Broglie]. Revue d’Histoire des Sciences, 38(3–4):365–367, juillet– décembre 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632504.

Blay:1985:I

[2940] Michel Blay. Informations. Revue d’Histoire des Sciences, 38(3–4): 369–374, juillet–décembre 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632505.

Postel:1985:AOB

[2941] Jacques Postel. Analyses d’ouvrages: Psicoanalisi e storia delle scienze. (^{≪}Gabinetto Scientifico Letteràrio G. P. Vieusseux ^{≫}) par Michele Ranchetti. Revue d’Histoire des Sciences, 38(3–4):375–377, juillet– décembre 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632506.

Blay:1985:AOBb

[2942] Michel Blay. Analyses d’ouvrages: The Show of Science, The Friends of the Bancroft Library, (^{≪}Series of Keepsakes issued par the Friends of the Bancroft Library ^{≫}, n^{o} 31) par Robin E. Rider. Revue d’Histoire des Sciences, 38(3–4):377–378, juillet–décembre 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632507.

Anonymous:1985:BMa

[2943] Anonyme. Back matter. Revue d’Histoire des Sciences, 38(1):??, January 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632609.

Anonymous:1985:BMb

[2944] Anonyme. Back matter. Revue d’Histoire des Sciences, 38(2):??, avril– juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632535.

Anonymous:1985:BMc

[2945] Anonyme. Back matter. Revue d’Histoire des Sciences, 38(3–4): ??, juillet–décembre 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632508.

Anonymous:1985:FMa

[2946] Anonyme. Front matter. Revue d’Histoire des Sciences, 38(1):??, January 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632595.

Anonymous:1985:FMb

[2947] Anonyme. Front matter. Revue d’Histoire des Sciences, 38(2):??, avril– juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632509.

Anonymous:1985:FMc

[2948] Anonyme. Front matter. Revue d’Histoire des Sciences, 38(3–4): ??, juillet–décembre 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632495.

Anonymous:1985:LCB

[2949] Anonyme. Liste des catalogues, bibliographies et ouvrages de référence reçus. Revue d’Histoire des Sciences, 38(2):??, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632534.

Nardi:1986:DFP

[2950] Antonio Nardi. Descartes ^{≪}presque ^{≫}galiléen : 18 février 1643. Revue d’Histoire des Sciences, 39(1):3–16, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632564.

Terrier:1986:LRV

[2951] Max Terrier. L’invention des ressorts de voiture. Revue d’Histoire des Sciences, 39(1):17–30, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632566.

Gaudant:1986:CAP

[2952] Jean Gaudant. Camille Arambourg (1885–1969), précurseur du ponctualisme. Revue d’Histoire des Sciences, 39(1):31–34, January 1986. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632568.

Bernes:1986:CRFa

[2953] Anne-Catherine Bernés and Pascal Lefébvre. La correspondance de René-François de Sluse: Essai de répertoire chronologique : première partie (1649–1664). Revue d’Histoire des Sciences, 39(1):35–69, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632570.

Berthon:1986:PCA

[2954] Pierre Berthon. Les plis cachetés de l’Académie des Sciences. (French) [The sealed envelopes of the Academy of Sciences]. Revue d’Histoire des Sciences, 39(1):71–78, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632572.

Theodorides:1986:PPG

[2955] Jean Théodoridès. Pierre-Paul Grassé (1895–1985). Revue d’Histoire des Sciences, 39(1):79–82, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632574.

Anonymous:1986:Ia

[2956] Anonyme. Informations. Revue d’Histoire des Sciences, 39(1):83–84, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632576.

Debru:1986:AOB

[2957] Claude Debru. Analyses d’ouvrages: Evolution from Molecules to Men par D. S. Bendall. Revue d’Histoire des Sciences, 39(1):85–87, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632579.

Mathieu:1986:AOBa

[2958] J.-P. Mathieu. Analyses d’ouvrages: Die Natur als arbeitende Maschine. Zur Entstehung des Energiebegriffs in der Physik (1840–1850), (^{≪}Campus Forschung ^{≫}, Bd. 296) par Herbert Breger. Revue d’Histoire des Sciences, 39(1):87, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632581.

Mathieu:1986:AOBb

[2959] J.-P. Mathieu. Analyses d’ouvrages: Werner Heisenberg. A bibliography of his writings, (^{≪}Berkeley Papers in History of Science ^{≫}, IX) par David Cassidy; Martha Baker. Revue d’Histoire des Sciences, 39(1):87–88, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632583.

Mathieu:1986:AOBc

[2960] J.-P. Mathieu. Analyses d’ouvrages: The beginnings of the Nobel institution. The Science Prizes, 1901–1915 par Elisabeth Crawford. Revue d’Histoire des Sciences, 39(1):88, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632585.

Grmek:1986:AOBa

[2961] Mirko D. Grmek. Analyses d’ouvrages: Disease in Ancient Man. An International Symposium par Gerald D. Hart. Revue d’Histoire des Sciences, 39(1):88–90, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632587.

Lagarde:1986:AOB

[2962] Lucie Lagarde. Analyses d’ouvrages: Les Portulans, cartes marines du XIII^{e} au XVII^{e} siècle … par Monique de la Roncière; Michel Mollat du Jourdin; M.-M. Azard; I. Raynaud-Nguyen; M.-A. Vannereau. Revue d’Histoire des Sciences, 39(1):90–91, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632589.

Grmek:1986:AOBb

[2963] M. D. Grmek. Analyses d’ouvrages: Richard Lower’s Vindicatio. A defence of the experimental method par Kenneth Dewhurst. Revue d’Histoire des Sciences, 39(1):92, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632590.

Bernhardt:1986:AOB

[2964] Jean Bernhardt. Analyses d’ouvrages: Correspondance du P. Marin Mersenne, religieux minime, t. XV, 1647 par Paul Tannery; Cornelis De Waard; Armand Beaulieu. Revue d’Histoire des Sciences, 39(1):92–93, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632591.

Thillaud:1986:AOB

[2965] P. L. Thillaud. Analyses d’ouvrages: M. Seoane. La introducción en España del sistema sanitario liberal (1791–1870), (coll. ^{≪}Textos clásicos españoles de la Salud pública ^{≫}, vol. 12) par J. M. Lopez Piñero. Revue d’Histoire des Sciences, 39(1):93–94, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632592.

Anonymous:1986:PRa

[2966] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 39(1):95– 96, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632593.

Dahan-Dalmedico:1986:TPM

[2967] Amy Dahan-Dalmédico. Un texte de philosophie mathématique de Gergonne (mémoire inédit déposé à l’Académie de Bordeaux). Revue d’Histoire des Sciences, 39(2):97–126, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632718.

Debarbat:1986:NHO

[2968] Suzanne Débarbat. Newton, Halley et l’Observatoire de Paris. Revue d’Histoire des Sciences, 39(2):127–154, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632719.

Bernes:1986:CRFb

[2969] Anne-Catherine Bernès and Pascal Lefèbvre. La correspondance de René-François de sluse: Essai de répertoire chronologique : deuxième partie (1665–1668). Revue d’Histoire des Sciences, 39(2):155–175, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632720.

Cazenobe:1986:EHR

[2970] Jean Cazenobe. Études sur l’histoire de la radio. Revue d’Histoire des Sciences, 39(2):177–180, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632721.

Anonymous:1986:Ib

[2971] Anonyme. Informations. Revue d’Histoire des Sciences, 39(2):181–182, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632722.

Mathieu:1986:AOBd

[2972] J.-P. Mathieu. Analyses d’ouvrages: Galilée, Newton lus par Einstein. Espace et relativité, (Coll. ^{≪}Philosophies ^{≫}) par Françoise Balibar. Revue d’Histoire des Sciences, 39(2):183, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632723.

deBeauregard:1986:AOBa

[2973] O. Costa de Beauregard. Analyses d’ouvrages: La construction de la notion de temps, 3 t. : 1. Genèse anthropologique de la notion; 2. Conditionnement bio-psychologique de la genèse représentative; 3. Constitution rationnelle du concept par Hervé Barreau. Revue d’Histoire des Sciences, 39(2):183–184, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632724.

Hulin:1986:AOB

[2974] Nicole Hulin. Analyses d’ouvrages: La presse d’éducation et d’enseignement, XVIII^{e} siècle-1940, (t. I : A-C, 1981 ; t. II : D-J, 1984) par P. Caspard. Revue d’Histoire des Sciences, 39(2):184, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632725.

Grmek:1986:AOBc

[2975] M.-D. Grmek. Analyses d’ouvrages: Histoire de la médecine aux armées, t. 2 : De la Révolution française au conflit mondial de 1914 par Albert Fabre. Revue d’Histoire des Sciences, 39(2):185, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632726.

Beaulieu:1986:AOB

[2976] A. Beaulieu. Analyses d’ouvrages: Le ^{≪}Giornale de’ Letterati ^{≫}de Rome (1668–1681), (Accademia Toscana di Scienze e Lettere ^{≪}La Colombaria ≫, Studi LXIX) par Jean-Michel Gardair; Leo S. Olschki Firenze. Revue d’Histoire des Sciences, 39(2):186–188, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632727.

Mathieu:1986:AOBe

[2977] J.-P. Mathieu. Analyses d’ouvrages: Inner Exile. Recollections of a life with Werner Heisenberg par Elisabeth Heisenberg; S. Cappellari; C. Morris; V. Weisskopf. Revue d’Histoire des Sciences, 39(2):188, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632728.

Theodorides:1986:AOBa

[2978] Jean Théodoridès. Analyses d’ouvrages: Felice Fontana, life and works par Peter K. Knoefel. Revue d’Histoire des Sciences, 39(2):189–190, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632729.

Mathieu:1986:AOBf

[2979] J.-P. Mathieu. Analyses d’ouvrages: Imagery in Scientific Thought creating 20th Century Physics. A ^{≪}Pro Scientia Viva ^{≫}title par Arthur I. Mller. Revue d’Histoire des Sciences, 39(2):190, avril–juin 1986. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632730.

Anonymous:1986:PRb

[2980] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 39(2):191– 192, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632731.

DeGandt:1986:SMB

[2981] François De Gandt. Le style mathématique des “Principia” de Newton. Revue d’Histoire des Sciences, 39(3):195–222 (1987), juillet–septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632481. Études sur

l’histoire du calcul infinitésimal.

Blay:1986:DMC

[2982] Michel Blay. Deux moments de la critique du calcul infinitésimal : Michel rolle et George Berkeley. Revue d’Histoire des Sciences, 39(3):223–253 (1987), juillet–septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/ 23632482. Études sur l’histoire du calcul infinitésimal.

Seidengart:1986:NDP

[2983] Jean Seidengart and Sylvain Zac. Noumène et différentielle dans la philosophie de Salomon Maïmon. (French) [Noumène and differential in the philosophy of Salomon Maïmon]. Revue d’Histoire des Sciences, 39 (3):255–272, juillet–septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632483.

Gillispie:1986:FSP

[2984] Charles C. Gillispie. Le fonds seguin à privas. Revue d’Histoire des Sciences, 39(3):273–275, juillet–septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632484.

Costabel:1986:NKV

[2985] Pierre Costabel. Nécrologie: Kurt Vogel (1888–1985). Revue d’Histoire des Sciences, 39(3):277–278, juillet–septembre 1986. CODEN RHSAAM.

ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632485.

Theodorides:1986:I

[2986] Jean Théodoridès. Informations. Revue d’Histoire des Sciences, 39(3): 279–280, juillet–septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632486.

Jacques:1986:AOBa

[2987] Jean Jacques. Analyses d’ouvrages: Justus von Liebig und August Wilhelm Hofmann in ihren Briefen (1841–1873) par W. H. Brock. Revue d’Histoire des Sciences, 39(3):281, juillet–septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632487.

Goupil:1986:AOB

[2988] Michelle Goupil. Analyses d’ouvrages: Amedeo Avogadro : A scientific Biography, (^{≪}Chemists and Chemistry ^{≫}) par Mario Morselli. Revue d’Histoire des Sciences, 39(3):282–283, juillet–septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632488.

Grmek:1986:AOBd

[2989] M. D. Grmek. Analyses d’ouvrages: Eine medizinische Erklärung der männlichen Homosexualität aus der Antike par P. H. Schrijvers. Revue d’Histoire des Sciences, 39(3):283, juillet–septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632489.

Russo:1986:AOB

[2990] François Russo. Analyses d’ouvrages: A Theory of Method par Husain Sarkar. Revue d’Histoire des Sciences, 39(3):283–284, juillet–septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632490.

Guipaud:1986:AOB

[2991] C. Guipaud and R. Laurent. Analyses d’ouvrages: I sei libri della prospettiva di Guidobaldo … marchesi del monte par Rocco Sinisgalli; Gaspare De Fiore. Revue d’Histoire des Sciences, 39(3):284–286, juillet–septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632491.

Merleau-Ponty:1986:AOB

[2992] Jacques Merleau-Ponty. Analyses d’ouvrages: Cosmology and Astrophysics, Essays in honor of Thomas Gold par Yervant Terzian; Elizabeth M. Bilson. Revue d’Histoire des Sciences, 39(3):286–287, juillet– septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632492.

Postel:1986:AOB

[2993] Jacques Postel. Analyses d’ouvrages: A generous confidence. Thomas Story Kirkbride and the art of Asylum-keeping, 1840–1883 (^{≪}Cambridge History of Medicine ^{≫}) par Nancy Tomes. Revue d’Histoire des Sciences, 39(3):287–288, juillet–septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632493.

Taton:1986:CHS

[2994] René Taton. Sur la ^{≪}comète de Halley ^{≫}et son ^{≪}retour ^{≫}en 1986. Revue d’Histoire des Sciences, 39(4):289–300, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632432.

Bailhache:1986:VAA

[2995] Patrice Bailhache. Valeur actuelle de l’acoustique musicale de Helmholtz. Revue d’Histoire des Sciences, 39(4):301–324, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632433.

Bernes:1986:CRF

[2996] Anne-Catherine Bernès and Pascal Lefèbvre. La correspondance de René-François de sluse: Essai de répertoire chronologique : troisième partie (1669–1685). Revue d’Histoire des Sciences, 39(4):325–344, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632434.

Fauque:1986:ESA

[2997] Danielle Fauque. Une élégante solution au problème des longitudes : les horloges marines de Ferdinand Berthoud. Revue d’Histoire des Sciences, 39(4):345–350, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632435.

Debarbat:1986:I

[2998] Suzanne Débarbat. Informations. Revue d’Histoire des Sciences, 39 (4):351–356, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632436.

Largeault:1986:AOB

[2999] J. Largeault. Analyses d’ouvrages: La philosophie des sciences aujourd’hui par J. Hamburger; René Thom; Anatole Abragam; Pierre Jacob; Gilles Granger; Daniel Andler; Antoine Danchin. Revue d’Histoire des Sciences, 39(4):357–358, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632437.

Debarbat:1986:AOBa

[3000] Suzanne Débarbat. Analyses d’ouvrages: Astrophysics and Twentieth-Century astronomy to 1950, Vol. 4 A de The General History of Astronomy par M. Hoskin; O. Gingerich. Revue d’Histoire des Sciences, 39(4):358–360, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632438.

deBeauregard:1986:AOBb

[3001] O. Costa de Beauregard. Analyses d’ouvrages: On Time. An investigation into scientific knowledge and human experience par Michael Shallis. Revue d’Histoire des Sciences, 39(4):361, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632439.

Mathieu:1986:AOBg

[3002] J.-P. Mathieu. Analyses d’ouvrages: Physical Sciences and History of Physics, (^{≪}Boston Studies in the Philosophy of Science ^{≫}, vol. 82) par R. S. Cohen; M. W. Wartofsky. Revue d’Histoire des Sciences, 39(4):362, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632440.

Blanckaert:1986:AOB

[3003] Claude Blanckaert. Analyses d’ouvrages: Archetypes and Ancestors : Palaeontology in Victorian London, 1850–1875 par Adrian Desmond. Revue d’Histoire des Sciences, 39(4):362–364, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632441.

Payen:1986:AOB

[3004] Jacques Payen. Analyses d’ouvrages: The Montgolfier brothers and the invention of aviation, 1783–1784 … par Charles Coulston Gillispie; La part du rêve : de la montgolfière au satellite, 1783–1983. Revue d’Histoire des Sciences, 39(4):364–366, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632442.

Debarbat:1986:AOBb

[3005] Suzanne Débarbat. Analyses d’ouvrages: Cecilia Payne-Gasposchkin, An autobiography and other recollections par Katherine Haramundanis. Revue d’Histoire des Sciences, 39(4):366–367, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632443.

deBeauregard:1986:AOBc

[3006] O. Costa de Beauregard. Analyses d’ouvrages: The creation of Quantum Mechanics and the Bohr–Pauli Dialogue, (^{≪}Studies in the History of Modern Science ^{≫}, 14) par John Hendry. Revue d’Histoire des Sciences, 39(4):367–368, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632444.

Mathieu:1986:AOBh

[3007] J.-P. Mathieu. Analyses d’ouvrages: Le noyau atomique. Physique nucléaire des énergies moyennes, (coll. ^{≪}Que sais-je? ^{≫}, 1826) par Théo Kahan. Revue d’Histoire des Sciences, 39(4):368, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632445.

Jacques:1986:AOBb

[3008] Jean Jacques. Analyses d’ouvrages: Boussingault, Chemist and Agriculturist par F. W. J. McCosh. Revue d’Histoire des Sciences, 39(4):369, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632446.

Dahan-Dalmedico:1986:AOB

[3009] Amy Dahan-Dalmedico. Analyses d’ouvrages: The Cauchy Method of Residues. Theory and Applications, (^{≪}Mathematics and its Applications. East European Series ^{≫}) par Dragoslav S. Mitrinović; Jovan D. Keckić. Revue d’Histoire des Sciences, 39(4):370, October 1986. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632447.

Aujac:1986:AOB

[3010] G. Aujac. Analyses d’ouvrages: Astronomy and History, Selected Essays … par O. Neugebauer. Revue d’Histoire des Sciences, 39(4):371, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632448.

Elena:1986:AOB

[3011] Alberto Elena. Analyses d’ouvrages: Scientific Writings and Astronomical Tables in Cracow : A Census of Manuscript Sources (XIVth–XVIth Centuries), (^{≪}Studia Copernicana ^{≫}, XXII) par Grazyna Rosińska. Revue d’Histoire des Sciences, 39(4):371–372, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632449.

Levy:1986:AOB

[3012] Jacques Lévy. Analyses d’ouvrages: Chronology of Eclipses and Comets, AD 1–1000 … par D. Justin Schove; Alan Fletcher. Revue d’Histoire des Sciences, 39(4):372–373, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632450.

V:1986:AOB

[3013] M. V. Analyses d’ouvrages: Enseignement et diffusion des sciences en France au XVIII^{e} siècle, (^{≪}Histoire de la pensée ^{≫}, XI) par René Taton; Ch. Bedel; A. Birembaut; P. Costabel; F. de Daninville. Revue d’Histoire des Sciences, 39(4):373, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632451.

Evans:1986:AOB

[3014] James Evans. Analyses d’ouvrages: Ptolemy’s Almagest par G. J. Toomer. Revue d’Histoire des Sciences, 39(4):373–375, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632452.

Theodorides:1986:AOBb

[3015] Jean Théodoridès. Analyses d’ouvrages: The zoological work of Petrus Camper (1722–1789) par Robert Paul Willem Visser. Revue d’Histoire des Sciences, 39(4):375–376, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632453.

Mathieu:1986:AOBi

[3016] J.-P. Mathieu. Analyses d’ouvrages: An introduction to science studies. The philosophical and social aspects of science and technology par John Ziman. Revue d’Histoire des Sciences, 39(4):376, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632454.

Anonymous:1986:BMa

[3017] Anonyme. Back matter. Revue d’Histoire des Sciences, 39(1):??, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632594.

Anonymous:1986:BMb

[3018] Anonyme. Back matter. Revue d’Histoire des Sciences, 39(2):??, avril– juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632732.

Anonymous:1986:BMc

[3019] Anonyme. Back matter. Revue d’Histoire des Sciences, 39(3):??, juillet– septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632494.

Anonymous:1986:BMd

[3020] Anonyme. Back matter. Revue d’Histoire des Sciences, 39(4):??, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632455.

Anonymous:1986:FMa

[3021] Anonyme. Front matter. Revue d’Histoire des Sciences, 39(1):??, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632562.

Anonymous:1986:FMb

[3022] Anonyme. Front matter. Revue d’Histoire des Sciences, 39(2):??, avril– juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632717.

Anonymous:1986:FMc

[3023] Anonyme. Front matter. Revue d’Histoire des Sciences, 39(3):??, juillet– septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632479.

Anonymous:1986:FMd

[3024] Anonyme. Front matter. Revue d’Histoire des Sciences, 39(4):??, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632431.

Blay:1986:EHD

[3025] Michel Blay. Études sur l’histoire du calcul infinitésimal. Revue d’Histoire des Sciences, 39(3):??, juillet–septembre 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632480.

Blay:1987:MPJ

[3026] Michel Blay. Mathématiques et philosophie : Jean Cavaillès; Albert Lautman. Revue d’Histoire des Sciences, 40(1):3, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632756.

Benis-Sinaceur:1987:SCE

[3027] Hourya Benis-Sinaceur. Structure et concept dans l’épistémologie mathématique de Jean Cavaillès. Revue d’Histoire des Sciences, 40(1): 5–30, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632757.

Mathématiques et philosophie: Jean Cavaillès, Albert Lautman.

Heinzmann:1987:PCP

[3028] Gerhard Heinzmann. La position de Cavaillès dans le problème des fondements en mathématiques, et sa différence avec celle de Lautman. Revue d’Histoire des Sciences, 40(1):31–47, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632758. Mathématiques et philosophie:

Jean Cavaillès, Albert Lautman.

Chevalley:1987:ALS

[3029] Catherine Chevalley. Albert Lautman et le souci logique. Revue d’Histoire des Sciences, 40(1):49–77, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www. jstor.org/stable/23632759. Mathématiques et philosophie: Jean Cavaillès, Albert Lautman.

Petitot:1987:RBI

[3030] Jean Petitot. Refaire le ^{≪}Timée ^{≫}: Introduction à la philosophie mathématique d’Albert Lautman. Revue d’Histoire des Sciences, 40 (1):79–115, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/ 23632760. Mathématiques et philosophie: Jean Cavaillès, Albert Lautman.

Benis-Sinaceur:1987:LIJ

[3031] Hourya Benis-Sinaceur. Lettres inédites de Jean Cavaillès à Albert Lautman. Revue d’Histoire des Sciences, 40(1):117–128, January 1987. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632761. Mathématiques et philoso-

phie: Jean Cavaillès, Albert Lautman.

Anonymous:1987:LIG

[3032] Anonyme. Lettre inédite de Gaston Bachelard à Albert Lautman. Revue d’Histoire des Sciences, 40(1):129, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632762.

Benis-Sinaceur:1987:LIG

[3033] Hourya Benis-Sinaceur. Lettre inédite de Gaston Bachelard à Albert Lautman. Revue d’Histoire des Sciences, 40(1):129, 1987. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). Mathématiques et philosophie: Jean Cavaillès, Albert Lautman.

Jaouiche:1987:I

[3034] K. Jaouiche and J. Payen. Informations. Revue d’Histoire des Sciences, 40(1):131–134, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632763.

Mathieu:1987:AOBa

[3035] J.-P. Mathieu. Analyses d’ouvrages: Otto Hahn, 1879–1968. Ein Forscherleben unserer Zeit ^{≪}Grosse Naturforscher ^{≫}, Bd. 45 par Walter Gerlach; Dietrich Hahn. Revue d’Histoire des Sciences, 40(1):135, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632764.

Sene:1987:AOB

[3036] Monique Sené. Analyses d’ouvrages: Cockroft and the Atom par Guy Hartcup; T. E. Allibone. Revue d’Histoire des Sciences, 40(1):135–136, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632765;
 http://www.persee.fr/doc/rhs_0151-4105_1987_num_40_1_4493_
t1_0135_0000_2.

Mathieu:1987:AOBb

[3037] J.-P. Mathieu. Analyses d’ouvrages: Cambridge Physics in the Thirties par John Hendry. Revue d’Histoire des Sciences, 40(1):136, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632766.

Simon:1987:AOB

[3038] Gérard Simon. Analyses d’ouvrages: Le secret du monde par Jean Kepler; Alain Segonds. Revue d’Histoire des Sciences, 40(1):136–138, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632767.

Jacquart:1987:AOB

[3039] Danielle Jacquart. Analyses d’ouvrages: ^{≪}Libreto de tute le cosse che se manzano ^{≫}, Un libro di dietetica di Michele Savonarola, medico padovano del secolo XV, vol. 1 par Jane Nystedt. Revue d’Histoire des Sciences, 40(1):138–139, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632768.

Darrigol:1987:AOB

[3040] Olivier Darrigol. Analyses d’ouvrages: Niels Bohr, 1885–1962; Leben und Werk eines Atomphysikers. ^{≪}Grosse Naturforscher ^{≫}, Bd. 47 par Ulrich Röseberg. Revue d’Histoire des Sciences, 40(1):139–140, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632769.

Taton:1987:AOB

[3041] René Taton. Analyses d’ouvrages: The mathematical work of John Wallis, DD, FRS (1616–1703) (2nd ed.) par J. F. Scott; E. N. da C. Andrade. Revue d’Histoire des Sciences, 40(1):140–141, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632770.

Dieudonne:1987:AOB

[3042] J. Dieudonné. Analyses d’ouvrages: A history of Algebra : From al-Khwarizmi to Emmy Noether par B. L. Van der Waerden. Revue d’Histoire des Sciences, 40(1):141–143, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632771.

Hulin:1987:AOB

[3043] Nicole Hulin. Analyses d’ouvrages: Le Personnel de l’enseignement supérieur en France aux XIX^{e} et XX^{e} siècle. Colloque organisé par l’Institut d’Histoire moderne et contemporaine et l’École des Hautes Études en Sciences sociales les 25 et 26 juin 1984 par Christophe Charle; Régine Ferré. Revue d’Histoire des Sciences, 40(1):143–144, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632772.

Langins:1987:EEE

[3044] Jānis Langins. Sur l’enseignement et les examens à l’École polytechnique sous le Directoire: à propos d’une lettre inédite de Laplace. Revue d’Histoire des Sciences, 40(2):145–177, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632739.

Dhombres:1987:PEE

[3045] Jean Dhombres. Les présupposés d’Euler dans l’emploi de la méthode fonctionnelle. Revue d’Histoire des Sciences, 40(2):179–202, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632740.

Bernhardt:1987:RCI

[3046] Jean Bernhardt. Le rôle des conceptions d’Isaac Beeckman dans la formation de Thomas Hobbes et dans l’élaboration de son “Short Tract”. Revue d’Histoire des Sciences, 40(2):203–215, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632741.

Meyer:1987:QFM

[3047] Fernand Meyer. Quelle ^{≪}médecine chinoise ^{≫}pour les praticiens occidentaux actuels? Revue d’Histoire des Sciences, 40(2):217–222, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632742.

Kubbinga:1987:ODA

[3048] H. H. Kubbinga. A l’occasion du 80^{e} anniversaire de Reyer Hooykaas. Revue d’Histoire des Sciences, 40(2):223–224, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632743.

Poulle:1987:AB

[3049] Emmanuel Poulle. Nécrologie. Alain Brieux (1922–1985). Revue d’Histoire des Sciences, 40(2):225–226, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632744.

Russo:1987:AB

[3050] François Russo. Arthur Birembaut (1905–1986). Revue d’Histoire des Sciences, 40(2):227–232, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632745.

Scriba:1987:I

[3051] C. J. Scriba and Józef Hurwic. Informations. Revue d’Histoire des Sciences, 40(2):233–236, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632746.

Payen:1987:AOBa

[3052] Jacques Payen. Analyses d’ouvrages: Ministère de la Culture, Direction du Patrimoine, Mission du Patrimoine ethnologique, Les moteurs hydrauliques et leurs applications industrielles en France (Rapport final d’activité. Rapport de synthèse) par Louis Bergeron. Revue d’Histoire des Sciences, 40(2):237–238, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632747.

Cohen:1987:AOB

[3053] Elliott Cohen. Analyses d’ouvrages: The Mathematics of Sonya Kovalevskaya par Roger Cooke. Revue d’Histoire des Sciences, 40(2):238, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632748.

Gourevitch:1987:AOB

[3054] Danielle Gourevitch. Analyses d’ouvrages: Santé et société en Grèce ancienne par Jean-Nicolas Corvisier; Jacques Dupâquier. Revue d’Histoire des Sciences, 40(2):239–242, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632749.

Lestienne:1987:AOB

[3055] Rémy Lestienne. Analyses d’ouvrages: La notion de temps : Equivalence avec l’espace. (^{≪}L’Histoire des sciences. Textes et Études ^{≫}) par Olivier Costa de Beauregard. Revue d’Histoire des Sciences, 40(2):242– 243, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632750.

Russo:1987:AOB

[3056] François Russo. Analyses d’ouvrages: L’Aventure des parallèles. Histoire de la géométrie non euclidienne : précurseurs et attardés par Jean-Claude Pont. Revue d’Histoire des Sciences, 40(2):243–245, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632751.

Fischer:1987:AOBa

[3057] J.-L. Fischer. Analyses d’ouvrages: Lamarck, sa vie et son œuvre par Yves Delange. Revue d’Histoire des Sciences, 40(2):245, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632752.

Fischer:1987:AOBb

[3058] J.-L. Fischer. Analyses d’ouvrages: Oltre il mito, Lamarck e le scienze naturali del suo tempo par Pietro Corsi. Revue d’Histoire des Sciences, 40(2):245, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633187.

Fischer:1987:AOBc

[3059] J.-L. Fischer. Analyses d’ouvrages: Emilt du Bois-Reymond (1818– 1896), Anton Dohrn (1840–1909), Briefwechsel par Christiane Groeben; Klaus Hierholzer; Ernst Florey. Revue d’Histoire des Sciences, 40(2):246, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633188.

Fischer:1987:AOBd

[3060] J.-L. Fischer. Analyses d’ouvrages: The correspondence of Charles Darwin, vol. 1 : 1821–1836 par Frederick Burkhardt; Sydney Smith; David Kohn; William Montgomery; Stephen V. Pocock; Charlotte Bowman; Janet Browne; Anne Secord; Anne Schlabach Burkhardt; Nora Carrol Stevenson. Revue d’Histoire des Sciences, 40(2):246–247, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633189.

Fischer:1987:AOBe

[3061] J.-L. Fischer. Analyses d’ouvrages: A history of embryology par T. J. Horder; J. A. Witkowski; C. G. Wylie. Revue d’Histoire des Sciences, 40(2):247–248, avril–juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633190.

Fischer:1987:AOBf

[3062] J.-L. Fischer. Analyses d’ouvrages: Le metafisiche dell’embrione, scienze deéllia vita e filosiofía da Malpighi a Spallanzani (1678–1793) par Walter Bernardi. Revue d’Histoire des Sciences, 40(2):248, avril–juin 1987. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633191.

Blay:1987:BNQ

[3063] Michel Blay. Les “Principia” de Newton questions et commentaires. Revue d’Histoire des Sciences, 40(3–4):249, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632901.

Costabel:1987:BNL

[3064] Pierre Costabel. Les “Principia” de Newton et leurs colonnes d’Hercule. Revue d’Histoire des Sciences, 40(3–4):251–271, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632902.

Barthelemy:1987:FAB

[3065] Georges Barthélemy. De la force accélératrice dans les “Principia”. Revue d’Histoire des Sciences, 40(3–4):273–280, juillet–décembre 1987. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632903.

DeGandt:1987:PIP

[3066] François De Gandt. Le problème inverse (prop. 39–41). Revue d’Histoire des Sciences, 40(3–4):281–309, juillet–décembre 1987. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632904.

Vilain:1987:NMM

[3067] Christiane Vilain. Newton et le modèle mécaniste de la réfraction. Revue d’Histoire des Sciences, 40(3–4):311–324, juillet–décembre 1987. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632905.

Blay:1987:TND

[3068] Michel Blay. Le traitement newtonien du mouvement des projectiles dans les milieux résistants. Revue d’Histoire des Sciences, 40(3–4): 325–355, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632906.

Greenberg:1987:INT

[3069] John L. Greenberg. Isaac Newton et la théorie de la figure de la Terre. Revue d’Histoire des Sciences, 40(3–4):357–366, juillet–décembre 1987. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632907.

Nicolaidis:1987:LQO

[3070] Efthymios Nicolaïdis. Les livres qui ont introduit les sciences dans le monde grec au siècle des Lumières. Revue d’Histoire des Sciences, 40(3– 4):367–376, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632908.

Chevalley:1987:LJC

[3071] Catherine Chevalley. L’optique des Jésuites et celle des médecins: À propos de deux ouvrages récents. Revue d’Histoire des Sciences, 40(3– 4):377–382, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632909.

Dhombres:1987:ODB

[3072] Jean Dhombres. Àl’occasion du bicentenaire de la mort de Leonhard Euler (1707–1783). Revue d’Histoire des Sciences, 40(3–4):383–386, juillet– décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632910.

Goupil:1987:I

[3073] Michelle Goupil. Informations. Revue d’Histoire des Sciences, 40(3–4): 387–389, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632911.

Mathieu:1987:AOBc

[3074] Jean-Paul Mathieu. Analyses d’ouvrages: Aller-retour pour l’espace par Joseph Allen; Russell Martin; J. Guiod. Revue d’Histoire des Sciences, 40(3–4):391, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632912.

Diara:1987:AOB

[3075] Anne Diara. Analyses d’ouvrages: L’esprit des protéines : Histoire et philosophies biochimiques par Claude Debru. Revue d’Histoire des Sciences, 40(3–4):391–394, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632913.

Payen:1987:AOBb

[3076] Jacques Payen. Analyses d’ouvrages: Aviation : An Historical Survey from its Origins to the end of World War II par Charles Harvard Gibbs-Smith. Revue d’Histoire des Sciences, 40(3–4):394–395, juillet– décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632914.

Mathieu:1987:AOBd

[3077] J.-P. Mathieu. Analyses d’ouvrages: Gesammelte Werke, Collected Works, Series A/Part I: Original scientific Papers. Wissenschaftliche Originalarbeiten par Werner Heisenberg; W. Blum; H.-P. Dürr; H. Rechenberg. Revue d’Histoire des Sciences, 40(3–4):395–396, juillet– décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632915.

Speziali:1987:AOB

[3078] Pierre Speziali. Analyses d’ouvrages: Els orígens i l’ensenyament de l’àlgebra simbòlica (1478–1545), vol. I par Antoni Malet; Jaume Paradís. Revue d’Histoire des Sciences, 40(3–4):396–398, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632916.

Guiot:1987:AOB

[3079] J.-P. Guiot. Analyses d’ouvrages: Mariotte, savant et philosophe (+ 1684) : Analyse d’une renommée. (^{≪}Histoire des Sciences. Textes et études ^{≫}). Revue d’Histoire des Sciences, 40(3–4):398–399, juillet– décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632917.

Meyer:1987:AOB

[3080] Fernand Meyer. Analyses d’ouvrages: Science and Civilisation in China. Vol. V : Chemistry and chemical technology, part 5 : Spagyrical discovery and invention : physiological alchemy par Joseph Needham. Revue d’Histoire des Sciences, 40(3–4):399–402, juillet–décembre 1987. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632918.

deBeauregard:1987:AOB

[3081] O. Costa de Beauregard. Analyses d’ouvrages: Le grand débat de la théorie quantique par Franco Selleri. Revue d’Histoire des Sciences, 40(3– 4):402–403, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632919.

Poulle:1987:AOB

[3082] Emmanuel Poulle. Analyses d’ouvrages: The Time Museum. Vol. 1 : Time Measuring Instruments.Part 1 : Astrolabes, Astrolabe Related Instruments par A. J. Turner. Revue d’Histoire des Sciences, 40(3–4): 403–405, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632920.

Debarbat:1987:AOB

[3083] Suzanne Débarbat. Analyses d’ouvrages: The Milky Way Galaxy. Proceedings of the 106th Symposium of the IAU. ^{≪}International Astronomical Union ^{≫}, Symposium n^{o} 106 par Hugo Van Woerden; Ronald J. Allen; W. Butler Burton. Revue d’Histoire des Sciences, 40(3–4): 405–406, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632921.

Anonymous:1987:PR

[3084] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 40(3–4): 407–408, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632922.

Anonymous:1987:BMa

[3085] Anonyme. Back matter. Revue d’Histoire des Sciences, 40(1):??, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632773.

Anonymous:1987:BMb

[3086] Anonyme. Back matter. Revue d’Histoire des Sciences, 40(2):??, avril– juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632753.

Anonymous:1987:BMc

[3087] Anonyme. Back matter. Revue d’Histoire des Sciences, 40(3–4): ??, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632923.

Anonymous:1987:FMa

[3088] Anonyme. Front matter. Revue d’Histoire des Sciences, 40(1):??, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632755.

Anonymous:1987:FMb

[3089] Anonyme. Front matter. Revue d’Histoire des Sciences, 40(2):??, avril– juin 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632738.

Anonymous:1987:FMc

[3090] Anonyme. Front matter. Revue d’Histoire des Sciences, 40(3–4): ??, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632900.

Six:1988:PNB

[3091] Jules Six. Pourquoi ni Bothe ni les Joliot–Curie n’ont découvert le neutron. (French) [Why neither [Walther] Bothe nor [Pierre] Joliot–Curie discovered the neutron]. Revue d’Histoire des Sciences, 41(1):3–24, January 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632789.

Rosmorduc:1988:ANP

[3092] Jean Rosmorduc. Arago et la naissance de la polarimétrie. Revue d’Histoire des Sciences, 41(1):25–38, January 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632790.

Taton:1988:DLB

[3093] René Taton. Le départ de Lagrange de Berlin et son installation à Paris en 1787. Revue d’Histoire des Sciences, 41(1):39–74, January 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632791.

Pestre:1988:SFP

[3094] Dominique Pestre. Sur la science en France, 1860–1940: À propos de deux ouvrages récents de mary jo nye et Harry w. Paul. Revue d’Histoire des Sciences, 41(1):75–83, January 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632792.

Anonymous:1988:Ia

[3095] Anonyme. Informations. Revue d’Histoire des Sciences, 41(1):85–86, January 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632793.

Payen:1988:AOB

[3096] Jacques Payen. Analyses d’ouvrages: Introduction à l’Histoire des Techniques. (^{≪}Bibliothèque scientifique Albert Blanchard ^{≫}) par François Russo. Revue d’Histoire des Sciences, 41(1):87–92, January 1988. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632794.

Evans:1988:AOBa

[3097] James Evans. Analyses d’ouvrages: Les Tables Alphonsines avec les canons de Jean de Saxe, (^{≪}Sources d’Histoire médiévale publ. par l’Institut de Recherche et d’Histoire des Textes ^{≫}) par Emmanuel Poulle. Revue d’Histoire des Sciences, 41(1):93–95, January 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632795.

Russo:1988:AOB

[3098] François Russo. Analyses d’ouvrages: L’Affaire Lumière : Du mythe à l’histoire. Enquête sur les origines du cinéma. (coll. ^{≪}Le cinéma et son histoire ^{≫}) par Léo Sauvage. Revue d’Histoire des Sciences, 41(1):95–96, January 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632796.

Benis-Sinaceur:1988:DMH

[3099] Hourya Benis-Sinaceur. Deux moments dans l’histoire du théorème d’algèbre de Ch. F. Sturm. Revue d’Histoire des Sciences, 41(2):99–132, avril–juin 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632982.

Dubucs:1988:JBT

[3100] Jacques-Paul Dubucs. L. J. E. Brouwer : topologie et constructivisme. Revue d’Histoire des Sciences, 41(2):133–155, avril–juin 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632983.

Salanskis:1988:LNS

[3101] Jean-Michel Salanskis. L’analyse non standard et la tradition de l’infini. Revue d’Histoire des Sciences, 41(2):157–207, avril–juin 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632984.

Anonymous:1988:Ib

[3102] Anonyme. Informations. Revue d’Histoire des Sciences, 41(2):209–212, avril–juin 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632985.

Crepel:1988:AOB

[3103] Pierre Crépel. Analyses d’ouvrages: Les manuscrits mathématiques de Marx par Alain Alcouffe. Revue d’Histoire des Sciences, 41(2):213–214, avril–juin 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632986.

Pueyo:1988:AOB

[3104] G. Pueyo. Analyses d’ouvrages: Histoire de la génétique et de l’évolutionnisme en France. (^{≪}Histoires ^{≫}) par Denis Buican. Revue d’Histoire des Sciences, 41(2):214–216, avril–juin 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632987.

Gourevitch:1988:AOB

[3105] D. Gourevitch. Analyses d’ouvrages: Studi su papiri greci di logica e medicina par W. Cavini; M. C. Donnini Macciò; M. S. Funghi; D. Manetti. Revue d’Histoire des Sciences, 41(2):216–217, avril–juin 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632988.

Martinez-Chavanz:1988:AOB

[3106] Régino Martínez-Chavanz. Analyses d’ouvrages: Understanding Relativity : Origin and impact of a Scientific Revolution par Stanley Goldberg. Revue d’Histoire des Sciences, 41(2):217–218, avril–juin 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632989.

Reyssat:1988:AOB

[3107] Eric Reyssat. Analyses d’ouvrages: Number theory. An approach through history : From Hammurapi to Legendre par André Weil. Revue d’Histoire des Sciences, 41(2):218–220, avril–juin 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632990.

Anonymous:1988:PR

[3108] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 41(2):221– 223, avril–juin 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632991.

Anonymous:1988:NR

[3109] Anonyme. Nouvelles des revues. Revue d’Histoire des Sciences, 41(2): 224, avril–juin 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632992.

Darrigol:1988:QEA

[3110] Olivier Darrigol. The quantum electrodynamical analogy in early nuclear theory or the roots of Yukawa’s theory. Revue d’Histoire des Sciences, 41(3–4):225–297, juillet–décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633104.

Roussel:1988:CMC

[3111] Fabrice Roussel. Le concept de mélancolie chez Aristote. Revue d’Histoire des Sciences, 41(3–4):299–330, juillet–décembre 1988. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633105.

Knobloch:1988:VOC

[3112] Eberhard Knobloch. Sur la vie et l’œuvre de Christophore Clavius (1538– 1612). Revue d’Histoire des Sciences, 41(3–4):331–356, juillet–décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633106.

Ricard:1988:ELM

[3113] Jean-Christian Ricard. Equitation, locomotion et mécanisme des allures au XIX^{e} siècle: de la méthode graphique à la chronophotographie. Revue d’Histoire des Sciences, 41(3–4):357–376, juillet–décembre 1988. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633107.

Larmore:1988:NCC

[3114] Charles Larmore. La notion de certitude chez Newton. Revue d’Histoire des Sciences, 41(3–4):377–384, juillet–décembre 1988. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633108.

Balibar:1988:EFD

[3115] Françoise Balibar. Sur l’ “Encyclopédie de la Fondation Diderot”. Revue d’Histoire des Sciences, 41(3–4):385–388, juillet–décembre 1988. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633109.

Corpus:1988:COP

[3116] Le Corpus. Le “corpus des œuvres de philosophie en langue française” et l’histoire des sciences. Revue d’Histoire des Sciences, 41(3–4):389– 393, juillet–décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633110.

Anonymous:1988:Ic

[3117] Anonyme. Informations. Revue d’Histoire des Sciences, 41(3–4):395– 397, juillet–décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633111.

Evans:1988:AOBb

[3118] James Evans. Analyses d’ouvrages: The Astronomy of Levi ben Gerson (1288–1344). A Critical Edition of Chapters 1–20 with Translation and Commentary (^{≪}Studies in the History of Mathematics and Physical Sciences ^{≫}, 11) par Bernard R. Goldstein; Theory and Observation in Ancient and Medieval Astronomy (^{≪}Collected Studies Series ^{≫}) par Bernard R. Goldstein. Revue d’Histoire des Sciences, 41(3–4):399–401, juillet– décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633112.

Caroti:1988:AOB

[3119] Stefano Caroti. Analyses d’ouvrages: Nicole Oresme and the Marvels of Nature. A Study of his De causis mirabilium with Critical Edition, Translation and Commentary (^{≪}Studies and Texts ^{≫}, 68) par Bert Hansen. Revue d’Histoire des Sciences, 41(3–4):401–403, juillet–décembre 1988. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633113.

Segonds:1988:AOB

[3120] A. Ph. Segonds. Analyses d’ouvrages: The Birth of History and Philosophy of Science. Kepler’s “A defence of Tycho against Ursus with essays on its provenance and significance” par N. Jardine. Revue d’Histoire des Sciences, 41(3–4):403–404, juillet–décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633114.

DeGandt:1988:AOB

[3121] François De Gandt. Analyses d’ouvrages: Histoire générale de la nature et théorie du ciel (1755), (^{≪}Bibliothèque des Textes philosophiques ^{≫}) par Emmanuel Kant; Pierre Kerszberg; Anne-Marie Roviello; Jean Seidengart. Revue d’Histoire des Sciences, 41(3–4):405–406, juillet–décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633115.

Emptoz:1988:AOB

[3122] G. Emptoz. Analyses d’ouvrages: Œuvres de Lavoisier. Correspondance, fasc. IV: 1784–1786 par A. Horeau; M. Goupil. Revue d’Histoire des Sciences, 41(3–4):406–407, juillet–décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633116.

Eckart:1988:AOB

[3123] Wolfgang U. Eckart. Analyses d’ouvrages: Der Schweizer Arzl und Philosoph Ignaz Paul Vital Troxler (1780–1866). Seine Philosophie, Anthropologie und Medizintheorie par Peter Heusser. Revue d’Histoire des Sciences, 41(3–4):407–408, juillet–décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633117.

Bernhardt:1988:AOB

[3124] Jean Bernhardt. Analyses d’ouvrages: Correspondance du P. Marin Mersenne, t. XVI : 1648 par Paul Tannery; Cornélis de Waard; Armand Beaulieu. Revue d’Histoire des Sciences, 41(3–4):409–410, juillet– décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633118.

Grmek:1988:AOB

[3125] M. D. Grmek. Analyses d’ouvrages: Galen’s Commentary on the Hippocratic Treatise Airs, Waters, Places in the Hebrew Translation of Solomon ha-Me’ati par Abraham Wasserstein; Galen on Bloodletting : A Study of the Origins, Development and Validity of His Opinions, with a Translation of the Three Works par Peter Brain. Revue d’Histoire des Sciences, 41(3–4):410–411, juillet–décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633119.

Jacquart:1988:AOB

[3126] Danielle Jacquart. Analyses d’ouvrages: Roger Bacon’s Philosophy of Nature : A Critical Edition, with English Translation, Introduction and Notes, of “De multiplicatione specierum and De speculis comburentibus” par David C. Lindberg. Revue d’Histoire des Sciences, 41(3–4): 411–412, juillet–décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633120.

Anonymous:1988:BMa

[3127] Anonyme. Back matter. Revue d’Histoire des Sciences, 41(1):??, January 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632797.

Anonymous:1988:BMb

[3128] Anonyme. Back matter. Revue d’Histoire des Sciences, 41(2):??, avril– juin 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632993.

Anonymous:1988:BMc

[3129] Anonyme. Back matter. Revue d’Histoire des Sciences, 41(3–4): ??, juillet–décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633121.

Anonymous:1988:FMa

[3130] Anonyme. Front matter. Revue d’Histoire des Sciences, 41(1):??, January 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632788.

Anonymous:1988:FMb

[3131] Anonyme. Front matter. Revue d’Histoire des Sciences, 41(2):??, avril– juin 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632980.

Anonymous:1988:FMc

[3132] Anonyme. Front matter. Revue d’Histoire des Sciences, 41(3–4): ??, juillet–décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633103.

Blay:1988:AAT

[3133] Michel Blay. Algèbre, analyse, topologie. questions d’histoire et d’interprétation. Revue d’Histoire des Sciences, 41(2):??, avril–juin 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632981.

Anonymous:1989:M

[3134] Anonyme. La Mathématisation 1780–1830. Revue d’Histoire des Sciences, 42(1–2):3, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633022.

Chabert:1989:GMM

[3135] Jean-Luc Chabert. Gauss et la méthode des moindres carrés. Revue d’Histoire des Sciences, 42(1–2):5–26, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www. jstor.org/stable/23633023. La mathématisation 1780–1830.

Bottazzini:1989:LPK

[3136] Umberto Bottazzini. Lagrange et le problème de Kepler. Revue d’Histoire des Sciences, 42(1–2):27–42, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/ stable/23633024. La mathématisation 1780–1830.

Dhombres:1989:TCS

[3137] Jean Dhombres. La théorie de la capillarité selon Laplace : mathématisation superficielle ou étendue? Revue d’Histoire des Sciences, 42(1–2):43–77, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/ 23633025. La mathématisation 1780–1830.

Dahan-Dalmedico:1989:NPM

[3138] Amy Dahan-Dalmédico. La notion de pression : de la métaphysique aux diverses mathématisations: Causalité et statut des hypothèses. Revue d’Histoire des Sciences, 42(1–2):79–108, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633026. La mathématisation 1780–1830.

Scholz:1989:RSC

[3139] Erhard Scholz. The rise of symmetry concepts in the atomistic and dynamistic schools of crystallography, 1815–1830. Revue d’Histoire des Sciences, 42(1–2):109–122, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/ stable/23633027. La mathématisation 1780–1830.

Blondel:1989:VPE

[3140] Christine Blondel. Vision physique ^{≪}éthérienne ^{≫}, mathématisation ≪laplacienne ^{≫}: l’électrodynamique d’Ampère. Revue d’Histoire des Sciences, 42(1–2):123–137, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/ stable/23633028. La mathématisation 1780–1830.

Pourprix:1989:MPG

[3141] Bernard Pourprix. La mathématisation des phénomènes galvaniques par G. S. Ohm (1825–1827). Revue d’Histoire des Sciences, 42(1–2):139– 154, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633029. La

mathématisation 1780–1830.

Picon:1989:IME

[3142] Antoine Picon. Les ingénieurs et la mathématisation. L’exemple du génie civil et de la construction. Revue d’Histoire des Sciences, 42(1–2):155– 172, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633030. La

mathématisation 1780–1830.

Kubbinga:1989:NCL

[3143] H. H. Kubbinga. Nouveau: le Catalogus [ …] librorum d’Isaac Beeckman. Revue d’Histoire des Sciences, 42(1–2):173–174, January 1989. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633031.

Anonymous:1989:I

[3144] Anonyme. Informations. Revue d’Histoire des Sciences, 42(1–2):177–178, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633032.

Diara:1989:AOB

[3145] A. Diara. Analyses d’ouvrages: Le même et l’autre : Recherches sur l’individualité dans les sciences de la vie (^{≪}Fondements des sciences ^{≫}) par H. Barreau. Revue d’Histoire des Sciences, 42(1–2):179–180, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633033.

Rulliere:1989:AOB

[3146] Roger Rullière. Analyses d’ouvrages: Ambroise Paré, chirurgien de quatre rois de France par Paule Dumaître. Revue d’Histoire des Sciences, 42(1–2):181, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633034.

Pont:1989:AOB

[3147] Jean-Claude Pont. Analyses d’ouvrages: La théorie des parallèles en pays d’Islam: Contribution à la préhistoire des géométries non-euclidiennes (^{≪}L’Histoire des Sciences. Textes et Études ^{≫}) par Khalil Jaouiche. Revue d’Histoire des Sciences, 42(1–2):181–182, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633035.

Grmek:1989:AOBa

[3148] M. D. Grmek. Analyses d’ouvrages: Index Hippocraticus, fasc. I: A–Δ, fasc. II: E–K, fasc. III: Δ — II. Cui elaborando interfuerunt sodales Thesauri Linguae Graecae Hamburgensis. Curas postremas adhibuerunt K. Alpers, A. Anastassiou, D. Irmer & V. Schmidt par Josef Hans Kühn; Ulrich Fleischer. Revue d’Histoire des Sciences, 42(1–2):182–183, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633036.

Laurent:1989:AOB

[3149] Goulven Laurent. Analyses d’ouvrages: ^{≪}Histoire des idées et théorie de l’évolution ^{≫}, Revue de synthèse, t. CVII : 3 (1986) par C. Devillers; S. J. Gould; J. Greene; F. Jacob; E. Mayr; J. Roger. Revue d’Histoire des Sciences, 42(1–2):183–184, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633037.

Grmek:1989:AOBb

[3150] M. D. Grmek. Analyses d’ouvrages: Nicolaus Steno (1638–1686). A reconsideration par Danish scientists par Jacob E. Poulsen; Egill Snorrason; Niccolo Stenone nella Firenze e nell’Europa del suo tempo : Mostra di documenti, manoscritti, opere par Stefano De Rosa. Revue d’Histoire des Sciences, 42(1–2):184–185, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633038.

Valentin:1989:AOB

[3151] Michel Valentin. Analyses d’ouvrages: Pasteur et la révolution pastorienne, (^{≪}Médecine et sociétés ^{≫}) par Claire Salomon-Bayet; Bernard Lécuyer; Jacques Léonard; Viviane Thévenin; Roger Carvais; Bruno Latour; André Lwoff. Revue d’Histoire des Sciences, 42(1–2):185–186, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633039.

Delorme:1989:AOB

[3152] Suzanne Delorme. Analyses d’ouvrages: Jean Rostand, un homme du futur par Andrée Tétry; Jacques Testart. Revue d’Histoire des Sciences, 42(1–2):186–188, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633040.

Delange:1989:AOB

[3153] Yves Delange. Analyses d’ouvrages: Histoire du concept d’espèce dans les sciences de la vie par Scott Atran; Richard W. Burkhardt; Pietro Corsi; Anne Diara; Bernardino Fantini; Jean-Louis Fischer; M. J. S. Hodge; Goulven Laurent; Antonello La Vergata; Anto Leikola; Pierre Louis; Ernst Mayr; Gerhard H. Müller; Roselyne Rey; Jacques Roger; Philipp R. Sloan; Étienne Wolff. Revue d’Histoire des Sciences, 42(1–2):188–189, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633041.

Gohau:1989:AOBa

[3154] G. Gohau. Analyses d’ouvrages: History of Geophysical Research in the Netherlands and its former Overseas Territories par J. Veldkamp. Revue d’Histoire des Sciences, 42(1–2):189–190, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633042.

Gohau:1989:AOBb

[3155] G. Gohau. Analyses d’ouvrages: Down to Earth : 150 Years of the British Geological Survey … par H. E. Wilson. Revue d’Histoire des Sciences, 42(1–2):190–191, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633043.

Mathieu:1989:AOBa

[3156] J.-P. Mathieu. Analyses d’ouvrages: Tabibito. Ein Wanderer, 1907– 1981. Erinnerungen eines Physikers, (^{≪}Grosse Naturforscher ^{≫}, Bd. 48) par Hideki Yukawa; Erwin Müller–Hartmann. Revue d’Histoire des Sciences, 42(1–2):191, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633044.

Mathieu:1989:AOBb

[3157] J.-P. Mathieu. Analyses d’ouvrages: Actes de la Journée ^{≪}Découvertes scientifiques et innovations techniques aux XIX^{e} et XX^{e} siècles ^{≫}, Bulletin d’histoire de l’électricité, 5 (juin 1985). Revue d’Histoire des Sciences, 42(1–2):191–192, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633045.

Levy:1989:LSC

[3158] Tony Lévy. L’étude des sections coniques dans la tradition médiévale hébraïque: Ses relations avec les traditions arabe et latine. Revue d’Histoire des Sciences, 42(3):193–239, juillet–septembre 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632799.

Verdet:1989:DH

[3159] Jean-Pierre Verdet. La diffusion de l’héliocentrisme. Revue d’Histoire des Sciences, 42(3):241–253, juillet–septembre 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632800.

Lerner:1989:PMC

[3160] Michel-Pierre Lerner. Le problème de la matière céleste après 1550 : aspects de la bataille des cieux fluides. Revue d’Histoire des Sciences, 42(3): 255–280, juillet–septembre 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632801.

Dupuigrenet-Desroussilles:1989:RSI

[3161] François Dupuigrenet-Desroussilles. Regards et savoirs : images du jardin botanique de l’Université de Padoue au XVI^{e} siècle. Revue d’Histoire des Sciences, 42(3):281–291, juillet–septembre 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632802.

Russo:1989:MAK

[3162] François Russo. Mémoire d’Alexandre Koyré. Revue d’Histoire des Sciences, 42(3):293–301, juillet–septembre 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632803.

Peiffer:1989:LNL

[3163] Jeanne Peiffer. Leibniz, Newton et leurs disciples. Revue d’Histoire des Sciences, 42(3):303–312, juillet–septembre 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632804.

Theodorides:1989:AD

[3164] Jean Théodoridès. André Doyon (1906–1988). Revue d’Histoire des Sciences, 42(3):313–314, juillet–septembre 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632805.

Pestre:1989:AOB

[3165] Dominique Pestre. Analyses d’ouvrages: The Birth of Particle Physics. Based on a Fermilab Symposium par Laurie M. Brown; Lillian Hoddeson. Revue d’Histoire des Sciences, 42(3):315–316, juillet–septembre 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632806.

Moulin:1989:AOB

[3166] Anne-Marie Moulin. Analyses d’ouvrages: Disease and Civilization. The cholera in Paris, 1832 par François Delaporte; Arthur Goldhammer. Revue d’Histoire des Sciences, 42(3):316–317, juillet–septembre 1989. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632807.

Russo:1989:AOBa

[3167] François Russo. Analyses d’ouvrages: Histoire de l’École polytechnique. ^{≪}Librairie du Bicentenaire de la Révolution française ^{≫}par Ambroise Fourcy; J. Dhombres. Revue d’Histoire des Sciences, 42(3):

317–318, juillet–septembre 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632808.

Russo:1989:AOBb

[3168] François Russo. Analyses d’ouvrages: La République avait besoin de savants : Les débuts de l’École polytechnique, L’École centrale des travaux publics et les cours révolutionnaires de l’an III. ^{≪}Librairie du Bicentenaire de la Révolution française ^{≫}par Janis Langins; Emmanuel Grison. Revue d’Histoire des Sciences, 42(3):318–319, juillet–septembre 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632809.

Debarbat:1989:AOB

[3169] Suzanne Débarbat. Analyses d’ouvrages: The analysis of starlight : One hundred and fifty years of astronomical spectroscopy par J. B. Hearnshaw. Revue d’Histoire des Sciences, 42(3):319–320, juillet–septembre 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632810.

Anonymous:1989:PHS

[3170] Anonyme. Problèmes d’histoire des sciences en Chine. Revue d’Histoire des Sciences, 42(4):321–322, October 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632775.

Gernet:1989:SRO

[3171] Jacques Gernet. Sciences et rationalité : l’originalité des données chinoises. Revue d’Histoire des Sciences, 42(4):323–332, October 1989. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632776.

Brenier:1989:SGS

[3172] Joël Brenier, Colette Diény, Jean-Claude Martzloff, and Wladyslaw de Wieclawik. Shen Gua (1031–1095) et les sciences. Revue d’Histoire des Sciences, 42(4):333–351, October 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632777.

Metailie:1989:HNH

[3173] Georges Métailié. Histoire naturelle et humanisme en Chine et en Europe au XVI^{e} siècle : Li Shizhen et Jacques Dalechamp. Revue d’Histoire des Sciences, 42(4):353–374, October 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632778.

Horiuchi:1989:PRE

[3174] Annick Horiuchi. Sur un point de rupture entre les traditions chinoise et japonaise des mathématiques. Revue d’Histoire des Sciences, 42(4):375–390, October 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/ 23632779. Problèmes d’histoire des sciences en Chine, I.

Jami:1989:CMS

[3175] Catherine Jami. Classification en mathématiques: la structure de l’encyclopédie “Yu Zhi Shu Li Jing Yun” (1723). Revue d’Histoire des Sciences, 42(4):391–406, October 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/ stable/23632780. Problèmes d’histoire des sciences en Chine, I.

Fichant:1989:YBF

[3176] Michel Fichant. Yvon Belaval (24 février 1908–19 novembre 1988). Revue d’Histoire des Sciences, 42(4):407–412, October 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632781.

Pueyo:1989:AOB

[3177] Guy Pueyo. Analyses d’ouvrages: Science and Civilisation in China, vol. 6 : Biology and biological technology. Part II : Agriculture par Joseph Needham; Francesca Bray. Revue d’Histoire des Sciences, 42(4):413–414, October 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632782.

Chemla:1989:AOB

[3178] Karine Chemla. Analyses d’ouvrages: Les premières théories planétaires chinoises, ^{≪}Mémoires de l’Institut des Hautes Études chinoises ^{≫}, vol. XXI par Michel Teboul; Jean-Claude Pecker. Revue d’Histoire des Sciences, 42(4):415–416, October 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632783.

Anonymous:1989:PR

[3179] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 42(4):417– 423, October 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632784.

Anonymous:1989:BMa

[3180] Anonyme. Back matter. Revue d’Histoire des Sciences, 42(1–2):??, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633046.

Anonymous:1989:BMb

[3181] Anonyme. Back matter. Revue d’Histoire des Sciences, 42(3):??, juillet– septembre 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632811.

Anonymous:1989:BMc

[3182] Anonyme. Back matter. Revue d’Histoire des Sciences, 42(4):??, October 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632785.

Anonymous:1989:FMa

[3183] Anonyme. Front matter. Revue d’Histoire des Sciences, 42(1–2):??, January 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633021.

Anonymous:1989:FMb

[3184] Anonyme. Front matter. Revue d’Histoire des Sciences, 42(3):??, juillet– septembre 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632798.

Anonymous:1989:FMc

[3185] Anonyme. Front matter. Revue d’Histoire des Sciences, 42(4):??, October 1989. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632774.

Anonymous:1990:PHS

[3186] Anonyme. Problèmes d’histoire des sciences en Chine. Revue d’Histoire des Sciences, 43(1):3–4, January 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633133.

Kalinowski:1990:CDR

[3187] Marc Kalinowski. Le calcul du rayon céleste dans la cosmographie chinoise. Revue d’Histoire des Sciences, 43(1):5–34, January 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633134.

Despeux:1990:CEP

[3188] Catherine Despeux and Frédéric Obringer. Conceptualisation d’un état pathologique dans la médecine chinoise traditionnelle : exemple de la toux. Revue d’Histoire des Sciences, 43(1):35–56, January 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633135.

Chemla:1990:DPE

[3189] Karine Chemla. Du parallélisme entre énoncés mathématiques. Analyse d’un formulaire rédigé en Chine au XIIIe siècle. Revue d’Histoire des Sciences, 43(1):57–80, January 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633136.

Martzloff:1990:EMC

[3190] Jean-Claude Martzloff. Un exemple de mathématiques chinoises non triviales: les formules sommatoires finies de Li Shan Lan (1811–1882). Revue d’Histoire des Sciences, 43(1):81–98, January 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633137.

Anonymous:1990:BCH

[3191] Anonyme. Bibliographie concernant l’histoire des sciences en Chine. Revue d’Histoire des Sciences, 43(1):99–107, January 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633138.

Pestre:1990:LAP

[3192] Dominique Pestre. Leviathan and the air-pump, Hobbes, Boyle, and the experimental life. Revue d’Histoire des Sciences, 43(1):109–116, January 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633139.

Anonymous:1990:Ia

[3193] Anonyme. Informations. Revue d’Histoire des Sciences, 43(1):117–118, January 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633140.

Payen:1990:AOBa

[3194] Jacques Payen. Analyses d’ouvrages: Histoire de l’ingénierie géomécanique jusqu’à 1700 par Jean Kerisel; Down to Earth. Foundations Past and Present : The Invisible Art of the Builder par Jean Kerisel. Revue d’Histoire des Sciences, 43(1):119–120, January 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633141.

Picolet:1990:AOB

[3195] Guy Picolet. Analyses d’ouvrages: The Correspondence of Henry Oldenburg. Vol. XII: October 1675–July 1676, Letters 2755–2940 par A. Rupert Hall; Marie Boas Hall; The Correspondence of Henry Oldenburg. Vol. XIII: July 1676–July 1681, Letters 2941–3139; Additions and Corrections to Vol. I–XI par A. Rupert Hall; Marie Boas Hall. Revue d’Histoire des Sciences, 43(1):120–123, January 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633142.

Poulle:1990:AOBa

[3196] Emmanuel Poulle. Analyses d’ouvrages: Horoscopes and history ≪Warburg Institute surveys and texts ^{≫}, XIII par John D. North. Revue d’Histoire des Sciences, 43(1):124–126, January 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633143.

Levy:1990:AOB

[3197] Jacques Lévy. Analyses d’ouvrages: Recueil des inscriptions gauloises (RIG). Vol. III: Les calendriers (Coligny, Villards-d’Héria), XLV^{e} supplément à Gallia par Paul-Marie Duval; Georges Pinault. Revue d’Histoire des Sciences, 43(1):126–128, January 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633144.

Durand:1990:GAS

[3198] Marie-José Durand. Genèse de l’algèbre symbolique en Angleterre: une influence possible de J. Locke. Revue d’Histoire des Sciences, 43(2–3): 129–180, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633075.

Gispert:1990:PAC

[3199] Hélène Gispert. Principes de l’analyse chez Darboux et Houël (1870– 1880) : textes et contextes. Revue d’Histoire des Sciences, 43(2–3): 181–220, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633076.

Sinaceur:1990:DPR

[3200] Mohammed-Allal Sinaceur. Dedekind et le programme de Riemann. Revue d’Histoire des Sciences, 43(2–3):221–236, avril–septembre 1990. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633077. Voir [3201].

Dedekind:1990:AUB

[3201] R. Dedekind. Analytische Untersuchungen zu Bernhard Riemann’s Abhandlungen Über die Hypothesen, welche der Geometrie zu Grunde liegen. Revue d’Histoire des Sciences, 43(2–3):237–296, 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). Texte en allemand avec traduction française. Voir aussi [3200].

Taton:1990:PC

[3202] René Taton. Pierre Costabel (1912–1989). Revue d’Histoire des Sciences, 43(2–3):297–324, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor. org/stable/23633078. Une liste des publications de Pierre Costabel se trouve aux pages 313–324.

Rav:1990:GCP

[3203] Yehuda Rav. Georg Cantor, 1845–1918 par Walter Purkert et Hans-Joachim Ilgauds. Revue d’Histoire des Sciences, 43(2–3):325–331, avril– septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633079.

Mathieu:1990:TA

[3204] Jean-Paul Mathieu. Sur le théorème d’Ampère. Revue d’Histoire des Sciences, 43(2–3):333–338, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633080.

Dhombres:1990:I

[3205] Jean Dhombres. Informations. Revue d’Histoire des Sciences, 43(2–3): 339–343, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633081.

Martzloff:1990:AOB

[3206] Jean-Claude Martzloff. Analyses d’ouvrages: Chinese Mathematics, a Concise History par Lĭ Yăn; Dù Shírán; traduction anglaise par John N. Crossley; Anthony W. C. Lun. Revue d’Histoire des Sciences, 43(2–3):

345–346, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633082.

Chemla:1990:AOB

[3207] Karine Chemla. Analyses d’ouvrages: Histoire des mathématiques chinoises par Jean-Claude Martzloff; J. Gernet; J. Dhombres. Revue d’Histoire des Sciences, 43(2–3):346–349, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633083.

Goupil:1990:AOBa

[3208] Michelle Goupil. Analyses d’ouvrages: Avis aux ouvriers en fer sur la fabrication de l’acier, publié par ordre du Comité de Salut public. Revue d’Histoire des Sciences, 43(2–3):350, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633085.

Grmek:1990:AOBa

[3209] Mirko D. Grmek. Analyses d’ouvrages: Introduction to the history of medical and veterinary mycology par G. C. Ainsworth. Revue d’Histoire des Sciences, 43(2–3):350, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633084.

Mathieu:1990:AOBa

[3210] J.-P. Mathieu. Analyses d’ouvrages: The wisdom of science : Its relevance to culture and religion par Hanbury Brown. Revue d’Histoire des Sciences, 43(2–3):351, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633086.

Grmek:1990:AOBb

[3211] Mirko D. Grmek. Analyses d’ouvrages: The history of scurvy and vitamin C par Kenneth J. Carpentier. Revue d’Histoire des Sciences, 43(2–3): 351–352, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633087.

Debarbat:1990:AOB

[3212] Suzanne Débarbat. Analyses d’ouvrages: Fasc. 1 : Manuscrits Delambre, in Notes scientifiques et techniques du Bureau des Longitudes, S008 par Michelle Chapront–Touzé. Revue d’Histoire des Sciences, 43(2–3):352, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633088.

Laissus:1990:AOB

[3213] Yves Laissus. Analyses d’ouvrages: Jean-Henri Fabre, l’homme qui aimait les insectes, naturaliste total et pédagogue du XIX^{e} siècle, (Paris– Genève : Champion-Slatkine, 1986) par Yves Delange; Hervé Harant; Marcelle Desurvire. Revue d’Histoire des Sciences, 43(2–3):353, avril– septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633089.

Guiot:1990:AOB

[3214] Jean-Paul Guiot. Analyses d’ouvrages: Die Entdeckung des Isomorphismus. Eine Fallstudie zur Geschichte der Mineralogie und der Chemie ≪Arbor Scientiarum ^{≫}, Reihe A, Bd. IX par Hans-Werner Schütt. Revue d’Histoire des Sciences, 43(2–3):353–354, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633090.

Poulle:1990:AOBb

[3215] Emmanuel Poulle. Analyses d’ouvrages: Imprimeurs & libraires parisiens du XVI^{e} siècle, fasc. Cavellat, Marnef et Cavellat. Revue d’Histoire des Sciences, 43(2–3):354–356, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633091.

Hulin:1990:AOB

[3216] Nicole Hulin. Analyses d’ouvrages: Les Inspecteurs généraux de l’Instruction publique. Dictionnaire biographique 1802–1914, ^{≪}Histoire biographique de l’enseignement ^{≫}par Guy Caplat. Revue d’Histoire des Sciences, 43(2–3):356, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633092.

Turner:1990:AOB

[3217] A. J. Turner. Analyses d’ouvrages: All the Astrolabes par Harold N. Saunders. Revue d’Histoire des Sciences, 43(2–3):357–358, avril– septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633093.

Jacques:1990:AOB

[3218] Jean Jacques. Analyses d’ouvrages: The Letters from Gerrit Jan Mulder to Justus Liebig (1838–1846). (Janus, Suppl. vol. IX) par H. A.

M. Snelders. Revue d’Histoire des Sciences, 43(2–3):359–360, avril– septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633094.

deBeauregard:1990:AOB

[3219] O. Costa de Beauregard. Analyses d’ouvrages: Foundations of Physics par P. Weingartner; G. Dorn. Revue d’Histoire des Sciences, 43(2–3): 360–361, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633095.

Darrigol:1990:AOBa

[3220] Olivier Darrigol. Analyses d’ouvrages: Models of scientific development and the case of nuclear magnetic resonance, ^{≪}Synthese Library ^{≫}, 184 par Henk Zandvoort. Revue d’Histoire des Sciences, 43(2–3):361–362, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633096.

Payen:1990:AOBb

[3221] Jacques Payen. Analyses d’ouvrages: Un canal … Des canaux. Exposition présentée à Paris du 7 mars au 8 juin 1986 dans le cadre de la Conciergerie. Revue d’Histoire des Sciences, 43(2–3):362–363, avril– septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633097.

Mathieu:1990:AOBb

[3222] J.-P. Mathieu. Analyses d’ouvrages: Die Symmetrie der Krystalle, von René-Just Haüy zur kristallographischen Schule in Zürich par Johann Jakob Burkhardt; Erhard Scholz. Revue d’Histoire des Sciences, 43 (2–3):363, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633098.

Blay:1990:AOB

[3223] Michel Blay and Jeanne Peiffer. Analyses d’ouvrages: Sämtliche Schriften und Briefe, hrsg. von der Akademie der Wissenschaften der ddr, Dritte Reihe : Mathematischer Naturwissenschaftlicher und Technischer Briefwechsel, Zweiter Band : 1676–1679 par Gottfried Wilhelm Leibniz; Heinz Jürgen Hess. Revue d’Histoire des Sciences, 43(2–3):364, avril– septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633099.

Polanco:1990:AOB

[3224] Xavier Polanco. Analyses d’ouvrages: Histoire des sciences et des savants depuis deux siècles, d’après l’opinion des principales académies ou sociétés scientifiques. ^{≪}Corpus des Œuvres de philosophie en langue française^{≫}par Alphonse de Candolle; Bruno Latour. Revue d’Histoire des Sciences, 43(2–3):365–366, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633100.

Armogathe:1990:AOB

[3225] Jean-Robert Armogathe. Analyses d’ouvrages: Exercices pour les éléments des solides : Essai en complément d’Euclide. Progymnasmata de solidorum elementis. ^{≪}Epiméthée ^{≫} par René Descartes; Pierre Costabel. Revue d’Histoire des Sciences, 43(2–3):366–368, avril–septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633101.

Anonymous:1990:LSA

[3226] Anonyme. L’enseignement scientifique au tournant des XIX^{e} et XX^{e} siècles. Revue d’Histoire des Sciences, 43(4):369–370, October 1990. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632930.

Belhoste:1990:LSF

[3227] Bruno Belhoste. L’enseignement secondaire français et les sciences au début du XX^{e} siècle: La réforme de 1902 des plans d’études et des programmes. Revue d’Histoire des Sciences, 43(4):371–400, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632931.

Hulin:1990:DDS

[3228] Nicole Hulin. Les doctorats dans les disciplines scientifiques au XIX^{e} siècle. Revue d’Histoire des Sciences, 43(4):401–426, October 1990. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632932.

Charle:1990:PFS

[3229] Christophe Charle. Les professeurs des facultés des sciences en France : une comparaison Paris-province (1880–1900). Revue d’Histoire des Sciences, 43(4):427–450, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632933.

Telkes:1990:PFS

[3230] Eva Telkes. Présentation de la Faculté des sciences et de son personnel, à Paris (1901–1939). Revue d’Histoire des Sciences, 43(4):451–476, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632934.

Hulin:1990:EHE

[3231] Nicole Hulin. Études sur l’histoire de l’enseignement. Revue d’Histoire des Sciences, 43(4):477–480, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632935.

Anonymous:1990:Ib

[3232] Anonyme. Informations. Revue d’Histoire des Sciences, 43(4):481–482, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632936.

Emptoz:1990:AOB

[3233] Gérard Emptoz. Analyses d’ouvrages: Berthelot, 1827–1907 : Autopsie d’une mythe. ^{≪}Un savant, une époque. ^{≫}par Jean Jacques; Jean Dhombres. Revue d’Histoire des Sciences, 43(4):483–485, October 1990. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632937.

Pestre:1990:AOB

[3234] Dominique Pestre. Analyses d’ouvrages: La matière première : la recherche des particules fondamentales et de leurs interactions par Michel Crozon. Revue d’Histoire des Sciences, 43(4):485–486, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632938.

Delaporte:1990:AOB

[3235] François Delaporte. Analyses d’ouvrages: Yellow fever in the north: the methods of early epidemiology par William Coleman; Finlay, el hombre y la verdad cientifica par José Lopez Sanchez. Revue d’Histoire des Sciences, 43(4):487–488, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632939.

Caveing:1990:AOB

[3236] Maurice Caveing. Analyses d’ouvrages: The Greek Cosmologists. Vol. 1 : The formation of the atomic theory and its earliest critics par David Furley. Revue d’Histoire des Sciences, 43(4):488–489, October 1990. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632940.

Youschkevitch:1990:AOB

[3237] A. P. Youschkevitch. Analyses d’ouvrages: Mathematik in Antike und Orient par Helmut Gericke. Revue d’Histoire des Sciences, 43(4):489– 490, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632941.

Darrigol:1990:AOBb

[3238] Olivier Darrigol. Analyses d’ouvrages: Atomi metafore paradossi: Niels Bohr e la costruzione di una nuova fisica par Sandro Petruccioli. Revue d’Histoire des Sciences, 43(4):490–492, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632942.

Sinaceur:1990:AOB

[3239] M. A. Sinaceur. Analyses d’ouvrages: Œuvres mathématiques, Algèbre et géométrie au XII^{e} siècle. ^{≪}Sciences et philosophie arabes. Textes et études ^{≫}par Sharaf al–Dîn al–Tusî; R. Rashed. Revue d’Histoire des Sciences, 43(4):492–493, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632943.

Grmek:1990:AOBc

[3240] Mirko D. Grmek. Analyses d’ouvrages: Edizione nazionale delle opere. Parte prima : Carteggi. vol. I (1984), vol. II (1984), vol. III (1985), vol. IV (1985), vol. V (1985) vol. VI (1986), vol. VII (1987), vol. VIII (1987), IX (1988), X (1988), XI (1989) par Lazzaro Spallanzani; Pericle Di Pietro. Revue d’Histoire des Sciences, 43(4):493–495, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632944.

Lerner:1990:AOB

[3241] M.-P. Lerner. Analyses d’ouvrages: Œuvres d’astronomie. ^{≪}Sciences et philosophie arabes. Texte et Études. ^{≫}par Thabit ibn Qurra; Régis Morelon. Revue d’Histoire des Sciences, 43(4):495–497, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632945.

Goupil:1990:AOBb

[3242] Michelle Goupil. Analyses d’ouvrages: François Broussais, Empéreur de la Médecine par Michel Valentin; Jean Bernard. Revue d’Histoire des Sciences, 43(4):497–498, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632946.

Mathieu:1990:AOBc

[3243] J.-P. Mathieu. Analyses d’ouvrages: Kelvin and Stokes: A comparative study in Victorian Physics par David B. Wilson. Revue d’Histoire des Sciences, 43(4):499, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632947.

Valentin:1990:AOB

[3244] Michel Valentin. Analyses d’ouvrages: Les historiens français de la médecine au XIX^{e} siècle et leur bibliographie, ^{≪}L’histoire des Sciences. Textes et Études ^{≫}par Isabelle Wohnlich-Despaigne; Pierre Chaunu. Revue d’Histoire des Sciences, 43(4):499–501, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632948.

Comte:1990:AOB

[3245] Claude Comte. Analyses d’ouvrages: Mach I, Mach II, Einstein und die Relativitätstheorie : Eine Fälschung und ihre Folgen par Gereib Wolters. Revue d’Histoire des Sciences, 43(4):501–502, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632949.

Anonymous:1990:BMa

[3246] Anonyme. Back matter. Revue d’Histoire des Sciences, 43(1):??, January 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633145.

Anonymous:1990:BMb

[3247] Anonyme. Back matter. Revue d’Histoire des Sciences, 43(2–3):??, avril– septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633102.

Anonymous:1990:BMc

[3248] Anonyme. Back matter. Revue d’Histoire des Sciences, 43(4):??, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632950.

Anonymous:1990:FMa

[3249] Anonyme. Front matter. Revue d’Histoire des Sciences, 43(1):??, January 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633132.

Anonymous:1990:FMb

[3250] Anonyme. Front matter. Revue d’Histoire des Sciences, 43(2–3):??, avril– septembre 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633074.

Anonymous:1990:FMc

[3251] Anonyme. Front matter. Revue d’Histoire des Sciences, 43(4):??, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632929.

Debru:1991:CME

[3252] Claude Debru. Créativité de la médecine expérimentale : sur quelques questions d’endocrinologie et de neuro-endocrinologie. Revue d’Histoire des Sciences, 44(1):3–21, January 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632814.

Petit:1991:ESC

[3253] Annie Petit. Enseignement scientifique et culture selon Ernest Renan. Revue d’Histoire des Sciences, 44(1):23–60, January 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632815.

Langins:1991:PEP

[3254] Janis Langins. La préhistoire de l’École polytechnique. Revue d’Histoire des Sciences, 44(1):61–89, January 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632816.

Festa:1991:GH

[3255] Egidio Festa. Galilée hérétique? Revue d’Histoire des Sciences, 44(1):91– 116, January 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632817.

Gaudant:1991:AGF

[3256] Jean Gaudant. Albert Gaudry (1827–1908) et les ^{≪}Enchaînements du Monde animal ^{≫}. Revue d’Histoire des Sciences, 44(1):117–127, January 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632818.

Mathieu:1991:AOBa

[3257] J.-P. Mathieu. Analyses d’ouvrages: Fifty years of neutron diffraction: The advent of neutron Scattering par G. E. Bacon. Revue d’Histoire des Sciences, 44(1):129, January 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632819.

Simon:1991:AOB

[3258] Gérard Simon. Analyses d’ouvrages: La Structure poétique du monde : Copernic, Kepler. ^{≪}Des Travaux ^{≫} par Fernand Hallyn. Revue d’Histoire des Sciences, 44(1):129–130, January 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632820.

Lagarde:1991:AOB

[3259] Lucie Lagarde. Analyses d’ouvrages: Cartography in France, 1660–1848 : Science, engineering and statecraft par Josef W. Konvitz; Emmanuel Le Roy Ladurie. Revue d’Histoire des Sciences, 44(1):131–132, January 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632821.

Blanckaert:1991:AOB

[3260] Claude Blanckaert. Analyses d’ouvrages: Paléontologie et évolution en France de 1800 à 1860. Une histoire des idées de Cuvier et Lamarck à Darwin, ^{≪}Mémoires de la Section d’histoire des sciences et des techniques ≫, 4 par Goulven Laurent; Charles C. Gillispie. Revue d’Histoire des Sciences, 44(1):132–134, January 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632822.

Rouvillois:1991:AOB

[3261] Bertrand Rouvillois. Analyses d’ouvrages: La Machine locomotive en France des origines au milieu du XIX^{e} siècle par Jacques Payen; Bernard Escudié; Jean-Marc Combe. Revue d’Histoire des Sciences, 44(1):134– 135, January 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632823.

Debarbat:1991:AOBa

[3262] Suzanne Débarbat. Analyses d’ouvrages: James Lick’s monument: The saga of captain Richard Floyd and the building of the Lick Observatory par Helen Wright. Revue d’Histoire des Sciences, 44(1):135–136, January 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632824.

Darrigol:1991:CCM

[3263] Olivier Darrigol. Cohérence et complétude de la mécanique quantique : l’exemple de ^{≪}Bohr–Rosenfeld ^{≫}. Revue d’Histoire des Sciences, 44(2):137–179, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633147.

Robinet:1991:CCM

[3264] André Robinet. La conquête de la chaire de mathématiques de Padoue par les leibniziens. Revue d’Histoire des Sciences, 44(2):181–201, avril– juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633148.

Duris:1991:QLI

[3265] Pascal Duris. Quatre lettres inédites de Jean-Henri Fabre à Léon Dufour. Revue d’Histoire des Sciences, 44(2):203–218, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633149.

Drouin:1991:SLS

[3266] Jean-Marc Drouin. Une société locale de sciences naturelles : la Société linnéenne de provence. Revue d’Histoire des Sciences, 44(2):219–234, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633150.

Caveing:1991:TLM

[3267] Maurice Caveing. Les traductions latines médiévales des “Éléments” d’Euclide: À propos de deux publications récentes. Revue d’Histoire des Sciences, 44(2):235–239, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633151.

Mesnard:1991:CAS

[3268] Jean Mesnard. Sur le chemin de l’Académie des sciences: le cercle du mathématicien Claude Mylon (1654–1660). Revue d’Histoire des Sciences, 44(2):241–251, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633152.

Anonymous:1991:Ia

[3269] Anonyme. Informations. Revue d’Histoire des Sciences, 44(2):253–254, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633153.

Balibar:1991:AOB

[3270] Françoise Balibar. Analyses d’ouvrages: The Collected Papers of Albert Einstein, vol. I : Early years, 1879–1902 par John Stachel. Revue d’Histoire des Sciences, 44(2):255–257, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633154.

Darrigol:1991:AOBa

[3271] Olivier Darrigol. Analyses d’ouvrages: Tributes to Paul Dirac par J. G. Taylor. Revue d’Histoire des Sciences, 44(2):258, avril–juin 1991. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633155.

Delaporte:1991:AOB

[3272] François Delaporte. Analyses d’ouvrages: Histoire du sida : début et origine d’une pandémie actuelle, ^{≪}Médecine et Sociétés Payot ^{≫} par Mirko D. Grmek. Revue d’Histoire des Sciences, 44(2):258–260, avril– juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633156.

Gabbey:1991:AOB

[3273] Alan Gabbey. Analyses d’ouvrages: Jean Picard et les débuts de l’astronomie de précision au XVII^{e} siècle par Guy Picolet. Revue d’Histoire des Sciences, 44(2):260–261, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633157.

Grmek:1991:AOBa

[3274] Mirko D. Grmek. Analyses d’ouvrages: Professional and popular medicine in France, 1770–1830 : The social world of medical practice, ^{≪}Cambridge History of Medicine ^{≫}par Matthew Ramsey. Revue d’Histoire des Sciences, 44(2):262, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633159.

Mathieu:1991:AOBb

[3275] Jean-Paul Mathieu. Analyses d’ouvrages: Physical Science in the Middle Ages, ^{≪}The Cambridge History of Science Series ^{≫}par Edward Grant.

Revue d’Histoire des Sciences, 44(2):262, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633158.

Gohau:1991:AOBa

[3276] Gabriel Gohau. Analyses d’ouvrages: Theory of Earth Science par Wolf von Engelhardt; Jörg Zimmermann; L. Fischer. Revue d’Histoire des Sciences, 44(2):263, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633160.

Gohau:1991:AOBb

[3277] Gabriel Gohau. Analyses d’ouvrages: Drifting continents and shifting theories : The modern revolution in geology and shifting theories par H. E. Le Grand. Revue d’Histoire des Sciences, 44(2):263–264, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633161.

Bernhardt:1991:AOB

[3278] Jean Bernhardt. Analyses d’ouvrages: Correspondance du P. Marin Mersenne, t. XVII : Supplément par Cornelis De Waard; B. Rochot; Armand Beaulieu; René Taton. Revue d’Histoire des Sciences, 44(2):265, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633162.

Goupil:1991:AOB

[3279] Michelle Goupil. Analyses d’ouvrages: De la richesse territoriale du Royaume de France par A.-L. Lavoisier; Jean-Claude Perrot. Revue d’Histoire des Sciences, 44(2):266, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633163.

Levy:1991:AOB

[3280] Jacques Lévy. Analyses d’ouvrages: Ces astronomes fous du ciel, ou l’histoire de l’observation des étoiles doubles par Paul Couteau. Revue d’Histoire des Sciences, 44(2):266–267, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633164.

Parmentier:1991:AOB

[3281] Marc Parmentier. Analyses d’ouvrages: Classical Probability in the Enlightenment par Lorraine Daston. Revue d’Histoire des Sciences, 44(2):267–269, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633165.

Debarbat:1991:AOBb

[3282] Suzanne Débarbat. Analyses d’ouvrages: Astronomical centers of the world par Kevin Krisciunas. Revue d’Histoire des Sciences, 44(2):269– 270, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633166.

Anonymous:1991:PR

[3283] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 44(2):271– 278, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633167.

Niderst:1991:I

[3284] Alain Niderst. Introduction. Revue d’Histoire des Sciences, 44(3–4): 279–280, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632862.

Rappaport:1991:FIE

[3285] Rhoda Rappaport. Fontenelle interprets the Earth’s history. Revue d’Histoire des Sciences, 44(3–4):281–300, juillet–décembre 1991. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632863.

Mortureux:1991:REM

[3286] Marie-Françoise Mortureux. De la ressemblance entre les mathématiques et l’amour dans les Entretiens de Fontenelle. Revue d’Histoire des Sciences, 44(3–4):301–311, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632864.

Niderst:1991:LAV

[3287] Alain Niderst. Les “lettres anglaises” de Voltaire, une vulgarisation méthodique et imprudente. Revue d’Histoire des Sciences, 44(3–4): 313–323, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632865.

Cohen:1991:BMD

[3288] Claudine Cohen. Benoît de Maillet et la diffusion de l’histoire naturelle à l’aube des lumières. Revue d’Histoire des Sciences, 44(3–4):

325–342, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632866.

Artigas-Menant:1991:VSB

[3289] Geneviève Artigas-Menant. La vulgarisation scientifique dans “Le Nouveau Magasin français” de Mme Leprince de Beaumont. Revue d’Histoire des Sciences, 44(3–4):343–357, juillet–décembre 1991. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632867.

Duranton:1991:DNH

[3290] Henri Duranton. La diffusion d’une nouvelle histoire : les avatars de Clio au XVIII^{e} siècle. Revue d’Histoire des Sciences, 44(3–4):359–374, juillet– décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632868.

Mayer:1991:DCP

[3291] Jean Mayer. Diderot et le calcul des probabilités dans l’ “Encyclopédie”. Revue d’Histoire des Sciences, 44(3–4):375–391, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632869.

Poulouin:1991:CDP

[3292] Claudine Poulouin and Claudine Pouloin. La connaissance du passé et la vulgarisation du débat sur les chronologies dans l’ “Encyclopédie”. Revue d’Histoire des Sciences, 44(3–4):393–411, juillet–décembre 1991. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632870.

Rey:1991:VMA

[3293] Roselyne Rey. La vulgarisation médicale au XVIII^{e} siècle : le cas des dictionnaires portatifs de santé. Revue d’Histoire des Sciences, 44(3–4): 413–433, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632871.

Debru:1991:HOA

[3294] Armelle Debru. Hérophile, ou l’art de la médecine dans l’Alexandrie antique. Revue d’Histoire des Sciences, 44(3–4):435–445, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632872.

Piepenbrock:1991:LMM

[3295] N. Piepenbrock, M. Dürr, V. Malquarti, and R. Toellner. L’enseignement médical dans les manuels scolaires en France aux XVII^{e} et XVIII^{e} siècles. Revue d’Histoire des Sciences, 44(3–4):447–456, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632873.

Gardies:1991:PDL

[3296] Jean-Louis Gardies. La proposition 14 du livre V dans l’économie des “Éléments” d’Euclide. Revue d’Histoire des Sciences, 44(3–4): 457–467, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/ 23632874. Voir remarques [3497].

Rey:1991:JRH

[3297] Roselyne Rey and Jean-Louis Fischer. Jacques Roger, historien des sciences (1920–1990). Revue d’Histoire des Sciences, 44(3–4):469– 478, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632875.

Anonymous:1991:Ib

[3298] Anonyme. Informations. Revue d’Histoire des Sciences, 44(3–4):479– 480, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632876.

Markovits:1991:AOB

[3299] Francine Markovits. Analyses d’ouvrages: Fontenelle. Actes du Colloque tenu à Rouen du 6 au 10 octobre 1987 par Alain Niderst; Jean Mesnard. Revue d’Histoire des Sciences, 44(3–4):481–483, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632877.

Rey:1991:AOB

[3300] Roselyne Rey. Analyses d’ouvrages: Fontenelle. ^{≪}Collection biographique ^{≫}par Alain Niderst. Revue d’Histoire des Sciences, 44(3–4): 483–484, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632878.

Gayon:1991:AOB

[3301] Jean Gayon. Analyses d’ouvrages: Science and Religion. Baden Powell and the Anglican debate, 1800–1860 par Pietro Corsi. Revue d’Histoire des Sciences, 44(3–4):484–485, juillet–décembre 1991. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632879.

Elie:1991:AOB

[3302] Maurice Elie. Analyses d’ouvrages: Optique. précédé de la Lumière aujourd’hui et suivi de Études sur l’optique newtonienne. Coll. ^{≪}Epistémé classiques ^{≫}par Isaac Newton; Jean-Paul Marat; Françoise Balibar; Michel Blay. Revue d’Histoire des Sciences, 44(3–4):485–487, juillet– décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632880.

Caveing:1991:AOBa

[3303] Maurice Caveing. Analyses d’ouvrages: The Ancient Tradition of Geometric Problems par Wilbur Richard Knorr. Revue d’Histoire des Sciences, 44(3–4):487–489, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632881.

Caveing:1991:AOBb

[3304] Maurice Caveing. Analyses d’ouvrages: The Mathematics of Plato’s Academy, A new reconstruction par D. H. Fowler. Revue d’Histoire des Sciences, 44(3–4):489–491, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632882.

Taton:1991:AOB

[3305] René Taton. Analyses d’ouvrages: Claude Perrault, 1613–1688, ou la curiosité d’un classique par Antoine Picon. Revue d’Histoire des Sciences, 44(3–4):491–494, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632883.

Russo:1991:AOB

[3306] François Russo. Analyses d’ouvrages: The History of Sciences: The French Debate, ^{≪}New Trends in French Social Sciences ^{≫}par Pietro Redondi; P. V. Pillai. Revue d’Histoire des Sciences, 44(3–4):494– 495, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632884.

Provost:1991:AOB

[3307] Sophie Provost. Analyses d’ouvrages: Broad and Narrow Interpretations of Philosophy of Technology. ^{≪}Official publications of the Society for Philosophy and Technology ^{≫}, vol. 7 par Paul T. Durbin. Revue d’Histoire des Sciences, 44(3–4):495–496, juillet–décembre 1991. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632885.

Richard:1991:AOB

[3308] Nathalie Richard. Analyses d’ouvrages: Boucher de Perthes : les origines romantiques de la préhistoire, ^{≪}Un savant, une époque ^{≫}par C. Cohen; J.-J. Hublin. Revue d’Histoire des Sciences, 44(3–4):496–497, juillet– décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632886.

Frey:1991:AOB

[3309] Anne Frey. Analyses d’ouvrages: Le Code alchimique dévoilé. Distillateurs, alchimistes et symbolistes par Suzanne Colnort-Bodet; Jean Glénisson; Jean Mesnard. Revue d’Histoire des Sciences, 44(3–4): 497–498, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632887.

Sinaceur:1991:AOB

[3310] Hourya Sinaceur. Analyses d’ouvrages: Boole, 1815–1864. L’oiseau de nuit en plein jour, ^{≪}Un savant, une époque ^{≫} par Souleymane Bachir Diagne; Marie–José Durand. Revue d’Histoire des Sciences, 44(3–4): 498–499, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632888.

Debru:1991:AOB

[3311] Claude Debru. Analyses d’ouvrages: Les causes de la mort. Histoire naturelle et facteurs de risque par Anne Fagot-Largeault; G. Canguilhem. Revue d’Histoire des Sciences, 44(3–4):500–501, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632889.

Grmek:1991:AOBb

[3312] Mirko Drazen Grmek. Analyses d’ouvrages: The uses of experiment. Studies in the natural sciences par David Gooding; Trevor Pinch; Simon Schaffer. Revue d’Histoire des Sciences, 44(3–4):501–502, juillet– décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632890.

Gourevitch:1991:AOBa

[3313] Danielle Gourevitch. Analyses d’ouvrages: Médecine et philosophie au temps d’Hippocrate. ^{≪}Histoires de science ^{≫}par Bernard Vitrac. Revue d’Histoire des Sciences, 44(3–4):502, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632891.

Gourevitch:1991:AOBb

[3314] Danielle Gourevitch. Analyses d’ouvrages: Das Prognostikon wurde nicht vor, sondern nach den Epidemienbüchern III und I verfasst. Zweiter Beitrag zur Chronologie der echten Hippokratischen Schriften, XIII. Hippokratische Studie. ^{≪}Études d’histoire de la médecine ^{≫}, VII par Charles Lichtenthaeler. Revue d’Histoire des Sciences, 44(3–4):502–503, juillet– décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632892.

Darrigol:1991:AOBb

[3315] Olivier Darrigol. Analyses d’ouvrages: Niels Bohr’s philosophy of physics, 2nd ed. par Dugald Murdoch. Revue d’Histoire des Sciences, 44 (3–4):503, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632893.

Kahn:1991:AOB

[3316] Didier Kahn. Analyses d’ouvrages: Johannes’ de Rupescissa ^{≪}Liber de consideratione quintae essentiae omnium rerum ^{≫}deutsch. Studien zur Alchemia medica des 15. bis 17. Jahrhunderts mit kritischer Edition des Textes. ^{≪}Heidelberger Studien zur Naturkunde der frühen Neuzeit ≫, Bd. 1 par Udo Benzenhöfer; Georg von Welling (1655–1727). Studien zu Leben und Werk. ^{≪}Heidelberger Studien zur Naturkunde der frühen Neuzeit ^{≫}, Bd. 2 par Petra Jungmayr. Revue d’Histoire des Sciences, 44(3–4):503–504, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632894.

Grmek:1991:AOBc

[3317] Mirko Drazen Grmek. Analyses d’ouvrages: The Polish School of philosophy of medicine. From Tytus Chalubinski (1820–1889) to Ludwik Fleck (1896–1961), ^{≪}Philosophy and Medicine ^{≫}, 37 par Ilana Löwy. Revue d’Histoire des Sciences, 44(3–4):504, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632895.

Pestre:1991:AOB

[3318] Dominique Pestre. Analyses d’ouvrages: La Physique des particules élémentaires, de sa naissance à sa maturité, 1930–1960 par Jeanne Laberrigue-Frolow. Revue d’Histoire des Sciences, 44(3–4): 504–505, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632896.

Acot:1991:AOB

[3319] Pascal Acot. Analyses d’ouvrages: Nature as the laboratory. Darwinian plant ecology in the German Empire, 1880–1900 par Eugene Cittadino. Revue d’Histoire des Sciences, 44(3–4):505, juillet–décembre 1991. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23632897.

Anonymous:1991:BMa

[3320] Anonyme. Back matter. Revue d’Histoire des Sciences, 44(1):??, January 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632825.

Anonymous:1991:BMb

[3321] Anonyme. Back matter. Revue d’Histoire des Sciences, 44(2):??, avril– juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633168.

Anonymous:1991:BMc

[3322] Anonyme. Back matter. Revue d’Histoire des Sciences, 44(3–4): ??, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632898.

Anonymous:1991:FMa

[3323] Anonyme. Front matter. Revue d’Histoire des Sciences, 44(1):??, January 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632813.

Anonymous:1991:FMb

[3324] Anonyme. Front matter. Revue d’Histoire des Sciences, 44(2):??, avril– juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633146.

Anonymous:1991:FMc

[3325] Anonyme. Front matter. Revue d’Histoire des Sciences, 44(3–4): ??, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632861.

Anonymous:1992:EC

[3326] Anonyme. Études sur Cauchy (1789–1857). Revue d’Histoire des Sciences, 45(1):3–4, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633049.

Truesdell:1992:CMM

[3327] Clifford A. Truesdell. Cauchy and the modern mechanics of continua. Revue d’Histoire des Sciences, 45(1):5–24, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633050. Études sur Cauchy (1789–1857).

Dhombres:1992:REF

[3328] Jean Dhombres. Le rôle des équations fonctionnelles dans l’Analyse algébrique de Cauchy. Revue d’Histoire des Sciences, 45(1):25–49, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633051. Études sur

Cauchy (1789–1857).

Sinaceur:1992:CSR

[3329] Hourya Sinaceur. Cauchy, Sturm et les racines des équations. Revue d’Histoire des Sciences, 45(1):51–67, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www. jstor.org/stable/23633052. Études sur Cauchy (1789–1857).

Gray:1992:CEA

[3330] Jeremy Gray. Cauchy: elliptic and Abelian integrals. Revue d’Histoire des Sciences, 45(1):69–81, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/ stable/23633053. Études sur Cauchy (1789–1857).

Dalmedico:1992:LED

[3331] Amy Dahan Dalmedico. L’intégration des équations aux dérivées partielles linéaires à coefficients constants dans les travaux de Cauchy (1821– 1830). Revue d’Histoire des Sciences, 45(1):83–114, January 1992. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633054. Études sur Cauchy (1789–

1857).

Laugwitz:1992:EDF

[3332] Detlef Laugwitz. Early delta functions and the use of infinitesimals in research. Revue d’Histoire des Sciences, 45(1):115–128, January 1992. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633055. Études sur Cauchy (1789–

1857).

Dhombres:1992:BCC

[3333] Jean Dhombres and Christian Gilain. Bibliographie concernant Cauchy (1974 à aujourd’hui). Revue d’Histoire des Sciences, 45(1):129–133, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633056. Études sur

Cauchy (1789–1857).

Chemla:1992:NIC

[3334] Karine Chemla. Des nombres irrationnels en Chine entre le premier et le troisième siècle. Revue d’Histoire des Sciences, 45(1):135–140, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633057. Études sur

Cauchy (1789–1857).

Anonymous:1992:Ia

[3335] Anonyme. Informations. Revue d’Histoire des Sciences, 45(1):141–143, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633058.

Blay:1992:AOB

[3336] Michel Blay. Analyses d’ouvrages: Logique des sciences de la culture, ≪Fondation critique ou fondation herméneutique des sciences de la culture? ^{≫} par Ernst Cassirer; Jean Carro; Joël Gaubert; Ernst Cassirer. De Marbourg à New York par Jean Seidengart. Revue d’Histoire des Sciences, 45(1):145, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633059.

Gayon:1992:AOB

[3337] Jean Gayon. Analyses d’ouvrages: Companion to the history of modern science par R. C. Olby; G. N. Cantor; J. R. R. Christie; M. J. S. Hodge. Revue d’Histoire des Sciences, 45(1):145–147, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633060.

Grmek:1992:AOBa

[3338] Mirko Drazen Grmek. Analyses d’ouvrages: The Medical Enlightenment of the eighteenth century par Andrew Cunningham; Roger French. Revue d’Histoire des Sciences, 45(1):147–148, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633061.

Kalinowski:1992:AOB

[3339] Marc Kalinowski. Analyses d’ouvrages: Dall’Europa alla Cina : contributi per una storia dell’ astronomia par Isaia Iannaccone; Adolfo Tamburello. Revue d’Histoire des Sciences, 45(1):148–149, January 1992. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633062.

Bourienne:1992:AOBa

[3340] Véronique Bourienne. Analyses d’ouvrages: Structures in Mathematical Theories. Reports of the San Sebastian International Symposium, September 25–29, 1990 par A. Diaz; J. Echeverria; A. Ibarra. Revue d’Histoire des Sciences, 45(1):149, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633063.

Darrigol:1992:AOBa

[3341] Olivier Darrigol. Analyses d’ouvrages: Experiment, right or wrong par Allan Franklin. Revue d’Histoire des Sciences, 45(1):149–150, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633064.

Goupil:1992:AOBa

[3342] Michelle Goupil. Analyses d’ouvrages: Genèse et diffusion du système métrique par Bernard Garnier; Jean-Claude Hocquet. Revue d’Histoire des Sciences, 45(1):150, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633065.

Boyer:1992:AOB

[3343] Alain Boyer. Analyses d’ouvrages: Karl Popper, science et philosophie, Coll. ^{≪}Science-histoire-philosophie ^{≫}. [Études rassemblées après le Colloque tenu à Strasbourg, 25–27 mars 1982.] par Renée Bouveresse; Hervé Barreau. Revue d’Histoire des Sciences, 45(1):150–152, January 1992.

CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633066.

Merleau-Ponty:1992:AOB

[3344] Jacques Merleau-Ponty. Analyses d’ouvrages: Modern Cosmology in retrospect par B. Bertotti; R. Balbinot; S. Bergia; A. Messina. Revue d’Histoire des Sciences, 45(1):153, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633067.

Poulle:1992:AOB

[3345] Emmanuel Poulle. Analyses d’ouvrages: Catalogue de l’horlogerie et des instruments de précision du début du XVI^{e} siècle au milieu du XVII^{e} par Château d’Ecouen; Adolphe Chapiro; Chantal Meslin-Perrier; Anthony Turner; Hervé Oursel. Revue d’Histoire des Sciences, 45(1):153–155, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633068.

deGandillac:1992:AOB

[3346] Maurice de Gandillac. Analyses d’ouvrages: Christus-Wahrheit des Denkens. Eine Untersuchung zur Schrift ^{≪}de docta ignorantia ^{≫}des Nikolaus von Kues. ^{≪}Beiträge zur Geschichte der Philosophie und Theologie des Mittelalters ^{≫}, Neue Folge, Bd. 33 par Ulrich Offermann. Revue d’Histoire des Sciences, 45(1):155, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633069.

Levy:1992:AOB

[3347] Jacques Lévy. Analyses d’ouvrages: Advances in geosciences [Selected papers from the symposia ^{≪}The history of geomagnetism and aeronomy ≫and ^{≪}Problems of uncertainties in geophysical time series ^{≫}of the Interdivisional commission on history of the IAGA during the IAGA general assembly, held in Exeter, 1989.] par Wilfried Schröder. Revue d’Histoire des Sciences, 45(1):155–157, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633070.

Drouin:1992:AOB

[3348] Jean-Marc Drouin. Analyses d’ouvrages: Les Origines de la Société de physique et d’histoire naturelle de Genève (1790–1822). La science genevoise face au modèle français. ^{≪}Mémoires de la SPHN ^{≫}, 45 (1990), fasc. 1 par René Sigrist; Histoire des sciences de la terre entre 1790 et 1815 vue à travers les documents inédits de la Société de physique et d’histoire naturelle de Genève. Trois grands protagonistes: Marc-Auguste Pictet, Guillaume-Antoine Deluc et Jean Tollot. ^{≪}Mémoires de la SPHN ≫, 45 (1990), fasc. 2 par Albert V. Carozzi. Revue d’Histoire des Sciences, 45(1):157, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633071.

Bourienne:1992:AOBb

[3349] Véronique Bourienne. Analyses d’ouvrages: L’histoire des sciences et des techniques, la philosophie des sciences, dans ^{≪}La Recherche ^{≫}, ^{≪}Pour la science ^{≫}, ^{≪}Revue d’histoire des sciences ^{≫}, ^{≪}Revue des questions scientifiques ^{≫}, 1937–1990. Bibliographie par Jean-François Stoffel. Revue d’Histoire des Sciences, 45(1):157–158, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633072.

Mayaud:1992:FNF

[3350] Pierre-Noël Mayaud. Une ^{≪}nouvelle ^{≫}affaire Galilée? Revue d’Histoire des Sciences, 45(2–3):161–230, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632995.

Souffrin:1992:HDC

[3351] Pierre Souffrin. Sur l’histoire du concept de vitesse d’Aristote à Galilée. Revue d’Histoire des Sciences, 45(2–3):231–267, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632996.

Gautero:1992:NDF

[3352] Jean-Luc Gautero and Pierre Souffrin. Note sur la démonstration “mécanique” du théorème de l’isochronisme des cordes du cercle dans les Discorsi de Galilée. Revue d’Histoire des Sciences, 45(2–3):269–280, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632997.

Gapaillard:1992:GPD

[3353] Jacques Gapaillard. Galilée et le principe du chasseur. Revue d’Histoire des Sciences, 45(2–3):281–306, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632998.

Festa:1992:NAI

[3354] Egidio Festa. La notion d’ ^{≪}agrégat d’indivisibles ^{≫}dans la constitution de la cinématique galiléenne : Cavalieri, Galilée, Torricelli. Revue d’Histoire des Sciences, 45(2–3):307–336, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632999.

Bailhache:1992:IBD

[3355] Patrice Bailhache. Isaac Beeckman a-t-il démontré la loi des cordes vibrantes selon laquelle la fréquence est inversement proportionnelle à la longueur? Revue d’Histoire des Sciences, 45(2–3):337–344, avril– septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633000.

Breny:1992:PEN

[3356] Henri Breny. Premières évaluations numériques pour une ^{≪}loi des gaz ^{≫}. Revue d’Histoire des Sciences, 45(2–3):345–359, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633001.

Minazzi:1992:FIL

[3357] Fabio Minazzi. L’ ^{≪}Institut Ludovico Geymonat pour la philosophie de la science, la logique et l’histoire de la science et de la technique ^{≫}. Revue d’Histoire des Sciences, 45(2–3):361–365, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633002.

Anonymous:1992:Ib

[3358] Anonyme. Informations. Revue d’Histoire des Sciences, 45(2–3):367–368, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633003.

Bos:1992:AOB

[3359] Henk J. M. Bos. Analyses d’ouvrages: La méthode de Fermat : Son statut et sa diffusion, Cahiers d’histoire et de philosophie des sciences, nouvelle série, n^{o} 33 par Giovanna Cleonice Cifoletti. Revue d’Histoire des Sciences, 45(2–3):369–371, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633004.

Corbellini:1992:AOB

[3360] Gilberto Corbellini. Analyses d’ouvrages: Le dernier langage de la médecine : Histoire de l’immunologie de Pasteur au Sida, ^{≪}Pratiques théoriques ^{≫} par Anne-Marie Moulin; Niels K. Jerne. Revue d’Histoire des Sciences, 45(2–3):371–372, avril–septembre 1992. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633005.

Darrigol:1992:AOBb

[3361] Olivier Darrigol. Analyses d’ouvrages: Chandra: A biography of S. Chandrasekhar par Kameshwar C. Wali. Revue d’Histoire des Sciences, 45 (2–3):373, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633006.

Gourevitch:1992:AOB

[3362] Danielle Gourevitch. Analyses d’ouvrages: Vol. 1 : Galen’s method of healing. Proceedings of the 1982 Galen Symposium, Vol. 2 : Hippocrates, Pseudepigraphic writings: Letters, Embassy, Speech from the Altar, Decree, Vol. 3 : Garum and Salsamenta: Production and commerce in materia medica par Fridolf Kudlien; Richard J. Durling; Wesley D. Smith; Robert I. Curtis. Revue d’Histoire des Sciences, 45(2–3):373–374, avril– septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633007.

Grmek:1992:AOBb

[3363] Mirko Drazen Grmek. Analyses d’ouvrages: On fracture of the skull or cranium. ^{≪}Transactions of the American Philosophy Society ^{≫}, vol. 80, part 4 par Berengario Da Carpi; L. R. Lind. Revue d’Histoire des Sciences, 45(2–3):374–375, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633008.

Moulin:1992:AOB

[3364] Anne-Marie Moulin. Analyses d’ouvrages: The Threat and the Glory: Reflections on Science and Scientists par Peter Medawar; David Pyke; Lewis Thomas. Revue d’Histoire des Sciences, 45(2–3):375–376, avril– septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633009.

Niderst:1992:AOB

[3365] Alain Niderst. Analyses d’ouvrages: Il secolo delle cose: Scienza e storia in Francesco Algarotti, ^{≪}Ricerche Studi Strumenti-Letteratura ^{≫}par Franco Arato. Revue d’Histoire des Sciences, 45(2–3):376–377, avril– septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633010.

Payen:1992:AOB

[3366] Jacques Payen. Analyses d’ouvrages: Histoire de l’hydraulique. L’eau conquise : les origines et le monde antique par Marcel Nordon. Revue d’Histoire des Sciences, 45(2–3):377, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633011.

Russo:1992:AOBa

[3367] François Russo. Analyses d’ouvrages: Science and religion: Some historical perspectives, ^{≪}Cambridge History of Science ^{≫}par John Hedley Brooke. Revue d’Histoire des Sciences, 45(2–3):378–379, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633012.

Theodorides:1992:AOB

[3368] Jean Théodoridès. Analyses d’ouvrages: The early evolution of Metazoa and the significance of problematic taxa, Proceedings of an International Symposium held at the University of Camerino, 27–31 March 1989 par Alberto M. Simonetta; Simon Conway Morris. Revue d’Histoire des Sciences, 45(2–3):379–380, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633013.

Darrigol:1992:AOBc

[3369] Olivier Darrigol. Analyses d’ouvrages: Dirac : A scientific biography par Helge S. Kragh. Revue d’Histoire des Sciences, 45(2–3):380–381, avril– septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633014.

Elie:1992:AOB

[3370] Maurice Élie. Analyses d’ouvrages: Goethe’s history of science par Karl J. Fink. Revue d’Histoire des Sciences, 45(2–3):381, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633015.

Darrigol:1992:AOBd

[3371] Olivier Darrigol. Analyses d’ouvrages: Niels Bohr’s times, in physics, philosophy, and polity par Abraham Pais. Revue d’Histoire des Sciences, 45(2–3):382–383, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633016.

Darrigol:1992:AOBe

[3372] Olivier Darrigol. Analyses d’ouvrages: Physique atomique et connaissance humaine, ^{≪}Folio/essais ^{≫}, 157 par Niels Bohr; Edmond Bauer; Roland Omnes; Catherine Chevalley. Revue d’Histoire des Sciences, 45 (2–3):383, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633017.

Rey:1992:EEB

[3373] Roselyne Rey. Espèces, espaces : la biogéographie sans frontières. Revue d’Histoire des Sciences, 45(4):385–388, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632961.

Laurent:1992:BPF

[3374] Goulven Laurent. Biogéographie et paléontologie en France au XIX^{e} siècle : acteurs et débats. Revue d’Histoire des Sciences, 45(4):389–417, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632962.

Castrillon:1992:AHG

[3375] Alberto Castrillon. Alexandre de Humboldt et la géographie des plantes. Revue d’Histoire des Sciences, 45(4):419–433, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632963.

Muller:1992:RBA

[3376] Gerhard H. Müller. Ratzel et la biogéographie en Allemagne dans la deuxième moitié du XIX^{e} siècle. Revue d’Histoire des Sciences, 45(4):435–452, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632964.

Browne:1992:SEB

[3377] Janet Browne. A science of empire: British biogeography before Darwin. Revue d’Histoire des Sciences, 45(4):453–475, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632965.

Deleage:1992:OSE

[3378] Jean-Paul Deléage. Aux origines de la science écologique : À propos de quelques ouvrages récents. Revue d’Histoire des Sciences, 45(4):477–490, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632966.

Nicolaidis:1992:LLF

[3379] Efthymios Nicolaïdis and Dimitrios Dialetis. L’influence des Lumières sur la formation scientifique grecque : une approche par l’analyse factorielle. Revue d’Histoire des Sciences, 45(4):491–499, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632967.

Anonymous:1992:Ic

[3380] Anonyme. Informations. Revue d’Histoire des Sciences, 45(4):501–502, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632968.

Horiuchi:1992:AOB

[3381] Annick Horiuchi. Analyses d’ouvrages: The formation of science in Japan: Building a research tradition par James R. Bartholomew. Revue d’Histoire des Sciences, 45(4):503, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632969.

Veuille:1992:AOB

[3382] Michel Veuille. Analyses d’ouvrages: Darwin, darwinisme, évolutionnisme. ≪Histoire des idées, théorie politique et recherche en sciences sociales ≫par Daniel Becquemont. Revue d’Histoire des Sciences, 45(4):504–505, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632970.

Kounelis:1992:AOB

[3383] Catherine Kounelis. Analyses d’ouvrages: Sunto di un corso di filosofia chimica par Stanislao Cannizzaro; Luigi Cerruti; Leonello Paoloni. Revue d’Histoire des Sciences, 45(4):505–506, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632971.

Chevalley:1992:AOB

[3384] Catherine Chevalley. Analyses d’ouvrages: Ludwig Boltzmann Principien der Naturfilosofi. Lectures on Natural Philosophy, 1903–1906 par Ilse M. Fasol-Boltzmann; S. G. Brusch; G. Fasol. Revue d’Histoire des Sciences, 45(4):506–507, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632972.

Dubucs:1992:AOB

[3385] Jacques Dubucs. Analyses d’ouvrages: Le raisonnement par l’absurde. ≪Bibliothèque d’histoire des sciences ^{≫}par Jean-Louis Gardies. Revue d’Histoire des Sciences, 45(4):507–508, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632973.

Goupil:1992:AOBb

[3386] Michelle Goupil. Analyses d’ouvrages: Eighteenth-century chemistry as an investigative enterprise. Five lectures delivered at the International Summer School in History of Science, Bologna, August 1988 par Frederic Lawrence Holmes. Revue d’Histoire des Sciences, 45(4):508, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632974.

Russo:1992:AOBb

[3387] François Russo. Analyses d’ouvrages: Regards historiques sur les sciences, Impact : Science et Société, I : n^{o} 159 (vol. 40, 1990, n^{o} 3), II : n^{o} 160 (vol. 40, 1990, n^{o} 4). Revue d’Histoire des Sciences, 45(4):509, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632975.

Vernant:1992:AOB

[3388] Denis Vernant. Analyses d’ouvrages: The mathematical philosophy of Bertrand Russel: Origins and development par Francisco A. Rodriguez-Consuegra. Revue d’Histoire des Sciences, 45(4):509–510, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23632976.

Laurent:1992:AOB

[3389] Goulven Laurent. Analyses d’ouvrages: Reading the Shape of Nature: Comparative Zoology at the Agassiz Museum. ^{≪}Science and its conceptual foundations ^{≫}par Mary P. Winsor. Revue d’Histoire des Sciences, 45(4):510–511, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632977.

Anonymous:1992:PR

[3390] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 45(4):513– 522, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23632978.

Anonymous:1992:BMa

[3391] Anonyme. Back matter. Revue d’Histoire des Sciences, 45(1):??, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633073.

Anonymous:1992:BMb

[3392] Anonyme. Back matter. Revue d’Histoire des Sciences, 45(2–3):??, avril– septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633018.

Anonymous:1992:BMc

[3393] Anonyme. Back matter. Revue d’Histoire des Sciences, 45(4):??, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632979.

Anonymous:1992:FMa

[3394] Anonyme. Front matter. Revue d’Histoire des Sciences, 45(1):??, January 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633048.

Anonymous:1992:FMb

[3395] Anonyme. Front matter. Revue d’Histoire des Sciences, 45(2–3):??, avril– septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632994.

Anonymous:1992:FMc

[3396] Anonyme. Front matter. Revue d’Histoire des Sciences, 45(4):??, October 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632960.

Blondel:1993:P

[3397] Christine Blondel, Patrice A. Carré, and Nicole Hulin. Présentation. Revue d’Histoire des Sciences, 46(1):3–6, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633739.

Hulin:1993:EBF

[3398] Nicole Hulin. Édouard Branly, la formation d’un physicien parmi d’autres. Revue d’Histoire des Sciences, 46(1):7–26, January 1993. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633740.

Cazenobe:1993:BEI

[3399] Jean Cazenobe. Branly est-il l’inventeur … du tube de Branly? Revue d’Histoire des Sciences, 46(1):27–37, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633741.

Blondel:1993:BFI

[3400] Christine Blondel. Branly face à l’innovation technique : un cas d’espèce? Revue d’Histoire des Sciences, 46(1):39–58, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633742.

Kleinert:1993:FBD

[3401] Andreas Kleinert. Ferdinand Braun et les débuts de la TSF en Allemagne. Revue d’Histoire des Sciences, 46(1):59–71, January 1993. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633743.

Brenni:1993:TLC

[3402] Paolo Brenni. Le tube à limaille de Calzecchi Onesti et la TSF de Marconi: Quelques réflexions sur les débuts de la TSF en Italie. Revue d’Histoire des Sciences, 46(1):73–82, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633744.

Pestre:1993:CAE

[3403] Dominique Pestre. Contrepoint autour d’Édouard Branly : trois propositions de révision. Revue d’Histoire des Sciences, 46(1):83–93, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633745.

Blondel:1993:BCE

[3404] Christine Blondel. Bibliographie concernant Édouard Branly. Revue d’Histoire des Sciences, 46(1):95–96, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633746.

Giraud:1993:DCS

[3405] Guy Giraud and Roland Faure. Du cohéreur à la science des milieux granulaires. Revue d’Histoire des Sciences, 46(1):97–103, January 1993. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633747.

Anonymous:1993:Ia

[3406] Anonyme. Informations. Revue d’Histoire des Sciences, 46(1):105–106, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633748.

Richard:1993:AOB

[3407] Nathalie Richard. Analyses d’ouvrages: Aux origines d’homo sapiens, ≪Nouvelle Encyclopédie Diderot ^{≫}par Jean-Jacques Hublin; Anne-Marie Tillier. Revue d’Histoire des Sciences, 46(1):107–108, January 1993. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633749.

Gourevitch:1993:AOBa

[3408] Michel Gourevitch. Analyses d’ouvrages: Citoyen fou (Le), ^{≪}Nouvelle Encyclopédie Diderot ^{≫} [Douze contributions. Contient un compte rendu de la table ronde ^{≪}Le citoyen fou ^{≫}du 12 février 1991 à la Cité des sciences et de l’industrie, à laquelle participaient 8 des 12 auteurs des contributions.] par Nathalie Robatel; Dominique Lecourt. Revue d’Histoire des Sciences, 46(1):108–109, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633750.

Levy:1993:AOBa

[3409] Jacques Lévy. Analyses d’ouvrages: Colloque André Danjon, 28–30 mai 1990 par N. Capitaine; S. Débarbat. Revue d’Histoire des Sciences, 46(1):109–110, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633751.

Darrigol:1993:AOBa

[3410] Olivier Darrigol. Analyses d’ouvrages: Theory construction and selection in modern physics. The S Matrix par James T. Cushing. Revue d’Histoire des Sciences, 46(1):110–111, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633752.

Chevalley:1993:AOB

[3411] Catherine Chevalley. Analyses d’ouvrages: Niels Bohr. His Heritage and Legacy: An Anti-realist View of Quantum Mechanics, ^{≪}Science and Philosophy ^{≫}, vol. 6 par Jan Faye; Niels Bohr’s Philosophical Background, ^{≪}Mat. Fys. Medd. Dan. Vid. Selsk. ^{≫}, 63 par David Favrholdt. Revue d’Histoire des Sciences, 46(1):111–113, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633753.

Monti:1993:AOB

[3412] Maria Teresa Monti. Analyses d’ouvrages: La naissance de la vie : Une anthologie. ^{≪}Agora. Les classiques ^{≫}, 92 par Jean-Louis Fischer; J.-L. Fischer. Revue d’Histoire des Sciences, 46(1):113–115, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633754.

Grison:1993:AOB

[3413] Emmanuel Grison. Analyses d’ouvrages: Du flou au clair? Histoire de l’affinité chimique de Cardan à Prigogine, ^{≪}Mémoires de la section d’histoire des sciences et des techniques ^{≫}, 7 par Michelle Goupil. Revue d’Histoire des Sciences, 46(1):115–117, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633755.

Festa:1993:AOB

[3414] Egidio Festa. Analyses d’ouvrages: La filosofia di Francesco Buonamici, professore di Galileo a Pisa par Mario Otto Helbing. Revue d’Histoire des Sciences, 46(1):117, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633756.

Veuille:1993:AOB

[3415] Michel Veuille. Analyses d’ouvrages: Histoire de la génétique. Pratiques, techniques et théories par Jean-Louis Fischer; William H. Schneider. Revue d’Histoire des Sciences, 46(1):118–119, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633757.

Grmek:1993:AOBa

[3416] Mirko D. Grmek. Analyses d’ouvrages: Antike Krankenbehandlung in historisch-vergleichender Sicht. Studien zum konkret-anschaulichen Denken, ^{≪}Bibliothek der klassischen Altertumswissenschaften ^{≫}, Reihe 2, Neue Folge, Bd. 81 par Günther Lorenz. Revue d’Histoire des Sciences, 46(1):119, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633758.

Gourevitch:1993:AOBb

[3417] Danielle Gourevitch. Analyses d’ouvrages: Hippocratic lives and legends, ^{≪}Studies in ancient medicine ^{≫}, vol. 4 par Jody Rubin Pinault.

Revue d’Histoire des Sciences, 46(1):119–120, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633759.

Grmek:1993:AOBb

[3418] Mirko D. Grmek. Analyses d’ouvrages: The Correspondence between Albrecht von Haller and Horace–Bénédict de Saussure, ^{≪}Studia Halleriana ≫, Bd. III par Otto Sonntag. Revue d’Histoire des Sciences, 46(1):120– 121, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633760.

Roux:1993:AOB

[3419] Sophie Roux. Analyses d’ouvrages: Newtons Entwurf “Über die Gravitation … ”: Ein Stück Entwicklungsgeschichte seiner Mechanik,^{≪}Boethius ≫, Bd. XXVI par Friedrich Steinle. Revue d’Histoire des Sciences, 46(1):121–123, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633761.

Levy:1993:AOBb

[3420] Tony Lévy. Analyses d’ouvrages: Apollonius: Conics Books V to VII. The Arabic Translation of the Lost Greek Original in the Version of the Banu Musa. I : Introd., text and translation, II : Commentary, fig. and indexes, ^{≪}Sources in the History of Mathematics and Physical Sciences ≫, 9 par Gerald J. Toomer; G. J. Toomer. Revue d’Histoire des Sciences, 46(1):123–124, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633762.

Bernhardt:1993:I

[3421] Jean Bernhardt. [introduction]. Revue d’Histoire des Sciences, 46(2–3): 129–130, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633817.

Bernhardt:1993:ERS

[3422] Jean Bernhardt. Empirisme rationnel et statut des “universalia”: le problème de la théorie de la science chez Hobbes. Revue d’Histoire des Sciences, 46(2–3):131–152, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633818.

Jesseph:1993:AIH

[3423] Douglas M. Jesseph. Of analytics and indivisibles: Hobbes on the methods of modern mathematics. Revue d’Histoire des Sciences, 46(2–3): 153–193, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633819.

Prins:1993:WPH

[3424] Jan Prins. Ward’s polemic with Hobbes on the sources of his optical theories. Revue d’Histoire des Sciences, 46(2–3):195–224, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633820.

Bernhardt:1993:QDV

[3425] Jean Bernhardt. La question du vide chez Hobbes. Revue d’Histoire des Sciences, 46(2–3):225–232, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633821.

Cotte:1993:LMD

[3426] Michel Cotte. L’approche mathématique du pont suspendu chez Marc Seguin, 1822–1826. Revue d’Histoire des Sciences, 46(2–3):233–257, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633822.

Saillard:1993:CCE

[3427] Michel Saillard and Yves Cortial. Calcul de la courbe d’efficacité lumineuse spectrale de l’œil effectué à partir des mesures des intensités des différentes couleurs du spectre solaire de Josef Fraunhofer (1817). Revue d’Histoire des Sciences, 46(2–3):259–272, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633823.

Kryzhanovsky:1993:BLE

[3428] Leonid Kryzhanovsky and Leonid Kryzhanovski. La bouteille de Leyde et l’électrophore au XVIII^{e} siècle : des répercussions russes. Revue d’Histoire des Sciences, 46(2–3):273–280, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633824.

Romano:1993:PMJ

[3429] Antonella Romano. À propos des mathématiques jésuites: notes et réflexions sur l’ouvrage d’Albert Krayer, “Mathematik im Studienplan der Jesuiten”. Revue d’Histoire des Sciences, 46(2–3):281–292, avril– septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633825.

Hachler:1993:I

[3430] Stefan Hächler, Kathrin Ott, and Urs Boschung. Informations. Revue d’Histoire des Sciences, 46(2–3):293–295, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633826.

Gourevitch:1993:AOBc

[3431] Danielle Gourevitch. Analyses d’ouvrages: The Greek Way of Life, from conception to old age par Robert Garland. Revue d’Histoire des Sciences, 46(2–3):297, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633827.

Narcy:1993:AOB

[3432] Michel Narcy. Analyses d’ouvrages: The Uses of Antiquity: The Scientific Revolution and the Classical Tradition, ^{≪}Australian Studies in History and Philosophy of Science ^{≫}, 10 par Stephen Gaukroger. Revue d’Histoire des Sciences, 46(2–3):297–298, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633828.

Dalmedico:1993:AOB

[3433] Amy Dahan Dalmedico. Analyses d’ouvrages: La France mathématique : La Société mathématique de France (1872–1914), Cahiers d’Histoire et de Philosophie des Sciences, 34 (1991) par Hélène Gispert; R. Bkouche; C. Gilain; C. Houzel; J.-P. Kahane; M. Zerner; J.-P. Bourguignon. Revue d’Histoire des Sciences, 46(2–3):299–300, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633829.

Parmentier:1993:AOBa

[3434] Marc Parmentier. Analyses d’ouvrages: The Taming of Chance. ^{≪}Ideas in context ^{≫}par Ian Hacking; The Empire of Chance: How probability changed science and everyday life, ^{≪}Ideas in context ^{≫}par Gerd Gigerenzer; Zeno Swijtink; Théodore Porter; Lorraine Daston; John Beatty; Lorenz Krüger. Revue d’Histoire des Sciences, 46(2–3):300–304, avril– septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633830.

Rey:1993:AOBa

[3435] Roselyne Rey. Analyses d’ouvrages: Les Professeurs de la faculté de médecine de Paris: Dictionnaire biographique, 1794–1939, coll. ≪Histoire biographique de l’enseignement ^{≫}par Françoise Huguet. Revue d’Histoire des Sciences, 46(2–3):304–305, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633831.

Darrigol:1993:AOBb

[3436] Olivier Darrigol. Analyses d’ouvrages: The Maxwellians ^{≪}Cornell History of Science Series ^{≫}par Bruce J. Hunt. Revue d’Histoire des Sciences, 46(2–3):305, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633832.

Parmentier:1993:AOBb

[3437] Marc Parmentier. Analyses d’ouvrages: Los Inicios de la Teoría de la Probabilidad : Siglos XVI y XVII par Marisol De Mora Charles. Revue d’Histoire des Sciences, 46(2–3):306, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633833.

Hugonnard-Roche:1993:AOB

[3438] Henri Hugonnard-Roche. Analyses d’ouvrages: Astronomica et astrologica cracoviensia ante annum 1550, ^{≪}Istituto nazionale di studi sul rinascimento. Studi e testi ^{≫}, XX par Mieczyslaw Markowski. Revue d’Histoire des Sciences, 46(2–3):306–307, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633834.

Mengal:1993:AOF

[3439] Paul Mengal. Analyses d’ouvrages: ^{≪}autour du cognitivisme ^{≫}, à propos des ouvrages suivants: Réflexions sur l’intelligence des machines : 25 ans de recherches, par Donald Michie; Neurophilosophie du rêve, ≪Savoir/sciences ^{≫}par Claude Debru; Historical foundations of cognitive science, ^{≪}Philosophical studies series ^{≫}, 46 par J.-C. Smith; Psychology of science: Contributions to metascience par Barry Gholson; William R. Shadish,; Robert A. Neimeyer; Arthur C. Houts. ^{≪}Manuels informatiques Masson ^{≫}. Revue d’Histoire des Sciences, 46(2–3):307–308, avril– septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633835.

Bitbol:1993:AOB

[3440] Michel Bitbol. Analyses d’ouvrages: Schrödinger, life and thought par Walter Moore. Revue d’Histoire des Sciences, 46(2–3):309–310, avril– septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633836.

Darrigol:1993:AOBc

[3441] Olivier Darrigol. Analyses d’ouvrages: Physics as a calling: Discipline and practice in the Königsberg Seminar for physics, ^{≪}Cornell History of Science Series ^{≫}par Kathryn M. Olesko. Revue d’Histoire des Sciences, 46(2–3):310–311, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633837.

Salanskis:1993:AOBa

[3442] Jean-Michel Salanskis. Analyses d’ouvrages: Mathématiques et existence: Ordres, fragments, empiètements.^{≪}Collection Milieux^{≫}par Daniel Parrochia; François Dagognet. Revue d’Histoire des Sciences, 46(2–3): 311–312, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633838.

Pellegrin:1993:AOB

[3443] Pierre Pellegrin. Analyses d’ouvrages: Die aristotelisch-scholastische Theorie der Bewegung. Studien zum Kommentar Alberts von Sachsen zur Physik des Aristoteles, ^{≪}Beiträge zur Geschichte der Philosophie und Theologie des Mittelalters ^{≫}, Neue Folge, Bd. 32 par Jürgen Sarnowsky. Revue d’Histoire des Sciences, 46(2–3):312–313, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633839.

Salanskis:1993:AOBb

[3444] Jean-Michel Salanskis. Analyses d’ouvrages: Rechnen mit dem Unendlichen. Beiträge zur Entwicklung eines kontroversen Gegenstandes par Detlef D. Spalt. Revue d’Histoire des Sciences, 46(2–3):313–314, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633840.

Anonymous:1993:PR

[3445] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 46(2–3): 315–327, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633841.

Fichant:1993:I

[3446] Michel Fichant. [introduction]. Revue d’Histoire des Sciences, 46(4):329– 331, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633295.

Fichant:1993:LLM

[3447] Michel Fichant. Leibniz lecteur de Mariotte. Revue d’Histoire des Sciences, 46(4):333–405 (1994), October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633296.

Alcantera:1993:CGL

[3448] Jean-Pascal Alcantéra and Gottfried Wilhelm Leibniz. La “Caractéristique géométrique” leibnizienne: travail du discernement et relations fondamentales. Revue d’Histoire des Sciences, 46(4):407–437 (1994), October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633297.

Parmentier:1993:CJP

[3449] Marc Parmentier. Concepts juridiques et probabilistes chez Leibniz. Revue d’Histoire des Sciences, 46(4):439–485, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633298.

Fichant:1993:BL

[3450] Michel Fichant. Bibliographie leibnizienne. Revue d’Histoire des Sciences, 46(4):487–492 (1994), October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633299.

Jullien:1993:EGL

[3451] Vincent Jullien. Les étendues géométriques et la ligne droite de Roberval. Revue d’Histoire des Sciences, 46(4):493–526, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633300.

Anonymous:1993:Ib

[3452] Anonyme. Informations. Revue d’Histoire des Sciences, 46(4):527–530, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633301.

Russo:1993:AOB

[3453] François Russo. Analyses d’ouvrages: La Naissance de la mécanique analytique : La science du mouvement au tournant des XVII^{e} et XVIII^{e} siècles, ^{≪}Bibliothèque d’histoire des sciences ^{≫} par Michel Blay. Revue d’Histoire des Sciences, 46(4):531, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633302.

Milanesi:1993:AOB

[3454] Claudio Milanesi. Analyses d’ouvrages: Dal Mito alla Storia. Biologia e medicina in Pasteur, ^{≪}Filosofia ^{≫} par Antonio Cadeddu. Revue d’Histoire des Sciences, 46(4):532–534, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633303.

Fischer:1993:AOBa

[3455] Jean-Louis Fischer. Analyses d’ouvrages: Correspondence of Charles Darwin (The), vol. 5 : 1851–1855 (1989), vol. 6 : 1856–1857 (1990), vol. 7 : 1858–1859 par Frederick Burkhardt; Sydney Smith; Charles Darwin’s Beagle Diary, 1831–1836 par Richard Darwin Keynes; Charles Darwin in Australia par F. W. Nicholas; J. M. Nicholas. Revue d’Histoire des Sciences, 46(4):535–537, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633304.

Grmek:1993:AOBc

[3456] Mirko D. Grmek. Analyses d’ouvrages: Dictionnaire des journaux, 1600– 1789 par Jean Sgard. Revue d’Histoire des Sciences, 46(4):538, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633305.

Fichant:1993:AOB

[3457] Michel Fichant. Analyses d’ouvrages: Leibniz et la méthode de la science, ^{≪}L’interrogation philosophique ^{≫} par François Duchesneau. Revue d’Histoire des Sciences, 46(4):538–539, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633306.

Beaulieu:1993:AOB

[3458] Armand Beaulieu. Analyses d’ouvrages: Lettres à Claude Saumaise et à son entourage (1620–1637) par Nicolas-Claude Fabri de Peireisc; Agnès Bresson. Revue d’Histoire des Sciences, 46(4):540, October 1993. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633307.

Beauvalet:1993:AOB

[3459] Scarlett Beauvalet. Analyses d’ouvrages: Patients, Power, and the Poor in Eighteenth Century Bristol, ^{≪}Cambridge History of Medicine ^{≫}par Mary E. Fissel. Revue d’Histoire des Sciences, 46(4):540–541, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633308.

Fischer:1993:AOBb

[3460] Jean-Louis Fischer. Analyses d’ouvrages: Histoire des connaissances zoologiques et ses rapports avec la zoologie, l’archéologie, la médecine vétérinaire, l’ethnologie (L’). Journée d’étude, univ. de Liège, 4 mars 1989. ^{≪}Colloques d’histoire des connaissances zoologiques ^{≫}, 1 par Liliane Bodson; Roland Libois. Revue d’Histoire des Sciences, 46(4):541– 542, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633309.

Debru:1993:AOB

[3461] Claude Debru. Analyses d’ouvrages: For the Love of Enzymes: The Odyssey of a Biochemist par Arthur Kornberg. Revue d’Histoire des Sciences, 46(4):542–543, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633310.

Gourevitch:1993:AOBd

[3462] Danielle Gourevitch. Analyses d’ouvrages: La Folie au Moyen Âge, XI^{e}– XIII^{e} siècles par Muriel Laharie; Jacques Le Goff. Revue d’Histoire des Sciences, 46(4):543–544, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633311.

Rey:1993:AOBb

[3463] Roselyne Rey. Analyses d’ouvrages: Artistes et mortels par Michel Lemire; B. Faye; E. A. Cabanis. Revue d’Histoire des Sciences, 46 (4):544–545, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633312.

Atten:1993:AOB

[3464] Michel Atten. Analyses d’ouvrages: Innovation in Maxwell’s electromagnetic theory. Molecular Vortices, Displacement Current, and Light par Daniel M. Siegel. Revue d’Histoire des Sciences, 46(4):545–547, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633313.

Anonymous:1993:BMa

[3465] Anonyme. Back matter. Revue d’Histoire des Sciences, 46(1):??, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633763.

Anonymous:1993:BMb

[3466] Anonyme. Back matter. Revue d’Histoire des Sciences, 46(2–3):??, avril– septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633842.

Anonymous:1993:BMc

[3467] Anonyme. Back matter. Revue d’Histoire des Sciences, 46(4):??, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633314.

Anonymous:1993:FMa

[3468] Anonyme. Front matter. Revue d’Histoire des Sciences, 46(1):??, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633738.

Anonymous:1993:FMb

[3469] Anonyme. Front matter. Revue d’Histoire des Sciences, 46(2–3):??, avril– septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633816.

Anonymous:1993:FMc

[3470] Anonyme. Front matter. Revue d’Histoire des Sciences, 46(4):??, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633294.

Grmek:1994:I

[3471] Mirko D. Grmek. [introduction]. Revue d’Histoire des Sciences, 47(1):3– 5, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633471.

Galperin:1994:VPC

[3472] Charles Galperin. Virus, provirus et cancer. Revue d’Histoire des Sciences, 47(1):7–56, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633472.

Gaudilliere:1994:CEI

[3473] Jean-Paul Gaudillière. Le cancer entre infection et hérédité: gènes, virus et souris au National Cancer Institute (1937–1977). Revue d’Histoire des Sciences, 47(1):57–89, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633473.

Brunet:1994:PHC

[3474] Guy Brunet. Populations et hérédité : à la croisée des pistes. Revue d’Histoire des Sciences, 47(1):91–106, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633474.

Monti:1994:CCI

[3475] Maria Teresa Monti. La chirurgie de la cataracte: Institutions, techniques et modèles scientifiques de Brisseau à Daviel. Revue d’Histoire des Sciences, 47(1):107–127, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633475.

Taylor:1994:NLG

[3476] Kenneth L. Taylor. New light on geological mapping in Auvergne during the eighteenth century: the Pasumot–Desmarest collaboration. Revue d’Histoire des Sciences, 47(1):129–136, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633476.

Braunstein:1994:YC

[3477] Jean-François Braunstein. Yvette Conry (1930–1992). Revue d’Histoire des Sciences, 47(1):137–139, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633477.

Theodorides:1994:GL

[3478] Jean Théodoridès. Georgette Legée (1914–1993). Revue d’Histoire des Sciences, 47(1):141, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633478.

Anonymous:1994:Ia

[3479] Anonyme. Informations. Revue d’Histoire des Sciences, 47(1):143–144, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633479.

Gourevitch:1994:AOBa

[3480] Danielle Gourevitch. Analyses d’ouvrages: Meanings of Sex Difference in the Middle Ages: Medicine, Science and Culture, ^{≪}Cambridge History of Medicine ^{≫}par Joan Cadden. Revue d’Histoire des Sciences, 47(1):145, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633480.

Morange:1994:AOBa

[3481] Michel Morange. Analyses d’ouvrages: The Laboratory revolution in medicine par Andrew Cunningham; Perry Williams. Revue d’Histoire des Sciences, 47(1):145–146, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633481.

Gourevitch:1994:AOBb

[3482] Danielle Gourevitch. Analyses d’ouvrages: Theophrastus of Eresus: Sources for his life, writings, thought and influence, Part one : Life, writings, various reports, logic, physics, Part two : Psychology, human physiology, living creatures, botany, ethics, religion, politics, rhetoric and poetics, music, miscellanea, ^{≪}Philosophia antiqua. A series of studies on ancient philosophy ^{≫}, vol. LIV, 1–2 par William W. Fortenbaugh. Revue d’Histoire des Sciences, 47(1):147, January 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633482.

Byl:1994:AOB

[3483] Simon Byl. Analyses d’ouvrages: Hippocrate par Jacques Jouanna. Revue d’Histoire des Sciences, 47(1):148–149, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633483.

Monnoyeur:1994:AOB

[3484] Françoise Monnoyeur. Analyses d’ouvrages: Philosophy, Science and Religion in England, 1640–1700 par Richard Kroll; Richard Ashcraft; Perez Zagorin. Revue d’Histoire des Sciences, 47(1):149–150, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633484.

Bensaude-Vincent:1994:AOB

[3485] Bernadette Bensaude-Vincent. Analyses d’ouvrages: Enlightenment Science in the Romantic Era: The chemistry of Berzelius and its cultural setting par Evan M. Melhado; Tore Frängsmyr. Revue d’Histoire des Sciences, 47(1):150–151, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633485.

Acot:1994:AOB

[3486] Pascal Acot. Analyses d’ouvrages: The State of Nature: Ecology, Community and American Social Thought, 1900–1950, ^{≪}Science and its conceptual foundations ^{≫}par Gregg Mitman. Revue d’Histoire des Sciences, 47(1):151–152, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633486.

Combe:1994:AOB

[3487] Jean-Marc Combe. Analyses d’ouvrages: André Chapelon, 1892–1978 : Le génie français de la vapeur par H. C. B. Rogers; Jacques Payen; Bernard Escudie. Revue d’Histoire des Sciences, 47(1):152–154, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633487.

Gourevitch:1994:AOBc

[3488] Danielle Gourevitch. Analyses d’ouvrages: Demokrit lachender Philosoph und sanguinischer Melancholiker : Eine pseudohippokratische Geschichte, ^{≪}Mnemosyne. Bibliotheca classica batava ^{≫}, Suppl., 118 par Thomas Rütten. Revue d’Histoire des Sciences, 47(1):154, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633488.

Blondel-Megrelis:1994:AOB

[3489] Marika Blondel-Megrelis. Analyses d’ouvrages: Eilhard Mitscherlich : Baumeister am Fundament der Chemie, ^{≪}Abhandlungen und Berichte ≫, Neue Folge, Bd. 8 par Hans Werner Schütt. Revue d’Histoire des Sciences, 47(1):155, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633489.

Milhaud-Cappe:1994:AOB

[3490] Danielle Milhaud-Cappe. Analyses d’ouvrages: Le Clinicien et le chercheur : Des grandes maladies de carence à la médecine moléculaire (1880–1980), ^{≪}Bibliothèque d’histoire des sciences^{≫}par Christiane Sinding. Revue d’Histoire des Sciences, 47(1):155–158, January 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633490.

Petroni:1994:AOB

[3491] Angelo M. Petroni. Analyses d’ouvrages: Sviluppi critici del pensiero scientifico moderno nella ricostruzione metodologica di Ernest Nagel, ≪Collana di studi epistemologici ^{≫}, 7 par Sandra Tugnoli Pattaro. Revue d’Histoire des Sciences, 47(1):158–159, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633491.

Caveing:1994:PGD

[3492] Maurice Caveing. La proportionnalité des grandeurs dans la doctrine de la nature d’Aristote. Revue d’Histoire des Sciences, 47(2):163–188, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633234.

Gardies:1994:LDL

[3493] Jean-Louis Gardies. L’organisation du livre XII des “Éléments” d’Euclide et ses anomalies. Revue d’Histoire des Sciences, 47(2):189–208, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633235.

Koelblen:1994:PCR

[3494] Sabine Koelblen. Une pratique de la composition des raisons dans un exercice de combinatoire. Revue d’Histoire des Sciences, 47(2):209–247, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633236.

Guillaumin:1994:SDC

[3495] Jean-Yves Guillaumin. La structure du chapitre 1,4 de l’Institution arithmétique de Boèce et le cours d’Ammonios sur Nicomaque. Revue d’Histoire des Sciences, 47(2):249–258, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633237.

Simon:1994:OTM

[3496] Gérard Simon. Aux origines de la théorie des miroirs: sur l’authenticité de la “Catoptrique” d’Euclide. Revue d’Histoire des Sciences, 47(2):259– 272, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633238.

Saito:1994:DPB

[3497] Ken Saïto. Debate: Proposition 14 of Book V of the Elements — a proposition that remained a local lemma. [Comment on: “Proposition 14 of Book V in the organization of Euclid’s Elements” [Rev. Histoire Sci. 44 (1991), no. 3–4, 457–467; MR1164428 (93g:01010)] by J.-L. Gardies]. Revue d’Histoire des Sciences, 47(2):273–282, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633239. See [3296] and response [3498].

Gardies:1994:RJL

[3498] Jean-Louis Gardies. Réponse de Jean-Louis Gardies. Revue d’Histoire des Sciences, 47(2):283–284, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633240. Voir [3497].

Caveing:1994:BVF

[3499] Maurice Caveing. Thèse. Bernard Vitrac, ^{≪}De quelques questions touchant au traitement de la proportionnalité dans “Les Éléments” d’Euclide ^{≫}. Revue d’Histoire des Sciences, 47(2):285–291, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633241.

Merleau-Ponty:1994:AOB

[3500] Jacques Merleau-Ponty. Analyses d’ouvrages: Ernst Mach — a deeper look: Documents and new perspectives, ^{≪}Boston Studies in the Philosophy of Science ^{≫}, vol. 143 par John Blackmore. Revue d’Histoire des Sciences, 47(2):293–295, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633242.

Chapront-Touze:1994:AOB

[3501] Michelle Chapront-Touzé. Analyses d’ouvrages: Models and Precision: The Quality of Ptolemy’s observations and parameters, ^{≪}Sources and Studies in the History and Philosophy of Classical Science ^{≫}, 1 par John Phillips Britton. Revue d’Histoire des Sciences, 47(2):295, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633243.

Thillaud:1994:AOB

[3502] Pierre L. Thillaud. Analyses d’ouvrages: Les Cellules folles : L’homme face au cancer de l’Antiquité à nos jours par Pierre Darmon. Revue d’Histoire des Sciences, 47(2):296–297, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633244.

Gourevitch:1994:AOBd

[3503] Danielle Gourevitch. Analyses d’ouvrages: Galenus Latinus II : Burgundio of Pisa’s translation of Galen’s Peri ton peponthoton topon ^{≪}de interioribus ^{≫}, ^{≪}Ars Medica, Abt. II : Griechisch-Lateinische Medizin ^{≫}, Bd. 6/2A und 6/2B par Richard J. Durling; R. J. Durling. Revue d’Histoire des Sciences, 47(2):297, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633245.

Caveing:1994:AOB

[3504] Maurice Caveing. Analyses d’ouvrages: Mathématiques dans l’Antiquité. ≪Mémoires du Centre Jean-Palerne ^{≫}, XI par Jean-Yves Guillaumin; J.- Y. Guillaumin. Revue d’Histoire des Sciences, 47(2):297–300, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633246.

Gourevitch:1994:AOBe

[3505] Danielle Gourevitch. Analyses d’ouvrages: Augenheilkunde im Rom der frühen Kaiserzeit, A. C. Celsus De medicina libri octo (Heidelberg : Kaden Verlag, 1992) par Ludwig Limmer; Günter K. Krieglstein. Revue d’Histoire des Sciences, 47(2):300–301, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633247.

Blanckaert:1994:AOB

[3506] Claude Blanckaert. Analyses d’ouvrages: History, humanity and evolution. Essays for John C. Greene par James R. Moore. Revue d’Histoire des Sciences, 47(2):301–304, avril–juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633248.

Scheidecker-Chevallier:1994:TMC

[3507] Myriam Scheidecker-Chevallier and Robert Locqueneux. La théorie mathématique de la combinaison chimique d’André-Marie Ampère. Revue d’Histoire des Sciences, 47(3–4):309–352, juillet–décembre 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633495.

Goudaroulis:1994:SND

[3508] Yorgos Goudaroulis. Searching for a name: the development of the concept of the critical point (1822–1869). Revue d’Histoire des Sciences, 47(3–4):353–379, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633496.

Medard:1994:LSP

[3509] Louis Médard. L’œuvre scientifique de Paul Vieille (1854–1934). Revue d’Histoire des Sciences, 47(3–4):381–404, juillet–décembre 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633497.

Roque:1994:CCN

[3510] Georges Roque. Les couleurs complémentaires : un nouveau paradigme. Revue d’Histoire des Sciences, 47(3–4):405–433, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633498.

Vilain:1994:PMD

[3511] Christiane Vilain. La proportionnalité de la masse et du poids dans la dynamique newtonienne. Revue d’Histoire des Sciences, 47(3–4): 435–473, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633499.

LeRu:1994:FAE

[3512] Véronique Le Ru. La force accélératrice: un exemple de définition contextuelle dans le Traité de dynamique de d’Alembert. Revue d’Histoire des Sciences, 47(3–4):475–494, juillet–décembre 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633500.

Anonymous:1994:Ib

[3513] Anonyme. Informations. Revue d’Histoire des Sciences, 47(3–4):495– 496, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633501.

Dubois:1994:AOB

[3514] Jacques Dubois. Analyses d’ouvrages: Newton mécanicien du cosmos, ≪Mathesis ^{≫}par Georges Barthélemy. Revue d’Histoire des Sciences, 47(3–4):497–498, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633502.

Duris:1994:AOB

[3515] Pascal Duris. Analyses d’ouvrages: Buffon 88. ^{≪}Science, histoire, philosophie ^{≫} par Jean-Claude Beaune; Serge Benoit; Jean Gayon; Jacques Roger; Denis Woronoff; Ernst Mayr; Georges Canguilhem. Revue d’Histoire des Sciences, 47(3–4):498–499, juillet–décembre 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633503.

Beaujouan:1994:AOB

[3516] Guy Beaujouan. Analyses d’ouvrages: Jordanus de Nemore, De elementis arithmetice artis: A medieval treatise on number theory. Part I: Text and paraphrase; Part II: Conspectus siglorum and critical apparatus, ^{≪}Boethius ^{≫}, Bd. XII, I und II par H. L. L. Busard. Revue d’Histoire des Sciences, 47(3–4):499–501, juillet–décembre 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633504.

Chevalley:1994:AOB

[3517] Catherine Chevalley and Olivier Darrigol. Analyses d’ouvrages: Uncertainty: The Life and Science of Werner Heisenberg par David C. Cassidy. Revue d’Histoire des Sciences, 47(3–4):501–503, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633505.

Mengal:1994:AOB

[3518] Paul Mengal. Analyses d’ouvrages: Cerveau et l’esprit (Le) par Hervé Barreau. Revue d’Histoire des Sciences, 47(3–4):503, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633506.

Bailhache:1994:AOB

[3519] Patrice Bailhache. Analyses d’ouvrages: Music and Science in the Age of Galileo ^{≪}The Univ. of Western Ontario Series in Philosophy of Science ^{≫}, vol. 51 par Victor Coelho. Revue d’Histoire des Sciences, 47 (3–4):504, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633507.

Dhombres:1994:AOB

[3520] Jean Dhombres. Analyses d’ouvrages: Science under Control: The French Academy of Sciences, 1795–1914 par Maurice Crosland. Revue d’Histoire des Sciences, 47(3–4):504–509, juillet–décembre 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633508.

Glasner:1994:AOB

[3521] Ruth Glasner. Analyses d’ouvrages: Studies on Gersonides: A fourteenth-century Jewish philosopher-scientist, ^{≪}Coll. de travaux de l’Académie internationale d’histoire des sciences ^{≫}, t. 36 par Gad Freudenthal. Revue d’Histoire des Sciences, 47(3–4):509–511, juillet– décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633509.

Clavelin:1994:AOB

[3522] Maurice Clavelin. Analyses d’ouvrages: Sidereus Nuncius/Le Messager céleste. ^{≪}Science et humanisme ^{≫} par Galileo Galilei; Isabelle Pantin. Revue d’Histoire des Sciences, 47(3–4):512, juillet–décembre 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633510.

Jullien:1994:AOBa

[3523] Vincent Jullien. Analyses d’ouvrages: Geometria e atomismo nella scuola galileana, ^{≪}Biblioteca di Nuncius, Studi e testi ^{≫}, X par M. Bucciantini; M. Torrini. Revue d’Histoire des Sciences, 47(3–4):512– 513, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633511.

Morange:1994:AOBb

[3524] Michel Morange. Analyses d’ouvrages: Styles of Scientific Thought: The German Genetics Community, 1900–1933, ^{≪}Science and its Conceptual Foundations Series ^{≫}par Jonathan Harwood. Revue d’Histoire des Sciences, 47(3–4):513–514, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633512.

Jullien:1994:AOBb

[3525] Vincent Jullien. Analyses d’ouvrages: Histoire de fractions, fractions d’histoire, ^{≪}Science Networks. Historical Studies ^{≫}, vol. 10 par Paul Benoit; Karine Chemla; Jim Ritter. Revue d’Histoire des Sciences, 47(3– 4):514–515, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633513.

Gourevitch:1994:AOBf

[3526] Danielle Gourevitch. Analyses d’ouvrages: The Cambridge World History of Human Disease par Kenneth F. Kiple. Revue d’Histoire des Sciences, 47(3–4):516–517, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633514.

Petitjean:1994:AOB

[3527] Patrick Petitjean. Analyses d’ouvrages: Colonialism and Science: Saint-Domingue in the Old Regime par James E. McClellan. Revue d’Histoire des Sciences, 47(3–4):517–518, juillet–décembre 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633515.

Debru:1994:AOB

[3528] Claude Debru. Analyses d’ouvrages: Classification, evolution and the nature of biology par Alec L. Panchen. Revue d’Histoire des Sciences, 47(3– 4):518–519, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633516.

Mondot:1994:AOB

[3529] Jean Mondot. Analyses d’ouvrages: Johann Gottfried Herder: La raison dans les peuples, ^{≪}Bibliothèque franco-allemande ^{≫}par Pierre Pénisson. Revue d’Histoire des Sciences, 47(3–4):519, juillet–décembre 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633517.

Weiner:1994:AOB

[3530] Dora B. Weiner. Analyses d’ouvrages: Penser la folie : Essais sur Michel Foucault, ^{≪}Débats ^{≫}. [Publ. de communications présentées lors du IX^{e} Colloque de la Soc. d’hist. de la psychiatrie et de la psychanalyse qui avait pour thème ^{≪}Histoire de la folie trente ans après ^{≫}.]. Revue d’Histoire des Sciences, 47(3–4):520–521, juillet–décembre 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633518.

Gourevitch:1994:AOBg

[3531] Danielle Gourevitch. Analyses d’ouvrages: Quid pro quo: Studies in the history of drugs par John M. Riddle. Revue d’Histoire des Sciences, 47 (3–4):521, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633519.

Grmek:1994:AOBa

[3532] Mirko D. Grmek. Analyses d’ouvrages: Explaining epidemics and other studies in the history of medicine par Charles E. Rosenberg. Revue d’Histoire des Sciences, 47(3–4):521–522, juillet–décembre 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633520.

Bonneuil:1994:AOB

[3533] Christophe Bonneuil. Analyses d’ouvrages: Science and Empires: Historical Studies about Scientific Development and European Expansion, ≪Boston Studies in the Philosophy of Science ^{≫}, vol. 136 par Patrick Petitjean; Catherine Jami; Anne-Marie Moulin. Revue d’Histoire des Sciences, 47(3–4):522–523, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633521.

Grmek:1994:AOBb

[3534] Mirko D. Grmek. Analyses d’ouvrages: Animals and Disease: An introduction to the history of comparative medicine par Lise Wilkinson. Revue d’Histoire des Sciences, 47(3–4):523–524, juillet–décembre 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633522.

Malherbe:1994:AOB

[3535] Michel Malherbe. Analyses d’ouvrages: Francis Bacon: History, Politics and Science, 1561–1626 par B. H. G. Wormald. Revue d’Histoire des Sciences, 47(3–4):524–525, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633523.

Elie:1994:AOB

[3536] Maurice Elie. Analyses d’ouvrages: Zur Edition naturwissenschaftlicher Texte der Goethezeit. Leopoldina-Meeting vom 22. bis 23. Mai 1992 in Halle (Saale). Acta Historica Leopoldina, Nr. 20 par Wolf von Engelhardt; Ilse Jahn; Dorothea Kuhn. Revue d’Histoire des Sciences, 47(3– 4):525–526, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633524.

Anonymous:1994:PR

[3537] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 47(3–4): 527–538, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633525.

Anonymous:1994:TGM

[3538] Anonyme. Table générale des matières du Tome XLVII. Revue d’Histoire des Sciences, 47(3–4):539–542, juillet–décembre 1994. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633526.

Anonymous:1994:BMa

[3539] Anonyme. Back matter. Revue d’Histoire des Sciences, 47(1):??, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633492.

Anonymous:1994:BMb

[3540] Anonyme. Back matter. Revue d’Histoire des Sciences, 47(2):??, avril– juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633249.

Anonymous:1994:BMc

[3541] Anonyme. Back matter. Revue d’Histoire des Sciences, 47(3–4): ??, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633527.

Anonymous:1994:E

[3542] Anonyme. Erratum. Revue d’Histoire des Sciences, 47(3–4):??, juillet– décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633494.

Anonymous:1994:FMa

[3543] Anonyme. Front matter. Revue d’Histoire des Sciences, 47(1):??, January 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633470.

Anonymous:1994:FMb

[3544] Anonyme. Front matter. Revue d’Histoire des Sciences, 47(2):??, avril– juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633232.

Anonymous:1994:FMc

[3545] Anonyme. Front matter. Revue d’Histoire des Sciences, 47(3–4): ??, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633493.

Anonymous:1994:P

[3546] Anonyme. Présentation. Revue d’Histoire des Sciences, 47(2):??, avril– juin 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633233.

Bret:1995:DCA

[3547] Patrice Bret. Débats et chantiers actuels autour de Lavoisier et de la révolution chimique. introduction. Revue d’Histoire des Sciences, 48(1– 2):3–8, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633791.

Holmes:1995:BLC

[3548] Frederic L. Holmes. The boundaries of Lavoisier’s chemical revolution. Revue d’Histoire des Sciences, 48(1–2):9–48, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633792.

Bensaude-Vincent:1995:EME

[3549] Bernadette Bensaude-Vincent. Eaux et mesures. Éclairages sur l’itinéraire intellectuel du jeune Lavoisier. Revue d’Histoire des Sciences, 48(1–2):49–69, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633793.

Beretta:1995:LRC

[3550] Marco Beretta. Lavoisier as a reader of chemical literature. Revue d’Histoire des Sciences, 48(1–2):71–94, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633794.

Mauskopf:1995:LIG

[3551] Seymour H. Mauskopf. Lavoisier and the improvement of gunpowder production. Revue d’Histoire des Sciences, 48(1–2):95–121, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633795.

Aceves:1995:LAN

[3552] Patricia Aceves. L’enseignement et l’application de la nouvelle chimie au Mexique au temps de Lavoisier. Revue d’Histoire des Sciences, 48 (1–2):123–132, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633796.

Kritsman:1995:SLW

[3553] Victor A. Kritsman and Brigitte Hoppe. The study of Lavoisier’s works par Russian scientists. Revue d’Histoire des Sciences, 48(1–2):133–142, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633797.

Fauque:1995:LDC

[3554] Danielle Fauque. Lavoisier, deux cents ans après: À propos des ouvrages du bicentenaire. Revue d’Histoire des Sciences, 48(1–2):143–168, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633798.

Bret:1995:TDE

[3555] Patrice Bret. Trois décennies d’études lavoisiennes: Supplément aux bibliographies de Duveen. Revue d’Histoire des Sciences, 48(1–2):169– 205, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633799.

Taton:1995:MG

[3556] René Taton. Michelle Goupil (1934–1993). Revue d’Histoire des Sciences, 48(1–2):207–218, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633800.

Eisenstaedt:1995:AOB

[3557] Jean Eisenstaedt. Analyses d’ouvrages: L’Univers sous le regard du temps : La cosmologie théorique moderne et ses racines, coll. ^{≪}de caelo ≫par Henri Andrillat; J. Merleau-Ponty. Revue d’Histoire des Sciences, 48(1–2):219, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633801.

Cotte:1995:AOB

[3558] Michel Cotte. Analyses d’ouvrages: Travaux publics de France : Un siècle d’entrepreneurs et d’entreprises (1883–1992) par Dominique Barjot. Revue d’Histoire des Sciences, 48(1–2):220, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633802.

Langins:1995:AOBa

[3559] Janis Langins. Analyses d’ouvrages: Histoire de la chimie, ^{≪}Histoire des sciences ^{≫}par Bernadette Bensaude-Vincent; Isabelle Stengers. Revue d’Histoire des Sciences, 48(1–2):220–221, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633803.

Langins:1995:AOBb

[3560] Janis Langins. Analyses d’ouvrages: The Enlightenment of Matter : The Definition of Chemistry from Agricola to Lavoisier par Marco Beretta. Revue d’Histoire des Sciences, 48(1–2):221–222, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633804.

Jullien:1995:AOB

[3561] Vincent Jullien. Analyses d’ouvrages: Les Raisons de l’infini : Du monde clos à l’univers mathématique, ^{≪}NRF Essais ^{≫}par Michel Blay. Revue d’Histoire des Sciences, 48(1–2):222–223, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633805.

Beaulieu:1995:AOB

[3562] Armand Beaulieu. Analyses d’ouvrages: Nel Segno di Galileo : Erudizione, filosofia e scienza a Firenze nel secolo XVII. I trattati accademici di Vincenzio Capponi par Stefano Caroti; G. Alessandri; M. L. Migliore. Revue d’Histoire des Sciences, 48(1–2):223–224, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633806.

Acot:1995:AOB

[3563] Pascal Acot. Analyses d’ouvrages: Linné et la France (1780–1850), ≪Histoire des idées et critique littéraire ^{≫}, vol. 318 par Pascal Duris. Revue d’Histoire des Sciences, 48(1–2):224–225, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633807.

Gourevitch:1995:AOBa

[3564] Danielle Gourevitch. Analyses d’ouvrages: Restless youth in ancient Rome par Emiel Eyben. Revue d’Histoire des Sciences, 48(1–2):225, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633808.

Nicolaidis:1995:AOBa

[3565] Efthymios Nicolaïdis. Analyses d’ouvrages: Science in Russia and the Soviet Union : A Short History, ^{≪}Cambridge History of Science ^{≫}par Loren R. Graham. Revue d’Histoire des Sciences, 48(1–2):225–226, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633809.

Clavelin:1995:AOB

[3566] Maurice Clavelin. Analyses d’ouvrages: Dissertatio cum Nuncio Sidereo (Discussion avec le messager céleste). Narratio de Observatis jovis satellitibus (Rapport sur l’observation des satellites de Jupiter). ^{≪}Science et humanisme ^{≫}par Johannes Kepler; Isabelle Pantin. Revue d’Histoire des Sciences, 48(1–2):226–227, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633810.

Duchesne:1995:AOB

[3567] Raymond Duchesne. Analyses d’ouvrages: Science and the Canadian Arctic : A century of exploration, 1818–1918 par Trevor H. Levere. Revue d’Histoire des Sciences, 48(1–2):227–228, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633811.

Pepe:1995:AOB

[3568] Luigi Pepe. Analyses d’ouvrages: Anton Maria Lorgna e l’Europa par Franco Piva. Revue d’Histoire des Sciences, 48(1–2):228, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633812.

Gohau:1995:AOBa

[3569] Gabriel Gohau. Analyses d’ouvrages: Les Scientifiques et la montagne. Actes du 116^{e} Congrès national des sociétés savantes (Chambéry, 29 avril–4 mai 1991), section d’histoire des sciences et des techniques ; Sciences naturelles et montagne. Actes du 116^{e} Congrès national des sociétés savantes. Revue d’Histoire des Sciences, 48(1–2):228–229, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633813.

Nicolaidis:1995:AOBb

[3570] Efthymios Nicolaïdis. Analyses d’ouvrages: Alexander A. Friedmann : The man who made the universe expand par Eduard A. Tropp; Viktor Ya. Frenkel; Arthur D. Chernin; Alexander Dron; Michael Burov. Revue d’Histoire des Sciences, 48(1–2):230, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633814.

Fischer:1995:RRH

[3571] Jean-Louis Fischer. Roselyne Rey (1951–1995), historienne des sciences du Siècle des lumières. Revue d’Histoire des Sciences, 48(3): 233–239, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633203.

Baratay:1995:ZEC

[3572] Éric Baratay. Zoologie et Église catholique dans la France du XVIII^{e} siècle (1670–1840) : une science au service de Dieu. Revue d’Histoire des Sciences, 48(3):241–265, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633204.

Cherni:1995:HBP

[3573] Amor Cherni. Haller et Buffon : à propos des “Réflexions”. Revue d’Histoire des Sciences, 48(3):267–305, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633205.

Lopez-Beltran:1995:FMH

[3574] Carlos López-Beltrán. ^{≪}Les maladies héréditaires ^{≫}: 18th century disputes in France. Revue d’Histoire des Sciences, 48(3):307–350, juillet– septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633206.

Gourevitch:1995:RR

[3575] Danielle Gourevitch, Anne-Marie Chouillet, and Anne Fagot-Largeault. Roselyne Rey (1951–1995). Revue d’Histoire des Sciences, 48(3):351– 363, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633207.

Anonymous:1995:Ia

[3576] Anonyme. Informations. Revue d’Histoire des Sciences, 48(3):365– 368, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633208.

Poirier:1995:AOB

[3577] Jacques Poirier. Analyses d’ouvrages: Atlas de la Révolution française, vol. 7 : Médecine et santé par Jean-Pierre Goubert; Roselyne Rey; Jacques Bertrand; Alexandra Laclau. Revue d’Histoire des Sciences, 48 (3):369–370, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633209.

Ben-Dov:1995:AOB

[3578] Yoav Ben-Dov. Analyses d’ouvrages: Erwin Schrödinger : Philosophy and the birth of quantum mechanics. [Recueil de 24 communications présentées au Colloque ^{≪}Erwin Schrödinger, philosophie et naissance de la mécanique quantique ^{≫}] par Michel Bitbol; Olivier Darrigol. Revue d’Histoire des Sciences, 48(3):370–371, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633210.

Bonah:1995:AOB

[3579] Christian Bonah. Analyses d’ouvrages: Preserve your love for science : Life of William A. Hammond, American neurologist, ^{≪}Cambridge history of medicine ^{≫}par Bonnie Ellen Blustein. Revue d’Histoire des Sciences, 48(3):372, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633211.

Lebedeva:1995:AOB

[3580] Natalia V. Lebedeva. Analyses d’ouvrages: The Laser in America, 1950– 1970 par Joan Lisa Bromberg. Revue d’Histoire des Sciences, 48(3): 372–374, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633212.

Darrigol:1995:AOB

[3581] Olivier Darrigol. Analyses d’ouvrages: Hermann von Helmholtz and the foundations of nineteenth-century science, ^{≪}California studies in the history of science ^{≫}par David Cahan. Revue d’Histoire des Sciences, 48(3): 374–375, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633213.

Loty:1995:AOB

[3582] Laurent Loty. Analyses d’ouvrages: Des sciences contre l’homme, vol. I: Classer, hiérarchiser, exclure, vol. II: Au nom du Bien, série ^{≪}Sciences en société ^{≫}, n^{os} 8 et 9 par Claude Blanckaert. Revue d’Histoire des Sciences, 48(3):375–377, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633214.

Panza:1995:AOB

[3583] Marco Panza. Analyses d’ouvrages: Archimede : Mito tradizione scienza, ≪Biblioteca di Nuncius. Studi e testi ^{≫}, 4 par Corrado Dollo. Revue d’Histoire des Sciences, 48(3):377–379, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633215.

Gourevitch:1995:AOBb

[3584] Danielle Gourevitch. Analyses d’ouvrages: A dictionary of medical terms in Galen, ^{≪}Studies in ancient medicine ^{≫}, vol. 5 par Richard J. Durling. Revue d’Histoire des Sciences, 48(3):379, juillet–septembre 1995. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633216.

Gohau:1995:AOBb

[3585] Gabriel Gohau. Analyses d’ouvrages: Histoire de la géologie, t. 2 : La Grande Eclosion et ses prémices, 1660–1810,^{≪}Petite collection d’histoire des sciences ^{≫}par François Ellenberger. Revue d’Histoire des Sciences, 48(3):379–380, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633217.

Jacquart:1995:AOBa

[3586] Danielle Jacquart. Analyses d’ouvrages: The Naming of the beasts : Natural history in the medieval bestiary par Wilma George; Brunsdon Yapp. Revue d’Histoire des Sciences, 48(3):380–381, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633218.

Jacquart:1995:AOBb

[3587] Danielle Jacquart. Analyses d’ouvrages: Healing and society in medieval England : A middle English translation of the pharmaceutical writings of Gilbertus Anglicus, ^{≪}Wisconsin publications in the history of science and medicine ^{≫}, no. 8 par Faye Marie Getz. Revue d’Histoire des Sciences, 48(3):381–382, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633219.

Sinding:1995:AOB

[3588] Christiane Sinding. Analyses d’ouvrages: Between biology and medicine : The formation of intermediary metabolism par Frederic Lawrence Holmes. Revue d’Histoire des Sciences, 48(3):382–383, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633220.

Gourevitch:1995:AOBc

[3589] Danielle Gourevitch. Analyses d’ouvrages: The Healing Past : Pharmaceuticals in the biblical and rabbinic world, ^{≪}Studies in ancient medicine ≫, vol. 7 par Irene Jacob; Walter Jacob. Revue d’Histoire des Sciences, 48(3):383–384, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633221.

Galperin:1995:AOB

[3590] C. Galperin. Analyses d’ouvrages: The Molecular Vision of life : Caltech, the Rockfeller Foundation, and the rise of the new biology, ^{≪}Monographs on the history and philosophy of biology ^{≫}par Lily E. Kay. Revue d’Histoire des Sciences, 48(3):384–385, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633222.

Barras:1995:AOP

[3591] Vincent Barras. Analyses d’ouvrages: Peter Keating, La Science du mal: L’institution de la psychiatrie au Québec, 1800–1914 (Québec: Boréal, 1993). Jean Garrabé (dir.), Philippe Pinel (Le Plessis-Robinson: Synthé labo, 1994). Revue d’Histoire des Sciences, 48(3):386–387, juillet– septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633223.

Castrillon:1995:AOB

[3592] Alberto Castrillon. Analyses d’ouvrages: Mundialización de la ciencia y cultura nacional. Actas del Congreso internacional ^{≪}Ciencia descubrimiento y mundo colonial ^{≫} par A. Lafuente; A. Elena; M. L. Ortega. Revue d’Histoire des Sciences, 48(3):387–388, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633224.

Grmek:1995:AOB

[3593] Mirko D. Grmek. Analyses d’ouvrages: Emerging Viruses par Stephen S. Morse. Revue d’Histoire des Sciences, 48(3):388–389, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633225.

Debru:1995:AOB

[3594] Claude Debru. Analyses d’ouvrages: Psychologie et cerveau, Symposium de l’Association de psychologie scientifique de langue française (Toulouse, 1987), ^{≪}Psychologie d’aujourd’hui ^{≫} par Xavier Seron. Revue d’Histoire des Sciences, 48(3):389–390, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633226.

Poulle:1995:AOBa

[3595] Emmanuel Poulle. Analyses d’ouvrages: Optique et mathématiques : Recherches sur l’histoire de la pensée scientifique en arabe, ^{≪}Collected studies series ^{≫}par Roshdi Rashed. Revue d’Histoire des Sciences, 48 (3):391, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633227.

Weiner:1995:AOB

[3596] Dora B. Weiner. Analyses d’ouvrages: Journal d’observations sur les principaux hôpitaux et sur quelques prisons d’Angleterre (1787), ^{≪}Publ. de la fac. des lettres de Clermont, univ. Blaise-Pascal ^{≫}, nouv. sér., fasc. 37 par Jacques Tenon. Revue d’Histoire des Sciences, 48(3): 391–393, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633228.

Moulin:1995:AOB

[3597] Anne-Marie Moulin. Analyses d’ouvrages: Medicine in society : Historical essays par Andrew Wear. Revue d’Histoire des Sciences, 48(3): 393–394, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633229.

DeGandt:1995:AOB

[3598] François De Gandt. Analyses d’ouvrages: The Life of Isaac Newton par Richard S. Westfall. Revue d’Histoire des Sciences, 48(3):

394–395, juillet–septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633230.

Kragh:1995:AMW

[3599] Helge Kragh. Arthur March, Werner Heisenberg, and the search for a smallest length. Revue d’Histoire des Sciences, 48(4):401–434, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633252.

Panza:1995:NEF

[3600] Marco Panza. De la nature épargnante aux forces généreuses : le principe de moindre action entre mathématiques et métaphysique. Maupertuis et Euler, 1740–1751. Revue d’Histoire des Sciences, 48(4):435–520, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633253.

Petit:1995:LDP

[3601] Annie Petit. L’héritage du positivisme dans la création de la chaire d’histoire générale des sciences au Collège de France. Revue d’Histoire des Sciences, 48(4):521–556, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633254.

Anonymous:1995:Ib

[3602] Anonyme. Informations. Revue d’Histoire des Sciences, 48(4):557, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633255.

Poulle:1995:AOBb

[3603] Emmanuel Poulle. Analyses d’ouvrages: Making instruments count : Essays on historical instruments presented to Gerard L’Estrange Turner par R. G. W. Anderson; J. A. Bennett; W. F. Ryan. Revue d’Histoire des Sciences, 48(4):559–560, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633256.

Bernhardt:1995:AOB

[3604] Jean Bernhardt. Analyses d’ouvrages: Claude Ptolémée, astronome, astrologue, géographe : Connaissance et représentation du monde habité par Germaine Aujac; Ptolemy’s universe : The natural philosophical and ethical foundations of Ptolemy’s astronomy par Liba Chaia Taub. Revue d’Histoire des Sciences, 48(4):560–562, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633257.

Locqueneux:1995:AOB

[3605] Robert Locqueneux. Analyses d’ouvrages: Robert Mayer and the conservation of energy par Kenneth L. Caneva. Revue d’Histoire des Sciences, 48(4):562–563, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633258.

Morange:1995:AOB

[3606] Michel Morange. Analyses d’ouvrages: Croisées biologiques : Systémique et analytique, écologie et biologie moléculaire en dialogue. ^{≪}Catalyses ≫par Bernard Feltz; Jean Ladrière. Revue d’Histoire des Sciences, 48(4):563–564, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633259.

Nicolaidis:1995:AOBc

[3607] Efthymios Nicolaïdis. Analyses d’ouvrages: Théodore Méliténiote, Tribiblos astronomique, Livre II, 2 vol., I : Edition critique et traduction, II : Commentaire, ^{≪}Corpus des astronomes byzantins ^{≫}, V et VI par Régine Leurquin. Revue d’Histoire des Sciences, 48(4):564–565, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633260.

Labarriere:1995:AOB

[3608] Jean-Louis Labarrière. Analyses d’ouvrages: Pour en finir avec les mentalités. ^{≪}Textes à l’appui ^{≫}, sér. histoire classique par G. E. R. Lloyd; Franz Regnot; Methods and problems in Greek science par G. E. R. Lloyd. Revue d’Histoire des Sciences, 48(4):566–567, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633261.

Sinaceur:1995:AOB

[3609] Hourya Sinaceur. Analyses d’ouvrages: Logique et mathématique chez Bernard Bolzano, ^{≪}L’histoire des sciences. Textes et études ^{≫}par Jan Sebestik. Revue d’Histoire des Sciences, 48(4):567–569, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633262.

Blanchard:1995:AOB

[3610] Alain Blanchard. Analyses d’ouvrages: Edwin Hubble, the discoverer of the big bang universe par Alexander S. Sharov; Igor D. Nivikov; Vitaly Kisin. Revue d’Histoire des Sciences, 48(4):569–570, October 1995. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633263.

Gourevitch:1995:AOBd

[3611] Danielle Gourevitch. Analyses d’ouvrages: Theophrastus, Metaphysics. ≪Mnémosyne, bibliotheca classica batava ^{≫}, suppl., 125 par Théophraste; Marlein Van Raalte. Revue d’Histoire des Sciences, 48(4):570, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633264.

Nicolaidis:1995:AOBd

[3612] Efthymios Nicolaïdis. Analyses d’ouvrages: Astronomie et astrophysique, ≪Textes essentiels ^{≫}par Jean-Pierre Verdet. Revue d’Histoire des Sciences, 48(4):570–571, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633265.

Anonymous:1995:PR

[3613] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 48(4):573– 587, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633266.

Anonymous:1995:TGM

[3614] Anonyme. Table générale des matières du Tome XLVIII. Revue d’Histoire des Sciences, 48(4):589–591, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633267.

Anonymous:1995:BMa

[3615] Anonyme. Back matter. Revue d’Histoire des Sciences, 48(1–2):??, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633815.

Anonymous:1995:BMb

[3616] Anonyme. Back matter. Revue d’Histoire des Sciences, 48(3):??, juillet– septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633231.

Anonymous:1995:BMc

[3617] Anonyme. Back matter. Revue d’Histoire des Sciences, 48(4):??, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633268.

Anonymous:1995:FMa

[3618] Anonyme. Front matter. Revue d’Histoire des Sciences, 48(1–2):??, January 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633790.

Anonymous:1995:FMb

[3619] Anonyme. Front matter. Revue d’Histoire des Sciences, 48(3):??, juillet– septembre 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633202.

Anonymous:1995:FMc

[3620] Anonyme. Front matter. Revue d’Histoire des Sciences, 48(4):??, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633251.

Hartmann:1996:RVD

[3621] Claude Hartmann. La révolution verte du Siècle des lumières : les deux sociétés savantes orléanaises de la fin de l’ancien Régime. Revue d’Histoire des Sciences, 49(1):5–21, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633529.

Duris:1996:LHN

[3622] Pascal Duris. L’enseignement de l’histoire naturelle dans les écoles centrales (1795–1802). Revue d’Histoire des Sciences, 49(1):23–52, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633530.

Schnitter:1996:DDM

[3623] Claude Schnitter. Le développement du Muséum national d’histoire naturelle de Paris au cours de la seconde moitié du XIX^{e} siècle; ^{≪}se transformer ou périr ^{≫}. Revue d’Histoire des Sciences, 49(1):53–97, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633531.

Matagne:1996:LFA

[3624] Patrick Matagne. L’écologie en France au XIX^{e} siècle : résistances et singularités. Revue d’Histoire des Sciences, 49(1):99–111, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633532.

Lamontagne:1996:LMS

[3625] Roland Lamontagne. L’influence de Maurepas sur les sciences : le botaniste Jean Prat à La Nouvelle–Orléans, 1735–1746. Revue d’Histoire des Sciences, 49(1):113–124, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633533.

Nicoud:1996:AOB

[3626] Marilyn Nicoud. Analyses d’ouvrages: Les Consilia médicaux. ≪Typologie des sources du Moyen Âge occidental ^{≫}, A-V.D.3, fasc. 69 par Jole Agrimi; Chiara Crisciani; Caroline Viola. Revue d’Histoire des Sciences, 49(1):127, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633534.

Montiel:1996:AOB

[3627] Luis Montiel. Analyses d’ouvrages: El Mito de Vesalio, ^{≪}Lecció oberta ≫par Juan José Barcia Goyanes. Revue d’Histoire des Sciences, 49 (1):127–128, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633535.

Dhombres:1996:AOBa

[3628] Jean Dhombres. Analyses d’ouvrages: Mersenne, le grand minime par Armand Beaulieu. Revue d’Histoire des Sciences, 49(1):128–130, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633536.

Lowy:1996:AOB

[3629] Ilana Löwy. Analyses d’ouvrages: The Uses of life: A history of biotechnology, [1^{re} éd., 1993.] par Robert Bud. Revue d’Histoire des Sciences, 49(1):130–131, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633537.

Grmek:1996:AOBa

[3630] Mirko D. Grmek. Analyses d’ouvrages: Science and the practice of medicine in the nineteenth century, ^{≪}Cambridge history of science ^{≫}par William F. Bynum. Revue d’Histoire des Sciences, 49(1):131–132, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633538.

Baertschi:1996:AOB

[3631] Bernard Baertschi. Analyses d’ouvrages: Charles Bonnet savant et philosophe (1720–1793). Actes du Colloque international de Genève (25– 27 novembre 1993), ^{≪}Mémoires de la Société de physique et d’histoire naturelle de Genève ^{≫}, vol. 47 par Marino Buscaglia; René Sigrist; Jacques Trembley; Jean Wüest. Revue d’Histoire des Sciences, 49(1):132–133, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633539.

Becquemont:1996:AOBa

[3632] Daniel Becquemont. Analyses d’ouvrages: Darwinism, war and history : The debate over the biology of war from the ^{≪}Origin of species ≫to the First World War par Paul Crook. Revue d’Histoire des Sciences, 49(1):133–134, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633540.

Sinding:1996:AOB

[3633] Christiane Sinding. Analyses d’ouvrages: Les Sciences biologiques et médicales en France, 1920–1950. Actes du Colloque de Dijon, 25–27 juin 1992, (Centre Gaston Bachelard de recherche sur l’imaginaire et la rationalité, univ. de Bourgogne et Institut d’histoire et de philosophie des sciences et des techniques, CNRS, Paris), ^{≪}Cahiers pour l’histoire de la recherche ^{≫}par Claude Debru; Jean Gayon; Jean-François Picard. Revue d’Histoire des Sciences, 49(1):134–135, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633541.

Becquemont:1996:AOBb

[3634] Daniel Becquemont. Analyses d’ouvrages: Huxley : The Devil’s disciple par Adrian Desmond. Revue d’Histoire des Sciences, 49(1):136–137, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633542.

Balan:1996:AOB

[3635] B. Balan. Analyses d’ouvrages: Histoire du concept de récapitulation : ontogenèse et phylogenèse en biologie et sciences humaines, ^{≪}Histoire des sciences^{≫}par Paul Mengal. Revue d’Histoire des Sciences, 49(1):137–138, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633543.

Grmek:1996:AOBb

[3636] Mirko D. Grmek. Analyses d’ouvrages: French Medical Culture in the nineteenth century, ^{≪}Clio medica ^{≫}, 25, ^{≪}The Wellcome Institute series in the history of medicine ^{≫}par Ann La Berge; Mordechaï Feingold. Revue d’Histoire des Sciences, 49(1):138–139, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633544.

Gourevitch:1996:AOBa

[3637] Danielle Gourevitch. Analyses d’ouvrages: Prolegomena : Questions to be settled before the study of an author, or a text, ^{≪}Philosophia antiqua. A series of studies on ancient philosophy ^{≫}, vol. LXI par Jaap Mansfeld. Revue d’Histoire des Sciences, 49(1):139–140, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633545.

Kleiche:1996:AOB

[3638] Mina Kleiche. Analyses d’ouvrages: Nature, the exotic and the science of French colonialism, ^{≪}Science, technology and society ^{≫}par Michael A. Osborne. Revue d’Histoire des Sciences, 49(1):140–141, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633546.

Hamraoui:1996:AOB

[3639] Eric Hamraoui. Analyses d’ouvrages: Histoire de la notion de vie par André Pichot. Revue d’Histoire des Sciences, 49(1):141–143, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633547.

Gourevitch:1996:AOBb

[3640] Danielle Gourevitch. Analyses d’ouvrages: Theophrastus of Eresus : Sources for his life, writings, thought and influence. Commentary, vol. 5 : Sources on biology (human physiology, living creatures, botany : texts 328–435), ^{≪}Philosophia antiqua. A series of studies on ancient philosophy ^{≫}, vol. LXIV par R. W. Sharples. Revue d’Histoire des Sciences, 49(1):143, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633548.

Joly:1996:QAE

[3641] Bernard Joly. Quand l’alchimie était une science: Introduction. Revue d’Histoire des Sciences, 49(2–3):147–157, avril–septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633454.

Newman:1996:FDP

[3642] William R. Newman. ^{≪}Decknamen or pseudochemical language ^{≫}? Eirenaeus Philalethes and Carl Jung. Revue d’Histoire des Sciences, 49(2– 3):159–188, avril–septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633455.

Viano:1996:AAG

[3643] Cristina Viano. Aristote et l’alchimie grecque : La transmutation et le modèle aristotélicien entre théorie et pratique. Revue d’Histoire des Sciences, 49(2–3):189–213, avril–septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633456.

Obrist:1996:ANA

[3644] Barbara Obrist. Art et nature dans l’alchimie médiévale. Revue d’Histoire des Sciences, 49(2–3):215–286, avril–septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633457.

Clericuzio:1996:APC

[3645] Antonio Clericuzio. Alchimie, philosophie corpusculaire et minéralogie dans la “Metallographia” de John Webster. Revue d’Histoire des Sciences, 49(2–3):287–304, avril–septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633458.

Joly:1996:LDU

[3646] Bernard Joly. L’alkahest, dissolvant universel ou quand la théorie rend pensable une pratique impossible. (French) [Alkahest, universal solvent or when theory makes an impossible practice thinkable]. Revue d’Histoire des Sciences, 49(2–3):305–344, avril–septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633459.

Joly:1996:B

[3647] Bernard Joly. Bibliographie. Revue d’Histoire des Sciences, 49(2–3): 345–354, avril–septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633460.

Joly:1996:AOBa

[3648] Bernard Joly. Analyses d’ouvrages: Ecrits alchimiques. ^{≪}Aux sources de la tradition ^{≫}par Nicolas Flamel; Didier Kahn; Nicolas Flamel sous investigation, suivi de l’édition annotée du ^{≪}Livre des figures hiéroglyphiques ^{≫}, ^{≪}Le lieu du Loup ^{≫}par Claude Gagnon; La Table d’émeraude et sa tradition alchimique par Hermès Trismégiste; Didier Kahn. Revue d’Histoire des Sciences, 49(2–3):355–356, avril–septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633461.

Debus:1996:AOB

[3649] Allen G. Debus. Analyses d’ouvrages: Robert Boyle reconsidered par Michael Hunter. Revue d’Histoire des Sciences, 49(2–3):356–357, avril– septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633462.

Freudentahl:1996:AOB

[3650] Gad Freudentahl and Pierre Lory. Analyses d’ouvrages: Alchemie, Ketzerei, Apokryphen im frühen Islam: Gesammelte Aufsätze par Paul Kraus; Rémi Brague; Names, natures and things: The alchemist Jabir ibn H .ayyan and his Kitab al-Ah.jar (Book of Stones). ^{≪}Boston Studies in the Philosophy of Science ^{≫}, vol. 158 par Syed Nomanul Haq; David E. Pingree. Revue d’Histoire des Sciences, 49(2–3):357–361, avril– septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633463.

Joly:1996:AOBb

[3651] Bernard Joly. Analyses d’ouvrages: The ^{≪}Summa perfectionis ^{≫}of Pseudo-Geber. A Critical Edition, Translation and Study, ^{≪}Travaux de l’Académie Internationale d’Histoire des Sciences ^{≫}par William R. Newman; Constantine of Pisa, ^{≪}The book of the secrets of alchemy ^{≫}. Introduction, critical edition, translation and commentary, ^{≪}Travaux de l’Académie Internationale d’Histoire des Sciences ^{≫}par Barbara Obrist. Revue d’Histoire des Sciences, 49(2–3):361–364, avril–septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633464.

Fontaine:1996:AOB

[3652] Resianne Fontaine. Analyses d’ouvrages: The Jewish Alchemists : A history and source book par Raphael Patai. Revue d’Histoire des Sciences, 49(2–3):364–365, avril–septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633465.

Joly:1996:AOBc

[3653] Bernard Joly. Analyses d’ouvrages: Alchemy and chemistry in the 16th and 17th centuries, ^{≪}Archives internationales d’histoire des idées/International Archives of the ideas ^{≫}par Piyo Rattansi; Antonio Clericuzio. Revue d’Histoire des Sciences, 49(2–3):365–367, avril– septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633466.

Bernhardt:1996:AOB

[3654] Jean Bernhardt. Analyses d’ouvrages: Leviathan et la pompe à air: Hobbes et Boyle entre science et politique. ^{≪}Textes à l’appui, Sér. anthropologie des sc. et techn. ^{≫}[ 1^{re} éd.] par Steven Shapin; Simon Schaffer; Thierry Plélat; Sylvie Barjansky. Revue d’Histoire des Sciences, 49(2– 3):368–369, avril–septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633467.

Joly:1996:AOBd

[3655] Bernard Joly. Analyses d’ouvrages: The Business of Alchemy : Science and culture in the Holy Roman Empire par Pamela H. Smith. Revue d’Histoire des Sciences, 49(2–3):369–370, avril–septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633468.

Grmek:1996:MB

[3656] Mirko D. Grmek. La méthodologie de Boscovich. Revue d’Histoire des Sciences, 49(4):379–400, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633765.

Pappas:1996:RJB

[3657] John Pappas. R. J. Boscovich et l’Académie des sciences de Paris. Revue d’Histoire des Sciences, 49(4):401–414, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633766.

Taton:1996:RER

[3658] René Taton. Les relations entre R. J. Boscovich et Alexis-Claude Clairaut (1759–1764). Revue d’Histoire des Sciences, 49(4):415–458, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633767.

McClaughlin:1996:WTE

[3659] Trevor McClaughlin. Was there an empirical movement in midseventeenth century France? Experiments in Jacques Rohault’s “Traité de physique”. Revue d’Histoire des Sciences, 49(4):459–481, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633768.

Forbes:1996:CAE

[3660] Eric G. Forbes and Jacques Gapaillard. La correspondance astronomique entre l’abbé Nicolas-Louis de Lacaille et Tobias Mayer. Revue d’Histoire des Sciences, 49(4):483–541, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633769.

Morando:1996:ETD

[3661] Bruno Morando and Denis Savoie. Étude de la théorie du Soleil des Tables pruténiques. Revue d’Histoire des Sciences, 49(4):543–567, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633770.

Granger:1996:JCH

[3662] Gilles Gaston Granger. Jean Cavaillès et l’histoire. Revue d’Histoire des Sciences, 49(4):569–582, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633771.

Balibar:1996:AOB

[3663] Françoise Balibar. Analyses d’ouvrages: Collected Papers of Albert Einstein (The), vol. 3 : The Swiss years : writings, 1909–1911. vol. 5. The Swiss years : correspondence, 1902–1914 par Martin J. Klein; A. J. Kox; Jürgen Renn; Robert Schulmann. Revue d’Histoire des Sciences, 49(4):583–584, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633772.

Labica:1996:AOB

[3664] Georges Labica. Analyses d’ouvrages: Philosophie et la Révolution française (La). Actes du Colloque de la Société française de philosophie, 31 mai, 1^{er} et 2 juin 1989, ^{≪}Biblioth. d’hist. de la philos. ^{≫}par B. Bourgeois; J. D’Hondt. Revue d’Histoire des Sciences, 49(4):585, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633773.

Vitrac:1996:AOB

[3665] Bernard Vitrac. Analyses d’ouvrages: La Sphère, instrument au service de la découverte du monde : D’Autolycos de Pitanè à Jean de Sacrobosco, ^{≪}Varia ^{≫}, n^{o} 11 par Germaine Aujac. Revue d’Histoire des Sciences, 49(4):585–586, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633774.

Goldstein:1996:AOB

[3666] Catherine Goldstein. Analyses d’ouvrages: Messengers of mathematics: European mathematical journals, 1800–1946 ^{≪}Ciencia ^{≫}par Elena Ausejo; Mariano Hormigon. Revue d’Histoire des Sciences, 49(4):586– 587, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633775.

Schappacher:1996:AOB

[3667] Norbert Schappacher. Analyses d’ouvrages: Briefwechsel zwischen Karl Weierstraß und Sofja Kowalewskaja par Reinhard Bölling. Revue d’Histoire des Sciences, 49(4):587–588, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633776.

Ribemont:1996:AOB

[3668] Bernard Ribémont. Analyses d’ouvrages: Science in Western and Eastern civilization in Carolingian Times par Paul Leo Butzer; Dietrich Lohrmann. Revue d’Histoire des Sciences, 49(4):588–590, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633777.

Dhombres:1996:AOBb

[3669] Jean Dhombres. Analyses d’ouvrages: Les Enjeux du mobile : Mathématiques, physique, philosophie. ^{≪}Des travaux ^{≫}par Gilles Châtelet; J. T. Desanti. Revue d’Histoire des Sciences, 49(4):590–593, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633778.

Guillerme:1996:AOB

[3670] André Guillerme. Analyses d’ouvrages: Un ingénieur des Lumières : Emiland-Marie Gauthey par Anne Coste; Antoine Picon; François Sidot. Revue d’Histoire des Sciences, 49(4):593–594, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633779.

Dhombres:1996:AOBc

[3671] Jean Dhombres. Analyses d’ouvrages: Réflexions sur la mesure par François Dagognet. Revue d’Histoire des Sciences, 49(4):594–595, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633780.

Festa:1996:AOB

[3672] Egidio Festa. Analyses d’ouvrages: Et pourtant elle tourne : Le mouvement de la Terre par Jacques Gapaillard. Revue d’Histoire des Sciences, 49(4):595–600, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633781.

Debarbat:1996:AOB

[3673] Suzanne Débarbat. Analyses d’ouvrages: The Man who sold the milky way : A biography of Bart Bok par David H. Levy. Revue d’Histoire des Sciences, 49(4):600–601, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633782.

Vilain:1996:AOB

[3674] Christiane Vilain. Analyses d’ouvrages: Penetralia Motus : La fondazione relativistica della meccanica in Christiaan Huygens, con l’edizione del codex Hugeniorum 7A, ^{≪}Pubblicazioni della facoltà di lettere e filosofia dell’universitàdi Milano^{≫}, CLIII par Gianfranco Mormino. Revue d’Histoire des Sciences, 49(4):601–602, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633783.

Hallyn:1996:AOB

[3675] Fernand Hallyn. Analyses d’ouvrages: Novelties in the Heavens : Rhetoric and Science in the Copernican controversy par Jean Dietz Moss. Revue d’Histoire des Sciences, 49(4):602–603, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633784.

Dhombres:1996:AOBd

[3676] Jean Dhombres. Analyses d’ouvrages: Un univers de propensions : Deux études sur la causalité et l’évolution. ^{≪}Collection tiré à part ^{≫}par Karl Popper; Alain Boyer. Revue d’Histoire des Sciences, 49(4):603, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633785.

Dhombres:1996:AOBe

[3677] Jean Dhombres. Analyses d’ouvrages: Les Mécanismes de la découverte scientifique : Une épistémologie interactionniste, ^{≪}Philosophica ^{≫}, 44 par Serge Robert. Revue d’Histoire des Sciences, 49(4):604–605, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633786.

Anonymous:1996:PR

[3678] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 49(4):607– 612, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633787.

Anonymous:1996:TGM

[3679] Anonyme. Table générale des matières du Tome 49. Revue d’Histoire des Sciences, 49(4):613–616, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633788.

Anonymous:1996:BMa

[3680] Anonyme. Back matter. Revue d’Histoire des Sciences, 49(1):??, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633549.

Anonymous:1996:BMb

[3681] Anonyme. Back matter. Revue d’Histoire des Sciences, 49(2–3):??, avril– septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633469.

Anonymous:1996:BMc

[3682] Anonyme. Back matter. Revue d’Histoire des Sciences, 49(4):??, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633789.

Anonymous:1996:FMa

[3683] Anonyme. Front matter. Revue d’Histoire des Sciences, 49(1):??, January 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633528.

Anonymous:1996:FMb

[3684] Anonyme. Front matter. Revue d’Histoire des Sciences, 49(2–3):??, avril– septembre 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633453.

Anonymous:1996:FMc

[3685] Anonyme. Front matter. Revue d’Histoire des Sciences, 49(4):??, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633764.

Pinault:1997:NDF

[3686] Michel Pinault. Naissance d’un dessein: Frédéric Joliot et le nucléaire français (août 1944–septembre 1945). Revue d’Histoire des Sciences, 50(1–2):3–47, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633270.

Bustamante:1997:JSP

[3687] Martha Cecilia Bustamante. Jacques Solomon (1908–1942) : profil d’un physicien théoricien dans la France des années trente. Revue d’Histoire des Sciences, 50(1–2):49–87, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633271.

Lelong:1997:PVJ

[3688] Benoit Lelong. Paul Villard, J.-J. Thomson et la composition des rayons cathodiques. Revue d’Histoire des Sciences, 50(1–2):89–130, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633272.

Belna:1997:NRF

[3689] Jean-Pierre Belna. Les nombres réels : Frege critique de Cantor et de Dedekind. Revue d’Histoire des Sciences, 50(1–2):131–158, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633273.

Scheidecker-Chevallier:1997:LAA

[3690] Myriam Scheidecker-Chevallier. L’hypothèse d’Avogadro (1811) et d’Ampère (1814) : la distinction atome/molécule et la théorie de la combinaison chimique. Revue d’Histoire des Sciences, 50(1–2):159–194, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633274.

Darrigol:1997:HPP

[3691] Olivier Darrigol. Histoires de pratiques, pratiques de l’histoire. Revue d’Histoire des Sciences, 50(1–2):195–206, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633275.

Pasdeloup:1997:RM

[3692] Maurice Pasdeloup. Regards sur la matière. Revue d’Histoire des Sciences, 50(1–2):207–211, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633276.

Darrigol:1997:AOBa

[3693] Olivier Darrigol. Analyses d’ouvrages: Weber’s electrodynamics, ≪Fundamental theories of physics ^{≫}, vol. 66 par André Koch Torres Assis. Revue d’Histoire des Sciences, 50(1–2):213, January 1997. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633277.

Atten:1997:AOB

[3694] Michel Atten. Analyses d’ouvrages: Restaging Coulomb: Usages, controverses et réplications autour de la balance de torsion, ^{≪}Biblioteca di Nuncius. Studi e testi ^{≫}, XV par Christine Blondel; Matthias Dörries. Revue d’Histoire des Sciences, 50(1–2):213–214, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633278.

Parmentier:1997:AOBa

[3695] Marc Parmentier. Analyses d’ouvrages: La Mesure de l’État : Administrateurs et géomètres au XVIII^{e} siècle, ^{≪}L’évolution de l’humanité ^{≫}par Eric Brian. Revue d’Histoire des Sciences, 50(1–2):215–216, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633279.

Darrigol:1997:AOBb

[3696] Olivier Darrigol. Analyses d’ouvrages: The Creation of scientific effects: Heinrich Hertz and electric waves, ^{≪}History and Philosophy of Science/Physics ^{≫}par Jed Z. Buchwald. Revue d’Histoire des Sciences, 50(1–2):216–218, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633280.

Belna:1997:AOB

[3697] Jean-Pierre Belna. Analyses d’ouvrages: Frege’s theory of sense and reference : Its origins and scope, ^{≪}Modern European Philosophy ^{≫}par Wolfgang Carl. Revue d’Histoire des Sciences, 50(1–2):218–219, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633281.

Parmentier:1997:AOBb

[3698] Marc Parmentier. Analyses d’ouvrages: Arithmétique politique : Textes rares ou inédits (1767–1789) par Condorcet; Bernard Bru; Pierre Crépel. Revue d’Histoire des Sciences, 50(1–2):219–221, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633282.

Merleau-Ponty:1997:AOBa

[3699] Jacques Merleau-Ponty. Analyses d’ouvrages: Niels Bohr and contemporary philosophy, ^{≪}Boston Studies in the Philos. of Sc. ^{≫}, vol. 153 par Jan Faye; Henry J. Folse. Revue d’Histoire des Sciences, 50(1–2):221–223, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633283.

Brenner:1997:AOB

[3700] Anastasios Brenner. Analyses d’ouvrages: Trends in the historiography of science, ^{≪}Boston Studies in the Philos. of Sc. ^{≫}, vol. 151. [32 contributions présentées au Colloque intern. organisé par la Soc. grecque d’hist. des sc. et des techn. en juin 1990 à Corfou.] par Kostas Gavroglu; Jean Christianidis; Efthymios Nicolaïdis. Revue d’Histoire des Sciences, 50(1– 2):223, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633284.

Chevalley:1997:AOB

[3701] Catherine Chevalley. Analyses d’ouvrages: Physics and our view of the world par Jan Hilgevoord. Revue d’Histoire des Sciences, 50(1–2):224, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633285.

Merleau-Ponty:1997:AOBb

[3702] Jacques Merleau-Ponty. Analyses d’ouvrages: Eddington’s search for a fundamental theory : A key to the universe par C. W. Kilmister. Revue d’Histoire des Sciences, 50(1–2):225–226, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633286.

Clavelin:1997:AOBa

[3703] Maurice Clavelin. Analyses d’ouvrages: Histoire et méthodologie des sciences : Programmes de recherche et reconstruction rationnelle. ^{≪}Biblioth. d’hist. des sc. ^{≫}. [The Methodology of scientific research programmes (Philosophical Papers, vol. I), chap. 1, 2 et 4] par Imre Lakatos; Catherine Malamoud; Jean-Fabien Spitz; Luce Giard; L. Giard. Revue d’Histoire des Sciences, 50(1–2):226–228, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633287.

Jullien:1997:AOBa

[3704] Vincent Jullien. Analyses d’ouvrages: Jean Le Rond D’Alembert philosophe. ^{≪}Mathésis ^{≫} par Véronique Le Ru; Maurice Clavelin. Revue d’Histoire des Sciences, 50(1–2):228–229, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633288.

Petit:1997:AOB

[3705] Annie Petit. Analyses d’ouvrages: Auguste Comte : An intellectual biography, vol. I par Mary Pickering. Revue d’Histoire des Sciences, 50(1–2):229–231, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633289.

Kleinert:1997:AOB

[3706] Andreas Kleinert. Analyses d’ouvrages: Science, technology and national socialism par Monika Renneberg; Mark Walker. Revue d’Histoire des Sciences, 50(1–2):232, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633290.

Anonymous:1997:PRa

[3707] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 50 (1–2):233–235, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633291.

Blay:1997:RJH

[3708] Michel Blay and Nicole Hulin. René-Just Haüy (1743–1822), physicien. introduction. Revue d’Histoire des Sciences, 50(3):241–242, juillet– septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633351.

Hulin:1997:RJH

[3709] Nicole Hulin. René-Just Haüy : Des leçons de l’an III au “Traité élémentaire de physique”. Revue d’Histoire des Sciences, 50(3):243– 263, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633352.

Blondel:1997:HED

[3710] Christine Blondel. Haüy et l’électricité : de la démonstration-spectacle à la diffusion d’une science newtonienne. Revue d’Histoire des Sciences, 50(3):265–282, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633353.

Blay:1997:LCT

[3711] Michel Blay. Lumière et couleurs dans le “Traité” de Haüy, ou ^{≪}la véritable méthode pour parvenir à l’explication^{≫}dans les sciences. Revue d’Histoire des Sciences, 50(3):283–292, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633354.

Curien:1997:CRJ

[3712] Hubert Curien. Le cristallographe René-Just Haüy. Revue d’Histoire des Sciences, 50(3):293–294, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633355.

Guinier:1997:CHJ

[3713] André Guinier. La cristallographie, de Haüy jusqu’aux Bragg et à leurs successeurs. Revue d’Histoire des Sciences, 50(3):295–301, juillet– septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633356.

Touret:1997:RPI

[3714] Lydie Touret. Relations parisiennes et internationales de René-Just Haüy. Revue d’Histoire des Sciences, 50(3):303–334, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633357.

Bilodeau-Guinamard:1997:MRJ

[3715] Bénédicte Bilodeau-Guinamard. Les manuscrits de René-Just Haüy conservés à la bibliothèque centrale du Muséum national d’histoire naturelle de Paris. Revue d’Histoire des Sciences, 50(3):335–353, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633358.

Bilodeau-Guinamard:1997:BCR

[3716] Bénédicte Bilodeau-Guinamard. Bibliographie complémentaire de René-Just Haüy, depuis 1944. Revue d’Histoire des Sciences, 50(3):355– 356, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633359.

Ickowicz:1997:DMD

[3717] Pierre Ickowicz. Découverte à Montbard de deux manuscrits mathématiques de Buffon. Revue d’Histoire des Sciences, 50(3):357–360, juillet– septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633360.

Gowing:1997:PVM

[3718] Ronald Gowing. Pierre Varignon and the measurement of time. Revue d’Histoire des Sciences, 50(3):361–368, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633361.

Nicolaidis:1997:GG

[3719] Efthymios Nicolaïdis. Georges Goudaroulis (1945–1996). Revue d’Histoire des Sciences, 50(3):369–373, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633362.

Zerner:1997:AOB

[3720] Martin Zerner. Analyses d’ouvrages: Études historico-mathématiques, t. 34 : Sur l’histoire de l’analyse mathématique. Revue d’Histoire des Sciences, 50(3):375–376, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633363.

Brenier:1997:AOB

[3721] Joël Brenier. Analyses d’ouvrages: Sous les nombres, le monde : Matériaux pour l’histoire du nombre en Chine ancienne, ^{≪}Extrême– Orient, Extrême–Occident. Cahiers de recherches comparatives ^{≫}, 16 par Alexeï Volkov. Revue d’Histoire des Sciences, 50(3):376–377, juillet– septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633364.

Seris:1997:AOBa

[3722] Jean-Pierre Séris. Analyses d’ouvrages: La Mesure : Instruments et philosophies, ^{≪}Milieux ^{≫}. [Actes du Colloque qui s’est tenu au Centre d’analyse des formes et systèmes de la fac. de philos. de l’univ. J.- Moulin-Lyon III, les 28–29 oct. 1993.] par Jean-Claude Beaune. Revue d’Histoire des Sciences, 50(3):377–378, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633365.

Licoppe:1997:AOB

[3723] Christian Licoppe. Analyses d’ouvrages: La Formation polytechnicienne, 1794–1994 par Bruno Belhoste; Amy Dahan Dalmedico; Antoine Picon. Revue d’Histoire des Sciences, 50(3):378, juillet–septembre 1997. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633366.

Morange:1997:AOB

[3724] Michel Morange. Analyses d’ouvrages: Tout ce que vous devriez savoir sur la science. ^{≪}Science ouverte ^{≫}. [Ed. orig. : The Golem] par Harry Collins; Trevor Pinch; Thierry Piélat. Revue d’Histoire des Sciences, 50 (3):379–381, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633367.

Jullien:1997:AOBb

[3725] Vincent Jullien. Analyses d’ouvrages: Les Éléments, trad. du texte de Heiberg, vol. II : Livres V–VI : Proportions et similitude, Livres VII– IX : Arithmétique. ^{≪}Bibliothèque d’histoire des sciences ^{≫} par Euclide d’Alexandrie; Bernard Vitrac. Revue d’Histoire des Sciences, 50(3): 381–382, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633368.

Seris:1997:AOBb

[3726] Jean-Pierre Séris. Analyses d’ouvrages: Technology, pessimism, and postmodernism, ^{≪}Sociology of the Sc. ^{≫}, A Yearbook, vol. XVII par Yaron Ezrahi; Everett Mendelsohn; Howard Segal. Revue d’Histoire des Sciences, 50(3):382–383, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633369.

Clavelin:1997:AOBb

[3727] Maurice Clavelin. Analyses d’ouvrages: Galileo: For Copernicanism and for the church. ^{≪}Studi galileiani ^{≫}, vol. 3. [Originally publ. in Italian in 1993 under the title Galileo: Per il copernicanesimo e per la chiesa.] par Annibale Fantoli; George V. Coyne. Revue d’Histoire des Sciences, 50(3): 383–384, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633370.

Passeron:1997:AOB

[3728] Irène Passeron. Analyses d’ouvrages: The Problem of the Earth’s shape from Newton to Clairaut: The rise of mathematical science in eighteenthcentury Paris and the fall of ^{≪}normal ^{≫}science par John L. Greenberg. Revue d’Histoire des Sciences, 50(3):384–386, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633371.

Jami:1997:AOB

[3729] Catherine Jami. Analyses d’ouvrages: Les Mathématiques japonaises à l’époque d’Edo (1600–1868) : Une étude des travaux de Seki Takakazu (?–1708) et de Takebe Katahiro (1664–1739), ^{≪}Mathesis ^{≫}par Annick Horiuchi. Revue d’Histoire des Sciences, 50(3):386, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633372.

Gayon:1997:AOBa

[3730] Jean Gayon. Analyses d’ouvrages: La Fourmi et le sociobiologiste par Pieere Jaisson. Revue d’Histoire des Sciences, 50(3):386–387, juillet– septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633373.

Friedelmeyer:1997:AOB

[3731] Jean-Pierre Friedelmeyer. Analyses d’ouvrages: Friedrich Wilhelm Bessel, 1784–1846. ^{≪}Vita mathematica ^{≫}, Bd. 9. [ Ed. orig. : Moscou, Isdatel’stvo ^{≪}Nauka ^{≫}, 1989.] par Kasimir Lawrynowicz; Katja Hansen-Matyssek; Heinz Matyssek. Revue d’Histoire des Sciences, 50(3):387– 388, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633374.

Rosaura:1997:AOB

[3732] Ruiz G. Rosaura. Analyses d’ouvrages: El Megaterio de Bru y el presidente Jefferson: Una relación insospechada en los albores de la paleontología, ^{≪}Cuadernos valenc. de hist. de la med. y de la ciencia ^{≫}, ser. A (monograf.), XLII par José María López Piñero; Thomas F. Glick. Revue d’Histoire des Sciences, 50(3):388–389, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633375.

Fauque:1997:AOB

[3733] Danielle Fauque. Analyses d’ouvrages: The Cross-staff : History and development of a navigational instrument par Willem F. J. Mörzer– Bruyns. Revue d’Histoire des Sciences, 50(3):389–390, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633376.

Poulle:1997:AOB

[3734] Emmanuel Poulle. Analyses d’ouvrages: Die astronomische Uhr des Strassburger Münsters : Funktion und Bedeutung eines Kosmos-Modells des 16. Jahrhunderts par Günther Oestmann. Revue d’Histoire des Sciences, 50(3):390–391, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633377.

Phaf:1997:AOB

[3735] Ineke Phaf. Analyses d’ouvrages: Las Primeras Noticias sobre plantas americanas en las relaciones de viajes y crónicas de Indias (1493–1553), ≪Cuadernos valenc. de hist. de la med. y de la ciencia ^{≫}, ser. A. (monograf.), XL par José Pardo Tomás; María Luz López Terrada; La Experiencia americana y la terapéutica en los Secretos de chirurgía (1567), de Pedro Arias de Benavides, ^{≪}Cuadermos valenc. de hist. de la med. y de la ciencia ^{≫}, ser. A (monograf.), XLI par José Luis Fresquet Febrer; Nuevos Materiales y noticias sobre la historia de las plantas de Nueva España, de Francisco Hernández, ^{≪}Cuadernos valenc. de hist. de la med. y de la ciencia ^{≫}, ser. A (monograf.), XLIV par José María López Piñero; José Pardo Tomás. Revue d’Histoire des Sciences, 50(3):391–394, juillet– septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633378.

Pellegrin:1997:AOB

[3736] Pierre Pellegrin. Analyses d’ouvrages: Recherches sur les plantes, t. III, livres V–VI. ^{≪}Coll. des univ. de France ^{≫}par Théophraste; Suzanne Amigues. Revue d’Histoire des Sciences, 50(3):394, juillet–septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633379.

Salmon:1997:MGM

[3737] Fernando Salmón. The many Galens of the medieval commentators on vision. Revue d’Histoire des Sciences, 50(4):397–419, October 1997. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633316.

Ratcliff:1997:CIP

[3738] Marc James Ratcliff. Le concept d’intensité dans la psychologie de Charles Bonnet. Revue d’Histoire des Sciences, 50(4):421–446, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633317.

Benaroyo:1997:RPB

[3739] Lazare Benaroyo. La renaissance de la pensée biologique de Rudolf Virchow dans l’œuvre de Ludwig Aschoff. Revue d’Histoire des Sciences, 50(4):447–460, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633318.

Acot:1997:LFI

[3740] Pascal Acot and Jean-Marc Drouin. L’introduction en France des idées de l’écologie scientifique américaine dans l’entre-deux-guerres. Revue d’Histoire des Sciences, 50(4):461–479, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633319.

Ribemont:1997:AOBa

[3741] Bernard Ribémont. Analyses d’ouvrages: Comprendre et maîtriser la nature au Moyen Âge : Mélanges d’histoire des sciences offerts à Guy Beaujouan. ^{≪}V. Hautes Études médiévales et modernes ^{≫}, 73. Revue d’Histoire des Sciences, 50(4):481–483, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633320.

Niderst:1997:AOB

[3742] Alain Niderst. Analyses d’ouvrages: Nature, histoire, société: Essais en hommage à Jacques Roger par Claude Blanckaert; Jean-Louis Fischer; Roselyne Rey. Revue d’Histoire des Sciences, 50(4):484, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633321.

Matagne:1997:AOB

[3743] Patrick Matagne. Analyses d’ouvrages: The Naturalist in Britain : A social history. 2nd ed. ^{≪}History and philosophy of science/Biological sciences ^{≫}. [1st ed.] par David Elliston Allen. Revue d’Histoire des Sciences, 50(4):484–485, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633322.

Gourevitch:1997:AOB

[3744] Danielle Gourevitch. Analyses d’ouvrages: Medicine, society, and faith in the ancient and medieval worlds par Darrel W. Amundsen. Revue d’Histoire des Sciences, 50(4):485–486, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633323.

Josset:1997:AOB

[3745] Patrice Josset. Analyses d’ouvrages: Les Papyrus médicaux de l’Egypte pharaonique. ^{≪}Penser la médecine ^{≫}par Thierry Bardinet. Revue d’Histoire des Sciences, 50(4):486, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633324.

Dhombres:1997:AOBa

[3746] Jean Dhombres. Analyses d’ouvrages: Les Sciences dans l’enseignement secondaire français, t. 1 : 1789–1914 par Bruno Belhoste; Claudette Balpe; Thierry Laporte. Revue d’Histoire des Sciences, 50(4):487–488, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633325.

Bathiard:1997:AOB

[3747] Arnaud Bathiard. Analyses d’ouvrages: Le Fondement social des découvertes scientifiques. ^{≪}Sciences, modernités, philosophies ^{≫}. [ The Social Basis of scientific discoveries] par Augustine Brannigan; Christian Puech. Revue d’Histoire des Sciences, 50(4):489–490, October 1997. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633326.

Becquemont:1997:AOB

[3748] Daniel Becquemont. Analyses d’ouvrages: Charles Darwin : voyaging : Volume I of a biography par Janet Browne. Revue d’Histoire des Sciences, 50(4):490–492, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633327.

Gohau:1997:AOB

[3749] Gabriel Gohau. Analyses d’ouvrages: Horace–Bénédict de Saussure : forerunner in glaciology. New manuscript evidence on the earliest explorations of the glaciers of Chamonix and the fundamental contribution of Horace–Bénédict de Saussure to the study of glaciers between 1760 and 1792, ^{≪}Mémoires de la Société de physique et d’histoire naturelle de Genève ^{≫}, 48 par Albert V. Carozzi; John K. Newman. Revue d’Histoire des Sciences, 50(4):492–493, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633328.

Tete:1997:AOB

[3750] Alain Tête. Analyses d’ouvrages: La Communication non verbale avant la lettre, ^{≪}Histoire des sciences humaines ^{≫}par Anne-Marie Drouin-Hans. Revue d’Histoire des Sciences, 50(4):493, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633329.

Gayon:1997:AOBb

[3751] Jean Gayon. Analyses d’ouvrages: Essays in the history and philosophy of science par Pierre Duhem; Roger Ariew; Peter Barker. Revue d’Histoire des Sciences, 50(4):493–494, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633330.

Bastit:1997:AOB

[3752] Michel Bastit. Analyses d’ouvrages: Self-motion : From Aristotle to Newton par Mary Louise Gill; James G. Lennox. Revue d’Histoire des Sciences, 50(4):494–495, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633331.

Grmek:1997:AOBa

[3753] Mirko D. Grmek. Analyses d’ouvrages: Bibliographie médicale des Antilles françaises : Imprimés médicaux dans les colonies françaises des Antilles sous l’Ancien Régime : Saint-Domingue, Guadeloupe, La Martinique, Sainte-Lucie, Grenade et Guyane, 1765–1805 par Francisco Guerra. Revue d’Histoire des Sciences, 50(4):495–496, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633332.

Lejbowicz:1997:AOB

[3754] Max Lejbowicz. Analyses d’ouvrages: Bede, the schools and the Computus. ^{≪}Collected Studies Series ^{≫}, 436 par Charles W. Jones; Wesley M. Stevens. Revue d’Histoire des Sciences, 50(4):496–498, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633333.

Espinosa:1997:AOB

[3755] Julieta Espinosa. Analyses d’ouvrages: Steno, life, science, philosophy : With Niels Stensen’s Prooemium or ^{≪}Preface to a demonstration in the Copenhagen Anatomical Theater in the year 1673 ^{≫}and Holger Jacoboeus ≪Niels Stensen’s anatomical demonstration no. XVI ^{≫}, ^{≪}Acta historica scientiarum naturalium et medicinalium ^{≫}, 42 par Troels Kardel; Emmanuel Collins; Paul Maquet. Revue d’Histoire des Sciences, 50(4):498– 499, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633334.

Lherminier:1997:AOB

[3756] Philippe Lherminier. Analyses d’ouvrages: Au nom de l’eugénisme: Génétique et politique dans le monde anglo-saxon. ^{≪}Science, histoire et société^{≫}. [Ed. orig.: In the name of eugenics, genetics and the uses of human heredity, 1985] par Daniel J. Kevles; Marcel Blanc. Revue d’Histoire des Sciences, 50(4):499–500, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633335.

Grmek:1997:AOBb

[3757] Mirko D. Grmek. Analyses d’ouvrages: Neuer Kommentar zu den ersten zwölf Krankengeschichten im III. Epidemienbuch des Hippokrates, XV: Hippokratische Studie. ^{≪}Hermes Einzelschriften ^{≫}, 65 par Charles Lichtenthaeler; Markwart Michler. Revue d’Histoire des Sciences, 50(4):500– 501, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633336.

Nicolaidis:1997:AOB

[3758] Efthymios Nicolaïdis. Analyses d’ouvrages: An almanac for Trebizond for the year 1336, ^{≪}Corpus des astronomes byzantins ^{≫}, VII par Raymond Mercier. Revue d’Histoire des Sciences, 50(4):501–502, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633337.

Ribemont:1997:AOBb

[3759] Bernard Ribémont. Analyses d’ouvrages: L’Enciclopedismo medievale, ≪Memoria del tempo ^{≫}. [Actes du Colloque ^{≪}L’enciclopedismo medievale ≫, San Gimignano, 8–10 oct. 1992.] par Michelangelo Picone. Revue d’Histoire des Sciences, 50(4):502–504, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633338.

Triadou:1997:AOB

[3760] Patrick Triadou. Analyses d’ouvrages: Su Wen : Les 11 premiers traités, ≪Variétés sinologiques^{≫}par Elisabeth Rochat de la Vallée; Claude Larre. Revue d’Histoire des Sciences, 50(4):504–505, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633339.

Dhombres:1997:AOBb

[3761] Jean Dhombres. Analyses d’ouvrages: Libres Propos sur l’histoire des sciences, ^{≪}Bibliothèque scientifique Albert Blanchard ^{≫}par François Russo. Revue d’Histoire des Sciences, 50(4):505–507, October 1997. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633340.

Lowy:1997:AOB

[3762] Ilana Löwy. Analyses d’ouvrages: The Immune Self : Theory or metaphor?. ^{≪}Cambridge Studies in philosophy and biology ^{≫}par Alfred I. Tauber. Revue d’Histoire des Sciences, 50(4):507–508, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633341.

Elie:1997:AOB

[3763] Maurice Elie. Analyses d’ouvrages: In the eye’s mind : Vision and Helmholtz-Hering controversy par Roy Steven Turner. Revue d’Histoire des Sciences, 50(4):508–510, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633342.

Moulin:1997:AOB

[3764] Anne-Marie Moulin. Analyses d’ouvrages: La Médecine islamique, ≪Islamiques ^{≫}. [Trad. de Islamic Medicine, 1978.] par Manfred Ullmann; Fabienne Hareau. Revue d’Histoire des Sciences, 50(4):510, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633343.

Sinding:1997:AOB

[3765] Christiane Sinding. Analyses d’ouvrages: Comment Fleming n’a pas inventé la pénicilline. ^{≪}Les Empêcheurs de penser en rond ^{≫}par Wai chen; Sophie Mayoux. Revue d’Histoire des Sciences, 50(4):511, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633344.

Dupont:1997:AOB

[3766] Jean-Claude Dupont. Analyses d’ouvrages: Julius Wagner-Jauregg (1857–1940) par Magda Whitrow. Revue d’Histoire des Sciences, 50 (4):511–512, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633345.

Balan:1997:AOB

[3767] Bernard Balan. Analyses d’ouvrages: Histoire du développement de la biologie, vol. III. ^{≪}Biologie ^{≫}. [Ed. orig. : Ontwikkelingsgeschiedenis van de biologie, 3 vol. vol. I et II] par Hendrik C. D. de Wit; A. Baudière. Revue d’Histoire des Sciences, 50(4):512–513, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633346.

Anonymous:1997:PRb

[3768] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 50(4):515– 525, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633347.

Anonymous:1997:TGM

[3769] Anonyme. Table générale des matières du Tome 50. Revue d’Histoire des Sciences, 50(4):527–531, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633348.

Anonymous:1997:BMa

[3770] Anonyme. Back matter. Revue d’Histoire des Sciences, 50(1–2):??, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633292.

Anonymous:1997:BMb

[3771] Anonyme. Back matter. Revue d’Histoire des Sciences, 50(3):??, juillet– septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633380.

Anonymous:1997:BMc

[3772] Anonyme. Back matter. Revue d’Histoire des Sciences, 50(4):??, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633349.

Anonymous:1997:FMa

[3773] Anonyme. Front matter. Revue d’Histoire des Sciences, 50(1–2):??, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633269.

Anonymous:1997:FMb

[3774] Anonyme. Front matter. Revue d’Histoire des Sciences, 50(3):??, juillet– septembre 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633350.

Anonymous:1997:FMc

[3775] Anonyme. Front matter. Revue d’Histoire des Sciences, 50(4):??, October 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633315.

Darrigol:1998:CEM

[3776] Olivier Darrigol. Aux confins de l’électrodynamique maxwellienne : Ions et électrons vers 1897. Revue d’Histoire des Sciences, 51(1):5–34, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633905.

Berlioz:1998:ELL

[3777] Dominique Berlioz and Filipe Drapeau Contim. Un essai logique de Leibniz, ^{≪}le calcul des ingrédients ^{≫}: Présentation et traduction. Revue d’Histoire des Sciences, 51(1):35–64, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633906.

Kury:1998:IVE

[3778] Lorelai Kury. Les instructions de voyage dans les expéditions scientifiques françaises (1750–1830). Revue d’Histoire des Sciences, 51(1):65–91, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633907.

Huta:1998:JSD

[3779] Carole Huta. Jean Senebier (1742–1809) : Un dialogue entre l’ombre et la lumière l’art d’observer à la fin du XVIII^{e} siècle. Revue d’Histoire des Sciences, 51(1):93–105, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633908.

Lerner:1998:DCF

[3780] Michel-Pierre Lerner. Le débat cosmologique à la fin du XVI^{e} siècle : Sur un ouvrage de Miguel A. Granada. Revue d’Histoire des Sciences, 51(1):107–114, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633909.

Clavelin:1998:GHC

[3781] Maurice Clavelin. Galilée, homme de cour : Sur un ouvrage de Mario Biagioli. Revue d’Histoire des Sciences, 51(1):115–126, January 1998. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633910.

Descotes:1998:GCL

[3782] Dominique Descotes. Genèse des Corollaires 1 et 2 de la “Lettre à Carcavy” de Blaise Pascal. Revue d’Histoire des Sciences, 51(1):127–138, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633911.

Toulmonde:1998:JPS

[3783] Michel Toulmonde. Jean Picard et les satellites de Saturne. Revue d’Histoire des Sciences, 51(1):139–143, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633912.

Malov:1998:DMT

[3784] Vladimir N. Malov. Découvert à Moscou: Le traité inédit d’un académicien des sciences de Paris sur les poudres (1720). Revue d’Histoire des Sciences, 51(1):145–149, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633913.

Bret:1998:AOB

[3785] Patrice Bret. Analyses d’ouvrages: Lavoisier, ex-libris : Une collection bordelaise Service inter-établissements de coopération documentaire, Bibliothèque univ. des sciences et techniques de Bordeaux. Revue d’Histoire des Sciences, 51(1):151–152, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633914.

Fauque:1998:AOB

[3786] Danielle Fauque. Analyses d’ouvrages: Lavoisier in European context : Negotiating a new language for chemistry, ^{≪}European Studies in science history and the arts ^{≫}, I par Bernadette Bensaude-Vincent; Ferdinando Abbri. Revue d’Histoire des Sciences, 51(1):152–153, January 1998. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633915.

Elie:1998:AOB

[3787] Maurice Elie. Analyses d’ouvrages: Les Figures de l’arc-en-ciel, ^{≪}Vues des sciences ^{≫}par Michel Blay. Revue d’Histoire des Sciences, 51(1):153– 154, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633916.

Guillerme:1998:AOB

[3788] André Guillerme. Analyses d’ouvrages: Science, industry, and the social order in post-revolutionary France,^{≪}Collected Studies series^{≫}par Robert Fox. Revue d’Histoire des Sciences, 51(1):154–155, January 1998. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633917.

Lagues:1998:AOBa

[3789] Michel Laguës. Analyses d’ouvrages: Fritz London : A scientific biography par Kostas Gavroglu. Revue d’Histoire des Sciences, 51(1):155, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633918.

Shea:1998:AOB

[3790] William Shea. Analyses d’ouvrages: Discours et démonstrations mathématiques concernant deux sciences nouvelles. ^{≪}Epiméthée ^{≫} par Galilée; Maurice Clavelin. Revue d’Histoire des Sciences, 51(1):155–156, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633919.

Dhombres:1998:AOB

[3791] Jean Dhombres. Analyses d’ouvrages: Traité des courbes spéciales remarquables planes et gauches, 3 t., réimpr. [1^{re} éd. française : 1908 (t. I), 1909 (t. II), 1915 (t. III).] par Francisco Gomes Teixeira. Revue d’Histoire des Sciences, 51(1):156–157, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633920.

Darrigol:1998:AOBa

[3792] Olivier Darrigol. Analyses d’ouvrages: The Scientific Letters and papers of James Clerk Maxwell, vol. II : 1862–1873 par P. M. Harman. Revue d’Histoire des Sciences, 51(1):157, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633921.

Brenner:1998:AOBa

[3793] Anastasios Brenner. Analyses d’ouvrages: Les Savants et l’épistémologie vers la fin du XIX^{e} siècle, ^{≪}Bibliothèque scientifique Albert Blanchard ≫par Marco Panza; Jean-Claude Pont. Revue d’Histoire des Sciences, 51(1):157–158, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633922.

Licoppe:1998:AOB

[3794] Christian Licoppe. Analyses d’ouvrages: The Diffident Naturalist : Robert Boyle and the philosophy of experiment, ^{≪}Science and its conceptual foundations ^{≫} par Rose-Mary Sargent. Revue d’Histoire des Sciences, 51(1):158–159, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633923.

Lejbowicz:1998:AOBa

[3795] Max Lejbowicz. Analyses d’ouvrages: Les Mathématiques dans l’Occident médiéval, ^{≪}Mathématiques et civilisation ^{≫}par Jean de Siebenthal. Revue d’Histoire des Sciences, 51(1):159–161, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633924.

Lejbowicz:1998:AOBb

[3796] Max Lejbowicz. Analyses d’ouvrages: Cycles of time and scientific learning in medieval Europe, ^{≪}Collected Studies series ^{≫}par Wesley M. Stevens. Revue d’Histoire des Sciences, 51(1):161–162, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633925.

Darrigol:1998:AOBb

[3797] Olivier Darrigol. Analyses d’ouvrages: From obscurity to enigma : The work of Oliver Heaviside, 1872–1889, ^{≪}Science networks. Historical studies ^{≫}par Ido Yavetz. Revue d’Histoire des Sciences, 51(1):162–163, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633926.

Serfati:1998:FPD

[3798] Michel Serfati. ^{≪}Pour Descartes ^{≫}: Mathématiques et physique cartésiennes. introduction. Revue d’Histoire des Sciences, 51(2–3): 171–182, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/ 23633586. “Pour Descartes”: mathématiques et physique cartésiennes (Paris, 1996).

Israel:1998:BB

[3799] Giorgio Israel. Des Regulæ à la Géométrie. Revue d’Histoire des Sciences, 51(2–3):183–236, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor. org/stable/23633587. “Pour Descartes”: mathématiques et physique cartésiennes (Paris, 1996).

Serfati:1998:DCE

[3800] Michel Serfati. Descartes et la constitution de l’écriture symbolique mathématique. Revue d’Histoire des Sciences, 51(2–3):237–289, avril– septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633588. “Pour

Descartes”: mathématiques et physique cartésiennes (Paris, 1996).

Bos:1998:SBD

[3801] Henk J. M. Bos. La structure de la Géométrie de Descartes. Revue d’Histoire des Sciences, 51(2–3):291–317, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:/

/www.jstor.org/stable/23633589. “Pour Descartes”: mathématiques

et physique cartésiennes (Paris, 1996).

Blay:1998:RCS

[3802] Michel Blay. Les règles cartésiennes de la science du mouvement dans Le monde ou traité de la lumière. Revue d’Histoire des Sciences, 51(2– 3):319–346, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/ 23633590. “Pour Descartes”: mathématiques et physique cartésiennes (Paris, 1996).

Rodis-Lewis:1998:QRQ

[3803] Geneviève Rodis-Lewis. Quelques remarques sur la question de la vitesse de la lumière chez Descartes. Revue d’Histoire des Sciences, 51(2–3): 347–354, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/ 23633591. “Pour Descartes”: mathématiques et physique cartésiennes (Paris, 1996).

Gregoire:1998:CED

[3804] Michèle Grégoire. La correspondance entre Descartes et Fermat. Revue d’Histoire des Sciences, 51(2–3):355–362, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:/

/www.jstor.org/stable/23633592. “Pour Descartes”: mathématiques

et physique cartésiennes (Paris, 1996).

Jullien:1998:DRR

[3805] Vincent Jullien. Descartes–Roberval, une relation tumultueuse. Revue d’Histoire des Sciences, 51(2–3):363–371, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:/

/www.jstor.org/stable/23633593. “Pour Descartes”: mathématiques

et physique cartésiennes (Paris, 1996).

Vilain:1998:DCS

[3806] Christiane Vilain. Descartes, correspondant scientifique de Constantyn Huygens. Revue d’Histoire des Sciences, 51(2–3):373–379, avril– septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633594. “Pour

Descartes”: mathématiques et physique cartésiennes (Paris, 1996).

Dubois:1998:AOB

[3807] Jacques Dubois. Analyses d’ouvrages: Les ^{≪}Principia ^{≫}de Newton, ≪Philosophies ^{≫}par Michel Blay. Revue d’Histoire des Sciences, 51(2– 3):381–382, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633595.

McClaughlin:1998:BRB

[3808] Trevor McClaughlin. Book review: Discipline & experience : The mathematical way in the scientific revolution, ^{≪}Science and its conceptual foundations ^{≫}by Peter Dear. Revue d’Histoire des Sciences, 51(2–3): 382–383, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633596.

Delon:1998:AOB

[3809] Michel Delon. Analyses d’ouvrages: Études sur le XVIII siècle: XXIII: Autour du père Castel et du clavecin oculaire par Roland Mortier; Hervé Hasquin. Revue d’Histoire des Sciences, 51(2–3):383–384, avril– septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633598.

Biard:1998:AOBa

[3810] Joël Biard. Analyses d’ouvrages: The Foundations of modern science in the Middle Ages : Their religious institutional and intellectual context, index, ^{≪}Cambridge history of science ^{≫}par Edward Grant. Revue d’Histoire des Sciences, 51(2–3):384, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633600.

Caveing:1998:AOB

[3811] Maurice Caveing. Analyses d’ouvrages: Éléments de géométrie de G. P. de Roberval, ^{≪}Mathesis ^{≫}par Vincent Jullien. Revue d’Histoire des Sciences, 51(2–3):384–388, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633603.

Poulle:1998:AOB

[3812] Emmanuel Poulle. Analyses d’ouvrages: Pélerin de Prusse on the astrolabe. ^{≪}Practique de astralabe ^{≫}, ^{≪}Medieval and Renaissance Texts and Studies^{≫}, vol. 127 par Edgar Laird; Robert Fischer. Revue d’Histoire des Sciences, 51(2–3):388–390, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633605.

Perfetti:1998:AOB

[3813] Amalia Perfetti. Analyses d’ouvrages: La Nave di ermete : La cosmologia di Girolamo Fracastoro, ^{≪}Istit. nazionale di studi sul Rinascimento, Quaderni di Rinascimento ^{≫}, XXXI par Enrico Peruzzi. Revue d’Histoire des Sciences, 51(2–3):390, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633607.

LeRu:1998:AOB

[3814] Véronique Le Ru. Analyses d’ouvrages: Descartes’s Imagination : Proportion, images, and the activity of thinking par Dennis L. Sepper. Revue d’Histoire des Sciences, 51(2–3):390–392, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633609.

Biard:1998:AOBb

[3815] Joël Biard. Analyses d’ouvrages: Richard Brinkley’s Obligationes : A late fourteenth century treatise on the logic of disputation, ^{≪}Beiträge zur Gesch. der Philos. und Theol. des Mittelalters ^{≫}, Neue Folge, Bd. 43 par Paul Vincent Spade; Gordon A. Wilson. Revue d’Histoire des Sciences, 51(2–3):392, avril–septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633611.

Pourprix:1998:AOB

[3816] Bernard Pourprix. Analyses d’ouvrages: Chladni und die Entwicklung der Akustik von 1750–1860, ^{≪}Science Networks, Historical Studies ^{≫}, 19 par Dieter Ullmann. Revue d’Histoire des Sciences, 51(2–3):393, avril– septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633613.

Drouin:1998:I

[3817] Jean-Marc Drouin and Nicole Hulin. Introduction. Revue d’Histoire des Sciences, 51(4):403–407, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633845.

Hulin:1998:PSN

[3818] Nicole Hulin. La place des sciences naturelles au sein de l’enseignement scientifique au XIX^{e} siècle. Revue d’Histoire des Sciences, 51(4):409–433, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633847.

Fischer:1998:CCE

[3819] Jean-Louis Fischer. La création de la chaire d’embryogénie comparée du Collège de France (1844): A la mémoire d’Étienne Wolff (1904–1996). Revue d’Histoire des Sciences, 51(4):435–456, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633849.

Blanckaert:1998:GAM

[3820] Claude Blanckaert. Les Gaulois au Muséum : Savoirs naturalistes et principe des nationalités à l’époque romantique. Revue d’Histoire des Sciences, 51(4):457–505, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633851.

Drouin-Hans:1998:IPA

[3821] Anne-Marie Drouin-Hans and Jean-Marc Drouin. Les idées pédagogiques d’Augustin-Pyramus et d’Alphonse de Candolle. Revue d’Histoire des Sciences, 51(4):507–534, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633853.

Anonymous:1998:I

[3822] Anonyme. Informations. Revue d’Histoire des Sciences, 51(4):535, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633855.

Rosmorduc:1998:AOBa

[3823] Jean Rosmorduc. Analyses d’ouvrages: Chronologie d’histoire des sciences, ^{≪}Le Temps déployé ^{≫}. Revue d’Histoire des Sciences, 51(4):537, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633857.

Ferrari:1998:AOB

[3824] Jean Ferrari. Analyses d’ouvrages: Le Bien, le mal et la science : Les dimensions éthiques de l’entreprise techno-scientifique. ^{≪}Thémis. Philosophie ^{≫}par Evandro Agazzi; d’Isolda Agazzi. Revue d’Histoire des Sciences, 51(4):537–538, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633859.

Rosmorduc:1998:AOBb

[3825] Jean Rosmorduc. Analyses d’ouvrages: Les Sciences au lycée : Un siècle de réformes des mathématiques et de la physique en France et à l’étranger par Bruno Belhoste; Hélène Gispert; Nicole Hulin. Revue d’Histoire des Sciences, 51(4):538–540, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633861.

Brenner:1998:AOBb

[3826] Anastasios Brenner. Analyses d’ouvrages: Ludwig Boltzmann, his later life and philosophy, 1900–1906, book one : A documentary history, ≪Boston Studies in the philosophy of science ^{≫}, vol. 168 par John Blackmore. Revue d’Histoire des Sciences, 51(4):540–541, October 1998. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633863.

Schnitter:1998:AOB

[3827] Claude Schnitter. Analyses d’ouvrages: Les Sciences physiologiques en Europe : Analyses comparées du XIX^{e} siècle. ^{≪}Science-Histoire-Philosophie ^{≫} par Christian Bonah; Frederic L. Holmes. Revue d’Histoire des Sciences, 51(4):541, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633865.

McClellan:1998:AOB

[3828] James E. McClellan III. Analyses d’ouvrages: Histoire et mémoire de l’Académie des sciences : Guide de recherches par Eric Brian; Christiane Demeulenaere–Douyère. Revue d’Histoire des Sciences, 51(4):542, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633867.

Pont:1998:AOB

[3829] Jean-Claude Pont. Analyses d’ouvrages: Teoria delle parallele secondo Lobacevskij : Nuova didattica delle geometrie non euclidee par Salvatore Cicenia; Antonino Drago. Revue d’Histoire des Sciences, 51(4):543–544, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633869.

Turner:1998:AOB

[3830] Anthony J. Turner. Analyses d’ouvrages: Ancient Egyptian Science, vol. II : Calendars, clocks, and astronomy, ^{≪}Memoirs of the American philosophical society ^{≫}, vol. 214 par Marshall Clagett. Revue d’Histoire des Sciences, 51(4):544–545, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633871.

Raichvarg:1998:AOBa

[3831] Daniel Raichvarg. Analyses d’ouvrages: La Transmission des savoirs scientifiques. ^{≪}Colloques du CTHS ^{≫}, 15 par Rémi Coutin; Henri Hudrisier; Marcel V. Locquin. Revue d’Histoire des Sciences, 51(4):545–546, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633873.

Drouin:1998:AOB

[3832] Jean-Marc Drouin. Analyses d’ouvrages: La Bretagne des savants et des ingénieurs : 1825–1900 par Jean Dhombres; Michel Cabaret; Nicole Dhombres. Revue d’Histoire des Sciences, 51(4):546–548, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633874.

Bessot:1998:AOB

[3833] Didier Bessot. Analyses d’ouvrages: Géométrie.^{≪}Sources du savoir^{≫}par Albrecht Dürer; Jeanne Peiffer. Revue d’Histoire des Sciences, 51(4):548– 549, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633875.

Gayon:1998:AOB

[3834] Jean Gayon. Analyses d’ouvrages: Le Chaudron de Médée : L’expérimentation sur le vivant dans l’Antiquité, ^{≪}Les Empêcheurs de penser en rond ^{≫}par Mirko D. Grmek. Revue d’Histoire des Sciences, 51(4):550, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633876.

Locqueneux:1998:AOB

[3835] Robert Locqueneux. Analyses d’ouvrages: André-Marie Ampère : Enlightenment and electrodynamics, ^{≪}Cambridge Science Biographies^{≫}par James R. Hofmann. Revue d’Histoire des Sciences, 51(4):551, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633877.

Bathiard:1998:AOB

[3836] Arnaud Bathiard. Analyses d’ouvrages: Misunderstanding science? The public reconstruction of science and technology par Alan Irwin; Brian Wynne. Revue d’Histoire des Sciences, 51(4):552–553, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633878.

Raichvarg:1998:AOBb

[3837] Daniel Raichvarg. Analyses d’ouvrages: De la science en littérature à la sciencefiction. ^{≪}Colloques du CTHS ^{≫}par Danielle Jacquart. Revue d’Histoire des Sciences, 51(4):553–554, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633879.

Greene:1998:BRB

[3838] John C. Greene. Book review: Cultures of natural history par Nicholas Jardine; James A. Secord; Emma C. Spary. Revue d’Histoire des Sciences, 51(4):554–556, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633880.

Gilain:1998:AOB

[3839] Christian Gilain. Analyses d’ouvrages: Mathematics of the 19th century : Geometry, analytic function theory par A. N. Kolmogorov; A. P. Yushkevich; Roger Cooke. Revue d’Histoire des Sciences, 51(4):556–557, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633881.

Lagues:1998:AOBb

[3840] Michel Laguës. Analyses d’ouvrages: After the breakthrough : The emergence of high-temperature superconductivity as a research field par Helga Nowotny; Ulrike Felt. Revue d’Histoire des Sciences, 51(4):557–558, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633882.

Gispert:1998:AOB

[3841] Hélène Gispert. Analyses d’ouvrages: Development of mathematics, 1900–1950 par Jean-Paul Pier. Revue d’Histoire des Sciences, 51(4):558– 559, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633883.

Lejbowicz:1998:AOBc

[3842] Max Lejbowicz. Analyses d’ouvrages: Astronomie planétaire au Moyen Âge latin, ^{≪}Collected Studies Series ^{≫}par Emmanuel Poulle. Revue d’Histoire des Sciences, 51(4):560, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633884.

Renault:1998:AOB

[3843] Emmanuel Renault. Analyses d’ouvrages: Festschriften der Versammlungen Deutscher Naturforscher und ´’arzte : 1822 bis 1920, ^{≪}Schrift. zur Gesch. der Versamml. Deutsch. Naturf. und ´’arzte ^{≫}, Bd 6 par Irene Scheifele. Revue d’Histoire des Sciences, 51(4):561, October 1998. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633885.

Merleau-Ponty:1998:AOB

[3844] Jacques Merleau-Ponty. Analyses d’ouvrages: Mgr Georges Lemaître, savant et croyant. La Physique d’Einstein, ^{≪}Réminisciences ^{≫}, 3 par Jean-François Stoffel; Georges Lemaître. Revue d’Histoire des Sciences, 51(4):561–562, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633886.

Anonymous:1998:PR

[3845] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 51(4):563– 570, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633887.

Anonymous:1998:TGM

[3846] Anonyme. Table générale des matières du Tome 51. Revue d’Histoire des Sciences, 51(4):571–575, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633888.

Anonymous:1998:BMa

[3847] Anonyme. Back matter. Revue d’Histoire des Sciences, 51(1):??, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633927.

Anonymous:1998:BMb

[3848] Anonyme. Back matter. Revue d’Histoire des Sciences, 51(2–3):??, avril– septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633616.

Anonymous:1998:BMc

[3849] Anonyme. Back matter. Revue d’Histoire des Sciences, 51(4):??, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633889.

Anonymous:1998:FMa

[3850] Anonyme. Front matter. Revue d’Histoire des Sciences, 51(1):??, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633904.

Anonymous:1998:FMb

[3851] Anonyme. Front matter. Revue d’Histoire des Sciences, 51(2–3):??, avril– septembre 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633585.

Anonymous:1998:FMc

[3852] Anonyme. Front matter. Revue d’Histoire des Sciences, 51(4):??, October 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633844.

Bergandi:1999:MOE

[3853] Donato Bergandi. Les métamorphoses de l’organicisme en écologie : de la communauté végétale aux écosystèmes. Revue d’Histoire des Sciences, 52(1):5–31, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633565.

Santesmases:1999:CIB

[3854] María Jesús Santesmases and Emilio Muñoz. La construction institutionnelle de la biochimie espagnole (1945–1970) : Rôle des échanges avec l’Europe du Nord et l’Amérique. Revue d’Histoire des Sciences, 52(1): 33–49, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633566.

Fournier:1999:HOM

[3855] Paul Fournier and Josette Fournier. Hasard ou mémoire dans la découverte de la radioactivité? Revue d’Histoire des Sciences, 52(1):

51–79, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633567.

Sabatier:1999:LSP

[3856] Xavier Sabatier. La logique dans la science : Place et statut de la logique dans la philosophie de Jean Cavaillès. Revue d’Histoire des Sciences, 52(1):81–106, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633568.

Sureson:1999:TEO

[3857] Claude Sureson. Théorie des ensembles ou ensemble de théories? Revue d’Histoire des Sciences, 52(1):107–137, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633569.

Torrini:1999:CGE

[3858] Maurizio Torrini. La correspondance de Galilée entre chronique et histoire des sciences. Revue d’Histoire des Sciences, 52(1):139–154, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633570.

Anonymous:1999:Ia

[3859] Anonyme. Informations. Revue d’Histoire des Sciences, 52(1):155, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633571.

Maio:1999:AOB

[3860] Giovanni Maio. Analyses d’ouvrages: The Codification of medical morality : Historical and philosophical studies of the formalization of Western medical morality in the eighteenth and nineteenth centuries, vol. 2 : Anglo-American medical ethics and medical jurisprudence in the nineteenth century, ^{≪}Philosophy and medicine ^{≫}, 49 par Robert Baker. Revue d’Histoire des Sciences, 52(1):157–159, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633572.

Ramunni:1999:AOB

[3861] G. Ramunni. Analyses d’ouvrages: User-driven Innovation : The world’s first business computer par David Caminer; John Aris; Peter Hermon; Frank Land. Revue d’Histoire des Sciences, 52(1):159–160, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633573.

Dagognet:1999:AOB

[3862] François Dagognet. Analyses d’ouvrages: Formes, figures, réalité, ≪Milieux ^{≫}par Gérard Chazal. Revue d’Histoire des Sciences, 52(1):160– 161, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633574.

Bourret:1999:AOB

[3863] Pascale Bourret. Analyses d’ouvrages: La Matérialité des sciences : Savoir-faire et instruments dans les sciences de la vie. ^{≪}Les Empêcheurs de penser en rond ^{≫}. [1^{re} ed. : The Right Tools for the job : At work in twentieth-century life science] par Adèle Clarke; Joan Fujimura; Françoise Bouillot; Emilie Hermant. Revue d’Histoire des Sciences, 52(1):161–162, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633575.

Lejbowicz:1999:AOB

[3864] Max Lejbowicz. Analyses d’ouvrages: The Measure of reality : Quantification and Western society, 1250–1600 (Cambridge-New York-Melbourne : Cambridge Univ. Press, 1997) par Alfred W. Crosby. Revue d’Histoire des Sciences, 52(1):162–163, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633576.

Walter:1999:AOB

[3865] Scott Walter. Analyses d’ouvrages: The Collected Papers, vol. 6 : The Berlin years : writings, 1914–1917 par Albert Einstein; A. J. Kox; Martin J. Klein; Robert Schulman. Revue d’Histoire des Sciences, 52(1):163– 164, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633577.

Belna:1999:AOB

[3866] Jean-Pierre Belna. Analyses d’ouvrages: From Dedekind to Gödel : Essays on the development of the foundations of mathematics, ^{≪}Synthese Library ^{≫}, vol. 251 par Jaakko Hintikka. Revue d’Histoire des Sciences, 52(1):164–165, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633578.

Gourevitch:1999:AOB

[3867] Danielle Gourevitch. Analyses d’ouvrages: La Science médicale occidentale entre deux renaissances (XII^{e} s.–XV^{e} s.), ^{≪}Collected Studies Series ^{≫}par Danielle Jacquart. Revue d’Histoire des Sciences, 52(1):166, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633579.

Karenberg:1999:AOB

[3868] Axel Karenberg. Analyses d’ouvrages: Charitable Knowledge : Hospital pupils and practitioners in eighteenth-century London, ^{≪}Cambridge History of medicine ^{≫} par Susan C. Lawrence. Revue d’Histoire des Sciences, 52(1):166–167, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633580.

Gaudilliere:1999:AOB

[3869] Jean-Paul Gaudillière. Analyses d’ouvrages: Between bench and bedside : Science, healing, and interleukin-2 in a cancer ward par Ilana Löwy. Revue d’Histoire des Sciences, 52(1):167–169, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633581.

Gattinara:1999:AOB

[3870] Enrico Castelli Gattinara. Analyses d’ouvrages: ^{≪}Tempo ^{≫}e ^{≪}Relatività ≫nella filosofia francese contemporanea, ^{≪}La Ruota ^{≫}, 26 par Silvestro Marcucci. Revue d’Histoire des Sciences, 52(1):169–170, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633582.

Radvanyi:1999:AOL

[3871] Pierre Radvanyi. Analyses d’ouvrages: Lise Meitner: A life in physics, ≪California Studies in the history of science ^{≫}, vol. 13 par Ruth Lewin Sime. Revue d’Histoire des Sciences, 52(1):170–173, January 1999. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633583.

Cassou-Nogues:1999:RHP

[3872] Pierre Cassou-Noguès. Recherches de Husserl pour une philosophie de la géométrie. Revue d’Histoire des Sciences, 52(2):179–206, avril–juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633974.

Weill-Parot:1999:CAF

[3873] Nicolas Weill-Parot. Causalité astrale et ^{≪}science des images ^{≫}au Moyen Âge : Éléments de réflexion. Revue d’Histoire des Sciences, 52(2):207– 240, avril–juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633975.

Balpe:1999:CEE

[3874] Claudette Balpe. Constitution d’un enseignement expérimental : La physique et chimie dans les écoles centrales. Revue d’Histoire des Sciences, 52(2):241–283, avril–juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633976.

Fauche:1999:EDC

[3875] Serge Fauché. Des exercices du corps à la guérison de l’esprit aux XVIII^{e} et XIX^{e} siècles. Revue d’Histoire des Sciences, 52(2):285–305, avril–juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633977.

Vitrac:1999:TMI

[3876] Bernard Vitrac. Le type mathématique de l’idéalité dans la pensée grecque : Sur un ouvrage de Maurice Caveing. Revue d’Histoire des Sciences, 52(2):307–314, avril–juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633978.

Salanskis:1999:BRT

[3877] Jean-Michel Salanskis. Nombres réels et théorie du continu: Sur un ouvrage édité par Philip Ehrlich. Revue d’Histoire des Sciences, 52(2):315– 320, avril–juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633979.

Anonymous:1999:Ib

[3878] Anonyme. Informations. Revue d’Histoire des Sciences, 52(2):321, avril– juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633980.

Pellegrin:1999:AOBa

[3879] Pierre Pellegrin. Analyses d’ouvrages: De la médecine, t. I, liv. I–II. ≪Universités de France ^{≫} par Celse; Guy Serbat. Revue d’Histoire des Sciences, 52(2):323, avril–juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633981.

Breard:1999:AOB

[3880] Andrea Bréard. Analyses d’ouvrages: Chinese Studies in the history and philosophy of science and technology. ^{≪}Boston Studies in the philosophy of science ^{≫}, vol. 179 par Dainian Fan; Robert Cohen; Kathleen Dugan; Mingshan Jiang. Revue d’Histoire des Sciences, 52(2):324–325, avril–juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633982.

Forest:1999:AOB

[3881] Denis Forest. Analyses d’ouvrages: Techniques du corps et traitements de l’esprit aux XVIII^{e} et XIX^{e} siècles, ^{≪}Conversciences ^{≫}Serge Fauché. Revue d’Histoire des Sciences, 52(2):325–326, avril–juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633983.

Caveing:1999:AOB

[3882] Maurice Caveing. Analyses d’ouvrages: L’Organisation des mathématiques grecques de Théétète à Archimède, ^{≪}Problèmes et controverses ^{≫} par Jean-Louis Gardies. Revue d’Histoire des Sciences, 52(2):327–329, avril– juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633984.

Pellegrin:1999:AOBb

[3883] Pierre Pellegrin. Analyses d’ouvrages: Airs, eaux, lieux. ^{≪}Universités de France ^{≫}par Hippocrate; Jacques Jouanna. Revue d’Histoire des Sciences, 52(2):330, avril–juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633985.

Poulle:1999:AOB

[3884] Emmanuel Poulle. Analyses d’ouvrages: Lexique de la langue scientifique (astrologie, mathématiques, médecine …): Matériaux pour le ^{≪}Dictionnaire du moyen français ^{≫} (DMF) par Danielle Jacquart; Claude Thomasset; Sylvie Bazin-Tacchella; Jean-Patrice Boudet; Thérèse Charmasson; Joëlle Ducos; Hervé L’Huillier. Revue d’Histoire des Sciences, 52(2):330–331, avril–juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633986.

Flament:1999:AOBa

[3885] Dominique Flament. Analyses d’ouvrages: Through a reporter’s eyes : The life of Stephan Banach par Roman Kaluza; Ann Kostant; Wojbor Woyczyński. Revue d’Histoire des Sciences, 52(2):332, avril–juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633987.

Flament:1999:AOBb

[3886] Dominique Flament. Analyses d’ouvrages: Hermann Günther Grassmann (1809–1877) : Visionary mathematician, scientist and neohumanist scholar. Papers from a sesquicentennial conference, ^{≪}Boston Studies in the philosophy of science ^{≫}, vol. 187 par Gert Schubring. Revue d’Histoire des Sciences, 52(2):333–334, avril–juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633988.

Sinaceur:1999:I

[3887] Hourya Sinaceur. Introduction. Revue d’Histoire des Sciences, 52(3–4): 339–341, juillet–décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633690.

Chihara:1999:FBR

[3888] Charles Chihara. Frege’s and Bolzano’s rationalist conceptions of arithmetic. Revue d’Histoire des Sciences, 52(3–4):343–361, juillet– décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633691.

Mathématique et logique chez Bolzano.

Proust:1999:BTR

[3889] Joëlle Proust. Bolzano’s theory of representation. Revue d’Histoire des Sciences, 52(3–4):363–383, juillet–décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www. jstor.org/stable/23633692. Mathématique et logique chez Bolzano.

Hafner:1999:BCI

[3890] Johannes Hafner. Bolzano’s criticism of indirect proofs. Revue d’Histoire des Sciences, 52(3–4):385–398, juillet–décembre 1999. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633693. Mathématique et logique

chez Bolzano.

Rusnock:1999:PMB

[3891] Paul Rusnock. Philosophy of mathematics : Bolzano’s responses to Kant and Lagrange. Revue d’Histoire des Sciences, 52(3–4):399– 427, juillet–décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/ 23633694. Mathématique et logique chez Bolzano.

Mancosu:1999:BCM

[3892] Paolo Mancosu. Bolzano and Cournot on mathematical explanation. Revue d’Histoire des Sciences, 52(3–4):429–455, juillet–décembre 1999. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633695. Mathématique et logique

chez Bolzano.

Sinaceur:1999:RMR

[3893] Hourya Sinaceur. Réalisme mathématique, réalisme logique chez Bolzano. Revue d’Histoire des Sciences, 52(3–4):457–477, juillet– décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633696.

Mathématique et logique chez Bolzano.

Sebestik:1999:FVD

[3894] Jan Sebestik. Forme, variation et déductibilité dans la logique de Bolzano. Revue d’Histoire des Sciences, 52(3–4):479–506, juillet– décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633697.

Mathématique et logique chez Bolzano.

Anonymous:1999:Ic

[3895] Anonyme. Informations. Revue d’Histoire des Sciences, 52(3–4):507– 510, juillet–décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633698.

Boulaine:1999:AOB

[3896] Jean Boulaine. Analyses d’ouvrages: Traditions agronomiques européennes: Élaboration et transmission depuis l’Antiquité par Marie-Claire Amouretti; François Sigaut. Revue d’Histoire des Sciences, 52(3– 4):511–512, juillet–décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633699.

Sebestik:1999:AOB

[3897] Jan Sebestik. Analyses d’ouvrages: Paradoxes de l’infini, ^{≪}Sources du savoir ^{≫} par Bernard Bolzano; Hourya Sinaceur. Revue d’Histoire des Sciences, 52(3–4):512–514, juillet–décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633700.

Bitbol:1999:AOB

[3898] Michel Bitbol. Analyses d’ouvrages: Interpreting the quantum world par Jeffrey Bub. Revue d’Histoire des Sciences, 52(3–4):514–515, juillet– décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633701.

Darrigol:1999:AOB

[3899] Olivier Darrigol. Analyses d’ouvrages: J. J. Thomson and the discovery of the electron par Edward A. Davis; Isobel J. Falconer. Revue d’Histoire des Sciences, 52(3–4):515–517, juillet–décembre 1999. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633702.

Belhoste:1999:AOB

[3900] Bruno Belhoste. Analyses d’ouvrages: Lazare Carnot par Jean Dhombres; Nicole Dhombres. Revue d’Histoire des Sciences, 52(3–4):517– 518, juillet–décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633703.

Bange:1999:AOB

[3901] Christian Bange. Analyses d’ouvrages: Sulla fina anatomia degli organi centrali del sistema nervoso. ^{≪}Biblioteca della scienza italiana ^{≫}, X par Camillo Golgi. Revue d’Histoire des Sciences, 52(3–4):518–520, juillet– décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633704.

Crozon:1999:AOB

[3902] Michel Crozon. Analyses d’ouvrages: The Rise of the standard model : Particle physics in the 1960s and 1970s par Lillian Hoddeson; Laurie Brown; Michael Riordan; Max Dresden. Revue d’Histoire des Sciences, 52(3–4):520–522, juillet–décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633705.

Bitbol-Hesperies:1999:AOBa

[3903] Annie Bitbol-Hespériès. Analyses d’ouvrages: The Collected Letters of Antoni Van Leeuwenhoek, vol. XIII : 1700–1701. The complete works of Van Leeuwenhoek, issued and annotated under the auspices of the Leeuwenhoek-Commission of the royal Netherlands Academy of arts and sciences par L. C. Palm. Revue d’Histoire des Sciences, 52(3–4): 522–523, juillet–décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633706.

Bitbol-Hesperies:1999:AOBb

[3904] Annie Bitbol-Hespériès. Analyses d’ouvrages: The Microscope in the Dutch Republic : The shaping of discovery par Edward G. Ruestow. Revue d’Histoire des Sciences, 52(3–4):523–525, juillet–décembre 1999. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633707.

Anonymous:1999:PR

[3905] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 52(3–4): 527–534, juillet–décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633708.

Anonymous:1999:TGM

[3906] Anonyme. Table générale des matières du Tome 52. Revue d’Histoire des Sciences, 52(3–4):535–540, juillet–décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633709.

Anonymous:1999:BMa

[3907] Anonyme. Back matter. Revue d’Histoire des Sciences, 52(1):??, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633584.

Anonymous:1999:BMb

[3908] Anonyme. Back matter. Revue d’Histoire des Sciences, 52(2):??, avril– juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633989.

Anonymous:1999:BMc

[3909] Anonyme. Back matter. Revue d’Histoire des Sciences, 52(3–4): ??, juillet–décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633710.

Anonymous:1999:FMa

[3910] Anonyme. Front matter. Revue d’Histoire des Sciences, 52(1):??, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633564.

Anonymous:1999:FMb

[3911] Anonyme. Front matter. Revue d’Histoire des Sciences, 52(2):??, avril– juin 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633973.

Anonymous:1999:FMc

[3912] Anonyme. Front matter. Revue d’Histoire des Sciences, 52(3–4): ??, juillet–décembre 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633689.

Salomon-Bayet:2000:P

[3913] Claire Salomon-Bayet. Présentation. Revue d’Histoire des Sciences, 53(1):5–8, January 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633642.

Braunstein:2000:CAC

[3914] Jean-François Braunstein. Canguilhem avant Canguilhem. Revue d’Histoire des Sciences, 53(1):9–26, January 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633643.

Lautman:2000:SC

[3915] Jacques Lautman. Un stoïcien chaleureux. Revue d’Histoire des Sciences, 53(1):27–45, January 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633644.

Grene:2000:PSG

[3916] Marjorie Grene. The philosophy of science of Georges Canguilhem : A transatlantic view. Revue d’Histoire des Sciences, 53(1):47–63, January 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633645.

Hodge:2000:CHB

[3917] Jonathan Hodge. Canguilhem and the history of biology. Revue d’Histoire des Sciences, 53(1):65–81, January 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633646.

Morange:2000:GCB

[3918] Michel Morange. Georges Canguilhem et la biologie du XX^{e} siècle. Revue d’Histoire des Sciences, 53(1):83–105, January 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633647.

Fragu:2000:IPO

[3919] Philippe Fragu. Interactions physiologie-outils thérapeutiques dans les constructions physiopathologiques du goitre exophtalmique (1860–1960). Revue d’Histoire des Sciences, 53(1):107–132, January 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633648.

Scheidecker-Chevallier:2000:MAG

[3920] Myriam Scheidecker-Chevallier. Marc-Antoine Gaudin, Alexandre-Édouard Baudrimont, Auguste Laurent et l’approche structurale en chimie. Revue d’Histoire des Sciences, 53(1):133–167, January 2000. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633649.

Pellegrin:2000:AOBa

[3921] Pierre Pellegrin. Analyses d’ouvrages: De morbis acutis et chroniis. ≪Studies in ancient medicine ^{≫}, vol. 12 par Anonymi medici; Ivan Garofalo; Brian Fuchs. Revue d’Histoire des Sciences, 53(1):169, January 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633650.

Duris:2000:AOB

[3922] Pascal Duris. Analyses d’ouvrages: Hommage à Théodore Monod, naturaliste d’exception. ^{≪}Archives ^{≫} par Roland Billard; Isabelle Jarry; Abdou Diouf. Revue d’Histoire des Sciences, 53(1):170, January 2000. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633651.

Perfetti:2000:AOB

[3923] Amalia Perfetti. Analyses d’ouvrages: La Diffusione del copernicanesimo in Italia (1543–1610), ^{≪}Biblioteca di “Nuncius”. Studi e testi ^{≫}, XXI par Massimo Bucciantini; Maurizio Torrini. Revue d’Histoire des Sciences, 53(1):170–172, January 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633652.

Lejbowicz:2000:AOB

[3924] Max Lejbowicz. Analyses d’ouvrages: Éclipses, comètes, autres phénomènes célestes et tremblements de terre au Moyen Âge : Enquête sur six siècles d’historiographie médiévale dans les limites de la Belgique actuelle (600–1200), ^{≪}Travaux de la fac. de philos. et lettres de l’univ. cathol. de Louvain ^{≫}, XXXVIII, section d’histoire, IX par Isabelle Draelants. Revue d’Histoire des Sciences, 53(1):172–173, January 2000. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633653.

Grene:2000:AOB

[3925] Marjorie Grene. Analyses d’ouvrages: Les Modèles du vivant de Descartes à Leibniz, ^{≪}Mathesis ^{≫} par François Duchesneau. Revue d’Histoire des Sciences, 53(1):173–174, January 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633654.

Valentin:2000:AOB

[3926] Michel Valentin. Analyses d’ouvrages: To see with a better eye : A life of R. T. H. Laennec par Jacalyn Duffin. Revue d’Histoire des Sciences, 53(1):174–175, January 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633655.

Debru:2000:AOBa

[3927] Claude Debru. Analyses d’ouvrages: Le Legs de Claude Bernard,^{≪}Penser la médecine ^{≫}par Mirko D. Grmek. Revue d’Histoire des Sciences, 53(1):175–176, January 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633656.

Massignat:2000:GEA

[3928] Corinne Massignat. Gassendi et l’élasticité de l’air: Une étape entre Pascal et la loi de Boyle–Mariotte. Revue d’Histoire des Sciences, 53(2):179– 203, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633991.

Despoix:2000:MDM

[3929] Philippe Despoix. Mesure du monde et représentation européenne au XVIII^{e} siècle : Le programme britannique de détermination de la longitude en mer. Revue d’Histoire des Sciences, 53(2):205–233, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633992.

Kawashima:2000:MLT

[3930] Keiko Kawashima. Madame Lavoisier et la traduction française de l’ “Essay on phlogiston” de Kirwan. Revue d’Histoire des Sciences, 53(2):235–263, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633993.

Fedi:2000:ACT

[3931] Laurent Fedi. Auguste Comte et la technique. Revue d’Histoire des Sciences, 53(2):265–293, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633994.

Herreman:2000:STM

[3932] Alain Herreman. Le sens d’un texte mathématique. (French) [The meaning of mathematical text. Critical review: Catherine Goldstein’s A theorem of Fermat and its readers [Presses Univ. Vincennes, Saint-Denis, 1995; MR1351497 (97c:01020)]]. Revue d’Histoire des Sciences, 53(2):295–301, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633995.

Gispert:2000:CDS

[3933] Hélène Gispert. La capture du social dans les mathématiques et leur histoire. (French) [finding the social in mathematics and its history. Critical review: Catherine Goldstein’s A theorem of Fermat and its reader [Presses Univ. Vincennes, Saint-Denis, 1995; MR1351497 (97c:01020)]]. Revue d’Histoire des Sciences, 53(2):303–306, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633996.

Callens:2000:AOB

[3934] Stéphane Callens. Analyses d’ouvrages: Actualité et universalité de la pensée scientifique d’Adolphe Quetelet. ^{≪}Classe des sciences ^{≫} par Philippe Roberts-Jones. Revue d’Histoire des Sciences, 53(2):307–308, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633997.

Gillispie:2000:AOB

[3935] Charles C. Gillispie. Analyses d’ouvrages: L’Invention scientifique de la Méditerranée : Égypte, Morée, Algérie, ^{≪}Recherches d’histoire et de sciences sociales^{≫}par Marie–Noëlle Bourguet; Bernard Lepetit; Daniel Nordman; Maroula Sinarellis. Revue d’Histoire des Sciences, 53(2):308– 310, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633998.

Fauque:2000:AOBa

[3936] Danielle Fauque. Analyses d’ouvrages: Lavoisier et l’Encyclopédie méthodique: Le manuscrit des régisseurs des Poudres et salpêtres pour le Dictionnaire de l’artillerie (1787), ^{≪}Biblioteca di Nuncius. Studi e testi ≫, XXVIII par Patrice Bret. Revue d’Histoire des Sciences, 53(2):310– 311, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633999.

Fauque:2000:AOBb

[3937] Danielle Fauque. Analyses d’ouvrages: Œuvres de Lavoisier : Correspondance, vol. VI : 1789–1791 par Patrice Bret; Henri Kagan. Revue d’Histoire des Sciences, 53(2):311–312, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634000.

Pariente:2000:AOB

[3938] Jean-Claude Pariente. Analyses d’ouvrages: Bergson et les neurosciences. Actes du colloque international de neurophilosophie (Fac. libre de médecine-Inst. de philosophie), ^{≪}Les Empêcheurs de penser en rond ^{≫}par Philippe Gallois; Gérard Forzy. Revue d’Histoire des Sciences, 53(2):313, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634001.

Barthelemy:2000:AOB

[3939] Georges Barthélemy. Analyses d’ouvrages: Pierre-Simon Laplace, 1749– 1827 : A life in exact science par Charles Coulston Gillispie; Robert Fox; Ivor Grattan-Guinness. Revue d’Histoire des Sciences, 53(2):314, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634002.

Debarbat:2000:AOB

[3940] Suzanne Débarbat. Analyses d’ouvrages: Victorian Telescope Makers : The lives and letters of Thomas and Howard Grubb par Ian S. Glass; Patrick Moore. Revue d’Histoire des Sciences, 53(2):314–315, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634003.

Fauque:2000:AOBc

[3941] Danielle Fauque. Analyses d’ouvrages: Ricerche fisiche intorno alla salubrità dell’aria : Con un articolo di Alessandro Volta sull’eudiometria. ≪Biblioteca della scienza italiana ^{≫}, VI par Marsilio Landriani; Marco Beretta. Revue d’Histoire des Sciences, 53(2):315–316, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23634004.

Balan:2000:AOB

[3942] Bernard Balan. Analyses d’ouvrages: Étienne Geoffroy Saint-Hilaire, 1772–1844 : Un naturaliste visionnaire, ^{≪}Un savant, une époque ^{≫}par Hervé Le Guyader. Revue d’Histoire des Sciences, 53(2):316–317, avril– juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634005.

Chazal:2000:AOB

[3943] Gérard Chazal. Analyses d’ouvrages: Pourquoi la science? Impacts et limites de la recherche, ^{≪}Milieux ^{≫} par Michel Meulders; Marc Crommelinck; Bernard Feltz. Revue d’Histoire des Sciences, 53(2):318, avril– juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634006.

Bitbol:2000:AOB

[3944] Michel Bitbol. Analyses d’ouvrages: The Interpretation of quantum mechanics and the measurement process par Peter Mittelstaedt. Revue d’Histoire des Sciences, 53(2):319–320, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634007.

Debru:2000:AOBb

[3945] Claude Debru. Analyses d’ouvrages: What is life? The next fifty years : Speculations on the future of biology par Michael P. Murphy; Luke A. J. O’Neill. Revue d’Histoire des Sciences, 53(2):320–321, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634008.

Peset:2000:AOB

[3946] José Luis Peset. Analyses d’ouvrages: Un reto para la sociedad madrileña: La epidemia de gripe de 1918–1919. ^{≪}Madrid en el tiempo ≫par María Isabel Porras Gallo. Revue d’Histoire des Sciences, 53 (2):321–322, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634009.

Gohau:2000:AOB

[3947] Gabriel Gohau. Analyses d’ouvrages: When geologists were historians, 1665–1750 par Rhoda Rappaport. Revue d’Histoire des Sciences, 53(2): 322, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634010.

Cotte:2000:AOB

[3948] Michel Cotte. Analyses d’ouvrages: Rhône–Alpes, terre d’industries à la Belle Époque : 1899–1914, ^{≪}Archives économiques du Crédit lyonnais ≫par Henri Morsel. Revue d’Histoire des Sciences, 53(2):323, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634011.

Dupont:2000:AOBa

[3949] Jean-Claude Dupont. Analyses d’ouvrages: 175 ans d’industrie pharmaceutique française : Histoire de Synthélabo par Michèle Ruffat. Revue d’Histoire des Sciences, 53(2):324, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634012.

Acot:2000:AOBa

[3950] Pascal Acot. Analyses d’ouvrages: Imaginaire, science et discipline. ≪Géographies en liberté^{≫}par Olivier Soubeyran; Bernardo Secchi; André Corboz. Revue d’Histoire des Sciences, 53(2):324–325, avril–juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23634013.

Galperin:2000:ADD

[3951] Charles Galperin. Aspects du développement dans la biologie du XX^{e} siècle: Introduction. Revue d’Histoire des Sciences, 53(3–4): 331–338, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633947.

LeGuyader:2000:CPO

[3952] Hervé Le Guyader. Le concept de plan d’organisation : Quelques aspects de son histoire. Revue d’Histoire des Sciences, 53(3–4):339–379, juillet– décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633948.

Schmitt:2000:LRG

[3953] Stéphane Schmitt. L’œuvre de Richard Goldschmidt : Une tentative de synthèse de la génétique, de la biologie du développement et de la théorie de l’évolution autour du concept d’homéose. Revue d’Histoire des Sciences, 53(3–4):381–399, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633949.

Rheinberger:2000:EPE

[3954] Hans-Jörg Rheinberger. Ephestia : Le projet expérimental d’une génétique de la physiologie du développement dans l’œuvre d’Alfred Kühn. Revue d’Histoire des Sciences, 53(3–4):401–446, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633950.

Fischer:2000:EWS

[3955] Jean-Louis Fischer. Étienne Wolff (1904–1996) : Ses débuts, ses cahiers de laboratoire (1932–1938). Revue d’Histoire des Sciences, 53(3–4): 447–474, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633951.

Gayon:2000:CRA

[3956] Jean Gayon. De la croissance relative à l’allométrie (1918–1936). Revue d’Histoire des Sciences, 53(3–4):475–498, juillet–décembre 2000. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633952.

Dupont:2000:RLM

[3957] Jean-Claude Dupont. Rita Levi-Montalcini et les débuts de la neuroembryologie. Revue d’Histoire des Sciences, 53(3–4):499–519, juillet– décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633953.

Morange:2000:CSM

[3958] Michel Morange. Le complexe t de la souris : Un mirage riche d’enseignements. Revue d’Histoire des Sciences, 53(3–4):521–554, juillet– décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633954.

Gilbert:2000:PSN

[3959] Scott F. Gilbert. Paradigm shifts in neural induction. Revue d’Histoire des Sciences, 53(3–4):555–579, juillet–décembre 2000. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633955.

Galperin:2000:EEG

[3960] Charles Galperin. De l’embryologie expérimentale à la génétique du développement: de Hans Spemann à Antonio García-Bellido. Revue d’Histoire des Sciences, 53(3–4):581–616, juillet–décembre 2000. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633956.

Gourevitch:2000:MG

[3961] Danielle Gourevitch. Mirko Grmek (1924–2000). Revue d’Histoire des Sciences, 53(3–4):617–621, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633957.

Pellegrin:2000:AOBb

[3962] Pierre Pellegrin. Analyses d’ouvrages: Alchimistes grecs (Les), t. IV, 1^{re} partie: Zosime de Panopolis: Mémoires authentiques. ^{≪}Universités de France ^{≫}par Michèle Mertens. Revue d’Histoire des Sciences, 53 (3–4):623, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633958.

Dupont:2000:AOBb

[3963] Jean-Claude Dupont. Analyses d’ouvrages: Annales d’histoire et de philosophie du vivant, vol. I. Revue d’Histoire des Sciences, 53(3–4): 623–624, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633959.

Pellegrin:2000:AOBc

[3964] Pierre Pellegrin. Analyses d’ouvrages: Le Corps respirant : La pensée scientifique chez Galien ^{≪}Studies in ancient medicine ^{≫}, vol. 13 par Armelle Debru. Revue d’Histoire des Sciences, 53(3–4):624–625, juillet– décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633960.

Karli:2000:AOB

[3965] Pierre Karli. Analyses d’ouvrages: Essays in the history of the physiological sciences. Proceedings of a network symposium of the European Association for the history of medicine and health held at the univ. Louis Pasteur, Strasbourg, on March 26–27th, 1993, ^{≪}Clio medica ^{≫}, 33 par Claude Debru. Revue d’Histoire des Sciences, 53(3–4):625–626, juillet– décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633961.

Laurent:2000:AOB

[3966] Goulven Laurent. Analyses d’ouvrages: Évolutionnisme et fixisme en France : Histoire d’un combat (1800–1882), ^{≪}CNRS Histoire des sciences ^{≫}par Cédric Grimoult. Revue d’Histoire des Sciences, 53(3–4): 626–628, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633962.

Weiner:2000:AOB

[3967] Dora B. Weiner. Analyses d’ouvrages: Public Health and social justice in the age of Chadwick : Britain, 1800–1854, ^{≪}Cambridge History of medicine ^{≫} par Christopher Hamlin. Revue d’Histoire des Sciences, 53 (3–4):628, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633963.

Tirard:2000:AOB

[3968] Stéphane Tirard. Analyses d’ouvrages: Bichat, la vie et la mort, ≪Philosophies ^{≫}, 111 par Philippe Huneman. Revue d’Histoire des Sciences, 53(3–4):629–631, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633964.

Blondel-Megrelis:2000:AOB

[3969] Marika Blondel-Mégrelis. Analyses d’ouvrages: Pharmaciens au Muséum : Chimistes et naturalistes. ^{≪}Archives ^{≫} par Philippe Jaussaud; Pierre Potier. Revue d’Histoire des Sciences, 53(3–4):631–632, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633965.

Kaeser:2000:AOB

[3970] Marc-Antoine Kaeser. Analyses d’ouvrages: Carl Vogt (1817–1895) : Science, philosophie et politique. Actes du Colloque de mai 1995, ≪Bibliothèque d’histoire des sciences ^{≫}, 1 par Jean-Claude Pont; Danièle Bui; Françoise Dubosson; Jan Lacki. Revue d’Histoire des Sciences, 53(3– 4):632–634, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633966.

Cherni:2000:AOB

[3971] Amor Cherni. Analyses d’ouvrages: Maupertuis : Un savant oublié. ≪L’Amateur averti^{≫}par Michel Valentin; Gabriel Richet; Jacques Ruffié. Revue d’Histoire des Sciences, 53(3–4):634, juillet–décembre 2000. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633967.

Acot:2000:AOBb

[3972] Pascal Acot. Analyses d’ouvrages: La Biosphère. ^{≪}Latitudes^{≫} par Vladimir Vernadsky; Jean-Paul Deléage; The Biosphere par Vladimir Ivanovitch Vernadsky; D. B. Langmuir; Jacques Grinevald. Revue d’Histoire des Sciences, 53(3–4):635–636, juillet–décembre 2000. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633968.

Anonymous:2000:PR

[3973] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 53(3–4): 637–647, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633969.

Anonymous:2000:TGM

[3974] Anonyme. Table générale des matières du Tome 53. Revue d’Histoire des Sciences, 53(3–4):649–654, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633970.

Anonymous:2000:BMa

[3975] Anonyme. Back matter. Revue d’Histoire des Sciences, 53(2):??, avril– juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634014.

Anonymous:2000:BMb

[3976] Anonyme. Back matter. Revue d’Histoire des Sciences, 53(3–4): ??, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633971.

Anonymous:2000:FMa

[3977] Anonyme. Front matter. Revue d’Histoire des Sciences, 53(1):??, January 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633641.

Anonymous:2000:FMb

[3978] Anonyme. Front matter. Revue d’Histoire des Sciences, 53(2):??, avril– juin 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633990.

Anonymous:2000:FMc

[3979] Anonyme. Front matter. Revue d’Histoire des Sciences, 53(3–4): ??, juillet–décembre 2000. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633946.

Guillerme:2001:LSP

[3980] André Guillerme. L’Électricité dans ses premières grandeurs (1760– 1820): Présentation. Revue d’Histoire des Sciences, 54(1):5–9, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633599.

Heilbron:2001:SCA

[3981] John Heilbron. Some connections among the heroes. Revue d’Histoire des Sciences, 54(1):11–28, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633601.

Beretta:2001:NVL

[3982] Marco Beretta. From Nollet to Volta : Lavoisier and electricity. Revue d’Histoire des Sciences, 54(1):29–52, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633602.

Bernardi:2001:CEA

[3983] Walter Bernardi. La controverse sur l’électricité animale dans l’Italie du XVIII^{e} siècle : Galvani, Volta et … d’autres. Revue d’Histoire des Sciences, 54(1):53–70, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633604.

Segala:2001:EAM

[3984] Marco Segala. Électricité animale, magnétisme animal, galvanisme universel : À la recherche de l’identité entre l’homme et la nature. Revue d’Histoire des Sciences, 54(1):71–84, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633606.

Ciardi:2001:FMA

[3985] Marco Ciardi. Forces et molécules : Amedeo Avogadro, l’électricité et l’hypothèse de 1811. Revue d’Histoire des Sciences, 54(1):85–97, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633608.

Roche:2001:EIS

[3986] Daniel Roche. Électricité et institution sociale de la science : Réflexions pour une conclusion. Revue d’Histoire des Sciences, 54(1):99–114, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633610.

Guillerme:2001:ASI

[3987] André Guillerme. Un accès aux sources imprimées de l’histoire de l’électricité et du magnétisme : Le conservatoire numérique du conservatoire national des arts et métiers (http://cnum.cnam.fr). Revue d’Histoire des Sciences, 54(1):115–117, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633612.

Souffrin:2001:AOB

[3988] Pierre Souffrin. Analyses d’ouvrages: Galileo on the world systems : A new abridged translation and guide par Galileo Galilei; Maurice A. Finocchiaro. Revue d’Histoire des Sciences, 54(1):119–122, January 2001. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633614.

Guillerme:2001:AOBa

[3989] André Guillerme. Analyses d’ouvrages: Mine et métal, 1780–1880 : Les non-ferreux et l’industrialisation, ^{≪}Art & Société ^{≫}par Anne-Françoise Garçon. Revue d’Histoire des Sciences, 54(1):122–123, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633617.

Diagne:2001:AOB

[3990] Souleymane Bachir Diagne. Analyses d’ouvrages: George Boole : Selected manuscripts on logic and its philosophy, ^{≪}Science networks. Historical studies ^{≫}, vol. 20 par Ivor Grattan-Guinness; Gérard Bornet. Revue d’Histoire des Sciences, 54(1):123–124, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633619.

Darrigol:2001:AOB

[3991] Olivier Darrigol. Analyses d’ouvrages: The Natural philosophy of James Clerk Maxwell par Peter M. Harman. Revue d’Histoire des Sciences, 54(1):124–125, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633621.

Crozon:2001:AOB

[3992] Michel Crozon. Analyses d’ouvrages: History of CERN, vol. III par John Krige. Revue d’Histoire des Sciences, 54(1):125–126, January 2001. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633623.

Guillerme:2001:AOBb

[3993] André Guillerme. Analyses d’ouvrages: La Bénédiction de Prométhée: Religion et technologie, XIX^{e}–XX^{e} siècle par Michel Lagrée. Revue d’Histoire des Sciences, 54(1):127–128, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633625.

Granada:2001:AOB

[3994] Miguel A. Granada. Analyses d’ouvrages: Matemáticas, cosmología y humanismo en la España del siglo XVI : Los Comentarios al segundo libro de la Historia natural de Plinio de Jerónimo Muñoz, ^{≪}Cuadernos valencianos de hist. de la med. y de la ciencia^{≫}, LIV, serie A (monografías) par Víctor Navarro Brotóns; Enrique Rodríguez Galdeano. Revue d’Histoire des Sciences, 54(1):128–130, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633627.

Simon:2001:AOB

[3995] Gérard Simon. Analyses d’ouvrages: L’hypothèse d’Oxford : Essai sur les origines de la perspective, ^{≪}Sociologies ^{≫}par Dominique Raynaud. Revue d’Histoire des Sciences, 54(1):130–133, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633629.

Poulle:2001:AOB

[3996] Emmanuel Poulle. Analyses d’ouvrages: Historic scientific instruments of the Adler Planetarium & astronomy museum, vol. I : Western astrolabes par Roderick Webster; Marjorie Webster; Sara Schechner Genuth. Revue d’Histoire des Sciences, 54(1):133–136, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633631.

Blay:2001:P

[3997] Michel Blay and Michel Serfati. Présentation. Revue d’Histoire des Sciences, 54(2):139–141, avril–juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633618.

Knobloch:2001:DEC

[3998] Eberhard Knobloch. Déterminants et élimination chez Leibniz. Revue d’Histoire des Sciences, 54(2):143–164, avril–juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633620. Mathématiques et physique leib-

niziennes, 1re Partie.

Serfati:2001:MPS

[3999] Michel Serfati. Mathématiques et pensée symbolique chez Leibniz. Revue d’Histoire des Sciences, 54(2):165–221, avril–juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633622. Mathématiques et physique leib-

niziennes, 1re Partie.

Bouveresse:2001:MLC

[4000] Jacques Bouveresse. Mathématiques et logique chez Leibniz. Revue d’Histoire des Sciences, 54(2):223–246, avril–juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633624. Mathématiques et physique leib-

niziennes, 1re Partie.

Niderst:2001:GRP

[4001] Alain Niderst. La géométrie et les réalités: À propos des “Éléments de la géométrie de l’infini” de Fontenelle. Revue d’Histoire des Sciences, 54(2):247–254, avril–juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633626.

Darrigol:2001:RCO

[4002] Olivier Darrigol. Revue critique. Sur l’ouvrage de Peter Galison, “Image and logic: A material culture of microphysics”. Revue d’Histoire des Sciences, 54(2):255–259, avril–juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633628.

Caveing:2001:AOB

[4003] Maurice Caveing. Analyses d’ouvrages: Sciences exactes et sciences appliquées à Alexandrie: III^{e} siècle av. J.-C. — I^{er} siècle apr. J.-C. Actes du colloque international de Saint-Étienne (6–8 juin, 1996), ^{≪}Centre Jean-Palerne. Mémoires ^{≫}, XVI par Gilbert Argoud; Jean-Yves Guillaumin. Revue d’Histoire des Sciences, 54(2):261–262, avril–juin 2001. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633630.

Castel:2001:AOB

[4004] Pierre-Henri Castel. Analyses d’ouvrages: Clinique de l’écriture : Une histoire du regard médical sur l’écriture, ^{≪}Les Empêcheurs de penser en rond ^{≫}par Philippe Artières. Revue d’Histoire des Sciences, 54(2):262– 263, avril–juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633632.

Fauque:2001:AOBa

[4005] Danielle Fauque and Alain Niderst. Analyses d’ouvrages: La Naissance de la science classique au XVII^{e} siècle, ^{≪}Histoire des sciences ^{≫}, 128 par Michel Blay. Revue d’Histoire des Sciences, 54(2):263–265, avril– juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633633.

Rommevaux:2001:AOB

[4006] Sabine Rommevaux. Analyses d’ouvrages: Johannes de Muris, De arte mensurandi : A geometrical handbook of the fourteenth century, ≪Boethius ^{≫}, Bd. 41 par H. L. L. Busard. Revue d’Histoire des Sciences, 54(2):265–266, avril–juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633635.

Pellegrin:2001:AOB

[4007] Pierre Pellegrin. Analyses d’ouvrages: Hippocrate, t. VIII : Plaies, nature des os, coeur, anatomie. ^{≪}Coll. des universités de France ^{≫}par Hippocrate; Marie-Paule Duminil. Revue d’Histoire des Sciences, 54(2):266, avril–juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633636.

Fauque:2001:AOBb

[4008] Danielle Fauque. Analyses d’ouvrages: Photométrie ou de la mesure et de la gradation de la lumière, des couleurs et de l’ombre (1760). ^{≪}La Philosophie en commun ^{≫}par Jean-Henri Lambert; J. Boye; J. Couty; M. Saillard. Revue d’Histoire des Sciences, 54(2):267–268, avril–juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633637.

Laurent:2001:AOBa

[4009] Roger Laurent. Analyses d’ouvrages: Épures d’architecture: De la coupe des pierres à la géométrie descriptive, XVI^{e}–XIX^{e} siècle, ^{≪}Science networks. Historical studies ^{≫}, SN 21 par Joël Sakarovitch. Revue d’Histoire des Sciences, 54(2):268–269, avril–juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633638.

Malherbe:2001:AOBa

[4010] Michel Malherbe. Analyses d’ouvrages: Objectivity in the making : Francis Bacon and the politics of inquiry par Julie Robin Solomon. Revue d’Histoire des Sciences, 54(2):269–271, avril–juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633639.

Thomas:2001:AOB

[4011] Jean-Paul Thomas. Analyses d’ouvrages: Histoire d’une grande peur: La masturbation, ^{≪}Les Empêcheurs de penser en rond ^{≫}. 1^{re} ed. par Jean Stengers; Anne Van Neck. Revue d’Histoire des Sciences, 54(2):271–272, avril–juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633640.

Parmentier:2001:DIP

[4012] Marc Parmentier. Démonstrations et infiniment petits dans la Quadratura arithmetica de Leibniz. Revue d’Histoire des Sciences, 54(3): 275–289, juillet–septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633712.

Blay:2001:ASF

[4013] Michel Blay. De l’apparition subreptice des futures formules de conservation à l’occasion de l’algorithmisation de la science du mouvement au tournant des XVII^{e} et XVIII^{e} siècles. Revue d’Histoire des Sciences, 54(3):291–301, juillet–septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633713.

Devillairs:2001:LDC

[4014] Laurence Devillairs. L’immutabilité divine comme fondement des lois de la nature chez Descartes et les éléments de la critique leibnizienne. Revue d’Histoire des Sciences, 54(3):303–324, juillet–septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633714.

Simon:2001:OPP

[4015] Gérard Simon. Optique et perspective : Ptolémée, Alhazen, Alberti. Revue d’Histoire des Sciences, 54(3):325–350, juillet–septembre 2001. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633715.

Fauque:2001:DBU

[4016] Danielle Fauque. Du bon usage de l’éloge : Cas de celui de Pierre Bouguer. Revue d’Histoire des Sciences, 54(3):351–382, juillet–septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633716.

Boistel:2001:DIP

[4017] Guy Boistel. Documents inédits des pères jésuites R. J. Boscovich et Esprit Pezenas sur les longitudes en mer. Revue d’Histoire des Sciences, 54(3):383–397, juillet–septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633717.

Fauque:2001:AOBc

[4018] Danielle Fauque. Analyses d’ouvrages: Instruments of science : An historical encyclopedia, ^{≪}Garland encyclopedias in the history of science ^{≫}, 2, ^{≪}Garland reference library of the social sciences ^{≫}, 936 par Robert Bud; Deborah Jean Warner. Revue d’Histoire des Sciences, 54(3): 399–400, juillet–septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633718.

Fragu:2001:AOB

[4019] Philippe Fragu. Analyses d’ouvrages: Histoire de la neurotransmission, ^{≪}Science, histoire et société ^{≫}par Jean-Claude Dupont; Claude Debru. Revue d’Histoire des Sciences, 54(3):400–401, juillet–septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633719.

Gilain:2001:AOB

[4020] Christian Gilain. Analyses d’ouvrages: Mathematics of the 19th century, vol. 3 : Function theory according to Chebyshev, ordinary differential equations, calculus of variations, theory of finite differences par A. N.

Kolmogorov; A. P. Yushkevich; Roger Cooke. Revue d’Histoire des Sciences, 54(3):401–402, juillet–septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633720.

Garcon:2001:AOB

[4021] Anne-Françoise Garçon. Analyses d’ouvrages: Le Chantier du canal de Suez (1859–1869) : Une histoire des pratiques techniques par Nathalie Montel. Revue d’Histoire des Sciences, 54(3):402–403, juillet–septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633721.

Duris:2001:AOB

[4022] Pascal Duris. Analyses d’ouvrages: Un naturaliste à Paris, 1834 : Notes sur mon premier séjour à Paris, en 1834, du 4 septembre au 23 octobre. ^{≪}Sens de l’histoire ^{≫}par Alfred Moquin-Tandon; Jean-Louis Fischer; Isabelle Dussert-Carbonne. Revue d’Histoire des Sciences, 54 (3):404, juillet–septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633722.

Sinaceur:2001:AOB

[4023] Hourya Sinaceur. Analyses d’ouvrages: Les Mathématiques infinitésimales du IX^{e} au XI^{e} siècle, vol. I: Fondateurs et commentateurs: Banu Musa, Ibn Qurra, Ibn Sinan, al-Khazin, al-Quhi, Ibn al-Samh., Ibn Hud (1996), vol. II: Ibn al-Haytham (1993) par Roshdi Rashed. Revue d’Histoire des Sciences, 54(3):405–409, juillet–septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633723.

Engel:2001:AOB

[4024] Pascal Engel. Analyses d’ouvrages: Philosophy and memory traces : Descartes to connectionism par John Sutton. Revue d’Histoire des Sciences, 54(3):409–410, juillet–septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633724.

Malherbe:2001:AOBb

[4025] Michel Malherbe. Analyses d’ouvrages: Francis Bacon, 1^{re} ed. par Perez Zagorin. Revue d’Histoire des Sciences, 54(3):410–411, juillet–septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633725.

Dagron:2001:AOB

[4026] Tristan Dagron. Analyses d’ouvrages: Systèmes de pensée précartésiens. Études d’aprés le colloque international organisé à Haïfa en 1994 par Ilana Zinguer; Heinz Schott; I. Zinguer; H. Schott. Revue d’Histoire des Sciences, 54(3):412, juillet–septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633726.

Sigrist:2001:LCR

[4027] René Sigrist. L’expérimentation comme rhétorique de la preuve : L’exemple du “Traité d’insectologie” de Charles Bonnet. Revue d’Histoire des Sciences, 54(4):419–449, October 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633891.

Denis:2001:PPT

[4028] Gilles Denis. Pratiques paysannes et théories savantes préagronomiques au XVIII^{e} siècle : Le cas des débats sur la transmission des maladies des grains de blé. Revue d’Histoire des Sciences, 54(4):451–494, October 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633892.

Schmitt:2001:TMM

[4029] Stéphane Schmitt. Type et métamorphose dans la morphologie de Goethe, entre classicisme et romantisme. Revue d’Histoire des Sciences, 54(4):495–521, October 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633893.

Matagne:2001:LNC

[4030] Patrick Matagne. Limites naturelles contre limites administratives, ou quand la géographie botanique croise la politique. Revue d’Histoire des Sciences, 54(4):523–541, October 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633894.

Cheung:2001:CPD

[4031] Tobias Cheung. Cuvier et la perfection du parfait. Revue d’Histoire des Sciences, 54(4):543–553, October 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633895.

Gayon:2001:AOB

[4032] Jean Gayon. Analyses d’ouvrages: L’Histoire des sciences ^{≪}Que sais-je? ≫, 3495 par Pascal Acot. Revue d’Histoire des Sciences, 54(4):555–556, October 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633896.

Dupont:2001:AOB

[4033] Jean-Claude Dupont. Analyses d’ouvrages: Leçons sur le corps, le cerveau et l’esprit : Les racines des sciences de la cognition au Collège de France ^{≪}Travaux du Collège de France ^{≫}par Alain Berthoz; Marie– Hélène Grosbras. Revue d’Histoire des Sciences, 54(4):556–557, October 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633897.

Laurent:2001:AOBb

[4034] Goulven Laurent. Analyses d’ouvrages: La Bretagne des savants et des ingénieurs : Le XX^{e} siècle par Jean Dhombres. Revue d’Histoire des Sciences, 54(4):557–558, October 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633898.

Samama:2001:AOBa

[4035] Évelyne Samama. Analyses d’ouvrages: Les Maladies dans l’art antique ^{≪}Penser la médecine ^{≫}par Mirko D. Grmek; Danielle Gourevitch. Revue d’Histoire des Sciences, 54(4):558–559, October 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633899.

Samama:2001:AOBb

[4036] Évelyne Samama. Analyses d’ouvrages: La Chirurgie dans l’Égypte grécoromaine d’après les papyrus littéraires grecs ^{≪}Studies in ancient medicine ^{≫}, vol. 17 par Marie–Hélène Marganne. Revue d’Histoire des Sciences, 54(4):559–560, October 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633900.

Anonymous:2001:PR

[4037] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 54(4):561– 567, October 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633901.

Anonymous:2001:TGM

[4038] Anonyme. Table générale des matières du Tome 54. Revue d’Histoire des Sciences, 54(4):569–574, October 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633902.

Anonymous:2001:BMa

[4039] Anonyme. Back matter. Revue d’Histoire des Sciences, 54(3):??, juillet– septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633727.

Anonymous:2001:BMb

[4040] Anonyme. Back matter. Revue d’Histoire des Sciences, 54(4):??, October 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633903.

Anonymous:2001:FMa

[4041] Anonyme. Front matter. Revue d’Histoire des Sciences, 54(1):??, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633597.

Anonymous:2001:FMb

[4042] Anonyme. Front matter. Revue d’Histoire des Sciences, 54(2):??, avril– juin 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633615.

Anonymous:2001:FMc

[4043] Anonyme. Front matter. Revue d’Histoire des Sciences, 54(3):??, juillet– septembre 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633711.

Anonymous:2001:FMd

[4044] Anonyme. Front matter. Revue d’Histoire des Sciences, 54(4):??, October 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633890.

Strasser:2002:TLM

[4045] Bruno J. Strasser. Totems de laboratoires, microscopes électroniques et réseaux scientifiques : L’émergence de la biologie moléculaire à Genève (1945–1960). Revue d’Histoire des Sciences, 55(1):5–43, January 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633658.

Borella:2002:EEP

[4046] Vincent Borella. Les écrits épistémologiques de Poincaré, obstacles à la diffusion de la relativité? Revue d’Histoire des Sciences, 55(1):45–81, January 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633659.

Sinaceur:2002:MMQ

[4047] Hourya Sinaceur. Modernité mathématique : Quelques invariants épistémologiques. Revue d’Histoire des Sciences, 55(1):83–100, January 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633660.

Hulin:2002:LSN

[4048] Nicole Hulin. L’enseignement des sciences naturelles au XIX^{e} siècle dans ses liens à d’autres disciplines. Revue d’Histoire des Sciences, 55(1):101– 120, January 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633661.

Gilain:2002:AOB

[4049] Christian Gilain. Analyses d’ouvrages: Correspondance entre Henri Poincaré et Gösta Mittag-Leffler (La) : Avec en annexes les lettres échangées par Poincaré avec Fredholm, Gyldén et Phragmén. ≪Publications des archives Henri Poincaré ^{≫}par Philippe Nabonnand. Revue d’Histoire des Sciences, 55(1):121–123, January 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633662.

Belna:2002:AOB

[4050] Jean-Pierre Belna. Analyses d’ouvrages: Idéographie. ^{≪}Bibliothèque des textes philosophiques ^{≫}par Gotlob Frege; Corine Besson; Jonathan Barnes. Revue d’Histoire des Sciences, 55(1):123–124, January 2002. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633663.

Simon:2002:AOB

[4051] Gérard Simon. Analyses d’ouvrages: La Mutation du visible: Essai sur la portée épistémologique des instruments d’optique au XVII^{e} siècle, vol. I: Du Sidereus nuncius de Galilée à la Dioptrique cartésienne, ^{≪}Histoire des sciences. Savoirs et systèmes de pensée ^{≫}par Philippe Hamou. Revue d’Histoire des Sciences, 55(1):124–126, January 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633664.

Parot:2002:AOBa

[4052] Françoise Parot. Analyses d’ouvrages: Histoire des sciences de l’homme (L’) : Trajectoires, enjeux et questions vives, ^{≪}Histoire des sciences humaines ^{≫}par Claude Blanckaert; Loïc Blondiaux; Laurent Loty; Marc Renneville; Nathalie Richard. Revue d’Histoire des Sciences, 55(1):126, January 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633665.

Elie:2002:AOB

[4053] Maurice Élie. Analyses d’ouvrages: Zur Personalisation des Vollzuges der Wissenschaftslehre J. G. Fichtes : Die systematische Funktion des Begriffes ^{≪}Hiatus irrationalis ^{≫}in den Vorlesungen zur Wissenschaftslehre in den Jahren 1804–1805 par Christoph A. Riedel. Revue d’Histoire des Sciences, 55(1):127–132, January 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633666.

Poulle:2002:AOB

[4054] Emmanuel Poulle. Analyses d’ouvrages: Astrolabes, cartes du ciel: Les comprendre et les construire par Jean-Noël Tardy. Revue d’Histoire des Sciences, 55(1):132–133, January 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633667.

Parot:2002:AOBb

[4055] Françoise Parot. Analyses d’ouvrages: Constructing scientific psychology : Karl Lashley’s mind-brain debates, ^{≪}Cambridge studies in the history of psychology ^{≫} par Nadine M. Weidman. Revue d’Histoire des Sciences, 55(1):133–134, January 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633668.

Engel:2002:AOB

[4056] Pascal Engel. Analyses d’ouvrages: Alfred Tarski and the Vienna Circle : Austro-Polish connections in logical empiricism,^{≪}Vienna Circle Institute yearbook ^{≫}, 6 par Jan Wolenski; Eckehart Köhler. Revue d’Histoire des Sciences, 55(1):135, January 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633669.

Perfetti:2002:I

[4057] Amalia Perfetti. Introduction. Revue d’Histoire des Sciences, 55(2):139– 142, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633672.

Kahn:2002:CTA

[4058] Didier Kahn. La condamnation des thèses d’Antoine de Villon et Étienne de Clave contre Aristote, Paracelse et les ^{≪}cabalistes ^{≫}(1624). Revue d’Histoire des Sciences, 55(2):143–198, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633673.

Ponzio:2002:PAG

[4059] Paolo Ponzio. À propos de l’atomisme de Galilée : Questions cosmologiques et problèmes théologiques. Revue d’Histoire des Sciences, 55(2):199–214, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633674.

Perfetti:2002:LAL

[4060] Amalia Perfetti. L’hypothèse atomistique dans “l’autre monde” de Cyrano de Bergerac. Revue d’Histoire des Sciences, 55(2):215–238, avril– juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633675.

Giglioni:2002:FGN

[4061] Guido Giglioni. Francis Glisson’s notion of “confœderatio naturæ” in the context of hylozoistic corpuscularianism. Revue d’Histoire des Sciences, 55(2):239–262, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633676.

Perfetti:2002:SSM

[4062] Amalia Perfetti. La syphilis sive de morbo gallico de Jérôme Fracastor : Un exemple de la diffusion de Lucrèce en Italie dans la première moitié du XVI^{e} siècle. Revue d’Histoire des Sciences, 55(2):263–271, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633677.

Ernst:2002:LRC

[4063] Germana Ernst, Eugenio Canone, and T. Campanella. Une lettre retrouvée : Campanella à peiresc, le 19 juin 1636. Revue d’Histoire des Sciences, 55(2):273–285, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633678.

Matton:2002:NQC

[4064] Sylvain Matton. Note sur quelques critiques oubliées de l’atomisme : À propos de la transsubstantiation eucharistique. Revue d’Histoire des Sciences, 55(2):287–294, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633679.

Gispert:2002:FPS

[4065] Hélène Gispert. ^{≪}Par la science, pour la patrie ^{≫}, l’Association française pour l’avancement des sciences (AFAS), 1872–1914: Un projet d’histoire des sciences et de leur diffusion. Revue d’Histoire des Sciences, 55(2):295– 300, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633680.

Mothu:2002:AOB

[4066] Alain Mothu. Analyses d’ouvrages: La Face cachée des Lumières : Recherches sur les manuscrits philosophiques clandestins de l’âge classique, ^{≪}Bibliographica ^{≫}, vol. 3 par Miguel Benítez. Revue d’Histoire des Sciences, 55(2):301–305, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633681.

Totaro:2002:AOB

[4067] Pina Totaro. Analyses d’ouvrages: Marcello Malpighi : Anatomist and physician, ^{≪}Biblioteca di Nuncius. Studi e testi ^{≫}, XXVII par Domenico Bertoloni Meli. Revue d’Histoire des Sciences, 55(2):305–307, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633682.

Serrapica:2002:AOBa

[4068] Salvatore Serrapica. Analyses d’ouvrages: D’Andrea atomista : L’ ≪Apologia ^{≫}e altri inediti nella polemica filosofica della Napoli di fine Seicento, ^{≪} Quaderni del dipartimento di filosofia e politica, Istituto universitario orientale ^{≫}par Antonio Borrelli. Revue d’Histoire des Sciences, 55(2):307–309, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633683.

vonWille:2002:AOB

[4069] Dagmar von Wille. Analyses d’ouvrages: Poemi filosofici latini : De triplici minimo et mensura — De monade, numero et figura — De innumerabilibus, immenso et infigurabili, ^{≪}Medium coeli ^{≫}, I par Giordano Bruno; Eugenio Canone. Revue d’Histoire des Sciences, 55(2):309–310, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633684.

Serrapica:2002:AOBb

[4070] Salvatore Serrapica. Analyses d’ouvrages: Atomismo e continuo nel XVII secolo. Atti del convegno internazionale ^{≪}Atomisme et continuum au XVII^{e} siècle ^{≫}, ^{≪}Biblioteca europea ^{≫}, 19 par Egidio Festa; Romano Gatto. Revue d’Histoire des Sciences, 55(2):311–312, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633685.

Ingegno:2002:AOB

[4071] Alfonso Ingegno. Analyses d’ouvrages: Géométrie, atomisme et vide dans l’école de Galilée par Egidio Festa; Vincent Jullien; Maurizio Torrini. Revue d’Histoire des Sciences, 55(2):312–318, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633686.

Perfetti:2002:AOB

[4072] Amalia Perfetti. Analyses d’ouvrages: Le Passioni degli atomi : Montanari e Rossetti, una polemica tra galileiani, ^{≪}Biblioteca di Nuncius. Studi e testi ^{≫}, XXIII par Susana Gomez Lopez. Revue d’Histoire des Sciences, 55(2):318–319, avril–juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633687.

Ghesquier:2002:CCD

[4073] Danièle Ghesquier. La centrifugation et la cellule : La déconstruction du protoplasme entre 1880 et 1910. Revue d’Histoire des Sciences, 55(3): 323–377, juillet–septembre 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633848.

Belna:2002:FGP

[4074] Jean-Pierre Belna. Frege et la géométrie projective: la Dissertation inaugurale de 1873. Revue d’Histoire des Sciences, 55(3):379–410, juillet– septembre 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633850.

Bailhache:2002:EDF

[4075] Patrice Bailhache. Un effet Doppler–Fizeau méconnu: Roemer et la vitesse de la lumière. Revue d’Histoire des Sciences, 55(3):411– 430, juillet–septembre 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633852.

Fauque:2002:AOB

[4076] Danielle Fauque. Analyses d’ouvrages: Atti del VII convegno nazionale di storia e fondamenti della chimica. Organizzato da università degli studi dell’Aquila, dipartimento di chimica, ingegneria chimica e materiali; Gruppo nazionale di fondamenti e storia della chimica; Accademia nazionale delle scienze detta dei XL, L’Aquila, 8–11 ottobre 1997. [Extr. du vol. 115 des Memorie di scienze fisiche e naturali, ^{≪}Rendiconti della Acc. naz. delle scienze detta dei XL ^{≫}, série V, vol. XXI, part. II, t. II, 1997 ] par Franco Calascibetta; Atti dell’VIII convegno nazionale di storia e fondamenti della chimica. Organizzato da università degli studi di Siena, facoltà di lettere e filosofia di Arezzo; Gruppo nazionale di fondamenti e storia della chimica; Accademia nazionale delle scienze detta dei XL; biblioteca città di Arezzo, Arezzo, 28–30 ottobre 1999. [Extr. du vol. 117 des Memorie di scienze fisiche e naturali, ^{≪}Rendiconti della Acc. naz. delle scienze detta dei XL ^{≫}, série V, vol. XXIII, part. II, t. II, 1999 ] par Ferdinando Abbri; Marco Ciardi. Revue d’Histoire des Sciences, 55(3):431–433, juillet–septembre 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633854.

Chamak:2002:AOB

[4077] Brigitte Chamak. Analyses d’ouvrages: The Practices of human genetics, ^{≪}Sociology of the sciences ^{≫}, A yearbook, vol. XXI, 1997 par Michael Fortun; Everett Mendelsohn. Revue d’Histoire des Sciences, 55(3): 433–435, juillet–septembre 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633856.

Clavelin:2002:AOB

[4078] Maurice Clavelin. Analyses d’ouvrages: Dialogo: Sopra i due massimi sistemi del mondo tolemaico e copernicano. ^{≪}Medioevo e umanesimo ≫[Vol. I: Testo; vol. II: Commento] par Galileo Galilei; Ottavio Besomi; Mario Helbing. Revue d’Histoire des Sciences, 55(3):435–437, juillet– septembre 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633858.

Bustamante:2002:AOB

[4079] Martha Cecilia Bustamante. Analyses d’ouvrages: Controversy and consensus : Nuclear beta decay 1911–1934. ^{≪}Historical studies. Science networks ^{≫}par Carsten Jensen; Finn Aaserud; Helge Kragh; Erik Rüdinger; Roger H. Stuewer. Revue d’Histoire des Sciences, 55(3):437–438, juillet– septembre 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633860.

Levy:2002:AOB

[4080] Jacques Lévy. Analyses d’ouvrages: Une histoire de la Méridienne : Textes, enjeux, débats et passions autour du Méridien de Paris, 1666– 1827 par Jean-Pierre Martin; Suzanne Débarbat. Revue d’Histoire des Sciences, 55(3):438–439, juillet–septembre 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633862.

Clericuzio:2002:AOB

[4081] Antonio Clericuzio. Analyses d’ouvrages: L’Alchimie à la cour de Côme I^{er} de Médicis: Savoirs, culture et politique ^{≪}Études et essais sur la Renaissance ^{≫}, XVI par Alfredo Perifano. Revue d’Histoire des Sciences, 55(3):439–440, juillet–septembre 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633864.

Granada:2002:AOB

[4082] Miguel A. Granada. Analyses d’ouvrages: The Unmaking of the medieval Christian cosmos, 1500–1760 : From solid heavens to boundless œther par W. G. L. Randles. Revue d’Histoire des Sciences, 55(3): 440–442, juillet–septembre 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633866.

Ratcliff:2002:AOB

[4083] Marc J. Ratcliff. Analyses d’ouvrages: Aspetti poco noti délia storia dell’anatomia patológica ira ’600 e ’700: William Harvey, Marcello Malpighi, Antonio Cocchi, Giovanni Maria Lancisi, Verso Morgagni, ≪Accademia toscana di scienze e lettere “La Colombaria”. Studi^{≫}, CLXI par Giorgio Weber. Revue d’Histoire des Sciences, 55(3):442–443, juillet– septembre 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633868.

Maillard:2002:AOB

[4084] Jean-François Maillard. Analyses d’ouvrages: Opuscule tres-eccelent de la vraye philosophie naturelle des metaulx. ^{≪}Textes et travaux de Chrysopoeia ^{≫}, 6 par D. Zecaire; Renan Crouvizier; Jean-Claude Margolin. Revue d’Histoire des Sciences, 55(3):443–444, juillet–septembre 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633870.

Carvallo:2002:CSH

[4085] Sarah Carvallo. Chimie et scepticisme : Héritage et ruptures d’une science. analyse du “chimiste sceptique”, 1661, Robert Boyle. Revue d’Histoire des Sciences, 55(4):451–492, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633929.

Franckowiak:2002:SNG

[4086] Rémi Franckowiak. Les sels neutres de Guillaume-François Rouelle. Revue d’Histoire des Sciences, 55(4):493–532, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633930.

Savaton:2002:MST

[4087] Pierre Savaton. De la minéralogie aux sciences de la Terre : La géologie dans l’enseignement secondaire de 1795 à 1988. Revue d’Histoire des Sciences, 55(4):533–558, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633931.

Caye:2002:AOB

[4088] Pierre Caye. Analyses d’ouvrages: Leon Battista Alberti : Architettura e cultura. Atti del Convegno internazionale, Mantova, 16–19 novembre 1994, ^{≪}Miscellanea ^{≫}, 7 Accademia nazionale virgiliana di scienze, lettere e arti. Revue d’Histoire des Sciences, 55(4):559–560, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23633932.

Xhayet:2002:AOB

[4089] Geneviève Xhayet. Analyses d’ouvrages: La Draperie au Moyen Âge : Essor d’une grande industrie européenne par Dominique Cardon. Revue d’Histoire des Sciences, 55(4):560–561, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633933.

Gohau:2002:AOBa

[4090] Gabriel Gohau. Analyses d’ouvrages: Manuscripts and publications of Horace–Bénédict de Saussure on the origin of basalt (1772–1797) /Manuscrits et publications de Horace–Bénédict de Saussure sur l’origine du basalte (1772–1797), bilingual volume par Albert V. Carozzi. Revue d’Histoire des Sciences, 55(4):561–563, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633934.

Gohau:2002:AOBb

[4091] Gabriel Gohau. Analyses d’ouvrages: Voyage à l’intérieur de la Terre : De la géographie antique à la géophysique moderne. Une histoire des idées par Vincent Deparis; Hilaire Legros; Jean-Paul Poirier. Revue d’Histoire des Sciences, 55(4):564–565, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633935.

Bensaude-Vincent:2002:AOB

[4092] Bernadette Bensaude-Vincent. Analyses d’ouvrages: Miroir de la chimie, ≪Science ouverte ^{≫}par Pierre Laszlo. Revue d’Histoire des Sciences, 55(4):565–566, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633936.

Dupuy:2002:AOB

[4093] Michel Dupuy. Analyses d’ouvrages: Von der Ökonomie der Natur zur Ökologie : Die Entwicklung ökologischen Denkens und seiner sprachlichen Ausdrucksformen ^{≪}Philologische Studien und Quellen ^{≫}, fasc. 160 par Erwin Morgenthaler. Revue d’Histoire des Sciences, 55(4):566–567, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633937.

Givernaud:2002:AOB

[4094] Nicolas Givernaud. Analyses d’ouvrages: La Fondation Rockefeller et la recherche médicale. ^{≪}Science, histoire et société ^{≫}par Jean-François Picard; Dominique Lecourt. Revue d’Histoire des Sciences, 55(4):568– 570, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633938.

Grene:2002:AOB

[4095] Marjorie Grene. Analyses d’ouvrages: Toward a history of epistemic things: Synthesizing proteins in the test tube, ^{≪}Writing science ^{≫}par Hans-Jörg Rheinberger; Experimentalsysteme und epistemische Dinge: Eine Geschichte der Proteinsynthese im Reagenzglas par Hans-Jörg Rheinberger. Revue d’Histoire des Sciences, 55(4):570–575, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633939.

Knobloch:2002:AOB

[4096] Eberhard Knobloch. Analyses d’ouvrages: La Contre–Réforme mathématique : Constitution et diffusion d’une culture mathématique jésuite à la Renaissance (1540–1640) ^{≪}Bibliothèque des Écoles françaises d’Athènes et de Rome ^{≫}, fasc. 306 par Antonella Romano. Revue d’Histoire des Sciences, 55(4):575–576, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633940.

Biard:2002:AOB

[4097] Joël Biard. Analyses d’ouvrages: Kontingenz und Wissen : Die Lehre von den futura contingentia bei Johannes Duns Scotus, ^{≪}Beiträge zur Gesch. der Phil. und Theol. des Mittelalters ^{≫}, neue Folge, Bd 49 par Joachim Roland Söder. Revue d’Histoire des Sciences, 55(4):577, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633941.

Ferrari:2002:AOB

[4098] Jean Ferrari. Analyses d’ouvrages: Sistema delle scienze naturali e unità della conoscenza nell’ultimo Kant, ^{≪}Biblioteca di storia della scienza ^{≫}, vol. 42 par Paola Vasconi. Revue d’Histoire des Sciences, 55(4):578, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633942.

Anonymous:2002:PR

[4099] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 55(4):579– 584, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633943.

Anonymous:2002:TGM

[4100] Anonyme. Table générale des matières du Tome 55. Revue d’Histoire des Sciences, 55(4):585–591, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633944.

Anonymous:2002:BMa

[4101] Anonyme. Back matter. Revue d’Histoire des Sciences, 55(1):??, January 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633670.

Anonymous:2002:BMb

[4102] Anonyme. Back matter. Revue d’Histoire des Sciences, 55(2):??, avril– juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633688.

Anonymous:2002:BMc

[4103] Anonyme. Back matter. Revue d’Histoire des Sciences, 55(3):??, juillet– septembre 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633872.

Anonymous:2002:BMd

[4104] Anonyme. Back matter. Revue d’Histoire des Sciences, 55(4):??, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633945.

Anonymous:2002:FMa

[4105] Anonyme. Front matter. Revue d’Histoire des Sciences, 55(1):??, January 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633657.

Anonymous:2002:FMb

[4106] Anonyme. Front matter. Revue d’Histoire des Sciences, 55(2):??, avril– juin 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633671.

Anonymous:2002:FMc

[4107] Anonyme. Front matter. Revue d’Histoire des Sciences, 55(3):??, juillet– septembre 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633846.

Anonymous:2002:FMd

[4108] Anonyme. Front matter. Revue d’Histoire des Sciences, 55(4):??, October 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633928.

Blay:2003:AL

[4109] Michel Blay. Avertissement aux lecteurs. Revue d’Histoire des Sciences, 56(1):2, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634521.

Chareix:2003:BAR

[4110] Fabien Chareix. “Experientia ac ratio”: l’œuvre de Christiaan Huygens. Revue d’Histoire des Sciences, 56(1):5–13, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634042. “Expérience et raison”, la science

chez Huygens (1629–1695).

Chareix:2003:DLD

[4111] Fabien Chareix. La découverte des lois du choc par Christiaan Huygens. Revue d’Histoire des Sciences, 56(1):15–58, January 2003. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634043. “Expérience et raison”, la

science chez Huygens (1629–1695).

Charrak:2003:HTM

[4112] André Charrak. Huygens et la théorie musicale. Revue d’Histoire des Sciences, 56(1):59–78, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634044.

Grave:2003:LSH

[4113] Patricia Radelet de Grave. L’Univers selon Huygens, le connu et l’imaginé. Revue d’Histoire des Sciences, 56(1):79–112, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23634045. “Expérience et rai-

son”, la science chez Huygens (1629–1695).

Mormino:2003:RDO

[4114] Gianfranco Mormino. Le rôle de Dieu dans l’œuvre scientifique et philosophique de Christiaan Huygens. Revue d’Histoire des Sciences, 56(1):113–133, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634046.

Yoder:2003:LCH

[4115] Joella G. Yoder. The letters of Christiaan Huygens. Revue d’Histoire des Sciences, 56(1):135–143, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor. org/stable/23634047. “Expérience et raison”, la science chez Huygens (1629–1695).

Mormino:2003:QPE

[4116] Gianfranco Mormino. Sur quelques problèmes éditoriaux concernant l’œuvre de Christiaan Huygens. Revue d’Histoire des Sciences, 56 (1):145–151, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634048.

Mormino:2003:BH

[4117] Gianfranco Mormino and Fabien Chareix. Bibliographie huguenienne. Revue d’Histoire des Sciences, 56(1):153–190, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634049. “Expérience et raison”, la science

chez Huygens (1629–1695).

Veron:2003:LTE

[4118] Philippe Véron. L’équatorial de la tour de l’est de l’observatoire de Paris. Revue d’Histoire des Sciences, 56(1):191–220, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634050.

Boistel:2003:ICO

[4119] Guy Boistel. Inventaire chronologique des œuvres imprimées et manuscrites du père Esprit Pezenas (1692–1776), jésuite, astronome et hydrographe marseillais. Revue d’Histoire des Sciences, 56(1):221–245, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634051.

LeRu:2003:AOB

[4120] Véronique Le Ru. Analyses d’ouvrages: Discours préliminaire de l’Encyclopédie. ^{≪}Textes et commentaires ^{≫}par [Jean Le Rond d’]Alembert; Michel Malherbe. Revue d’Histoire des Sciences, 56(1):247– 248, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634052.

Kahn:2003:AOB

[4121] Didier Kahn. Analyses d’ouvrages: Geheimnisse der Alchemie par Manuel Bachmann; Thomas Hofmeier. Revue d’Histoire des Sciences, 56(1):248–249, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634053.

Souffrin:2003:AOB

[4122] Pierre Souffrin. Analyses d’ouvrages: All’alba della scienza galileiana : Michel Varro e il suo De motu tractatus. Un importante capitolo nella storia della meccanica di fine cinquecento, ^{≪}Storia del pensiero scientifico ^{≫}, 2 par Michele Camerota; Mario Otto Helbing. Revue d’Histoire des Sciences, 56(1):249–252, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634054.

Bitbol-Hesperies:2003:AOBa

[4123] Annie Bitbol-Hespériès. Analyses d’ouvrages: Écrits physiologiques et médicaux. ^{≪}Épiméthée ^{≫} par Descartes; Vincent Aucante. Revue d’Histoire des Sciences, 56(1):252–254, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634055.

Simon:2003:AOB

[4124] Gérard Simon. Analyses d’ouvrages: La Mutation du visible : Essai sur la portée épistémologique des instruments d’optique au XVII^{e} siècle, vol. 2 : Microscopes et télescopes en Angleterre, de Bacon à Hooke, ≪Histoire des sciences ^{≫} par Philippe Hamou. Revue d’Histoire des Sciences, 56(1):255–257, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634056.

Clavelin:2003:AOB

[4125] Maurice Clavelin. Analyses d’ouvrages: Der ungebändigte Galilei : Beiträge zu einem Symposion, ^{≪}Sudhoffs Archiv. Beihefte ^{≫}, 44 par Michael Segre; Eberhard Knobloch. Revue d’Histoire des Sciences, 56(1):257–258, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634057.

Rommevaux:2003:RBE

[4126] Sabine Rommevaux. La réception des “Éléments” d’Euclide au Moyen Âge et à la Renaissance introduction. Revue d’Histoire des Sciences, 56 (2):267–273, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634017.

Vitrac:2003:SGB

[4127] Bernard Vitrac. Les scholies grecques aux “Éléments” d’Euclide. Revue d’Histoire des Sciences, 56(2):275–292, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634018.

Djebbar:2003:QES

[4128] Ahmed Djebbar. Quelques exemples de scholies dans la tradition arabe des “Éléments” d’Euclide. Revue d’Histoire des Sciences, 56(2): 293–321, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634019.

Solere:2003:LAC

[4129] Jean-Luc Solère. L’ordre axiomatique comme modèle d’écriture philosophique dans l’Antiquité et au Moyen Âge. Revue d’Histoire des Sciences, 56(2):323–345, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634020.

Lejbowicz:2003:PTS

[4130] Max Lejbowicz. Le premier témoin scolaire des “Éléments” arabo-latins d’Euclide: Thierry de Chartres et l’ “Heptateuchon”. Revue d’Histoire des Sciences, 56(2):347–368, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634021.

Murdoch:2003:TUE

[4131] John E. Murdoch. Transmission into use: the evidence of marginalia in the medieval Euclides latinus. Revue d’Histoire des Sciences, 56(2): 369–382, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634022.

Biard:2003:MPB

[4132] Joël Biard. Mathématiques et philosophie dans les Questions de Blaise de Parme sur le Traité des rapports de Thomas Bradwardine. Revue d’Histoire des Sciences, 56(2):383–400, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634023.

Rommevaux:2003:LDD

[4133] Sabine Rommevaux. L’irrationalité de la diagonale et du côté d’un même carré dans les Questions de Blaise de Parme sur le Traité des rapports de Bradwardine. Revue d’Histoire des Sciences, 56(2):401–418, juillet– décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634024.

Celeyrette:2003:MDP

[4134] Jean Celeyrette and Edmond Mazet. Le mouvement du point de vue de la cause et le mouvement du point de vue de l’effet dans le Traité des rapports d’Albert de Saxe. Revue d’Histoire des Sciences, 56(2): 419–437, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634025.

Mehl:2003:EFR

[4135] Édouard Mehl. Euclide et la fin de la Renaissance : Sur le scholie de la proposition XIII. 18. Revue d’Histoire des Sciences, 56(2):439– 455, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634026.

Esteve:2003:TEP

[4136] Maria Rosa Massa Esteve. La théorie euclidienne des proportions dans les “Geometriæ speciosæ elementa” (1659) de Pietro Mengoli. Revue d’Histoire des Sciences, 56(2):457–474, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634027.

Mandosio:2003:FMV

[4137] Jean-Marc Mandosio. Des ^{≪}mathématiques vulgaires ^{≫}à la ^{≪}monade hiéroglyphique ^{≫}: Les Éléments d’Euclide vus par John Dee. Revue d’Histoire des Sciences, 56(2):475–491, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634028.

Savoie:2003:PDO

[4138] Denis Savoie. Problèmes de datation d’une occultation observée par Aristote. Revue d’Histoire des Sciences, 56(2):493–504, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634029.

Mascre:2003:AOB

[4139] David Mascré. Analyses d’ouvrages: Cantor, ^{≪}Figures du savoir ^{≫}par Jean-Pierre Belna. Revue d’Histoire des Sciences, 56(2):505–507, juillet– décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634030.

Ramunni:2003:AOB

[4140] Girolamo Ramunni. Analyses d’ouvrages: La Technologie au risque de l’histoire par Robert Belot; Michel Cotte; Pierre Lamard. Revue d’Histoire des Sciences, 56(2):507–509, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634031.

Elie:2003:AOB

[4141] Maurice Élie. Analyses d’ouvrages: Lumières sur les couleurs : Le regard du physicien ^{≪}L’Esprit des sciences ^{≫}, 11 par Michel Blay. Revue d’Histoire des Sciences, 56(2):509–511, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634032.

Balibar:2003:AOB

[4142] Françoise Balibar. Analyses d’ouvrages: Œuvres, XX: La Théorie de la relativité d’Einstein: Éléments pour une théorie de la connaissance. ≪Passages ^{≫}par Ernst Cassirer; Jean Seidengart. Revue d’Histoire des Sciences, 56(2):511–513, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634033.

Bitbol-Hesperies:2003:AOBb

[4143] Annie Bitbol-Hespériès. Analyses d’ouvrages: Spirits and clocks : Machine and organism in Descartes par Dennis Des Chene. Revue d’Histoire des Sciences, 56(2):513–515, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634034.

Brenner:2003:AOB

[4144] Anastasios Brenner. Analyses d’ouvrages: Karl Popper, the formative years, 1902–1945 : Politics and philosophy in interwar Vienna par Haim Malachi Hacohen. Revue d’Histoire des Sciences, 56(2):515–516, juillet– décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634035.

Serfati:2003:AOB

[4145] Michel Serfati. Analyses d’ouvrages: La Topologie et ses signes : Éléments pour une histoire sémiotique des mathématiques par Alain Herreman. Revue d’Histoire des Sciences, 56(2):516–520, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634036.

Pellegrin:2003:AOBa

[4146] Pierre Pellegrin. Analyses d’ouvrages: Hippocrate, t. IV, 3^{e} partie : Épidémies V et VII. ^{≪}Coll. des Universités de France ^{≫}par Hippocrate; Jacques Jouanna; Mirko D. Grmek. Revue d’Histoire des Sciences, 56 (2):520–521, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634037.

Pellegrin:2003:AOBb

[4147] Pierre Pellegrin. Analyses d’ouvrages: Maladies des femmes, t. IV, livre IV. ^{≪}Coll. des Universités de France ^{≫}par Soranos d’Éphèse; Paul Burgière; Danielle Gourevitch; Yves Malinas. Revue d’Histoire des Sciences, 56(2):521, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634038.

Anonymous:2003:TGM

[4148] Anonyme. Table générale des matières du Tome 56. Revue d’Histoire des Sciences, 56(2):523–527, juillet–décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634039.

Anonymous:2003:BMa

[4149] Anonyme. Back matter. Revue d’Histoire des Sciences, 56(1):??, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634058.

Anonymous:2003:BMb

[4150] Anonyme. Back matter. Revue d’Histoire des Sciences, 56(2):??, juillet– décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634040.

Anonymous:2003:FMa

[4151] Anonyme. Front matter. Revue d’Histoire des Sciences, 56(1):??, January 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634041.

Anonymous:2003:FMb

[4152] Anonyme. Front matter. Revue d’Histoire des Sciences, 56(2):??, juillet– décembre 2003. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634016.

Forest:2004:CPH

[4153] Denis Forest. Le concept de proprioception dans l’histoire de la sensibilité interne. Revue d’Histoire des Sciences, 57(1):5–31, January 2004. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634126.

Parot:2004:HIS

[4154] Françoise Parot. Honorer l’incertain : La science positive du XIX^{e} enfante le spiritisme. Revue d’Histoire des Sciences, 57(1):33–63, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634127.

Thomas:2004:NTI

[4155] Marion Thomas. De nouveaux territoires d’introduction du mendélisme en France : Louis Blaringhem (1878–1958), un généticien néolamarckien sur le terrain agricole. Revue d’Histoire des Sciences, 57(1):65–100, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634128.

Chamak:2004:SPO

[4156] Brigitte Chamak. Un scientifique pendant l’occupation : Le cas d’Antoine Lacassagne. Revue d’Histoire des Sciences, 57(1):101–133, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634129.

Barthelemy:2004:NDD

[4157] Georges Barthélemy. Newton découvreur du poids de toutes choses. Revue d’Histoire des Sciences, 57(1):135–159, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634130.

Mandea:2004:GNP

[4158] Mioara Mandea and Pierre-Noël Mayaud. Guillaume Le Nautonier, un précurseur dans l’histoire du géomagnétisme. Revue d’Histoire des Sciences, 57(1):161–173, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634131.

Kahn:2004:ADA

[4159] Didier Kahn. Autour du 500^{e} anniversaire de Paracelse (1993): Quelques ouvrages récents. Revue d’Histoire des Sciences, 57(1):175–188, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634132.

Fauque:2004:PEP

[4160] Danielle Fauque. Un projet européen pour l’histoire de la chimie sous l’égide de l’European Science Foundation (1992–1997). Revue d’Histoire des Sciences, 57(1):189–201, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634133.

Scheidecker-Chevallier:2004:AOB

[4161] Myriam Scheidecker-Chevallier. Analyses d’ouvrages: La Chimica e le tecnologie chimiche nel veneto dell’800. Atti del settimo seminario di storia delle scienze e delle tecniche nell’ottocento veneto, Venezia, 9 e 10 ottobre 1998 (^{≪}Collana dei seminari di storia delle scineze e delle tecniche ≫), 7 par Angelo Bassani. Revue d’Histoire des Sciences, 57(1):203–205, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634134.

Segal:2004:AOB

[4162] Jérôme Segal. Analyses d’ouvrages: Technogenèse de la télévision : Le diable en histoire des machines. ^{≪}Épistémologie et philosophie des sciences^{≫}par Jean Cazenobe. Revue d’Histoire des Sciences, 57(1):205–207, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634135.

Schmitt:2004:AOB

[4163] Stéphane Schmitt. Analyses d’ouvrages: Épistémologie de la transparence : Sur l’embryologie de A. von Haller. ^{≪}Mathesis ^{≫}par Amor Cherni; François Dagognet. Revue d’Histoire des Sciences, 57(1):207– 209, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634136.

Simon:2004:AOB

[4164] Gérard Simon. Analyses d’ouvrages: Une histoire de la science arabe : Introduction à la connaissance du patrimoine scientifique des pays d’Islam. ^{≪}Points-Sciences ^{≫}, 144 par Ahmed Djebbar; Jean Rosmorduc. Revue d’Histoire des Sciences, 57(1):209–210, January 2004. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634137.

Brouzeng:2004:AOB

[4165] Paul Brouzeng. Analyses d’ouvrages: Chimie et industrie en Europe: L’apport des sociétés savantes industrielles du XIX^{e} siècle à nos jours. ≪Histoire des sciences, des techniques et de la médecine ^{≫}par Ulrike Fell. Revue d’Histoire des Sciences, 57(1):210–211, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634138.

Ruelle:2004:AOB

[4166] Charles-Édouard Ruelle. Analyses d’ouvrages: Ainsi s’achèvent les expériences : La place des expériences dans la physique du XX^{e} siècle. ≪Textes à l’appui. Anthropologie des sciences et des techniques ^{≫}. [Éd. originale : How experiments end (1987).] par Peter Galison; Bernard Nicquevert. Revue d’Histoire des Sciences, 57(1):211–213, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23634139.

Moulin:2004:AOB

[4167] Anne-Marie Moulin. Analyses d’ouvrages: The Introduction of modern science and technology to Turkey and Japan par Feza Günergun; Shigehisa Kuriyama. Revue d’Histoire des Sciences, 57(1):213–214, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634140.

Barberousse:2004:AOB

[4168] Anouk Barberousse. Analyses d’ouvrages: Philosophie : Le manuscrit de 1942. ^{≪}Sources du savoir ^{≫}par Werner Heisenberg; Catherine Chevalley. Revue d’Histoire des Sciences, 57(1):214–216, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634141.

Thomas:2004:AOB

[4169] Marion Thomas. Analyses d’ouvrages: Au carrefour de la chimie et de l’agriculture: Les sciences agronomiques en France et en Allemagne, 1840–1914. ^{≪}Histoire des sciences, des techniques et de la médecine ≫par Nathalie Jas; Jean-Luc Mayaud. Revue d’Histoire des Sciences, 57(1):216–218, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634142.

Clavelin:2004:AOB

[4170] Maurice Clavelin. Analyses d’ouvrages: Entrer en matière : Les atomes expliquent-ils le monde? ^{≪}Science ouverte ^{≫}par Pablo Jensen. Revue d’Histoire des Sciences, 57(1):218, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634143.

Grange:2004:AOB

[4171] Juliette Grange. Analyses d’ouvrages: Philosophie des sciences de la nature, ^{≪}L’Interrogation philosophique ^{≫} par Angèle Kremer-Marietti. Revue d’Histoire des Sciences, 57(1):218–220, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634144.

Radvanyi:2004:AOL

[4172] Pierre Radvanyi. Analyses d’ouvrages: Lise Meitner and the dawn of the nuclear age. [1^{re} éd. 1990, en allemand] par Patricia Rife; John Archibald Wheeler. Revue d’Histoire des Sciences, 57(1):220–224, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634145.

Castel:2004:AOB

[4173] Pierre-Henri Castel. Analyses d’ouvrages: Lire le délire : Aliénisme, rhétorique et littérature en France au XIX^{e} siècle par Juan Rigoli. Revue d’Histoire des Sciences, 57(1):224–226, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634146.

Engel:2004:AOB

[4174] Pascal Engel. Analyses d’ouvrages: Le Concept de preuve à la lumière de l’intelligence artificielle, ^{≪}Nouvelle encyclopédie Diderot ^{≫}par Jean Sallantin; Jean-Jacques Szczeciniarz. Revue d’Histoire des Sciences, 57(1):226–227, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634147.

Biard:2004:AOB

[4175] Joël Biard. Analyses d’ouvrages: Kontingenz und Wissen : Die Lehre von den futura contingentia bei Johannes Duns Scotus, ^{≪}Beiträge zur Geschichte der Philosophie und Theologie des Mittelalters ^{≫}, Neue Folge, Bd 49 par Joachim Roland Söder. Revue d’Histoire des Sciences, 57(1):227–228, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634148.

Lefeve:2004:AOB

[4176] Céline Lefeve. Analyses d’ouvrages: Maine de Biran et l’évolutionnisme en psychologie. ^{≪}Les empêcheurs de penser en rond ^{≫}par Gabriel Tarde; Éric Alliez; Anne Devarieux. Revue d’Histoire des Sciences, 57(1):228– 230, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634149.

Nabonnand:2004:AOB

[4177] Philippe Nabonnand. Analyses d’ouvrages: Mathesis : Festschrift zum siebzigsten Geburtstag von Matthias Schramm par Rüdiger Thiele. Revue d’Histoire des Sciences, 57(1):230–231, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634150.

Laurent:2004:AOB

[4178] Goulven Laurent. Analyses d’ouvrages: Smithsonian Institution Secretary Charles Doolittle Walcott par Ellis L. Yochelson. Revue d’Histoire des Sciences, 57(1):231–233, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634151.

Brenner:2004:AOB

[4179] Anastasios Brenner. Analyses d’ouvrages: Essai d’épistémologie réaliste. ≪Mathesis ^{≫}par Élie Zahar; Alain Boyer. Revue d’Histoire des Sciences, 57(1):233–234, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634152.

Anonymous:2004:PRa

[4180] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 57(1):235– 240, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634153.

Armatte:2004:MMN

[4181] Michel Armatte and Amy Dahan Dalmedico. Modèles et modélisations, 1950–2000 : Nouvelles pratiques, nouveaux enjeux. Revue d’Histoire des Sciences, 57(2):243–303 (2005), juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634062.

Bissell:2004:MBB

[4182] Chris Bissell. Models and ^{≪}black boxes ^{≫}: Mathematics as an enabling technology in the history of communications and control engineering. Revue d’Histoire des Sciences, 57(2):305–338, juillet–décembre 2004. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634063.

Morgan:2004:SBT

[4183] Mary S. Morgan. Simulation : The birth of a technology to create ≪evidence ^{≫}in economics. Revue d’Histoire des Sciences, 57(2):339– 375 (2005), juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634064.

Kieken:2004:RMP

[4184] Hubert Kieken. RAINS : Modéliser les pollutions atmosphériques pour la négociation internationale. Revue d’Histoire des Sciences, 57(2):377– 406 (2005), juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634065.

MacKenzie:2004:MMF

[4185] Donald MacKenzie. Models of markets : Finance theory and the historical sociology of arbitrage. Revue d’Histoire des Sciences, 57(2):407– 431 (2005), juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634066.

Armatte:2004:BG

[4186] Michel Armatte and Amy Dahan Dalmedico. Bibliographie générale. Revue d’Histoire des Sciences, 57(2):433–440, juillet–décembre 2004. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634067.

Armatte:2004:RBM

[4187] Michel Armatte and Amy Dahan Dalmedico. Revue bibliographique : Modèles, un panorama réflexif. Revue d’Histoire des Sciences, 57(2): 441–447, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634068.

Nastasi:2004:AOB

[4188] Pietro Nastasi. Analyses d’ouvrages: Changing images in mathematics: From the French Revolution to the new millennium ^{≪}Studies in the history of science, technology, and medicine ^{≫}, 13 par Umberto Bottazzini; Amy Dahan Dalmedico. Revue d’Histoire des Sciences, 57(2): 449–451, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634069.

Aubin:2004:AOB

[4189] David Aubin. Analyses d’ouvrages: Philosophies des mathématiques et de la modélisation : Du chercheur à l’ingénieur par Nicolas Bouleau; La Mathématisation du réel : Essai sur la modélisation mathématique, ≪Science ouverte ^{≫}par Giorgio Israel. Revue d’Histoire des Sciences, 57 (2):451–454, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634070.

Walliser:2004:AOB

[4190] Bernard Walliser. Analyses d’ouvrages: Models as mediators : Perspectives on natural and social science, ^{≪}Ideas in context ^{≫}par Mary S. Morgan; Margaret Morrison. Revue d’Histoire des Sciences, 57(2): 454–455, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634071.

Pinet:2004:PP

[4191] Patrice Pinet. Pasteur et les philosophes. Revue d’Histoire des Sciences, 57(2):457–471, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634072.

Milhaud-Cappe:2004:LHB

[4192] Danielle Milhaud-Cappe and Anastasios Brenner. Lettres d’Henri Bergson à Gaston Milhaud. Revue d’Histoire des Sciences, 57(2):473– 491, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634073.

Fauque:2004:NLA

[4193] Danielle Fauque. De Nicolas Lemery à Adolphe Wurtz : Sur quelques ouvrages d’histoire de la chimie. Revue d’Histoire des Sciences, 57(2): 493–508, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634074.

Wunsch:2004:ECI

[4194] Danielle Wünsch and Emmanuelle Wiss. Einstein et la commission internationale de coopération intellectuelle. Revue d’Histoire des Sciences, 57 (2):509–520, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634075.

Fauque:2004:AOB

[4195] Danielle Fauque. Analyses d’ouvrages: Enseigner la physique au collège et au lycée : Une approche historique, ^{≪}Didact sciences ^{≫}par Claudette Balpe. Revue d’Histoire des Sciences, 57(2):521–523, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634076.

Dorlin:2004:AOB

[4196] Elsa Dorlin. Analyses d’ouvrages: Récit véritable de la naissance de Messeigneurs et Dames les enfants de France et Instruction à ma fille, ≪Textes littéraires français ^{≫}, 530 par Louise Boursier; Bourgeois Louise; épouse Boursier; François Rouget; Colette H. Winn; Lana Caprina : Une controverse médicale sur l’Utérus pensant à l’université de Bologne en 1771–1772 par Giacomo Casanova; Paul Mengal; Roberto Poma. Revue d’Histoire des Sciences, 57(2):523–525, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634077.

Soler:2004:AOB

[4197] Léna Soler. Analyses d’ouvrages: Histories of the electron : The birth of microphysics. ^{≪}Dibner Institute studies in the history of science and technology ^{≫} par Jed Z. Buchwald; Andrew Warwick. Revue d’Histoire des Sciences, 57(2):525–528, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634078.

Vitrac:2004:AOB

[4198] Bernard Vitrac. Analyses d’ouvrages: Johannes de Tinemue’s redaction of Euclid’s “Elements”, the so-called Adelard III version: vol. I : Introduction, sigla and descriptions of the manuscripts, editorial remarks, Euclides, Elementa, vol. II : Conspectus siglorum, apparatus criticus, addenda, ^{≪}Boethius ^{≫}, Bd. 45 par H. L. L. Busard. Revue d’Histoire des Sciences, 57(2):528–529, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634079.

Magnard:2004:AOB

[4199] Pierre Magnard. Analyses d’ouvrages: De l’acédie monastique à l’anxio-dépression : Histoire philosophique de la transformation d’un vice en pathologie, ^{≪}Les empêcheurs de penser en rond ^{≫}par Bernard Forthomme. Revue d’Histoire des Sciences, 57(2):529–530, juillet– décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634080.

Bourel:2004:AOB

[4200] Dominique Bourel. Analyses d’ouvrages: Wissenschaft und Öffentlichkeit in Berlin, 1870–1930 par Constantin Goschler. Revue d’Histoire des Sciences, 57(2):530–531, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634081.

Chaline:2004:AOB

[4201] Jean Chaline. Analyses d’ouvrages: Le Sentier de la guerre : Visages de la violence préhistorique par Jean Guilaine; Jean Zammit. Revue d’Histoire des Sciences, 57(2):531–532, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634082.

Chollet:2004:AOB

[4202] Jacques Chollet and Amor Cherni. Analyses d’ouvrages: Commentarius de formatione cordis in ovo incubato, ^{≪}Studia Halleriana ^{≫}, VI par Albrecht von Haller; Maria Teresa Monti. Revue d’Histoire des Sciences, 57 (2):533–535, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634083.

Gilain:2004:AOB

[4203] Christian Gilain. Analyses d’ouvrages: Mathematics of the 19th century : Mathematical logic, algebra, number theory, probability theory, 2nd revised ed. [Éd. orig. : Izdatel’stvo ^{≪}Nauka ^{≫}, Moskva, 1978.] par A. N. Kolmogorov; A. P. Yushkevich; A. Shenitzer; H. Grant; O. B. Sheinin. Revue d’Histoire des Sciences, 57(2):535, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634084.

Parmentier:2004:AOB

[4204] Marc Parmentier. Analyses d’ouvrages: Hauptschriften zur Versicherungsund Finanzmathematik par Gottfried Wilhelm Leibniz; Eberhard Knobloch; J.-Matthias Graf von der Schulenburg; Ivo Schneider; Edgar Neuburger; Walter Karten; Klaus Luig. Revue d’Histoire des Sciences, 57 (2):536–538, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634085.

Ghermani:2004:AOB

[4205] Laïla Ghermani. Analyses d’ouvrages: Représentations scientifiques et images poétiques en Angleterre au XVII^{e} siècle: À la recherche de l’invisible ^{≪}CNRS Littérature ^{≫} par Margaret Llasera. Revue d’Histoire des Sciences, 57(2):538–539, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634086.

Kottek:2004:AOB

[4206] Samuel S. Kottek. Analyses d’ouvrages: The Tabula antidotarii of Armengaud Blaise and its Hebrew translation, ^{≪}Transactions of the American philosophical society ^{≫}vol. 90 part 6 par Michael McVaugh; Lola Ferre. Revue d’Histoire des Sciences, 57(2):539–540, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634087.

Kleinert:2004:AOB

[4207] Andreas Kleinert. Analyses d’ouvrages: Romanticism in science : Science in Europe 1790–1840, ^{≪}Boston Studies in the philosophy of science ≫vol. 152 par Stefano Poggi; Maurizio Bossi. Revue d’Histoire des Sciences, 57(2):540–541, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634088.

Rignol:2004:AOB

[4208] Loïc Rignol. Analyses d’ouvrages: Le Langage des crânes : Histoire de la phrénologie. ^{≪}Les empêcheurs de penser en rond ^{≫} par Marc Renneville; Georges Lantéri-Laura. Revue d’Histoire des Sciences, 57(2): 541–546, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634089.

Sicard:2004:AOB

[4209] Monique Sicard. Analyses d’ouvrages: Science and the visual image in the Enlightenment, ^{≪}European studies in science history and the arts ^{≫}, IV par William R. Shea. Revue d’Histoire des Sciences, 57(2): 546–548, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634090.

Ramunni:2004:AOB

[4210] Girolamo Ramunni. Analyses d’ouvrages: Technological innovation as an evolutionary process par John Ziman. Revue d’Histoire des Sciences, 57(2):548, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634091.

Anonymous:2004:PRb

[4211] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 57(2): 549–552, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634092.

Anonymous:2004:TGM

[4212] Anonyme. Table générale des matières du Tome 57. Revue d’Histoire des Sciences, 57(2):553–559, juillet–décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634093.

Anonymous:2004:BMa

[4213] Anonyme. Back matter. Revue d’Histoire des Sciences, 57(1):??, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634154.

Anonymous:2004:BMb

[4214] Anonyme. Back matter. Revue d’Histoire des Sciences, 57(2):??, juillet– décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634094.

Anonymous:2004:FMa

[4215] Anonyme. Front matter. Revue d’Histoire des Sciences, 57(1):??, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634125.

Anonymous:2004:FMb

[4216] Anonyme. Front matter. Revue d’Histoire des Sciences, 57(2):??, juillet– décembre 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634061.

Devillairs:2005:LBP

[4217] Laurence Devillairs. La lecture des “Principia philosophiae” de Descartes. Présentation. Revue d’Histoire des Sciences, 58(1):5–7, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634096.

Chevalley:2005:CQP

[4218] Catherine Chevalley. Ce que Pascal doit à la physique des “Principia”. Revue d’Histoire des Sciences, 58(1):9–27, January 2005. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634097.

Carraud:2005:ATP

[4219] Vincent Carraud. Approfondir trop et parler de tout: Les “Principia philosophiae” dans les “Pensées” (note complémentaire sur ≪disproportion de l’homme ^{≫}). Revue d’Histoire des Sciences, 58(1): 29–52, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634098.

Moreau:2005:BS

[4220] Pierre-François Moreau. Les “Principia” de Spinoza. Revue d’Histoire des Sciences, 58(1):53–66, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634099.

Scribano:2005:IIR

[4221] Emanuela Scribano. Invention et interprétation dans la réécriture spinozienne de la métaphysique des “Principia philosophiae”. Revue d’Histoire des Sciences, 58(1):67–85, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634100.

Devillairs:2005:IDM

[4222] Laurence Devillairs. Immutabilité de Dieu et mérites des hommes : À propos du scolie de la proposition XIII de la deuxième partie des “principes de la philosophie de Descartes”. Revue d’Histoire des Sciences, 58(1):87–103, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634101.

deBuzon:2005:ROM

[4223] Frédéric de Buzon. Repos ou mouvement conspirant: Leibniz et les articles 54 et 55 de la partie II des “Principia philosophiae”. Revue d’Histoire des Sciences, 58(1):105–122, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634102.

Belgioioso:2005:RSM

[4224] Giulia Belgioioso. Les racines suaréziennes de la métaphysique cartésienne : La lecture des “Principia” de Francesco Maria Spinelli (1733). Revue d’Histoire des Sciences, 58(1):123–144, January 2005. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634103.

Roveda:2005:LAD

[4225] Lyndia Roveda. Lamarck et l’art des distinctions. Revue d’Histoire des Sciences, 58(1):145–168, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634104.

Aragon:2005:RIS

[4226] Santiago Aragón. Le rayonnement international de la Société zoologique d’acclimatation : Participation de l’Espagne entre 1854 et 1861. Revue d’Histoire des Sciences, 58(1):169–206, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634105.

Gonzalez:2005:VT

[4227] Solange Gonzalez. Varignon et la transsubstantiation. Revue d’Histoire des Sciences, 58(1):207–223, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634106.

Charmasson:2005:AMW

[4228] Thérèse Charmasson, Stéphanie Méchine, and Marc Petit. Archives et manuscrits de Wolfgang Doeblin. Revue d’Histoire des Sciences, 58(1):225–236, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634107.

Parmentier:2005:AOB

[4229] Marc Parmentier. Analyses d’ouvrages: Locke : Idées, langage et connaissance, ^{≪}Philo ^{≫}par Geneviève Brykman. Revue d’Histoire des Sciences, 58(1):237–238, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634108.

Dupont:2005:AOB

[4230] Jean-Claude Dupont. Analyses d’ouvrages: Pharmacopoeias and related literature in Britain and America, 1618–1847, ^{≪}Variorum collected studies series ^{≫}, CS700 par David L. Cowen. Revue d’Histoire des Sciences, 58(1):238–240, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634109.

Marietti:2005:AOB

[4231] Angèle Kremer Marietti. Analyses d’ouvrages: Intrusions spiritualistes et impostures intellectuelles en sciences par Jean Dubessy; Guillaume Lecointre; Jacques Bouveresse. Revue d’Histoire des Sciences, 58(1):240– 242, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634110.

Castelao-Lawless:2005:AOB

[4232] Teresa Castelão-Lawless. Analyses d’ouvrages: A epistemologia de Claude Bernard, ^{≪}Coleçã Cle ^{≫}, 33 par Luiz Henrique; Araujo Dutra. Revue d’Histoire des Sciences, 58(1):242–244, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634111.

Brenner:2005:AOBa

[4233] Anastasios Brenner. Analyses d’ouvrages: Les Philosophies françaises et la science : Dialogue avec Kant par Laurent Fedi; Jean-Michel Salanskis. Revue d’Histoire des Sciences, 58(1):244–245, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634112.

Galonnier:2005:AOB

[4234] Alain Galonnier. Analyses d’ouvrages: Rectitudo Wahrheit und Freiheit bei Anselm von Canterbury : Eine philosophische Untersuchung seines Denkansatzes, ^{≪}Beiträge zur Geschichte der Philosophie und Theologie des Mittelalters ^{≫}, Neue Folge, Bd 56 by Bernd Goebel. Revue d’Histoire des Sciences, 58(1):246–247, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634113.

Schmitt:2005:AOB

[4235] Stéphane Schmitt. Analyses d’ouvrages: Embryos in wax : Models from the Ziegler studio, with a reprint of ^{≪}Embryological wax models ≫ par Nick Hopwood; Friedrich Ziegler. Revue d’Histoire des Sciences, 58(1):248–249, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634114.

Duris:2005:AOB

[4236] Pascal Duris. Analyses d’ouvrages: Les Coquillages de Léonard: Réflexions sur l’histoire naturelle. ^{≪}Science ouverte ^{≫}. [Éd. orig.: Leonardo’s mountain of clams and the diet of worms] par Stephen Jay Gould; Marcel Blanc. Revue d’Histoire des Sciences, 58(1):249–251, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634115.

Sinding:2005:AOB

[4237] Christiane Sinding. Analyses d’ouvrages: L’Avènement de la médecine clinique moderne en Europe, 1750–1815 : Politiques, institutions et savoirs, ^{≪}Bibliothèque d’histoire de la médecine et de la santé ^{≫}par Othmar Keel. Revue d’Histoire des Sciences, 58(1):251–253, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634116.

Serfati:2005:AOB

[4238] Michel Serfati. Analyses d’ouvrages: The Ciphers of the monks: A forgotten number notation of the Middle Ages, ^{≪}Boethius ^{≫}, 44 par David A. King. Revue d’Histoire des Sciences, 58(1):253–255, January 2005. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634117.

Panza:2005:AOB

[4239] Marco Panza. Analyses d’ouvrages: Icona e diagramma : Il segno matematico in Charles Sanders Peirce, ^{≪}Il filarete ^{≫}, n^{o} 203 par Susanna Marietti. Revue d’Histoire des Sciences, 58(1):255–257, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634118.

Bienvenu:2005:AOB

[4240] Alexis Bienvenu. Analyses d’ouvrages: Helmholtz : Des lumières aux neurosciences par Michel Meulders. Revue d’Histoire des Sciences, 58(1):257–259, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634119.

Maury:2005:AOB

[4241] Liliane Maury. Analyses d’ouvrages: De Darwin à Piaget : Pour une histoire de la psychologie de l’enfant,^{≪}Histoire des sciences^{≫}par Dominique Ottavi. Revue d’Histoire des Sciences, 58(1):259–260, January 2005. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634120.

Brenner:2005:AOBb

[4242] Anastasios Brenner. Analyses d’ouvrages: Scientific opportunism / L’Opportunisme scientifique : An anthology, ^{≪}Publications des Archives Henri Poincaré ^{≫} par Henri Poincaré; Louis Rougier; Laurent Rollet. Revue d’Histoire des Sciences, 58(1):261–262, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634121.

Cherni:2005:AOB

[4243] Amor Cherni. Analyses d’ouvrages: Repertorium zu Albrecht von Hallers Korrespondenz, 1724–1777, ^{≪}Studia Halleriana ^{≫}, VII/1, VII/2 par Urs Boschung; Barbara Braun-Bücher; Stefan Hächler; Anne Kathrin Ott; Hubert Steinke; Martin Stuber. Revue d’Histoire des Sciences, 58(1):262– 264, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634122.

Fauque:2005:MRT

[4244] Danielle Fauque. In memoriam René Taton (1915–2004). Revue d’Histoire des Sciences, 58(2):267–303, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634158.

Brenner:2005:P

[4245] Anastasios Brenner. Présentation. Revue d’Histoire des Sciences, 58(2): 305–310, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634159.

Bensaude-Vincent:2005:PLH

[4246] Bernadette Bensaude-Vincent. Paul Langevin : L’histoire des sciences comme remède à tout dogmatisme. Revue d’Histoire des Sciences, 58(2): 311–327, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634160.

Petit:2005:LPA

[4247] Annie Petit. L’enseignement positiviste : Auxiliaire ou obstacle pour l’histoire des sciences? Revue d’Histoire des Sciences, 58(2):329– 365, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634161.

Braunstein:2005:DDH

[4248] Jean-François Braunstein. Daremberg et les débuts de l’histoire de la médecine en France. Revue d’Histoire des Sciences, 58(2):367– 387, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634162.

Hulin:2005:HSE

[4249] Nicole Hulin. Histoire des sciences et enseignement scientifique au lycée sous la Troisième République. Revue d’Histoire des Sciences, 58(2): 389–405, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634163.

Locqueneux:2005:LHP

[4250] Robert Locqueneux. L’intérêt de l’histoire de la physique pour la formation des physiciens selon Henri Bouasse. Revue d’Histoire des Sciences, 58 (2):407–431, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634164.

Brenner:2005:RSL

[4251] Anastasios Brenner. Réconcilier les sciences et les lettres: Le rôle de l’histoire des sciences selon Paul Tannery, Gaston Milhaud et Abel Rey. Revue d’Histoire des Sciences, 58(2):433–454, juillet–décembre 2005. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634165.

Guillo:2005:QCF

[4252] Jean-Marie Guillo and Robert Locqueneux. Quelques contributions de Félix Savary et Jean-Firmin Demonferrand à l’électrodynamique (1822– 1823). Revue d’Histoire des Sciences, 58(2):455–504, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634166.

Blay:2005:EC

[4253] Michel Blay. Ernest Coumet (1933–2003). Revue d’Histoire des Sciences, 58(2):505–507, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634167.

Bontems:2005:AOB

[4254] Vincent Bontems. Analyses d’ouvrages: Philosophy of science : Contemporary readings, ^{≪}Routledge contemporary readings in philosohy ^{≫}par Yuri Balashov; Alex Rosenberg. Revue d’Histoire des Sciences, 58(2): 509–511, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634168.

Fauque:2005:AOBa

[4255] Danielle Fauque. Analyses d’ouvrages: L’Émergence d’une science des manuels: Les livres de chimie en France (1789–1852), ^{≪}Histoire des sciences, des techniques et de la médecine ^{≫}par Bernadette Bensaude-Vincent; Antonio García Belmar; José Ramón Bertomeu Sánchez. Revue d’Histoire des Sciences, 58(2):512–513, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634169.

Swinford:2005:AOB

[4256] Dean Swinford. Analyses d’ouvrages: Imagining the unimaginable : The poetics of early modern astronomy, ^{≪}Internationale Forschungen zur Allgemeinen und Vergleichenden ^{≫}par Ladina Bezzola Lambert. Revue d’Histoire des Sciences, 58(2):513–515, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634170.

Fauque:2005:AOBb

[4257] Danielle Fauque. Analyses d’ouvrages: Des chimistes de A à Z par Éric Brown. Revue d’Histoire des Sciences, 58(2):515–516, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634171.

Clavelin:2005:AOBa

[4258] Maurice Clavelin. Analyses d’ouvrages: Le Mythe Galilée, ^{≪}Science, histoire et société ^{≫} par Fabien Chareix. Revue d’Histoire des Sciences, 58 (2):516–517, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634172.

Fauque:2005:AOBc

[4259] Danielle Fauque. Analyses d’ouvrages: Mémoires de lycées: Archives et patrimoine par Thérèse Charmasson; Armelle Le Goff. Revue d’Histoire des Sciences, 58(2):518, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634173.

Eisenstaedt:2005:AOB

[4260] Jean Eisenstaedt. Analyses d’ouvrages: Pushing gravity : New perspectives on Le Sage’s theory of gravitation par Matthew R. Edwards. Revue d’Histoire des Sciences, 58(2):519–520, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634174.

Dhombres:2005:AOB

[4261] Jean Dhombres. Analyses d’ouvrages: Histoire de l’université de Nantes, 1460–1993 par Gérard Emptoz; Guy Saupin; Anne-Claire Déré; Jean-Paul Molinari. Revue d’Histoire des Sciences, 58(2):521–523, juillet– décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634175.

Raichvarg:2005:AOB

[4262] Daniel Raichvarg. Analyses d’ouvrages: Un pubblico per la scienza : La divulgazione scientifica nell’Italia in formazione par Paola Govoni. Revue d’Histoire des Sciences, 58(2):523–524, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634176.

Brenner:2005:AOBc

[4263] Anastasios Brenner. Analyses d’ouvrages: Historical ontology par Ian Hacking. Revue d’Histoire des Sciences, 58(2):524–525, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634177.

Balibar:2005:AOB

[4264] Françoise Balibar. Analyses d’ouvrages: Les Femmes et l’enseignement scientifique, ^{≪}Science, histoire et société ^{≫}par Nicole Hulin; Bénédicte Bilodeau; Claudine Hermann. Revue d’Histoire des Sciences, 58(2): 525–528, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634178.

Stenger:2005:AOB

[4265] Gerhardt Stenger. Analyses d’ouvrages: Sciences, musiques, lumières : Mélanges offerts à Anne-Marie Chouillet. ^{≪}Publications du Centre international d’étude du XVIIIe siècle ^{≫}, 11 par Ulla Kölving; Irène Passeron.

Revue d’Histoire des Sciences, 58(2):528–529, juillet–décembre 2005. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634179.

Brenner:2005:AOBd

[4266] Anastasios Brenner. Analyses d’ouvrages: Auguste Comte, trajectoires positivistes, 1798–1998, ^{≪}Épistémologie et philosophie des sciences ^{≫}par Annie Petit. Revue d’Histoire des Sciences, 58(2):530–532, juillet– décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634180.

Fauque:2005:AOBd

[4267] Danielle Fauque. Analyses d’ouvrages: Antoine d’Abbadie, ^{≪}Mémoire de la science ^{≫}, 2 par Jean-Paul Poirier; Antony Turner. Revue d’Histoire des Sciences, 58(2):532–533, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634181.

Vieille-Blanchard:2005:AOB

[4268] Élodie Vieille-Blanchard. Analyses d’ouvrages: The Origins of theoretical population genetics, second edition with a new afterword par William B. Provine. Revue d’Histoire des Sciences, 58(2):533–535, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634182.

Laurent:2005:AOB

[4269] Goulven Laurent. Analyses d’ouvrages: The New science of geology : Studies in the earth sciences in the age of Revolution, ^{≪}Variorum collected studies series ^{≫}, 789 par Martin J. S. Rudwick. Revue d’Histoire des Sciences, 58(2):535–537, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634183.

Fauque:2005:AOBe

[4270] Danielle Fauque. Analyses d’ouvrages: Lavoisier in Italia : La communità scientifica italiana e la rivoluzione chimica, ^{≪}Biblioteca di storia della scienza ^{≫} par Raffaella Seligardi. Revue d’Histoire des Sciences, 58 (2):537–539, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634184.

Clavelin:2005:AOBb

[4271] Maurice Clavelin. Analyses d’ouvrages: Domingo de Soto and the early Galileo : Essays on intellectual history, ^{≪}Variorum collected studies series ^{≫}, 783 par William A. Wallace. Revue d’Histoire des Sciences, 58 (2):539–541, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634185.

Anonymous:2005:PR

[4272] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 58(2): 543–546, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634186.

Anonymous:2005:TGM

[4273] Anonyme. Table générale des matières du Tome 58. Revue d’Histoire des Sciences, 58(2):547–552, juillet–décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634187.

Anonymous:2005:BMa

[4274] Anonyme. Back matter. Revue d’Histoire des Sciences, 58(1):??, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634123.

Anonymous:2005:BMb

[4275] Anonyme. Back matter. Revue d’Histoire des Sciences, 58(2):??, juillet– décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634188.

Anonymous:2005:FMa

[4276] Anonyme. Front matter. Revue d’Histoire des Sciences, 58(1):??, January 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634095.

Anonymous:2005:FMb

[4277] Anonyme. Front matter. Revue d’Histoire des Sciences, 58(2):??, juillet– décembre 2005. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634157.

Perru:2006:ORS

[4278] Olivier Perru. Aux origines des recherches sur la symbiose vers 1868– 1883. Revue d’Histoire des Sciences, 59(1):5–27, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634217.

Guedj:2006:DCT

[4279] Muriel Guedj. Du concept de travail vers celui d’énergie : L’apport de Thomson. Revue d’Histoire des Sciences, 59(1):29–50, January 2006. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634218.

LeGars:2006:JRC

[4280] Stéphane Le Gars and Laetitia Maison. Janssen, Rayet, Cornu : Trois parcours exemplaires dans la construction de l’astronomie physique en France (1860–1890). Revue d’Histoire des Sciences, 59(1):51–81, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634219.

Lamy:2006:TUO

[4281] Jérôme Lamy. Transmissions et usages des outils scientifiques : L’observatoire de Toulouse et ses instruments (XVIII^{e}–XIX^{e} siècles). Revue d’Histoire des Sciences, 59(1):83–95, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634220.

Laboulais-Lesage:2006:SEE

[4282] Isabelle Laboulais-Lesage. Les systèmes : Un enjeu épistémologique de la géographie des lumières. Revue d’Histoire des Sciences, 59(1):97–125, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634221.

Pequignot:2006:PED

[4283] Amandine Péquignot. Une peau entre deux feuilles, l’usage de l’^{≪}herbier ≫en taxidermie aux XVIII^{e} et XIX^{e} siècles en France. Revue d’Histoire des Sciences, 59(1):127–136, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634222.

Bontems:2006:LEH

[4284] Vincent Bontems. L’actualité de l’épistémologie historique. Revue d’Histoire des Sciences, 59(1):137–147, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634223.

Pellegrin:2006:AOB

[4285] Pierre Pellegrin. Analyses d’ouvrages: Études de botanique antique. ≪Mémoires de l’Académie des inscriptions et Belles Lettres^{≫}par Suzanne Amigues; Pierre Quézel. Revue d’Histoire des Sciences, 59(1):149, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634224.

Fabre:2006:AOB

[4286] François Fabre. Analyses d’ouvrages: La Sedacina ou l’Œuvre au crible. ≪Textes et travaux de Chrysopoeia ^{≫}, 8, vol. I: L’Alchimie de Guillaume Sedacer, carme catalan de la fin du XIV^{e} siècle: Études et outils, vol. II: L’Alchimie de Guillaume Sedacer, carme catalan de la fin du XIV^{e} siècle: Sedacina, édition critique et traduction, suivie du Liber alterquinus par Pascale Barthélemy; Guy Beaujouan. Revue d’Histoire des Sciences, 59(1):149–151, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634225.

Flament:2006:AOB

[4287] Dominique Flament. Analyses d’ouvrages: Les Constructions des nombres réels dans le mouvement d’arithmétisation de l’analyse, ≪Comprendre les mathématiques par les textes historiques ^{≫} par Jacqueline Boniface. Revue d’Histoire des Sciences, 59(1):151–153, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634226.

Stenger:2006:AOB

[4288] Gerhardt Stenger. Analyses d’ouvrages: Diderot : L’ordre et le devenir, ^{≪}Travaux d’histoire éthico-politique ^{≫}, LIX par Amor Cherni. Revue d’Histoire des Sciences, 59(1):153–155, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634227.

Fauque:2006:AOB

[4289] Danielle Fauque. Analyses d’ouvrages: Les Collèges français, XVI^{e}– XVIII^{e} siècles: Répertoire 3 — Paris par Marie-Madeleine Compère. Revue d’Histoire des Sciences, 59(1):155–156, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634228.

Nicolaidis:2006:AOB

[4290] Efthymios Nicolaidis. Analyses d’ouvrages: Correspondence (The) of John Flamsteed, the first astronomer Royal, (1703–1719), vol. 3 par Flamsteed John; Eric G. Forbes; Maria Forbes; Lesley Murdin; Frances Willmoth. Revue d’Histoire des Sciences, 59(1):156–157, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23634229.

Poulle:2006:AOB

[4291] Emmanuel Poulle. Analyses d’ouvrages: Cueillir les étoiles: Autour des astrolabes de Strasbourg par Francis Debeauvais; Paul-André Béfort; Agnès Acker; William Shea. Revue d’Histoire des Sciences, 59(1):158– 160, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634230.

Laurent:2006:AOB

[4292] Goulven Laurent. Analyses d’ouvrages: Jules Marcou (1824–1898), précurseur français de la géologie nord-américaine. ^{≪}Les Acteurs de la science ^{≫}par Michel Durand-Delga; Richard Moreau; Marie-Claude Fortier. Revue d’Histoire des Sciences, 59(1):160–161, January 2006. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634231.

Sinding:2006:AOB

[4293] Christiane Sinding. Analyses d’ouvrages: Philosophie du progrès en cardiologie par Éric Hamraoui. Revue d’Histoire des Sciences, 59(1):161– 162, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634232.

Pantin:2006:AOB

[4294] Isabelle Pantin. Analyses d’ouvrages: God’s two books : Copernican cosmology and biblical interpretation in early modern science par Kenneth J. Howell. Revue d’Histoire des Sciences, 59(1):163–165, January 2006. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634233.

Locqueneux:2006:AOB

[4295] Robert Locqueneux and Myriam Scheidecker-Chevallier. Analyses d’ouvrages: L’Histoire du concept de ^{≪}molécule ^{≫} par Henk Kubbinga. Revue d’Histoire des Sciences, 59(1):165–167, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634234.

Schmitt:2006:AOBa

[4296] Stéphane Schmitt. Analyses d’ouvrages: Tintinnabulum naturæ : Rêveries d’un individu semi homme semi bête engendré d’une négresse et d’un orang-outang suivi de Pensées métaphisiques lancées dans le tourbillon et de quelques poésies et pièces fugitives, par un solitaire de Champagne. ^{≪}Anecdota. Collection de textes inédits pour servir à l’histoire des idées ^{≫}, 3 par Sylvain Matton; Miguel Benitez; Alain Mothu; Alain Niderst; Charles Porset. Revue d’Histoire des Sciences, 59(1):168–169, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634235.

Chappert:2006:AOB

[4297] André Chappert. Analyses d’ouvrages: Les Ondes en physique, de Pythagore à nos jours : Vibrations, ondes, impulsions par Georges Mourier. Revue d’Histoire des Sciences, 59(1):169–170, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23634236.

Elie:2006:AOB

[4298] Maurice Elie. Analyses d’ouvrages: Philosophie chimique : Hegel et la science dynamiste de son temps,^{≪}Études culturelles^{≫}par Emmanuel Renault. Revue d’Histoire des Sciences, 59(1):171–172, January 2006. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634237.

Caye:2006:P

[4299] Pierre Caye and Thierry Gontier. [Présentation]. Revue d’Histoire des Sciences, 59(2):181–186, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634316.

Napolitani:2006:MPS

[4300] Pier Daniele Napolitani and Jean-Luc Gautero. En mémoire de Pierre Souffrin. Revue d’Histoire des Sciences, 59(2):187–196, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634317.

Furlan:2006:MAD

[4301] Francesco Furlan. In margine all’edizione degli “ex ludis rerum mathematicarum” : Ossia osservazioni e note per l’edizione di un testo scientifico e delle sue figure. (Italian) [on the publication of Ex ludis rerum mathematicarum, and remarks about publishing scientific texts and their illustrations]. Revue d’Histoire des Sciences, 59(2):197–217, juillet– décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634318.

Vengeon:2006:MCH

[4302] Frédéric Vengeon. Mathématiques, création et humanisme chez Nicolas de Cues. Revue d’Histoire des Sciences, 59(2):219–243, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634319.

Caye:2006:SSA

[4303] Pierre Caye. Scientia sine arte nihil est …: Architecture et mathématiques palladiennes II. Revue d’Histoire des Sciences, 59(2): 245–263, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634320.

Angelini:2006:AOD

[4304] Annarita Angelini. ^{≪}Un autre ordre du monde ^{≫}: Science et mathématiques d’après les commentateurs de Proclus au Cinquecento. Revue d’Histoire des Sciences, 59(2):265–283, juillet–décembre 2006. CO-DEN RHSAAM. ISBN 2-200-92108-X. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634321.

Gontier:2006:MSU

[4305] Thierry Gontier. Mathématiques et science universelle chez Bacon et chez Descartes. Revue d’Histoire des Sciences, 59(2):285–312, juillet– décembre 2006. CODEN RHSAAM. ISBN 2-200-92108-X. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634322.

Fredette:2006:PSB

[4306] Raymond Fredette. Pierre Souffrin, bref regard sur sa pratique du métier d’historien des idées scientifiques. Revue d’Histoire des Sciences, 59(2): 313–316, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634323.

Anonymous:2006:NRM

[4307] Anonyme. Un numéro de la revue “Medioevo” à la mémoire de Pierre Souffrin. Revue d’Histoire des Sciences, 59(2):317, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23634324.

Belna:2006:OPD

[4308] Jean-Pierre Belna. Objectivité et principe de dualité : Le paragraphe 26 des Fondements de l’arithmétique de Frege. Revue d’Histoire des Sciences, 59(2):319–344, juillet–décembre 2006. CODEN RHSAAM. ISBN 2-200-92108-X. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634325.

Gillispie:2006:AOB

[4309] Charles C. Gillispie. Analyses d’ouvrages: Montgolfiere, scienze e lumi nel tardo Settecento : Cultura accademica e conoscenze tecniche dalla vigilia della Rivoluzione francese all’età napoleonica, ^{≪}Collana di storia della scienze ^{≫}par Davide Arecco. Revue d’Histoire des Sciences, 59(2): 345–347, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634326.

Barthelemy:2006:AOB

[4310] Georges Barthélemy. Analyses d’ouvrages: Le Problème des objets dans la pensée mathématique, coll. ^{≪}Problèmes et controverses ^{≫}par Maurice Caveing. Revue d’Histoire des Sciences, 59(2):347–349, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634327.

Crozon:2006:AOB

[4311] Michel Crozon. Analyses d’ouvrages: Les Sciences pour la guerre, 1940– 1960, coll. ^{≪}Civilisations et sociétés ^{≫}, 120 par Amy Dahan; Dominique Pestre. Revue d’Histoire des Sciences, 59(2):349–351, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634328.

Bret:2006:AOB

[4312] Patrice Bret. Analyses d’ouvrages: Lavoisier et la naissance de la chimie moderne par Danielle Fauque. Revue d’Histoire des Sciences, 59(2): 351–352, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634329.

Guilhaumou:2006:AOB

[4313] Jacques Guilhaumou. Analyses d’ouvrages: L’Encyclopédie ou la création des disciplines par Martine Groult. Revue d’Histoire des Sciences, 59(2):352–355, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634330.

Brenner:2006:AOB

[4314] Anastasios Brenner. Analyses d’ouvrages: Les Structures rhétoriques de la science : De Kepler à Maxwell, ^{≪}Des travaux ^{≫}par Fernand Hallyn.

Revue d’Histoire des Sciences, 59(2):355–357, juillet–décembre 2006. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634331.

Eisenstaedt:2006:AOB

[4315] Jean Eisenstaedt. Analyses d’ouvrages: The Herschel partnership as viewed par Caroline par Michael Anthony Hoskin; Caroline Herschel’s autobiographies par Michael Anthony Hoskin. Revue d’Histoire des Sciences, 59(2):357–358, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634332.

Joly:2006:AOB

[4316] Bernard Joly. Analyses d’ouvrages: Affinity, that elusive dream : A genealogy of the chemical revolution par Mi Gyung Kim. Revue d’Histoire des Sciences, 59(2):358–360, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634333.

Schmitt:2006:AOBb

[4317] Stéphane Schmitt. Analyses d’ouvrages: Rhetoric and natural history : Buffon in polemical and literary context, ^{≪}Studies on Voltaire and the eighteenth century ^{≫}, 2001/3 par Jeff Loveland. Revue d’Histoire des Sciences, 59(2):361–362, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634334.

Savoie:2006:AOB

[4318] Denis Savoie. Analyses d’ouvrages: Astronomy in Orient and Occident: Selected papers on its cultural and scientific history, ^{≪}Collectanea ^{≫}, vol. XLIV par Yasukatsu Maeyama. Revue d’Histoire des Sciences, 59 (2):363, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634335.

Mazauric:2006:AOB

[4319] Simone Mazauric. Analyses d’ouvrages: Autour de la ^{≪}grande expérience ≫de Pascal : Trois médecins philosophes du XVII^{e} siècle (Pierre Mosnier, G.-B. de Saint-Romain, Guillaume Lamy). ^{≪}Libre pensée et littérature clandestine ^{≫} par Sylvain Matton; Jean Mesnard. Revue d’Histoire des Sciences, 59(2):363–366, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634336.

Dubucs:2006:AOB

[4320] Jacques-Paul Dubucs. Analyses d’ouvrages: New essays on semantic externalism and self-knowledge par Susana Nuccetelli. Revue d’Histoire des Sciences, 59(2):366–368, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634337.

Galuzzi:2006:AOB

[4321] Massimo Galuzzi. Analyses d’ouvrages: Abel’s proof : An essay on the sources and meaning of mathematical unsolvability par Peter Pesic. Revue d’Histoire des Sciences, 59(2):368–369, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634338.

Anonymous:2006:TG

[4322] Anonyme. Table générale. Revue d’Histoire des Sciences, 59(2):371– 379, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634339.

Anonymous:2006:BMa

[4323] Anonyme. Back matter. Revue d’Histoire des Sciences, 59(1):??, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634238.

Anonymous:2006:BMb

[4324] Anonyme. Back matter. Revue d’Histoire des Sciences, 59(2):??, juillet– décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634340.

Anonymous:2006:FMa

[4325] Anonyme. Front matter. Revue d’Histoire des Sciences, 59(1):??, January 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634216.

Anonymous:2006:FMb

[4326] Anonyme. Front matter. Revue d’Histoire des Sciences, 59(2):??, juillet– décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634315.

Joly:2007:P

[4327] Bernard Joly and Sabine Rommevaux. [Présentation]. Revue d’Histoire des Sciences, 60(1):7–8, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634243.

Simon:2007:HOR

[4328] Anne Simon. Histoire de l’optique et recherche littéraire : Le rayon visuel chez proust. Revue d’Histoire des Sciences, 60(1):9–24, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634244.

Hamou:2007:LBG

[4329] Philippe Hamou. L’optique des “Voyages de Gulliver”. Revue d’Histoire des Sciences, 60(1):25–45, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634245.

Rashed:2007:BCU

[4330] Roshdi Rashed. “The Configuration of the universe” : A book par al-Hasan ibn al-Haytham? Revue d’Histoire des Sciences, 60(1):47–63, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634246.

Smith:2007:BAA

[4331] A. Mark Smith. Le “De aspectibus” d’Alhazen : Révolutionnaire ou réformiste? Revue d’Histoire des Sciences, 60(1):65–81, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23634247.

Celeyrette:2007:AIC

[4332] Jean Celeyrette. Apparences et imaginations chez Nicole Oresme : Question III.1 sur la “physique” et question sur l’apparence d’une chose. Revue d’Histoire des Sciences, 60(1):83–100, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634248.

Vescovini:2007:MLP

[4333] Graziella Federici Vescovini. De la métaphysique de la lumière à la physique de la lumière dans la perspective des XIII^{e} et XIV^{e} siècles. Revue d’Histoire des Sciences, 60(1):101–118, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634249.

Blay:2007:VLQ

[4334] Michel Blay. La vue et la lumière : Sur quelques aspects de l’histoire de la lumière. Revue d’Histoire des Sciences, 60(1):119–132, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634250.

Dupont:2007:PVS

[4335] Jean-Claude Dupont. La psychologie de la vision spatiale : Gérard Simon et après. Revue d’Histoire des Sciences, 60(1):133–150, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634251.

Rommevaux:2007:FST

[4336] Sabine Rommevaux. ^{≪}La sphère touche le plan en un point seulement ≫: Un problème mathématique? Revue d’Histoire des Sciences, 60(1): 151–165, January 2007. CODEN RHSAAM. ISBN 2-200-92108-X. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634252.

Joly:2007:PPD

[4337] Bernard Joly. À propos d’une prétendue distinction entre la chimie et l’alchimie au XVII^{e} siècle : Questions d’histoire et de méthode. Revue d’Histoire des Sciences, 60(1):167–183, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634253.

Pourprix:2007:RPA

[4338] Bernard Pourprix. De la reconstitution de la physique allemande du XIX^{e} siècle: Les exemples de Georg Simon Ohm et Hermann Helmholtz. Revue d’Histoire des Sciences, 60(1):185–202, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634254.

Caveing:2007:SSS

[4339] Maurice Caveing. Savoirs et sciences selon Gérard Simon. Revue d’Histoire des Sciences, 60(1):203–216, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634255.

Macherey:2007:HSE

[4340] Pierre Macherey. Histoire des savoirs et épistémologie. Revue d’Histoire des Sciences, 60(1):217–236, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634256.

Simon:2007:PIL

[4341] Gérard Simon. Le problème inverse de Lévi–Strauss. Revue d’Histoire des Sciences, 60(1):237–246, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634257.

Anonymous:2007:BGS

[4342] Anonyme. Bibliographie de Gérard Simon. Revue d’Histoire des Sciences, 60(1):247–250, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634258.

Forest:2007:AOB

[4343] Denis Forest. Analyses d’ouvrages: Le Laboratoire du cerveau psychologique : Histoire et modèles, ^{≪}CNRS histoire des sciences ^{≫}par Bernard Andrieu. Revue d’Histoire des Sciences, 60(1):251–252, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634259.

Clavelin:2007:AOBa

[4344] Maurice Clavelin. Analyses d’ouvrages: Galileo e Keplero : Filosofia, cosmologia e teologia nell’Età della Controriforma, ^{≪}Biblioteca di cultura storica ^{≫}, vol. 243 par Massimo Bucciantini. Revue d’Histoire des Sciences, 60(1):252–254, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634260.

Clavelin:2007:AOBb

[4345] Maurice Clavelin. Analyses d’ouvrages: Galileo Galilei e la cultura scientifica nell’età della Controriforma, ^{≪}Profili ^{≫}, nuova serie, 35 par Michele Camerota. Revue d’Histoire des Sciences, 60(1):254–256, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634261.

Perru:2007:AOB

[4346] Olivier Perru. Analyses d’ouvrages: Expliquer la vie : Modèles, métaphores et machines en biologie du développement, ^{≪}Bibliothèque des sciences humaines ^{≫}par Evelyn Fox Keller; Stéphane Schmitt. Revue d’Histoire des Sciences, 60(1):256–258, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634262.

Lamy:2007:AOB

[4347] Jérôme Lamy. Analyses d’ouvrages: L’Empire du temps: Les horloges d’Einstein et les cartes de Poincaré par Peter Galison. Revue d’Histoire des Sciences, 60(1):258–260, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634263.

Pont:2007:AOB

[4348] Jean-Claude Pont. Analyses d’ouvrages: János Bolyai, non-Euclidean geometry, and the nature of space. With Bolyai’s Appendix, an introduction; the Appendix, a facsimile; and George Bruce Halsted’s translation of The Science absolute of space, ^{≪}Burndy Library Publications, new series ^{≫}, I par Jeremy J. Gray. Revue d’Histoire des Sciences, 60(1):260– 261, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634264.

Locqueneux:2007:AOB

[4349] Robert Locqueneux. Analyses d’ouvrages: Henry Le Chatelier (1850– 1936) ou la science appliquée à l’industrie, ^{≪}Collection Carnot ^{≫}par Michel Letté. Revue d’Histoire des Sciences, 60(1):261–263, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634265.

Parot:2007:AOB

[4350] Françoise Parot. Analyses d’ouvrages: La Naissance de la psychologie, ^{≪}Généalogie de la psychologie ^{≫}par Paul Mengal. Revue d’Histoire des Sciences, 60(1):264–265, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634266.

Clavelin:2007:AOBc

[4351] Maurice Clavelin. Analyses d’ouvrages: Carteggio (1896–1914), ≪Archivio della corrispondenza degli scienziati italiani ^{≫}, 16 par Giuseppe Peano; Louis Couturat; Erika Luciano; Clara Silvia Roero. Revue d’Histoire des Sciences, 60(1):265–266, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634267.

Clavelin:2007:AOBd

[4352] Maurice Clavelin. Analyses d’ouvrages: Sociologie des controverses scientifiques, ^{≪}Sociologies ^{≫}par Dominique Raynaud. Revue d’Histoire des Sciences, 60(1):267–268, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634268.

Dupont:2007:P

[4353] Jean-Claude Dupont and Denis Forest. [Présentation]. Revue d’Histoire des Sciences, 60(2):277–279, juillet–décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634295.

Clauzade:2007:SEE

[4354] Laurent Clauzade. De la science de l’esprit à l’étude du caractère : Alexander Bain et la psychologie des différences individuelles. Revue d’Histoire des Sciences, 60(2):281–301, juillet–décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634296.

Becquemont:2007:RBF

[4355] Daniel Becquemont. Les réticences de Bain face à la théorie des localisations cérébrales de Jackson et de ferrier. Revue d’Histoire des Sciences, 60 (2):303–325, juillet–décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634297.

Dupont:2007:SIP

[4356] Jean-Claude Dupont. Sources et implications physiologiques du discours psychologique chez Bain. Revue d’Histoire des Sciences, 60(2): 327–356, juillet–décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634298.

Forest:2007:BTC

[4357] Denis Forest. Bain et les théories centralistes de l’action et de la conscience d’agir. Revue d’Histoire des Sciences, 60(2):357–374, juillet– décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634299.

Dupont:2007:IBB

[4358] Jean-Claude Dupont and Denis Forest. Indications bio-bibliographiques sur Alexander Bain (1818–1903). Revue d’Histoire des Sciences, 60(2): 375–386, juillet–décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634300.

Baskevitch:2007:LNV

[4359] François Baskevitch. L’élaboration de la notion de vibration sonore : Galilée dans les “discorsi”. Revue d’Histoire des Sciences, 60(2): 387–418, juillet–décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634301.

Delord:2007:AOB

[4360] Julien Delord. Analyses d’ouvrages: Histoire du climat, ^{≪}Pour l’histoire ≫par Pascal Acot. Revue d’Histoire des Sciences, 60(2):419–420, juillet– décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634302.

Schmitt:2007:AOB

[4361] Stéphane Schmitt. Analyses d’ouvrages: Schelling in Rußland : Die frühe naturphilosophischen Schriften von Daniil Michajlovic Vellanskij (1774–1847), ^{≪}Die Deutsche Bibliothek ^{≫}par Thomas Bach. Revue d’Histoire des Sciences, 60(2):421–422, juillet–décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634303.

Laboulais:2007:AOB

[4362] Isabelle Laboulais. Analyses d’ouvrages: Les Instructions scientifiques pour les voyageurs (XVII^{e}–XIX^{e} siècle), ^{≪}Histoire des sciences humaines ≫ par Silvia Collini; Antonella Vannoni. Revue d’Histoire des Sciences, 60 (2):422–424, juillet–décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634304.

Perrin:2007:AOB

[4363] Michel Jean-Louis Perrin. Analyses d’ouvrages: Signe et prédiction dans l’Antiquité. Actes du colloque international interdisciplinaire de Créteil et de Paris, 22, 23 et 24 mai 2003, ^{≪}Centre Jean-Palerne. Mémoires ≫, XXIX par José Kany-Turpin. Revue d’Histoire des Sciences, 60(2): 424–425, juillet–décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634305.

Simon:2007:AOB

[4364] Jonathan Simon. Analyses d’ouvrages: Bacchic medicine : Wine and alcohol therapies from Napoleon to the French paradox, ^{≪}Clio medica ^{≫}, 64 par Harry W. Paul. Revue d’Histoire des Sciences, 60(2):425–427, juillet– décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634306.

Bourdeau:2007:AOB

[4365] Michel Bourdeau. Analyses d’ouvrages: How the cold war transformed philosophy of science : To the icy slopes of logic par George A. Reisch. Revue d’Histoire des Sciences, 60(2):427–430, juillet–décembre 2007. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634307.

Maury:2007:AOB

[4366] Liliane Maury. Analyses d’ouvrages: Ribot, Lévy-Bruhl, Durkheim, numéro spécial de la Revue philosophique de la France et de l’étranger, CXCV/3 (juillet–septembre 2005). Revue d’Histoire des Sciences, 60(2): 431–432, juillet–décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634308.

Chamayou:2007:AOB

[4367] Grégoire Chamayou. Analyses d’ouvrages: L’Invention dans les techniques: Cours et conférences, ^{≪}Traces écrites ^{≫}par Gilbert Simondon; Jean-Yves Château. Revue d’Histoire des Sciences, 60(2):432– 434, juillet–décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634309.

Clavelin:2007:AOBe

[4368] Maurice Clavelin. Analyses d’ouvrages: Galilée, Le ^{≪}Dialogue sur les deux grands systèmes du monde ^{≫}: Rhétorique, dialectique et démonstration, ^{≪}Philosophies ^{≫}par Marta Spranzi. Revue d’Histoire des Sciences, 60(2):434–436, juillet–décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634310.

Dupont:2007:AOB

[4369] Jean-Claude Dupont. Analyses d’ouvrages: Les Sciences de l’âme, XVI^{e}– XVIII^{e} siècle, ^{≪}Les dix-huitièmes siècles ^{≫}, 95 par Fernando Vidal. Revue d’Histoire des Sciences, 60(2):436–438, juillet–décembre 2007. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634311.

Anonymous:2007:PR

[4370] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 60(2): 439–447, juillet–décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634312.

Anonymous:2007:TG

[4371] Anonyme. Table générale. Revue d’Histoire des Sciences, 60(2):449– 456, juillet–décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634313.

Anonymous:2007:BMa

[4372] Anonyme. Back matter. Revue d’Histoire des Sciences, 60(1):??, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634269.

Anonymous:2007:BMb

[4373] Anonyme. Back matter. Revue d’Histoire des Sciences, 60(2):??, juillet– décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634314.

Anonymous:2007:FMa

[4374] Anonyme. Front matter. Revue d’Histoire des Sciences, 60(1):??, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634242.

Anonymous:2007:FMb

[4375] Anonyme. Front matter. Revue d’Histoire des Sciences, 60(2):??, juillet– décembre 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634294.

Anonymous:2008:Ia

[4376] Anonyme. [illustration]. Revue d’Histoire des Sciences, 61(1):2, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634192.

Chapront-Touze:2008:I

[4377] Michelle Chapront-Touzé and Pierre Crépel. [introduction]. Revue d’Histoire des Sciences, 61(1):7–23, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634193.

Dumont:2008:FST

[4378] Simone Dumont and Suzanne Débarbat. Fouchy et ses travaux en astronomie. Revue d’Histoire des Sciences, 61(1):25–39, January 2008. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634194.

Savoie:2008:LGO

[4379] Denis Savoie. L’aspect gnomonique de l’œuvre de Fouchy : La méridienne de temps moyen. Revue d’Histoire des Sciences, 61(1):41–61, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634195.

Fauque:2008:NIP

[4380] Danielle Fauque. Un nouvel instrument pour observer les hauteurs, inventé par M. Grandjean de Fouchy. Revue d’Histoire des Sciences, 61(1): 63–88, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634196.

Chapront-Touze:2008:AOV

[4381] Michelle Chapront-Touzé. Aspects de l’œuvre et de la vie de Pierre-Charles Le Monnier, astronome et académicien, collègue de Grandjean de Fouchy. Revue d’Histoire des Sciences, 61(1):89–103, January 2008. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634197.

Sigrist:2008:QAD

[4382] René Sigrist. Quand l’astronomie devint un métier : Grandjean de Fouchy, Jean III, Bernoulli et la ^{≪}république astronomique ^{≫}, 1700– 1830. Revue d’Histoire des Sciences, 61(1):105–132, January 2008. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634198.

Candaux:2008:GFC

[4383] Jean-Daniel Candaux. Grandjean de Fouchy en correspondance académique avec deux savants genevois. Revue d’Histoire des Sciences, 61(1):133–146, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634199.

Mazauric:2008:PFG

[4384] Simone Mazauric. Parallèle de Fontenelle et de Grandjean de Fouchy. Revue d’Histoire des Sciences, 61(1):147–163, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634200.

Passeron:2008:GFD

[4385] Irène Passeron. Grandjean de Fouchy, D’Alembert et Condorcet : Tracasseries et arrangements des secrétaires perpétuels. Revue d’Histoire des Sciences, 61(1):165–180, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634201.

Crepel:2008:PGF

[4386] Pierre Crépel. Le projet de Grandjean de Fouchy sur l’histoire de l’Académie des sciences. Revue d’Histoire des Sciences, 61(1):181–196, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634202.

Crepel:2008:LJP

[4387] Pierre Crépel, Nicole Dyonet, and Michelle Chapront-Touzé. Une lettre de Jean-Paul Grandjean de Fouchy à Daniel Jousse. Revue d’Histoire des Sciences, 61(1):197–201, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634203.

Courcelle:2008:GFS

[4388] Olivier Courcelle. Grandjean de Fouchy et la Société des arts à Stockholm. Revue d’Histoire des Sciences, 61(1):203–204, January 2008. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634204.

Jacob:2008:FLS

[4389] Marie Jacob. Le fonctionnement de l’Académie des sciences pendant le secrétariat de Fouchy. Revue d’Histoire des Sciences, 61(1):205–210, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634205.

Bensaude-Vincent:2008:AOB

[4390] Bernadette Bensaude-Vincent. Analyses d’ouvrages: L’Encyclopédie méthodique (1782–1832) : Des Lumières au positivisme, ^{≪}Bibliothèque des Lumières ^{≫}, vol. LXVIII par Claude Blanckaert; Michel Porret; Fabrice Brandli. Revue d’Histoire des Sciences, 61(1):211–213, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23634206.

Rommevaux:2008:AOB

[4391] Sabine Rommevaux. Analyses d’ouvrages: Campanus of Novara and Euclid’s “Elements”, ^{≪}Boethius ^{≫}, Band 51 par Hubert L. L. Busard. Revue d’Histoire des Sciences, 61(1):213–215, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634207.

Lamy:2008:AOBa

[4392] Jérôme Lamy. Analyses d’ouvrages: La Grande lunette de Meudon : Les yeux de la découverte, ^{≪}CNRS histoire des sciences ^{≫}par Audouin Dollfus; Paul Couteau; Daniel Egret. Revue d’Histoire des Sciences, 61(1):215–217, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634208.

Dhombres:2008:AOB

[4393] Jean Dhombres. Analyses d’ouvrages: Les Racines sociales et économiques des “Principia” de Newton : Une rencontre entre Newton et Marx à Londres en 1931 par Boris Hessen; Serge Guérout; Christopher A. J. Chilvers. Revue d’Histoire des Sciences, 61(1):217–218, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634209.

Savaton:2008:AOB

[4394] Pierre Savaton. Analyses d’ouvrages: Le Concept de semence dans les théories de la matière à la Renaissance : De Marsile Ficin à Pierre Gassendi, ^{≪}de diversis artibus ^{≫}, t. 72 (N. S. 35) par Hiro Hirai. Revue d’Histoire des Sciences, 61(1):218–220, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634210.

Thomas:2008:AOB

[4395] Marion Thomas. Analyses d’ouvrages: L’Intelligence des animaux selon Charles-Georges Leroy (1723–1789) par Charles-Georges Leroy; Boris Cyrulnik; François Sigaut; Jean-Luc Renck; Véronique Servais; Robinet. Revue d’Histoire des Sciences, 61(1):220–223, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634211.

Lamy:2008:AOBb

[4396] Jérôme Lamy. Analyses d’ouvrages: La Science au Parlement : Les débuts d’une politique des recherches scientifiques en France, ^{≪}CNRS histoire ^{≫}par Michel Pinault. Revue d’Histoire des Sciences, 61(1):223– 224, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634212.

Nicolaidis:2008:AOB

[4397] Efthymios Nicolaïdis. Analyses d’ouvrages: Les astronomes parisiens au XIV^{e} siècle et l’astronomie alphonsine, in Histoire littéraire de la France, t. 43, fasc. 1 par Emmanuel Poulle. Revue d’Histoire des Sciences, 61(1):225–227, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634213.

Jouary:2008:AOB

[4398] Jean-Paul Jouary. Analyses d’ouvrages: Émergence, complexité et dialectique : Sur les systèmes dynamiques non linéaires par Lucien Sève; Janine Guespin-Michel; Roland Charlionet; Philippe Gascuel. Revue d’Histoire des Sciences, 61(1):227–229, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634214.

Anonymous:2008:EI

[4399] Anonyme. Erratum: [illustration]. Revue d’Histoire des Sciences, 61 (2):237, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634273.

Joly:2008:P

[4400] Bernard Joly and Pierre-François Moreau. [Présentation]. Revue d’Histoire des Sciences, 61(2):239–244, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634274.

Hirai:2008:LSC

[4401] Hiro Hirai. Les “logoi spermatikoi” et le concept de semence dans la minéralogie et la cosmogonie de Paracelse. Revue d’Histoire des Sciences, 61(2):245–264, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634275.

Barker:2008:SAA

[4402] Peter Barker. Stoic alternatives to Aristotelian cosmology : Pena, Rothmann and Brahe. Revue d’Histoire des Sciences, 61(2):265–286, juillet– décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634276.

Joly:2008:PST

[4403] Bernard Joly. Présences stoïciennes dans les théories des marées aux XVI^{e} et XVII^{e} siècles. Revue d’Histoire des Sciences, 61(2):287– 311, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634277.

Giglioni:2008:CSU

[4404] Guido Giglioni. The cosmoplastic system of the universe : Ralph Cudworth on stoic naturalism. Revue d’Histoire des Sciences, 61(2): 313–331, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634278.

Chamayou:2008:QTT

[4405] Grégoire Chamayou. La querelle des têtes tranchées : Les médecins, la guillotine et l’anatomie de la conscience au lendemain de la Terreur. Revue d’Histoire des Sciences, 61(2):333–365, juillet–décembre 2008. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634279.

Popov:2008:OVD

[4406] Igor Popov. Orthogenesis versus Darwinism : The Russian case. Revue d’Histoire des Sciences, 61(2):367–397, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634280.

Pizarroso:2008:PHV

[4407] Noemí Pizarroso. La psychologie historique vue par la psychologie expérimentale : Analyse d’une rencontre manquée. Revue d’Histoire des Sciences, 61(2):399–434, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634281.

Anonymous:2008:Ib

[4408] Anonyme. Informations. Revue d’Histoire des Sciences, 61(2):435– 440, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634282.

Aragon:2008:AOB

[4409] Santiago Aragon. Analyses d’ouvrages: Actas del IX Congreso de la Sociedad española de historia de las ciencias y de las técnicas (27, 28, 29 et 30 septembre 2005), vol. I, vol. II. Revue d’Histoire des Sciences, 61 (2):441–442, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634283.

Bontems:2008:AOB

[4410] Vincent Bontems. Analyses d’ouvrages: Contingence et nécessité des lois de la nature au XVIII^{e} siècle : La philosophie seconde des Lumières, ≪Histoire de la philosophie ^{≫}par André Charrak. Revue d’Histoire des Sciences, 61(2):442–448, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634284.

Crepel:2008:AOB

[4411] Pierre Crépel. Analyses d’ouvrages: La Bibliothèque de l’Académie royale des sciences au XVIII^{e} siècle. coll. ^{≪}Sciences ^{≫}, 5 par Annie Chassagne; Pierre Messmer; Jean Dercourt; Nicole Le Douarin. Revue d’Histoire des Sciences, 61(2):448–449, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634285.

Missa:2008:AOB

[4412] Jean-Noël Missa. Analyses d’ouvrages: Histoire des aphasies : Une anatomie de l’expression ^{≪}Pratiques théoriques ^{≫} par Denis Forest. Revue d’Histoire des Sciences, 61(2):449–450, juillet–décembre 2008. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634286.

Fox:2008:AOB

[4413] Robert Fox. Analyses d’ouvrages: Entre l’État et l’usine: L’École des mines de Saint-Étienne au XIX^{e} siècle ^{≪}Histoire ^{≫}par Anne-Françoise Garçon. Revue d’Histoire des Sciences, 61(2):450–451, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634287.

Lamy:2008:AOBc

[4414] Jérôme Lamy. Analyses d’ouvrages: Lieux de savoir : Espaces et communautés par Christian Jacob. Revue d’Histoire des Sciences, 61(2): 452–454, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634288.

Fauque:2008:AOBa

[4415] Danielle Fauque. Analyses d’ouvrages: La Juste mesure : Quantifier, évaluer, mesurer, entre Orient et Occident (VIII^{e}–XVIII^{e} siècle), ≪Temps et espaces ^{≫} par Laurence Moulinier; Line Sallmann; Catherine Verna; Nicolas Weill-Parot. Revue d’Histoire des Sciences, 61(2): 455–457, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634289.

Fauque:2008:AOBb

[4416] Danielle Fauque. Analyses d’ouvrages: Henri Sainte-Claire Deville : Chimie, recherche et industrie ^{≪}Hommes de sciences, hommes de vérité ^{≫}par Catherine Paquot. Revue d’Histoire des Sciences, 61(2): 457–459, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634290.

Michel:2008:AOB

[4417] Jacques Michel. Analyses d’ouvrages: La Tradition scientifique : Méthode et histoire par Mohammad K. Salhab. Revue d’Histoire des Sciences, 61(2):459–461, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634291.

Anonymous:2008:TG

[4418] Anonyme. Table générale. Revue d’Histoire des Sciences, 61(2):463– 471, juillet–décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634292.

Anonymous:2008:BMa

[4419] Anonyme. Back matter. Revue d’Histoire des Sciences, 61(1):??, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634215.

Anonymous:2008:BMb

[4420] Anonyme. Back matter. Revue d’Histoire des Sciences, 61(2):??, juillet– décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634293.

Anonymous:2008:FMa

[4421] Anonyme. Front matter. Revue d’Histoire des Sciences, 61(1):??, January 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634191.

Anonymous:2008:FMb

[4422] Anonyme. Front matter. Revue d’Histoire des Sciences, 61(2):??, juillet– décembre 2008. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634272.

Beaulieu:2009:EDC

[4423] Liliane Beaulieu. Entre deux citations. Revue d’Histoire des Sciences, 62 (1):5–8, January 2009. CODEN RHSAAM. ISBN 2-200-92491-7. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634486.

Beaulieu:2009:RMF

[4424] Liliane Beaulieu. Regards sur les mathématiques en France entre les deux guerres. Introduction. Revue d’Histoire des Sciences, 62(1):9–38, January 2009. CODEN RHSAAM. ISBN 2-200-92491-7. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634487.

Gispert:2009:PMF

[4425] Hélène Gispert and Juliette Leloup. Des patrons des mathématiques en France dans l’entre-deux-guerres. Revue d’Histoire des Sciences, 62(1): 39–117, January 2009. CODEN RHSAAM. ISBN 2-200-92491-7. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634488.

Demidov:2009:RMF

[4426] Sergej S. Demidov. Les relations mathématiques franco-russes entre les deux guerres mondiales. Revue d’Histoire des Sciences, 62(1):119–142, January 2009. CODEN RHSAAM. ISBN 2-200-92491-7. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/ 23634489. With an appendix 2 by Liliane Beaulieu.

Goldstein:2009:TNF

[4427] Catherine Goldstein. La théorie des nombres en France dans l’entredeux-guerres : de quelques effets de la première guerre mondiale. Revue d’Histoire des Sciences, 62(1):143–175, January 2009. CODEN RHSAAM. ISBN 2-200-92491-7. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634490.

Guillaume:2009:LMF

[4428] Marcel Guillaume. La logique mathématique en France entre les deux guerres mondiales : Quelques repères. Revue d’Histoire des Sciences, 62(1):177–219, January 2009. CODEN RHSAAM. ISBN 2-200-92491-7. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634491.

Nabonnand:2009:NHC

[4429] Philippe Nabonnand. La notion d’holonomie chez Élie Cartan. Revue d’Histoire des Sciences, 62(1):221–245, January 2009. CODEN RHSAAM. ISBN 2-200-92491-7. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634492.

Siegmund-Schultze:2009:IHP

[4430] Reinhard Siegmund-Schultze. The Institute Henri Poincaré and mathematics in France between the wars. Revue d’Histoire des Sciences, 62(1): 247–283, January 2009. CODEN RHSAAM. ISBN 2-200-92491-7. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634493.

Gandon:2009:TRC

[4431] Sébastien Gandon. La théorie des rapports chez Augustus De Morgan. Revue d’Histoire des Sciences, 62(1):285–311, January 2009. CODEN RHSAAM. ISBN 2-200-92491-7. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634494.

Jullien:2009:AOB

[4432] Vincent Jullien. Analyses d’ouvrages: Spinoza: La science mathématique du salut, ^{≪}CNRS Philosophie ^{≫}par Françoise Barbaras. Revue d’Histoire des Sciences, 62(1):313–317, January 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634495.

Basso:2009:AOBa

[4433] Elisabetta Basso. Analyses d’ouvrages: L’Histoire des sciences : Méthodes, styles et controverses, ^{≪}Textes clés de l’histoire des sciences ^{≫} par Jean-François Braunstein; J.-F. Braunstein; V. Guillin; A. Zielinska. Revue d’Histoire des Sciences, 62(1):317–320, January 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634496.

Crepel:2009:AOB

[4434] Pierre Crépel. Analyses d’ouvrages: Les Archives des scientifiques, XVI^{e}–XX^{e} siècle: Guide des fonds conservés en France par Thérèse Charmasson; Marion Daniel; Catherine Gaziello; Carole Rota-Treguier. Revue d’Histoire des Sciences, 62(1):320–321, January 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634497.

Fauque:2009:AOBa

[4435] Danielle Fauque. Analyses d’ouvrages: Atti del X convegno nazionale di storia e fondamenti della chimica. [Extr. du vol. 121 des Memorie di scienze fisiche e naturali, ^{≪}Rendiconti della Acc. naz. delle scienze detta dei XL ^{≫}, sér. V, vol. XXVII, part. II, t. II, 2003.] par Marco Ciardi; Franco Giudice; Atti del XI convegno nazionale di storia e fondamenti della chimica. [Extr. du vol. 123 des Memorie di scienze fisiche e naturali, ^{≪}Rendiconti della Acc. naz. delle scienze detta dei XL ^{≫}, sér. V, vol. XXIX, part. II, t. II, 2005.] par Luigi Cerruti; Francesca Turco. Revue d’Histoire des Sciences, 62(1):322–324, January 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634498.

Dupont:2009:AOBa

[4436] Jean-Claude Dupont. Analyses d’ouvrages: La Matrice de la race : Généalogie sexuelle et coloniale de la nation française. ^{≪}Textes à l’appui/Genre et sexualité ^{≫}par Elsa Dorlin; Joan W. Scott; Éric Fassin. Revue d’Histoire des Sciences, 62(1):324–326, January 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634499.

Dupont:2009:AOBb

[4437] Jean-Claude Dupont. Analyses d’ouvrages: Le Siècle du gène par Evelyn Fox Keller; Stéphane Schmitt; François Jacob. Revue d’Histoire des Sciences, 62(1):326–328, January 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634500.

Dupont:2009:AOBc

[4438] Jean-Claude Dupont. Analyses d’ouvrages: L’Éternel retour de l’eugénisme. ^{≪}Science, histoire et société ^{≫}par Jean Gayon; Daniel Jacobi; Marie-Claude Lorne. Revue d’Histoire des Sciences, 62(1):328–330, January 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634501.

Knittel:2009:AOB

[4439] Fabien Knittel. Analyses d’ouvrages: À quoi sert l’histoire des sciences? Conférence prononcée le 26 octobre 2006, coll. ^{≪}Sciences en questions ≫ par Michel Morange. Revue d’Histoire des Sciences, 62(1):330–332, January 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634502.

Savaton:2009:AOB

[4440] Pierre Savaton. Analyses d’ouvrages: The Earth sciences in the Enlightenment : Studies on the early development of geology, ^{≪}Variorum collected studies series ^{≫} par Kenneth L. Taylor. Revue d’Histoire des Sciences, 62(1):332–333, January 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634503.

Goldberg:2009:AOB

[4441] Sylvie-Anne Goldberg. Analyses d’ouvrages: L’Université hébraïque de Jérusalem à travers ses acteurs : La première génération de professeurs (1925–1948), ^{≪}Bibliothèque d’études juives ^{≫}, 21, ^{≪}Série Histoire ^{≫}, XVII par Eva Telkes-Klein. Revue d’Histoire des Sciences, 62(1):333– 335, January 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634504.

Anonymous:2009:PR

[4442] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 62(1):337– 342, January 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23634505.

Sackur:2009:LPE

[4443] Jérôme Sackur. L’introspection en psychologie expérimentale. Revue d’Histoire des Sciences, 62(2):349–372, juillet–décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634397.

Klein:2009:PSA

[4444] Alexandre Klein. La philosophie scientifique d’Alfred Binet. Revue d’Histoire des Sciences, 62(2):373–393, juillet–décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634398.

dHombres:2009:FOE

[4445] Emmanuel d’Hombres. ^{≪}Un organisme est une société, et réciproquement? ≫La délimitation des champs d’extension des sciences de la vie et des sciences sociales chez Alfred Espinas (1877). Revue d’Histoire des Sciences, 62(2):395–422, juillet–décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634399.

Grimoult:2009:DPF

[4446] Cédric Grimoult. Le débat sur le ponctualisme en France : Le colloque international de Dijon en 1982. Revue d’Histoire des Sciences, 62(2): 423–454, juillet–décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634400.

Bracco:2009:EMR

[4447] Christian Bracco and Jean-Pierre Provost. De l’électromagnétisme à la mécanique: le rôle de l’action dans le Mémoire de Poincaré de 1905. Revue d’Histoire des Sciences, 62(2):457–493, juillet–décembre 2009. CO-DEN RHSAAM. ISBN 2-200-92599-9. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634401.

Fauque:2009:RCS

[4448] Danielle Fauque. Revue critique. un siècle de physique à Oxford (1839– 1939). Revue d’Histoire des Sciences, 62(2):495–510, juillet–décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634402.

Hamou:2009:MGS

[4449] Philippe Hamou. In memoriam: Gérard Simon (1931–2009). Revue d’Histoire des Sciences, 62(2):511–516, juillet–décembre 2009. CODEN RHSAAM. ISBN 2-200-92599-9. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634403.

Basso:2009:AOBb

[4450] Elisabetta Basso. Analyses d’ouvrages: La Philosophie de la médecine d’Auguste Comte : Vaches carnivores, Vierge mère et morts vivants, ≪Science, histoire et société ^{≫} par Jean-François Braunstein. Revue d’Histoire des Sciences, 62(2):517–520, juillet–décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634404.

Guillin:2009:AOB

[4451] Vincent Guillin. Analyses d’ouvrages: Canguilhem : Histoire des sciences et politique du vivant ^{≪}Débats philosophiques ^{≫}par Jean-François Braunstein. Revue d’Histoire des Sciences, 62(2):520–522, juillet– décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634405.

Pourprix:2009:AOB

[4452] Bernard Pourprix. Analyses d’ouvrages: L’Édification au XIX^{e} siècle d’une science du phénomène lumineux : Propagation de la lumière et conception du phénomène lumineux du début du XIX^{e} siècle à la naissance de la théorie de la relativité, ^{≪}Mathesis ^{≫}par André Chappert. Revue d’Histoire des Sciences, 62(2):522–524, juillet–décembre 2009. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634406.

Aragon:2009:AOBa

[4453] Santiago Aragon. Analyses d’ouvrages: Museología médica, aspectos teóricos y cuestiones prácticas, vol. 1 : vol. 2 par Felip Cid. Revue d’Histoire des Sciences, 62(2):524–526, juillet–décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634407.

Maury:2009:AOB

[4454] Liliane Maury. Analyses d’ouvrages: Rousseau l’animal et l’homme : L’animalité dans l’horizon anthropologique des Lumières, ^{≪}La nuit surveillée ^{≫} par Jean-Luc Guichet. Revue d’Histoire des Sciences, 62(2): 526–527, juillet–décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634408.

Schmitt:2009:AOB

[4455] Stéphane Schmitt. Analyses d’ouvrages: Buffon : Histoire naturelle et philosophie, ^{≪}Les dix-huitièmes siècles ^{≫}, 92 par Thierry Hoquet. Revue d’Histoire des Sciences, 62(2):527–529, juillet–décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634409.

Lamy:2009:AOB

[4456] Jérôme Lamy. Analyses d’ouvrages: British university observatories, 1772–1939, ^{≪}Science, technology and culture, 1700–1945 ^{≫}par Roger Hutchins. Revue d’Histoire des Sciences, 62(2):529–531, juillet–décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634410.

Fauque:2009:AOBb

[4457] Danielle Fauque. Analyses d’ouvrages: Volta : Science and culture in the Age of Enlightenment, 2^{e} éd. [1^{re} éd., 2003] par Giuliano Pancaldi. Revue d’Histoire des Sciences, 62(2):531–533, juillet–décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634411.

Aragon:2009:AOBb

[4458] Santiago Aragon. Analyses d’ouvrages: Science in Latin America: A history, ^{≪}Latin American studies: History of science ^{≫}par Juan José Saldaña; Bernabé Madrigal. Revue d’Histoire des Sciences, 62(2): 534–536, juillet–décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634412.

Anonymous:2009:TG

[4459] Anonyme. Table générale. Revue d’Histoire des Sciences, 62(2):537– 543, juillet–décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634413.

Anonymous:2009:BMa

[4460] Anonyme. Back matter. Revue d’Histoire des Sciences, 62(1):??, January 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634506.

Anonymous:2009:BMb

[4461] Anonyme. Back matter. Revue d’Histoire des Sciences, 62(2):??, juillet– décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634414.

Anonymous:2009:FMa

[4462] Anonyme. Front matter. Revue d’Histoire des Sciences, 62(1):??, January 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634485.

Anonymous:2009:FMb

[4463] Anonyme. Front matter. Revue d’Histoire des Sciences, 62(2):??, juillet– décembre 2009. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634396.

Fauque:2010:PBFa

[4464] Danielle Fauque. Pierre Bouguer, figure emblématique ou savant singulier? Revue d’Histoire des Sciences, 63(1):5–21, January 2010. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634435.

Fauque:2010:PBFb

[4465] Danielle Fauque. Pierre Bouguer et l’^{≪}affaire du jaugeage ^{≫}, 1721– 1726. Revue d’Histoire des Sciences, 63(1):23–66, January 2010. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634436.

Mayrargue:2010:MOE

[4466] Arnaud Mayrargue. De la méthode d’observer exactement sur mer la hauteur des astres. Revue d’Histoire des Sciences, 63(1):67–91, January 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634437.

Ferreiro:2010:PBS

[4467] Larrie D. Ferreiro, Michel Blay, and Alexandre Guilbaud. Pierre Bouguer and the solid of least resistance. Revue d’Histoire des Sciences, 63(1): 93–119, January 2010. CODEN RHSAAM. ISBN 2-200-92659-6. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/ stable/23634438. With a letter in French by Michel Blay and Alexandre Guilbaud to Ferreiro, dated September 2008, concerning a solid of least resistance.

Boistel:2010:PBC

[4468] Guy Boistel. Pierre Bouguer, commissaire pour la marine et expert pour les longitudes : Un opposant au développement de l’horlogerie de marine au XVIII^{e} siècle? Revue d’Histoire des Sciences, 63(1):121–159, January 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634439.

Alfonsi:2010:SBE

[4469] Liliane Alfonsi. Un successeur de Bouguer : Étienne Bézout (1730– 1783), commissaire et expert pour la marine. Revue d’Histoire des Sciences, 63(1):161–187, January 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634440.

Fauque:2010:RCA

[4470] Danielle Fauque. Revue critique. de l’art de naviguer à la science nautique au Siècle des lumières. Revue d’Histoire des Sciences, 63(1):189–219, January 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634441.

Colin:2010:JSI

[4471] Cécile Colin. Jacques Solomon et l’interprétation de la théorie quantique. Revue d’Histoire des Sciences, 63(1):221–246, January 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634442.

Telkes-Klein:2010:GIR

[4472] Eva Telkes-Klein. La genèse d’“Identité et réalité” (1908) à travers une lettre d’Émile Meyerson à sa sœur. Revue d’Histoire des Sciences, 63(1):247–297, January 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634443.

Brenner:2010:EMS

[4473] Anastasios Brenner. Sur Émile Meyerson et son œuvre. revue critique. Revue d’Histoire des Sciences, 63(1):299–302, January 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634444.

Karela:2010:AOB

[4474] Catherine Karela. Analyses d’ouvrages: Gödel, ^{≪}Figures du savoir ^{≫}par Pierre Cassou-Noguès. Revue d’Histoire des Sciences, 63(1):303–305, January 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634445.

Brenner:2010:AOB

[4475] Anastasios Brenner. Analyses d’ouvrages: Le Traité de navigation par Jean-Baptiste Denoville; Ships and science : The birth of naval architecture in the scientific revolution, 1600–1800 par Larrie D. Ferreiro; Essere, apparire e interpretare : Saggio sul pensiero di Duhem (1861–1916), ≪Epistemologia ^{≫}par Mirella Fortino; Verificazione e olismo par Pierre Duhem; Mirella Fortino. Revue d’Histoire des Sciences, 63(1):305–307, January 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634446.

Rommevaux:2010:AOB

[4476] Sabine Rommevaux. Analyses d’ouvrages: L’Épistémologie d’Émile Meyerson: Une anthropologie de la connaissance par Frédéric Fruteau de Laclos; Regiomontano e il rinnovamento del sapere matematico e astronomico nel Quattrocento, ^{≪}Collana di storia della scienza ^{≫}par Michela Malpangotto. Revue d’Histoire des Sciences, 63(1):307–308, January 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634447.

Basso:2010:AOB

[4477] Elisabetta Basso. Analyses d’ouvrages: Lettres françaises, ^{≪}Cahiers du Centre de recherche français de Jérusalem. Série Hommes & sociétés ≫par Émile Meyerson; Bernadette Bensaude-Vincent; Eva Telkes-Klein; Naissance de la psychiatrie biologique : Histoire des traitements des maladies mentales au XX^{e} siècle, ^{≪}Science, histoire et société ^{≫}par Jean-Noël Missa. Revue d’Histoire des Sciences, 63(1):308–312, January 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23634448.

Fauque:2010:AOB

[4478] Danielle Fauque. Analyses d’ouvrages: Sextants at Greenwich : A catalogue of the mariner’s quadrants, mariner’s astrolabes, cross-staffs, backstaffs, octants, sextants, quintants, reflecting circles, and artificial horizons in the National Maritime Museum, Greenwich par Willem F. J. Mörzer Bruyns; Richard Dunn; The Quest for the true figure of the Earth : Ideas and expeditions in four centuries of geodesy, ^{≪}Science, technology and culture, 1700–1945 ^{≫}par Michael Rand Hoare. Revue d’Histoire des Sciences, 63(1):312–316, January 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634449.

Barras:2010:I

[4479] Vincent Barras and Jean-Claude Dupont. [introduction]. Revue d’Histoire des Sciences, 63(2):325–329, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634363.

Mendelzweig:2010:PCC

[4480] Marion Droz Mendelzweig. La ploasticité cérébrale de Cajal à Kandel : Cheminement d’une notion constitutive du sujet cérébral. Revue d’Histoire des Sciences, 63(2):331–367, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634364.

Barbara:2010:RMS

[4481] Jean-Gaël Barbara. Relations médecine — sciences dans l’individualisation des maladies nerveuses à la Salpêtriére à la fin du XIX^{e} siècle. Revue d’Histoire des Sciences, 63(2):369–407, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634365.

Cherici:2010:DEC

[4482] Céline Cherici. La définition d’une entité clinique entre développements techniques et spécialisation médicale : Épilepsie et épileptologie au XX^{e} siècle. Revue d’Histoire des Sciences, 63(2):409–437, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634366.

Pidoux:2010:ECE

[4483] Vincent Pidoux. Expérimentation et clinique électroencéphalographiques entre physiologie, neurologie et psychiatrie (Suisse, 1935–1965). Revue d’Histoire des Sciences, 63(2):439–472, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634367.

Aragon:2010:PTI

[4484] Santiago Aragón and Miguel Villena. Les premières tentatives d’institutionnalisation de la zoologie en Espagne. Revue d’Histoire des Sciences, 63(2):473–499, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634368.

Canabal:2010:DSO

[4485] Manuela Canabal. De la découverte scientifique à l’occultation du savoir : Isaac-Bénédict Prévost et la carie du blé (1798–1807). Revue d’Histoire des Sciences, 63(2):501–527, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634369.

Nicolaidis:2010:AOBa

[4486] Efthymios Nicolaïdis. Analyses d’ouvrages: Médecins et ingénieurs ottomans à l’âge des nationalismes, ^{≪}Institut français d’études anatoliennes ^{≫}par Méropi Anastassiadou-Dumont. Revue d’Histoire des Sciences, 63(2):529–531, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634370.

Nicolaidis:2010:AOBb

[4487] Efthymios Nicolaïdis. Analyses d’ouvrages: Jérôme Lalande (1732–1807) : Une trajectoire scientifique, ^{≪}Histoire ^{≫}par Guy Boistel; Jérôme Lamy; Colette Le Lay. Revue d’Histoire des Sciences, 63(2):531–532, juillet– décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634371.

Celeyrette:2010:AOB

[4488] Jean Celeyrette. Analyses d’ouvrages: Traité des rapports entre les rapidités dans les mouvements, Sur les rapports de rapports. ^{≪}Sagesses médiévales ^{≫}par Thomas Bradwardine; Nicole Oresme; Sabine Rommevaux. Revue d’Histoire des Sciences, 63(2):533–534, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634372.

Simon:2010:AOB

[4489] Jonathan Simon. Analyses d’ouvrages: Scientific institutions and practice in France and Britain, c. 1700– c. 1870, ^{≪}Variorum collected studies series ^{≫} par Maurice Crosland. Revue d’Histoire des Sciences, 63(2): 534–535, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634373.

Maury:2010:AOB

[4490] Liliane Maury. Analyses d’ouvrages: Genèse et développement d’un fait scientifique. ^{≪}Médecine et sciences humaines ^{≫}par Ludwig Fleck; Nathalie Jas; Ilana Löwy; Bruno Latour. Revue d’Histoire des Sciences, 63(2):535–538, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634374.

Fabre:2010:AOB

[4491] François Fabre. Analyses d’ouvrages: Philosophie et alchimie à la Renaissance et à l’Âge classique, 1 : Scolastique et alchimie (XVI^{e}–XVII^{e} siècles), ^{≪}Textes et travaux de Chrysopoeia ^{≫}, 10 par Sylvain Matton. Revue d’Histoire des Sciences, 63(2):538–540, juillet–décembre 2010. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634375.

Jullien:2010:AOBa

[4492] Vincent Jullien. Analyses d’ouvrages: Géométries de Port-Royal, ≪Sources classiques ^{≫}, 100 par Blaise Pascal; Antoine Arnauld; François de Nonancourt; Dominique Descotes. Revue d’Histoire des Sciences, 63 (2):540–542, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634376.

Sasaki:2010:AOB

[4493] Chikara Sasaki. Analyses d’ouvrages: Les Mathématiques infinitésimales du IX^{e} au XI^{e} siècle, vol. V : Ibn al-Haytham : Astronomie, géométrie sphérique et trigonométrie par Roshdi Rashed. Revue d’Histoire des Sciences, 63(2):542–546, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634377.

Jullien:2010:AOBb

[4494] Vincent Jullien. Analyses d’ouvrages: Physique nouvelle (1667), ≪Anecdota ^{≫}, 6 par Jacques Rohault. Revue d’Histoire des Sciences, 63 (2):547–548, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634378.

Verdet:2010:AOB

[4495] Cyril Verdet. Analyses d’ouvrages: La Loi de la gravitation universelle. Newton, Euler et Laplace : Le cheminement d’une révolution scientifique vers une science normale par Prosper Schroeder. Revue d’Histoire des Sciences, 63(2):548–550, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634379.

Anonymous:2010:PR

[4496] Anonyme. Publications reçues. Revue d’Histoire des Sciences, 63(2): 551–556, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634380.

Anonymous:2010:TG

[4497] Anonyme. Table générale. Revue d’Histoire des Sciences, 63(2):557– 563, juillet–décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634381.

Anonymous:2010:BMa

[4498] Anonyme. Back matter. Revue d’Histoire des Sciences, 63(1):??, January 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634450.

Anonymous:2010:BMb

[4499] Anonyme. Back matter. Revue d’Histoire des Sciences, 63(2):??, juillet– décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634382.

Anonymous:2010:FMa

[4500] Anonyme. Front matter. Revue d’Histoire des Sciences, 63(1):??, January 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634434.

Anonymous:2010:FMb

[4501] Anonyme. Front matter. Revue d’Histoire des Sciences, 63(2):??, juillet– décembre 2010. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634362.

Diemer:2011:LPA

[4502] Arnaud Diemer and Hervé Guillemin. L’économie politique au miroir de la physique: Adam Smith et Isaac Newton. Revue d’Histoire des Sciences, 64(1):5–26, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634417.

Frobert:2011:TCS

[4503] Ludovic Frobert. Théorie cellulaire, science économique et République dans l’œuvre de François-Vincent Raspail autour de 1830. Revue d’Histoire des Sciences, 64(1):27–58, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634418.

Eisenstaedt:2011:AVL

[4504] Jean Eisenstaedt and Michel Combes. Arago et la vitesse de la lumière (1806–1810), un manuscrit inédit, une nouvelle analyse. Revue d’Histoire des Sciences, 64(1):59–120, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634419.

Leclercq:2011:BPC

[4505] Frédéric Leclercq. Biot, la polarisation chromatique et la théorie des accès. Revue d’Histoire des Sciences, 64(1):121–156, January 2011. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634420.

Bremond:2011:RGP

[4506] Alain Brémond. La recherche sur les galaxies : La place de la France de 1910 à 1950. Revue d’Histoire des Sciences, 64(1):157–187, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634421.

Darrigol:2011:AOB

[4507] Olivier Darrigol. Analyses d’ouvrages: Quantum theory at the crossroads : Reconsidering the 1927 Solvay conference par Guido Bacciagaluppi; Antony Valentini. Revue d’Histoire des Sciences, 64(1):189–191, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634422.

Basso:2011:AOBa

[4508] Elisabetta Basso. Analyses d’ouvrages: Bachelard, ^{≪}Figures du savoir ≫par Vincent Bontems. Revue d’Histoire des Sciences, 64(1):191–194, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634423.

Verdet:2011:AOB

[4509] Cyril Verdet. Analyses d’ouvrages: Science, histoire et philosophie selon Gaston Milhaud : La constitution d’un champ disciplinaire sous la Troisième République, ^{≪}Cahiers d’histoire et de philosophie des sciences ^{≫}par Anastasios Brenner; Annie Petit. Revue d’Histoire des Sciences, 64(1):194–195, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634424.

Nicolaidis:2011:AOB

[4510] Efthymios Nicolaïdis. Analyses d’ouvrages: Philosophie et sciences à Byzance de 1204 à 1453 : Les textes, les doctrines et leur transmission. Actes de la table ronde organisée au XX^{e} congrès international d’études byzantines, Paris, 2001. ^{≪}Orientalia Lovaniensia Analecta ^{≫}, 146 par Michel Cacouros; Marie–Hélène Congourdeau; Jean Irigoin. Revue d’Histoire des Sciences, 64(1):195–198, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634425.

Basso:2011:AOBb

[4511] Elisabetta Basso. Analyses d’ouvrages: Il Valore dei concetti : Filosofia e critica, ^{≪}Epistemologia ^{≫} par Antonella Cutro. Revue d’Histoire des Sciences, 64(1):198–201, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634426.

Belna:2011:AOBa

[4512] Jean-Pierre Belna. Analyses d’ouvrages: Cantor et la France: Correspondance du mathématicien allemand avec les Français à la fin du XIX^{e} siècle, ^{≪}Philosophie, épistémologie ^{≫} par Anne-Marie Décaillot. Revue d’Histoire des Sciences, 64(1):201–203, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634427.

Belna:2011:AOBb

[4513] Jean-Pierre Belna. Analyses d’ouvrages: Cantor et le transfini : Mathématique et ontologie, ^{≪}Ouverture philosophique^{≫}par Philippe Lauria. Revue d’Histoire des Sciences, 64(1):203–205, January 2011. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634428.

Fauque:2011:AOB

[4514] Danielle Fauque. Analyses d’ouvrages: Creating networks in chemistry : The founding and early history of chemical societies in Europe par Anita Kildebæk Nielsen; Sona Strbánova. Revue d’Histoire des Sciences, 64(1):205–207, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634429.

Basso:2011:AOBc

[4515] Elisabetta Basso. Analyses d’ouvrages: Lo Sguardo psichiatrico : Studi e materiali dalle cartelle cliniche tra Otto e Novecento, ^{≪}Ricerca ^{≫}par Riccardo Panattoni. Revue d’Histoire des Sciences, 64(1):207–209, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634430.

Gilain:2011:AOB

[4516] Christian Gilain. Analyses d’ouvrages: Collected works of Charles-François Sturm par Jean-Claude Pont; Flavia Padovani. Revue d’Histoire des Sciences, 64(1):209–211, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634431.

Giroux:2011:CHE

[4517] Élodie Giroux. Contribution à l’histoire de l’épidémiologie des facteurs de risque. Revue d’Histoire des Sciences, 64(2):219–224, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634467.

Morabia:2011:EPE

[4518] Alfredo Morabia. Épidémiologie politique des études cas-témoins (1926– 1950). Revue d’Histoire des Sciences, 64(2):225–242, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634468.

Parascandola:2011:ETC

[4519] Mark Parascandola. The epidemiologic transition and changing concepts of causation and causal inference. Revue d’Histoire des Sciences, 64(2): 243–262, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634469.

Aronowitz:2011:FHS

[4520] Robert A. Aronowitz. The Framingham heart study and the emergence of the risk factor approach to coronary heart disease, 1947–1970. Revue d’Histoire des Sciences, 64(2):263–295, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634470.

Giroux:2011:OEP

[4521] Élodie Giroux. Origines de l’étude prospective de cohorte : Épidémiologie cardio-vasculaire américaine et étude de Framingham. Revue d’Histoire des Sciences, 64(2):297–318, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634471.

Giroux:2011:IBH

[4522] Élodie Giroux. Indications bibliographiques sur l’histoire de l’épidémiologie. Revue d’Histoire des Sciences, 64(2):319–322, juillet–décembre 2011. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634472.

Quin:2011:GSI

[4523] Grégory Quin. Genèse et structure d’un interchamp orthopédique (première moitié du XIX^{e} siècle) : Contribution à l’histoire de l’institutionnalisation d’un champ scientifique. Revue d’Histoire des Sciences, 64(2):323–347, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634473.

Barroux:2011:SAE

[4524] Gilles Barroux. La santé des animaux et l’émergence d’une médecine vétérinaire au XVIII^{e} siècle. Revue d’Histoire des Sciences, 64(2): 349–376, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634474.

Lamy:2011:PCS

[4525] Jérôme Lamy and Arnaud Saint-Martin. Pratiques et collectifs de la science en régimes note critique. Revue d’Histoire des Sciences, 64(2): 377–389, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634475.

Barroux:2011:AOB

[4526] Gilles Barroux. Analyses d’ouvrages: Quand les sciences dialoguent avec la métaphysique par Pascal Charbonnat; Francine Markovits-Pessel. Revue d’Histoire des Sciences, 64(2):391–393, juillet–décembre 2011. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634476.

Clavelin:2011:AOB

[4527] Maurice Clavelin. Analyses d’ouvrages: On sunspots, ^{≪}Science/History ≫par Galileo Galilei; Christoph Scheiner; Eileen Reeves; Albert Van Helden. Revue d’Histoire des Sciences, 64(2):393–395, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634477.

Lamy:2011:AOB

[4528] Jérôme Lamy. Analyses d’ouvrages: Lieux de savoir, t. 2 : Les mains de l’intellect par Christian Jacob. Revue d’Histoire des Sciences, 64(2): 395–400, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634478.

Maury:2011:AOB

[4529] Liliane Maury. Analyses d’ouvrages: Cours de philosophie de 1830 à l’École normale suivi de Du suicide, ^{≪}Anecdota ^{≫}, 4 par Théodore Jouffroy; Sylvain Matton; Serge Nicolas. Revue d’Histoire des Sciences, 64 (2):400–403, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634479.

Niderst:2011:AOB

[4530] Alain Niderst. Analyses d’ouvrages: Émilie Du Châtelet : Éclairages et documents nouveaux, ^{≪}Publications du Centre international d’étude du XVIII^{e} siècle ^{≫}, 21 par Ulla Kölving; Olivier Courcelle. Revue d’Histoire des Sciences, 64(2):403–404, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634480.

Knittel:2011:AOB

[4531] Fabien Knittel. Analyses d’ouvrages: Parmentier, Chaptal, Chevreul : Trois grands pionniers de la chimie alimentaire par Paul Mazliak. Revue d’Histoire des Sciences, 64(2):404–407, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634481.

LeRu:2011:AOB

[4532] Véronique Le Ru. Analyses d’ouvrages: L’Arbre et le labyrinthe : Descartes selon l’ordre des Lumières par Mariafranca Spallanzani; Jean Dagen. Revue d’Histoire des Sciences, 64(2):407–409, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634482.

Anonymous:2011:TG

[4533] Anonyme. Table générale. Revue d’Histoire des Sciences, 64(2):411– 416, juillet–décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634483.

Anonymous:2011:BMa

[4534] Anonyme. Back matter. Revue d’Histoire des Sciences, 64(1):??, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634432.

Anonymous:2011:BMb

[4535] Anonyme. Back matter. Revue d’Histoire des Sciences, 64(2):??, juillet– décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634484.

Anonymous:2011:FMa

[4536] Anonyme. Front matter. Revue d’Histoire des Sciences, 64(1):??, January 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634416.

Anonymous:2011:FMb

[4537] Anonyme. Front matter. Revue d’Histoire des Sciences, 64(2):??, juillet– décembre 2011. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634466.

Carteret:2012:MAA

[4538] Xavier Carteret. Michel Adanson au Sénégal (1749–1754) : Un grand voyage naturaliste et anthropologique du Siècle des lumières. Revue d’Histoire des Sciences, 65(1):5–25, January 2012. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23634385.

Tesio:2012:CMA

[4539] Stéphanie Tésio. Climatologie et médecine au Canada au XVIII^{e} siècle. Revue d’Histoire des Sciences, 65(1):27–59, January 2012. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634386.

Loison:2012:PDN

[4540] Laurent Loison. Le projet du néolamarckisme français (1880–1910). Revue d’Histoire des Sciences, 65(1):61–79, January 2012. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634387.

Beaulieu:2012:ANW

[4541] Alain Beaulieu. Alfred North Whitehead précurseur des théories de l’auto-création. Revue d’Histoire des Sciences, 65(1):81–101, January 2012. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634388.

Bergandi:2012:PNA

[4542] Donato Bergandi and Patrick Blandin. De la protection de la nature au développement durable : Genèse d’un oxymore éthique et politique. Revue d’Histoire des Sciences, 65(1):103–142, January 2012. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634389.

Burnand:2012:SEL

[4543] Léonard Burnand. Le savant et l’éditeur : Les lettres de Charles Bonnet à Fortunato Bartolomeo De Felice (1766–1769). Revue d’Histoire des Sciences, 65(1):143–158, January 2012. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634390.

Triaire:2012:LBP

[4544] Dominique Triaire. Lalande, Bernoulli, Poczobut … lettres de savants de l’Ouest à des astronomes de l’Est. Revue d’Histoire des Sciences, 65(1):159–180, January 2012. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634391.

Guillaume:2012:TLA

[4545] Marcel Guillaume. Traité de logique algorithmique, ^{≪}publications des Archives Henri–Poincaré, textes et travaux, approches philosophiques en logique ^{≫}par Louis Couturat; Oliver Schlaudt; Mohsen Sakhri. Revue d’Histoire des Sciences, 65(1):181–184, January 2012. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634392.

Basso:2012:AOB

[4546] Elisabetta Basso. Analyses d’ouvrages: Sur le sommeil: Ou plutôt sur les rêves, et autres textes inédits, ^{≪}Sources en perspective ^{≫}par Gabriel Tarde; Jacqueline Carroy; Louise Salmon. Revue d’Histoire des Sciences, 65(1):184–185, January 2012. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634393.

Fauque:2012:OCO

[4547] Danielle Fauque. Aux origines de la chimie organique : Méthodes et pratiques des pharmaciens et des chimistes (1785–1835), coll. ^{≪}Carnot ≫par sacha tomic. Revue d’Histoire des Sciences, 65(1):186–188, January 2012. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634394.

Malpangotto:2012:PMB

[4548] Michela Malpangotto. Les premiers manuscrits des Theoricæ novæ planetarum de Georg Peurbach: présentation, description, évolution d’un ouvrage. Revue d’Histoire des Sciences, 65(2):339–380, 2012. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

Bakhouche:2013:TVC

[4549] Béatrice Bakhouche. La théorie de la vision chez calcidius (IV^{e} siècle) entre géométrie, médecine et philosophie. Revue d’Histoire des Sciences, 66(1):5–31, January 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634452.

Gnassounou:2013:CPL

[4550] Bruno Gnassounou. Causalité, puissance et lois de la nature chez Leibniz. Revue d’Histoire des Sciences, 66(1):33–72, January 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634453.

Morel:2013:MBE

[4551] Thomas Morel. Mathématiques et Naturphilosophie: l’exemple de la controverse entre Johann Jakob Wagner et Johann Schön (1803–1804). Revue d’Histoire des Sciences, 66(1):73–105, January 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23634454.

Segala:2013:JBB

[4552] Marco Segala. Jean-Baptiste Biot collaborateur du Mercure de France : Vulgarisation et analyse philosophique des sciences. Revue d’Histoire des Sciences, 66(1):107–136, January 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634455.

Bracco:2013:PVP

[4553] Christian Bracco and Jean-Pierre Provost. Les points de vue de Poincaré sur la ^{≪}mécanique nouvelle ^{≫}et leurs rapports à l’enseignement et à sa pratique scientifique. Revue d’Histoire des Sciences, 66(1):137–165, January 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634456.

Varenne:2013:TMC

[4554] Franck Varenne. Théorie mathématique des catégories en biologie et notion d’équivalence naturelle chez Robert Rosen. Revue d’Histoire des Sciences, 66(1):167–197, January 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634457.

Gasnier:2013:RSA

[4555] Adeline Gasnier. Riolan, savant en alchimie? Revue d’Histoire des Sciences, 66(1):199–211, January 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634458.

Maury:2013:PSF

[4556] Liliane Maury. Penser Spencer, ^{≪}La philosophie hors de soi ^{≫}par Daniel Becquemont; Dominique Ottavi. Revue d’Histoire des Sciences, 66(1):213–215, January 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634459.

Bakhouche:2013:AOB

[4557] Béatrice Bakhouche. Analyses d’ouvrages: Monks, manuscripts and sundials: The navicula in medieval England, ^{≪}History of science and medicine library ^{≫}, 13, ^{≪}Medieval and early modern science ^{≫}, 11 par Catherine Eagleton. Revue d’Histoire des Sciences, 66(1):215–217, January 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634460.

Verdet:2013:BSP

[4558] Cyril Verdet. Énergie, science et philosophie, au tournant des XIX^{e} et XX^{e} siècles, vol. 1 : L’Émergence de l’énergie dans les sciences de la nature par Danièle Ghesquier-Pourcin; Muriel Gued; Gabriel Gohau; Michel Paty. Revue d’Histoire des Sciences, 66(1):217–219, January 2013. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23634461.

Petit:2013:LDS

[4559] Caroline Petit. Les limites du soi : Immunologie et identité biologique, ≪analytiques ^{≫}, 18. par Thomas pradeu. Revue d’Histoire des Sciences, 66(1):220–222, January 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634462.

Jullien:2013:BKD

[4560] Vincent Jullien. D’al-Khwarizmi à Descartes: Études sur l’histoire des mathématiques classiques par Roshdi Rashed. Revue d’Histoire des Sciences, 66(1):222–224, January 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634463.

Morel:2013:DPS

[4561] Thomas Morel. The disordered police state : German cameralism as science and practice par Andre Wakefield. Revue d’Histoire des Sciences, 66(1):224–226, January 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23634464.

Bodenmann:2013:GLC

[4562] Siegfried Bodenmann and Anne-Lise Rey. La guerre en lettres: La controverse scientifique dans les correspondances des Lumières. Revue d’Histoire des Sciences, 66(2):233–248, juillet–décembre 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/42771774.

Goldstein:2013:RCM

[4563] Catherine Goldstein. Routine controversies: Mathematical challenges in Mersenne’s correspondence. Revue d’Histoire des Sciences, 66(2): 249–273, juillet–décembre 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

42771775.

Rey:2013:LPA

[4564] Anne-Lise Rey. L’expérience de pensée au péril de la fiction: Le cas de la correspondance entre Leibniz et Papin. Revue d’Histoire des Sciences, 66 (2):275–298, juillet–décembre 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

42771776.

Schandeler:2013:VSV

[4565] Jean-Pierre Schandeler. La violence sous le verbe académique: Un débat sur les traditions historiques à l’Académie royale des inscriptions et belles-lettres au XVIII^{e} siècle. Revue d’Histoire des Sciences, 66(2): 299–323, juillet–décembre 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

42771777.

Steinke:2013:AHP

[4566] Hubert Steinke. Albrecht von Haller, patron dans son réseau: Le rôle de la correspondance dans les controverses scientifiques. Revue d’Histoire des Sciences, 66(2):325–359, juillet–décembre 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/42771778.

Bodenmann:2013:FAV

[4567] Siegfried Bodenmann. Faux accords et vrais désaccords: Rhétoriques et stratégies de la controverse dans le réseau de correspondance de Leonhard Euler. Revue d’Histoire des Sciences, 66(2):361–394, juillet–décembre 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/42771779.

Leclercq:2013:ABF

[4568] Frédéric Leclercq. Arago, Biot et Fresnel expliquent la polarisation rotatoire. Revue d’Histoire des Sciences, 66(2):395–416, juillet–décembre 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/42771780.

Basso:2013:AOB

[4569] Elisabetta Basso. Analyses d’ouvrages: Foucault’s archaeology: Science and transformation par David Webb. Revue d’Histoire des Sciences, 66 (2):417–420, juillet–décembre 2013. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

42771781.

Rommevaux:2014:AAD

[4570] Sabine Rommevaux. Un auteur anonyme du XIVe siècle, à Oxford, lecteur de Pierre de Maricourt. (French) [An anonymous author of the 14th Century in Oxford, reader of Pierre de Maricourt]. Revue d’Histoire des Sciences, 67(1):5–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.cairn.info/

article.php?ID_ARTICLE=RHS_671_0005.

Descamps:2014:LMO

[4571] Pascal Descamps. La ligne méridienne de l’Observatoire de Paris : Une analyse des registres des observations de Cassini II de 1730 à 1755. (French) [The meridian line of the Paris Observatory: An analysis of the registers of observations of Cassini II from 1730 to 1755]. Revue d’Histoire des Sciences, 67(1):35–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.cairn.info/

article.php?ID_ARTICLE=RHS_671_0035.

Balan:2014:NTM

[4572] Bernard Balan. La naissance de la tectonique moderne en Pennsylvanie au XIXe siècle. (French) [The birth of modern tectonics in Pennsylvania in the nineteenth century]. Revue d’Histoire des Sciences, 67 (1):71–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.cairn.info/article.php?
ID_ARTICLE=RHS_671_0071.

Seignan:2014:QPD

[4573] Gérard Seignan. Quand la psychologie détaille les forces du psychisme: Épistémologie et histoire du concept d’ <<effort volontaire >>du début du XIXe siècle au milieu du XXe. (French) [When psychology details the forces of the psyche: epistomology and the history of the concept of voluntary effort from the beginning of the nineteenth century to the middle of the twentieth century]. Revue d’Histoire des Sciences, 67 (1):93–114, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.cairn.info/article.php?
ID_ARTICLE=RHS_671_0093.

Aucouturier:2014:ATE

[4574] Etienne Aucouturier. Auguste Trillat : Épidémiologie et immunologie de guerre. (French) [Auguste Trillat: Epidemiology and immunology of war]. Revue d’Histoire des Sciences, 67(1):115–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.cairn.info/article.php?ID_ARTICLE=RHS_671_0115.

Debru:2014:RCI

[4575] Claude Debru. Revue critique. l’implicite cognitif. (French) [Critical review. The cognitive implicit]. Revue d’Histoire des Sciences, 67(1): 151–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.cairn.info/article.php?
ID_ARTICLE=RHS_671_0151.

Anonymous:2014:AOa

[4576] Anonymous. Analyses d’ouvrages. (French) [Reviews of books]. Revue d’Histoire des Sciences, 67(1):165–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

cairn.info/article.php?ID_ARTICLE=RHS_671_0165.

Bourdeau:2014:I

[4577] Michel Bourdeau and Anastasios Brenner. Introduction. Revue d’Histoire des Sciences, 67(2):175–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.cairn.info/

article.php?ID_ARTICLE=RHS_672_0175.

Marion:2014:LSD

[4578] Mathieu Marion and Amirouche Moktefi. La logique symbolique en débat à Oxford à la fin du dix-neuvième siècle: Les disputes logiques de Lewis Carroll et John Cook Wilson. (French) [Symbolic logic in debate in Oxford at the end of the Nineteenth Century: The logic disputes of Lewis Carroll and John Cook Wilson]. Revue d’Histoire des Sciences, 67(2):185–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.cairn.info/article.php?
ID_ARTICLE=RHS_672_0185.

Peckhaus:2014:ESP

[4579] Volker Peckhaus. Ernst Schröder on pasigraphy. Revue d’Histoire des Sciences, 67(2):207–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.cairn.info/

article.php?ID_ARTICLE=RHS_672_0207.

Brenner:2014:RDL

[4580] Anastasios Brenner. La réception du logicisme en France en réaction à la controverse Poincaré–Russell. (French) [The reception of logic in France in reaction to the Poincaré–Russell controversy]. Revue d’Histoire des Sciences, 67(2):231–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.cairn.info/

article.php?ID_ARTICLE=RHS_672_0231.

VanAtten:2014:OQP

[4581] Mark Van Atten, Göran Sundholm, Michel Bourdeau, and Vanessa Van Atten. <<Que les principes de la logique ne sont pas fiables >>: Nouvelle traduction française annotée et commentée de l’article de 1908 de L. E. J. Brouwer. (French) [“That the principles of logic are not reliable”: new annotated and commented French translation of the 1908 article by L. E. J. Brouwer]. Revue d’Histoire des Sciences, 67 (2):257–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.cairn.info/article.php?
ID_ARTICLE=RHS_672_0257.

Roy:2014:OTN

[4582] Jean-Michel Roy. <<there is no such thing as philosophic logic >>: Le problème de l’exception wittgensteinienne. (French) [The problem of the Wittgenstein exception]. Revue d’Histoire des Sciences, 67(2): 283–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.cairn.info/article.php?
ID_ARTICLE=RHS_672_0283.

Bourdeau:2014:LCC

[4583] Michel Bourdeau. La logique à la croisée des chemins: La controverse Goblot–Rougier sur la nature de la démonstration et du raisonnement déductif (1907–1921). (French) [Logic at the crossroads: The Goblot–Rougier controversy on the nature of demonstration and deductive reasoning (1907–1921)]. Revue d’Histoire des Sciences, 67(2): 311–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.cairn.info/article.php?
ID_ARTICLE=RHS_672_0311.

Mazliak:2014:ADR

[4584] Laurent Mazliak and Marc Sage. Au-delà des réels : Émile Borel et l’approche probabiliste de la réalité. (French) [Beyond realities: Émile Borel and the probabilistic approach of reality]. Revue d’Histoire des Sciences, 67(2):331–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.cairn.info/

article.php?ID_ARTICLE=RHS_672_0331.

Anonymous:2014:AOb

[4585] Anonymous. Analyses d’ouvrages. (French) [Reviews of books]. Revue d’Histoire des Sciences, 67(2):359–??, ???? 2014. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

cairn.info/article.php?ID_ARTICLE=RHS_672_0359.

LeNoxaic:2015:CBP

[4586] Armand Le Noxaïc. Comment Blaise Pascal a pu envisager et réaliser l’expérience des liqueurs de Rouen. (French) [How Blaise Pascal was able to envisage and realize the liqueur experiment of Rouen]. Revue d’Histoire des Sciences, 68(1):5–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.cairn.

info/article.php?ID_ARTICLE=RHS_681_0005.

Loison:2015:PRT

[4587] Laurent Loison. Pourquoi refuser la théorie cellulaire ? le projet d’une anatomie chimique chez Charles Robin (1821–1885). (French) [Why deny cell theory? The chemical anatomy project of Charles Robin (1821– 1885)]. Revue d’Histoire des Sciences, 68(1):23–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.cairn.info/article.php?ID_ARTICLE=RHS_681_0023.

Descamps:2015:DNE

[4588] Pascal Descamps. La découverte de Neptune: Entre triomphe et camouflet. (French) [The discovery of Neptune: Between triumph and explosive device]. Revue d’Histoire des Sciences, 68(1):47–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.cairn.info/article.php?ID_ARTICLE=RHS_681_0047.

Schut:2015:LDM

[4589] Pierre-Olaf Schut and Matthieu Delalandre. L’échec d’une discipline: Montée et déclin de la spéléologie en France (1888–1978). (French) [The failure of a discipline: the rise and decline of speleology in France (1888– 1978)]. Revue d’Histoire des Sciences, 68(1):81–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.cairn.info/article.php?ID_ARTICLE=RHS_681_0081.

Bracco:2015:ESD

[4590] Christian Bracco. L’environnement scientifique du jeune Albert Einstein : La période milanaise (1899–1901). (French) [The scientific environment of the young Albert Einstein: the Milan period (1899–1901)]. Revue d’Histoire des Sciences, 68(1):109–144, January 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.cairn.info/article.php?ID_ARTICLE=RHS_681_0109.

Savaton:2015:EPS

[4591] Pierre Savaton. Les exercices pratiques de sciences naturelles de 1902 aux années 1930 : Des textes aux pratiques. (French) [Practical exercises in the natural sciences from 1902 to the 1930s: From texts to practices]. Revue d’Histoire des Sciences, 68(1):145–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.cairn.info/article.php?ID_ARTICLE=RHS_681_0145.

Lamy:2015:SHS

[4592] Jérôme Lamy and Arnaud Saint-Martin. La sociologie historique des sciences et techniques : Essai de généalogie conceptuelle et d’histoire configurationnelle. (French) [The historical sociology of science and technology: An essay on conceptual genealogy and conflict history]. Revue d’Histoire des Sciences, 68(1):175–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.cairn.info/

article.php?ID_ARTICLE=RHS_681_0175.

Kahn:2015:NJT

[4593] Didier Kahn. Nécrologie : Joachim Telle (1939–2013). (French) [Obituary: Joachim Telle (1939–2013)]. Revue d’Histoire des Sciences, 68 (1):215–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.cairn.info/article.php?
ID_ARTICLE=RHS_681_0215.

Anonymous:2015:AOa

[4594] Anonymous. Analyses d’ouvrages. (French) [Reviews of books]. Revue d’Histoire des Sciences, 68(1):225–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

cairn.info/article.php?ID_ARTICLE=RHS_681_0225.

Bruneau:2015:I

[4595] Olivier Bruneau and Irène Passeron. Introduction. Revue d’Histoire des Sciences, 68(2):259–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.cairn.info/

article.php?ID_ARTICLE=RHS_682_0259.

Schmit:2015:DJJ

[4596] Christophe Schmit. Les dynamiques de Jean-Jacques Dortous de Mairan. (French) [The dynamics of Jean-Jacques Dortous de Mairan]. Revue d’Histoire des Sciences, 68(2):281–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

cairn.info/article.php?ID_ARTICLE=RHS_682_0281.

LeGars:2015:DMT

[4597] Stéphane Le Gars. Dortous de Mairan et la théorie des aurores polaires : Trajectoire et circulation d’une idée, de 1733 à 1933. (French) [Dortous of Mairan and the theory of the polar aurora: Trajectory and circulation of an idea, from 1733 to 1933]. Revue d’Histoire des Sciences, 68 (2):311–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.cairn.info/article.php?
ID_ARTICLE=RHS_682_0311.

Baskevitch:2015:PDS

[4598] François Baskevitch. La propagation du son chez Dortous de Mairan (1737) : Des particules d’air de différentes élasticités. (French) [Sound propagation in Dortous de Mairan (1737): Air particles of different elasticities]. Revue d’Histoire des Sciences, 68(2):335–??, ???? 2015. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.cairn.info/article.php?ID_ARTICLE=RHS_682_0335.

LeLay:2015:DG

[4599] Colette Le Lay and Frédérique Rémy. La dissertation sur la glace (1749). (French) [The dissertation on ice (1749)]. Revue d’Histoire des Sciences, 68(2):359–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.cairn.info/article.php?
ID_ARTICLE=RHS_682_0359.

Chabot:2015:REO

[4600] Hugues Chabot and Nicolas Breteil. La réception épistémologique de l’œuvre scientifique de Dortous de Mairan dans l’Encyclopédie. (French) [The epistemological reception of the scientific work of Dortous de Mairan in the booktitle Encyclopedia]. Revue d’Histoire des Sciences, 68(2):

375–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.cairn.info/article.php?
ID_ARTICLE=RHS_682_0375.

Kachur:2015:BDM

[4601] Iryna Kachur. La bibliothèque de Dortous de Mairan et ses livres retrouvés. (French) [The library of Dortous de Mairan and its rediscovered books]. Revue d’Histoire des Sciences, 68(2):405–??, ???? 2015. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.cairn.info/article.php?ID_ARTICLE=RHS_682_0405.

Frank:2015:CCQ

[4602] Martin Frank. The curious case of QP.6: The Reception of Archimedes’ Mechanics by Federico Commandino and Guidobaldo dal Monte. Revue d’Histoire des Sciences, 68(2):419–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

cairn.info/article.php?ID_ARTICLE=RHS_682_0419.

Parmentier:2015:LV

[4603] Marc Parmentier. Leibniz et le virtuel. (French) [Leibniz and the virtual]. Revue d’Histoire des Sciences, 68(2):447–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.cairn.info/article.php?ID_ARTICLE=RHS_682_0447.

VilaValls:2015:ECP

[4604] Adrien Vila Valls. Entre causalisme et positivisme : La réception épistémologique de la mécanique quantique par les théoriciens français de la génération de l’entre-deux-guerres. (French) [Between causalism and positivism: The epistemological reception of quantum mechanics by the French theorists of the inter-war generation]. Revue d’Histoire des Sciences, 68(2):475–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.cairn.info/

article.php?ID_ARTICLE=RHS_682_0475.

Anonymous:2015:AOb

[4605] Anonymous. Analyses d’ouvrages. (French) [Reviews of books]. Revue d’Histoire des Sciences, 68(2):509–??, ???? 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

cairn.info/article.php?ID_ARTICLE=RHS_682_0509.

Anonymous:2016:FMa

[4606] Anonymous. Front matter. Revue d’Histoire des Sciences, 69(1):??, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL https://www.jstor.org/stable/26633446.

Fauque:2016:IOM

[4607] Danielle Fauque. Introduction: Aux origines de la Maison de la Chimie. (French) [Introduction: The origins of the House of Chemistry]. Revue d’Histoire des Sciences, 69(1):5–18, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https:

//www.jstor.org/stable/26633447.

Lette:2016:CCR

[4608] Michel Letté. Chimie, chimistes et rationalisation sous les auspices du ministre du Commerce et de l’Industrie Étienne Clémentel (1917–1919). (French) [Chemistry, chemists and rationalization under the auspices of Minister of Trade and Industry Étienne Clément (1917–1919)]. Revue d’Histoire des Sciences, 69(1):19–40, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https:

//www.jstor.org/stable/26633448.

Fauque:2016:DAC

[4609] Danielle Fauque. La documentation au cœur de la réorganisation de la chimie dans l’entre-deux-guerres: Rôle des sociétés savantes et institutions françaises dans le contexte international. (French) [Documentation at the heart of the reorganization of chemistry in the interwar period: Role of French learned societies and institutions in the international context]. Revue d’Histoire des Sciences, 69(1):41–76, January/ June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL https://www.jstor.org/stable/26633449.

Fox:2016:SCD

[4610] Robert Fox. Science, celebrity, diplomacy: The Marcellin Berthelot centenary, 1927. Revue d’Histoire des Sciences, 69(1):77–116, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL https://www.jstor.org/stable/26633450.

Langlinay:2016:AHC

[4611] Erik Langlinay. Albin Haller, chimie et politique (1918–1925). (French) [Albin Haller, chemistry and politics (1918–1925)]. Revue d’Histoire des Sciences, 69(1):117–136, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.jstor.

org/stable/26633451.

Benoit:2016:CCS

[4612] Serge Benoit and Gérard Emptoz. Les collections de chimie de la Société d’encouragement pour l’industrie nationale. (French) [The chemistry collections of the Society for Encouragement of National Industry]. Revue d’Histoire des Sciences, 69(1):137–152, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https:

//www.jstor.org/stable/26633452.

Pelucchi:2016:CDC

[4613] Stéphane Pelucchi. La collection du cabinet d’histoire naturelle de Lavoisier: Sa place dans son œuvre scientifique. (French) [The collection of Lavoisier’s natural history cabinet: Its place in its scientific work]. Revue d’Histoire des Sciences, 69(1):153–170, January/June 2016. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

https://www.jstor.org/stable/26633453.

BenSaad:2016:BRL

[4614] Meyssa Ben Saâd. Book review: Le Livre des animaux d’Al–Jâhiz. ^{≪} Acteurs de la science ^{≫} by Ahmed Aarab, Philippe Lherminier. Revue d’Histoire des Sciences, 69(1):171–172, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https:

//www.jstor.org/stable/26633454.

Chabot:2016:BRO

[4615] Hugues Chabot. Book review: Œuvres complètes, série V : ^{≪} Correspondance générale ^{≫}, vol. 2 : Correspondance générale, 1741–1752 by Jean Le Rond d’Alembert, Irène Passeron, Jean-Daniel Candaux, Alain Cernuschi, Frédéric Chambat. Revue d’Histoire des Sciences, 69(1):172–174, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL https://www.jstor.org/stable/26633455.

Chandelier:2016:BRL

[4616] Cédric Chandelier. Book review: L’Épistémologie française, 1830–1970, 2e édition [1re éd.] ^{≪} Sciences et philosophie ^{≫} by Michel Bitbol, Jean Gayon, Paula Quinon, Manuel Quinon. Revue d’Histoire des Sciences, 69(1):174–176, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.jstor.org/stable/

26633456.

Chabert:2016:BRA

[4617] Jean-Luc Chabert and Christian Gilain. Book review: André-Louis Cholesky: Mathematician, topographer and army officer by Claude Brezinski, Dominique Tournès. Revue d’Histoire des Sciences, 69(1): 176–177, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.jstor.org/stable/

26633457.

Poirier:2016:BRL

[4618] Jacques Poirier. Book review: L’Exploration cérébrale : Histoire récente et nouveaux outils coll.^{≪} Histoire des sciences^{≫} by Jean-Claude Dupont, Céline Cherici. Revue d’Histoire des Sciences, 69(1):178–179, January/ June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL https://www.jstor.org/stable/26633458.

Leduc:2016:BRL

[4619] Christian Leduc. Book review: Les Lumières et Leibniz : Avant la publication des Nouveaux essais sur l’entendement humain ^{≪} Les Dixhuitièmes siècles ^{≫} by Claire Fauvergue. Revue d’Histoire des Sciences, 69(1):179–180, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.jstor.org/stable/

26633459.

Tarfaoui:2016:BRF

[4620] Sakina Tarfaoui. Book review: Femmes de sciences de l’Antiquité au XIXe siècle : Réalités et représentations, coll. ^{≪} Histoire et philosophie des sciences ^{≫} by Adeline Gargam, Adeline Gargam, Patrice Bret, Judith Zinsser. Revue d’Histoire des Sciences, 69(1):181–182, January/ June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL https://www.jstor.org/stable/26633460.

dHombres:2016:BRP

[4621] Emmanuel d’Hombres. Book review: Précis de philosophie de la biologie, ≪ Philosophie des sciences^{≫} by Thierry Hoquet, Francesca Merlin. Revue d’Histoire des Sciences, 69(1):182–185, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https:

//www.jstor.org/stable/26633461.

Teissier:2016:BRO

[4622] Pierre Teissier. Book review: Objects of chemical inquiry by Ursula Klein, Carsten Reinhardt. Revue d’Histoire des Sciences, 69(1):185–187, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL https://www.jstor.org/stable/26633462.

Bertin:2016:BRI

[4623] Pascal Bertin. Book review: Infini, logique, géométrie, coll. ^{≪} Mathesis ≫ by Paolo Mancosu. Revue d’Histoire des Sciences, 69(1):187–189, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL https://www.jstor.org/stable/26633463.

Cherici:2016:BRD

[4624] Céline Cherici. Book review: Les Déséquilibres de l’amour : La genèse du concept de perversion sexuelle de la Révolution française à Freud, coll. ^{≪} Philosophie, anthropologie, psychologie ^{≫} by Julie Mazaleigue-Labaste. Revue d’Histoire des Sciences, 69(1):190–191, January/June 2016. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

https://www.jstor.org/stable/26633464.

Kozluk:2016:BRN

[4625] Magdalena Koźluk. Book review: Nicolas Dortoman et Balaruc : La médecine thermale à la Renaissance, ^{≪} Série Histoire ^{≫} by Jean Meyers, Brigitte Pérez–Jean. Revue d’Histoire des Sciences, 69(1):191–193, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL https://www.jstor.org/stable/26633465.

Bimbot:2016:BRM

[4626] René Bimbot. Book review: Molecular beam epitaxy : A short history by John W. Orton, Tom Foxon. Revue d’Histoire des Sciences, 69 (1):193–195, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.jstor.org/stable/

26633466.

Bellis:2016:BRS

[4627] Delphine Bellis. Book review: From sight to light : The passage from ancient to modern optics by A. Mark Smith. Revue d’Histoire des Sciences, 69(1):195–197, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.jstor.org/stable/

26633467.

Fauque:2016:BRH

[4628] Danielle Fauque. Book review: Une histoire de la chimie du solide : Synthèses, formes, identités, coll. ^{≪} Intersciences ^{≫} by Pierre Tessier, Jacques Livage. Revue d’Histoire des Sciences, 69(1):197–200, January/ June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL https://www.jstor.org/stable/26633468.

Nicolaidis:2016:BRL

[4629] Efthymios Nicolaïdis. Book review: L’Aventure spatiale française : De 1945 à la naissance d’Ariane, coll. ^{≪} Histoire des sciences ^{≫} by Philippe Varnoteaux, Jacques Blamont. Revue d’Histoire des Sciences, 69(1): 200–202, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.jstor.org/stable/

26633469.

Anonymous:2016:BMa

[4630] Anonymous. Back matter. Revue d’Histoire des Sciences, 69(1):??, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL https://www.jstor.org/stable/26633470.

Anonymous:2016:Aa

[4631] Anonymous. Accessibility. Revue d’Histoire des Sciences, 69(1):??, January/June 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

Anonymous:2016:FMb

[4632] Anonymous. Front matter. Revue d’Histoire des Sciences, 69(2):??, July/ December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL https://www.jstor.org/stable/26384005.

Maffioli:2016:LVS

[4633] Cesare S. Maffioli. Leonardo da Vinci et le savoir des ingénieurs : Aménagement et science des eaux à Milan aux environs de 1500. (French) [Leonardo da Vinci and the knowledge of engineers: Planning and water science in Milan around 1500]. Revue d’Histoire des Sciences, 69 (2):209–243, July/December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.jstor.org/stable/

26384006.

Taussig:2016:VFS

[4634] Sylvie Taussig. Vrai et faux soleil(s) chez Gassendi. (French) [True and false sun(s) in Gassendi]. Revue d’Histoire des Sciences, 69(2):245–258, July/December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL https://www.jstor.org/stable/26384007.

Rabouin:2016:MUA

[4635] David Rabouin. Mathesis universalis et algèbre générale dans les Regulae ad directionem ingenii de Descartes. (French) [Mathesis universalis and general algebra in the Regulae ad directionem ingenii of Descartes]. Revue d’Histoire des Sciences, 69(2):259–309, July/December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https:

//www.jstor.org/stable/26384008.

Dubouclez:2016:LOG

[4636] Olivier Dubouclez. L’^{≪} ordre géométrique ^{≫} et le dispositif argumentatif des méditations métaphysiques de René Descartes. (French) [“Geometrical order” and the argumentative device of the metaphysical meditations of René Descartes]. Revue d’Histoire des Sciences, 69(2):311–334, July/ December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL https://www.jstor.org/stable/26384009.

Dayrat:2016:HLD

[4637] Benoît Dayrat. Henri de Lacaze-Duthiers and the creation of the Laboratoire de zoologie expérimentale, Roscoff, France. (French) [Henri de Lacaze-Duthiers and the creation of the Laboratory of Experimental Zoology, Roscoff, France]. Revue d’Histoire des Sciences, 69(2):335–368, July/December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL https://www.jstor.org/stable/26384010.

Herring:2016:ESM

[4638] Emily Herring. Des évolutionnismes sans mécanisme : Les néolamarckismes métaphysiques d’Albert Vandel (1894–1980) et Pierre-Paul Grassé (1895–1985). (French) [Evolutionaries without mechanism: The metaphysical neolamarckisms of Albert Vandel (1894–1980) and Pierre-Paul Grassé (1895–1985)]. Revue d’Histoire des Sciences, 69(2):369–398, July/ December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL https://www.jstor.org/stable/26384011.

Lautman:2016:ACS

[4639] Jacques Lautman. Adieu à Claire Salomon-Bayet (1932–2016). (French) [Farewell to Claire Salomon-Bayet (1932–2016)]. Revue d’Histoire des Sciences, 69(2):399–401, July/December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.jstor.

org/stable/26384012.

Lorenat:2016:BRE

[4640] Jemma Lorenat. Book review: Éléments d’une biographie de l’espace géométrique, ^{≪} Histoires de géométries ^{≫} by Lise Bioesmat-Martagon. Revue d’Histoire des Sciences, 69(2):403–405, July/December 2016. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

https://www.jstor.org/stable/26384013.

Ginoux:2016:BRT

[4641] Jean-Marc Ginoux. Book review: Les Textes fondateurs de l’épistémologie française : Duhem, Poincaré, Brunschvicg et autres philosophes. ^{≪} Hermann Philosophie ^{≫} by Anastasios Brenner. Revue d’Histoire des Sciences, 69(2):405–407, July/December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.jstor.

org/stable/26384014.

Lancelot:2016:BRO

[4642] Mathilde Lancelot. Book review: Œuvres complètes, t. IV: Résistance, philosophie biologique et histoire des sciences (1940–1965). ^{≪} Bibliothèque des textes philosophiques ^{≫} by Georges Canguilhem, Camille Limoges. Revue d’Histoire des Sciences, 69(2):407–409, July/December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL https://www.jstor.org/stable/26384015.

Mehl:2016:BRR

[4643] Édouard Mehl. Book review: De revolutionibus orbium cœlestium/Des révolutions des orbes célestes, coll.^{≪} Science et humanisme^{≫}, n o 11; vol. I : Introduction, vol. II : Texte et traduction, vol. III : Notes, appendice, dossier iconographique et index général by Nicolas Copernic, Michel-Pierre Lerner, Alain-Philippe Segonds, Concetta Luna, Isabelle Pantin, Denis Savoie, M.-P. Lerner, A.-P. Segonds, Jean-Pierre Verdet, C. Luna, J.-P. Verdet, D. Savoie, Michel Toulmonde. Revue d’Histoire des Sciences, 69(2):409–411, July/December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.jstor.

org/stable/26384016.

Ducourant:2016:BRQ

[4644] Samuel Ducourant. Book review: Quelle révolution scientifique? Les sciences de la vie dans la querelle des Anciens et des Modernes, XVI e–XVIIIe siècles, ^{≪} Les collections de la République des lettres. Études ≫ by Pascal Duris. Revue d’Histoire des Sciences, 69(2):411–413, July/ December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL https://www.jstor.org/stable/26384017.

Debru:2016:BRA

[4645] Claude Debru. Book review: 350 ans de l’Académie des sciences: Une compagnie en son siècle by Pascal Griset, Florence Greffe, Catherine Bréchignac, Jean-François Bach. Revue d’Histoire des Sciences, 69(2): 413–414, July/December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.jstor.org/stable/

26384018.

Moatti:2016:BRD

[4646] Alexandre Moatti. Book review: Dompter Prométhée : Technologies et socialismes à l’âge romantique (1820–1870), ^{≪} Les Cahiers de la MSHE Ledoux. Archives de l’imaginaire social ^{≫} by François Jarrige. Revue d’Histoire des Sciences, 69(2):414–416, July/December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https:

//www.jstor.org/stable/26384019.

Peiffer:2016:BRJ

[4647] Jeanne Peiffer. Book review: Les Jombert : Une famille de libraires parisiens dans l’Europe des Lumières, 1680–1824, ^{≪} Bibliothèque des Lumières ^{≫}, vol. 84 by Greta Kaucher. Revue d’Histoire des Sciences, 69 (2):416–418, July/December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.jstor.org/stable/

26384020.

Anonymous:2016:TG

[4648] Anonymous. Table générale. Revue d’Histoire des Sciences, 69(2): 419–426, July/December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.jstor.org/stable/

26384021.

Anonymous:2016:BMb

[4649] Anonymous. Back matter. Revue d’Histoire des Sciences, 69(2):??, July/ December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL https://www.jstor.org/stable/26384022.

Anonymous:2016:Ab

[4650] Anonymous. Accessibility. Revue d’Histoire des Sciences, 69(2):??, July/ December 2016. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic).

Jami:2017:IF

[4651] Catherine Jami. Introduction. Revue d’Histoire des Sciences, 70(1): 7–13, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL https://www.cairn.info/article.php?ID_
ARTICLE=RHS_701_0007.

Chu:2017:JIC

[4652] Longfei Chu. From the Jesuits’ treatises to the imperial compendium: The appropriation of the Tychonic system in seventeenth and eighteenthcentury China. Revue d’Histoire des Sciences, 70(1):15–46, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL https://www.cairn.info/article.php?ID_ARTICLE=
RHS_701_0015.

Cams:2017:CIS

[4653] Mario Cams. Converging interests and scientific circulation between Paris and Beijing (1685–1735): The path towards a new Qing cartographic practice. Revue d’Histoire des Sciences, 70(1):47–78, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https:

//www.cairn.info/article.php?ID_ARTICLE=RHS_701_0047.

Vigouroux:2017:SAP

[4654] Mathias Vigouroux. The surgeon’s acupuncturist: Philipp Franz von Siebold’s encounter with Ishizaka Sōtetsu and nineteenth century Japanese acupuncture. Revue d’Histoire des Sciences, 70(1):79–108, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL https://www.cairn.info/article.php?ID_ARTICLE=
RHS_701_0079.

Cousin:2017:MGJ

[4655] Marion Cousin. Les manuels de géométrie dans le Japon de l’ère Meiji (1868–1912): Des témoins des révolutions du livre et de l’enseignement des mathématiques. (French) [The textbooks of geometry in Japan of the Meiji era (1868–1912): Witnesses to the revolutions of the book and the teaching of mathematics]. Revue d’Histoire des Sciences, 70 (1):109–145, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL https://www.cairn.info/article.php?
ID_ARTICLE=RHS_701_0109.

Terrail-Lormel:2017:EAT

[4656] Sarah Terrail-Lormel. De l’éreuthophobie au taijinkyōfushō: Circulation et élaboration d’un diagnostic psychiatrique entre Europe, États-Unis et Japon (décennies 1890–1930). (French) [From ereuthophobia to taijinkyōfushō: Circulation and development of a psychiatric diagnosis between Europe, United States and Japan (decades 1890–1930)]. Revue d’Histoire des Sciences, 70(1):147–174, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.

cairn.info/article.php?ID_ARTICLE=RHS_701_0147.

Fedi:2017:PES

[4657] Laurent Fedi. La psychologie de l’esprit scientifique chez Bachelard et ses prédécesseurs. (French) [The psychology of the scientific spirit in Bachelard and his predecessors]. Revue d’Histoire des Sciences, 70 (1):175–216, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL https://www.cairn.info/article.php?
ID_ARTICLE=RHS_701_0175.

LeRoux:2017:LSD

[4658] Benjamin Le Roux. ^{≪} lui seul doit être glorifié ^{≫}, science et religion dans l’œuvre d’Henri Devaux (1862–1956). (French) [“He alone must be glorified”, science and religion in the work of Henri Devaux (1862– 1956)]. Revue d’Histoire des Sciences, 70(1):217–236, ???? 2017. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

https://www.cairn.info/article.php?ID_ARTICLE=RHS_701_0217.

Anonymous:2017:AOFa

[4659] Anonymous. Analyses d’ouvrages. (French) [Analysis of works]. Revue d’Histoire des Sciences, 70(1):237–248, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.

cairn.info/article.php?ID_ARTICLE=RHS_701_0237.

Basso:2017:IPS

[4660] Elisabetta Basso and Mireille Delbraccio. Introduction : La psychiatrie en ses archives, entre histoire et épistémologie. (French) [Introduction: Psychiatry in its archives, between history and epistemology]. Revue d’Histoire des Sciences, 70(2):255–273, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.

cairn.info/article.php?ID_ARTICLE=RHS_702_0255.

Engstrom:2017:APP

[4661] Eric J. Engstrom, Wolfgang Burgmair, and Matthias M. Weber. Archiving psychiatry’s past: Emil Kraepelin’s archive and papers. Revue d’Histoire des Sciences, 70(2):275–297, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.

cairn.info/article.php?ID_ARTICLE=RHS_702_0275.

Priani:2017:AAA

[4662] Egidio Priani. Les archives de l’ancien asile psychiatrique de San Servolo (Venise), 1840–1877: Trames, classifications, sujets. (French) [The archives of the former psychiatric asylum of San Servolo (Venice), 1840– 1877: Patterns, classifications, subjects]. Revue d’Histoire des Sciences, 70(2):299–326, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL https://www.cairn.info/article.php?
ID_ARTICLE=RHS_702_0299.

Delille:2017:UMN

[4663] Emmanuel Delille. Usages et mésusages de la notion de précurseur en histoire des sciences et des techniques, en particulier en histoire de la psychiatrie: À propos d’un échange épistolaire entre Henri Ellenberger et Georges Canguilhem (1967). (French) [Uses and misconceptions of the notion of precursor in the history of science and technology, in particular in the history of psychiatry: about a pistol exchange between Henri Ellenberger and Georges Canguilhem (1967)]. Revue d’Histoire des Sciences, 70(2):327–350, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL https://www.cairn.info/article.php?
ID_ARTICLE=RHS_702_0327.

Delille:2017:APH

[4664] Emmanuel Delille. Une archive pour l’histoire des sociabilités savantes au xxe siècle : Georges Lantéri-Laura (1930–2004) et l’Encyclopédie médico-chirurgicale. (French) [An archive for the history of learned societies in the twentieth century: Georges Lantéri-Laura (1930–2004) and the Medico-surgical Encyclopaedia]. Revue d’Histoire des Sciences, 70 (2):351–377, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL https://www.cairn.info/article.php?
ID_ARTICLE=RHS_702_0351.

Queyroux:2017:CDD

[4665] Nathalie Queyroux. Le centre documentaire du CAPHÉS : Un lieu d’archives. (French) [The CAPHÉS documentary center: a place of archives]. Revue d’Histoire des Sciences, 70(2):379–396, ???? 2017. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

https://www.cairn.info/article.php?ID_ARTICLE=RHS_702_0379.

Stamenkovic:2017:RPP

[4666] Philippe Stamenkovic. Retour sur la philosophie de la physique quantique d’ernst cassirer. (French) [Back to Ernst Cassirer’s philosophy of quantum physics]. Revue d’Histoire des Sciences, 70(2):397–426, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL https://www.cairn.info/article.php?ID_ARTICLE=
RHS_702_0397.

Anonymous:2017:AOFb

[4667] Anonymous. Analyses d’ouvrages. (French) [Analysis of works]. Revue d’Histoire des Sciences, 70(2):427–451, ???? 2017. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.

cairn.info/article.php?ID_ARTICLE=RHS_702_0427.

Ribas:2018:APA

[4668] Marie-Noëlle Ribas. Aristote, père de l’astrophysique scientifique. (French) [Aristotle, father of scientific astrophysics]. Revue d’Histoire des Sciences, 71(1):5–24, ???? 2018. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.cairn.info/

article.php?ID_ARTICLE=RHS_711_0005.

Boarini:2018:LPV

[4669] Serge Boarini. L’inoculation de la petite vérole à Lyon : de la condamnation à l’acceptation (1779–1811). (French) [The inoculation of the little woman in Lyon: from condemnation to acceptance (1779–1811)]. Revue d’Histoire des Sciences, 71(1):25–48, ???? 2018. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.

cairn.info/article.php?ID_ARTICLE=RHS_711_0025.

daSilva:2018:CCC

[4670] Matheus Alves Duarte da Silva. Compétition, controverses et cultures de microbes: le développement du sérum antipesteux entre Paris et l’Inde (1894–1899). (French) [Competition, controversies and cultures of microbes: the development of the anti-plague serum between Paris and India (1894–1899)]. Revue d’Histoire des Sciences, 71(1):49– 77, ???? 2018. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL https://www.cairn.info/article.php?ID_
ARTICLE=RHS_711_0049.

Dolbeault:2018:EUB

[4671] Joël Dolbeault. Einstein et l’univers-bloc. (French) [Einstein and the block universe]. Revue d’Histoire des Sciences, 71(1):79–109, ???? 2018. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL https://www.cairn.info/article.php?ID_ARTICLE=
RHS_711_0079.

Anonymous:2018:AOFa

[4672] Anonymous. Analyses d’ouvrages. (French) [Analysis of works]. Revue d’Histoire des Sciences, 71(1):111–140, ???? 2018. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.

cairn.info/article.php?ID_ARTICLE=RHS_711_0111.

Loison:2018:IBP

[4673] Laurent Loison. Introduction : Biologie, philosophie de la biologie et histoire de la biologie chez Georges Canguilhem (1904–1995). (French) [Introduction: Biology, philosophy of biology and history of biology in Georges Canguilhem (1904–1995)]. Revue d’Histoire des Sciences, 71 (2):147–154, ???? 2018. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL https://www.cairn.info/article.php?
ID_ARTICLE=RHS_712_0147.

Gabel:2018:BRH

[4674] Isabel Gabel. La biologie, la réflexivité et l’histoire : Réinscrire Canguilhem dans son milieu. (French) [Biology, reflexivity and history: Reregister Canguilhem in its midst]. Revue d’Histoire des Sciences, 71 (2):155–177, ???? 2018. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL https://www.cairn.info/article.php?
ID_ARTICLE=RHS_712_0155.

Diez:2018:CAG

[4675] Ivan Moya Diez. Canguilhem avec Goldstein : De la normativité de la vie à la normativité de la connaissance. (French) [Canguilhem with Goldstein: From the normativity of life to the normativity of knowledge]. Revue d’Histoire des Sciences, 71(2):179–204, ???? 2018. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

https://www.cairn.info/article.php?ID_ARTICLE=RHS_712_0179.

Methot:2018:GCF

[4676] Pierre-Olivier Méthot. Georges Canguilhem et le ^{≪} problème de l’évolution ^{≫} dans le normal et le pathologique. (French) [Georges Canguilhem and the “problem of evolution” in the normal and the pathologic]. Revue d’Histoire des Sciences, 71(2):205–241, ???? 2018. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https:

//www.cairn.info/article.php?ID_ARTICLE=RHS_712_0205.

Vagelli:2018:SMC

[4677] Matteo Vagelli. De la science des monstres: Canguilhem et la tératologie expérimentale d’Étienne Wolff. (French) [From the science of monsters: Canguilhem and the experimental teratology of Étienne Wolff]. Revue d’Histoire des Sciences, 71(2):243–270, ???? 2018. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.

cairn.info/article.php?ID_ARTICLE=RHS_712_0243.

Loison:2018:EPG

[4678] Laurent Loison. Un enthousiasme paradoxal ? Georges Canguilhem et la biologie moléculaire (1966–1973). (French) [Paradoxical enthusiasm? Georges Canguilhem and molecular biology (1966–1973)]. Revue d’Histoire des Sciences, 71(2):271–300, ???? 2018. CODEN RHSAAM.

ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.

cairn.info/article.php?ID_ARTICLE=RHS_712_0271.

Debru:2018:PBC

[4679] Claude Debru. Philosophie et biologie : La connaissance de la vie et les enseignements de Canguilhem à la faculté des lettres de Strasbourg (1941–1948). (French) [Philosophy and biology: Knowledge of life and the teachings of Canguilhem at the Faculty of Letters of Strasbourg (1941– 1948)]. Revue d’Histoire des Sciences, 71(2):301–310, ???? 2018. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

https://www.cairn.info/article.php?ID_ARTICLE=RHS_712_0301.

Huneman:2018:JGF

[4680] Philippe Huneman. Jean Gayon (1949–2018). Revue d’Histoire des Sciences, 71(2):311–317, ???? 2018. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.cairn.info/

article.php?ID_ARTICLE=RHS_712_0311.

Anonymous:2018:AOFb

[4681] Anonymous. Analyses d’ouvrages. (French) [Analysis of works]. Revue d’Histoire des Sciences, 71(2):319–350, ???? 2018. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.

cairn.info/article.php?ID_ARTICLE=RHS_712_0319.

DeRisi:2019:I

[4682] Vincenzo De Risi and David Rabouin. Introduction. Revue d’Histoire des Sciences, 72(1):5–9, ???? 2019. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.cairn.info/

article.php?ID_ARTICLE=RHS_721_0005.

Garber:2019:DLE

[4683] Daniel Garber. La dynamique de Leibniz est-elle compatible avec sa monadologie ?. (French) [Is Leibniz’s dynamics compatible with his monadology?]. Revue d’Histoire des Sciences, 72(1):11–30, ???? 2019. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https:

//www.cairn.info/article.php?ID_ARTICLE=RHS_721_0011.

Rey:2019:EMD

[4684] Anne-Lise Rey. Les enjeux métaphysiques de la dynamique. (French) [The metaphysical challenges of dynamics]. Revue d’Histoire des Sciences, 72(1):31–37, ???? 2019. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.cairn.info/

article.php?ID_ARTICLE=RHS_721_0031.

Duchesneau:2019:RPA

[4685] François Duchesneau. Le recours aux principes architectoniques dans la Dynamica de Leibniz. (French) [The use of architectural principles in Leibniz’s Dynamica]. Revue d’Histoire des Sciences, 72(1):39–62, ???? 2019. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL https://www.cairn.info/article.php?ID_ARTICLE=
RHS_721_0039.

Adomaitis:2019:EHG

[4686] Laurynas Adomaitis. Equivalence of hypotheses and Galilean censure in Leibniz: A conspiracy or a way to moderate censure? Revue d’Histoire des Sciences, 72(1):63–85, ???? 2019. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.cairn.info/

article.php?ID_ARTICLE=RHS_721_0063.

Gibier:2019:DPL

[4687] Mathieu Gibier. Dynamique et pneumatique : Leibniz face aux expériences paradoxales de Huygens. (French) [Dynamics and pneumatics: Leibniz facing the paradoxical experiences of Huygens]. Revue d’Histoire des Sciences, 72(1):87–135, ???? 2019. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.

cairn.info/article.php?ID_ARTICLE=RHS_721_0087.

Costa:2019:LCD

[4688] Andrea Costa and Enrico Pasini. L’édition critique de la Dynamica de potentia seu de legibus naturæ corporeæ de G. W. Leibniz. (French) [The critical edition of Dynamica de potentia seu de legibus naturæ corporeæ by G. W. Leibniz]. Revue d’Histoire des Sciences, 72(1):137– 161, ???? 2019. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL https://www.cairn.info/article.php?ID_
ARTICLE=RHS_721_0137.

Loison:2019:OCG

[4689] Laurent Loison. Les Œuvres complètes de Georges Canguilhem. (French) [The complete works of Georges Canguilhem]. Revue d’Histoire des Sciences, 72(1):163–169, ???? 2019. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.cairn.info/

article.php?ID_ARTICLE=RHS_721_0163.

Anonymous:2019:AOF

[4690] Anonymous. Analyses d’ouvrages. (French) [Analysis of works]. Revue d’Histoire des Sciences, 72(1):171–220, ???? 2019. CODEN RHSAAM.

REFERENCES 894 ISSN 0151-4105 (print), 1969-6582 (electronic). URL https://www.

cairn.info/article.php?ID_ARTICLE=RHS_721_0171.